The MITSFS Titledex  September 24, 2007  Page 1
TITLE  AUTHOR  CODE
"A" COMME ANDROMEDE [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  FP
A FOR ANDROMEDA [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  H:2,P
A FOR ANYTHING  KNIGHT, DAMON  C/P,SFWA-P,P,H
A LA LUNA EN UN SEGUNDO [ENVIADO SECRETO 81]  CAUDETT, FRANK  FP
"A" MEN  LE PAGE, RAND  SR-P
A. MERRITT: REFLECTIONS IN THE MOON POOL  MOSKOWITZ, SAM  R/REF-H
AARDE VALT AAN!, DE (DUTCH) [@PERRY RHODAN #4]  SCHEER, K. H./DARLTON, CLARK  FPR
AARDVARK AFFAIR, THE [@BIG BRAIN #1]  BRANDNER, GARY  C/P:2,P
AARN MUNRO IL GIOVIANO  CAMPBELL, JOHN W. JR.  FH
ABACAR THE WIZAR [ABACAR 1]  ERENBERGER, TIMOTHY  LP
ABACAR THE WIZARD AND THE DARK LORD OF DRAGONS [ABACAR 2]  ERENBERGER, TIMOTHY  LP
ABANDON GALAXY!  SOMERS, BART  C/P:2,P
ABANDON IN PLACE  OLTION, JERRY  LP
ABC OF SCIENCE FICTION, AN  BOARDMAN, TOM  PA,C/PA

ABDUCTION, THE [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #1] KING, J. ROBERT

P

ABDUCTION  PHILBRICK, RODMAN/HARNETT, LYNN  P
ABDUCTION (CASSETTE)  MACK, JOHN E., M. D.  SR-MM
ABDUCTORS: CONSPIRACY, THE  FRAKES, JONATHAN/SMITH, DEAN WESLEY H,P
ABHORSEN TRILOGY, THE [ABHORSEN 1,2,3]  NIX, GARTH  H
ABILITY QUOTIENT  REYNOLDS, MACK  C/P:2,P
ABISMO DE CHICAGO, O  BRADBURY, RAY  FP
ABODE OF LIFE, THE [STAR TREK]  CORREY, LEE  C/P:3,P
ABOMINABLE EARTHMAN, THE  POHL, FREDERIK  C/P:2,P
ABOMINATION, THE [OMEN 5]  MCGILL, GORDON  C/P,P
ABOMINATIONS OF YONDO, THE  SMITH, CLARK ASHTON  SR-P
ABOUT TIME  FINNEY, JACK  LP
ABOVE THE HUMAN LANDSCAPE  MCNELLY, WILLIS E./STOVER, LEON E.  LPA
ABOVE THE LOWER SKY  DEITZ, TOM  C/P,P,H
ABRAXAS MARVEL CIRCUS, THE  LEIGH, STEPHEN  C/P:2,P
ABSOLUTE AT LARGE, THE  CAPEK, KAREL  LP
ABSOLUTE MAGNITUDE  LAPINE, WARREN/PAGEL, STEPHEN  HA,LPA
ABSOLUTE SANDMAN, VOL. 1 [SANDMAN]  GAIMAN, NEIL  CX
ABSOLUTE UNCERTAINTY  SUSSEX, LUCY  LP
ABSOLUTE VALUES  MENARD, ANDREW R.  H
ABSOLUTELY BRILLIANT IN CHROME  OLEXA, KEITH  LPA
ABSOLUTION GAP [REVELATION SPACE 3]  REYNOLDS, ALASTAIR  H
ABTEILUNG PSI [@PERRY RHODAN #131]  MAHR, KURT  FPR
ABYSS, THE  CARD, ORSON SCOTT  C/P:2,P
ABYSS [STAR TREK DS9 SECTION 31]  WEDDLE, DAVID/LANG, JEFFREY  P
ABYSS  WILHELM, KATE  C/P,SFWA,P,H
ABYSS OF LIGHT, AN [GAMANT 1]  O’NEAL, KATHLEEN M.  P
ABYSS OF WONDERS, THE  SHEEHAN, PERLEY POORE  SR-H
ABYSSAL WARRIORS [@PLANESCAPE BLOOD WARS #2]  KING, J. ROBERT  P
ABYSSOS  HARDIE, RAYMOND  SFWA-P
AC3  OLIVER, STANLEY  H
ACACIA [WAR WITH THE MEIN 1]  DURHAM, DAVID ANTHONY  H
ACACIA (PROOF) [WAR WITH THE MEIN 1]  DURHAM, DAVID ANTHONY  SR-LP
ACADEM’S FURY [@CODEX ALERA #2]  BUTCHER, JIM  H
ACADEMY OF TERROR [@PRIVATE SCHOOL #2]  CHARLES, STEVEN  P
ACCELERANDO  STROSS, CHARLES  H
ACCELERATED LEARNING  ROSE, COLIN  [C/RAN-H]
ACCEPTABLE TIME, AN  L’ENGLE, MADELEINE  C/P,P,H
ACCESS DENIED [TURING HOPPER 3]  ANDREWS, DONNA  P
ACCIDENTAL CREATURES  HARRIS, ANNE  LP
ACCIDENTAL EARTH, THE  KELLEY, LEO P.  C/P:2,P
ACCIDENTAL GODDESS, AN  SINCLAIR, LINNEA  P
ACCURSED [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  LP
ACCUSATIONS [@BABYLON 5 #2]  TILTON, LOIS  C/P,P
ACE! [DOCTOR WHO]  ALDRED, SOPHIE/TUCKER, MIKE  R/REF-H
ACE ALPHA-NUMERIC ANNOTATED SCIENCE FICTION CHECKLIST, AN  DILLON, PETER C.  R/IND
ACE IN THE HOLE [@WILD CARDS #6]  MARTIN, GEORGE R. R.  C/PA,PA
ACE OF THE WHITE DEATH [@G-8 #3]  HOGAN, ROBERT J.  SR-P
ACE SCIENCE FICTION READER  WOLLHEIM, DONALD A.  C/PA,PA
ACES ABROAD [@WILD CARDS #4]  MARTIN, GEORGE R. R.  C/PA,PA
ACES HIGH [@WILD CARDS #2]  MARTIN, GEORGE R. R.  C/PA,PA
ACHILLES’ CHOICE  NIVEN, LARRY/BARNES, STEVEN  H:2,P,SFWA
ACKERMANTHOLOGY!  ACKERMAN, FORREST J.  LPA
ACORNA [ACORNA 1]  MCCAFFREY, ANNE/BALL, MARGARET  H

ACORNA’S PEOPLE [ACORNA 3] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H,P

ACORNA’S QUEST [ACORNA 2]  MCCAFFREY, ANNE/BALL, MARGARET  C/P,P,H

ACORNA’S REBELS [ACORNA 6] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN H

The MITSFS Titledex  September 24, 2007  Page 2
TITLE  AUTHOR  CODE

ACORNA’S REBELS (PROOF) [ACORNA 6] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

SR-LP

ACORNA’S SEARCH [ACORNA 5]  MCCAFFREY, ANNE  H

ACORNA’S TRIUMPH [ACORNA 7] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

P

ACORNA’S TRIUMPH (PROOF) [ACORNA 7] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

SR-LP

ACORNA’S WORLD [ACORNA 4] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

ACROSS A BILLION YEARS  SILVERBERG, ROBERT  C/P,SFWA-P,P,H
ACROSS A WINE-DARK SEA  BRYAN, JESSICA  P
ACROSS PARIS AND OTHER STORIES  AYME, MARCEL  H
ACROSS REALTIME  VINGE, VERNOR  H:2,SFWA-P,P
ACROSS THE ACHERON  WITTIG, MONIQUE  P
ACROSS THE AGES  STATTEN, VARGO  SR-P
ACROSS THE DARK METROPOLIS [BORRIBLES 3]  DE LARRABEITI, MICHAEL  P
ACROSS THE FAR MOUNTAIN [@WILDERNESS OF FOUR #1]  HANCOCK, NIEL  C/P,P
ACROSS THE NIGHTINGALE FLOOR [@OTORI #1]  HEARN, LIAN  H:2
ACROSS THE NIGHTINGALE FLOOR (PROOF) [@OTORI #1]  HEARN, LIAN  SR-LP
ACROSS THE SEA OF SUNS (ORIGINAL) [GALACTIC CENTER 2]  BENFORD, GREGORY  H:2

ACROSS THE SEA OF SUNS (REVISED: EXTRA CHAPTER AT END)

[GALACTIC CENTER 2] BENFORD, GREGORY

C/P,P

ACROSS THE THLASSA MEY [THLASSA MEY 4]  MCCARTY, DENNIS  C/P,P
ACROSS THE UNIVERSE [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P,P
ACROSS THE WALL [ABHORSEN]  NIX, GARTH  H
ACROSS THE ZODIAC  GREG, PERCY  LP
ACROSS TIME  GRINNELL, DAVID  DD286,P,C/P:2
ACT OF GOD  ASHBY, RICHARD  SR-P,C/P
ACT OF GOD [ISLAND WORLDS 1]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  P

ACT OF PROVIDENCE BRENNAN, JOSEPH PAYNE/ GRANT, DONALD M.

SR-H

ACTION: DIVISION 3 [@PERRY RHODAN #94]  ACE  C/PA,PA
ACTION STATIONS [@WING COMMANDER 6]  FORSTCHEN, WILLIAM R.  C/P,P
ACTION STORIES  ODYSSEY PUBLICATIONS  SR-LPA
ACTIVE-ENZYME LEMON-FRESHENED JUNIOR HIGH SCHOOL WITCH, THE  HILDICK, E. W.  LP
ACTIVE MEASURES  MORRIS, JANET/DRAKE, DAVID  C/P,P
ACTOR, THE  JACKSON, ANITA  SR-P
ACTS OF CONSCIENCE  BARTON, WILLIAM  P
ACTS OF GOD [@CHRIST CLONE #3]  BEAUSEIGNEUR, JAMES  H
ACTS OF KING ARTHUR AND HIS NOBLE KNIGHTS, THE  STEINBECK, JOHN  C/P:2,LP
ACTS OF THE APOSTLES  SUNDMAN, JOHN F. X.  H:2,LP
A.D. 999  BELL, JADRIEN  P
ADAM AND EVE  ERSKINE, JOHN  H
ADAM AND HIS WOMEN  CLIFFORD, SARAH  P
ADAM CHASERS, THE  BOWER, B. M.  H
ADAM EXPERIMENT, THE  SIMMONS, GEOFFREY  P
ADAM LINK — ROBOT  BINDER, EANDO  C/P:2,P
ADAM M-1  ANDERSON, WILLIAM C.  H
ADAM PROJECT, THE  WOOLFOLK, WILLIAM  P
ADAM’S COMMON  WISEMAN, DAVID  H
ADAO E EVA ROBOTS  BINDER, EANDO  FP
ADDAMS FAMILY, THE [ADDAMS FAMILY]  SHARKEY, JACK  C/P:2,P
ADDAMS FAMILY VALUES [ADDAMS FAMILY]  HODGMAN, ANN  P
ADDING A DIMENSION  ASIMOV, ISAAC  [C/RAN-P]
ADDRESS: CENTAURI  WALLACE, F. L.  H:2,P
ADDRESS UNKNOWN  PHILLPOTTS, EDEN  H
ADELE & THE BEAST  TARDI, JACQUES  CX

ADEPT, THE [@ADEPT #1] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P:2,P

ADIAMANTE  MODESITT, L. E. JR.  P
ADOLESCENCE OF P-1, THE  RYAN, THOMAS J.  H:2,LP
ADRIFT IN A BONEYARD  TAYLOR, ROBERT LEWIS  C/P,P,H
ADRIFT IN SPACE  ELWOOD, ROGER  SR-HA
ADRIFT IN THE PACIFIC  VERNE, JULES  H
ADRIFT IN THE STRATOSPHERE  LOW, PROF. A. M.  H
ADRIFT IN THE UNKNOWN  COOK, WALLACE WILLIAM  SR-P
ADS INFINITUM  WAYMAN, TONY RUSSELL  P
ADULT VERSION OF THE ESCAPADES OF CLEOPATRA, THE  ANONYMOUS  SR-L
ADULTHOOD RITES [XENOGENESIS 2]  BUTLER, OCTAVIA E.  H:2,P
ADVANCE AND RETREAT [WAR BETWEEN THE PROVINCES 3]  TURTLEDOVE, HARRY  H
ADVENTURE GIRLS IN THE AIR, THE  BLANK, CLAIR  H
ADVENTURE IN FORGOTTEN VALLEY  FREWER, GLYN  H
ADVENTURE OF THE PEERLESS PEER, THE  FARMER, PHILIP JOSE  SR-H,C/P,P
ADVENTURER, THE  OSBOURNE, LLOYD  H

The MITSFS Titledex  September 24, 2007  Page 3
TITLE  AUTHOR  CODE
ADVENTURES [LUCIFER JONES 1]  RESNICK, MIKE  P
ADVENTURES IN THE DREAM TRADE  GAIMAN, NEIL  SR-H
ADVENTURES IN THE FAR FUTURE  WOLLHEIM, DONALD A.  DD073

ADVENTURES IN THE LIADEN UNIVERSE 1-5 [LIADEN UNIVERSE 0/ LIADEN ADVENTURES 1,2,3,4,5] LEE, SHARON/MILLER, STEVE

SR-H

ADVENTURES IN THE SUPERNORMAL  GARRETT, EILEEN J.  [C/RAN-P:2]
ADVENTURES IN THE TWILIGHT ZONE  SERLING, CAROL  PA
ADVENTURES IN TIME AND SPACE  HEALY, RAYMOND J./MCCOMAS, J. FRANCIS HA,SR-HA
ADVENTURES IN TIME AND SPACE (SELECTIONS)  HEALY, RAYMOND J./MCCOMAS, J. FRANCIS PA,C/PA

ADVENTURES IN TIME AND SPACE WITH MAX MERRIWELL [MAX MERRIWELL 4] MURPHY, PAT

H

ADVENTURES IN TOMORROW  CROSSEN, KENDELL FOSTER  C/PA,PA
ADVENTURES IN UNHISTORY  DAVIDSON, AVRAM  H
ADVENTURES OF ALYX, THE  RUSS, JOANNA  C/P:2,P
ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  DICKS, TERRANCE  H
ADVENTURES OF JULES DE GRANDIN, THE  QUINN, SEABURY  C/P,P

ADVENTURES OF K9 AND OTHER MECHANICAL CREATURES [DOCTOR WHO] DICKS, TERRANCE

P

ADVENTURES OF MYHR, THE  ELROD, P. N.  P
ADVENTURES OF PHOEBE ZEIT-GEIST, THE  O’DONOGHUE, MICHAEL/SPRINGER, FRANK CX
ADVENTURES OF SAMURAI CAT, THE [SAMURAI CAT 1]  ROGERS, MARK E.  SFWA,LP
ADVENTURES OF SHERLOCK HOLMES, THE  DOYLE, ARTHUR CONAN  C/P,P
ADVENTURES OF TERRA TARKINGTON, THE  WEBB, SHARON  C/P:2,P
ADVENTURES OF THE FAITHFUL COUNSELOR, THE  SHELDON, ANNE  LP
ADVENTURES OF THE RAT FAMILY  VERNE, JULES  H

ADVENTURES OF THE STAINLESS STEEL RAT, THE [STAINLESS STEEL RAT 1,2,3] HARRISON, HARRY

C/P,P,H

ADVENTURES OF TOM BOMBADIL, THE  TOLKIEN, J. R. R.  H
ADVENTURES OF TOM BOMBADIL, THE (BRITISH)  TOLKIEN, J. R. R.  SR-H
ADVENTURES ON OTHER PLANETS  WOLLHEIM, DONALD A.  PA,C/PA
ADVENTURES ON STRANGE ISLANDS  FINLAY, ROGER T.  H
ADVENTURES TO COME  ESENWEIN, V. BERG  SR-HA
ADVENTURESS OF HENRIETTA STREET, THE [DOCTOR WHO]  MILES, LAWRENCE  P
ADVERSARY, THE [PLIOCENE EXILE 4]  MAY, JULIAN  H:2,P
ADVERSARY  RHODES, DANIEL  SFWA-P
AEGYPT [AEGYPT 1]  CROWLEY, JOHN  H,LP
AELITA  TOLSTOI, ALEXEI  H
AENEID, THE  VIRGIL  P
AEON  WILBURN, STEPHEN PAUL  LP
AERIE [DRAGON JOUSTERS 4]  LACKEY, MERCEDES  H
AEROMANCER [MANCER 4]  CALLANDER, DON  P
AESOP’S FABLES  AESOP  H
AESTIVAL TIDE [WINTERLONG 2]  HAND, ELIZABETH  C/P,P
AFFAIR OF SORCERERS, AN [MONGO FREDERICKSON 3]  CHESBRO, GEORGE C.  H:2,P
AFFAIR OF THE GENTLE SABOTEUR, THE [MAN FROM UNCLE]  KEITH, BRANDON  H
AFFAIR OF THE GUNRUNNER’S GOLD, THE [MAN FROM UNCLE]  KEITH, BRANDON  H
AFFECTIONISM SOCIETY, THE  SKOLSKY, SID  SR-P
AFFINITY TRAP, THE [STRUCTURE 1]  SKETCHLEY, MARTIN  LP
AFFIRMATION, THE  PRIEST, CHRISTOPHER  SFWA
AFRICA ZERO  ASHER, NEAL  LP
AFRICAN MERCENARY, THE [@CASCA #12]  SADLER, BARRY  P
AFRICAN WITCH, THE  CARY, JOYCE  H
AFRIKORPS [@AFRIKORPS #1]  DOLAN, BILL  P
AFTER 12,000 YEARS  COBLENTZ, STANTON A.  SR-H,SR-LP
AFTER DARK [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
AFTER DOOMSDAY  ANDERSON, POUL  C/P:2,P
AFTER LONDON  JEFFERIES, RICHARD  LP
AFTER LONG SILENCE  TEPPER, SHERI S.  C/P:2,P,H
AFTER MAN  DIXON, DOUGAL  A
AFTER MANY A SUMMER DIES THE SWAN  HUXLEY, ALDOUS  SR-P,C/P
AFTER MIDNIGHT  GRANT, CHARLES L.  PA
AFTER SOME TOMORROW  REYNOLDS, MACK  C/P:2,P
AFTER THE APOCALYPSE  FOX, W. RANDOLPH  P
AFTER THE ATOM  LA SALLE, VICTOR  SR-P
AFTER THE BOMB  MIKLOWITZ, GLORIA  C/P,P
AFTER THE DARK [@DARK ANGEL]  COLLINS, MAX ALLAN  P
AFTER THE FACT [PILGRIM 2]  SABERHAGEN, FRED  C/P,SFWA-P,P
AFTER THE FALL [STAR TREK NEW FRONTIER 15]  DAVID, PETER  H
AFTER THE FALL  SHECKLEY, ROBERT  PA
AFTER THE FLAMES [ALIEN STARS 2]  MITCHELL, ELIZABETH  PA
AFTER THE GOOD WAR  BREGGIN, PETER  P
AFTER THE KING  GREENBERG, MARTIN H.  C/PA,PA
AFTER THE LONG GOODBYE [GHOST IN THE SHELL 2]  YAMADA, MASAKI  H
AFTER THE RAIN  BOWEN, JOHN  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 4
TITLE  AUTHOR  CODE
AFTER THE ZAP  ARMSTRONG, MICHAEL  C/P,P
AFTER THINGS FELL APART  GOULART, RON  C/P:3,P
AFTER UTOPIA  REYNOLDS, MACK  C/P,P
AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  BALMER, EDWIN/WYLIE, PHILIP  H
AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  WYLIE, PHILIP/BALMER, EDWIN  C/P,P,H:2
AFTER WORLD’S END/LEGION OF TIME, THE  WILLIAMSON, JACK  SR-P,C/P:2
AFTERAGE  NAVARRO, YVONNE  P
AFTERBURN [BIO RESCUE 2]  VIEHL, S. L.  H
AFTERIMAGE [@BUFFY]  ASKEGREN, PIERCE  P

AFTERIMAGE RUSCH, KRISTINE KATHRYN/ ANDERSON, KEVIN J.

P

AFTERIMAGE AFTERSHOCK RUSCH, KRISTINE KATHRYN/ ANDERSON, KEVIN J.

LP

AFTERLIVES  SARGENT, PAMELA/WATSON, IAN  PA
AFTERMATH [@THIEVES’ WORLD #10]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
AFTERMATH [WASTEWORLD 1]  BARTON, JAMES  P
AFTERMATH  BURTON, LEVAR  C/P,P
AFTERMATH [STAR TREK S.C.E. 8]  POCKET BOOKS  LPA
AFTERMATH  SHEFFIELD, CHARLES  C/P:2,P
AFTERSHOCK [@SHADOWRUN]  RABE, JEAN/HELFERS, JOHN  P
AFTERSHOCK [@STAR TREK STARFLEET ACADEMY #2]  VORNHOLT, JOHN  LP
AFTERWAR  MORRIS, JANET  PA
AGAIN: ATLAN! [@PERRY RHODAN #46]  ACE  PA,C/PA
AGAIN, DANGEROUS VISIONS  ELLISON, HARLAN  HA,SFWA
AGAIN, DANGEROUS VISIONS 2  ELLISON, HARLAN  PA,C/PA
AGAINST A DARK BACKGROUND  BANKS, IAIN M.  C/P:2,P
AGAINST ALL ENEMIES [PAUL SINCLAIR 4]  HEMRY, JOHN G.  P
AGAINST ARCTURUS  PUTNEY, SUSAN K.  D00990
AGAINST INFINITY  BENFORD, GREGORY  C/P:2,P
AGAINST THE FALL OF NIGHT  CLARKE, ARTHUR C.  C/P:2,P
AGAINST THE GIANTS [@GREYHAWK]  EMERSON, RU  P
AGAINST THE HORDE [DRENAI 1 (ORIGINAL)]  GEMMELL, DAVID  P
AGAINST THE NIGHT, THE STARS  HOLLOW, JOHN  R/REF-H
AGAINST THE ODDS [HERIS SERRANO 7]  MOON, ELIZABETH  H,P
AGAINST THE PRINCE OF HELL [@RED SONJA #5]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
AGAINST THE TIDE [DRAGONS 3]  RINGO, JOHN  H
AGAINST THE TIDE OF YEARS [NANTUCKET 2]  STIRLING, S. M.  P
AGAINST THE WIND [@BLOOD OF TEN CHIEFS #4]  PINI, RICHARD  C/PA,LPA,SFWA
AGAINST TIME’S ARROW: THE HIGH CRUSADE OF POUL ANDERSON  MIESEL, SANDRA  R/REF-P
AGAINST TOMORROW  HOSKINS, ROBERT  PA
AGATHA HETERODYNE AND THE AIRSHIP CITY [@GIRL GENIUS #2]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGATHA HETERODYNE AND THE BEETLEBURG CLANK [@GIRL GENIUS #1]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGATHA HETERODYNE AND THE CIRCUS OF DREAMS [@GIRL GENIUS #4]  FOGLIO, PHIL/FOGLIO, KAJA  CX

AGATHA HETERODYNE AND THE CLOCKWORK PRINCESS [@GIRL GENIUS #5] FOGLIO, PHIL/FOGLIO, KAJA

CX

AGATHA HETERODYNE AND THE MONSTER ENGINE [@GIRL GENIUS #3]  FOGLIO, PHIL/FOGLIO, KAJA  CX
AGE OF AQUARIUMS [@SABRINA THE TEENAGE WITCH #20]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
AGE OF CHAOS, THE [@UNBALANCED EARTH #3]  WYLIE, JONATHAN  C/P,P
AGE OF CHIVALRY, THE/LEGENDS OF CHARLEMAGNE, THE  BULFINCH, THOMAS  [C/RAN-P]
AGE OF MIRACLES  BRUNNER, JOHN  C/P:2,P
AGE OF REASON, THE  ROTH, KURT  SR-HA
AGE OF RUIN, THE  FAUCETTE, JOHN M.  DH103
AGE OF SINATRA, THE  OHLE, DAVID  LP
AGE OF THE PUSSYFOOT, THE  POHL, FREDERIK  C/P,P,H
AGE OF THE TAIL, THE  SMITH, H. ALLEN  H:2,P
AGE OF WONDERS, THE  DWIGHT, JEFFRY  LPA
AGE OF WONDERS  HARTWELL, DAVID G.  R/REF-H,C/REF-P
AGENT OF BYZANTIUM  TURTLEDOVE, HARRY  H,P
AGENT OF BYZANTIUM (REVISED)  TURTLEDOVE, HARRY  C/P,SFWA-P,P
AGENT OF CHANGE [LIADEN UNIVERSE 1]  LEE, SHARON/MILLER, STEVE  P
AGENT OF CHAOS [FLEETS OF THE GREAT BATTLE 2]  ETHERIDGE, RUTLEDGE  P
AGENT OF CHAOS  SPINRAD, NORMAN  C/P,P,H
AGENT OF DESTRUCTION [FLEETS OF THE GREAT BATTLE 1]  ETHERIDGE, RUTLEDGE  P
AGENT OF ENTROPY  SIEGEL, MARTIN  P

AGENT OF THE TERRAN EMPIRE (ORIGINAL) (BRITISH)

[DOMINIC FLANDRY] ANDERSON, POUL

SR-P

AGENT OF THE TERRAN EMPIRE (REVISED) [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
AGENT OF THE UNKNOWN  SAINT CLAIR, MARGARET  DD150
AGENT OF VEGA  SCHMITZ, JAMES H.  C/P,P,H
AGENT OF VEGA AND OTHER STORIES  SCHMITZ, JAMES H.  P
AGENT SPATIAL NO. 1  RICHARD-BESSIERE, F.  FP
AGENT TO THE STARS  SCALZI, JOHN  SR-H
AGENTE PLANETARIO X  REYNOLDS, MACK  FP

AGENTS OF CHAOS [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS 1,2] LUCENO, JAMES H

The MITSFS Titledex  September 24, 2007  Page 5
TITLE  AUTHOR  CODE
AGENTS OF DESTRUCTION [PERRY RHODAN 134]  BRAND, KURT  SR-LP
AGENTS OF INSIGHT  KLAPER, STEVEN  P,SFWA-P
AGENTS OF LIGHT AND DARKNESS [NIGHTSIDE 2]  GREEN, SIMON R.  P
AGES IN CHAOS  VELIKOVSKY, IMMANUEL  [C/RAN-H]
AGGRESSOR SIX [AGGRESSOR SIX 1]  MCCARTHY, WIL  C/P,P
AGNES DAY [QUEST FOR THE WHITE DUCK 3]  FENN, LIONEL  C/P,SFWA-P,P
AGOG! FANTASTIC FICTION  SPARKS, CAT  LPA
AGOG! SMASHING STORIES  SPARKS, CAT  LPA
AGOG! TERRIFIC TALES  SPARKS, CAT  HA
AGUAS PROFUNDAS  HODGSON, WILLIAM HOPE  FP
AGVIQ  ARMSTRONG, MICHAEL  C/P,P
AGYAR  BRUST, STEVEN  H,P
AHEAD OF TIME  HARRISON, HARRY/GORDON, THEODORE J.  SFWA
AHEAD OF TIME  KUTTNER, HENRY  C/P:2,P
AI WAR, THE [BIOFAB 3]  BERRY, STEPHEN AMES  SFWA-P
AIR  RYMAN, GEOFF  LP
AIR ESCAPE AND EVASION  SUNDERMAN, JAMES F.  [C/RAN-H]
AIR FORCE!  HARVEY, FRANK  C/P,P
AIR MONSTER  GREEN, EDWIN  H
AIR OF MARS, THE  GINSBURG, MIRRA  HA
AIR TRUST, THE  ENGLAND, GEORGE ALLAN  LP
AIRBORN [MATT CRUSE 1]  OPPEL, KENNETH  H,P
AIRFORCE ONE IS HAUNTED  SERLING, ROBERT  H,P
AIRFRAME  CRICHTON, MICHAEL  H:2
AIRS BENEATH THE MOON  BISHOP, TOBY  P
A.I.S  DANN, JACK/DOZOIS, GARDNER  PA
AISLING [INDIGO 8]  COOPER, LOUISE  C/P:2,P
AKA  JONES, TRISTAN  H,P
A.K.A.: A COSMIC FABLE  SWIGART, ROB  LP
AKHUNATON  STEWART, DANIEL BLAIR  LP
AKIMA’S STORY [@TITAN A.E.]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
AKU-AKU  HEYERDAHL, THOR  [C/RAN-P]
ALABASTER STAFF, THE [@FORGOTTEN REALMS ROGUES]  BOLME, EDWARD  P
ALADDIN: MASTER OF THE LAMP  RESNICK, MIKE/GREENBERG, MARTIN H.  C/PA,PA
ALADORE  NEWBOLT, SIR HENRY  LP
ALAMUT [ALAMUT 1]  TARR, JUDITH  C/P,LP
ALANNA [SONG OF THE LIONESS 1]  PIERCE, TAMORA  P
ALANYA TO ALANYA [@MARQ’SSAN #1]  DUCHAMP, L. TIMMEL  LP
ALARM IN DE ATOOMCENTRALE  DEL REY, LESTER  FP
ALAS, BABYLON  FRANK, PAT  C/P:2,P,H
ALAS, THAT GREAT CITY  ASHTON, FRANCIS  H
ALASTOR [ALASTOR]  VANCE, JACK  H
ALBANIAN-ENGLISH AND ENGLISH-ALBANIAN DICTIONARY  DRIZARI, NELO  DICT
ALBINO KNIFE, THE [MATADOR 5]  PERRY, STEVE  C/P:2,SFWA-P:2,P
ALBINO’S DANCER, THE [TIME HUNTER 9]  SMITH, DALE  LP
ALCHEMICAL MARRIAGE OF ALISTAIR CROMPTON, THE  SHECKLEY, ROBERT  P
ALCHEMIST, THE  BOYD, DONNA  LP
ALCHEMIST, THE  WHITTEN, LESLIE H.  C/P,P
ALCHEMIST OF TIME, THE [@SLOW WORLD #3]  RIPLEY, KAREN  P
ALCHEMISTS, THE [AUTUMN WORLD MOSAIC 1]  GRAVEL, GEARY  C/P:2,P
ALCHEMIST’S APPRENTICE, THE  DUNCAN, DAVE  LP
ALCHEMIST’S DOOR, THE  GOLDSTEIN, LISA  H
ALCHEMY AND ACADEME  MCCAFFREY, ANNE  C/PA,PA
ALCHEMY UNLIMITED [CORWYN THE ALCHEMIST 1]  CLARK, DOUGLAS W.  C/P,P
ALCHYMIST’S CAT, THE [@DEPTFORD HISTORIES #2]  JARVIS, ROBIN  LP
ALDAIR, ACROSS THE MISTY SEA [ALDAIR 3]  BARRETT, NEAL JR.  P
ALDAIR IN ALBION [ALDAIR 1]  BARRETT, NEAL JR.  P
ALDAIR, MASTER OF SHIPS [ALDAIR 2]  BARRETT, NEAL JR.  C/P,P
ALDAIR: THE LEGION OF BEASTS [ALDAIR 4]  BARRETT, NEAL JR.  P
ALDEBARAN CAMPAIGN, THE [SEEDS OF WAR 2]  RANDLE, KEVIN/CORNETT, ROBERT  C/P,P
ALDISS UNBOUND  MATHEWS, RICHARD  R/REF-P
ALECTOR’S CHOICE [COREAN 4]  MODESITT, L. E. JR.  H
ALEJANDRA VARIATIONS, THE  COOK, PAUL  P
ALEPH AND OTHER STORIES, THE (1933–1969)  BORGES, JORGE LUIS  H,P
ALEXANDRIAN RING, THE [GAMESTER WARS 1]  FORSTCHEN, WILLIAM R.  C/P:2,P
ALGEBRA OF ICE, THE [DOCTOR WHO PAST DOCTORS]  ROSE, LLOYD  P
ALGEBRAIC FANTASIES AND REALISTIC ROMANCES  STABLEFORD, BRIAN M.  R/REF-P
ALGEBRAIST, THE  BANKS, IAIN M.  H
ALGERNON BLACKWOOD  ASHLEY, MIKE  R/IND
ALGERNON, CHARLIE AND I  KEYES, DANIEL  R/REF-H
ALGORITHM  GAWRON, JEAN MARK  P

ALICE IN WONDERLAND (STAGE ADAPTATION)

CARROLL, LEWIS/LE GALLIENNE, EVA/ FRIEBUS, FLORIDA

SR-P

ALICE’S WORLD  LUNDWALL, SAM J.  D58880

The MITSFS Titledex  September 24, 2007  Page 6
TITLE  AUTHOR  CODE
ALICIA AUSTIN’S AGE OF DREAMS  AUSTIN, ALICIA  A
ALICIA II  THURSTON, ROBERT  H,P
ALIEN, THE  BESAW, VICTOR  P
ALIEN, THE  DAVIES, L. P.  H,SFWA
ALIEN [ALIENS]  FOSTER, ALAN DEAN  H:2,P
ALIEN, THE  JONES, RAYMOND F.  C/P:2,GN1.06,P
ALIEN  LEONARD, GEORGE H.  C/P,P
ALIEN  MULLER, JOHN E.  SR-P
ALIEN 3 [ALIENS]  FOSTER, ALAN DEAN  C/P,P
ALIEN (CHINESE)  RAYJEAN, MAX-ANDRE  FP
ALIEN ABDUCTORS, THE  BEVIS, H. U.  H
ALIEN AFFAIR, AN [@MISSION EARTH #4]  HUBBARD, L. RON  H,P
ALIEN ART  DICKSON, GORDON R.  H:2,P
ALIEN ART/ARCTURUS LANDING  DICKSON, GORDON R.  C/P,P
ALIEN ATLAS  ALEXANDER, C. M.  P
ALIEN BLOOD [CAT 1,2]  VINGE, JOAN D.  H:2
ALIEN BLUES [ELAKI 1]  HIGHTOWER, LYNN S.  P,SFWA-P
ALIEN BODIES [DOCTOR WHO EIGHTH DOCTOR]  MILES, LAWRENCE  C/P:2,P
ALIEN BOOTLEGGER  ORE, REBECCA  LP
ALIEN BOOTLEGGER AND OTHER STORIES  ORE, REBECCA  H
ALIEN BOUNTY [SAM MCCADE 3]  DIETZ, WILLIAM C.  C/P,P
ALIEN CARNIVAL  LIEBSCHER, WALT  SR-P
ALIEN CITADEL [HUNTSMAN 3]  HILL, DOUGLAS  H
ALIEN CONDITION, THE  GOLDIN, STEPHEN  C/PA,PA
ALIEN DARK, THE  GALLAGHER, DIANA G.  C/P,P
ALIEN DARKNESS, AN  CASTRO, ADAM-TROY  LP
ALIEN DEATH FLEET [@STAR FRONTIER #1]  HUDSON, EDWARD S.  C/P:2,P
ALIEN DEBT, THE [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
ALIEN DREAMS  SEGRIFF, LARRY  P
ALIEN DUST  TUBB, E. C.  H
ALIEN EARTH, THE  ELDER, MICHAEL  C/P,P
ALIEN EARTH  LINDHOLM, MEGAN  C/P,P
ALIEN EARTH AND OTHER STORIES  ELWOOD, ROGER/MOSKOWITZ, SAM  C/PA,PA
ALIEN EMBASSY  WATSON, IAN  C/P,P
ALIEN EMERGENCIES [SECTOR GENERAL 4,5,6]  WHITE, JAMES  LP
ALIEN EMPIRE  O’TOOLE, CHRISTOPHER  [C/RAN-H]
ALIEN ENCOUNTER  DEVIN, FLANNA  P
ALIEN ENCOUNTERS  FINDER, JAN HOWARD  SR-HA
ALIEN ENCOUNTERS  ROSE, MARK  R/REF-P
ALIEN EYES [ELAKI 2]  HIGHTOWER, LYNN S.  C/P,P
ALIEN FACTOR, THE  LEE, STAN/TIMMONS, STAN  C/P:2,P
ALIEN FLESH  QUINN, SEABURY  SR-H
ALIEN FROM ARCTURUS  DICKSON, GORDON R.  DD139
ALIEN FROM THE STARS  FANTHORPE, R. L.  H
ALIEN HARVEST [ALIENS 5]  SHECKLEY, ROBERT  C/P,P
ALIEN HEAT [ELAKI 3]  HIGHTOWER, LYNN S.  P
ALIEN HEAT, AN [DANCERS AT THE END OF TIME 1]  MOORCOCK, MICHAEL  H:2,P,SFWA
ALIEN HORIZONS  NOLAN, WILLIAM F.  C/P:2,P
ALIEN IN A BOTTLE  MACKEL, KATHRYN  H
ALIEN INFLUENCES  RUSCH, KRISTINE KATHRYN  P
ALIEN INVADERS [@SPACE HAWKS #2]  PACKARD, EDWARD  P
ALIEN ISLAND [@SPACE PRECINCT #3]  BISCHOFF, DAVID F.  C/P,P
ALIEN ISLAND  SHERRED, T. L.  C/P:2,P

ALIEN LANDSCAPES HOLDSTOCK, ROBERT P./ EDWARDS, MALCOLM

A

ALIEN LIFE  TUBB, E. C.  LP
ALIEN LIGHT, AN  KRESS, NANCY  H:2
ALIEN LOVE  RODGERS, ALAN  LP
ALIEN MAIN  SHERRED, T. L./BIGGLE, LLOYD JR.  H:2,SFWA
ALIEN MEETINGS  STEIGER, BRAD  [C/RAN-P]
ALIEN MINDS  EVANS, E. EVERETT  SR-H
ALIEN MINDS  LAUMER, KEITH  C/P,P
ALIEN NATION [ALIEN NATION]  FOSTER, ALAN DEAN  C/P:2,P
ALIEN ONES, THE  BRETT, LEO  SR-P,C/P
ALIEN PERSPECTIVE  HOUSTON, DAVID  P
ALIEN PETS  LITTLE, DENISE  PA
ALIEN PLANET  PRATT, FLETCHER  C/P,P,H
ALIEN PLOT  ANTHONY, PIERS  P

ALIEN PREGNANT BY ELVIS FRIESNER, ESTHER M./

GREENBERG, MARTIN H.

C/PA,PA

ALIEN PROBE, THE [@NEW TOM SWIFT #3]  APPLETON, VICTOR  H,LP
ALIEN REALMS [LORD TEDRIC 4]  SMITH, E. E.  SR-P
ALIEN RESONANCE  EASTON, THOMAS A.  H,LP
ALIEN RESURRECTION [ALIENS]  CRISPIN, A. C./O’MALLEY, KATHLEEN  P

TITLE  AUTHOR  CODE
ALIEN RITES [ELAKI 4]  HIGHTOWER, LYNN S.  P
ALIEN SALUTE [@SAND WARS #4]  INGRID, CHARLES  C/P,P
ALIEN SEA  RACKHAM, JOHN  DH040
ALIEN SEA, THE [@DRAGONLANCE CHAMPIONS]  SOULBAN, LUCIEN  P
ALIEN SECRETS  KLAUSE, ANNETTE CURTIS  H
ALIEN SEED [SPACE: 1999 7]  TUBB, E. C.  C/P:2,P
ALIEN SEX  DATLOW, ELLEN  SR-HA,C/PA,PA
ALIEN STARS [ALIEN STARS 1]  MITCHELL, ELIZABETH  PA
ALIEN SWORDMASTER, THE [V]  SUCHARITKUL, SOMTOW  C/P:2,SFWA-P,P
ALIEN TASTE [UKIAH OREGON 1]  SPENCER, WEN  C/P,P
ALIEN TONGUE [NEXT WAVE 2]  LEIGH, STEPHEN  C/P:2,P
ALIEN TRACE, THE [@ALIEN TRACE #1]  MAJOR, H. M.  P
ALIEN UPSTAIRS, THE  SARGENT, PAMELA  P,SFWA
ALIEN WAY, THE  DICKSON, GORDON R.  C/P:2,SFWA-TD31,P
ALIEN WEB, THE [@MASTERS OF SPACE #2]  VARDEMAN, ROBERT E.  P
ALIEN WITHIN, THE  BOLTON, JOHANNA M.  P
ALIEN WITHIN, THE [@VOYAGERS #2]  BOVA, BEN  C/P,SFWA-P,P,H:2
ALIEN WITHIN, THE  DE BOLT, ADRIANA  P
ALIEN WORLD  LUKENS, ADAM  H
ALIEN WORLDS  COLLINS, PAUL  SR-HA
ALIEN WORLDS  ELWOOD, ROGER  PA,C/PA
ALIEN YEARS, THE  SILVERBERG, ROBERT  H:2,P
ALIENIST [GERALD KNAVE]  JANIFER, LAURENCE M.  LP
ALIENS!  DOZOIS, GARDNER/DANN, JACK  C/PA,PA
ALIENS [ALIENS]  FOSTER, ALAN DEAN  P
ALIENS, THE  LEINSTER, MURRAY  C/P:2,P
ALIENS AMONG US  DANN, JACK/DOZOIS, GARDNER  C/PA:2,PA
ALIENS AMONG US, THE  WHITE, JAMES  C/P:2,P
ALIENS AND ENEMIES [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
ALIENS ATE MY HOMEWORK [ROD ALLBRIGHT 1]  COVILLE, BRUCE  P
ALIENS COLONIAL MARINES TECHNICAL MANUAL [@ALIENS]  BRIMMICOMBE-WOOD, LEE  VLP
ALIENS FOUR  STURGEON, THEODORE  C/P:2,P
ALIENS FROM ANALOG  SCHMIDT, STANLEY  HA,SFWA
ALIENS FROM SPACE  KEYHOE, DONALD E.  SFWA
ALIENS FROM SPACE  OSBORNE, DAVID  H
ALIENS IN SPACE  CALDWELL, STEVEN  A
ALIENS, OF EARTH, THE  KRESS, NANCY  H
ALIENS, TRAVELERS, AND OTHER STRANGERS  MACMILLAN PUBLISHING  HA
ALIGHT IN THE VOID  ANDERSON, POUL  C/P,P
ALL ABOARD  OLSON, MIKE  LPA
ALL ABOUT BROTHER BIRD  PLUMMER, CATHARINE  P
ALL ABOUT ROCKETS AND JETS  PRATT, FLETCHER  [C/RAN-H]
ALL ABOUT VENUS  ALDISS, BRIAN W.  C/PA,PA
ALL-AMERICAN ALIEN BOY  STEELE, ALLEN  C/P,P
ALL-CONSUMING FIRE [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY  C/P:2,P
ALL DARKNESS MET [DREAD EMPIRE 3/RAGNARSON 5]  COOK, GLEN  C/P:2,P
ALL FLESH IS GRASS  SIMAK, CLIFFORD D.  C/P,P,H
ALL FOOLS’ DAY  COOPER, EDMUND  P
ALL GOOD THINGS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  H:2,P
ALL HALLOW’S EVE  WILLIAMS, CHARLES  SR-P,C/P,H
ALL HELL BREAKING LOOSE  GREENBERG, MARTIN H.  PA
ALL I REALLY NEED TO KNOW I LEARNED FROM WATCHING STAR TREK  MARINACCIO, DAVE  [C/RAN-H:2]
ALL IN COLOR FOR A DIME  LUPOFF, RICHARD A./THOMPSON, DON  R/REF-P
ALL IN GOOD TIME  ORMONDROYD, EDWARD  P
ALL JUDGMENT FLED  WHITE, JAMES  H:2,P
ALL MY SINS REMEMBERED  HALDEMAN, JOE  H:2,P
ALL NIGHT AWAKE [WILLIAM SHAKESPEARE 2]  HOYT, SARAH A.  P
ALL OF AN INSTANT  GARFINKLE, RICHARD  H,LP
ALL ONE UNIVERSE  ANDERSON, POUL  LP
ALL OUR YESTERDAYS  WARNER, HARRY JR.  R/REF-H
ALL RIGHT, EVERYBODY OFF THE PLANET!  OTTUM, BOB  C/P,P,H
ALL SHADOWS FLED [@FORGOTTEN REALMS SHADOW OF AVATAR #3]  GREENWOOD, ED  P
ALL-STAR ZEPPELIN STORIES  MOLES, DAVID/LAKE, JAY  LPA
ALL THAT GLITTERS [@SABRINA THE TEENAGE WITCH #12]  GARTON, RAY  P
ALL THE BELLS ON EARTH  BLAYLOCK, JAMES P.  P
ALL THE COLORS OF DARKNESS [JAN DARZEK]  BIGGLE, LLOYD JR.  C/P,P,H:2
ALL THE DEVILS ARE HERE  DEYO, DAVID D. JR.  SR-LPA
ALL THE GODS OF EISERNON  LANG, SIMON  C/P:2,P
ALL THE MYRIAD WAYS  NIVEN, LARRY  C/P:2,P
ALL THE RAGE [REPAIRMAN JACK]  WILSON, F. PAUL  H
ALL THE SHATTERED WORLDS  VANCE, STEVE  P
ALL THE TRAPS OF EARTH AND OTHER STORIES  SIMAK, CLIFFORD D.  H:2,P
ALL THE VISIONS  RUCKER, RUDY  OVD-2
ALL THE WAY TO THE GALLOWS  DRAKE, DAVID  C/P,P

The MITSFS Titledex  September 24, 2007  Page 8
TITLE  AUTHOR  CODE
ALL THE WEYRS OF PERN [PERN]  MCCAFFREY, ANNE  H:2,P
ALL THE WONDERS WE SEEK  MARTI-IBANEZ, FELIX  SR-H
ALL THESE EARTHS  BUSBY, F. M.  C/P:2,P
ALL THINGS UNDER THE MOON [TEDDY LONDON,P.I.]  MORGAN, ROBERT  P
ALL TIMES POSSIBLE  EKLUND, GORDON  C/P,P
ALL TOGETHER DEAD [SOUTHERN VAMPIRES 7]  HARRIS, CHARLAINE  H:2
ALL TOMORROW’S PARTIES  GIBSON, WILLIAM  H,LP
ALL YOU NEED IS A LOVE SPELL [@SABRINA THE TEENAGE WITCH #7]  REISFELD, RANDI  P
ALLAMANDA [ARCADY 2]  WILLIAMS, MICHAEL  H,LP
ALLAN’S WIFE [ALLAN QUATERMAIN]  HAGGARD, H. RIDER  LP
ALLEGIANCE [@STAR WARS]  ZAHN, TIMOTHY  H:3
ALLELUIA FILES, THE [SAMARIA]  SHINN, SHARON  LP
ALLEY GOD, THE  FARMER, PHILIP JOSE  C/P:2,P
ALLIANCE [ROBOTS AND ALIENS 4]  OLTION, JERRY  P,SFWA-P:2
ALLIANCES [@STARGATE SG-1 #8]  MILLER, KAREN  P
ALLIANCES [@DRAGONLANCE ELVEN EXILES #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
ALLIES AND ALIENS  ALLEN, ROGER MACBRIDE  P
ALLIES OF ANTARES [DRAY PRESCOT 26]  AKERS, ALAN BURT  P
ALLY [WESS’HAR WARS 5]  TRAVISS, KAREN  P
AL’MANAKH NAUCHNOI FANTASTIKI 1963, VYPUSK 1  MALININA  FH
AL’MANAKH NAUCHNOI FANTASTIKI 1964, VYPUSK 2  ANDREEV, K.  FH
ALMAS DE METAL  ASHLEY, MIKE  FP
ALMURIC  HOWARD, ROBERT E.  C/P:2,P
ALOHA  CHRISTENSEN, MARK  H
ALONE AGAINST TOMORROW  ELLISON, HARLAN  C/P,P,H:2
ALONE BY NIGHT  CONGDON, DON/CONGDON, MICHAEL  C/PA,PA
ALONE IN THE ASHES [ASHES 5]  JOHNSTONE, WILLIAM W.  P
ALONGSIDE NIGHT  SCHULMAN, J. NEIL  C/P:2,P
ALPH  MAINE, CHARLES ERIC  H:2,P
ALPHA [CHARON 2]  ASARO, CATHERINE  H
ALPHA 1  SILVERBERG, ROBERT  C/PA,PA
ALPHA 2  SILVERBERG, ROBERT  C/PA,PA
ALPHA 3  SILVERBERG, ROBERT  C/PA,PA
ALPHA 4  SILVERBERG, ROBERT  C/PA,PA
ALPHA 5  SILVERBERG, ROBERT  C/PA,PA
ALPHA 6  SILVERBERG, ROBERT  C/PA,PA
ALPHA 7  SILVERBERG, ROBERT  C/PA,PA
ALPHA 8  SILVERBERG, ROBERT  PA,C/PA
ALPHA 9  SILVERBERG, ROBERT  C/PA,PA
ALPHA CENTAURI  BARTON, WILLIAM/CAPOBIANCO, MICHAEL C/P,LP,H
ALPHA CENTAURI  SIEGEL, ROBERT  LP
ALPHA CENTAURI — OR DIE!  BRACKETT, LEIGH  C/P:2,DF187
ALPHA-II  HUBSCHMAN, THOMAS  P
ALPHA TRANSIT  MCSWEEGAN, EDWARD  LP
ALPHA TRAP, THE  BYRNE, STUART J.  P
ALPHA YES, TERRA NO!  PETAJA, EMIL  DM121
ALPHABET OF THORN  MCKILLIP, PATRICIA A.  LP
ALQUA DREAMS  POLLACK, RACHEL  H
ALREADY DEAD  HUSTON, CHARLIE  LP
ALSO PEOPLE, THE [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN  C/P:2,P
ALTA [DRAGON JOUSTERS 2]  LACKEY, MERCEDES  H,P
ALTAR ON ASCONEL, THE  BRUNNER, JOHN  DM123
ALTER EGO  WATSON, PATRICK  H
ALTER EVIL [@STARSHIP ORPHEUS #3]  JADE, SYMON  P
ALTERATION, THE  AMIS, KINGSLEY  H,P
ALTERED CARBON [TAKESHI KOVACS 1]  MORGAN, RICHARD  H
ALTERED EGO, THE  SOHL, JERRY  H:2,P

ALTERED I:

URSULA K. LE GUIN’S SCIENCE FICTION WRITING WORKSHOP, THE HARDING, LEE

H,SR-LP

ALTERED STATES  CHAYEFSKY, PADDY  C/P,P,H

ALTERNATE AMERICAS [@WHAT MIGHT HAVE BEEN #4] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA

ALTERNATE ASIMOVS, THE  ASIMOV, ISAAC  H,P

ALTERNATE EMPIRES [@WHAT MIGHT HAVE BEEN #1] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA,SFWA-P

ALTERNATE GENERALS  TURTLEDOVE, HARRY  PA,C/PA
ALTERNATE GENERALS II  TURTLEDOVE, HARRY  HA
ALTERNATE GENERALS III  TURTLEDOVE, HARRY/GREEN, ROLAND J.  HA
ALTERNATE GERROLDS  GERROLD, DAVID  LP

ALTERNATE GETTYSBURGS THOMSEN, BRIAN M./

GREENBERG, MARTIN H.

HA

ALTERNATE HEROES [@WHAT MIGHT HAVE BEEN #2] BENFORD, GREGORY/

GREENBERG, MARTIN H.

C/PA,PA

ALTERNATE KENNEDYS  RESNICK, MIKE  PA
ALTERNATE MARTIANS, THE  CHANDLER, A. BERTRAM  DM129

The MITSFS Titledex  September 24, 2007  Page 9
TITLE  AUTHOR  CODE
ALTERNATE ORBITS [GRIMES]  CHANDLER, A. BERTRAM  D13783
ALTERNATE OUTLAWS  RESNICK, MIKE  PA
ALTERNATE PRESIDENTS  RESNICK, MIKE  C/PA,PA
ALTERNATE SKIFFY  RESNICK, MIKE/HAYDEN, PATRICK NIELSEN LP
ALTERNATE TYRANTS  RESNICK, MIKE  LPA
ALTERNATE WARRIORS  RESNICK, MIKE  PA
ALTERNATE WORLDCONS AND AGAIN, ALTERNATE WORLDCONS  RESNICK, MIKE  SR-LPA
ALTERNATE WORLDS  GUNN, JAMES E.  R/REF-H,C/REF-H:3
ALTERNATING CURRENTS  POHL, FREDERIK  C/P:2,P
ALTERNATIVES  ADAMS, ROBERT  SFWA-P
ALTERNITIES  GERROLD, DAVID  C/PA,PA
ALTERNITIES  KUBE-MCDOWELL, MICHAEL P.  H,SFWA,SFWA-P,P
ALTMUTANTEN, DIE [@PERRY RHODAN SILBERBANDE #65]  PABEL MOEWIG KG  FPR

ALTON GIFT, THE [DARKOVER] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

ALTON’S UNGUESSABLE  SUTTON, JEFF  D76096
ALUMINUM MAN, THE  EDMONDSON, G. C.  C/P:2,P
ALVIN JOURNEYMAN [ALVIN MAKER 4]  CARD, ORSON SCOTT  C/P:2,P
ALVIN STEADFAST ON VERNACULAR ISLAND  JACOBS, FRANK  H
ALVIN WANDERING [ALVIN MAKER 4,5]  CARD, ORSON SCOTT  H
ALWAYS COMING HOME  LE GUIN, URSULA K.  C/P:2,SFWA,P
ALWAYS COMING HOME (BOXED HARDCOVER AND CASSETTE)  LE GUIN, URSULA K.  H,SR-MM:3
ALWAYS COMING HOME (COMPANION CASSETTE)  LE GUIN, URSULA K.  SR-MM
ALWAYS DARKEST  TIMMONS, STAN  LP
ALWAYS FOREVER [AGE OF MISRULE 3]  CHADBOURN, MARK  LP
ALWAYS THE BLACK KNIGHT  HOFFMAN, LEE  C/P:2,P
AM RAND DES BLAUEN NEBELS [@PERRY RHODAN #8]  KNEIFEL, HANS  FPR
AMATEURS IN ALCHEMY  DEEGAN, JON J.  SR-P
AMAZING DR. DARWIN, THE  SHEFFIELD, CHARLES  H
AMAZING MAURICE AND HIS EDUCATED RODENTS, THE [DISCWORLD 27]  PRATCHETT, TERRY  H
AMAZING POWER OF ASHUR FINE, THE  SOBOL, DONALD J.  H
AMAZING SPIDER-MAN, THE [MARVEL COMICS]  LEE, STAN  SCX
AMAZING STORIES, THE [STAR TREK]  STAR TREK BOOKS  LPA
AMAZING STORIES VISIONS OF OTHER WORLDS  GREENBERG, MARTIN H.  LPA
AMAZON PLANET  REYNOLDS, MACK  C/P:2,P
AMAZONS [MIND MASTERS 4]  ROSS, IAN  C/P,P
AMAZONS!  SALMONSON, JESSICA AMANDA  C/PA,PA
AMAZONS II  SALMONSON, JESSICA AMANDA  PA,C/PA
AMAZONS OF SOMELON, THE  KAMINSKI, RAYMOND  P
AMBASSADOR, THE [@WARHAMMER VON VELTEN 1]  MCNEILL, GRAHAM  P
AMBASSADOR OF PROGRESS  WILLIAMS, WALTER JON  C/P,SFWA-P,P
AMBASSADORS FROM AURIGEL, THE [@PERRY RHODAN #64]  ACE  PA,C/PA
AMBASSADORS OF DEATH, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
AMBER AND ASHES [@DRAGONLANCE DARK DISCIPLE #1]  WEIS, MARGARET  H
AMBER AND IRON [@DRAGONLANCE DARK DISCIPLE #2]  WEIS, MARGARET  H
AMBER CITADEL, THE [JEWELFIRE 1]  WARRINGTON, FREDA  P
AMBER DREAMS: A ROGER ZELAZNY BILIOGRAPHY  LEVACK, DANIEL J. H.  R/IND
AMBER EFFECT, THE  PRATHER, RICHARD S.  SFWA-P
AMBER ENCHANTRESS, THE [@DARK SUN PRISM #3]  DENNING, TROY  C/P,P
AMBER GODS, THE  SPOFFORD, HARRIET PRESCOTT  LP
AMBER SPYGLASS, THE [HIS DARK MATERIALS 3]  PULLMAN, PHILIP  C/P,P,H

AMBER SPYGLASS, THE (LIMITED EDITION) (CD)

[HIS DARK MATERIALS 1,2,3] PULLMAN, PHILIP

SR-MM

AMBER WIZARD, THE [OSSERIAN 1]  FORBES, DAVID  P
AMBIENT [DRYCO 1]  WOMACK, JACK  H,P
AMBIVALENT MAGICIAN, THE  HAWKE, SIMON  P
AMBROV KEON [SIME/GEN 7]  LORRAH, JEAN  C/P:2,P
AMBULANCE SHIP [SECTOR GENERAL 4]  WHITE, JAMES  C/P:2,P
AMBUSH AT CORELLIA [STAR WARS CORELLIAN 1]  ALLEN, ROGER MACBRIDE  C/P:2,P
AMBUSH AT OSIRAK  CROWDER, HERBERT  P
AMBUSH OF SHADOWS, AN [PELBAR 5]  WILLIAMS, PAUL O.  C/P:2,P
AMEACA DE ANDROMEDA [ANDROMEDA 1]  HOYLE, FRED/ELLIOT, JOHN  FP
AMERICA 2040 [AMERICA 2040 1]  INNES, EVAN  P
AMERICAN BOOK OF THE DEAD, THE  BILLIAS, STEPHEN  P
AMERICAN DEFIANCE [@DOOMSDAY WARRIOR #7]  STACY, RYDER  P
AMERICAN EDEN [@DOOMSDAY WARRIOR #11]  STACY, RYDER  P
AMERICAN EMPEROR, AN  TRACY, LOUIS  H
AMERICAN FANTASY AND SCIENCE FICTION  TYMN, MARSHALL B.  SRL/IND
AMERICAN FANTASY TRADITION, THE  THOMSEN, BRIAN M.  LPA
AMERICAN FRONT [CONFEDERATE STATES: GREAT WAR 1]  TURTLEDOVE, HARRY  C/P,P
AMERICAN GLORY [@DOOMSDAY WARRIOR #8]  STACY, RYDER  P
AMERICAN GODS  GAIMAN, NEIL  C/P:2,P
AMERICAN GOTHIC  BLOCH, ROBERT  SFWA-P
AMERICAN MEAT [@DARK FUTURE AMERICAN MEAT 1]  MOORE, STUART  P

The MITSFS Titledex  September 24, 2007  Page 10
TITLE  AUTHOR  CODE
AMERICAN REBELLION [@DOOMSDAY WARRIOR #6]  STACY, RYDER  P
AMERICAN ZONE, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  LP
AMERICA’S LAST DECLARATION [@DOOMSDAY WARRIOR #5]  STACY, RYDER  P
AMERICA’S SECRET SERVICE ACE  CARR, NICK  R/REF-P
AMERICA’S ZERO HOUR [@DOOMSDAY WARRIOR #9]  STACY, RYDER  P
AMERIKA  POUNS, BRAUNA E.  P
AMES DEVOREES, LES  JETER, K. W.  FP
AMES HOLBROOK, DEITY  RUNYON, CHARLES W.  C/P,P
AMITYVILLE HORROR, THE  ANSON, JAY  [C/RAN-P:2]
AMMONITE  GRIFFITH, NICOLA  C/P:2,P
AMNESIA MOON  LETHEM, JONATHAN  H
AMOK TIME [STAR TREK]  STURGEON, THEODORE  SR-S
AMONG MADMEN  STARLIN, JIM/GRAZIUNAS, DAINA  C/P,P
AMONG THE CANNIBALS  VERNE, JULES  H
AMONG THE DEAD  BRYANT, EDWARD  H,SFWA,P
AMONG THE FOREST PEOPLE  PIERSON, CLARA D.  H
AMORALITY TALE [DOCTOR WHO PAST DOCTORS]  BISHOP, DAVID  P
AMORE A QUATTRO DIMENSIONI  GALASSIA  FP
AMOROUS PHILANDRE  DE BIBIENA, JEAN-GALLI  SR-P
AMOROUS UMBRELLA, THE  KAYE, MARVIN  H,SFWA
AMOTINADOS DO POLAR LION, OS  ROSHWALD, MORDECAI  FP
AMOUNT TO CARRY, THE  SCHOLZ, CARTER  LP
AMPHIBIAN, THE  BELIAEV, ALEXANDER  H
AMPHIBIANS, THE  WRIGHT, S. FOWLER  C/P:2,GN1.04
AMSIRS AND THE IRON THORN, THE  BUDRYS, ALGIS  C/P:2,P
AMULET, THE  HALL, S. C.  SR-L
AMULET, THE  MORLAN, A. R.  P
AMULET OF SAMARKAND, THE [@BARTIMAEUS #1]  STROUD, JONATHAN  H
AMULET OF TARV, THE  KENSETT, PERCY F.  H
ANACHRONISMS  HINZ, CHRISTOPHER  P
ANACHROPHOBIA [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, JONATHAN  P
ANACKIRE [VIS 2]  LEE, TANITH  C/P:2,P
ANAL ASTRONAUT  GOLD, BOB  SR-L
ANALOG 1  CAMPBELL, JOHN W. JR.  PA,HA
ANALOG 2  CAMPBELL, JOHN W. JR.  C/PA,PA,HA
ANALOG 3  CAMPBELL, JOHN W. JR.  HA
ANALOG 4  CAMPBELL, JOHN W. JR.  HA
ANALOG 5  CAMPBELL, JOHN W. JR.  HA
ANALOG 6  CAMPBELL, JOHN W. JR.  C/PA,PA
ANALOG 7  CAMPBELL, JOHN W. JR.  HA,PA
ANALOG 8  CAMPBELL, JOHN W. JR.  HA,SFWA
ANALOG 9  BOVA, BEN  HA
ANALOG ANNUAL  BOVA, BEN  PA,C/PA
ANALOG: WRITER’S CHOICE II  SCHMIDT, STANLEY  SFWA
ANALOG YEARBOOK  BOVA, BEN  HA,LPA
ANALOG YEARBOOK II  SCHMIDT, STANLEY  C/PA,PA
ANALOGUE MEN  KNIGHT, DAMON  C/P:2,P
ANANSI BOYS, THE  GAIMAN, NEIL  H:2
ANARCH LORDS, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
ANARCHAOS  CLARK, CURT  C/P:2,P
ANARCHISTIC COLOSSUS, THE  VAN VOGT, A. E.  C/P:2,P
ANARCHY IN THE ASHES [ASHES 3]  JOHNSTONE, WILLIAM W.  P
ANARCTICA  ROBINSON, KIM STANLEY  P
ANASAZI  ING, DEAN  C/P:2,P
ANATOMY OF WONDER: SCIENCE FICTION  BARRON, NEIL  R/REF-P,C/REF-H
ANAZ-VOOHRI [OPERATION PLEIADES]  SCHARTZ, VIJAYA  P
ANCESTOR CELL, THE [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER/COLE, STEPHEN  P
ANCESTOR’S WORLD [STARBRIDGE 6]  CRISPIN, A. C./KING, T. JACKSON  P
ANCESTRAL BLOOD  HALLOWELL, MADELYNE  LP
ANCESTRAL HUNGERS  BAKER, SCOTT  LP
ANCIENT, THE [@SEAQUEST DSV 3]  BISCHOFF, DAVID F.  C/P,P
ANCIENT BLOOD [STAR TREK DAY OF HONOR 1/STAR TREK TNG]  CAREY, DIANE  C/P:2,P
ANCIENT DESIGNS  HAGE, RICHARD  LP
ANCIENT ECHOES  HOLDSTOCK, ROBERT P.  C/P,P

ANCIENT ENCHANTRESSES MASSIE-FERCH, KATHLEEN M./ GREENBERG, MARTIN H./ GILLIAM, RICHARD

PA

ANCIENT ENEMY, THE [ARNA 1]  ROWLEY, CHRISTOPHER  P
ANCIENT ENEMY, THE  THOMPSON, DONALD  P
ANCIENT EVENINGS  MAILER, NORMAN  H
ANCIENT EXHUMATIONS +2  SARGENT, STANLEY C.  LP
ANCIENT GAMES [@ELVEN WAYS #2]  CIENCIN, SCOTT  P
ANCIENT HEAVENS  VARDEMAN, ROBERT E.  P
ANCIENT IMAGES  CAMPBELL, RAMSEY  SFWA-P

TITLE  AUTHOR  CODE
ANCIENT IRISH TALES  CROSS, TOM PEETE/SLOVER, CLARK HARRIS HA
ANCIENT LEGACY, THE [MATHEW LEWIN 3]  GRAHAM, MITCHELL  P
ANCIENT LIGHT [ORTHE 2]  GENTLE, MARY  C/P,P,H
ANCIENT LIGHTS  GRUBB, DAVIS  LP
ANCIENT, MY ENEMY  DICKSON, GORDON R.  H:2,P,SFWA
ANCIENT MYSTERIES READER, THE  HAINING, PETER  SFWA
ANCIENT OF DAYS  BISHOP, MICHAEL  H,P
ANCIENT OF DAYS, THE  GREENFIELD, IRVING A.  P
ANCIENT OF DAYS [@CONFLUENCE #2]  MCAULEY, PAUL J.  C/P,P,H
ANCIENT ONE, THE  BARRON, T. A.  P
ANCIENT PUZZLES  OLIVASTRO, DOMINIC  [C/RAN-H]
ANCIENT SHORES  MCDEVITT, JACK  C/P,P,H:2
ANCIENT SOLITARY REIGN, THE  HOCKE, MARTIN  H
ANCIENT SORCERIES  BLACKWOOD, ALGERNON  VLP
AND AFTERWARD, THE DARK  COPPER, BASIL  H
AND ALL BETWEEN [GREEN-SKY 2]  SNYDER, ZILPHA KEATLEY  C/P,SFWA-P,P
AND ALL THE STARS A STAGE  BLISH, JAMES  C/P:2,P,H
AND CALL MY KILLER...MODOK! [IRON MAN]  ROTSLER, WILLIAM  P
AND CHAOS DIED  RUSS, JOANNA  C/P:2,P
AND DISREGARDS THE REST  VOERMANS, PAUL  P
AND ETERNITY [INCARNATIONS OF IMMORTALITY 7]  ANTHONY, PIERS  C/P:2,P,H
AND HAVING WRIT...  BENSEN, DONALD R.  C/P,P,H
AND LOVE SURVIVED  CHETWYND-HAYES, R.  P
AND LOVING IT! [GET SMART 6]  JOHNSTON, WILLIAM  P
AND NOT MAKE DREAMS YOUR MASTER  GOLDIN, STEPHEN  C/P:2,P
AND NOW THE NEWS... [COMPLETE STORIES OF THEODORE STURGEON 9] STURGEON, THEODORE  H
...AND OTHERS SHALL BE BORN  LONG, FRANK BELKNAP  SR-DB0809
AND PEACE SHALL SLEEP [MAGIC: THE GATHERING]  LYRIS, SONIA ORIN  P
"...AND SOME WERE HUMAN"  DEL REY, LESTER  C/P:2,P
AND STILL THE EARTH  BRANDAO, IGNACIO DELOYOLA  P
AND STRANGE AT ECBATAN THE TREES  BISHOP, MICHAEL  H:2,SFWA,SFWA-TD16
AND THE ANGELS SING  WILHELM, KATE  H
AND THE DEVIL WILL DRAG YOU UNDER  CHALKER, JACK L.  C/P:2,P
AND THE DEVIL WILL DRAG YOU UNDER (PROOF)  CHALKER, JACK L.  SR-LP
AND THE GODS LAUGHED  BROWN, FREDRIC  H
AND THE STARS REMAIN  BERRY, BRYAN  SR-P
AND THEN THE TOWN TOOK OFF  WILSON, RICHARD  DD437
AND THEN THERE’LL BE FIREWORKS [OZARK 3]  ELGIN, SUZETTE HADEN  C/P,SFWA,P,H
"...AND THEN WE’LL GET HIM!"  WILSON, GAHAN  CX
AND WALK NOW GENTLY THROUGH THE FIRE  ELWOOD, ROGER  SR-HA,HA
ANDERE SEITE DES HIMMELS, DIE  CLARKE, ARTHUR C.  FP
ANDERSEN’S FAIRY TALES  ANDERSEN, HANS CHRISTIAN  H,P
ANDRE NORTON: A PRIMARY AND SECONDARY BIBLIOGRAPHY  SCHLOBIN, ROGER C./HARRISON, IRENE R.  R/IND
ANDROID  ZEIGFREID, KARL  SR-P
ANDROID ARMAGGEDON  TRALINS, ROBERT  P
ANDROID AT ARMS  NORTON, ANDRE  C/P:2,P
ANDROID AVENGER  WHITE, TED  DM123
ANDROID PLANET [SPACE: 1999 8]  RANKINE, JOHN  C/P,P
ANDROIDS ARE COMING, THE  SILVERBERG, ROBERT  LPA
ANDROID’S DREAM, THE  SCALZI, JOHN  H
ANDROIDS, TIME MACHINES AND BLUE GIRAFFES  ELWOOD, ROGER/GHIDALIA, VIC  HA
ANDROMEDA  YEFREMOV, IVAN  SR-P
ANDROMEDA 1  WESTON, PETER  SR-PA,HA
ANDROMEDA 2  WESTON, PETER  SR-PA
ANDROMEDA 3  WESTON, PETER  SR-PA
ANDROMEDA BREAKTHROUGH [ANDROMEDA 2]  HOYLE, FRED/ELLIOT, JOHN  C/P,P,H
ANDROMEDA GUN  BOYD, JOHN  C/P,P,H
ANDROMEDA STRAIN, THE  CRICHTON, MICHAEL  C/P,P,H
ANDY BUCKRAM’S TIN MEN  BRINK, CAROL RYRIE  P
ANDY WARHOL’S DRACULA  NEWMAN, KIM  MB002
ANGADO [DUMAREST 29]  TUBB, E. C.  C/P:2,P
ANGEL  KILWORTH, GARRY  P
ANGEL AND THE SWORD, THE  HOLLAND, CECELIA  LP
ANGEL AT APOGEE  LEWITT, SHARIANN  C/P:2,SFWA-P,P
ANGEL CHRONICLES, THE, VOL. 1 [BUFFY]  HOLDER, NANCY  C/P,P
ANGEL CHRONICLES, THE, VOL. 2 [BUFFY]  TANKERSLEY, RICHIE  C/P,P
ANGEL CHRONICLES, THE, VOL. 3 [BUFFY]  HOLDER, NANCY  C/P,P
ANGEL FIRE  GREELEY, ANDREW M.  SFWA-P
ANGEL FIRE EAST [VOID & WORD 3]  BROOKS, TERRY  H
ANGEL ISLAND  GILLMORE, INEZ HAYNES  LP
ANGEL OF DARKNESS  DE LINT, CHARLES  H,LP
ANGEL OF DESTRUCTION [JUDICIARY]  MATTHEWS, SUSAN R.  P
ANGEL OF PAIN, THE [DAVID LYDYARD 2]  STABLEFORD, BRIAN M.  H
ANGEL OF THE REVOLUTION, THE  GRIFFITH, GEORGE  LP

The MITSFS Titledex  September 24, 2007  Page 12
TITLE  AUTHOR  CODE
ANGEL SCENE  KADREY, RICHARD  ERD2
ANGEL-SEEKER [SAMARIA]  SHINN, SHARON  H,P
ANGEL SOULS AND DEVIL HEARTS [SHADOWS 2]  GOLDEN, CHRISTOPHER  C/P,P
ANGEL STATION  WILLIAMS, WALTER JON  C/P,SFWA-P,P,H
ANGEL STATION (ERRATA SHEET INCLUDED)  WILLIAMS, WALTER JON  SR-H,H,SFWA
ANGEL WITH THE SWORD  CHERRYH, C. J.  H:2,P
ANGELIC AVENGERS, THE  ANDREZEL, PIERRE  P
ANGELICA [SAMARIA]  SHINN, SHARON  H
ANGELMASS  ZAHN, TIMOTHY  H:2,P
ANGELS!  DANN, JACK/DOZOIS, GARDNER  PA
ANGELS & VISITATIONS  GAIMAN, NEIL  SR-H,H:2
ANGELS AND SPACESHIPS  BROWN, FREDRIC  H
ANGEL’S COMMAND, THE [CASTAWAYS OF THE FLYING DUTCHMAN]  JACQUES, BRIAN  P
ANGELS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,SFWA-P
ANGELS OF DARKNESS [@WARHAMMER 40K]  THORPE, GAV  P
ANGELS UNAWARE [QUANTUM LEAP]  STORM, L. ELIZABETH  P
ANGRY ANGEL, THE [SISTERS OF THE NIGHT]  YARBRO, CHELSEA QUINN  H,LP
ANGRY CANDY  ELLISON, HARLAN  H,LP
ANGRY ESPERS, THE  BIGGLE, LLOYD JR.  DD485
ANGRY GHOST, THE [DOC SAVAGE #86]  ROBESON, KENNETH  @P
ANGRY LEAD SKIES [GARRETT 10]  COOK, GLEN  C/P,P
ANGRY PLANET, THE  CROSS, JOHN KEIR  H:2
ANGUISHED DAWN, THE [CRADLE OF SATURN 2]  HOGAN, JAMES P.  P
ANHALTER DURCH DIE GALAXIS, PER  ADAMS, DOUGLAS  FP
ANIARA  MARTINSON, HARRY  LP

ANIMAL BRIGADE 3000 GREENBERG, MARTIN H./ WAUGH, CHARLES G.

PA

ANIMAL DOCTOR, THE  JERSILD, P. C.  H,SFWA
ANIMAL FARM  ORWELL, GEORGE  C/P,P,H
ANIMAL FORUM  LEVY, PHIL  LP
ANIMAL PEOPLE, THE  COBLENTZ, STANTON A.  C/P:2,P
ANIMAL WIFE, THE  THOMAS, ELIZABETH MARSHALL  P
ANIMALS  SKIPP, JOHN/SPECTOR, CRAIG  C/P:2,P
ANIMAUX DENATURES, LES  VERCORS  FP
ANIMIST  FORWARD, EVE  H
ANITA  ROBERTS, KEITH  P
ANN, THE GENTLE [@PHENWICK #20]  KIMBROUGH, KATHERYN  P
ANNA TO THE INFINITE POWER  AMES, MILDRED  P
ANNALS OF KLEPSIS  LAFFERTY, R. A.  P
ANNALS OF THE BLACK COMPANY [BLACK COMPANY 1,2,3]  COOK, GLEN  H:2
ANNALS OF THE HEECHEE, THE [HEECHEE 4]  POHL, FREDERIK  H:2,P,SFWA
ANNALS OF THE TIME PATROL  ANDERSON, POUL  H:2
ANNE RICE  ROBERTS, BETTE B.  R/REF-H
ANNE RICE TRIVIA BOOK, THE  RAMSLAND, KATHERINE  R/REF-P,C/REF-P
ANNI DELLA TURBINA, GLI  JAKES, JOHN  FH
ANNIHILATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #5]  ATHANS, PHILIP  P
ANNIHILATION [MORTAL KOMBAT 2]  PREISLER, JEROME  P
ANNIHILATION!  STATTEN, VARGO  SR-P
ANNIHILATION FACTOR  BAYLEY, BARRINGTON J.  D33710
ANNIHILATION SQUAD [WARHAMMER 40K LAST CHANCERS 3]  THORPE, GAV  P
ANNIHILIST [DOC SAVAGE #31]  ROBESON, KENNETH  C/P:2,@P
ANNO DRACULA [ANNO DRACULA 1]  NEWMAN, KIM  H,P
ANNOTATED ALICE, THE  CARROLL, LEWIS  SR-VLH,H
ANNOTATED BIBLIOGRAPHY OF RECURSIVE SCIENCE FICTION, AN  LEWIS, ANTHONY R.  R/IND
ANNOTATED DRACULA, THE  STOKER, BRAM  SR-VLH
ANNOTATED GUIDE TO ROBERT E. HOWARD’S SWORD AND SORCERY, THE  WEINBERG, ROBERT  R/REF-H
ANNOTATED GULLIVER’S TRAVELS, THE  SWIFT, JONATHAN  H
ANNOTATED HOBBIT, THE  TOLKIEN, J. R. R./ANDERSON, DOUGLAS A.  R/REF-H
ANNOTATED JULES VERNE: FROM THE EARTH TO THE MOON, THE  VERNE, JULES  SR-VLH

ANNOTATED JULES VERNE:

TWENTY THOUSAND LEAGUES UNDER THE SEA, THE VERNE, JULES

SR-VLP,H

ANNOTATED SHERLOCK HOLMES 1, THE  BARING-GOULD, WILLIAM S.  SR-VLH
ANNOTATED SHERLOCK HOLMES 2, THE  BARING-GOULD, WILLIAM S.  SR-VLH
ANNOTATED WIZARD OF OZ, THE  BAUM, L. FRANK  SR-H,H
1972 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1973 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1974 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA,HA
1975 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1976 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA,HA
1977 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1978 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA,HA
1979 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1980 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1981 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1982 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA

The MITSFS Titledex  September 24, 2007  Page 13
TITLE  AUTHOR  CODE
1983 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1984 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  HA,PA
1985 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1986 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
1987 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  HA,PA
1988 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A.  HA
1989 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  HA,PA
1990 ANNUAL WORLD’S BEST SF, THE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  C/PA,PA
ANNUNCIATE, THE  PARK, SEVERNA  H
ANOMALOUS PHENOMENA  VERNE, JULES  H
ANOMALY, THE  SOHL, JERRY  C/P,P
ANOME, THE  VANCE, JACK  C/P:2,P
ANONYMOUS REX  GARCIA, ERIC  LP
ANOTHER DAY, ANOTHER DUNGEON [CUPS AND SORCERY 1]  COSTIKYAN, GREG  P,SFWA-P
ANOTHER END  KING, VINCENT  C/P:2,P
ANOTHER FINE MYTH... [MYTH ADVENTURES 1]  ASPRIN, ROBERT LYNN  C/P:2,LP

ANOTHER GIRL, ANOTHER PLANET [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] DAY, MARTIN/BEECH, LEN

C/P,P

ANOTHER HEAVEN, ANOTHER EARTH  HOOVER, H. M.  LP
ANOTHER KIND  OLIVER, CHAD  SR-H,C/P:2,P
ANOTHER MAN’S HELL  KEMPER, WILLIAM E. JR.  SR-P
ANOTHER ORPHAN  KESSEL, JOHN  TD6,SFWA-TD6
ANOTHER PART OF THE GALAXY  CONKLIN, GROFF  PA,C/PA
ANOTHER ROUND AT THE SPACEPORT BAR  SCITHERS, GEORGE/SCHWEITZER, DARRELL PA,C/PA
ANOTHER SPACE — ANOTHER TIME  CAMPBELL, H. J.  SR-P
ANOTHER WORLD  DOZOIS, GARDNER  HA
ANSWER, THE  WYLIE, PHILIP  C/P,P
ANT MEN, THE  NORTH, ERIC  H,P
ANT-MEN OF TIBET AND OTHER STORIES, THE  PRINGLE, DAVID  LPA
ANTAGONIST [CHILDE CYCLE]  DICKSON, GORDON R./WIXON, DAVID W.  H
ANTARES DAWN [ANTARES 1]  MCCOLLUM, MICHAEL  C/P:2,P
ANTARES PASSAGE [ANTARES 2]  MCCOLLUM, MICHAEL  C/P,P
ANTARES VICTORY [ANTARES 3]  MCCOLLUM, MICHAEL  LP
ANTARKTOS CYCLE, THE  PRICE, ROBERT M.  LP
ANTENNA SYNDROME, THE  MARKS, ALAN  C/P,P
ANTES DEL FIN  CHILSON, ROBERT  FP
ANTHEM  RAND, AYN  C/P:2,P
ANTHOLOGY [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  WIECK, STEWART  PA
ANTHOLOGY: THREE NOVELS  CAMPBELL, JOHN W. JR.  SFWA
ANTHONOLOGY  ANTHONY, PIERS  C/P,SFWA-P,P,H
ANTHRAX MUTATION, THE  SCOTT, ALAN  C/P:2,P
ANTHROPOL  TRIMBLE, LOUIS  DH059
ANTHROPOS-SPECTER BEAST, THE  KONWICKI, TADEUSZ  H
ANTI-GRAV UNLIMITED  LONG, DUNCAN  C/P,P
ANTI-HOMMES, LES  RAYJEAN, MAX-ANDRE  FP
ANTI-ICE  BAXTER, STEPHEN  C/P,P
ANTI-MAN  KOONTZ, DEAN R.  P
ANTI-SOCIAL REGISTER  HITCHCOCK, ALFRED  PA
ANTI-UNIVERSUM [@PERRY RHODAN SILBERBANDE #68]  PABEL MOEWIG KG  FPR
ANTIBODIES [X-FILES]  ANDERSON, KEVIN J.  H
ANTIBODIES  SKAL, DAVID J.  H,P
ANTICHRIST  MCBIRNIE, W. S.  [C/RAN-P]
ANTICIPATIONS  PRIEST, CHRISTOPHER  C/PA,SFWA
ANTIGRAV: COSMIC COMEDIES BY SF MASTERS  STRICK, PHILIP  HA
ANTIMATTER [STAR TREK DS9]  VORNHOLT, JOHN  C/P:2,P
ANTINOMY  ROBINSON, SPIDER  C/P:2,P
ANTIPOPE, THE [BRENTFORD 1]  RANKIN, ROBERT  P
ANTIQUE VINTAGE, AN  PELAN, JOHN  B
ANTOLOGIA DE FICCAO CIENTIFICA 1  BERTASO, JOSE  FP
ANTOLOGIA DE FICCAO CIENTIFICA 2  BERTASO, JOSE  FP
ANTOLOGIA DE FICCAO CIENTIFICA 3  BERTASO, JOSE  FP
ANTOLOGIA ESPANOLA DE CIENCIA FICCION 1  TORRES, RAUL  FP
ANTON YORK, IMMORTAL  BINDER, EANDO  C/P:2,P
ANTON YORK, INMORTAL (SPANISH)  BINDER, EANDO  FP
ANTRAX [VOYAGE OF THE JERLE SHANNARA 2/SHANNARA JERLE]  BROOKS, TERRY  H,P
ANTRO, THE LIFE-GIVER  DEEGAN, JON J.  SR-P,H
ANUBIS GATES, THE  POWERS, TIM  C/P:2,P
ANUBIS MURDERS, THE [@DANGEROUS JOURNEYS 1]  GYGAX, GARY  C/P,SFWA-P,P
ANVIL, THE [GENERAL 3]  STIRLING, S. M./DRAKE, DAVID  C/P:2,SFWA-P,P
ANVIL  VAN PALLANDT, NICHOLAS  P
ANVIL OF ICE, THE [@WINTER OF THE WORLD #1]  ROHAN, MICHAEL SCOTT  C/P,P
ANVIL OF STARS [FORGE OF GOD 2]  BEAR, GREG  H:2,P
ANVIL OF THE HEART  HOLMES, BRUCE T.  P
ANVIL OF THE SUN [@CLOAK AND DAGGER #1]  GROELL, ANNE LESLEY  P

The MITSFS Titledex  September 24, 2007  Page 14
TITLE  AUTHOR  CODE
ANVIL OF THE WORLD, THE  BAKER, KAGE  H
ANY MAN SO DARING  HOYT, SARAH A.  P
ANYTHING BOX, THE  HENDERSON, ZENNA  C/P,P,H
ANYTHING TREE, THE  RACKHAM, JOHN  D89250
ANYTHING YOU CAN DO...  GARRETT, RANDALL  P
ANYTHING YOU CAN DO...  LANGART, DARREL T.  H:2
ANYTIME RINGS, THE  FARADAY, ROBERT  LP
ANYWHEN  BLISH, JAMES  H:2,P
ANYWHERE, ANYWHEN  ENGDAHL, SYLVIA LOUISE  SR-HA
ANYWHERE BUT HERE  OLTION, JERRY  H
APE AND ESSENCE  HUXLEY, ALDOUS  C/P,P,H
APEMAN, SPACEMAN  STOVER, LEON E./HARRISON, HARRY  PA
APHILE [@PERRY RHODAN SILBERBANDE #81]  PABEL MOEWIG KG  FPR
APHRODISIAC GIRL  KOONTZ, DEAN R./KOONTZ, GERDA  SR-L
APOCALYPSE [@MAGIC: THE GATHERING INVASION 3]  KING, J. ROBERT  C/P,P
APOCALYPSE  SPRINGER, NANCY  C/P:2,SFWA-P,P
APOCALYPSE ARRAY [ARCHANGEL 4]  MOREHOUSE, LYDA  P
APOCALYPSE DOOR, THE  MACDONALD, JAMES D.  H
APOCALYPSE MEMORIES [@BUFFY]  BURNS, LAURA J./METZ, MELINDA  C/P,P
APOCALYPSE TROLL, THE  WEBER, DAVID  C/P,P,H
APOCALYPSES  LAFFERTY, R. A.  C/P,P
APOCALYPSO  RANKIN, ROBERT  P
APOLLO AT GO  SUTTON, JEFF  H:2,P
APOLLO LEGACY, THE  BARKER, ALBERT  P
APOLLO 13  LOVELL, JIM/KLUGER, JEFFREY  P
APOLLO’S WAR [@BATTLESTAR GALACTICA 1978 #13]  LARSON, GLEN A./THURSTON, ROBERT  P,SFWA-P
APOLOGY FOR RAIN, AN  GAWRON, JEAN MARK  H,SFWA
APOSTATE, THE  LONARDO, PAUL  LP
APOSTLE  LOVIN, ROGER  H,LP
APOSTROPHES & APOCALYPSES  BARNES, JOHN  H,LP
APPARITION TRAIL, THE  SMEDMAN, LISA  LP
APPLESEED  CLUTE, JOHN  LP
APPOINTMENT AT BLOODSTAR [FAMILY D’ALEMBERT #5]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
APPOINTMENT WITH YESTERDAY  DORN, FRANK  P
APPRENTICE, THE  TALMADGE-BICKMORE, DEBORAH  P
APPRENTICE FANTASTIC  GREENBERG, MARTIN H./DAVIS, RUSSELL  C/PA,PA
APPROACHING OBLIVION  ELLISON, HARLAN  H:2,SFWA,P
APPROACHING STORM, THE [@STAR WARS]  FOSTER, ALAN DEAN  H,P
AQUAMANCER [MANCER 2]  CALLANDER, DON  P,SFWA-P
AQUARIAN ATTACK, THE [SEEDS OF WAR 3]  RANDLE, KEVIN/CORNETT, ROBERT  C/P:2,P
AQUARIUS  MARK, JAN  SFWA
AQUARIUS MISSION  CAIDIN, MARTIN  C/P,P
AQUATECH WARRIORS [@TOM SWIFT (ARCHWAY) #6]  APPLETON, VICTOR  P
AQUILA AND THE IRON HORSE [@AQUILIAD #2]  SOMTOW, S. P.  P
AQUILA AND THE SPHINX [@AQUILIAD #3]  SOMTOW, S. P.  P
AQUILA IN THE NEW WORLD [@AQUILIAD #1]  SOMTOW, S. P.  C/P,P
AQUILIAD, THE  SUCHARITKUL, SOMTOW  C/P:2,P
ARABESQUES  SHWARTZ, SUSAN  PA,C/PA
ARABESQUES 2  SHWARTZ, SUSAN  C/PA,PA
ARABIAN NIGHTS  CAMPBELL, JOSEPH  PA
ARABIAN NIGHTS, THE  WIGGIN, KATE DOUGLAS/SMITH, NORA A.  LP
ARACHNE  MASON, LISA  H,SFWA,P
ARAFEL’S SAGA [ARAFEL 1,2]  CHERRYH, C. J.  H:2
ARAMAYA  ROUTLEY, JANE  P
ARAMINTA STATION [CADWAL 1]  VANCE, JACK  C/P,SFWA-P,P,H
ARAMINTA STATION (VIE) [@VANCE INTEGRAL EDITION #39/CADWAL 1]  VANCE, JACK  SR-H
ARA’S FIELD  MARKS, LAURIE J.  P

ARBOR HOUSE TREASURY OF GREAT SCIENCE FICTION SHORT NOVELS, THE SILVERBERG, ROBERT/

GREENBERG, MARTIN H.

HA

ARC OF INFINITY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
ARC OF THE DREAM [RADIX 3]  ATTANASIO, A. A.  C/P,P
ARC RIDERS [ARC RIDERS]  DRAKE, DAVID/MORRIS, JANET  C/P:2,P
ARCADE  MAXXE, ROBERT  H
ARCADY [ARCADY 1]  WILLIAMS, MICHAEL  C/P,P
ARCANE  SHERRELL, CARL  C/P:2,P
ARCHANGEL  CONNER, MICHAEL  P
ARCHANGEL [SAMARIA 1]  SHINN, SHARON  C/P:2,P
ARCHANGEL BLUES [@HELLFLOWER #3]  BES SHAHAR, ELUKI  P
ARCHANGEL PROTOCOL [ARCHANGEL 1]  MOREHOUSE, LYDA  P
ARCHER’S GOON  JONES, DIANA WYNNE  P,SFWA-P
ARCHFORM: BEAUTY [ARCHFORM BEAUTY 1]  MODESITT, L. E. JR.  H
ARCHITECT OF SLEEP, THE  BOYETT, STEVEN R.  P
ARCHITECTS OF DREAMS  BAILEY, ROBIN WAYNE  HA,LPA
ARCHITECTS OF EMORTALITY [EMORTALITY]  STABLEFORD, BRIAN M.  H

The MITSFS Titledex  September 24, 2007  Page 15
TITLE  AUTHOR  CODE
ARCHITECTS OF HYPERSPACE, THE  MCDONOUGH, THOMAS R.  C/P:2,P
ARCHITECTURE OF DESIRE, THE  GENTLE, MARY  P,SFWA
ARCHITECTURE OF FEAR, THE  CRAMER, KATHRYN/PAUTZ, PETER D.  SFWA
ARCHIVES OF HAVEN, THE [LEMMUS #3]  SAVARIN, JULIAN JAY  @SR-P,H
ARCHON CONSPIRACY, THE  HUNT, DAVE  LP
ARCHONS OF THE STARS, THE [@DRAGON THRONE #3]  BAIRD, ALISON  LP
ARCTURUS LANDING  DICKSON, GORDON R.  C/P:2,SFWA-P,P
ARDNEH’S WORLD [EMPIRE OF THE EAST 3]  SABERHAGEN, FRED  P,SFWA-P
ARDOR ON AROS  OFFUTT, ANDREW J.  C/P,P
ARDREYS, THE  VERNON, KAY  P
ARE ALL THE GIANTS DEAD?  NORTON, MARY  LP
ARE WE HAVING FUN YET?  SANDERS, WILLIAM  LP
AREA 51 [AREA 51 1]  DOHERTY, ROBERT  P
AREA 51: EXCALIBUR [AREA 51 6]  DOHERTY, ROBERT  P
AREA 51: LEGEND [AREA 51 9]  DOHERTY, ROBERT  C/P,P
AREA 51: NOSFERATU [@AREA 51 8]  DOHERTY, ROBERT  P
AREA 51: NOSFERATU (PROOF) [AREA 51 8]  DOHERTY, ROBERT  SR-P
AREA 51: THE GRAIL [AREA 51 5]  DOHERTY, ROBERT  P
AREA 51: THE MISSION [AREA 51 3]  DOHERTY, ROBERT  P
AREA 51: THE REPLY [AREA 51 2]  DOHERTY, ROBERT  P
AREA 51: THE SPHINX [AREA 51 4]  DOHERTY, ROBERT  P
AREA 51: THE TRUTH [@AREA 51 7]  DOHERTY, ROBERT  P
ARENA [MAGIC: THE GATHERING]  FORSTCHEN, WILLIAM R.  C/P:2,P
ARENA OF ANTARES [DRAY PRESCOT 7]  AKERS, ALAN BURT  P
ARENA: SPORTS SF  FERMAN, EDWARD L./MALZBERG, BARRY N. HA,SFWA
ARES EXPRESS  MCDONALD, IAN  P
ARGENTIS  SHAW, BRIAN  SR-P
ARGONAUT  SCHMIDT, STANLEY  H
ARGONAUT AFFAIR, THE [TIMEWARS 7]  HAWKE, SIMON  P
ARGUS GAMBIT, THE [DREAMERS OF THE DAY 1]  ROSS, DAVID D.  H,SFWA
ARIADNE’S WEB [GODS 2]  SABERHAGEN, FRED  P
ARIEL  BICKHAM, JACK M.  P,SFWA-P
ARIEL  BOYETT, STEVEN R.  C/P:2,P
ARIOSTO  YARBRO, CHELSEA QUINN  C/P:2,P
ARISIA ’00 (PROGRAM BOOK)  ARISIA 2000 CONVENTION COMMITTEE  VLPA
ARISIA ’98 (PROGRAM BOOK)  ARISIA 1998 CONVENTION COMMITTEE  VLPA
ARISTOI  WILLIAMS, WALTER JON  P
ARISTOKIA  PEZET, A. WASHINGTON  SR-H
ARK, THE [DOCTOR WHO]  ERICKSON, PAUL  SR-P,C/P
ARK, THE  SZYDLOW, JARL  P
ARK LIBERTY  BRADLEY, WILL  C/P,SFWA-P,P
ARK OF DOOM [MAN FROM ATLANTIS 4]  WOODLEY, RICHARD  P
ARK OF MARS, THE  BRACKETT, LEIGH  B
ARK OF THE STARS [@PERRY RHODAN LEMURA #1]  BORSCH, FRANK  P
ARK OF VENUS  CLASON, CLYDE B.  H
ARK TWO [@NEW TOM SWIFT #7]  APPLETON, VICTOR  H,LP
ARKADIANS, THE (PROOF)  ALEXANDER, LLOYD  SR-LP
ARKHAM TALES  JONES, WILLIAM  LPA
ARM OF THE STARFISH, THE  L’ENGLE, MADELEINE  C/P,P
ARM OF THE STONE, THE  STRAUSS, VICTORIA  P
ARMADA  JAHN, MICHAEL  C/P:2,P
ARMADA OF ANTARES [DRAY PRESCOT 11]  AKERS, ALAN BURT  C/P,P
ARMADA OF THE AIR  BENTLY, NORMAN S.  H
ARMADILHA EM ZARKASS  WUL, STEFAN  FP
ARMAGEDDON  BOLIN, M. C.  P
ARMAGEDDON  DRAKE, DAVID/MOSIMAN, BILLIE SUE  C/PA,PA
ARMAGEDDON [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/GOLDEN, CHRISTOPHER  P
ARMAGEDDON! [@THERE WILL BE WAR #8]  POURNELLE, JERRY  PA,SFWA-P
ARMAGEDDON 2000 [OMEN 4]  MCGILL, GORDON  P
ARMAGEDDON 2419 A.D. [BUCK ROGERS]  NOWLAN, PHILIP FRANCIS  C/P,P,H

ARMAGEDDON 2419 A.D. (REVISED) [BUCK ROGERS/ BUCK ROGERS (NIVEN & POURNELLE’S) 1] NOWLAN, PHILIP FRANCIS

C/P,P

ARMAGEDDON BLUES, THE  MORAN, DANIEL KEYS  C/P,P
ARMAGEDDON BOX, THE  WEINBERG, ROBERT  P
ARMAGEDDON CRAZY, THE  FARREN, MICK  C/P,P
ARMAGEDDON INHERITANCE, THE [COLIN MACINTYRE 2]  WEBER, DAVID  C/P,SFWA-P,P
ARMAGEDDON OFF VESTA [BUCK ROGERS MARTIAN WARS 3]  MURDOCK, M. S.  C/P,P
ARMAGEDDON RAG, THE  MARTIN, GEORGE R. R.  C/P,P,H
ARMAGEDDON SKY [STAR TREK DAY OF HONOR 2/STAR TREK DS9]  GRAF, L. A.  C/P:2,P
ARMAGEDDON: THE MUSICAL [ARMAGEDDON 1]  RANKIN, ROBERT  LP
ARMAGEDDON: THE MUSICAL (PROOF) [ARMAGEDDON 1]  RANKIN, ROBERT  SR-LP
ARMAGEDDONS  DANN, JACK/DOZOIS, GARDNER  PA
ARMAGEDDON’S CHILDREN  BROOKS, TERRY  H
ARMED CAMPS  REED, KIT  C/P,P

The MITSFS Titledex  September 24, 2007  Page 16
TITLE  AUTHOR  CODE
ARMED MEMORY  YOUNG, JIM  C/P,P
ARMIES OF DAYLIGHT, THE [DARWATH 3]  HAMBLY, BARBARA  C/P:2,P
ARMIES OF ELFLAND, THE  ANDERSON, POUL  C/P:2,P
ARMIES OF LIGHT AND DARK [@BABYLON 5 LEGIONS OF FIRE #2]  DAVID, PETER  C/P,P
ARMIES OF MEMORY, THE [THOUSAND CULTURES 4]  BARNES, JOHN  H
ARMLESS MAIDEN, THE  WINDLING, TERRI  LPA
ARMLET OF THE GODS, THE [LAND BEYOND THE GATE 2]  ESHBACH, LLOYD ARTHUR  P
ARMOR  STEAKLEY, JOHN  C/P:2,P
ARMOR OF LIGHT, THE  SCOTT, MELISSA/BARNETT, LISA A.  C/P,SFWA-P,P,H
ARMORY OF SWORDS, AN [SWORDS]  SABERHAGEN, FRED  PA
ARMOUR OF CONTEMPT, THE [@WARHAMMER 40K]  ABNETT, DAN  H
ARMS OF HERCULES, THE [GODS 3]  SABERHAGEN, FRED  C/P,LP

ARMS OF THE KRAKEN, THE (NOVELLA FROM DRAGON#305)

[SONG OF ICE AND FIRE 4.01] MARTIN, GEORGE R. R.

H

ARMY OF DARKNESS [MAGIRA 2]  WALKER, HUGH  C/P,P
ARMY OF THE UNDEAD [INVADERS 3]  BERNARD, RAFE  P
AROUND THE WORLD IN EIGHTY DAYS  VERNE, JULES  C/P:2,P
ARRIVAL [BUCK ROGERS]  TSR  PA
ARRIVE AT EASTERWINE  LAFFERTY, R. A.  C/P,SFWA,P,H
ARROGANT HISTORY OF WHITE BEN, THE  DANE, CLEMENCE  H
ARROW BOOK OF GHOST STORIES  KRAMER, NORA  PA
ARROW FROM EARTH  BUSBY, F. M.  P
ARROW’S FALL [HERALDS OF VALDEMAR 3]  LACKEY, MERCEDES  C/P:2,P
ARROW’S FLIGHT [HERALDS OF VALDEMAR 2]  LACKEY, MERCEDES  C/P:2,P
ARROWS OF HERCULES, THE  DE CAMP, L. SPRAGUE  P
ARROWS OF THE QUEEN [HERALDS OF VALDEMAR 1]  LACKEY, MERCEDES  C/P:2,P
ARROWS OF THE SUN [@AVARYAN RISING #4]  TARR, JUDITH  H:2,P
ARS MAGICA  TARR, JUDITH  H:2,P
ARSENAL, THE [@SURVIVALIST #16]  AHERN, JERRY  P
ARSENAL OF MIRACLES, THE  FOX, GARDNER F.  P,DF299
ARSENAL OUT OF TIME, THE  MCDANIEL, DAVID  C/P:2,P
ARSENE LUPIN  LEBLANC, MAURICE  H
ARSLAN  ENGH, M. J.  C/P,SFWA,SFWA-P,P
ART IN THE BLOOD [@VAMPIRE FILES #4]  ELROD, P. N.  C/P,P
ART OF ARROW CUTTING, THE [MAGE AND TAKUMO 1]  DEDMAN, STEPHEN  LP
ART OF DESTRUCTION, THE [@DOCTOR WHO]  COLE, STEPHEN  H

ART OF GEORGE R.R. MARTIN’S A SONG OF ICE AND FIRE, THE [SONG OF ICE AND FIRE] WOOD, BRIAN

A

ART OF MICHAEL WHELAN, THE  WHELAN, MICHAEL  A
ART OF STAR WARS, THE [STAR WARS]  TITELMAN, CAROL  A
ART OF STAR WARS: EPISODE V, THE [STAR WARS]  CALL, DEBORAH  A
ART OF STAR WARS: EPISODE VI, THE [STAR WARS]  LUCASFILM LTD.  A
ART OF THE BROTHERS HILDEBRANDT, THE  SUMMERS, IAN  A
ART OF THE IMPOSSIBLE, THE [@STAR TREK LOST ERA (2328-46 AD)]  DECANDIDO, KEITH R. A.  P
ART OF THE SWORD, THE [@NIGHT-THREADS 5]  EMERSON, RU  C/P,P
ARTEMIS FOWL [ARTEMIS FOWL 1]  COLFER, EOIN  C/P,LP
ARTEMIS FOWL: THE ARCTIC INCIDENT [ARTEMIS FOWL 2]  COLFER, EOIN  LP
ARTEMIS FOWL: THE ETERNITY CODE [ARTEMIS FOWL 3]  COLFER, EOIN  H
ARTEMIS FOWL: THE LOST COLONY [ARTEMIS FOWL 5]  COLFER, EOIN  H
ARTEMIS FOWL: THE OPAL DECEPTION [ARTEMIS FOWL 4]  COLFER, EOIN  H
ARTERY OF FIRE  SCORTIA, THOMAS N.  H:2,SFWA,P
ARTHUR [@PENDRAGON #3]  LAWHEAD, STEPHEN R.  P
ARTHUR AND THE FORBIDDEN CITY  BESSON, LUC  LP
ARTHUR C. CLARKE  MCALEER, NEIL  R/REF-H

ARTHUR C. CLARKE OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

R/REF-H

ARTHUR C. CLARKE  RABKIN, ERIC S.  R/REF-P
ARTHUR C. CLARKE’S JULY 20, 2019  CLARKE, ARTHUR C.  H:2
ARTHUR, KING  ANDERSON, DENNIS LEE  C/P,P
ARTHUR REX  BERGER, THOMAS  P
ARTHUR WAR LORD [ARTHUR WAR LORD 1]  AB HUGH, DAFYDD  P
ARTICLE 23 [JUSTIN BELL 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
ARTICLES OF THE FEDERATION [STAR TREK]  DECANDIDO, KEITH R. A.  P
ARTIFACT  BENFORD, GREGORY  C/P:2,SFWA-P:2,P,H
ARTIFACT, THE  GEAR, W. MICHAEL  C/P,P
ARTIFACT OF EVIL [@GREYHAWK #2]  GYGAX, GARY  C/P:2,P
ARTIFACT OF THE SYSTEM [@BLACK HOLE TRAVEL AGENCY #2]  MCKINNEY, JACK  C/P:2,P
ARTIFACTS  JONES, WILLIAM  LP
ARTIFICIAL INTELLIGENCE: HOW MACHINES THINK  PEAT, F. DAVID  SFWA-P
ARTIFICIAL KID, THE  STERLING, BRUCE  C/P,SFWA-P,H,P,SFWA
ARTIFICIAL MAN, THE  DAVIES, L. P.  H:2,P
ARTIFICIAL THINGS  FOWLER, KAREN JOY  C/P:2,P
ARTS AND BEYOND, THE  MONTELEONE, THOMAS F.  HA,SFWA
AS DREAM AND SHADOW  PETAJA, EMIL  SR-H

TITLE  AUTHOR  CODE
AS IT IS ON MARS (REVISED SECOND EDITION) [AS IT IS ON MARS 1]  CRONIN, THOMAS W.  H
AS ON A DARKLING PLAIN  BOVA, BEN  C/P:2,SFWA-P,H,P,SFWA
AS PUCK WOULD HAVE IT [@CHARMED]  RUDITIS, PAUL  P
AS SHE CLIMBED ACROSS THE TABLE  LETHEM, JONATHAN  H
AS THE CURTAIN FALLS  CHILSON, ROBERT  C/P:2,P
AS THE GREEN STAR RISES [GREEN STAR 4]  CARTER, LIN  C/P,P
ASCENDANCE [DEMONWARS II 1]  SALVATORE, R. A.  H,P
ASCENDANCIES  COMPTON, D. G.  H,SFWA,P
ASCENDANT SUN [SKOLIAN EMPIRE 5]  ASARO, CATHERINE  H,P
ASCENDING [LEAGUE OF PEOPLES 4]  GARDNER, JAMES ALAN  H,P
ASCENDING, THE  WRIGHT, T. M.  H
ASCENSION [@WATER #1]  DALKEY, KARA  LP
ASCENSION [@WARHAMMER 40K DAWN OF WAR 2]  GOTO, C. S.  P
ASCENSION  GRANT, CHARLES L.  P
ASCENSION FACTOR, THE [SHIP 4]  HERBERT, FRANK/RANSOM, BILL  C/P,SFWA-P,P,H
ASCENSION WARRIOR, THE [@HORIZON WAR #2/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P
ASCENT OF WONDER, THE  HARTWELL, DAVID G./CRAMER, KATHRYN  HA
ASCENTS OF WONDER  GERROLD, DAVID  C/PA,PA
ASGARD RUN, THE  VANCE, STEVE  C/P,P
ASH OCK [PARATWA 2]  HINZ, CHRISTOPHER  H,P
ASH OF STARS  SALLIS, JAMES  R/REF-P
ASH STAFF, THE [ASH STAFF 1]  FISHER, PAUL R.  C/P,P
ASH WEDNESDAY  WILLIAMSON, CHET  SFWA-P

ASHES AND ANGEL WINGS [@WORLD OF DARKNESS DEMON THE FALLEN #1] STOLZE, GREG

P

ASHES AND STARS  ZEBROWSKI, GEORGE  C/P,P
ASHES, ASHES  BARJAVEL, RENE  H,P
ASHES OF EDEN, THE [STAR TREK ODYSSEY 1]  SHATNER, WILLIAM  H,P
ASHES OF TAMAR, THE  WADE, ELIZABETH  P
ASHES OF THE SUN [MAGIC: THE GATHERING]  BRADDOCK, HANOVI  C/P,P
ASHES OF VICTORY [HONOR HARRINGTON 9]  WEBER, DAVID  H:2,P
ASHES TO ASHES: THE DRAGON’S FLAME [@BLACKSTONE #3]  SAUL, JOHN  P
ASHLING [OBERNEWTYN 3]  CARMODY, ISOBELLE  H
ASHTON HORROR, THE [@DARK FORCES 12]  BRIDGES, LAURIE  P
ASIMOV ANALYZED  GOBLE, NEIL  R/REF-H,C/REF-H
ASIMOV CHRONICLES 1, THE  ASIMOV, ISAAC  C/P,P
ASIMOV CHRONICLES 2, THE  ASIMOV, ISAAC  P
ASIMOV CHRONICLES 3, THE  ASIMOV, ISAAC  P
ASIMOV CHRONICLES 4, THE  ASIMOV, ISAAC  C/P:2,P
ASIMOV CHRONICLES 5, THE  ASIMOV, ISAAC  C/P,P
ASIMOV ON NUMBERS  ASIMOV, ISAAC  [C/RAN-H]
ASIMOV SELECCION 1  ASIMOV, ISAAC  FP
ASIMOV SELECCION 3  ASIMOV, ISAAC  FP
ASIMOV’S CHOICE: ASTRONAUTS AND ANDROIDS  SCITHERS, GEORGE  PA
ASIMOV’S CHOICE: BLACK HOLES AND BUG-EYED MONSTERS  SCITHERS, GEORGE  PA
ASIMOV’S CHOICE: COMETS AND COMPUTERS  SCITHERS, GEORGE  PA
ASIMOV’S CHOICE: DARK STARS AND DRAGONS  SCITHERS, GEORGE  PA
ASIMOV’S CHOICE: EXTRATERRESTRIALS AND ECLIPSES  SCITHERS, GEORGE  PA
ASIMOV’S CHRONOLOGY OF SCIENCE AND DISCOVERY  ASIMOV, ISAAC  R/REF-H
ASIMOV’S MYSTERIES  ASIMOV, ISAAC  H:2,P
ASIMOV’S SCIENCE FICTION 30TH ANNIVERSARY ANTHOLOGY  WILLIAMS, SHEILA  LPA
ASSAMITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
ASSAMITE [@WORLD OF DARKNESS DARK AGES CLAN #2]  PETRUCHA, STEFAN  P
ASSASSIN, THE [@CASCA #13]  SADLER, BARRY  P

ASSASSIN FANTASTIC GREENBERG, MARTIN H./ POTTER, ALEXANDER

PA

ASSASSIN GAMBIT, THE [GAMESTER WARS 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
ASSASSIN OF GOR [GOR 5]  NORMAN, JOHN  C/P:2,P
ASSASSINATION AFFAIR, THE [@MAN FROM UNCLE #10]  HOLLY, JOAN HUNTER  C/P:2,P
ASSASSIN’S APPRENTICE [FARSEER 1]  HOBB, ROBIN  C/P:3,P
ASSASSIN’S BLADE [@MAGIC: THE GATHERING LEGENDS TWO 1]  MCGOUGH, SCOTT  P
ASSASSIN’S EDGE, THE [EINARINN 5]  MCKENNA, JULIET E.  P
ASSASSINS FROM TOMORROW [MIND BROTHERS 2]  HEATH, PETER  C/P:2,P
ASSASSINS OF TAMURIN, THE  TOWER, S. D.  H
ASSASSINS OF TAMURIN, THE (PROOF)  TOWER, S. D.  SR-LP
ASSASSIN’S QUEST [FARSEER 3]  HOBB, ROBIN  H,P
ASSAULT AT SELONIA [STAR WARS CORELLIAN 2]  ALLEN, ROGER MACBRIDE  C/P:2,P
ASSAULT ON ATLANTIS [ATLANTIS 5]  DONEGAN, GREG  P
ASSAULT ON OMEGA 4  HORVAT, DILWYN  P
ASSAULT ON THE GODS  GOLDIN, STEPHEN  P,SFWA
ASSEMBLERS OF INFINITY  ANDERSON, KEVIN J./BEASON, DOUG  C/P,P,H
ASSIGNMENT: ETERNITY [STAR TREK]  COX, GREG  C/P,P
ASSIGNMENT: HELLHOLE [@SPACEWAYS #14]  CLEVE, JOHN  P
ASSIGNMENT IN ETERNITY  HEINLEIN, ROBERT A.  SR-H,C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 18
TITLE  AUTHOR  CODE
ASSIGNMENT IN ETERNITY (ABRIDGED)  HEINLEIN, ROBERT A.  SR-P
ASSIGNMENT IN NOWHERE [IMPERIUM 3]  LAUMER, KEITH  C/P:2,P
ASSIGNMENT IN SPACE WITH RIP FOSTER  SAVAGE, BLAKE  H
ASSIGNMENT IN TOMORROW  POHL, FREDERIK  HA,PA
ASSIGNMENT — MOON GIRL  AARONS, EDWARD S.  P
ASSIGNMENT NOR’DYREN  VAN SCYOC, SYDNEY J.  C/P:2,P
ASSUMPTION OF RISK [@BATTLETECH 12]  STACKPOLE, MICHAEL A.  C/P,P
ASTERISK DESTINY  BLACK, CAMPBELL  P
ASTERIX THE LEGIONARY  GOSCINNY/UDERZO  [C/RAN-H]
ASTEROID FORMA  LE PAGE, RAND  SR-P
ASTEROID MAN  FANTHORPE, R. L.  H
ASTEROID WARS, THE [ASTEROID WARS 1,2,3]  BOVA, BEN  H
ASTEROIDE DEI PARIA, L  PLATT, CHARLES  FP
ASTOUNDING-ANALOG READER 1, THE  HARRISON, HARRY/ALDISS, BRIAN W.  HA,SFWA
ASTOUNDING-ANALOG READER 2, THE  HARRISON, HARRY/ALDISS, BRIAN W.  HA,SFWA
ASTOUNDING DAYS  CLARKE, ARTHUR C.  R/REF-P,C/REF-H
ASTOUNDING — JOHN W. CAMPBELL MEMORIAL ANTHOLOGY  HARRISON, HARRY  C/PA,PA,SFWA
ASTOUNDING SCIENCE FICTION ANTHOLOGY, THE  CAMPBELL, JOHN W. JR.  PA,HA
ASTOUNDING TALES OF SPACE AND TIME  CAMPBELL, JOHN W. JR.  PA,C/PA
ASTRA AND FLONDRIX  CULLEN, SEAMUS  C/P,LP
ASTRAL FORTRESS, THE [@NEW TOM SWIFT #5]  APPLETON, VICTOR  C/P,LP
ASTRAL MIRROR, THE  BOVA, BEN  C/P:2,SFWA-P,P
ASTRAL QUEST [SPACE: 1999 6]  RANKINE, JOHN  P
ASTRO RACE  LANG, KING  SR-P
ASTRONAUT, THE  BLUMGARTEN, JAMES  C/P,P
ASTRONAUT, THE  SEARLS, HANK  C/P:2,P
ASTRONAUTS AND ANDROIDS  DAVIS PUBLICATIONS  C/PA,PA
ASTRONAUTS MUST NOT LAND, THE  BRUNNER, JOHN  DF227
ASTRONAVE PIRATA, A  VAN VOGT, A. E.  FP
ASTRONEF PIRATE, L  LEINSTER, MURRAY  FP
ASTROPILOTS  MIXON, LAURA J.  C/P,P
ASUTRA, THE  VANCE, JACK  C/P:2,P
ASYLUM [DOCTOR WHO PAST DOCTORS]  DARVILL-EVANS, PETER  P
ASYLUM  JOHNSTON, WILLIAM  P
ASYLUM [@BLACKSTONE #6]  SAUL, JOHN  P
ASYLUM AND CIRCUS  BARRETT, MICHAEL DENNIS  P
ASYLUM EARTH  ELLIOTT, BRUCE  C/P:2,P
ASYLUM WORLD, THE  JAKES, JOHN  C/P:2,P
AT ALL COSTS [HONOR HARRINGTON 11]  WEBER, DAVID  H
AT ALL COSTS DISK, THE (CD) [HONOR HARRINGTON]  BAEN PUBLISHING ENTERPRISES  SR-MM
AT AMBERLEAF FAIR  KARR, PHYLLIS ANN  C/P,LP
AT ANY PRICE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
AT SWORD’S POINT [@KNIGHTS OF THE BLOOD]  MACMILLAN, SCOTT  C/P,P
AT SWORDS’ POINTS  NORTON, ANDRE  H:2
AT THE BACK OF THE NORTH WIND  MACDONALD, GEORGE  H,P
AT THE CITY LIMITS OF FATE  BISHOP, MICHAEL  SR-LP
AT THE EARTH’S CORE [PELLUCIDAR #1]  BURROUGHS, EDGAR RICE  H:2,@P

AT THE EARTH’S CORE [PELLUCIDAR 1,2,3]/ PELLUCIDAR [PELLUCIDAR 1,2,3]/ TANAR OF PELLUCIDAR [PELLUCIDAR 1,2,3] BURROUGHS, EDGAR RICE

LP

AT THE EDGE OF SPACE [HANAN REBELLION 1,2]  CHERRYH, C. J.  P
AT THE EDGE OF THE WORLD  DUNSANY, LORD  SR-P
AT THE EYE OF THE OCEAN  SCHENCK, HILBERT  P
AT THE MOUNTAINS OF MADNESS  LOVECRAFT, H. P.  H:2,SR-H,P
AT THE NARROW PASSAGE [ERIC MATHERS 1]  MEREDITH, RICHARD C.  C/P,P,H
AT THE NORTH POLE  VERNE, JULES  H
AT THE SEVENTH LEVEL [COYOTE JONES 3]  ELGIN, SUZETTE HADEN  C/P,P
AT WINTER’S END [WINTER 1]  SILVERBERG, ROBERT  H:2,P
ATHELING, THE  CHETWIN, GRACE  H,P
ATHENA, ACT 1 [ATHENA 1]  HSIEH, DEAN  SCX
ATHENA, ACT 2 [ATHENA 2]  HSIEH, DEAN  SCX
ATHYRA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,SFWA-P:2,P
ATLAN [ATLAN]  GASKELL, JANE  C/P,P,H
ATLAN IN DANGER [@PERRY RHODAN #82]  ACE  C/PA,PA
ATLANTEAN CHRONICLES  EICHNER, HENRY M.  R/REF-H
ATLANTIC ABOMINATION, THE  BRUNNER, JOHN  C/P:2,DD465,P
ATLANTIDA  BENOIT, PIERRE  P

ATLANTIS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

ATLANTIS [ATLANTIS 1]  DONEGAN, GREG  C/P,P
ATLANTIS: DEVIL’S SEA [ATLANTIS 3]  DONEGAN, GREG  P
ATLANTIS ENDGAME [TIME TRADERS]  NORTON, ANDRE/SMITH, SHERWOOD  P
ATLANTIS FOUND  GARCIA Y ROBERTSON, R.  P
ATLANTIS RISING  DAY, ALYSSA  P

The MITSFS Titledex  September 24, 2007  Page 19
TITLE  AUTHOR  CODE
ATLANTIS RISING  STEIGER, BRAD  [C/RAN-P]
ATLANTIS STATION [@STAR TREK TNG STARFLEET ACADEMY #5]  MITCHELL, V. E.  H,LP
ATLANTIS: THREE TALES  DELANY, SAMUEL R.  H
ATLAS OF FANTASY, AN  POST, J. B.  A
ATLAS OF MIDDLE-EARTH, THE [LORD OF THE RINGS]  FONSTAD, KAREN WYNN  R/REF-H
ATLAS OF PERN, THE [PERN]  FONSTAD, KAREN WYNN  R/REF-H,C/REF-H:3
ATLAS SHRUGGED  RAND, AYN  C/P:2,P
ATOM  AYLETT, STEVE  H
ATOM BOMB BLUES [DOCTOR WHO PAST DOCTORS]  CARTMEL, ANDREW  P
ATOM CONSPIRACY, THE  SUTTON, JEFF  C/P,P,H
ATOM CURTAIN, THE  WILLIAMS, NICK BODDIE  DD139
ATOM FIRE ON MECHANICA [PERRY RHODAN 121]  DARLTON, CLARK  SR-LP
ATOM HELL OF GRAUTIER, THE [@PERRY RHODAN #71]  ACE  PA
ATOMIC NEMESIS  ZEIGFREID, KARL  SR-P
ATOMIC STORY, THE  CAMPBELL, JOHN W. JR.  R/REF-H
ATOMS AND EVIL  BLOCH, ROBERT  C/P:2,P
ATOMS IN ACTION  SHELDON, ROY  SR-P,H
ATOOM-ALARM! (DUTCH) [@PERRY RHODAN #5]  MAHR, KURT  FPR
ATROCITY ARCHIVES, THE [ATROCITY ARCHIVES 1]  STROSS, CHARLES  H
ATTA  BELLAMY, FRANCIS RUFUS  DD079,P
ATTACK FROM ATLANTIS  DEL REY, LESTER  H:2,P
ATTACK FROM THE UNSEEN [@PERRY RHODAN #50]  ACE  PA,C/PA
ATTACK OF THE CLONES [STAR WARS MOVIE 2]  SALVATORE, R. A.  H
ATTACK OF THE JAZZ GIANTS  FROST, GREGORY  H
ATTACK OF THE MUTANT [GOOSEBUMPS]  STINE, R. L.  LP
ATTAR’S REVENGE [@ATTAR #1]  GRAHAM, ROBERT  C/P,P
ATTITUDE OF SILENCE, THE [ANDROMEDA (GENE RODDENBERRY’S) 5]  MARIOTTE, JEFF  H
AU SEUIL DU FUTUR  FAST, HOWARD  FP
AUBADE FOR GAMELON  WILLETT, JOHN  C/P,P
AUDIO EXTRACTS (CD)  PEERY, NORRIS RAY  SR-MM
AUDIO SAMPLER 1 (CASSETTE)  SPINE-TINGLING PRESS  SR-MM
AUDREY ROSE  DE FELITTA, FRANK  H,P
AUF UNBEKANNTEM STERN  KOLNBERGER, ANTON  FH
AUGUR’S TEACHER [EARTH FINAL CONFLICT]  SMITH, SHERWOOD  H,LP
AUGUSTA, THE FIRST [@PHENWICK #1]  KIMBROUGH, KATHERYN  P
AUGUSTA, THE SECOND [@PHENWICK #28]  KIMBROUGH, KATHERYN  P
AUNT MARIA  JONES, DIANA WYNNE  LP
AURIAN [AURIAN 1]  FUREY, MAGGIE  C/P,P
AURORA [ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  P

AURORA: BEYOND EQUALITY MCINTYRE, VONDA N./

ANDERSON, SUSAN JANICE

C/PA,PA

AURORA QUEST [@EARTHBLOOD]  AXLER, JAMES  P
AUSTRALIAN SCIENCE FICTION  IKIN, VAN  HA,LPA
AUSTRALIAN SCIENCE FICTION INDEX 1939–1962  STONE, GRAHAM  R/IND:2

AUSTRALIEN ABSURDITIES MCKENZIE, CHUCK/

ROBERTS, TANSY RAYNER

LPA

AUTHENTIC TOUCH, THE  WODHAMS, JACK  C/P:2,P
AUTHOR INDEX TO THE DOC SAVAGE MAGAZINE, AN  CLARK, WILLIAM J.  R/IND
AUTOCRACY OF MR. PARHAM, THE  WELLS, H. G.  H
AUTOCRISI  PROSPERI, PIERO  FP
AUTOMATED ALICE  NOON, JEFF  P
AUTOMATED GOLIATH, THE  TEMPLE, WILLIAM F.  DF129
AUTOPSY FOR A COSMONAUT  HAY, JACOB/KESHISHIAN, JOHN M.  C/P,P
AUTUMN ACCELERATOR, THE  LESLIE, PETER  P
AUTUMN ANGELS  COVER, ARTHUR BYRON  C/P,P
AUTUMN MIST [DOCTOR WHO EIGHTH DOCTOR]  MCINTEE, DAVID A.  P
AVALANCHE SOLDIER  MATTHEWS, SUSAN R.  C/P:2,P
AVALON  LAWHEAD, STEPHEN R.  P
AVARYAN RESPLENDENT [AVARYAN RISING 4,5,6]  TARR, JUDITH  LP
AVATAR, THE  ANDERSON, POUL  C/P,P,H
AVATAR [@INDIGO #6]  COOPER, LOUISE  C/P:2,P
AVATAR [@ANGEL]  PASSARELLA, JOHN  P
AVATAR, BOOK ONE [@STAR TREK DS9]  PERRY, S. D.  P
AVATAR, BOOK TWO [@STAR TREK DS9]  PERRY, S. D.  P
AVATARS, THE [@PRIMAL RAGE]  VORNHOLT, JOHN  P
AVENGER, THE [CHAOS GATE 3]  COOPER, LOUISE  C/P:2,P
AVENGER [STAR TREK ODYSSEY 3]  SHATNER, WILLIAM  P
AVENGER OF ANTARES [DRAY PRESCOT 10]  AKERS, ALAN BURT  P
AVENGERS, THE  KAEWERT, JULIE  P

AVENGERS AND THE THUNDERBOLTS, THE [AVENGERS (MARVEL COMICS CHARACTERS)] ASKEGREN, PIERCE

C/P,P

AVENGERS BATTLE THE EARTH-WRECKER, THE [AVENGERS (MARVEL COMICS CHARACTERS)] BINDER, OTTO

C/P,P

AVENGERS OF CARRIG, THE  BRUNNER, JOHN  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 20
TITLE  AUTHOR  CODE
AVENGING MARTIAN, THE  STATTEN, VARGO  SR-P
AVENGING RAY, THE  SMALL, AUSTIN J.  H
AVENGING SPIRIT, THE [TWILIGHT 10]  STEVENSON, E.  P
AVENTINE  KILLOUGH, LEE  C/P:2,P
AVENTURES D’ALICE AU PAYS DES MERVEILLES  CARROLL, LEWIS  FP

AVON FANTASY READER, THE WOLLHEIM, DONALD A./ ERNSBERGER, GEORGE

PA,C/PA

2ND AVON FANTASY READER, THE WOLLHEIM, DONALD A./ ERNSBERGER, GEORGE

C/PA,PA

AVON GHOST READER  AVON  SR-PA
AVONTUUR DER NATUURWETENSCHAP, HET  GIRARDEAU, EMILE  FP
AVONTUUR IN HET VERLEDEN  ANDERSON, POUL  FP
AVP: ALIEN VS. PREDATOR [ALIENS/PREDATOR]  CERASINI, MARC  P
AVRAM DAVIDSON TREASURY, THE  DAVIDSON, AVRAM  H
AWAKEN ME DARKLY [ALIEN HUNTRESS 1]  SHOWALTER, GENA  LP
AWAKENERS, THE [AWAKENERS 1,2]  TEPPER, SHERI S.  H:2,LP
AWAKENING, THE [@SURVIVALIST #10]  AHERN, JERRY  P
AWAKENING, THE [@MAGELORD 1]  MARTIN, THOMAS K.  P
AWAKENING, THE [DOCTOR WHO]  PRINGLE, ERIC  SR-P,C/P
AWAKENING [@GUNDAM MOBILE SUIT #1]  TOMINO, YOSHIYUKI  C/P,P
AWAKENING DREAM, THE [WHISPERING HILLS 5]  CAMERON, KATE  P
AWAKENING OF ZOJAS, THE  MICHELSON, MIRIAM  H
AWARD SCIENCE FICTION READER  NORTON, ALDEN H.  PA,C/PA
AWARE, THE [@ISLES OF GLORY #1]  LARKE, GLENDA  C/P,P
AWAY AND BEYOND  VAN VOGT, A. E.  C/P:2,P
AWAY FROM THE HERE AND NOW  HARRIS, CLARE WINGER  H
AWFUL EGG, THE [DOC SAVAGE #92]  ROBESON, KENNETH  @P
AXIOMATIC  EGAN, GREG  H
AXMAN COMETH, THE  FARRIS, JOHN  SFWA-P
AYES OF TEXAS, THE [TEXAS 1]  DA CRUZ, DANIEL  C/P,P
AYESHA: THE RETURN OF SHE [AYESHA 2]  HAGGARD, H. RIDER  C/P,P
AZAZEL  ASIMOV, ISAAC  C/P,P,H:2
AZRIEL UPRISING, THE  THOMPSON, ALLYN  P
AZTECS, THE [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P,C/P:2
AZURE BONDS [FORGOTTEN REALMS FINDER’S STONE 1]  NOVAK, KATE/GRUBB, JEFF  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 21
TITLE  AUTHOR  CODE
BAAA  MACAULAY, DAVID  H
BAAL  MCCAMMON, ROBERT R.  P
BAALBAK QUEST, THE [KILLSTAR 1]  KELLY, DAVID J.  SR-P
BABEL-17  DELANY, SAMUEL R.  C/P:2,P
BABEL-17 (ITALIAN)  DELANY, SAMUEL R.  FP
BABY  LIEBERMAN, ROBERT  H
BABY IS THREE [COMPLETE STORIES OF THEODORE STURGEON 6]  STURGEON, THEODORE  H
BABY IS THREE/AND MY FEAR IS GREAT  STURGEON, THEODORE  SR-P,C/P
BABY SQUAD, THE  NEIDERMAN, ANDREW  P
BABYLON FALLING  KEENE, BRIAN  BD00001
BABYLON GATE, THE  BYERS, EDWARD A.  C/P,P
BABYLON RISING  URBANCIK, JOHN  BD00001
BABYLON SISTERS  DI FILIPPO, PAUL  LP
BACHMAN BOOKS, THE  KING, STEPHEN  C/P,LP,H
BACK FOR MORE  WRIGHTSON, BERNI  CX
BACK FROM HELL [CAR WARRIORS 3]  FARREN, MICK  P
BACK FROM THE DEAD [@NECROMUNDA]  KYME, NICK  P
BACK HOME AT THE O.R.G.Y. [MAN FROM ORGY]  MARK, TED  SR-P
BACK TO THE FUTURE [BACK TO THE FUTURE 1]  GIPE, GEORGE  C/P,P
BACK TO THE FUTURE PART 2 [BACK TO THE FUTURE 2]  GARDNER, CRAIG SHAW  P
BACK TO THE FUTURE PART 3 [BACK TO THE FUTURE 3]  GARDNER, CRAIG SHAW  P
BACK TO THE MOON  HICKAM, HOMER H. JR.  P
BACK TO THE STONE AGE [PELLUCIDAR #5]  BURROUGHS, EDGAR RICE  C/P,@P,H
BACK TO THE TIME TRAP  LAUMER, KEITH  C/P,P
BACKBLAST [CONEK HAYDEN]  MCKEONE, LEE  P
BACKDROP OF STARS  HARRISON, HARRY  PA
BACKFLASH [@RACK #3]  JAMES, LAURENCE  P
BACKLASH MISSION, THE [BLACKCOLLAR 2]  ZAHN, TIMOTHY  C/P,P
BACKSHOT [@STARFIST FORCE RECON 1]  SHERMAN, DAVID/CRAGG, DAN  P
BACKWARD IN TIME  KELLEY, LEO P.  SR-P
BACTRA WAR, THE [@STAR WARS X-WING 4]  STACKPOLE, MICHAEL A.  P
BAD BARGAIN [@BUFFY]  GALLAGHER, DIANA G.  P
BAD DAY FOR ALI BABA, A [OTHER SINBAD 2]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
BAD DREAMS  NEWMAN, KIM  H
BAD MEDICINE [CHARLIE MOON 1]  YARBRO, CHELSEA QUINN  P
BAD MOON RISING [@JUDGE DREDD]  BISHOP, DAVID  P
BAD MOON RISING  DISCH, THOMAS M.  HA,SFWA
BAD PRINCE CHARLIE  MOORE, JOHN  P

BAD RONALD [@VANCE INTEGRAL EDITION #12]/ DARK OCEAN, THE [@VANCE INTEGRAL EDITION #12] VANCE, JACK

SR-H

BAD RONALD  VANCE, JOHN HOLBROOK  P
BAD SPELL IN YURT, A [WIZARD OF YURT 1]  BRITTAIN, C. DALE  P,SFWA-P
BAD THERAPY [DOCTOR WHO NEW ADVENTURES]  JONES, MATTHEW  C/P,P
BAD TIMING [STRONTIUM DOG]  LEVENE, REBECCA  P
BAD TIMING AND OTHER STORIES  BROWN, MOLLY  LP
BAD VOLTAGE  LITTELL, JONATHAN  P
BADGE OF INFAMY  DEL REY, LESTER  D76960
BADLAND [@WARLORD #3]  FROST, JASON  P
BADLANDS 1, THE [STAR TREK]  WRIGHT, SUSAN  C/P:2,P
BADLANDS 2, THE [STAR TREK]  WRIGHT, SUSAN  C/P,P
BAG OF BONES  KING, STEPHEN  H
BAGOMBO SNUFF BOX  VONNEGUT, KURT JR.  H
BAIT OF DREAMS, A  CLAYTON, JO  C/P:2,P
BAKER’S BOY, THE [BOOK OF WORDS 1]  JONES, J. V.  C/P,LP
BALAK  RAINEY, STEPHEN MARK  LP
BALANCE OF POWER [STAR TREK TNG]  AB HUGH, DAFYDD  C/P:2,P
BALANCE OF POWER [DAEDALUS 5]  STABLEFORD, BRIAN M.  C/P,P
BALANCE OF POWER, THE  WILLIAMS, MICHAEL  P
BALANCE OF TERROR [STAR TREK]  SCHNEIDER, PAUL  SR-S
BALANCE OF TRADE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  H
BALANCE POINT [@STAR WARS NEW JEDI ORDER]  TYERS, KATHY  P
BALDUR’S GATE [@FORGOTTEN REALMS BALDUR’S GATE I]  ATHANS, PHILIP  P
BALEFIRES  DRAKE, DAVID  H
BALLAD OF BETA-2, THE  DELANY, SAMUEL R.  DM121,C/P:2,P
BALLAD OF BETA-2, THE/EMPIRE STAR  DELANY, SAMUEL R.  C/P:2,P
BALLAD OF BILLY BADASS AND THE ROSE OF TURKESTAN, THE  SANDERS, WILLIAM  LP

BALLAD OF THE STARS ALTOV, GENRIKH/

ZHURAVLYOVA, VALENTINA

H

BALLADE DE BETA 2, LA  DELANY, SAMUEL R.  FP
BALLANTINE TEACHERS’ GUIDE TO SCIENCE FICTION, THE  ALLEN, L. DAVID  R/REF-P,C/REF-P
BALLROOM OF THE SKIES  MACDONALD, JOHN D.  C/P,P,H
BALLS!  ROHMER, RICHARD  P
BALOOK  ANTHONY, PIERS  P
BALOR OF THE EVIL EYE  CRENSHAW, NADINE  P

The MITSFS Titledex  September 24, 2007  Page 22
TITLE  AUTHOR  CODE
BALSHAZZAR’S SERPENT [THREE KINGS 1]  CHALKER, JACK L.  H
BALTICON 33 PROGRAM BOOK: DAVID WEBER GOH  BSFS  HA,VLPA
BALTIMORE CATECHISM 2  HORAN, ELLAMAY  [C/RAN-H]
BAMBOO DEMONS, THE [@CHILL #3]  SHERMAN, JORY  P
BANANA 9000 OPERATING SYSTEM (MACINTOSH)  BANANA COMPANY  SR-MM
BAND OF BROTHERS, A [LOST REGIMENT #7]  FORSTCHEN, WILLIAM R.  P
BANDER SNATCH  O’DONNELL, KEVIN JR.  P
BANDIT OF HELL’S BEND, THE  BURROUGHS, EDGAR RICE  P
BANE OF KANTHOS, THE  DAIN, ALEX  D42800
BANE OF LORD CALADON, THE  MILLS, CRAIG  P
BANE OF NIGHTMARES [@DREAM LORDS #3]  COLE, ADRIAN  P
BANE OF THE BLACK SWORD, THE [ELRIC 5]  MOORCOCK, MICHAEL  C/P:2,P
BANEWREAKER [@SUNDERING #1]  CAREY, JACQUELINE  H
BANGS AND WHIMPERS  FRENKEL, JAMES  HA
BANISHED, THE  WILLIAMSON, J. N.  P
BANNERS OF THE SA’YEN  STATEHAM, B. R.  P
BANQUETS OF THE BLACK WIDOWERS  ASIMOV, ISAAC  C/P:2,P
BANQUO LEGACY, THE [DOCTOR WHO EIGHTH DOCTOR]  LANE, ANDY/RICHARDS, JUSTIN  P
BANTAN — GOD-LIKE ISLANDER  GARDNER, MAURICE B.  H
BANTAN VALIANT  GARDNER, MAURICE B.  H
BAPHOMET’S METEOR  BARBET, PIERRE  C/P,P
BARBARA, THE VALIANT [@PHENWICK #16]  KIMBROUGH, KATHERYN  P
BARBARELLA  FOREST, JEAN-CLAUDE  CX
BARBARIAN, THE [@CASCA #5]  SADLER, BARRY  P
BARBARIAN OF WORLD’S END, THE [GONDWANE 4]  CARTER, LIN  P

BARBARIANS ADAMS, ROBERT/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

BARBARIANS OF MARS [MICHAEL KANE 3]  BRADBURY, EDWARD P.  SR-P,C/P:2
BARBARIANS OF MARS [MICHAEL KANE 3]  MOORCOCK, MICHAEL  P
BARBARY  MCINTYRE, VONDA N.  C/P,SFWA-P,H,P,SFWA
BARBED COIL, THE  JONES, J. V.  H:2,P
BARBED ROSE, THE [ONE ROSE 2]  DAYTON, GAIL  LP
BARBIE MURDERS, THE  VARLEY, JOHN  C/P:2,P
BARD  LLYWELYN, MORGAN  SFWA-P
BARD [@BARD #1]  TAYLOR, KEITH  C/P:2,P
BARD II [@BARD #2]  TAYLOR, KEITH  C/P,P
BARD OF SORCERY, THE  HOUARNER, GERARD DANIEL  P
BAREFOOT IN THE HEAD  ALDISS, BRIAN W.  C/P:2,P
BARGAIN, THE [DARK FORCES 5]  SPARGER, REX  C/P,P

BARLOWE’S GUIDE TO EXTRATERRESTRIALS BARLOWE, WAYNE DOUGLAS/

SUMMERS, IAN/MEACHAM, BETH

R/REF-H

BARLY FIELDS, THE  NATHAN, ROBERT  H
BARNABAS COLLINS  ROSS, MARILYN  C/P,P
BARNABAS COLLINS AND QUENTIN’S DEMON  ROSS, MARILYN  C/P,P
BARNABAS COLLINS AND THE GYPSY WITCH  ROSS, MARILYN  P
BARNABAS COLLINS AND THE MYSTERIOUS GHOST  ROSS, MARILYN  P
BARNABAS COLLINS IN A FUNNY VEIN  PAPERBACK LIBRARY  [C/RAN-P]
BARNABAS COLLINS VS. THE WARLOCK  ROSS, MARILYN  C/P:2,P
BARNABAS, QUENTIN AND DOCTOR JEKYLL’S SON  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE AVENGING GHOST  ROSS, MARILYN  C/P,P
BARNABAS, QUENTIN AND THE BODY SNATCHERS  ROSS, MARILYN  C/P,P
BARNABAS, QUENTIN AND THE CRYSTAL COFFIN  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE FRIGHTENED BRIDE  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE GRAVE ROBBERS  ROSS, MARILYN  P
BARNABAS, QUENTIN, AND THE HAUNTED CAVE  ROSS, MARILYN  C/P:2,P
BARNABAS, QUENTIN AND THE MAD MAGICIAN  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE MAGIC POTION  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE MUMMY’S CURSE  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE NIGHTMARE ASSASSIN  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE SCORPIO CURSE  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE SEA GHOST  ROSS, MARILYN  C/P:2,P
BARNABAS, QUENTIN AND THE SERPENT  ROSS, MARILYN  P
BARNABAS, QUENTIN AND THE WITCH’S CURSE  ROSS, MARILYN  P
BARNACLE BILL THE SPACER AND OTHER STORIES  SHEPARD, LUCIUS  LP
BARNARD’S PLANET  BOYD, JOHN  C/P,P,H
BARNSTORMER IN OZ, A  FARMER, PHILIP JOSE  LP
BARONESS OF BLOOD [RAVENLOFT]  BERGSTROM, ELAINE  P
BARONS OF BEHAVIOR, THE  PURDOM, TOM  C/P:2,P
BAROQUE FABLE, A  YARBRO, CHELSEA QUINN  P
BARRAYAR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 2]  BUJOLD, LOIS MCMASTER  C/P:2,SFWA-P,P,H
BARREL OF LAUGHS, A VALE OF TEARS, A  FEIFFER, JULES  LP
BARRENLANDS [DUN LADY’S JESS 4]  DURGIN, DORANNA  P
BARRENS AND OTHERS, THE  WILSON, F. PAUL  LP
BARRIER 346  ZEIGFREID, KARL  SR-P,C/P

The MITSFS Titledex  September 24, 2007  Page 23
TITLE  AUTHOR  CODE
BARRIER UNKNOWN  MERAK, A. J.  SR-P
BARRIER WORLD  CHARBONNEAU, LOUIS  C/P:2,P
BARROW  DEAKINS, JOHN  P
BARRY TROTTER AND THE SHAMELESS PARODY  GERBER, MICHAEL  H
BARSOOM: EDGAR RICE BURROUGHS AND THE MARTIAN VISION  LUPOFF, RICHARD A.  R/REF-H
BARSOOM PROJECT, THE [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  C/P,SFWA-P,P,H
BART SCIENCE FICTION TRIPLET #1  BART BOOKS  PA
BASE LUNAR  TUBB, E. C.  FP
BASES SUR VENUS (FRENCH) [@PERRY RHODAN #4]  SCHEER, K. H.  FPR
BASIC WRITINGS OF BERTRAND RUSSELL, THE  RUSSELL, BERTRAND  [C/RAN-H]
BASIL RATHBONE SELECTS STRANGE TALES  RATHBONE, BASIL  PA
BASILISK  KUSHNER, ELLEN  C/PA,PA
BASTAARD REBELLION, THE [@WARBOTS #3]  STINE, G. HARRY  C/P:2,P
BASTARD KING, THE [@SCEPTER OF MERCY #1]  CHERNENKO, DAN  LP
BASTARD PRINCE, THE [DERYNI HEIRS 3]  KURTZ, KATHERINE  H:2,P
BASTION OF DARKNESS [YNIS AIELLE 3]  SALVATORE, R. A.  P
BAT STAFFEL, THE [@G-8 #1]  HOGAN, ROBERT J.  SR-P
BATALHA DO VACUO, A  MAINE, CHARLES ERIC  FP
BATMAN [BATMAN 1]  GARDNER, CRAIG SHAW  C/P:2,P
BATMAN: DIGITAL JUSTICE  MORENO, PEPE  CX
BATMAN FOREVER  DAVID, PETER  P
BATMAN MURDERS, THE [BATMAN 2]  GARDNER, CRAIG SHAW  P
BATMAN RETURNS  GARDNER, CRAIG SHAW  C/P,P

BATMAN RETURNS O’NEIL, DENNIS/ERWIN, STEVE/ GARCIA-LOPEZ, JOSE LUIS

CX

BATMAN: THE LAST ANGEL  LUSTBADER, ERIC VAN  CX
BATMAN: TO STALK A SPECTER  HAWKE, SIMON  P
BATS IN THE BELFRY  MATSON, NORMAN  H
BATS OUT OF HELL  SMITH, GUY N.  P
*BATTERIES NOT INCLUDED  DREW, WEYLAND  P
BATTERIES NOT INCLUDED [NOT QUITE HUMAN 1]  MCEVOY, SETH  P
BATTLE AT ZERO POINT [@STARHAWK 4]  MALONEY, MACK  P
BATTLE CIRCLE [BATTLE CIRCLE 1,2,3]  ANTHONY, PIERS  C/P:2,P
BATTLE CRY [@ROBOTECH #2]  MCKINNEY, JACK  P
BATTLE FOR ATLANTIS [ATLANTIS 6]  DONEGAN, GREG  P
BATTLE FOR TERRA TWO, THE [BIOFAB 2]  BERRY, STEPHEN AMES  C/P,P
BATTLE FOR THE PLANET OF THE APES  GERROLD, DAVID  C/P:2,P
BATTLE FOR THE STARS  HAMILTON, EDMOND  P,H:2,SFWA-TD8,TD8
BATTLE HYMN [@LOST REGIMENT 5]  FORSTCHEN, WILLIAM R.  P
BATTLE LINES [@STAR TREK VOYAGER]  GALANTER, DAVE/BRODEUR, GREG  C/P:2,P
BATTLE MAGIC  GREENBERG, MARTIN H./SEGRIFF, LARRY  C/PA,PA
BATTLE OF BETAZED, THE [STAR TREK TNG]  DOUGLAS, CHARLOTTE/KEARNEY, SUSAN  C/P,P
BATTLE OF CORRIN, THE [DUNE JIHAD 3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H
BATTLE OF DISNEYLAND, THE  KEYES, THOM  H
BATTLE OF DORKING CONTROVERSY, THE  CORNMARKET REPRINTS  HA
BATTLE OF EVERNIGHT, THE [BITTERBYNDE 3]  DART-THORNTON, CECILIA  C/P,P
BATTLE OF EVERNIGHT, THE (PROOF) [BITTERBYNDE 3]  DART-THORNTON, CECILIA  SR-LP
BATTLE OF FOREVER, THE  VAN VOGT, A. E.  C/P:2,P
BATTLE OF THE RING [STARWOLVES 2]  GUNNARSSON, THORARINN  P,SFWA-P
BATTLE OF THE SEXES IN SCIENCE FICTION, THE  LARBALESTIER, JUSTINE  R/REF-H
BATTLE OF ZORN, THE  MCDANIEL, LURLENE  P
BATTLE ON! [XENA]  COX, GREG  R/REF-P
BATTLE ON MERCURY  VAN LHIN, ERIK  H:2
BATTLE ON VENUS  TEMPLE, WILLIAM F.  D76380,DF195
BATTLE STATION  BOVA, BEN  C/P,P
BATTLE SURGEONS [@STAR WARS CLONE WARS MEDSTAR 1]  REAVES, MICHAEL/PERRY, STEVE  P
BATTLEDRAGON [BAZIL BROKETAIL 4]  ROWLEY, CHRISTOPHER  P
BATTLEFIELD [DOCTOR WHO]  PLATT, MARC  C/P,P
BATTLEFIELD EARTH  HUBBARD, L. RON  H,P
BATTLEHYMN [@ROBOTECH #4]  MCKINNEY, JACK  P
BATTLESPACE [@LEGACY #2]  DOUGLAS, IAN  P
BATTLESTAR GALACTICA [BATTLESTAR GALACTICA 1978 #1]  LARSON, GLEN A./THURSTON, ROBERT  H:2,@P
BATTLESTATION [@BATTLESTATION #1]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
BATTLESTATIONS! [STAR TREK]  CAREY, DIANE  C/P:2,P
BAVARIAN GATE, THE [CURTIS MACURDY 2]  DALMAS, JOHN  P
BAY CITY BURNOUT [@OUTRIDER #4]  HARDING, RICHARD  P
BAZIL BROKETAIL [BAZIL BROKETAIL 1]  ROWLEY, CHRISTOPHER  C/P,SFWA-P,P
B.C. GREAT ZOT, I’M BEAUTIFUL  HART, JOHNNY  [C/RAN-P]
BE BURIED IN THE RAIN  MICHAELS, BARBARA  P
BEACHHEAD  WILLIAMSON, JACK  H
BEACHHEAD PLANET  WILLIAMS, ROBERT MOORE  C/P:2,P
BEACHHEADS IN SPACE  DERLETH, AUGUST  C/PA,PA
BEACON, THE  FREIREICH, VALERIE J.  C/P,P
BEAKER’S DOZEN  KRESS, NANCY  H,LP

TITLE  AUTHOR  CODE
BEAMRIDERS!  CAIDIN, MARTIN  C/P:2,SFWA-P,P
BEANSTALK  RACKHAM, JOHN  C/P:2,P
BEAR DAUGHTER  BERMAN, JUDITH  LP
BEARING AN HOURGLASS [INCARNATIONS OF IMMORTALITY 2]  ANTHONY, PIERS  C/P,P,H
BEARS DISCOVER FIRE  BISSON, TERRY  H
BEAR’S FANTASIES  BEAR, GREG  SR-H
BEARWALKER (PROOF)  BRUCHAC, JOSEPH  SR-LP
BEAST  BENCHLEY, PETER  P
B.E.A.S.T.  MAINE, CHARLES ERIC  C/P:2,P
BEAST, THE  STALLMAN, ROBERT  C/P,P
BEAST, THE  VAN VOGT, A. E.  C/P,P,H
BEAST MASTER, THE [HOSTEEM STORM 1]  NORTON, ANDRE  C/P,P,H,DD509
BEAST MASTER TEAM [HOSTEEM STORM 3,4]  NORTON, ANDRE/MCCONCHIE, LYN  H
BEAST MASTER’S ARK [HOSTEEM STORM 3]  NORTON, ANDRE/MCCONCHIE, LYN  P
BEAST MASTER’S CIRCUS [HOSTEEM STORM 4]  NORTON, ANDRE/MCCONCHIE, LYN  P
BEAST MASTER’S QUEST [HOSTEEM STORM 5]  NORTON, ANDRE/MCCONCHIE, LYN  H
BEAST OF HEAVEN, THE  KELLEHER, VICTOR  H
BEAST THAT SHOUTED LOVE AT THE HEART OF THE WORLD, THE  ELLISON, HARLAN  H:2,P
BEAST WITH FIVE FINGERS, THE  HARVEY, WILLIAM FRYER  H
BEAST WITH THE RED HANDS, THE  STUART, SIDNEY  P
BEAST WITHIN, THE  LEVY, EDWARD  P
BEAST WITHIN, THE [WORLD OF DARKNESS VAMPIRE]  WIECK, STEWART  PA
BEASTCHILD  KOONTZ, DEAN R.  P
BEASTMAKER  SMITH, JAMES V. JR.  P
BEASTMARKS  ATTANASIO, A. A.  H
BEASTNIGHTS  YARBRO, CHELSEA QUINN  P
BEASTS  CROWLEY, JOHN  H:2,SFWA,P
BEASTS BELOW, THE [PERRY RHODAN 128]  MAHR, KURT  SR-LP
BEASTS FROM BEYOND, THE  WELLMAN, MANLY WADE  SR-P
BEASTS IN VELVET [WARHAMMER GENEVIEVE 3]  YEOVIL, JACK  P
BEASTS OF ANTARES [DRAY PRESCOT 23]  AKERS, ALAN BURT  P
BEASTS OF BARAKHAI [@BARAKHAI #1]  REICHERT, MICKEY ZUCKER  H,P
BEASTS OF GOR [GOR 12]  NORMAN, JOHN  C/P:2,P
BEASTS OF HADES, THE [EMPIRE PRINCESS 4]  DIAMOND, GRAHAM  P
BEASTS OF KOHL, THE  RACKHAM, JOHN  DG592
BEASTS OF TARZAN, THE [TARZAN #3]  BURROUGHS, EDGAR RICE  C/P,@P,H
BEASTS OF THE MIST, THE [SWORDS OF RAEMLLYN 5]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P:2,P
BEASTS OF VALHALLA, THE [MONGO FREDERICKSON 4]  CHESBRO, GEORGE C.  H,P
BEASTSLAYER [WARHAMMER GOTREK & FELIX 5]  KING, WILLIAM  P
BEAUTY  MCKINLEY, ROBIN  C/P,P,H
BEAUTY  TEPPER, SHERI S.  C/P,P,LP,H
BEAUTY AND THE BEAST [BEAUTY AND THE BEAST]  HAMBLY, BARBARA  C/P,P
BEAUTY’S PUNISHMENT [BEAUTY 2]  ROQUELAURE, A. N.  SR-L
BEAUTY’S RELEASE [BEAUTY 3]  ROQUELAURE, A. N.  H,SR-L
BECKONING, THE  WEBSTER, JOSH  P
BECOME THE HUNTED  SWYCAFFER, JEFFERSON P.  P
BECOMING, THE [ANNA STRONG 1]  STEIN, JEANNE C.  P
BECOMING ALIEN [BECOMING ALIEN 1]  ORE, REBECCA  H:2,P
BECOMING DEATH [@ENEMIES OF SOCIETY #1]  DAVID, JOHN  LP
BECOMING HUMAN [HARMONY OF WORLDS]  FREIREICH, VALERIE J.  P
BECOMING HUMAN (PROOF)  THOMPSETT, NEIL LEE  SR-P
BEDFORD INCIDENT, THE  RASCOVICH, MARK  P
BEDLAM BOYZ [ERIC BANYON]  GUON, ELLEN  C/P:2,SFWA-P,P
BEDLAM PLANET  BRUNNER, JOHN  C/P:2,P
BEDLAM’S BARD [ERIC BANYON 1,2]  LACKEY, MERCEDES/GUON, ELLEN  H,P
BEDLAM’S EDGE  LACKEY, MERCEDES/EDGHILL, ROSEMARY  HA
BEELZEBUB BUSINESS, THE [@BIG BRAIN #2]  BRANDNER, GARY  P
BEETLE IN THE ANTHILL  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H
BEFORE... 12:01 ...AND AFTER  LUPOFF, RICHARD  H
BEFORE ADAM  LONDON, JACK  H:2,P
BEFORE ARMAGEDDON, VOLUME 1  MOORCOCK, MICHAEL  PA
BEFORE THE DAWN [@DARK ANGEL]  COLLINS, MAX ALLAN  C/P,P
BEFORE THE DAWN  TAINE, JOHN  SR-H
BEFORE THE GOLDEN AGE  ASIMOV, ISAAC  SFWA,HA
BEFORE THE GOLDEN AGE 1  ASIMOV, ISAAC  PA,C/PA
BEFORE THE GOLDEN AGE 2  ASIMOV, ISAAC  PA,C/PA
BEFORE THE GOLDEN AGE 3  ASIMOV, ISAAC  PA,C/PA
BEFORE THE INVID STORM [@ROBOTECH #21]  MCKINNEY, JACK  C/P,P
BEFORE THE MASK [@DRAGONLANCE VILLIANS #1]  WILLIAMS, MICHAEL/WILLIAMS, TERI  C/P,P
BEFORE THE STORM [@STAR WARS BLACK FLEET CRISIS 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
BEFORE THE UNIVERSE  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
BEGGAR QUEEN, THE [WESTMARK 3]  ALEXANDER, LLOYD  P
BEGGARS AND CHOOSERS [BEGGARS 2]  KRESS, NANCY  H:2,P
BEGGAR’S BANQUET [FANUILH 3]  HOOD, DANIEL  P

The MITSFS Titledex  September 24, 2007  Page 25
TITLE  AUTHOR  CODE
BEGGARS IN SPAIN [BEGGARS 1]  KRESS, NANCY  H:2,P
BEGGARS RIDE [BEGGARS 3]  KRESS, NANCY  C/P:2,P
BEGINNING, THE [@DARK CHRONICLES #1]  SOROKA, CYNTHIA  P
BEGINNING AND THE END, THE  ASIMOV, ISAAC  SFWA,R/REF-P
BEGINNING AT DUSK: AN INTERLUDE  WILLIAMSON, THAMES  H
BEGINNING OF UNBELIEF, THE  JONES, ROBIN D.  SFWA
BEGINNING PLACE, THE  LE GUIN, URSULA K.  C/P,SFWA,P
BEGINNINGS [STAR TREK TNG]  CARLIN, MIKE  CX
BEGINNINGS  DICKSON, GORDON R.  C/P:2,SFWA-P,P
BEGINNINGS OF FOREVER, THE  SIROIS, A. L.  LP
BEGUILEMENT [@SHARING KNIFE 1]  BUJOLD, LOIS MCMASTER  H
BEGUM’S FORTUNE, THE  VERNE, JULES  H
BEHEMOTH: B-MAX [RIFTERS 3.0]  WATTS, PETER  H
BEHEMOTH: SEPPUKU [RIFTERS 3.1]  WATTS, PETER  H
BEHIND ENEMY LINES [@STAR TREK DOMINION WAR #1/STAR TREK TNG]  VORNHOLT, JOHN  P
BEHIND THE ATTIC WALL  CASSEDY, SYLVIA  LP
BEHIND THE FLYING SAUCERS  SCULLY, FRANK  [C/RAN-H]
BEHIND THE WALLS OF TERRA [WORLD OF TIERS 4]  FARMER, PHILIP JOSE  C/P:2,P
BEHIND TIME [EMMA MERRIGAN 2]  ABBEY, LYNN  C/P,P
BEHOLD THE MAN [KARL GLOGAUER 1]  MOORCOCK, MICHAEL  C/P:2,P
BEHOLD THE STARS  BULMER, KENNETH  DM131
BEHOLDER’S EYE [WEB SHIFTERS 1]  CZERNEDA, JULIE E.  C/P,P
BEIGE PLANET MARS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  PARKIN, LANCE/CLAPHAM, MARK  C/P,P
BEING  REDFINN, MICHAEL  P
BEING A GREEN MOTHER [INCARNATIONS OF IMMORTALITY 5]  ANTHONY, PIERS  C/P,P,H
BEING ALIEN [BECOMING ALIEN 2]  ORE, REBECCA  H:2,SFWA-P
BEING GARDENER DOZOIS  SWANWICK, MICHAEL  R/REF-H
BEING OF TWO MINDS  SERVICE, PAMELA  SFWA
BEING THERE  KOSINSKI, JERZY  C/P,P
BELARUS  HOGAN, LEE  P
BELDAN’S FIRE [QUEEN’S QUARTER 3]  SNYDER, MIDORI  C/P:2,P
BELGARATH THE SORCERER [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  H:2,P
BELGARIAD 1, THE [BELGARIAD 1,2,3]  EDDINGS, DAVID  H:2
BELGARIAD 2, THE [BELGARIAD 4,5]  EDDINGS, DAVID  H:2
BELIEVE.  SHATNER, WILLIAM/TOBIAS, MICHAEL  C/P,SFWA-P:2,P
BELIEVERS’ WORLD  LOWNDES, ROBERT A. W.  H
BELL FROM INFINITY, THE  WILLIAMS, ROBERT MOORE  C/P:2,P
BELL OF NENDRUM, THE  ANDREWS, J. S.  P
BELLADONNA [EPHEMERA]  BISHOP, ANNE  H
BELLE TERRE [@STAR TREK NEW EARTH #2]  SMITH, DEAN WESLEY/CAREY, DIANE  C/P,P
BELLMAKER, THE [REDWALL]  JACQUES, BRIAN  H,P
BELLWETHER  WILLIS, CONNIE  C/P,P
BELLY OF THE BOW, THE [FENCER 2]  PARKER, K. J.  LP
BELLY OF THE WOLF, THE [@LENS OF THE WORLD #3]  MACAVOY, R. A.  C/P,P
BELOVED EXILE  GODWIN, PARKE  H,P
BELOVED OF THE DEAD [@EXALTED SECOND AGE #2]  DANSKY, RICHARD E.  P
BELOVED SON [ETHICAL SOCIETY 1]  TURNER, GEORGE  C/P,P
BELOW THE ROOT [GREEN-SKY 1]  SNYDER, ZILPHA KEATLEY  P,SFWA-P
BELOW THE THRESHOLD [V]  WOLD, ALLEN L.  C/P,SFWA-P,P
BELTEMPEST [DOCTOR WHO EIGHTH DOCTOR]  MORTIMORE, JIM  P
BELUTHAHATCHIE AND OTHER STORIES  DUNCAN, ANDY  H
BEN  RALSTON, GILBERT A.  P
BEN HARDY’S FLYING MACHINE  WEBSTER, FRANK V.  H
BEN THERE, DONE THAT [@SABRINA THE TEENAGE WITCH #6]  LOCKE, JOSEPH  P
BENDING THE LANDSCAPE: FANTASY  GRIFFITH, NICOLA/PAGEL, STEPHEN  LPA
BENDING THE LANDSCAPE: SCIENCE FICTION  GRIFFITH, NICOLA/PAGEL, STEPHEN  HA,LPA
BENEATH A SAPPHIRE SEA  BRYAN, JESSICA  P
BENEATH AN OPAL MOON [SUNSET WARRIOR 4]  LUSTBADER, ERIC VAN  H,SFWA,P
BENEATH THE GATED SKY  REED, ROBERT  H
BENEATH THE MOORS  LUMLEY, BRIAN  SR-H
BENEATH THE PLANET OF THE APES  AVALLONE, MICHAEL  C/P:2,P
BENEATH THE SHATTERED MOONS  BISHOP, MICHAEL  P
BENEATH THE THIRTEEN MOONS  KENNEDY, KATHRYNE  H
BENEATH THE TREE OF HEAVEN [@CHUNG KUO #5]  WINGROVE, DAVID  C/P:2,LP
BENEATH THE VAULTED HILLS [RIVER INTO DARKNESS 1]  RUSSELL, SEAN  C/P,P
BENEATH THE WEB [WOODEN SWORD 2]  ABBEY, LYNN  P
BENEFITS  FAIRBAIRNS, ZOE  P
BENIGHTED  WHITFIELD, KIT  LP

BEOWULF’S CHILDREN [HEOROT 2] NIVEN, LARRY/POURNELLE, JERRY/

BARNES, STEVEN

H:2,P

BERBORA [@BERBORA #1]  SIROTA, MIKE  C/P,P
BERKLEY SHOWCASE 1, THE  SCHOCHET, VICTORIA/SILBERSACK, JOHN  PA
BERKLEY SHOWCASE 2, THE  SCHOCHET, VICTORIA/SILBERSACK, JOHN  PA
BERKLEY SHOWCASE 3, THE  SCHOCHET, VICTORIA/SILBERSACK, JOHN  C/PA,PA

The MITSFS Titledex  September 24, 2007  Page 26
TITLE  AUTHOR  CODE
BERKLEY SHOWCASE 4, THE  SCHOCHET, VICTORIA/SILBERSACK, JOHN  PA
BERKLEY SHOWCASE 5, THE  SCHOCHET, VICTORIA/SINGER, MELISSA  C/PA,PA
BERNHARD THE CONQUEROR  LUNDWALL, SAM J.  C/P:2,P
BERSERKER [@ALIENS 9]  PERRY, S. D.  P
BERSERKER [BERSERKER 1]  SABERHAGEN, FRED  C/P:2,P
BERSERKER ATTACK, THE  SABERHAGEN, FRED  C/P:2,P

BERSERKER BASE [BERSERKER 2] SABERHAGEN, FRED SFWA,SFWA-P,HA:2,

LPA

BERSERKER: BLUE DEATH [BERSERKER 3]  SABERHAGEN, FRED  C/P:2,SFWA,SFWA-P,LP
BERSERKER FURY [BERSERKER 7]  SABERHAGEN, FRED  H,P
BERSERKER KILL [BERSERKER 6]  SABERHAGEN, FRED  H:2
BERSERKER LIES [BERSERKER]  SABERHAGEN, FRED  C/P,P
BERSERKER MAN [BERSERKER]  SABERHAGEN, FRED  C/P:2,P
BERSERKER PRIME [BERSERKER]  SABERHAGEN, FRED  P
BERSERKER THRONE, THE [BERSERKER 4]  SABERHAGEN, FRED  H:2,SFWA-P,LP
BERSERKER WARS, THE [BERSERKER 1]  SABERHAGEN, FRED  C/P,SFWA-P,P
BERSERKERS, THE  ELWOOD, ROGER  SR-HA
BERSERKER’S HORSE, THE  MAXWELL, LISA  P
BERSERKER’S PLANET [BERSERKER 5]  SABERHAGEN, FRED  C/P:2,P
BERSERKER’S STAR [BERSERKER]  SABERHAGEN, FRED  P
BERSERKERS: THE BEGINNING [BERSERKER]  SABERHAGEN, FRED  C/P,P
BERSERKERS: THE ULTIMATE ENEMY [BERSERKER 4]  SABERHAGEN, FRED  C/P,P

BERTREM’S GUIDE TO THE WAR OF SOULS, VOL. ONE [@DRAGONLANCE WAR OF SOULS] CROOK, JEFF/GRUBBER, JOHN/

HERBERT, MARY H./

BERBERICK, NANCY VARIAN

P

BEST ALTERNATE HISTORY STORIES OF THE 20TH CENTURY, THE TURTLEDOVE, HARRY/

GREENBERG, MARTIN H.

LPA

BEST AND THE BRIGHTEST, THE [STAR TREK]  WRIGHT, SUSAN  P
BEST AUSTRALIAN SCIENCE FICTION WRITING, THE  GERRAND, ROB  LPA
BEST DESTINY [STAR TREK]  CAREY, DIANE  C/P,P,H
BEST FROM AMAZING STORIES, THE  WHITE, TED  PA,C/PA
BEST FROM FANTASTIC, THE  WHITE, TED  C/PA,PA
BEST FROM FANTASY & SCIENCE FICTION, THE  FERMAN, EDWARD L./VAN GELDER, GORDON HA
BEST FROM FANTASY AND SCIENCE FICTION 3, THE  BOUCHER, ANTHONY/MCCOMAS, J. FRANCIS C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 4, THE  BOUCHER, ANTHONY  HA,PA
BEST FROM FANTASY AND SCIENCE FICTION 5, THE  BOUCHER, ANTHONY  HA,PA
BEST FROM FANTASY AND SCIENCE FICTION 6, THE  BOUCHER, ANTHONY  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 7, THE  BOUCHER, ANTHONY  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 8, THE  BOUCHER, ANTHONY  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 9, THE  MILLS, ROBERT P.  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 10, THE  MILLS, ROBERT P.  PA,C/PA
BEST FROM FANTASY AND SCIENCE FICTION 11, THE  MILLS, ROBERT P.  PA,C/PA
BEST FROM FANTASY AND SCIENCE FICTION 12, THE  DAVIDSON, AVRAM  HA,PA
BEST FROM FANTASY AND SCIENCE FICTION 13, THE  DAVIDSON, AVRAM  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 14, THE  DAVIDSON, AVRAM  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 15, THE  FERMAN, EDWARD L.  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 16, THE  FERMAN, EDWARD L.  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 17, THE  FERMAN, EDWARD L.  C/PA,PA
BEST FROM FANTASY AND SCIENCE FICTION 18, THE  FERMAN, EDWARD L.  HA
BEST FROM FANTASY AND SCIENCE FICTION 19, THE  FERMAN, EDWARD L.  C/PA:2,PA,SFWA
BEST FROM FANTASY AND SCIENCE FICTION 20, THE  FERMAN, EDWARD L.  C/PA,PA,SFWA
BEST FROM FANTASY AND SCIENCE FICTION 22, THE  FERMAN, EDWARD L.  C/PA,PA,SFWA
BEST FROM FANTASY AND SCIENCE FICTION 23, THE  FERMAN, EDWARD L.  PA,SFWA
BEST FROM FANTASY AND SCIENCE FICTION 24, THE  FERMAN, EDWARD L.  HA,PA,SFWA

BEST FROM FANTASY AND SCIENCE FICTION:

A SPECIAL 25TH ANNIVERSARY ANTHOLOGY, THE FERMAN, EDWARD L.

HA,PA,SFWA

BEST FROM GALAXY 1, THE  GALAXY  C/PA,PA
BEST FROM GALAXY 2, THE  GALAXY  C/PA,PA
BEST FROM GALAXY III, THE  BAEN, JAMES  C/PA:2,PA
BEST FROM GALAXY IV, THE  BAEN, JAMES  PA,C/PA
BEST FROM IF 1, THE  IF EDITORS  PA
BEST FROM IF 2, THE  IF EDITORS  C/PA,PA
BEST FROM IF III, THE  BAEN, JAMES  C/PA,PA
BEST FROM ORBIT, THE  KNIGHT, DAMON  C/PA,PA
BEST FROM STARTLING STORIES, THE  MINES, SAMUEL  HA
BEST FROM THE REST OF THE WORLD, THE  WOLLHEIM, DONALD A.  HA,PA,SFWA
BEST FROM UNIVERSE, THE  CARR, TERRY  SFWA
BEST GHOST STORIES  LE FANU, J. SHERIDAN  LPA
BEST GHOST STORIES OF ALGERNON BLACKWOOD  BLACKWOOD, ALGERNON  LP
BEST GHOST STORIES OF M. R. JAMES  JAMES, M. R.  H
BEST HORROR FROM FANTASY TALES, THE  JONES, STEPHEN/SUTTON, DAVID  LPA

BEST HORROR STORIES FORM THE MAGAZINE OF FANTASY AND SCIENCE FICTION 1, THE FERMAN, EDWARD L./JORDAN, ANNE

PA

BEST-IN-BOOKS (CONTAINS "DANDELION WINE" BY RAY BRADBURY)  DOUBLEDAY BOOKS  HA
BEST-IN-BOOKS (CONTAINS "THE FAR TRAVELLER" BY MANNING COLES)  DOUBLEDAY BOOKS  HA

The MITSFS Titledex  September 24, 2007  Page 27
TITLE  AUTHOR  CODE
BEST KNOWN WORKS OF VOLTAIRE, THE  VOLTAIRE  H
BEST LAID SCHEMES, THE  EISENBERG, LARRY  H,P

BEST MILITARY SCIENCE FICTION OF THE 20TH CENTURY, THE TURTLEDOVE, HARRY/

GREENBERG, MARTIN H.

HA,LPA

BEST MYSTERIES OF ISAAC ASIMOV, THE  ASIMOV, ISAAC  C/P:2,SFWA,P
BEST NEW FANTASY  WALLACE, SEAN  LPA
BEST NEW HORROR 3  JONES, STEPHEN/CAMPBELL, RAMSEY  LPA
BEST OF A. E. VAN VOGT, THE  VAN VOGT, A. E.  C/P,P
BEST OF ALL POSSIBLE WARS [@MAN-KZIN WARS]  NIVEN, LARRY  PA
BEST OF ALL POSSIBLE WORLDS, THE  DE ROSA, PETER  SR-P
BEST OF ALL POSSIBLE WORLDS, THE  ROBINSON, SPIDER  C/PA,PA
BEST OF AMAZING, THE  ROSS, JOSEPH  PA
BEST OF ANALOG, THE  BOVA, BEN  C/PA,LPA
BEST OF ARTHUR C. CLARKE: 1937–1955, THE  CLARKE, ARTHUR C.  SR-P,C/P
BEST OF ASTOUNDING, THE  LEWIS, TONY  LPA
BEST OF AVRAM DAVIDSON, THE  DAVIDSON, AVRAM  H,SFWA
BEST OF BARRY N. MALZBERG, THE  MALZBERG, BARRY N.  C/P:2,P
BEST OF BEAUMONT  BEAUMONT, CHARLES  P
BEST OF C. M. KORNBLUTH, THE  KORNBLUTH, C. M.  H:2,P
BEST OF CEMETARY DANCE, THE, VOLUME ONE  CHIZMAR, RICHARD T.  LPA
BEST OF CEMETARY DANCE, THE, VOLUME TWO  CHIZMAR, RICHARD T.  LPA
BEST OF C.L. MOORE, THE  MOORE, C. L.  H:2,P
BEST OF CLIFFORD D. SIMAK, THE  SIMAK, CLIFFORD D.  SR-P
BEST OF CORDWAINER SMITH, THE  SMITH, CORDWAINER  H:2,P
BEST OF CRANK!, THE  CHOLFIN, BRYAN  HA,LPA
BEST OF CREEPY, THE  WARREN, JAMES  SCX
BEST OF DAMON KNIGHT, THE  KNIGHT, DAMON  C/P,P,H:2
BEST OF DESTINIES 1980, THE  BAEN, JAMES  PA
BEST OF DREAMS OF DECADENCE, THE  KESSLER, ANGELA  C/PA,PA
BEST OF E. E. "DOC" SMITH, THE  SMITH, E. E.  SR-P,C/P
BEST OF EDMOND HAMILTON, THE  HAMILTON, EDMOND  C/P,P,H:2
BEST OF ERIC FRANK RUSSELL, THE  RUSSELL, ERIC FRANK  C/P:2,P
BEST OF FREDERIK POHL, THE  POHL, FREDERIK  H:2,P
BEST OF FREDRIC BROWN, THE  BROWN, FREDRIC  H:2,P
BEST OF FRITZ LEIBER, THE  LEIBER, FRITZ  H:2,P

BEST OF H. P. LOVECRAFT:

BLOODCURDLING TALES OF HORROR AND THE MACABRE, THE LOVECRAFT, H. P.

LP

BEST OF HAL CLEMENT, THE  CLEMENT, HAL  C/P:2,P
BEST OF HARRY HARRISON, THE  HARRISON, HARRY  C/P,P
BEST OF HENRY KUTTNER, THE  KUTTNER, HENRY  H:2,P
BEST OF INTERZONE, THE  PRINGLE, DAVID  HA
BEST OF ISAAC ASIMOV, THE  ASIMOV, ISAAC  C/P,P,H
BEST OF JACK VANCE, THE  VANCE, JACK  C/P:2,P
BEST OF JACK WILLIAMSON, THE  WILLIAMSON, JACK  C/P:2,P
BEST OF JAMES BLISH, THE  BLISH, JAMES  C/P,P
BEST OF JAMES H. SCHMITZ, THE  SCHMITZ, JAMES H.  SR-H,H
BEST OF JIM BAEN’S UNIVERSE, THE (2006)  FLINT, ERIC  HA
BEST OF JOHN BRUNNER, THE  BRUNNER, JOHN  P
BEST OF JOHN JAKES, THE  JAKES, JOHN  C/P,P
BEST OF JOHN SLADEK, THE  SLADEK, JOHN  P
BEST OF JOHN W. CAMPBELL, THE  CAMPBELL, JOHN W. JR.  C/P:2,P,H
BEST OF JUDITH MERRIL, THE  MERRIL, JUDITH  C/P:2,P
BEST OF JULES VERNE  VERNE, JULES  H
BEST OF KEITH LAUMER, THE  LAUMER, KEITH  C/P:2,P
BEST OF L. SPRAGUE DE CAMP, THE  DE CAMP, L. SPRAGUE  H:2,P
BEST OF LEIGH BRACKETT, THE  BRACKETT, LEIGH  H:2,P
BEST OF LESTER DEL REY, THE  DEL REY, LESTER  C/P,P,H
BEST OF MACK REYNOLDS, THE  REYNOLDS, MACK  C/P:2,P
BEST OF MARION ZIMMER BRADLEY, THE  BRADLEY, MARION ZIMMER  C/P,P
BEST OF MARION ZIMMER BRADLEY’S FANTASY MAGAZINE, THE  BRADLEY, MARION ZIMMER  PA

BEST OF MARION ZIMMER BRADLEY’S FANTASY MAGAZINE, THE, VOL. II BRADLEY, MARION ZIMMER/ WATERS, ELISABETH

C/PA,PA

BEST OF MURRAY LEINSTER, THE  LEINSTER, MURRAY  C/P:2,P
BEST OF NEW DIMENSIONS, THE  SILVERBERG, ROBERT  PA
BEST OF OMNI SCIENCE FICTION, THE  BOVA, BEN/MYRUS, DON  VLPA
BEST OF OMNI SCIENCE FICTION NO. 4, THE  BOVA, BEN/MYRUS, DON  VLPA
BEST OF PAMELA SARGENT, THE  SARGENT, PAMELA  P
BEST OF PHILIP K. DICK, THE  DICK, PHILIP K.  P
BEST OF PIRATE WRITINGS 1, THE  MCFADDEN, EDWARD J. III  LPA
BEST OF PLANET STORIES #1, THE  BRACKETT, LEIGH  PA,C/PA
BEST OF POUL ANDERSON, THE  ANDERSON, POUL  C/P,P
BEST OF PULPHOUSE, THE  RUSCH, KRISTINE KATHRYN  HA
BEST OF RANDALL GARRETT, THE  GARRETT, RANDALL  C/P:2,P
BEST OF RAYMOND Z. GALLUN, THE  GALLUN, RAYMOND Z.  C/P,P
BEST OF ROBERT BLOCH, THE  BLOCH, ROBERT  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 28
TITLE  AUTHOR  CODE
BEST OF ROBERT SILVERBERG, THE  SILVERBERG, ROBERT  C/P:3,P
BEST OF SAKI, THE  SAKI  LP
BEST OF SCIENCE FICTION, THE (1946 EDITION; COMPLETE)  CONKLIN, GROFF  HA
BEST OF SHADOWS, THE  GRANT, CHARLES L.  HA
BEST OF STANLEY G. WEINBAUM, THE  WEINBAUM, STANLEY G.  C/P:2,P
BEST OF STAR TREK, THE NEXT GENERATION, THE  FRIEDMAN, MICHAEL JAN/DELANCIE, JOHN CX
BEST OF SYDNEY J. BOUNDS, VOLUME 1, THE  BOUNDS, SYDNEY J.  LP
BEST OF SYDNEY J. BOUNDS, VOLUME 2, THE  BOUNDS, SYDNEY J.  LP
BEST OF TALES, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
BEST OF TALES II, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  C/PA,PA
BEST OF THE BEST, THE  DOZOIS, GARDNER  LPA
BEST OF THE BEST 2, THE  DOZOIS, GARDNER  LPA
BEST OF THE BEST OF TREK, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF THE BEST OF TREK II, THE  IRWIN, WALTER/LOVE, G. B.  SFWA,R/REF-P
BEST OF THE BUSHEL, THE  SHAW, BOB  R/REF-P
BEST OF THE HORROR SHOW  SILVA, DAVID B.  LPA
BEST OF THE REALMS, BOOK II, THE [@FORGOTTEN REALMS]  GREENWOOD, ED  P
BEST OF THE REALMS BOOK III, THE [@FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
BEST OF THE REALMS BOOK I, THE [@FORGOTTEN REALMS]  WIZARDS OF THE COAST  PA
BEST OF THE REST 2  YOUMANS, BRIAN  LPA
BEST OF THE REST 3  YOUMANS, BRIAN  LPA
BEST OF THOMAS N. SCORTIA, THE  SCORTIA, THOMAS N.  H,SFWA
BEST OF TREK, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #2, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #3, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #4, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #5, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #6, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #7, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #8, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #9, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #10, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #11, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #12, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #13, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #15, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF TREK #16, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P
BEST OF TREK #18, THE  IRWIN, WALTER/LOVE, G. B.  R/REF-P,C/REF-P
BEST OF WALTER M. MILLER JR., THE  MILLER, WALTER M. JR.  C/P,P

BEST OF WEIRD TALES: 1923, THE KAYE, MARVIN/

BETANCOURT, JOHN GREGORY

LPA

BEST OF WILSON TUCKER, THE  TUCKER, WILSON  C/P,P

BEST ROOTIN’ TOOTIN’ SHOOTIN’ GUNSLINGER IN THE WHOLE DAMNED GALAXY, THE [GALACTIC MIDWAY 4] RESNICK, MIKE

C/P:2,P

BEST SCIENCE FICTION AND FANTASY OF THE YEAR VOLUME 1, THE  STRAHAN, JONATHAN  LPA
BEST SCIENCE FICTION FOR 1972  POHL, FREDERIK  PA,C/PA
BEST SCIENCE FICTION FOR 1973  ACKERMAN, FORREST J.  PA
BEST SCIENCE FICTION FROM IF  POHL, FREDERIK  SR-PA,C/PA
BEST SCIENCE FICTION FROM WORLDS OF TOMORROW  POHL, FREDERIK  SR-PA,C/PA
BEST SCIENCE FICTION NOVELLAS OF THE YEAR #1, THE  CARR, TERRY  C/PA,PA
BEST SCIENCE FICTION NOVELLAS OF THE YEAR #2, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF E. C. TUBB, THE  TUBB, E. C.  H
BEST SCIENCE FICTION OF ISAAC ASIMOV, THE  ASIMOV, ISAAC  SFWA
BEST SCIENCE FICTION OF J. G. BALLARD  BALLARD, J. G.  P
BEST SCIENCE FICTION OF THE YEAR, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF THE YEAR #2, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF THE YEAR #3, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF THE YEAR #4, THE  CARR, TERRY  C/PA:2,PA
BEST SCIENCE FICTION OF THE YEAR #5, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF THE YEAR #6, THE  CARR, TERRY  PA,C/PA
BEST SCIENCE FICTION OF THE YEAR #7, THE  CARR, TERRY  PA
BEST SCIENCE FICTION OF THE YEAR #8, THE  CARR, TERRY  HA,PA
BEST SCIENCE FICTION OF THE YEAR #9, THE  CARR, TERRY  PA
BEST SCIENCE FICTION OF THE YEAR #10, THE  CARR, TERRY  C/PA,PA
BEST SCIENCE FICTION OF THE YEAR #11, THE  CARR, TERRY  C/PA,PA
BEST SCIENCE FICTION OF THE YEAR #12, THE  CARR, TERRY  C/PA,PA
BEST SCIENCE FICTION OF THE YEAR #13, THE  CARR, TERRY  C/PA,PA
BEST SCIENCE FICTION STORIES, 1949, THE  BLEILER, EVERETT F./DIKTY, T. E.  HA
BEST SCIENCE FICTION STORIES, 1951, THE  BLEILER, EVERETT F./DIKTY, T. E.  HA
BEST SCIENCE FICTION STORIES, 1952, THE  BLEILER, EVERETT F./DIKTY, T. E.  HA
BEST SCIENCE FICTION STORIES AND NOVELS 1956, THE  DIKTY, T. E.  HA
BEST SCIENCE FICTION STORIES AND NOVELS, NINTH SERIES, THE  DIKTY, T. E.  HA
BEST SCIENCE FICTION STORIES OF CLIFFORD D. SIMAK  SIMAK, CLIFFORD D.  H:2,P
BEST SCIENCE FICTION STORIES OF H. G. WELLS  WELLS, H. G.  H,LP
BEST SCIENCE FICTION STORIES OF THE YEAR 1  DEL REY, LESTER  HA,PA

The MITSFS Titledex  September 24, 2007  Page 29
TITLE  AUTHOR  CODE
BEST SCIENCE FICTION STORIES OF THE YEAR 2  DEL REY, LESTER  HA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 3  DEL REY, LESTER  HA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 4  DEL REY, LESTER  C/PA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 5  DEL REY, LESTER  C/PA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 6  DOZOIS, GARDNER  C/PA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 7  DOZOIS, GARDNER  HA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 8  DOZOIS, GARDNER  HA,PA
BEST SCIENCE FICTION STORIES OF THE YEAR 9  DOZOIS, GARDNER  HA
BEST SF: 1967  HARRISON, HARRY  PA,C/PA
BEST SF: 1968  HARRISON, HARRY/ALDISS, BRIAN W.  C/PA,PA
BEST SF: 1969  HARRISON, HARRY/ALDISS, BRIAN W.  C/PA,PA
BEST SF: 1970  HARRISON, HARRY/ALDISS, BRIAN W.  HA
BEST SF: 1971  HARRISON, HARRY/ALDISS, BRIAN W.  PA
BEST SF: 1972  HARRISON, HARRY/ALDISS, BRIAN W.  C/PA,PA
BEST SF: 1973  HARRISON, HARRY/ALDISS, BRIAN W.  C/PA,PA
BEST SF: 1974  HARRISON, HARRY/ALDISS, BRIAN W.  HA
BEST SF: 1975  HARRISON, HARRY/ALDISS, BRIAN W.  HA
BEST SF STORIES FROM NEW WORLDS, THE  MOORCOCK, MICHAEL  PA,C/PA
BEST SF STORIES FROM NEW WORLDS 2, THE  MOORCOCK, MICHAEL  PA,C/PA
BEST SF STORIES FROM NEW WORLDS 3, THE  MOORCOCK, MICHAEL  C/PA,PA
BEST SF STORIES FROM NEW WORLDS 4, THE  MOORCOCK, MICHAEL  PA,C/PA
BEST SF STORIES FROM NEW WORLDS 5, THE  MOORCOCK, MICHAEL  C/PA,PA
BEST SF STORIES FROM NEW WORLDS 6, THE  MOORCOCK, MICHAEL  PA,C/PA
BEST SF STORIES FROM NEW WORLDS 8, THE  MOORCOCK, MICHAEL  SR-PA
BEST SF THREE  CRISPIN, EDMUND  SR-LPA
BEST SHORT NOVELS 2004  STRAHAN, JONATHAN  HA
BEST SHORT NOVELS 2005  STRAHAN, JONATHAN  HA
BEST SHORT NOVELS 2006  STRAHAN, JONATHAN  HA
BEST SHORT STORIES  BERGER, ERIC  C/PA,PA
BEST SHORT STORIES OF J. G. BALLARD, THE  BALLARD, J. G.  H,P
BEST STORIES OF H. G. WELLS  WELLS, H. G.  C/P:2,P
BEST STORIES OF WALTER DE LA MARE  DE LA MARE, WALTER J.  SR-H
BEST SUPERNATURAL TALES OF ALGERNON BLACKWOOD, THE  BLACKWOOD, ALGERNON  H
BEST TALES OF HOFFMANN, THE  HOFFMANN, E. T. A.  LP
BEST "THINKING MACHINE" DETECTIVE STORIES  FUTRELLE, JACQUES  LP
BEST TIME TRAVEL STORIES OF ALL TIME, THE  MALZBERG, BARRY N.  PA

BEST TIME TRAVEL STORIES OF THE 20TH CENTURY, THE TURTLEDOVE, HARRY/

GREENBERG, MARTIN H.

LPA

BEST YE BREED, THE  REYNOLDS, MACK  C/P:2,P
BESTIARY!  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
BESTIARY MOUNTAIN [BESTIARY 1]  FORRESTER, JOHN  H,P
BESTSELLERS GUARANTEED  LANSDALE, JOE R.  P
BETA COLONY  ENSTROM, ROBERT  H,SFWA
BETCHA CAN’T READ JUST ONE  FOSTER, ALAN DEAN  C/PA,PA
BETRAYAL [STAR WARS LEGACY OF THE FORCE 1]  ALLSTON, AARON  H
BETRAYAL [@DRAGONLANCE DHAMON #2]  RABE, JEAN  C/P,P
BETRAYAL [STAR TREK DS9]  TILTON, LOIS  C/P:2,P
BETRAYALS  FREEMANTLE, BRIAN  SFWA-P
BETRAYALS [@BLENDING #4]  GREEN, SHARON  C/P,P
BETRAYALS [@BABYLON 5 #6]  STIRLING, S. M.  C/P,P
BETTER ANGELS  HENDRIX, HOWARD V.  H,LP
BETTER IN THE DARK [SAINT-GERMAIN 8]  YARBRO, CHELSEA QUINN  H,LP
BETTER MAN, THE [STAR TREK]  WEINSTEIN, HOWARD  C/P:2,P
BETTER MANTRAP, A  SHAW, BOB  P
BETTER PART OF VALOR, THE [TORIN KERR 2]  HUFF, TANYA  P
BETTER THAN LIFE [@RED DWARF 2]  NAYLOR, GRANT  C/P:2,P
BETTER THAN ONE  KNIGHT, DAMON/WILHELM, KATE  SR-H
BETTYANN  NEVILLE, KRIS  P
BETWEEN, THE  DUE, TANANARIVE  LP
BETWEEN DARKNESS AND LIGHT [SHOLAN ALLIANCE 7]  NORMAN, LISANNE  P
BETWEEN DARKNESS AND LIGHT (PROOF) [SHOLAN ALLIANCE 7]  NORMAN, LISANNE  SR-LP
BETWEEN PLANETS  HEINLEIN, ROBERT A.  C/P,P,H,SR-LP
BETWEEN THE DARKNESS AND THE FIRE  DWIGHT, JEFFRY  LPA
BETWEEN THE DEVIL AND THE DEEP  MORGAN, J. M.  P
BETWEEN THE GALAXIES [PERRY RHODAN 119]  MAHR, KURT  SR-LP
BETWEEN THE RIVERS  TURTLEDOVE, HARRY  H,P
BETWEEN THE STARS [ISLAND WORLDS 3]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P:2,SFWA-P,P
BETWEEN THE STROKES OF NIGHT  SHEFFIELD, CHARLES  C/P:2,P
BETWEEN THE STROKES OF NIGHT (REVISED)  SHEFFIELD, CHARLES  P
BETWEEN TIME AND TIMBUKTU, OR PROMETHEUS-5  VONNEGUT, KURT JR.  H,LP
BETWEEN TWO WORLDS  CARR, TERRY  SR-DNESFA1
BETWEEN TWO WORLDS  FODOR, NANDOR  [C/RAN-P]
BETWEEN WORLDS  SILVERBERG, ROBERT  HA
BETWEEN WORLDS  SMITH, GARRETT  SR-LP
BEWARE MORE BEASTS  GHIDALIA, VIC/ELWOOD, ROGER  PA

The MITSFS Titledex  September 24, 2007  Page 30
TITLE  AUTHOR  CODE
BEWARE THE BEASTS  GHIDALIA, VIC/ELWOOD, ROGER  PA
BEWARE THE METAL CHILDREN [@OUTER LIMITS]  PEEL, JOHN  LP
BEWARE THE MICROBOTS [@PERRY RHODAN #35]  ACE  PA,C/PA
BEWARE TOMORROW  ELLIS, WALTER  P
BEWARE WHAT YOU WISH [@CHARMED]  GALLAGHER, DIANA G.  P
BEWITCHED  HINE, AL  P
BEWITCHED  O’DAY-FLANNERY, CONSTANCE  P
BEWITCHED AND BEWILDERED  BORISOFF, NORMAN  P
BEWITCHED CAVERNS, THE  RIENOW, LEONA  H
BEWITCHED: THE OPPOSITE UNCLE  JOHNSTON, WILLIAM  H
BEWITCHING FAMILIAR  BURNES, CAROLINE  P
BEWITCHMENTS OF LOVE AND HATE, THE [WRAETHTHU 2]  CONSTANTINE, STORM  P
BEYOND  BERKLEY BOOKS  PA,C/PA
BEYOND, THE [WELL-BUILT CITY 3]  FORD, JEFFREY  H
BEYOND  STURGEON, THEODORE  C/P:2,P

BEYOND (2003 CALENDAR)

SCIENCE APPLICATIONS INTERNATIONAL CORP.

CAL

BEYOND ANOTHER SUN  GODWIN, TOM  P
BEYOND APOLLO  MALZBERG, BARRY N.  H:2,P

BEYOND ARMAGEDDON MILLER, WALTER M. JR./ GREENBERG, MARTIN H.

HA,PA,SFWA

BEYOND BELIEF  HURLEY, R. J.  PA,C/PA
BEYOND CAPELLA  RACKHAM, JOHN  D05595
BEYOND CONTROL  SILVERBERG, ROBERT  HA,PA
BEYOND EARTH’S GATES  PADGETT, LEWIS/MOORE, C. L.  SR-DD069
BEYOND EDEN  DUNCAN, DAVID  C/P,P
BEYOND FLESH  DANN, JACK/DOZOIS, GARDNER  PA
BEYOND FUTURE SHOCK  ALANIZ, ALEX  LP
BEYOND GRAVITY [SALLY FORTH 3]  LEIBER, JUSTIN  C/P:2,SFWA-P,P
BEYOND HEAVEN’S RIVER  BEAR, GREG  P,SFWA-P
BEYOND HUMAN KEN  MERRIL, JUDITH  C/PA,PA
BEYOND HUMANITY [SALLY FORTH 2]  LEIBER, JUSTIN  C/P:2,P
BEYOND INFINITY [FALL OF NIGHT]  BENFORD, GREGORY  H
BEYOND INFINITY  CARR, ROBERT SPENCER  C/P:2,P
BEYOND LICH GATE [ELIXIR OF LIFE 1]  FRANKOS, STEVEN  P
BEYOND MIDNIGHT  MCCAULEY, KIRBY  PA
BEYOND RAGNAROK [RENSHAI CHRONICLES 1]  REICHERT, MICKEY ZUCKER  P,SFWA
BEYOND REALITY  CARR, TERRY  HA
BEYOND REJECTION [SALLY FORTH 1]  LEIBER, JUSTIN  C/P,P,H
BEYOND SANCTUARY [THIEVES’ WORLD TEMPUS 1]  MORRIS, JANET  H:2,P
BEYOND SILENCE  CAMERON, ELEANOR  P
BEYOND SINGULARITY  DANN, JACK/DOZOIS, GARDNER  PA
BEYOND SPACE AND TIME: AN ESP CASEBOOK  EBON, MARTIN  [C/RAN-H]
BEYOND STAR TREK  KRAUSS, LAWRENCE M.  R/REF-P,R/REF-H
BEYOND THE BARRIER  KNIGHT, DAMON  H:2,P
BEYOND THE BARRIER OF SPACE  TORRO, PEL  P
BEYOND THE BARRIERS OF SPACE AND TIME  MERRIL, JUDITH  HA
BEYOND THE BEYOND  ANDERSON, POUL  SR-H,H:2,P
BEYOND THE BLACK ENIGMA  SOMERS, BART  C/P:2,P
BEYOND THE BLUE EVENT HORIZON [HEECHEE 2]  POHL, FREDERIK  H:2,P,SFWA
BEYOND THE BLUE MOON  GREEN, SIMON R.  C/P,P
BEYOND THE BORDERS 7  HOWARD, ROBERT E.  P
BEYOND THE BURNING LANDS [LUKE 2]  CHRISTOPHER, JOHN  H,P
BEYOND THE CURTAIN OF DARK  HAINING, PETER  PA
BEYOND THE DAR AL-HARB  DICKSON, GORDON R.  C/P:2,SFWA-P,P
BEYOND THE DOOR  BLACKWOOD, GARY L.  SFWA
BEYOND THE DRAAK’S TEETH  BENNETT, MARCIA J.  C/P:2,P
BEYOND THE DRAGON PORTAL  HABER, MELISSA GLENN  H
BEYOND THE END OF TIME  POHL, FREDERIK  PA,C/PA
BEYOND THE FALL OF NIGHT  CLARKE, ARTHUR C./BENFORD, GREGORY  H:2,P
BEYOND THE FARTHEST STAR  BURROUGHS, EDGAR RICE  C/P:2,P
BEYOND THE FIELDS WE KNOW  DUNSANY, LORD  SR-P
BEYOND THE GALACTIC LENS [@CAP KENNEDY #16]  KERN, GREGORY  P
BEYOND THE GALACTIC RIM [GRIMES]  CHANDLER, A. BERTRAM  DF237
BEYOND THE GAP [OPENING OF THE WORLD]  TURTLEDOVE, HARRY  H
BEYOND THE GASLIGHT: SCIENCE IN POPULAR FICTION 1895–1905  EVANS, HILARY/EVANS, DIK  HA
BEYOND THE GATE [GOLDEN QUEEN 2]  WOLVERTON, DAVE  H,SFWA:2,P

BEYOND THE GATE OF WORLDS SILVERBERG, ROBERT/BRUNNER, JOHN/

YARBRO, CHELSEA QUINN

PA,SFWA-P

BEYOND THE GATES  WELLS, CATHERINE  C/P,P
BEYOND THE GATES OF DREAM  CARTER, LIN  C/P:2,P
BEYOND THE GOLDEN STAIR  BOK, HANNES  P
BEYOND THE HANGING WALL  DOUGLASS, SARA  H
BEYOND THE HIGH ROAD [@FORGOTTEN REALMS CORMYR 2]  DENNING, TROY  P
BEYOND THE IMPERIUM [IMPERIUM 2,3]  LAUMER, KEITH  C/P:2,P

TITLE  AUTHOR  CODE
BEYOND THE LAST STAR  SMITH, SHERWOOD  LPA
BEYOND THE MOON  HAMILTON, EDMOND  C/P,P
BEYOND THE MOONS [@SPELLJAMMER CLOAKMASTER #1]  COOK, DAVID  C/P:2,P
BEYOND THE NORTH WIND  BRADSHAW, GILLIAN  H
BEYOND THE OUTER MIRR [LEMMUS #2]  SAVARIN, JULIAN JAY  @SR-P,H
BEYOND THE PALE [LAST RUNE 1]  ANTHONY, MARK  P
BEYOND THE PALE [SHADOWRUN 30/DRAGON HEART 3]  KOKE, JAK  P
BEYOND THE RESURRECTION  EKLUND, GORDON  H,SFWA
BEYOND THE SAFE ZONE  SILVERBERG, ROBERT  C/P:2,P
BEYOND THE SEALED WORLD  VALE, RENA  C/P:2,P
BEYOND THE SHROUD [WORLD OF DARKNESS WRAITH]  HAUTALA, RICK  H
BEYOND THE SILVER SKY  BULMER, KENNETH  DD507
BEYOND THE SPECTRUM  THOMAS, MARTIN  C/P:2,P
BEYOND THE STARS  CUMMINGS, RAY  C/P:2,P
BEYOND THE STARS [QUEST FOR TOMORROW 4]  SHATNER, WILLIAM  H
BEYOND THE STRANGE  FATE, EDITORS OF  [C/RAN-P]
BEYOND THE SUMMERLAND [BINDING OF THE BLADE 1]  GRAHAM, L. B.  LP
BEYOND THE SUN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  JONES, MATTHEW  C/P,P
BEYOND THE THIRTEENTH SUN  WALDEN, MARK  P
BEYOND THE TOMORROW MOUNTAINS [NOREN 2]  ENGDAHL, SYLVIA LOUISE  H
BEYOND THE VANISHING POINT  CUMMINGS, RAY  DD331
BEYOND THE VEIL [THIEVES’ WORLD TEMPUS 2]  MORRIS, JANET  C/P,P,H
BEYOND THE VEIL  THANET, NEIL  SR-P
BEYOND THE VEIL OF STARS  REED, ROBERT  C/P,P,H
BEYOND THE VISIBLE  CAMPBELL, H. J.  SR-P,H
BEYOND THE VOID [@STARHAWKS #3]  DALTON, SEAN  C/P,P
BEYOND THE VOID  MULLER, JOHN E.  SR-P
BEYOND THE ’ZONE [SHATTERZONE 1]  STARK, ED  P
BEYOND THESE SUNS  LE PAGE, RAND  SR-P,C/P
BEYOND THIS HORIZON  CARRELL, CHRISTOPHER  VLPA
BEYOND THIS HORIZON  HEINLEIN, ROBERT A.  C/P,P,H
BEYOND TIME  JOHNS, MARSTON  H
BEYOND TIME  LEY, SANDRA  C/PA,PA
BEYOND TIME  MULLER, JOHN E.  SR-P
BEYOND TIME AND SPACE  DERLETH, AUGUST  PA,C/PA
BEYOND TIME AND SPACE  POTTER, ROBERT R.  LP
BEYOND TOMORROW  KNIGHT, DAMON  C/PA,PA
BEYOND UHURA [STAR TREK]  NICHOLS, NICHELLE  R/REF-H
BEYOND VARALLAN [STARDOC 2]  VIEHL, S. L.  C/P,P
BEYOND WIZARDWALL [THIEVES’ WORLD TEMPUS]  MORRIS, JANET  H:2,SFWA-P,P
BEYOND WORLD’S END [ERIC BANYON 3]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H:2
BEYOND ZOASTER  CHARLES, NEIL  SR-P
BEYONDERS, THE  WELLMAN, MANLY WADE  C/P,P
BHAGAVAD-GITA  GOD  PA
BHAGAVATI [BLOOD OF THE GODDESS 3]  DALKEY, KARA  H
BIBBLINGS  PAUL, BARBARA  C/P,P
BIBLIOGRAPH OF A. E. VAN VOGT  DENGROVE, RICHARD A.  SRL/IND
BIBLIOHOLISM  RAABE  [C/RAN-P]
BICENTENNIAL MAN AND OTHER STORIES, THE  ASIMOV, ISAAC  H:2,P,SFWA

BICYCLING THROUGH SPACE AND TIME [BICYCLING THROUGH SPACE AND TIME 1] SIROTA, MIKE

C/P,SFWA-P:2,P

BID TIME RETURN  MATHESON, RICHARD  SR-P
BIG BALL OF WAX, THE  MEAD, SHEPHERD  C/P:2,P
BIG BANG  GOULART, RON  P
BIG BLACK MARK, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
BIG BOOK OF SCIENCE FICTION, THE  CONKLIN, GROFF  PA,C/PA
BIG BOOK OF SPACE, THE  HURST, EARL OLIVER  [C/RAN-H]
BIG EMPTY, THE  LUCENO, JAMES  C/P,P
BIG EYE, THE  EHRLICH, MAX  C/P,SR-P,H:2
BIG FOUR, THE  CHRISTIE, AGATHA  P
BIG GAME [PREDATOR]  SCHOFIELD, SANDY  P
BIG GAME, THE [STAR TREK DS9]  SCHOFIELD, SANDY  C/P:2,P
BIG GAME, THE (PROOF) [STAR TREK DS9]  SCHOFIELD, SANDY  SR-VLP
BIG HUNT, THE [BERNICE SUMMERFIELD]  PARKIN, LANCE  H
BIG JUMP, THE  BRACKETT, LEIGH  C/P:2,DD103,P
BIG LIFTERS, THE  ING, DEAN  C/P,SFWA-P,P,H
BIG LOSERS, THE  TILLEY, ROBERT J.  P
BIG PLANET  VANCE, JACK  DD295,SFWA-P,H:2,P
BIG SHOW, THE  LAUMER, KEITH  C/P,P
BIG TIME, THE  LEIBER, FRITZ  C/P,DD491,SFWA-P,P,H
BIG U, THE  STEPHENSON, NEAL  H,LP
BIG WIN, THE  MILLER, JIMMY  C/P,P
BIG X, THE  SEARLS, HANK  H,P
BIGFOOT STALKS THE COAST OF MAINE  EASTON, THOMAS A.  LP

The MITSFS Titledex  September 24, 2007  Page 32
TITLE  AUTHOR  CODE
BIJAPUR [BLOOD OF THE GODDESS 2]  DALKEY, KARA  P
BIKINI PLANET  GARNETT, DAVID  P
BILBO LE HOBBIT  TOLKIEN, J. R. R.  FP
BILBO’S LAST SONG  TOLKIEN, J. R. R.  H:2
BILI THE AXE [HORSECLANS 10]  ADAMS, ROBERT  C/P:2,P
BILL, THE GALACTIC HERO [BILL 0]  HARRISON, HARRY  C/P,P,H
BILL, THE GALACTIC HERO ON THE PLANET OF BOTTLED BRAINS [BILL 2] HARRISON, HARRY/SHECKLEY, ROBERT  C/P:2,P

BILL, THE GALACTIC HERO ON THE PLANET OF TASTELESS PLEASURE [BILL 3] HARRISON, HARRY/BISCHOFF, DAVID F.

C/P:2,P

BILL, THE GALACTIC HERO ON THE PLANET OF TEN THOUSAND BARS [BILL 5] HARRISON, HARRY/BISCHOFF, DAVID F.

C/P,P

BILL, THE GALACTIC HERO ON THE PLANET OF ZOMBIE VAMPIRES [BILL 4] HARRISON, HARRY/HALDEMAN, JACK C. II

C/P:2,P

BILL, THE GALACTIC HERO: THE FINAL INCOHERENT ADVENTURE [BILL 6] HARRISON, HARRY/HARRIS, DAVID  C/P,SFWA-P,P
BILL THE GALACTIC HERO: THE PLANET OF THE ROBOT SLAVES [BILL 1]  HARRISON, HARRY  C/P,P
BILLENIUM  BALLARD, J. G.  C/P:2,P
BILLIBUB BADDINGS AND THE CASE OF THE SINGING SWORD  MORRIS, TEE  LP
BILLION DAYS OF EARTH, A  PISERCHIA, DORIS  C/P:2,P
BILLION DOLLAR BOY, THE  SHEFFIELD, CHARLES  C/P:2,P
BILLION DOLLAR BRAIN, THE  DEIGHTON, LEN  P

BILLION YEAR SPREE ALDISS, BRIAN W.

SFWA,R/REF-H, C/REF-P:2

BIMBOS OF THE DEATH SUN  MCCRUMB, SHARYN  C/P:2,P
BINARY  LANGE, JOHN  C/P:2,P
BINARY DIVINE  HARTRIDGE, JON  H:2,P
BINDER’S ROAD, THE [ILLUMINATION 2]  MCGARRY, TERRY  P
BINDING FORCE [@BATTLETECH 32]  COLEMAN, LOREN L.  P
BINDING STONE, THE [@EBERRON DRAGON BELOW #1]  BASSINGTHWAITE, DON  P
BINDINGS [@BOOKS OF MAGIC #2]  JABLONSKI, CARLA  P
BIO BLITZ [@ENFORCER #5]  SUGAR, ANDREW  P
BIO-FUTURES  SARGENT, PAMELA  C/PA:2,PA
BIO-MUTON  ELLIOT, LEE  SR-P

BIO OF AN OGRE ANTHONY, PIERS R/REF-P,SFWA,R/REF-H, C/REF-P:2,SFWA-P

BIO RESCUE [BIO RESCUE 1]  VIEHL, S. L.  H
BIOBLAST  GALLUN, RAYMOND Z.  C/P,P
BIODROIDS 2300  ALEXANDER, C. M.  P
BIOFAB WAR, THE [BIOFAB 1]  BERRY, STEPHEN AMES  P

BIOGRAPHICAL DICTIONARY OF SCIENCE FICTION AND FANTASY ARTISTS, A WEINBERG, ROBERT

R/REF-H

BIOGRAPHY OF THE EARTH  GAMOW, GEORGE  [C/RAN-P]
BIOLOGY "A"  KENT, BRAD  SR-P
BIOS  WILSON, ROBERT CHARLES  H
BIOTA RISK, THE  JERGENS, PHILLIP  P
BIPOHL  POHL, FREDERIK  C/P:2,P
BIRD OF TIME, THE  EFFINGER, GEORGE ALEC  H:2
BIRD OF TIME, THE  WEST, WALLACE  C/P,P,H
BIRDS-OF-A-FEATHER AFFAIR, THE [@GIRL FROM UNCLE #1]  AVALLONE, MICHAEL  P
BIRDS OF PREY  DRAKE, DAVID  C/P,SFWA-P,P,H
BIRTH AND DEATH OF THE SUN, THE  GAMOW, GEORGE  [C/RAN-P]
BIRTH OF AN AGE [@CHRIST CLONE #2]  BEAUSEIGNEUR, JAMES  P
BIRTH OF AN AGE (PROOF) [CHRIST CLONE 2]  BEAUSEIGNEUR, JAMES  SR-LP
BIRTH OF FIRE  POURNELLE, JERRY  C/P:2,P
BIRTH OF FLUX AND ANCHOR, THE [@SOUL RIDER #4]  CHALKER, JACK L.  C/P:2,SFWA-P,P
BIRTH OF THE BLADE, THE [THLASSA MEY 5]  MCCARTY, DENNIS  C/P,P
BIRTH OF THE PEOPLE’S REPUBLIC OF ANTARCTICA, THE  BATCHELOR, JOHN CALVIN  H,SFWA,LP
BIRTHDAY OF THE WORLD, THE  LE GUIN, URSULA K.  LP
BIRTHGRAVE, THE  LEE, TANITH  C/P:2,P
BIRTHPYRE  BRAND, LARRY  P
BIRTHRIGHT [STAR TREK TNG]  BRAGA, BRANDON  SR-S
BIRTHRIGHT [@DIABLO SIN WAR #1]  KNAAK, RICHARD A.  P
BIRTHRIGHT [DOCTOR WHO NEW ADVENTURES]  ROBINSON, NIGEL  C/P:2,P
BIRTHRIGHT  SKY, KATHLEEN  SR-P,C/P
BIRTHRIGHT: THE BOOK OF MAN  RESNICK, MIKE  C/P:2,P
BISHOP’S HEIR, THE [DERYNI KELSON 1]  KURTZ, KATHERINE  C/P,P,H,SFWA
BISHOP’S JAEGERS, THE  SMITH, THORNE  H

BITE HAMILTON, LAURELL K./ HARRIS, CHARLAINE/ DAVIDSON, MARYJANICE/ KNIGHT, ANGELA/TAYLOR, VICKIE

P

BITE OF MONSTERS, THE  O’NEIL, DENNIS  C/P:2,P
BITING THE SUN  LEE, TANITH  P
BITTEN [WOMEN OF THE OTHERWORLD 1]  ARMSTRONG, KELLEY  P
BITTER BLOOD [VAMPIRE LEGACY 2]  TAYLOR, KAREN E.  P

The MITSFS Titledex  September 24, 2007  Page 33
TITLE  AUTHOR  CODE
BITTER ENDS  BLOCH, ROBERT  PA
BITTER FRUIT [@DEATHLANDS]  AXLER, JAMES  P
BITTER GOLD HEARTS [GARRETT 2]  COOK, GLEN  C/P:2,P
BITTER WATERS [UKIAH OREGON 3]  SPENCER, WEN  C/P,P
BIZARRE SIGHTS & ODD VISIONS  BOND, SIMON  [C/RAN-P]
BLACK ABYSS  POWERS, J. L.  SR-P
BLACK AMBROSIA  ENGSTROM, ELIZABETH  SFWA-P
BLACK AND BLUE MAGIC  SNYDER, ZILPHA KEATLEY  LP
BLACK AS BLOOD  CHILSON, ROBERT  P
BLACK ATLANTIC [@JUDGE DREDD]  JOWETT, SIMON/EVANS, PETER J.  P
BLACK AURA  SLADEK, JOHN  P
BLACK AVENGERS, THE  STATTEN, VARGO  SR-P
BLACK BARGAIN  STATTEN, VARGO  SR-P
BLACK BEAST, THE [BOOKS OF THE ISLE 4]  SPRINGER, NANCY  C/P,P
BLACK BLADE, THE  BALLOTTI, MATTHEW  LP
BLACK BODY  TURK, H. C.  H
BLACK BOUQUET, THE [@FORGOTTEN REALMS ROGUES]  BYERS, RICHARD LEE  P
BLACK BOX, THE  SHIEL, M. P.  H
BLACK BRILLION [ARCHONATE 3]  HUGHES, MATTHEW  H
BLACK CANAAN  HOWARD, ROBERT E.  P
BLACK CASTLE, THE [DON SEBASTIAN 1]  DANIELS, LES  H,P
BLACK CAT  RUSSO, JOHN  P
BLACK CATS AND BROKEN MIRRORS  GREENBERG, MARTIN H./HELFERS, JOHN  C/PA,PA
BLACK CAULDRON, THE [PRYDAIN 2]  ALEXANDER, LLOYD  C/P:2,LP,H
BLACK CHALICE, THE  JAKOBER, MARIE  LP
BLACK CHARIOTS [@AVENGER #30]  ROBESON, KENNETH  C/P:2,P
BLACK CLOUD, THE  HOYLE, FRED  H:2,P
BLACK COCKTAIL  CARROLL, JONATHAN  SFWA,LP
BLACK COMPANY, THE [BLACK COMPANY 1]  COOK, GLEN  C/P:2,SFWA-P,P

BLACK COMPANY GOES SOUTH, THE [BLACK COMPANY 4,5/ BLACK COMPANY SOUTH 1,2] COOK, GLEN

H

BLACK COMPANY, THE: GLITTERING STONE I [BLACK COMPANY GLITTERING STONE 1,2] COOK, GLEN

H

BLACK COMPANY, THE: GLITTERING STONE II [BLACK COMPANY GLITTERING STONE 3,4] COOK, GLEN

H

BLACK CORRIDOR, THE  MOORCOCK, MICHAEL  H:2,P
BLACK CRUSADE  MACHEN, ARTHUR  P
BLACK DEATH, THE  CRAVENS, GWYNETH/MARR, JOHN S.  C/P,P
BLACK DEATH, THE [@AVENGER #22]  ROBESON, KENNETH  C/P,P
BLACK DIAMONDS, THE  SMITH, CLARK ASHTON  LP
BLACK DIAMONDS  VERNE, JULES  H
BLACK DOG, THE  GOFF, GEORGENA  P
BLACK DRAGON, THE [@TOM SWIFT (ARCHWAY) #1]  APPLETON, VICTOR  P
BLACK DRAGON [@BATTLETECH 29]  MILAN, VICTOR  C/P,P
BLACK EAGLE RISING [@ELITA #3]  JACOBY, KATE  P
BLACK EASTER  BLISH, JAMES  C/P,P,H
BLACK FIRE [STAR TREK]  COOPER, SONNI  C/P:3,P
BLACK FLAME, THE  ABBEY, LYNN  C/P,LP,H
BLACK FLAME, THE  WEINBAUM, STANLEY G.  SR-H,C/P:2,P
BLACK GALAXY, THE  LEINSTER, MURRAY  C/P,GN4.20
BLACK GENESIS [@MISSION EARTH #2]  HUBBARD, L. RON  C/P,SFWA-P,H,P
BLACK GOD’S SHADOW  MOORE, C. L.  SR-H
BLACK GRAIL, THE  BRODERICK, DAMIEN  C/P:2,P
BLACK GRYPHON, THE [MAGE WARS 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P,SFWA,P,H
BLACK HARVEST [@DEATHLANDS]  AXLER, JAMES  P
BLACK HEART, IVORY BONES [COMPLEMENTARY FAIRY COLORS 6]  DATLOW, ELLEN/WINDLING, TERRI  LPA
BLACK HOLE, THE  FOSTER, ALAN DEAN  C/P,P,H
BLACK HOLE OF CARCOSA, THE [KAMUS OF KADIZHAR]  SHIRLEY, JOHN  C/P,P
BLACK HOLE PLANET  PEIRCE, HAYFORD  H
BLACK HOLES  POURNELLE, JERRY  PA,C/PA
BLACK HOLES: THE END OF THE UNIVERSE?  TAYLOR, JOHN G.  SFWA
BLACK HORDE, THE  LEWIS, RICHARD  C/P,P
BLACK HORSES FOR THE KING  MCCAFFREY, ANNE  H,LP
BLACK IN TIME  JAKES, JOHN  C/P:2,P
BLACK INFINITY  BRETT, LEO  SR-P
BLACK INFINITY  CAMERON, BERL  SR-P
BLACK IS THE COLOR  BRUNNER, JOHN  C/P:2,P
BLACK JADE ROAD, THE [TEN THOUSAND WILLOWS 2]  GRANT, KATHRYN  C/P,P
BLACK JEWELS TRILOGY, THE [BLACK JEWELS 1,2,3]  BISHOP, ANNE  LP
BLACK JUICE  LANAGAN, MARGO  H,P
BLACK KING, THE [BLACK THRONE 2]  RUSCH, KRISTINE KATHRYN  P
BLACK KNIGHT OF THE IRON SPHERE [@LORD TEDRIC #3]  EKLUND, GORDON  H,LP,P
BLACK LEGION OF CALLISTO [JANDAR 2]  CARTER, LIN  C/P:2,P
BLACK LIGHT  HAND, ELIZABETH  H

The MITSFS Titledex  September 24, 2007  Page 34
TITLE  AUTHOR  CODE
BLACK LIGHTNING  SAUL, JOHN  P
BLACK LYNX, THE [SKYLA 2]  BOYER, ELIZABETH H.  P
BLACK MADONNA [@SHADOWRUN 20]  SARGENT, CARL/GASCOIGNE, MARC  P
BLACK MAGIC  WARD, DON  C/PA,PA
BLACK MAGICIAN, THE  EDMONDSON, G. C./KOTLAN, C. M.  P
BLACK MARIA  JONES, DIANA WYNNE  LP
BLACK MEDICINE  BURKS, ARTHUR J.  SR-H
BLACK MILK  REED, ROBERT  C/P:2,P
BLACK MIST  CARD, ORSON SCOTT/FERRELL, KEITH  PA
BLACK MOON, THE [CABAL 2]  DUNN, PHILIP  P
BLACK MOON INN [ANITA BLAKE 7,8]  HAMILTON, LAURELL K.  H
BLACK MOUNTAINS, THE [EMPIRE OF THE EAST 2]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
BLACK ON BLACK  WENTWORTH, K. D.  P
BLACK ORCHID [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P,C/P
BLACK ORCHID  GAIMAN, NEIL/MCKEAN, DAVE  CX
BLACK OXEN  ATHERTON, GERTRUDE  SR-H
BLACK POCKETS  ZEBROWSKI, GEORGE  HA
BLACK POWDER WAR [TEMERAIRE 3]  NOVIK  C/P:2,P
BLACK PROJECTS, WHITE KNIGHTS [COMPANY]  BAKER, KAGE  H
BLACK QUEEN, THE [BLACK THRONE 1]  RUSCH, KRISTINE KATHRYN  P
BLACK RAVEN, THE [DEVERRY 10/DRAGON MAGE 2]  KERR, KATHARINE  H,LP
BLACK ROAD, THE [@DIABLO #2]  ODOM, MEL  C/P,P
BLACK ROADS, THE  HENSLEY, J. L.  SR-P
BLACK ROOD, THE [CELTIC CRUSADES 2]  LAWHEAD, STEPHEN R.  P
BLACK SABBAT  HERMAN, J. B.  P
BLACK SEAS OF INFINITY  LOVECRAFT, H. P.  H:2
BLACK SHIELDS, THE [STORMLANDS 2]  ROBERTS, JOHN MADDOX  C/P:2,P
BLACK SHIP, THE  ROWLEY, CHRISTOPHER  C/P:2,P
BLACK SHORE, THE [STAR TREK VOYAGER]  COX, GREG  C/P:2,P
BLACK SNOW DAYS  O’KEEFE, CLAUDIA  C/P,P
BLACK SPOT, THE [DOC SAVAGE #76]  ROBESON, KENNETH  C/P:2,@P
BLACK STAR, THE  CARTER, LIN  C/P:2,P
BLACK STAR CHRONICLES  ARNESON, D. J./TALLARICO, TONY  P
BLACK STAR PASSES, THE  CAMPBELL, JOHN W. JR.  SR-H,C/P:2,P
BLACK STAR RISING  POHL, FREDERIK  C/P,P,H
BLACK STEEL [MATADOR 6]  PERRY, STEVE  C/P,SFWA-P:2,P
BLACK STORM RISING [ORIANUS CREATION 1]  BURTON, D. SHANE  LP
BLACK STRANGER, THE  HOWARD, ROBERT E.  LP
BLACK SUN  LEININGER, ROBERT  P
BLACK SUN  TARABILDA, JAMES  P
BLACK SUN, THE  WILLIAMSON, JACK  H,P
BLACK SUN RISING [COLDFIRE 1]  FRIEDMAN, C. S.  H:2,P
BLACK SWAN, THE  LACKEY, MERCEDES  C/P,P,H
BLACK SWAN, WHITE RAVEN [COMPLEMENTARY FAIRY COLORS 4]  DATLOW, ELLEN/WINDLING, TERRI  HA
BLACK THORN, WHITE ROSE [COMPLEMENTARY FAIRY COLORS 2]  DATLOW, ELLEN/WINDLING, TERRI  C/PA,PA
BLACK THRONE, THE  ZELAZNY, ROGER/SABERHAGEN, FRED  C/P,SFWA-P,P,H
BLACK TOWER, THE [DUNGEON #1]  LUPOFF, RICHARD A.  C/P:2,P

BLACK TRILLIUM [TRILLIUM 1] BRADLEY, MARION ZIMMER/MAY, JULIAN/

NORTON, ANDRE

C/P,P,H

BLACK TRUMP [@WILD CARDS NEW CYCLE 3]  MARTIN, GEORGE R. R.  C/PA,PA,SFWA-P
BLACK UNICORN, THE [@LANDOVER 2]  BROOKS, TERRY  C/P,P,H:2
BLACK UNICORN [TANAQUIL 1]  LEE, TANITH  H:2,SFWA,P
BLACK VESSEL, THE [MYSTARA]  SIMON, MORRIS  P
BLACK VULMEA’S VENGEANCE  HOWARD, ROBERT E.  C/P,VLP
BLACK WHEEL, THE  MERRITT, A./BOK, HANNES  C/P,P
BLACK WIND  WILSON, F. PAUL  H,SFWA-P,P
BLACK WINE  DORSEY, CANDAS JANE  H,LP
BLACK WINE  WINTER, DOUGLAS E.  HA
BLACK WING, THE [@DRAGONLANCE VILLAINS #2]  KIRCHOFF, MARY  C/P,P
BLACK WING OF MARS  STATTEN, VARGO  SR-P
BLACK WIZARDS [@FORGOTTEN REALMS MOONSHAE #2]  NILES, DOUGLAS  C/P,P
BLACK WOLF, THE  ELFLANDSSON, GALAD  SR-H
BLACK WOLF [@FORGOTTEN REALMS SEMBIA]  GROSS, DAVE  P
BLACKBIRD IN SILVER, A  WARRINGTON, FREDA  LP
BLACKCOLLAR, THE [BLACKCOLLAR 1]  ZAHN, TIMOTHY  C/P:2,P
BLACKCOLLAR: THE JUDAS SOLUTION [BLACKCOLLAR 2]  ZAHN, TIMOTHY  H
BLACKGOD, THE [CHILDREN OF THE CHANGELING 2]  KEYES, J. GREGORY  H,P
BLACKMAN’S BURDEN  REYNOLDS, MACK  D06612
BLACKMANTLE [KELTIAD MISC 1]  KENNEALY-MORRISON, PATRICIA  H:2,P
BLACKOUT [@BUFFY]  DECANDIDO, KEITH R. A.  P
BLACKSTAFF [FORGOTTEN REALMS WIZARDS]  SCHEND, STEVEN E  P
BLACKWOOD FARM [LESTAT]  RICE, ANNE  H:2
BLADE  ODOM, MEL  P
BLADE DANCER [STARDOC JORY RASK]  VIEHL, S. L.  H,P

The MITSFS Titledex  September 24, 2007  Page 35
TITLE  AUTHOR  CODE
BLADE OF CONAN, THE [CONAN]  DE CAMP, L. SPRAGUE  R/REF-P,C/REF-P
BLADE OF FORTRIU [@BRIDEI #2]  MARILLIER, JULIET  H
BLADE OF THE POISONER  HILL, DOUGLAS  P
BLADE OF TYSHALE [HARI MICHAELSON 2]  STOVER, MATTHEW WOODRING  P
BLADE RUNNER  DICK, PHILIP K.  C/P:2,P
BLADE TRINITY  RHODES, NATASHA  P
BLADERUNNER, THE  NOURSE, ALAN E.  C/P,P,H
BLADES FROM THE WILLOWS  HUANZHULOUZHU  LP
BLADES OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 2]  THORPE, GAV  P
BLADES OF MARS [MICHAEL KANE 2]  BRADBURY, EDWARD P.  SR-P,C/P:2
BLADES OF MARS [MICHAEL KANE 2]  MOORCOCK, MICHAEL  P
BLADES OF THE TIGER [@DRAGONLANCE TALADAS #1]  PIERSON, CHRIS  P
BLADESINGER [@FORGOTTEN REALMS FIGHTERS]  STROHM, KEITH FRANCIS  P
BLADESMAN OF ANTARES [DRAY PRESCOT 9]  AKERS, ALAN BURT  P
BLAEDUD THE BIRDMAN  CHAPMAN, VERA  C/P:2,P
BLAKELY’S ARK  MACMILLAN, IAN  P
BLAKE’S PROGRESS  NELSON, RAY FARADAY  SR-P
BLAKE’S SEVEN [BLAKE’S 7]  HOYLE, TREVOR  C/P,P
BLAL, THE  VAN VOGT, A. E.  C/P,P
BLAMELESS IN ABADDON [CORPUS DEI 2]  MORROW, JAMES  H
BLANK SLATE  MCGARRY, MARK J.  P
BLANKET OF THE DARK, THE  BUCHAN, JOHN  H
BLAST OFF INTO SPACE [LOST IN SPACE]  PACKER, PETER  SR-S
BLAZE OF GLORY [STAR TREK TNG]  HAWKE, SIMON  C/P:2,P
BLAZE OF WRATH [HARRY BORG 3]  HAWKINS, WARD  C/P,P
BLAZING AFFAIR, THE [@GIRL FROM UNCLE #2]  AVALLONE, MICHAEL  P
BLAZING SUN [@PERRY RHODAN #86]  ACE  C/PA:2,PA
BLAZON  BULMER, KENNETH  P

BLEAK SEASONS [@BLACK COMPANY GLITTERING STONE 1/ BLACK COMPANY 6] COOK, GLEN

H,P

BLEEDING CHALICE, THE [@WARHAMMER 40K SOUL DRINKERS 2]  COUNTER, BEN  P
BLEEDING FROM THE ROMAN  ROMILLY, ERIC  H
BLESS THE BEASTS [STAR TREK VOYAGER]  HABER, KAREN  P
BLESSING PAPERS, THE  BARNWELL, WILLIAM  P
BLESSINGS AND CURSES [@CRAFTERS #2]  STASHEFF, CHRISTOPHER/FAWCETT, BILL  C/PA,PA,SFWA-P:2
BLIND AMBITIONS  SIROIS, A. L.  LP
BLIND ARCHER, THE  BETANCOURT, JOHN GREGORY  C/P,P
BLIND FIRE [@ZONE #2]  ROUCH, JAMES  P
BLIND GEOMETER, THE  ROBINSON, KIM STANLEY  SFWA-TD13,TD13
BLIND JUSTICE  LEWITT, SHARIANN  P,SFWA-P:2
BLIND KNIGHT, THE  VAN ASTEN, GAIL  P
BLIND LAKE  WILSON, ROBERT CHARLES  H
BLIND MEN AND THE ELEPHANT, THE  GRIFFIN, RUSSELL M.  P,SFWA-P
BLIND PROPHET  DAVIS, BART  H
BLIND SPOT, THE  HALL, AUSTIN/FLINT, HOMER EON  SR-H,C/P:2,P
BLIND VOICES  REAMY, TOM  H:2,P
BLIND WAVES  GOULD, STEVEN C.  C/P,P,H
BLIND WORM, THE  STABLEFORD, BRIAN M.  D06707
BLINDFOLD  ANDERSON, KEVIN J.  P
BLINDSIGHT  STEWART, MICHAEL  H
BLINDSIGHT  WATTS, PETER  H:2
BLINK OF THE MIND [TWILIGHT #5]  FRANCIS, DOROTHY BRENNER  P
BLITZKRIEG GALACTICA [PERRY RHODAN 129]  BRAND, KURT  SR-LP
BLIZZARD  STONE, GEORGE  H,P
BLOB, THE  BISCHOFF, DAVID F.  P
BLOCKADE: LEPSO [@PERRY RHODAN #101]  ACE  PA
BLOND BARBARIANS AND NOBLE SAVAGES  DE CAMP, L. SPRAGUE  R/REF-P
BLONDE GODDESS, THE  RICHARDS, R. P. J.  H
BLOOD  DEE, RON  P
BLOOD [SECOND ETHER 1]  MOORCOCK, MICHAEL  H
BLOOD & IVORY [JAMETHIEL DREAM-WEAVER]  HODGELL, P. C.  LP
BLOOD & RUST [CLEVELAND UNDEAD 1,2]  SWINIARSKI, S. A.  P
BLOOD & STEEL [@WARHAMMER BRUNNER 2]  WERNER, C. L.  P
BLOOD AND BONE [EAGLEHEART 3]  WESTCOTT, C. T.  C/P:2,P
BLOOD AND BURNING  BUDRYS, ALGIS  C/P,P
BLOOD AND DREAMS [PARSIVAL 4]  MONACO, RICHARD  P
BLOOD AND FIRE  GERROLD, DAVID  LP
BLOOD AND FOG [@BUFFY]  HOLDER, NANCY  P
BLOOD AND GOLD [LESTAT]  RICE, ANNE  H
BLOOD AND HONOR [@EBERRON WAR-TORN #4]  DAVIS, GRAEME  P
BLOOD AND HONOR  GREEN, SIMON R.  C/P,P
BLOOD AND HOPE [DOCTOR WHO]  MCLAUGHLIN, IAIN  SR-H
BLOOD AND IRON [PROMETHEAN AGE 1]  BEAR, ELIZABETH  LP
BLOOD AND IRON [@THERE WILL BE WAR #3]  POURNELLE, JERRY  C/PA,PA

The MITSFS Titledex  September 24, 2007  Page 36
TITLE  AUTHOR  CODE
BLOOD AND IRON [CONFEDERATE STATES AMERICAN EMPIRE 1]  TURTLEDOVE, HARRY  C/P,P,H
BLOOD AND JUDGMENT  WALKER, LARS  P
BLOOD AND MEMORY [@QUICKENING #2]  MCINTOSH, FIONA  C/P,LP
BLOOD AND WAR [@HARRIERS #2]  DICKSON, GORDON R.  C/PA,PA
BLOOD AUTUMN  PTACEK, KATHRYN  SFWA-P
BLOOD AVATAR [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P
BLOOD BOUND [MERCY THOMPSON 2]  BRIGGS, PATRICIA  C/P,P
BLOOD BOWL [@BLOOD BOWL 1]  FORBECK, MATT  P
BLOOD BROTHERS  BARNES, STEVEN  P
BLOOD BROTHERS OF GOR [GOR 18]  NORMAN, JOHN  C/P,P
BLOOD CANTICLE [LESTAT]  RICE, ANNE  H:2
BLOOD CHILD  BUTLER, OCTAVIA E.  H
BLOOD CONTACT [@STARFIST 4]  SHERMAN, DAVID/CRAGG, DAN  P
BLOOD COUNTESS, THE [@AVENGER #33]  ROBESON, KENNETH  C/P:2,P
BLOOD COUNTY  SELBY, CURT  P
BLOOD DEBT [VICTORY NELSON 5]  HUFF, TANYA  C/P:2,P
BLOOD DEBT, THE [@CATACLYSM #2]  WILLIAMS, SEAN  H
BLOOD DRIVE [ANNA STRONG 2]  STEIN, JEANNE C.  P
BLOOD ENEMY [@UNDERWORLD 0]  COX, GREG  P
BLOOD FEUD  SICILIANO, SAM  P

BLOOD FEUDS [@WAR WORLD] STIRLING, S. M./TARR, JUDITH/

SHWARTZ, SUSAN/TURTLEDOVE, HARRY

C/P,SFWA-P,P

BLOOD FOLLOWS [MALAZAN]  ERIKSON, STEVEN  SR-LP
BLOOD FOUNTAIN [SWORDS OF RAEMLLYN 3]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P:2,P
BLOOD GAMES [SAINT-GERMAIN 3]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
BLOOD HARVEST [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P:2,P
BLOOD HEAT [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P,P
BLOOD HERITAGE  TEPPER, SHERI S.  C/P:2,P
BLOOD HIGHWAY [@OUTRIDER #3]  HARDING, RICHARD  P
BLOOD HOSTAGES [@PLANESCAPE BLOOD WARS #1]  KING, J. ROBERT  P
BLOOD HUNT [GARRETH MIKAELIAN 1]  KILLOUGH, LEE  C/P,SFWA-P,P
BLOOD IN THE ASHES [ASHES 4]  JOHNSTONE, WILLIAM W.  P
BLOOD IS NOT ENOUGH  DATLOW, ELLEN  C/PA,PA
BLOOD JAGUAR, THE  PAYNE, MICHAEL H.  H,P
BLOOD KNIGHT, THE [KINGDOMS OF THORN AND BONE 3]  KEYES, J. GREGORY  H
BLOOD LEGACY [@BATTLETECH #21/BLOOD OF KERENSKY 2]  STACKPOLE, MICHAEL A.  P
BLOOD LINES  BURKETT, WILLIAM R. JR.  P
BLOOD LINES [VICTORY NELSON 3]  HUFF, TANYA  C/P:2,P
BLOOD LINES [LILY YU 3]  WILKS, EILEEN  P
BLOOD LUST  DEE, RON  P
BLOOD MONEY [@ARHAMMER BRUNNER 1]  WERNER, C. L.  P
BLOOD MOON  DIVONO, SHARMAN  P
BLOOD MUSIC  BEAR, GREG  H:2,P,SFWA
BLOOD OATH [@BABYLON 5 #3]  VORNHOLT, JOHN  C/P,P
BLOOD OF A DRAGON, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
BLOOD OF AMBER [AMBER 7]  ZELAZNY, ROGER  C/P,H,@P,SFWA
BLOOD OF HEROES [@BATTLETECH 11]  KEITH, ANDREW  C/P,P
BLOOD OF MYSTERY [LAST RUNE 4]  ANTHONY, MARK  P

BLOOD OF TEN CHIEFS, THE [@BLOOD OF TEN CHIEFS #1] PINI, RICHARD/ASPRIN, ROBERT LYNN/

ABBEY, LYNN

SFWA-P,C/PA,PA

BLOOD OF THE COLYN MUR  ZIMMER, PAUL EDWIN/DE CLES, JON  P
BLOOD OF THE DRAGON [@WARHAMMER BRUNNER 3]  WERNER, C. L.  P
BLOOD OF THE FOLD [SWORD OF TRUTH 3]  GOODKIND, TERRY  C/P:2,P,H
BLOOD OF THE HEROES  WHITE, STEVE  H
BLOOD OF THE IMPALER  SACKETT, JEFFREY  P,SFWA-P
BLOOD OF THE ISLE [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
BLOOD OF THE LAMB, THE  MONTELEONE, THOMAS F.  H,P
BLOOD OF THE LAND, THE [RILSIN 2]  BRENNAN, NOEL-ANNE  P
BLOOD OF THE TIGER [LOST LANDS 1]  ESTES, ROSE  P
BLOOD OF WOLVES [@AGE OF CONAN KERN #1]  COLEMAN, LOREN L.  P
BLOOD ON THE MOON  COHEN, BARNEY  P
BLOOD ON THE SUN [WORLD OF DARKNESS VAMPIRE]  HERBERT, BRIAN/LANDIS, MARIE  P
BLOOD ON THE WATER [@VAMPIRE FILES #6]  ELROD, P. N.  C/P:2,P
BLOOD PACT [VICTORY NELSON 4]  HUFF, TANYA  C/P:2,P
BLOOD PRICE [VICTORY NELSON 1]  HUFF, TANYA  C/P:2,P
BLOOD QUEST [SURVIVAL 2000 1]  MCPHEE, JAMES  P
BLOOD RED ANGEL  COLE, ADRIAN  P
BLOOD RED ROSES [TWILIGHT 8]  ARMSTRONG, SARAH  P
BLOOD RELATIONS [@WORLD OF DARKNESS VAMPIRE]  MURRAY, DOUG  P
BLOOD RELATIONS  SIROIS, A. L.  LP
BLOOD RELATIVE [@ROGUE TROOPER]  SWALLOW, JAMES  P
BLOOD RING, THE [@AVENGER #6]  ROBESON, KENNETH  C/P:2,P
BLOOD RITES  BERGSTROM, ELAINE  P
BLOOD RITES [@DRESDEN FILES #6]  BUTCHER, JIM  C/P,P

The MITSFS Titledex  September 24, 2007  Page 37
TITLE  AUTHOR  CODE
BLOOD RIVER DOWN [QUEST FOR THE WHITE DUCK 1]  FENN, LIONEL  P,SFWA-P
BLOOD ROSES [SAINT-GERMAIN 12]  YARBRO, CHELSEA QUINN  LP
BLOOD ROYAL [@NECROMUNDA]  MCDERMOTT, WILL/RENNIE, GORDON  P
BLOOD SECRETS [VAMPIRE LEGACY 1]  TAYLOR, KAREN E.  P
BLOOD SPORT  JONES, ROBERT F.  P
BLOOD SPORT  SCOTT, R. C.  P
BLOOD SPORT [SHADOWRUN 29]  SMEDMAN, LISA  P
BLOOD STONES, THE [BALZAN 1]  MOORE, WALLACE  P
BLOOD TIDE [@OMEGA SUB #4]  CAMERON, J. D.  P
BLOOD TIES [@THIEVES’ WORLD #9]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
BLOOD TIES [@TIMECOP 3]  PARKINSON, DAN  P
BLOOD TO BLOOD  BERGSTROM, ELAINE  P
BLOOD TRAIL [VICTORY NELSON 2]  HUFF, TANYA  C/P:2,P
BLOOD TRILLIUM [TRILLIUM 2]  MAY, JULIAN  H:2,P

BLOOD VENGEANCE [@WAR WORLD] STIRLING, S. M./TARR, JUDITH/

SHWARTZ, SUSAN/TURTLEDOVE, HARRY

P,SFWA-P

BLOOD WEDDING [@VAMPIRELLA #4]  GOULART, RON  P
BLOODCIRCLE [@VAMPIRE FILES #3]  ELROD, P. N.  C/P,P
BLOODED [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
BLOODED ON ARACHNE  BISHOP, MICHAEL  C/P,P
BLOODFANG [BLOODFANG 3]  WEAVER, MICHAEL D.  P
BLOODHYPE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
BLOODLETTER [STAR TREK DS9]  JETER, K. W.  C/P:2,P
BLOODLINES [@MAGIC: THE GATHERING ARTIFACTS 4]  COLEMAN, LOREN L.  C/P,P
BLOODLINES [@STAR WARS LEGACY OF THE FORCE 2]  TRAVISS, KAREN  P
BLOODLINKS [GARRETH MIKAELIAN 2]  KILLOUGH, LEE  P,SFWA-P
BLOODLIST [@VAMPIRE FILES #1]  ELROD, P. N.  C/P:2,P
BLOODLOVER [POWER OF THE BOOD 4]  KILPATRICK, NANCY  LP
BLOODMIST  TREMAYNE, PETER  P,SFWA-P
BLOODNAME [@BATTLETECH #2/JADE PHOENIX 2]  THURSTON, ROBERT  C/P,P
BLOODRIGHT  TREMAYNE, PETER  H,P
BLOODRIGHTS  WOOD, N. LEE  P
BLOODRING  HUNTER, FAITH  LP
BLOODSONGS [FROST 3]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
BLOODSPORT  BURKETT, WILLIAM R. JR.  P
BLOODSTALK [@VAMPIRELLA #1]  GOULART, RON  C/P,P
BLOODSTONE [STONES OF POWER 5]  GEMMELL, DAVID  C/P,P
BLOODSTONE  WAGNER, KARL EDWARD  C/P:2,SFWA-P,P
BLOODSTORM [@WARHAMMER MALUS DARKBLADE 2]  ABNETT, DAN/LEE, MIKE  P
BLOODSTORM [NAGA TEOT 2]  GLADNEY, HEATHER  C/P:2,P
BLOODSUCKING FIENDS: A LOVE STORY  MOORE, CHRISTOPHER  LP
BLOODTHIRST [STAR TREK]  DILLARD, J. M.  C/P:2,P
BLOODTIDE  BURGESS, MELVIN  H,P
BLOODWALK [FORGOTTEN REALMS WIZARDS 5]  DAVIS, JAMES P  P
BLOODWALK [GARRETH MIKAELIAN 1,2]  KILLOUGH, LEE  LP

BLOODWAR [@WORLD OF DARKNESS VAMPIRE MASQUERADE RED DEATH 1] WEINBERG, ROBERT

P

BLOODWIND, THE  GRANT, CHARLES L.  C/P,SFWA-P,P
BLOODWINTER [ERON 1]  DEITZ, TOM  LP
BLOODWORLD  JANIFER, LAURENCE M.  C/P:2,P
BLOODY AMERICA [@DOOMSDAY WARRIOR #4]  STACY, RYDER  P
BLOODY BONES [ANITA BLAKE 5]  HAMILTON, LAURELL K.  C/P:2,P
BLOODY CROWN OF CONAN, THE [CONAN]  HOWARD, ROBERT E.  H
BLOODY EYE, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
BLOODY RED BARON, THE [ANNO DRACULA 2]  NEWMAN, KIM  P

BLOODY SEPTEMBER [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #3] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

BLOODY SUN, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
BLOODY SUN, THE (EXPANDED) [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
BLOOM  MCCARTHY, WIL  C/P,P
BLOOP BOX, THE [NANNY AND THE PROFESSOR 3]  JOHNSTON, WILLIAM  P
BLOSSOM CULP AND THE SLEEP OF DEATH [BLOSSOM CULP 4]  PECK, RICHARD  P
BLOSSOM OF ERDA, THE  TAYLOR, L. A.  H,SFWA,P
BLOTE ZON, DE  ASIMOV, ISAAC  FP
BLOWN [HERALD CHILDE 2]  FARMER, PHILIP JOSE  SR-L
BLOWN DEAD [HORN]  SLOANE, BEN  P
BLUE ADEPT [APPRENTICE ADEPT #2]  ANTHONY, PIERS  C/P,@P,H
BLUE ANGEL, THE [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL/HOAD, JEREMY  P
BLUE ASP  LE PAGE, RAND  SR-P
BLUE ATOM, THE  WILLIAMS, ROBERT MOORE  DD322
BLUE BAMBOO  DAZAI, OSAMU  P
BLUE BOX [DOCTOR WHO PAST DOCTORS]  ORMAN, KATE  P
BLUE CHAIR, THE  THOMPSON, JOYCE  C/P,P
BLUE CHAMPAGNE  VARLEY, JOHN  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 38
TITLE  AUTHOR  CODE
BLUE CORDON  CARTER, DEE  SR-P
BLUE DWARFS, THE [@PERRY RHODAN #54]  ACE  PA,C/PA
BLUE FACE  EDMONDSON, G. C.  C/P:2,P
BLUE FAIRY BOOK, THE  LANG, ANDREW  SR-HA
BLUE GHOST MYSTERY, THE [@RICK BRANT #15]  BLAINE, JOHN  H
BLUE GIRL, THE  DE LINT, CHARLES  H
BLUE GROTTO TERROR, THE  CLAUDY, CARL H.  H:2
BLUE HAWK, THE  DICKINSON, PETER  C/P:2,P
BLUE ICE PILOT, THE  FISHER, LOU  C/P,P
BLUE JUGGERNAUT  FANE, BRON  H
BLUE LAB  JONES, J. A.  P
BLUE LIGHT  MOSLEY, WALTER  H
BLUE LIGHT RED LIGHT VOL. 2 #1  PARKER, JOY  SFWA
BLUE LIMBO  GREEN, TERENCE M.  H,P
BLUE MAGIC [SOUL DRINKER 2]  CLAYTON, JO  C/P:2,P
BLUE MARS [MARS 3]  ROBINSON, KIM STANLEY  C/P,P,H
BLUE MOON [ANITA BLAKE 8]  HAMILTON, LAURELL K.  C/P:2,P
BLUE MOON RISING  GREEN, SIMON R.  C/P:2,SFWA-P:2,P
BLUE MOTEL  CROWTHER, PETER  PA
BLUE NOON [@MIDNIGHTERS #3]  WESTERFELD, SCOTT  H
BLUE NOON (PROOF) [@MIDNIGHTERS #3]  WESTERFELD, SCOTT  SR-LP
BLUE STAR, THE  PRATT, FLETCHER  C/P:2,P
BLUE SWORD, THE  MCKINLEY, ROBIN  H:2,P
BLUE SYSTEM, THE [@PERRY RHODAN #99]  ACE  C/PA,PA
BLUE WALL, THE  CHILD, RICHARD WASHBURN  H
BLUE WORLD  MCCAMMON, ROBERT R.  C/P:2,P
BLUE WORLD, THE  VANCE, JACK  C/P:2,P
BLUE WORLD (VIE), THE [@VANCE INTEGRAL EDITION #16]  VANCE, JACK  SR-H
BLUEBEARD  VONNEGUT, KURT JR.  H,P
BLUEBLOODS [STARHUNTERS 3]  DRAKE, DAVID  C/PA,PA,SFWA-P
BLUEHEART  SINCLAIR, ALISON  P
BLUEPRINT FOR YESTERDAY  WETHERELL, JUNE  H,SFWA
BLUESONG  VAN SCYOC, SYDNEY J.  C/P:2,LP
BOAT OF A MILLION YEARS, THE  ANDERSON, POUL  H:2,SFWA-P,P
BOATS OF THE "GLEN CARRIG", THE  HODGSON, WILLIAM HOPE  C/P:2,LP
BOBBY’S GIRL  AUSTIN, J. D.  P
BODDEKKER’S DEMONS [PEMBROKE HALL 2]  FAUST, JOE CLIFFORD  P
BODELAN WAY, THE  TRIMBLE, LOUIS  C/P,P
BODE’S CARTOON CONCERT  BODE, VAUGHN  SCX
BODY AND SOUL [ALIEN NATION 3]  DAVID, PETER  C/P:2,P

BODY ARMOR: 2000 HALDEMAN, JOE/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA

BODY MORTGAGE  ENGLING, RICHARD  P
BODY SNATCHERS, THE  FINNEY, JACK  C/P:2,P
BODYGUARD  DIETZ, WILLIAM C.  C/P:2,P
BODYGUARD  GOLD, HORACE L.  PA,C/PA
BODYGUARD AND FOUR OTHER SHORT NOVELS FROM GALAXY  GOLD, HORACE L.  HA
BODYGUARD OF LIGHTNING [ORCS FIRST BLOOD 1]  NICHOLLS, STAN  LP
BODYSNATCHERS, THE [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, MARK  C/P:2,P
BOG, THE  TALBOT, MICHAEL  H
BOGEY MEN  BLOCH, ROBERT  C/P:2,P
BOGEYMAN, THE  O’CALLAGHAN, MAXINE  SFWA-P
BOGGART, THE  COOPER, SUSAN  SFWA
BOHACK: SYMBIOTIC WORLDS  LINEHAN, L.  P
BOHR MAKER, THE  NAGATA, LINDA  P
BOK  DE LA REE, GERRY  A
BOLD AS LOVE [ROCK AND ROLL REICH 1]  JONES, GWYNETH  LP
BOLDLY LIVE AS YOU’VE NEVER LIVED BEFORE  RABEN, RICHARD/COHEN, HIYAGUHA  R/REF-H
BOLO [BOLOS]  LAUMER, KEITH  C/P:2,P
BOLO BRIGADE [BOLOS]  KEITH, WILLIAM H. JR.  C/P,P
BOLO RISING [BOLOS]  KEITH, WILLIAM H. JR.  C/P,P
BOLO STRIKE [BOLOS]  KEITH, WILLIAM H. JR.  H
BOMB THAT FAILED, THE  CLARK, RONALD  H
BOMBS FROM THE MURDER WOLVES [@G-8 #4]  HOGAN, ROBERT J.  SR-P
BOMBS IN ORBIT  SUTTON, JEFF  C/P:2,P
BONDS OF ETERNITY, THE [@PERRY RHODAN #69]  ACE  PA
BONDS OF VENGEANCE [WINDS OF THE FORELANDS 3]  COE, DAVID B.  H
BONE DANCE  BULL, EMMA  H:2,SFWA-P:2,P
BONE DOLL’S TWIN, THE [TOBIN OF SKALA 1]  FLEWELLING, LYNN  P
BONE FOREST, THE  HOLDSTOCK, ROBERT P.  P,SFWA-P

BONE GNAWERS [WORLD OF DARKNESS WEREOLF 4]/ STARGAZERS [WORLD OF DARKNESS WEREOLF 4] BRIDGES, BILL/ACHILLI, JUSTIN

PA

BONE WARS  DAVIS, BRETT  C/P,P
BONEHUNTERS, THE [MALAZAN 6]  ERIKSON, STEVEN  LP

The MITSFS Titledex  September 24, 2007  Page 39
TITLE  AUTHOR  CODE
BONES, THE  TEPPER, SHERI S.  P,SFWA-P
BONES OF GOD, THE  LEIGH, STEPHEN  C/P,P
BONES OF HAVEN, THE [HAWK & FISHER 6]  GREEN, SIMON R.  P
BONES OF THE EARTH  SWANWICK, MICHAEL  H,P
BONES OF THE EARTH (PROOF)  SWANWICK, MICHAEL  SR-LP
BONES OF THE MOON  CARROLL, JONATHAN  H:2,P
BONES OF THE PAST [FAIA RISSEDOTTE 2]  LISLE, HOLLY  C/P:2,SFWA-P,P
BONES OF THE WORLD  ROGERS, BRUCE HOLLAND  LPA
BONES OF TIME, THE  GOONAN, KATHLEEN ANN  C/P,P

BONES OF ZORA, THE [KRISHNA] DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

H,P

BONES WIZARD, THE  RYAN, ALAN  H,SFWA
BOOGEYMEN [STAR TREK TNG]  GILDEN, MEL  C/P:2,P
BOOK DRAGON, A  KUSHNER, DONN  C/P:2,P
BOOK OF AIR, THE [@DRAGON QUARTET #4]  KELLOGG, MARJORIE BRADLEY  C/P,P
BOOK OF ANDRE NORTON, THE  NORTON, ANDRE  C/P:2,P
BOOK OF ARTHUR, THE  MATTHEWS, JOHN  H
BOOK OF ATHYRA, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  LP
BOOK OF ATRIX WOLFE, THE  MCKILLIP, PATRICIA A.  P
BOOK OF BEING, THE  WATSON, IAN  P
BOOK OF BRIAN ALDISS, THE  ALDISS, BRIAN W.  C/P:2,P
BOOK OF DEAD DAYS, THE  SEDGWICK, MARCUS  H
BOOK OF DREAMS, THE [DEMON PRINCES 5]  VANCE, JACK  C/P:2,P

BOOK OF DREAMS, THE (VIE) [@VANCE INTEGRAL EDITION #26/ DEMON PRINCES 5] VANCE, JACK

SR-H

BOOK OF DZYAN, THE  MARONEY, TIM  R/REF-H
BOOK OF EARTH, THE [@DRAGON QUARTET #1]  KELLOGG, MARJORIE BRADLEY  C/P,P
BOOK OF EIBON, THE  PRICE, ROBERT M.  LPA
BOOK OF ELLISON, THE  PORTER, ANDREW  R/REF-P

BOOK OF ENCHANTMENT, A ADAMS, KATHLEEN/

ATCHINSON, FRANCES ELIZABETH

H

BOOK OF ENCHANTMENTS  WREDE, PATRICIA C.  H
BOOK OF FANTASY, THE  BORGES, JORGE LUIS  LPA
BOOK OF FIRE, THE [@DRAGON QUARTET #3]  KELLOGG, MARJORIE BRADLEY  C/P:2,P
BOOK OF FLYING, THE  MILLER, KEITH  H
BOOK OF FOURS, THE [@BUFFY]  HOLDER, NANCY  H:2,P
BOOK OF FRANK HERBERT, THE  HERBERT, FRANK  C/P:2,P
BOOK OF FREEDOMS, THE [BOTANY 1,2,3]  MCCAFFREY, ANNE  H
BOOK OF FRITZ LEIBER, THE  LEIBER, FRITZ  C/P:2,P
BOOK OF GORDON DICKSON, THE  DICKSON, GORDON R.  C/P:2,P
BOOK OF GUINEVERE, THE  HOPKINS, ANDREA  H
BOOK OF HYPERBOREA, THE  SMITH, CLARK ASHTON  SR-P
BOOK OF IAN WATSON, THE  WATSON, IAN  H

BOOK OF IMAGINARY BEINGS, THE BORGES, JORGE LUIS/

GUERRERO, MARGARITA

C/P,LP

BOOK OF JHEREG, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H
BOOK OF JOHN BRUNNER, THE  BRUNNER, JOHN  C/P,P
BOOK OF KANTELA, THE [THRONE OF SHERRAN 1]  GREEN, ROLAND J./MURRAY, FRIEDA A.  C/P,SFWA-P,LP
BOOK OF KELLS, THE  MACAVOY, R. A.  C/P:2,P
BOOK OF KINGS, THE  GILLIAM, RICHARD/GREENBERG, MARTIN H. C/PA,PA
BOOK OF LIES, THE  MOLONEY, JAMES  H
BOOK OF LOST TALES 1, THE [HISTORY OF MIDDLE-EARTH 1]  TOLKIEN, J. R. R.  H:2,LP
BOOK OF LOST TALES 2, THE [HISTORY OF MIDDLE-EARTH 2]  TOLKIEN, J. R. R.  H:2,P
BOOK OF LOST THINGS, THE  CONNOLLY, JOHN  H
BOOK OF MERLYN, THE  WHITE, T. H.  C/P,P,H
BOOK OF MORDRED, THE  HANRATTY, PETER  P
BOOK OF MORGAINE, THE [MORGAINE 1,2,3]  CHERRYH, C. J.  H:2
BOOK OF MORMON, THE  MORMON  PA
BOOK OF NIGHT WITH MOON, THE [FELINE WIZARDS 1]  DUANE, DIANE  C/P,LP
BOOK OF PARADOX, THE  COOPER, LOUISE  C/P,P
BOOK OF PHILIP JOSE FARMER, THE  FARMER, PHILIP JOSE  C/P,P
BOOK OF PHILIP K. DICK, THE  DICK, PHILIP K.  C/P,P
BOOK OF POUL ANDERSON, THE  ANDERSON, POUL  P
BOOK OF PTATH, THE  VAN VOGT, A. E.  C/P:2,P,H
BOOK OF RACK THE HEALER, THE  HUGHES, ZACH  C/P:2,P
BOOK OF REVELATIONS, THE  SWIGART, ROB  H
BOOK OF ROBERT E. HOWARD, THE  HOWARD, ROBERT E.  P
BOOK OF SABERHAGEN, THE  SABERHAGEN, FRED  C/P:2,P
BOOK OF SAND, THE  BORGES, JORGE LUIS  LP
BOOK OF SHADOWS  KILLUS, JAMES  C/P,P
BOOK OF SHADOWS  OLDEN, MARC  C/P,P
BOOK OF SHAI, THE  WALTHER, DANIEL  P
BOOK OF SILENCE, THE [LORDS OF DUS 4]  WATT-EVANS, LAWRENCE  C/P:2,P
BOOK OF SKAITH, THE [SKAITH 1,2,3]  BRACKETT, LEIGH  H:2

The MITSFS Titledex  September 24, 2007  Page 40
TITLE  AUTHOR  CODE
BOOK OF SKULLS, THE  SILVERBERG, ROBERT  H:2,P
BOOK OF SORROWS, THE [DUN COW 2]  WANGERIN, WALTER JR.  H
BOOK OF STIER, THE  SANBORN, ROBIN  C/P:2,P
BOOK OF STONES  JAMES, L. DEAN  C/P,P
BOOK OF SUNS, THE  SPRINGER, NANCY  C/P:2,P
BOOK OF TALTOS, THE [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H:2
BOOK OF THE BEAST, THE [PARADYS 2]  LEE, TANITH  H
BOOK OF THE BLACK SUN, THE  THOMAS, G. W.  LP
BOOK OF THE CAULDRON, THE [@HALLOWED ISLE #3]  PAXSON, DIANA L.  LP
BOOK OF THE DAMNED, THE  FORT, CHARLES  [C/RAN-P:2]
BOOK OF THE DAMNED [PARADYS 1]  LEE, TANITH  LP
BOOK OF THE DEAD, THE [PARADYS 3]  LEE, TANITH  H,LP
BOOK OF THE DEAD [@ANGEL]  MCCONNELL, ASHLEY  P
BOOK OF THE DEAD  SKIPP, JOHN/SPECTOR, CRAIG  PA
BOOK OF THE DRAGON, THE  SANT, MONTE  A
BOOK OF THE DUN COW, THE [DUN COW 1]  WANGERIN, WALTER JR.  C/P:2,P
BOOK OF THE MAD, THE [PARADYS 4]  LEE, TANITH  H,LP
BOOK OF THE NEW SUN, THE [NEW SUN 1,2,3,4]  WOLFE, GENE  H
BOOK OF THE NIGHT, THE  LERMAN, RHODA  P
BOOK OF THE RIVER, THE  WATSON, IAN  P
BOOK OF THE SHORT SUN, THE [BOOK OF THE SHORT SUN 1,2,3]  WOLFE, GENE  H:2
BOOK OF THE SPEAR, THE [@HALLOWED ISLE #2]  PAXSON, DIANA L.  LP
BOOK OF THE STARS, THE  WATSON, IAN  P
BOOK OF THE STILL, THE [DOCTOR WHO EIGHTH DOCTOR]  EBBS, PAUL  P
BOOK OF THE STONE, THE [@HALLOWED ISLE #4]  PAXSON, DIANA L.  H,LP

BOOK OF THE SUBGENIUS:

LUNATIC PROPHECIES FOR THE COMING WEIRD TIMES, THE SUBGENIUS FOUNDATION

VLP

BOOK OF THE SWORD, THE [@HALLOWED ISLE #1]  PAXSON, DIANA L.  LP
BOOK OF THE WAR, THE [@FACTION PARADOX 0]  MILES, LAWRENCE  LPA
BOOK OF THREE, THE [PRYDAIN 1]  ALEXANDER, LLOYD  C/P:2,LP,H
BOOK OF ULTIMATE TRUTHS, THE [CORNELIUS MURPHY 1]  RANKIN, ROBERT  P
BOOK OF VAN VOGT, THE  VAN VOGT, A. E.  C/P:2,P
BOOK OF VIRGIL FINLAY, THE  DE LA REE, GERRY  A
BOOK OF WATER, THE [@DRAGON QUARTET #2]  KELLOGG, MARJORIE BRADLEY  C/P,P
BOOK OF WEIRD, THE  BYFIELD, BARBARA NINDE  R/REF-H
BOOK OF WEREWOLVES, THE  BARING-GOULD, SABINE  R/REF-H,C/REF-H
BOOK OF WONDER, THE  DUNSANY, LORD  SR-L,LP
BOOKMAN’S FANTASY, A  LERNER, FRED  R/REF-P
BOOKS OF GREAT ALTA, THE [JENNA 1,2]  YOLEN, JANE  H
BOOKS OF MAGIC  GAIMAN, NEIL  CX
BOOKS OF ROBERT E. HOWARD, THE  HOWARD, ROBERT E.  P
BOOKS OF THE BLACK CURRENT, THE  WATSON, IAN  H:2
BOOKS OF THE GODS, THE, PART ONE [GODS 1,2,3]  SABERHAGEN, FRED  H
BOOKS OF THE KEEPERS, THE [SPELLKEY 3]  DOWNER, ANN  H
BOOSTED MAN, THE [@HOOK #2]  ZETFORD, TULLY  C/P,P

BOOTCAMP 3000 WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA

BORDER, THE  FITCH, MARINA  P
BORDER, BREED NOR BIRTH  REYNOLDS, MACK  D06612
BORDER DISPUTE [INDIW 2]  KERNS, DANIEL R.  P
BORDER WAR [@TRAVELER #6]  DRUMM, D. B.  P
BORDERLAND [BORDERLANDS 1]  WINDLING, TERRI/ARNOLD, MARK ALAN  PA,C/PA
BORDERLANDS  MONTELEONE, THOMAS F.  PA
BORDERLANDS 1 (INCLUDES "THE SCENT OF VINEGAR")  MONTELEONE, THOMAS F.  C/PA,PA
BORDERLANDS 2  MONTELEONE, THOMAS F.  C/PA,PA
BORDERLANDS 3  MONTELEONE, THOMAS F.  C/PA,PA

BORDERLANDS 4 MONTELEONE, ELIZABETH/ MONTELEONE, THOMAS F.

C/PA,PA

BORDERLANDS OF SCIENCE  SHEFFIELD, CHARLES  R/REF-H
BORDERS JUST BEYOND, THE  BRENNAN, JOSEPH PAYNE  SR-H
BORDERS OF INFINITY [MILES VORKOSIGAN 3]  BUJOLD, LOIS MCMASTER  C/P:2,SFWA-P,P
BORDERS OF LIFE, THE  KATHRYNS, G. A.  C/P,P
BORDERTOWN [BORDERLANDS 2]  WINDLING, TERRI/ARNOLD, MARK ALAN  C/PA,PA
BORED OF THE RINGS  HARVARD LAMPOON  C/P:3,SFWA-P,P
BORGEL  PINKWATER, DANIEL  SFWA
BORIS KARLOFF’S FAVORITE HORROR STORIES  KARLOFF, BORIS  PA
BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2003  VALLEJO, BORIS/BELL, JULIE  CAL
BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2004  VALLEJO, BORIS/BELL, JULIE  CAL
BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2005  VALLEJO, BORIS/BELL, JULIE  CAL
BORIS VALLEJO CLASSIC 1997 CALENDAR  VALLEJO, BORIS  CAL
BORIS VALLEJO’S FANTASY CALENDAR 1995  VALLEJO, BORIS  CAL
BORIS VALLEJO’S FANTASY CALENDAR 1996  VALLEJO, BORIS  CAL
BORIS VALLEJO’S FANTASY CALENDAR 1998  VALLEJO, BORIS  CAL
BORIS VALLEJO’S FANTASY CALENDAR 2000  VALLEJO, BORIS  CAL

The MITSFS Titledex  September 24, 2007  Page 41
TITLE  AUTHOR  CODE
BORIS VALLEJO’S FANTASY CALENDAR 2001  VALLEJO, BORIS  CAL
BORIS VALLEJO’S MYTHOLOGY CALENDAR 1989  VALLEJO, BORIS  CAL
BORIS VALLEJO’S MYTHOLOGY CALENDAR 1992  VALLEJO, BORIS  CAL
BORIS VALLEJO’S MYTHOLOGY CALENDAR 1993  VALLEJO, BORIS  CAL
BORIS VALLEJO’S MYTHOLOGY CALENDAR 1994  VALLEJO, BORIS  CAL
BORN IN CAPTIVITY  BERRY, BRYAN  H
BORN IN DEATH [EVE DALLAS]  ROBB, J. D.  H
BORN IN FIRE [RISING STARS 1]  COVER, ARTHUR BYRON  P
BORN LEADER  MCINTOSH, J. T.  H:2
BORN OF BLOOD [WARHAMMER DARKBLADE #1]  ABNETT, DAN  CX
BORN OF LUNA  STATTEN, VARGO  SR-P
BORN TO EXILE  EISENSTEIN, PHYLLIS  C/P,P,H
BORN TO RUN [@SHADOWRUN KELLAN COLT 1]  KENSON, STEPHEN  P
BORN TO RUN [SERRATED EDGE 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,SFWA-P,P
BORN UNDER MARS  BRUNNER, JOHN  C/P:2,P

BORN WITH THE DEAD SILVERBERG, ROBERT C/P,SFWA-TD3,H,P,TD3,

SFWA

BORRIBLES, THE [BORRIBLES 1]  DE LARRABEITI, MICHAEL  C/P,P
BORRIBLES GO FOR BROKE, THE [BORRIBLES 2]  DE LARRABEITI, MICHAEL  P
BORROBIL  DICKINSON, WILLIAM GROFT  P
BORROWED TIDES  LEVINSON, PAUL  H
BORROWED TIME  HRUSKA, ALAN  C/P,P
BOSS IN THE WALL, A TREATISE ON THE HOUSE DEVIL, THE  DAVIDSON, AVRAM/DAVIS, GRANIA  SR-LP
BOSS OF TERROR, THE [DOC SAVAGE #85]  ROBESON, KENNETH  @P
BOTSCHAFT AUS DEM ALL  BRUNNER, JOHN  FP
BOTSCHAFTER VON DEN STERNEN  CLEMENT, HAL  FP
BOUNCING OFF THE MOON [DINGILLIAD 2]  GERROLD, DAVID  H
BOUND BY IRON [@EBERRON INQUISITIVES]  BOLME, EDWARD  P
BOUNDARIES  WRIGHT, T. M.  SFWA-P
BOUNDARY  FLINT, ERIC/SPOOR, RYK E.  H
BOUNTY HUNTER WARS, THE [STAR WARS BOUNTY HUNTER WARS 1,2,3]  JETER, K. W.  H:2
BOW DOWN TO NUL  ALDISS, BRIAN W.  C/P:2,DD443,P
BOWL OF BAAL, THE  BENNET, ROBERT AMES  H
BOWMAN TEST, THE  ELIAS, ALBERT J.  C/P,P
BOX OF DELIGHTS, THE  MASEFIELD, JOHN  P
BOX OF NOTHING, A  DICKINSON, PETER  H,P
BOXEN  LEWIS, C. S.  H
BOY AND HIS TANK, A  FRANKOWSKI, LEO  H,LP
BOY BEYOND THE MOON  ALLUM, TOM  H
BOY FROM THE BURREN, THE [PAINTER 1]  GILLULY, SHEILA  P
BOY INVENTORS’ FLYING SHIP  BONNER, RICHARD  H
BOY WHO COULD FLY, THE  NEWMAN, ROBERT  P
BOY WHO REVERSED HIMSELF, THE  SLEATOR, WILLIAM  P
BOY WHO SAVED EARTH, THE  SLATER, JIM  H,P
BOY WHO SAW TOMORROW, THE  NIALL, IAN  H
BOY WHO WAS THROWN AWAY, THE [SNOW-EYES 2]  SMITH, STEPHANIE A.  C/P,P
BOY WHO WOULD LIVE FOREVER, THE [HEECHEE 6]  POHL, FREDERIK  H
BOYS ARE BACK IN TOWN, THE  GOLDEN, CHRISTOPHER  LP
BOYS ARE BACK IN TOWN, THE (PROOF)  GOLDEN, CHRISTOPHER  SR-LP
BOYS FROM BRAZIL, THE  LEVIN, IRA  H,SFWA,P
BOY’S LIFE  MCCAMMON, ROBERT R.  P
BOYS’ LIFE BOOK OF OUTER SPACE STORIES, THE  BOYS’ LIFE EDITORS  HA
BOYS WHO VANISHED, THE  CARSON, JOHN F.  H
BRADAMANT: THE IRON TEMPEST  MILLER, RON  H,LP
BRADBURY: AN ILLUSTRATED LIFE  WEIST, JERRY  R/REF-H
BRADBURY CHRONICLES, THE  NOLAN, WILLIAM F./GREENBERG, MARTIN H. SFWA:2
BRADBURY CHRONICLES, THE  WELLER, SAM  R/REF-H
BRADBURY SPEAKS  BRADBURY, RAY  R/REF-H
BRADBURY STORIES  BRADBURY, RAY  LP
BRAIDED WORLD, THE [MAXIMUM ICE 2]  KENYON, KAY  P
BRAIN  COOK, ROBIN  H,P
BRAIN 2000  GANN, ERNEST K.  H
BRAIN CHILD  TURNER, GEORGE  SFWA-P,H:2,P
BRAIN MACHINE, THE  SMITH, GEORGE O.  C/P:2,P
BRAIN PALAEO  ROMILUS, ARN  SR-P
BRAIN PLAGUE [ELYSIUM]  SLONCZEWSKI, JOAN  H
BRAIN-STEALERS, THE  LEINSTER, MURRAY  SR-P,DD079
BRAIN TWISTER  GARRETT, RANDALL/JANIFER, LAURENCE M. P
BRAIN TWISTER  PHILLIPS, MARK  C/P:2,P
BRAIN ULTIMATE  CAMPBELL, H. J.  SR-P
BRAIN WAVE  ANDERSON, POUL  C/P:2,P
BRAIN WORLD  REYNOLDS, MACK  C/P,P
BRAINCHILD  SAUL, JOHN  C/P,P
BRAINFIRE  BLACK, CAMPBELL  P
BRAINRACK  PEDLER, KIT/DAVIS, GERRY  P

The MITSFS Titledex  September 24, 2007  Page 42
TITLE  AUTHOR  CODE
BRAINROSE  KRESS, NANCY  C/P,P
BRAINS OF EARTH, THE  VANCE, JACK  DM141
BRAINSTORM  KRAUZER, STEVEN M.  P
BRAINSTORM  MYERS, WALTER DEAN  P
BRAINZ, INC.  GOULART, RON  C/P:2,P
BRAK THE BARBARIAN [BRAK 1]  JAKES, JOHN  P
BRAK THE BARBARIAN (JAPANESE) [BRAK 1]  JAKES, JOHN  FP
BRAK THE BARBARIAN VERSUS THE MARK OF THE DEMONS [BRAK 2]  JAKES, JOHN  C/P:2,P
BRAK THE BARBARIAN VERSUS THE SORCERESS [BRAK 3]  JAKES, JOHN  C/P,P
BRAM STOKER BEDSIDE COMPANION, THE  STOKER, BRAM  H
BRAN MAK MORN [BRAN MAK MORN 1]  HOWARD, ROBERT E.  SR-P,C/P
BRAN MAK MORN: THE LAST KING [BRAN MAK MORN]  HOWARD, ROBERT E.  H
BRANCH, THE  RESNICK, MIKE  P
BRANCH AND CROWN [@WATER! #3]  BAUDINO, GAEL  C/P,P
BRANCH POINT  CLEE, MONA  C/P:2,P
BRAND NEW CHERRY FLAVOR  GRIMSON, TODD  H,LP
BRAND NEW WORLD, A  CUMMINGS, RAY  C/P:2,P
BRAND OF THE WEREWOLF [DOC SAVAGE #5]  ROBESON, KENNETH  C/P:2,@P
BRANDER’S BOOK [WINTER WORLD 4]  MILLS, C. J.  C/P,P
BRANDYJACK  FUNNELL, AUGUSTINE  SR-P
BRASS BOTTLE, THE  ANSTEY, F.  H
BRASS BUTTERFLY, THE  GOLDING, WILLIAM  SR-P
BRASS DRAGON, THE  BRADLEY, MARION ZIMMER  C/P:2,P,D37250
BRASS MAN [POLITY GRIDLINKED 3]  ASHER, NEAL  P
BRASYL  MCDONALD, IAN  H
BRAVE AND THE BOLD, BOOK ONE [@STAR TREK]  DECANDIDO, KEITH R. A.  P
BRAVE AND THE BOLD, BOOK TWO [@STAR TREK]  DECANDIDO, KEITH R. A.  P
BRAVE FREE MEN, THE  VANCE, JACK  C/P:2,P
BRAVE LITTLE TOASTER GOES TO MARS, THE  DISCH, THOMAS M.  H
BRAVE NEW WORDS  PRUCHER, JEFF  R/REF-H
BRAVE NEW WORLD  HUXLEY, ALDOUS  H:2,SR-P,P
BRAVE NEW WORLD/BRAVE NEW WORLD REVISITED  HUXLEY, ALDOUS  LP
BRAVE NEW WORLD REVISITED  HUXLEY, ALDOUS  [C/RAN-H,C/RAN-P:2]
BRAZEN GAMBIT, THE [@DARK SUN ATHAS #1]  ABBEY, LYNN  P
BRAZEN RULE, THE [MATTHEWS 1]  BURGAUER, STEVEN  H:2,LP
BREACH IN THE WATERSHED, A [WATERSHED 1]  NILES, DOUGLAS  H,LP
BREAKAWAY [CASSANDRA KRESNOV 2]  SHEPHERD, JOEL  LP
BREAKAWAY [SPACE: 1999 1]  TUBB, E. C.  C/P:2,P
BREAKAWAY [@STAR TREK TNG STARFLEET ACADEMY #12]  WEISS, BOBBI JG/WEISS, DAVID CODY  H,P

BREAKDOWNS [@STAR TREK S.C.E. #7] CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ DILMORE, KEVIN/JARMAN, HEATHER/ WARD, DAYTON

P

BREAKFAST IN THE RUINS  MALZBERG, BARRY N.  R/REF-H
BREAKFAST IN THE RUINS [KARL GLOGAUER 2]  MOORCOCK, MICHAEL  H,P
BREAKFAST OF CHAMPIONS  VONNEGUT, KURT JR.  H:2,P
BREAKFAST WITH THE ONES YOU LOVE  FINTUSHEL, ELIOT  LP
BREAKING EARTH, THE  LAUMER, KEITH  P,SFWA-P
BREAKING OF NORTHWALL, THE [PELBAR 1]  WILLIAMS, PAUL O.  C/P:2,P
BREAKING OF THE SEALS, THE  ASHTON, FRANCIS  H,LP
BREAKING POINT  GUNN, JAMES E.  C/P,SFWA,P,H
BREAKING STRAIN [@VENUS PRIME #1]  PREUSS, PAUL  C/P:2,P
BREAKTHROUGH  COWPER, RICHARD  C/P:2,P
BREAKTHROUGH [@FLEET #3]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P
BREAKTHROUGH  IGGULDEN, JOHN  SR-P
BREAKTHROUGHS [CONFEDERATE STATES: GREAT WAR 3]  TURTLEDOVE, HARRY  H,P
BREATH OF SNOW AND ASHES, A [OUTLANDER (GABALDON) 6]  GABALDON, DIANA  H
BREATH OF SUSPENSION, THE  JABLOKOV, ALEXANDER  C/P,P
BREATHE  RUSSON, PENNI  H
BREATHE DEEPLY  BASSINGTHWAITE, DON  C/P,P
BREATHING SPACE ONLY  WHITEFORD, WYNNE  SR-P,C/P
BREATHMOSS AND OTHER EXHALATIONS  MACLEOD, IAN R.  H
BRED FOR WAR [@BATTLETECH 16]  STACKPOLE, MICHAEL A.  P
BRED TO KILL  HOFFMAN, LEE  P
BREED TO COME  NORTON, ANDRE  C/P,P,H
BREEDS OF MAN, THE  BUSBY, F. M.  C/P:2,P
BREGA PATH, THE [@SILVER CALL #2/MITHGAR 11]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
BRENNENDEN SCHIFFE, DIE [@PERRY RHODAN ANDROMEDA 1]  ANTON, UWE  FPR
BRENTFORD TRIANGLE, THE [BRENTFORD 2]  RANKIN, ROBERT  P
BRENTWOOD WITCHES, THE  JORDAN, BRENDA  P
BRETTA MARTYN [HENRY MARTYN 2]  SMITH, L. NEIL  H,P
BREWING STORM, THE [@CHARMED]  RUDITIS, PAUL  P
BRIAR KING, THE [KINGDOMS OF THORN AND BONE 1]  KEYES, J. GREGORY  H
BRIAR ROSE  YOLEN, JANE  C/P,P,H
BRIAR’S BOOK [@CIRCLE OF MAGIC #4]  PIERCE, TAMORA  C/P,P

TITLE  AUTHOR  CODE
BRIDAL BEDLAM [@SABRINA THE TEENAGE WITCH #23]  GALLAGHER, DIANA G.  P
BRIDE, THE  MCINTYRE, VONDA N.  P
BRIDE OF FRANKENSTEIN, THE  DREADSTONE, CARL  P
BRIDE OF FU MANCHU, THE  ROHMER, SAX  C/P,P
BRIDE OF STONE, THE  WILLIAMS, THOMAS  LP
BRIDE OF THE CASTLE [CASTLE 8]  DE CHANCIE, JOHN  C/P,P
BRIDE OF THE DARK CASTLE  MOORE, DOREEN  P
BRIDE OF THE DARK CASTLE  MOORE, DORINE  P

BRIDE OF THE FAT WHITE VAMPIRE [FAT WHITE VAMPIRE (JULES DUCHON) 2] FOX, ANDREW

LP

BRIDE OF THE RAT GOD  HAMBLY, BARBARA  C/P:2,P
BRIDE OF THE SLIME MONSTER [CINEVERSE 2]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P:2,P
BRIDE OF THE VAMPYRE [@VARNEY #4]  RYMER, JAMES  LP
BRIDES OF DRACULA, THE  OWEN, DEAN  SR-P
BRIDGE, THE  BANKS, IAIN M.  C/P,P,H
BRIDGE, THE  MANO, D. KEITH  C/P,P,H
BRIDGE, THE  SKIPP, JOHN/SPECTOR, CRAIG  P
BRIDGE, THE  YOUNG, JANINE ELLEN  P
BRIDGE OF ASHES  ZELAZNY, ROGER  C/P:2,P
BRIDGE OF BIRDS [MASTER LI 1]  HUGHART, BARRY  C/P,P,H
BRIDGE OF DAWN, THE [WINDAMEIR CIRCLE 4]  HANCOCK, NIEL  P
BRIDGE OF LIGHT, THE  VERRILL, A. HYATT  SR-H
BRIDGE OF LOST DESIRE, THE  DELANY, SAMUEL R.  C/P,P,H
BRIDGE OF SOULS [QUICKENING 3]  MCINTOSH, FIONA  LP
BRIDGE OF THE SEPARATOR [VIDESSOS]  TURTLEDOVE, HARRY  H
BRIDGE OF TIME, THE  WARNER, WILLIAM HENRY  H
BRIDGE OF VALOR [@CLOAK AND DAGGER #2]  GROELL, ANNE LESLEY  P
BRIDGE OF YEARS, A  WILSON, ROBERT CHARLES  C/P,LP,H
BRIDGE TO YESTERDAY  ARCH, E. L.  H
BRIDGEHEAD  DRAKE, DAVID  C/P:2,SFWA-P,P
BRIDGET LOVES BERNIE  FAIRMAN, PAUL W.  P
BRIDLING CHAOS [JANNA BRILL 1,2,3]  KILLOUGH, LEE  LP
BRIEF LIVES [@SANDMAN #7]  GAIMAN, NEIL  H:2,CX
BRIGANDS OF THE MOON  CUMMINGS, RAY  C/P,P,H
BRIGHT AND SHINING TIGER  EDWARDS, CLAUDIA J.  P
BRIGHT AND THE DARK, THE [FIVE COUNTRIES]  WELCH, MICHELLE M.  C/P,P
BRIGHT ANGEL  BLAIR, JOHN  P
BRIGHT COMPANION, THE  LLEWELLYN, EDWARD  P
BRIGHT ISLANDS IN A DARK SEA  DOUGLAS, L. WARREN  P
BRIGHT MESSENGERS [RAMA]  LEE, GENTRY  C/P,P
BRIGHT NEW UNIVERSE  WILLIAMSON, JACK  C/P:2,P
BRIGHT OF THE SKY [ENTIRE AND ROSE 1]  KENYON, KAY  H
BRIGHT PHOENIX, THE  MEAD, HAROLD  SR-H,C/P:2,P
BRIGHTLY BURNING [VALDEMAR]  LACKEY, MERCEDES  H:2,P
BRIGHTNESS FALLS FROM THE AIR  TIPTREE, JAMES JR.  H:2,SFWA-P,P
BRIGHTNESS REEF [UPLIFT 4]  BRIN, DAVID  H:2,P
BRIGHTONOMICON, THE  RANKIN, ROBERT  LP
BRIGHTSUIT MACBEAR  SMITH, L. NEIL  C/P,P
BRILLIANCE OF THE MOON [@OTORI #3]  HEARN, LIAN  H
BRING BACK YESTERDAY [GRIMES]  CHANDLER, A. BERTRAM  DD517
BRING IT ON [RETRIEVERS]  GILMAN, LAURA ANNE  LP
BRING ME THE HEAD OF PRINCE CHARMING [AZZIE 1]  ZELAZNY, ROGER/SHECKLEY, ROBERT  H:2,LP,P

BRING ON THE BAD GUYS: ORIGINS OF THE MARVEL COMICS VILLAINS [MARVEL COMICS] LEE, STAN

CX

BRING THE JUBILEE  MOORE, WARD  H:2,P
BRINGER OF STORMS [BINDING OF THE BLADE 2]  GRAHAM, L. B.  LP
BRINGERS OF DEATH [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
BRINGING DOWN THE MOON  ANDERSON, JANI  SR-LP
BRINKMAN  GOULART, RON  H,SFWA
BRISINGAMEN  PAXSON, DIANA L.  C/P,P
BRISTLING WOOD, THE [DEVERRY 3/DEVERRY PROPER 3]  KERR, KATHARINE  C/P:2,P
BRITISH SUMMERTIME  CORNELL, PAUL  LP
BRITTLE INNINGS  BISHOP, MICHAEL  H,P
BROCA’S BRAIN  SAGAN, CARL  R/REF-P
BROECKER SAMPLER, THE  BROECKER, RANDY  B
BROKE DOWN ENGINE  GOULART, RON  C/P,P,H
BROKEDOWN PALACE [DRAGAERA FENARIO]  BRUST, STEVEN  C/P:3,P
BROKEN [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  P
BROKEN ANGELS [TAKESHI KOVACS 2]  MORGAN, RICHARD  H,LP
BROKEN BLADE, THE [@DARK SUN ATHAS #3]  HAWKE, SIMON  P
BROKEN BLADE [RUNE BLADE 3]  MARSTON, ANN  P
BROKEN BOW [@STAR TREK ENTERPRISE]  CAREY, DIANE  P
BROKEN CHALICE, THE [BOOK OF MAN 2]  WELCH, JANE  P
BROKEN CITADEL, THE [BROKEN CITADEL 1]  GREGORIAN, JOYCE BALLOU  C/P,P

TITLE  AUTHOR  CODE
BROKEN CLAW [JAMES BOND 23]  GARDNER, JOHN (BRITISH, 1926-)  P
BROKEN COUNTRY, THE [@GAMMALAW 3]  DALEY, BRIAN  C/P,P
BROKEN CRESCENT  SWANN, S. ANDREW  P
BROKEN CROWN, THE [@SUN SWORD #1]  WEST, MICHELLE  P
BROKEN CYCLE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
BROKEN GOD, THE [REQUIEM FOR HOMO SAPIENS 2]  ZINDELL, DAVID  C/P:2,P
BROKEN GODDESS, THE  BEMMANN, HANS  P,SFWA,SFWA-P
BROKEN KINGS, THE [MERLIN CODEX 3]  HOLDSTOCK, ROBERT P.  H
BROKEN LANCE, THE [@WARHAMMER BLACK HEARTS 2]  LONG, NATHAN  P
BROKEN LAND, THE  MCDONALD, IAN  C/P,LP,P
BROKEN LANDS, THE [EMPIRE OF THE EAST 1]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
BROKEN PLACES, THE [ANDROMEDA (GENE RODDENBERRY’S)]  VARE, ETHLIE ANN/MORRIS, DANIEL  H
BROKEN SPHERE, THE [@SPELLJAMMER CLOAKMASTER #5]  FINDLEY, NIGEL D.  P
BROKEN STONE  MONACO, RICHARD  P
BROKEN SUNRISE [@BUFFY WICKED WILLOW #3]  NAVARRO, YVONNE  C/P,P
BROKEN SWORD, THE  ANDERSON, POUL  C/P:2,SFWA-P,P
BROKEN SWORD, THE [FOREVER KING 2]  COCHRAN, MOLLY/MURPHY, WARREN  C/P,P,H
BROKEN SYMMETRIES  PREUSS, PAUL  C/P,P,H
BROKEN SYMMETRIES (PROOF)  PREUSS, PAUL  SR-LP
BROKEN TIME  THOMAS, MAGGY  P
BROKEN WHEEL, THE [@CHUNG KUO #2]  WINGROVE, DAVID  C/P:2,P
BROKEN WORLDS, THE  HARRIS, RAYMOND  P
BROMELIAD TRILOGY, THE [BROMELIAD 1,2,3]  PRATCHETT, TERRY  H
BROMIUS PHENOMENON, THE  RANKINE, JOHN  P
BRONTOMEK!  CONEY, MICHAEL G.  SR-P,C/P
BRONWYN’S BANE [ARGONIA 3]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
BRONZE AXE, THE [@BLADE #1]  LORD, JEFFREY  C/P:2,P
BRONZE GOD OF RHODES, THE  DE CAMP, L. SPRAGUE  C/P,P
BRONZE KING, THE [VALENTINE MARSH 1]  CHARNAS, SUZY MCKEE  H,SFWA
BRONZE OF EDDARTA, THE [GANDALARA 3]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
BROOD OF HELIOS  BERTIN, JACK  H
BROOD OF THE WITCH-QUEEN  ROHMER, SAX  C/P,P
BROOM-ADELAIDE  FREEMAN, BARBARA E.  H
BROTHEL IN ROSENSTRASSE, THE [VON BEK]  MOORCOCK, MICHAEL  LP
BROTHER ASSASSIN [BERSERKER 2]  SABERHAGEN, FRED  C/P:2,P
BROTHER ASTRONOMER  CONSOLMAGNO, GUY  R/REF-H
BROTHER BERSERKER [BERSERKER 2]  SABERHAGEN, FRED  H
BROTHER DEATH [MATADOR 7]  PERRY, STEVE  P,SFWA-P:2
BROTHER ESAU  ORGILL, DOUGLAS/GRIBBIN, JOHN  SFWA-P
BROTHER JOHN  ETHERIDGE, RUTLEDGE  P
BROTHER JONATHAN  KILIAN, CRAWFORD  P
BROTHER NIGHT  KELLEHER, VICTOR  H
BROTHER OF THE DRAGON [@DRAGONLANCE BARBARIANS #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
BROTHER TERMITE  ANTHONY, PATRICIA  C/P,P,H
BROTHER TO DEMONS, BROTHER TO GODS  WILLIAMSON, JACK  C/P,P,H
BROTHER TO DRAGONS  SHEFFIELD, CHARLES  C/P,P
BROTHER TO DRAGONS, COMPANION TO OWLS  LINDSKOLD, JANE  P
BROTHER TO SHADOWS  NORTON, ANDRE  C/P,P,H
BROTHER TO THE LION [LOST LANDS 2]  ESTES, ROSE  P
BROTHERHOOD OF DIABLO, THE [@VISION OF BEASTS #3]  LOVEJOY, JACK  P,SFWA-P
BROTHERHOOD OF THE STARS  GREENE, KIRBY  P
BROTHERHOOD OF THE STRANGE  MACKLIN, JOHN  [C/RAN-P]
BROTHERHOOD OF THE WOLF [RUNELORDS 2]  FARLAND, DAVID  H:2,P
BROTHERS  BOVA, BEN  H,P
BROTHERS IN ARMS [MILES VORKOSIGAN 2]  BUJOLD, LOIS MCMASTER  C/P:2,P
BROTHERS IN ARMS [SCOTT ST. ANDREW 1]  WEAVER, BEN  P
BROTHERS IN ARMS [@DRAGONLANCE RAISTLIN 2]  WEIS, MARGARET/PERRIN, DON  P
BROTHERS MAJERE [DRAGONLANCE PRELUDES 3]  STEIN, KEVIN  C/P:2,P
BROTHERS OF EARTH [HANAN REBELLION 1]  CHERRYH, C. J.  H:2,P
BROTHERS OF THE DRAGON [BROTHERS OF THE DRAGON 1]  BAILEY, ROBIN WAYNE  C/P:2,P
BROTHERS OF THE SNAKE [@WARHAMMER 40K]  ABNETT, DAN  H
BROTHER’S PRICE, A  SPENCER, WEN  P
BROTHERS’ WAR, THE [@MAGIC: THE GATHERING ARTIFACTS 1]  GRUBB, JEFF  P
BROWN GIRL IN THE RING  HOPKINSON, NALO  H,LP
BRRRR!  CONKLIN, GROFF  PA
BRUCE COVILLE’S BOOK OF ALIENS  COVILLE, BRUCE  LPA
BRUCE COVILLE’S BOOK OF MAGIC  COVILLE, BRUCE  LPA
BRUJA [@ANGEL]  ODOM, MEL  P
BRUJO  RELLING, WILLIAM JR.  SFWA-P
BRUTE FORCE [GUARDIANS 7]  AUSTIN, RICHARD  P
BRUTE ORBITS  ZEBROWSKI, GEORGE  P
BUCHANAN CAMPAIGN, THE [SECOND COMMONWEALTH 1]  SHELLEY, RICK  P
BUCK ALICE AND THE ACTOR-ROBOT  KOENIG, WALTER  P
BUCK ROGERS AND THE DOOM COMET  NOWLAN, PHILIP FRANCIS  BLB

TITLE  AUTHOR  CODE
BUCK ROGERS AND THE PLANETOID PLOT  NOWLAN, PHILIP FRANCIS  BLB
BUCK ROGERS IN THE 25TH CENTURY [@BUCK ROGERS #1]  STEELE, ADDISON E.  P
BUCK ROGERS: THE FIRST 60 YEARS IN THE 25TH CENTURY  WILLIAMS, LORRAINE DILLE  CX
BUCKAROO BANZAI  RAUCH, EARL MAC  C/P:2,P
BUDAYEEN NIGHTS [MARID AUDRAN]  EFFINGER, GEORGE ALEC  H
BUDDY HOLLY IS ALIVE AND WELL ON GANYMEDE  DENTON, BRADLEY  C/P,P
BUDRYS’ INFERNO  BUDRYS, ALGIS  C/P:2,P
BUDSPY  DVORKIN, DAVID  H,SFWA
BUFFALO GALS  LE GUIN, URSULA K.  C/P,P
BUFFY THE VAMPIRE SLAYER [BUFFY]  CUSICK, RICHIE TANKERSLEY  P
BUG-EYED MONSTERS  PRONZINI, BILL/MALZBERG, BARRY N.  HA,LP
BUG JACK BARRON  SPINRAD, NORMAN  C/P:2,P
BUG LIFE CHRONICLES, THE  JENNINGS, PHILLIP C.  C/P,P
BUG PARK  HOGAN, JAMES P.  C/P,P,H
BUG WARS, THE  ASPRIN, ROBERT LYNN  H,SFWA-P:2,P
BUGGED!  GLUT, DONALD F.  P
BUGS  ROSZAK, THEODORE  H:2,P
BUILDING HARLEQUIN’S MOON  NIVEN, LARRY/COOPER, BRENDA  H
BULFINCH’S MYTHOLOGY: THE AGE OF FABLE  BULFINCH, THOMAS  R/REF-H
BULL AND THE SPEAR, THE [CORUM 4]  MOORCOCK, MICHAEL  C/P:2,P
BULL CHIEF, THE [@BERSERKER SWIFTAXE 2]  CARLSEN, CHRIS  P
BULL FROM THE SEA, THE  RENAULT, MARY  P
BULL GOD  GELLIS, ROBERTA  P
BULLARD OF THE SPACE PATROL  JAMESON, MALCOLM  H:2
BULLET TIME [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  P
BULPINGTON OF BLUP, THE  WELLS, H. G.  H
BUMPER CROP  LANSDALE, JOE R.  H
BUMSIDER  MACAPP, C. C.  C/P:2,P
BUNDUKI  EDSON, J. T.  C/P,P
BUNKER SOLDIERS [DOCTOR WHO PAST DOCTORS]  DAY, MARTIN  P
BUNNICULA [BUNNICULA 1]  HOWE, DEBORAH/HOWE, JAMES  LP
BUNYIP ARCHIVES, THE  SCHUTTE, JAMES E.  H
BURBANK THE NORTHERNER  VERNE, JULES  H
BURDEN OF PROOF [PAUL SINCLAIR 2]  HEMRY, JOHN G.  P
BUREAU 13 [BUREAU 13 1]  POLLOTTA, NICK  P
BURIED AGE, THE [@STAR TREK TNG]  BENNETT, CHRISTOPHER L.  P
BURIED DEEP [RETRIEVAL ARTIST 4]  RUSCH, KRISTINE KATHRYN  C/P,P
BURIED PYRAMID, THE  LINDSKOLD, JANE  P
BURIED WORLD, THE  DAY, LIONEL  B
BURN  KELLY, JAMES PATRICK  H
BURN  RANSOM, BILL  C/P,P
BURN, WITCH, BURN  MERRITT, A.  SR-LP
BURNDIVE  LOWACHEE, KARIN  C/P,P
BURNING  CHAMBERS, JANE  P
BURNING, THE  FAIN, JEFF  P
BURNING, THE  GUNN, JAMES E.  C/P:2,P
BURNING, THE [@BUFFY ANGEL UNSEEN 1]  HOLDER, NANCY/MARIOTTE, JEFF  P
BURNING, THE  MASTERTON, GRAHAM  SFWA
BURNING, THE [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
BURNING BRIDGES [RETRIEVERS]  GILMAN, LAURA ANNE  LP
BURNING BRIGHT [@SHADOWRUN 15]  DOWD, TOM  P
BURNING BRIGHT  RUSSELL, JAY  LP
BURNING BRIGHT  SCOTT, MELISSA  H:2,P
BURNING CHROME  GIBSON, WILLIAM  C/P,SFWA-P,P,H,SFWA
BURNING CITY, THE [MAGIC 4]  NIVEN, LARRY/POURNELLE, JERRY  H:2,P
BURNING COURT, THE  CARR, JOHN DICKSON  P
BURNING CROWN, THE [SERPENT’S EGG 2]  MCCURDY, J. FITSGERALD  LP
BURNING DREAMS [STAR TREK]  BONANNO, MARGARET WANDER  P
BURNING EYE, THE [@WAR WORLD #1]  POURNELLE, JERRY  PA,C/PA,SFWA-P
BURNING GODDESS, THE [@SHADOW WORLD 1]  HAMMELL, IAN  P
BURNING GROUND, THE [SHADOWSONG 2]  CLAYTON, JO  C/P:2,P
BURNING HEART [DOCTOR WHO MISSING ADVENTURES]  STONE, DAVE  C/P,P
BURNING HEART OF NIGHT, THE  CAT, IVAN  P
BURNING LAND, THE [@FIREBRATS #1]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
BURNING LAND, THE  STRAUSS, VICTORIA  P
BURNING LAND, THE (PROOF)  STRAUSS, VICTORIA  SR-LP
BURNING MOUNTAIN, THE  COPPEL, ALFRED  H
BURNING REALM, THE  REAVES, MICHAEL  C/P:2,P
BURNING RING, THE  BURDEKIN, KATHARINE  H
BURNING SHORE, THE [@WARHAMMER]  EARL, ROBERT  P
BURNING STONE, THE [CROWN OF STARS 3]  ELLIOTT, KATE  C/P,P,H
BURNING STONE, THE [MAGES OF GARILLON 1]  HARRIS, DEBORAH TURNER  C/P,SFWA,SFWA-P,P
BURNING TEARS OF SASSURUM [NAPHAR 3]  BAKER, SHARON  P
BURNING THE ICE  MIXON, LAURA J.  H

The MITSFS Titledex  September 24, 2007  Page 46
TITLE  AUTHOR  CODE
BURNING TIME, THE [@SHADOWRUN 40]  KENSON, STEPHEN  P
BURNING TOWER [MAGIC 5]  NIVEN, LARRY/POURNELLE, JERRY  H:2
BURNING VOID  MAGROON, VECTOR  SR-P
BURNING WATER  LACKEY, MERCEDES  C/P:2,P
BURNING WORLD, THE  BALLARD, J. G.  C/P:2,P
BURNT LANDS, THE  ELLIOTT, RICHARD  P
BURNT OFFERINGS [ANITA BLAKE 7]  HAMILTON, LAURELL K.  C/P:2,P
BURR IN THE GARDEN OF EDEN, THE [MACNAIR OF MACNAIR 2]  PEIRCE, HAYFORD  LP
BURROWERS BENEATH, THE  LUMLEY, BRIAN  C/P,P
BURSTER  CAPOBIANCO, MICHAEL  P
BURY HIM DARKLY  BLACKBURN, JOHN  H
BURYING THE SHADOW  CONSTANTINE, STORM  LP
BUSHIDO INCIDENT, THE  CRAWFORD, BETTY ANNE  C/P:2,P
BUSINESS OF FERRETS, A  HILGARTNER, BETH  LP
BUSINESS UNUSUAL [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  C/P:2,P
BUSINESSMAN, THE  DISCH, THOMAS M.  H,SFWA
BUSY BODIES  MARTIN, ED  SR-L
BUT GENTLY DAY  NATHAN, ROBERT  H
BUT WE ARE NOT OF EARTH  KARL, JEAN E.  P
BUT WHAT OF EARTH?  ANTHONY, PIERS  C/P:2,SFWA-P,P
BUT WHAT OF EARTH?  ANTHONY, PIERS/COULSON, ROBERT  SR-P
BUTCHER’S BILL, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
BUTTERFLY AND HELLFLOWER [BUTTERFLY ST. CYR 1,2,3]  BES SHAHAR, ELUKI  H
BUTTERFLY KID, THE  ANDERSON, CHESTER  C/P:2,P
BUY JUPITER AND OTHER STORIES  ASIMOV, ISAAC  H:2,P,SFWA
BUYING TIME  HALDEMAN, JOE  H:2,P
BWANA  RESNICK, MIKE  TD33
BWANA/BULLY!  RESNICK, MIKE  C/P:3,P
BY AIR EXPRESS TO VENUS  ROCKWOOD, ROY  H
BY ANY OTHER FAME  RESNICK, MIKE/GREENBERG, MARTIN H.  C/PA,PA
BY ANY OTHER NAME  ROBINSON, SPIDER  P
BY BALLOON TO THE SAHARA  TERMAN, DOUGLAS  P
BY BLOOD ALONE [LEGION OF THE DAMNED 3]  DIETZ, WILLIAM C.  C/P,P
BY CHAOS CURSED [BIFROST GUARDIANS 5]  REICHERT, MICKEY ZUCKER  P
BY DAYBREAK THE EAGLE  DORER, NANCY/DORER, FRANCES  P
BY FORCE OF ARMS [LEGION OF THE DAMNED 4]  DIETZ, WILLIAM C.  P
BY FURIES POSSESSED  WHITE, TED  C/P:2,P
BY HONOR BETRAY’D [MAGEWORLDS 3]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
BY SLANDEROUS TONGUES  LACKEY, MERCEDES/GELLIS, ROBERTA  H
BY SPACE SHIP TO SATURN  ROCKWOOD, ROY  H
BY SPACE SHIP TO THE MOON  COGGINS, JACK/PRATT, FLETCHER  [C/RAN-H]

BY TEMPTATIONS AND BY WAR [@MECHWARRIOR DARK AGE/ BATTLETECH] COLEMAN, LOREN L.

LP

BY TEMPTATIONS AND BY WAR (PROOF) [@MECHWARRIOR DARK AGE/ BATTLETECH] COLEMAN, LOREN L.

SR-LP

BY THE BOOK [STAR TREK ENTERPRISE] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

BY THE LIGHT OF THE GREEN STAR [GREEN STAR 3]  CARTER, LIN  C/P:2,P
BY THE LIGHT OF THE MOON  KOONTZ, DEAN R.  H
BY THE SWORD [MAGIC OF THE PLAINS #1]  COSTIKYAN, GREG  C/P,@H:2,P
BY THE SWORD [KEROWYN’S TALE]  LACKEY, MERCEDES  C/P:2,P
BY THE TIME I GET TO NASHVILLE [DIEGO 2]  FENN, LIONEL  P
BYPASS TO OTHERNESS  KUTTNER, HENRY  C/P:2,P
BYTE BEAUTIFUL  TIPTREE, JAMES JR.  H
BYWORLDER, THE  ANDERSON, POUL  C/P:2,SFWA-P,P
BYZANTIUM  LAWHEAD, STEPHEN R.  H,P
BYZANTIUM! [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH  P
BYZANTIUM’S CROWN [HEIRS TO BYZANTIUM 1]  SHWARTZ, SUSAN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 47
TITLE  AUTHOR  CODE
CABAL  BARKER, CLIVE  H
CABAL, THE [CABAL 1]  DUNN, PHILIP  P
CABINET OF LIGHT, THE [DOCTOR WHO/TIME HUNTER 0]  O’MAHONY, DANIEL  SR-H
CABU  RUSSELL, JOHN ROBERT  C/P:2,P
CACHALOT [HUMANX]  FOSTER, ALAN DEAN  H:2,P
CACHE, THE  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
CACHE FROM OUTER SPACE  FARMER, PHILIP JOSE  DF165
CACOPHONY (CASSETTE) [STAR TREK]  MOLLOY, J. J.  SR-MM
CADAVER OF GIDEON WYCK, THE  LAING, ALEXANDER  H
CADET KIRK [@STAR TREK STARFLEET ACADEMY #3]  CAREY, DIANE  LP
CADMIAN’S CHOICE [COREAN 5]  MODESITT, L. E. JR.  H
CADRE LUCIFER [CADRE 2]  O’RIORDAN, ROBERT  C/P,SFWA-P,P
CADRE MESSIAH [CADRE 3]  O’RIORDAN, ROBERT  C/P:2,P
CADRE ONE [CADRE 1]  O’RIORDAN, ROBERT  C/P:2,P
C.A.D.S. [@C.A.D.S. #1]  SIEVERT, JOHN  P
CADUCEUS WILD  MOORE, WARD/BRADFORD, ROBERT  C/P,P
CAESAR DIES  MUNDY, TALBOT  C/P,P
CAESAR’S BICYCLE [@TIMELINE WARS #3]  BARNES, JOHN  C/P,P
CAFE PURGATORIUM  TOR BOOKS  SR-HA
CAGE, THE [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,SFWA-P:2,P
CAGE A MAN [DEMU 1]  BUSBY, F. M.  H:2,P
CAGEBIRD  LOWACHEE, KARIN  C/P,P
CAHENA  WELLMAN, MANLY WADE  H,SFWA
CAILLOUX DANS LE CIEL  ASIMOV, ISAAC  FP
CAIN FACTOR, THE  LEWIS, CHARLES  H
CAIN’S LAND [VERESHCHAGIN 3]  FREZZA, ROBERT  P
CALAMITY’S CHILD [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  SR-LP
CALCULATED MAGIC, A [LOGICAL MAGICIAN 2]  WEINBERG, ROBERT  C/P,P
CALCULATING GOD  SAWYER, ROBERT J.  H,P
CALCULUS OF ANGELS, A [AGE OF UNREASON 2]  KEYES, J. GREGORY  H:2,LP
CALDE OF THE LONG SUN [LONG SUN 3]  WOLFE, GENE  C/P,P,H
CALEDON OF THE MISTS [CALEDON 1]  HARRIS, DEBORAH TURNER  C/P,SFWA-P,P
CALE’S STORY [@TITAN A.E.]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
CALIBAN [CALIBAN 1]  ALLEN, ROGER MACBRIDE  H,SFWA:2,P
CALIBAN COVE [@RESIDENT EVIL #2]  PERRY, S. D.  P
CALIBAN LANDING  POPKES, STEVEN  H,P
CALIBAN’S HOUR  WILLIAMS, TAD  C/P,P,H
CALIBRATED ALLIGATOR AND OTHER SCIENCE FICTION STORIES, THE  SILVERBERG, ROBERT  H
CALIFIA’S DAUGHTERS  RICHARDS, LEIGH  P
CALIFORNIA COVEN PROJECT, THE  STICKGOLD, BOB  P
CALIFORNIA ICEBERG, THE  HARRISON, HARRY  H
CALIFORNIA SORCERY  NOLAN, WILLIAM F./SCHAFER, WILLIAM  C/PA,PA
CALIFORNIA VOODOO GAME, THE [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  C/P,P,H
CALIFORNIOS, THE  L’AMOUR, LOUIS  P
CALIX STAY [@CIRCLE OF LIGHT #3]  HANCOCK, NIEL  C/P,P
CALL FOR AN EXORCIST  BIXBY, JEROME  SR-L
CALL FROM A DISTANT SHORE  BURNS, STEPHEN L.  P

CALL OF CHAOS, THE [WARHAMMER] GASCOIGNE, MARC/JONES, ANDY/

DUNN, CHRISTIAN

CX

CALL OF DUTY [BATTLETECH]  PARDOE, BLAINE LEE  P
CALL OF EARTH, THE [HOMECOMING 2]  CARD, ORSON SCOTT  C/P,P,H
CALL OF MADNESS [CAITHAN CRUSADE 1]  SMITH, JULIE DEAN  C/P,P
CALL TO ARMS... [@STAR TREK DOMINION WAR #2/STAR TREK DS9]  CAREY, DIANE  C/P,P
CALL TO ARMS, A [MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
CALL TO ARMS, A [@DAMNED #1]  FOSTER, ALAN DEAN  C/P,P,H
CALL TO ARMS [STAR COMMANDOS 9]  GRIFFIN, P. M.  C/P,P
CALL TO ARMS, A [TWO-EDGED SWORD 3]  MARTIN, THOMAS K.  P
CALL TO ARMS, A [@BABYLON 5]  SHECKLEY, ROBERT  C/P,P
CALL TO BATTLE [WORLD OF DARKNESS RAGE JAY NO-NAME 1]  MURRAY, DOUG  P
CALL TO BATTLE! [@THERE WILL BE WAR #7]  POURNELLE, JERRY  PA,SFWA-P
CALL TO DARKNESS, A [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
CALLAHAN TOUCH, THE [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,P,H
CALLAHAN’S CON [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H
CALLAHAN’S CROSSTIME SALOON [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
CALLAHAN’S KEY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H,P
CALLAHAN’S LADY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,SFWA-P:2,P,H,SFWA
CALLAHAN’S LEGACY [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  H,P
CALLAHAN’S SECRET [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
CALLER FROM ETERNITY [@PERRY RHODAN #106]  ACE  PA
CALLER OF THE BLACK, THE  LUMLEY, BRIAN  SR-H
CALLING, THE [@IMMORTALS]  ASHLEY, JENNIFER  P
CALLING CAPTAIN FUTURE  HAMILTON, EDMOND  C/P,P
CALLING DOCTOR KILL! [@ENFORCER #2]  SUGAR, ANDREW  P
CALLING DR. PATCHWORK  GOULART, RON  P

The MITSFS Titledex  September 24, 2007  Page 48
TITLE  AUTHOR  CODE
CALLING OF BARA, THE  SULLIVAN, SHEILA  C/P,P
CALLING OF THE THREE, THE [NIGHT-THREADS 1]  EMERSON, RU  SFWA-P:2
CALLING ON DRAGONS [ENCHANTED FOREST 3]  WREDE, PATRICIA C.  P
CALLING STATION E-A-R-T-H  WILLIAMS, GURNEY III  LP
CALLIPYGIA [TERRA MAGICA 4]  CARTER, LIN  P
CALLY’S WAR  RINGO, JOHN/COCHRANE, JULIE  H
CALTRAPS OF TIME, THE  MASSON, DAVID I.  SR-P
CALVIN COOLIDGE HOME FOR DEAD COMEDIANS, THE  DENTON, BRADLEY  LP
CALY  COMBES, SHARON  P
CAMBER OF CULDI [DERYNI CAMBER 1]  KURTZ, KATHERINE  C/P,P,H
CAMBER THE HERETIC [DERYNI CAMBER 3]  KURTZ, KATHERINE  C/P:2,P,H
CAMELOT 30K  FORWARD, ROBERT L.  C/P:2,P
CAMELOT CAPER, THE  PETERS, ELIZABETH  SFWA-P
CAMELOT CHRONICLES, THE  ASHLEY, MIKE  LPA

CAMELOT FANTASTIC SCHIMEL, LAWRENCE/

GREENBERG, MARTIN H.

PA

CAMELOT IN ORBIT [CAMELOT 2]  LANDIS, ARTHUR H.  C/P,P
CAMERA OBSCURA [DOCTOR WHO EIGHTH DOCTOR]  ROSE, LLOYD  P
CAMINO A CASA, UN  STURGEON, THEODORE  FP
CAMOUFLAGE  HALDEMAN, JOE  H
CAMP CONCENTRATION  DISCH, THOMAS M.  C/P,P,H
CAMP OF THE SAINTS, THE  RASPAIL, JEAN  P
CAMPAIGN [DOCTOR WHO PAST DOCTORS]  MORTIMORE, JIM  SR-P
CAMPBELL WOOD  SARRANTONIO, AL  H,SFWA
CAMPFIRE COLLECTION, THE  HYAMS, GINA  LPA
CAMPO ARCHIMEDE  DISCH, THOMAS M.  FP
CAN YOU FEEL ANYTHING WHEN I DO THIS?  SHECKLEY, ROBERT  H:2,SFWA,P
CANBY’S LEGION  BALDWIN, BILL  C/P:2,P
CANCRO 2000  JONES, GONNER  FP
CANDLE  BARNES, JOHN  H,P
CANDLE, THE  HOPKINS, LINTON C.  SR-L
CANDLE FOR D’ARTAGNAN, A [SAINT-GERMAIN OLIVIA 3]  YARBRO, CHELSEA QUINN  H:2,LP
CANDLE IN A BOTTLE  GILMAN, CAROLYN IVES  LP
CANDLE IN THE SUN  WELLS, ROBERT  C/P:2,P
CANDLE OF DISTANT EARTH, THE  FOSTER, ALAN DEAN  H
CANDY MAN  KING, VINCENT  C/P:2,P
CANNON’S ORB  DOUGLAS, L. WARREN  P
CANOPY OF TIME, THE  ALDISS, BRIAN W.  C/P,P
CANTICLE [@FORGOTTEN REALMS CLERIC #1]  SALVATORE, R. A.  P
CANTICLE FOR LEIBOWITZ, A [LEIBOWITZ 1]  MILLER, WALTER M. JR.  H:2,LP
CANYONS  CACEK, P. D.  H
CAPACITY [RECURSION 2]  BALLANTYNE, TONY  P
CAPELLA’S GOLDEN EYES  EVANS, CHRISTOPHER  P
CAPITAL OFFENSE [@MERCHANT PRINCE #3]  SHIMERMAN, ARMIN  P
CAPITOL  CARD, ORSON SCOTT  C/P:2,LP
CAPITOL HILL CLONES  TABLER, JOSEPH  P
CAPO’S REVENGE [@CYBERNARC #4]  CAIN, ROBERT  P
CAPRICORN GAMES  SILVERBERG, ROBERT  H
CAPRICORN ONE  GOULART, RON  C/P,P
CAPTAIN, THE [@TELNARIAN 2]  NORMAN, JOHN  C/P:2,P
CAPTAIN [DIRIGENT MERCENARY CORPS 3]  SHELLEY, RICK  C/P,P
CAPTAIN EMPIRICAL  NICHOLSON, SAM  C/P,P
CAPTAIN FUTURE AND THE SPACE EMPEROR  HAMILTON, EDMOND  P
CAPTAIN FUTURE’S CHALLENGE  HAMILTON, EDMOND  C/P,P
CAPTAIN JACK ZODIAC  KANDEL, MICHAEL  C/P:2,P
CAPTAIN MIDNIGHT AND SHEIK JOMAK KHAN  BERKE, HELEN  BLB
CAPTAIN NEMO  ANDERSON, KEVIN J.  H,P
CAPTAIN NICE  JOHNSTON, WILLIAM  P

CAPTAIN’S BLOOD [STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

H

CAPTAIN’S DAUGHTER, THE [STAR TREK]  DAVID, PETER  C/P:2,P

CAPTAIN’S GLORY [STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

H

CAPTAIN’S HONOR, THE [STAR TREK TNG]  DVORKIN, DAVID/DVORKIN, DANIEL  C/P:2,P
CAPTAINS OUTRAGEOUS [CAPTAINS OUTRAGEOUS 1]  YOUNG, ROY V.  P

CAPTAIN’S PERIL [@STAR TREK] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

CAPTIVE, THE  STALLMAN, ROBERT  C/P:2,P
CAPTIVE OF GOR [GOR 7]  NORMAN, JOHN  C/P:2,P
CAPTIVE ON THE FLYING SAUCERS  FINN, RALPH L.  SR-P
CAPTIVE SCORPIO [DRAY PRESCOT 17]  AKERS, ALAN BURT  P

TITLE  AUTHOR  CODE
CAPTIVE SOUL, THE [@HIGHLANDER]  SHERMAN, JOSEPHA  P
CAPTIVE UNIVERSE  HARRISON, HARRY  C/P,P,H
CAPTIVES IN SPACE [@DIG ALLEN #2]  GREENE, JOSEPH  P,@H
CAPTIVES OF BLUE MOUNTAIN [@COMPLETE ELFQUEST #3]  PINI, WENDY/PINI, RICHARD  CX
CAPTIVES OF RAAG, THE [LONE WOLF 14]  DEVER, JOE  P
CAPTIVES OF THE FLAME  DELANY, SAMUEL R.  DF199
CAPTIVES OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 3]  LORRAH, JEAN  C/P,P
CAPTIVITY [OUTLANDER (SALTERBERG) 1]  SALTERBERG, B. J.  LP
CAPTURE THE FLAG [@STAR TREK TNG STARFLEET ACADEMY #4]  VORNHOLT, JOHN  LP
CAR, THE  SHRYACK, DENNIS/BUTLER, MICHAEL  P

CAR SINISTER SILVERBERG, ROBERT/

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

PA

CARAVAN  GOLDIN, STEPHEN  SR-P
CARAVAN OF SHADOWS [WORLD OF DARKNESS WRAITH]  BYERS, RICHARD LEE  P
CARBON-COPY KILLER  BROWNE, HOWARD  GRD15
CARD MASTER [FANTASY ADVENTURES]  EMERY, CLAYTON  P
CARD SHARKS [@WILD CARDS NEW CYCLE 1]  MARTIN, GEORGE R. R.  PA,C/PA,SFWA-P

CARDASSIA: LOTUS FLOWER, THE [STAR TREK DS9 WORLDS 1]/ ANDOR: PARADIGM [STAR TREK DS9 WORLDS 1] MCCORMACK, UNA/JARMAN, HEATHER

C/P,P

CARDER’S PARADISE  LEVENE, MALCOLM  H
CARDINAL CRIMSON [@NECROMUNDA]  MCDERMOTT, WILL  P
CARDINAL OF THE KREMLIN, THE [JACK RYAN]  CLANCY, TOM  P
CARDS OF GRIEF  YOLEN, JANE  C/P,P
CAREFULLY CONSIDERED RAPE OF THE WORLD, THE  MEAD, SHEPHERD  P
CARESS OF TWILIGHT, A [MEREDITH GENTRY 2]  HAMILTON, LAURELL K.  H,P
CARETAKER [STAR TREK VOYAGER]  GRAF, L. A.  C/P,P
CARLUCCI’S EDGE [FRANK CARLUCCI 2]  RUSSO, RICHARD PAUL  P
CARLUCCI’S HEART [FRANK CARLUCCI 3]  RUSSO, RICHARD PAUL  P
CARMEN DOG  EMSHWILLER, CAROL  LP
CARMEN MIRANDA’S GHOST IS HAUNTING SPACE STATION THREE  SAKERS, DON  PA,C/PA,SFWA-P
CARMICHAEL’S DOG  KOSTER, R. M.  H
CARMILLA/HAUNTED BARONET, THE  LE FANU, J. SHERIDAN  P
CARNACKI THE GHOST-FINDER  HODGSON, WILLIAM HOPE  SR-P,C/P
CARNADYNE HORDE, THE [@SPACEWAYS #17]  CLEVE, JOHN  P
CARNELIAN CUBE, THE  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  C/P:2,P
CARNELIAN THRONE, THE [SILISTRA 4]  MORRIS, JANET  C/P,P
CARNIVAL  BEAR, ELIZABETH  P
CARNIVAL OF DESTRUCTION, THE [DAVID LYDYARD 3]  STABLEFORD, BRIAN M.  P
CARNIVAL OF FEAR [RAVENLOFT]  KING, J. ROBERT  P
CARNIVAL OF SOULS [@BUFFY]  HOLDER, NANCY  P
CARNIVORES OF LIGHT AND DARKNESS [JOURNEYS OF THE CATECHIST 1] FOSTER, ALAN DEAN  H
CARNOSAUR  KNIGHT, HARRY ADAM  P
CAROL, THE PURSUED [@PHENWICK #29]  KIMBROUGH, KATHERYN  P
CAROLINE  NORTON, ANDRE/CUSHING, ENID  P
CARPATHIAN CASTLE  VERNE, JULES  H,P
CARPE DEMON  KENNER, JULIE  LP
CARPE DIEM [LIADEN UNIVERSE 3]  LEE, SHARON/MILLER, STEVE  C/P,P
CARPE JUGULUM [DISCWORLD 23/DISCWORLD WITCHES 6]  PRATCHETT, TERRY  H:2,P
CARPET MAKERS, THE  ESCHBACH, ANDREAS  H
CARRIE  KING, STEPHEN  C/P:2,P
CARRION COMFORT  SIMMONS, DAN  C/P,P
CARRYING THE FIRE  COLLINS, MICHAEL (ASTRONAUT)  R/REF-P
CARSON OF VENUS [VENUS #3]  BURROUGHS, EDGAR RICE  C/P:2,@P
CART AND CWIDDER [DALEMARK 1]  JONES, DIANA WYNNE  C/P:2,LP
CARTHAGE ASCENDANT [@ASH #2]  GENTLE, MARY  C/P,P
CARTOMANCY [AGE OF DISCOVERY 2]  STACKPOLE, MICHAEL A.  LP
CARTOON CRIMES, THE [@AVENGER #31]  ROBESON, KENNETH  C/P,P

CARTOON HISTORY OF THE UNIVERSE 1:

THE EVOLUTION OF EVERYTHING, THE GONICK, LARRY

SR-VLP

CARTOON HISTORY OF THE UNIVERSE 2: STICKS AND STONES, THE  GONICK, LARRY  SR-VLP
CARTOON HISTORY OF THE UNIVERSE 3: RIVER REALMS, THE  GONICK, LARRY  SR-VLP

CARTOON HISTORY OF THE UNIVERSE 4:

PART OF THE OLD TESTAMENT, THE GONICK, LARRY

SR-VLP

CARTOON HISTORY OF THE UNIVERSE 5: BRAINS AND BRONZE, THE  GONICK, LARRY  SR-VLP

CARTOON HISTORY OF THE UNIVERSE 6:

WHO ARE THESE ATHENIANS?, THE GONICK, LARRY

SR-VLP

CARVE THE SKY  JABLOKOV, ALEXANDER  C/P,P,H
CASCA THE DEFIANT [CASCA]  DENGELEGI, PAUL  P
CASCA THE LIBERATOR [CASCA]  DENGELEGI, PAUL  P
CASCADE POINT  ZAHN, TIMOTHY  TD2,SFWA-TD2
CASCADE POINT (AND OTHER STORIES)  ZAHN, TIMOTHY  C/P,P,H
CASE AGAINST TOMORROW, THE  POHL, FREDERIK  C/P:2,P
CASE AND THE DREAMER  STURGEON, THEODORE  H:2,P

TITLE  AUTHOR  CODE
CASE OF CHARLES DEXTER WARD, THE  LOVECRAFT, H. P.  C/P:2,P
CASE OF CONSCIENCE, A  BLISH, JAMES  H:2,P
CASE OF THE COLONIST’S CORPSE, THE [@STAR TREK SAMUEL COGLEY]  INGERSOLL, BOB/ISABELLA, TONY  P
CASE OF THE LITTLE GREEN MEN, THE  REYNOLDS, MACK  H
CASE OF THE SOLID KEY, THE  BOUCHER, ANTHONY  P
CASE OF THE TOXIC SPELL DUMP, THE  TURTLEDOVE, HARRY  C/P:2,SFWA-P,P
CASE OF THE VANISHING BOY, THE  KEY, ALEXANDER  C/P,P
CASEBOOK OF JULES DE GRANDIN, THE  QUINN, SEABURY  P
CASEBOOK OF LUCIUS LEFFING, THE  BRENNAN, JOSEPH PAYNE  SR-H
CASEBOOK OF THE BLACK WIDOWERS  ASIMOV, ISAAC  C/P:2,P
CASEY AGONISTES AND OTHER SCIENCE FICTION AND FANTASY STORIES  MCKENNA, RICHARD  H,P
CASINO ROYALE [JAMES BOND 1]  FLEMING, IAN  C/P,P
CASSANDRA COMPLEX, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H
CASSANDRA CROSSING, THE  KATZ, ROBERT  P
CASSANDRA RISING  LAURANCE, ALICE  HA,SFWA
CASSILEE [LIVING PLANET 2]  COON, SUSAN  C/P,P
CASSINI DIVISION, THE [FALL REVOLUTION]  MACLEOD, KEN  H,P
CASSIOPEIA AFFAIR, THE  ZERWICK, CHLOE/BROWN, HARRISON  P
CAST A COLD EYE  RYAN, ALAN  P
CAST DOWN THE STARS  JONES, MCCLURE  H
CAST IN COURTLIGHT [ELANTRA 2]  SAGARA, MICHELLE  LP
CAST IN SECRET [ELANTRA 3]  SAGARA, MICHELLE  LP
CAST IN SHADOW [ELANTRA 1]  SAGARA, MICHELLE  P
CAST OF CORBIES, A [BARDIC CHOICES]  LACKEY, MERCEDES/SHERMAN, JOSEPHA  C/P:2,P
CAST OF FATE [DRAGON DICE]  VARNEY, ALLEN  SR-P
CASTAWAY FROM SPACE  BRACK, VEKTIS  SR-P
CASTAWAYS IN LILLIPUT  WINTERFELD, HENRY  H
CASTAWAYS IN TIME [@CASTAWAYS IN TIME #1]  ADAMS, ROBERT  C/P,LP
CASTAWAYS OF TANAGAR, THE  STABLEFORD, BRIAN M.  C/P,P
CASTAWAYS’ WORLD  BRUNNER, JOHN  DF242
CASTELLO D’ACCIAIO, IL  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  FP
CASTING FORTUNE [LIAVEK 6]  FORD, JOHN M.  H,SFWA-P
CASTING SHADOWS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #1]  CAVELOS, JEANNE  C/P,P
CASTLE DEATH [LONE WOLF 7]  DEVER, JOE  P
CASTLE DREAMS [CASTLE 6]  DE CHANCIE, JOHN  C/P:2,P
CASTLE FANTASTIC  DE CHANCIE, JOHN/GREENBERG, MARTIN H. PA
CASTLE FOR RENT [CASTLE 2]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
CASTLE HOHENFELS  MOORE, DORINE  P
CASTLE IN THE AIR [CASTLE 2]  JONES, DIANA WYNNE  H
CASTLE IN THE ATTIC, THE  WINTHROP, ELIZABETH  LP
CASTLE KEEPS, THE  OFFUTT, ANDREW J.  C/P,P
CASTLE KIDNAPPED [CASTLE 3]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
CASTLE MURDERS [CASTLE 5]  DE CHANCIE, JOHN  C/P:2,P
CASTLE OF CROSSED DESTINIES, THE  CALVINO, ITALO  LP
CASTLE OF DAYS  WOLFE, GENE  LP
CASTLE OF DECEPTION  FITCH, ED  LP
CASTLE OF DECEPTION [BARD’S TALE 1]  LACKEY, MERCEDES/SHERMAN, JOSEPHA  C/P:2,SFWA-P,P
CASTLE OF HAPE, THE [YNELL 3]  MURPHY, SHIRLEY ROUSSEAU  H,P
CASTLE OF IRON, THE  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  C/P,P,H
CASTLE OF LLYR, THE [PRYDAIN 3]  ALEXANDER, LLOYD  C/P:3,LP
CASTLE OF OTRANTO, THE  WALPOLE, HORACE  P
CASTLE OF THE OTTER, THE  WOLFE, GENE  R/REF-H,C/REF-H
CASTLE OF THE SILVER WHEEL, THE [CELYDONN GWENLLIANT 1]  EDGERTON, TERESA  SFWA-P:2
CASTLE OF WIZARDRY [BELGARIAD 4]  EDDINGS, DAVID  C/P:3,P
CASTLE PERILOUS [CASTLE 1]  DE CHANCIE, JOHN  C/P:2,P
CASTLE ROOGNA [@XANTH #3]  ANTHONY, PIERS  C/P:2,P
CASTLE ROOGNA (PROOF) [XANTH 3]  ANTHONY, PIERS  SR-LP
CASTLE SPELLBOUND [CASTLE 7]  DE CHANCIE, JOHN  C/P,P
CASTLE TERROR  BRADLEY, MARION ZIMMER  C/P,P
CASTLE WAR! [CASTLE 4]  DE CHANCIE, JOHN  C/P:2,SFWA-P,P
CASTLEDOWN [BROKEN CITADEL 2]  GREGORIAN, JOYCE BALLOU  P
CASTLES MADE OF SAND [ROCK AND ROLL REICH 2]  JONES, GWYNETH  LP
CASTLEVIEW  WOLFE, GENE  H,SFWA-P,P
CASTROVALVA [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,C/P
CASUALTIES OF WAR [DOCTOR WHO EIGHTH DOCTOR]  EMMERSON, STEVE  C/P,P
CAT-A-LYST  FOSTER, ALAN DEAN  C/P,SFWA-P,P,H
CAT AND THE CURATE  GILSON, CHARLES  H
CAT BY THE TAIL [@SABRINA THE TEENAGE WITCH SALEM #7]  VERNEY, SARAH J.  P
CAT CRIMES 2  GREENBERG, MARTIN H./GORMAN, ED  PA
CAT FROM OUTER SPACE, THE  KEY, TED  C/P:2,P
CAT HOUSE  PEAK, MICHAEL  P
CAT KARINA  CONEY, MICHAEL G.  C/P:2,P
CAT MAGIC  STRIEBER, WHITLEY  SFWA-P
CAT O’ FIVE TAILS  WHIP, LEATHER  WRD

TITLE  AUTHOR  CODE
CAT OF SILVERY HUE, A [HORSECLANS 4]  ADAMS, ROBERT  C/P:2,P
CAT ON THE EDGE [JOE GREY 1]  MURPHY, SHIRLEY ROUSSEAU  P
CAT RAISE THE DEAD [JOE GREY 3]  MURPHY, SHIRLEY ROUSSEAU  P
CAT SCRATCH FEVER  HARPER, TARA K.  C/P:2,P
CAT THAT WASN’T THERE, THE  WHITELAW, STELLA  P
CAT TV [@SABRINA THE TEENAGE WITCH SALEM #1]  DUBOWSKI, MARK  P
CAT UNDER FIRE [JOE GREY 2]  MURPHY, SHIRLEY ROUSSEAU  C/P,P
CAT WHO WALKED THROUGH TIME, THE [DOCTOR WHO]  KELLY, TOM/KELLY, ALRYSSA  SR-VLP
CAT WHO WALKS THROUGH WALLS, THE  HEINLEIN, ROBERT A.  C/P,P,H
CAT WHO WENT TO HEAVEN, THE  COATSWORTH, ELIZABETH  P
CATACLISMO SOLAR  BALLARD, J. G.  FP
CATACLYSM  PENDLETON, DON  C/P:2,P
CATACLYSM, THE  SHERRIFF, R. C.  SR-P
CATACLYSM  STATTEN, VARGO  SR-P
CATACLYSM, THE [@DRAGONLANCE TALES II #2]  TSR  C/PA,PA
CATACOMB YEARS  BISHOP, MICHAEL  H,SFWA,P
CATACOMBS  FARRIS, JOHN  SFWA-P
CATALINA  MAUGHAM, W. SOMERSET  P
CATALOG OF FANTASTIC THINGS  CARELMAN, JACQUES  [C/RAN-H]
CATALOGO GENERALE DELLA FANTASCIENZA  BERTONI, ALFIO/MISSIAJA, GIANLUIGI  R/IND
CATALYST, THE  HARNESS, CHARLES L.  C/P,P
CATALYST  HOFFMAN, NINA KIRIKI  LP
CATALYST  KENT, BRAD  SR-P
CATALYST, THE  STATTEN, VARGO  SR-P
CATALYST OF SORROWS [@STAR TREK LOST ERA (2360 AD)]  BONANNO, MARGARET WANDER  P
CATALYST, THE (PROOF)  HARNESS, CHARLES L.  SR-LP
CATAMOUNT  PEAK, MICHAEL  P
CATARACT  HARPER, TARA K.  C/P,P
CATASTROPHE PLANET  LAUMER, KEITH  C/P:2,P
CATASTROPHEA [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  C/P,P
CATASTROPHE’S SPELL [DANCE OF GODS 1]  BRENNER, MAYER ALAN  C/P,P
CATCH A FALLING STAR  BRUNNER, JOHN  C/P:2,P
CATCH THE LIGHTNING [SKOLIAN EMPIRE 2]  ASARO, CATHERINE  C/P:2,P
CATCH THE STAR WINDS  CHANDLER, A. BERTRAM  C/P:2,P
CATCH TRAP, THE  BRADLEY, MARION ZIMMER  P
CATCHING MIDNIGHT  HOLLY, EMMA  H
CATCHWORLD  BOYCE, CHRIS  H,SFWA,P
CATERPILLARIAN CHRONICLES, THE  PEMBROKE, LYNNE  P
CATERPILLAR’S QUESTION, THE  ANTHONY, PIERS/FARMER, PHILIP JOSE  H,SFWA:2,P
CATFANTASTIC  NORTON, ANDRE/GREENBERG, MARTIN H.  C/PA,PA
CATFANTASTIC II  NORTON, ANDRE/GREENBERG, MARTIN H.  C/PA,PA
CATFANTASTIC III  NORTON, ANDRE/GREENBERG, MARTIN H.  C/PA,PA
CATFANTASTIC IV  NORTON, ANDRE/GREENBERG, MARTIN H.  C/PA,PA
CATFANTASTIC V  NORTON, ANDRE/GREENBERG, MARTIN H.  PA
CATHARSIS CENTRAL  ALBAN, ANTONY  C/P,P
CATHEDRAL [@STAR TREK DS9 MISSION GAMMA #3]  MARTIN, MICHAEL A./MANGELS, ANDY  C/P,P
CATHEDRAL OF THORNS [ELIXIR OF LIFE 2]  FRANKOS, STEVEN  P
CATHEDRAL OPTION, THE  MONTANA, RON  P
CATHERDRAL OF VAMPIRES  MITCHELL, MARY ANN  P
CATHOUSE [MAN-KZIN WARS]  ING, DEAN  C/P:2,SFWA-P,P
CATIVEIRO HUMANO  WUL, STEFAN  FP
CATS, THE  PHIPSON, JOAN  P
CAT’S CLAW [CHARLIE MOON 4]  YARBRO, CHELSEA QUINN  P
CAT’S CRADLE  VONNEGUT, KURT JR.  H:2,P
CAT’S CRADLE: TIME’S CRUCIBLE [DOCTOR WHO NEW ADVENTURES]  PLATT, MARC  C/P:2,P
CAT’S CRADLE: WARHEAD [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
CAT’S CRADLE: WITCH MARK [DOCTOR WHO NEW ADVENTURES]  HUNT, ANDREW  C/P:2,P
CAT’S EYE  ATWOOD, MARGARET  LP
CAT’S GAMBIT [ORIANI 2]  GADALLAH, LESLIE  C/P,P
CATS HAVE NO LORD  SHETTERLY, WILL  C/P:2,P
CATS IN CYBERSPACE  HILGARTNER, BETH  LP
CATS IN SPACE AND OTHER PLACES  FAWCETT, BILL  PA,C/PA,SFWA-P
CATS OF SEROSTER, THE  WESTALL, ROBERT  H
CAT’S PAJAMAS, THE  BRADBURY, RAY  H
CAT’S PAJAMAS & OTHER STORIES, THE  MORROW, JAMES  H
CAT’S PAW  TAYLOR, L. A.  C/P,P
CAT’S PAWN [ORIANI 1]  GADALLAH, LESLIE  P
CATSEYE [SHATTERED LIGHT 3]  FORSTCHEN, WILLIAM R./DEMAREST, JAKI  P
CATSEYE [DIPPLE 1]  NORTON, ANDRE  C/P,P,H
CATSPAW [CAT 2]  VINGE, JOAN D.  C/P:2,P
CATSWOLD PORTAL, THE  MURPHY, SHIRLEY ROUSSEAU  C/P,SFWA:2,P,H
CATTLE MUTILATORS, THE  DALTON, JOHN J.  P
CATWINGS  LE GUIN, URSULA K.  LP
CATWOMAN  ABBEY, LYNN/ASPRIN, ROBERT LYNN  P

The MITSFS Titledex  September 24, 2007  Page 52
TITLE  AUTHOR  CODE
CAUGHT IN CRYSTAL [LYRA 1]  WREDE, PATRICIA C.  C/P:2,SFWA-P,P
CAUGHT IN THE CROSSFIRE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P
CAULDRON OF FEAR, THE [LONE WOLF 9]  DEVER, JOE  P
CAULDRON OF INIQUITY [@CLOAK AND DAGGER #3]  GROELL, ANNE LESLEY  P
CAULDRON OF PLENTY, THE [DANANS 2]  TAYLOR, KEITH  C/P,P
CAUSE OF DEATH, THE [BUREAU OF SPECIAL INVESTIGATIONS]  ALLEN, ROGER MACBRIDE  P
CAUTION! INFLAMMABLE!  SCORTIA, THOMAS N.  H,SFWA,P
CAUTIONARY TALES  YARBRO, CHELSEA QUINN  C/P,SFWA,P
CAVALCADE OF COMEDY  KRONENBERGER, LOUIS  HA
CAVE AND THE ROCK, THE  FAURE, RAOUL C.  H
CAVE GIRL, THE  BURROUGHS, EDGAR RICE  H:2,SR-P,P
CAVE OF SPEARS, THE  DEL REY, LESTER  H:2
CAVE OF STARS  ZEBROWSKI, GEORGE  P
CAVE OF TIME, THE  PACKARD, EDWARD  C/P,P
CAVERN OF BLACK ICE, A [SWORD OF SHADOWS 1]  JONES, J. V.  H,P
CAVERN OF DOOM, THE [ZORK 3]  MERETZKY, S. ERIC  C/P,P
CAVERN OF THE DAMNED  STEELE, CURTIS  SR-LPA
CAVERNICOLES DE WOLF, LES  BARBET, PIERRE  FP
CAVERNS [@JOURNEYS OF MCGILL FEIGHAN #1]  O’DONNELL, KEVIN JR.  C/P:2,P
CAVERNS OF KALTE, THE [LONE WOLF 3]  DEVER, JOE/CHALK, GARY  C/P,P
CAVERNS OF SOCRATES  MCKIERNAN, DENNIS L.  P
CAVES OF ANDROZANI, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
CAVES OF BUDA, THE  CUTTER, LEAH R.  P
CAVES OF FEAR, THE [@RICK BRANT #8]  BLAINE, JOHN  H
CAVES OF FIRE AND ICE [YNELL 4]  MURPHY, SHIRLEY ROUSSEAU  C/P,P
CAVES OF ICE [@WARHAMMER 40K CIAPHAS CAIN 2]  MITCHELL, SANDY  P
CAVES OF KARST, THE  HOFFMAN, LEE  C/P,P
CAVES OF KLYDOR, THE [COLSEC 2]  HILL, DOUGLAS  C/P,P,H
CAVES OF MADNESS, THE [BALZAN 2]  MOORE, WALLACE  P
CAVES OF MARS, THE  PETAJA, EMIL  DM133
CAVES OF MARS (PROOF) [FALCON GOLD 1]  WOLFF, ROBERT SHERMAN  SR-LP
CAVES OF NIGHT, THE  CHRISTOPHER, JOHN  P
CAVES OF REGLATHIUM, THE [@DANNUS #3]  SIROTA, MIKE  C/P,P
CAVES OF STEEL, THE [LIJE BALEY 1]  ASIMOV, ISAAC  C/P:2,P,H
CAVES OF THE DRUUFS [@PERRY RHODAN #72]  ACE  C/PA,PA
CAVIAR  STURGEON, THEODORE  SR-H,C/P:2,P
CD OF AMAZINGLY NEAT STUFF! (CD)  NEO-OPSIS SCIENCE FICTION MAGAZINE  SR-MM
CEE-TEE MAN  MORGAN, DAN  SR-P
CELEBRATED CASES OF JUDGE DEE  VAN GULIK, ROBERT  LP
CELEBRITY VAMPIRES  GREENBERG, MARTIN H.  PA
CELERY STALKS AT MIDNIGHT, THE [BUNNICULA 3]  HOWE, JAMES  LP
CELESTIAL BLUEPRINT AND OTHER STORIES, THE  FARMER, PHILIP JOSE  DF165
CELESTIAL CHESS  BONTLY, THOMAS  P
CELESTIAL DEBRIS  WATT-EVANS, LAWRENCE  LP
CELESTIAL HIT LIST [@SAND WARS #3]  INGRID, CHARLES  C/P,P
CELESTIAL MATTERS  GARFINKLE, RICHARD  H,LP
CELESTIAL OMNIBUS AND OTHER STORIES, THE  FORSTER, E. M.  C/P,P
CELESTIAL PURSUITS IN THE HUB OF THE UNIVERSE  CARBERRY, C. A.  LP
CELESTIAL STEAM LOCOMOTIVE, THE [SONG OF EARTH 1]  CONEY, MICHAEL G.  H:2,SFWA
CELESTIAL TOYMAKER, THE [DOCTOR WHO]  DAVIS, GERRY/BINGEMAN, ALISON  SR-P
CELESTIS  PARK, PAUL  LP
CELL, THE  CASE, DAVID  H
CELLAR, THE  LAYMON, RICHARD  P
CELLULOID GANGS  TOLNAY, TOM  SFWA
CELTIC FAIRY TALES  JACOBS, JOSEPH  LP
CELTIC MYTHS AND LEGENDS  ROLLESTON, T. W.  R/REF-H
CELTIKA [MERLIN CODEX 1]  HOLDSTOCK, ROBERT P.  P
CEMETERY WORLD  SIMAK, CLIFFORD D.  H:2,P
CENOTAPH ROAD [CENOTAPH ROAD 1]  VARDEMAN, ROBERT E.  P
CENTAUR AISLE [@XANTH #4]  ANTHONY, PIERS  C/P,P,H
CENTAUR IN THE GARDEN, THE  SCLIAR, MOACYR  LP
CENTAURI DAWN [@ALPHA CENTAURI #1]  ELY, MICHAEL  P
CENTAURI DEVICE, THE  HARRISON, M. JOHN  H:2,SFWA:2,SR-P
CENTAURI EXPRESS 1 #3 (CASSETTE)  CENTAURI COMPANY  SR-MM
CENTAURUS  HARTWELL, DAVID G./BRODERICK, DAMIEN HA

CENTER CANNOT HOLD, THE [CONFEDERATE STATES AMERICAN EMPIRE 2] TURTLEDOVE, HARRY

H

CENTER OF THE CIRCLE, THE [SERVANTS OF ARK 2]  WYLIE, JONATHAN  C/P,P
CENTERFORCE  WATERS, T. A.  P
CENTRAL HEAT  DVORKIN, DAVID  C/P:2,P
CENTRAL PASSAGE  SCHOONOVER, LAWRENCE  H:2,P
CENTRIFUGAL RICKSHAW DANCER, THE  WATKINS, WILLIAM JOHN  P
CENTURION’S EMPIRE, THE  MCMULLEN, SEAN  P
CENTURY OF FANTASY: 1980-1989, A  SILVERBERG, ROBERT  HA

TITLE  AUTHOR  CODE
CENTURY OF GREAT SHORT SCIENCE FICTION NOVELS, A  KNIGHT, DAMON  HA,PA
CENTURY OF PROGRESS, A  SABERHAGEN, FRED  C/P:2,SFWA-P,P
CENTURY OF SCIENCE FICTION, A  KNIGHT, DAMON  C/PA,PA
CENTURY OF SCIENCE FICTION: 1950-1959, A  SILVERBERG, ROBERT  HA
CENTURY OF THE MANIKIN  TUBB, E. C.  C/P,P
CENTURY RAIN [REVELATION SPACE]  REYNOLDS, ALASTAIR  H:2
CENTURY’S END  GRIFFIN, RUSSELL M.  P
C’ERA UNA VOLTA L’AMERICA  BARRETT, NEAL JR.  FP
CERBERUS: A WOLF IN THE FOLD [@FOUR LORDS OF THE DIAMOND #2]  CHALKER, JACK L.  C/P:2,P
CEREMONIES, THE  KLEIN, T. E. D.  P
CEREMONY [@DARKWAR #3]  COOK, GLEN  P
CERES SOLUTION, THE  SHAW, BOB  C/P,P
CERES STORM  HERTER, DAVID  H,P
CERULEAN SINS [ANITA BLAKE 11]  HAMILTON, LAURELL K.  C/P,H
CERULEAN STORM, THE [@DARK SUN PRISM #5]  DENNING, TROY  C/P,P
CERVEAUX SOUS CONTROLE  RICHARD-BESSIERE, F.  FP
CESTUS DECEPTION, THE [@STAR WARS CLONE WARS]  BARNES, STEVEN  P
CESTUS DEI  ROBERTS, JOHN MADDOX  P
CETAGANDA [MILES VORKOSIGAN 6]  BUJOLD, LOIS MCMASTER  C/P,P,H
CETTE LUEUR QUI VENAIT DES TENEBRES  RICHARD-BESSIERE, F.  FP
CHAIN OF ATTACK [STAR TREK]  DEWEESE, GENE  C/P:3,P
CHAIN OF CHANCE, THE  LEM, STANISLAW  C/P:2,P,H
CHAIN OF COMMAND [JEFFERSON’S WAR]  RANDLE, KEVIN  P
CHAINER’S TORMENT [@MAGIC: THE GATHERING ODYSSEY 2]  MCGOUGH, SCOTT  P
CHAINFIRE [SWORD OF TRUTH 9]  GOODKIND, TERRY  H
CHAINING THE LADY [CLUSTER #2]  ANTHONY, PIERS  C/P:2,@P
CHAINMAIL [STAR TREK GATEWAYS #2/STAR TREK CHALLENGER]  CAREY, DIANE  P
CHAINS OF COMMAND [STAR TREK TNG]  MCCAY, BILL/FLOOD, ELOISE  C/P:3,P
CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  SAGARA, MICHELLE  P
CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  WEST, MICHELLE  LP
CHAINS OF DESTRUCTION [CHAINS 2]  ROSEN, SELINA  LP
CHAINS OF FREEDOM [CHAINS 1]  ROSEN, SELINA  LP
CHAINS OF GOLD  SPRINGER, NANCY  H:2,SFWA,P
CHAINS OF LIGHT  THOMAS, QUENTIN  P,SFWA-P
CHAINS OF THE SEA  SILVERBERG, ROBERT  HA,PA
CHAINS THAT YOU REFUSE, THE  BEAR, ELIZABETH  LP
CHALCHIUHITE DRAGON, THE  MORRIS, KENNETH  H
CHALET BOUGY-VILLARS  MARVIN, SUSAN  P
CHALET DIABOLIQUE [LUCIFER COVE 5]  COFFMAN, VIRGINIA  P
CHALICE, THE [@SWORD,RING AND CHALICE 3]  CHESTER, DEBORAH  P
CHALICE AND THE BLADE, THE  MCREYNOLDS, GLENNA  P
CHALK GIANTS, THE  ROBERTS, KEITH  C/P,P
CHALLENGE  BULMER, H. K.  SR-P
CHALLENGE MET [@SAND WARS #6]  INGRID, CHARLES  C/P:2,P
CHALLENGE OF CLIMATE, THE  SILVERBERG, ROBERT  [C/RAN-H]
CHALLENGE OF THE CLANS [FINN MACCUMHAL 1]  FLINT, KENNETH C.  C/P,P
CHALLENGE OF THE SPACESHIP, THE  CLARKE, ARTHUR C.  [C/RAN-P:2]
CHALLENGE OF THE UNKNOWN [@PERRY RHODAN #32]  ACE  C/PA,PA
CHALLENGE THE HELLMAKER  RICHMOND, WALT/RICHMOND, LEIGH  C/P:2,P
CHALLENGE TO DRACULA [@DRACULA #9]  LORY, ROBERT  P
CHALLENGER [@STAR TREK NEW EARTH #6]  CAREY, DIANE  C/P,P
CHALLENGER’S HOPE [NICHOLAS SEAFORT 2]  FEINTUCH, DAVID  C/P:2,P
CHALLENGERS OF THE UNKNOWN  GOULART, RON  P
CHALLENGES  BOVA, BEN  H,P
CHALLENGES [@BLENDING #3]  GREEN, SHARON  C/P,P
CHAMBER OF HORRORS, A  HADFIELD, JOHN  HA
CHAMELEON CORPS AND OTHER SHAPE CHANGERS, THE  GOULART, RON  H:2,P
CHAMIEL  PEARSON, EDWARD  C/P,P
CHAMPION OF GARATHORM, THE [CASTLE BRASS 2]  MOORCOCK, MICHAEL  C/P,P
CHAMPION OF THE GODS [@BLADE #21]  LORD, JEFFREY  P
CHAMPION OF THE LAST BATTLE [HORSECLANS 11]  ADAMS, ROBERT  C/P:2,P
CHAMPIONS OF THE FORCE [@STAR WARS JEDI ACADEMY 3]  ANDERSON, KEVIN J.  C/P:2,P
CHAMPIONS OF THE SIDHE [SIDHE 2]  FLINT, KENNETH C.  C/P:2,P
CHAMPION’S TRIAL [@MAGIC: THE GATHERING LEGENDS TWO 3]  MCGOUGH, SCOTT  P
CHANCE AND OTHER GESTURES OF THE HAND OF FATE  SPRINGER, NANCY  C/P:2,SFWA-P,P
CHANCE CHILD, A  WALSH, JILL PATON  P
CHANCELLOR, THE  VERNE, JULES  H
CHANGE, THE  FOSTER, GEORGE C.  SR-P
CHANGE, THE [@ALIEN NATION #4]  LONGYEAR, BARRY B.  C/P:2,P
CHANGE  MAXWELL, ANN  C/P,P
CHANGE OF COMMAND [HERIS SERRANO 6]  MOON, ELIZABETH  H:2,P
CHANGE SONG  HOFFMAN, LEE  H,SFWA
CHANGE THE SKY AND OTHER STORIES  SAINT CLAIR, MARGARET  C/P,P
CHANGED MAN, THE  CARD, ORSON SCOTT  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 54
TITLE  AUTHOR  CODE
CHANGELING [OTHERWORLD]  GALENORN, YASMINE  P
CHANGELING [@SHADOWRUN #5]  KUBASIK, CHRIS  C/P,SFWA-P,P
CHANGELING [LIADEN UNIVERSE 0/LIADEN ADVENTURES 6]  LEE, SHARON/MILLER, STEVE  SR-LP
CHANGELING [ROBOTS AND ALIENS 1]  LEIGH, STEPHEN  C/P:2,P
CHANGELING, THE [@FEY 2]  RUSCH, KRISTINE KATHRYN  C/P,P
CHANGELING  SHERMAN, DELIA  H
CHANGELING, THE  VAN VOGT, A. E.  C/P:2,P
CHANGELING [POL DETSON 1]  ZELAZNY, ROGER  C/P,H,LP
CHANGELING EARTH [EMPIRE OF THE EAST 3]  SABERHAGEN, FRED  C/P:2,P
CHANGELING GARDEN, THE  ELZE, WINIFRED  H
CHANGELING PLACES [@CHARMED]  OSTOW, MICOL  P
CHANGELING PLAGUE, THE  MITCHELL, SYNE  P
CHANGELING PRINCE, THE  VELDE, VIVIAN VANDE  C/P,P
CHANGELING SEA, THE  MCKILLIP, PATRICIA A.  C/P:2,P
CHANGELING WAR, THE [@CHANGELING #1]  GARRISON, PETER  C/P,LP
CHANGELING WORLDS, THE  BULMER, KENNETH  DD369

CHANGELINGS [TWINS OF PETAYBEE 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2

CHANGEOVER  JONES, DIANA WYNNE  LP
CHANGEOVER, THE  MAHY, MARGARET  C/P,P
CHANGER [ATHANOR 1]  LINDSKOLD, JANE  P
CHANGER OF DAYS [ANGHARA 2]  ALEXANDER, ALMA  P
CHANGER OF WORLDS [HONOR HARRINGTON/WORLDS OF HONOR 3]  WEBER, DAVID  HA
CHANGER’S MOON [DUEL OF SORCERY 3]  CLAYTON, JO  C/P:2,P
CHANGES  BISHOP, MICHAEL/WATSON, IAN  PA
CHANGES, THE [CHANGES 1,2,3]  DICKINSON, PETER  H,LP
CHANGES  HOWARTH, MATT  CX
CHANGESPELL [DUN LADY’S JESS 2]  DURGIN, DORANNA  P
CHANGESPELL LEGACY [DUN LADY’S JESS 3]  DURGIN, DORANNA  P
CHANGEWAR  LEIBER, FRITZ  P
CHANGEWEAVER  BALL, MARGARET  C/P:2,SFWA-P,P
CHANGEWINDS, THE [CHANGEWINDS 1,2,3]  CHALKER, JACK L.  LP
CHANGING FATE  WATERS, ELISABETH  P
CHANGING LAND, THE  ZELAZNY, ROGER  C/P,P,H
CHANGING MAZE, THE  SNYDER, ZILPHA KEATLEY  SFWA
CHANGING PLANES  LE GUIN, URSULA K.  H
CHANGING THE PAST  BERGER, THOMAS  H
CHANGING VISION [@WEB SHIFTERS #2]  CZERNEDA, JULIE E.  C/P:2,P
CHANGING WOMEN [@WOMAN OF THE IRON PEOPLE #2]  ARNASON, ELEANOR  P,SFWA-P
CHANNELING CLEOPATRA  SCARBOROUGH, ELIZABETH ANN  H
CHANNEL’S DESTINY [SIME/GEN 5]  LORRAH, JEAN/LICHTENBERG, JACQUELINE C/P:2,SFWA,P
CHANTRY GUILD, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,SFWA-P,P,H
CHANUR’S HOMECOMING [CHANUR 4]  CHERRYH, C. J.  H:2,P
CHANUR’S LEGACY [CHANUR 5]  CHERRYH, C. J.  C/P,SFWA,P,H:2
CHANUR’S VENTURE [CHANUR 2]  CHERRYH, C. J.  H:2,P
CHAOS  BELL, THORNTON  SR-P
CHAOS AND AMBER [ROGER ZELAZNY’S DAWN OF AMBER 2]  BETANCOURT, JOHN GREGORY  H
CHAOS AND ORDER [GAP 4]  DONALDSON, STEPHEN R.  H:2,P
CHAOS BALANCE, THE [RECLUCE 7]  MODESITT, L. E. JR.  H:2,P
CHAOS BLEEDS [@BUFFY]  MOORE, JAMES A.  P
CHAOS CHILD [WARHAMMER INQUISITION WAR 3]  WATSON, IAN  P
CHAOS COME AGAIN  BAIRD, WILHELMINA  P
CHAOS CURSE, THE [@FORGOTTEN REALMS CLERIC #5]  SALVATORE, R. A.  C/P,P
CHAOS FIGHTERS, THE  WILLIAMS, ROBERT MOORE  C/P:2,P
CHAOS GATE, THE [BARD’S TALE 5]  SHERMAN, JOSEPHA  P
CHAOS IN LAGRANGIA  REYNOLDS, MACK  C/P,P
CHAOS KID, THE [@PSI-MAN #4]  DAVID, PETER  C/P,P
CHAOS KING, THE [GEORGIE AND BUG 2]  RUBY, LAURA  SR-LP
CHAOS MODE [MODE 3]  ANTHONY, PIERS  C/P,P,H
CHAOS WEAPON, THE  KAPP, COLIN  C/P:2,P
CHAPAYECA  EDMONDSON, G. C.  H,SFWA
CHAPTER WAR [@WARHAMMER 40K]  COUNTER, BEN  P
CHAPTERHOUSE: DUNE [DUNE 6]  HERBERT, FRANK  C/P,SFWA-P,P,H
CHARIOT INTO TIME  ZEIGFREID, KARL  SR-P
CHARIOT OF FIRE  HALES, E. E. Y.  P
CHARIOT TO THE STARS  MILLER, STEVE  SR-LP
CHARIOTS OF FIRE  PARRY, MICHEL  P
CHARIOTS OF RA, THE [KEYS TO THE DIMENSIONS 7]  BULMER, KENNETH  D10293
CHARIOTS OF THE GODS?  VON DANIKEN, ERICH  [C/RAN-P:2]
CHARISMA  BARNES, STEVEN  H,P
CHARISMA  CONEY, MICHAEL G.  SR-P,C/P
CHARLATAN, THE [RELAMS OF ARKANIA]  KIESOW, ULRICH  P
CHARLEMAGNE’S CHAMPION [ROLAND 1]  VAN ASTEN, GAIL  C/P,P
CHARLES FORT NEVER MENTIONED WOMBATS  DEWEESE, GENE/COULSON, ROBERT  H,SFWA

The MITSFS Titledex  September 24, 2007  Page 55
TITLE  AUTHOR  CODE
CHARLES KEEPING’S BOOK OF CLASSIC GHOST STORIES  KEEPING, CHARLES  HA
CHARLIEUNCLENORFOLKTANGO  WHITE, TONY  LP
CHARMED AGAIN [@CHARMED]  LENHARD, ELIZABETH  P
CHARMED DEATH, A [@BEWITCHING MYSTERY 2]  ALT, MADELYN  P
CHARMED DESTINIES  SILHOUETTE BOOKS  PA
CHARMED LIFE [CHRESTOMANCI 1]  JONES, DIANA WYNNE  C/P:2,P
CHARMED SPHERE, THE  ASARO, CATHERINE  LP
CHARNEL HOUSE  MASTERTON, GRAHAM  SFWA-P
CHARNEL PRINCE, THE [KINGDOMS OF THORN AND BONE 2]  KEYES, J. GREGORY  H
CHARON: A DRAGON AT THE GATE [@FOUR LORDS OF THE DIAMOND #3]  CHALKER, JACK L.  C/P:2,P
CHARON’S ARK  GAUGER, RICK  C/P,P
SAINT AND THE LAST HERO, THE  CHARTERIS, LESLIE  SR-P
CHARWOMAN’S SHADOW, THE  DUNSANY, LORD  C/P,P,H
CHASE INTO SPACE [@CYBORG COMMANDO #2]  O’NEILL, PAMELA/MOHAN, KIM  P
CHASE THE MORNING [STEPHEN FISHER 1]  ROHAN, MICHAEL SCOTT  P,SFWA
CHASM CITY [REVELATION SPACE]  REYNOLDS, ALASTAIR  H
CHASM OF DOOM, THE [LONE WOLF 4]  DEVER, JOE  P
CHASSE A L’IMPONDERABLE, LA  LE MAY, J./LE MAY, D.  FP
CHATTERING GODS  CRAWLEY, RAYBURN  H
CHAUVINISTO  MERWIN, SAM JR.  P
CHEAP COMPLEX DEVICES  SUNDMAN, JOHN COMPTON  HA,LPA
CHEAP THRILLS  GOULART, RON  R/REF-H
CHECKLIST OF FANTASTIC LITERATURE, THE  BLEILER, EVERETT F.  SRL/IND
CHECKLIST OF GERMAN "DIME-NOVELS"  JOHNSON, KENNETH R.  R/IND
CHECKLIST OF KIM STANLEY ROBINSON, A  JOYCE, TOM/STEPHENS, CHRISTOPHER P.  R/IND

CHECKLIST OF SCIENCE-FICTION AND SUPERNATURAL FICTION 1800–1948, THE BLEILER, EVERETT F.

R/IND

CHECKLIST OF SCIENCE FICTION ANTHOLOGIES, A  COLE, W. R.  SRL/IND
CHECKLIST TO FANTASTIC UNIVERSE  WOOD, EDWARD  SRL/IND
CHECKMATE  BAKER, ERIC T.  C/P,P
CHECKMATE: UNIVERSE [@PERRY RHODAN #74]  ACE  C/PA,PA
CHECKPOINT LAMBDA  LEINSTER, MURRAY  C/P,P
CHEIM MANUSCRIPT, THE  PRATHER, RICHARD S.  SFWA-P
CHEKHOV’S JOURNEY  WATSON, IAN  H,P
CHEKOV’S ENTERPRISE  KOENIG, WALTER  R/REF-P
CHENGTU STRAIN, THE  KAY, KENNETH/GOLDBERG, MARSHALL  P
CHEON OF WELTANLAND I: THE FOUR WISHES  STONE, CHARLOTTE  C/P,P
CHERNAGOR PIRATES, THE [@SCEPTER OF MERCY #2]  CHERNENKO, DAN  LP
CHERNEVOG  CHERRYH, C. J.  H:2,P
CHERNOBYL  POHL, FREDERIK  C/P:2,P
CHERNOBYL SYNDROME, THE  ING, DEAN  SFWA-P
CHERRON  COMBES, SHARON  P
CHESS WITH A DRAGON  GERROLD, DAVID  C/P,P,H
CHESSBOARD PLANET  PADGETT, LEWIS  GN5.26
CHESSBOARD QUEEN, THE [GUINEVERE 2]  NEWMAN, SHARAN  LP
CHESSMEN OF MARS, THE [BARSOOM #5]  BURROUGHS, EDGAR RICE  C/P:2,@P
CHI [VIRTUAL REALITY 3]  BESHER, ALEXANDER  H
CHICAGO CONVERSION, THE [V]  PROCTOR, GEO. W.  C/P,P
CHICAGO RED  MELUCH, R. M.  C/P:2,P
CHICK IS IN THE MAIL, THE [CHICKS IN CHAINMAIL 5]  FRIESNER, ESTHER M.  PA,C/PA
CHICKS IN CHAINMAIL [CHICKS IN CHAINMAIL 1]  FRIESNER, ESTHER M.  PA
CHICKS ’N CHAINED MALES [CHICKS IN CHAINMAIL 3]  FRIESNER, ESTHER M.  HA,PA
CHIEFTAIN, THE [@TELNARIAN 1]  NORMAN, JOHN  C/P,P
CHIEFTAIN OF ANDOR  OFFUTT, ANDREW J.  C/P:2,P
CHIEN DES BASKERVILLE, LE  DOYLE, ARTHUR CONAN  FP
CHILD ACROSS THE SKY, A  CARROLL, JONATHAN  P
CHILD BUYER, THE  HERSEY, JOHN  C/P,P,H
CHILD CHRISTOPHER AND GOLDILIND THE FAIR  MORRIS, WILLIAM  LP
CHILD GARDEN, THE  RYMAN, GEOFF  H
CHILD OF A RAINLESS YEAR  LINDSKOLD, JANE  P
CHILD OF EARTH [SEA OF GRASS 1]  GERROLD, DAVID  LP
CHILD OF ELVISH, A [JEWELS OF ELVISH 2]  BERBERICK, NANCY VARIAN  C/P:2,P
CHILD OF FLAME [CROWN OF STARS 4]  ELLIOTT, KATE  H:2
CHILD OF FORTUNE  SPINRAD, NORMAN  H:2,P
CHILD OF SATURN [CELYDONN GREEN LION 1]  EDGERTON, TERESA  C/P:2,SFWA-P,P
CHILD OF THE DARK PROPHECY [@GREAT TREE OF AVALON 1]  BARRON, T. A.  P
CHILD OF THE EAGLE  FRIESNER, ESTHER M.  C/P,P
CHILD OF THE GROVE  HUFF, TANYA  C/P,P
CHILD OF THE HUNT [@BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
CHILD OF THE JOURNEY [MADAGASCAR MANIFESTO 2]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
CHILD OF THE LIGHT [MADAGASCAR MANIFESTO 1]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
CHILD OF THE PROPHECY [SEVENWATERS 3]  MARILLIER, JULIET  H
CHILD OF THE RIVER [@CONFLUENCE #1]  MCAULEY, PAUL J.  H,P
CHILD OF THUNDER [RENSHAI 3]  REICHERT, MICKEY ZUCKER  C/P,P

The MITSFS Titledex  September 24, 2007  Page 56
TITLE  AUTHOR  CODE
CHILD OF TOMORROW [TIME KEEPER 2]  BARTHOLOMEW, BARBARA  P
CHILD OF VENUS [VENUS OF DREAMS 3]  SARGENT, PAMELA  P
CHILD QUEEN, THE [CHILD QUEEN 1]  MCKENZIE, NANCY  P
CHILDE MORGAN [DERYNI]  KURTZ, KATHERINE  H
CHILDE ROLANDE  LEE, SAMANTHA  LP
CHILDHOOD’S END  CLARKE, ARTHUR C.  H:3,SR-H,P
CHILDREN OF AMARID [LON TOBYN 1]  COE, DAVID B.  C/P:2,P
CHILDREN OF ANTHI, THE [ANTHI]  BLAKENEY, JAY D.  P
CHILDREN OF ARABLE  BELDEN, DAVID  C/P,P
CHILDREN OF ASHGAROTH, THE [FARADAWN 3]  FORD, RICHARD  P
CHILDREN OF CHAOS [CHILDREN OF CHAOS 1]  DUNCAN, DAVE  H
CHILDREN OF CTHULHU, THE  PELAN, JOHN/ADAMS, BENJAMIN  SR-LPA
CHILDREN OF DUNE [DUNE 3]  HERBERT, FRANK  C/P:2,P,H
CHILDREN OF ENCHANTMENT [POWER AND THE PATTERN 2]  BUSH, ANNE KELLEHER  P
CHILDREN OF FLUX AND ANCHOR [@SOUL RIDER #5]  CHALKER, JACK L.  C/P:2,SFWA-P,P

CHILDREN OF GAIA [WORLD OF DARKNESS WEREOLF 5]/ UKTENA [WORLD OF DARKNESS WEREOLF 5] BYERS, RICHARD LEE/PETRUCHA, STEFAN

PA

CHILDREN OF GOD  RUSSELL, MARY DORIA  LP
CHILDREN OF GREEN KNOWE, THE [GREEN KNOWE 1]  BOSTON, L. M.  H,LP
CHILDREN OF HAMLIN, THE [STAR TREK TNG]  CARTER, CARMEN  C/P:3,P
CHILDREN OF HASTUR [DARKOVER]  BRADLEY, MARION ZIMMER  H:2
CHILDREN OF HATE, THE  CREASEY, JOHN  H
CHILDREN OF HOPE [NICHOLAS SEAFORT 7]  FEINTUCH, DAVID  H
CHILDREN OF INFINITY  ELWOOD, ROGER  HA,SFWA
CHILDREN OF LLYR, THE [MABINOGI 2]  WALTON, EVANGELINE  C/P:2,P
CHILDREN OF MAGIC  GREENBERG, MARTIN H./HUGHES, KERRIE  PA
CHILDREN OF MEN, THE  JAMES, P. D.  H,P
CHILDREN OF MORROW  HOOVER, H. M.  P
CHILDREN OF MU, THE  CHURCHWARD, JAMES  [C/RAN-P:2]
CHILDREN OF SHINY MOUNTAIN, THE  DVORKIN, DAVID  P
CHILDREN OF THE ATOM  SHIRAS, WILMAR H.  H:2,P
CHILDREN OF THE BLOOD [SUNDERED 2]  SAGARA, MICHELLE  P,LP
CHILDREN OF THE COMPANY, THE [COMPANY 6]  BAKER, KAGE  H:2
CHILDREN OF THE DARK  VELEY, CHARLES  P
CHILDREN OF THE DRAGON [@EXALTED SECOND AGE #3]  DANSKY, RICHARD E.  P
CHILDREN OF THE DRAGON  ESTES, ROSE  P
CHILDREN OF THE DRAGON  ROBINSON, FRANK S.  P
CHILDREN OF THE DRAKE [DRAGONREALM ORIGINS 2]  KNAAK, RICHARD A.  P
CHILDREN OF THE DUSK [MADAGASCAR MANIFESTO 3]  BERLINER, JANET/GUTHRIDGE, GEORGE  P
CHILDREN OF THE DUST  LAWRENCE, LOUISE  H
CHILDREN OF THE JEDI [STAR WARS]  HAMBLY, BARBARA  H:2,P
CHILDREN OF THE LENS [LENSMEN 6]  SMITH, E. E.  C/P,P,H
CHILDREN OF THE LIGHT  WESTON, SUSAN B.  H,SFWA,P
CHILDREN OF THE LION [@CHEYSULI #5,6]  ROBERSON, JENNIFER  P
CHILDREN OF THE MIND [ENDER 4]  CARD, ORSON SCOTT  H:3,P
CHILDREN OF THE NIGHT  BLACKBURN, JOHN  P
CHILDREN OF THE NIGHT [SPACE MAVERICKS 2]  KRING, MICHAEL  P
CHILDREN OF THE NIGHT [DIANA TREGARDE]  LACKEY, MERCEDES  C/P:2,P
CHILDREN OF THE NIGHT  SIMMONS, DAN  C/P,P,H
CHILDREN OF THE PLAINS [@DRAGONLANCE BARBARIANS #1]  THOMPSON, PAUL B./COOK, TONYA C.  P
CHILDREN OF THE RUNE [MONTE COOK’S ARCANA EVOLVED]  COOK, SUE WEINLEIN  C/PA,PA
CHILDREN OF THE SERPENT GATE [TEARS OF ARTAMON 3]  ASH, SARAH  H
CHILDREN OF THE SHAMAN  RYDILL, JESSICA  P
CHILDREN OF THE THUNDER  BRUNNER, JOHN  C/P:2,P
CHILDREN OF THE VOID  DEXTER, WILLIAM  C/P:2,P
CHILDREN OF THE WIND, THE [SEVEN CITADELS 2]  HARRIS, GERALDINE  P
CHILDREN OF THE WIND  SHIEL, M. P.  H
CHILDREN OF THE WIND  WILHELM, KATE  H,SFWA-P,LP
CHILDREN OF TOMORROW  VAN VOGT, A. E.  C/P:2,P
CHILDREN OF WONDER  TENN, WILLIAM  HA
CHILDREN OF WRATH, THE [RENSHAI CHRONICLES 3]  REICHERT, MICKEY ZUCKER  P
CHILDREN STAR, THE [ELYSIUM]  SLONCZEWSKI, JOAN  H:2,P
CHILDREN’S CRUSADE, THE [@TRAVELER #11]  DRUMM, D. B.  P
CHILDREN’S HOUR, THE [MAN-KZIN WARS]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
CHILEKINGS, THE  WEST, JESSAMYN  C/P:2,P
CHILL [@CHILL #2]  SHERMAN, JORY  P
CHILL AND THE KILL, THE  FLEMING, JOAN  P
CHILL FACTOR [WEATHER WARDEN 3]  CAINE, RACHEL  C/P,P
CHILL IN THE BLOOD, A [@VAMPIRE FILES 7]  ELROD, P. N.  H
CHILLED TO THE BONE  GARCIA, ROBERT T.  SR-LPA
CHILLER  BLAKE, STERLING  H
CHIMAERA’S COPPER [KELVIN HACKLEBERRY 3]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P,P,H
CHIMERA  BARTH, JOHN  C/P,P
CHIMERA  ROSENBLUM, MARY  P

The MITSFS Titledex  September 24, 2007  Page 57
TITLE  AUTHOR  CODE
CHIMERA  SHETTERLY, WILL  H
CHIMERA [@ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  LP
CHIN MUSIC  MCMANUS, JAMES  H
CHINA MOUNTAIN ZHANG  MCHUGH, MAUREEN F.  C/P,P,H
CHINDI [PRISCILLA HUTCHINS 3]  MCDEVITT, JACK  H:2
CHINESE AGENDA, THE  POYER, JOE  H
CHINESE FAIRY STORIES  PITMAN, NORMAN H.  H
CHIVALRY  CABELL, JAMES BRANCH  H
CHIY-UNE [LIVING PLANET 4]  COON, SUSAN  P
CHLOROPLASM  CARTER, DEE  SR-P
CHOCKY  WYNDHAM, JOHN  H:2,P
CHOCOLATE, THE CONSUMING PASSION  BOYNTON, SANDRA  [C/RAN-H]
CHOICE CUTS  BOILEAU, PIERRE/NARCEJAC, THOMAS  P
CHOICE OF DESTINIES, A  SCOTT, MELISSA  C/P,P
CHOICE OF GODS, A  SIMAK, CLIFFORD D.  C/P,P,H
CHOICE OF THE CAT [@VAMPIRE EARTH #2]  KNIGHT, E. E.  P
CHOIR OF ILL CHILDREN, A  PICCIRILLI, TOM  P
CHOIR OF ILL CHILDREN, A (PROOF)  PICCIRILLI, TOM  SR-P

CHOOSE YOUR ENEMIES CAREFULLY [SHADOWRUN 2/ SECRETS OF POWER 2] CHARRETTE, ROBERT N.

C/P:2,P

CHOOSING NAMES [@MAN-KZIN WARS #8]  NIVEN, LARRY  C/PA,PA
CHORALE  MALZBERG, BARRY N.  H,SFWA
CHORUS SKATING [SPELLSINGER 8]  FOSTER, ALAN DEAN  C/P,P
CHOSEN [@BUFFY]  HOLDER, NANCY  P
CHOSEN, THE [STONE DANCE OF THE CHAMELEON 1]  PINTO, RICARDO  C/P,P,H
CHOSEN, THE [GENERAL]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P

CHOSEN, THE [STARGATE ATLANTIS #3] WHITELAW, SONNY/

CHRISTENSEN, ELIZABETH

P

CHOSEN/PLACEBO EFFECT, THE  WINTERS, K. C.  P
CHOSEN OF MIDA [JALAV 3]  GREEN, SHARON  C/P,P
CHOSEN OF THE GODS [@DRAGONLANCE KINGPRIEST #1]  PIERSON, CHRIS  P
CHOSEN THE SUN [@EXALTED SECOND AGE #1]  DANSKY, RICHARD E.  P
CHRESTOMANCI QUARTET, THE [CHRESTOMANCI 1,2,3,4]  JONES, DIANA WYNNE  H
CHRESTOMATHY  LAUMER, KEITH  C/P:2,P
CHRISTENING QUEST, THE [ARGONIA 4]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
CHRISTINE  KING, STEPHEN  P

CHRISTMAS BESTIARY GREENBERG, ROSALIND M./ GREENBERG, MARTIN H.

PA,C/PA

CHRISTMAS CAROL AND OTHER VICTORIAN FAIRY TALES, A  KNOEPFLMACHER, U. C.  PA
CHRISTMAS FOREVER  HARTWELL, DAVID G.  HA
CHRISTMAS GHOSTS  CRAMER, KATHRYN/HARTWELL, DAVID G.  SFWA
CHRISTMAS GHOSTS  RESNICK, MIKE/GREENBERG, MARTIN H.  PA
CHRISTMAS MAGIC  HARTWELL, DAVID G.  PA
CHRISTMAS ON A RATIONAL PLANET [DOCTOR WHO NEW ADVENTURES]  MILES, LAWRENCE  P
CHRISTMAS ON GANYMEDE  GREENBERG, MARTIN H.  C/PA,PA
CHRISTMAS SLAUGHTER  GRANT, MARK  P
CHRISTMAS STARS  HARTWELL, DAVID G.  C/PA,LPA
CHRISTMAS TREE THAT ATE MY MOTHER, THE  MARNEY, DEAN  LP
CHROME  NADER, GEORGE  H,P
CHROME BORNE, THE [SERRATED EDGE 1,4]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P
CHROME CIRCLE [SERRATED EDGE 4]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,P
CHROMOSOMAL CODE, THE  WATT-EVANS, LAWRENCE  C/P,P
CHROMOSOME 6  COOK, ROBIN  P
CHRONICLE  ZOSS, JOEL  C/P:2,P

CHRONICLES OF A COMER AND OTHER RELIGIOUS SCIENCE FICTION STORIES ELWOOD, ROGER

LPA

CHRONICLES OF AMBER 1, THE [AMBER 1,2,3]  ZELAZNY, ROGER  H:2
CHRONICLES OF AMBER 2, THE [AMBER 4,5]  ZELAZNY, ROGER  H:2
CHRONICLES OF CHRESTOMANCI, VOLUME II [CHRESTOMANCI 3,2]  JONES, DIANA WYNNE  P
CHRONICLES OF CORUM, THE [CORUM 1,2,3]  MOORCOCK, MICHAEL  P
CHRONICLES OF ER-DA: BOOK ONE [ER-DA 1]  REDMOND, JEFFREY  LP
CHRONICLES OF ER-DA: BOOK TWO [ER-DA 2]  REDMOND, JEFFREY  LP
CHRONICLES OF NARNIA CALENDAR, THE (1982)  CALENDAR  CAL
CHRONICLES OF RIDDICK, THE  FOSTER, ALAN DEAN  P
CHRONICLES OF SCAR, THE [SCAR 1]  SARTI, RON  C/P,P
CHRONICLES OF THE DERYNI, THE [DERYNI CHRONICLES 1,2,3]  KURTZ, KATHERINE  H
CHRONICLES OF THE RAVEN 1 [CHRONICLES OF THE RAVEN 1,2]  BARCLAY, JAMES  H
CHRONO CRUSADE, VOL. 1 [CHRONO CRUSADE 1]  MORIYAMA, DAISUKE  SCX
CHRONOCULES  COMPTON, D. G.  C/P:2,P
CHRONOLITHS, THE  WILSON, ROBERT CHARLES  C/P,P,H
CHRONOLYSIS  JEURY, MICHAEL  H:2
CHRONOMASTER  LINDSKOLD, JANE  P
CHRONOPOLIS AND OTHER STORIES  BALLARD, J. G.  H:2
CHRONOPOLIS: THE SCIENCE FICTION OF J. G. BALLARD  BALLARD, J. G.  C/P:2,P

TITLE  AUTHOR  CODE
CHRONOSEQUENCE  SCHENCK, HILBERT  C/P,SFWA-P,P,H
CHRONOSPACE  STEELE, ALLEN  H
CHRYSALIDS, THE  WYNDHAM, JOHN  C/P:2,P
CHRYSALIS  TORGESON, ROY  PA
CHRYSALIS [@STAR TREK VOYAGER]  WILSON, DAVID NIALL  P
CHRYSALIS 2  TORGESON, ROY  PA,C/PA
CHRYSALIS 3  TORGESON, ROY  PA
CHRYSALIS 4  TORGESON, ROY  C/PA,PA
CHRYSALIS 5  TORGESON, ROY  PA
CHRYSALIS 6  TORGESON, ROY  PA
CHRYSALIS 7  TORGESON, ROY  PA
CHRYSALIS 8  TORGESON, ROY  HA,PA
CHRYSALIS 9  TORGESON, ROY  HA,PA,SFWA
CHRYSALIS 10  TORGESON, ROY  HA,SFWA
CHRYSALIS OF DEATH  ROBINSON, ELEANOR  P
CHTHON [CHTHON #1]  ANTHONY, PIERS  C/P:2,@P
CHYMICAL WEDDING, THE  CLARKE, LINDSAY  P

CIAPHAS CAIN: HERO OF THE IMPERIUM [@WARHAMMER 40K CIAPHAS CAIN 1,2,3] MITCHELL, SANDY

LP

CIARA’S SONG [WITCH WORLD]  NORTON, ANDRE/MCCONCHIE, LYN  P
CIMMERIAN RAGE [@AGE OF CONAN KERN #2]  COLEMAN, LOREN L.  P
CINDER  DE MARINIS, RICK  P
CINDERBLOCK  YOUNG, JANINE ELLEN  P
CINEMA OF THE FANTASTIC  STEINBRUNNER, CHRIS/GOLDBLATT, BURT  R/REF-H
CINEVERSE CYCLE, THE [CINEVERSE 1,2,3]  GARDNER, CRAIG SHAW  H
CINNABAR  BRYANT, EDWARD  C/P,P,H
CINNABAR SHADOWS [@DARK SUN ATHAS #4]  ABBEY, LYNN  P
CINQ SEMAINES AU BALLON  VERNE, JULES  FP
CIRCLE AT CENTER [SEVEN CIRCLES 1]  NILES, DOUGLAS  LP
CIRCLE, CRESCENT, STAR [KANTMORIE 2]  DIBELL, ANSEN  C/P:2,P
CIRCLE IN THE SEA, A  SENN, STEVE  H
CIRCLE OF ONE [JENNY SIXA 1]  FULLILOVE, ERIC JAMES  C/P,P
CIRCLE OF THE MOON [SISTERS OF THE RAVEN 2]  HAMBLY, BARBARA  LP
CIRCLE WAR [WINGMAN 2]  MALONEY, MACK  P
CIRCLES AND STANDING STONES  HADINGHAM, EVAN  SFWA
CIRCUIT [CIRCUIT 1]  SNODGRASS, MELINDA M.  C/P:2,P
CIRCUIT BREAKER [CIRCUIT 2]  SNODGRASS, MELINDA M.  P,SFWA-P
CIRCUIT OF HEAVEN [CIRCUIT OF HEAVEN 1]  DANVERS, DENNIS  H,P
CIRCUMPOLAR!  LUPOFF, RICHARD A.  H,P
CIRCUS OF DOCTOR LAO, THE  BRADBURY, RAY  C/PA,PA
CIRCUS OF DR. LAO, THE  FINNEY, CHARLES G.  C/P:2,SR-LP,LP
CIRCUS OF FEAR  ESTES, ROSE  P
CIRCUS OF HELLS, A [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
CIRCUS OF THE DAMNED [ANITA BLAKE 3]  HAMILTON, LAURELL K.  C/P:2,P
CIRCUS WORLD [CIRCUS WORLD 3]  LONGYEAR, BARRY B.  C/P,P,H
CIRQUE  CARR, TERRY  C/P,SFWA,P,H
CITADEL, THE [@DRAGONLANCE CLASSICS]  KNAAK, RICHARD A.  P
CITADEL, THE [GALACTIC MI]  RANDLE, KEVIN  P
CITADEL OF DREAMS [DOCTOR WHO]  STONE, DAVE  SR-H
CITADEL OF FEAR, THE  STEVENS, FRANCIS  C/P:2,P
CITADEL OF SHADOWS [ELUNDIUM 4]  JEFFERIES, MIKE  C/P,P
CITADEL OF THE AUTARCH, THE [NEW SUN 4]  WOLFE, GENE  C/P,P,H
CITADELS ON EARTH [@FLASH GORDON #6]  ANONYMOUS  C/P,P
CITADELS UNDER ATTACK [@FLASH GORDON #5]  ANONYMOUS  P
CITIES  CROWTHER, PETER  HA
CITIES IN FLIGHT [CITIES IN FLIGHT 1,2,3,4]  BLISH, JAMES  C/P,H:2,P
CITIES IN FLIGHT, VOL. 1 [CITIES IN FLIGHT 1,2]  BLISH, JAMES  C/P:2,SFWA-P,P
CITIES IN FLIGHT, VOL. 2 [CITIES IN FLIGHT 3,4]  BLISH, JAMES  P
CITIES IN SPACE [ENDLESS FRONTIER 3]  POURNELLE, JERRY/CARR, JOHN F.  PA
CITIES OF THE RED NIGHT  BURROUGHS, WILLIAM S.  LP
CITIES OF WONDER  KNIGHT, DAMON  PA,C/PA
CITIZEN IN SPACE  SHECKLEY, ROBERT  C/P,P,H
CITIZEN OF THE GALAXY  HEINLEIN, ROBERT A.  C/P:2,P
CITIZEN PHAID [PHAID 2]  FARREN, MICK  C/P:2,P
CITIZEN VAMPIRE [DON SEBASTIAN 3]  DANIELS, LES  P
CITY, THE [ATLAN]  GASKELL, JANE  C/P,P,H
CITY [@SMALLVILLE]  GRAYSON, DEVIN  P
CITY  SIMAK, CLIFFORD D.  C/P,SFWA-P,P,H

CITY: 2000 A.D.: URBAN LIFE THROUGH SCIENCE FICTION, THE CLEM, RALPH/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

C/PA,PA

CITY (INCLUDES "EPILOG")  SIMAK, CLIFFORD D.  P
CITY AND THE STARS, THE  CLARKE, ARTHUR C.  H:2,P
CITY AT THE EDGE OF TIME [THIEVES’ WORLD TEMPUS 5]  MORRIS, JANET/MORRIS, CHRIS  C/P:2,SFWA-P,P
CITY AT WORLD’S END [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  P

TITLE  AUTHOR  CODE
CITY AT WORLD’S END  HAMILTON, EDMOND  SR-P,C/P:2,GN3.18,P
CITY BEYOND THE CLOUDS, THE  ROCKWOOD, ROY  H
CITY COME A-WALKIN’  SHIRLEY, JOHN  P
CITY DESTROYER, THE [@SPIDER #3]  STOCKBRIDGE, GRANT  P
CITY IN THE AUTUMN STARS, THE [VON BEK 2]  MOORCOCK, MICHAEL  C/P,SFWA,SFWA-P,P
CITY IN THE GLACIER, THE [WAR OF POWERS 2]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
CITY IN THE MIST [AMERICA 2040 3]  INNES, EVAN  P
CITY IN THE NORTH, A  RANDALL, MARTA  C/P:2,P
CITY IN THE SAHARA, THE  VERNE, JULES  H,P
CITY IN THE SEA, THE  TUCKER, WILSON  SR-P,C/P,GN2.11
CITY IN THE SKY  SIODMAK, CURT  H:2,P
CITY IN THE STARS, THE [@NEW TOM SWIFT #1]  APPLETON, VICTOR  H,LP
CITY IN WINTER, A  HELPRIN, MARK  H
CITY MACHINE, THE  TRIMBLE, LOUIS  C/P:2,P
CITY, NOT LONG AFTER, THE  MURPHY, PAT  C/P,P
CITY OF [@ANGEL]  HOLDER, NANCY  P
CITY OF A MILLION LEGENDS [KREN 2]  LICHTENBERG, JACQUELINE  C/P:2,P
CITY OF A THOUSAND SUNS [FALL OF THE TOWERS 3]  DELANY, SAMUEL R.  C/P:2,P
CITY OF ASSASSINS [@SHADOW WORLD 3]  HAMMELL, IAN  P
CITY OF BARABOO [CIRCUS WORLD 1]  LONGYEAR, BARRY B.  C/P,SFWA,P,H
CITY OF BONES  WELLS, MARTHA  H,P
CITY OF BRASS  HOCH, EDWARD D.  C/P,P
CITY OF CAIN  WILHELM, KATE  C/P:2,P
CITY OF CRYSTAL SHADOW [QUESTIONER 2]  SCHMIDT, DENNIS  P
CITY OF DARKNESS  BOVA, BEN  P
CITY OF DARKNESS: UNSEEN [WORLD OF DARKNESS]  KELLY, ERIN/WIECK, STEWART  C/PA,PA
CITY OF DIAMOND  EMERSON, JANE  P
CITY OF ENDLESS NIGHT  HASTINGS, MILO  H,LP
CITY OF EXILE, THE  HARRIS, DEBORAH TURNER  C/P,P
CITY OF FEAR [@OMEGA SUB #3]  CAMERON, J. D.  P
CITY OF FIRE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
CITY OF FORTUNE [@DRAGONLANCE NEW ADVENTURES ELIDOR #3]  SOESBEE, REE  LP
CITY OF GLASS  LOOMIS, NOEL  SR-P
CITY OF GOLD AND LEAD, THE [TRIPODS 2]  CHRISTOPHER, JOHN  C/P:2,P
CITY OF GOLDEN SHADOW [@OTHERLAND #1]  WILLIAMS, TAD  H,P
CITY OF HAWKS [GORD THE ROGUE]  GYGAX, GARY  C/P:2,SFWA-P,P
CITY OF ILLUSIONS [HAINISH 3]  LE GUIN, URSULA K.  C/P:2,SFWA,P
CITY OF IRON [@SEARCHERS #1]  WILLIAMSON, CHET  P
CITY OF PEARL [WESS’HAR WARS 1]  TRAVISS, KAREN  P
CITY OF PEARL (PROOF) [WESS’HAR WARS 1]  TRAVISS, KAREN  SR-P
CITY OF PILLARS  PELOSO, DOMINIC  LP
CITY OF RAVENS, THE [@FORGOTTEN REALMS CITIES]  BAKER, RICHARD  P
CITY OF SAINTS AND MADMEN [AMBERGRIS]  VANDERMEER, JEFF  LP
CITY OF SORCERY [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
CITY OF SPLENDORS, THE [@FORGOTTEN REALMS CITIES]  GREENWOOD, ED/CUNNINGHAM, ELAINE  P
CITY OF THE BEAST, THE [MICHAEL KANE 1]  MOORCOCK, MICHAEL  C/P:2,P
CITY OF THE CHASCH  VANCE, JACK  C/P:2,P
CITY OF THE DEAD [@RESIDENT EVIL #3]  PERRY, S. D.  P
CITY OF THE DEAD, THE [DOCTOR WHO EIGHTH DOCTOR]  ROSE, LLOYD  P
CITY OF THE DEAD [@STAR WARS GALAXY OF FEAR #2]  WHITMAN, JOHN  P
CITY OF THE GODS [@STARGATE SG-1 #4]  WHITELAW, SONNY  P
CITY OF THE LIVING DEAD [@BLADE #26]  LORD, JEFFREY  P
CITY OF THE LOST [@DRAGONLANCE LINSHA #1]  HERBERT, MARY H.  P
CITY OF THE SINGING FLAME, THE  SMITH, CLARK ASHTON  C/P,P
CITY OF THE SORCERERS [DARK HORSE 3]  HERBERT, MARY H.  P
CITY OF THE SUN [DAEDALUS 4]  STABLEFORD, BRIAN M.  C/P,P
CITY OF TOWERS, THE [@EBERRON DREAMING DARK #1]  BAKER, KEITH  P
CITY OF TRUTH  MORROW, JAMES  H:2,LP
CITY OF WHISPERING STONE [MONGO FREDERICKSON 2]  CHESBRO, GEORGE C.  C/P,P,H
CITY OF WONDER  VIVIAN, E. CHARLES  P
CITY ON FIRE [AIAH 2]  WILLIAMS, WALTER JON  H,P
CITY ON THE EDGE OF FOREVER, THE [STAR TREK]  ELLISON, HARLAN  SR-S
CITY ON THE EDGE OF FOREVER (FOTONOVEL) [STAR TREK]  RODDENBERRY, GENE  [C/RAN-P]
CITY ON THE MOON  LEINSTER, MURRAY  DD277,H
CITY OUTSIDE THE WORLD, THE [MAN WHO LOVED MARS 3]  CARTER, LIN  C/P,P
CITY UNDER GROUND, THE  MARTEL, SUZANNE  H
CITY UNDER THE SEA  BULMER, KENNETH  DD255
CITY UNDER THE SEA  FAIRMAN, PAUL W.  P
CITY WARS  PALUMBO, DENNIS  C/P,P
CITY WHO FOUGHT, THE [SHELLPEOPLE 4]  MCCAFFREY, ANNE/STIRLING, S. M.  H:2,P
CIUDAD DE ILUSIONES  LE GUIN, URSULA K.  FP
CIUDAD, LA  LEVRERO, MARIO  FP
CIUDAD QUE NO EXISTIA, LA  EBLY, P.  FH
CIVIL CAMPAIGN, A [MILES VORKOSIGAN 9]  BUJOLD, LOIS MCMASTER  H,P

TITLE  AUTHOR  CODE
CIVIL WAR FANTASTIC  GREENBERG, MARTIN H.  PA
CIVIL WAR II  BRITAIN, DAN  P
CIVIL WAR SECRET AGENT  PERRY, STEVE  P
CIVILIZATION’S LAST HURRAH  COHEN, GARY G.  P
CIVILWARLAND IN BAD DECLINE  SAUNDERS, GEORGE  LP
CLADE  BUDZ, MARK  C/P,P
CLADE (PROOF)  BUDZ, MARK  H,SR-LP
CLAIDI JOURNALS, THE [CLAIDI 1,2,3]  LEE, TANITH  H
CLAIM ON FORTY MILE CREEK, THE  VERNE, JULES  H
CLAIMED  STEVENS, FRANCIS  P
CLAIMED BY SHADOW [CASSANDRA PALMER]  CHANCE, KAREN  P
CLAIMING OF SLEEPING BEAUTY, THE [BEAUTY 1]  ROQUELAURE, A. N.  H,SR-L
CLAIRVOYANT COUNTESS, THE  GILMAN, DOROTHY  P
CLAN AND CROWN [@JANISSARIES 2]  POURNELLE, JERRY/GREEN, ROLAND  C/P,LP,H
CLAN CORPORATE, THE [MERCHANT PRINCES 3]  STROSS, CHARLES  H:2
CLAN GROUND [RATHA 2]  BELL, CLARE  H
CLAN OF THE CATS, THE [HORSECLANS 18]  ADAMS, ROBERT  P
CLAN OF THE CAVE BEAR, THE [EARTH’S CHILDREN 1]  AUEL, JEAN M.  C/P:2,P
CLAN OF THE WARLORD, THE [SKYLA 1]  BOYER, ELIZABETH H.  P
CLANDESTINE CIRCLE, THE [@DRAGONLANCE CHRONICLES]  HERBERT, MARY H.  P
CLANS OF THE ALPHANE MOON  DICK, PHILIP K.  C/P:2,P
CLAREMONT TALES II  LUPOFF, RICHARD  H
CLARET, SANDWICHES AND SIN  DUKE, MADELAINE  P
CLARION  GREENLEAF, WILLIAM  C/P,P
CLARION  WILSON, ROBIN SCOTT  PA,C/PA
CLARION AWARDS, THE  KNIGHT, DAMON  HA,SFWA
CLARION II  WILSON, ROBIN SCOTT  C/PA,PA
CLARION III  WILSON, ROBIN SCOTT  C/PA,PA
CLARION SF  WILHELM, KATE  C/PA,PA
CLARKE COUNTY, SPACE  STEELE, ALLEN  C/P:2,SFWA-P:2,P
CLARK’S LAW [@BABYLON 5 #4]  MORTIMORE, JIM  C/P:2,P
CLASH BY NIGHT [CROW]  WILLIAMSON, CHET  LP
CLASH OF ANGELS  DANIELS, JONATHAN  H
CLASH OF EAGLES  RUTMAN, LEO  P
CLASH OF KINGS, A [SONG OF ICE AND FIRE 2]  MARTIN, GEORGE R. R.  C/P,LP,H:2
CLASH OF KINGS, A (BRITISH) [SONG OF ICE AND FIRE 2]  MARTIN, GEORGE R. R.  H
CLASH OF STAR-KINGS  DAVIDSON, AVRAM  DG576
CLASH OF SYMBOLS: THE TRIUMPH OF JAMES BLISH, A  STABLEFORD, BRIAN M.  R/REF-P
CLASH OF THE TITANS  FOSTER, ALAN DEAN  C/P,P
CLASS DIS-MYTHED [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  H
CLASS G-ZERO  HENDRICKSON, WALTER B. JR.  P
CLASS SIX CLIMB  COCHRANE, WILLIAM E.  P
CLASSIC BOOK OF SCIENCE FICTION, THE  CONKLIN, GROFF  HA
CLASSIC PHILIP JOSE FARMER 1952–1964, THE  FARMER, PHILIP JOSE  H
CLASSIC PHILIP JOSE FARMER 1964–1973, THE  FARMER, PHILIP JOSE  H
CLASSIC SCIENCE FICTION: THE FIRST GOLDEN AGE  CARR, TERRY  SFWA
CLAW OF THE CONCILIATOR, THE [NEW SUN 2]  WOLFE, GENE  C/P,P,H
CLAWS OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 1]  THORPE, GAV  P
CLAY’S ARK  BUTLER, OCTAVIA E.  C/P,SFWA,P,H
CLD: COLLECTIVE LANDING DETACHMENT  MILAN, VICTOR  P
CLEFT, THE  TABORI, PAUL  C/P:2,P
CLEFT AND OTHER ODD TALES, THE  WILSON, GAHAN  LP
CLEOPATRA  HAGGARD, H. RIDER  C/P,P
CLEOPATRA CRISIS, THE [TIMEWARS 11]  HAWKE, SIMON  P
CLES DE L’UNIVERS, LES  RAYJEAN, MAX-ANDRE  FP
CLEWISTON TEST, THE  WILHELM, KATE  C/P:2,P
CLICK HERE FOR MURDER [TURING HOPPER 2]  ANDREWS, DONNA  P
CLICKWHISTLE  WATKINS, WILLIAM JOHN  SFWA
CLIFFS [@JOURNEYS OF MCGILL FEIGHAN #4]  O’DONNELL, KEVIN JR.  C/P,P
CLIMACTICON, THE  LIVINGSTON, HAROLD  C/P:2,P
CLIMB THE WIND  SARGENT, PAMELA  P
CLIMBING OLYMPUS  ANDERSON, KEVIN J.  P
CLIPJOINT [CRASHCOURSE 2]  BAIRD, WILHELMINA  P
CLIPPER OF THE CLOUDS, THE  VERNE, JULES  H
CLIQUE  YERMAKOV, NICHOLAS  P
CLIVE BARKER’S BOOKS OF BLOOD 1  BARKER, CLIVE  P
CLIVE BARKER’S BOOKS OF BLOOD 2  BARKER, CLIVE  P
CLIVE BARKER’S BOOKS OF BLOOD 3  BARKER, CLIVE  P
CLOAK [STAR TREK SECTION 31]  PERRY, S. D.  C/P,P
CLOAK AND DAGGER [@STAR TREK VOYAGER DARK MATTERS #1]  GOLDEN, CHRISTIE  P
CLOAK OF AESIR  CAMPBELL, JOHN W. JR.  C/P:2,P
CLOAK OF DECEPTION [@STAR WARS]  LUCENO, JAMES  P
CLOAK OF ILLUSION [CINGULUM 2]  ROBERTS, JOHN MADDOX  P,SFWA-P
CLOAK OF NIGHT AND DAGGERS, THE [TWELVE TREASURES 3]  EDGHILL, ROSEMARY  P

TITLE  AUTHOR  CODE
CLOAK OF SHADOWS [@FORGOTTEN REALMS SHADOW OF AVATAR #2]  GREENWOOD, ED  P
CLOAKED IN SHADOW  HORNER, W. H.  LPA
CLOCK OF DREAMS, THE [TITUS CROW]  LUMLEY, BRIAN  P
CLOCK STRIKES SWORD [@SHADOW WORLD 2]  HAMMELL, IAN  P
CLOCK STRIKES TWELVE, THE  WAKEFIELD, H. R.  H
CLOCKS OF IRAZ, THE [RELUCTANT KING 2]  DE CAMP, L. SPRAGUE  C/P:2,P
CLOCKWISE MAN, THE [DOCTOR WHO NINTH DOCTOR]  RICHARDS, JUSTIN  H
CLOCKWORK ASYLUM [@SHADOWRUN DRAGON HEART 2]  KOKE, JAK  P
CLOCKWORK ORANGE, A  BURGESS, ANTHONY  C/P:2,LP
CLOCKWORK TRAITOR, THE [FAMILY D’ALEMBERT #3]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
CLOCKWORK UNIVERSE OF ANTHONY BURGESS, THE  MATHEWS, RICHARD  R/REF-P
CLOCKWORK WOMAN, THE [@TIME HUNTER 3]  BOTT, CLAIRE  LP
CLOCKWORK’S PIRATES  GOULART, RON  D11182
CLONE  COWPER, RICHARD  H:2,P,SFWA
CLONE, THE  THOMAS, THEODORE L./WILHELM, KATE  C/P:2,P
CLONE CRISIS, THE  MCKEONE, LEE  P
CLONE REBELLION, THE  LIEF, EVELYN  P
CLONE REBELLION, THE (PROOF)  LIEF, EVELYN  SR-LP
CLONE REPUBLIC, THE [CLONE REPUBLIC 1]  KENT, STEVEN L.  P
CLONED LIVES  SARGENT, PAMELA  C/P:2,P
CLONES  DANN, JACK/DOZOIS, GARDNER  PA
CLONES, THE  OLEMY, P. T.  P
CLONES, THE  SHAFFER, EUGENE CARL  P
CLONING  SHEAR, DAVID  C/P,SFWA,P,H
CLONING OF JOANNA MAY, THE  WELDON, FAY  H
CLONING OF MAN, THE  EBON, MARTIN  [C/RAN-P]
CLOSE ENCOUNTERS OF THE THIRD KIND  SPIELBERG, STEVEN  H:2,P
CLOSE ENCOUNTERS OF THE THIRD KIND: A DOCUMENT OF THE FILM  DURWOOD, THOMAS  R/REF-H,C/REF-H
CLOSE ENCOUNTERS OF THE THIRD KIND: THE SPECIAL EDITION  SPIELBERG, STEVEN  C/P,P
CLOSE ENCOUNTERS WITH THE DEITY  BISHOP, MICHAEL  LP
CLOSE QUARTERS [@BATTLETECH 15]  MILAN, VICTOR  P
CLOSE TO CRITICAL  CLEMENT, HAL  C/P:2,P
CLOSE TO THE GROUND [@ANGEL]  MARIOTTE, JEFF  C/P,P
CLOSED CORPORATION, THE  RIDGEWAY, JAMES  [C/RAN-P]
CLOSED SYSTEM  HUGHES, ZACH  C/P:2,P
CLOSED WORLDS, THE [@STARWOLF #2]  HAMILTON, EDMOND  C/P,P
CLOUD ATLAS  MITCHELL, DAVID  LP
CLOUD CASTLES [STEPHEN FISHER 3]  ROHAN, MICHAEL SCOTT  P
CLOUD CHAMBER  MYERS, HOWARD L.  P
CLOUD GATHERER, THE [HARAN OF OAS 1]  TAFF, JOHN F. D.  LP
CLOUD PEOPLE, THE  KELLY, ROBERT B.  P
CLOUD WALKER, THE  COOPER, EDMUND  C/P,P
CLOUD WARRIOR [AMTRAK WARS 1]  TILLEY, PATRICK  C/P:2,P
CLOUD WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
CLOUDBEARER’S SHADOW [SWORD IN EXILE 1]  MARSTON, ANN  P
CLOUDCRY  VAN SCYOC, SYDNEY J.  H:2,P
CLOUDS END  STEWART, SEAN  C/P,P,H
CLOUDS OF MAGELLAN [TIMELAPSE 2]  NIGHBERT, DAVID E.  H,SFWA
CLOUDS OF SATURN, THE  MCCOLLUM, MICHAEL  C/P:2,P
CLOUD’S RIDER  CHERRYH, C. J.  C/P,P,H
CLOVEN HOOVES  LINDHOLM, MEGAN  C/P,P
CLUB DEAD [SOUTHERN VAMPIRES 3]  HARRIS, CHARLAINE  P
CLUB VAMPYRE [ANITA BLAKE 1,2,3]  HAMILTON, LAURELL K.  H
CLUSTER [CLUSTER #1]  ANTHONY, PIERS  C/P:2,@P
CLUSTER COMMAND [CRISIS OF EMPIRE 2]  DIETZ, WILLIAM C./DRAKE, DAVID  P,SFWA-P
CLUTCH OF VAMPIRES, A  MCNALLY, RAYMOND T.  HA,PA
CLYPSIS [@ALIEN SPEEDWAY 1]  CARVER, JEFFREY A.  C/P:2,P
COACHMAN RAT, THE  WILSON, DAVID HENRY  C/P,SFWA-P,P,H
COALESCENT [DESTINY’S CHILDREN 1]  BAXTER, STEPHEN  H,P
COAST OF CORAL, THE  CLARKE, ARTHUR C.  [C/RAN-P]
COBRA [COBRA 1]  ZAHN, TIMOTHY  C/P:2,P
COBRA BARGAIN [COBRA 3]  ZAHN, TIMOTHY  C/P:3,P
COBRA STRIKE [COBRA 2]  ZAHN, TIMOTHY  C/P:2,P
COBRA TRILOGY, THE [COBRA 1,2,3]  ZAHN, TIMOTHY  H
COBRAS TWO [COBRA 1,2]  ZAHN, TIMOTHY  C/P:2,SFWA-P,P
COBWEB, THE  STEPHENSON, NEAL/GEORGE, J. FREDERICK LP
COCKATRICE BOYS, THE  AIKEN, JOAN  H,LP
COCOON [COCOON 1]  SAPERSTEIN, DAVID  C/P,P
C.O.D. MARS  TUBB, E. C.  DH040
CODE BLUE — EMERGENCY [SECTOR GENERAL 7]  WHITE, JAMES  C/P:2,P
CODE DUELLO  REYNOLDS, MACK  D11650,DH103
CODE NAME: JUDAS [@MISSION IMPOSSIBLE #2]  WALKER, MAX  P
CODE NAME PEREGRINE [@STARHAWKS #2]  DALTON, SEAN  C/P,P
CODE NAME: RAPIER [@MISSION IMPOSSIBLE #3]  WALKER, MAX  P

The MITSFS Titledex  September 24, 2007  Page 62
TITLE  AUTHOR  CODE
CODE NOIR  DE PIERRES, MARIANNE  P
CODE OF CONDUCT [JANI KILIAN 1]  SMITH, KRISTINE  C/P,P,H
CODE OF THE LIFEMAKER  HOGAN, JAMES P.  H:2,P
CODE THREE  RAPHAEL, RICK  C/P:2,P
CODE: VERONICA [@RESIDENT EVIL #6]  PERRY, S. D.  P
CODENAME WOLVERINE [X-MEN]  GOLDEN, CHRISTOPHER  H
CODEX DERYNIANUS II [DERYNI]  KURTZ, KATHERINE/REGINALD, ROBERT  R/REF-H
CODGERSPACE  FOSTER, ALAN DEAN  C/P,SFWA-P:2,P,H,SFWA
CODOMINIUM: REVOLT ON WAR WORLD  POURNELLE, JERRY  C/PA,PA,SFWA-P
COELURA, THE  MCCAFFREY, ANNE  C/P,SFWA-P,P,H,SR-H
COHESION [STAR TREK VOYAGER STRING THEORY 1]  LANG, JEFFREY  P
COHORT OF THE DAMNED [@FIFTH FOREIGN LEGION #3]  KEITH, ANDREW  C/P,P
COILS  SABERHAGEN, FRED/ZELAZNY, ROGER  H:2
COILS  ZELAZNY, ROGER/SABERHAGEN, FRED  C/P:2,SFWA-P,LP
COILS OF TIME, THE  CHANDLER, A. BERTRAM  DM107,P
COIN OF EL DIABLO AFFAIR, THE [MAN FROM UNCLE]  GIBSON, WALTER B.  VLP
COINS OF MURPH, THE  KELLEY, LEO P.  C/P:2,P
COINSPINNER’S STORY [@LOST SWORDS #5]  SABERHAGEN, FRED  C/P,SFWA-P,P,H
COLCHICINE FACTOR, THE  BRYCE, ROBERT  P
COLD ALLIES  ANTHONY, PATRICIA  C/P,P,H
COLD AS ICE [BATTACHARIYA 1]  SHEFFIELD, CHARLES  C/P:2,P,H
COLD ASYLUM [@DEATHLANDS]  AXLER, JAMES  P
COLD CASH WAR, THE [COLD CASH WAR 1]  ASPRIN, ROBERT LYNN  C/P,SFWA-P:2,P,H
COLD CASH WARRIOR [COLD CASH WAR 2]  ASPRIN, ROBERT LYNN/FAWCETT, BILL  P
COLD CHILLS  BLOCH, ROBERT  C/P,SFWA,P,H
COLD COPPER TEARS [GARRETT 3]  COOK, GLEN  C/P:2,P
COLD DEATH [DOC SAVAGE #21]  ROBESON, KENNETH  C/P,@P
COLD EQUATIONS, THE  GODWIN, TOM  P
COLD FIRE  KOONTZ, DEAN R.  P
COLD FUSION [DOCTOR WHO MISSING ADVENTURES]  PARKIN, LANCE  C/P,P
COLD HAND IN MINE  AICKMAN, ROBERT  H,P
COLD HAND OF BETRAYAL, THE [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  P
COLD IRON  MICHAELS, MELISA C.  P
COLD MOONS, THE  CLEMENT, AERON  LP
COLD NIGHT’S DEATH, A  HARRISON, BARBARA  P
COLD PRINT  CAMPBELL, RAMSEY  SFWA-P
COLD SEA RISING  MORAN, RICHARD  H
COLD STEEL [@BOLOS #6]  FAWCETT, BILL  PA
COLD STEEL  SHIEL, M. P.  H
COLD VICTORY  ANDERSON, POUL  C/P,P
COLD WAR [@PREDATOR 2]  ARCHER, NATHAN  C/P,P
COLD WAR IN A COUNTRY GARDEN [MATTHEW DILKE 1]  GUTTERIDGE, LINDSAY  C/P:2,P
COLD WARS [STAR TREK GATEWAYS #6/STAR TREK NEW FRONTIER]  DAVID, PETER  P
COLD WIND FROM ORION, A  ASNIN, SCOTT  C/P,P
COLDHEART [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
COLDNESS IN THE BLOOD, A [DRACULA]  SABERHAGEN, FRED  H
COLL AND HIS WHITE PIG  ALEXANDER, LLOYD  LP
COLLAPSIUM, THE [QUEENDOM OF SOL 1]  MCCARTHY, WIL  H
COLLECTED EDITORIALS FROM ANALOG  CAMPBELL, JOHN W. JR.  R/REF-H
COLLECTED FANTASIES  DAVIDSON, AVRAM  P
COLLECTED FEGHOOT, THE  BRIARTON, GRENDEL  LP
COLLECTED GHOST STORIES  WILKINS-FREEMAN, MARY  SR-H
COLLECTED GHOST STORIES OF OLIVER ONIONS, THE  ONIONS, OLIVER  LP
COLLECTED JORKENS, THE, VOLUME TWO [JORKENS]  DUNSANY, LORD  H
COLLECTED JORKENS, THE, VOLUME THREE [JORKENS]  DUNSANY, LORD  H
COLLECTED NOVELS  VERNE, JULES  H
COLLECTED SHORT FICTION OF C.J. CHERRYH, THE  CHERRYH, C. J.  H
COLLECTED SHORT STORIES  FORSTER, E. M.  SR-P
COLLECTED STORIES OF ARTHUR C. CLARKE  CLARKE, ARTHUR C.  H
COLLECTED STORIES OF GREG BEAR, THE  BEAR, GREG  H
COLLECTED STORIES OF PHILIP K. DICK VOLUME 1, THE  DICK, PHILIP K.  LP
COLLECTED STORIES OF PHILIP K. DICK VOLUME 2, THE  DICK, PHILIP K.  LP
COLLECTED STORIES OF PHILIP K. DICK VOLUME 3, THE  DICK, PHILIP K.  LP
COLLECTED STORIES OF PHILIP K. DICK VOLUME 5, THE  DICK, PHILIP K.  LP
COLLECTED STORIES OF VERNOR VINGE, THE  VINGE, VERNOR  H
COLLECTED STORIES, VOLUME ONE  MATHESON, RICHARD  LP
COLLECTED STORIES, VOLUME TWO  MATHESON, RICHARD  LP

COLLECTED WORKS OF BUCK ROGERS IN THE 25TH CENTURY, THE [BUCK ROGERS] DILLE, ROBERT C.

CX

COLLECTION OF SCIENCE FICTION (CHINESE)  YI HUANG  FP
COLLECTOR’S BOOK OF SCIENCE FICTION BY H.G. WELLS, THE  WELLS, H. G.  H:2
COLLECTOR’S POUL ANDERSON, THE  STEVER, DAVID/WHYTE, ANDREW ADAMS  R/IND:2
COLLEGE OF MAGICS, A [COLLEGE OF MAGICS 1]  STEVERMER, CAROLINE  H
COLLINS SPANISH-ENGLISH ENGLISH-SPANISH DICTIONARY  GONZALEZ, MIKE  DICT

The MITSFS Titledex  September 24, 2007  Page 63
TITLE  AUTHOR  CODE
COLLISION COURSE  BAYLEY, BARRINGTON J.  C/P:2,P
COLLISION COURSE  SILVERBERG, ROBERT  DF123,C/P:2,SFWA-P
COLLISION COURSE [SPACE: 1999 4]  TUBB, E. C.  C/P,P
COLLISION WITH CHRONOS  BAYLEY, BARRINGTON J.  P
COLONEL [DIRIGENT MERCENARY CORPS 6]  SHELLEY, RICK  C/P,P
COLONEL SUN [JAMES BOND 14]  AMIS, KINGSLEY  P
COLONIAL SURVEY  LEINSTER, MURRAY  H
COLONIES IN SPACE  HEPPENHEIMER, T. A.  [C/RAN-P]
COLONISTS OF SPACE  CARR, CHARLES  SR-P
COLONIZATION: AFTERSHOCKS [WORLDWAR 7]  TURTLEDOVE, HARRY  H,P
COLONIZATION: DOWN TO EARTH [WORLDWAR 6]  TURTLEDOVE, HARRY  C/P,P
COLONIZATION: SECOND CONTACT [WORLDWAR 5]  TURTLEDOVE, HARRY  H
COLONY  BOVA, BEN  C/P:2,SFWA-P,P
COLONY  GRANT, ROB  P
COLONY, THE  LALLEY, PAUL  P
COLONY [@PLANET OF THE APES]  QUICK, W. T.  P
COLONY, THE  VIGLIANTE, MARY  P
COLONY: EARTH  MOONEY, RICHARD E.  [C/RAN-H,C/RAN-P]
COLONY FLEET  MATTHEWS, SUSAN R.  C/P,P
COLONY IN PERIL [STAR COMMANDOS 2]  GRIFFIN, P. M.  P,SFWA-P
COLONY OF LIES, THE [DOCTOR WHO PAST DOCTORS]  BRAKE, COLIN  P
COLOR OF DISTANCE, THE  THOMSON, AMY  LP
COLOR OF EVIL, THE [DARK DESCENT 1]  HARTWELL, DAVID G.  PA
COLOR OF HER PANTIES, THE [@XANTH #15]  ANTHONY, PIERS  C/P,SFWA-P,P,H
COLOR OF NEANDERTHAL EYES, THE  TIPTREE, JAMES JR.  SFWA-TD16
COLOR OUT OF TIME, THE  SHEA, MICHAEL  C/P,P
COLORS OF CHAOS [RECLUCE 9]  MODESITT, L. E. JR.  H:2,P
COLORS OF CHAOS [@WEAPONS OF CHAOS #3]  VARDEMAN, ROBERT E.  P
COLORS OF MAGIC, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA
COLORS OF SPACE, THE  BRADLEY, MARION ZIMMER  C/P:2,P
COLOSSUS [COLOSSUS 1]  JONES, D. F.  H:2,P
COLOSSUS AND THE CRAB [COLOSSUS 3]  JONES, D. F.  C/P,P
COLOUR OF MAGIC, THE [DISCWORLD 1/DISCWORLD RINCEWIND 1]  PRATCHETT, TERRY  C/P,P,H:2
COLOUR OUT OF SPACE AND OTHERS, THE  LOVECRAFT, H. P.  C/P:3,P
COLOURS IN THE STEEL [FENCER 1]  PARKER, K. J.  LP
COLSEC REBELLION [COLSEC 3]  HILL, DOUGLAS  H,P
COLUMBUS AFFAIR, THE [@PERRY RHODAN #80]  ACE  PA
COLUMBUS OF SPACE, A  SERVISS, GARRETT P.  LP
COMA  COOK, ROBIN  C/P:2,P
COMBAT ROCK [DOCTOR WHO PAST DOCTORS]  LEWIS, MICK  P
COMBAT SF  DICKSON, GORDON R.  C/PA,PA,SFWA
COME ARMAGEDDON  PERRY, ANNE  H
COME AWAY HOME  SMITH, ALISON  SFWA
COME BE MY O.R.G.Y. [MAN FROM ORGY]  MARK, TED  SR-P,C/P
COME ENDLESS DARKNESS [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
COME TWILIGHT [SAINT-GERMAIN 14]  YARBRO, CHELSEA QUINN  H
COMEBACK TOUR [DARK FUTURE]  YEOVIL, JACK  LP
COMES THE BLIND FURY  SAUL, JOHN  C/P,P,H
COMET, THE  CHARLES, ROBERT  P,SFWA-P
COMET  SAGAN, CARL/DRUYAN, ANN  [C/RAN-H]
COMET  WHITE, JANE  H,SFWA
COMET HALLEY  HOYLE, FRED  H
COMET KINGS, THE  HAMILTON, EDMOND  P
COMETEERS, THE  WILLIAMSON, JACK  SR-H,C/P:2,P

COMETS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

COMET’S CURSE, THE [GALAHAD 1]  TESTA, DOM  LP
COMING, THE  HALDEMAN, JOE  C/P,P,H
COMING ATTRACTIONS  GREENBERG, MARTIN  HA
COMING BACK: THE SCIENCE OF REINCARNATION  BHAKTIVEDANTA BOOK TRUST  [C/RAN-P:2]
COMING DARK AGE, THE  VACCA, ROBERTO  SFWA
COMING EVENT, THE [DUMAREST 26]  TUBB, E. C.  C/P,P
COMING OF AGE, A  ZAHN, TIMOTHY  C/P:2,SFWA,P
COMING OF CONAN THE CIMMERIAN, THE [CONAN]  HOWARD, ROBERT E.  H,LP
COMING OF JONATHAN SMITH, THE  LUDLAM, HARRY  SR-P
COMING OF STEELEYE, THE [STEELEYE 1]  DUNN, SAUL  SR-P
COMING OF THE AMAZONS, THE  JOHNSON, OWEN  H
COMING OF THE DEMONS, THE  HOOD, GWENYTH  H
COMING OF THE HORSECLANS, THE [HORSECLANS 1]  ADAMS, ROBERT  C/P:2,P
COMING OF THE KING, THE  TOLSTOY, NIKOLAI  C/P,P,H:2
COMING OF THE QUANTUM CATS, THE  POHL, FREDERIK  C/P,P,H
COMING OF THE RATS, THE  SMITH, GEORGE H.  SR-P
COMING OF THE ROBOTS, THE  MOSKOWITZ, SAM  PA,C/PA
COMING OF THE SPACE AGE, THE  CLARKE, ARTHUR C.  [C/RAN-H]

The MITSFS Titledex  September 24, 2007  Page 64
TITLE  AUTHOR  CODE
COMING OF THE STRANGERS, THE  LYMINGTON, JOHN  P
COMING OF THE TERRANS, THE  BRACKETT, LEIGH  C/P:2,P
COMING OF WISDOM, THE [SEVENTH SWORD 2]  DUNCAN, DAVE  C/P:2,P
COMING SELF-DESTRUCTION OF THE UNITED STATES OF AMERICA, THE  SEYMOUR, ALAN  SR-P
COMMAND DECISION [@OMEGA SUB #2]  CAMERON, J. D.  P
COMMAND DECISION [VATTA’S WAR 4]  MOON, ELIZABETH  H:2
COMMANDER-1  GEORGE, PETER  C/P:2,P
COMMANDER TOAD AND THE PLANET OF THE GRAPES  YOLEN, JANE  LP

COMMANDO BRIGADE 3000 GREENBERG, MARTIN H./ WAUGH, CHARLES G.

PA

COMMANDO DU NON-RETOUR  RANDA, PETER  FP
COMMENCEMENT  JAMES, ROBY  P
COMMITMENT  JAMES, ROBY  P
COMMITMENT HOUR  GARDNER, JAMES ALAN  H,P
COMMITTED MEN, THE  HARRISON, M. JOHN  H,SFWA
COMMITTEE, THE  BUCKLAND, RAYMOND  P
COMMODORE AT SEA, THE [GRIMES]/SPARTAN PLANET [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
COMMON CLAY  ALDISS, BRIAN W.  H:2
COMMON TIME  BLISH, JAMES  B
COMMUNE 2000 A. D.  REYNOLDS, MACK  C/P:2,P
COMMUNICATION WITH EXTRATERRESTRIAL INTELLIGENCE  SAGAN, CARL  [C/RAN-P]
COMMUNION BLOOD [SAINT-GERMAIN 13]  YARBRO, CHELSEA QUINN  H
COMMUNIPATH WORLDS [COYOTE JONES 1,2,3]  ELGIN, SUZETTE HADEN  C/P:2,P
COMMUNIPATHS, THE [COYOTE JONES 1]  ELGIN, SUZETTE HADEN  D11560
COMPANHEIROS DO INFINITO, OS  BARBET, PIERRE  FP
COMPANION PIECE [DOCTOR WHO]  PERRY, ROBERT/TUCKER, MIKE  SR-H
COMPANIONS, THE [@DRAGONLANCE MEETINGS #6]  DANIELL, TINA  C/P,P
COMPANIONS [DOCTOR WHO]  HOWE, DAVID J./STAMMERS, MARK  R/REF-H
COMPANIONS [@LAWS OF THE BLOOD 3]  SIZEMORE, SUSAN  C/P,P
COMPANIONS, THE  TEPPER, SHERI S.  H,P
COMPANIONS OF THE NIGHT  VELDE, VIVIAN VANDE  P
COMPANIONS ON THE ROAD  LEE, TANITH  P
COMPANIONS ON THE ROAD/WINTER PLAYERS, THE  LEE, TANITH  H
COMPANY MAN, THE  FAUST, JOE CLIFFORD  P
COMPANY OF GLASS, THE [EVERIEN 1]  LEITH, VALERY  C/P,LP
COMPANY OF GLORY, THE  PANGBORN, EDGAR  P
COMPANY OF STARS, A [@STARSHIP TROUPERS #1]  STASHEFF, CHRISTOPHER  C/P,P,H
COMPASS OF THE SOUL, THE [RIVER INTO DARKNESS 2]  RUSSELL, SEAN  P
COMPASS REACH [SECANTIS 1]  TIEDEMANN, MARK W.  H,LP
COMPASS ROSE, THE [ONE ROSE 1]  DAYTON, GAIL  LP
COMPASS ROSE, THE  LE GUIN, URSULA K.  C/P:2,P,H
COMPETITIONS [@BLENDING #2]  GREEN, SHARON  C/P,P
COMPKILL  PAULSEN, GARY  C/P,P
COMPLEAT BOLO, THE [BOLOS]  LAUMER, KEITH  C/P,SFWA-P,P
COMPLEAT BOUCHER, THE  BOUCHER, ANTHONY  H
COMPLEAT CROW, THE [TITUS CROW]  LUMLEY, BRIAN  LP
COMPLEAT DYING EARTH, THE [DYING EARTH 1,2,3,4]  VANCE, JACK  H:2

COMPLEAT ENCHANTER: THE MAGICAL MISADVENTURES OF HAROLD SHEA, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/PRATT, FLETCHER

C/P,P,H

COMPLEAT FEGHOOT, THE  BRIARTON, GRENDEL  SR-P,C/P:2
COMPLEAT TRAVELLER IN BLACK, THE  BRUNNER, JOHN  C/P,LP
COMPLEAT WEREWOLF, THE  BOUCHER, ANTHONY  H:2,P
COMPLETE ADVENTURES OF LUCKY STARR, THE [LUCKY STARR 1,2,3,4,5,6] ASIMOV, ISAAC  H
COMPLETE AMBER SOURCEBOOK, THE [AMBER]  KRULIK, THEODORE  R/REF-H
COMPLETE BOOK OF SWORDS, THE [@SWORDS #1,2,3]  SABERHAGEN, FRED  H:2
COMPLETE BOOK OF WITCHCRAFT, THE  AHMED, ROLLO  [C/RAN-P]
COMPLETE CHECKLIST OF SCIENCE-FICTION MAGAZINES, THE  DAY, BRADFORD M.  R/IND:2
COMPLETE COMPLEAT ENCHANTER, THE [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  C/P:2,SFWA-P,P

COMPLETE GUIDE TO MIDDLE-EARTH:

FROM THE HOBBIT TO THE SILMARILLION, THE FOSTER, ROBERT

R/REF-P,C/REF-P

COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 1 [HAMMER’S SLAMMERS] DRAKE, DAVID

H

COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 2 [HAMMER’S SLAMMERS] DRAKE, DAVID

H

COMPLETE INDEX TO ASTOUNDING/ANALOG, THE  ASHLEY, MIKE  R/IND
COMPLETE IVORY, THE [IVORY 1,2,3]  EGAN, DORIS  P
COMPLETE PARATIME, THE [PARATIME POLICE]  PIPER, H. BEAM  H,LP
COMPLETE PEGANA, THE  DUNSANY, LORD  LP
COMPLETE POEMS OF RAY BRADBURY, THE  BRADBURY, RAY  [C/RAN-P]
COMPLETE ROBOT, THE  ASIMOV, ISAAC  H,SFWA
COMPLETE SCIENCE FICTION TREASURY OF H. G. WELLS, THE  WELLS, H. G.  H:2
COMPLETE SHORT STORIES OF AMBROSE BIERCE, VOL. 1, THE  BIERCE, AMBROSE  P
COMPLETE SHORT STORIES OF AMBROSE BIERCE, VOL. 2, THE  BIERCE, AMBROSE  P
COMPLETE STORIES AND POEMS OF EDGAR ALLAN POE  POE, EDGAR ALLAN  H

The MITSFS Titledex  September 24, 2007  Page 65
TITLE  AUTHOR  CODE
COMPLETE STORIES, THE, VOLUME 1  ASIMOV, ISAAC  LP
COMPLETE STORIES, THE, VOLUME 2  ASIMOV, ISAAC  H
COMPLETE VENUS EQUILATERAL, THE  SMITH, GEORGE O.  C/P:2,P
COMPLETE WAR OF THE WORLDS, THE  HOLMSTEN, BRIAN/LUBERTOZZI, ALEX  SR-MM
COMPLEX MAN  FARCA, MARIE C.  H,SFWA
COMPLICITY  BANKS, IAIN M.  P
COMPOUNDED INTERESTS  REYNOLDS, MACK  SR-H
COMPUTER CONNECTION, THE  BESTER, ALFRED  C/P,P,H
COMPUTER WAR  REYNOLDS, MACK  DH034,D11650
COMPUTERWORLD  VAN VOGT, A. E.  C/P:2,P

CONAN [CONAN 1] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P:2,P

CONAN AND THE AMAZON [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN AND THE DEATH LORD OF THANZA [CONAN]  GREEN, ROLAND J.  P
CONAN AND THE EMERALD LOTUS [CONAN 38]  HOCKING, JOHN C.  LP
CONAN AND THE GODS OF THE MOUNTAIN [CONAN]  GREEN, ROLAND J.  P
CONAN AND THE GRIM GREY GOD [CONAN]  MOORE, SEAN A.  LP
CONAN AND THE MANHUNTERS [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN AND THE MISTS OF DOOM [@CONAN]  GREEN, ROLAND J.  P,SFWA-P
CONAN AND THE SHAMAN’S CURSE [CONAN]  MOORE, SEAN A.  P
CONAN AND THE SORCERER [CONAN]  OFFUTT, ANDREW J.  C/P,LP
CONAN AND THE SPIDER GOD [BANTAM CONAN 5]  DE CAMP, L. SPRAGUE  C/P,P
CONAN AND THE TREASURE OF PYTHON [CONAN]  ROBERTS, JOHN MADDOX  LP
CONAN AT THE DEMON’S GATE [CONAN]  GREEN, ROLAND J.  LP
CONAN OF AQUILONIA [CONAN 11]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P

CONAN OF CIMMERIA [CONAN 2] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P:2,P

CONAN OF THE ISLES [CONAN 12]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN OF THE RED BROTHERHOOD [CONAN]  CARPENTER, LEONARD  P
CONAN OF VENARIUM [CONAN]  TURTLEDOVE, HARRY  P
CONAN READER, THE [CONAN]  DE CAMP, L. SPRAGUE  R/REF-H
CONAN, SCOURGE OF THE BLOODY COAST [CONAN]  CARPENTER, LEONARD  P
CONAN THE ADVENTURER [CONAN 5]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P

CONAN THE AVENGER [CONAN 10] HOWARD, ROBERT E./NYBERG, BJORN/

DE CAMP, L. SPRAGUE

C/P:2,P

CONAN THE BARBARIAN [CONAN]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE BOLD [CONAN]  ROBERTS, JOHN MADDOX  P,SFWA-P
CONAN THE BUCCANEER [CONAN 6]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE CHAMPION [CONAN]  ROBERTS, JOHN MADDOX  P
CONAN THE CONQUEROR [CONAN 9]  HOWARD, ROBERT E.  C/P:2,SR-DD036,P
CONAN THE DEFENDER [CONAN]  JORDAN, ROBERT  C/P,SFWA-P,LP
CONAN THE DEFIANT [CONAN]  PERRY, STEVE  SFWA-P,LP
CONAN THE DESTROYER [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE FEARLESS [CONAN]  PERRY, STEVE  H,SFWA-P
CONAN: THE FLAME KNIFE [CONAN]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE P
CONAN THE FORMIDABLE [CONAN] PERRY, STEVE C/P,SFWA-P,LP,P
CONAN THE FREEBOOTER [CONAN 3]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONAN THE FREELANCE [CONAN]  PERRY, STEVE  P,SFWA-P
CONAN THE GLADIATOR [CONAN]  CARPENTER, LEONARD  P
CONAN THE GREAT [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE GUARDIAN [CONAN]  GREEN, ROLAND J.  P,SFWA-P
CONAN THE HERO [CONAN]  CARPENTER, LEONARD  C/P,P
CONAN THE INDOMITABLE [CONAN]  PERRY, STEVE  C/P,SFWA,SFWA-P,LP
CONAN THE INVINCIBLE [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE LIBERATOR [BANTAM CONAN 2]  DE CAMP, L. SPRAGUE/CARTER, LIN  C/P:2,P
CONAN THE MAGNIFICENT [CONAN]  JORDAN, ROBERT  C/P:2,P
CONAN THE MARAUDER [CONAN]  ROBERTS, JOHN MADDOX  P,SFWA-P
CONAN THE MERCENARY [CONAN]  OFFUTT, ANDREW J.  P
CONAN THE OUTCAST [CONAN]  CARPENTER, LEONARD  P
CONAN THE RAIDER [CONAN]  CARPENTER, LEONARD  P,SFWA,SFWA-P,LP
CONAN THE REBEL [BANTAM CONAN 6]  ANDERSON, POUL  C/P,P
CONAN THE RELENTLESS  GREEN, ROLAND J.  P
CONAN THE RENEGADE [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE ROGUE [CONAN]  ROBERTS, JOHN MADDOX  H,LP
CONAN THE SAVAGE [CONAN]  CARPENTER, LEONARD  P

CONAN THE SWORDSMAN [BANTAM CONAN 1] DE CAMP, L. SPRAGUE/CARTER, LIN/

NYBERG, BJORN

C/P,P

CONAN THE TRIUMPHANT [CONAN]  JORDAN, ROBERT  C/P:2,SFWA-P,P
CONAN THE UNCONQUERED [CONAN]  JORDAN, ROBERT  C/P:2,SFWA-P,P
CONAN THE USURPER [CONAN 8]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONAN THE VALIANT [CONAN]  GREEN, ROLAND J.  P,SFWA-P:2
CONAN THE VALOROUS [CONAN]  ROBERTS, JOHN MADDOX  SFWA,LP
CONAN THE VICTORIOUS [CONAN]  JORDAN, ROBERT  C/P,SFWA,SFWA-P:2,P,H

The MITSFS Titledex  September 24, 2007  Page 66
TITLE  AUTHOR  CODE

CONAN THE WANDERER [CONAN 4] HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/

CARTER, LIN

C/P,P

CONAN THE WARLORD [CONAN]  CARPENTER, LEONARD  P,SFWA-P
CONAN THE WARRIOR [CONAN 7]  HOWARD, ROBERT E./DE CAMP, L. SPRAGUE C/P:2,P
CONAN’S WORLD AND ROBERT E. HOWARD  SCHWEITZER, DARRELL  R/REF-P
CONCORDANCE TO CORDWAINER SMITH (PRELIMINARY EDITION)  LEWIS, ANTHONY R.  R/REF-H
CONCORDANCE TO CORDWAINER SMITH (THIRD EDITION)  LEWIS, ANTHONY R.  R/REF-P
CONCRETE ISLAND  BALLARD, J. G.  P
CONCRETE JUNGLE [@PREDATOR 1]  ARCHER, NATHAN  P
CONDEMNATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #3]  BAKER, RICHARD  H,P
CONDITIONALLY HUMAN  MILLER, WALTER M. JR.  C/P:2,P
CONEHEAD  FOX, GARDNER F.  C/P:2,P
CONEHEADS: THE LIFE AND TIMES OF BELDAR CONEHEAD  DAVIS, TOM/AYKROYD, DAN  LP

CONFEDERACY OF THE DEAD GILLIAM, RICHARD/GREENBERG, MARTIN H./

KRAMER, EDWARD E.

PA

CONFEDERATION HANDBOOK, THE [NIGHT’S DAWN]  HAMILTON, PETER F.  P
CONFEDERATION MATADOR  BONE, J. F.  LP
CONFEDERATION OF VALOR, A [TORIN KERR 1,2]  HUFF, TANYA  P
CONFESSIONS OF A CRAP ARTIST  DICK, PHILIP K.  P
CONFESSIONS OF A TEENAGE FROG [@SAMANTHA SLADE #2]  SMITH, SUSAN  P
CONFIDENCE GAME [FIVE COUNTRIES]  WELCH, MICHELLE M.  P
CONFIDENCE GAME (PROOF) [FIVE COUNTRIES]  WELCH, MICHELLE M.  SR-P
CONFLAGRATION ARTIST, A  DENTON, BRADLEY  LP
CONFLICT  ANDERSON, POUL  C/P:2,P
CONFLICT CENTER: NAATOR [@PERRY RHODAN #77]  ACE  C/PA,PA
CONFLICT OF HONORS [LIADEN UNIVERSE 2]  LEE, SHARON/MILLER, STEVE  C/P,P
CONFLUENCE [CONFLUENCE 1,2,3]  MCAULEY, PAUL J.  H
CONFRONTATION [@GUNDAM MOBILE SUIT #3]  TOMINO, YOSHIYUKI  C/P,P
CONFUSION, THE [BAROQUE CYCLE 2]  STEPHENSON, NEAL  H,LP
CONGLOMEROID COCKTAIL PARTY, THE  SILVERBERG, ROBERT  C/P:2,P,H
CONGO  CRICHTON, MICHAEL  C/P,P,H
CONJURE WIFE  LEIBER, FRITZ  SR-P,C/P:2,TD36
CONJURER PRINCESS, THE  VELDE, VIVIAN VANDE  C/P,P
CONJURERS, THE  HARRIS, MARILYN  C/P,P
CONNECTICUT YANKEE IN KING ARTHUR’S COURT, A  TWAIN, MARK  C/P,P,H
CONNOISSEUR’S SCIENCE FICTION  BOARDMAN, TOM  SR-PA,C/PA
CONQUERED, THE [@STAR TREK DS9 REBELS #1]  AB HUGH, DAFYDD  P
CONQUERED PLACE, THE  SHAFER, ROBERT  H
CONQUERING SWORD OF CONAN, THE [CONAN]  HOWARD, ROBERT E.  LP

CONQUERING THE SUN’S EMPIRE ORDWAY, FREDERICK I. III/ WAKEFORD, RONALD C.

[C/RAN-H]

CONQUEROR (PROOF) [GENERAL 3,4,5]  STIRLING, S. M./DRAKE, DAVID  SR-LP
CONQUEROR FANTASTIC  SARGENT, PAMELA  PA
CONQUEROR WORMS, THE  KEENE, BRIAN  P
CONQUEROR’S CHILD, THE [HOLDFAST 4]  CHARNAS, SUZY MCKEE  H,LP
CONQUERORS FROM THE DARKNESS  SILVERBERG, ROBERT  C/P:2,P
CONQUERORS’ HERITAGE [CONQUERORS 2]  ZAHN, TIMOTHY  C/P,P,H
CONQUERORS’ LEGACY [CONQUERORS 3]  ZAHN, TIMOTHY  H,P
CONQUEROR’S MOON [BOREAL MOON 1]  MAY, JULIAN  H
CONQUERORS OF REGLATHIUM, THE [@DANNUS #2]  SIROTA, MIKE  P
CONQUERORS OF VENUS  KENNEDY, EDGAR REES  SR-P
CONQUERORS’ PRIDE [CONQUERORS 1]  ZAHN, TIMOTHY  C/P,P,H
CONQUEST [@UNICORN AND DRAGON #2]  ABBEY, LYNN  LP
CONQUEST [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #1]  KEYES, J. GREGORY  P
CONQUEST AT QUENDOR [ZORK 4]  MERETZKY, S. ERIC  SFWA-P
CONQUEST OF ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 2]  GREEN, JONATHAN  P
CONQUEST OF EARTH  BANISTER, MANLY  C/P:2,P
CONQUEST OF LIFE  LUKENS, ADAM  H
CONQUEST OF SPACE, THE  LEY, WILLY/BONESTELL, CHESLEY  R/REF-H
CONQUEST OF THE PLANET OF THE APES  JAKES, JOHN  C/P:2,P
CONQUEST OF THE SPACE SEA  WILLIAMS, ROBERT MOORE  DD099
CONQUISTADOR, THE [@CASCA #10]  SADLER, BARRY  P
CONQUISTADOR  STIRLING, S. M.  H
CONRAD’S FATE [CHRESTOMANCI 5]  JONES, DIANA WYNNE  H
CONRAD’S QUEST FOR RUBBER [CONRAD STARGARD 6]  FRANKOWSKI, LEO  C/P,P
CONRAD’S TIME MACHINE [CONRAD STARGARD 0]  FRANKOWSKI, LEO  H
CONSCIENCE INTERPLANETARY  GREEN, JOSEPH  C/P,SFWA,P,H
CONSCIENCE OF THE BEAGLE  ANTHONY, PATRICIA  C/P,SR-LP
CONSCIENCE PLACE  THOMPSON, JOYCE  P,SFWA
CONSCIOUSNESS PLAGUE, THE [PHIL D’AMATO 2]  LEVINSON, PAUL  H
CONSEQUENCES [RETRIEVAL ARTIST 3]  RUSCH, KRISTINE KATHRYN  P
CONSIDER HER WAYS AND OTHERS  WYNDHAM, JOHN  SR-P,C/P
CONSIDER PHLEBAS [CULTURE]  BANKS, IAIN M.  H,P
CONSPICUOUS CONSUMPTION [@WORLD OF DARKNESS WEREWOLF 2]  VON ALLMEN, STEWART  P

TITLE  AUTHOR  CODE
CONSPIRACIES [REPAIRMAN JACK]  WILSON, F. PAUL  H
CONSPIRACY FILES, THE  GREENBERG, MARTIN H./URBAN, SCOTT H.  PA
CONSPIRACY GAME  FEEHAN, CHRISTINE  P
CONSTANTINE  SHIRLEY, JOHN  P
CONSTELLATION OF CATS, A  LITTLE, DENISE  PA
CONSTELLATIONS  CROWTHER, PETER  PA
CONSTELLATIONS [STAR TREK]  PALMIERI, MARCO  LPA
CONSTITUTION [@STAR TREK MY BROTHER’S KEEPER #2]  FRIEDMAN, MICHAEL JAN  P
CONTACT  KEYES, NOEL  PA,C/PA
CONTACT  SAGAN, CARL  C/P,P,H:2
CONTACT AND COMMUNE [ERIS 1]  SMITH, L. NEIL  P
CONTACT IMMINENT [JANI KILIAN 4]  SMITH, KRISTINE  C/P,P
CONTACT IMMINENT (PROOF) [JANI KILIAN 4]  SMITH, KRISTINE  SR-P
CONTAMINATION [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
CONTEMPORARY SCIENCE FICTION, FANTASY, AND HORROR POETRY  GREEN, SCOTT E.  R/REF-H
CONTEXT [NULAPEIRON 2]  MEANEY, JOHN  H
CONTINENT MAKERS AND OTHER TALES OF THE VIAGENS, THE  DE CAMP, L. SPRAGUE  H:2,P
CONTINENT OF LIES, THE  MORROW, JAMES  C/P,P,H
CONTINUOUS KATHERINE MORTENHOE, THE  COMPTON, D. G.  P
CONTINUUM 1  ELWOOD, ROGER  C/PA,PA,SFWA
CONTINUUM 2  ELWOOD, ROGER  HA,PA,SFWA
CONTINUUM 3  ELWOOD, ROGER  C/PA,PA,SFWA
CONTINUUM 4  ELWOOD, ROGER  C/PA,PA
CONTRABAND  FOY, GEORGE  H,P
CONTRABAND FROM OTHERSPACE [GRIMES]  CHANDLER, A. BERTRAM  DG609
CONTRABAND ROCKET  CORREY, LEE  DD146
CONTRARYWISE [CONTRARYWISE 1]  GREENHALGH, ZOHRA  P
CONTROL  GOLDMAN, WILLIAM  P
CONUNDRUM [@DRAGONLANCE AGE OF MORTALS]  CROOK, JEFF  P
CONUNDRUM [DOCTOR WHO NEW ADVENTURES]  LYONS, STEVE  C/P:2,P
CONVENTIONS OF WAR [@DREAD EMPIRE’S FALL 3]  WILLIAMS, WALTER JON  P
CONVERGENCE [@BLENDING #1]  GREEN, SHARON  C/P,P
CONVERGENCE, THE [@LAST ON EARTH #2]  KAYE, MARILYN  P
CONVERGENCE [HERITAGE UNIVERSE 4]  SHEFFIELD, CHARLES  H,P
CONVERGENT SERIES  NIVEN, LARRY  C/P:2,P
CONVERGENT SERIES [HERITAGE UNIVERSE (REVISED) 1,2]  SHEFFIELD, CHARLES  P
CONVERSATIONAL KLINGON (CASSETTE) [STAR TREK]  OKRAND, MARC  SR-MM
CONVERSATIONS  MALZBERG, BARRY N.  H
CONVERSE AND CONFLICT [ERIS 2]  SMITH, L. NEIL  P
COOL WAR, THE  POHL, FREDERIK  C/P,P,H
COPENHAGEN AFFAIR, THE [@MAN FROM UNCLE #3]  ORAM, JOHN  C/P:2,P
COPERNICK’S REBELLION  FRANKOWSKI, LEO  P
COPPA PIENA DI STELLE, UNA  YOUNG, ROBERT F.  FP
COPPER CROWN, THE [KELTIAD 1]  KENNEALY-MORRISON, PATRICIA  C/P:2,SFWA,P
COPPERHEAD  HENDERSON, JAMES  H,P
COPY SHOP, THE  SMITH, EVELYN E.  SFWA
CORAL KINGDOM, THE [@FORGOTTEN REALMS DRUIDHOME #2]  NILES, DOUGLAS  C/P,P
CORALINE  GAIMAN, NEIL  H:2
CORDELIA COLLECTION, THE, VOL. 1 [@BUFFY]  KRULIK, NANCY  P
CORDELIA’S HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1,2]  BUJOLD, LOIS MCMASTER  LP
CORDWAINER SMITH CHECKLIST, A  BENNETT, MIKE  R/IND
CORE  PREUSS, PAUL  C/P:2,P
CORFU AFFAIR, THE [@MAN FROM UNCLE #20]  PHILLIFENT, JOHN T.  P
CORIOLANUS, THE CHARIOT!  YATES, ALAN  P
CORMAC MAC ART  HOWARD, ROBERT E.  P
CORMORANT, THE  GREGORY, STEPHEN  P
CORMYR [@FORGOTTEN REALMS CORMYR #1]  GREENWOOD, ED/GRUBB, JEFF  P
CORNELIUS CHRONICLES, THE [CORNELIUS 1,2,3,4]  MOORCOCK, MICHAEL  C/P,P
CORNELIUS CHRONICLES II, THE [CORNELIUS 5,6]  MOORCOCK, MICHAEL  P
CORNELIUS CHRONICLES III, THE [CORNELIUS 7,8]  MOORCOCK, MICHAEL  P
COROBITE MINES, THE  ROBERTS, PETER  P
CORONA [STAR TREK]  BEAR, GREG  C/P:3,P
CORPORATE MENTALITY  PIZIKS, STEVEN  C/P,P
CORPSE MARKER [DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  P
CORPSEMAN  SHERMAN, JOEL HENRY  P
CORPUS EARTHLING  CHARBONNEAU, LOUIS  SR-P,C/P
CORRESPONDENCE  THOMAS, SUE  H,LP
CORRIDOR, THE  SMITH, A. D.  LP
CORRIDORS OF TIME, THE  ANDERSON, POUL  C/P,H:2,P
CORRUPTED, THE [@WARHAMMER]  EARL, ROBERT  P
CORRUPTING DR. NICE  KESSEL, JOHN  H,LP
CORSAIR  BUNCH, CHRIS  P
CORTEZ ON JUPITER  HOGAN, ERNEST  P,SFWA-P
CORUNDUM’S WOMAN [@SPACEWAYS #2]  CLEVE, JOHN  P

The MITSFS Titledex  September 24, 2007  Page 68
TITLE  AUTHOR  CODE
COSM  BENFORD, GREGORY  C/P:2,P
COSMIC BOMBER, THE [DR. BONES]  WU, WILLIAM F.  P
COSMIC CARNAGE [@STARSHIP ORPHEUS #2]  JADE, SYMON  P
COSMIC CARNIVAL OF STANISLAW LEM, THE  LEM, STANISLAW  H,LP
COSMIC CHECKMATE  DE VET, CHARLES V./MACLEAN, KATHERINE DF149
COSMIC COCKTAILS  LITTLE, DENISE  PA
COSMIC COMPUTER, THE  PIPER, H. BEAM  C/P:2,P
COSMIC CONNECTION, THE  SAGAN, CARL  [C/RAN-P:2]
COSMIC CONQUEST  BLAIR, ADRIAN  SR-P
COSMIC CRUSADE  SAXON, RICHARD  H
COSMIC CRUSADERS  BARBET, PIERRE  C/P:2,P
COSMIC DECOY, THE [@PERRY RHODAN #21]  ACE  PA,C/PA
COSMIC ENCOUNTER  VAN VOGT, A. E.  H,SFWA
COSMIC ENGINEERS  SIMAK, CLIFFORD D.  SR-H,C/P:2,P
COSMIC EXODUS  HOLT, CONRAD G.  SR-P
COSMIC EYE, THE  REYNOLDS, MACK  C/P:2,P
COSMIC FLAME, THE  STATTEN, VARGO  SR-P
COSMIC FORCES OF MU, THE  CHURCHWARD, JAMES  [C/RAN-P]
COSMIC GASH, THE  KAINEN, RAY  C/P,SR-L
COSMIC GEOIDS, THE  TAINE, JOHN  SR-H
COSMIC KALEIDOSCOPE  SHAW, BOB  C/P,SFWA,P

COSMIC KNIGHTS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

COSMIC MANHUNT  DE CAMP, L. SPRAGUE  SR-DD061
COSMIC PUPPETS, THE  DICK, PHILIP K.  DD249
COSMIC RAPE, THE  STURGEON, THEODORE  C/P:2,P
COSMIC TALES: ADVENTURES IN SOL SYSTEM  WEISSKOPF, T. K. F.  PA
COSMIC THUNDER  WATKINS, WILLIAM JOHN  P
COSMIC TRAITOR [@PERRY RHODAN #26]  ACE  PA,C/PA
COSMIC TRIGGER/THE FINAL SECRET OF THE ILLUMINATI  WILSON, ROBERT ANTON  P
COSMICOMICS  CALVINO, ITALO  C/P:2,P
COSMONAUT KEEP [ENGINES OF LIGHT 1]  MACLEOD, KEN  H

COSMOS LATINOS BELL, ANDREA L./

MOLINA-GAVILAN, YOLANDA

LPA

COSMOZOIDS, THE  TRALINS, ROBERT  C/P:2,P
COST OF HONOR, THE [@STARGATE SG-1 #5]  MALCOLM, SALLY  P
COSTIGAN’S NEEDLE  SOHL, JERRY  H:2,P
COUNCIL OF BLADES, THE [@FORGOTTEN REALMS NOBLES #6]  KIDD, PAUL  P
COUNCIL TO SAVE THE PLANET, THE  LAWRENCE, CONRAD  LP
COUNT BRASS [CASTLE BRASS 1]  MOORCOCK, MICHAEL  C/P,P
COUNT GEIGER’S BLUES  BISHOP, MICHAEL  H,LP
COUNT SCAR  BRITTAIN, C. DALE/BOUCHARD, ROBERT A. C/P,P
COUNT ZERO [NEUROMANCER 2]  GIBSON, WILLIAM  SFWA-P,H:2,P,SFWA
COUNTDOWN FOR CINDY  ENGLE, ELOISE  C/P:2,P
COUNTDOWN TO MIDNIGHT  FRANKLIN, H. BRUCE  PA
COUNTER-CLOCK WORLD  DICK, PHILIP K.  C/P:2,P
COUNTERATTACK [@FLEET #2]  DRAKE, DAVID/FAWCETT, BILL  PA,C/PA,SFWA-P
COUNTERCOMMANDMENT  CAMPBELL, JOHN W. JR.  C/PA,PA
COUNTERFEIT HEINLEIN, THE  JANIFER, LAURENCE M.  LP
COUNTERFEIT KINGS  CONNELL, ADAM  LP
COUNTERFEIT MAN, THE  NOURSE, ALAN E.  H:2,P
COUNTERFEITS, THE  KELLEY, LEO P.  P
COUNTERMEASURES [@FORBIDDEN BORDERS #3]  GEAR, W. MICHAEL  C/P:2,P
COUNTERPARTS  ROYLE, NICHOLAS  P
COUNTERPROBE [PROBE 2]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P
COUNTERSOLAR!  LUPOFF, RICHARD A.  H,SFWA,P
COUNTESS DRACULA  PARRY, MICHEL  P
COUNTING HEADS  MARUSEK, DAVID  H
COUNTING THE COST [HAMMER’S SLAMMERS]  DRAKE, DAVID  P,SFWA-P
COUNTING THE EONS  ASIMOV, ISAAC  SFWA
COUNTING UP COUNTING DOWN  TURTLEDOVE, HARRY  LP
COUNTRY LOVE AND POISON RAIN  TATE, PETER  H:2,SFWA
COUNTRY MATTERS  DERLETH, AUGUST  LP
COUNTRY OF THE BLIND AND OTHER STORIES, THE  WELLS, H. G.  P
COURAGE OF FALCONS [SECRET TEXTS 3]  LISLE, HOLLY  H,LP
COURAGEOUS, THE [@STAR TREK DS9 REBELS #2]  AB HUGH, DAFYDD  C/P,P
COURIER  BRAMSTRUP, M. A.  SR-P
COURSE OF EMPIRE, THE  FLINT, ERIC/WENTWORTH, K. D.  H
COURSE OF EMPIRE, THE (PROOF)  FLINT, ERIC/WENTWORTH, K. D.  SR-LP
COURT-MARTIAL OF GEORGE ARMSTRONG CUSTER, THE  JONES, DOUGLAS C.  P
COURT OF A THOUSAND SUNS, THE [STEN 3]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
COURTS OF CHAOS, THE [AMBER 5]  ZELAZNY, ROGER  C/P,@P,H,SFWA
COURTS OF SORCERY, THE [DEMON WARS 3]  MCCONNELL, ASHLEY  P
COURTSHIP OF PRINCESS LEIA, THE [STAR WARS]  WOLVERTON, DAVE  H:2,P

The MITSFS Titledex  September 24, 2007  Page 69
TITLE  AUTHOR  CODE
COURTSHIP RITE  KINGSBURY, DONALD  H:2,SFWA,P
COVEN  RIMMER, STEVEN WILLIAM  P
COVEN, THE  SAINT JOHN, DAVID  H
COVENANT OF JUSTICE, A [PHAESTOR 2]  GERROLD, DAVID  C/P,P
COVENANT OF THE CROWN, THE [STAR TREK]  WEINSTEIN, HOWARD  C/P,P,H
COVENANT OF THE FORGE, THE [@DRAGONLANCE DWARVEN NATIONS #1]  PARKINSON, DAN  C/P,P
COVENANT RISING, THE [DREAMTIME 1]  NICHOLLS, STAN  H,LP
COVENANTS [BORDERLANDS (IVERSTERRE)]  FREEMAN, LORNA  P
COWBOY FENG’S SPACE BAR AND GRILLE  BRUST, STEVEN  C/P:2,P
COWBOY HEAVEN  GOULART, RON  H:2,SFWA
COWL  ASHER, NEAL  H
COYOTE [COYOTE 1]  STEELE, ALLEN  H
COYOTE BLUE  MOORE, CHRISTOPHER  LP
COYOTE COWGIRL  ANTIEAU, KIM  H
COYOTE DREAMS [@WALKER #3]  MURPHY, C. E.  LP
COYOTE FRONTIER [COYOTE 3]  STEELE, ALLEN  H
COYOTE KINGS OF THE SPACE-AGE BACHELOR PAD, THE  FAUST, MINISTER  LP
COYOTE MOON [BUFFY]  VORNHOLT, JOHN  P
COYOTE RISING [COYOTE 2]  STEELE, ALLEN  H
CRAB, THE [FIVE RINGS CLAN WAR 5]  BROWN, STAN  P
CRABS’ MOON [CRABS 2]  SMITH, GUY N.  P
CRABS ON THE RAMPAGE [CRABS 3]  SMITH, GUY N.  C/P,P
CRACHE  BUDZ, MARK  P
CRACHE (PROOF)  BUDZ, MARK  SR-P
CRACK IN SPACE, THE  DICK, PHILIP K.  C/P:2,P
CRACK IN THE SKY, THE  LUPOFF, RICHARD A.  P
CRACKPOT  GOULART, RON  H:2,SFWA
CRADLE  CLARKE, ARTHUR C./LEE, GENTRY  C/P,P,H
CRADLE OF SATURN [CRADLE OF SATURN 1]  HOGAN, JAMES P.  H:2,P
CRADLE OF SPLENDOR  ANTHONY, PATRICIA  P
CRADLE OF THE SUN  STABLEFORD, BRIAN M.  D12140
CRAFT OF LIGHT, THE [@NIGHT-THREADS 4]  EMERSON, RU  P
CRAFT OF SCIENCE FICTION, THE  BRETNOR, REGINALD  SFWA
CRAFTERS, THE  STASHEFF, CHRISTOPHER/FAWCETT, BILL  PA,C/PA,SFWA-P
CRAGHOLD LEGACY, THE  NOONE, EDWINA  P
CRANE, THE [FIVE RINGS CLAN WAR 3]  SOESBEE, REE  P
CRASH  BALLARD, J. G.  P
CRASH DELUXE [PARRISH PLESSIS 3]  DE PIERRES, MARIANNE  P
CRASH GO THE CHARIOTS  WILSON, CLIFFORD  [C/RAN-P]
CRASH LANDING ON IDUNA  TOFTE, ARTHUR  SR-P,C/P
CRASH OF ’79, THE  ERDMAN, PAUL E.  P
CRASH OF 2086, THE  HERSHMAN, MORRIS  P
CRASH OF EMPIRE, THE [@IMPERIAL STARS #3]  POURNELLE, JERRY  SFWA-P
CRASHCOURSE [CRASHCOURSE 1]  BAIRD, WILHELMINA  C/P:2,SFWA-P,P
CRASHING SUNS  HAMILTON, EDMOND  C/P:2,P
CRASHLANDER [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
CRATER LAKE [@DEATHLANDS]  AXLER, JAMES  P
CRATER OF MYSTERY [@NEW TOM SWIFT #8]  APPLETON, VICTOR  LP

CRAVINGS HAMILTON, LAURELL K./

DAVIDSON, MARYJANICE/WILKS, EILEEN/ YORK, REBECCA

H,P

CRAWLING BETWEEN HEAVEN AND EARTH  HOYT, SARAH A.  SR-LP
CRAZY TIME  WILHELM, KATE  C/P,P,H
CREAM OF THE JEST, THE  CABELL, JAMES BRANCH  P
CREATION DESCENDING [@VISION OF BEASTS #1]  LOVEJOY, JACK  C/P,P
CREATION OF DINO DE LAURENTIIS’ KING KONG, THE  BAHRENBURG, BRUCE  R/REF-P,C/REF-P
CREATION OF THE UNIVERSE, THE  GAMOW, GEORGE  [C/RAN-P]
CREATION OF TOMORROW, THE  CARTER, PAUL A.  R/REF-H
CREATOR, THE  HINKEMEYER, M. THOMAS  P
CREATOR  LEVEN, JEREMY  C/P,P,H
CREATURE!  PRONZINI, BILL  LP
CREATURE  SAUL, JOHN  P
CREATURE EPARSE, LA  BRUSS, B. R.  FP
CREATURE FANTASTIC  LITTLE, DENISE  PA
CREATURE FEATURES, UPDATED EDITION  STANLEY, JOHN  R/REF-H
CREATURE FROM BEYOND INFINITY, THE  KUTTNER, HENRY  C/P:2,P
CREATURE FROM THE BLACK LAGOON  DREADSTONE, CARL  P
CREATURES AND DEMONS [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
CREATURES FROM BEYOND  CARR, TERRY  HA
CREATURES OF LIGHT AND DARKNESS  ZELAZNY, ROGER  C/P:3,P
CREATURES OF MAN, THE  MYERS, HOWARD L.  P
CREATURES OF THE ABYSS  LEINSTER, MURRAY  C/P:2,P
CREED FOR THE THIRD MILLENNIUM, A  MCCULLOUGH, COLLEEN  H,P
CREEP, SHADOW!  MERRITT, A.  H

The MITSFS Titledex  September 24, 2007  Page 70
TITLE  AUTHOR  CODE
CREEP, SHADOW CREEP!  MERRITT, A.  SR-P
CREEPING SHROUD, THE  WRIGHT, LAN  SR-P
CREEPS BY NIGHT  HAMMETT, DASHIELL  C/PA,PA
CREEPSHOW  KING, STEPHEN  CX
CRESCENT CITY RHAPSODY [NANOTECH]  GOONAN, KATHLEEN ANN  H,P
CRESCENT IN THE SKY [MECHANICAL SKY 1]  MOFFITT, DONALD  C/P,P
CRETACEOUS SEA [CRETACEOUS SEA 1]  HUBBELL, WILL  P
CREW OF THE MEKONG, THE  VOISKUNSKY, E./LUKODYANOV, I.  P
CREWEL LYE [@XANTH #8]  ANTHONY, PIERS  C/P:2,P
CRIME CAMPAIGN  KUPPERBERG, PAUL  P
CRIME PREVENTION IN THE 30TH CENTURY  SANTESSON, HANS STEFAN  HA
CRIMSON CAPSULE, THE  COBLENTZ, STANTON A.  H:2
CRIMSON CHALICE, THE  CANNING, VICTOR  P
CRIMSON CLAW, THE [@ALIEN CHRONICLES #2]  CHESTER, DEBORAH  P
CRIMSON FLASH, THE  SNELL, ROY J.  H
CRIMSON GODDESS, THE [ARRIS 3]  COOKE, CATHERINE  C/P:2,P
CRIMSON GOLD, THE [@FORGOTTEN REALMS ROGUES]  WHITNEY-ROBINSON, VORONICA  P
CRIMSON LEGION, THE [@DARK SUN PRISM #2]  DENNING, TROY  C/P,P
CRIMSON SERPENT, THE [DOC SAVAGE #78]  ROBESON, KENNETH  C/P:2,@P
CRIMSON SHADOW, THE [CRIMSON SHADOW 1,2,3]  SALVATORE, R. A.  LP

CRIMSON SKIES NYLUND, ERIC S./LEE, MICHAEL B./

BERMAN, NANCY/TRAUTMANN, ERIC S.

PA

CRIMSON SKY, THE [@KEEPERS OF THE HIDDEN WAYS #3]  ROSENBERG, JOEL  P
CRIMSON SPEAR  FROST, GREGORY  LP
CRIMSON SPELL, THE [@CHARMED]  GOLDSBOROUGH, F.  P
CRIMSON SWORD, THE [ASAHIEL 1]  THOMPSON, ELDON  P
CRIMSON TALISMAN, THE [@EBERRON WAR-TORN #1]  COLE, ADRIAN  P
CRIMSON TEARS [@WARHAMMER 40K SOUL DRINKERS 3]  COUNTER, BEN  P
CRIMSON UNIVERSE [@PERRY RHODAN #67]  ACE  PA
CRIMSON WITCH, THE  KOONTZ, DEAN R.  SR-P,C/P:2
CRISIS [@FLEET #6]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
CRISIS  ELWOOD, ROGER  SR-HA
CRISIS!  GUNN, JAMES E.  C/P,SFWA-P,P
CRISIS AT CRYSTAL REEF [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
CRISIS AT STARLIGHT [@BIOWARRIORS #2]  VARDEMAN, ROBERT E.  C/P,P
CRISIS IN 2140  PIPER, H. BEAM/MCGUIRE, JOHN J.  DD227
CRISIS ON CENTAURUS [STAR TREK]  FERGUSON, BRAD  C/P:2,P
CRISIS ON CHEIRON  COULSON, JUANITA  DH027
CRISIS ON CITADEL II [@FLASH GORDON #3]  ANONYMOUS  P
CRISIS ON DOONA [DOONA]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P:2,SFWA-P,P
CRISIS ON VULCAN [@STAR TREK STARFLEET ACADEMY #1]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
CRISPAN MAGICKER  LOWENTHAL, MARK M.  C/P,P
CRISSCROSS [REPAIRMAN JACK]  WILSON, F. PAUL  P
CRITICAL MASS  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
CRITICAL THRESHOLD [DAEDALUS 2]  STABLEFORD, BRIAN M.  C/P:2,P
CRIVIT EXPERIMENT, THE [V]  WOLD, ALLEN L.  C/P,P
CROCK OF GOLD, THE  STEPHENS, JAMES  H,LP
CROISET THE CLAIRVOYANT  POLLACK, JACK HARRISON  [C/RAN-P]
CROMM  FLINT, KENNETH C.  H,P
CROMPTON DIVIDED  SHECKLEY, ROBERT  C/P,P,H
CRONICI MARTIENE  BRADBURY, RAY  FP
CROOK FACTORY, THE  SIMMONS, DAN  P
CROOKED COMPUTER, THE  SHINGLER, WILLIAM G. JR.  P
CROOKED HOUSE, THE  DUNCAN, DAVE  C/P,P

CROOKED HOUSE MONTELEONE, THOMAS F./ DE CHANCIE, JOHN

SFWA-P

CROOKED LETTER, THE [CATACLYSM 1]  WILLIAMS, SEAN  H
CROOKED TREE  WILSON, ROBERT C.  P
CROOKED WORLD, THE [DOCTOR WHO EIGHTH DOCTOR]  LYONS, STEVE  P
CROQUET PLAYER, THE  WELLS, H. G.  H:2
CROSS AND CRESCENT  SHWARTZ, SUSAN  H,LP
CROSS-CURRENTS  ASPRIN, ROBERT LYNN/ABBEY, LYNN  HA
CROSS OF BLOOD [@ALIEN NATION #8]  JETER, K. W.  P
CROSS OF EMPIRE  LAMPTON, CHRISTOPHER  SR-P
CROSS OF FIRE, THE  MALZBERG, BARRY N.  C/P,P
CROSS OF FRANKENSTEIN, THE  MYERS, ROBERT J.  H,P
CROSS OF GOLD AFFAIR, THE [MAN FROM UNCLE 14]  DAVIES, FREDRIC  C/P:2,P
CROSS THE STARS  DRAKE, DAVID  C/P:2,P
CROSS-TIME ENGINEER, THE [CONRAD STARGARD 1]  FRANKOWSKI, LEO  C/P:2,P
CROSSFIRE [@WARHAMMER 40K SHIRA CALPURNIA 1]  FARRER, MATTHEW  P
CROSSFIRE [CROSSFIRE 1]  KRESS, NANCY  H
CROSSFIRE [@STAR TREK TNG STARFLEET ACADEMY #11]  VORNHOLT, JOHN  LP
CROSSING INFINITY  HABER, KAREN  H
CROSSING THE LINE [WESS’HAR WARS 2]  TRAVISS, KAREN  P

TITLE  AUTHOR  CODE
CROSSINGS [@BUFFY]  ODOM, MEL  P
CROSSOVER [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  P
CROSSOVER [@RIFTWORLD 1]  MCCAY, BILL  C/P:2,SFWA-P,P
CROSSOVER [CASSANDRA KRESNOV 1]  SHEPHERD, JOEL  LP
CROSSROAD [STAR TREK]  HAMBLY, BARBARA  C/P:2,P
CROSSROADS  COX, F. BRET/DUNCAN, ANDY  HA
CROSSROADS [SHADOWRUN 36]  KENSON, STEPHEN  P
CROSSROADS [VALDEMAR]  LACKEY, MERCEDES  PA
CROSSROADS IN TIME  CONKLIN, GROFF  C/PA,PA
CROSSROADS OF TIME, THE [CROSSTIME]  NORTON, ANDRE  C/P:2,DD164,P
CROSSROADS OF TWILIGHT [WHEEL OF TIME 10]  JORDAN, ROBERT  H
CROSSTIME TRAFFIC  WATT-EVANS, LAWRENCE  C/P,P
CROSSWAYS [@DEATHLANDS]  AXLER, JAMES  P
CROUCHING TIGER HIDDEN DRAGON, VOL. 5  SETO, ANDY  SCX
CROWN AND THE SWORD, THE [@DRAGONLANCE RISE OF SOLAMNIA #2]  NILES, DOUGLAS  P
CROWN DISOWNED, A [OAK,YEW,ASH,AND ROWAN 3]  NORTON, ANDRE/MILLER, SASHA  H
CROWN DUEL  SMITH, SHERWOOD  P
CROWN JEWELS, THE [DRAKE MAIJSTRAL 1]  WILLIAMS, WALTER JON  C/P:2,P
CROWN OF DALEMARK, THE [DALEMARK 4]  JONES, DIANA WYNNE  C/P,P
CROWN OF EMPIRE [CRISIS OF EMPIRE 4]  YARBRO, CHELSEA QUINN  P
CROWN OF FIRE [@FORGOTTEN REALMS HARPERS #9/SHANDRIL 2]  GREENWOOD, ED  C/P:2,P
CROWN OF FIRE [FIREBIRD 3]  TYERS, KATHY  LP
CROWN OF HORN  COOPER, LOUISE  P
CROWN OF INFINITY  FAUCETTE, JOHN M.  DH051
CROWN OF KINGS, THE [@SORCERY! #4]  JACKSON, STEVE  [C/RAN-P]
CROWN OF SHADOWS [COLDFIRE 3]  FRIEDMAN, C. S.  H,P
CROWN OF SILENCE, THE [MAGRAVANDIAS 2]  CONSTANTINE, STORM  H:2
CROWN OF SLAVES [HONOR HARRINGTON]  WEBER, DAVID/FLINT, ERIC  H
CROWN OF STARS [CROWN OF STARS 7]  ELLIOTT, KATE  H
CROWN OF STARS  TIPTREE, JAMES JR.  C/P,SFWA-P,P,H
CROWN OF SWORDS, A [WHEEL OF TIME 7]  JORDAN, ROBERT  H:2,P
CROWN OF THE SERPENT  WOLD, ALLEN L.  C/P,P
CROWN OF THE SWORD GOD  NORVIL, MANNING  P
CROWN OF THIEVES [@DRAGONLANCE NEW ADVENTURES ELIDOR #1]  SOESBEE, REE  LP
CROWN ROSE, THE  AVERY, FIONA  H
CROYD  WALLACE, IAN  C/P:2,P
CRUACHAN AND THE KILLANE, THE  CRISTABEL  P
CRUCIBLE [FOX 2]  CHASE, ROBERT R.  P
CRUCIBLE [CROSSFIRE 2]  KRESS, NANCY  H
CRUCIBLE [@ROGUE TROOPER]  RENNIE, GORDON  P
CRUCIBLE OF POWER, THE [@COLLECTED STORIES JACK WILLIAMSON #5]  WILLIAMSON, JACK  H
CRUCIBLE OF TIME [@DEATHLANDS]  AXLER, JAMES  C/P,P
CRUCIBLE OF TIME, THE  BRUNNER, JOHN  C/P:2,P,H
CRUCIBLE OF WAR [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA

CRUCIBLE: THE TRIAL OF CYRIC THE MAD [@FORGOTTEN REALMS AVATAR 5] DENNING, TROY

P

CRUEL MIRACLES  CARD, ORSON SCOTT  P
CRUISER DREAMS [KERRION EMPIRE 2]  MORRIS, JANET  H:2,P
CRUNCH BUNCH, THE  BISCHOFF, DAVID F.  P
CRUSADE [FORGOTTEN REALMS EMPIRES 3]  LOWDER, JAMES  C/P:2,P
CRUSADE [STARFIRE 2]  WEBER, DAVID/WHITE, STEVE  C/P:2,SFWA-P,P
CRUSADE FOR ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 1]  GREEN, JONATHAN  P
CRUSADE OF FIRE  KURTZ, KATHERINE  PA
CRUSADER [WAYFARER REDEMPTION 6]  DOUGLASS, SARA  H
CRUSADERS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA
CRUSADER’S TORCH [SAINT-GERMAIN OLIVIA 2]  YARBRO, CHELSEA QUINN  C/P,SFWA-P,P,H:2
CRUSADING WIZARD, THE [WIZARD IN RHYME 7]  STASHEFF, CHRISTOPHER  C/P,P
CRY FOR THE STRANGERS  SAUL, JOHN  P
CRY HORROR!  LOVECRAFT, H. P.  C/P,P
CRY OF THE BEAST  MEYERS, RICHARD S.  P
CRY OF THE ONLIES, THE [STAR TREK]  KLASS, JUDY  C/P:2,P
CRY PLAGUE!  DRACHMAN, THEODORE S.  SR-DD013
CRY REPUBLIC [PROCURATOR 3]  MITCHELL, KIRK  C/P,SFWA-P,P
CRY SILVER BELLS  SWANN, THOMAS BURNETT  C/P:2,P
CRY WOLF  CHRONISTER, ALAN B.  P
CRY WOLF [@K-9 CORPS 3]  VON GUNDEN, KENNETH  P
CRYGENDER  THOMAS, THOMAS T.  C/P,SFWA-P,P
CRYING OF LOT 49, THE  PYNCHON, THOMAS  C/P:2,P
CRYING SHAME, A  JOHNSTONE, WILLIAM W.  P
CRYPT OF THE SHADOWKING [@FORGOTTEN REALMS HARPERS #6]  ANTHONY, MARK  P
CRYPTOGRAM, THE  VERNE, JULES  H
CRYPTONOMICON  STEPHENSON, NEAL  H:3,LP
CRYPTOZOIC!  ALDISS, BRIAN W.  H:2,P
CRYSTAL AGE, A  HUDSON, W. H.  H

The MITSFS Titledex  September 24, 2007  Page 72
TITLE  AUTHOR  CODE
CRYSTAL BALL, THE  SPICER, DOROTHY  P
CRYSTAL BUCEPHALUS, THE [DOCTOR WHO MISSING ADVENTURES]  HINTON, CRAIG  C/P:2,P
CRYSTAL CAVE, THE [ARTHURIAN 1]  STEWART, MARY  C/P:2,P,LP
CRYSTAL CHALICE, THE [@DRAGONLANCE NEW ADVENTURES ELIDOR #2]  SOESBEE, REE  LP
CRYSTAL CITY, THE [ALVIN MAKER 6]  CARD, ORSON SCOTT  H
CRYSTAL CLOUDS, THE [@JADE DEMONS #3]  VARDEMAN, ROBERT E.  P
CRYSTAL CROWN, THE [AVERIDAN 1]  CLOUGH, BRENDA W.  C/P:2,P
CRYSTAL DESERT, THE [@GUARDIAN #3]  GRAY, JULIA  P
CRYSTAL DRAGON, THE [DRAGONREALM 5]  KNAAK, RICHARD A.  C/P,P
CRYSTAL DRAGON [LIADEN UNIVERSE GREAT MIGRATION 2]  LEE, SHARON/MILLER, STEVE  H
CRYSTAL EMPIRE, THE  SMITH, L. NEIL  H,P
CRYSTAL EXPRESS  STERLING, BRUCE  C/P:2,SFWA-P:2,P
CRYSTAL EYE, THE [@ALIEN CHRONICLES #3]  CHESTER, DEBORAH  P
CRYSTAL GORGE [@DREAMERS #3]  EDDINGS, DAVID/EDDINGS, LEIGH  H
CRYSTAL GRYPHON, THE [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE  C/P:2,SFWA-P,P
CRYSTAL HEAT [@SHADOWSONG #3]  CLAYTON, JO  C/P,P
CRYSTAL HORDE, THE  TAINE, JOHN  SR-H
CRYSTAL LINE [CRYSTAL SINGER 3]  MCCAFFREY, ANNE  H:2,P
CRYSTAL MAN, THE  MITCHELL, EDWARD PAGE  H,SFWA
CRYSTAL MEMORY, THE  LEIGH, STEPHEN  H:2,P
CRYSTAL PALACE, THE [CRAY ORMORU 2]  EISENSTEIN, PHYLLIS  C/P:2,P
CRYSTAL PHOENIX  BERLYN, MICHAEL  P
CRYSTAL PRINCE, THE [@ATLAN #4]/WAR OF THE GHOSTS [ATLAN 5]  ACE  PA
CRYSTAL PRISON, THE [DEPTFORD MICE 2]  JARVIS, ROBIN  LP
CRYSTAL RAIN  BUCKELL, TOBIAS S.  H
CRYSTAL ROSE, THE [MERI 3]  BOHNHOFF, MAYA KAATHRYN  P
CRYSTAL SAGE  DALKEY, KARA  C/P,P
CRYSTAL SEAS, THE [@BLADE #16]  LORD, JEFFREY  P
CRYSTAL SHARD, THE [@FORGOTTEN REALMS ICEWIND DALE #1]  SALVATORE, R. A.  C/P:2,P
CRYSTAL SHIP, THE  SILVERBERG, ROBERT  C/PA,PA
CRYSTAL SINGER [CRYSTAL SINGER 1]  MCCAFFREY, ANNE  C/P:2,P,H:3
CRYSTAL SOLDIER [LIADEN UNIVERSE GREAT MIGRATION 1]  LEE, SHARON/MILLER, STEVE  H
CRYSTAL SORCERORS, THE [CRYSTAL WARRIORS 2]  FORSTCHEN, WILLIAM R./MORRISON, GREG P
CRYSTAL STAR, THE [@STAR WARS]  MCINTYRE, VONDA N.  C/P:2,P
CRYSTAL SWORD, THE  MARTINE-BARNES, ADRIENNE  C/P,P
CRYSTAL WARRIORS, THE [CRYSTAL WARRIORS 1]  FORSTCHEN, WILLIAM R./MORRISON, GREG C/P:2,P
CRYSTAL WITNESS  TYERS, KATHY  C/P,P
CRYSTAL WORLD, THE  BALLARD, J. G.  H:2,P
CRYSTAL WORLD  HAIBLUM, ISIDORE  P
CRYSTALS OF AIR AND WATER [PARSINA 3]  GOLDIN, STEPHEN  P
CRYSTALS OF MIDA, THE [JALAV 1]  GREEN, SHARON  C/P,P
CTHULHU 2000  TURNER, JIM  HA,LPA
CTHULHU MYTHOS, THE  DERLETH, AUGUST  H
CTHULHU: THE MYTHOS AND KINDRED HORRORS  HOWARD, ROBERT E.  C/P:2,P
CTZ PARADIGM, THE  FRANCOIS, YVES REGIS  H,SFWA
CUBE ROOT OF UNCERTAINTY, THE  SILVERBERG, ROBERT  C/P,P,H
CUCKOO TREE, THE  AIKEN, JOAN  P
CUCKOO’S BOYS, THE  REED, ROBERT  H
CUCKOO’S EGG  CHERRYH, C. J.  H:2,P
CUGEL THE CLEVER [@VANCE INTEGRAL EDITION #15]  VANCE, JACK  SR-H

CUGEL: THE SKYBREAK [@VANCE INTEGRAL EDITION #35]/ SPATTERLIGHT [@VANCE INTEGRAL EDITION #35] VANCE, JACK

SR-H

CUGEL’S SAGA [DYING EARTH 3]  VANCE, JACK  C/P,P,H
CUJO  KING, STEPHEN  C/P:2,P
CULAI HERITAGE, THE [TALES OF THE BARD 1,2,3]  SCOTT, MICHAEL  LP
CULT OF LOVING KINDNESS, THE [STARBRIDGE CHRONICLES 3]  PARK, PAUL  H,P
CULTURAL BREAKS  ALDISS, BRIAN W.  H
CULTURES BEYOND THE EARTH [MATT CRUSE 1]  MARUYAMA, MAGOROH/HARKINS, ARTHUR R/REF-P
CUNNING BLOOD, THE  DUNTEMANN, JEFF  H
CUNNINGHAM EQUATIONS, THE  EDMONDSON, G. C./KOTLAN, C. M.  C/P:2,P
CUP OF CLAY [TALISWOMAN 1]  DOUGLAS, CAROLE NELSON  H:2,P
CUP OF MORNING SHADOWS, THE [TWELVE TREASURES 2]  EDGHILL, ROSEMARY  C/P:2,P
CUP OF THANATOS, THE [PAUL HOLTON 2]  HUNT, CHARLOTTE  P
CUPFUL OF SPACE, A  CLINGERMAN, MILDRED  C/P:2,P
CURE FOR CANCER, A [CORNELIUS 2]  MOORCOCK, MICHAEL  H
CURE FOR DEATH  VALENTINE, VICTOR  SR-P
CURIOUS QUESTS OF BRIGADIER FFELLOWES, THE  LANIER, STERLING E.  SR-H,H
CURRANT EVENTS [@XANTH #28]  ANTHONY, PIERS  H
CURRENTS OF SPACE, THE [GALACTIC EMPIRE 2]  ASIMOV, ISAAC  C/P:2,P,H
CURSE, THE [@DARK MOON #1]  BLAIR, CYNTHIA  P
CURSE [@SMALLVILLE]  GRANT, ALAN  P
CURSE, THE  GRANT, CHARLES L.  P
CURSE, THE  WEINBERG, LARRY  P
CURSE OF ARKADY, THE [@MAGICKERS #2]  DRAKE, EMILY  P

The MITSFS Titledex  September 24, 2007  Page 73
TITLE  AUTHOR  CODE
CURSE OF CHALION, THE [CHALION 1]  BUJOLD, LOIS MCMASTER  H:2
CURSE OF COLLINWOOD, THE  ROSS, MARILYN  C/P,P
CURSE OF LEO, THE [@HORRORSCOPE #3]  LORY, ROBERT  P
CURSE OF LOCH NESS, THE  TREMAYNE, PETER  SR-P
CURSE OF PLANET KUZ  VAN LODEN, ERLE  SR-P
CURSE OF QUINTANA ROO, THE  GARDNER, MATT  P
CURSE OF RATHLAW, THE [GUARDIANS (SAXON) 0]  SAXON, PETER  C/P,P
CURSE OF SAGAMORE, THE [SAGAMORE 1]  DALKEY, KARA  P
CURSE OF SLAGFID, THE [WIZARD’S WAR 2]  BOYER, ELIZABETH H.  P
CURSE OF THE BLACK HERON [BARD’S TALE 8]  LISLE, HOLLY  C/P:2,P
CURSE OF THE BLUE FIGURINE, THE  BELLAIRS, JOHN  H,P
CURSE OF THE GIANT HOGWEED, THE [PROFESSOR SHANDY 5]  MACLEOD, CHARLOTTE  P
CURSE OF THE LOST GROVE [@KNIGHTS OF THE SILVER DRAGON #10]  GRAHAM, DENISE B.  LP
CURSE OF THE MISTWRAITH [WARS OF LIGHT AND SHADOW 1]  WURTS, JANNY  C/P:3,SFWA,P
CURSE OF THE OBELISK, THE  GOULART, RON  C/P:2,P
CURSE OF THE SHADOWMAGE [@FORGOTTEN REALMS HARPERS #11]  ANTHONY, MARK  P
CURSE OF THE UNDEAD, THE  CARTER, M. L.  PA
CURSE OF THE VAMPIRE  CAINE, GEOFFREY  P
CURSE OF THE WITCH-QUEEN, THE  VOLSKY, PAULA  C/P:2,P
CURSE OF WHISPERING HILLS, THE [WHISPERING HILLS 2]  CAMERON, KATE  P
CURSE THE DARK [RETRIEVERS]  GILMAN, LAURA ANNE  H,LP
CURSED, THE  DUNCAN, DAVE  P
CURSED [@BUFFY/ANGEL]  ODOM, MEL  C/P,P
CURSED, THE [@CASCA #18]  SADLER, BARRY  P
CURSED LAND, THE [MAGIC: THE GATHERING]  MCLAREN, TERI  P
CURSED TOWERS, THE [WITCHES OF EILEANAN 3]  FORSYTH, KATE  C/P:2,P

CURSES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

CURSOR’S FURY [@CODEX ALERA #3]  BUTCHER, JIM  H:2
CURVE OF THE SNOWFLAKE, THE  WALTER, W. GREY  H
CUSP  METZGER, ROBERT A.  H
CUTTHROAT, THE [@WARLORD #2]  FROST, JASON  P
CUTTHROAT CANNIBALS, THE [@LAST RANGER #8]  SARGENT, CRAIG  P
CUTTING EDGE, THE [HANDFUL OF MEN 1]  DUNCAN, DAVE  C/P:2,P
CUTTING EDGE  ETCHISON, DENNIS  SFWA
CV [CV 1]  KNIGHT, DAMON  C/P,SFWA-P,P,H
CYBELE, WITH BLUEBONNETS  HARNESS, CHARLES L.  H
CYBER WAY  FOSTER, ALAN DEAN  C/P:2,SFWA-P:2,P
CYBERBOOKS  BOVA, BEN  C/P,SFWA-P,P,H
CYBERIAD, THE  LEM, STANISLAW  H:2,SFWA:2,P
CYBERNETIC BRAINS, THE  JONES, RAYMOND F.  C/P,P,H
CYBERNETIC CONTROLLER  CLARKE, A. V./BULMER, H. K.  SR-P
CYBERNETIC IMAGINATION IN SCIENCE FICTION, THE  WARRICK, PATRICIA S.  R/REF-H
CYBERNETIC JUNGLE  LEWITT, SHARIANN  C/P,P
CYBERNETIC SAMURAI, THE [CYBERNETIC SAMURAI 1]  MILAN, VICTOR  H:2,SFWA,P
CYBERNETIC SHOGUN, THE [CYBERNETIC SAMURAI 2]  MILAN, VICTOR  H
CYBERNETIC WALRUS, THE [@WONDERLAND GAMBIT #1]  CHALKER, JACK L.  C/P,SFWA-P,LP,H:2
CYBERNIA  CAMERON, LOU  C/P:2,P
CYBERPUNK HANDBOOK  ST. JUDE/SIRIUS, R. U./NAGEL, BART  [C/RAN-P]
CYBERSONG [STAR TREK VOYAGER]  LEWITT, SHARIANN  P
CYBERSTEALTH [CARGO AND GHOSTER 1]  LEWITT, SHARIANN  P
CYBERWEB  MASON, LISA  C/P,P
CYBORG [@CYBORG #1]  CAIDIN, MARTIN  C/P:2,P
CYBORG [@ROBOT CITY #3]  WU, WILLIAM F.  C/P:2,SFWA-P,P
CYBORG AND THE SORCERERS, THE [WAR SURPLUS 1]  WATT-EVANS, LAWRENCE  C/P:2,P
CYBORG FROM EARTH, THE  SHEFFIELD, CHARLES  C/P,P,H
CYBORG IV [SIX MILLION DOLLAR MAN 6]  CAIDIN, MARTIN  H,@P
CYBORG KICKBOXER [@TOM SWIFT (ARCHWAY) #3]  APPLETON, VICTOR  P
CYBORG KING, THE [STAR HAWKS 2]  GOULART, RON  P
CYCLE OF FIRE  CLEMENT, HAL  C/P:2,P
CYCLE OF FIRE, THE [CYCLE OF FIRE 1,2,3]  WURTS, JANNY  LP
CYCLE OF HATRED [WORLD OF WARCRAFT]  DECANDIDO, KEITH R. A.  P
CYCLE OF NEMESIS  BULMER, KENNETH  C/P:2,P
CYCLE OF THE WEREWOLF  KING, STEPHEN  SR-LP
CYCLOPS  CUSSLER, CLIVE  P
CYGNET AND THE FIREBIRD, THE [CYGNET 2]  MCKILLIP, PATRICIA A.  SFWA
CYLON DEATH MACHINE, THE [@BATTLESTAR GALACTICA 1978 #2]  LARSON, GLEN A./THURSTON, ROBERT  C/P,P
CYLONS’ SECRET, THE [@BATTLESTAR GALACTICA 2003]  GARDNER, CRAIG SHAW  LP
CYLON’S SECRET, THE (PROOF) [@BATTLESTAR GALACTICA 2003]  GARDNER, CRAIG SHAW  SR-LP
CYRION  LEE, TANITH  C/P,P
CYTEEN [UNION/ALLIANCE CYTEEN 1,2,3]  CHERRYH, C. J.  H
CYTEEN: THE BETRAYAL [UNION/ALLIANCE CYTEEN 1]  CHERRYH, C. J.  C/P:2,P
CYTEEN: THE REBIRTH [UNION/ALLIANCE CYTEEN 2]  CHERRYH, C. J.  C/P:2,P
CYTEEN: THE VINDICATION [UNION/ALLIANCE CYTEEN 3]  CHERRYH, C. J.  C/P:2,P

TITLE  AUTHOR  CODE
CYTHERA  CALDER, RICHARD  H
CZAR OF FEAR, THE [DOC SAVAGE #22]  ROBESON, KENNETH  C/P:2,@P

The MITSFS Titledex  September 24, 2007  Page 75
TITLE  AUTHOR  CODE
D-99  FYFE, H. B.  C/P:2,P
D.A.  WILLIS, CONNIE  H
DA VINCI CODE, THE  BROWN, DAN  H
DA VINCI LEGACY, THE  PERDUE, LEWIS  P
DA VINCI MACHINE, THE  CONRAD, EARL  C/P,P,H
DACOBRA, OR THE WHITE PRIESTS OF AHRIMAN  BURLAND, HARRIS  H
DAD’S NUKE  LAIDLAW, MARC  H
DAEDALUS’S CHILDREN, PART TWO [STAR TREK ENTERPRISE]  STERN, DAVE  P
DAEMON WORLD [@WARHAMMER 40K]  COUNTER, BEN  P
DAEMONMANIA  CROWLEY, JOHN  H
DAEMON’S CURSE, THE [@WARHAMMER MALUS DARKBLADE 1]  ABNETT, DAN/LEE, MIKE  P
DAEMONS, INC. [KEVIN BRADLEY AND EVAN DAVIS 1,2]  BACON-SMITH, CAMILLE  H
DAEMONSLAYER [WARHAMMER GOTREK & FELIX 3]  KING, WILLIAM  P
DAGGER [THIEVES’ WORLD]  DRAKE, DAVID  C/P,SFWA-P,P,H
DAGGER AFFAIR, THE [@MAN FROM UNCLE #4]  MCDANIEL, DAVID  C/P:2,P
DAGGER AND THE CROSS, THE [ALAMUT 2]  TARR, JUDITH  H,P
DAGGER IN THE SKY, THE [DOC SAVAGE #40]  ROBESON, KENNETH  C/P:2,@P

DAGGER MAGIC [ADEPT 4] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,P,H

DAGGER OF DOOM, THE [@KNIGHTS OF THE SILVER DRAGON #6]  ROBERTS, KERRY DANIEL  LP
DAGGER OF THE MIND  SHAW, BOB  C/P,P
DAGGER POINT [@BATTLETECH]  GRESSMAN, THOMAS S.  P
DAGGER’S EDGE [SHADOW 4]  LOGSTON, ANNE  P
DAGGER’S POINT [SHADOW 5]  LOGSTON, ANNE  P
DAGGERSPELL [DEVERRY 1/DEVERRY PROPER 1]  KERR, KATHARINE  H:2,SFWA,P
DAGGERSPELL (REVISED) [DEVERRY 1/DEVERRY PROPER 1]  KERR, KATHARINE  C/P:2,P
DAGON AND OTHER MACABRE TALES  LOVECRAFT, H. P.  SR-H,H
DAHUT [@KING OF YS 3]  ANDERSON, POUL/ANDERSON, KAREN  C/P,SFWA-P,P
DAI-SAN [SUNSET WARRIOR 3]  LUSTBADER, ERIC VAN  C/P,SFWA,P,H
DAI-SHO  OLDEN, MARC  H

DAIMON (LATIN)

CASE, RICHARD D./GESS, NICHOLAS M./ ROGERSON, EDWARD G./ HIRSCHLER, PHILIP/ SALTONSTALL, WILLIAM L. JR.

FH

DAJA’S BOOK [@CIRCLE OF MAGIC #3]  PIERCE, TAMORA  P
DAKOTA PROJECT, THE  BEECHING, JACK  P
DALAMAR THE DARK [@DRAGONLANCE CLASSICS]  BERBERICK, NANCY VARIAN  P
DALEK FACTOR, THE [DOCTOR WHO]  CLARK, SIMON  SR-H
DALETH EFFECT, THE  HARRISON, HARRY  C/P,P,H
DAM OF DEATH [INVADERS]  PEARL, JACK  H
DAMAGED GOODS [DOCTOR WHO NEW ADVENTURES]  DAVIES, RUSSELL T.  C/P,P
DAMIA [ROWAN 2]  MCCAFFREY, ANNE  C/P,SFWA-P:2,P,H
DAMIANO [DAMIANO 1]  MACAVOY, R. A.  C/P:2,P
DAMIANO’S LUTE [DAMIANO 2]  MACAVOY, R. A.  C/P:2,P
DAMIA’S CHILDREN [ROWAN 3]  MCCAFFREY, ANNE  C/P,SFWA-P,P,H
DAMIEN — OMEN II [OMEN 2]  HOWARD, JOSEPH  P
DAMNATION ALLEY  ZELAZNY, ROGER  C/P:2,P
DAMNATION GAME, THE  BARKER, CLIVE  H:2,P
DAMNATIONS  LEMING, R. L.  SR-VLPA
DAMNBANNA  SPRINGER, NANCY  SR-H
DAMNED, THE [@CASCA #7]  SADLER, BARRY  P
DAMNED DISCIPLES, THE [@LAST RANGER #9]  SARGENT, CRAIG  P
DAMNED NATION [@BUREAU 13]  POLLOTTA, NICK  LP
DAMSEL UNDER STRESS [KATIE CHANDLER 3]  SWENDSON, SHANNA  LP
DAN BARRY’S DAUGHTER  BRAND, MAX  P
DAN OWEN AND THE ANGEL JOE  MITCHELL, RONALD  H
DANCE BAND ON THE TITANIC  CHALKER, JACK L.  C/P:2,P
DANCE DOWN THE STARS [@DANCER #3]  CLAYTON, JO  C/P:2,P
DANCE OF DEATH [JONATHAN BARRETT 4]  ELROD, P. N.  C/P,P
DANCE OF DEATH [TWILIGHT 19]  KASEM, LOU  P
DANCE OF DEMONS [GORD THE ROGUE]  GYGAX, GARY  C/P,P
DANCE OF KNIVES  MCMAHON, DONNA  LP
DANCE OF THE APOCALYPSE  EKLUND, GORDON  SR-P
DANCE OF THE DEAD [@RAVENLOFT]  GOLDEN, CHRISTIE  P
DANCE OF THE GODS [CIRCLE 2]  ROBERTS, NORA  H
DANCE OF THE HAG  LEIGH, STEPHEN  C/P:2,P
DANCE OF THE TIGER  KURTEN, BJORN  C/P:2,P

DANCE OF THE VOODOO HANDBAG, THE [TRILOGY THAT DARE NOT SPEAK ITS NAME 2] RANKIN, ROBERT

P

DANCE OF TIME, THE [BELISARIUS 6]  FLINT, ERIC/DRAKE, DAVID  H
DANCE TO THE SUN  ESRAC, WILLIAM  P
DANCE WITH A VAMPIRE (PROOF) [@VAMPIRE KISSES #4]  SHCREIBER, ELLEN  SR-LP
DANCE WITH THE DEVIL [DARK-HUNTER]  KENYON, SHERRILYN  P
DANCER FROM ATLANTIS, THE  ANDERSON, POUL  H:2,P

The MITSFS Titledex  September 24, 2007  Page 76
TITLE  AUTHOR  CODE
DANCER OF GOR [GOR 22]  NORMAN, JOHN  P
DANCER OF THE SIXTH  CREAN, MICHELLE SHIREY  P
DANCER’S ILLUSION  MAXWELL, ANN  C/P:2,P
DANCERS IN THE AFTERGLOW  CHALKER, JACK L.  C/P:2,P
DANCER’S LUCK  MAXWELL, ANN  C/P:2,P
DANCERS OF ARUN, THE [TORNOR 2]  LYNN, ELIZABETH A.  C/P,P,H,SFWA
DANCERS OF NOYO, THE  SAINT CLAIR, MARGARET  C/P,P
DANCER’S RISE [@DANCER #1]  CLAYTON, JO  C/P:2,P
DANCING AT THE EDGE OF THE WORLD  LE GUIN, URSULA K.  R/REF-H
DANCING BEARS  SABERHAGEN, FRED  H:2,P
DANCING FLOOR, THE  BUCHAN, JOHN  H
DANCING JACK  MARKS, LAURIE J.  P
DANCING NAKED  TENN, WILLIAM  H
DANCING ON AIR  KRESS, NANCY  LP
DANCING THE CODE [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P:2,P
DANCING VAC [CARGO AND GHOSTER 2]  LEWITT, SHARIANN  P
DANCING WITH MYSELF  SHEFFIELD, CHARLES  C/P:2,P
DANDELION WINE  BRADBURY, RAY  C/P,P,H
DANGER BELOW! [@RICK BRANT #23]  BLAINE, JOHN  H
DANGER: DINOSAURS!  MARSTEN, RICHARD  H:2
DANGER FROM VEGA  RACKHAM, JOHN  DG576
DANGER IN DEEP SPACE [TOM CORBETT 2]  ROCKWELL, CAREY  H
DANGER IN THE ASHES [ASHES 8]  JOHNSTONE, WILLIAM W.  P
DANGER PLANET  STERLING, BRETT  C/P:2,P
DANGER...DINOSAURES!  MARSTEN, RICHARD  FP
DANGEROUS GAMES [RILEY JENSON 4]  ARTHUR, KERI  P
DANGEROUS GAMES [@FORGOTTEN REALMS ARCANE AGE NETHERIL #2]  EMERY, CLAYTON  P
DANGEROUS GAMES  RANDALL, MARTA  C/P:2,P
DANGEROUS INTERFACES [@TIME GATE #2]  SILVERBERG, ROBERT/FAWCETT, BILL  PA,C/PA,SFWA-P
DANGEROUS MAGIC, A  LITTLE, DENISE  PA
DANGEROUS SPACE  ESKRIDGE, KELLEY  LP

DANGEROUS VEGETABLES WAUGH, CHARLES G./

GREENBERG, MARTIN H.

C/PA,PA

DANGEROUS VISIONS  ELLISON, HARLAN  HA,PA
DANGEROUS VISIONS #1  ELLISON, HARLAN  PA
DANGEROUS VISIONS #2  ELLISON, HARLAN  PA
DANGEROUS VISIONS #3  ELLISON, HARLAN  PA
DANIEL’S VEIL  STAVIS, R. H.  P
DANILOV THE VIOLIST  ORLOV, VLADIMIR  H
DANNY DUNN AND THE ANTI-GRAVITY PAINT [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN AND THE AUTOMATIC HOUSE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE FOSSIL CAVE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN AND THE HEAT RAY [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE HOMEWORK MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE SMALLIFYING MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE SWAMP MONSTER [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE UNIVERSAL GLUE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN AND THE VOICE FROM SPACE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN AND THE WEATHER MACHINE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P,H
DANNY DUNN, INVISIBLE BOY [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN ON A DESERT ISLAND [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANNY DUNN ON THE OCEAN FLOOR [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN, SCIENTIFIC DETECTIVE [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P:2,P
DANNY DUNN, TIME TRAVELER [DANNY DUNN]  WILLIAMS, JAY/ABRASHKIN, RAYMOND  C/P,P
DANSE MACABRE [ANITA BLAKE 14]  HAMILTON, LAURELL K.  H:2
DANSE MACABRE  KING, STEPHEN  R/REF-P,C/REF-H
DANTE’S DISCIPLES  CROWTHER, PETER/KRAMER, EDWARD E.  PA
DANTE’S EQUATION  JENSEN, JANE  LP
DANUBE PILOT, THE  VERNE, JULES  H
DANZATORI DEL CREPUSCOLO, I (ITALIAN)  CHALKER, JACK L.  FP
DAPHNE DU MAURIER’S CLASSICS OF THE MACABRE  DU MAURIER, DAPHNE  SFWA
DARE  FARMER, PHILIP JOSE  SR-P,C/P:2
DARE TO GO A-HUNTING [MOON SINGER 4]  NORTON, ANDRE  H:2,SFWA-P,P
DAREDEVIL  COX, GREG  H
DAREDEVILS, LTD. [EXCHAMELEON 1]  GOULART, RON  C/P:2,P
DARGONESTI, THE [@DRAGONLANCE LOST HISTORIES #3]  THOMPSON, PAUL B./COOK, TONYA C.  P
DARK, THE  DATLOW, ELLEN  LPA
DARK, THE  FRANKLIN, MAX  P
DARK, THE  HERBERT, JAMES  P
DARK ABYSS, THE [DUNGEON 2]  COVILLE, BRUCE  C/P:2,P
DARK ADEPTUS [@WARHAMMER 40K GREY KNIGHTS 2]  COUNTER, BEN  P
DARK ALLIES [@STAR TREK NEW FRONTIER #8]  DAVID, PETER  P
DARK AND HUNGRY GOD ARISES, A [GAP 3]  DONALDSON, STEPHEN R.  H:2,@P
DARK ANDROMEDA  MERAK, A. J.  SR-P

TITLE  AUTHOR  CODE
DARK APPRENTICE [@STAR WARS JEDI ACADEMY 2]  ANDERSON, KEVIN J.  C/P:2,P
DARK ARARAT [EMORTALITY]  STABLEFORD, BRIAN M.  H
DARK AS DAY [BATTACHARIYA 3]  SHEFFIELD, CHARLES  H
DARK ASCENT, THE  HUNT, WALTER H.  P
DARK ATLANTIS  CRAIGIE, DAVID  H
DARK BEASTS, THE  LONG, FRANK BELKNAP  C/P:2,P
DARK BETWEEN THE STARS, THE  ANDERSON, POUL  H,P
DARK BEYOND THE STARS, THE  ROBINSON, FRANK M.  H:2,P
DARK BIBLE  MANACHINO, ALBERT J.  B
DARK BOUNDARIES  LORRAINE, PAUL  SR-P
DARK BRIGHT WATER, THE  WRIGHTSON, PATRICIA  C/P,P
DARK CASTLE, WHITE HORSE  LEE, TANITH  C/P,P
DARK CELEBRATION [CARPATHIAN 17]  FEEHAN, CHRISTINE  H
DARK CHAMBER, THE  CLINE, LEONARD  P
DARK CHANNEL  GARTON, RAY  H
DARK CITIES UNDERGROUND  GOLDSTEIN, LISA  H,LP
DARK CITY  LAURIA, FRANK  P
DARK COMPANION  NORTON, ANDRE  H
DARK COMPANIONS  CAMPBELL, RAMSEY  H,P
DARK CONTINUUM  MULLER, JOHN E.  SR-P
DARK COUNTRY, THE  ETCHISON, DENNIS  SR-H,H
DARK CRUSADE, THE [XOR 3]  HUNT, WALTER H.  P
DARK CRUSADE  WAGNER, KARL EDWARD  C/P:2,SFWA-P,P
DARK CRYSTAL, THE  SMITH, A. C. H.  C/P,P
DARK DANCE [BLOOD OPERA 1]  LEE, TANITH  H
DARK DECEMBER  COPPEL, ALFRED  C/P:2,P
DARK DEMON [CARPATHIAN 16]  FEEHAN, CHRISTINE  P
DARK DESCENT, THE  HARTWELL, DAVID G.  LPA
DARK DESIGN, THE [RIVERWORLD 3]  FARMER, PHILIP JOSE  C/P,P,H
DARK DESIGN, THE (PROOF) [RIVERWORLD 3]  FARMER, PHILIP JOSE  SR-LP
DARK DESIRE [CARPATHIAN 2]  FEEHAN, CHRISTINE  P
DARK DESTINY [CARPATHIAN 13]  FEEHAN, CHRISTINE  P
DARK DESTINY III: CHILDREN OF DRACULA  KRAMER, EDWARD E.  PA
DARK DESTROYERS, THE  WELLMAN, MANLY WADE  DD443
DARK DIMENSIONS, THE [GRIMES]  CHANDLER, A. BERTRAM  D13783
DARK DIMENSIONS, THE [GRIMES]/RIM GODS, THE [GRIMES]  CHANDLER, A. BERTRAM  P
DARK DIVIDE [RUNESWORD 5]  ACRES, MARK  C/P,P
DARK DOMAIN, THE  GRABINSKI, STEFAN  LP
DARK DOMINION  DUNCAN, DAVID  C/P:2,P
DARK DOOR, THE  WILHELM, KATE  P
DARK DOOR OPENS, THE [@LEGENDS OF LONE WOLF #2]  DEVER, JOE/GRANT, JOHN  P
DARK DREAMS [CHILDREN OF THE NIGHT 1]  HODGMAN, ANN  P
DARK DRUID, THE [FINN MACCUMHAL 3]  FLINT, KENNETH C.  P
DARK ELF TRILOGY, THE [@FORGOTTEN REALMS DARK ELF]  SALVATORE, R. A.  LP
DARK EMBLEM [@DEATHLANDS]  AXLER, JAMES  P
DARK FORCE RISING [STAR WARS]  ZAHN, TIMOTHY  H:2,P
DARK FORCES  MCCAULEY, KIRBY  SR-HA,C/PA,PA
DARK FOUNDATIONS, THE [@LAMB AMONG THE STARS #2]  WALLEY, CHRIS  H
DARK FUTURES [@TERMINATOR 2 NEW JOHN CONNOR 1]  BLACKFORD, RUSSELL  P
DARK GENESIS [@BABYLON 5 PSI CORPS 1]  KEYES, J. GREGORY  C/P,P
DARK GLORY WAR, THE [DRAGONCROWN WAR]  STACKPOLE, MICHAEL A.  P
DARK GODS  KLEIN, T. E. D.  H,P
DARK HALF, THE  KING, STEPHEN  C/P,P,H
DARK HAND OF MAGIC, THE [SUNWOLF AND STARHAWK 3]  HAMBLY, BARBARA  C/P,P,H
DARK HEART [@DRAGONLANCE MEETINGS #3]  DANIELL, TINA  C/P:2,P
DARK HEART  WEIS, MARGARET/BALDWIN, DAVID  P
DARK HEART OF TIME, THE [@TARZAN]  FARMER, PHILIP JOSE  P
DARK HEAVENS  LEVY, ROGER  LP
DARK HORIZON [@ALIEN NATION #2]  JETER, K. W.  P
DARK HORN BLOWING, A  IPCAR, DAHLOV  P
DARK HORSE [DARK HORSE 1]  HERBERT, MARY H.  P
DARK HOURS [@BLOOD OF TEN CHIEFS #5]  PINI, RICHARD  LPA
DARK IMAGININGS  BOYER, ROBERT H./ZAHORSKI, KENNETH J. HA,LPA
DARK IMPERIUM [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
DARK INTRUDER AND OTHER STORIES, THE  BRADLEY, MARION ZIMMER  DF273
DARK IS RISING, THE [DARK IS RISING 2]  COOPER, SUSAN  C/P,LP,H:2
DARK IS THE MOON [@VIEW FROM THE MIRROR #3]  IRVINE, IAN  P
DARK IS THE SUN  FARMER, PHILIP JOSE  H:2,P
DARK JOURNEY [@STAR WARS NEW JEDI ORDER]  CUNNINGHAM, ELAINE  P
DARK JOURNEY  MORLAN, A. R.  P
DARK KNIGHT OF KARAMEIKOS [MYSTARA]  BROWN, TIMOTHY  P
DARK LADY, THE  RESNICK, MIKE  C/P:2,SFWA-P,P
DARK LAGOON, THE  SHINEROCK, SIMON M.  P
DARK LEGACY [@MAGIC: THE GATHERING]  VARDEMAN, ROBERT E.  P

TITLE  AUTHOR  CODE
DARK LIGHT [ENGINES OF LIGHT 2]  MACLEOD, KEN  H
DARK LIGHT [SHATTERED LIGHT 2]  YARBRO, CHELSEA QUINN  P
DARK LIGHT-YEARS, THE  ALDISS, BRIAN W.  C/P:2,P
DARK LORD, THE [OATH OF EMPIRE 4]  HARLAN, THOMAS  H
DARK LORD OF DERKHOLM [DERKHOLM 1]  JONES, DIANA WYNNE  H
DARK LORD OF PENGERSICK, THE  CARLYON, RICHARD  P
DARK LORD: THE RISE OF DARTH VADER [STAR WARS]  LUCENO, JAMES  H
DARK MAGIC [CARPATHIAN 4]  FEEHAN, CHRISTINE  P
DARK MAGIC [@GODWARS #2]  WELLS, ANGUS  C/P:2,P
DARK MATTER  REEVES-STEVENS, GARFIELD  LP
DARK MATTER  THOMAS, SHEREE R.  LPA
DARK MESSIAH [@PHOENIX #1]  ALEXANDER, DAVID  P
DARK MESSIAH [MESSIAH STONE 2]  CAIDIN, MARTIN  C/P:2,P
DARK MILLENNIUM, THE  MERAK, A. J.  C/P:2,P,H
DARK MIND, DARK HEART  DERLETH, AUGUST  SR-PA,C/PA
DARK MIRROR [STAR TREK TNG]  DUANE, DIANE  H:2,P
DARK MIRROR [@ANGEL]  GARDNER, CRAIG SHAW  C/P,P
DARK MIRROR, THE [@BRIDEI #1]  MARILLIER, JULIET  H
DARK MIRROR, DARK DREAMS [ALEXIA AND TIRAN 2/HIDDEN REALMS 2] GREEN, SHARON  P
DARK MONDAYS  BAKER, KAGE  H
DARK MOON  GEMMELL, DAVID  P
DARK MOON, THE [@GUARDIAN #1]  GRAY, JULIA  P
DARK MOON DEFENDER [TWELVE HOUSES]  SHINN, SHARON  H
DARK MUSIC [CHILDREN OF THE NIGHT 2]  HODGMAN, ANN  P
DARK MUSIC AND OTHER SPECTRAL TALES  SNOW, JACK  SR-H
DARK NADIR [SHOLAN ALLIANCE 5]  NORMAN, LISANNE  P
DARK NEST [@STAR WARS DARK NEST 1,2,3]  DENNING, TROY  H
DARK OF THE MOON  DERLETH, AUGUST  HA
DARK OF THE MOON [JAMETHIEL DREAM-WEAVER 2]  HODGELL, P. C.  H:2,SFWA-P,P
DARK OF THE SUN [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
DARK OF THE WOODS  KOONTZ, DEAN R.  D13793
DARK OTHER, THE  WEINBAUM, STANLEY G.  SR-H
DARK PARADISE [QUESTIONER 3]  SCHMIDT, DENNIS  P
DARK PASSAGE  PODRUG, JUNIUS  P
DARK PASSIONS BOOK ONE [STAR TREK]  WRIGHT, SUSAN  C/P,P
DARK PASSIONS BOOK TWO [STAR TREK]  WRIGHT, SUSAN  C/P,P
DARK PATH, THE [XOR 2]  HUNT, WALTER H.  H
DARK PATH, THE [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
DARK PIPER  NORTON, ANDRE  C/P:2,P,H
DARK PLANET, THE  HOLLY, JOAN HUNTER  H,P
DARK PLANET  RACKHAM, JOHN  D13805
DARK POWERS [@SENTINELS #2/ROBOTECH 14]  MCKINNEY, JACK  P
DARK PRIESTESS  COULSON, JUANITA  P
DARK PRINCE [MACEDON 2]  GEMMELL, DAVID  LP
DARK PRINCE [WORLD OF DARKNESS VAMPIRE]  HERBER, KEITH  C/P,P
DARK PROGENY [DOCTOR WHO EIGHTH DOCTOR]  EMMERSON, STEVE  P

DARK PROPHECY [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #3] FLEMING, GHERBOD

P

DARK QUEEN, THE [@DRAGONLANCE VILLAINS #6]  WILLIAMS, MICHAEL/WILLIAMS, TERI  P
DARK REALM, THE [@TORG POSSIBILITY WARS #2]  KAUFMAN, DOUGLAS  C/P,P
DARK REMAINS, THE [LAST RUNE 3]  ANTHONY, MARK  P
DARK RENDEZVOUS [@STAR WARS CLONE WARS YODA]  STEWART, SEAN  P
DARK RIVER, THE [@FOURTH REALM #2]  TWELVE HAWKS, JOHN  H
DARK SATANIC  BRADLEY, MARION ZIMMER  C/P:2,SFWA-P,P
DARK SEAS OF MALTERN MANOR, THE  VERNON, KAY  P
DARK SECRET [CARPATHIAN 15]  FEEHAN, CHRISTINE  H
DARK SEEKER  JETER, K. W.  SFWA-P
DARK SHADOWS  ROSS, MARILYN  P
DARK SHADOWS BOOK OF VAMPIRES AND WEREWOLVES, THE  COLLINS, BARNABAS/COLLINS, QUENTIN  PA
DARK SHADOWS COOKBOOK, THE  MALIS, JODY CAMERON  [C/RAN-P]
DARK SHORE, THE [@DOMINIONS OF IRTH #1]  LEE, ADAM  C/P,P
DARK SIDE, THE  HUGHES, ZACH  C/P,P
DARK SIDE, THE  KNIGHT, DAMON  HA
DARK SIDE OF THE EARTH, THE  BESTER, ALFRED  C/P:2,P
DARK SIDE OF THE MOON, THE  VANCE, JACK  H
DARK SIDE OF THE SUN [@FARSCAPE]  DYMOND, ANDREW  C/P,P
DARK SIDE OF THE SUN, THE  PRATCHETT, TERRY  H
DARK SIDE OF VENUS  MACARTNEY, CLEM  SR-P
DARK SINS, DARK DREAMS  MALZBERG, BARRY N./PRONZINI, BILL  HA,SFWA
DARK SISTER  JOYCE, GRAHAM  H
DARK SKY LEGION  BARTON, WILLIAM  C/P:3,SFWA-P,P
DARK SLEEP, THE [VAMPIRE FILES 8]  ELROD, P. N.  H:2
DARK SLEEPER [WESTERN LIGHTS 1]  BARLOUGH, JEFFREY E.  LP
DARK SONATA [IRA VEDDER 4]  BOND, EVELYN  P

The MITSFS Titledex  September 24, 2007  Page 79
TITLE  AUTHOR  CODE
DARK STAR  FOSTER, ALAN DEAN  C/P:2,P
DARK STARS  SILVERBERG, ROBERT  PA,C/PA
DARK STARS AND OTHER ILLUMINATIONS  MONTELEONE, THOMAS F.  H,SFWA
DARK STRAITS OF REGLATHIUM, THE [@DANNUS #4]  SIROTA, MIKE  P
DARK SWORD’S LOVER, THE [ROLAND 2]  VAN ASTEN, GAIL  C/P,P
DARK SYMPHONY [CARPATHIAN 10]  FEEHAN, CHRISTINE  P
DARK SYMPHONY, THE  KOONTZ, DEAN R.  C/P,P
DARK THANE [@DRAGONLANCE AGE OF MORTALS]  CROOK, JEFF  P
DARK THANE [@LAST CLANSMAN #2]  LAU, MILLER  P
DARK THINGS  DERLETH, AUGUST  SR-HA
DARK TIDE  FORREST, ELIZABETH  P
DARK TIDE, THE [@IRON TOWER #1/MITHGAR 7]  MCKIERNAN, DENNIS L.  C/P:2,SFWA,P
DARK TIDE [@STAR WARS NEW JEDI ORDER DARK TIDE 1,2]  STACKPOLE, MICHAEL A.  H
DARK TIMES [@BUFFY LOST SLAYER #2]  GOLDEN, CHRISTOPHER  C/P
DARK TOWER, THE [@DARK TOWER #7]  KING, STEPHEN  H:2
DARK TOWER AND OTHER STORIES, THE  LEWIS, C. S.  H,LP
DARK TRAVELING, A  ZELAZNY, ROGER  C/P,P,H
DARK UNIVERSE  GALOUYE, DANIEL F.  C/P:2,P

DARK VALLEY DESTINY: THE LIFE OF ROBERT E. HOWARD DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK/ GRIFFIN, JANE WHITTINGTON

R/REF-H

DARK VENGEANCE [@CHARMED]  GALLAGHER, DIANA G.  P

DARK VICTORY [STAR TREK MIRROR 2] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

DARK VISIONS  O’CALLAGHAN, MAXINE  SFWA-P
DARK VOICES 4  SUTTON, DAVID/JONES, STEPHEN  SR-PA
DARK WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
DARK WATERS [CHRONICLES UNBINDING 2]  STONE, ROBERT S.  P
DARK WATER’S EMBRACE  LEIGH, STEPHEN  P
DARK WAY TO GLORY, A [@OUTREMER #3]  BRENCHLEY, CHAZ  P
DARK WING, THE [XOR 1]  HUNT, WALTER H.  C/P,P
DARK WING  WEST, CARL/MACLEAN, KATHERINE  H
DARK WORLD, THE  KUTTNER, HENRY  C/P:2,P
DARKANGEL, THE [DARKANGEL 1]  PIERCE, MEREDITH ANN  C/P,P
DARKANGEL TRILOGY, THE [DARKANGEL 1,2,3]  PIERCE, MEREDITH ANN  H
DARKCHILD  VAN SCYOC, SYDNEY J.  C/P:2,LP
DARKE CRUSADE, THE [LONE WOLF 15]  DEVER, JOE  P
DARKENHEIGHT [WATERSHED 2]  NILES, DOUGLAS  P
DARKENING, THE  MCGREW, CHANDLER  P
DARKENING, THE [@BUFFY WICKED WILLOW #1]  NAVARRO, YVONNE  C/P,P
DARKENING, THE [@IMMORTALS]  POPP, ROBIN T.  P
DARKENING FLOOD, THE  FLINT, KENNETH C.  P
DARKENING ISLAND  PRIEST, CHRISTOPHER  H,P
DARKER GEOMETRY, A [KNOWN SPACE]  MARTIN, MARK O./BENFORD, GREGORY  P
DARKER JEWELS [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H,LP
DARKER MAGIC, A  BEDARD, MICHAEL  P
DARKER THAN YOU THINK  WILLIAMSON, JACK  SR-H,C/P:2,P
DARKEST DAY, THE [@IRON TOWER #3/MITHGAR 9]  MCKIERNAN, DENNIS L.  C/P:2,SFWA,P
DARKEST HEART [SONJA BLUE]  COLLINS, NANCY A.  H
DARKEST HOUR [AGE OF MISRULE 2]  CHADBOURN, MARK  P
DARKEST KNIGHT [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
DARKEST NIGHT, THE  SAXON, PETER  P
DARKEST ROAD, THE [FIONAVAR TAPESTRY 3]  KAY, GUY GAVRIEL  H:2,SFWA,P
DARKFALL [WORLD OF ADEN THUNDERSCAPE]  HENSLEY, SHANE LACY  P
DARKFALL  KOONTZ, DEAN R.  P
DARKFELL  CORRAN, MARY  P
DARKFEVER  MONING, KAREN MARIE  H
DARKHOLD  BROWN, JERRY EARL  C/P,P
DARKLIGHT GRIMOIRE, THE [WHEEL 3]  FRANKOS, STEVEN  P
DARKLING, THE  KESTERTON, DAVID  H
DARKLING HILLS, THE  MARTIN, LORI  C/P,P
DARKLING PLAIN, A [HUNGRY CITY 4]  REEVE, PHILIP  H
DARKLING PLAIN, A (PROOF) [@HUNGRY CITY 4]  REEVE, PHILIP  SR-LP
DARKLING WIND, THE [INQUESTOR 4]  SUCHARITKUL, SOMTOW  C/P:2,P
DARKLOOM  OSBORNE, CARY  C/P,P
DARKLOST [VICTOR RENQUIST 1]  FARREN, MICK  H
DARKMAGE [WINDROSE 1,2]  HAMBLY, BARBARA  H:2
DARKMAN  BOYLL, RANDALL  P
DARKNESS & LIGHT [DRAGONLANCE PRELUDES 1]  THOMPSON, PAUL B./CARTER, TONYA R.  C/P:2,P
DARKNESS AND DAWN  ENGLAND, GEORGE ALLAN  H,LP
DARKNESS AND DAWN  NORTON, ANDRE  P
DARKNESS AND THE LIGHT  STAPLEDON, OLAF  H,LP
DARKNESS AT SETHANON, A [RIFTWAR 4]  FEIST, RAYMOND E.  C/P,P,H,SFWA

The MITSFS Titledex  September 24, 2007  Page 80
TITLE  AUTHOR  CODE
DARKNESS BEFORE THE DAWN, THE [DARK SUN ATHAS #2]  HUGHES, RYAN  P
DARKNESS BEFORE TOMORROW, THE  WILLIAMS, ROBERT MOORE  DF141
DARKNESS DESCENDING [INTO THE DARKNESS 2]  TURTLEDOVE, HARRY  C/P,P,H
DARKNESS DESCENDS [@MYSTIC REBEL]  SYVERTSEN, RYDER  P
DARKNESS IN MY SOUL, A  KOONTZ, DEAN R.  C/P:2,P
DARKNESS OF GOD, THE [@SHADOW WARRIOR #3]  BUNCH, CHRIS  P
DARKNESS ON DIAMONDIA, THE  VAN VOGT, A. E.  C/P:2,P
DARKNESS THAT COMES BEFORE, THE [@PRINCE OF NOTHING #1]  BAKKER, R. SCOTT  H
DARKNESS UPON THE ICE, A [ICE 3]  FORSTCHEN, WILLIAM R.  C/P,P
DARKNESS VISIBLE  GOLDING, WILLIAM  P
DARKNESS WEAVES  WAGNER, KARL EDWARD  P
DARKNESS WEAVES WITH MANY SHADES...  WAGNER, KARL EDWARD  SR-P
DARKNESSES [COREAN 2]  MODESITT, L. E. JR.  P
DARKOVER CONCORDANCE, THE  BREEN, WALTER  R/REF-H
DARKOVER LANDFALL [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
DARKSABER [@STAR WARS]  ANDERSON, KEVIN J.  H:2,P
DARKSIDE  ETCHISON, DENNIS  C/P:2,P
DARKSONG RISING [SPELLSONG 3]  MODESITT, L. E. JR.  C/P,P,H
DARKSPAWN  TILTON, LOIS  LP
DARKSPELL [DEVERRY 2/DEVERRY PROPER 2]  KERR, KATHARINE  C/P,SFWA,P,H
DARKSPELL (REVISED) [DEVERRY 2/DEVERRY PROPER 2]  KERR, KATHARINE  C/P,P
DARKSTEEL EYE, THE [@MAGIC: THE GATHERING MIRRODIN 2]  LEBOW, JESS  P
DARKTHUNDER’S WAY [OTHERWORLD 3]  DEITZ, TOM  C/P:2,P
DARKTRADERS [HELLFLOWER 2]  BES SHAHAR, ELUKI  C/P:2,P
DARKVISION [@FORGOTTEN REALMS WIZARDS]  CORDELL, BRUCE R.  P
DARKWALKER ON MOONSHAE [@FORGOTTEN REALMS MOONSHAE #1]  NILES, DOUGLAS  C/P:2,P
DARKWATER  ALEXANDER, JAN  P
DARKWELL [@FORGOTTEN REALMS MOONSHAE #3]  NILES, DOUGLAS  C/P:2,P
DARKWITCH RISING [TROY GAME 3]  DOUGLASS, SARA  H
DARKWORLD DETECTIVE  REAVES, MICHAEL  C/P:2,P
DARTH MAUL: SHADOW HUNTER [STAR WARS]  REAVES, MICHAEL  P
DARWINIA  WILSON, ROBERT CHARLES  C/P,P,H
DARWIN’S BLADE  SIMMONS, DAN  H
DARWIN’S CHILDREN [SHEVA 2]  BEAR, GREG  H
DARWIN’S RADIO [SHEVA 1]  BEAR, GREG  H:2,P

DARWIN’S WATCH [@SCIENCE OF DISCWORLD #3/DISCWORLD] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK

H,P

DARYA OF THE BRONZE AGE [ERIC CARSTAIRS 4]  CARTER, LIN  P
D.A.R.Y.L.  KLEINBAUM, N. H.  C/P,P
DASTARD, THE [XANTH #24]  ANTHONY, PIERS  C/P,P,H
DATE ME, BABY, ONE MORE TIME  ROWE, STEPHANIE  P
DATE WITH DEATH [@CHARMED]  LENHARD, ELIZABETH  P

DATES FROM HELL HARRISON, KIM/SANDS, LYNSAY/

ARMSTRONG, KELLEY/HANDELAND, LORI

P

DAUGHTER OF ANCIENTS [BRIDGE OF D’ARNATH 4]  BERG, CAROL  P
DAUGHTER OF DESTINY  SALSITZ, R. A. V.  C/P,P
DAUGHTER OF ELYSIUM [ELYSIUM]  SLONCZEWSKI, JOAN  C/P,P
DAUGHTER OF EXILE  GLASS, ISABEL  P
DAUGHTER OF FU MANCHU  ROHMER, SAX  H:2,P
DAUGHTER OF HOUNDS  KIERNAN, CAITLIN R.  LP
DAUGHTER OF INDEPENDENCE [KYDAN 3]  BROWN, SIMON  P
DAUGHTER OF IS  DAVIDSON, MICHAEL  C/P:2,P
DAUGHTER OF MAGIC [WIZARD OF YURT 5]  BRITTAIN, C. DALE  P
DAUGHTER OF NIGHT  OBUKHOVA, LYDIA  H,P
DAUGHTER OF PROPHECY [POWER AND THE PATTERN 1]  BUSH, ANNE KELLEHER  C/P,P
DAUGHTER OF PROPHECY  OWENS, MILES  LP
DAUGHTER OF REGALS AND OTHER TALES  DONALDSON, STEPHEN R.  C/P,P,H
DAUGHTER OF THE BEAR KING  ARNASON, ELEANOR  P
DAUGHTER OF THE BLOOD [BLACK JEWELS 1]  BISHOP, ANNE  C/P,P
DAUGHTER OF THE BRIGHT MOON  ABBEY, LYNN  C/P:2,LP
DAUGHTER OF THE DESERT  BRENNAN, NOEL-ANNE  P
DAUGHTER OF THE DRAGON [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P

DAUGHTER OF THE DROW [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 1] CUNNINGHAM, ELAINE

P

DAUGHTER OF THE EMPIRE [MARA 1]  FEIST, RAYMOND E./WURTS, JANNY  C/P,P,H
DAUGHTER OF THE FOREST [SEVENWATERS 1]  MARILLIER, JULIET  H
DAUGHTER OF THE LION [CHEYSULI 6]  ROBERSON, JENNIFER  C/P,P
DAUGHTER OF THE NIGHT  BERGSTROM, ELAINE  P
DAUGHTER OF THE SHAMAN  GUTTENBERG, ELYSE  P
DAUGHTER OF WITCHES [LYRA 4]  WREDE, PATRICIA C.  C/P:2,P
DAUGHTER OF ZIR [HORSE GODDESS PEOPLE 2]  TARR, JUDITH  LP
DAUGHTERS OF A CORAL DAWN [MATERNAS 1]  FORREST, KATHERINE V.  LP
DAUGHTERS OF AN AMBER MOON [MATERNAS 2]  FORREST, KATHERINE V.  LP
DAUGHTERS OF AN EMERALD DUSK [MATERNAS 3]  FORREST, KATHERINE V.  LP

The MITSFS Titledex  September 24, 2007  Page 81
TITLE  AUTHOR  CODE
DAUGHTERS OF EARTH  LARBALESTIER, JUSTINE  LPA
DAUGHTERS OF EARTH  MERRIL, JUDITH  C/P:2,P
DAUGHTERS OF THE DOLPHIN [DOLPHIN 2]  MEYERS, ROY  C/P:2,P
DAUGHTERS OF THE SUNSTONE  VAN SCYOC, SYDNEY J.  H:2
DAUNTLESS [@LOST FLEET 1]  CAMPBELL, JACK  P
DAVID BLAIZE AND THE BLUE DOOR  BENSON, E. F.  H
DAVID COPPERFIELD’S BEYOND IMAGINATION  COPPERFIELD, DAVID  PA
DAVID COPPERFIELD’S TALES OF THE IMPOSSIBLE  COPPERFIELD, DAVID/BERLINER, JANET  HA,PA
DAVID STARR, SPACE RANGER [LUCKY STARR 1]  ASIMOV, ISAAC  C/P:2,P

DAVID STARR, SPACE RANGER [LUCKY STARR 1,2]/ LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 1,2] ASIMOV, ISAAC

P

DAVID’S SLING  STIEGLER, MARC  C/P,SFWA-P,P
DAVY  PANGBORN, EDGAR  C/P:2,SFWA-P,P

DAW 30TH ANNIVERSARY FANTASY WOLLHEIM, ELIZABETH R./ GILBERT, SHEILA E.

PA

DAW 30TH ANNIVERSARY SCIENCE FICTION WOLLHEIM, ELIZABETH R./ GILBERT, SHEILA E.

HA

DAW SCIENCE FICTION READER, THE  WOLLHEIM, DONALD A.  PA,C/PA
DAWN [XENOGENESIS 1]  BUTLER, OCTAVIA E.  H:2,P
DAWN  WRIGHT, S. FOWLER  H
DAWN FOR A DISTANT EARTH [FOREVER HERO 1]  MODESITT, L. E. JR.  C/P:2,SFWA-P,P
DAWN OF A DARK AGE [BOOK OF MAN 1]  WELCH, JANE  P
DAWN OF AMBER, THE [ROGER ZELAZNY’S DAWN OF AMBER 1]  BETANCOURT, JOHN GREGORY  H:2
DAWN OF DARKNESS  DEL MARTIA, ASTRON  SR-P
DAWN OF MAN, THE  LEWIS, ELBERT JR.  LP
DAWN OF NIGHT [@FORGOTTEN REALMS EREVIS CALE #2]  KEMP, PAUL S.  P
DAWN OF THE DEAD  ROMERO, GEORGE A./SPARROW, SUSANNA  H,LP
DAWN OF THE DRAGONS [LONE WOLF]  DEVER, JOE  P
DAWN OF THE ICE BEAR [@AGE OF CONAN MARAUDERS #3]  MARIOTTE, JEFF  P
DAWN OF THE SPACE AGE  HARPER, HARRY  [C/RAN-H]
DAWN OVER THE AMAZON  BEALS, CARLETON  H
DAWN SONG  GREEN, SHARON  C/P,P
DAWN STAG, THE [DALRIADA 2]  WATSON, JULES  H
DAWN STAR, THE [MEL DAWNFIELD 2]  ASARO, CATHERINE  LP
DAWNING LIGHT, THE  RANDALL, ROBERT  C/P,P,H
DAWNING OF A NEW AGE, THE [@DRAGONLANCE FIFTH AGE 1]  RABE, JEAN  P

DAWNING SHADOWS: LIGHT ON THE SOUND, THE (REVISED)

[INQUESTOR 1] SUCHARITKUL, SOMTOW

C/P:2,P

DAWNING SHADOWS: THRONE OF MADNESS, THE (REVISED)

[INQUESTOR 2] SUCHARITKUL, SOMTOW

P

DAWNMAN PLANET  REYNOLDS, MACK  DG580
DAWN’S UNCERTAIN LIGHT  BARRETT, NEAL JR.  P
DAWNTHIEF [CHRONICLES OF THE RAVEN 1]  BARCLAY, JAMES  C/P:2,P
DAY AFTER DOOMSDAY, THE  VALE, RENA  C/P:2,P
DAY AFTER JUDGMENT, THE  BLISH, JAMES  C/P,SFWA,SR-P,H
DAY AFTER TOMORROW, THE  HEINLEIN, ROBERT A.  C/P:2,P
DAY AFTER TOMORROW  REYNOLDS, MACK  C/P,P
DAY BEFORE FOREVER, THE/THUNDERHEAD  LAUMER, KEITH  C/P,P,H
DAY BEFORE TOMORROW, THE  KLEIN, GERARD  C/P:2,P
DAY BY NIGHT  LEE, TANITH  C/P:2,P
DAY DARK AS NIGHT, A [@EXALTED #1]  BOWEN, CARL  P
DAY DARK, NIGHT BRIGHT  LEIBER, FRITZ  SR-H
DAY FOR DAMNATION, A [WAR AGAINST THE CHTORR 2]  GERROLD, DAVID  C/P:2,P

DAY FOR DAMNATION, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 2] GERROLD, DAVID

C/P,LP

DAY HE DIED, THE  PADGETT, LEWIS  SR-P
DAY IN THE LIFE, A  DOZOIS, GARDNER  HA,PA
DAY MILLION  POHL, FREDERIK  C/P:2,P
DAY NEW YORK TREMBLED, THE  LEWIS, IRWIN  C/P:2,P
DAY NEW YORK WENT DRY, THE  EINSTEIN, CHARLES  C/P:2,P
DAY OF CREATION, THE  BALLARD, J. G.  H,P

DAY OF DESCENT, THE [@ALIEN NATION #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

DAY OF RECKONING: THE STEREOSCOPE [@BLACKSTONE #5]  SAUL, JOHN  C/P,P
DAY OF THE BEASTS  MULLER, JOHN E.  C/P,P
DAY OF THE BROWN HORDE, THE  TOOKER, RICHARD  H
DAY OF THE BURNING, THE  MALZBERG, BARRY N.  C/P,P
DAY OF THE CHEETAH [MAC MCLANAHAN 2]  BROWN, DALE  H

DAY OF THE DINOSAUR, THE DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

[C/RAN-P]

DAY OF THE DISSONANCE, THE [SPELLSINGER 3]  FOSTER, ALAN DEAN  C/P:2,P
DAY OF THE DOLPHIN, THE  MERLE, ROBERT  C/P,P,H
DAY OF THE DRAGON [@WARCRAFT #1]  KNAAK, RICHARD A.  P

DAY OF THE DRAGONSTAR [DRAGONSTAR 1] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 82
TITLE  AUTHOR  CODE
DAY OF THE DRONES, THE  LIGHTNER, A. M.  C/P:2,P
DAY OF THE GIANTS  DEL REY, LESTER  C/P,P,H
DAY OF THE KLESH, THE [LER 3]  FOSTER, M. A.  C/P:2,P
DAY OF THE MINOTAUR  SWANN, THOMAS BURNETT  C/P:2,P
DAY OF THE NESS, THE  NORTON, ANDRE/GILBERT, MICHAEL  H,P
DAY OF THE SHIELD, THE  ALBAN, ANTONY  C/P:2,P
DAY OF THE STAR CITIES, THE  BRUNNER, JOHN  C/P:2,P
DAY OF THE STARWIND [LAST LEGIONARY 3]  HILL, DOUGLAS  H,P
DAY OF THE TEMPEST, THE [@DRAGONLANCE FIFTH AGE]  RABE, JEAN  P
DAY OF THE TIMESTOP, THE  FARMER, PHILIP JOSE  P
DAY OF THE TRIFFIDS, THE (AMERICAN) [TRIFFIDS 1]  WYNDHAM, JOHN  C/P,P,H
DAY OF THE TRIFFIDS, THE (BRITISH) [TRIFFIDS 1]  WYNDHAM, JOHN  SR-P,C/P
DAY OF THE TYRANT [@THERE WILL BE WAR #4]  POURNELLE, JERRY  PA,C/PA,SFWA-P
DAY OF THEIR RETURN, THE  ANDERSON, POUL  H:2,P
DAY OF WRATH [DIES IRAE 3]  STABLEFORD, BRIAN M.  C/P:2,P
DAY ONE  MACCABEE, JOHN  P
DAY SEVEN  BICKHAM, JACK M.  P
DAY STAR, THE  GESTON, MARK S.  C/P:2,P
DAY THE COMMUNISTS TOOK OVER AMERICA, THE  MOORE, ISABEL  SR-P
DAY THE EARTH CAUGHT FIRE, THE  WELLS, BARRY  C/P:2,P
DAY THE EARTH FROZE, THE  HATCH, GERALD  C/P:2,P
DAY THE GODS DIED, THE  ERNSTING, WALTER  P
DAY THE MACHINES STOPPED, THE  ANVIL, CHRISTOPHER  P
DAY THE MAGIC STOPPED, THE  STASHEFF, CHRISTOPHER  PA
DAY THE MARTIANS CAME, THE  POHL, FREDERIK  H:2,P
DAY THE OCEANS OVERFLOWED, THE  FONTENAY, CHARLES L.  C/P,P
DAY THE SUN STOOD STILL, THE  DEL REY, LESTER  PA,C/PA
DAY THE WORLD DIED, THE  MULLER, JOHN E.  SR-P,C/P
DAY THE WORLD ENDED, THE  ROHMER, SAX  C/P:2,P
DAY THE WORLD STOPPED, THE  COBLENTZ, STANTON A.  H:2
DAY THEY H-BOMBED LOS ANGELES, THE  WILLIAMS, ROBERT MOORE  C/P:2,P
DAY THEY INVADED NEW YORK, THE  LEWIS, IRWIN  C/P:2,P
DAY WATCH [NIGHTWATCH 2]  LUKYANENKO, SERGEI  H,LP
DAYBREAK — 2250 A.D.  NORTON, ANDRE  C/P:2,SR-DD069,P
DAYDREAMS UNDERTAKEN  ANTCZAK, STEPHEN L.  LP
DAYMARES  BROWN, FREDRIC  C/P:2,P
DAYS AFTER TOMORROW, THE  SANTESSON, HANS STEFAN  HA
DAYS BETWEEN STATIONS  ERICKSON, STEVE  LP
DAYS OF AIR AND DARKNESS [DEVERRY 8/WESTLANDS 4]  KERR, KATHARINE  C/P:2,LP
DAYS OF ATONEMENT  WILLIAMS, WALTER JON  C/P,P,H,SFWA
DAYS OF BITTER STRENGTH [@CHUNG KUO #7]  WINGROVE, DAVID  C/P,LP,H
DAYS OF BLOOD AND FIRE [DEVERRY 7/WESTLANDS 3]  KERR, KATHARINE  C/P:2,LP
DAYS OF CAIN  DUNN, J. R.  C/P,P
DAYS OF GLORY, THE [DIES IRAE 1]  STABLEFORD, BRIAN M.  C/P:2,P
DAYS OF GRASS  LEE, TANITH  P
DAYS OF MAGIC, NIGHTS OF WAR [ABARAT 2]  BARKER, CLIVE  H
DAYSTAR AND SHADOW  JOHNSON, JAMES B.  P
DAYWORLD [DAYWORLD 1]  FARMER, PHILIP JOSE  C/P,P,H
DAYWORLD BREAKUP [DAYWORLD 3]  FARMER, PHILIP JOSE  H:2,SFWA-P,P
DAYWORLD REBEL [DAYWORLD 2]  FARMER, PHILIP JOSE  C/P,SFWA-P,P,H
DAZZLE OF DAY, THE  GLOSS, MOLLY  LP

DE CAMP: AN L. SPRAGUE DE CAMP BIBLIOGRAPHY LAUGHLIN, CHARLOTTE/ LEVACK, DANIEL J. H.

R/IND

DE LA TERRE A LA LUNE  VERNE, JULES  FP
DEAD AIR [@SHADOWRUN 22]  KOKE, JAK  P
DEAD AND BURIED  YARBRO, CHELSEA QUINN  C/P,P,H
DEAD AND LOVING IT  DAVIDSON, MARYJANICE  LP
DEAD AND THE DAMNED, THE [WARHAMMER]  GREEN, JONATHAN  P
DEAD AS A DOORNAIL [SOUTHERN VAMPIRES 5]  HARRIS, CHARLAINE  H
DEAD ASTRONAUT, THE  PLAYBOY BOOKS  PA,C/PA
DEAD BALL [@BLOOD BOWL 2]  FORBECK, MATT  P
DEAD BEAT [DRESDEN FILES 7]  BUTCHER, JIM  H
DEAD BY DAY [SOUTHERN VAMPIRES 4,5]  HARRIS, CHARLAINE  H
DEAD EASY [HALFLIFE 4]  SIMMONS, WM. MARK  H
DEAD GIRLS [DEAD 1]  CALDER, RICHARD  P
DEAD GIRLS’ DANCE, THE [@MORGANVILLE VAMPIRES 2]  CAINE, RACHEL  P
DEAD GIRLS DEAD BOYS DEAD THINGS [DEAD 1,2,3]  CALDER, RICHARD  LP
DEAD GOD DANCING, A  MAXWELL, ANN  C/P,P
DEAD GROUND  AMIES, CHRIS  LP
DEAD IN DIXIE [SOUTHERN VAMPIRES 1,2,3]  HARRIS, CHARLAINE  H
DEAD KINGDOM, THE [SEVEN CITADELS 3]  HARRIS, GERALDINE  P
DEAD KNOWLEDGE  CAMPBELL, JOHN W. JR.  B
DEAD LINES  SKIPP, JOHN/SPECTOR, CRAIG  P
DEAD LIVE, THE [@PERRY RHODAN #48]  ACE  PA,C/PA

The MITSFS Titledex  September 24, 2007  Page 83
TITLE  AUTHOR  CODE
DEAD MAN ON THE MOON [LCSI]  HARPER, STEVEN  LP
DEAD MAN RISING [DANTE VALENTINE 2]  SAINTCROW, LILITH  P
DEAD MAN’S GOLD  HOFFMAN, LEE  P
DEAD MAN’S HAND  COLLINS, NANCY A.  LP
DEAD MAN’S HAND [@WILD CARDS #7]  MARTIN, GEORGE R. R./MILLER, JOHN J.  C/P:2,P
DEAD MOON  KELLEY, LEO P.  SR-P
DEAD MORN  ANTHONY, PIERS/FUENTES, ROBERTO  C/P:2,SFWA-P,P
DEAD OF NIGHT  HAINING, PETER  HA
DEAD OF WINTER, THE [@THIEVES’ WORLD #7]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
DEAD ON MY FEET [HALFLIFE 2]  SIMMONS, WM. MARK  H
DEAD ROMANCE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  MILES, LAWRENCE  P

DEAD ROMANCE (WITH ADDITIONAL STORIES)

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] MILES, LAWRENCE

LP

DEAD SKY, BLACK SUN [@WARHAMMER 40K ULTRAMARINES 3]  MCNEILL, GRAHAM  P
DEAD TO THE WORLD [SOUTHERN VAMPIRES 4]  HARRIS, CHARLAINE  H
DEAD UNTIL DARK [SOUTHERN VAMPIRES 1]  HARRIS, CHARLAINE  P
DEAD WITCH WALKING [RACHEL MORGAN 1]  HARRISON, KIM  P
DEAD WORLD, THE  HARRIS, CLARE WINGER/BREUER, MILES J.  B
DEAD ZONE [@STAR TREK TNG MAXIMUM WARP #1]  GALANTER, DAVE/BRODEUR, GREG  P
DEAD ZONE, THE  KING, STEPHEN  H:2,SR-H,P
DEADALUS [STAR TREK ENTERPRISE]  STERN, DAVE  P
DEADEYE DICK  VONNEGUT, KURT JR.  H,P
DEADFALL [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RUSSELL, GARY  C/P,P
DEADHOUSE GATES [MALAZAN 2]  ERIKSON, STEVEN  C/P,LP
DEADLIEST SHOW IN TOWN, THE [MATHEW SWAIN 3]  MCQUAY, MIKE  P
DEADLINE TO PLUTO  STATTEN, VARGO  SR-P
DEADLY CYBORGS, THE [JOHN EAGLE EXPEDITER]  EDWARDS, PAUL  P
DEADLY DEEP, THE  MESSMANN, JON  P
DEADLY DUTCHMAN, THE [@RICK BRANT #22]  BLAINE, JOHN  H
DEADLY DWARF, THE [DOC SAVAGE #28]  ROBESON, KENNETH  C/P:2,@P
DEADLY GAME, A  JACKSON, ANITA  SR-P
DEADLY IMAGE  COOPER, EDMUND  C/P:2,P
DEADLY LITTER  WHITE, JAMES  C/P:2,P
DEADLY MESSIAH, THE  HILL, DAVID CAMPBELL/HILL, ALBERT FAY  C/P,P
DEADLY NIGHTSHADE  HAINING, PETER  HA
DEADLY QUICKSILVER LIES [GARRETT 7]  COOK, GLEN  P,SFWA-P
DEADLY RELATIONS [@BABYLON 5 PSI CORPS 2]  KEYES, J. GREGORY  P
DEADLY REUNION [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE/LETTS, BARRY  P
DEADLY SILENTS  KILLOUGH, LEE  C/P,P
DEADLY SKY, THE  JORGENSEN, IVAR  P
DEADLY SKY, THE  PISERCHIA, DORIS  P
DEADLY SLEEP [TWILIGHT 1]  COWAN, DALE  P
DEADLY STREETS, THE  ELLISON, HARLAN  C/P,P
DEADLY YEARS, THE (FOTONOVEL) [STAR TREK]  HARMON, DAVID P.  [C/RAN-P]
DEADMAN SWITCH  ZAHN, TIMOTHY  C/P:2,P
DEADMEN SHOULDN’T DIE [PERRY RHODAN 132]  DARLTON, CLARK  SR-LP
DEADSPAWN [NECROSCOPE 5]  LUMLEY, BRIAN  P
DEADSPEAK [NECROSCOPE 4]  LUMLEY, BRIAN  P,SFWA-P
DEADSTOCK [PUNKTOWN]  THOMAS, JEFFREY  P
DEADSTONE MEMORIAL, THE [@DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
DEADWALK [@VAMPIRELLA #3]  GOULART, RON  P
DEAKEN’S WAR  FREEMANTLE, BRIAN  SFWA-P
DEAL WITH THE DEVIL  DICHIARA, ROBERT  P
DEALER’S CHOICE [@WILD CARDS #11]  MARTIN, GEORGE R. R.  PA,C/PA
DEALERS IN LIGHT AND DARKNESS  WILDER, CHERRY  LP
DEALING IN FUTURES  HALDEMAN, JOE  H:2,P
DEALING WITH DRAGONS [ENCHANTED FOREST 1]  WREDE, PATRICIA C.  P
DEALS WITH THE DEVIL  DAVENPORT, BASIL  PA,C/PA

DEALS WITH THE DEVIL RESNICK, MIKE/GREENBERG, MARTIN H./

ESTLEMAN, LOREN D.

PA,C/PA

DEAN KOONTZ COMPANION, THE GREENBERG, MARTIN H./GORMAN, ED/

MUNSTER, BILL

SFWA

DEAR ABBEY  BISSON, TERRY  SR-LP
DEAR DEAD DAYS  HOCH, EDWARD D.  SFWA
DEATH  SCHIFF, STUART DAVID  PA
DEATH AMERICAN STYLE [@DOOMSDAY WARRIOR #12]  STACY, RYDER  P
DEATH AND DAMNATION [WORLD OF DARKNESS WRAITH]  KRAUSE, STALEY/WIECK, STEWART  PA
DEATH AND DIPLOMACY [DOCTOR WHO NEW ADVENTURES]  STONE, DAVE  C/P,P
DEATH AND LIFE OF SUPERMAN, THE [DC COMICS]  STERN, ROGER  C/P,P,H
DEATH AND THE MAIDEN [JONATHAN BARRETT 2]  ELROD, P. N.  C/P:2,P
DEATH AND THE SPIDER [@SPIDER #4]  STOCKBRIDGE, GRANT  P
DEATH AND THRAXAS [THRAXAS]  SCOTT, MARTIN  P
DEATH ANGEL’S SHADOW  WAGNER, KARL EDWARD  C/P,P
DEATH ARMS  JETER, K. W.  H

The MITSFS Titledex  September 24, 2007  Page 84
TITLE  AUTHOR  CODE
DEATH-BRINGER [@AMTRAK WARS #5]  TILLEY, PATRICK  P
DEATH BY DREAMING  WHITE, JON MANCHIP  P
DEATH CELL  GOULART, RON  C/P:2,P
DEATH CHANTS  STRETE, CRAIG  SFWA
DEATH CLOUD  MANNION, MICHAEL  P
DEATH-COACH  WILLIAMSON, J. N.  P
DEATH COUNT [STAR TREK]  GRAF, L. A.  C/P:2,P
DEATH DEALERS, THE  ASIMOV, ISAAC  P
DEATH DISCIPLE, THE  MOORE, ROBIN/GRIFFIN, GERALD G.  P
DEATH DIVE [@OMEGA SUB #5]  CAMERON, J. D.  P
DEATH DOLLS OF LYRA, THE  HOLLY, JOAN HUNTER  P
DEATH DRAWS FIVE [@WILD CARDS #17]  MILLER, JOHN J.  H
DEATH DREAM  BOVA, BEN  C/P,P,H
DEATH DREAM  MASTERTON, GRAHAM  SFWA-P
DEATH DREAMS  KATZ, WILLIAM  P
DEATH FROM THE SEA  ANTWORTH, TERRANCE  LP
DEATH GOD’S CITADEL, THE  COULSON, JUANITA  C/P:2,P
DEATH GOD’S DOOM [MALKAR 1]  TUBB, E. C.  LP
DEATH HUNT ON A DYING PLANET  RUSE, GARY ALAN  C/P,P
DEATH IN DELHI [@DANGEROUS JOURNEYS 3]  GYGAX, GARY  P
DEATH IN FLORENCE  EFFINGER, GEORGE ALEC  H,SFWA
DEATH IN SILVER [DOC SAVAGE #26]  ROBESON, KENNETH  C/P:2,@P
DEATH IN SLOW MOTION [@AVENGER #18]  ROBESON, KENNETH  C/P:2,P
DEATH IN THE SPIRIT HOUSE  STRETE, CRAIG  H
DEATH IN WINTER [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  H
DEATH IS A DREAM  TUBB, E. C.  H,DH034
DEATH IS A LONELY BUSINESS  BRADBURY, RAY  P
DEATH JAG  ELLIS, ALBERT C./SLATEN, JEFF  P
DEATH MACHINE, THE [@AVENGER #32]  ROBESON, KENNETH  P
DEATH MASKS [@DRESDEN FILES #5]  BUTCHER, JIM  C/P,P
DEATH MASQUE [JONATHAN BARRETT 3]  ELROD, P. N.  C/P,P
DEATH MASTER, THE  APPEL, BENJAMIN  P
DEATH MATCH [@BLOOD BOWL 3]  FORBECK, MATT  P
DEATH OF A DARKLORD [@RAVENLOFT]  HAMILTON, LAURELL K.  C/P,LP
DEATH OF A LEGEND, THE [HORSECLANS 8]  ADAMS, ROBERT  C/P:2,P
DEATH OF A NEUTRON STAR [STAR TREK VOYAGER]  KOTANI, ERIC/SMITH, DEAN WESLEY  C/P:2,P
DEATH OF A REGIMENT [JEFFERSON’S WAR 5]  RANDLE, KEVIN  C/P,P
DEATH OF A WORLD  FARJEON, J. JEFFERSON  H

DEATH OF AN ADEPT [ADEPT 5] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

H,P

DEATH OF ART, THE [DOCTOR WHO NEW ADVENTURES]  BUCHER-JONES, SIMON  C/P,P
DEATH OF CHAOS, THE [RECLUCE 5]  MODESITT, L. E. JR.  H:2,P
DEATH OF DOCTOR ISLAND, THE  WOLFE, GENE  SFWA-TD25,TD25
DEATH OF HONOR, A  FAUST, JOE CLIFFORD  C/P,P,H
DEATH OF PRINCES, THE [STAR TREK TNG]  PEEL, JOHN  C/P:2,P
DEATH OF SLEEP, THE [PLANET PIRATES 2]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P,P,H
DEATH OF THE DRAGON [@FORGOTTEN REALMS CORMYR #3]  GREENWOOD, ED/DENNING, TROY  P
DEATH OF THE FUHRER, THE  PUCCETTI, ROLAND  P
DEATH OF THE MOON  PHILLIPS, ALEXANDER M.  B
DEATH OF THE NECROMANCER, THE  WELLS, MARTHA  P
DEATH OF THE PHOENIX [MISSION CRITICAL]  CHAFE, PAUL C.  P
DEATH OF THE RAVEN [LADY BRANWEN 2]  MEANEY, DEE MORRISON  P
DEATH ON A WARM WIND  WARNER, DOUGLAS  C/P,P
DEATH OR GLORY [@WARHAMMER 40K CIAPHAS CAIN 4]  MITCHELL, SANDY  P
DEATH PLANET [STAR COMMANDOS 4]  GRIFFIN, P. M.  P
DEATH QUAKE [@TOM SWIFT (ARCHWAY) #12]  APPLETON, VICTOR  P
DEATH QUEST [@MISSION EARTH #6]  HUBBARD, L. RON  H,P
DEATH RAY, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
DEATH REIGN OF THE VAMPIRE KING [@SPIDER #1]  STOCKBRIDGE, GRANT  P
DEATH RIDERS OF HEL [HEL 2]  DRAKE, ASA  P
DEATH SCOUTS [MAN FROM ATLANTIS 2]  WOODLEY, RICHARD  P
DEATH SENTENCE  KELLEY, LEO P.  SR-P
DEATH: THE HIGH COST OF LIVING  GAIMAN, NEIL  CX
DEATH: THE TIME OF YOUR LIFE  GAIMAN, NEIL  CX
DEATH TIDE [V]  CRISPIN, A. C./MARSHALL, DEBORAH A.  C/P,P
DEATH TO THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DEATH TOLLS  STITH, JOHN E.  P
DEATH TOWER, THE  GRANT, MAXWELL  P
DEATH TREK  WALLMANN, JEFFREY N.  P
DEATH WAITS IN SEMISPACE [@PERRY RHODAN #61]  ACE  C/PA,PA
DEATH WORLD [@WARHAMMER 40K]  LYONS, STEVE  P
DEATHBEAST  GERROLD, DAVID  C/P:2,P
DEATHBELL  SMITH, GUY N.  P
DEATHBIRD STORIES  ELLISON, HARLAN  SR-H,C/P,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 85
TITLE  AUTHOR  CODE
DEATHBIRD STORIES (CORRECTED)  ELLISON, HARLAN  H:2
DEATHDAY [DEATHDAY 1]  DIETZ, WILLIAM C.  C/P,P
DEATHGAME [@VAMPIRELLA #5]  GOULART, RON  P
DEATHGIFT  ZEDDIES, ANN TONSOR  P
DEATHHUNTER  WATSON, IAN  C/P,P,H
DEATHKILLER [MINDKILLER 1,2]  ROBINSON, SPIDER  C/P,P
DEATHKNIGHT  OFFUTT, ANDREW J.  P
DEATHLESS, THE [@BUFFY]  DECANDIDO, KEITH R. A.  P
DEATHLORD OF IXIA, THE [LONE WOLF 17]  DEVER, JOE  P
DEATHMASQUES [JUDGE DREDD]  STONE, DAVE  P
DEATH’S ACOLYTE [SWORDS OF RAEMLLYN 4]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  P
DEATH’S ANGEL [STAR TREK]  SKY, KATHLEEN  C/P:2,P
DEATH’S CITY [@WARHAMMER BLOOD ON THE REIK 2]  MITCHELL, SANDY  P
DEATH’S DEPUTY  HUBBARD, L. RON  SR-H,P
DEATH’S GRAY LAND [@DESTINY MAKERS #4]  SHUPP, MIKE  C/P,P
DEATH’S HEAD REBELLION [@WAR WORLD #2]  POURNELLE, JERRY  PA,C/PA
DEATH’S LAW [TALES OF THE BARD #3]  SCOTT, MICHAEL  LP
DEATH’S MASTER [FLAT EARTH #2]  LEE, TANITH  C/P,@P
DEATH’S MESSENGER [@WARHAMMER BLOOD ON THE REIK 1]  MITCHELL, SANDY  C/P,P
DEATHSCAPE [@PSI-MAN #2]  DAVID, PETER  P
DEATHSTALKER [DEATHSTALKER 1]  GREEN, SIMON R.  C/P:2,P
DEATHSTALKER CODA [DEATHSTALKER]  GREEN, SIMON R.  H
DEATHSTALKER DESTINY [DEATHSTALKER 5]  GREEN, SIMON R.  C/P,P
DEATHSTALKER HONOR [DEATHSTALKER 4]  GREEN, SIMON R.  P
DEATHSTALKER LEGACY [DEATHSTALKER]  GREEN, SIMON R.  H
DEATHSTALKER REBELLION [DEATHSTALKER 2]  GREEN, SIMON R.  C/P,P
DEATHSTALKER RETURN [DEATHSTALKER]  GREEN, SIMON R.  P
DEATHSTALKER WAR [DEATHSTALKER 3]  GREEN, SIMON R.  C/P,P
DEATHSTAR VOYAGE  WALLACE, IAN  C/P,P,H:2
DEATHWEAVE  OSBORNE, CARY  C/P,P
DEATHWING [WARHAMMER 40K]  JONES, NEIL/PRINGLE, DAVID  PA
DEATHWING OVER VEYNAA [LAST LEGIONARY 2]  HILL, DOUGLAS  H,P
DEATHWISH WORLD  REYNOLDS, MACK/ING, DEAN  C/P:2,P
DEATHWORLD  HARRISON, HARRY  C/P:2,P
DEATHWORLD 2  HARRISON, HARRY  C/P,P
DEATHWORLD 3  HARRISON, HARRY  C/P:2,P
DEATHWORLD TRILOGY, THE  HARRISON, HARRY  C/P,P,H
DEATHWORMS OF KRATOS, THE [EXPENDABLES 1]  AVERY, RICHARD  C/P:2,P
DEBRACY’S DRUG  GRIDBAN, VOLSTED  SR-P
DEBT, THE  ROBINSON, G. P.  H
DEBT OF AGES  WHITE, STEVE  P
DEBT OF BONES [SWORD OF TRUTH 10]  GOODKIND, TERRY  H
DEBT OF HONOR [STAR TREK]  CLAREMONT, CHRIS  CX
DEBTORS’ PLANET [STAR TREK TNG]  THOMPSON, W. R.  C/P:2,P
DECADE OF FANTASY AND SCIENCE FICTION, A  MILLS, ROBERT P.  C/PA,PA
DECADE: THE 1940’S  ALDISS, BRIAN W./HARRISON, HARRY  HA
DECADE: THE 1950’S  ALDISS, BRIAN W./HARRISON, HARRY  HA

DECADENT RETURN OF THE HI-FI QUEEN AND HER EMBRYONIC REPTILE INFECTION, THE LOGAN, SIMON

ERD3

DECALOG [DOCTOR WHO] STAMMERS, MARK/

WALKER, STEPHEN JAMES

C/PA,PA

DECALOG 2: LOST PROPERTY [DOCTOR WHO] STAMMERS, MARK/

WALKER, STEPHEN JAMES

C/PA:2,PA

DECALOG 3: CONSEQUENCES [DOCTOR WHO]  LANE, ANDY/RICHARDS, JUSTIN  C/PA,PA
DECALOG 4 — RE:GENERATIONS [DOCTOR WHO]  LANE, ANDY/RICHARDS, JUSTIN  PA
DECALOG 5 — WONDERS [DOCTOR WHO]  LEONARD, PAUL/MORTIMORE, JIM  PA
DECEIT [DOCTOR WHO NEW ADVENTURES]  DARVILL-EVANS, PETER  C/P:2,P
DECEIVER, THE [CHAOS GATE 1]  COOPER, LOUISE  C/P,P
DECEIVERS, THE  BESTER, ALFRED  H:2,SFWA-P,LP
DECEPTION WELL  NAGATA, LINDA  P
DECEPTIONS [@BLENDING ENTHRONED #2]  GREEN, SHARON  P
DECEPTIONS [@LAWS OF THE BLOOD 4]  SIZEMORE, SUSAN  C/P:2,P
DECHRONIZATION OF SAM MAGRUDER, THE  SIMPSON, GEORGE GAYLORD  H
DECIPHER  PAVLOU, STEL  P
DECISION AT DOONA [DOONA]  MCCAFFREY, ANNE  C/P:2,P
DECISION AT THUNDER RIFT [@BATTLETECH 6]  KEITH, WILLIAM H. JR.  P
DECLARE  POWERS, TIM  H,P
DECLIN DE LA SAGESSE, LE  MARCEL, GABRIEL  FP
DECOY PRINCESS, THE [DECOY PRINCESS 1]  COOK, DAWN  C/P,P
DECREATION  STATTEN, VARGO  SR-P
DECTRA CHAIN [@DEATHLANDS]  AXLER, JAMES  P
DEDALE 2  PLANCHAT, HENRY-LUC  FP
DEDALUS BOOK OF AUSTRIAN FANTASY: 1890-2000, THE  MITCHELL, MIKE  LPA
DEDALUS BOOK OF FINNISH FANTASY, THE  SINISALO, JOHANNA  LPA

The MITSFS Titledex  September 24, 2007  Page 86
TITLE  AUTHOR  CODE
DEDALUS BOOK OF POLISH FANTASY, THE  POWAGA, WIESIEK  LPA
DEED OF PAKSENARRION, THE [PAKSENARRION 1,2,3]  MOON, ELIZABETH  H,SFWA
DEEP, THE  CROWLEY, JOHN  C/P,SFWA,P,H
DEEP BLUE [DOCTOR WHO PAST DOCTORS]  MORRIS, MARK  C/P,P
DEEP DOMAIN [STAR TREK]  WEINSTEIN, HOWARD  C/P:3,P
DEEP FIRE RISING (PROOF)  DU BRUL, JACK  SR-LP
DEEP FIX, THE  COLVIN, JAMES  SR-P
DEEP FREEZE  HUGHES, ZACH  C/P,P
DEEP FREEZE  WHYTE, H. WALTER  C/P,P
DEEP FUTURE  BAXTER, STEPHEN  [C/RAN-P:2]
DEEP FUTURE  BROWN, ERIC  H
DEEP GODS, THE  MASON, DAVID  C/P:2,P
DEEP QUARRY  STITH, JOHN E.  P
DEEP RANGE, THE  CLARKE, ARTHUR C.  C/P,SR-H,P,H
DEEP SECRET  JONES, DIANA WYNNE  H
DEEP SPACE  RUSSELL, ERIC FRANK  C/P,P,H
DEEP SPACE  SILVERBERG, ROBERT  HA,PA
DEEP STRIKE [SPEC OPS SQUAD 2]  SHELLEY, RICK  C/P,P
DEEP TREK [@EARTHBLOOD]  AXLER, JAMES  P
DEEP, VERY DEEP SPACE  NATHENSON, JOSEPH  C/P,P
DEEP WATER [BUFFY]  GILMAN, LAURA ANNE/SHERMAN, JOSEPHA P
DEEP WATERS  HODGSON, WILLIAM HOPE  SR-H
DEEP WIZARDRY  DUANE, DIANE  H
DEEPDRIVE  JABLOKOV, ALEXANDER  C/P:2,P
DEEPER SEA, A  JABLOKOV, ALEXANDER  C/P,SFWA,P,H
DEEPER THAN THE DARKNESS  BENFORD, GREGORY  P
DEEPEST SEA, THE  BARNITZ, CHARLES  P
DEEPNESS IN THE SKY, A [DEEP 2]  VINGE, VERNOR  C/P,P,H
DEEPSIX [PRISCILLA HUTCHINS 2]  MCDEVITT, JACK  H
DEEPWATER DREAMS  VAN SCYOC, SYDNEY J.  C/P:2,P
DEEPWOOD [KARAVANS 2]  ROBERSON, JENNIFER  H
DEER’S CRY, THE [KELTIAD]  KENNEALY-MORRISON, PATRICIA  C/P,P,H
DEERSKIN  MCKINLEY, ROBIN  P,SFWA
DEFENDER [ATEVI 5]  CHERRYH, C. J.  H:2,P
DEFENDERS, THE [HELMSMAN 5]  BALDWIN, BILL  C/P:2,P
DEFENDERS OF AR [GUARDIANS OF THE THREE 4]  LOVEJOY, JACK  C/P:2,P
DEFIANCE, THE [HELMSMAN 7]  BALDWIN, BILL  P
DEFIANT AGENTS, THE [TIME TRADERS 3]  NORTON, ANDRE  C/P,P,H

DEFIANT AGENTS, THE (REVISED) [TIME TRADERS 3,4]/ KEY OUT OF TIME (REVISED) [TIME TRADERS 3,4] NORTON, ANDRE

P

DEFINITELY DEAD [SOUTHERN VAMPIRES 6]  HARRIS, CHARLAINE  H:2
DEFINITELY MAYBE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P:2,P,H
DEITIES AND DEMIGODS  TSR GAMES  [C/RAN-H]
DEITY-FATHER, THE [@SPACE PRECINCT #1]  BISCHOFF, DAVID F.  P
DEJA DEMAIN  KUTTNER, HENRY/MOORE, C. L.  FP
DELAN THE MISLAID  MARKS, LAURIE J.  P
DELANY INTERSECTION, THE  SLUSSER, GEORGE EDGAR  R/REF-P,C/REF-P
DELETE ALL SUSPECTS [TURING HOPPER 4]  ANDREWS, DONNA  H
DELIA OF VALLIA [DRAY PRESCOT 28]  AKERS, ALAN BURT  P
DELICATE APE, THE  HUGHES, DOROTHY B.  SR-P,H
DELIGHTS OF TURKEY, THE  RODITI, EDOUARD  LP
DELIKON, THE  HOOVER, H. M.  C/P,P
DELIRIUM’S MISTRESS [@FLAT EARTH #4]  LEE, TANITH  P
DELIVERER [ATEVI 9]  CHERRYH, C. J.  H
DELON CITY [@OZ CHRONICLES #2]  RIDLEY, R. W.  LP
DELTA, THE [SPIDER WORLD 4]  WILSON, COLIN  P
DELTA AND THE BANNERMEN (MISSPELLED ON SPINE) [DOCTOR WHO]  KOHLL, MALCOLM  SR-P,C/P
DELTA CITY [HUMILITY GARDEN 2]  SAVAGE, FELICITY  P
DELTA GREEN: DENIED TO THE ENEMY [DELTA GREEN]  DETWILLER, DENNIS  LP
DELTA PAVONIS [ISLAND WORLDS 4]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P,P
DELTA SEARCH [@QUEST FOR TOMORROW 1]  SHATNER, WILLIAM  P
DELUGE, THE  DA VINCI, LEONARDO  SR-P
DELUGE  WRIGHT, S. FOWLER  H
DELUGE (PROOF)  WRIGHT, S. FOWLER  SR-LP
DELUGE DRIVERS, THE [HUMANX ICERIGGER 3]  FOSTER, ALAN DEAN  C/P:2,P
DELUSION WORLD  DICKSON, GORDON R.  DF119
DELUSION’S MASTER [@FLAT EARTH #3]  LEE, TANITH  C/P:2,P
DELUSIONS OF GRANDEUR [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
DEMANDS OF HONOR [@STAR TREK ERRAND OF FURY #2]  RYAN, KEVIN  P
DEMENTIA [@MUTANT CHRONICLES APOSTLE OF INSANITY 3]  STACKPOLE, MICHAEL A.  P
DEMOGORGON  LUMLEY, BRIAN  P
DEMOLISHED MAN, THE  BESTER, ALFRED  H:2,LP
DEMOLITION WINTER [SPACE: ABOVE AND BEYOND 2]  TELEP, PETER  C/P,P
DEMON, THE  SACKETT, JEFFREY  P

The MITSFS Titledex  September 24, 2007  Page 87
TITLE  AUTHOR  CODE
DEMON [GAEA 3]  VARLEY, JOHN  H:2,LP
DEMON  YOUNGER, JACK  P
DEMON-4  MACE, DAVID  C/P,P
DEMON AND THE CITY, THE [DETECTIVE INSPECTOR CHEN 2]  WILLIAMS, LIZ  LP
DEMON APOSTLE, THE [DEMONWARS 3]  SALVATORE, R. A.  C/P,P,H
DEMON AWAKENS, THE [DEMONWARS 1]  SALVATORE, R. A.  C/P,SR-LP,H
DEMON BLADE [DEMON BLADE 1]  GARLAND, MARK A./MCGRAW, CHARLES G. C/P,P
DEMON BLUES [HERE BE DEMONS 2]  FRIESNER, ESTHER M.  P
DEMON BREED, THE  SCHMITZ, JAMES H.  C/P:2,P,H
DEMON CAT  CLARK, LYDIA BENSON  P
DEMON DOPPELGANGERS [@CHARMED]  ELLIOT, GREG  P
DEMON DOWNLOAD [DARK FUTURE DEMON DOWNLOAD 1]  YEOVIL, JACK  LP
DEMON DRUMS  SEVERANCE, CAROL  C/P,P
DEMON HAND, THE [@GREYHAWK #5]  ESTES, ROSE  C/P:2,P
DEMON IN THE MIRROR [WAR OF THE WIZARDS 1]  OFFUTT, ANDREW J./LYON, RICHARD K.  C/P,P
DEMON IN THE SKULL  POHL, FREDERIK  C/P:2,P
DEMON ISLAND [@AVENGER #36]  ROBESON, KENNETH  C/P:2,P
DEMON KIND  ELWOOD, ROGER  PA
DEMON KING, THE [DAMASTES 2]  BUNCH, CHRIS  LP
DEMON KNIGHT  BOYLL, RANDALL  P
DEMON KNIGHT [LONGDIRK 3]  HOOD, KEN  P
DEMON LORD, THE [BOOK OF YEARS 2]  MORWOOD, PETER  C/P,P
DEMON LORD OF KARANDA [MALLOREON 3]  EDDINGS, DAVID  H:2,P
DEMON LOVER, THE  VALDEMI, MARIA  P
DEMON LOVERS AND STRANGE SEDUCTIONS  CARTER, M. L.  PA,C/PA
DEMON MOON  WILLIAMSON, JACK  H,P
DEMON NIGHT [@RED SONJA #2]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
DEMON OF BARNABAS COLLINS, THE  ROSS, MARILYN  C/P,P
DEMON OF CAWNPORE, THE  VERNE, JULES  H,P

DEMON OF SCATTERY, THE ANDERSON, POUL/

BROXON, MILDRED DOWNEY

H:2,LP

DEMON OF THE DARK ONES [WAR OF POWERS 6]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
DEMON OF THE EIFFEL TOWER, THE  TARDI, JACQUES  CX
DEMON OF THE NIGHT  LAKIN, RITA  P
DEMON OF UNDOING  ALTON, ANDREA I.  P
DEMON PIG  BRUSH, KAREN A.  C/P,P
DEMON PRINCES, THE [DEMON PRINCES 1,2,3,4,5]  VANCE, JACK  H
DEMON RIDER [LONGDIRK 2]  HOOD, KEN  P
DEMON SAMURAI, THE  GRANT, CLAY  P
DEMON SEED  KOONTZ, DEAN R.  C/P:2,P
DEMON SHADOWS  SIROTA, MIKE  P
DEMON SOUL, THE [@WARCRAFT WAR OF THE ANCIENTS #2]  KNAAK, RICHARD A.  P
DEMON SPIRIT, THE [DEMONWARS 2]  SALVATORE, R. A.  C/P,P
DEMON SWORD [LONGDIRK 1]  HOOD, KEN  P
DEMON TREE [TWILIGHT 9]  DANIEL, COLIN  P
DEMON WING [@SPACE PRECINCT #2]  BISCHOFF, DAVID F.  P
DEMON WITCH [@RAVENSCLIFF #2]  HUNTINGTON, GEOFFREY  P
DEMON YOU KNOW, THE [OTHERS (WARREN)]  WARREN, CHRISTINE  P
DEMONS [@WITCHBLADE]  BARON, MIKE  P
DEMONS, THE  BULMER, KENNETH  SR-P
DEMONS!  DANN, JACK/DOZOIS, GARDNER  PA
DEMONS [STAR TREK]  DILLARD, J. M.  C/P:3,P

DEMONS AT RAINBOW BRIDGE, THE [QUINTARA 1] CHALKER, JACK L.

C/P,SFWA-P:2,H:2,P,

SFWA

DEMONS BY DAYLIGHT  CAMPBELL, RAMSEY  SR-H,P
DEMON’S DAUGHTER, THE  HOLLY, EMMA  H
DEMONS DON’T DREAM [@XANTH #16]  ANTHONY, PIERS  H:2,P
DEMONS ET MERVEILLES  LOVECRAFT, H. P.  FP
DEMON’S GATE  WHITE, STEVE  P
DEMON’S GATE (PROOF)  WHITE, STEVE  SR-LP
DEMONS IN THE GREEN, THE  DEITZ, TOM  P
DEMON’S MIRROR, THE  WALLERSTEIN, JAMES S.  H

DEMONS OF AIR AND DARKNESS [STAR TREK GATEWAYS #4/ STAR TREK DS9] DECANDIDO, KEITH R. A.

P

DEMONS OF EDEN [@DEATHLANDS]  AXLER, JAMES  P
DEMONS OF SANDORRA, THE  TABORI, PAUL  P
DEMONS OF THE DANCING GODS [DANCING GODS 2]  CHALKER, JACK L.  C/P:2,P
DEMONS OF ZAMMAR, THE [@RO-LAN #4]  SIROTA, MIKE  P
DEMONS’ WORLD  BULMER, KENNETH  DF289
DEMONSTORM [LEGENDS OF THE RAVEN 3]  BARCLAY, JAMES  P
DEMONTAGE [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  C/P,P
DEMU TRILOGY, THE [DEMU 1,2,3]  BUSBY, F. M.  C/P:2,P
DENNER’S WRECK  WATT-EVANS, LAWRENCE  C/P,P,H:2
DENVER IS MISSING  JONES, D. F.  C/P,P,H

TITLE  AUTHOR  CODE
DENY THY FATHER [@STAR TREK LOST ERA (2355-57 AD)]  MARIOTTE, JEFF  P
DEPARTURES  TURTLEDOVE, HARRY  C/P:2,P
DEPRESSION OR BUST/DAWNMAN PLANET  REYNOLDS, MACK  C/P:2,P
DEPRIVERS  ALTMAN, STEVEN-ELLIOT  LP
DEPTHS, THE  CREASEY, JOHN  P
DEPTHS OF MADNESS [@FORGOTTEN REALMS DUNGEONS]  DE BIE, ERIK SCOTT  P
DEPTHS OF TIME, THE [SOLACE 1]  ALLEN, ROGER MACBRIDE  H,LP
DEPUTY SHERIFF OF COMMANCHE COUNTY, THE  BURROUGHS, EDGAR RICE  P
DERAI [DUMAREST 2]  TUBB, E. C.  C/P,DH077,D89301
DERELICT  HOVORKA, ROBERT L.  C/P,P
DERELICT FOR TRADE [SOLAR QUEEN 5]  NORTON, ANDRE/SMITH, SHERWOOD  C/P,P
DERIK’S BANE [WYNDHAM WEREWOLF]  DAVIDSON, MARYJANICE  P
DERRICK DEVIL, THE [DOC SAVAGE #74]  ROBESON, KENNETH  C/P:2,@P
DERVISH DAUGHTER [TRUE GAME 5]  TEPPER, SHERI S.  C/P,SFWA-P,P
DERVISH IS DIGITAL [KONSTANTIN 2]  CADIGAN, PAT  H
DERYNI ARCHIVES, THE [DERYNI]  KURTZ, KATHERINE  C/P,P,H
DERYNI CHALLENGE [DERYNI]  BILLIAS, STEPHEN  P,SFWA-P
DERYNI CHECKMATE [DERYNI CHRONICLES 2]  KURTZ, KATHERINE  C/P:3,P
DERYNI MAGIC [DERYNI]  KURTZ, KATHERINE  C/P:2,P
DERYNI RISING [DERYNI CHRONICLES 1]  KURTZ, KATHERINE  C/P:3,P
DERYNI TALES [DERYNI]  KURTZ, KATHERINE  C/PA,PA
DESCENT [STAR TREK TNG]  CAREY, DIANE  C/P,P
DESCENT, THE  LONG, JEFF  P
DESCENT [@DESCENT 1]  TELEP, PETER  P
DESCENT INTO HELL  WILLIAMS, CHARLES  H,LP
DESCENT INTO THE DEPTHS OF THE EARTH [@GREYHAWK]  KIDD, PAUL  P
DESCENT OF ANANSI, THE  NIVEN, LARRY/BARNES, STEVEN  C/P:2,P
DESERT, THE [SPIDER WORLD 1]  WILSON, COLIN  P
DESERT MERCENARY [@CASCA #16]  SADLER, BARRY  P
DESERT OF DEATH’S DOMAIN [@PERRY RHODAN #100]  ACE  C/PA,PA
DESERTED CITIES OF THE HEART  SHINER, LEWIS  C/P,P,H
DESFILE DE LA VICTORIA, EL  DIAZ-PLAJA, FERNANDO  FP
DESIERTO DE NIEBLA Y CENIZAS  TRIGO, JOAN  FP
DESIGN FOR GREAT-DAY  FOSTER, ALAN DEAN/RUSSELL, ERIC FRANK C/P:2,P
DESIGNATED TARGETS [AXIS OF TIME 2]  BIRMINGHAM, JOHN  LP
DESOLATION ROAD [GUARDIANS 8]  AUSTIN, RICHARD  P
DESOLATION ROAD  MCDONALD, IAN  C/P:2,P
DESPATCHES FROM THE FRONTIERS OF THE FEMALE MIND  GREEN, JEN/LEFANU, SARAH  SR-PA
DESPERATE, DECAYING DARKNESS, A  CASTRO, ADAM-TROY  LP
DESPERATE GAMES  BOULLE, PIERRE  H
DESPERATE MEASURES [ANGEL’S LUCK 1]  FAUST, JOE CLIFFORD  C/P,P,H
DESPERATION  KING, STEPHEN  C/P,P
DESTINATION ALPHA  CAMERON, BERL  SR-P
DESTINATION: AMALTHEIA  DIXON, RICHARD  SR-PA,C/PA
DESTINATION: BRAIN  MCEVOY, SETH  P
DESTINATION: INFINITY  KUTTNER, HENRY  C/P:2,P
DESTINATION MARS  BROWN, GEORGE SHELDON  SR-P
DESTINATION MOON  KENTON, L. P.  SR-P
DESTINATION: MUTINY [@STARHAWKS #5]  DALTON, SEAN  C/P:2,P
DESTINATION: SATURN  GRINNELL, DAVID/CARTER, LIN  DH085
DESTINATION: TERRA  HASSLER, KENNETH W.  H
DESTINATION: TERRE!  WUPI  CX
DESTINATION: UNIVERSE!  VAN VOGT, A. E.  C/P,P,H
DESTINATION UNKNOWN  CROWTHER, PETER  LPA
DESTINATION: VOID [SHIP 1]  HERBERT, FRANK  C/P:2,P
DESTINAZIONE UOMO  GALASSIA  FP
DESTINED QUEEN, THE [MAURA WOODBURY 2]  HALE, DEBORAH  LP
DESTINIES 1 #1  BAEN, JAMES  PA,C/PA
DESTINIES 1 #2  BAEN, JAMES  C/PA,PA
DESTINIES 1 #3  BAEN, JAMES  C/PA,PA
DESTINIES 1 #4  BAEN, JAMES  C/PA,PA
DESTINIES 1 #5  BAEN, JAMES  PA,C/PA
DESTINIES 2 #1  BAEN, JAMES  PA
DESTINIES 2 #2  BAEN, JAMES  PA
DESTINIES 2 #3  BAEN, JAMES  C/PA,PA
DESTINIES 2 #4  BAEN, JAMES  PA
DESTINIES 3 #1  BAEN, JAMES  C/PA,PA
DESTINIES 3 #2  BAEN, JAMES  PA,C/PA
DESTINY [@BLENDING ENTHRONED #3]  GREEN, SHARON  C/P,P
DESTINY [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  P
DESTINY [SYMPHONY OF AGES 3]  HAYDON, ELIZABETH  H,P
DESTINY  HUGHES, RUPERT  H
DESTINY AND THE DOLPHINS [DOLPHIN 3]  MEYERS, ROY  C/P:2,P
DESTINY DICE, THE [GAMING MAGI 1]  BISCHOFF, DAVID F.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 89
TITLE  AUTHOR  CODE
DESTINY DOLL  SIMAK, CLIFFORD D.  C/P:2,P
DESTINY MAKERS, THE  TURNER, GEORGE  H:2,SFWA,P
DESTINY MASK, THE [STRUCTURE 2]  SKETCHLEY, MARTIN  LP
DESTINY OF THE SWORD, THE [SEVENTH SWORD 3]  DUNCAN, DAVE  C/P:2,P
DESTINY STONE, THE [WAR OF POWERS 3]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
DESTINY TIMES THREE  LEIBER, FRITZ  C/P,GN5.28,BS1
DESTINY’S END  SULLIVAN, TIM  C/P:2,P
DESTINY’S FORGE [@MAN-KZIN WARS]  CHAFE, PAUL C.  H
DESTINY’S ORBIT  GRINNELL, DAVID  H,DF161
DESTINY’S ROAD  NIVEN, LARRY  C/P,P,H
DESTINY’S SHIELD [BELISARIUS 3]  FLINT, ERIC/DRAKE, DAVID  C/P:2,P
DESTINY’S WAY [@STAR WARS NEW JEDI ORDER]  WILLIAMS, WALTER JON  P
DESTROYER [ATEVI 7]  CHERRYH, C. J.  H:2
DESTROYER GODDESS, THE [IN FIRE FORGED 2]  RESNICK, LAURA  P
DESTROYERS, THE  MANNING, P. L.  SR-P
DESTROYING ANGEL [FRANK CARLUCCI 1]  RUSSO, RICHARD PAUL  C/P,P
DESTROYING WORLDS [@ENEMIES OF SOCIETY #2]  DAVID, JOHN  LP
DESTRUCTION OF ILLUSIONS [@ANDROMEDA (GENE RODDENBERRY’S) #1]  DECANDIDO, KEITH R. A.  P
DESTRUCTION OF THE BOOKS, THE [ROVER 2]  ODOM, MEL  H
DESTRUCTION OF THE TEMPLE, THE  MALZBERG, BARRY N.  C/P,P
DETACHED RETINA, THE  ALDISS, BRIAN W.  R/REF-P
DETAILS OF THE HUNT  BAUMBACH, LAURA  LP
DETOUR  WILSON, WILLIAM  P
DEUS ENCARMINE [@WARHAMMER 40K BLOOD ANGELS 1]  SWALLOW, JAMES  P
DEUS EX MACHINA  BRUMMELS, J. V.  C/P:2,P
DEUS IRAE  DICK, PHILIP K./ZELAZNY, ROGER  C/P,SFWA,P,H
DEUS LE VOLT [TIME HUNTER 8]  MILLER, JON DE BURGH  SR-H
DEUS MACHINE, THE  OUELLETTE, PIERRE  C/P,P,H
DEUS SANGUINIUS [@WARHAMMER 40K BLOOD ANGELS 2]  SWALLOW, JAMES  P
DEUS X  SPINRAD, NORMAN  P
DEUX ROBOTS, LES  BRUCKNER, KARL  FH
DEVIANT STRAIN, THE [DOCTOR WHO NINTH DOCTOR]  RICHARDS, JUSTIN  H
DEVIATES, THE  JONES, RAYMOND F.  SR-P,C/P:2
DEVICE OF DEATH, A [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P
DEVIL & DAN COOLEY, THE [SYMPATHY FOR THE DEVIL 2]  LISLE, HOLLY  C/P,P
DEVIL AND DEEP SPACE, THE [JUDICIARY 4]  MATTHEWS, SUSAN R.  P
DEVIL AND MAX DEVLIN, THE  GROSSBACH, ROBERT  P
DEVIL AND W. KASPAR, THE  APPEL, BENJAMIN  C/P,P
DEVIL DELIVERED, THE  ERIKSON, STEVEN  SR-H
DEVIL GENGHIS, THE [DOC SAVAGE #79]  ROBESON, KENNETH  C/P:2,@P
DEVIL GOBLINS FROM NEPTUNE, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH/DAY, MARTIN  C/P:2,P
DEVIL HIS DUE, THE  HILL, DOUGLAS  C/PA,PA
DEVIL IN A FOREST, THE  WOLFE, GENE  C/P:2,P
DEVIL IN IRON, THE  HOWARD, ROBERT E.  VLP
DEVIL IN THE DARK, THE [STAR TREK]  COON, GENE L.  SR-S
DEVIL IN THE DUST, THE [@OUTREMER #1]  BRENCHLEY, CHAZ  P
DEVIL IN THE SKY [STAR TREK DS9]  COX, GREG/BETANCOURT, JOHN GREGORY  C/P:2,P
DEVIL IN VELVET, THE  CARR, JOHN DICKSON  P
DEVIL IS DEAD, THE  LAFFERTY, R. A.  C/P:2,P
DEVIL IS NOT MOCKED, THE  WELLMAN, MANLY WADE  H
DEVIL-MAY-CARE  PETERS, ELIZABETH  SFWA-P
DEVIL OF PEI-LING, THE  ASBURY, HERBERT  H
DEVIL ON LAMMAS NIGHT, THE  HOWATCH, SUSAN  P
DEVIL ON MY BACK [ARC ONE 1]  HUGHES, MONICA  P
DEVIL ON THE MOON [DOC SAVAGE #50]  ROBESON, KENNETH  C/P:2,@P
DEVIL ON THE ROAD, THE  WESTALL, ROBERT  P
DEVIL RIDES OUT, THE  WHEATLEY, DENNIS  C/P,P
DEVIL WIND [@DARK FORCES 4]  BRIDGES, LAURIE/ALEXANDER, PAUL  P
DEVIL WIVES OF LI FONG, THE  PRICE, E. HOFFMANN  C/P:2,P
DEVIL WORLD [STAR TREK]  EKLUND, GORDON  C/P:2,P
DEVIL YOU SAY, THE  BREIG, JOSEPH A.  H
DEVIL YOU SAY, THE  DECARLO, ELISA  C/P,P
DEVILDAY  HALL, ANGUS  P

DEVILS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

DEVIL’S ADVOCATE, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #1] FLEMING, GHERBOD

P

DEVILS AND DEMONS  KAYE, MARVIN  HA,SFWA
DEVIL’S ARMOR [EYES OF GOD 2]  MARCO, JOHN  P
DEVIL’S BARGAIN  TARR, JUDITH  LP
DEVIL’S BRIDE, THE  QUINN, SEABURY  P
DEVIL’S CHILDREN, THE [CHANGES 1]  DICKINSON, PETER  H
DEVIL’S CHILDREN, THE  PARRY, MICHEL  PA
DEVIL’S DAY, THE  BLISH, JAMES  C/P:2,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 90
TITLE  AUTHOR  CODE
DEVIL’S DICTIONARY, THE  BIERCE, AMBROSE  H:2,LP
DEVIL’S ENGINE [DEVIL’S WEST 2]  SUMNER, MARK  P
DEVIL’S GAMBLE  SLAUGHTER, FRANK G.  H,P
DEVIL’S GAME, THE  ANDERSON, POUL  C/P:2,P
DEVIL’S GENERATION, THE  GHIDALIA, VIC  C/PA,PA
DEVIL’S GUARD, THE  MUNDY, TALBOT  P
DEVIL’S HAND, THE [@SENTINELS #1/ROBOTECH 13]  MCKINNEY, JACK  C/P,P
DEVIL’S HEART, THE [STAR TREK TNG]  CARTER, CARMEN  C/P,P,H
DEVIL’S HORNS, THE [@AVENGER #4]  ROBESON, KENNETH  C/P:2,P
DEVIL’S MISTRESS, THE  BRODIE-INNES, J. W.  P
DEVIL’S MISTRESS, THE [LUCIFER COVE 1]  COFFMAN, VIRGINIA  P
DEVIL’S MODE, THE  BURGESS, ANTHONY  P
DEVILS OF D-DAY, THE  MASTERTON, GRAHAM  P

DEVILS OF THE DEEP [DOC SAVAGE #123]/ HEADLESS MEN, THE [DOC SAVAGE 124] ROBESON, KENNETH

@P

DEVIL’S PLAYGROUND, THE [DOC SAVAGE #25]  ROBESON, KENNETH  C/P:2,@P
DEVIL’S RAIN, THE  WILLIS, MAUD  P
DEVIL’S TANGO [DEVIL DANCES 2]  JUBERT, HERVE  H
DEVIL’S TOWER [DEVIL’S WEST 1]  SUMNER, MARK  C/P,P
DEVIL’S VIRGIN, THE [LUCIFER COVE 3]  COFFMAN, VIRGINIA  P
DEVIL’S WORK, THE  BRUNNER, JOHN  H
DEVINE WAR  CARO, DENNIS R.  H,SFWA
DEVLIN’S HONOR [SWORD OF CHANGE 2]  BRAY, PATRICIA  C/P,P
DEVLIN’S HONOR (PROOF) [SWORD OF CHANGE 2]  BRAY, PATRICIA  SR-P
DEVLIN’S JUSTICE [SWORD OF CHANGE 3]  BRAY, PATRICIA  C/P,P
DEVLIN’S LUCK [SWORD OF CHANGE 1]  BRAY, PATRICIA  C/P,P
DEVOLUTIONIST, THE/EMANCIPATRIX, THE  FLINT, HOMER EON  C/P:2,P
DEVOURING FIRE  STATTEN, VARGO  SR-P
DEVOURING VOID, THE [@BLOOD OF THE LAMB #2]  ROGERS, MARK E.  P
DEXTA [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 1]  RYAN, C. J.  P
DEZRA’S QUEST [@DRAGONLANCE BRIDGES OF TIME]  PIERSON, CHRIS  P
DHALGREN  DELANY, SAMUEL R.  C/P,P,H
DHAMPIR [NOBLE DEAD SERIES 1 1]  HENDEE, BARB/HENDEE, J. C.  P
DHAMPIRE  BAKER, SCOTT  C/P:2,P
DHIAMMARA [AURIAN 4]  FUREY, MAGGIE  P
DIA DE LOS MUERTOS, EL  HOPKINS, BRIAN A.  SR-H
DIABOLIST, THE  FAIRMAN, PAUL W.  P
DIABOLS, THE  MACKELWORTH, R. W.  C/P:2,P
DIABOLUS  SAINT JOHN, DAVID  P
DIADEM FROM THE STARS [DIADEM 1]  CLAYTON, JO  C/P:3,P
DIALING THE WIND  GRANT, CHARLES L.  SFWA-P
DIALOGUE WITH DARKNESS  ANDERSON, POUL  C/P:2,P
DIALOGUES WITH THE DEVIL  CALDWELL, TAYLOR  C/P,P
DIAMOND AGE, THE  STEPHENSON, NEAL  C/P,P,H
DIAMOND CONTESSA, THE [KEYS TO THE DIMENSIONS 8]  BULMER, KENNETH  P
DIAMOND DOGS, TURQUOISE DAYS [REVELATION SPACE]  REYNOLDS, ALASTAIR  H
DIAMOND MASK [GALACTIC MILIEU 2]  MAY, JULIAN  C/P,P,H
DIAMOND MOON, THE [@VENUS PRIME #5]  PREUSS, PAUL  C/P:2,P
DIAMOND THRONE, THE [ELENIUM 1]  EDDINGS, DAVID  H:2,P
DIAMONDS ARE FOREVER [JAMES BOND 4]  FLEMING, IAN  P

DIANA WYNNE JONES ROSENBERG, TEYA/HIXON, MARTHA P./

SCAPPLE, SHARON M./WHITE, DONNA R.

R/REF-H

DIANETICS  HUBBARD, L. RON  [C/RAN-H]
DIAPASON  SULLIVAN, THOMAS  P
DIASPORA  EGAN, GREG  C/P,P,H
DIASPORAH  YATES, W. R.  C/P,P
DICING WITH DRAGONS  LIVINGSTONE, IAN  [C/RAN-P]
DICK TRACY MEETS THE NIGHT CRAWLER  GOULD, CHESTER  H
DICKSON!  DICKSON, GORDON R.  SR-H
DICTATOR [@ROBOTS IN TIME 4]  WU, WILLIAM F.  C/P,P
DICTIONARY OF IMAGINARY PLACES, THE  MANGUEL, ALBERTO/GUADALUPI, GIANNI  R/REF-H,C/REF-H
DICTIONARY OF IMAGINARY PLACES, THE (EXPANDED EDITION)  MANGUEL, ALBERTO/GUADALUPI, GIANNI  R/REF-H
DICTIONARY OF THE KHAZARS (MALE EDITION)  PAVIC, MILORAD  LP
DID MAN GET HERE BY EVOLUTION OR BY CREATION?  WATCHTOWER  [C/RAN-H]
DID SHE FALL?  SMITH, THORNE  H
DID YOU SAY CHICKS?! [CHICKS IN CHAINMAIL 2]  FRIESNER, ESTHER M.  C/PA,PA
DIDO AND PA  AIKEN, JOAN  H
"DIE, CHAMELEON!" [@BATTLESTAR GALACTICA 1978 #12]  LARSON, GLEN A./THURSTON, ROBERT  P
DIE FOR LOVE  PETERS, ELIZABETH  SFWA-P
DIE MACHT DER SCHATTEN [@PERRY RHODAN #133]  EWERS, H. G.  FPR
DIE TODESMASCHINE [@PERRY RHODAN #119]  EWERS, H. G.  FPR
DIE WELT DER GLUCKSELIGEN [@PERRY RHODAN #65]  EWERS, H. G.  FPR
DIE ZEITSPRINGER [@PERRY RHODAN #7]  EWERS, H. G.  FPR
DIES THE FIRE [EMBERVERSE 1]  STIRLING, S. M.  H

TITLE  AUTHOR  CODE
DIEU COULEUR DE NUIT, LE  LIMAT, MAURICE  FP
DIFERENTE ESPACIO, DIFERENTE TIEMPO  CAMPBELL, H. J.  FP
DIFFERENCE ENGINE, THE  GIBSON, WILLIAM/STERLING, BRUCE  H:2,P
DIFFERENT FLESH, A  TURTLEDOVE, HARRY  C/P:2,P
DIFFERENT LIGHT, A  LYNN, ELIZABETH A.  C/P:2,P
DIFFERENT SEASONS  KING, STEPHEN  C/P:2,P
DIFFICULTY WITH DWARVES, A [WUNTVOR 1]  GARDNER, CRAIG SHAW  C/P:2,P
DIG, THE  FOSTER, ALAN DEAN  C/P,P,H
DIG IN TIME, A  GRIFFIN, PENI R.  SFWA
DIGGING LEVIATHAN, THE  BLAYLOCK, JAMES P.  P
DIGITAL DEAD, THE  BALFOUR, BRUCE  P
DIGITAL EFFECT, THE  PERRY, STEVE  C/P,P
DIGITAL KNIGHT  SPOOR, RYK E.  C/P,P
DIGITS AND DASTARDS  POHL, FREDERIK  C/P:2,P
DILATION EFFECT  MASON, DOUGLAS R.  C/P:2,P
DILEMMAS  HOUGHTON MIFFLIN  LP
DILVISH, THE DAMNED  ZELAZNY, ROGER  C/P,P,H

DIME STORE MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1] ARMSTRONG, KELLEY

C/P:2,P

DIMENSION 4  CONKLIN, GROFF  PA,C/PA
DIMENSION A  DAVIES, L. P.  C/P:2,P
DIMENSION OF DREAMS [@BLADE #11]  LORD, JEFFREY  C/P,P
DIMENSION OF HORROR [@BLADE #30]  LORD, JEFFREY  P
DIMENSION OF MIRACLES  SHECKLEY, ROBERT  C/P:3,P
DIMENSION RIDERS, THE [DOCTOR WHO NEW ADVENTURES]  BLYTHE, DANIEL  C/P:2,P
DIMENSION SEARCH [@PERRY RHODAN #60]  ACE  PA
DIMENSION THIRTEEN  SILVERBERG, ROBERT  C/P:2,P
DIMENSION X  KNIGHT, DAMON  HA
DIMENSION X [@WEB WARRIORS #2]  LUCENO, JAMES  P
DIMENSIONEERS, THE  PISERCHIA, DORIS  C/P,P
DIMENSIONS  SHAW, BOB  P
DIMENSIONS OF SHECKLEY  SHECKLEY, ROBERT  H
DINNER AT DEVIANT’S PALACE  POWERS, TIM  C/P,P,H
DINO DE LAURENTIIS PRODUCTION OF KING KONG, THE  SEMPLE, LORENZO JR.  P
DINOSAUR BEACH  LAUMER, KEITH  H:2,SFWA,P
DINOSAUR CONQUEST [RAY BRADBURY’S DINOSAURS 6]  LEIGH, STEPHEN  P
DINOSAUR EMPIRE [RAY BRADBURY’S DINOSAURS 5]  LEIGH, STEPHEN/MILLER, JOHN J.  P
DINOSAUR FANTASTIC  RESNICK, MIKE/GREENBERG, MARTIN H.  HA,PA
DINOSAUR NEXUS  GRIMES, LEE  P
DINOSAUR PARK  PEIRCE, HAYFORD  P
DINOSAUR PLANET [RAY BRADBURY’S DINOSAURS 2]  LEIGH, STEPHEN  P,SFWA-P
DINOSAUR PLANET [DINOSAUR PLANET 1]  MCCAFFREY, ANNE  C/P:3,P
DINOSAUR PLANET SURVIVORS [DINOSAUR PLANET 2]  MCCAFFREY, ANNE  C/P:2,P
DINOSAUR SAMURAI [RAY BRADBURY’S DINOSAURS 3]  LEIGH, STEPHEN/MILLER, JOHN L.  P
DINOSAUR SUMMER  BEAR, GREG  P
DINOSAUR THAT FOLLOWED ME HOME, THE [CAMP HAUNTED HILL 3]  COVILLE, BRUCE  LP
DINOSAUR WARRIORS [RAY BRADBURY’S DINOSAURS 4]  LEIGH, STEPHEN  P
DINOSAUR WORLD [RAY BRADBURY’S DINOSAURS 1]  LEIGH, STEPHEN  P
DINOSAURS!  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
DINOSAURS  WATSON, JANE WERNER  [C/RAN-H]
DINOSAURS II  DANN, JACK/DOZOIS, GARDNER  PA
DINOTOPIA CALENDAR, THE (1994)  GURNEY, JAMES  CAL
DINOTOPIA LOST [DINOTOPIA]  FOSTER, ALAN DEAN  H
DINOVERSE  CIENCIN, SCOTT  H
DIONYSUS: THE ULTIMATE EXPERIMENT  RUBEN, WILLIAM S.  P
DIPLOIDS, THE  MACLEAN, KATHERINE  C/P:2,P
DIPLOMACY GUILD, THE [@ISAAC’S UNIVERSE #1]  GREENBERG, MARTIN H.  C/PA,PA
DIPLOMACY OF WOLVES [SECRET TEXTS 1]  LISLE, HOLLY  C/P:2,LP
DIPLOMATIC ACT  JURASIK, PETER/KEITH, WILLIAM H. JR.  P
DIPLOMATIC IMMUNITY [MILES VORKOSIGAN 10]  BUJOLD, LOIS MCMASTER  H,P
DIPLOMATIC IMPLAUSIBILITY [STAR TREK TNG]  DECANDIDO, KEITH R. A.  P
DIRDIR, THE  VANCE, JACK  C/P:2,P
DIRECT DESCENT  HERBERT, FRANK  C/P:2,P
DIRECTORY OF FANDOM (1969)  NESFA  SR-L
DIRGE [HUMANX FOUNDING COMMONWEALTH 2]  FOSTER, ALAN DEAN  H
DIRGE FOR SABIS, A [SWORD OF KNOWLEDGE #1]  CHERRYH, C. J./FISH, LESLIE  C/P:2,SFWA-P,P
DIRK GENTLY’S HOLISTIC DETECTIVE AGENCY [DIRK GENTLY 1]  ADAMS, DOUGLAS  H:2,P
DIRTY, ROTTEN, DEPRIVING RAY, THE [@MAN FROM TOMCAT #4]  KNIGHT, MALLORY T.  P
DIRTY TRICKS  EFFINGER, GEORGE ALEC  H,SFWA
DIRTY WORK  CADIGAN, PAT  H
DIRTY WORK [JASON COSMO 3]  MCGIRT, DAN  C/P,P
DISAGREEMENT WITH DEATH, A [WUNTVOR 3]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
DISAPPEARANCE, THE  WYLIE, PHILIP  C/P,P
DISAPPEARED, THE [RETRIEVAL ARTIST 1]  RUSCH, KRISTINE KATHRYN  P

The MITSFS Titledex  September 24, 2007  Page 92
TITLE  AUTHOR  CODE
DISAPPEARING ACT  BALL, MARGARET  H
DISAPPEARING DWARF, THE  BLAYLOCK, JAMES P.  C/P,P
DISAPPEARING NIGHTLY [ESTHER DIAMOND 1]  RESNICK, LAURA  LP
DISASTER [@MISSION EARTH #8]  HUBBARD, L. RON  C/P,P,H
DISASTER AREA, THE  BALLARD, J. G.  SR-P
DISCIPLES OF CTHULHU, THE  BERGLUND, EDWARD P.  C/PA,PA
DISCIPLES OF DREAD  CAVE, HUGH B.  P
DISCLOSURE  CRICHTON, MICHAEL  H:2
DISCLOSURES IN SCARLET  JACOBI, CARL  SR-H
DISCOVERY OF KING ARTHUR, THE  ASHE, GEOFFREY  SFWA
DISCWORLD COMPANION, THE [DISCWORLD]  PRATCHETT, TERRY/BRIGGS, STEPHEN  R/REF-P

DISINHERITED, THE [STAR TREK] DAVID, PETER/FRIEDMAN, MICHAEL JAN/

GREENBERGER, ROBERT

C/P:2,P

DISINHERITED, THE [DISINHERITED 1]  WHITE, STEVE  P,SFWA-P
DISPOSABLE PEOPLE  GOLDBERG, MARSHALL/KAY, KENNETH  P
DISPOSSESSED, THE  LE GUIN, URSULA K.  C/P,P,H,SFWA
DISSENSION [@MAGIC: THE GATHERING RAVNICA 3]  HERNDON, CORY J.  P
DISSOLUTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #1]  BYERS, RICHARD LEE  P
DISTANCE HAZE  NASIR, JAMIL  P
DISTANT FRIENDS AND OTHERS  ZAHN, TIMOTHY  C/P,SFWA-P,P
DISTANT PLANES [MAGIC: THE GATHERING]  ICE, KATHY  HA,PA
DISTANT SHORES [STAR TREK VOYAGER]  PALMIERI, MARCO  LPA
DISTANT STARS  DELANY, SAMUEL R.  H,LP
DISTANT SUNS, THE  MOORCOCK, MICHAEL/JAMES, PHILIP  H,SR-LP
DISTANT WORLDS  COLLINS, PAUL  SR-HA
DISTRACTION  STERLING, BRUCE  H,P
DISTRESS  EGAN, GREG  C/P,P
DISUNITED STATES OF AMERICA, THE [@CROSSTIME TRAFFIC #4]  TURTLEDOVE, HARRY  H
DIUTURNITY’S DAWN [HUMANX FOUNDING COMMONWEALTH 3]  FOSTER, ALAN DEAN  H
DIVERGENCE [RECURSION 3]  BALLANTYNE, TONY  P
DIVERGENCE [HERITAGE UNIVERSE 2]  SHEFFIELD, CHARLES  H,P
DIVERSITY ALLIANCE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
DIVIDE, THE  OVERGARD, WILLIAM  P
DIVIDE, THE  WILSON, ROBERT CHARLES  C/P,LP
DIVIDE AND RULE  DE CAMP, L. SPRAGUE  SR-H,C/P:2,P
DIVIDED, THE  WAITMAN, KATIE  LP
DIVIDED ALLEGIANCE [PAKSENARRION 2]  MOON, ELIZABETH  C/P:3,SFWA-P,P
DIVIDED CROWN, THE  GLASS, ISABEL  H
DIVIDED IN DEATH [EVE DALLAS]  ROBB, J. D.  H
DIVIDED LOYALTIES [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  P
DIVINE ENCOUNTERS  SITCHIN, ZECHARIA  R/REF-P
DIVINE ENDURANCE [DIVINE ENDURANCE 1]  JONES, GWYNETH  H,SFWA-P,P
DIVINE HAMMER [@DRAGONLANCE KINGPRIEST #2]  PIERSON, CHRIS  C/P,P
DIVINE INTERVENTION  WHARTON, KEN  P
DIVINE INTERVENTION (CD)  ECKLAR, JULIA  SR-MM
DIVINE INVASION, THE  DICK, PHILIP K.  C/P,P,H
DIVINE QUEEN, THE [DOOM-QUEST OF ARA-KARN 2]  CORBY, ADAM  P
DIVINE RIGHT [@MEROVINGEN NIGHTS #5]  CHERRYH, C. J.  PA
DIVING DAMES AFFAIR, THE [@MAN FROM UNCLE #9]  LESLIE, PETER  C/P:2,P
DIVISION OF THE SPOILS [STARCRUISER SHENANDOAH 2]  GREEN, ROLAND J.  P
DJINN, THE  MASTERTON, GRAHAM  SFWA-P
DNA DISASTER, THE [@TOM SWIFT (ARCHWAY) #4]  APPLETON, VICTOR  P
DO ANDROIDS DREAM OF ELECTRIC SHEEP?  DICK, PHILIP K.  C/P:3,P
DO COMETS DREAM? [@STAR TREK TNG]  SOMTOW, S. P.  P
DOADORES DE MUNDOS  BRUNNER, JOHN  FP
DOC – FIRST GALACTIC ROAMER  LUCCHETTE, STEPHEN C.  R/REF-H
DOC SAVAGE FILES, THE [DOC SAVAGE]  DENT, LESTER  R/REF-P

DOC SAVAGE: HIS APOCALYPTIC LIFE FARMER, PHILIP JOSE SFWA,R/REF-P,

C/REF-H:3,C/REF-P

DOC SAVAGE OMNIBUS 1 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 3 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 4 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 5 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 6 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 7 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 8 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 9 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 10 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 11 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE OMNIBUS 12 [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
DOC SAVAGE OMNIBUS 13 [DOC SAVAGE]  ROBESON, KENNETH  P
DOC SAVAGE: REFLECTIONS IN BRONZE  MURRAY, WILL  R/REF-P
DOC SAVAGE, SUPREME ADVENTURER  NANOVIC, JOHN L.  B
DOC SIDHE [DOC SIDHE 1]  ALLSTON, AARON  P

TITLE  AUTHOR  CODE
DR. ADDER  JETER, K. W.  H,SFWA,P
DOCTOR AND THE ENTERPRISE, THE [DOCTOR WHO/STAR TREK]  AIREY, JEAN  SR-LP
DR. BLOODMONEY, OR HOW WE GOT ALONG AFTER THE BOMB  DICK, PHILIP K.  C/P:2,P
DR. CYCLOPS  GARTH, WILL  C/P,P
DR. CYCLOPS  POPULAR LIBRARY  PA
DR. DIMENSION [DOCTOR DIMENSION]  DE CHANCIE, JOHN/BISCHOFF, DAVID F.  C/P:2,P
DR. DIMENSION: MASTERS OF SPACETIME [DOCTOR DIMENSION]  DE CHANCIE, JOHN/BISCHOFF, DAVID F.  P
DOCTOR DOLITTLE’S CIRCUS [DOCTOR DOLITTLE]  LOFTING, HUGH  H,P
DOCTOR DOLITTLE’S GARDEN [DOCTOR DOLITTLE]  LOFTING, HUGH  H
DOCTOR DOLITTLE’S POST OFFICE [DOCTOR DOLITTLE]  LOFTING, HUGH  H,P
DOCTOR DOLITTLE’S RETURN [DOCTOR DOLITTLE]  LOFTING, HUGH  H
DOCTOR DOLITTLE’S ZOO [DOCTOR DOLITTLE]  LOFTING, HUGH  P
DOCTOR FAUSTUS  MARLOWE, CHRISTOPHER  LP
DOCTOR FOGG  MATSON, NORMAN  H
DR. FUTURITY  DICK, PHILIP K.  DD421,D15697
DOCTOR JEKYLL AND MR. HYDE  STEVENSON, ROBERT LOUIS  C/P:2,P
DOCTOR JEKYLL AND MR. HYDE AND OTHER STORIES (POCKET)  STEVENSON, ROBERT LOUIS  P
DOCTOR JEKYLL AND MR. HYDE AND OTHER STORIES (SCHOLASTIC)  STEVENSON, ROBERT LOUIS  P
DR. KRASINSKI’S SECRET  SHIEL, M. P.  H
DOCTOR MIRABILIS  BLISH, JAMES  SR-H,P
DR. NIKOLA RETURNS  BOOTHBY, GUY  LP
DOCTOR NO [JAMES BOND 6]  FLEMING, IAN  C/P:2,P
DOCTOR ORIENT  LAURIA, FRANK  P
DOCTOR ORPHEUS  WALLACE, IAN  H:2,P
DOCTOR OX AND OTHER STORIES  VERNE, JULES  H
DR. PHIBES  GOLDSTEIN, WILLIAM  C/P,P
DR. PHIBES RISES AGAIN!  GOLDSTEIN, WILLIAM  P
DR. QUAKE [DESTROYER 5]  SAPIR, RICHARD BEN/MURPHY, WARREN  P
DOCTOR RAT  KOTZWINKLE, WILLIAM  H,P
DR. SATAN  WEINBERG, ROBERT  SR-LPA
DR. SCOFFLAW  GOULART, RON  BS3
DOCTOR STRANGE: MASTER OF THE MYSTIC ARTS [MARVEL COMICS]  LEE, STAN  CX

DR. STRANGELOVE OR:

HOW I LEARNED TO STOP WORRYING AND LOVE THE BOMB GEORGE, PETER

C/P,P

DOCTOR TIME [@AVENGER #28]  ROBESON, KENNETH  C/P,P
DOCTOR TO THE STARS [MED]  LEINSTER, MURRAY  C/P:2,P
DOCTOR WHO [DOCTOR WHO]  RUSSELL, GARY  C/P,P
DOCTOR WHO 1985 SUMMER SPECIAL CLASSIC [DOCTOR WHO]  MARVEL COMICS  CX
DOCTOR WHO: A CELEBRATION  HAINING, PETER  R/REF-H,C/REF-H:2
DOCTOR WHO AND AN UNEARTHLY CHILD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE ABOMINABLE SNOWMEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE ANDROID INVASION [DOCTOR WHO]  DICKS, TERRANCE  C/P,P
DOCTOR WHO AND THE ANDROIDS OF TARA [DOCTOR WHO]  DICKS, TERRANCE  SR-P,P
DOCTOR WHO AND THE ARK IN SPACE [DOCTOR WHO]  MARTER, IAN  SR-P
DOCTOR WHO AND THE ARMAGEDDON FACTOR [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE AUTON INVASION [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE BRAIN OF MORBIUS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,H,P
DOCTOR WHO AND THE CARNIVAL OF MONSTERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE CAVE-MONSTERS [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE CLAWS OF AXOS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE CREATURE FROM THE PIT [DOCTOR WHO]  FISHER, DAVID  SR-P,C/P
DOCTOR WHO AND THE CRUSADERS [DOCTOR WHO]  WHITAKER, DAVID  SR-P,C/P:2
DOCTOR WHO AND THE CURSE OF PELADON [DOCTOR WHO]  HAYLES, BRIAN  SR-P,C/P:2
DOCTOR WHO AND THE CYBERMEN [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE DAEMONS [DOCTOR WHO]  LETTS, BARRY  SR-P,C/P:2
DOCTOR WHO AND THE DALEK INVASION OF EARTH [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:3
DOCTOR WHO AND THE DALEKS [DOCTOR WHO]  WHITAKER, DAVID  SR-P,P
DOCTOR WHO AND THE DAY OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE DEADLY ASSASSIN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE DESTINY OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE DINOSAUR INVASION [DOCTOR WHO]  HULKE, MALCOLM  P
DOCTOR WHO AND THE DOOMSDAY WEAPON [DOCTOR WHO]  HULKE, MALCOLM  C/P:2,P
DOCTOR WHO AND THE ENEMY OF THE WORLD [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
DOCTOR WHO AND THE FACE OF EVIL [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE GENESIS OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE GIANT ROBOT [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE GREEN DEATH [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE HAND OF FEAR [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE HORNS OF NIMON [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE HORROR OF FANG ROCK [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE ICE WARRIORS [DOCTOR WHO]  HAYLES, BRIAN  SR-P,C/P
DOCTOR WHO AND THE IMAGE OF THE FENDAHL [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE INVASION OF TIME [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE INVISIBLE ENEMY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2

TITLE  AUTHOR  CODE
DOCTOR WHO AND THE KEEPER OF TRAKEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE KEYS OF MARINUS [DOCTOR WHO]  HINCHCLIFFE, PHILIP  SR-P,C/P:2
DOCTOR WHO AND THE LEISURE HIVE [DOCTOR WHO]  FISHER, DAVID  SR-P
DOCTOR WHO AND THE LOCH NESS MONSTER [DOCTOR WHO]  DICKS, TERRANCE  C/P,P
DOCTOR WHO AND THE MASQUE OF MANDRAGORA [DOCTOR WHO]  HINCHCLIFFE, PHILIP  C/P,P
DOCTOR WHO AND THE MONSTER OF PELADON [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE MUTANTS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE NIGHTMARE OF EDEN [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PLANET OF EVIL [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PLANET OF THE DALEKS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE PLANET OF THE SPIDERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE POWER OF KROLL [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE PYRAMIDS OF MARS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE REVENGE OF THE CYBERMEN [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE RIBOS OPERATION [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
DOCTOR WHO AND THE ROBOTS OF DEATH [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE SEA-DEVILS [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE SEEDS OF DOOM [DOCTOR WHO]  HINCHCLIFFE, PHILIP  C/P:2,P
DOCTOR WHO AND THE SONTARAN EXPERIMENT [DOCTOR WHO]  MARTER, IAN  SR-P
DOCTOR WHO AND THE SPACE WAR [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P
DOCTOR WHO AND THE STATE OF DECAY [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE STONES OF BLOOD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE SUNMAKERS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
DOCTOR WHO AND THE TALONS OF WENG-CHIANG [DOCTOR WHO]  DICKS, TERRANCE  C/P:2,P
DOCTOR WHO AND THE TENTH PLANET [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE TERROR OF THE AUTONS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE THREE DOCTORS [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
DOCTOR WHO AND THE TIME MONSTER, THE [DOCTOR WHO]  DICKS, TERRANCE  P
DOCTOR WHO AND THE TIME WARRIOR [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE TOMB OF THE CYBERMEN [DOCTOR WHO]  DAVIS, GERRY  SR-P,C/P:2
DOCTOR WHO AND THE UNDERWORLD [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE VISITATION [DOCTOR WHO]  SAWARD, ERIC  SR-P
DOCTOR WHO AND THE WAR GAMES [DOCTOR WHO]  HULKE, MALCOLM  SR-P,C/P:2
DOCTOR WHO AND THE WEB OF FEAR [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
DOCTOR WHO AND THE ZARBI [DOCTOR WHO]  STRUTTON, BILL  SR-P,C/P:2
DOCTOR WHO AND WARRIORS’ GATE [DOCTOR WHO]  LYDECKER, JOHN  SR-P,C/P:2
DOCTOR WHO ANNUAL 1985 [DOCTOR WHO]  BBC TV  HA
DOCTOR WHO ANNUAL 1986 [DOCTOR WHO]  BBC TV  HA

DOCTOR WHO ANNUAL (YEAR UNKNOWN, COPYRIGHT 1981)

[DOCTOR WHO] BBC TV

HA

DOCTOR WHO: ATTACK OF THE CYBERMEN [DOCTOR WHO]  SAWARD, ERIC  P
DOCTOR WHO CROSSWORD BOOK, THE [DOCTOR WHO]  ROBINSON, NIGEL  [C/RAN-P]
DOCTOR WHO CYBERMEN [DOCTOR WHO]  BANKS, DAVID  R/REF-H

DOCTOR WHO IN AN EXCITING ADVENTURE WITH THE DALEKS [DOCTOR WHO] WHITAKER, DAVID

C/P:2,P

DOCTOR WHO: JOURNEY THROUGH TIME [DOCTOR WHO]  BBC TV  H

DOCTOR WHO PROGRAMME GUIDE 1: THE PROGRAMMES, THE [DOCTOR WHO] LOFFICIER, JEAN MARC

R/REF-P,C/REF-P:2

DOCTOR WHO PROGRAMME GUIDE 2:

WHAT’S WHAT AND WHO’S WHO, THE [DOCTOR WHO] LOFFICIER, JEAN MARC

R/REF-P,C/REF-P:2

DOCTOR WHO PROGRAMME GUIDE, THE (2ND EDITION) [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P,C/REF-P

DOCTOR WHO TECHNICAL MANUAL, THE (OFFICIAL ANNIVERSARY VOLUME) [DOCTOR WHO] HARRIS, MARK

R/REF-H,C/REF-H

DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME TWO [DOCTOR WHO]  BIG FINISH  HA
DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME THREE [DOCTOR WHO]  BIG FINISH  HA
DOCTOR WHO: THE CHASE [DOCTOR WHO]  PEEL, JOHN  P

DOCTOR WHO: THE DALEK’S MASTERPLAN - PART I:

MISSION TO THE UNKNOWN [DOCTOR WHO] PEEL, JOHN

P

DOCTOR WHO: THE INSIDE STORY [DOCTOR WHO]  RUSSELL, GARY  R/REF-H
DOCTOR WHO: THE KEY TO TIME [DOCTOR WHO]  HAINING, PETER  R/REF-H
DOCTOR WHO: THE LEGEND CONTINUES [DOCTOR WHO]  RICHARDS, JUSTIN  R/REF-H
DOCTOR WHO: THE SCRIPT OF THE FILM [DOCTOR WHO]  JACOBS, MATTHEW  LP
DOCTOR WHO: THE UNFOLDING TEXT [DOCTOR WHO]  TULLOCH, JOHN/ALVARADO, MANUEL  R/REF-P,C/REF-H
DOCTOR WHO: THE WAR MACHINES [DOCTOR WHO]  BLACK, IAN STUART  P
DOCTOR WHO: THE WHEEL IN SPACE [DOCTOR WHO]  DICKS, TERRANCE  P
DOCTOR WHO YEARBOOK (1991) [DOCTOR WHO]  MARVEL COMICS  HA
DOCTOR’S ORDERS [STAR TREK]  DUANE, DIANE  C/P:3,P

DOCUMENTS RELATING TO THE SENTIMENTAL AGENTS IN THE VOLYEN EMPIRE [CANOPUS IN ARGOS 5] LESSING, DORIS

H,LP

DODD, MEAD GALLERY OF HORROR, THE  GRANT, CHARLES L.  HA
DOES ANYONE ELSE HAVE SOMETHING FURTHER TO ADD?  LAFFERTY, R. A.  SFWA
DOG AND THE WOLF, THE [@KING OF YS 4]  ANDERSON, POUL/ANDERSON, KAREN  P,SFWA-P
DOG CALLED DEMOLITION, A  RANKIN, ROBERT  P
DOG DAY AFTERNOON [@SABRINA THE TEENAGE WITCH SALEM #5]  GALLAGHER, DIANA G.  P

The MITSFS Titledex  September 24, 2007  Page 95
TITLE  AUTHOR  CODE
DOG DAYS OF ARTHUR CANE, THE  BETHANCOURT, T. ERNESTO  P
DOG EAT DOG [INGERSOL 2]  CARROLL, JERRY JAY  H,LP
DOG FROM HELL, THE [@STAR RISK,LTD. 4]  BUNCH, CHRIS  P
DOG IN THE SKY  CORWIN, NORMAN  H
DOG WARRIOR [UKIAH OREGON 4]  SPENCER, WEN  C/P:2,P
DOG WIZARD [WINDROSE 3]  HAMBLY, BARBARA  C/P,P,H
DOGE’S RING, THE  VARE, DANIELE  H
DOGLAND  SHETTERLY, WILL  H
DOG’S HEAD, A  DUTOURD, JEAN  SR-P
DOG’S LIFE, A [@SABRINA THE TEENAGE WITCH #9]  DUBOWSKI, CATHY EAST  P
DOGS OF TRUTH  REED, KIT  LP
DOGS OF WAR  DRAKE, DAVID  PA
DOGSBODY  JONES, DIANA WYNNE  P
DOGTALES!  DANN, JACK/DOZOIS, GARDNER  PA

DOGTOWN TOURIST AGENCY, THE [@VANCE INTEGRAL EDITION #32]/ FREITZKE’S TURN [@VANCE INTEGRAL EDITION #32] VANCE, JACK

SR-H

DOLFIJNENEILAND  CLARKE, ARTHUR C.  FP
DOLL, THE [DARK FORCES 3]  SPARGER, REX  P
DOLL AND ONE OTHER, THE  BLACKWOOD, ALGERNON  H
DOLL MAKER, THE  SARBAN  P
DOLL WHO ATE HIS MOTHER, THE  CAMPBELL, RAMSEY  H,P
DOLL WITH OPAL EYES, THE  DEWEESE, JEAN  P
DOLLARVILLE  DAVIES, PETE  H
DOLL’S HOUSE, THE [@SANDMAN #2]  GAIMAN, NEIL  H:2,CX
DOLPHIN AND THE DEEP, THE  SWANN, THOMAS BURNETT  C/P:2,P
DOLPHIN BOY [DOLPHIN 1]  MEYERS, ROY  C/P:2,P
DOLPHIN ISLAND  CLARKE, ARTHUR C.  C/P:2,P
DOLPHIN’S BELL, THE [PERN]  MCCAFFREY, ANNE  H:2
DOLPHINS OF ALTAIR, THE  SAINT CLAIR, MARGARET  C/P:2,P
DOLPHINS OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
DOMAIN  ALTEN, STEVE  H
DOMAIN  HERBERT, JAMES  P
DOMAINS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
DOMAINS OF KORYPHON, THE [@VANCE INTEGRAL EDITION #28]  VANCE, JACK  SR-H
DOME  REAVES, MICHAEL/PERRY, STEVE  P,SFWA-P
DOME AROUND AMERICA  WILLIAMSON, JACK  DD118
DOME IN THE FOREST, THE [PELBAR 3]  WILLIAMS, PAUL O.  C/P:3,P
DOME WORLD  MCLAUGHLIN, DEAN  C/P:2,P
DOMES OF FIRE [TAMULI 1]  EDDINGS, DAVID  H:2,LP
DOMINANT SPECIES  WARREN, GEORGE  H,LP
DOMINATION  CECILIONE, MICHAEL  P
DOMINATION, THE [DRAKA (REVISED) 1,2,3]  STIRLING, S. M.  H
DOMINATORS, THE [DOCTOR WHO]  MARTER, IAN  SR-P,C/P:2
DOMINION [DRACULA]  SABERHAGEN, FRED  C/P:2,P
DOMINION [WARHAMMER]  SAVILE, STEVEN  P
DOMINION [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  C/P,P
DOMINION’S REACH [UNIFIED WORLDS 3]  THORNLEY, DIANN  H,P
DOMINO EFFECT, THE [DOCTOR WHO EIGHTH DOCTOR]  BISHOP, DAVID  P
DOMNEI  CABELL, JAMES BRANCH  C/P:2,P
DON OF THE DEAD [PEPPER MARTIN]  DANIELS, CASEY  P
DON RODRIGUEZ: CHRONICLES OF SHADOW VALLEY  DUNSANY, LORD  P
DONA FLOR AND HER TWO HUSBANDS  AMADO, JORGE  P
DONNERJACK  ZELAZNY, ROGER/LINDSKOLD, JANE  H:2,P
DONOVAN’S BRAIN  SIODMAK, CURT  C/P:2,P
DON’T BET ON THE PRINCE  ZIPES, JACK  LP
DON’T BITE THE SUN  LEE, TANITH  P
DON’T FORGET YOUR SPACESUIT, DEAR  NYE, JODY LYNN  C/PA,PA
DON’T OPEN THIS BOOK!  KAYE, MARVIN  HA
DON’T PANIC [HITCHHIKERS]  GAIMAN, NEIL  R/REF-P
DOOM BRIGADE, THE [@DRAGONLANCE CHAOS WAR 1]  WEIS, MARGARET/PERRIN, DON  P
DOOM CITY [GREYSTONE BAY 2]  GRANT, CHARLES L.  PA,SFWA-P
DOOM COMMANDER [@C.A.D.S. #7]  SIEVERT, JOHN  P
DOOM OF THE DARKSWORD [@DARKSWORD #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DOOM OF THE GREEN PLANET [GREEN PLANET 2]  PETAJA, EMIL  DH070
DOOM OF THREE PLANETS  POHLE, ROBERT  P
DOOM STAR [DOOM STAR 1]  MEYERS, RICHARD S.  SR-P,C/P:2
DOOM STAR NUMBER TWO [DOOM STAR 2]  MEYERS, RICHARD S.  SR-P
DOOM STONE, THE  WOOLRICH, CORNELL  P
DOOM THAT CAME TO SARNATH, THE  LOVECRAFT, H. P.  C/P:2,P
DOOM WIND (PROOF)  GUNTHER, MAX  SR-LP
DOOMED NATION OF THE SKIES  FUTURE, STEVE  SR-P
DOOMED PLANET, THE [@MISSION EARTH #10]  HUBBARD, L. RON  H
DOOMED PLANET  SHELDON, LEE  H
DOOMFARERS OF CORAMONDE, THE [CORAMONDE 1]  DALEY, BRIAN  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 96
TITLE  AUTHOR  CODE
DOOM’S BREAK [@ARNA #3]  ROWLEY, CHRISTOPHER  P
DOOMSAYER, THE [@SURVIVALIST #4]  AHERN, JERRY  P
DOOMSDAY [@ROBOTECH #6]  MCKINNEY, JACK  P
DOOMSDAY [FANTASTIC FOUR]  WOLFMAN, MARV  C/P,P
DOOMSDAY, 1999  MACTYRE, PAUL  C/P:2,P
DOOMSDAY AFFAIR, THE [@MAN FROM UNCLE #2]  WHITTINGTON, HARRY  C/P:2,P
DOOMSDAY BOOK [DUNWORTHY & FINCH 1]  WILLIS, CONNIE  C/P,LP,H:2
DOOMSDAY BRUNETTE, THE [ZACHARY NIXON JOHNSON 2]  ZAKOUR, JOHN/GANEM, LAWRENCE  C/P,P
DOOMSDAY DECK [BUFFY]  GALLAGHER, DIANA G.  P
DOOMSDAY EFFECT, THE  WREN, THOMAS  P
DOOMSDAY EVE  WILLIAMS, ROBERT MOORE  DD215
DOOMSDAY EXAM [BUREAU 13 2]  POLLOTTA, NICK  C/P,P
DOOMSDAY EXHIBIT, THE  FAIRMAN, PAUL W.  C/P,P
DOOMSDAY GENE, THE  BOYD, JOHN  H
DOOMSDAY MEN, THE  BULMER, KENNETH  H:2,P
DOOMSDAY MEN, THE  PRIESTLEY, J. B.  H,P
DOOMSDAY MORNING  MOORE, C. L.  H:2,P
DOOMSDAY ON AJIAT [PROFESSOR JAMESON 5]  JONES, NEIL R.  C/P:2,P
DOOMSDAY PLANET, THE  VINCENT, HARL  C/P:2,P
DOOMSDAY PLUS TWELVE  FORMAN, JAMES D.  H
DOOMSDAY SPIRAL, THE  LAND, JOHN  P
DOOMSDAY WARRIOR [@DOOMSDAY WARRIOR #1]  STACY, RYDER  P
DOOMSDAY WING  SMITH, GEORGE H.  C/P:2,P

DOOMSDAY WORLD [STAR TREK TNG] CARTER, CARMEN/

DAVID, PETER/FRIEDMAN, MICHAEL JAN/ GREENBERGER, ROBERT

C/P:2,P

DOOMSMAN  ELLISON, HARLAN  SR-DB0779,SR-DB0244
DOOMSTALKER [@DARKWAR #1]  COOK, GLEN  P
DOOMSTAR  HAMILTON, EDMOND  C/P:2,P
DOOMTIME  PISERCHIA, DORIS  P
DOOON MODE [MODE 4]  ANTHONY, PIERS  H
DOOR IN THE HEDGE, THE  MCKINLEY, ROBIN  C/P:2,P
DOOR INTO FIRE, THE [TALE OF THE FIVE 1]  DUANE, DIANE  P
DOOR INTO FIRE, THE (REVISED) [TALE OF THE FIVE 1]  DUANE, DIANE  C/P,SFWA-P,LP
DOOR INTO OCEAN, A [ELYSIUM]  SLONCZEWSKI, JOAN  H:2,P,SFWA
DOOR INTO SHADOW, THE [TALE OF THE FIVE 2]  DUANE, DIANE  H:2,SFWA-P,LP
DOOR INTO SUMMER, THE  HEINLEIN, ROBERT A.  C/P,P,H
DOOR INTO SUNSET, THE [TALE OF THE FIVE 3]  DUANE, DIANE  SR-P,H
DOOR NUMBER THREE  O’LEARY, PATRICK  H,LP
DOOR THROUGH SPACE, THE  BRADLEY, MARION ZIMMER  C/P:2,DF117,P
DOOR THROUGH WASHINGTON SQUARE, THE  BERGSTROM, ELAINE  P
DOOR TO ALTERNITY [@BUFFY ANGEL UNSEEN 2]  HOLDER, NANCY/MARIOTTE, JEFF  C/P,P
DOOR TO AMBERMERE, THE [AMBERMERE 1]  PIERCE, J. CALVIN  C/P,P
DOOR TO ANYWHERE  POHL, FREDERIK  PA
DOOR TO DECEMBER, THE  KOONTZ, DEAN R.  P
DOOR TO SATURN, THE  SMITH, CLARK ASHTON  H
DOORS INTO CHAOS [STAR TREK GATEWAYS #3/STAR TREK TNG]  GREENBERGER, ROBERT  P
DOORS OF DEATH AND LIFE [PEARL OF IMMORTALITY 2]  CLOUGH, BRENDA W.  H

DOORS OF HIS FACE, THE LAMPS OF HIS MOUTH, AND OTHER STORIES, THE ZELAZNY, ROGER

C/P,SFWA,P,H

DOORS OF THE UNIVERSE, THE [NOREN 3]  ENGDAHL, SYLVIA LOUISE  H
DOORWAY INTO TIME  MOSKOWITZ, SAM  PA,C/PA
DOORWAY TO ETERNITY  WILLIAMS, SEAN  SR-LP
DOORWAYS IN THE SAND  ZELAZNY, ROGER  C/P,P,H,SFWA
DOORWAYS TO SPACE  WELLS, BASIL  SR-LP
DOPPELGANGER [MIREI 1]  BRENNAN, MARIE  P
DOPPELGANGER GAMBIT, THE [JANNA BRILL 1]  KILLOUGH, LEE  C/P,P
DOPPELGANGERS  HEARD, H. F.  C/P:2,P
DORELLA  GARLAND, MARK A./MCCRAW, CHARLES G. C/P:2,SFWA-P,P
DORK COVENANT  KOVALIC, JOHN  CX
DOROTHEA DREAMS  CHARNAS, SUZY MCKEE  P,SFWA,SFWA-P
DOROTHY AND THE WIZARD IN OZ [OZ #4]  BAUM, L. FRANK  H,@P,@C/P
DOROTHY, THE TERRIFIED [@PHENWICK #19]  KIMBROUGH, KATHERYN  P
DORSAI! [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
DORSAI COMPANION, THE [CHILDE CYCLE]  DICKSON, GORDON R.  H,LP

DORSAI’S COMMAND [@COMBAT COMMAND CHILDE CYCLE] DICKSON, GORDON R./DENNING, TROY/

GLABERSON, CORY

P

DOSADI EXPERIMENT, THE  HERBERT, FRANK  C/P,P,H
DOUBLE-BLIND [@BATTLETECH 31]  COLEMAN, LOREN L.  P
DOUBLE BLIND [KURTIS:P385XL47:RUTH 3]  STAMEY, SARA  P
DOUBLE CONTACT [SECTOR GENERAL 12]  WHITE, JAMES  C/P,P
DOUBLE DISAPPEARANCE OF WALTER FOZBEK, THE  SENN, STEVE  LP
DOUBLE, DOUBLE  BRUNNER, JOHN  C/P:2,P
DOUBLE, DOUBLE [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P

TITLE  AUTHOR  CODE
DOUBLE EAGLE [@WARHAMMER 40K]  ABNETT, DAN  P
DOUBLE EXPOSURE [APPRENTICE ADEPT 1,2,3]  ANTHONY, PIERS  H:2
DOUBLE EXPOSURE  FLEISCHMAN, THEO  H,P
DOUBLE FEATURE  BULL, EMMA/SHETTERLY, WILL  SR-H,H
DOUBLE FULL MOON NIGHT [RAMA]  LEE, GENTRY  H,P
DOUBLE HELIX FALL  FERGUSON, NEIL  LP
DOUBLE IDENTITY  GALLUN, RAYMOND Z.  B
DOUBLE IN SPACE  PRATT, FLETCHER  H:2
DOUBLE INVADERS, THE  RACKHAM, JOHN  DG623
DOUBLE JEOPARDY [CAR WARRIORS 2]  ALLSTON, AARON  P
DOUBLE JEOPARDY  PRATT, FLETCHER  C/P,H,GN5.30
DOUBLE MAN, THE  BINDER, EANDO  C/P:2,P
DOUBLE MEANING  KNIGHT, DAMON  TD34
DOUBLE MOBIUS SPHERE  NIM, P. S.  C/P:2,P
DOUBLE NOCTURNE  FELICE, CYNTHIA  H:2,P
DOUBLE OR NOTHING [@STAR TREK TNG DOUBLE HELIX 5]  DAVID, PETER  P
DOUBLE OR NOTHING [@QUANTUM LEAP]  HENDERSON, C. J.  P
DOUBLE PHOENIX  BALLANTINE BOOKS  PA,C/PA
DOUBLE PLANET  GRIBBIN, JOHN/CHOWN, MARCUS  C/P,P
DOUBLE SHADOW, THE  SMITH, CLARK ASHTON  LP
DOUBLE SHADOW, A  TURNER, FREDERICK  H,P
DOUBLE SOLITAIRE [@WILD CARDS #10]  SNODGRASS, MELINDA M.  C/P:2,P
DOUBLE STAR  HEINLEIN, ROBERT A.  C/P:2,P
DOUBLE STAR (CHINESE)  HEINLEIN, ROBERT A.  FP
DOUBLECROSS PROGRAM, THE [STAR RISK,LTD. 3]  BUNCH, CHRIS  P
DOUBLES, THE  MIKHANOVSKY, VLADIMIR  P
DOUGLAS ADAMS’S STARSHIP TITANIC  JONES, TERRY  LP
DOUGLAS CONVOLUTION, THE  LLEWELLYN, EDWARD  P
DOVE ISABEAU  YOLEN, JANE  H
DOVE LOOKED IN, THE [@WATER! #2]  BAUDINO, GAEL  P
DOVER BEACH  BOWKER, RICHARD  P
DOWN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  MILES, LAWRENCE  C/P,P
DOWN AMONG THE DEAD MEN  GREEN, SIMON R.  C/P,P
DOWN AND DIRTY [@WILD CARDS #5]  MARTIN, GEORGE R. R.  C/PA,PA
DOWN AND OUT IN THE MAGIC KINGDOM  DOCTOROW, CORY  H
DOWN BOUND TRAIN  GARNETT, BILL  H,SFWA,P
DOWN BY THE OLD BLOODSTREAM  HITCHCOCK, ALFRED  C/PA,PA
DOWN HERE IN THE DREAM QUARTER  MALZBERG, BARRY N.  H,SFWA
DOWN IN FLAMES (OUTLINE) [KNOWN SPACE]  NIVEN, LARRY  [C/RAN-H:3]
DOWN IN THE BLACK GANG  FARMER, PHILIP JOSE  H:2,P
DOWN IN THE BOTTOMLANDS (AND OTHER PLACES)  TURTLEDOVE, HARRY/DE CAMP, L. SPRAGUE P
DOWN THE AMAZON  VERNE, JULES  H
DOWN THE BRIGHT WAY  REED, ROBERT  C/P:2,P
DOWN THE STREAM OF STARS [CHANGELING STAR 2]  CARVER, JEFFREY A.  C/P:2,P
DOWN THERE IN DARKNESS  TURNER, GEORGE  LP
DOWN THESE DARK SPACEWAYS  RESNICK, MIKE  HA
DOWN TIME [EMMA MERRIGAN 4]  ABBEY, LYNN  P
DOWN TO A SUNLESS SEA [MAN WHO LOVED MARS 4]  CARTER, LIN  C/P,P
DOWN TO A SUNLESS SEA  GRAHAM, DAVID  H:2,P
DOWN TO EARTH  CAPON, PAUL  H
DOWN TO EARTH  CHARBONNEAU, LOUIS  C/P:2,P
DOWN TO EARTH  WRIGHTSON, PATRICIA  P
DOWN TO THE SEA [@LOST REGIMENT 9]  FORSTCHEN, WILLIAM R.  C/P,P
DOWN TOWN  POLIKARPUS, VIIDO/KING, TAPPAN  P,SFWA,SFWA-P
DOWNBELOW STATION [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
DOWNFALL [@DRAGONLANCE DHAMON #1]  RABE, JEAN  H,P
DOWNFALL MATRIX, THE [@PATTERNS OF CHAOS #3]  INGRID, CHARLES  P
DOWNS-LORD DAWN [DOWNS-LORD 1]  WHITBOURN, JOHN  P
DOWNS-LORD DAY [DOWNS-LORD 2]  WHITBOURN, JOHN  P
DOWNSTAIRS ROOM, THE  WILHELM, KATE  C/P:2,P
DOWNTIME  FELICE, CYNTHIA  H:2
DOWNTIME [DOCTOR WHO MISSING ADVENTURES]  PLATT, MARC  C/P:2,P
DOWNTIMING THE NIGHT SIDE  CHALKER, JACK L.  C/P:2,SFWA-P,P
DOWNWARD TO THE EARTH  SILVERBERG, ROBERT  H:2,P
DOZEN BLACK ROSES, A [SONJA BLUE]  COLLINS, NANCY A.  LP
DR. TIM BOOK ONE  VARIAN, CHRISTOPHER  CX
DRACHENFELS [WARHAMMER GENEVIEVE 1]  YEOVIL, JACK  P
DRACO [WARHAMMER INQUISITION WAR 1]  WATSON, IAN  P
DRACO TAVERN, THE  NIVEN, LARRY  H:2
DRACONIAN MEASURES [@DRAGONLANCE KANG’S REGIMENT #2]  PERRIN, DON/WEIS, MARGARET  P
DRACULA  STOKER, BRAM  H:2,P
DRACULA AND THE VIRGINS OF THE UNDEAD  AUBIN, ETIENNE  SR-P
DRACULA ARCHIVES, THE  RUDORFF, RAYMOND  C/P,P
DRACULA: ASYLUM  WITCOVER, PAUL  P

The MITSFS Titledex  September 24, 2007  Page 98
TITLE  AUTHOR  CODE
DRACULA BEGAN  KIMBERLY, GAIL  P
DRACULA BOOK OF GREAT VAMPIRE STORIES, THE  SHEPARD, LESLIE  C/PA,LPA
DRACULA CAPER, THE [TIMEWARS 8]  HAWKE, SIMON  C/P,P
DRACULA IN LONDON  ELROD, P. N.  LPA
DRACULA IN LOVE  SHIRLEY, JOHN  C/P,P
DRACULA: PRINCE OF DARKNESS  GREENBERG, MARTIN H.  PA
DRACULA RETURNS [@DRACULA #1]  LORY, ROBERT  P
DRACULA TAPE, THE [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
DRACULA: THE VAMPIRE PLAY  DEANE, HAMILTON/BALDERSTON, JOHN L.  H
DRACULA UNBOUND  ALDISS, BRIAN W.  C/P,P
DRACULA’S BROTHERS [@DRACULA #3]  LORY, ROBERT  C/P,P
DRACULA’S DAUGHTER  DREADSTONE, CARL  P
DRACULA’S DISCIPLE [@DRACULA #8]  LORY, ROBERT  C/P,P
DRACULA’S GOLD [@DRACULA #4]  LORY, ROBERT  P
DRAG QUEEN OF ELFLAND  SCHIMEL, LAWRENCE  P
DRAGON [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,P,H
DRAGON, THE [ATLAN]  GASKELL, JANE  C/P:2,P
DRAGON [@SMALLVILLE]  GRANT, ALAN  P
DRAGON, THE [FIVE RINGS CLAN WAR 6]  SOESBEE, REE  P
DRAGON AMERICA  RESNICK, MIKE  LP
DRAGON AND HERDSMAN [DRAGONBACK 4]  ZAHN, TIMOTHY  H
DRAGON AND PHOENIX [DRAGONLORD LINDEN 2]  BERTIN, JOANNE  H:2
DRAGON AND PHOENIX (PROOF) [DRAGONLORD LINDEN 2]  BERTIN, JOANNE  SR-LP
DRAGON AND THE DJINN, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  C/P,P
DRAGON AND THE FAIR MAID OF KENT, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  H
DRAGON AND THE GEORGE, THE [DRAGON KNIGHT 1]  DICKSON, GORDON R.  C/P,P,H:2
DRAGON AND THE GNARLY KING, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  C/P,P,H
DRAGON AND THE UNICORN, THE  ATTANASIO, A. A.  C/P:2,LP
DRAGON AND THIEF [DRAGONBACK 1]  ZAHN, TIMOTHY  H
DRAGON AT WAR, THE [DRAGON KNIGHT 4]  DICKSON, GORDON R.  H,SFWA,SFWA-P:2,P
DRAGON AT WORLD’S END, A [BAZIL BROKETAIL 5]  ROWLEY, CHRISTOPHER  P
DRAGON AVENGER [@AGE OF FIRE #2]  KNIGHT, E. E.  LP
DRAGON BLADE [OAK,YEW,ASH,AND ROWAN 4]  NORTON, ANDRE/MILLER, SASHA  H
DRAGON BLOOD [WARD OF HUROG 2]  BRIGGS, PATRICIA  C/P,P
DRAGON BONES [WARD OF HUROG 1]  BRIGGS, PATRICIA  C/P,P
DRAGON BURNING [DRAGON CIRCLE 3]  GARDNER, CRAIG SHAW  H,@P
DRAGON CAULDRON [DRAGON OF THE LOST SEA 3]  YEP, LAURENCE  SFWA
DRAGON CHAMPION [@AGE OF FIRE #1]  KNIGHT, E. E.  LP
DRAGON CHARMER, THE [FERN CHAPEL 2]  SIEGEL, JAN  P
DRAGON CHRONICLES, THE  MEIKLEJOHN GRAPHICS  H:2
DRAGON CIRCLE, THE  KRENSKY, STEPHEN  LP
DRAGON CIRCLE, THE [@STARGODS #2]  RADFORD, IRENE  P
DRAGON COMPANION [DRAGON COMPANION 1]  CALLANDER, DON  C/P:2,P

DRAGON DAY [@DRAGONLANCE NEW ADVENTURES #6/ DRAGONLANCE DRAGON QUARTET 2] BROWN, STAN

LP

DRAGON DEATH [DRAGONSWORD 3]  BAUDINO, GAEL  C/P,SFWA-P,P
DRAGON DELASANGRE, THE [DRAGON DELASANGRE 1]  TROOP, ALAN F.  P
DRAGON DELASANGRE, THE (PROOF) [DRAGON DELASANGRE 1]  TROOP, ALAN F.  SR-LP

DRAGON FANTASTIC GREENBERG, MARTIN H./

GREENBERG, ROSALIND M.

C/PA,PA

DRAGON FEAST  ELLIOTT, JOHN  P
DRAGON GUARD, THE [@MAGICKERS #3]  DRAKE, EMILY  P
DRAGON HOARD, THE  LEE, TANITH  P
DRAGON HUNT [@WARCRAFT SUNWELL 1]  KNAAK, RICHARD A.  SCX
DRAGON IN DANGER  MANNING, ROSEMARY  P
DRAGON IN LYONESSE, THE [DRAGON KNIGHT]  DICKSON, GORDON R.  H
DRAGON IN THE SEA, THE  HERBERT, FRANK  H:2,P
DRAGON IN THE SWORD, THE [EREKOSE 3/ETERNAL CHAMPION 3]  MOORCOCK, MICHAEL  C/P,SFWA-P,P,H
DRAGON ISLES, THE [@DRAGONLANCE CROSSROADS]  SULLIVAN, STEPHEN D.  P
DRAGON KEEPERS, THE  HUGHES, RODNEY  P
DRAGON KING, THE [CRIMSON SHADOW 3]  SALVATORE, R. A.  P
DRAGON KNIGHT, THE [DRAGON KNIGHT 2]  DICKSON, GORDON R.  H:2,SFWA,P

DRAGON KNIGHT [@DRAGONLANCE NEW ADVENTURES #7/ DRAGONLANCE DRAGON QUARTET 3] WILLIS, DAN

LP

DRAGON LENSMAN, THE [LENSMEN 7]  KYLE, DAVID A.  C/P:2,P
DRAGON LORD, THE  DRAKE, DAVID  C/P,SFWA-P,H,P,SFWA
DRAGON LORD [BOOK OF YEARS 3]  MORWOOD, PETER  P
DRAGON LORD OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 2]  LORRAH, JEAN  C/P,P
DRAGON LORD, THE (PROOF)  DRAKE, DAVID  SR-LP
DRAGON LOVERS  SIGNET ECLIPSE  LPA
DRAGON-LOVER’S TREASURY OF THE FANTASTIC, A  WEIS, MARGARET  C/PA,LPA
DRAGON-LOVER’S TREASURY OF THE FANTASTIC, A (2 CASSETTES)  WEIS, MARGARET  SR-MM
DRAGON MAGIC  NORTON, ANDRE  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 99
TITLE  AUTHOR  CODE

DRAGON MASTERS, THE VANCE, JACK C/P,H,DF185,P,D16640,

SFWA-TD28

DRAGON MOON [DRAGON DELASANGRE 2]  TROOP, ALAN F.  P
DRAGON NEVER SLEEPS, THE  COOK, GLEN  C/P:2,P
DRAGON OF DESPAIR, THE [FIREKEEPER 3]  LINDSKOLD, JANE  P
DRAGON OF MISHBIL, THE [AVERIDAN 4]  CLOUGH, BRENDA W.  C/P:2,P
DRAGON OF THE ISHTAR GATE, THE  DE CAMP, L. SPRAGUE  C/P:2,P
DRAGON ON A PEDESTAL [@XANTH #7]  ANTHONY, PIERS  C/P,P,H
DRAGON ON THE BORDER, THE [DRAGON KNIGHT 3]  DICKSON, GORDON R.  H:2,SFWA-P:2,P
DRAGON PRECINCT  DECANDIDO, KEITH R. A.  P
DRAGON PRINCE [DRAGON PRINCE 1]  RAWN, MELANIE  C/P,P
DRAGON QUEEN, THE [TALES OF GUINEVERE 1]  BORCHARDT, ALICE  H,P
DRAGON QUINTET, THE  KAYE, MARVIN  HA
DRAGON REBORN, THE [WHEEL OF TIME 3]  JORDAN, ROBERT  C/P,P,H:2
DRAGON REBORN, THE (PROOF) [WHEEL OF TIME 3]  JORDAN, ROBERT  SR-LP
DRAGON REFORGED [GONARD 2]  CYMRI, CHRYS  P
DRAGON RESCUE [DRAGON COMPANION 2]  CALLANDER, DON  C/P,P
DRAGON REVENANT, THE [DEVERRY 4/DEVERRY PROPER 4]  KERR, KATHARINE  C/P,LP,H
DRAGON RIDER  FUNKE, CORNELIA  H
DRAGON RIGGER [STAR DRAGONS 2]  CARVER, JEFFREY A.  C/P,P,H
DRAGON RISES, THE  MARTINE-BARNES, ADRIENNE  P
DRAGON RISING [@MECHWARRIOR DARK AGE/BATTLETECH]  BICK, ILSA J.  P
DRAGON SEASON  CASSUTT, MICHAEL  C/P,P
DRAGON SLEEPING [DRAGON CIRCLE 1]  GARDNER, CRAIG SHAW  H:2,SFWA,P
DRAGON SOCIETY, THE [OBSIDIAN CHRONICLES 2]  WATT-EVANS, LAWRENCE  H
DRAGON SORCERER, THE  BETANCOURT, JOHN GREGORY  P

DRAGON SPELL [@DRAGONLANCE NEW ADVENTURES #8/ DRAGONLANCE DRAGON QUARTET 4] SAMPSON, JEFF

LP

DRAGON STONES  PALMER, JOSEPH  P
DRAGON SUN [@ALPHA CENTAURI #2]  ELY, MICHAEL  P

DRAGON SWORD [@DRAGONLANCE NEW ADVENTURES #5/ DRAGONLANCE DRAGON QUARTET 1] SOESBEE, REE

LP

DRAGON TALES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

DRAGON TEARS  KOONTZ, DEAN R.  H:2,P
DRAGON TEMPEST  CALLANDER, DON  P
DRAGON, THE EARL, AND THE TROLL, THE [DRAGON KNIGHT 5]  DICKSON, GORDON R.  H,P
DRAGON TOKEN, THE [DRAGON STAR 2]  RAWN, MELANIE  H:2,P
DRAGON TOME [DRAGONREALM ORIGINS 3]  KNAAK, RICHARD A.  P
DRAGON ULTIMATE [BAZIL BROKETAIL 7]  ROWLEY, CHRISTOPHER  P
DRAGON VENOM [OBSIDIAN CHRONICLES 3]  WATT-EVANS, LAWRENCE  H
DRAGON WAITING, THE  FORD, JOHN M.  C/P:2,P,H
DRAGON WAKING [DRAGON CIRCLE 2]  GARDNER, CRAIG SHAW  C/P,P,H
DRAGON WAR [LAND BETWEEN THE RIVERS 2]  ACRES, MARK  P
DRAGON WEATHER [OBSIDIAN CHRONICLES 1]  WATT-EVANS, LAWRENCE  H

DRAGON WELL, THE [@DRAGONLANCE NEW ADVENTURES #3/ DRAGONLANCE SPELLBINDER QUARTET 3] WILLIS, DAN

LP

DRAGON WING [@DEATH GATE #1]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
DRAGON WINTER  HANCOCK, NIEL  C/P:2,P
DRAGONBACK BARGAIN [DRAGONBACK 1,2,3]  ZAHN, TIMOTHY  H
DRAGONBARDS, THE  MURPHY, SHIRLEY ROUSSEAU  P
DRAGONBONE CHAIR, THE [MEMORY,SORROW AND THORN 1]  WILLIAMS, TAD  C/P,H:2,P
DRAGONBOUND  MILLER, CARL  P
DRAGONCHARM [FORTUNE 1]  EDWARDS, GRAHAM  C/P,P
DRAGONDOOM [MITHGAR 5]  MCKIERNAN, DENNIS L.  C/P:2,P
DRAGONDRUMS [PERN HARPER HALL 3]  MCCAFFREY, ANNE  H:2,P
DRAGONFALL 5 AND THE HIJACKERS  EARNSHAW, BRIAN  SR-P
DRAGONFALL 5 AND THE ROYAL BEAST  EARNSHAW, BRIAN  SR-P
DRAGONFALL 5 AND THE SPACE COWBOYS  EARNSHAW, BRIAN  SR-P
DRAGONFIELD AND OTHER STORIES  YOLEN, JANE  C/P,P
DRAGONFIRE  NYE, JODY LYNN  C/P,SFWA-P,P
DRAGONFLIGHT [PERN DRAGONRIDERS 1]  MCCAFFREY, ANNE  C/P:2,P
DRAGONFLY  DURBIN, FREDERIC C.  H
DRAGONFLY IN AMBER [OUTLANDER (GABALDON) 2]  GABALDON, DIANA  C/P,P
DRAGONFRIGATE WIZARD HALCYON BLITHE [HALCYON BLITHE 2]  WARD, JAMES M.  H
DRAGONHARPER  NYE, JODY LYNN  C/P:2,SFWA-P,P
DRAGONHEART  POGUE, CHARLES EDWARD  P
DRAGONHOLDER  MCCAFFREY, TODD  R/REF-H,C/REF-H
DRAGONKIN  BAILEY, ROBIN WAYNE  H
DRAGONKING OF MYSTARA [@MYSTARA DRAGONLORD 2]  GUNNARSSON, THORARINN  P
DRAGONKNIGHT [DRAGONSPELL 3]  PAUL, DONITA K.  LP
DRAGONLANCE AND OTHER WORLDS (1996)  CALENDAR  CAL
DRAGONLANCE: THE ANNOTATED CHRONICLES [DRAGONLANCE 1,2,3]  WEIS, MARGARET/HICKMAN, TRACY  LP

TITLE  AUTHOR  CODE

DRAGONLANCE: THE ANNOTATED LEGENDS [DRAGONLANCE CHRONICLES 1,2,3] WEIS, MARGARET/HICKMAN, TRACY

LP

DRAGONLORD OF MYSTARA [@MYSTARA DRAGONLORD 1]  GUNNARSSON, THORARINN  P
DRAGONLOVER’S GUIDE TO PERN, THE [PERN]  NYE, JODY LYNN/MCCAFFREY, ANNE  R/REF-H
DRAGONMAGE OF MYSTARA [@MYSTARA DRAGONLORD 3]  GUNNARSSON, THORARINN  P
DRAGONNE’S EG [UNEXPECTED DRAGON 3]  BROWN, MARY  P
DRAGONQUEST [PERN DRAGONRIDERS 2]  MCCAFFREY, ANNE  C/P:2,P
DRAGONQUEST [DRAGONSPELL 2]  PAUL, DONITA K.  LP
DRAGONRANK MASTER [BIFROST GUARDIANS 3]  REICHERT, MICKEY ZUCKER  C/P,P
DRAGONRIDERS OF PERN [PERN DRAGONRIDERS 1,2,3]  MCCAFFREY, ANNE  H:2
DRAGONROUGE [TERRA MAGICA 2]  CARTER, LIN  C/P,P
DRAGONS  BLANPIED, PAMELA WHARTON  C/P,P
DRAGONS!  DANN, JACK/DOZOIS, GARDNER  PA
DRAGONS, THE [@DRAGONLANCE LOST HISTORIES #6]  NILES, DOUGLAS  P

DRAGONS & MYSTICS: THE ART OF MICHAEL WHELAN 2005 CALENDAR (2005)

WHELAN, MICHAEL

CAL

DRAGONS AND NIGHTMARES  BLOCH, ROBERT  C/P:2,P
DRAGON’S ASCENSION, A [BAND OF FOUR]  GREENWOOD, ED  C/P,P
DRAGONS AT WAR, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGON’S BAIT  VELDE, VIVIAN VANDE  P
DRAGON’S BLOOD [SCORPIO 4]  MCDONOUGH, ALEX  C/P,P
DRAGON’S BLOOD [PIT DRAGONS 1]  YOLEN, JANE  C/P,P
DRAGON’S BLUFF [@DRAGONLANCE CROSSROADS]  HERBERT, MARY H.  P
DRAGONS CAN ONLY RUST [GONARD 1]  CYMRI, CHRYS  P
DRAGON’S CARBUNCLE, THE [WIZARD’S WAR 3]  BOYER, ELIZABETH H.  C/P,P
DRAGON’S CLAW [SCORPIO 6]  MCDONOUGH, ALEX  C/P,P
DRAGON’S DAGGER, THE [SPEARWIELDER 2]  SALVATORE, R. A.  C/P,P
DRAGONS DISK, THE (CD)  BAEN PUBLISHING ENTERPRISES  SR-MM
DRAGON’S DOMAIN  GUNNARSSON, THORARINN  P
DRAGON’S DOOM [BAND OF FOUR]  GREENWOOD, ED  H
DRAGON’S EGG  FORWARD, ROBERT L.  C/P:2,SFWA,P,H
DRAGON’S EGG (PROOF)  FORWARD, ROBERT L.  SR-LP
DRAGONS, ELVES, AND HEROES  CARTER, LIN  PA
DRAGON’S EYE, THE  CHAMPETIER, JOEL  LP
DRAGON’S EYE [STONEFORT,MAINE 1]  HETLEY, JAMES A.  LP
DRAGON’S EYE [SCORPIO 5]  MCDONOUGH, ALEX  P
DRAGON’S EYE  STASHEFF, CHRISTOPHER  PA
DRAGON’S FIRE [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  H
DRAGON’S GOLD [KELVIN HACKLEBERRY 1]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:2,SFWA-P,P
DRAGON’S HONOR [STAR TREK TNG]  JOHNSON, KIJ/COX, GREG  C/P:2,P
DRAGONS IN THE ARCHIVES [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS IN THE STARS [STAR DRAGONS 1]  CARVER, JEFFREY A.  C/P,P
DRAGONS IN THE WATERS  L’ENGLE, MADELEINE  C/P,P
DRAGON’S ISLAND  WILLIAMSON, JACK  C/P:2,P
DRAGON’S KIN [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  H
DRAGON’S KIN (PROOF) [PERN]  MCCAFFREY, ANNE/MCCAFFREY, TODD  SR-LP
DRAGON’S KNIGHT  DENNIS, CAROL L.  P
DRAGON’S MILK  FLETCHER, SUSAN  LP
DRAGONS OF A FALLEN SUN [DRAGONLANCE WAR OF SOULS 1]  WEIS, MARGARET/HICKMAN, TRACY  H
DRAGONS OF A LOST STAR [DRAGONLANCE WAR OF SOULS 2]  WEIS, MARGARET/HICKMAN, TRACY  H
DRAGONS OF A VANISHD MOON [DRAGONLANCE WAR OF SOULS 3]  WEIS, MARGARET/HICKMAN, TRACY  H
DRAGONS OF ARGONATH [BAZIL BROKETAIL 6]  ROWLEY, CHRISTOPHER  P
DRAGONS OF AUTUMN TWILIGHT [@DRAGONLANCE CHRONICLES #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF CHAOS, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF DARKNESS  CARD, ORSON SCOTT  C/PA,LPA
DRAGONS OF EDEN, THE  SAGAN, CARL  [C/RAN-P:2]
DRAGONS OF ENGLOR, THE [@BLADE #24]  LORD, JEFFREY  P
DRAGONS OF KOMAKO, THE [DR. BONES]  BETANCOURT, JOHN GREGORY  P
DRAGONS OF KRYNN, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF LIGHT  CARD, ORSON SCOTT  C/PA,PA,SR-LPA
DRAGONS OF MAGIC, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  PA
DRAGONS OF NORTH CHITTENDON, THE  SCHAEFFER, SUSAN FROMBERG  H,LP
DRAGONS OF SPRING DAWNING [@DRAGONLANCE CHRONICLES #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF SPRINGSPACE, THE  REED, ROBERT  H
DRAGONS OF SUMMER FLAME [DRAGONLANCE CHRONICLES 4]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
DRAGONS OF THE CUYAHOGA, THE [DRAGONS OF THE CUYAHOGA 1]  SWANN, S. ANDREW  C/P,P

DRAGONS OF THE DWARVEN DEPTHS [@DRAGONLANCE LOST CHRONICLES #1] WEIS, MARGARET/HICKMAN, TRACY

H

DRAGONS OF THE HIGHLORD SKIES [@DRAGONLANCE LOST CHRONICLES #2] WEIS, MARGARET/HICKMAN, TRACY

H

DRAGONS OF THE RHINE, THE [WODAN’S CHILDREN 2]  PAXSON, DIANA L.  C/P,P
DRAGONS OF TIME [DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
DRAGONS OF WAR [BAZIL BROKETAIL 3]  ROWLEY, CHRISTOPHER  C/P,P
DRAGONS OF WINTER NIGHT [@DRAGONLANCE CHRONICLES #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 101
TITLE  AUTHOR  CODE
DRAGONS ON THE TOWN [CALIFORNIA DRAGONS 3]  GUNNARSSON, THORARINN  P
DRAGONS OVER ENGLAND [TORG POSSIBILITY WARS]  WEST END BOOKS  PA
DRAGONS PAST [DRAGON WARS 2]  GENTILE, GARY  C/P,P
DRAGON’S PAWN  DENNIS, CAROL L.  C/P,P
DRAGON’S QUEEN  DENNIS, CAROL L.  P
DRAGON’S RETURN, THE [MONTE COOK’S ARCANA EVOLVED]  COOK, SUE WEINLEIN  PA
DRAGONS REVEALED [KNIGHTS OF THE DRAGONS REVELATIONS 2]  FORBECK, MATT  LP
DRAGON’S SISTER, THE/TIMOTHY TRAVELS  STOREY, MARGARET  C/P,LP
DRAGON’S TEETH [STONEFORT,MAINE 2]  HETLEY, JAMES A.  LP
DRAGON’S TEETH [JANNA BRILL 3]  KILLOUGH, LEE  C/P:2,P
DRAGON’S TEETH  MILES, KEITH  C/P,P
DRAGON’S TOMB, THE [@DUNGEONS & DRAGONS PENHALIGON #2]  HEINRICH, D. J.  P
DRAGON’S TONGUE [DEMON-BOUND 1]  UNDERWOOD, LAURA J.  LP
DRAGON’S TOUCHSTONE, THE [DRAGON NIMBUS HISTORY 1]  RADFORD, IRENE  C/P,P
DRAGON’S TREASURE [DRAGON’S COUNTY 2]  LYNN, ELIZABETH A.  H
DRAGON’S WINTER [DRAGON’S COUNTY 1]  LYNN, ELIZABETH A.  C/P,P
DRAGONS: WORLD AFIRE  WIZARDS OF THE COAST  H
DRAGONS’ WRATH [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  C/P,P
DRAGONSBANE [WINTERLANDS 1]  HAMBLY, BARBARA  C/P,P,H
DRAGONSBLOOD [PERN]  MCCAFFREY, TODD  H
DRAGONSDAWN [PERN]  MCCAFFREY, ANNE  H:2,P
DRAGONSEYE [PERN]  MCCAFFREY, ANNE  H:2,P
DRAGONSHADOW [WINTERLANDS 2]  HAMBLY, BARBARA  H
DRAGONSINGER [PERN HARPER HALL 2]  MCCAFFREY, ANNE  H:2,P
DRAGONSLAYER  DREW, WAYLAND  H:2,P
DRAGONSLAYER [WARHAMMER GOTREK & FELIX 4]  KING, WILLIAM  P
DRAGONSLAYERS, THE  COVILLE, BRUCE  LP
DRAGONSLAYER’S RETURN [SPEARWIELDER 3]  SALVATORE, R. A.  P
DRAGONSONG [PERN HARPER HALL 1]  MCCAFFREY, ANNE  H:2,P
DRAGONSPAWN [LAND BETWEEN THE RIVERS 1]  ACRES, MARK  P
DRAGONSPELL [DRAGONSPELL 1]  PAUL, DONITA K.  LP
DRAGONSTAR [WINTERLANDS 4]  HAMBLY, BARBARA  H

DRAGONSTAR DESTINY [DRAGONSTAR 3] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P,P

DRAGONSTONE, THE [MITHGAR 1]  MCKIERNAN, DENNIS L.  P
DRAGONSTORM [FORTUNE 2]  EDWARDS, GRAHAM  P
DRAGONSWORD [DRAGONSWORD 1]  BAUDINO, GAEL  C/P:2,P
DRAGONWALL [FORGOTTEN REALMS EMPIRES 2]  DENNING, TROY  C/P:2,P
DRAGONWORLD  PREISS, BYRON/REAVES, J. MICHAEL  H:2,P
DRAKAS! [DRAKA]  STIRLING, S. M.  C/PA,PA
DRAKON [DRAKA 4]  STIRLING, S. M.  C/P:2,P
DRAMA IN LIVONIA  VERNE, JULES  H
DRAMA READER, THE  CONGDON, S. PERRY II  LP
DRAMATURGES OF YAN, THE  BRUNNER, JOHN  C/P:2,P
DRAMOCLES  SHECKLEY, ROBERT  H:2
DRASTIC DRAGON OF DRACO, TEXAS, THE  SCARBOROUGH, ELIZABETH ANN  P
DRAUGHT OF THE BLUE, A  BAIN, F. W.  H
DRAW ONE IN THE DARK  HOYT, SARAH A.  H
DRAWING OF THE DARK, THE  POWERS, TIM  P
DRAWING OF THE THREE, THE [DARK TOWER 2]  KING, STEPHEN  C/P:2,@H:2,P
DREAD BRASS SHADOWS [GARRETT 5]  COOK, GLEN  C/P,P
DREAD COMPANION  NORTON, ANDRE  H:2,P
DREAD VISITOR  BERRY, BRYAN  SR-P
DREADFUL FUTURE OF BLOSSOM CULP, THE [BLOSSOM CULP 3]  PECK, RICHARD  C/P,P
DREADFUL HOLLOW  KARLOVA, IRINA  P
DREADFUL SANCTUARY (REVISED)  RUSSELL, ERIC FRANK  C/P:2,P
DREADNOUGHT! [STAR TREK]  CAREY, DIANE  C/P:3,P
DREAM A LITTLE DREAM  ANTHONY, PIERS/BRADY, JULIE  P
DREAM BABY  MCALLISTER, BRUCE  LP
DREAM BABY  MCCALLISTER, BRUCE  P
DREAM CATCHER, THE [ARC ONE 2]  HUGHES, MONICA  SFWA
DREAM CHARIOTS  NORVIL, MANNING  C/P,P
DREAM COMPASS, THE  BREDENBERG, JEFF  C/P,P
DREAM COUNTRY [@SANDMAN #3]  GAIMAN, NEIL  H,CX
DREAM DANCER [KERRION EMPIRE 1]  MORRIS, JANET  H:2,SFWA,P
DREAM DAYS  GRAHAME, KENNETH  LP
DREAM DOCTOR, THE  REEVE, ARTHUR B.  H
DREAM GAMES  HANSEN, KARL  C/P,P
DREAM HUNTERS, THE [@SANDMAN]  GAIMAN, NEIL  CX
DREAM MAKER  HARBINSON, W. A.  SFWA
DREAM MASTER, THE  ZELAZNY, ROGER  C/P:2,SFWA-P,P
DREAM MILLENNIUM, THE  WHITE, JAMES  C/P:2,P
DREAM OF DRACULA, A  WOLF, LEONARD  R/REF-P
DREAM OF GLASS  GAWRON, JEAN MARK  H:2

TITLE  AUTHOR  CODE
DREAM OF KINSHIP, A [BIRD OF KINSHIP 2]  COWPER, RICHARD  C/P,P
DREAM OF THE DEAD  SETH, MARIE  SR-P
DREAM OF THINGS THAT WERE, A  ALEXANDER, KATHERINE  P
DREAM OF WESSEX, A  PRIEST, CHRISTOPHER  P
DREAM OF X, THE  HODGSON, WILLIAM HOPE  SR-VLH
DREAM, OR THE SIMIAN MAID  WRIGHT, S. FOWLER  H
DREAM PALACE, THE  STEPHENS, BRYNNE  P
DREAM PARK [DREAM PARK]  NIVEN, LARRY/BARNES, STEVEN  H:3,P
DREAM-QUEST OF UNKNOWN KADATH, THE  LOVECRAFT, H. P.  C/P:2,P
DREAM SCIENCE  PALMER, THOMAS  H
DREAM SPHERES [@FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
DREAM SQUAD, THE  HASSLER, KENNETH W.  H
DREAM STORIES AND FANTASIES  O’BRIEN, FITZ-JAMES  SFWA
DREAM THIEF  LAWHEAD, STEPHEN R.  LP
DREAM VESSEL, THE  BREDENBERG, JEFF  P,SFWA-P
DREAM WARRIOR, THE  CONWAY, D. J.  LP
DREAM YEARS, THE  GOLDSTEIN, LISA  H:2,P
DREAMBUILDER [SOULSMITH 2]  DEITZ, TOM  P,SFWA-P
DREAMCASTLE [WEB,THE]  BOWKETT, STEPHEN  P
DREAMCATCHER  KING, STEPHEN  H:2
DREAMER [SILENT EMPIRE 1]  HARPER, STEVEN  P
DREAMER  QUINN, DANIEL  SFWA-P
DREAMER IN DISCORD, THE  MILLS, CRAIG  P
DREAMERS, THE  GUNN, JAMES E.  H,SFWA
DREAMERS, THE  MANVELL, ROGER  P
DREAMER’S TALES AND OTHER STORIES, A  DUNSANY, LORD  SR-P,H
DREAMFALL [CAT 3]  VINGE, JOAN D.  H:2
DREAMFIELDS, THE  JETER, K. W.  SR-P,C/P
DREAMHUNTER [DREAMHUNTER 1]  KNOX, ELIZABETH  H
DREAMING CITY, THE [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
DREAMING DOWN UNDER  DANN, JACK/WEBB, JANEEN  HA,LPA
DREAMING DRAGONS, THE  BRODERICK, DAMIEN  C/P,P
DREAMING EARTH, THE  BRUNNER, JOHN  C/P:3,P
DREAMING IN SMOKE  SULLIVAN, TRICIA  P
DREAMING JEWELS, THE  STURGEON, THEODORE  SR-H,C/P,LP,H
DREAMING JEWELS, THE/COSMIC RAPE, THE/VENUS PLUS X  STURGEON, THEODORE  H
DREAMING METAL  SCOTT, MELISSA  LP
DREAMING TREE, THE [ARAFEL (REVISED) 1,2]  CHERRYH, C. J.  P
DREAMLAND CHRONICLES [DREAMLAND 1,2,3]  SIMMONS, WM. MARK  LP
DREAMQUAKE [DREAMHUNTER 2]  KNOX, ELIZABETH  H
DREAMRIDER  MIESEL, SANDRA  C/P,P
DREAMS  JACKSON, ANITA  SR-P
DREAMS AND DELIGHTS  BECK, L. ADAMS  H
DREAM’S EDGE  CARR, TERRY  HA,LPA
DREAMS MADE FLESH  BISHOP, ANNE  LP
DREAMS OF AN UNSEEN PLANET  PLOWRIGHT, TERESA  H,SFWA
DREAMS OF DAWN  STEUSSY, MARTI  C/P,P,H
DREAMS OF EMPIRE [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
DREAMS OF FLESH AND SAND  QUICK, W. T.  C/P,P
DREAMS OF GODS AND MEN  QUICK, W. T.  P
DREAMS OF STEEL [BLACK COMPANY 5/BLACK COMPANY SOUTH 2]  COOK, GLEN  C/P:2,SFWA-P,P
DREAMS OF STONE [@UNBALANCED EARTH #1]  WYLIE, JONATHAN  P
DREAMS OF THE RAVEN [STAR TREK]  CARTER, CARMEN  C/P:2,P
DREAMS OF THE SEA  VONARBURG, ELISABETH  LP
DREAMS THAT BURN IN THE NIGHT  STRETE, CRAIG  H,SFWA
DREAMS UNDERFOOT  DE LINT, CHARLES  C/P,P,H
DREAMSEEKER’S ROAD [OTHERWORLD 6]  DEITZ, TOM  H,P
DREAMSHIPS  SCOTT, MELISSA  H:2,SFWA,P
DREAMSIDE  JOYCE, GRAHAM  H,LP
DREAMSNAKE  MCINTYRE, VONDA N.  C/P,H:2,P
DREAMSPY  LICHTENBERG, JACQUELINE  H
DREAMSTONE, THE [ARAFEL 1]  CHERRYH, C. J.  C/P:2,P
DREAMSTONE, THE [RUNESWORD 3]  RIVKIN, J. F.  C/P,P
DREAMSTONE MOON [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
DREAMTHIEF’S DAUGHTER, THE  MOORCOCK, MICHAEL  H,P
DREAMWATCHER  ROSZAK, THEODORE  SFWA
DREAMWEAVER’S DILEMMA  BUJOLD, LOIS MCMASTER  H
DREAMWRIGHT [@MIGHT AND MAGIC #1]  GRAVEL, GEARY  P
DRIFT [DOCTOR WHO PAST DOCTORS]  FORWARD, SIMON A.  P
DRIFT, THE  KROPP, LLOYD  H,P
DRIFTER [DRIFTER 1]  DIETZ, WILLIAM C.  C/P:2,SFWA-P:2,P
DRIFTER’S RUN [DRIFTER 2]  DIETZ, WILLIAM C.  C/P,P
DRIFTER’S WAR [DRIFTER 3]  DIETZ, WILLIAM C.  C/P:2,SFWA-P,P
DRIFTGLASS  DELANY, SAMUEL R.  H:2,P

TITLE  AUTHOR  CODE
DRINK DOWN THE MOON [JACKY ROWAN 2]  DE LINT, CHARLES  C/P:2,SFWA-P:2,P
DRINK THE FIRE FROM THE FLAMES [ASHLU]  BAKER, SCOTT  P,SFWA-P
DRINKER OF SOULS [SOUL DRINKER 1]  CLAYTON, JO  C/P:2,P
DRINKING MIDNIGHT WINE  GREEN, SIMON R.  LP
DRINKING SAPPHIRE WINE  LEE, TANITH  P
DRITTE MACHT, DIE [@PERRY RHODAN SILBERBANDE #1]  PABEL MOEWIG KG  FPR
DRIVE-IN, THE [@DRIVE-IN #1]  LANSDALE, JOE R.  C/P,P
DRIVE-IN 2, THE [@DRIVE-IN #2]  LANSDALE, JOE R.  C/P,P
DRIVE TO THE EAST [@SETTLING ACCOUNTS 2]  TURTLEDOVE, HARRY  H
DRIVEN TO KILL!  GARTH, G. G.  P
DRIVING BLIND  BRADBURY, RAY  P
DROPPED FROM THE CLOUDS  VERNE, JULES  H
DROPS OF CORRUPTION [SHADOWRUN]  HARDY, JASON M.  P
DROUGHT, THE  BALLARD, J. G.  C/P,P
DROWNED AMMET [DALEMARK 2]  JONES, DIANA WYNNE  C/P,P
DROWNED QUEEN, THE [AVENGERS 6]  LAUMER, KEITH  P
DROWNED WEDNESDAY [@KEYS TO THE KINGDOM #3]  NIX, GARTH  LP
DROWNED WORLD, THE  BALLARD, J. G.  C/P:2,VLP
DROWNED WORLD, THE/WIND FROM NOWHERE, THE  BALLARD, J. G.  H:2
DROWNING TOWERS  TURNER, GEORGE  H
DROWNING WORLD [HUMANX]  FOSTER, ALAN DEAN  H
DROWNTIDE  VAN SCYOC, SYDNEY J.  C/P,SFWA-P,P
DROWSY  MITCHELL, J. A.  H
D.R.T. [@BATTLETECH 14]  LONG, JAMES D.  P
DRUG OF CHOICE  LANGE, JOHN  P
DRUID OF SHANNARA, THE [SHANNARA HERITAGE 2]  BROOKS, TERRY  H:2,P
DRUID QUEEN, THE [@FORGOTTEN REALMS DRUIDHOME #3]  NILES, DOUGLAS  C/P,P
DRUID STONE, THE  MAJORS, SIMON  C/P:2,P
DRUIDS  LLYWELYN, MORGAN  P
DRUID’S BLOOD  FRIESNER, ESTHER M.  P
DRUID’S SWORD [TROY GAME 4]  DOUGLASS, SARA  H
DRUM CALLS [DRUMS OF CHAOS 2]  CLAYTON, JO  P
DRUM INTO SILENCE [DRUMS OF CHAOS 3]  CLAYTON, JO/MURPHY, KEVIN ANDREW  H,P
DRUM WARNING [DRUMS OF CHAOS 1]  CLAYTON, JO  C/P,P
DRUMS OF DARKNESS  BRADLEY, MARION ZIMMER  P
DRUMS OF DRACULA [@DRACULA #5]  LORY, ROBERT  P
DRUMS OF FU MANCHU, THE  ROHMER, SAX  C/P,P,H
DRUMS OF TAPAJOS  MEEK, COLONEL S. P.  H
DRUMS OF THE DARK GODS  BALLINGER, W. A.  P
DRUNE, THE  PALMER, JANE  LP
DRUNKARD’S WALK  POHL, FREDERIK  C/P:2,P
DRY PILGRIMAGE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  LEONARD, PAUL/WALTERS, NICK  C/P,P
DRY WATER  NYLUND, ERIC S.  LP
DRYAD, THE  MCCARTHY, JUSTIN HUNTLY  H
DRYLANDS, THE  ROSENBLUM, MARY  P
D’SHAI [D’SHAI 1]  ROSENBERG, JOEL  C/P,SFWA-P:2,P,H
DUAL TO THE DEATH [@BATMAN ANIMATED SERIES 2]  GRAVEL, GEARY  P
DUBIOUS HILLS, THE  DEAN, PAMELA  H,P
DUCHESS OF KNEEDEEP, THE  NOEL, ATANIELLE ANNYN  C/P:2,P
DUDGEON & DRAGONS [@KEDRIGERN CHRONICLES #2]  MORRESSY, JOHN  LP
DUDLEY AND GILDEROY  BLACKWOOD, ALGERNON  C/P,P,H
DUEL IN NIGHTMARE WORLDS  FLACKES, B.  SR-P
DUEL OF DRAGONS [DRAGONSWORD 2]  BAUDINO, GAEL  C/P,SFWA-P:2,P
DUEL UNDER THE DOUBLE SUN [@PERRY RHODAN #108]  ACE  PA
DUELING MACHINE, THE  BOVA, BEN  SR-H,C/P:2,P
DUENDE HISTORY OF THE SHADOW MAGAZINE, THE  MURRAY, WILL  R/REF-H
DUENDE MEADOW  COOK, PAUL  P
DUENO DEL MUNDO  MILK, LOUIS G.  FP
DUJONIAN’S HOARD [STAR TREK CAPTAIN’S TABLE 2/STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P,P
DUKE OF SUMAVA, THE  WRENCH, SARA J.  P
DUKE OF URANIUM, THE [JAK JINNAKA 1]  BARNES, JOHN  C/P:2,P
DUKE’S BALLAD, THE  NORTON, ANDRE/MCCONCHIE, LYN  H
DULUTH  VIDAL, GORE  H
DUMB GODS SPEAK, THE  OPPENHEIM, E. PHILLIPS  H
DUN LADY’S JESS [DUN LADY’S JESS 1]  DURGIN, DORANNA  P
DUNCTON WOOD  HORWOOD, WILLIAM  P
DUNE [DUNE 1]  HERBERT, FRANK  C/P:3,SFWA-P,P,H
DUNE ENCYCLOPEDIA, THE  MCNELLY, WILLIS E.  R/REF-H,C/REF-H
DUNE: HOUSE ATREIDES [DUNE 0.1]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE: HOUSE CORRINO [DUNE 0.3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE: HOUSE CORRINO (6 CASSETTES, BOXED) [DUNE 0.3]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-MM
DUNE: HOUSE HARKONNEN [DUNE 0.2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2,P
DUNE MESSIAH [DUNE 2]  HERBERT, FRANK  C/P,P,H
DUNE: THE BUTLERIAN JIHAD [DUNE JIHAD 1]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2

The MITSFS Titledex  September 24, 2007  Page 104
TITLE  AUTHOR  CODE
DUNE: THE MACHINE CRUSADE [DUNE JIHAD 2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H:2
DUNE: THE MACHINE CRUSADE (PROOF) [DUNE JIHAD 2]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-LP
DUNE: WHIPPING MEK [DUNE JIHAD]  HERBERT, BRIAN/ANDERSON, KEVIN J.  B
DUNE: WHIPPING MEK (BOOKLET AND CD) [DUNE JIHAD]  HERBERT, BRIAN/ANDERSON, KEVIN J.  SR-MM:2
DUNES OF PRADAI  WAYMAN, TONY RUSSELL  P
DUNGEON OF DREAD  ESTES, ROSE  [C/RAN-P:2]
DUNGEONS & DRAGONS THE MOVIE [@DUNGEONS & DRAGONS]  BARRETT, NEAL JR.  P
DUNGEONS OF KUBA [EMPIRE PRINCESS 2]  DIAMOND, GRAHAM  P
DUNGEONS OF TORGAR, THE [LONE WOLF 10]  DEVER, JOE  P
DUNN’S CONUNDRUM  LEE, STANLEY R.  C/P,P
CURSE OF THE WISE WOMAN, THE  DUNSANY, LORD  SR-P
DUNWICH HORROR, THE  LOVECRAFT, H. P.  SR-H,H
DUNWICH HORROR, THE (BART)  LOVECRAFT, H. P.  SR-P
DUNWICH HORROR, THE (LANCER)  LOVECRAFT, H. P.  C/P:2,P
DUPLICATE, THE  SLEATOR, WILLIAM  H,P
DUPLICATE LOVERS  BURCH, RALPH  SR-L
DUPLICATED MAN, THE  BLISH, JAMES/LOWNDES, ROBERT A. W.  C/P:2,P
DUPLICATORS, THE  LEINSTER, MURRAY  DF275
DURDANE [@VANCE INTEGRAL EDITION #27]  VANCE, JACK  SR-H
DUSHAU [DUSHAU 1]  LICHTENBERG, JACQUELINE  C/P,P
DUSK  LEBBON, TIM  LP
DUST  PELLEGRINO, CHARLES  H
DUST DESTROYERS, THE  STATTEN, VARGO  SR-P
DUST OF DEATH [DOC SAVAGE #32]  ROBESON, KENNETH  C/P:2,@P
DUST OF FAR SUNS  VANCE, JACK  P
DUSTLAND [JUSTICE 2]  HAMILTON, VIRGINIA  C/P,P
DUTCHMAN  KNAAK, RICHARD A.  P
DUTY CALLS [@WARHAMMER 40K]  MITCHELL, SANDY  P
DUTY, HONOR, REDEMPTION [@STAR TREK]  MCINTYRE, VONDA N.  LP
DWARVES OF WHISKEY ISLAND [DRAGONS OF THE CUYAHOGA 2]  SWANN, S. ANDREW  P
DWELLER ON THE THRESHOLD, THE  HICHENS, ROBERT  H
DWELLER ON TWO PLANETS, A  PHYLOS  H
DWELLERS IN THE CRUCIBLE [STAR TREK]  BONANNO, MARGARET WANDER  C/P:3,P
DWELLERS IN THE MIRAGE  MERRITT, A.  C/P:2,SFWA-P,P
DWELLERS OF THE DEEP  O’DONNELL, K. M.  D27400
DYDEETOWN WORLD  WILSON, F. PAUL  C/P:2,SFWA-P,P
DYING DAYS, THE [DOCTOR WHO NEW ADVENTURES]  PARKIN, LANCE  C/P,P
DYING EARTH, THE [DYING EARTH 1]  VANCE, JACK  C/P:2,P
DYING FOR TOMORROW  MOORCOCK, MICHAEL  C/P,P
DYING IN THE SUN [DOCTOR WHO PAST DOCTORS]  MILLER, JON DE BURGH  P
DYING INSIDE  SILVERBERG, ROBERT  C/P,SFWA,P,H

DYING KINGDOM, THE [@DRAGONLANCE NEW ADVENTURES #2/ DRAGONLANCE SPELLBINDER QUARTET 2] SULLIVAN, STEPHEN D.

LP

DYING LIGHT  CHANDLER, EVAN  P
DYING OF PARADISE  COUPER, STEPHEN  SR-P
DYING OF THE LIGHT  MARTIN, GEORGE R. R.  H:2,SFWA-P,P
DYING TIME, THE [@MECHWARRIOR]  GRESSMAN, THOMAS S.  P
DYKSTRA’S WAR  KOOISTRA, JEFFREY D.  C/P,P
DYNO-DEPRESSANT, THE  GRIDBAN, VOLSTED  SR-P
DYNOSTAR MENACE, THE  PEDLER, KIT/DAVIS, GERRY  H
DYNTERYX  MURDOCK, M. S.  P

DYSON SPHERE [@STAR TREK TNG] PELLEGRINO, CHARLES/ ZEBROWSKI, GEORGE

C/P:2,P

DYSTOPIAN VISIONS  ELWOOD, ROGER  SR-HA
DZIENNIKI GWIAZDOWE  LEM, STANISLAW  FP
DZUR [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  H
DZURLORD  FAWCETT, BILL  SFWA-P

TITLE  AUTHOR  CODE
E PLURIBUS UNICORN  STURGEON, THEODORE  C/P:2,P
EAGLE, THE [CAMULOD 9]  WHYTE, JACK  H:2
EAGLE & THE NIGHTINGALES, THE [@BARDIC VOICES #3]  LACKEY, MERCEDES  H,P
EAGLE AGAINST THE STARS  WHITE, STEVE  C/P,P
EAGLE AND THE SWORD, THE  ATTANASIO, A. A.  LP
EAGLE AND THE SWORD, THE  SCHREIBER, HARVEY K.  P
EAGLE-SAGE [LON TOBYN 3]  COE, DAVID B.  H,P
EAGLES’ BROOD, THE [CAMULOD 3]  WHYTE, JACK  P
EAR, THE  JACKSON, ANITA  SR-P
EARL AUBEC AND OTHER STORIES [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  P
EARLY ASIMOV, THE  ASIMOV, ISAAC  H:2,SFWA
EARLY ASIMOV 1, THE  ASIMOV, ISAAC  C/P:2,P
EARLY ASIMOV 2, THE  ASIMOV, ISAAC  C/P:2,P
EARLY DEL REY  DEL REY, LESTER  H:2,SFWA
EARLY DEL REY 1, THE  DEL REY, LESTER  C/P,P
EARLY DEL REY 2, THE  DEL REY, LESTER  C/P:2,P
EARLY HARVEST  BEAR, GREG  SR-H
EARLY LONG, THE  LONG, FRANK BELKNAP  H:2,SFWA
EARLY POHL, THE  POHL, FREDERIK  H:2,SFWA
EARLY WILLIAMSON, THE  WILLIAMSON, JACK  H,SFWA,P
EARTH  BRIN, DAVID  H:2,P
EARTH  FARCA, MARIE C.  H:2,SFWA
EARTH 2 [EARTH 2]  CRANDALL, MELISSA  C/P,P
EARTH ABIDES  STEWART, GEORGE R.  H:2,P
EARTH AGAIN REDEEMED, THE  GREEN, MARTIN  H
EARTH AND ELSEWHERE  MACMILLAN PUBLISHING  HA
EARTH BOOK OF STORMGATE, THE  ANDERSON, POUL  H:2,P
EARTH DESCENDED  SABERHAGEN, FRED  C/P:2,SFWA-P,P
EARTH DIES, THE [@PERRY RHODAN #41]  ACE  PA,C/PA
EARTH DREAMS [KERRION EMPIRE 3]  MORRIS, JANET  C/P,P,H
EARTH ENSLAVED [@CAP KENNEDY #9]  KERN, GREGORY  P
EARTH FACTOR X (ORIGINALLY THE SECRET GALACTICS)  VAN VOGT, A. E.  C/P,P
EARTH FIRE [@SURVIVALIST #9]  AHERN, JERRY  P
EARTH GODS ARE COMING, THE  BULMER, KENNETH  DD453
EARTH GONE MAD, AN  DEE, ROGER  DD084
EARTH HAS BEEN FOUND  JONES, D. F.  C/P,P
EARTH HERALD  CLARK, JAN  P
EARTH HIVE [@ALIENS #1]  PERRY, STEVE  C/P,P
EARTH IN PERIL, THE  WOLLHEIM, DONALD A.  DD205
EARTH IN TRANSIT  SCHWARTZ, SHEILA  C/PA,PA
EARTH IN TWILIGHT  PISERCHIA, DORIS  P
EARTH IN UPHEAVAL  VELIKOVSKY, IMMANUEL  [C/RAN-P:2]
EARTH INVADER  GARRETT, RANDALL  C/P:2,P
EARTH INVASION BATTALION, THE  HUGHES, DENIS  SR-P
EARTH IS ALL THAT LASTS, THE [CHILDREN OF EARTH 1]  WELLS, CATHERINE  P
EARTH IS HEAVEN [DUMAREST 27]  TUBB, E. C.  C/P,P
EARTH IS NEAR, THE  PESEK, LUDEK  C/P,P
EARTH IS ROOM ENOUGH  ASIMOV, ISAAC  C/P,P,H

EARTH IS THE ALIEN PLANET:

J. G. BALLARD’S FOUR-DIMENSIONAL NIGHTMARE PRINGLE, DAVID

R/REF-P

EARTH IS THE STRANGEST PLANET  SILVERBERG, ROBERT  HA
EARTH LIES SLEEPING [@RACK #1]  JAMES, LAURENCE  P
EARTH LOGIC [ELEMENTAL LOGIC 2]  MARKS, LAURIE J.  H
EARTH LORDS, THE  DICKSON, GORDON R.  C/P:2,SFWA-P,P
EARTH MADE OF GLASS [THOUSAND CULTURES 2]  BARNES, JOHN  C/P,P,H
EARTH MAGIC  PANSHIN, ALEXEI/PANSHIN, CORY  C/P:2,P
EARTH MAN ON VENUS, AN  FARLEY, RALPH M.  SR-P
EARTH OF FIRE, SKY OF ICE [SHALRODAN SAGA 1]  HOTCHKISS, ROBERT R.  LP
EARTH: OUR CROWDED SPACESHIP  ASIMOV, ISAAC  [C/RAN-P]
EARTH REMEMBERS, THE  OLAN, SUSAN TORIAN  C/P,P
EARTH SATELLITES  MOORE, PATRICK  [C/RAN-H]
EARTH SATELLITES  STINE, G. HARRY  [C/RAN-P]
EARTH SAVER, THE [CHILDREN OF EARTH 3]  WELLS, CATHERINE  P
EARTH SET FREE  TUBB, E. C.  LP
EARTH-SHAKER, THE [ZARKON 4]  CARTER, LIN  H,SFWA
EARTH SHIP AND STAR SONG  SHEDLEY, ETHAN I.  C/P,P
EARTH SONG [EARTH SONG 2]  WEBB, SHARON  C/P,P,H
EARTH STATION SEX  KAINEN, RAY  SR-L
EARTH STRIKES BACK, THE  CHIZMAR, RICHARD T.  PA
EARTH TIMES TWO  REYNOLD, PAMELA  H
EARTH TRIPPER, THE  KELLEY, LEO P.  C/P:2,P
EARTH TWO  KELLEY, LEO P.  SR-P
EARTH UNAWARE  REYNOLDS, MACK  C/P:2,P
EARTH WAR, THE  REYNOLDS, MACK  C/P:2,P
EARTH WILL SHAKE, THE [@HISTORICAL ILLUMINATUS #1]  WILSON, ROBERT ANTON  C/P:2,LP,H

The MITSFS Titledex  September 24, 2007  Page 106
TITLE  AUTHOR  CODE
EARTH WINTER [NEW ICE AGE 2]  MORAN, RICHARD  P
EARTH WITCH, THE  LAWRENCE, LOUISE  H,P
EARTHBLOOD [@EARTHBLOOD]  AXLER, JAMES  P
EARTHBLOOD  LAUMER, KEITH/BROWN, ROSEL GEORGE  H:2,P
EARTHBORN [HOMECOMING 5]  CARD, ORSON SCOTT  C/P,P,H:2
EARTHBORN, THE [EARTHBORN WARS 1]  COLLINS, PAUL  H
EARTHBOUND  LESSER, MILTON  H:2
EARTHBOUND  MATHESON, RICHARD  P
EARTHCHILD  PISERCHIA, DORIS  C/P,P
EARTHCHILD [EARTH SONG 1]  WEBB, SHARON  C/P,P,H
EARTHCLAN [UPLIFT 2,3]  BRIN, DAVID  H:2
EARTHCORE  SIGLER, SCOTT  LP
EARTHFALL  BROWN, JERRY EARL  C/P,P
EARTHFALL [HOMECOMING 4]  CARD, ORSON SCOTT  H:2,P
EARTHGRIP  TURTLEDOVE, HARRY  C/P:2,P
EARTHJACKET  HARTRIDGE, JON  H
EARTHLIGHT  CLARKE, ARTHUR C.  C/P,P,H
EARTHLING  DANIEL, TONY  H,LP
EARTHLING SAMPLER, THE  EARTHLING PUBLICATIONS  B
EARTHLOVE  MCALEER, NEIL  LP
EARTHLY CROWN, AN [@SWORD OF HEAVEN #1]  ELLIOTT, KATE  C/P:2,P
EARTHMAN, COME HOME  BLISH, JAMES  H:2,P
EARTHMAN, GO HOME!  ANDERSON, POUL  DD479,P
EARTHMAN, GO HOME!  ELLISON, HARLAN  C/P,P
EARTHMAN’S BURDEN  ANDERSON, POUL/DICKSON, GORDON R.  C/P:2,P
EARTHMEN AND STRANGERS  SILVERBERG, ROBERT  PA,C/PA
EARTHQUAKE MACHINE, THE  MITCHELSON, AUSTIN/UTECHIN, NICHOLAS P
EARTHQUAKE-MAKER, THE  ROBERTS, MURRAY  B
EARTHQUAKE WEATHER [FAULT LINES 2]  POWERS, TIM  P
EARTHRIM  KAMIN, NICK  D66160
EARTHRISE [DEATHDAY 2]  DIETZ, WILLIAM C.  P
EARTH’S LAST CITADEL  MOORE, C. L./KUTTNER, HENRY  C/P:2,P
EARTH’S LAST FORTRESS  VAN VOGT, A. E.  DD431
EARTH’S OTHER SHADOW  SILVERBERG, ROBERT  C/P,P,H
EARTH’S SECRET INHABITANTS  ROGO, D. SCOTT/CLARK, JEROME  [C/RAN-P]
EARTHSEED  SARGENT, PAMELA  C/P:2,LP
EARTHSHOCK [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
EARTHSONG [@NATIVE TONGUE #3]  ELGIN, SUZETTE HADEN  C/P:2,P
EARTHSTONE, THE [WESTRIA 4]  PAXSON, DIANA L.  P,SFWA-P
EARTHSTRINGS  RACKHAM, JOHN  D10293
EARTHWALK  SLATER, PHILIP  [C/RAN-P]
EARTHWEB  STIEGLER, MARC  C/P,P
EARTHWIND  HOLDSTOCK, ROBERT P.  C/P:2,P
EARTHWORKS  ALDISS, BRIAN W.  H:2,P
EARTHWORLD [DOCTOR WHO EIGHTH DOCTOR]  RAYNER, JACQUELINE  P
EARTHWRECK!  SCORTIA, THOMAS N.  C/P:2,P
EAST COAST CRISIS [V]  WEINSTEIN, HOWARD/CRISPIN, A. C.  C/P,P
EAST OF EALING [BRENTFORD 3]  RANKIN, ROBERT  P
EAST OF MIDNIGHT  LEE, TANITH  C/P,P,H
EAST OF THE SUN, WEST OF THE MOON [DRAGONS 4]  RINGO, JOHN  H
EAST WIND COMING, AN  COVER, ARTHUR BYRON  C/P,P
EASTERCON SPEECHES, THE  SHAW, BOB  R/REF-P
EASTERN STANDARD TRIBE  DOCTOROW, CORY  H
EASY TRAVEL TO OTHER PLANETS  MOONEY, TED  P
EATEN ALIVE [@STAR WARS GALAXY OF FEAR #1]  WHITMAN, JOHN  P
EATER  BENFORD, GREGORY  H,P
EATER OF WASPS [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
EATER OF WORLDS, THE [@CAP KENNEDY #8]  KERN, GREGORY  C/P,P
EATERS OF THE DEAD  CRICHTON, MICHAEL  C/P:2,P
EATING MEMORIES  ANTHONY, PATRICIA  C/P,LP
EBONY EYE, THE [@DRAGONLANCE NEW ADVENTURES SUNCATCHER #2]  SAMPSON, JEFF  LP
ECCE AND OLD EARTH [CADWAL 2]  VANCE, JACK  C/P,P,H
ECCE AND OLD EARTH (VIE) [@VANCE INTEGRAL EDITION #40/CADWAL 2]  VANCE, JACK  SR-H
ECCENTRIC CIRCLES  LICKISS, REBECCA  P
ECHEC AUX FUSEES  BUD, BUSTER  FP
ECHELON  CONVISIER, JOSH  LP
ECHO IN THE SKULL  BRUNNER, JOHN  DD385
ECHO OF ETERNITY [SHADOWLEAGUE 3]  FUREY, MAGGIE  C/P,P
ECHO ROUND HIS BONES  DISCH, THOMAS M.  C/P:2,P
ECHO X  BARZMAN, BEN  C/P:2,P
ECHOES [@TIME HUNTER 6]  MCLAUGHLIN, IAIN/BARTLETT, CLAIRE  LP

ECHOES [@STAR TREK VOYAGER] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN/ HOFFMAN, NINA KIRIKI C/P,P

TITLE  AUTHOR  CODE
ECHOES FROM AN IRON HARP  HOWARD, ROBERT E.  SR-H
ECHOES IN TIME [TIME TRADERS 6]  NORTON, ANDRE/SMITH, SHERWOOD  H
ECHOES OF CHAOS, THE [@WEAPONS OF CHAOS #1]  VARDEMAN, ROBERT E.  C/P:2,SFWA-P,P
ECHOES OF EARTH [ECHOES OF EARTH 1]  WILLIAMS, SEAN/DIX, SHANE  C/P,P
ECHOES OF EVIL [WHISPERING HILLS 6]  CAMERON, KATE  P
ECHOES OF HONOR [HONOR HARRINGTON 8]  WEBER, DAVID  H:2,P
ECHOES OF ISSEL [UNIFIED WORLDS 2]  THORNLEY, DIANN  P
ECHOES OF THE FOURTH MAGIC [YNIS AIELLE 1]  SALVATORE, R. A.  P
ECHOES OF THE GREAT SONG  GEMMELL, DAVID  P

ECHOES OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 1] CHALKER, JACK L.

C/P,LP,H

ECHOES OF THUNDER  DUNN, JACK/HALDEMAN, JACK C. II  TD32
ECHOES OF VALOR  WAGNER, KARL EDWARD  PA,SFWA-P
ECHOES OF VALOR 2  WAGNER, KARL EDWARD  HA,PA,SFWA-P
ECHOES OF VALOR 3  WAGNER, KARL EDWARD  C/PA,PA
ECHOING WORLDS, THE  BURKE, JONATHAN  SR-P
ECLIPSE [SONG CALLED YOUTH 1]  SHIRLEY, JOHN  C/P,LP
ECLIPSE AT NOON [@DEATHLANDS]  AXLER, JAMES  P
ECLIPSE CORONA [SONG CALLED YOUTH 3]  SHIRLEY, JOHN  C/P,P
ECLIPSE EXPRESS, THE  STATTEN, VARGO  SR-P
ECLIPSE OF DAWN, THE  EKLUND, GORDON  C/P,P
ECLIPSE OF THE KAI [@LEGENDS OF LONE WOLF #1]  DEVER, JOE/GRANT, JOHN  P
ECLIPSE PENUMBRA [SONG CALLED YOUTH 2]  SHIRLEY, JOHN  P
ECLIPSES  FELICE, CYNTHIA  C/P:2,P
ECLIPSING BINARIES [FAMILY D’ALEMBERT #8]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
ECO-FICTION  STADLER, JOHN  PA

ECODEATH WATKINS, WILLIAM JOHN/ SNYDER, EUGENE V.

H,SFWA

ECOLITAN ENIGMA, THE [ECOLITAN 4]  MODESITT, L. E. JR.  H,P
ECOLITAN OPERATION, THE [ECOLITAN 2]  MODESITT, L. E. JR.  SFWA-P
ECOLITAN PRIME [ECOLITAN 1,4]  MODESITT, L. E. JR.  LP
ECOLOG, THE  NELSON, RAY FARADAY  SR-P
ECOLOGIC ENVOY, THE [ECOLITAN 1]  MODESITT, L. E. JR.  C/P,P
ECOLOGIC SECESSION, THE [ECOLITAN 3]  MODESITT, L. E. JR.  C/P:2,P
ECOTOPIA  CALLENBACH, ERNEST  C/P:2,LP
ECOTOPIA EMERGING  CALLENBACH, ERNEST  P
ECSTASIA  BLOCK, FRANCESCA LIA  SFWA
EDDIE AND THE ARCHANGEL MIKE  BENEFIELD, BARRY  H
EDDIE BEAR, PRIVATE DETECTIVE [EDDIE BEAR 1,2]  RANKIN, ROBERT  H
EDDIE’S BLUE-WINGER DRAGON  ADLER, C. S.  P
EDEN  LEM, STANISLAW  FH
EDEN CYCLE, THE  GALLUN, RAYMOND Z.  C/P,P
EDENBORN  SAGAN, NICK  LP
EDGAR CAYCE ON ATLANTIS  CAYCE, EDGAR  [C/RAN-P]
EDGAR RICE BURROUGHS: A BIBLIOGRAPHY  DAY, BRADFORD M.  R/IND
EDGAR RICE BURROUGHS: MASTER OF ADVENTURE  LUPOFF, RICHARD A.  R/REF-P
EDGAR RICE BURROUGHS: THE MAN WHO CREATED TARZAN  PORGES, IRWIN  R/REF-H
EDGE, THE  BEAUMONT, CHARLES  SR-P
EDGE IN MY VOICE, AN  ELLISON, HARLAN  R/REF-H,C/REF-H:2
EDGE OF DESTRUCTION, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,C/P
EDGE OF ETERNITY  MULLER, JOHN E.  SR-P
EDGE OF FOREVER, THE  OLIVER, CHAD  H
EDGE OF HUMAN, THE [@BLADE RUNNER #2]  JETER, K. W.  P
EDGE OF INFINITY  GROTEBOER, JEFFREY  P
EDGE OF NEVER, THE  HOSKINS, ROBERT  HA,PA
EDGE OF RUNNING WATER, THE  SLOANE, WILLIAM  H:2,SR-L,P
EDGE OF THE EARTH, THE  HILL, RUSSELL  P
EDGE OF THE INFINITE, THE [@SPACE: 1999 YEAR 2 #6]  BUTTERWORTH, MICHAEL  P
EDGE OF THE SWORD, THE [@STAR TREK ERRAND OF VENGEANCE #1]  RYAN, KEVIN  P
EDGE OF THE WORLD, THE [FIRST AMERICANS]  SARABONDE, WILLIAM  P
EDGE OF TIME  GRINNELL, DAVID  C/P:2,DD362,P
EDGE OF TOMORROW, THE  ASIMOV, ISAAC  H:2,SFWA-P,P
EDGE OF TOMORROW, THE  FAST, HOWARD  C/P:2,P
EDGE OF VICTORY [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY 1,2]  KEYES, J. GREGORY  H
EDGES  LE GUIN, URSULA K./KIDD, VIRGINIA  C/PA,PA
EDICT, THE  EHRLICH, MAX  H:2,P
EDISON’S CONQUEST OF MARS  SERVISS, GARRETT P.  SR-H
EDNA’S FRUIT HAT AND OTHER STORIES  PUDNEY, JOHN  H
EDUCATION OF JENNIFER PARRISH, THE  MARTIN, RUSS  SFWA-P
EDUCATION OF OVERSOUL 7, THE  ROBERTS, JANE  C/P,P,H
EDWARD MARET  KATZ, ROBERT I.  LP
EEVALU  LUKENS, ADAM  H
EFFENDI [EL ISKANDRYIA 2]  GRIMWOOD, JON COURTENAY  LP
EFFICIENCY IN HADES  VALE, ROBERT B.  SR-H

The MITSFS Titledex  September 24, 2007  Page 108
TITLE  AUTHOR  CODE
EFFIGIES  WELLS, WILLIAM K.  H
EGG-SHAPED THING, THE  HODDER-WILLIAMS, CHRISTOPHER  H:2
EGGCHILD, THE  BAXTER, LORNA  P
EGGHEADS  DEVENPORT, EMILY  C/P,P
EGIL’S BOOK [WINTER WORLD 2]  MILLS, C. J.  C/P,P

E.GODZ ASPRIN, ROBERT LYNN/ FRIESNER, ESTHER M.

H:2

EGYPTIAN CAT MYSTERY, THE [@RICK BRANT #16]  BLAINE, JOHN  H:2
EHOES OF AN ALIEN SKY  HOGAN, JAMES P.  H
EIFELHEIM  FLYNN, MICHAEL  H
EIGHT, THE  NEVILLE, KATHERINE  P
EIGHT AGAINST UTOPIA  MASON, DOUGLAS R.  C/P:2,P
EIGHT DAYS OF LUKE  JONES, DIANA WYNNE  H
EIGHT DIME NOVELS  BLEILER, EVERETT F.  VLPA
EIGHT DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  DICKS, TERRANCE  C/P:2,P
EIGHT FANTASMS AND MAGICS  VANCE, JACK  C/P:2,P
EIGHT HUNDRED LEAGUES ON THE AMAZON  VERNE, JULES  H
EIGHT KEYS TO EDEN  CLIFTON, MARK  H:2,P
EIGHT SKILLED GENTLEMEN [MASTER LI 3]  HUGHART, BARRY  C/P,LP,H
EIGHT SPELLS A WEEK [SABRINA THE TEENAGE WITCH]  POCKET BOOKS  PA
EIGHT STORIES FROM THE REST OF THE ROBOTS  ASIMOV, ISAAC  C/P:3,P
EIGHT STRANGE TALES  GHIDALIA, VIC  PA
18 GREATEST SCIENCE FICTION STORIES  JANIFER, LAURENCE M.  PA
EIGHTH DAY OF THE WEEK, THE  HLASKO, MAREK  H
EIGHTH RANK, THE [DREAMERS OF THE DAY 2]  ROSS, DAVID D.  H,SFWA
EIGHTH STAGE OF FANDOM, THE  BLOCH, ROBERT  R/REF-P

EIGHTIES, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-H

’80’S: A LOOK BACK AT THE TUMULTUOUS DECADE 1980–1989, THE HENDRA, TONY/CERF, CHRISTOPHER/

ELBLING, PETER

SR-VLHA

EIGHTY MINUTE HOUR, THE  ALDISS, BRIAN W.  C/P,SFWA,P,H
EINSTEIN INTERSECTION, THE  DELANY, SAMUEL R.  C/P,P,H
EINSTEIN’S BRAIN  OLSHAKER, MARK  P
EINSTEIN’S BRIDGE  CRAMER, JOHN  C/P,LP
EISENHORN [@WARHAMMER 40K EISENHORN 1,2,3]  ABNETT, DAN  H,LP
ELANTRIS  SANDERSON, BRANDON  H
ELDAR PROPHECY [@WARHAMMER 40K]  GOTO, C. S.  P
ELDER, THE  FREDDI, CHRIS  H
ELDER GODS, THE [@DREAMERS #1]  EDDINGS, DAVID/EDDINGS, LEIGH  H:2
ELDER GODS, THE (PROOF) [DREAMERS 1]  EDDINGS, DAVID/EDDINGS, LEIGH  SR-LP
ELDEST [INHERITANCE 2]  PAOLINI, CHRISTOPHER  H
ELDRIE THE HEALER [@BASTARD PRINCESS #1]  EDWARDS, CLAUDIA J.  P
ELDRITCH BLUE  O’BRIEN, KEVIN L.  LPA
ELECTION DAY 2084  ASIMOV, ISAAC/GREENBERG, MARTIN H.  HA
ELECTRIC CROCODILE, THE  COMPTON, D. G.  P
ELECTRIC FOREST  LEE, TANITH  H:2,P
ELECTRIC GENE MACHINE, THE [ORGANIC FUTURE]  EASTON, THOMAS A.  LP
ELECTRIC SWORD-SWALLOWERS, THE  BULMER, KENNETH  D05595
ELECTRONIC LULLABY MEAT MARKET, THE  SPENCER, JOHN  P
ELECTRONIC MIND READER, THE [@RICK BRANT #12]  BLAINE, JOHN  H:2

ELEGANT NIGHTMARES:

THE ENGLISH GHOST STORY FROM LE FANU TO BLACKWOOD SULLIVAN, JACK

R/REF-H

ELEGY FOR A LOST STAR [SYMPHONY OF AGES 5]  HAYDON, ELIZABETH  H,P
ELEGY FOR ANGELS AND DOGS  WILLIAMS, WALTER JON  TD24
ELEKTRICITEIT ZONDER DYNAMO  GARDNER, JOHN W.  FP
ELEKTRON UNION  HUNT, GILL  SR-P
ELEKTRONICA  ENGLEBARDT, STANLEY L.  FP
ELEMENT 79  HOYLE, FRED  H:2,P
ELEMENT 115  DUKE, CYNTHIA J.  P
ELEMENT OF FIRE, THE [HIGHLANDER]  HENDERSON, JASON  P
ELEMENT OF FIRE, THE  WELLS, MARTHA  H
ELEMENTAL  SAVILE, STEVEN/ONTIS, ALTHEA  HA
ELEMENTALS, THE  LLYWELYN, MORGAN  H,P
ELEMENTALS, THE  MCDOWELL, MICHAEL  P
ELEPHANT, THE  MROZEK, SLAWOMIR  H
ELEPHANT AND THE KANGAROO, THE  WHITE, T. H.  H
ELEPHANT FOR ARISTOTLE, AN  DE CAMP, L. SPRAGUE  P
ELEPHANT SONG [CIRCUS WORLD 2]  LONGYEAR, BARRY B.  C/P:2,P
ELEPHANTASM  LEE, TANITH  P
ELEVEN GREAT HORROR STORIES  OWEN, BETTY M  PA
ELEVEN MILLION MILE HIGH DANCER, THE  HILL, CAROL  H:2,LP
ELEVENTH COMMANDMENT, THE  DEL REY, LESTER  C/P:2,P
11TH PLAGUE, THE  PETERS, L. T.  C/P,P,H
ELEVENTH SEMIANNUAL REPORT TO CONGRESS 1964  NASA  [C/RAN-H]
ELEVENTH TIGER, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  P

TITLE  AUTHOR  CODE
ELF DEFENCE  FRIESNER, ESTHER M.  C/P:2,P
ELF FANTASTIC  GREENBERG, MARTIN H.  PA
ELF MAGIC  GREENBERG, MARTIN H.  C/PA,PA
ELF QUEEN & THE KING [RUIN MIST]  STANEK, ROBERT  LP
ELF QUEEN & THE KING II [RUIN MIST]  STANEK, ROBERT  LP
ELF QUEEN OF SHANNARA, THE [SHANNARA HERITAGE 3]  BROOKS, TERRY  H:2,P
ELF QUEEN’S QUEST [RUIN MIST DARK PATH 1]  STANEK, ROBERT  LP
ELFIN SHIP, THE  BLAYLOCK, JAMES P.  C/P:2,P
ELFQUEST 1  PINI, WENDY/PINI, RICHARD  CX
ELFQUEST 1 (FIRST HALF)  PINI, WENDY/PINI, RICHARD  SCX
ELFQUEST 2  PINI, WENDY/PINI, RICHARD  CX
ELFQUEST 3  PINI, WENDY/PINI, RICHARD  CX
ELFQUEST 4  PINI, WENDY/PINI, RICHARD  CX
ELFQUEST: JOURNEY TO SORROW’S END  PINI, WENDY/PINI, RICHARD  C/P:2,LP
ELFSHADOW [@FORGOTTEN REALMS HARPERS #2]  CUNNINGHAM, ELAINE  C/P:2,P
ELFSONG [@FORGOTTEN REALMS HARPERS #8]  CUNNINGHAM, ELAINE  C/P,P
ELFSORROW [LEGENDS OF THE RAVEN 1]  BARCLAY, JAMES  LP
ELFSTONES OF SHANNARA, THE [SHANNARA 2]  BROOKS, TERRY  C/P,LP,H
ELFWIN  WRIGHT, S. FOWLER  H
ELFWOOD  ESTES, ROSE  P,SFWA-P
GRAND JUBILEE, THE [OZARK 2]  ELGIN, SUZETTE HADEN  C/P,SFWA,P,H
ELIDOR  GARNER, ALAN  C/P:2,P
ELIMINATION  CAMPBELL, JOHN W. JR.  B
ELIZABETH  HAMILTON, JESSICA  P
ELLA ENCHANTED  LEVINE, GAIL CARSON  LP
ELLEANDER MORNING  YULSMAN, JERRY  C/P,P,H
ELLERY QUEEN’S LOST MEN  SULLIVAN, ELEANOR  SFWA
ELLISON WONDERLAND  ELLISON, HARLAN  C/P,P,H
ELLUVON GIFT, THE  LANG, SIMON  C/P:2,P
ELMINSTER IN HELL [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  P
ELMINSTER IN MYTH DRANNOR [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  H

ELMINSTER: THE MAKING OF A MAGE [@FORGOTTEN REALMS ELMINSTER] GREENWOOD, ED

C/P,P,H

ELMINSTER’S DAUGHTER [@FORGOTTEN REALMS ELMINSTER]  GREENWOOD, ED  H
ELOISE [DUMAREST 12]  TUBB, E. C.  C/P:2,P
ELPHAME’S CHOICE  CAST, P. C.  LP
ELRIC AT THE END OF TIME [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
ELRIC OF MELNIBONE [ELRIC 1]  MOORCOCK, MICHAEL  P
ELRIC OF MELNIBONE (OMNIBUS EDITION) [ELRIC]  MOORCOCK, MICHAEL  P
ELRIC SAGA PART I, THE [ELRIC 1,2,3]  MOORCOCK, MICHAEL  H:2
ELRIC SAGA PART II, THE [ELRIC 4,5,6]  MOORCOCK, MICHAEL  H:2
ELRIC SAGA PART III, THE [ELRIC]  MOORCOCK, MICHAEL  H
ELSEWHERE [BORDERLANDS 5]  SHETTERLY, WILL  H:2,P
ELSEWHERE  WINDLING, TERRI/ARNOLD, MARK ALAN  C/PA,PA
ELSEWHERE AND ELSEWHEN  CONKLIN, GROFF  PA,C/PA
ELSEWHERE, ELSEWHEN, ELSEHOW  DEFORD, MIRIAM ALLEN  H
ELSEWHERE II  WINDLING, TERRI/ARNOLD, MARK ALAN  C/PA,PA
ELSEWHERE III  WINDLING, TERRI/ARNOLD, MARK ALAN  C/PA,PA
ELVEN STAR [@DEATH GATE #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H
ELVENBANE, THE [HALFBLOOD 1]  NORTON, ANDRE/LACKEY, MERCEDES  H:2,P
ELVENBLOOD [HALFBLOOD 2]  NORTON, ANDRE/LACKEY, MERCEDES  H:2
ELVENBORN [HALFBLOOD 3]  NORTON, ANDRE/LACKEY, MERCEDES  H:2
ELVENDUDE  SHEPHERD, MARK  P
ELVES AND THE OTTERSKIN, THE [ALFAR 2]  BOYER, ELIZABETH H.  C/P:2,P
ELVISSEY [DRYCO 4]  WOMACK, JACK  H:2,LP
ELYSIUM  CARLSON, WILLIAM K.  H,SFWA
ELYSIUM COMMISSION, THE  MODESITT, L. E. JR.  H
EMBASSY [BRION BAYARD 1]  LAUMER, KEITH  P
EMBEDDING, THE  WATSON, IAN  C/P,SFWA,P,H
EMBER FROM THE SUN  CANTER, MARK  H
EMBRACE, THE [WORLD OF DARKNESS VAMPIRE DIARY]  WEINBERG, ROBERT/REIN-HAGEN, MARK  SR-H
EMBRYO  CHARBONNEAU, LOUIS  P
EMERALD CAVERN, THE [MATHEW LEWIN 2]  GRAHAM, MITCHELL  P
EMERALD CITY OF OZ, THE [OZ #6]  BAUM, L. FRANK  H:2,@P,@C/P:2
EMERALD ELEPHANT GAMBIT, THE [@AGENT OF TERRA #3]  MADDOCK, LARRY  C/P:2,P
EMERALD EYES [CONTINUING TIME 1]  MORAN, DANIEL KEYS  P
EMERALD FIRE [@DEATHLANDS]  AXLER, JAMES  P
EMERALD HOUSE RISING  KERR, PEG  C/P:2,P
EMERALD MAGIC  GREELEY, ANDREW M.  PA

EMERALD SCEPTER, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #3] REID, THOMAS M.

P

EMERALD SEA [DRAGONS 2]  RINGO, JOHN  H
EMERGENCE  PALMER, DAVID R.  C/P:2,P
EMERGENCE  SAN SOUCI, ROBERT D.  P

The MITSFS Titledex  September 24, 2007  Page 110
TITLE  AUTHOR  CODE
EMILE AND THE DUTCHMAN  ROSENBERG, JOEL  C/P,P
EMISSARY [STAR TREK DS9]  DILLARD, J. M.  C/P:2,P
EMISSARY (2 CASSETTES) [STAR TREK DS9]  DILLARD, J. M.  SR-MM
EMOTIONAL CHEMISTRY [DOCTOR WHO EIGHTH DOCTOR]  FORWARD, SIMON A.  P
EMPEROR [@TIME’S TAPESTRY #1]  BAXTER, STEPHEN  H
EMPEROR [@ROBOTS IN TIME 5]  WU, WILLIAM F.  P
EMPEROR AND CLOWN [MAN OF HIS WORD 4]  DUNCAN, DAVE  C/P,P
EMPEROR AND THE MONSTER, THE [@PERRY RHODAN #107]  ACE  PA
EMPEROR FOR THE LEGION, AN [VIDESSOS 2]  TURTLEDOVE, HARRY  C/P:2,P
EMPEROR FU MANCHU  ROHMER, SAX  P
EMPEROR MAGE [@IMMORTALS (PIERCE) #3]  PIERCE, TAMORA  P
EMPEROR OF AMERICA, THE  ROHMER, SAX  H
EMPEROR OF ANSALON [@DRAGONLANCE VILLAINS #3]  NILES, DOUGLAS  P
EMPEROR OF DAWN [BASIL CASTELLAN 2]  WHITE, STEVE  P
EMPEROR OF DREAMS: A CLARK ASHTON SMITH BIBLIOGRAPHY  SIDNEY-FRYER, DONALD  R/IND
EMPEROR OF ERIDANUS, THE  BARBET, PIERRE  P
EMPEROR OF EVERYTHING, THE [@EMANCIPATOR #2]  ALDRIDGE, RAY  C/P,P
EMPEROR OF GONDWANALAND, THE  DI FILIPPO, PAUL  LP
EMPEROR OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
EMPEROR OF THE LAST DAYS, THE  GOULART, RON  P
EMPEROR, SWORDS, PENTACLES  GOTLIEB, PHYLLIS  C/P:2,P
EMPEROR’S FIST [@MAGIC: THE GATHERING LEGENDS TWO 2]  MCGOUGH, SCOTT  P
EMPERORS OF THE TWILIGHT [MOREAU 2]  SWANN, S. ANDREW  C/P:2,P
EMPEROR’S PLAGUE, THE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
EMPERY [TRIGON DISUNITY 3]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,SFWA-P,P
EMPHYRIO  VANCE, JACK  C/P:2,P
EMPHYRIO (VIE) [@VANCE INTEGRAL EDITION #20]  VANCE, JACK  SR-H
EMPIRE  PIPER, H. BEAM  C/P:2,P
EMPIRE  SIMAK, CLIFFORD D.  C/P:2,GN2.07
EMPIRE & ECOLITAN [ECOLITAN 2,3]  MODESITT, L. E. JR.  LP
EMPIRE 99 [STAR HAWKS 1]  GOULART, RON  C/P,P
EMPIRE BUILDERS [PRIVATEERS 2]  BOVA, BEN  P
EMPIRE DREAMS  MCDONALD, IAN  C/P:2,P
EMPIRE FROM THE ASHES [COLIN MACINTYRE 1,2,3]  WEBER, DAVID  H,SR-LP
EMPIRE OF BLOOD [@DRAGONLANCE MINOTAUR WARS #3]  KNAAK, RICHARD A.  H
EMPIRE OF BLOOD [@BLADE #23]  LORD, JEFFREY  P
EMPIRE OF BONES  WILLIAMS, LIZ  P
EMPIRE OF CHAOS  BULMER, H. K.  SR-P
EMPIRE OF DEATH [DOCTOR WHO PAST DOCTORS]  BISHOP, DAVID  P
EMPIRE OF DREAMS AND MIRACLES  CARD, ORSON SCOTT/OLEXA, KEITH  HA,LPA
EMPIRE OF DUST [@SEARCHERS #2]  WILLIAMSON, CHET  C/P,P
EMPIRE OF FEAR, THE  STABLEFORD, BRIAN M.  C/P,P
EMPIRE OF GLASS, THE [DOCTOR WHO MISSING ADVENTURES]  LANE, ANDY  C/P:2,P
EMPIRE OF ICE, THE [NEW ICE AGE 1]  MORAN, RICHARD  H
EMPIRE OF ICE CREAM, THE  FORD, JEFFREY  H
EMPIRE OF THE ANTS  WERBER, BERNARD  P
EMPIRE OF THE ANTS  WEST, LINDSAY  C/P,P
EMPIRE OF THE ANTS (AND OTHER STORIES), THE  WELLS, H. G.  P
EMPIRE OF THE ATOM  VAN VOGT, A. E.  C/P,P,H,DD242,SFWA
EMPIRE OF THE EAGLE  NORTON, ANDRE/SHWARTZ, SUSAN  P
EMPIRE OF THE EAST [EMPIRE OF THE EAST 1,2,3]  SABERHAGEN, FRED  H:2,SFWA-P,P
EMPIRE OF THE SENSELESS  ACKER, KATHY  H,LP
EMPIRE OF THE STARS, THE [@DRAGON THRONE #2]  BAIRD, ALISON  LP
EMPIRE OF TIME, THE [CHRONOPLANE WARS 3]  KILIAN, CRAWFORD  C/P,P
EMPIRE OF TWO WORLDS  BAYLEY, BARRINGTON J.  P
EMPIRE OF UNREASON [AGE OF UNREASON 3]  KEYES, J. GREGORY  LP
EMPIRE STAR  DELANY, SAMUEL R.  DM139
EMPIRE STONE, THE  BUNCH, CHRIS  P
EMPIRE STRIKES BACK, THE [STAR WARS MOVIE 5]  GLUT, DONALD F.  H:2,P
EMPIRE STRIKES BACK COLLECTORS EDITION  LUCASFILM LTD.  A
EMPIRE STRIKES BACK NOTEBOOK, THE  ATTIAS, DIANA/SMITH, LINDSAY  R/REF-H
EMPIRE STRIKES BACK STORYBOOK  STENEMAN, SHEP  SR-VLH

EMPIRE STRIKES BACK, THE:

NATIONAL PUBLIC RADIO DRAMATIZATION, THE [STAR WARS] DALEY, BRIAN

LP

EMPIRE STRIKES BACK THE SCREENPLAY, THE [STAR WARS MOVIE 5]  BRACKETT, LEIGH/KASDAN, LAWRENCE  SR-S
EMPIRE’S DAUGHTER [KYDAN 1]  BROWN, SIMON  C/P,P
EMPIRE’S END [STEN 8]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
EMPIRE’S HORIZON  BRIZZOLARA, JOHN  P
EMPIRE’S LEGACY, THE [@CONCORDAT #1]  SWYCAFFER, JEFFERSON P.  C/P,P
EMPIRES OF FLUX AND ANCHOR [@SOUL RIDER #2]  CHALKER, JACK L.  C/P:2,P
EMPRESS AND THE ACOLYTE, THE [LYREMOUTH 3]  FLETCHER, JANE  LP
EMPRESS OF EARTH, THE [SILENCE LEIGH 3]  SCOTT, MELISSA  C/P:2,SFWA-P,P
EMPRESS OF LIGHT [SHANJI INTERSTELLAR EMPIRE 2]  GLASS, JAMES C.  C/P,P
EMPRESS OF OUTER SPACE  CHANDLER, A. BERTRAM  DM129

TITLE  AUTHOR  CODE
EMPRESS OF THE ENDLESS DREAM [@OROKON #5]  ARDEN, TOM  P
EMPRESS UNBORN [SAVAGE EMPIRE 7]  LORRAH, JEAN  C/P,P
EMPRISE [TRIGON DISUNITY 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
EMPTY CHAIR, THE [STAR TREK RIHANNSU 5]  DUANE, DIANE  P
EMPTY CITIES OF THE FULL MOON  HENDRIX, HOWARD V.  LP
EMPTY PEOPLE, THE  O’DONNELL, K. M.  C/P:2,P
EMPTY THRONE, THE [@XENA]  EMERSON, RU  P
EMSHWILLER: INFINITY X TWO  ORTIZ, LUIS  R/REF-H
ENCHANTED, THE  COATSWORTH, ELIZABETH  P
ENCHANTED CUP, THE  ROBERTS, DOROTHY JAMES  H
ENCHANTED FORESTS  KERR, KATHARINE/GREENBERG, MARTIN H. C/PA,PA
ENCHANTED, INC. [KATIE CHANDLER 1]  SWENDSON, SHANNA  LP
ENCHANTED ISLES, THE [@GODS OF IRELAND #2]  FLYNN, CASEY  P
ENCHANTED PILGRIMAGE  SIMAK, CLIFFORD D.  H:2,P
ENCHANTED PLANET, THE  BARBET, PIERRE  P
ENCHANTER  BAILEY, ROBIN WAYNE  C/P:2,P
ENCHANTER [WAYFARER REDEMPTION 2]  DOUGLASS, SARA  H

ENCHANTER COMPLEATED, THE (ORIGINALLY WALL OF SERPENTS)

[INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/PRATT, FLETCHER

SR-P

ENCHANTER COMPLETED, THE  TURTLEDOVE, HARRY  C/PA,PA

ENCHANTER REBORN, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/

STASHEFF, CHRISTOPHER

PA,C/PA,SFWA-P

ENCHANTERS’ END GAME [BELGARIAD 5]  EDDINGS, DAVID  C/P:2,P
ENCHANTMENT  CARD, ORSON SCOTT  H
ENCHANTMENT  VALLEJO, DORIS/VALLEJO, BORIS  A
ENCHANTMENTS OF FLESH AND SPIRIT, THE [WRAETHTHU 1]  CONSTANTINE, STORM  SFWA-P
ENCHANTRESS, THE  PARKER, EDGAR  H
ENCHANTRESS, THE  SUYIN, HAN  P
ENCHANTRESS FROM THE STARS  ENGDAHL, SYLVIA LOUISE  LP
ENCHANTRESS OF WORLD’S END, THE [GONDWANE 2]  CARTER, LIN  C/P,P
ENCORES IN FADE  LAMBERT, W. III  P
ENCOUNTER [@DEATHLANDS]  AXLER, JAMES  P
ENCOUNTER  HOLLY, JOAN HUNTER  C/P,P,H
ENCOUNTER AT FARPOINT [STAR TREK TNG]  GERROLD, DAVID  C/P:2,P
ENCOUNTER NEAR VENUS  WIBBERLEY, LEONARD  H:2
ENCOUNTER PROGRAM  ENSTROM, ROBERT  H,SFWA
ENCOUNTER THREE  CAIDIN, MARTIN  C/P,P
ENCOUNTER WITH TIBER  ALDRIN, BUZZ/BARNES, JOHN  P
ENCYCLOPEDIA OF ANCIENT AND FORBIDDEN KNOWLEDGE, THE  ZOLAR  [C/RAN-P]
ENCYCLOPEDIA OF FANTASY, THE  CLUTE, JOHN/GRANT, JOHN  R/REF-H
ENCYCLOPEDIA OF SCIENCE FICTION, THE  CLUTE, JOHN/NICHOLLS, PETER  R/REF-H
ENCYCLOPEDIA OF SCIENCE FICTION  HOLDSTOCK, ROBERT P.  R/REF-H

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 1:

WHO’S WHO, A–L, THE TUCK, DONALD H.

R/IND

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 2:

WHO’S WHO, M–Z, THE TUCK, DONALD H.

R/IND

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 3:

MISCELLANEOUS, THE TUCK, DONALD H.

R/IND

ENCYCLOPEDIA OF SCIENCE FICTION MOVIES, THE  HARDY, PHIL  R/REF-H
ENCYCLOPEDIA OF XANTH  NYE, JODY LYNN  P,SFWA-P
ENCYCLOPEDIE DE L’UTOPIE ET DE LA SCIENCE FICTION  VERSINS, PIERRE  R/REF-H
END AS A HERO  LAUMER, KEITH  C/P,P
END BRINGERS, THE  MASON, DOUGLAS R.  C/P:2,P
END GAME [@STAR TREK NEW FRONTIER #4]  DAVID, PETER  C/P,P

END GAMES [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #4] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

END IS COMING, THE [@SURVIVALIST #8]  AHERN, JERRY  P
END OF AFFLUENCE, THE  EHRLICH, PAUL R.  [C/RAN-P]
END OF ALL ROADS, THE [@OUTREMER #6]  BRENCHLEY, CHAZ  P
END OF ALL SONGS, THE [DANCERS AT THE END OF TIME 3]  MOORCOCK, MICHAEL  H:2,P,SFWA
END OF AN ERA  SAWYER, ROBERT J.  C/P:2,LP
END OF DAYS [CIRCUIT OF HEAVEN 2]  DANVERS, DENNIS  C/P,P
END OF ETERNITY, THE  ASIMOV, ISAAC  C/P:2,P
END-OF-EVERYTHING MAN, THE [@KING’S TRAMP #2]  DE HAVEN, TOM  C/P,P,H
END OF FAME, THE  ADAMS, BILL/BROOKS, CECIL  P
END OF MAN?, THE  JOHANNESSON, OLOF  P
END OF SUMMER, THE  MALZBERG, BARRY N./PRONZINI, BILL  C/PA,PA
END OF THE BEGINNING  TURTLEDOVE, HARRY  H
END OF THE CIRCLE  GOMERY, PERCY  H
END OF THE CIRCLE, THE [@ROBOTECH #18]  MCKINNEY, JACK  C/P,P
END OF THE DREAM, THE  WYLIE, PHILIP  C/P,P,H
END OF THE DREAMS, THE  GUNN, JAMES E.  H:2
END OF THE EMPIRE, THE  GILLILAND, ALEXIS A.  C/P,P
END OF THE GAME, THE [TRUE GAME 4,5,6]  TEPPER, SHERI S.  H
END OF THE JOURNEY  VERNE, JULES  H

The MITSFS Titledex  September 24, 2007  Page 112
TITLE  AUTHOR  CODE
END OF THE MATTER, THE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
END OF THE STORY, THE  SMITH, CLARK ASHTON  H
END OF THE THIRD WORLD, THE [LOST WORLD 2]  SOUZA, MARCIO  H,LP
END OF THE WORLD  OWEN, DEAN  P
END OF THE WORLD, THE  WOLLHEIM, DONALD A.  PA,C/PA
END OF THE WORLD NEWS, THE  BURGESS, ANTHONY  H
END OF THIS DAY’S BUSINESS, THE  BURDEKIN, KATHERINE  LP

END RUN [@WING COMMANDER 2] STASHEFF, CHRISTOPHER/ FORSTCHEN, WILLIAM R.

P,SFWA-P

ENDANGERED SPECIES [@ANGEL]  HOLDER, NANCY/MARIOTTE, JEFF  P
ENDANGERED SPECIES  WOLFE, GENE  C/P:2,SFWA-P,P
ENDER’S GAME (ORIGINAL) [ENDER 1]  CARD, ORSON SCOTT  P
ENDER’S GAME (REVISED 1991 EDITION) [ENDER 1]  CARD, ORSON SCOTT  C/P:3,SFWA-P,P
ENDER’S SHADOW [ENDER 1.1]  CARD, ORSON SCOTT  H:2,P

ENDER’S WAR (CONTAINS ORIGINAL NOVEL VERSION OF ENDER’S GAME)

[ENDER 1,2] CARD, ORSON SCOTT

H:2

ENDGAME [@DOOM #4]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P:2,P
ENDGAME [STAR TREK VOYAGER]  CAREY, DIANE  LP
ENDGAME [@MEROVINGEN NIGHTS #7]  CHERRYH, C. J.  C/PA,PA
ENDGAME [@BATTLETECH]  COLEMAN, LOREN L.  P
ENDGAME [DOCTOR WHO EIGHTH DOCTOR]  DICKS, TERRANCE  P
ENDGAME [@LOST SWORDS #7,8]  SABERHAGEN, FRED  H:2
ENDGAME (PROOF) [JANI KILIAN 5]  SMITH, KRISTINE  SR-P
ENDGAME ENIGMA  HOGAN, JAMES P.  C/P,P,H
ENDITHOR’S DAUGHTER [@RED SONJA #4]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
ENDLESS FRONTIER, THE  POURNELLE, JERRY  PA
ENDLESS FRONTIER II, THE  POURNELLE, JERRY/CARR, JOHN F.  C/PA,PA
ENDLESS KNOT, THE [SONG OF ALBION #3]  LAWHEAD, STEPHEN R.  H,@P
ENDLESS SHADOW  BRUNNER, JOHN  DF299
ENDLESS THINGS [AEGYPT]  CROWLEY, JOHN  H
ENDLESS UNIVERSE  BRADLEY, MARION ZIMMER  C/P:2,P
ENDLESS VOYAGE  BRADLEY, MARION ZIMMER  C/P:2,P
ENDLOSE ALPTRAUM, DER [@PERRY RHODAN #67]  VLCEK, ERNST  FPR
ENDS  DICKSON, GORDON R.  C/P:2,SFWA-P,P
ENDS OF THE CIRCLE, THE [PELBAR 2]  WILLIAMS, PAUL O.  C/P:2,P
ENDS OF THE EARTH, THE  SHEPARD, LUCIUS  P
ENDS OF TIME, THE  SILVERBERG, ROBERT  C/PA,PA
ENDURANCE [STARDOC 3]  VIEHL, S. L.  C/P,P
ENDYMION [HYPERION 3]  SIMMONS, DAN  H:2,P
ENDYMION SPRING  SKELTON, MATTHEW  H
ENEMIES  HOGAN, LEE  P
ENEMIES FROM BEYOND [INVADERS 2]  LAUMER, KEITH  P
ENEMIES OF THE SYSTEM  ALDISS, BRIAN W.  P,SFWA
ENEMY AT GREEN KNOWE, AN [GREEN KNOWE 5]  BOSTON, L. M.  H,LP
ENEMY GLORY [ENEMY GLORY 1]  MICHALSON, KAREN  H,P
ENEMY IN THE DARK [@PERRY RHODAN #85]  ACE  PA
ENEMY LINES [@STAR WARS NEW JEDI ORDER ENEMY LINES 1,2]  ALLSTON, AARON  H
ENEMY MINE  LONGYEAR, BARRY B./GERROLD, DAVID  TD6,SFWA-TD6,P
ENEMY OF MY ENEMY, THE  DAVIDSON, AVRAM  C/P:2,P
ENEMY OF MY ENEMY [@STAR TREK VOYAGER SPIRIT WALK 2]  GOLDEN, CHRISTIE  P
ENEMY OF MY ENEMY [TERRA NOVA]  OHLANDER, BEN/DRAKE, DAVID  C/P,P
ENEMY OF THE STATE, AN [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,SFWA,P,H
ENEMY PAPERS, THE  LONGYEAR, BARRY B.  H
ENEMY REBORN, AN [REALMS OF CHAOS 2]  STACKPOLE, MICHAEL A.  P
ENEMY STARS, THE  ANDERSON, POUL  SR-H,H:2,P
ENEMY TERRITORY [@STAR TREK I.K.S. GORKON 3]  DECANDIDO, KEITH R. A.  P
ENEMY UNSEEN [STAR TREK]  MITCHELL, V. E.  C/P:3,P
ENEMY WITHIN, THE [@WARHAMMER]  BYERS, RICHARD LEE  P
ENEMY WITHIN, THE [@PRIVATE SCHOOL #5]  CHARLES, STEVEN  P
ENEMY WITHIN, THE [@RAVENLOFT]  GOLDEN, CHRISTIE  P
ENEMY WITHIN, THE [@MISSION EARTH #3]  HUBBARD, L. RON  H,SFWA-P,P
ENEMY WITHIN THE SKULL [@CAP KENNEDY #4]  KERN, GREGORY  P
ENERGY ZERO [@BIG BRAIN #3]  BRANDNER, GARY  P
ENFORCER, THE [@ENFORCER #1]  SUGAR, ANDREW  P
ENGAGING THE ENEMY [VATTA’S WAR 3]  MOON, ELIZABETH  H:2
ENGINE CITY [ENGINES OF LIGHT 3]  MACLEOD, KEN  H
ENGINE OF RECALL, THE  SCHROEDER, KARL  H
ENGINE SUMMER  CROWLEY, JOHN  C/P,SFWA,P,H
ENGINEER, THE  ASHER, NEAL  LP
ENGINEER, THE  FURNAS, C. C./MCCARTHY, JOE  [C/RAN-H]
ENGINEER RECONDITIONED, THE  ASHER, NEAL  LP
ENGINES  DE CAMP, L. SPRAGUE  [C/RAN-P]
ENGINES OF DAWN, THE  COOK, PAUL  C/P,P
ENGINES OF DESTINY [@STAR TREK]  DEWEESE, GENE  P

TITLE  AUTHOR  CODE
ENGINES OF GOD, THE [PRISCILLA HUTCHINS 1]  MCDEVITT, JACK  C/P:2,P
ENGINES OF THE NIGHT, THE  MALZBERG, BARRY N.  SFWA,R/REF-H,C/REF-H
ENGLAND SWINGS SF  MERRIL, JUDITH  C/PA,PA
ENGLISH ASSASSIN, THE [CORNELIUS 3]  MOORCOCK, MICHAEL  H
ENGLISH WAY OF DEATH, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
ENIGMA [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
ENIGMA [TRIGON DISUNITY 2]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
ENIGMA FROM TANTALUS  BRUNNER, JOHN  DM115
ENIGMA VARIATIONS, THE  ROBERTS, JOHN MADDOX  P
ENLIGHTENMENT [DOCTOR WHO]  CLEGG, BARBARA  SR-P,C/P:2
ENNEAD, THE  MARK, JAN  P
ENORMOUS EGG, THE  BUTTERWORTH, OLIVER  P
ENORMOUS HOURGLASS, THE  GOULART, RON  P
ENORMOUS ROOM, THE  GOLD, HORACE L./KREPPS, ROBERT W.  B
ENORMOUS SHADOW, THE  HARLING, ROBERT  H
ENQUIRIES OF DOCTOR ESZTERHAZY, THE  DAVIDSON, AVRAM  P
ENSIGN FLANDRY [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P
ENSLAVED BRAINS  BINDER, EANDO  H
ENTANGLEMENT [@STARGATE ATLANTIS #6]  WELLS, MARTHA  P
ENTER THE ENCHANTED [@EVERWORLD #3]  APPLEGATE, K. A.  P
ENTERING TENEBREA [@TENEBREA #1]  DAWSON, ROXANN/GRAHAM, DANIEL  P
ENTERPRISE [@STAR TREK MY BROTHER’S KEEPER #3]  FRIEDMAN, MICHAEL JAN  P
ENTERPRISE STARDUST [@PERRY RHODAN #1]  ACE  PA,C/PA
ENTERPRISE: THE EXPANSE [STAR TREK ENTERPRISE]  DILLARD, J. M.  LP
ENTERPRISE: THE FIRST ADVENTURE [STAR TREK]  MCINTYRE, VONDA N.  C/P:3,P
ENTERTAINMENT  BUDRYS, ALGIS  LP
ENTITIES  RUSSELL, ERIC FRANK  H
ENTITY, THE  DE FELITTA, FRANK  P
ENTOMBED  SMITH, GUY N.  P
ENTOVERSE [GIANTS 4]  HOGAN, JAMES P.  H:2,P
ENTROPIA: A COLLECTION OF UNUSUALLY RARE STAMPS  SCHEURER, CHRISTIAN LORENZ  A
ENTROPY EFFECT, THE [STAR TREK]  MCINTYRE, VONDA N.  C/P:3,P
ENTRY, THE  HOLZER, HANS  P
ENTRY TO ELSEWHEN  BRUNNER, JOHN  C/P:2,P
ENVISAGED WORLDS  COLLINS, PAUL  SR-HA
ENVOY TO NEW WORLDS [RETIEF]  LAUMER, KEITH  C/P:2,DF223,P
ENVOY TO THE DOG STAR  SHAW, FREDERICK L. JR.  DG614
EON [EON 1]  BEAR, GREG  C/P,SFWA-P:2,H:2,P
EONS OF THE NIGHT [@HOWARD LIBRARY #5]  HOWARD, ROBERT E.  P
EOS READER 2004-2005, THE  EOS BOOKS  H:2,SR-LP
EPIC OF GILGAMESH, THE  SANDERS, N. K.  P
EPICUREAN, THE  MOORE, THOMAS  H
EPIDEMIC!  SLAUGHTER, FRANK G.  P
EPIDEMIC CENTER: ARALON [@PERRY RHODAN #37]  ACE  C/PA,PA
EPIPHANY [@STAR TREK VULCAN’S SOUL #3]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  H
EPIPHANY  YERMAKOV, NICHOLAS  C/P,P
EPIPHANY OF THE LONG SUN [LONG SUN 3,4]  WOLFE, GENE  H
EPISODES BEFORE THIRTY  BLACKWOOD, ALGERNON  SR-H
EPITAPH IN RUST  POWERS, TIM  SR-P
EPITAPH IN RUST, AN  POWERS, TIM  SR-H
EPOCH  ELWOOD, ROGER/SILVERBERG, ROBERT  C/PA,PA
EQUAL RITES [DISCWORLD 3/DISCWORLD WITCHES 1]  PRATCHETT, TERRY  C/P:2,P,H
EQUALITY  BELLAMY, EDWARD  SR-H
EQUALITY: IN THE YEAR 2000  REYNOLDS, MACK  C/P,P
EQUATIONS OF CHAOS [@WEAPONS OF CHAOS #2]  VARDEMAN, ROBERT E.  P
EQUATOR AND SEGREGATION  ALDISS, BRIAN W.  P
EQUINOX [STAR TREK VOYAGER]  CAREY, DIANE  P
EQUINOX [@DESCENT 3]  TELEP, PETER  P
EQUINOX ZERO [@OUTLANDERS]  AXLER, JAMES  P
ER-DAN STORIES, THE [ER-DA]  REDMOND, JEFFREY  LP
ERAGON [INHERITANCE 1]  PAOLINI, CHRISTOPHER  H
ERASING SHERLOCK [FACTION PARADOX 5]  HALE, KELLY  LP
ERASMUS MAGISTER  SHEFFIELD, CHARLES  C/P,P
ERBE DER YULOCS, DAS [@PERRY RHODAN SILBERBANDE #71]  PABEL MOEWIG KG  FPR
ERDE GEHT NICHT UNTER, DIE  KOCH, R.  FH
EREWHON  BUTLER, SAMUEL  P
ERIC [DISCWORLD 9/DISCWORLD RINCEWIND 4]  PRATCHETT, TERRY  C/P:3,SFWA-P,VLP
ERIC BRIGHTEYES [ERIC BRIGHTEYES 1]  HAGGARD, H. RIDER  C/P,LP
ERIC JOHN STARK: OUTLAW OF MARS  BRACKETT, LEIGH  P
ERIC OF ZANTHODON [ERIC CARSTAIRS 5]  CARTER, LIN  P
ERIDAHN  YOUNG, ROBERT F.  C/P,P
ERLING’S WORD  WALKER, LARS  P
ERNST ELLERT RETURNS! [@PERRY RHODAN #83]  ACE  PA
EROS ASCENDING [VELVET COMET #1]  RESNICK, MIKE  C/P,@P,H

TITLE  AUTHOR  CODE
EROS AT NADIR [VELVET COMET #4]  RESNICK, MIKE  C/P,@P
EROS AT ZENITH [VELVET COMET #2]  RESNICK, MIKE  C/P:2,@P,H
EROS DESCENDING [VELVET COMET #3]  RESNICK, MIKE  C/P,@P
EROS IN ORBIT  ELDER, JOSEPH  HA,PA,SFWA
ERRAND OF MERCY [STAR TREK]  COON, GENE L.  SR-S

ERRAND OF MERCY [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #4] MOORE, ROGER E.

P

ERROR DE LA MONTANA, EL  GILMAR, XANDRA  FP
ERSATZ NATION  KENYON, TIM  LP
ERSTE BUCH DER SCHWERTER, DAS  SABERHAGEN, FRED  FP
ERTHRING CYCLE, THE [ERTHRING 1,2,3]  DREW, WAYLAND  H:2
ES, ROBOTS (LATVIAN)  ASIMOV, ISAAC  FH
ESBAE: A WINTER’S TALE  HALDEMAN, LINDA  P
ESCALA NO TEMPO  HEINLEIN, ROBERT A.  FP
ESCALATION [@GUNDAM MOBILE SUIT #2]  TOMINO, YOSHIYUKI  C/P:2,P
ESCAPE!  BOVA, BEN  SR-H,P

ESCAPE, THE [STAR TREK VOYAGER] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

ESCAPE ACROSS THE COSMOS  FOX, GARDNER F.  C/P:2,P
ESCAPE CARTHUS  BROWNE, C. J.  P
ESCAPE FROM CHINA [TRIO]  MONTGOMERY, R. A.  P
ESCAPE FROM EARTH  DANN, JACK/DOZOIS, GARDNER  HA
ESCAPE FROM FLATLAND [FLATLAND]  HAVERSINE, THEDA  SR-P
ESCAPE FROM KATHMANDU  ROBINSON, KIM STANLEY  C/P,SFWA-P,P,H
ESCAPE FROM LOKI [DOC SAVAGE]  FARMER, PHILIP JOSE  P
ESCAPE FROM MACHO [@SPACEWAYS #3]  CLEVE, JOHN  C/P:2,P
ESCAPE FROM NEW YORK  MCQUAY, MIKE  P
ESCAPE FROM ROBOTROPOLIS [ROBOT ODYSSEY 1]  D’IGNAZIO, FRED  H:2
ESCAPE FROM ROKSAMUR [BARD’S TALE 4]  SHEPHERD, MARK  P

ESCAPE FROM TERROR LAGOON [@RETURN TO THE PLANET OF THE APES #2] ARROW, WILLIAM

P

ESCAPE FROM THE PLANET OF THE APES  POURNELLE, JERRY  C/P:2,P
ESCAPE FROM TOMORROW  LANNING, SERETA  SR-P
ESCAPE FROM UNDERMOUNTAIN [@FORGOTTEN REALMS NOBLES #3]  ANTHONY, MARK  P
ESCAPE INTO SPACE  TUBB, E. C.  H
ESCAPE ON VENUS [VENUS #4]  BURROUGHS, EDGAR RICE  C/P:2,@P
ESCAPE ORBIT, THE  WHITE, JAMES  C/P:2,P
ESCAPE PLUS  BOVA, BEN  C/P,P
ESCAPE TO EARTH  HOWARD, IVAN  PA,C/PA
ESCAPE TO INFINITY  ZEIGFREID, KARL  SR-P
ESCAPE TO THE WIND  DIMARCO, JENNIFER  LP
ESCAPE TO TOMORROW [@PLANET OF THE APES #2]  EFFINGER, GEORGE ALEC  C/P,P
ESCAPE TO VENUS [@PERRY RHODAN #15]  ACE  PA,C/PA
ESCAPE TO WITCH MOUNTAIN  KEY, ALEXANDER  C/P,P
ESCAPE VELOCITY [DOCTOR WHO EIGHTH DOCTOR]  BRAKE, COLIN  P
ESCAPE VELOCITY [WARLOCK 0]  STASHEFF, CHRISTOPHER  C/P:2,P
ESCAPES OF MR. HONEY, THE  ATKEY, BERTRAM  H
ESKIMO INVASION, THE  HOWARD, HAYDEN  C/P:2,P
ESP ORGY  WILLIAMS, J. X.  SR-L
E.S.P. WORM, THE  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:4,SFWA-P,P
ESP: YOUR SIXTH SENSE  STEIGER, BRAD  [C/RAN-P:2]
ESPACE NOIR, L  BRUSS, B. R.  FP
ESPER  BLISH, JAMES  C/P:2,P
ESPER GREIFEN EIN, DIE  MAGAZINE OF F & SF  FP
ESPER TRANSFER, THE  PROCTOR, GEO. W.  P
ESPIONAGE  DOXEY, WILLIAM S.  P
ESSENCE OF EVIL, THE [ANGEL’S LUCK 3]  FAUST, JOE CLIFFORD  H,P
ESSENTIAL BORDERTOWN, THE [BORDERLAND]  WINDLING, TERRI/SHERMAN, DELIA  LP
ESSENTIAL CLIVE BARKER, THE  BARKER, CLIVE  H
ESSENTIAL ELLISON, THE (REVISED, EXPANDED)  ELLISON, HARLAN  LP
ESSENTIAL HAL CLEMENT 1, THE: TRIO FOR SLIDE RULE & TYPEWRITER  CLEMENT, HAL  H
ESSENTIAL HAL CLEMENT 2, THE: MUSIC OF MANY SPHERES  CLEMENT, HAL  H

ESSENTIAL HAL CLEMENT 3, THE:

VARIATIONS ON A THEME BY SIR ISAAC NEWTON CLEMENT, HAL

H

ESTABLISHMENT IS ALIVE & WELL IN WASHINGTON, THE  BUCHWALD, ART  [C/RAN-P]
E.T. 101  MISSION CONTROL/JHO, ZOEV  H
E.T.: THE BOOK OF THE GREEN PLANET  KOTZWINKLE, WILLIAM  C/P:2,P
E.T.: THE EXTRA-TERRESTRIAL  KOTZWINKLE, WILLIAM  C/P,P,H
E.T.: THE EXTRA-TERRESTRIAL STORYBOOK  KOTZWINKLE, WILLIAM  H
ET TU, BABE  LEYNER, MARK  SFWA
ETCHED CITY, THE  BISHOP, K. J.  LP
ETERNAL CHAMPION, THE [EREKOSE 1/ETERNAL CHAMPION 1]  MOORCOCK, MICHAEL  C/P,SFWA,SFWA-P,P
ETERNAL CHAMPION, THE (OMNIBUS) [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  LP
ETERNAL CITY, THE  DRAKE, DAVID  P,SFWA-P:2

TITLE  AUTHOR  CODE
ETERNAL ENEMY, THE  BERLYN, MICHAEL  H,SFWA
ETERNAL FLAME [@GREAT TREE OF AVALON #3]  BARRON, T. A.  H
ETERNAL FOOTMAN, THE [CORPUS DEI 3]  MORROW, JAMES  H
ETERNAL FRONTIER  SCHMITZ, JAMES H.  LP
ETERNAL FRONTIERS, THE  SCHMITZ, JAMES H.  H:2,P
ETERNAL ICE, THE [@MAGIC: THE GATHERING ICE AGE 2]  GRUBB, JEFF  P
ETERNAL LIFE  NEWLOVE, DONALD  LP
ETERNAL LIGHT  MCAULEY, PAUL J.  C/P,SFWA,P,H
ETERNAL MERCENARY, THE [@CASCA #1]  SADLER, BARRY  C/P,P
ETERNAL SAVAGE, THE  BURROUGHS, EDGAR RICE  C/P:2,P
ETERNALLY BAD  ROBBINS, TRINA  H
ETERNITY [EON 2]  BEAR, GREG  SFWA-P,H:2,P
ETERNITY  REYNOLDS, MACK/ING, DEAN  C/P:2,P
ETERNITY ARTIFACT, THE  MODESITT, L. E. JR.  H
ETERNITY BRIGADE, THE  GOLDIN, STEPHEN  C/P,P
ETERNITY BRIGADE, THE (PROOF)  GOLDIN, STEPHEN  SR-P
ETERNITY ROAD  MCDEVITT, JACK  H
ETERNITY ROW [@STARDOC]  VIEHL, S. L.  C/P,P
ETERNITY STONE, THE  ROMINE, ADEN FOSTER/ROMINE, MARY COX P
ETERNITY UNBOUND [@SPACE: 1999]  LATHAM, WILLIAM  LP
ETERNITY WEEPS [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P,P
ETERNITY’S END  CARVER, JEFFREY A.  H:2
ETHAN OF ATHOS  BUJOLD, LOIS MCMASTER  C/P,P,H
ETHER ORE  TURK, H. C.  P,SFWA-P
ETHOS EFFECT, THE  MODESITT, L. E. JR.  P
ETIDORHPA  LLOYD, JOHN URI  SR-H,P
ETOILE DU NEANT, L  BARBET, PIERRE  FP
ETRANGE PLANETE ORGA, L  BRUSS, B. R.  FP
ETRUSCANS  LLYWELYN, MORGAN/SCOTT, MICHAEL  H,P
EUDAEMONIC PIE, THE  BASS, THOMAS  [C/RAN-H]
EUGENICS WARS, VOLUME 1 [STAR TREK]  COX, GREG  P
EUGENICS WARS, VOLUME 2 [STAR TREK]  COX, GREG  H,P
EUNUCH’S HEIR, THE [DURHYS FAMILY 2]  ISAAK, ELAINE  LP
EUREKA YEARS, THE  MCCOMAS, ANNETTE PELTZ  PA
EUROPA STRIKE [@HERITAGE #3]  DOUGLAS, IAN  C/P,P
EURYALE  DALKEY, KARA  P
EVANGELIST, THE [CABAL 3]  DUNN, PHILIP  P
EVE OF MIDSUMMER, THE  SHACKLEFORD, JACK D.  SR-P
EVE OF THE MAELSTROM, THE [@DRAGONLANCE FIFTH AGE]  RABE, JEAN  P
EVELYN, THE AMBITIOUS [@PHENWICK #23]  KIMBROUGH, KATHERYN  P
EVEN MORE NASTY STORIES  MCNAUGHTON, BRIAN  LP
EVEN VAMPIRES GET THE BLUES  MACALISTER, KATIE  P
EVENING AT JOE’S, AN [HIGHLANDER]  HORVATH, GILLIAN  LPA
EVENING’S EMPIRE  HERTER, DAVID  H,LP
EVENOR  MACDONALD, GEORGE  C/P,P
EVENT HORIZON  MCDONALD, STEVEN E.  P
EVENT HORIZON [@BLACK HOLE TRAVEL AGENCY #1]  MCKINNEY, JACK  C/P:2,P
EVENTS IN SPACE  LEY, WILLY  [C/RAN-P:2]
EVER: THE WAR IN THE WASTE [EVER 1]  SAVAGE, FELICITY  LP
EVERARD’S RIDE  JONES, DIANA WYNNE  SR-H
EVERFREE [IDLEWILD 3]  SAGAN, NICK  H
EVERGENCE [EVERGENCE 1,2,3]  WILLIAMS, SEAN/DIX, SHANE  H
EVERGENCE: A DARK IMBALANCE [EVERGENCE 3]  WILLIAMS, SEAN/DIX, SHANE  P
EVERGENCE: THE DYING LIGHT [EVERGENCE 2]  WILLIAMS, SEAN/DIX, SHANE  P
EVERGENCE: THE PRODIGAL SUN [EVERGENCE 1]  WILLIAMS, SEAN/DIX, SHANE  P
EVERMEET: ISLAND OF ELVES [FORGOTTEN REALMS]  CUNNINGHAM, ELAINE  P
EVERVILLE [ART 2]  BARKER, CLIVE  H,LP
EVERY INCH A KING  TURTLEDOVE, HARRY  H
EVERY WHICH WAY BUT DEAD [RACHEL MORGAN 3]  HARRISON, KIM  C/P,P
EVERYBODY COMES TO COSMO’S [EXCHAMELEON 3]  GOULART, RON  C/P,P
EVERYBODY COMES TO THE NIGHTSIDE [NIGHTSIDE 1,2,3]  GREEN, SIMON R.  H
EVERYBODY’S FAVORITE DUCK  WILSON, GAHAN  C/P,P
EVERYTHING BUT LOVE  MIR PUBLISHERS  SR-PA
EVERYTHING’S EVENTUAL  KING, STEPHEN  H
EVE’S SECOND APPLE  DOGBOLT, BARNABY  H
EVIDENCE [GARRISON 2]  JOHNSON, KARMELA  LP
EVIL, THE  CAVE, HUGH B.  P
EVIL ABYSS [@OUTLANDERS]  AXLER, JAMES  P
EVIL ASCENDING [DARK CONSPIRACY 2]  STACKPOLE, MICHAEL A.  P
EVIL EARTHS  ALDISS, BRIAN W.  C/PA,PA
EVIL, EVIL  DUBREUIL, LINDA  P
EVIL GNOME, THE [DOC SAVAGE #82]  ROBESON, KENNETH  C/P,@P
EVIL HOUR, AN [@TERMINATOR 2 NEW JOHN CONNOR #2]  BLACKFORD, RUSSELL  P
EVIL IS LIVE SPELLED BACKWARDS  OFFUTT, ANDREW J.  P

The MITSFS Titledex  September 24, 2007  Page 116
TITLE  AUTHOR  CODE
EVIL OF THE DALEKS [DOCTOR WHO]  PEEL, JOHN  P
EVIL ON THE BAYOU [TWILIGHT #21]  CUSICK, RICHIE TANKERSLEY  P
EVIL PEOPLE, THE  HAINING, PETER  PA
EVIL THAT MEN DO, THE  BRUNNER, JOHN  SR-DB1010
EVIL THAT MEN DO, THE [@BUFFY]  HOLDER, NANCY  C/P,P
EVIL TRIUMPHANT [DARK CONSPIRACY 3]  STACKPOLE, MICHAEL A.  P
EVIL UNDER THE WATER, THE  GURNEY, DAVID  P
EVOLUTION  BAXTER, STEPHEN  H:2
EVOLUTION [@STAR TREK VOYAGER STRING THEORY #3]  JARMAN, HEATHER  P
EVOLUTION [DOCTOR WHO MISSING ADVENTURES]  PEEL, JOHN  C/P:2,P
EVOLUTION IN ACTION  HUXLEY, JULIAN  [C/RAN-P]
EVOLUTION MAN, THE  LEWIS, ROY  SR-P
EVOLUTION VERSUS THE NEW WORLD  WATCHTOWER  [C/RAN-P]
EVOLUTION’S DARLING  WESTERFELD, SCOTT  LP
EVOLUTION’S SHORE [CHAGA 1]  MCDONALD, IAN  C/P,P,H
EX MACHINA [@STAR TREK]  BENNETT, CHRISTOPHER L.  P
EXALTATION OF STARS, AN  CARR, TERRY  PA,HA

EXCALIBUR GILLIAM, RICHARD/GREENBERG, MARTIN H./

KRAMER, EDWARD E.

HA,LPA

EXCALIBUR  LAUBENTHAL, SANDERS ANNE  C/P,P
EXCALIBUR ALTERNATIVE, THE [RANKS OF BRONZE]  WEBER, DAVID  H,P
EXCALIBUR: RENAISSANCE [@STAR TREK NEW FRONTIER #10]  DAVID, PETER  P
EXCALIBUR: REQUIEM [@STAR TREK NEW FRONTIER #9]  DAVID, PETER  P
EXCALIBUR: RESTORATION [@STAR TREK NEW FRONTIER #11]  DAVID, PETER  P
EXCESS OF ENCHANTMENTS, AN [WUNTVOR 2]  GARDNER, CRAIG SHAW  C/P:2,P
EXCESSION [CULTURE]  BANKS, IAIN M.  C/P,P,H
EXCHANGE OF GIFTS, AN  MCCAFFREY, ANNE  H,LP
EXCHANGE OF HOSTAGES, AN [JUDICIARY 1]  MATTHEWS, SUSAN R.  C/P,P
EXECUTION HOUR [WARHAMMER 40K]  RENNIE, GORDON  P
EXECUTIONER [@WORLD OF DARKNESS PREDATOR & PREY #6]  FLEMING, GHERBOD  P
EXECUTIVE [@BIO OF A SPACE TYRANT #4]  ANTHONY, PIERS  C/P:2,P
EXERCISE FOR MADMEN, AN  PAUL, BARBARA  C/P,P
EXERCISER, THE  WEINSTEIN, SOL/ALBRECHT, HOWARD  P
EXHALTATION OF LARKS, AN  REED, ROBERT  LP
EXILE [@STAR WARS LEGACY OF THE FORCE 4]  ALLSTON, AARON  H,P
EXILE  KUBE-MCDOWELL, MICHAEL P.  H,SFWA,SFWA-P:2
EXILE  LOGSTON, ANNE  P
EXILE [@PHOTON #5]  PETERS, DAVID  P
EXILE [@FORGOTTEN REALMS DARK ELF #2]  SALVATORE, R. A.  C/P,P
EXILE [@FIVE WORLDS #1]  SARRANTONIO, AL  P
EXILE AND OTHER TALES OF FANTASY  CUMMINGS, M. A.  P
EXILE AND THE SORCERER, THE [LYREMOUTH 1]  FLETCHER, JANE  LP
EXILE FROM XANADU  WRIGHT, LAN  DM103
EXILE KISS, THE [MARID AUDRAN 3]  EFFINGER, GEORGE ALEC  C/P,LP,H
EXILE OF ELLENDON, THE  MARDEN, WILLIAM  H,SFWA
EXILE OF TIME, THE  CUMMINGS, RAY  C/P:2,P
EXILE ON VLAHIL  MAYHAR, ARDATH  H,SFWA
EXILE WAITING, THE  MCINTYRE, VONDA N.  C/P,P,H
EXILE WAITING, THE (REVISED)  MCINTYRE, VONDA N.  C/P,P
EXILED FROM EARTH  BOVA, BEN  H:2,LP
EXILED IN SPACE  TORRO, PEL  H
EXILES [STAR TREK VULCAN’S SOUL 2]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  H
EXILES [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:2,P
EXILES AT THE WELL OF SOULS [WELL WORLD 2]  CHALKER, JACK L.  C/P:2,P
EXILE’S CHALLENGE [EXILES 2]  WELLS, ANGUS  LP
EXILE’S CHILDREN [EXILES 1]  WELLS, ANGUS  LP
EXILE’S GATE [MORGAINE 4]  CHERRYH, C. J.  H:2,P
EXILE’S HONOR [VALDEMAR]  LACKEY, MERCEDES  H,P
EXILES IN TIME  DEEGAN, JON J.  SR-P
EXILES OF COLSEC [COLSEC 1]  HILL, DOUGLAS  C/P,P,H
EXILES OF THE RYNTH  DOUGLAS, CAROLE NELSON  C/P:2,P
EXILES OF THE STARS [MOON SINGER 2]  NORTON, ANDRE  C/P,P,H
EXILES OF TIME  BOND, NELSON  C/P:2,P
EXILE’S QUEST  MEADE, RICHARD  P
EXILE’S RETURN [CONCLAVE OF SHADOWS 3]  FEIST, RAYMOND E.  H
EXILES’ RETURN [@GHATTI’S TALE #3]  GREENO, GAYLE  C/P,P
EXILE’S RETURN [@ELITA 1]  JACOBY, KATE  P
EXILE’S SONG [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
EXILES TO GLORY  POURNELLE, JERRY  C/P:2,SFWA-P,P
EXILES TO GLORY (REVISED)  POURNELLE, JERRY  P
EXILES TRILOGY, THE  BOVA, BEN  C/P,P
EXILE’S VALOR [VALDEMAR]  LACKEY, MERCEDES  H:2
EXILE’S VALOR (PROOF) [VALDEMAR]  LACKEY, MERCEDES  SR-LP
EXIT EARTH  CAIDIN, MARTIN  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 117
TITLE  AUTHOR  CODE
EXIT HUMANITY  BRETT, LEO  SR-P
EXIT INTO ETERNITY  EDDY, C. M. JR.  SR-LP
EXIT STRATEGY [INCONSTANT MOON 3]  ASKEGREN, PIERCE  P
EXIT TO REALITY  FORBES, EDITH  H
EXODUS [@HELLGATE LONDON #1]  ODOM, MEL  P
EXODUS [@STAR TREK VULCAN’S SOUL #1]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
EXODUS [STARFIRE 5]  WHITE, STEVE/MEIER, SHIRLEY  H
EXODUS FROM THE LONG SUN [LONG SUN 4]  WOLFE, GENE  C/P,P,H
EXODUS ROAD [@BATTLETECH #33/TWILIGHT OF THE CLANS 1]  PARDOE, BLAINE LEE  C/P,P

EXOGENESIS [@STARGATE ATLANTIS #5] WHITELAW, SONNY/

CHRISTENSEN, ELIZABETH

P

EXORCISM  NATAS, ETH  P
EXORCIST, THE [EXORCIST 1]  BLATTY, WILLIAM PETER  C/P:2,P
EXORCISTS, THE  MULLER, JOHN E.  SR-P

EXOTIC ENCHANTER, THE [INCOMPLETE ENCHANTER] DE CAMP, L. SPRAGUE/

STASHEFF, CHRISTOPHER

PA

EXPANDED UNIVERSE  HEINLEIN, ROBERT A.  C/P,H,LP
EXPATRIA  BROOKE, KEITH  LP
EXPECTED ONE, THE [BLOOD OF THE LAMB 1]  ROGERS, MARK E.  P
EXPECTING BEOWULF  HOLT, TOM  LP
EXPECTING SOMEONE TALLER  HOLT, TOM  C/P:3,SFWA-P:2,P
EXPEDITION  BARLOWE, WAYNE DOUGLAS  VLP
EXPEDITION TO EARTH  CLARKE, ARTHUR C.  C/P,P,H
EXPENDABLE [LEAGUE OF PEOPLES 1]  GARDNER, JAMES ALAN  H:3,P
EXPERIMENT, THE  DARNTON, JOHN  H
EXPERIMENT, THE  SETLOWE, RICHARD  H:2
EXPERIMENT IN TERRA [@BATTLESTAR GALACTICA 1978 #9]  LARSON, GLEN A./GOULART, RON  C/P,P
EXPERIMENT WITH TIME, AN  DUNNE, J. W.  [C/RAN-P]
EXPERIMENTS IN AUTOBIOGRAPHY  WELLS, H. G.  [C/RAN-H]
EXPERT DREAMERS, THE  POHL, FREDERIK  C/PA,PA
EXPIRATION DATE [FAULT LINES 1]  POWERS, TIM  H,P
EXPLOITS OF EBENEZUM, THE [EBENEZUM 1,2,3]  GARDNER, CRAIG SHAW  H
EXPLOITS OF MOOMINPAPPA, THE  JANSSON, TOVE  SR-P

EXPLOITS OF SHERLOCK HOLMES, THE DOYLE, ADRIAN CONAN/ CARR, JOHN DICKSON

P

EXPLORATION OF MARS, THE  LEY, WILLY/VON BRAUN, WERNHER  R/REF-H,C/REF-H
EXPLORATION OF SPACE, THE  CLARKE, ARTHUR C.  [C/RAN-H,C/RAN-P:2]
EXPLORATIONS  ANDERSON, POUL  C/P:2,P
EXPLORER [ATEVI 6]  CHERRYH, C. J.  H:2
EXPLORER  CZERNEDA, JULIE E.  LPA
EXPLORERS  DOZOIS, GARDNER  LPA
EXPLORERS  GIPE, GEORGE  P
EXPLORERS, THE  KORNBLUTH, C. M.  C/P:2,P
EXPLORERS IN HELL [HEROES IN HELL]  DRAKE, DAVID/MORRIS, JANET  P,SFWA-P
EXPLORERS INTO INFINITY  CUMMINGS, RAY  H
EXPLORERS OF GOR [GOR 13]  NORMAN, JOHN  C/P:2,P
EXPLORERS OF SPACE  SILVERBERG, ROBERT  HA
EXPLORERS OF THE INFINITE  MOSKOWITZ, SAM  R/REF-H
EXPLORING CORDWAINER SMITH  ALGOL PRESS  R/REF-P
EXPLORING FANTASY WORLDS  SCHWEITZER, DARRELL  R/REF-P
EXPLORING OTHER WORLDS  MOSKOWITZ, SAM  PA,C/PA
EXPLORING THE OCCULT  HUNT, DOUGLAS  [C/RAN-P]
EXPLORING THE UNIVERSE (1987)  CALENDAR  CAL
EXTERMINATOR!  BURROUGHS, WILLIAM S.  LP
EXTINCTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #4]  SMEDMAN, LISA  P
EXTINCTION BOMBER  HOUGH, S. B.  H
EXTRA DAY, THE  BLACKWOOD, ALGERNON  SR-H
EXTRA MAN, THE  GREY, CHARLES  SR-P
EXTRA MAN, THE  TUBB, E. C.  LP
EXTRA(ORDINARY) PEOPLE  RUSS, JOANNA  H,LP
EXTRACURRICULAR ACTIVITIES [BIONIC WOMAN 2]  LOTTMAN, EILEEN  C/P:2,P
EXTRAPOLASIS  MALEC, ALEXANDER  C/P:3,P
EXTRATERRESTRIAL  SHOCK, JULIAN  C/P,P
EXTRATERRESTRIAL CIVILIZATIONS  TOVMASYAN, G. M.  [C/RAN-H]
EXTRATERRITORIAL, THE  MORRESSY, JOHN  SR-P
EXTREME PREJUDICE [@ALIEN NATION #7]  GRAF, L. A.  C/P,P
EXTREMES [RETRIEVAL ARTIST 2]  RUSCH, KRISTINE KATHRYN  P
EXTRO  BESTER, ALFRED  P
EXULTANT [DESTINY’S CHILDREN 2]  BAXTER, STEPHEN  H
EXXONERATION  ROHMER, RICHARD  H
EYAS  KILIAN, CRAWFORD  C/P:2,P
EYE  HERBERT, FRANK  H:2,SFWA-P,P
EYE AMONG THE BLIND  HOLDSTOCK, ROBERT P.  C/P,SFWA,P,H
EYE FOR AN EYE: THE DOLL, AN [@BLACKSTONE #1]  SAUL, JOHN  P

The MITSFS Titledex  September 24, 2007  Page 118
TITLE  AUTHOR  CODE
EYE FOR EYE  CARD, ORSON SCOTT  SFWA-TD27,TD27
EYE IN THE PYRAMID, THE [ILLUMINATUS! 1]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P,P
EYE IN THE SKY  DICK, PHILIP K.  C/P:2,P
EYE IN THE STONE, THE  WOLD, ALLEN L.  P
EYE OF CAT  ZELAZNY, ROGER  H:2,SFWA,P
EYE OF CHARON, THE [@AGE OF CONAN SOLDIER’S QUEST #2]  KNAAK, RICHARD A.  P
EYE OF FORTUNE [@KNIGHTS OF THE SILVER DRAGON #4]  GRAHAM, DENISE B.  LP

EYE OF GEHENNA, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #2] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

EYE OF HEAVEN [DOCTOR WHO PAST DOCTORS]  MORTIMORE, JIM  C/P,P
EYE OF ISTAR, THE  LE QUEUX, WILLIAM  SR-H
EYE OF KARNAK, THE  MULLER, JOHN E.  SR-P
EYE OF NIGHT, THE  ALAMA, PAULINE J.  C/P,P
EYE OF TERROR [WARHAMMER 40K]  BAYLEY, BARRINGTON J.  P
EYE OF THE DAEMON [KEVIN BRADLEY AND EVAN DAVIS 1]  BACON-SMITH, CAMILLE  C/P,P
EYE OF THE EAGLE [GLOBAL 2000 1]  BASILE, GLORIA VITANZA  P
EYE OF THE GIANT, THE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P
EYE OF THE HERON, THE  LE GUIN, URSULA K.  C/P:2,P
EYE OF THE HUNTER, THE [MITHGAR 12]  MCKIERNAN, DENNIS L.  C/P,SFWA,LP
EYE OF THE LABYRINTH [SECOND SONS 2]  FALLON, JENNIFER  P
EYE OF THE LENS, THE  JONES, LANGDON  H,P
EYE OF THE MONSTER  NORTON, ANDRE  DF147,C/P
EYE OF THE QUEEN, THE  MANN, PHILLIP  H:2,P
EYE OF THE RAM, THE [@HERCULES THE LEGENDARY JOURNEYS]  BOGGS, TIMOTHY  C/P,P
EYE OF THE SERPENT [AELWYN 2]  CHARRETTE, ROBERT N.  C/P,P
EYE OF THE SUN  CONNER, MIKE  P
EYE OF THE TYGER, THE [DOCTOR WHO]  MCAULEY, PAUL J.  SR-H
EYE OF THE VULTURE [GYPSY 2/WEIRD HEROES 7]  GOULART, RON  P
EYE OF THE WORLD, THE [WHEEL OF TIME 1]  JORDAN, ROBERT  C/P:2,H,LP,SFWA
EYE OF THE WORLD, THE (CHAPTERS 1-18) [@WHEEL OF TIME 1]  JORDAN, ROBERT  C/P,P
EYE OF THE ZODIAC [DUMAREST 13]  TUBB, E. C.  C/P,P
EYE TO EYE  PENDLETON, DON  H,P
EYES  SUBIAGA, ROBERT JR.  LP
EYES HAVE IT, THE [GREYHAWK]  ESTES, ROSE  P
EYES OF AMBER  VINGE, JOAN D.  C/P:2,P
EYES OF BOLSK, THE  LORY, ROBERT  D77710,P
EYES OF CROW  SMITH-READY, JERI  LP
EYES OF DARKNESS, THE  NICHOLS, LEIGH  P
EYES OF FIRE  BISHOP, MICHAEL  C/P,P
EYES OF FIRE (PROOF)  BISHOP, MICHAEL  SR-LP
EYES OF GOD, THE [PINCH 2]  KREIGHBAUM, MARK  P
EYES OF GOD [EYES OF GOD 1]  MARCO, JOHN  H,P
EYES OF HEISENBERG, THE  HERBERT, FRANK  C/P:2,P
EYES OF HORNS  GRANT, JOAN  P
EYES OF KID MIDAS, THE  SHUSTERMAN, NEAL  P
EYES OF LAURA MARS  GILMOUR, H. B.  P
EYES OF LIGHT AND DARKNESS, THE  CAT, IVAN  P
EYES OF SARSIS, THE [WAR OF THE WIZARDS 2]  OFFUTT, ANDREW J./LYON, RICHARD K.  P
EYES OF SILVER  STACKPOLE, MICHAEL A.  C/P,P
EYES OF THE BEHOLDERS, THE [STAR TREK TNG]  CRISPIN, A. C.  C/P:2,P
EYES OF THE CALCULOR [GREAT MACHINE 2]  MCMULLEN, SEAN  H
EYES OF THE DRAGON, THE  KING, STEPHEN  C/P,P,H
EYES OF THE EMPRESS [KEVIN BRADLEY AND EVAN DAVIS 2]  BACON-SMITH, CAMILLE  P
EYES OF THE OVERWORLD, THE [DYING EARTH 2]  VANCE, JACK  SR-H,C/P:2,P
EYES OF THE TAROT [DARK FORCES 9]  COVILLE, BRUCE  P
EYRE AFFAIR, THE [THURSDAY NEXT 1]  FFORDE, JASPER  LP

The MITSFS Titledex  September 24, 2007  Page 119
TITLE  AUTHOR  CODE
’F’ CERTIFICATE, THE  GURNEY, DAVID  H
F-CUBED  DA CRUZ, DANIEL  C/P,P
FABLES & REFLECTIONS [@SANDMAN #6]  GAIMAN, NEIL  H,CX
FABLES AND FAIRY TALES  TOLSTOY, LEO  LP
FABULIST, THE  VORNHOLT, JOHN  P,SFWA-P
FABULOUS FORMLESS DARKNESS, A  HARTWELL, DAVID G.  PA
FABULOUS HARBORS [SECOND ETHER 2]  MOORCOCK, MICHAEL  H,LP
FABULOUS RIVERBOAT, THE [RIVERWORLD 2]  FARMER, PHILIP JOSE  C/P,P,H
FACE, THE  KOONTZ, DEAN R.  SR-LP
FACE, THE [DEMON PRINCES 4]  VANCE, JACK  C/P:2,P
FACE-EATER, THE [DOCTOR WHO EIGHTH DOCTOR]  MESSINGHAM, SIMON  P
FACE IN THE ABYSS, THE  MERRITT, A.  C/P,SR-LP,P,H
FACE IN THE FROST, THE  BELLAIRS, JOHN  C/P,SFWA-P,P,H
FACE IN THE MIRROR, THE  BAKER, DENYS VAL  H
FACE IN THE NIGHT  BRETT, LEO  SR-P
FACE OF APOLLO, THE [GODS 1]  SABERHAGEN, FRED  C/P,LP
FACE OF CHAOS, THE [@THIEVES’ WORLD #5]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
FACE OF FEAR, THE  KOONTZ, DEAN R.  C/P,P
FACE OF FEAR, THE  TORRO, PEL  SR-P
FACE OF HIM, THE  GREENFIELD, IRVING A.  P
FACE OF THE DEEP, THE  YOUNG, JIM  P
FACE OF THE DEEP (PROOF)  YOUNG, JIM  SR-LP
FACE OF THE ENEMY  FAWKES, RICHARD  C/P,P
FACE OF THE ENEMY, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P:2,P
FACE OF THE WATERS, THE  SILVERBERG, ROBERT  H:2,P
FACE OF TIME, THE  BACON-SMITH, CAMILLE  P
FACE OF X, THE  ROBERTS, LIONEL  SR-P
FACE/OFF  CARLTON, CLARK  P
FACE THAT MUST DIE, THE  CAMPBELL, RAMSEY  C/P,SR-P,H
FACE, THE (VIE) [@VANCE INTEGRAL EDITION #25/DEMON PRINCES 4]  VANCE, JACK  SR-H
FACE WITHOUT A HEART, A  REED, RICK R.  LP
FACELESS MAN, THE  VANCE, JACK  P
FACELESS ONES, THE [@DOCTOR WHO]  DICKS, TERRANCE  P
FACELESS PLANET  BRETT, LEO  SR-P
FACES  KENNEDY, LEIGH  H,SFWA
FACES IN THE FLAMES  TATE, PETER  H,SFWA
FACES OF CETI, THE  CARAKER, MARY  H
FACES OF DECEPTION [@FORGOTTEN REALMS LOST EMPIRES]  DENNING, TROY  P
FACES OF FANTASY, THE  PERRET, PATTI  R/REF-H
FACES OF FEAR  WINTER, DOUGLAS E.  R/REF-H
FACES OF FIRE [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P
FACES OF INFINITY [GATES OF TIME 2]  PARKINSON, DAN  P
FACES OF SCIENCE FICTION, THE  PERRET, PATTI  R/REF-H
FACES OF THE FUTURE  ASH, BRIAN  R/REF-H
FACES UNDER WATER [@SECRET BOOKS OF VENUS #1]  LEE, TANITH  H
FACETS  WILLIAMS, WALTER JON  C/P,SFWA-P,P,H
FACIAL JUSTICE  HARTLEY, L. P.  H
FACING THE FLAG  VERNE, JULES  H
FACT AND FANCY  ASIMOV, ISAAC  R/REF-H,C/REF-P:2
FACT AND FANCY (ARABIC)  ASIMOV, ISAAC  FP
FACTORING HUMANITY  SAWYER, ROBERT J.  H,P
FACTS ON THE FLYING SAUCER CONTROVERSY, THE  CHAMBERS, HOWARD V.  [C/RAN-H]
FADE  CORMIER, ROBERT  H,P
FADE-OUT  TILLEY, PATRICK  H,P
FADE TO BLACK [@SHADOWRUN 13]  SMITH, NYX  P
FADED STEEL HEAT [GARRETT 9]  COOK, GLEN  C/P,P
FADED SUN, THE [FADED SUN 1,2,3]  CHERRYH, C. J.  C/P:2,P
FADED SUN: KESRITH, THE [FADED SUN 1]  CHERRYH, C. J.  H:2,P
FADED SUN: KUTATH, THE [FADED SUN 3]  CHERRYH, C. J.  H:2,P
FADED SUN: SHON’JIR, THE [FADED SUN 2]  CHERRYH, C. J.  H:2,P
FAERIE QUEENE I, THE  SPENSER, EDMUND  SR-P
FAERIE QUEENE II, THE  SPENSER, EDMUND  SR-P
FAERIE TALE  FEIST, RAYMOND E.  C/P,P,H
FAERIE TALE (BRITISH)  FEIST, RAYMOND E.  SR-H
FAERIE TALES  GREENBERG, MARTIN H./DAVIS, RUSSELL  PA
FAERIE WARS  BRENNAN, HERBIE  LP

FAERIES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

FAERIES (TWENTY-FIFTH ANNIVERSARY EDITION)  FROUD, BRIAN/LEE, ALAN  A
FAERY!  WINDLING, TERRI  C/PA,PA
FAERY CONVENTION, THE  DAVIS, BRETT  P
FAERY IN SHADOW  CHERRYH, C. J.  C/P,P,H
FAERY LANDS FORLORN [MAN OF HIS WORD 2]  DUNCAN, DAVE  C/P:2,P
FAERY REEL, THE  DATLOW, ELLEN/WINDLING, TERRI  HA

The MITSFS Titledex  September 24, 2007  Page 120
TITLE  AUTHOR  CODE
FAHRENHEIT 451  BRADBURY, RAY  SR-P,H:2,P
FAHRENHEIT 666  HARMAN, ANDREW  P
FAIL-SAFE  BURDICK, EUGENE/WHEELER, HARVEY  C/P:2,P
FAINT ECHOES, DISTANT STARS  BOVA, BEN  R/REF-H
FAIR FOLK, THE  KAYE, MARVIN  HA
FAIR IN EMAIN MACHA, THE  DE LINT, CHARLES  SFWA-TD19,TD19
FAIR PERIL  SPRINGER, NANCY  C/P,P
FAIRY GODMOTHER, THE [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  P
FAIRYLAND  MCAULEY, PAUL J.  C/P,LP,H
FAITH & FIRE [@WARHAMMER 40K SISTERS OF BATTLE 1]  SWALLOW, JAMES  P
FAITH OF TAROT [TAROT #3]  ANTHONY, PIERS  C/P:2,@P
FAITH OF THE FALLEN [SWORD OF TRUTH 6]  GOODKIND, TERRY  C/P,P,H
FAITH OF THE UNFORGOTTEN [@FOUNDATIONS OF HOPE #1]  GRODY, LEATHEL  H
FAITH TRIALS, THE, VOL. 1 [BUFFY]  LAURENCE, JAMES  P
FAITHFUL TRAITOR, THE [WIZARD AND DRAGON 2]  HUGHES, ROBERT DAN  C/P,P
FAL THE DRAGONHARPER  PATTERSON, PETER  LP
FALCON  BULL, EMMA  C/P:2,SFWA-P,P
FALCON GUARD [@BATTLETECH #3/JADE PHOENIX 3]  THURSTON, ROBERT  C/P:2,P
FALCON OF EDEN, THE [EMPIRE PRINCESS 3]  DIAMOND, GRAHAM  P
FALCON RISES, THE  STAUDINGER, MICHAEL C.  P
FALCON RISING [@BATTLETECH #40/TWILIGHT OF THE CLANS 8]  THURSTON, ROBERT  P
FALCONS OF NARABEDLA  BRADLEY, MARION ZIMMER  P,DF273
FALKENBERG’S LEGION [FALKENBERG’S LEGION]  POURNELLE, JERRY  C/P,P
FALL, THE [@PLANET OF THE APES]  QUICK, W. T.  P
FALL GIRL [INCONSTANT MOON 2]  ASKEGREN, PIERCE  C/P,P
FALL INTO DARKNESS  YERMAKOV, NICHOLAS  P
FALL OF A KINGDOM [@FARSALA #1]  BELL, HILARI  P
FALL OF A NATION, THE  DIXON, THOMAS  H
FALL OF ANGELS [RECLUCE 6]  MODESITT, L. E. JR.  H:2,P

FALL OF ATLANTA, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #1] WIECK, STEWART/FLEMING, GHERBOD/

GRIFFIN, ERIC

LP

FALL OF ATLANTIS, THE [WEB 1,2]  BRADLEY, MARION ZIMMER  C/P:2,P
FALL OF CHRONOPOLIS, THE  BAYLEY, BARRINGTON J.  C/P:2,P
FALL OF COLOSSUS, THE [COLOSSUS 2]  JONES, D. F.  H:2,P
FALL OF HYPERION, THE [HYPERION 2]  SIMMONS, DAN  C/P:2,P,H
FALL OF MAGIC, THE [@DUNGEONS & DRAGONS PENHALIGON #3]  HEINRICH, D. J.  C/P,P
FALL OF MOONDUST, A  CLARKE, ARTHUR C.  C/P,P,H

FALL OF NESKAYA, THE [@CLINGFIRE #1/DARKOVER] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H,P

FALL OF NEW YORK, THE  DONIS, MILES  P
FALL OF PRINCES, A [@AVARYAN RISING #3]  TARR, JUDITH  C/P,SFWA-P,P,H
FALL OF ROME, THE  LAFFERTY, R. A.  SFWA
FALL OF SIRIUS, THE [AGGRESSOR SIX 2]  MCCARTHY, WIL  P

FALL OF TEROK NOR [@STAR TREK DS9 MILLENIUM #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

FALL OF THE DREAM MACHINE, THE  KOONTZ, DEAN R.  D22600
FALL OF THE FAMILIES, THE [GARDNER 2]  MANN, PHILLIP  H
FALL OF THE KINGS, THE [SWORDSPOINT 2]  KUSHNER, ELLEN/SHERMAN, DELIA  C/P,LP
FALL OF THE NIBELUNGS, THE  MACDOUGALL, W. B.  A
FALL OF THE REPUBLIC, THE [CHRONOPLANE WARS 1]  KILIAN, CRAWFORD  C/P,P
FALL OF THE REPUBLIC (STORY TREATMENT) [STAR WARS]  FLYNN, JOHN L.  SR-S
FALL OF THE RUSSIAN EMPIRE, THE  JAMES, DONALD  P
FALL OF THE SHELL, THE [PELBAR 4]  WILLIAMS, PAUL O.  C/P:2,P
FALL OF THE SKY LORDS, THE [SKY LORDS 3]  BROSNAN, JOHN  P
FALL OF THE TOWERS, THE [FALL OF THE TOWERS 1,2,3]  DELANY, SAMUEL R.  C/P:2,P
FALL OF THE WHITE SHIP AVATAR [HOBART FLOYT]  DALEY, BRIAN  C/P:2,P
FALL OF WINTER, THE  HALDEMAN, JACK C. II  C/P,P
FALL OF WORLDS, THE [@AREIA DARENGA #1]  MEZO, FRANCINE  C/P:2,P
FALL OF YQUATINE, THE [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  P
FALL REVOLUTION, THE [FALL REVOLUTION]  MACLEOD, KEN  H
FALLBACK  NIESEWAND, PETER  H,P
FALLEN, THE  BAILEY, DALE  H
FALLEN ANGELS [@SHADOWRUN KELLAN COLT 2]  KENSON, STEPHEN  P

FALLEN ANGELS NIVEN, LARRY/POURNELLE, JERRY/

FLYNN, MICHAEL

C/P,P,H

FALLEN COUNTRY, THE  SUCHARITKUL, SOMTOW  C/P,P
FALLEN DRAGON  HAMILTON, PETER F.  H
FALLEN FORTRESS, THE [@FORGOTTEN REALMS CLERIC #4]  SALVATORE, R. A.  P
FALLEN GODS [DOCTOR WHO]  BLUM, JONATHAN/ORMAN, KATE  SR-H
FALLEN HEROES [STAR TREK DS9]  AB HUGH, DAFYDD  C/P:2,P
FALLEN HEROES (PROOF) [STAR TREK DS9]  AB HUGH, DAFYDD  SR-VLP
FALLEN HOST [ARCHANGEL 2]  MOREHOUSE, LYDA  P
FALLEN ONES, THE [WAR OF POWERS 4]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
FALLEN SPACEMAN, THE  HARDING, LEE  H,LP

TITLE  AUTHOR  CODE
FALLEN STAR  BLISH, JAMES  SR-P,C/P:2,H
FALLIBLE FIEND, THE  DE CAMP, L. SPRAGUE  C/P:2,P
FALLING ANGEL  HJORTSBERG, WILLIAM  P
FALLING ASTRONAUTS, THE  MALZBERG, BARRY N.  P
FALLING FREE  BUJOLD, LOIS MCMASTER  C/P:2,SFWA-P,P
FALLING SIDEWAYS  HOLT, TOM  H
FALLING STARS [STARS 4]  FLYNN, MICHAEL  H
FALLING TORCH, THE  BUDRYS, ALGIS  C/P,P
FALLING TORCH (COMPLETE)  BUDRYS, ALGIS  C/P,P
FALLING TOWARD FOREVER  EKLUND, GORDON  SR-P,C/P
FALLING WOMAN, THE  MURPHY, PAT  H,SFWA-P:2,P
FALLOUT [CRAIG KREIDENT 2]  ANDERSON, KEVIN J./BEASON, DOUG  P
FALLS THE SHADOW [DOCTOR WHO NEW ADVENTURES]  O’MAHONY, DANIEL  C/P:2,P
FALLWAY  DOWNING, PAULA E.  P
FALSE COLORS [@WING COMMANDER 7]  FORSTCHEN, WILLIAM R./KEITH, ANDREW  C/P,P
FALSE DAWN  YARBRO, CHELSEA QUINN  H:2,SFWA,P
FALSE FRONT [@PERRY RHODAN #103]  ACE  PA
FALSE GODS [@WARHAMMER 40K HORUS HERESY 2]  MCNEILL, GRAHAM  P
FALSE HOUSE, THE [EVENMERE 2]  STODDARD, JAMES  P
FALSE MIRROR, THE [@DAMNED #2]  FOSTER, ALAN DEAN  C/P,P,H
FALSE NIGHT  BUDRYS, ALGIS  SR-P,C/P
FALSE NOTES [CHARLIE MOON 2]  YARBRO, CHELSEA QUINN  P
FAMILIAR DRAGON, THE [FANUILH 1,2,3]  HOOD, DANIEL  H
FAMILIAR SPIRIT  TUTTLE, LISA  SFWA-P
FAMILIARS  LITTLE, DENISE  PA
FAMILY CRYPT [TWILIGHT 20]  TRAINOR, JOSEPH  P
FAMILY TRADE, THE [MERCHANT PRINCES 1]  STROSS, CHARLES  H,P
FAMILY TREE, THE  TEPPER, SHERI S.  C/P:2,LP
FAMINE, THE  CREASEY, JOHN  H
FAMINE  MASTERTON, GRAHAM  P
FAMOUS FANTASTIC CLASSICS #1  FAX PUBLISHERS  SR-LP
FAMOUS FANTASTIC CLASSICS #2  FAX PUBLISHERS  SR-LP
FAMOUS GHOST STORIES  CERF, BENNETT  PA
FANCESTRAL VOICES  SPEER, JACK  R/REF-H
FANCIES AND GOODNIGHTS  COLLIER, JOHN  C/P,SR-P,LP,H
FANCYCLOPEDIA II  ENEY, DICK  R/REF-H
FANDOM DIRECTORY 1  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 3  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 4  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 5  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 6  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 8  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 9  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 10  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 11  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 12  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 13  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 14  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 15  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 16  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 17  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 18  HOPKINS, MARIANE S.  SR-L
FANDOM DIRECTORY 19  HOPKINS, MARIANE S.  R/REF-P,C/REF-H
FANDOM OF THE OPERATOR, THE  RANKIN, ROBERT  P
FANE  ALEXANDER, DAVID M.  C/P,P
FANG THE GNOME  CONEY, MICHAEL G.  H
FANGLITH  DALMAS, JOHN  C/P,P
FANGS OF THE HOODED DEMON, THE [LINCOLN BLACKTHORNE 4]  MARSH, GEOFFREY  SFWA-P
FANTASCIENZA, LA  ALDANI, LINO  R/REF-P
FANTASIA MATHEMATICA  FADIMAN, CLIFTON  H,LP
FANTASIES OF ROBERT A. HEINLEIN, THE  HEINLEIN, ROBERT A.  LP
FANTASIES TWO  GRIFFITH, WILLIAM JR.  H
FANTASISTS ON FANTASY  BOYER, ROBERT H./ZAHORSKI, KENNETH J. R/REF-P
FANTASMA MONTELEONE, THOMAS F. SFWA-P

FANTASMS 2.1:

A BIBLIOGRAPHY OF THE PUBLISHED FICTION OF JACK VANCE COCKRUM, KURT/LEVACK, DANIEL J. H./

UNDERWOOD, TIM

R/IND

FANTASMS: A BIBLIOGRAPHY OF THE LITERATURE OF JACK VANCE  LEVACK, DANIEL J. H./UNDERWOOD, TIM  R/IND
FANTASTIC ADVENTURES OF ROBIN HOOD, THE  GREENBERG, MARTIN H.  C/PA,PA
FANTASTIC ALICE  WEIS, MARGARET  LP
FANTASTIC ART OF FRANK FRAZETTA, THE  FRAZETTA, FRANK  A
FANTASTIC CIVIL WAR, THE  MCSHERRY, FRANK JR.  PA,SFWA-P
FANTASTIC IMAGINATION, THE  BOYER, ROBERT H./ZAHORSKI, KENNETH J. C/PA,PA
FANTASTIC ISLAND, THE [DOC SAVAGE #14]  ROBESON, KENNETH  C/P,@P
FANTASTIC KINGDOM, THE  LARKIN, DAVID  A
FANTASTIC LIVES  GREENBERG, MARTIN H.  R/REF-H

The MITSFS Titledex  September 24, 2007  Page 122
TITLE  AUTHOR  CODE
FANTASTIC MIRROR, THE  APPEL, BENJAMIN  R/REF-H
FANTASTIC PULPS, THE  HAINING, PETER  LP
FANTASTIC SAINT, THE  CHARTERIS, LESLIE  SFWA
FANTASTIC SCIENCE FICTION ART 1926–1954  DEL REY, LESTER  A
FANTASTIC STORIES OF CORNELL WOOLRICH, THE  WOOLRICH, CORNELL  H
FANTASTIC SWORDSMEN, THE  DE CAMP, L. SPRAGUE  C/PA,PA
FANTASTIC TELEVISION  GERANI, GARY/SCHULMAN, PAUL H.  R/REF-H
FANTASTIC UNIVERSE OMNIBUS, THE  SANTESSON, HANS STEFAN  C/PA,PA
FANTASTIC VOYAGE  ASIMOV, ISAAC  H:3,P
FANTASTIC VOYAGE II: DESTINATION BRAIN  ASIMOV, ISAAC  C/P,P,H,SFWA
FANTASTIC VOYAGE MICROCOSM  ANDERSON, KEVIN J.  P
FANTASTIC WORLD WAR II, THE  MCSHERRY, FRANK JR.  PA,SFWA-P
FANTASTIC WORLDS  RABKIN, ERIC S.  LPA
FANTASTIKA 73–74  MIKHAILOVA, S.  FH
FANTASTIKA 1964 GOD  SMIRNOV, G.  FH
FANTASTIKA 1965 VYPUSK I  VOSKRESENSKAYA, E./KLYUEVA, V.  FH
FANTASTIKA 1965 VYPUSK II  STRUGATSKY, ARKADY  FH
FANTASTIKA 1991  MOLODAYA GVARDIYA  FH
FANTASY  ANDERSON, POUL  C/P:2,P
FANTASY ADVENTURES #1  HARBOTTLE, PHILIP  LPA
FANTASY ADVENTURES #2  HARBOTTLE, PHILIP  LPA
FANTASY ADVENTURES 3  HARBOTTLE, PHILIP  LPA
FANTASY ANNUAL 4  HARBOTTLE, PHILIP/WALLACE, SEAN  LPA
FANTASY ANNUAL 5  HARBOTTLE, PHILIP/WALLACE, SEAN  LPA
FANTASY ANNUAL III  CARR, TERRY  PA
FANTASY ANNUAL IV  CARR, TERRY  HA,PA
FANTASY ANNUAL V  CARR, TERRY  PA
FANTASY ART OF KEITH PARKINSON CALENDAR, THE (2005)  PARKINSON, KEITH  CAL
FANTASY BOOK, THE  ROTTENSTEINER, FRANZ  R/REF-H
FANTASY FICTION DECIMAL CLASSIFICATION  SPEER, JACK  R/IND-F

FANTASY GONE WRONG GREENBERG, MARTIN H./ KOREN, BRITTANY A.

PA

FANTASY HALL OF FAME, THE (SILVERBERG & GREENBERG)

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

HA

FANTASY HALL OF FAME, THE (SILVERBERG ONLY)  SILVERBERG, ROBERT  HA,LPA
FANTASY LIFE  RUSCH, KRISTINE KATHRYN  P
FANTASY SHOWCASE TAROT, THE [TAROT DECK]  PELZ, BRUCE  SR-PA
FANTASY THE BEST OF 2001  SILVERBERG, ROBERT/HABER, KAREN  PA
FANTASY THE BEST OF 2004  HABER, KAREN/STRAHAN, JONATHAN  PA
FANTASY: THE BEST OF THE YEAR 2006  HORTON, RICH  LPA
FANTASY: THE BEST OF THE YEAR 2007  HORTON, RICH  HA,PA
FANTASY THE BEST 100 BOOKS  CAWTHORN, JAMES/MOORCOCK, MICHAEL  R/REF-H
FANTASY WORLDS OF PETER S. BEAGLE, THE  BEAGLE, PETER S.  H:2,LP
FANTASY WORLDS OF PETER STONE AND OTHER FABLES, THE  BOYD, MALCOLM  P
FANTASY WRITER’S ASSISTANT, THE  FORD, JEFFREY  H
FANTOME DE L’OPERA, LE  LEROUX, GASTON  FP
FANUILH [FANUILH 1]  HOOD, DANIEL  C/P,P
FANZINE INDEX  PAVLAT, BOB/EVANS, BILL  R/IND
FAR AND AWAY  BOUCHER, ANTHONY  C/P:2,P
FAR ARENA, THE  SAPIR, RICHARD BEN  P
FAR AS HUMAN EYE COULD SEE  ASIMOV, ISAAC  SFWA
FAR BELOW AND OTHER HORRORS  WEINBERG, ROBERT  HA
FAR BEYOND THE BLUE  AMPER, DRAX  SR-P
FAR BEYOND THE STARS [@STAR TREK DS9]  BARNES, STEVEN  P
FAR BEYOND THE WAVE [@ARTHUR WAR LORD #2]  AB HUGH, DAFYDD  P
FAR CALL, THE  DICKSON, GORDON R.  H:2,SFWA-P,P
FAR COUNTRY [@BATTLETECH 13]  RICE, PETER L.  C/P,P
FAR EDGE OF DARKNESS  EVANS, LINDA  C/P,P
FAR ENDS OF TIME AND EARTH, THE  ASIMOV, ISAAC  H
FAR FROM HOME  TEVIS, WALTER  H
FAR FROM THIS EARTH  OLIVER, CHAD  H
FAR FRONTIER, THE  ROTSLER, WILLIAM  P
FAR FRONTIERS  GREENBERG, MARTIN H./SEGRIFF, LARRY  C/PA,PA
FAR FRONTIERS  POURNELLE, JERRY/BAEN, JIM  PA,C/PA
FAR FRONTIERS II  POURNELLE, JERRY/BAEN, JIM  PA
FAR FRONTIERS III  POURNELLE, JERRY/BAEN, JIM  PA
FAR FRONTIERS IV  POURNELLE, JERRY/BAEN, JIM  PA
FAR FRONTIERS V  POURNELLE, JERRY/BAEN, JIM  PA
FAR FRONTIERS VI  POURNELLE, JERRY/BAEN, JIM  PA
FAR FRONTIERS VII  POURNELLE, JERRY/BAEN, JIM  PA
FAR FUTURE CALLING  STAPLEDON, OLAF  SR-H
FAR FUTURES  BENFORD, GREGORY  HA
FAR HARBOR  MICHAELS, MELISA C.  P
FAR HORIZONS  SILVERBERG, ROBERT  C/PA,PA
FAR KINGDOMS, THE  COLE, ALLAN/BUNCH, CHRIS  C/P,P

TITLE  AUTHOR  CODE
FAR OUT  KNIGHT, DAMON  H:2,P
FAR-OUT PEOPLE, THE  HOSKINS, ROBERT  PA,C/PA
FAR-OUT WORLDS OF A.E. VAN VOGT, THE  VAN VOGT, A. E.  C/P:2,P
FAR RAINBOW/SECOND INVASION FROM MARS, THE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P,LP,H
FAR-SEER [DINOSAUR EMPIRE 1]  SAWYER, ROBERT J.  C/P:2,SFWA-P:2,P
FAR SHORE OF TIME, THE [ESCHATON 3]  POHL, FREDERIK  C/P,P,H
FAR SIDE OF EVIL, THE  ENGDAHL, SYLVIA LOUISE  H:2,LP
FAR SIDE OF FOREVER, THE [@FAR SIDE OF FOREVER #1]  GREEN, SHARON  C/P,P
FAR SIDE OF THE STARS, THE [RCN 3]  DRAKE, DAVID  H:2
FAR SIDE OF THE STARS, THE (PROOF) [RCN 3]  DRAKE, DAVID  SR-LP
FAR SIDE OF TIME, THE  ELWOOD, ROGER  SR-HA,HA
FAR STARS WAR, THE [@WAR YEARS #1]  FAWCETT, BILL  PA
FAR SUNSET, A  COOPER, EDMUND  C/P:2,P
FAR TRAVELER, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
FARADAY’S ORPHANS  WOOD, N. LEE  C/P,P
FARAGON FAIRINGAY [@CIRCLE OF LIGHT #2]  HANCOCK, NIEL  C/P:2,P
FARAWAY LURS, THE  BEHN, HARRY  P
FARCE TO BE RECKONED WITH, A [AZZIE 3]  ZELAZNY, ROGER/SHECKLEY, ROBERT  H,LP
FAREWELL, EARTH’S BLISS  COMPTON, D. G.  C/P:2,P
FAREWELL HORIZONTAL  JETER, K. W.  H:2
FAREWELL, NIKOLA  BOOTHBY, GUY  H
FAREWELL TO YESTERDAY’S TOMORROW  PANSHIN, ALEXEI  H:2,SFWA,P
FARFETCH [DUSHAU 2]  LICHTENBERG, JACQUELINE  C/P,P
FARMER GILES OF HAM  TOLKIEN, J. R. R.  H
FARMER IN THE SKY  HEINLEIN, ROBERT A.  C/P,P,H
FARNHAM’S FREEHOLD  HEINLEIN, ROBERT A.  H:2,SFWA-P,P
FARSCAPE: THE ILLUSTRATED COMPANION [FARSCAPE]  SIMPSON, PAUL/HUGHES, DAVID  R/REF-P
FARSEEKERS, THE [OBERNEWTYN 2]  CARMODY, ISOBELLE  H,P
FARSEER [WARHAMMER 40K ELDAR]  KING, WILLIAM  P
FARSIDE CANNON  ALLEN, ROGER MACBRIDE  C/P:2,SFWA-P,P
FARSLAYER’S STORY [@LOST SWORDS #4]  SABERHAGEN, FRED  H:2,SFWA-P,P
FARTHER SHORE, THE [@STAR TREK VOYAGER HOMECOMING 2]  GOLDEN, CHRISTIE  P
FARTHEST REACH [@FORGOTTEN REALMS LAST MYTHAL #2]  BAKER, RICHARD  P
FARTHEST REACHES, THE  ELDER, JOSEPH  C/PA,PA
FARTHEST SEEKING, THE [@GHATTENS’ GAMBIT #2]  GREENO, GAYLE  C/P,P
FARTHEST SHORE, THE [EARTHSEA 3]  LE GUIN, URSULA K.  C/P,P,H,SFWA
FARTHEST SHORES OF URSULA K. LE GUIN, THE  SLUSSER, GEORGE EDGAR  R/REF-P
FARTHEST STAR  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
FARTHING  WALTON, JO  H
FAST FORWARD 1  ANDERS, LOU  LPA
FASTER THAN LIGHT  DANN, JACK/ZEBROWSKI, GEORGE  C/PA,PA,SFWA
FASTER THAN LIGHT [@SPACE HAWKS #1]  PACKARD, EDWARD  P
FASTWALKER  VALLEE, JACQUES/TORME, TRACY  LP
FAT CHANCE  LAUMER, KEITH  P
FAT WHITE VAMPIRE BLUES [FAT WHITE VAMPIRE (JULES DUCHON) 1]  FOX, ANDREW  LP
FATA MORGANA, THE  FRANKOWSKI, LEO  C/P,LP
FATAL ATTRACTION [TWILIGHT 4]  HOWE, IMOGEN  P
FATAL EXPOSURE  TOBIAS, MICHAEL  P
FATAL KISS MYSTERY, THE  KING, RUFUS  SR-P
FATAL LAW, THE  KENT, BRAD  SR-P
FATAL SHADOWS  GEORGE, SARA  P
FATE OF MICE, THE  PALWICK, SUSAN  LP
FATE OF THE PHOENIX, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
FATE OF THE VAMPYRE, THE [@VARNEY THE VAMPIRE #5]  RYMER, JAMES  LP
FATE WORSE THAN DRAGONS, A  MOORE, JOHN  P
FATEFUL LIGHTNING [@LOST REGIMENT 4]  FORSTCHEN, WILLIAM R.  P
FATES OF THE PRINCES OF DYFED, THE  MORRIS, KENNETH  LP
FATE’S TRICK  COSTELLO, MATTHEW J.  P,SFWA-P
FATHER CHRISTMAS LETTERS, THE  TOLKIEN, J. R. R.  VLP
FATHER MALACHY’S MIRACLE  MARSHALL, BRUCE  P
FATHER OF DRAGONS [@BINDING OF THE BLADE #4]  GRAHAM, L. B.  LP
FATHER OF LIES  BRUNNER, JOHN  SR-DB0081
FATHER TIME [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
FATHER TO THE MAN  GRIBBIN, JOHN  P,SFWA-P
FATHER TO THE STARS  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
FATHERLAND  HARRIS, ROBERT  H:2,P
FATHOMLESS CAVES, THE [WITCHES OF EILEANAN 6]  FORSYTH, KATE  C/P,P
FAULT LINES [FAULT LINES 1,2]  POWERS, TIM  H
FAULT LINES  REIDA, ALVAH  P
FAULT LINES  WILHELM, KATE  P
FAUN & GAMES [XANTH #21]  ANTHONY, PIERS  H,@P,@C/P
FAUN AND THE WOODCUTTER’S DAUGHTER, THE  PICARD, BARBARA LEONIE  H
FAUST AMONG EQUALS  HOLT, TOM  P
FEAR  HUBBARD, L. RON  P,GN5.29

The MITSFS Titledex  September 24, 2007  Page 124
TITLE  AUTHOR  CODE
FEAR/TYPEWRITER IN THE SKY  HUBBARD, L. RON  P
FEAR/ULTIMATE ADVENTURE, THE  HUBBARD, L. RON  C/P:2,P
FEAR AND TREMBLING  BLOCH, ROBERT  SFWA-P
FEAR CAY [DOC SAVAGE #11]  ROBESON, KENNETH  C/P:2,@P
FEAR ITSELF [DOCTOR WHO PAST DOCTORS]  WALLACE, NICK  P
FEAR ITSELF: THE HORROR FICTION OF STEPHEN KING  UNDERWOOD, TIM/MILLER, CHUCK  R/REF-H
FEAR OF GOD, THE [FEAR 2]  CHEPAITIS, B. A.  C/P,P
FEAR OF THE DARK [DOCTOR WHO PAST DOCTORS]  BAXENDALE, TREVOR  P
FEAR PLANET, THE  BLOCH, ROBERT  SR-H
FEAR PRINCIPLE, THE [FEAR 1]  CHEPAITIS, B. A.  P
FEAR THAT MAN  KOONTZ, DEAN R.  D23140
FEAR TODAY — GONE TOMORROW  BLOCH, ROBERT  P
FEARFUL ROCK  WELLMAN, MANLY WADE  H
FEARFUL SUMMONS, THE [STAR TREK]  FLINN, DENNY MARTIN  C/P:2,P
FEARFUL SYMMETRIES [MOREAU 4]  SWANN, S. ANDREW  P
FEARFUL SYMMETRY, A  LUCENO, JAMES  C/P,P
FEARLESS [@LOST FLEET 2]  CAMPBELL, JACK  P
FEARLESS [@ANGEL]  DURGIN, DORANNA  P
FEARLESS MASTER OF THE JUNGLE  EDSON, J. T.  SR-P
FEARS  GRANT, CHARLES L.  C/PA,PA
FEARSOME ISLAND, THE  KINROSS, ALBERT  SR-P
FEAST FOR CROWS, A [SONG OF ICE AND FIRE 4]  MARTIN, GEORGE R. R.  H
FEAST IN EXILE, A [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
FEAST OF FEAR  GHIDALIA, VIC  PA
FEAST OF FREEDOM, A  WIBBERLEY, LEONARD  P
FEAST OF ST. DIONYSUS  SILVERBERG, ROBERT  H:2,SFWA,P
FEAST OF SOULS [@MAGISTER #1]  FRIEDMAN, C. S.  H
FEAST OF THE DROWNED, THE [DOCTOR WHO TENTH DOCTOR]  COLE, STEPHEN  H
FEAST OF THE KING’S SHADOW [@OUTREMER #4]  BRENCHLEY, CHAZ  P
FEAST OF THE KING’S SHADOW (COMPLETE) [OUTREMER 3,4]  BRENCHLEY, CHAZ  P
FEAST UNKNOWN, A  FARMER, PHILIP JOSE  C/P,SR-L,P
FEATHER & BONE  SMITH, GUS  LP
FEATHER STROKE  VAN SCYOC, SYDNEY J.  C/P,P
FEATHERED DRAGON, THE [@FORGOTTEN REALMS MAZTICA #3]  NILES, DOUGLAS  P
FEATHERED OCTOPUS, THE [DOC SAVAGE #48]  ROBESON, KENNETH  C/P:2,@P
FEDERATION  PIPER, H. BEAM  C/P,LP,H

FEDERATION [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P,H

FEDERATION WORLD  WHITE, JAMES  C/P:2,P
FEE, FEI, FO, FUM  AYLESWORTH, JOHN  C/P:2,P
FEELIES, THE  FARREN, MICK  P
FEELING VERY STRANGE  KELLY, JAMES PATRICK/KESSEL, JOHN  LPA
FEERSUM ENDJINN  BANKS, IAIN M.  C/P,LP,H
FEESTERS IN THE LAKE  LEMAN, BOB  SR-H
FEET OF CLAY [DISCWORLD 19/DISCWORLD WATCHMEN 3]  PRATCHETT, TERRY  H:2,P
FELAHEEN [EL ISKANDRYIA 3]  GRIMWOOD, JON COURTENAY  H
FELINE FETISHES  CORWIN  SR-L
FELINE WIZARD, THE [WIZARD IN RHYME 8]  STASHEFF, CHRISTOPHER  C/P,P
FELL CARGO [@WARHAMMER]  ABNETT, DAN  P
FELLOW TRAVELER  BARTON, WILLIAM/CAPOBIANCO, MICHAEL C/P:2,P
FELLOWSHIP OF THE HAND, THE  HOCH, EDWARD D.  H,SFWA,P
FELLOWSHIP OF THE RING, THE [@LORD OF THE RINGS #1]  TOLKIEN, J. R. R.  SR-H,C/P,P,H
FELLOWSHIP OF THE RING, THE (ACE) [@LORD OF THE RINGS #1]  TOLKIEN, J. R. R.  SR-P
FELLOWSHIP OF THE STARS  CARR, TERRY  SR-HA,SFWA,HA
FELLOWSHIP OF THE TALISMAN, THE  SIMAK, CLIFFORD D.  H:2,P
FEMALE MAN, THE  RUSS, JOANNA  C/P:2,P
FEMALE WAR, THE [@ALIENS #3]  PERRY, STEVE/PERRY, STEPHANI  C/P,P
FEMININE EYE, THE  STAICAR, TOM  R/REF-H
FEMINIST FAIRY TALES  WALKER, BARBARA G.  LP
FEMINISTS, THE  COOPER, PARLEY J.  P
FENRIS DEVICE, THE [GRAINGER 5]  STABLEFORD, BRIAN M.  C/P,P
FERAL CELL  BOWES, RICHARD  P
FERAL DARKNESS, A  DURGIN, DORANNA  P
FERENGI RULES OF ACQUISITION, THE [STAR TREK DS9]  BEHR, IRA STEVEN  R/REF-P,C/REF-P

FERENGINAR: SATISFACTION IS NOT GUARANTEED [STAR TREK DS9 WORLDS 3]/ DOMINION, THE: OLYMPUS DESCENDING [STAR TREK DS9 WORLDS 3] DECANDIDO, KEITH R. A./ GEORGE, DAVID R. III

C/P,P

FERMAN’S DEVILS [PEMBROKE HALL 1]  FAUST, JOE CLIFFORD  P
FERRET FANTASY’S CHRISTMAS ANNUAL FOR 1973  LOCKE, GEORGE  R/IND
FERRY ROCKET  KINLEY, GEORGE  SR-P
FESTIVAL MOON [@MEROVINGEN NIGHTS #1]  CHERRYH, C. J.  C/PA,PA
FESTIVAL OF DEATH [DOCTOR WHO PAST DOCTORS]  MORRIS, JONATHAN  C/P,P
FEVER  COOK, ROBIN  C/P,P
FEVER IN URBICAND  SCHUITEN, FRANCOIS/PEETERS, BENOIT  CX

The MITSFS Titledex  September 24, 2007  Page 125
TITLE  AUTHOR  CODE
FEVER SEASON [@MEROVINGEN NIGHTS #2]  CHERRYH, C. J.  C/PA,PA
FEVRE DREAM  MARTIN, GEORGE R. R.  H:2,P
FEW LAST WORDS, A  SALLIS, JAMES  H
FEW WERE LEFT  REIN, HAROLD  H
FIASCO  LEM, STANISLAW  LP
FICKLE FINGER OF FATE, THE  KEEL, JOHN A.  P
FICTION FACTORY, THE  DANN, JACK  H
FIDDLER FAIR  LACKEY, MERCEDES  C/P,P
FIEBRE DEL HENO, LA  LEM, STANISLAW  FP
FIELD GUIDE TO THE LITTLE PEOPLE, A  ARROWSMITH, NANCY/MOORE, GEORGE  R/REF-P
FIELD OF DISHONOR [HONOR HARRINGTON 4]  WEBER, DAVID  C/P:2,P
FIELDS OF HONOR [RUIN MIST 3]  STANEK, ROBERT  LP
FIELDS OF SLEEP  VIVIAN, E. CHARLES  SR-H
FIEND, THE  PLAYBOY BOOKS  C/PA,PA
FIEND FOLIO  TSR GAMES  [C/RAN-H]
FIENDS  FARRIS, JOHN  SFWA-P
FIENDS AND CREATURES  KAYE, MARVIN  C/PA,PA
FIFTEEN HOURS [@WARHAMMER 40K]  SCANLON, MITCHEL  P
FIFTEENTH PELICAN, THE  RIOS, TERE  P
FIFTH DAWN, THE [@MAGIC: THE GATHERING MIRRODIN #3]  HERNDON, CORY J.  P
FIFTH DOCTOR, THE [DOCTOR WHO]  HOWE, DAVID J./WALKER, STEPHEN JAMES  R/REF-P,C/REF-P
FIFTH ELEMENT, THE  BISSON, TERRY  P
FIFTH ELEPHANT, THE [DISCWORLD 24/DISCWORLD WATCHMEN 5]  PRATCHETT, TERRY  H:2,P
FIFTH HEAD OF CERBERUS, THE  WOLFE, GENE  C/P,SFWA,P,H
FIFTH HORSEMAN, THE  COLLINS, LARRY/LAPIERRE, DOMINIQUE  C/P,P
FIFTH HORSEMAN, THE  SHERBANIUK, RICHARD  H
FIFTH HORSEMAN, THE [BLOODSHADOWS]  STARK, ED  LP
FIFTH PLANET  HOYLE, FRED/HOYLE, GEOFFREY  H:2,P

FIFTH QUADRANT, THE [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS #2] RYAN, C. J.

P

FIFTH QUARTER [QUARTER SINGERS 2]  HUFF, TANYA  C/P:2,P
FIFTH RING, THE [MATHEW LEWIN 1]  GRAHAM, MITCHELL  P
FIFTH SACRED THING, THE  STARHAWK  H
FIFTH SORCERESS, THE [BLOOD AND STONE 1]  NEWCOMB, ROBERT  H
FIFTY DEGREES BELOW [FORTY SIGNS OF RAIN 2]  ROBINSON, KIM STANLEY  H:2
FIFTY EXTREMELY SHORT FICTION STORIES  BASTRAW, MICHAEL  SR-LPA
50 GREAT GHOST STORIES  CANNING, JOHN  C/PA
50 IN 50  HARRISON, HARRY  H
FIFTY-METER MONSTERS AND OTHER HORRORS, THE  ELWOOD, ROGER  SR-PA
FIFTY-MINUTE HOUR, THE  LINDNER, ROBERT  [C/RAN-P]
FIFTY-ONE TALES  DUNSANY, LORD  LP
’57 T-BIRD  JACKSON, ANITA  SR-P
57TH FRANZ KAFKA, THE  RUCKER, RUDY  P
FIFTY SHORT SCIENCE FICTION TALES  ASIMOV, ISAAC/CONKLIN, GROFF  PA,C/PA
FIGHT FOR LIFE  LEINSTER, MURRAY  SR-P
FIGHTING MAN OF MARS, A [BARSOOM #7]  BURROUGHS, EDGAR RICE  C/P,@P,H
FIGHTING SLAVE OF GOR [GOR 14]  NORMAN, JOHN  C/P:2,P
FIGURE IN THE FROST [@KNIGHTS OF THE SILVER DRAGON #5]  PEREZ, LANA  LP
FIGURE IN THE SHADOWS, THE  BELLAIRS, JOHN  LP
FIGURES OF EARTH  CABELL, JAMES BRANCH  C/P,P,H
FINAL ASSAULT [BIOFAB 4]  BERRY, STEPHEN AMES  P,SFWA-P
FINAL BATTLE, THE [LEGION OF THE DAMNED 2]  DIETZ, WILLIAM C.  C/P,P
FINAL BATTLE, THE [DUNGEON #6]  LUPOFF, RICHARD A.  C/P:2,P
FINAL BLACKOUT  HUBBARD, L. RON  SR-H,P
FINAL CIRCLE OF PARADISE, THE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P,P
FINAL CIRCUIT [CIRCUIT 3]  SNODGRASS, MELINDA M.  P
FINAL COMMAND [DOUBLE SPIRAL WAR 3]  NORWOOD, WARREN  P
FINAL CONFLICT, THE [OMEN 3]  MCGILL, GORDON  P
FINAL COUNTDOWN, THE  CAIDIN, MARTIN  P
FINAL CUT, THE [@JUDGE DREDD]  SMITH, MATTHEW  P
FINAL DIAGNOSIS [SECTOR GENERAL 10]  WHITE, JAMES  H
FINAL ENCOUNTER  POHL, FREDERIK  C/PA,PA
FINAL ENCYCLOPEDIA, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H
FINAL FRONTIER [STAR TREK]  CAREY, DIANE  C/P:2,P
FINAL FURY, THE [STAR TREK INVASION! 4/STAR TREK VOYAGER]  AB HUGH, DAFYDD  C/P,P
FINAL GATE [@FORGOTTEN REALMS LAST MYTHAL #3]  BAKER, RICHARD  P
FINAL IMPACT [AXIS OF TIME 3]  BIRMINGHAM, JOHN  LP
FINAL IMPACT  NAVARRO, YVONNE  C/P,P
FINAL KEY, THE [SKOLIAN TRIAD 2]  ASARO, CATHERINE  H
FINAL NEXUS, THE [STAR TREK]  DEWEESE, GENE  C/P:2,P
FINAL NIGHTMARE, THE [@ROBOTECH #9]  MCKINNEY, JACK  C/P,P
FINAL PLANET, THE  GREELEY, ANDREW M.  H,SFWA-P,P
FINAL PROGRAMME, THE [CORNELIUS 1]  MOORCOCK, MICHAEL  C/P:2,P
FINAL PROPHECY, THE [@STAR WARS NEW JEDI ORDER]  KEYES, J. GREGORY  C/P,P

TITLE  AUTHOR  CODE
FINAL QUEST, THE [PARSIVAL 3]  MONACO, RICHARD  C/P,SFWA,P
FINAL RECKONING [@BABYLON 5 PSI CORPS 3]  KEYES, J. GREGORY  P
FINAL RECKONINGS  BLOCH, ROBERT  LP
FINAL REFLECTION, THE [STAR TREK]  FORD, JOHN M.  C/P:3,P
FINAL SACRIFICE [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
FINAL SANCTION, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  P
FINAL SHADOWS  GRANT, CHARLES L.  HA
FINAL SOLUTION  PECK, RICHARD E.  H,SFWA
FINAL STAGE  FERMAN, EDWARD L./MALZBERG, BARRY N. SR-PA,C/PA
FINAL STORM, THE [WINGMAN 6]  MALONEY, MACK  P
FINAL WAR, THE  TRACY, LOUIS  H
FINAL WAR AND OTHER FANTASIES  O’DONNELL, K. M.  D23775
FIND THE CHANGELING  BENFORD, GREGORY/EKLUND, GORDON  C/P,P
FIND THE FEATHERED SERPENT  HUNTER, EVAN  H:2
FIND YOUR OWN TRUTH [SHADOWRUN 3/SECRETS OF POWER 3]  CHARRETTE, ROBERT N.  C/P,P
FINDER [BORDERLANDS 4]  BULL, EMMA  C/P:2,P
FINDER’S BANE [@FORGOTTEN REALMS HARPERS]  NOVAK, KATE/GRUBB, JEFF  P
FINDERS KEEPERS  SINCLAIR, LINNEA  P
FINDER’S SEEKERS [@GHATTI’S TALE #1]  GREENO, GAYLE  C/P,P
FINE AND PRIVATE PLACE, A  BEAGLE, PETER S.  C/P:2,P
FINE FRIGHTS  CAMPBELL, RAMSEY  SFWA-P
FINE PREY  WESTERFELD, SCOTT  P
FINEST CHOICE, THE  RABE, JEAN  H
FINEST CREATION, THE  RABE, JEAN  H
FINGER IN THE SKY AFFAIR, THE [@MAN FROM UNCLE #23]  LESLIE, PETER  P
FINGER IN THE SKY AFFAIR, THE (BRITISH) [@MAN FROM UNCLE #5]  LESLIE, PETER  SR-P
FINISH LINE  GOLDIN, STEPHEN  SR-P
FINITY  BARNES, JOHN  C/P,P,H
FINITY’S END [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
FINN AND HENGEST: THE FRAGMENT AND THE EPISODE  TOLKIEN, J. R. R.  R/REF-H
FINN MAC COOL  LLYWELYN, MORGAN  H,P
FIRE  RODGERS, ALAN  P
FIRE AND FLIGHT [@COMPLETE ELFQUEST #1]  PINI, WENDY/PINI, RICHARD  CX
FIRE AND FOG [CENOTAPH ROAD 5]  VARDEMAN, ROBERT E.  C/P,P
FIRE AND HEMLOCK  JONES, DIANA WYNNE  H,P
FIRE AND ICE [@OUTRIDER #2]  HARDING, RICHARD  P
FIRE AND ICE [@MOUNTAIN MADE OF LIGHT #2]  MYERS, EDWARD  P,SFWA-P
FIRE AND SWORD [@KEYS OF POWER #2]  BROWN, SIMON  P
FIRE AND THE ROSE, THE [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
FIRE ANGELS [DION HOLYHANDS 1]  ROUTLEY, JANE  H,LP
FIRE AT THE CENTER  PROCTOR, GEO. W.  C/P,P
FIRE BELOW [@SEAQUEST DSV 2]  COSTELLO, MATTHEW J.  P
FIRE BIKER [@TOM SWIFT (ARCHWAY) #9]  APPLETON, VICTOR  P
FIRE, BURN!  CARR, JOHN DICKSON  C/P,P
FIRE CROSSING [NETWORK-CONSORTIUM 2]  FRANKLIN, CHERYL J.  C/P:2,P
FIRE DANCER  MAXWELL, ANN  C/P:2,P
FIRE DRAGON, THE [DEVERRY 11/DRAGON MAGE 3]  KERR, KATHARINE  P
FIRE DUKE, THE [@KEEPERS OF THE HIDDEN WAYS 1]  ROSENBERG, JOEL  P
FIRE-EATER, THE  GOULART, RON  C/P:2,P
FIRE FROM THE WINE-DARK SEA  SUCHARITKUL, SOMTOW  LP
FIRE GET [TAORMIN 1]  FRANKLIN, CHERYL J.  C/P:2,P
FIRE-HUNTER  KJELGAARD, JIM  C/P,P
FIRE IN A FARAWAY PLACE [VERESHCHAGIN 2]  FREZZA, ROBERT  P
FIRE IN HIS HANDS, THE [KING WITHOUT A THRONE 1/RAGNARSON 1]  COOK, GLEN  C/P,P
FIRE IN THE ABYSS  GORDON, STUART  C/P,P
FIRE IN THE ASHES [ASHES 2]  JOHNSTONE, WILLIAM W.  P
FIRE IN THE BLOOD [@VAMPIRE FILES #5]  ELROD, P. N.  C/P:2,P
FIRE IN THE HEAVENS  SMITH, GEORGE O.  DD375
FIRE IN THE MIST [FAIA RISSEDOTTE 1]  LISLE, HOLLY  C/P:2,SFWA-P,P
FIRE IN THE SKY [@SHADOWSONG #1]  CLAYTON, JO  C/P:2,P
FIRE IN THE SKY, A  KENDRICK, WALTER  P
FIRE IN THE SKY [@PLANET BUILDERS #10]  TALLIS, ROBYN  P
FIRE IN THE SKY  WOLF, CHRIS L./MAIKOWSKI, MICHAEL F.  C/P,P
FIRE IN THE SUN, A [MARID AUDRAN 2]  EFFINGER, GEORGE ALEC  C/P,P,H
FIRE LANCE  MACE, DAVID  C/P,P
FIRE LOGIC [ELEMENTAL LOGIC 1]  MARKS, LAURIE J.  H
FIRE LORD [TAORMIN 2]  FRANKLIN, CHERYL J.  C/P:2,P
FIRE MARGINS [SHOLAN ALLIANCE 3]  NORMAN, LISANNE  P
FIRE ME UP [AISLING GREY 2]  MACALISTER, KATIE  C/P,P
FIRE MUSIC [TIGUAFAYA 2]  GRAY, JULIA  C/P,P
FIRE OF THE WITCHES  LONG, LYDA BELKNAP  P
FIRE ON HIGH [@STAR TREK NEW FRONTIER #6]  DAVID, PETER  P
FIRE ON THE BORDER  O’DONNELL, KEVIN JR.  C/P,P
FIRE ON THE MOUNTAIN  BISSON, TERRY  H,P

TITLE  AUTHOR  CODE
FIRE ON THE WATER [LONE WOLF 2]  DEVER, JOE  P
FIRE OPAL, THE [LOST CONTINENT 2]  ASARO, CATHERINE  LP
FIRE PAST THE FUTURE  MAINE, CHARLES ERIC  C/P:2,P
FIRE PATTERN  SHAW, BOB  C/P,P
FIRE PLANET [STAR COMMANDOS 7]  GRIFFIN, P. M.  P
FIRE QUEEN, THE  HOLLAND, JACK  P,SFWA-P
FIRE ROSE, THE  LACKEY, MERCEDES  C/P:2,P
FIRE SANCTUARY [NUALA 1]  KIMBRIEL, KATHERINE ELISKA  C/P,P
FIRE SEA [@DEATH GATE #3]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
FIRE SHIP [STAR TREK CAPTAIN’S TABLE 4/STAR TREK VOYAGER]  CAREY, DIANE  C/P,P
FIRE! SURVIVAL AND PREVENTION  BLAIR, WILLIAM  [C/RAN-P]
FIRE SWORD, THE  MARTINE-BARNES, ADRIENNE  C/P:2,P
FIRE THRONE MOUNTAIN  BIXBY, E. REW  P
FIRE TIME  ANDERSON, POUL  H:2,P,SFWA:2
FIRE UPON THE DEEP, A [DEEP 1]  VINGE, VERNOR  H:2,P
FIRE WARRIOR [@WARHAMMER 40K]  SPURRIER, SIMON  P
FIRE WATCH  WILLIS, CONNIE  C/P:2,SFWA:2,P
FIRE WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
FIRE WHEN IT COMES, THE  GODWIN, PARKE  H,SFWA
FIREBALL  MAYHEW, VIC/LONG, DOUG  P
FIREBIRD  HARNESS, CHARLES L.  H,P
FIREBIRD  LACKEY, MERCEDES  C/P,P,H
FIREBIRD [FIREBIRD 1]  TYERS, KATHY  C/P,P
FIREBIRD TRILOGY, THE [FIREBIRD 1,2,3]  TYERS, KATHY  H
FIREBIRDS  NOVEMBER, SHARON  HA
FIREBIRDS RISING  NOVEMBER, SHARON  HA
FIREBIRD’S VENGEANCE, THE [ISAVALTA 3]  ZETTEL, SARAH  H
FIREBRAND, THE  BRADLEY, MARION ZIMMER  H:2,P
FIREBRINGER TRILOGY, THE [FIREBRINGER 1,2,3]  PIERCE, MEREDITH ANN  H
FIREBUG  BLOCH, ROBERT  SFWA-P
FIRECHILD  WILLIAMSON, JACK  C/P,SFWA-P,P,H
FIRECLOWN, THE  MOORCOCK, MICHAEL  SR-P,C/P:2
FIRECODE  YARBRO, CHELSEA QUINN  C/P,P
FIREDANCE [ASHLU]  BAKER, SCOTT  C/P,P
FIREDANCE [AUBRY KNIGHT]  BARNES, STEVEN  H
FIREDRAKE [DRAGONREALM 1]  KNAAK, RICHARD A.  C/P:2,P
FIREFIGHT 2000  ING, DEAN  C/P,P
FIREFLOOD AND OTHER STORIES  MCINTYRE, VONDA N.  C/P,P,H:2
FIREFLY  ANTHONY, PIERS  H,P
FIREFLY  STABLEFORD, BRIAN M.  LP
FIREFOX  THOMAS, CRAIG  P
FIREFOX DOWN!  THOMAS, CRAIG  P
FIREHAND [TIME TRADERS 5]  NORTON, ANDRE/GRIFFIN, P. M.  H:2
FIRELORD  GODWIN, PARKE  C/P,P,H
FIREMASK [LAST LEGION 2]  BUNCH, CHRIS  C/P,P
FIRES OF AGGAR [AGGAR 2]  WOLFE, CHRIS ANNE  LP
FIRES OF AZEROTH [MORGAINE 3]  CHERRYH, C. J.  C/P:2,P
FIRES OF COVENTRY, THE [SECOND COMMONWEALTH 2]  SHELLEY, RICK  P
FIRES OF EDEN  SIMMONS, DAN  C/P,P,H
FIRES OF FOREVER  OLIVER, CHAD  B
FIRES OF HEAVEN, THE [WHEEL OF TIME 5]  JORDAN, ROBERT  C/P,SFWA,P,H:2
FIRES OF LAN-KERN, THE  TREMAYNE, PETER  H
FIRES OF NUALA [NUALA 2]  KIMBRIEL, KATHERINE ELISKA  C/P:2,P
FIRES OF PARATIME, THE  MODESITT, L. E. JR.  C/P:2,P,H
FIRES OF SCORPIO [DRAY PRESCOT 29]  AKERS, ALAN BURT  P
FIRES OF THE FAITHFUL [ELIANA 1]  KRITZER, NAOMI  P
FIRES OF THE PAST  JORDAN, ANNE DEVEREAUX  SFWA
FIRES OF WINDAMEIR, THE [WINDAMEIR CIRCLE 1]  HANCOCK, NIEL  P
FIRE’S STONE, THE [HUFF 4]  HUFF, TANYA  C/P,P
FIRESHAPER’S DOOM [OTHERWORLD 2]  DEITZ, TOM  C/P:2,P
FIRESHIP  VINGE, JOAN D.  C/P,P,H:2
FIRESHIP/MOTHER AND CHILD  VINGE, JOAN D.  P
FIRESHIPS [REACHES 3]  DRAKE, DAVID  C/P,P,H
FIRESIGN THEATRE’S BIG BOOK OF PLAYS, THE  FIRESIGN THEATRE  VLP
FIRESIGN THEATRE’S BIG MYSTERY JOKE BOOK, THE  FIRESIGN THEATRE  VLP
FIRESTAR [STARS 1]  FLYNN, MICHAEL  H:2,P
FIRESTARTER  KING, STEPHEN  H:2,P
FIRESTORM [WEATHER WARDEN 5]  CAINE, RACHEL  P
FIRESTORM [STAR TREK]  GRAF, L. A.  C/P:2,P
FIRESTORM [@STARFIST 12]  SHERMAN, DAVID/CRAGG, DAN  H
FIRETHORN  MICKLEM, SARAH  LP
FIREWALK  LOGSTON, ANNE  P
FIREWORKS  CARTER, ANGELA  C/P,LP
FIREWORKS  MOORE, JAMES A.  LP

The MITSFS Titledex  September 24, 2007  Page 128
TITLE  AUTHOR  CODE
FIRST ACTION [@STARSEA INVADERS 1]  STINE, G. HARRY  C/P,P
FIRST AMENDMENT, THE [@STARGATE SG-1 (MCCONNELL) 3]  MCCONNELL, ASHLEY  P
FIRST AMERICAN INTO SPACE  SILVERBERG, ROBERT  [C/RAN-P]
FIRST AND FINAL RITES  LEWITT, SHARIANN  C/P:2,P
FIRST AND ONLY [WARHAMMER 40K GAUNT’S GHOSTS 1]  ABNETT, DAN  P
FIRST BATTLE [SKYRIDER 2]  MICHAELS, MELISA C.  P,SFWA-P
FIRST BETRAYAL, THE  BRAY, PATRICIA  P
FIRST BLOOD [ORCS FIRST BLOOD 1,2,3]  NICHOLLS, STAN  H
FIRST BOOK OF SWORDS, THE [@SWORDS #1]  SABERHAGEN, FRED  C/P,LP,H
FIRST BOY ON THE MOON  HICKS, CLIFFORD B.  H:2
FIRST CASUALTY, THE [SOCIETY OF HUMANITY/UNITY PARTY 1]  MOSCOE, MIKE  C/P,P
FIRST CASUALTY, THE [@HERCULES THE LEGENDARY JOURNEYS]  SEIDMAN, DAVID L.  P
FIRST CHANNEL [SIME/GEN 3]  LORRAH, JEAN/LICHTENBERG, JACQUELINE C/P,P,H,SFWA
FIRST CHRONICLES OF DRUSS THE LEGEND, THE [DRENAI]  GEMMELL, DAVID  C/P,P
FIRST CITIZEN  THOMAS, THOMAS T.  C/P,SFWA-P,P
FIRST CONTACT  KNIGHT, DAMON  C/PA,PA
FIRST CONTACT?  WALTERS, HUGH  H
FIRST CONTACTS  LEINSTER, MURRAY  H
FIRST CONTRACT  COSTIKYAN, GREG  H
FIRST CYCLE  PIPER, H. BEAM/KURLAND, MICHAEL  C/P:2,P
FIRST DAWN [@LOST MILLENNIUM 1]  MOSCOE, MIKE  C/P,P

FIRST DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-P,C/REF-P

FIRST DUELIST, THE  ETHERIDGE, RUTLEDGE  SFWA-P
FIRST EDITIONS OF ANDRE NORTON, THE  TURNER, DAVID G.  R/IND
FIRST FALL [PERN]  MCCAFFREY, ANNE  C/P,P,H:2
FIRST FAMILY, THE [AMTRAK WARS 2]  TILLEY, PATRICK  C/P:2,P
FIRST FLIGHT  KNIGHT, DAMON  PA,C/PA
FIRST FRONTIER [STAR TREK]  CAREY, DIANE/KIRKLAND, DR. JAMES I.  C/P:2,P
FIRST FRONTIER [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
FIRST HE DIED  SIMAK, CLIFFORD D.  C/P:2,P
FIRST HEROES, THE  TURTLEDOVE, HARRY/DOYLE, NOREEN  LPA
FIRST IMMORTAL, THE  HALPERIN, JAMES L.  H,P
FIRST KING OF SHANNARA [SHANNARA 0]  BROOKS, TERRY  H:2,P
FIRST LANDING  ZUBRIN, ROBERT  H
FIRST LENSMAN [LENSMEN 2]  SMITH, E. E.  H:2,P
FIRST LIGHT  ACKROYD, PETER  LP
FIRST MEETINGS IN THE ENDERVERSE [ENDER]  CARD, ORSON SCOTT  H
FIRST MEN IN THE MOON, THE  WELLS, H. G.  C/P,SR-P,H
FIRST MEN TO THE MOON  VON BRAUN, WERNHER  H
FIRST NAME, THE [@TWILIGHT OF THE GODS #1]  SCHMIDT, DENNIS  C/P:2,P
FIRST NAMED, THE [SERVANTS OF ARK 1]  WYLIE, JONATHAN  C/P,P
FIRST OMNI BOOK OF SCIENCE FICTION, THE  DATLOW, ELLEN  C/PA,PA
FIRST ON MARS  GORDON, REX  C/P:2,P
FIRST ON THE MOON  SUTTON, JEFF  C/P:2,P
FIRST ON THE MOON  WALTERS, HUGH  C/P:2,P
FIRST PERSON, PECULIAR  SHERRED, T. L.  C/P:2,P
FIRST POWER PLAY [BUCK ROGERS INNER PLANETS 1]  MILLER, JOHN  C/P:2,P
FIRST PROTECTOR, THE [EARTH FINAL CONFLICT]  WHITE, JAMES  H,P
FIRST RIDER’S CALL [GREEN RIDER 2]  BRITAIN, KRISTEN  H,P
FIRST STEP OUTWARD  HOSKINS, ROBERT  C/PA,PA
FIRST STONE, THE [LAST RUNE 6]  ANTHONY, MARK  P
FIRST STRIKE [STAR TREK INVASION! 1]  CAREY, DIANE  C/P,P
FIRST TEST [@PROTECTOR OF THE SMALL 1]  PIERCE, TAMORA  H
FIRST THROUGH TIME  GORDON, REX  C/P:2,P
FIRST TO AWAKEN, THE  HICKS, GRANVILLE/BENNETT, RICHARD M.  H
FIRST TO FIGHT [@STARFIST 1]  SHERMAN, DAVID/CRAGG, DAN  C/P:2,P
FIRST TO THE STARS  GORDON, REX  C/P:2,P
FIRST TRIAD, THE [@LOST SWORDS #1,2,3]  SABERHAGEN, FRED  H:2
FIRST TRUTH [FIRST TRUTH 1]  COOK, DAWN  P
FIRST TWO LIVES OF LUK AS-KASHA, THE  ALEXANDER, LLOYD  P

FIRST VIRTUE, THE [@STAR TREK TNG DOUBLE HELIX 6] FRIEDMAN, MICHAEL JAN/ GOLDEN, CHRISTIE

P

FIRST VOYAGES KNIGHT, DAMON/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

PA

FIRST WARNING [ACORNA’S CHILDREN 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

FIRST WORLD FANTASY AWARDS  WILSON, GAHAN  SFWA
FIRST WORLD FANTASY CONVENTION: THREE AUTHORS REMEMBER, THE  MICHAUD, MARC A.  R/REF-P
FIRST WORLD OF IF, THE  QUINN, JAMES/WULFF, EVE  SR-PA
FIRST, YOU FIGHT [@TRAVELER #1]  DRUMM, D. B.  P
FIRSTBORN [@STARCRAFT DARK TEMPLAR #1]  GOLDEN, CHRISTIE  P
FIRSTBORN [@DRAGONLANCE ELVEN NATIONS #1]  THOMPSON, PAUL B./CARTER, TONYA R.  P
FIRSTFLIGHT [NICOLE SHEA 1]  CLAREMONT, CHRIS  C/P:2,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 129
TITLE  AUTHOR  CODE
FISH DINNER IN MEMISON, A [ZIMIAMVIA 2]  EDDISON, E. R.  SR-H,C/P:2,P
FISHERMAN OF THE INLAND SEA, A  LE GUIN, URSULA K.  P
FISHERMAN OF THE INLAND SEA, A (PRE-PUBLICATION EDITION)  LE GUIN, URSULA K.  SR-H,H
FISHERMAN’S CURSE, THE  EASTON, M. COLEMAN  C/P,P
FISHERMAN’S HOPE [NICHOLAS SEAFORT 4]  FEINTUCH, DAVID  C/P:2,P
FISSION  HUNT, GILL  SR-P,C/P
FISTANDANTILUS REBORN [@DRAGONLANCE LOST LEGENDS #2]  NILES, DOUGLAS  P
FISTFUL OF CHARMS, A [RACHEL MORGAN 4]  HARRISON, KIM  C/P,P
FISTFUL OF DATA, A [@SHADOWRUN]  DEDMAN, STEPHEN  P
FISTFUL OF DIGITS, A  HODDER-WILLIAMS, CHRISTOPHER  SR-P,C/P
FISTFUL OF SKY, A  HOFFMAN, NINA KIRIKI  P
FITCHER’S BRIDES  FROST, GREGORY  LP
FITTEST, THE  MCINTOSH, J. T.  H:2
FITZPATRICK’S WAR  JUDSON, THEODORE  H
FIVE AGAINST ARLANE  PURDOM, TOM  DH022
FIVE AGAINST VENUS  LATHAM, PHILIP  H:2
FIVE CHILDREN AND IT  NESBIT, E.  C/P:2,P
FIVE DOCTORS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
FIVE FATES  DOUBLEDAY BOOKS  HA
FIVE FATES  WARNER BOOKS  C/PA,PA
FIVE FOR INFINITY  BARKER, THOMAS W.  P
5 GALAXY SHORT NOVELS  GOLD, HORACE L.  C/PA,PA
FIVE GOLD BANDS, THE  VANCE, JACK  C/P,DF185,P,D16640
FIVE HUNDRED YEARS AFTER [DRAGAERA KHAAVREN 2]  BRUST, STEVEN  C/P,P,H
FIVE-ODD  CONKLIN, GROFF  PA,C/PA
FIVE OF CUPS  KIERNAN, CAITLIN R.  H
FIVE SCIENCE FICTION NOVELS  GREENBERG, MARTIN  HA
FIVE STEPS TO TOMORROW  BINDER, EANDO  C/P,P
5 TALES FROM TOMORROW  DIKTY, T. E.  PA,C/PA
FIVE THOUSAND MILES UNDERGROUND  ROCKWOOD, ROY  H
FIVE TO TWELVE  COOPER, EDMUND  H:2,P
FIVE-TWELFTHS OF HEAVEN [SILENCE LEIGH 1]  SCOTT, MELISSA  C/P:2,P
5 UNEARTHLY VISIONS  CONKLIN, GROFF  PA,C/PA
FIVE VICTORIAN GHOST NOVELS  BLEILER, EVERETT F.  LPA
FIVE WAY SECRET AGENT, THE/MERCENARY FROM TOMORROW  REYNOLDS, MACK  C/P:2,P
FIVE WEEKS IN A BALLOON  FOX, GARDNER F.  P
FIVE WEEKS IN A BALLOON  VERNE, JULES  H
FIZE OF THE GABRIEL RATCHETS [@WINDHOVER TAPES #3]  NORWOOD, WARREN  C/P:2,P
FLAG FULL OF STARS, A [STAR TREK]  FERGUSON, BRAD  C/P:3,P
FLAG IN EXILE [HONOR HARRINGTON 5]  WEBER, DAVID  C/P,P,H
FLAME AND THE SHADOW-EATER  WEAVER, HENRIETTA  SR-H
FLAME BREATHERS, THE [@AVENGER #12]  ROBESON, KENNETH  C/P:2,P
FLAME IN BYZANTIUM, A [SAINT-GERMAIN OLIVIA 1]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P

FLAME IN HALI, A [@CLINGFIRE #3] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

FLAME IS GREEN, THE  LAFFERTY, R. A.  H
FLAME KEY, THE [KEYS TO PARADISE 1]  MORAN, DANIEL  C/P,P
FLAME OF IRIDAR, THE  CARTER, LIN  SR-DB0759
FLAME TREE PLANET  ELWOOD, ROGER  SR-PA
FLAME UPON THE ICE, THE [ICE 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
FLAME WINDS  PAGE, NORVELL W.  C/P:2,P
FLAMES, THE  STAPLEDON, OLAF  SR-H
FLAMES OF THE DRAGON [BROTHERS OF THE DRAGON 2]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
FLAMESONG [TEKUMEL 2]  BARKER, M. A. R.  C/P,P
FLAMEWEAVER  BALL, MARGARET  C/P,SFWA-P,P
FLAMING ARROW, THE [@STAR TREK NEW EARTH #4]  OLTION, KATHY/OLTION, JERRY  C/P,P
FLAMING FALCONS, THE [DOC SAVAGE #30]  ROBESON, KENNETH  C/P:2,@P
FLAMING MOUNTAIN, THE [@RICK BRANT #17]  BLAINE, JOHN  H
FLANDRY [DOMINIC FLANDRY]  ANDERSON, POUL  P,SFWA-P
FLANDRY OF TERRA [DOMINIC FLANDRY]  ANDERSON, POUL  C/P,P,H
FLARE  ZELAZNY, ROGER/THOMAS, THOMAS T.  C/P:2,SFWA-P,P
FLARE STAR  DOWNING, PAULA E.  P
FLASH [ARCHFORM BEAUTY 2]  MODESITT, L. E. JR.  H
FLASH GORDON  COVER, ARTHUR BYRON  P
FLASHBACK [STAR TREK VOYAGER]  CAREY, DIANE  C/P:2,P
FLASHFIRE [@STARFIST 11]  SHERMAN, DAVID/CRAGG, DAN  H
FLASHFORWARD  SAWYER, ROBERT J.  C/P,P,H
FLASHING SWORDS! #1  CARTER, LIN  PA,C/PA
FLASHING SWORDS! #2  CARTER, LIN  C/PA,PA
FLASHING SWORDS! #3: WARRIORS AND WIZARDS  CARTER, LIN  PA,C/PA
FLASHING SWORDS! #4: BARBARIANS AND BLACK MAGICIANS  CARTER, LIN  PA,C/PA
FLASHING SWORDS! #5: DEMONS AND DAGGERS  CARTER, LIN  PA,HA
FLASHMAN [FLASHMAN]  FRASER, GEORGE MACDONALD  P
FLASHMAN AT THE CHARGE [FLASHMAN]  FRASER, GEORGE MACDONALD  P

The MITSFS Titledex  September 24, 2007  Page 130
TITLE  AUTHOR  CODE
FLASHPOINT [@BATTLETECH 30]  COLEMAN, LOREN L.  P
FLATLAND [FLATLAND]  ABBOTT, EDWIN A.  H:2,LP
FLATLANDER [KNOWN SPACE]  NIVEN, LARRY  P
FLATTERLAND  STEWART, IAN  H
FLEDGLING  BUTLER, OCTAVIA E.  H
FLEET, THE [@FLEET #1]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P
FLEET ACTION [@WING COMMANDER 3]  FORSTCHEN, WILLIAM R.  P,SFWA-P
FLEET OF STARS, THE [HARVEST OF STARS 4]  ANDERSON, POUL  C/P:2,P
FLEET OF THE DAMNED [STEN 4]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
FLEET OF THE SPRINGERS, THE [@PERRY RHODAN #22]  ACE  PA,C/PA
FLESH  FARMER, PHILIP JOSE  C/P,SR-P,H
FLESH/LORD TYGER  FARMER, PHILIP JOSE  P
FLESH  LAYMON, RICHARD  SFWA-P
FLESH AND GOLD [FLESH AND GOLD 1]  GOTLIEB, PHYLLIS  H,LP
FLESH AND SILVER  BURNS, STEPHEN L.  C/P:2,P
FLESH AND SPIRIT  BERG, CAROL  LP
FLESH HUNTERS [@UFO #1]  MIALL, ROBERT  P
FLESH HUNTERS, THE  WYATT, LEE  SR-L
FLESH IN THE FURNACE, THE  KOONTZ, DEAN R.  C/P:2,P

FLESH MASK, THE [@VANCE INTEGRAL EDITION #10]/ STRANGE PEOPLE, QUEER NOTIONS [@VANCE INTEGRAL EDITION #10]/ BIRD ISLAND [@VANCE INTEGRAL EDITION #10] VANCE, JACK

SR-H

FLESHPOTS OF SANSATO, THE  TEMPLE, WILLIAM F.  H
FLESHWORKS [@NECROMUNDA]  SOULBAN, LUCIEN  P
FLEXING THE WARP [@WINDHOVER TAPES #2]  NORWOOD, WARREN  C/P:2,P
FLIERS OF ANTARES [DRAY PRESCOT 8]  AKERS, ALAN BURT  P
FLIES FROM THE AMBER  MCCARTHY, WIL  P
FLIES OF MEMORY, THE  WATSON, IAN  H
FLIGHT  BONTA, VANNA  LP
FLIGHT, THE  RUNYAN, C. F.  LP
FLIGHT FROM BERBORA [@BERBORA #2]  SIROTA, MIKE  P
FLIGHT FROM NEVERYON  DELANY, SAMUEL R.  P
FLIGHT FROM REBIRTH  MCINTOSH, J. T.  C/P:2,P

FLIGHT FROM TARKIHL [@ATLAN #2]/ MENACE OF ATOMIGEDDON (SPECIAL) [PERRY RHODAN] ACE

PA

FLIGHT FROM THE DARK [LONE WOLF 1]  DEVER, JOE/CHALK, GARY  P
FLIGHT FROM THE GRAVE [@G-8 #6]  HOGAN, ROBERT J.  P
FLIGHT FROM TIME ONE  ROMANO, DEANE  H,SFWA
FLIGHT FROM YESTERDAY  WILLIAMS, ROBERT MOORE  DF223
FLIGHT IN YIKTOR [MOON SINGER 3]  NORTON, ANDRE  H:2,P
FLIGHT INTO FEAR [DOC SAVAGE]  ROBESON, KENNETH  P
FLIGHT INTO SPACE  LEONARD, JONATHAN EDWARD  [C/RAN-P:2]
FLIGHT INTO SPACE  WOLLHEIM, DONALD A.  SR-PA,C/PA
FLIGHT INTO THE UNKNOWN  BARTHOLOMEW, BARBARA  P
FLIGHT INTO YESTERDAY  HARNESS, CHARLES L.  H
FLIGHT OF EXILES  BOVA, BEN  H
FLIGHT OF FEAR [LAND OF THE GIANTS]  RATHJEN, CARL HENRY  H
FLIGHT OF HONOR [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  P
FLIGHT OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 2]  TEPPER, SHERI S.  C/P:2,P
FLIGHT OF MICHAEL MCBRIDE, THE  SNYDER, MIDORI  C/P,P
FLIGHT OF THE BAMBOO SAUCER, THE  GORDON, FRITZ  C/P,P
FLIGHT OF THE BAT  GORDON, DONALD  H
FLIGHT OF THE DRAGONFLY, THE  FORWARD, ROBERT L.  C/P:2,SFWA,LP
FLIGHT OF THE DYING SUN [@EBERRON HEIRS OF ASH #2]  WULF, RICH  P
FLIGHT OF THE EISENSTEIN, THE [@WARHAMMER 40K HORUS HERESY 4]  SWALLOW, JAMES  P
FLIGHT OF THE ENDEAVOR, THE  FARRELL, JACKSON T.  P
FLIGHT OF THE FALCON [MECHWARRIOR DARK AGE/BATTLETECH]  MILAN, VICTOR  P
FLIGHT OF THE FALLEN [@DRAGONLANCE LINSHA #2]  HERBERT, MARY H.  P
FLIGHT OF THE HESPER  HAY, GEORGE  SR-P
FLIGHT OF THE HORSE, THE  NIVEN, LARRY  C/P:2,P
FLIGHT OF THE NIGHTHAWKS [DARKWAR 1]  FEIST, RAYMOND E.  H
FLIGHT OF THE OLD DOG [MAC MCLANAHAN 1]  BROWN, DALE  P
FLIGHT OF THE RAVEN [CHEYSULI 7]  ROBERSON, JENNIFER  C/P:2,P
FLIGHT OF THE STONEMAN’S SON, THE  MUNSEY, TERENCE  P

FLIGHT OF VENGEANCE [WITCH WORLD TURNING 2] NORTON, ANDRE/GRIFFIN, P. M./

SCHAUB, MARY H.

P

FLIGHT TO FRANCE  VERNE, JULES  H
FLIGHT TO LUCIFER, THE  BLOOM, HAROLD P.  H,LP
FLIGHT TO OPAR  FARMER, PHILIP JOSE  C/P,P
FLIGHT TO TERROR  ELDER, MICHAEL  C/P,P
FLIGHT TO THE LONESOME PLACE  KEY, ALEXANDER  H
FLIGHT TO THE SAVAGE EMPIRE [SAVAGE EMPIRE 4]  LORRAH, JEAN/HOWLETT, WINSTON A.  C/P,P
FLIGHT TO THLASSA MEY [THLASSA MEY 1]  MCCARTY, DENNIS  C/P,P
FLIGHT, VOLUME ONE  IMAGE COMICS  CX

The MITSFS Titledex  September 24, 2007  Page 131
TITLE  AUTHOR  CODE
FLIGHTLESS FALCON  REICHERT, MICKEY ZUCKER  H
FLIGHTS  SARRANTONIO, AL  HA

FLIGHTS OF FANTASY HAUPT, HANNA BEATE/HESTON, LILLA/ LITTELL, JOY/SOLOTAROFF, SARAH

LP

FLIGHTS OF FANTASY  LACKEY, MERCEDES  C/PA,PA

FLIGHTS OF FANTASY LITTELL, J./HESTON, L./HAUPT, H. B./

SOLOTAROFF, S.

HA

FLINT THE KING [@DRAGONLANCE PRELUDES II 2]  KIRCHOFF, MARY/NILES, DOUGLAS  C/P,P
FLINX IN FLUX [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
FLINX’S FOLLY [HUMANX]  FOSTER, ALAN DEAN  P
FLINX’S FOLLY (PROOF) [HUMANX FLINX]  FOSTER, ALAN DEAN  SR-LP
FLIP SIDE OF MEMORY  TABLER, MADELINE  P
FLOATER FACTOR [SKYRIDER 5]  MICHAELS, MELISA C.  P,SFWA-P
FLOATING CASTLE, THE [@KING’S QUEST 1]  MILLS, CRAIG  P
FLOATING CITY, A  VERNE, JULES  H
FLOATING DRAGON  STRAUB, PETER  H,P
FLOATING GAME, THE [@AVENGERS 1]  GARFORTH, JOHN  C/P:2,P
FLOATING GODS, THE [VIRICONIUM]  HARRISON, M. JOHN  P
FLOATING ISLAND  HAYDON, ELIZABETH  H
FLOATING WORLDS  BOWEN, GARY  B
FLOATING WORLDS  HOLLAND, CECELIA  C/P,P,H
FLOATING ZOMBIE, THE  JONES, D. F.  C/P,P
FLOOD, THE  CREASEY, JOHN  P
FLOOD AND FLAME  VERNE, JULES  H
FLOOD TIDE [@MEROVINGEN NIGHTS #6]  CHERRYH, C. J.  C/PA,PA
FLOODGATE, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #2]  CUNNINGHAM, ELAINE  P
FLORIANS, THE [DAEDALUS 1]  STABLEFORD, BRIAN M.  C/P,P
FLORIDA PROJECT, THE [V]  SULLIVAN, TIM  C/P,P
FLOUNDER, THE  GRASS, GUNTER  P
FLOW MY TEARS, THE POLICEMAN SAID  DICK, PHILIP K.  C/P,P,H,SFWA:2
FLOWER OF DORADIL  RACKHAM, JOHN  D24100
FLOWER OF ETERNITY, THE  HILTON, MARGERY  P
FLOWERCRASH  PALMER, STEPHEN  LP
FLOWERDUST [DIVINE ENDURANCE 2]  JONES, GWYNETH  LP
FLOWERS FOR ALGERNON  KEYES, DANIEL  C/P:2,P
FLOWERS FOR ALGERNON AND OTHER STORIES  MILLS, ROBERT P.  C/PA,PA
FLOWERS FROM THE MOON AND OTHER LUNACIES  BLOCH, ROBERT  H
FLUCHT DER ANDROIDEN [@PERRY RHODAN #147]  KNEIFEL, HANS  FPR
FLUGER, THE  PISERCHIA, DORIS  P
FLUKE  HERBERT, JAMES  P
FLUKE  MOORE, CHRISTOPHER  LP
FLUTE SONG MAGIC  SHETTLE, ANDREA  P
FLUX [XEELEE]  BAXTER, STEPHEN  C/P,P,H
FLUX  CARD, ORSON SCOTT  P
FLUX  GOULART, RON  C/P:2,P
FLY BY NIGHT  HARDINGE, FRANCES  H
FLYER  KIMBERLY, GAIL  C/P,P
FLYING DUTCH  HOLT, TOM  H,P
FLYING EYES, THE  HOLLY, JOAN HUNTER  C/P,P
FLYING FISH, THE  BERESFORD, LESLIE  B
FLYING GOBLIN, THE [DOC SAVAGE #90]  ROBESON, KENNETH  C/P,@P
FLYING IN PLACE  PALWICK, SUSAN  H
FLYING SAUCER CONSPIRACY, THE  KEYHOE, DONALD E.  [C/RAN-H]
FLYING SAUCER FROM MARS  ALLINGHAM, CEDRIC  [C/RAN-H]
FLYING SAUCER GAMBIT, THE [@AGENT OF TERRA #1]  MADDOCK, LARRY  C/P:2,P

FLYING SAUCERS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

FLYING SAUCERS — HERE AND NOW!  EDWARDS, FRANK  [C/RAN-H]
FLYING SAUCERS: HOAX OR REALITY?  STANTON, JEROME  [C/RAN-P]
FLYING SAUCERS IN FACT AND FICTION  SANTESSON, HANS STEFAN  PA
FLYING SAUCERS — SERIOUS BUSINESS  EDWARDS, FRANK  [C/RAN-H,C/RAN-P]

FLYING SAUCERS:

THE STARTLING EVIDENCE OF THE INVASION FROM OUTER SPACE LORENZEN, CORAL E.

[C/RAN-P]

FLYING SORCERERS, THE  GERROLD, DAVID/NIVEN, LARRY  C/P:2,P
FLYING SORCERERS, THE  HAINING, PETER  PA
FLYING STINGAREE, THE [@RICK BRANT #18]  BLAINE, JOHN  H
FLYING TO VALHALLA  PELLEGRINO, CHARLES  C/P,P,H
FLYING TOWARDS PLUTO (CHINESE)  CHI CHING CHOU  FP
FLYING VISIT, THE  FLEMING, PETER  H
FLYING WARLORD, THE [CONRAD STARGARD 4]  FRANKOWSKI, LEO  C/P:2,P
FLYING WINDMILL, THE  ADAMS, EUSTACE L.  H
FLYING YORKSHIREMAN, THE  KNIGHT, ERIC  C/P,P,H
FOCUS ON THE HORROR FILM  HUSS, ROY/ROSS, T. J.  R/REF-H
FOCUS ON THE SCIENCE FICTION FILM  JOHNSON, WILLIAM  R/REF-H

TITLE  AUTHOR  CODE
FOE OF BARNABAS COLLINS, THE  ROSS, MARILYN  P
FOG, THE  ETCHISON, DENNIS  P
FOG, THE  HERBERT, JAMES  P
FOLD A BANANA  ERSKINE, JIM/MORAN, GEORGE  [C/RAN-H]
FOLD IN THE TENT OF THE SKY, A  HALE, MICHAEL  H
FOLK OF THE AIR, THE  BEAGLE, PETER S.  C/P:2,SFWA,P
FOLK OF THE FRINGE, THE  CARD, ORSON SCOTT  C/P:2,P,H
FOLKLORE IN THE ENGLISH AND SCOTTISH BALLADS  WIMBERLY, LOWRY CHARLES  [C/RAN-H]
FOLLOW THE DRUM  NORTON, ANDRE  P
FOLLOW THE WHALES  BIEMILLER, CARL L.  SFWA
FOLSOM FLINT, THE  KELLER, DAVID H.  SR-H
FOND FAREWELL TO DYING, A  LOGSDON, SYD  P
FONDAZIONE E TERRA (ITALIAN) [FOUNDATION 5]  ASIMOV, ISAAC  FH
FONTANA BOOK OF GREAT GHOST STORIES, THE  AICKMAN, ROBERT  PA
3RD FONTANA BOOK OF GREAT GHOST STORIES, THE  AICKMAN, ROBERT  SR-PA,C/PA
4TH FONTANA BOOK OF GREAT GHOST STORIES, THE  AICKMAN, ROBERT  PA
8TH FONTANA BOOK OF GREAT GHOST STORIES, THE  AICKMAN, ROBERT  SR-PA
FONTANA BOOK OF GREAT HORROR STORIES, THE  BERNARD, CHRISTINE  SR-PA
SECOND FONTANA BOOK OF GREAT HORROR STORIES, THE  BERNARD, CHRISTINE  C/PA,PA
THIRD FONTANA BOOK OF GREAT HORROR STORIES, THE  BERNARD, CHRISTINE  PA
FOURTH FONTANA BOOK OF GREAT HORROR STORIES, THE  BERNARD, CHRISTINE  PA
FIFTH FONTANA BOOK OF GREAT HORROR STORIES, THE  DANBY, MARY  C/PA,PA
FOOD FOR DEMONS  EVANS, E. EVERETT  SR-LP
FOOD OF DEATH: FIFTY-ONE TALES, THE  DUNSANY, LORD  LP
FOOD OF THE GODS, THE  WELLS, H. G.  C/P,P,H
FOOL FOR AN AGENT  CURTIS, RICHARD  R/REF-H
FOOL ME TWICE [ARCHONATE 2]  HUGHES, MATTHEW  C/P,P
FOOL MOON [@DRESDEN FILES #2]  BUTCHER, JIM  C/P:2,P
FOOL ON THE HILL  RUFF, MATT  P,SFWA
FOOLS  CADIGAN, PAT  C/P,P
FOOL’S ERRAND [TAWNY MAN 1]  HOBB, ROBIN  C/P,P
FOOLS ERRANT [ARCHONATE 1]  HUGHES, MATTHEW  P
FOOL’S FATE [TAWNY MAN 3]  HOBB, ROBIN  H
FOOLS OF TIME, THE  BARRETT, WILLIAM E.  C/P,P,H
FOOL’S RUN  MCKILLIP, PATRICIA A.  C/P,P,H
FOOL’S WAR  ZETTEL, SARAH  C/P:2,P
FOOP!  GENOA, CHRIS  LP
FOOTFALL  NIVEN, LARRY/POURNELLE, JERRY  C/P:2,P,H,SFWA
FOOTPRINTS OF GOD, THE  ILES, GREG  H
FOOTPRINTS OF THE DEAD [TWILIGHT 14]  CALLAHAN, JAY  P
FOOTPRINTS OF THUNDER [FOOTPRINTS OF THUNDER 1]  DAVID, JAMES F.  P
FOR A FEW DEMONS MORE [RACHEL MORGAN 5]  HARRISON, KIM  H:2
FOR CAMELOT’S HONOR  ZETTEL, SARAH  LP
FOR CROWN AND KINGDOM [SWORDS OF RAEMLLYN 6]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  P
FOR FEAR OF LITTLE MEN  BLACKBURN, JOHN  SR-P
FOR FEAR OF THE NIGHT  GRANT, CHARLES L.  SFWA-P
FOR KING & COUNTRY [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  H,P
FOR LOVE AND GLORY  ANDERSON, POUL  H:2
FOR LOVE OF EVIL [INCARNATIONS OF IMMORTALITY 6]  ANTHONY, PIERS  C/P:2,P
FOR LOVE OF MOTHER-NOT [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
FOR MORE THAN GLORY [LEGION OF THE DAMNED 5]  DIETZ, WILLIAM C.  H,P
FOR SPECIAL SERVICES [JAMES BOND 16]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
FOR TEXAS AND ZED  HUGHES, ZACH  C/P,P
FOR THE EMPEROR [@WARHAMMER 40K CIAPHAS CAIN 1]  MITCHELL, SANDY  P
FOR THE FLAG  VERNE, JULES  C/P,P,H
FOR THE GLORY [@PHOTON #1]  PETERS, DAVID  P
FOR THE TIME BEING  DESJARDIN, MARIE  LP
FOR THE WITCH OF THE MISTS [BRAN MAK MORN 3]  SMITH, DAVID C./TIERNEY, RICHARD L.  C/P,P
FOR THOSE WHO FELL [LEGION OF THE DAMNED 6]  DIETZ, WILLIAM C.  H
FOR US, THE LIVING  HEINLEIN, ROBERT A.  H
FOR WANT OF A NAIL  SOBEL, ROBERT  H
FOR YOUR EYES ONLY [JAMES BOND 8]  FLEMING, IAN  C/P:2,P
FOR YOUR SIGHS ONLY  ALLISON, CLYDE  SR-L
FORAGERS  OBERNDORF, CHARLES  C/P:2,P
FORBIDDEN, THE  BRETT, LEO  SR-P
FORBIDDEN [SWORD & SORCERY DEAD GOD 3]  BYERS, RICHARD LEE  P
FORBIDDEN AREA  FRANK, PAT  H:2,P
FORBIDDEN CITY, THE [@LONE WOLF #2]  PAGE, IAN/DEVER, JOE  P
FORBIDDEN GARDEN, THE  TAINE, JOHN  SR-H,H:2
FORBIDDEN GROVE, THE [@COMPLETE ELFQUEST #2]  PINI, WENDY/PINI, RICHARD  CX
FORBIDDEN KNOWLEDGE [GAP 2]  DONALDSON, STEPHEN R.  H:2,P
FORBIDDEN LAND, THE [WITCHES OF EILEANAN 4]  FORSYTH, KATE  C/P,P
FORBIDDEN MAGIC  MCCRAY, CHEYENNE  P
FORBIDDEN MAGIC [@GODWARS #1]  WELLS, ANGUS  P

TITLE  AUTHOR  CODE
FORBIDDEN OBJECTS  DAVIS, MAGGIE  P
FORBIDDEN PLANET  STUART, W. J.  C/P:2,P
FORBIDDEN PLANETS  CROWTHER, PETER  PA
FORBIDDEN PLANETS  KAYE, MARVIN  HA
FORBIDDEN REGION [@TIME WARRIORS #2]  NORTH, DAVID  C/P,P
FORBIDDEN SANCTUARY  BOWKER, RICHARD  C/P:2,P
FORBIDDEN TOWER, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
FORBIDDEN WORLD  WHITE, TED/BISCHOFF, DAVID F.  C/P:2,P
FORBIDDEN ZONE, THE  STRIEBER, WHITLEY  H
FORCE 97X  TORRO, PEL  SR-P
FORCE HERETIC [@STAR WARS NEW JEDI ORDER FORCE HERETIC 1,2,3]  WILLIAMS, SEAN/DIX, SHANE  H
FORCE OF ARMS [@ROBOTECH #5]  MCKINNEY, JACK  P
FORCE OF ARMS [@WARBOTS #8]  STINE, G. HARRY  P
FORCE RED  BASS, MILTON R.  H,P
FORCES FROM THE FEDERATION [@FLASH GORDON #4]  ANONYMOUS  P
FORCES OF KRILL, THE [ZORK 1]  MERETZKY, S. ERIC  P
FORECAST 1955  GERNSBACK, HUGO  B
FORECAST 1956  GERNSBACK, HUGO  B
FORECAST 1957  GERNSBACK, HUGO  B
FORECAST 1958  GERNSBACK, HUGO  B
FORECAST 1959  GERNSBACK, HUGO  B
FORECAST 1960  GERNSBACK, HUGO  B
FORECAST 1961  GERNSBACK, HUGO  B
FORECAST 1963  GERNSBACK, HUGO  B
FORECAST 1964  GERNSBACK, HUGO  B
FOREIGN BODIES  DEDMAN, STEPHEN  H
FOREIGN DEVILS [DOCTOR WHO]  CARTMEL, ANDREW  SR-H,H
FOREIGN FOES [STAR TREK TNG]  GALANTER, DAVE/BRODEUR, GREG  C/P:2,P
FOREIGN LEGIONS [RANKS OF BRONZE]  DRAKE, DAVID  HA,PA
FOREIGNER [ATEVI 1]  CHERRYH, C. J.  H:2,P
FOREIGNER [DINOSAUR EMPIRE 3]  SAWYER, ROBERT J.  P
FOREKNOWLEDGE [QUANTUM LEAP]  DEFILIPPIS  P
FORERUNNER [FORERUNNER 4/DIPPLE 4]  NORTON, ANDRE  C/P:2,SFWA-P,P
FORERUNNER FORAY [FORERUNNER 3]  NORTON, ANDRE  H:2,P
FORERUNNER: THE SECOND VENTURE [FORERUNNER 5]  NORTON, ANDRE  H:2,SFWA,SFWA-P,P
FOREST HOUSE, THE [MISTS OF AVALON 2]  BRADLEY, MARION ZIMMER  H,LP
FOREST LORD, A  WILLIAMS, MICHAEL  P
FOREST MAGE [SOLDIER SON 2]  HOBB, ROBIN  H
FOREST OF FOREVER, THE  SWANN, THOMAS BURNETT  C/P:2,P
FOREST OF PELDAN, THE  BAYLEY, BARRINGTON J.  P
FOREST OF STARS, A [@SEVEN SUNS #2]  ANDERSON, KEVIN J.  H,P
FOREST OF STARS, A (PROOF) [@SEVEN SUNS #2]  ANDERSON, KEVIN J.  SR-LP
FOREST OF THE NIGHT [RIVERRUN 2]  SOMTOW, S. P.  P
FOREST OF THE NIGHT  STEUSSY, MARTI  C/P:2,P
FOREST OF TIME AND OTHER STORIES, THE  FLYNN, MICHAEL  LP
FORESTS OF GLEOR, THE [@BLADE #22]  LORD, JEFFREY  P
FORESTS OF THE HEART [NEWFORD]  DE LINT, CHARLES  H
FORESTS OF THE NIGHT [MOREAU 1]  SWANN, S. ANDREW  C/P,SFWA-P,P
FOREVER AFTER  BAEN BOOKS  C/PA:2,PA
FOREVER AND THE NIGHT  MILLER, LINDA LAEL  H
FOREVER CITY, THE  LUPOFF, RICHARD A.  H
FOREVER DARK [@STAR TREK TNG MAXIMUM WARP #2]  GALANTER, DAVE/BRODEUR, GREG  P
FOREVER DRUG, THE  PERRY, STEVE  C/P:2,P
FOREVER DRUG, THE [SHADOWRUN 37]  SMEDMAN, LISA  P
FOREVER EMBER  EARLE, RICHARD  SR-L
FOREVER FORMULA, THE  BONHAM, FRANK  C/P,P,H
FOREVER FREE  HALDEMAN, JOE  H,P
FOREVER HERO, THE [FOREVER HERO 1,2,3]  MODESITT, L. E. JR.  H
FOREVER KING, THE [FOREVER KING 1]  COCHRAN, MOLLY/MURPHY, WARREN  H:2,P
FOREVER MACHINE, THE  CLIFTON, MARK/RILEY, FRANK  SR-P,C/P:2
FOREVER MAN, THE  DICKSON, GORDON R.  C/P,SFWA-P,P,H
FOREVER PEACE  HALDEMAN, JOE  C/P,P,H
FOREVER WAR, THE  HALDEMAN, JOE  C/P,P,H
FORGE, THE [@GENERAL #1]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P
FORGE IN THE FOREST, THE [@WINTER OF THE WORLD #2]  ROHAN, MICHAEL SCOTT  P
FORGE OF GOD, THE [FORGE OF GOD 1]  BEAR, GREG  H:2,SFWA-P,P
FORGE OF HEAVEN [GENE WARS 2]  CHERRYH, C. J.  H:2
FORGE OF MARS, THE  BALFOUR, BRUCE  P
FORGE OF THE ELDERS [ERIS 1,2]  SMITH, L. NEIL  P
FORGE OF THE MINDSLAYERS [@EBERRON BLADE OF THE FLAME #2]  WAGGONER, TIM  P
FORGE OF THE TITANS  WHITE, STEVE  H
FORGE OF VIRTUE, THE [ULTIMA]  ABBEY, LYNN  C/P,P
FORGED BONDS  TAN, CECILIA  SR-L
FORGED BY FIRE [@DRAGON TEMPLE #3]  CROSS, JANINE  LP

The MITSFS Titledex  September 24, 2007  Page 134
TITLE  AUTHOR  CODE
FORGED IN BATTLE [@WARHAMMER]  HUNTER, JUSTIN  P
FORGETTING PLACES  SOMTOW, S. P.  H
FORGING OF THE DRAGON, THE [WIZARD AND DRAGON 1]  HUGHES, ROBERT DON  C/P:2,P
FORGING OF THE SHADOWS, THE [@LIGHTBRINGER #1]  JOHNSON, OLIVER  H,LP
FORGING THE DARKSWORD [@DARKSWORD #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
FORGING THE RUNES [PRINCE OF THE SIDHE 2]  SHERMAN, JOSEPHA  P
FORGOTTEN BEASTS OF ELD, THE  MCKILLIP, PATRICIA A.  C/P,P,H
FORGOTTEN DOOR, THE  KEY, ALEXANDER  C/P:2,LP
FORGOTTEN LIFE  ALDISS, BRIAN W.  P
FORGOTTEN PLANET, THE  JENKINS, WILL F.  H
FORGOTTEN PLANET, THE  LEINSTER, MURRAY  H:2,P,DD146
FORGOTTEN RACE, THE  NEWTON, JULIUS P.  H
FORGOTTEN REALM, THE [DOC SAVAGE]  ROBESON, KENNETH  P
FORGOTTEN REALMS ATLAS, THE [FORGOTTEN REALMS]  FONSTAD, KAREN WYNN  A
FORGOTTEN STAR, THE [DIG ALLEN #1]  GREENE, JOSEPH  P,@H
FORGOTTEN TRUTH [FIRST TRUTH 3]  COOK, DAWN  P
FORGOTTEN WAR, THE [@STAR TREK TNG]  FORSTCHEN, WILLIAM R.  C/P,P
FORLORN, THE  FREER, DAVE  C/P,P
FORLORN HOPE, THE  DRAKE, DAVID  C/P:2,SFWA-P:2,P
FORMER KING, THE [DOOM-QUEST OF ARA-KARN 1]  CORBY, ADAM  C/P,P
FORMS OF HEAVEN  BARKER, CLIVE  H
FORMULA 29X  TORRO, PEL  SR-P
FORNAX [MATTHEWS 3]  BURGAUER, STEVEN  H,LP
FORSAKE THE SKY  POWERS, TIM  C/P:2,P
FORSAKEN [SWORD & SORCERY DEAD GOD 1]  BYERS, RICHARD LEE  P
FORSAKEN, THE [@SPACE: 1999]  MUIR, JOHN KENNETH  P
FORSAKEN EARTH, THIS [@SEA BEGGARS 2]  KEARNEY, PAUL  LP
FORSAKEN HOUSE [@FORGOTTEN REALMS LAST MYTHAL #1]  BAKER, RICHARD  P
FORSWORN [SWORD & SORCERY DEAD GOD 2]  BYERS, RICHARD LEE  P
FORT AT RIVER’S BEND, THE [CAMULOD 5]  WHYTE, JACK  LP
FORT PRIVILEGE  REED, KIT  P,SFWA
FORTEC CONSPIRACY, THE  GARVIN, RICHARD M./ADDEO, EDMOND G.  C/P,P
FORTECKNING PA JULES VERNE MAGASINET  HJALMARSSON, KJELL  R/IND
FORTERESSE DES SIX LUNES, LA (FRENCH) [@PERRY RHODAN #7]  SCHEER, K. H./DARLTON, CLARK  FPR
FORTRESS  DRAKE, DAVID  C/P:2,SFWA,SFWA-P,P,H
FORTRESS, THE [SPIDER WORLD 3]  WILSON, COLIN  P
FORTRESS AND THE FIRE, THE  FRIEDMAN, MICHAEL JAN  C/P:2,P
FORTRESS ATLANTIS [@PERRY RHODAN #52]  ACE  C/PA:2,PA
FORTRESS DRACONIS [DRAGONCROWN WAR 1]  STACKPOLE, MICHAEL A.  LP
FORTRESS IN THE EYE OF TIME [TRISTEN 1]  CHERRYH, C. J.  H:2,P
FORTRESS IN THE EYE OF TIME (PROOF) [TRISTEN 1]  CHERRYH, C. J.  SR-LP
FORTRESS IN TIME [PERRY RHODAN 123]  MAHR, KURT  SR-LP
FORTRESS OF DRAGONS [TRISTEN 4]  CHERRYH, C. J.  C/P,P,H
FORTRESS OF EAGLES [TRISTEN 2]  CHERRYH, C. J.  H:2,P
FORTRESS OF ETERNITY, THE  WHITMORE, ANDREW  C/P,P
FORTRESS OF FROST AND FIRE [BARD’S TALE 2]  LACKEY, MERCEDES/EMERSON, RU  C/P:2,SFWA-P,P
FORTRESS OF GLASS, THE [CROWN OF THE ISLES 1]  DRAKE, DAVID  H
FORTRESS OF GREY ICE, A [SWORD OF SHADOWS 2]  JONES, J. V.  H
FORTRESS OF ICE [TRISTEN 5]  CHERRYH, C. J.  H
FORTRESS OF LIES [@MECHWARRIOR DARK AGE/BATTLETECH]  YORK, J. STEVEN  P
FORTRESS OF OWLS [TRISTEN 3]  CHERRYH, C. J.  C/P,P,H
FORTRESS OF SOLITUDE [DOC SAVAGE #23]  ROBESON, KENNETH  @P
FORTRESS OF THE PEARL, THE [ELRIC]  MOORCOCK, MICHAEL  H:2,SFWA,SFWA-P:2,P
FORTRESS OF THE SIX MOONS [@PERRY RHODAN #7]  ACE  PA,C/PA
FORTRESS ON THE SUN  COOK, PAUL  C/P,P
FORTRESS REPUBLIC [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
FORTUNATE FALL, THE  CARTER, RAPHAEL  LP
FORTUNE, THE [@BLACKWATER #5]  MCDOWELL, MICHAEL  P
FORTUNE FOR KREGEN, A [DRAY PRESCOT 21]  AKERS, ALAN BURT  P
FORTUNE OF FEAR [@MISSION EARTH #5]  HUBBARD, L. RON  H,P
FORTUNE OF FEAR (PROOF) [@MISSION EARTH #5]  HUBBARD, L. RON  SR-LP

FORTUNE TELLER, THE SCHIMEL, LAWRENCE/

GREENBERG, MARTIN H.

PA

FORTUNE’S FOOL [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  H
FORTUNE’S LIGHT [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
FORTUNES OF BRAK, THE [BRAK 5]  JAKES, JOHN  P
FORTUNE’S STROKE [BELISARIUS 4]  FLINT, ERIC/DRAKE, DAVID  H:2
FORTUNE’S WHEEL [SHOLAN ALLIANCE 2]  NORMAN, LISANNE  P
’48  HERBERT, JAMES  H
40-MINUTE WAR, THE  MORRIS, JANET/MORRIS, CHRIS  H,P
FORTY SIGNS OF RAIN [FORTY SIGNS OF RAIN 1]  ROBINSON, KIM STANLEY  P
FORTY THOUSAND IN GEHENNA [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
43,000 YEARS LATER  COON, HORACE  C/P:2,P
FORWARD!  DICKSON, GORDON R.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 135
TITLE  AUTHOR  CODE
FORWARD IN TIME  BOVA, BEN  C/P,P,H,SFWA
FORWARD THE FOUNDATION [FOUNDATION]  ASIMOV, ISAAC  C/P,P,H:2
FORWARD THE MAGE [JOE’S WORLD 2]  FLINT, ERIC/ROACH, RICHARD  C/P,P,H
FOSSIL [ISAAC’S UNIVERSE]  CLEMENT, HAL  C/P:2,P
FOSSIL HUNTER [DINOSAUR EMPIRE 2]  SAWYER, ROBERT J.  P,SFWA-P
FOUCAULT’S PENDULUM  ECO, UMBERTO  C/P,P,H
FOUND WANTING  CARTER, LIN  C/P:2,P
FOUNDATION [FOUNDATION 1]  ASIMOV, ISAAC  C/P:2,P
FOUNDATION (REVISED) [FOUNDATION 1]  ASIMOV, ISAAC  P
FOUNDATION AND CHAOS [SECOND FOUNDATION 2]  BEAR, GREG  C/P,P,H
FOUNDATION AND EARTH [FOUNDATION 5]  ASIMOV, ISAAC  H:3,SFWA,P
FOUNDATION AND EARTH PAPER BLOCK  CHUNK, GEORGE  SFWA-VLP
FOUNDATION AND EMPIRE [FOUNDATION 2]  ASIMOV, ISAAC  C/P:2,P,SR-H,H
FOUNDATION TRILOGY, THE [FOUNDATION 1,2,3]  ASIMOV, ISAAC  H:2,LP
FOUNDATIONS [@STAR TREK S.C.E. 5]  WARD, DAYTON/DILMORE, KEVIN  P
FOUNDATION’S EDGE [FOUNDATION 4]  ASIMOV, ISAAC  C/P,P,H:2,SFWA
FOUNDATION’S FEAR [SECOND FOUNDATION 1]  BENFORD, GREGORY  H:2,P
FOUNDATION’S FRIENDS [FOUNDATION]  GREENBERG, MARTIN H.  C/PA,SFWA-P,LPA
FOUNDATION’S TRIUMPH [SECOND FOUNDATION 3]  BRIN, DAVID  C/P,P,H
FOUNDER, THE  ROWLEY, CHRISTOPHER  C/P,P
FOUNDING, THE [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  LP
FOUNDLING, THE [PRYDAIN 6]  ALEXANDER, LLOYD  H,P
FOUNTAINHEAD, THE  RAND, AYN  P
FOUNTAINS OF MIRLACCA, THE [DEMON WARS 1]  MCCONNELL, ASHLEY  P
FOUNTAINS OF PARADISE, THE  CLARKE, ARTHUR C.  H:2,P
FOUNTAINS OF YOUTH, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H,LP
FOUR AND TWENTY BLACKBIRDS [BARDIC VOICES 4]  LACKEY, MERCEDES  H:2
FOUR AND TWENTY BLACKBIRDS [EDEN MOORE]  PRIEST, CHERIE  LP
FOUR CAME BACK  CAIDIN, MARTIN  C/P:2,SFWA-P,P,H
4-D FUNHOUSE, THE  EMERY, CLAYTON/WAJENBERG, EARL  P
FOUR-DAY PLANET  PIPER, H. BEAM  H
FOUR-DAY PLANET/LONE STAR PLANET  PIPER, H. BEAM  C/P:2,P
FOUR DAY WEEKEND, THE  SMITH, GEORGE H.  C/P:2,P
FOUR-DIMENSIONAL NIGHTMARE, THE  BALLARD, J. G.  SR-P
4 FOR THE FUTURE  CONKLIN, GROFF  PA,C/PA
FOUR FOR TOMORROW  ZELAZNY, ROGER  C/P:2,SFWA-P,P
FOUR FORGES, THE [ELVEN WAYS 1]  RHODES, JENNA  H
FOUR FROM PLANET 5  LEINSTER, MURRAY  C/P,P,H
FOUR FROM THE WITCH WORLD  NORTON, ANDRE  C/PA,PA,SFWA-P
FOUR FUTURES  HAWTHORN BOOKS  HA
4 FUTURES  MANOR BOOKS  C/PA,PA
FOUR HUNDRED BILLION STARS  MCAULEY, PAUL J.  C/P,P

FOUR LORDS OF THE DIAMOND, THE [FOUR LORDS OF THE DIAMOND 1,2,3,4] CHALKER, JACK L.

H:2

FOUR MAGICS, THE  FORSTCHEN, WILLIAM R./SEGRIFF, LARRY  P
FOUR MOONS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
FOUR PAST MIDNIGHT  KING, STEPHEN  H:2,P
FOUR-SIDED TRIANGLE  TEMPLE, WILLIAM F.  GN2.09,SR-H
FOUR TO DOOMSDAY [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
FOUR WAYS TO FORGIVENESS  LE GUIN, URSULA K.  C/P,P,H
FOUR WAYS TO FORGIVENESS (PROOF)  LE GUIN, URSULA K.  SR-LP
FOURS CROSSING  GARDEN, NANCY  P
FOURSIGHT  CROWTHER, PETER  HA
14 GREAT TALES OF ESP  STONE, IDELLA PURNELL  PA
FOURTH BOOK OF JORKENS, THE [JORKENS]  DUNSANY, LORD  H
FOURTH DIMENSION, THE  RUCKER, RUDY  [C/RAN-H]

FOURTH DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-P

FOURTH EMPIRE, THE [@STARHAWK 3]  MALONEY, MACK  P
FOURTH GHOST BOOK, THE  TURNER, JAMES  SR-PA
FOURTH GUARDIAN, THE [@ETERNAL GUARDIANS #1]  CROSS, RONALD ANTHONY  LP
FOURTH HEMISPHERE, THE  LAKE, DAVID J.  SR-P
FOURTH MANSIONS  LAFFERTY, R. A.  C/P:2,P
FOURTH OMNI BOOK OF SCIENCE FICTION, THE  DATLOW, ELLEN  PA
FOURTH PLANET FROM THE SUN  VAN GELDER, GORDON  HA,LPA
FOURTH ’R’, THE  SMITH, GEORGE O.  C/P:2,P
FOURTH ROME, THE [ARC RIDERS]  DRAKE, DAVID/MORRIS, JANET  C/P,P
FOURTH WORLD, THE  DANVERS, DENNIS  H
FOX AND EMPIRE [GERIN THE FOX 3]  TURTLEDOVE, HARRY  P
FOX RUN, THE [@ENDWORLD 1]  ROBBINS, DAVID  P

FOX VALLEY MURDERS, THE [@VANCE INTEGRAL EDITION #13]/ PLEASANT GROVE MURDERS, THE [@VANCE INTEGRAL EDITION #13] VANCE, JACK

SR-H

FOX WOMAN, THE  JOHNSON, KIJ  H,LP
FOX WOMAN AND OTHER STORIES, THE  MERRITT, A.  P

TITLE  AUTHOR  CODE
FOXES OF FIRSTDARK, THE  KILWORTH, GARRY  H
FOXGLOVE HOLLOW  BROWN, DENNIS PHILLIP  P
FOXMASK  MARILLIER, JULIET  LP
FRACAS FACTOR, THE  REYNOLDS, MACK  C/P,P
FRACTAL MODE [MODE 2]  ANTHONY, PIERS  H:2,SFWA-P:2,P
FRACTAL PAISLEYS  DI FILIPPO, PAUL  H
FRAGILE THINGS  GAIMAN, NEIL  H
FRAME OF REFERENCE  OLTION, JERRY  C/P,P
FRAMESHIFT  SAWYER, ROBERT J.  H,P
FRANCIS GOES TO WASHINGTON  STERN, DAVID  H
FRANK FRAZETTA CALENDAR, THE (1977)  CALENDAR  CAL
FRANK HERBERT  O’REILLY, TIMOTHY  R/REF-H
FRANK KELLY FREAS: THE ART OF SCIENCE FICTION  FREAS, FRANK KELLY  A
FRANKENSTEIN  SHELLEY, MARY  C/P,P,H:2
FRANKENSTEIN ’69  MARTIN, ED  SR-L
FRANKENSTEIN FACTORY, THE  HOCH, EDWARD D.  P
FRANKENSTEIN OMNIBUS, THE  HAINING, PETER  HA
FRANKENSTEIN PAPERS, THE  SABERHAGEN, FRED  C/P:2,P
FRANKENSTEIN READER, THE  BECK, CALVIN  SR-PA,C/PA
FRANKENSTEIN: THE MONSTER WAKES  GREENBERG, MARTIN H.  PA
FRANKENSTEIN UNBOUND  ALDISS, BRIAN W.  C/P,SFWA,P,H
FRANKENSTEIN WHEEL, THE  FAIRMAN, PAUL W.  P
FRANKENSTEINS AND FOREIGN DEVILS  WILLIAMS, WALTER JON  SR-H
FRATERNITY OF THE WEIRD  HERVEY, MICHAEL  [C/RAN-P]
FRAYED [DOCTOR WHO]  SAMMS, TARA  SR-H
FREAK SHOW, THE  HAINING, PETER  SR-PA
FREAKS AGAINST SUPERMEN  FINN, RALPH L.  SR-P
FREAKS’ AMOUR  DE HAVEN, TOM  LP
FRECKLED SHARK, THE [DOC SAVAGE #67]  ROBESON, KENNETH  C/P:2,@P
FREDDY AND THE FLYING SAUCER PLANS  BROOKS, WALTER R.  H
FREDDY’S BOOK  GARDNER, JOHN (AMERICAN, 1933-1982)  P
FREE AMAZONS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
FREE FLIGHT  TERMAN, DOUGLAS  P
FREE LANCERS [ALIEN STARS 4]  MITCHELL, ELIZABETH  C/PA,PA
FREE LIVE FREE  WOLFE, GENE  C/P,SFWA-P,P,H
FREE LUNCH, THE  ROBINSON, SPIDER  H,P
FREE RADICALS [@BLACK HOLE TRAVEL AGENCY #3]  MCKINNEY, JACK  C/P:2,P
FREE SPACE  LINAWEAVER, BRAD/KRAMER, EDWARD E.  LPA
FREE ZONE  PLATT, CHARLES  C/P,P
FREEBIRTH [@BATTLETECH #36/TWILIGHT OF THE CLANS 4]  THURSTON, ROBERT  C/P,P
FREEDOM!  TIER, MARK/GREENBERG, MARTIN H.  LPA
FREEDOM & NECESSITY  BRUST, STEVEN/BULL, EMMA  C/P:2,P
FREEDOM BEACH  KELLY, JAMES PATRICK/KESSEL, JOHN  SR-H,SFWA-P,LP
FREEDOM EXPRESS [WINGMAN 7]  MALONEY, MACK  P
FREEDOM FLIGHT [WING COMMANDER 1]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
FREEDOM LOST [@DEATHLANDS]  AXLER, JAMES  P
FREEDOM’S APPRENTICE [DEAD RIVERS 2]  KRITZER, NAOMI  P
FREEDOM’S CHALLENGE [BOTANY 3]  MCCAFFREY, ANNE  H:2,P
FREEDOM’S CHOICE [BOTANY 2]  MCCAFFREY, ANNE  C/P:2,P,H
FREEDOM’S GATE [DEAD RIVERS 1]  KRITZER, NAOMI  C/P,P
FREEDOM’S LANDING [BOTANY 1]  MCCAFFREY, ANNE  H:3,P
FREEDOM’S RANGERS [@FREEDOM’S RANGERS 1]  ANDREWS, KEITH WILLIAM  C/P,P
FREEDOM’S RANSOM [BOTANY 4]  MCCAFFREY, ANNE  H:2
F.R.E.E.FALL [F.R.E.E.LANCERS 2]  ODOM, MEL  P
FREEHOLD  DIETZ, WILLIAM C.  P
FREEHOLD [FREEHOLD WAR 1]  WILLIAMSON, MICHAEL Z.  P
F.R.E.E.LANCERS [F.R.E.E.LANCERS 1]  ODOM, MEL  P
FREEMAN, THE  AHERN, JERRY/AHERN, SHARON  P
FREEMASTER  JENSEN, KRIS  C/P:2,P
FREEWARE [WARE 3]  RUCKER, RUDY  LP
FREEZE FRAMES  KERR, KATHARINE  P
FREEZING DOWN  BODELSEN, ANDERS  H:2,P
FREK AND THE ELIXER  RUCKER, RUDY  LP
FRENCHMAN, THE [@MILLENIUM]  HAND, ELIZABETH  C/P,P
FRENZIED FICTION  LEACOCK, STEPHEN  H
FRENZY [@MUTANT CHRONICLES APOSTLE OF INSANITY 2]  PRICE, JOHN-ALLEN  P
FRESCO, THE  TEPPER, SHERI S.  H
FRIDAY  HEINLEIN, ROBERT A.  C/P,H,LP
FRIEND TO MANKIND [@PERRY RHODAN #91]  ACE  PA
FRIENDS COME IN BOXES  CONEY, MICHAEL G.  C/P,P,H
FRIGHT  COLLINS, CHARLES M.  PA
FRIGHT NIGHT  SKIPP, JOHN/SPECTOR, CRAIG  P
FRIGHTENED FISH, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
FRIGHTENERS, THE  JAHN, MICHAEL  P

TITLE  AUTHOR  CODE
FRIGHTS  MCCAULEY, KIRBY  PA
FRITZ LEIBER  FRANE, JEFF  R/REF-P
FRITZ LEIBER: A BIBLIOGRAPHY 1934–1979  MORGAN, CHRIS  R/IND
FROGS OF WAR, THE  HARMAN, ANDREW  P
FROM A BUICK 8  KING, STEPHEN  H
FROM A CHANGELING STAR [CHANGELING STAR 1]  CARVER, JEFFREY A.  H:2,P
FROM ALIEN TO THE MATRIX  KAVENY, ROZ  R/REF-H
FROM DREAM TO DREAM  GOLDS, ARLENE  H
FROM EARTH TO HEAVEN  ASIMOV, ISAAC  [C/RAN-P:2]
FROM ELFLAND TO POUGHKEEPSIE  LE GUIN, URSULA K.  R/REF-P
FROM MIND TO MIND  SCHMIDT, STANLEY  SFWA
FROM OUTER SPACE  CLEMENT, HAL  C/P:2,P
FROM OUTER SPACE  MENGER, HOWARD  [C/RAN-P]
FROM PRUSSIA WITH LOVE [CASTLE FALKENSTEIN]  DE CHANCIE, JOHN  P
FROM REALMS BEYOND  BRETT, LEO  SR-P
FROM RUSSIA, WITH LOVE [JAMES BOND 5]  FLEMING, IAN  C/P,P
FROM SATAN, WITH LOVE [LUCIFER COVE 6]  COFFMAN, VIRGINIA  P
FROM THE ARCHIVES OF EVIL #2  LEE, CHRISTOPHER/PARRY, MICHEL  PA

FROM THE BLOB TO STAR WARS:

THE SCIENCE FICTION MOVIE QUIZ BOOK ANDREWS, BART

R/REF-P,C/REF-P

FROM THE BORDERLANDS MONTELEONE, ELIZABETH/ MONTELEONE, THOMAS F.

PA

FROM THE DEPTHS [STAR TREK]  MILAN, VICTOR  C/P:2,P
FROM THE DUST RETURNED  BRADBURY, RAY  H:2
FROM THE EARTH TO THE MOON  VERNE, JULES  C/P,P
FROM THE EARTH TO THE MOON/TOUR OF THE MOON, A  VERNE, JULES  SR-H
FROM THE EARTH TO THE MOON/TRIP AROUND IT, A  VERNE, JULES  C/P,P,H
FROM THE END OF THE TWENTIETH CENTURY  FORD, JOHN M.  SR-H,H
FROM THE FILES OF THE TIME RANGERS  BOWES, RICHARD  H
FROM THE HEART OF DARKNESS  DRAKE, DAVID  P
FROM THE HEART OF THE STORM [FALLEN CLOUD 4]  GIAMBASTIANI, KURT R. A.  P
FROM THE JAWS OF VICTORY  FAIR, CHARLES  [C/RAN-H]
FROM THE LAND OF FEAR  ELLISON, HARLAN  C/P:2,P
FROM THE LEGEND OF BIEL  STATON, MARY  C/P:2,P
FROM THE OCEAN, FROM THE STARS  CLARKE, ARTHUR C.  SR-H,H
FROM THE "S" FILE  PLAYBOY BOOKS  PA,C/PA
FROM THE SEA TO THE STARS  NORTON, ANDRE  LP
FROM THE TEETH OF ANGELS  CARROLL, JONATHAN  P
FROM THE TWILIGHT ZONE  SERLING, ROD  H
FROM THESE ASHES  BROWN, FREDRIC  H
FROM THIS DAY FORWARD  BRUNNER, JOHN  C/P:2,SFWA,P,H:2
FROM TIME TO TIME [TIME AND AGAIN 2]  FINNEY, JACK  H,LP
FROM TOLKIEN TO OZ  MCGUIRE, WILLIAM  A
FROM UNKNOWN WORLDS  CAMPBELL, JOHN W. JR.  HA
FROM WHAT FAR STAR?  BERRY, BRYAN  H
FRONTERA  SHINER, LEWIS  P
FRONTIER  ANSARA, MICHAEL  H
FRONTIER EARTH [@FRONTIER EARTH 1]  BOXLEITNER, BRUCE  H
FRONTIER OF THE DARK  CHANDLER, A. BERTRAM  C/P,P
FRONTIER WORLDS [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER  P
FRONTIER WORLDS  COLLINS, PAUL  SR-PA
FRONTIERS IN SPACE  BLEILER, EVERETT F./DIKTY, T. E.  PA,C/PA
FRONTIERS OF THE UNKNOWN  MACKENZIE, ANDREW  [C/RAN-P]
FRONTIERS PAST AND FUTURE  ABBOTT, CARL  R/REF-H
FRONTIOS [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,P
FROOMB!  LYMINGTON, JOHN  H:2,P
FROST [FROST 1]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
FROST [DEMON BLADE 2]  GARLAND, MARK A./MCGRAW, CHARLES G. P
FROST AND FIRE  ZELAZNY, ROGER  C/P:2,P,H
FROST-HAIRED VIXEN, THE [ZACHARY NIXON JOHNSON 4]  ZAKOUR, JOHN  P
FROST ON THE WINDOW  WISMAN, KEN  LP
FROSTED DEATH, THE [@AVENGER #5]  ROBESON, KENNETH  C/P:2,P
FROSTFELL [@FORGOTTEN REALMS WIZARDS]  SEHESTEDT, MARK  P
FROSTFLOWER AND THORN [FROSTFLOWER]  KARR, PHYLLIS ANN  C/P:2,P
FROSTFLOWER AND WINDBOURNE [FROSTFLOWER]  KARR, PHYLLIS ANN  C/P:2,P
FROSTWING  KNAAK, RICHARD A.  C/P,P
FROSTWORLD AND DREAMFIRE  MORRESSY, JOHN  C/P,SFWA,P,H:2
FROZEN FIRE [@SURVIVAL 2000 3]  MCPHEE, JAMES  P
FROZEN LIMIT  GRIDBAN, VOLSTED  SR-P
FROZEN PIRATE, THE  RUSSELL, W. CLARK  H
FROZEN PLANET, THE  MACFADDEN BOOKS  PA,C/PA
FROZEN PLANET, THE  MERAK, A. J.  P
FROZEN PLANET  TORRO, PEL  H
FROZEN WAVES, THE [@JADE DEMONS #2]  VARDEMAN, ROBERT E.  P

The MITSFS Titledex  September 24, 2007  Page 138
TITLE  AUTHOR  CODE
FROZEN YEAR, THE  BLISH, JAMES  C/P,P,H
F.T.L. [@EXPLORATION #3]  RANDLE, KEVIN  P
FTL: FURTHER THAN LIFE  WILLIAMS, MICHAEL LINDSAY  C/P,P
FUDOKI  JOHNSON, KIJ  LP
FUGITIVE IN TRANSIT  LLEWELLYN, EDWARD  C/P,P
FUGITIVE OF THE STARS  HAMILTON, EDMOND  DM111

FUGITIVE PRINCE [WARS OF LIGHT AND SHADOW 3/ ALLIANCE OF LIGHT 1] WURTS, JANNY

H,P

FUGITIVE STARS, THE  RANSOM, DANIEL  P
FUGITIVE WORLDS, THE [TOLLER MARAQUINE 3]  SHAW, BOB  P
FUGITIVES OF CHAOS [ORPHANS OF CHAOS 2]  WRIGHT, JOHN C.  H
FUGUE STATE  FORD, JOHN M.  SFWA-TD25,TD25
FULFILMENTS OF FATE AND DESIRE, THE [WRAETHTHU 3]  CONSTANTINE, STORM  P,SFWA-P
FULL CIRCLE  ARISS, BRUCE  H
FULL CIRCLE [DOCTOR WHO]  SMITH, ANDREW  SR-P,C/P
FULL MOON/HIGH PRIEST OF CALIFORNIA  MUNDY, TALBOT/WILLEFORD, CHARLES  SR-P
FULL MOON RISING [RILEY JENSON 1]  ARTHUR, KERI  P
FULL MOONSTER [BUREAU 13 3]  POLLOTTA, NICK  P
FULL SPECTRUM  ARONICA, LOU/MCCARTHY, SHAWNA  PA,C/PA

FULL SPECTRUM 2 ARONICA, LOU/MCCARTHY, SHAWNA/ STOUT, AMY/LOBRUTTO, PAT

C/PA,PA

FULL SPECTRUM 3 ARONICA, LOU/STOUT, AMY/

MITCHELL, BETSY

C/PA,PA,SR-LPA

FULL SPECTRUM 4 ARONICA, LOU/STOUT, AMY/

MITCHELL, BETSY

C/PA,LPA

FULL SPECTRUM 4 (PROOF)

ARONICA, LOU/STOUT, AMY/ MITCHELL, BETSY

SR-LPA

FULL SPECTRUM 5 HERSHEY, JENNIFER/DUPRE, TOM/

SILVERSTEIN, JANNA

LPA

FULL TIDE OF NIGHT  DUNN, J. R.  H,P
FULLY AUTOMATED LOVE LIFE OF HENRY KEANRIDGE, THE  PLAYBOY BOOKS  C/PA,PA
FULLY DRESSED AND IN HIS RIGHT MIND  FESSIER, MICHAEL  SR-P,C/P
FUN WITH YOUR NEW HEAD  DISCH, THOMAS M.  H:2,SFWA,P
FUNCO FILE, THE  COLE, BURT  H:2,P
FUNDACION E IMPERIO  ASIMOV, ISAAC  FP
FUNDAMENTAL DISCH  DISCH, THOMAS M.  C/P:2,P
FUNERAL CUSTOMS IN TOLKIEN’S TRILOGY  ROCKOW, KAREN  R/REF-P
FUNERAL FOR THE EYES OF FIRE, A  BISHOP, MICHAEL  C/P:2,P
FUNF ZAUBERSTEIN ZU BINDEN FUNF VER-SCHIEDNE WELTEN  CHALKER, JACK L.  FP
FUNGI FROM YUGGOTH AND OTHER POEMS  LOVECRAFT, H. P.  C/P,P
FUNGUS, THE  KNIGHT, HARRY ADAM  H
FUNHOUSE, THE  APPEL, BENJAMIN  C/P,P
FUR COUNTRY OR 70 NORTH LATITUDE  VERNE, JULES  SR-H
FUR MAGIC  NORTON, ANDRE  C/P:2,P
FURIES, THE [HOLDFAST 3]  CHARNAS, SUZY MCKEE  C/P,P,H
FURIES, THE  ROBERTS, KEITH  C/P:2,P
FURIES OF CALDERON [CODEX ALERA 1]  BUTCHER, JIM  H
FURIOSO  LESTIENNE, VOLDEMAR  P
FURIOUS GULF [GALACTIC CENTER 5]  BENFORD, GREGORY  C/P,P,H
FURLOUGH FROM HEAVEN  DREIFUSS, JEROME  H
FURRY FANTASTIC  RABE, JEAN/THOMSEN, BRIAN M.  PA
FURTHER ADVENTURES OF BATMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  C/PA,PA
FURTHER ADVENTURES OF BATMAN 2, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF BATMAN 3, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  DICKS, TERRANCE  H
FURTHER ADVENTURES OF HALLEY’S COMET, THE  BATCHELOR, JOHN CALVIN  P
FURTHER ADVENTURES OF SUPERMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF THE JOKER, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF WONDER WOMAN, THE [DC COMICS]  GREENBERG, MARTIN H.  PA
FURTHER ADVENTURES OF XENA WARRIOR PRINCESS [XENA]  GREENBERG, MARTIN H.  PA
FURTHEST [COYOTE JONES 2]  ELGIN, SUZETTE HADEN  C/P:2,P
FURTHEST HORIZON, THE  DOZOIS, GARDNER  LPA
FURY  COYNE, JOHN  SFWA
FURY, THE  FARRIS, JOHN  C/P:2,P
FURY  KUTTNER, HENRY  C/P,P,H
FURY FROM EARTH, THE  MCLAUGHLIN, DEAN  C/P:2,P
FURY FROM THE DEEP [DOCTOR WHO]  PEMBERTON, VICTOR  SR-P
FURY OUT OF TIME, THE  BIGGLE, LLOYD JR.  C/P:2,P
FURY SCORNED, A [STAR TREK TNG]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P:2,P
FUSE POINT [TIME WARRIORS #1]  NORTH, DAVID  C/P,P
FUSION [STAR TREK VOYAGER STRING THEORY 2]  BEYER, KIRSTEN  P
FUSION FIRE [FIREBIRD 2]  TYERS, KATHY  C/P,P
FUTURE AS NIGHTMARE, THE  HILLEGAS, MARK R.  R/REF-H
FUTURE BOSTON  SMITH, DAVID ALEXANDER  HA

The MITSFS Titledex  September 24, 2007  Page 139
TITLE  AUTHOR  CODE
FUTURE CITY  ELWOOD, ROGER  C/PA,PA
FUTURE CORRUPTION  ELWOOD, ROGER  PA
FUTURE CRIME  BOVA, BEN  C/P:2,P
FUTURE CRIME  MANSON, CYNTHIA/ARDAI, CHARLES  HA
FUTURE CRIMES  DANN, JACK/DOZOIS, GARDNER  PA
FUTURE CRIMES  GREENBERG, MARTIN H./HELFERS, JOHN  C/PA,PA
FUTURE FEMALES: A CRITICAL ANTHOLOGY  BARR, MARLENE S.  R/REF-P
FUTURE FOR SALE  SAXON, RICHARD  SR-P
FUTURE FORWARD [REWIND TO YESTERDAY 2]  PFEFFER, SUSAN BETH  H
FUTURE GLITTER  VAN VOGT, A. E.  C/P:2,P

FUTURE I, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

C/PA,PA

FUTURE IMPERFECT [@STAR TREK JANUS GATE #2]  GRAF, L. A.  P
FUTURE IMPERFECT  GUNN, JAMES E.  C/P:2,P

FUTURE IN QUESTION, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

PA

FUTURE INDEFINITE [@GREAT GAME #3]  DUNCAN, DAVE  P
FUTURE IS NOW, THE  NOLAN, WILLIAM F.  HA,PA
FUTURE IS QUEER, THE  LABONTE, RICHARD/SCHIMEL, LAWRENCE  LPA
FUTURE MAKERS, THE  HAINING, PETER  HA,PA
FUTURE NET  GREENBERG, MARTIN H./SEGRIFF, LARRY  PA
FUTURE NOW, THE  HOSKINS, ROBERT  C/PA,PA
FUTURE ON FIRE  CARD, ORSON SCOTT  C/PA,PA,SFWA-P
FUTURE ON ICE  CARD, ORSON SCOTT  LPA
FUTURE PERFECT  FRANKLIN, H. BRUCE  LPA
FUTURE POWER  DANN, JACK/DOZOIS, GARDNER  SR-HA,SFWA,HA
FUTURE PRIMITIVE  ROBINSON, KIM STANLEY  LPA
FUTURE QUARTET  AVONOVA BOOKS  HA,PA
FUTURE QUEST  ELWOOD, ROGER  PA,C/PA
FUTURE SANCTUARY  HARDING, LEE  SR-P
FUTURE SHOCK  TOFFLER, ALVIN  [C/RAN-P]
FUTURE SIGHT [@MAGIC: THE GATHERING TIME SPIRAL #3]  MCGOUGH, SCOTT/DELANEY, JOHN  P
FUTURE SPORTS  DANN, JACK/DOZOIS, GARDNER  PA
FUTURE TENSE  CROSSEN, KENDELL FOSTER  SR-HA
FUTURE TENSE  CURTIS, RICHARD  PA,C/PA
FUTURE TENSE  VANCE, JACK  C/P:2,P
FUTURE TENSE: THE CINEMA OF SCIENCE FICTION  BROSNAN, JOHN  R/REF-H
FUTURE TIMES THREE  BARJAVEL, RENE  C/P:2,P
FUTURE TOOK US, THE  SEVERN, DAVID  SR-P,C/P:2
FUTURE WAR  DANN, JACK/DOZOIS, GARDNER  PA
FUTURE WAR  WODHAMS, JACK  SR-P
FUTURE WARS  GREENBERG, MARTIN H./SEGRIFF, LARRY  PA
FUTURE WASHINGTON  LILLEY, ERNEST  LPA
FUTURE WEAPONS OF WAR  HALDEMAN, JOE/GREENBERG, MARTIN H.  HA
FUTURE WITHOUT FUTURE  STERNBERG, JACQUES  H,SFWA
FUTURELAND  MOSLEY, WALTER  H,P
FUTURELOVE  ELWOOD, ROGER  SR-HA,HA
FUTURES  CROWTHER, PETER  PA
FUTURES IMPERFECT  WILLIS, CONNIE  H
FUTURES PAST  DANN, JACK/DOZOIS, GARDNER  PA
FUTURES PAST  VAN VOGT, A. E.  LP
FUTURES PAST  WHITE, JAMES  C/P:2,P
FUTURES TO INFINITY  MOSKOWITZ, SAM  PA,C/PA
FUTURES UNLIMITED  NORTON, ALDEN H.  C/PA,PA
FUTURESHOCKS  ANDERS, LOU  LPA
FUTUREWORLD  HALL, JOHN RYDER  P
FUTURIANS, THE  KNIGHT, DAMON  R/REF-H
FUTURO AL ROGO (ITALIAN)  GUNN, JAMES E.  FP
FUTUROLOGICAL CONGRESS, THE  LEM, STANISLAW  C/P,P,H
FUZZIES AND OTHER PEOPLE [FUZZIES 3]  PIPER, H. BEAM  H,P
FUZZY BONES [FUZZIES]  TUNING, WILLIAM  C/P:2,P
FUZZY PAPERS, THE [FUZZIES 1,2]  PIPER, H. BEAM  H:2,P
FUZZY SAPIENS [FUZZIES 2]  PIPER, H. BEAM  P

TITLE  AUTHOR  CODE
G-BOMB, THE  STATTEN, VARGO  SR-P
GAAL THE CONQUEROR  WHITE, JOHN  LP
GABRIEL  TUTTLE, LISA  SFWA-P
GABRIEL’S GHOST  SINCLAIR, LINNEA  P
GADGET MAN  GOULART, RON  C/P,SFWA,P,H
GADGET STORIES [@VANCE INTEGRAL EDITION #3]  VANCE, JACK  SR-H
GAHAN WILSON’S THE ULTIMATE HAUNTED HOUSE  COLLINS, NANCY A.  LPA
GAIA WAR, THE  LEON, MARK  C/P,P
GAIA WEBSTERS, THE  ANTIEAU, KIM  LP
GAIAN EXPEDIENT [ERTHRING 2]  DREW, WAYLAND  C/P:2,P
GAIA’S TOYS  ORE, REBECCA  H:2,SFWA,LP
GALACTIAD, THE  KERN, GREGORY  P
GALACTIC ALARM [@PERRY RHODAN #3]  ACE  PA,C/PA
GALACTIC ALIENS  FRANK, ALAN  A
GALACTIC ARENA, THE  DANE, CHRISTOPHER  P
GALACTIC BOUNTY [SAM MCCADE 1]  DIETZ, WILLIAM C.  P
GALACTIC BREED, THE  BRACKETT, LEIGH  DD099
GALACTIC CLUSTER  BLISH, JAMES  C/P:3,P
GALACTIC CONVOY [HELMSMAN 2]  BALDWIN, BILL  C/P:2,P
GALACTIC DERELICT [TIME TRADERS 2]  NORTON, ANDRE  C/P:2,P
GALACTIC DIPLOMAT [RETIEF]  LAUMER, KEITH  C/P,P
GALACTIC DREAMERS  SILVERBERG, ROBERT  SFWA
GALACTIC DREAMS  HARRISON, HARRY  H
GALACTIC DREAMS (PROOF)  HARRISON, HARRY  SR-LP
GALACTIC EFFECTUATOR  VANCE, JACK  C/P,P
GALACTIC EMPIRES 1  ALDISS, BRIAN W.  C/PA,PA
GALACTIC EMPIRES 2  ALDISS, BRIAN W.  PA,C/PA
GALACTIC GAMBIT  DUDLEY, ROY C.  H
GALACTIC GOURMET, THE [SECTOR GENERAL 9]  WHITE, JAMES  H,P
GALACTIC INTRIGUE  BULMER, H. K.  SR-P
GALACTIC INVADERS  BERRY, JAMES R.  SR-P
GALACTIC MEDAL OF HONOR  REYNOLDS, MACK  C/P:2,P
GALACTIC MI [GALACTIC MI]  RANDLE, KEVIN  P
GALACTIC NORTH  REYNOLDS, ALASTAIR  H
GALACTIC ODYSSEY  LAUMER, KEITH  C/P:2,SFWA-P,P
GALACTIC PATROL [LENSMEN 3]  SMITH, E. E.  C/P,P,H
GALACTIC POT-HEALER  DICK, PHILIP K.  H:2,P
GALACTIC RAPTURE  FLYNN, TOM  LP
GALACTIC REJECTS, THE  OFFUTT, ANDREW J.  H,P
GALACTIC RIDDLE, THE [@PERRY RHODAN #8]  ACE  PA,C/PA
GALACTIC RIFT  WARNER, R. D./WARNER, MICHAEL  P
GALACTIC SIBYL SUE BLUE  BROWN, ROSEL GEORGE  C/P:2,P
GALACTIC SILVER STAR, THE [JEFFERSON’S WAR 1]  RANDLE, KEVIN  P
GALACTIC STORM  HUNT, GILL  SR-P
GALACTIC TROUBADOURS, THE  LIGHTNER, A. M.  H
GALACTIC WARLORD [LAST LEGIONARY 1]  HILL, DOUGLAS  H:2,P
GALACTIC WARRIORS [STARHOUNDS 2]  BISCHOFF, DAVID F.  C/P:2,P
GALACTIC WHIRLPOOL, THE [STAR TREK]  GERROLD, DAVID  C/P:3,P
GALACTICA DISCOVERS EARTH [@BATTLESTAR GALACTICA 1978 #5]  LARSON, GLEN A./RESNICK, MIKE  P
GALAPAGOS  VONNEGUT, KURT JR.  H:2,P
GALATEA 2.2  POWERS, RICHARD  LP
GALATEA IN 2-D  ALLSTON, AARON  P,SFWA-P
GALAXIES  MALZBERG, BARRY N.  C/P,P
GALAXIES LIKE GRAINS OF SAND  ALDISS, BRIAN W.  C/P:2,P
GALAXY 5 (TEACHER’S GUIDE)  FEARON-PITMAN  R/REF-P
GALAXY 666  TORRO, PEL  SR-P,C/P
GALAXY BUILDER, THE [LAFAYETTE O’LEARY 4]  LAUMER, KEITH  C/P:2,P
GALAXY FOUR [DOCTOR WHO]  EMMS, WILLIAM  SR-P,C/P
GALAXY IN FLAMES [@WARHAMMER 40K HORUS HERESY 3]  COUNTER, BEN  P
GALAXY JANE  GOULART, RON  C/P:2,P
GALAXY MAGAZINE: THE DARK AND THE LIGHT YEARS  ROSHEIM, DAVID L.  R/REF-H
GALAXY MISSION  HAMILTON, EDMOND  C/P,P
GALAXY OF GHOULS  MERRIL, JUDITH  SR-PA,C/PA
GALAXY OF STRANGERS, A  BIGGLE, LLOYD JR.  H,SFWA
GALAXY OF THE LOST [@CAP KENNEDY #1]  KERN, GREGORY  C/P:2,P
GALAXY PRIMES, THE  SMITH, E. E.  C/P:2,P
GALAXY QUEST  BISSON, TERRY  P
SECOND GALAXY READER OF SCIENCE FICTION, THE  GOLD, HORACE L.  HA
THIRD GALAXY READER, THE  GOLD, HORACE L.  C/PA,PA
FOURTH GALAXY READER, THE  GOLD, HORACE L.  C/PA,PA
FIFTH GALAXY READER, THE  GOLD, HORACE L.  C/PA,PA
SIXTH GALAXY READER, THE  GOLD, HORACE L.  HA
SEVENTH GALAXY READER, THE  POHL, FREDERIK  HA
NINTH GALAXY READER, THE  POHL, FREDERIK  PA,C/PA

The MITSFS Titledex  September 24, 2007  Page 142
TITLE  AUTHOR  CODE
ELEVENTH GALAXY READER, THE  POHL, FREDERIK  HA
GALAXY: THE BEST OF MY YEARS  BAEN, JAMES  C/PA,PA

GALAXY: THIRTY YEARS OF INNOVATIVE SCIENCE FICTION POHL, FREDERIK/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

SR-HA,HA

GALAXY: VOLUME 2 POHL, FREDERIK/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

PA

GALAXY’S END  LUPOFF, RICHARD A.  P
GALEN BEKNIGHTED [@DRAGONLANCE HEROES II #3]  WILLIAMS, MICHAEL  C/P,P
GALILEE  BARKER, CLIVE  H
GALILEO’S CHILDREN  DOZOIS, GARDNER  HA
GALLAGHER’S GLACIER  RICHMOND, WALT/RICHMOND, LEIGH  P,D27235
GALLATIN DIVERGENCE, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P,P
GALLERY OF HORROR  GRANT, CHARLES L.  PA
GALLERY OF THE SAINTS, A  GARRETT, RANDALL  [C/RAN-P]
GALLICENAE [@KING OF YS 2]  ANDERSON, POUL/ANDERSON, KAREN  C/P,SFWA-P,P
GALLIFREY CHRONICLES, THE [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
GALVESTON [RESURRECTION MAN 2]  STEWART, SEAN  LP
GAMBLER’S FORTUNE, THE [EINARINN 3]  MCKENNA, JULIET E.  P
GAME, THE [DARK FORCES 1]  LOGAN, LES  P
GAME BEYOND, THE  SCOTT, MELISSA  C/P,P
GAME OF EMPIRE, THE [DOMINIC FLANDRY]  ANDERSON, POUL  C/P,P,H
GAME OF FOX AND LION, THE [FOX 1]  CHASE, ROBERT R.  P
GAME OF THRONES, A [SONG OF ICE AND FIRE 1]  MARTIN, GEORGE R. R.  C/P:2,P,H
GAME OF UNIVERSE, A  NYLUND, ERIC S.  P
GAME OF WORLDS, THE [@OUT OF TIME #3]  ALLEN, ROGER MACBRIDE  P
GAME OF X, THE  SHECKLEY, ROBERT  P
GAME OF YOU, A [@SANDMAN #5]  GAIMAN, NEIL  H,CX
GAME-PLAYERS OF TITAN, THE  DICK, PHILIP K.  C/P,P,H
GAMEARTH [GAMEARTH 1]  ANDERSON, KEVIN J.  P
GAMEPLAY [GAMEARTH 2]  ANDERSON, KEVIN J.  P
GAMEPLAYERS OF ZAN, THE [LER 1]  FOSTER, M. A.  C/P:2,P
GAME’S END [HIDDEN REALMS 3]  GREEN, SHARON  P
GAMES OF COMMAND  SINCLAIR, LINNEA  P
GAMES OF NEITH, THE  SAINT CLAIR, MARGARET  DD453
GAMES PEOPLE PLAY  BERNE, ERIC  [C/RAN-H]
GAMES PSYBORGS PLAY  BARBET, PIERRE  C/P,P
GAMESMAN, THE  MALZBERG, BARRY N.  C/P,P
GANDALARA CYCLE I, THE [GANDALARA 1,2,3]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
GANDALARA CYCLE II, THE [GANDALARA 4,5,6]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
GANGDOM’S DOOM  GRANT, MAXWELL  P
GANGLION AND OTHER STORIES  WIGHTMAN, WAYNE  SR-H
GANGREL [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
GANGSTER CONSPIRACY, THE [@STAR RISK,LTD.]  PERRY, STEVE/PERRY, DAL  P
GANWOLD’S CHILD [UNIFIED WORLDS 1]  THORNLEY, DIANN  H,P
GANYMEDE CLUB, THE [BATTACHARIYA 2]  SHEFFIELD, CHARLES  H,P
GANYMEDE TAKEOVER, THE  DICK, PHILIP K./NELSON, RAY FARADAY  C/P:2,P
GARAN THE ETERNAL  NORTON, ANDRE  C/P,P,H
GARBAGE CHRONICLES, THE  HERBERT, BRIAN  C/P,P
GARBAGE WORLD  PLATT, CHARLES  C/P:2,P
GARDEN, THE [STAR TREK VOYAGER]  SCOTT, MELISSA  P
GARDEN OF EDEN, THE  BRAND, MAX  P
GARDEN OF EVIL [@CHARMED]  HARRISON, EMMA  P
GARDEN OF EVIL, THE  STOKER, BRAM  P
GARDEN OF FEAR, THE  CRAWFORD PUBLICATIONS  SR-PA
GARDEN OF PARADISE, THE  WEIGALL, ARTHUR  H
GARDEN OF RAMA, THE [RAMA 3]  CLARKE, ARTHUR C./LEE, GENTRY  H:2,P
GARDEN OF SALT [HUMILITY GARDEN 1,2]  SAVAGE, FELICITY  H
GARDEN OF THE SHAPED, THE [SHAPER EXILE 1]  FINCH, SHEILA  C/P:2,P
GARDEN OF THE STONE, THE  STRAUSS, VICTORIA  C/P,P
GARDEN OF UNEARTHLY DELIGHTS, THE  RANKIN, ROBERT  P
GARDEN OF WINTER, THE  EKLUND, GORDON  P
GARDEN ON THE MOON  BOULLE, PIERRE  C/P,P,H
GARDENS OF DELIGHT, THE  WATSON, IAN  C/P:2,P
GARDENS OF FEAR  HOWARD, ROBERT E.  H
GARDENS OF THE MOON [MALAZAN 1]  ERIKSON, STEVEN  C/P,P,H:2
GARDENS ONE TO FIVE  TATE, PETER  H,SFWA
GARGANTUA AND PANTAGRUEL  RABELAIS  P
GARGOYLE, THE  LOVISI, GARY  B
GARGOYLE, THE (REVISED)  LOVISI, GARY  LP
GARIN DEATH RAY, THE  TOLSTOI, ALEXEI  H:2
GARMENTS OF CAEAN, THE (U.S. EDITION)  BAYLEY, BARRINGTON J.  C/P,SFWA,P,H
GARRETT FILES, THE [GARRETT 1,2,3]  COOK, GLEN  H
GARRETT INVESTIGATES [GARRETT 7,8,9]  COOK, GLEN  H
GARRETT ON THE CASE [GARRETT 10,11]  COOK, GLEN  H

TITLE  AUTHOR  CODE
GARRETT, P.I. [GARRETT 4,5,6]  COOK, GLEN  H
GARTH OF IZAR [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  P
GARUKAN BLOOD [DR. BONES]  WYLDE, THOMAS  P
GAS!  HIRSCHFELD, BURT  C/P,P
GASHLYCRUMB TIMES 2000 CALENDAR, THE  GOREY, EDWARD  CAL
GASP, THE  GARY, ROMAIN  C/P,P
GATE, THE [@EXPLORATION #4]  RANDLE, KEVIN  P
GATE OF DARKNESS, CIRCLE OF LIGHT [HUFF 3]  HUFF, TANYA  C/P:2,P
GATE OF FIRE, THE [OATH OF EMPIRE 2]  HARLAN, THOMAS  H,P
GATE OF GODS, THE [FALL OF ILE-RIEN 3]  WELLS, MARTHA  H
GATE OF IVORY, THE [IVORY 1]  EGAN, DORIS  C/P:2,P
GATE OF IVORY, GATE OF HORN [RYHOPE WOOD 3]  HOLDSTOCK, ROBERT P.  P
GATE OF IVREL [MORGAINE 1]  CHERRYH, C. J.  C/P:2,P
GATE OF THE CAT, THE [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P,SFWA,P,H
GATE OF TIME, THE  FARMER, PHILIP JOSE  C/P:2,P
GATE OF WORLDS, THE  SILVERBERG, ROBERT  C/P:2,P
GATE TO WOMEN’S COUNTRY, THE  TEPPER, SHERI S.  C/P,SR-LP,P,H:2
GATEKEEPERS, THE  GRAHAM, DANIEL, JR.  H,P
GATES OF CREATION, THE [WORLD OF TIERS 2]  FARMER, PHILIP JOSE  C/P:2,P
GATES OF DAWN, THE [@BLOOD AND STONE #2]  NEWCOMB, ROBERT  C/P,P,H
GATES OF DAWN, THE (PROOF) [BLOOD AND STONE 2]  NEWCOMB, ROBERT  H,SR-LP
GATES OF EDEN, THE  STABLEFORD, BRIAN M.  C/P,P
GATES OF HADES, THE  BETANCOURT, JOHN GREGORY  P
GATES OF HEAVEN, THE [SEVEN BROTHERS 3]  BENJAMIN, CURT  H
GATES OF HEAVEN, THE  PREUSS, PAUL  C/P:2,P
GATES OF HELL, THE [HEROES IN HELL]  CHERRYH, C. J./MORRIS, JANET  C/P,P,H
GATES OF NIGHT, THE [@EBERRON DREAMING DARK #3]  BAKER, KEITH  P
GATES OF NOON, THE [STEPHEN FISHER 2]  ROHAN, MICHAEL SCOTT  C/P,SFWA,P
GATES OF SLEEP, THE [ELEMENTAL MASTERS 2]  LACKEY, MERCEDES  H
GATES OF THE UNIVERSE  COULSON, ROBERT/DE WEESE, GENE  SR-P,C/P:2
GATES OF THORBARDIN, THE [@DRAGONLANCE HEROES II #2]  PARKINSON, DAN  C/P:2,P
GATES OF TIME, THE  BARRETT, NEAL JR.  D27400
GATES OF TWILIGHT, THE  VOLSKY, PAULA  LP
GATES OF VENSUNOR, THE [@SHUNLAR #1]  HELLER, CAROL  P
GATES OF WINTER, THE [LAST RUNE 5]  ANTHONY, MARK  P
GATES OF WINTER, THE (PROOF) [LAST RUNE 5]  ANTHONY, MARK  SR-LP
GATES TO TOMORROW  NORTON, ANDRE/DONALDY, ERNESTINE  HA
GATEWAY [HEECHEE 1]  POHL, FREDERIK  C/P,P,H
GATEWAY TO DOOM [@NEW TOM SWIFT #9]  APPLETON, VICTOR  LP
GATEWAY TO ELSEWHERE  LEINSTER, MURRAY  SR-DD053
GATEWAY TO HELL  WHEATLEY, DENNIS  P
GATEWAY TO LIMBO  LAMPTON, CHRISTOPHER  H:2,SFWA
GATEWAY TO NEVER, THE [GRIMES]  CHANDLER, A. BERTRAM  D37062
GATEWAY TO REMEMBRANCE  CRADOCK, PHYLLIS  H
GATEWAY TO THE STARS [@TRAVELLER 1]  ASKEGREN, PIERCE  P
GATEWAY TO XANADU [DIANA SANTEE 2]  GREEN, SHARON  P
GATEWAY TRIP, THE [HEECHEE 5]  POHL, FREDERIK  C/P,P,H
GATEWAYS  GREENBERG, MARTIN H.  PA
GATEWAYS TO FOREVER  ASHLEY, MIKE  R/REF-H
GATHER, DARKNESS!  LEIBER, FRITZ  C/P:2,SFWA,LP
GATHER IN THE HALL OF THE PLANETS  O’DONNELL, K. M.  D27415
GATHERER OF CLOUDS [INITIATE BROTHER 2]  RUSSELL, SEAN  C/P:2,SFWA-P,P
GATHERING, THE [JUSTICE 3]  HAMILTON, VIRGINIA  C/P,P
GATHERING DARK, THE [SHADOW SAGA]  GOLDEN, CHRISTOPHER  C/P,P
GATHERING DARK, THE [@MAGIC: THE GATHERING ICE AGE 1]  GRUBB, JEFF  P
GATHERING DARKNESS [@MOONRUNNER 2]  TOOMBS, JANE  P
GATHERING EVIL, A [DARK CONSPIRACY 1]  STACKPOLE, MICHAEL A.  P
GATHERING FLAME, THE [MAGEWORLDS 4]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
GATHERING OF GARGOYLES, A [DARKANGEL 2]  PIERCE, MEREDITH ANN  P,SFWA-P
GATHERING OF HEROES, A [@DARK BORDER]  ZIMMER, PAUL EDWIN  C/P,P
GATHERING OF STARS, A [MECHANICAL SKY 2]  MOFFITT, DONALD  P
GATHERING OF STONES, A [SOUL DRINKER 3]  CLAYTON, JO  C/P:2,P
GATHERING OF THE SINISTER SIX, THE [SPIDER-MAN]  CASTRO, ADAM-TROY  P
GATHERING OF WIDOWMAKERS, A [WIDOWMAKER 4]  RESNICK, MIKE  H
GATHERING STORM, THE [CROWN OF STARS 5]  ELLIOTT, KATE  H
GAUDY SHADOWS, THE  BRUNNER, JOHN  SR-P,C/P:2
GAUNTLET [@STAR TREK TNG STARGAZER #1]  FRIEDMAN, MICHAEL JAN  P
GAUNTLET OF MALICE, THE [MAGES OF GARILLON 2]  HARRIS, DEBORAH TURNER  C/P,SFWA,SFWA-P,LP
GEDANKENNETZ, DAS  FRANKE, HERBERT W.  FH
GEEK SPACE NINE [@STAR WRECK #6]  REWOLINSKI, LEAH  P
GEHIRN IN FESSELN [@PERRY RHODAN SILBERBANDE #70]  PABEL MOEWIG KG  FPR
GEIS OF THE GARGOYLE [@XANTH #18]  ANTHONY, PIERS  C/P,P
GELIEBTER ROBOTER  ASIMOV, ISAAC  FP
GEMINI [STAR TREK]  BARR, MIKE W.  P

TITLE  AUTHOR  CODE
GEMINI RISING  FILBRUN, J. S.  P
GEMINI SMILE, GEMINI KILL [@HORRORSCOPE #4]  LORY, ROBERT  C/P,P
GEMSTONES  CAMPBELL, TYREE  B
GEMWORLD, BOOK ONE [@STAR TREK TNG]  VORNHOLT, JOHN  C/P,P
GEMWORLD, BOOK TWO [@STAR TREK TNG]  VORNHOLT, JOHN  C/P,P
GENDER GENOCIDE  COOPER, EDMUND  P
GENDERFLEX  TAN, CECILIA  LPA
GENE RODDENBERRY: THE LAST CONVERSATION  FERN, YVONNE  R/REF-P
GENE RODDENBERRY: THE MYTH AND THE MAN BEHIND STAR TREK  ENGEL, JOEL  R/REF-H
GENE WOLFE’S BOOK OF DAYS  WOLFE, GENE  H,SFWA
GENELLAN: FIRST VICTORY [@GENELLAN #3]  GIER, SCOTT G.  P
GENELLAN: IN THE SHADOW OF THE MOON [@GENELLAN #2]  GIER, SCOTT G.  P
GENELLAN: PLANETFALL [@GENELLAN #1]  GIER, SCOTT G.  P
GENERAL ZAPPED AN ANGEL, THE  FAST, HOWARD  C/P,P,H
GENERAL’S PRESIDENT, THE  DALMAS, JOHN  C/P,SFWA-P,P
GENERATION  GERROLD, DAVID  PA,C/PA
GENERATION GAP, THE [@STAR WRECK #1]  REWOLINSKI, LEAH  C/P,P
GENERATION REMOVED, A  WOLF, GARY K.  H,SFWA
GENERATION WARRIORS [PLANET PIRATES 3]  MCCAFFREY, ANNE/MOON, ELIZABETH  C/P:2,SFWA-P,P
GENESIS  ANDERSON, POUL  C/P,P
GENESIS  HARBINSON, W. A.  C/P:2,P
GENESIS [@ROBOTECH #1]  MCKINNEY, JACK  C/P:2,P
GENESIS FIVE  ALLEN, HENRY WILSON  C/P:2,P
GENESIS FORCE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS MACHINE, THE  HOGAN, JAMES P.  C/P,SR-P,H
GENESIS PROTOCOL, THE  WARD, DAYTON  LP
GENESIS QUEST, THE [GENESIS 1]  MOFFITT, DONALD  C/P:2,P
GENESIS SHIELD, THE  SPRUILL, STEVEN G.  C/P,P
GENESIS TWO  DAVIES, L. P.  C/P,P,H
GENESIS WAVE BOOK ONE, THE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS WAVE BOOK TWO, THE [STAR TREK TNG]  VORNHOLT, JOHN  P
GENESIS WAVE BOOK THREE, THE [STAR TREK TNG]  VORNHOLT, JOHN  H,P
GENESIS WEB, THE [BUCK ROGERS INVADERS OF CHARON 1]  BRENNAN, C. M.  P
GENETIC BOMB  OFFUTT, ANDREW J./BERRY, D. BRUCE  C/P,P
GENETIC BUCCANEER, THE [@CAP KENNEDY #12]  KERN, GREGORY  C/P,P
GENETIC GENERAL, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,DD449,P
GENETIC SOLDIER  TURNER, GEORGE  H:2,P
GENETOPIA  BROOKE, KEITH  H
GENEVIEVE UNDEAD [WARHAMMER GENEVIEVE 2]  YEOVIL, JACK  P
GENIUS LOCI  SMITH, CLARK ASHTON  H
GENIUS LOCI AND OTHER TALES  SMITH, CLARK ASHTON  SR-P
GENIUS UNLIMITED  PHILLIFENT, JOHN T.  C/P:2,P
GENOCIDAL HEALER, THE [SECTOR GENERAL 8]  WHITE, JAMES  H:2,P
GENOCIDE [@ALIENS 4]  BISCHOFF, DAVID F.  P
GENOCIDE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
GENOCIDES, THE  DISCH, THOMAS M.  C/P:2,P
GENOMETRY  DANN, JACK/DOZOIS, GARDNER  PA
GENPEI  DALKEY, KARA  H
GENTLE GIANTS OF GANYMEDE, THE [GIANTS 2]  HOGAN, JAMES P.  C/P:2,SR-P
GENTLE INVADERS  SANTESSON, HANS STEFAN  C/PA,PA
GENTLE SEDUCTION, THE  STIEGLER, MARC  P,SFWA-P
GENTLEMAN FROM ANGELL STREET, THE  EDDY, MURIEL E./EDDY, C. M. JR.  R/REF-P
GENTLEMAN INTO GOOSE  WARD, CHRISTOPHER  SR-H
GENTLEMAN JUNKIE  ELLISON, HARLAN  P
GENTLEMEN IN HADES  KUMMER, FREDERIC ARNOLD  H
GENUS HOMO  DE CAMP, L. SPRAGUE/MILLER, P. SCHUYLER SR-H,C/P:2,P
GEODESIC DREAMS  DOZOIS, GARDNER  P
GEODESICA ASCENT [GEODESICA 1]  WILLIAMS, SEAN/DIX, SHANE  P
GEODESICA DESCENT [GEODESICA 2]  WILLIAMS, SEAN/DIX, SHANE  P
GEOMANCER [MANCER 3]  CALLANDER, DON  C/P,SFWA-P:2,P
GEORGE MACDONALD  REIS, RICHARD H.  R/REF-H
GEORGIA ON MY MIND AND OTHER PLACES  SHEFFIELD, CHARLES  LP
GERALD’S GAME  KING, STEPHEN  P
GERFALCON [NEUSTRIAN CYCLE 1]  BARRINGER, LESLIE  LP
GERNSBACK DAYS, THE  ASHLEY, MIKE/LOWNDES, ROBERT A. W.  R/REF-H
GET OFF MY WORLD  BINDER, EANDO  C/P,P
GET OFF MY WORLD!  LEINSTER, MURRAY  C/P:2,P
GET OFF THE UNICORN  MCCAFFREY, ANNE  C/P:2,P
GET OUT OF MY SKY  MARGULIES, LEO  PA,C/PA:2
GET SMART! [GET SMART 1]  JOHNSTON, WILLIAM  P
GET SMART ONCE AGAIN! [GET SMART 3]  JOHNSTON, WILLIAM  C/P,P
GETAWAY SPECIAL, THE  OLTION, JERRY  H
GETAWAY WORLD [FAMILY D’ALEMBERT #4]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
GETTING BACK  DIETRICH, WILLIAM  H

The MITSFS Titledex  September 24, 2007  Page 145
TITLE  AUTHOR  CODE
GETTING HOME  BUSBY, F. M.  C/P,P
GETTING INTO DEATH  DISCH, THOMAS M.  H
GETTING INTO DEATH AND OTHER STORIES  DISCH, THOMAS M.  C/P,P
GETTYSBURG [GETTYSBURG 1]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
GEVANGENEN VAN ATLANTIS, DE  ANDERSON, POUL  FP
GHOLAN GATE, THE [@CAP KENNEDY #7]  KERN, GREGORY  P
GHOST  ANTHONY, PIERS  C/P,SFWA-P,P,H
GHOST AND MRS. MUIR, THE  DENHAM, ALICE  P
GHOST AND MRS. MUIR, THE  DICK, R. A.  H
GHOST ARMY, THE [@ORON #5]  SMITH, DAVID C.  P
GHOST BELONGED TO ME, THE [BLOSSOM CULP 1]  PECK, RICHARD  P
GHOST BOOK, THE  ASQUITH, CYNTHIA  PA
GHOST BREAKER  GOULART, RON  D11182
GHOST BRIGADES, THE [OLD MAN’S WAR 2]  SCALZI, JOHN  H
GHOST DANCE [@STAR TREK VOYAGER DARK MATTERS #2]  GOLDEN, CHRISTIE  P
GHOST DANCE  NORMAN, JOHN  P
GHOST DANCE  PTACEK, KATHRYN  SFWA-P
GHOST DANCERS [@TRAVELER #13]  DRUMM, D. B.  P
GHOST DEVICES [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  BUCHER-JONES, SIMON  C/P,P
GHOST FROM THE GRAND BANKS, THE  CLARKE, ARTHUR C.  C/P,P,H
GHOST GARDEN, THE  RIVES, AMELIE  H
GHOST HOUSE [GHOST HOUSE 1]  MCNALLY, CLAIRE  P
GHOST HOUSE REVENGE [GHOST HOUSE 2]  MCNALLY, CLAIRE  P
GHOST IN THE SHELL  SHIROW, MASAMUNE  CX
GHOST KING [STONES OF POWER 1]  GEMMELL, DAVID  C/P,P
GHOST LEGION, THE [DOC SAVAGE 3]  ROBESON, KENNETH  H
GHOST LEGION [@STAR OF THE GUARDIANS #4]  WEIS, MARGARET  C/P:2,P
GHOST LIGHT, THE  LEIBER, FRITZ  P,LP
GHOST MAKERS, THE  GRANT, MAXWELL  P
GHOST MANSION  WILLIAMSON, J. N.  P
GHOST OF A CHANCE [STAR TREK VOYAGER]  GARLAND, MARK A./MCGRAW, CHARLES G. C/P:2,P
GHOST OF A CHANCE  NYE, JODY LYNN  C/P:2,SFWA-P,P
GHOST OF DIBBLE HOLLOW, THE  WALLACE, MAY NICKERSON  P
GHOST OF THE JEDI [@STAR WARS GALAXY OF FEAR #5]  WHITMAN, JOHN  P
GHOST OF THE REVELATOR, THE [JOHAN ESBACH 2]  MODESITT, L. E. JR.  H,P
GHOST OF THE WALL [@AGE OF CONAN MARAUDERS #1]  MARIOTTE, JEFF  P
GHOST OF THE WELL OF SOULS [WELL WORLD]  CHALKER, JACK L.  P
GHOST OF THE WHITE NIGHTS [JOHAN ESBACH 3]  MODESITT, L. E. JR.  P
GHOST OF WINTER [@MECHWARRIOR]  KENSON, STEPHEN  P
GHOST PIRATES, THE  HODGSON, WILLIAM HOPE  SR-P,LP
GHOST PIRATES AND OTHER REVENANTS OF THE SEA, THE  HODGSON, WILLIAM HOPE  H
GHOST ROADS [@BUFFY GATEKEEPER 2]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P:2,P
GHOST SHADOW [NETWORK-CONSORTIUM 4]  FRANKLIN, CHERYL J.  P
GHOST SHIP [STAR TREK TNG]  CAREY, DIANE  C/P:2,P
GHOST SHIP [DOCTOR WHO]  TOPPING, KEITH  SR-H
GHOST SISTER, THE  WILLIAMS, LIZ  C/P,P
GHOST SONG  DANIELS, DOROTHY  P
GHOST STORIES  FURMAN, A. L.  PA
GHOST STORIES  OPAR PRESS  R/IND
GHOST STORIES OF AN ANTIQUARY  JAMES, M. R.  LP
GHOST STORY  STRAUB, PETER  P
GHOST TRAIN  LAWS, STEPHEN  SFWA-P
GHOST-WALKER [STAR TREK]  HAMBLY, BARBARA  C/P:3,P
GHOST WAR [@MECHWARRIOR DARK AGE 1/BATTLETECH]  STACKPOLE, MICHAEL A.  P
GHOSTBOAT  SIMPSON, GEORGE E./BURGER, NEAL R.  P
GHOSTBUSTERS  MUELLER, RICHARD  P
GHOSTBUSTERS 2 [GHOSTBUSTERS]  NAHA, ED  P
GHOSTBUSTERS: THE RETURN [GHOSTBUSTERS]  FISCH, SHOLLY  P
GHOSTCOUNTRY’S WRATH  DEITZ, TOM  P
GHOSTHUNT [DIADEM 7]  CLAYTON, JO  C/P:2,P
GHOSTHUNTERS, THE  TOFTE, ARTHUR  P
GHOSTLIGHT  BRADLEY, MARION ZIMMER  H:2,LP
GHOSTLY TALES OF HENRY JAMES  JAMES, HENRY  LP
GHOSTMAKER [WARHAMMER 40K GAUNT’S GHOSTS 2]  ABNETT, DAN  P

GHOSTS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

GHOSTS  MAYNE, WILLIAM  HA
GHOSTS AND THINGS  CANTOR, HAL  PA
GHOSTS, CASTLES, AND VICTIMS: TALES OF GOTHIC HORROR  WOLF, JACK C./WOLF, BARBARA  PA
GHOSTS I HAVE BEEN [BLOSSOM CULP 2]  PECK, RICHARD  C/P:2,P
GHOSTS IN IRISH HOUSES  REYNOLDS, JAMES  [C/RAN-P]
GHOSTS IN THE SNOW [DUBRIC BYERLY 1]  JONES, TAMARA SILER  C/P,P
GHOSTS OF EPIDORIS, THE [@CAP KENNEDY #14]  KERN, GREGORY  C/P,P
GHOSTS OF GOL, THE [@PERRY RHODAN #10]  ACE  PA,C/PA

The MITSFS Titledex  September 24, 2007  Page 146
TITLE  AUTHOR  CODE
GHOSTS OF HARREL, THE  ROBERTS, WILLO DAVIS  P
GHOSTS OF MANACLE, THE  FINNEY, CHARLES G.  C/P:2,P
GHOSTS OF N-SPACE, THE [DOCTOR WHO MISSING ADVENTURES]  LETTS, BARRY  C/P,P
GHOSTS OF ONYX [@HALO]  NYLUND, ERIC S.  P
GHOSTS OF SLEATH, THE  HERBERT, JAMES  H,P
GHOSTS OF SLEATH, THE (PROOF)  HERBERT, JAMES  SR-LP
GHOSTS OF THE HEAVISIDE LAYER AND OTHER FANTASMS, THE  DUNSANY, LORD  H
GHOSTS WHO WENT TO SCHOOL  SPEARING, JUDITH  LP
GHOSTWALKER [@FORGOTTEN REALMS FIGHTERS]  DEBIE, ERIK SCOTT  P
GHOSTWORLD  GREEN, SIMON R.  P
GHOUL KEEPERS, THE  MARGULIES, LEO  PA,C/PA
GHOUL LOVER  TRALINS, ROBERT  P
GHOUL TROUBLE [BUFFY]  PASSARELLA, JOHN  P
GHOULS, THE  HAINING, PETER  HA,PA
GHOULS IN MY GRAVE  RAY, JEAN  SR-P
GIANT AMONG US, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #2]  DENNING, TROY  P
GIANT ANTHOLOGY OF SCIENCE FICTION, THE  MARGULIES, LEO/FRIEND, OSCAR J.  HA
GIANT LEAP, THE  BERRY, ADRIAN  R/REF-H
GIANT LIZARDS FROM ANOTHER STAR  MACLEOD, KEN  H
GIANT OF WORLD’S END [GONDWANE]  CARTER, LIN  P
GIANT RAT OF SUMATRA, THE  BOYER, RICHARD L.  P

GIANTS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

GIANTS FROM ETERNITY  WELLMAN, MANLY WADE  H
GIANTS IN THE DUST  OLIVER, CHAD  C/P,P
GIANTS NOVELS, THE [GIANTS 1,2,3]  HOGAN, JAMES P.  C/P,P,H
GIANT’S PARTNER, THE [@PERRY RHODAN #33]  ACE  PA,C/PA
GIANTS’ STAR [GIANTS 3]  HOGAN, JAMES P.  C/P:2,P
GIANTS UNLEASHED  CONKLIN, GROFF  PA,C/PA
GIANTSLAYER [WARHAMMER GOTREK & FELIX 7]  KING, WILLIAM  P
GIBBON’S DECLINE AND FALL  TEPPER, SHERI S.  C/P,P,H
GIBRALTAR EARTH  MCCOLLUM, MICHAEL  LP
GIDEON [@NIGHTWALKERS]  FRANK, JACQUELYN  P
GIDEON’S WALL  KURZAWA, GREG  LP
GIFT, THE  O’LEARY, PATRICK  H,LP
GIFT FROM EARTH, A [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
GIFT FROM THE STARS  GUNN, JAMES E.  LP
GIFT OF DRAGONS, A [PERN]  MCCAFFREY, ANNE  H
GIFT OF MAGIC, A  DUNCAN, LOIS  C/P:2,P
GIFT OF PROPHECY, A  MONTGOMERY, RUTH  [C/RAN-P]
GIFT OF THE GODS, THE  JONES, RAYMOND F.  B
GIFT OF THE MANTI  BONE, J. F./MYERS, ROY  SR-P
GIFT UPON THE SHORE, A  WREN, M. K.  P
GIFTS  LE GUIN, URSULA K.  P
GIFTS OF ASTI AND OTHER STORIES OF SCIENCE FICTION, THE  ELWOOD, ROGER  HA
GIFTS OF BLOOD  PETREY, SUSAN  P,SFWA-P
GIFTS OF THE GORBODUC VANDAL, THE  WILLIAMS, PAUL O.  C/P:2,P
GIFTWISH  MARTIN, GRAHAM DUNSTAN  H
GIGGLING GHOSTS, THE [DOC SAVAGE #56]  ROBESON, KENNETH  C/P:2,@P
GILDED CHAIN, THE [KING’S BLADES]  DUNCAN, DAVE  H,P
GILDED MAN, THE  SMYTH, CLIFFORD  H
GILDED SARCOPHAGUS, THE [PAUL HOLTON 1]  HUNT, CHARLOTTE  P
GILDEN-FIRE  DONALDSON, STEPHEN R.  H:2,SR-H
GILEAD BOMB, THE  SINCLAIR, DAVID  C/P,P
GILEAD’S BLOOD [WARHAMMER GILEAD]  ABNETT, DAN/VINCENT, NIK  P
GILES GOAT-BOY  BARTH, JOHN  P
GILFEATHER [@ISLES OF GLORY #2]  LARKE, GLENDA  P
GILGAMESH  GRUNDY, STEPHAN  H
GILGAMESH THE KING [GILGAMESH]  SILVERBERG, ROBERT  C/P,P,H:2
GILPIN’S SPACE  BRETNOR, REGINALD  P
GIL’S ALL FRIGHT DINER  MARTINEZ, A. LEE  LP
GINGER STAR, THE [SKAITH 1]  BRACKETT, LEIGH  C/P:2,P
GIORNO PRIMA DELL’ETERNITA, IL  LAUMER, KEITH  FP
GIOVANNI [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  ACHILLI, JUSTIN  P
GIRL FACTORY, THE [GIRL FACTORY 1]  MURPHY, ROBERT FRANKLIN  C/P,P
GIRL FROM EROSPHERE, THE [YOLANDA 1]  VERSEAU, DOMINIQUE  SR-L
GIRL FROM FARRIS’S, THE  BURROUGHS, EDGAR RICE  SR-VLP
GIRL FROM MARS, THE  WILLIAMSON, JACK/BREUER, DR. MILES J.  B,GRD17
GIRL FROM NOWHERE, THE [IRA VEDDER 2]  BOND, EVELYN  P
GIRL FROM NOWHERE, THE  MINTON, PAULA  P
GIRL FROM THE EMERALINE ISLAND, THE  BLUM, ROBERT S.  C/P:3,P
GIRL FROM TOMORROW, THE  ZEIGFREID, KARL  SR-P
GIRL IN A SWING, THE  ADAMS, RICHARD  C/P,P
GIRL IN LANDSCAPE  LETHEM, JONATHAN  LP

The MITSFS Titledex  September 24, 2007  Page 147
TITLE  AUTHOR  CODE
GIRL IN THE GOLDEN ATOM, THE  CUMMINGS, RAY  H,LP
GIRL IN THE HAIRY PAW, THE  GOTTESMAN, RONALD/GEDULD, HARRY  R/REF-H
GIRL OF MANY PARTS, A  SHARMAT, MITCHELL  P
GIRL, THE GOLD WATCH AND EVERYTHING, THE  MACDONALD, JOHN D.  C/P,P
GIRL WHO HEARD DRAGONS, THE  MCCAFFREY, ANNE  H:2,SFWA-P,P
GIRL WHO LOVED TOM GORDON, THE  KING, STEPHEN  P
GIRL WHO WAS PLUGGED IN, THE  TIPTREE, JAMES JR.  TD7,SFWA-TD7
GIRL WHO WAS POSSESSED, THE  BROWN, CARTER  P
GIRL WITH A SYMPHONY IN HER FINGERS, THE  CONEY, MICHAEL G.  H
GIRL WITH SOMETHING EXTRA, THE  FAIRMAN, PAUL W.  P
GIRL WITH THE HUNGRY EYES, THE  AVON  SR-PA
GIRL WITH THE JADE GREEN EYES, THE  BOYD, JOHN  SR-P
GIRL WITH THE SILVER EYES, THE  ROBERTS, WILLO DAVIS  P
GIRLS FROM PLANET 5, THE  WILSON, RICHARD  SR-H,C/P:2,P
GIRL’S GUIDE TO VAMPIRES, A  MACALISTER, KATIE  P
GISMO FROM OUTER SPACE, THE  LAZARUS, KEO FELKER  LP
GIST HUNTER, THE [ARCHONATE/HENGHIS HAPTHORN]  HUGHES, MATTHEW  H
GIVE ME LIBERTY  GREENBERG, MARTIN H./TIER, MARK  PA
GIVE ME LIBERTY  MILLER, FRANK/GIBBONS, DAVE  CX
GIVE US THIS MARS [AS IT IS ON MARS 2]  CRONIN, THOMAS W.  H
GIVE WARNING TO THE WORLD  BRUNNER, JOHN  C/P,P
GIVER, THE  LOWRY, LOIS  P
GIVER OF ROSES [@GUARDIANS OF GADIEL]  MORGAN, KATHLEEN  LP
GLADIATOR, THE  EARLS, BILL  P
GLADIATOR  WYLIE, PHILIP  H,LP,P
GLADIATOR-AT-LAW  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
GLADIATORS OF HAPANU [@BLADE #31]  LORD, JEFFREY  P
GLAHMIAN SHOCK  CONRAD, ALLAN  P
GLAICE [ARBITER 1]  DOUGLAS, L. WARREN  P
GLAIVE, THE [IROSHI 2]  OSBORNE, CARY  P
GLAMOUR, THE  PRIEST, CHRISTOPHER  H:2,SFWA
GLASS AND AMBER  CHERRYH, C. J.  SR-H
GLASS BEES, THE  JUENGER, ERNST  LP
GLASS CAGE, THE  LUKENS, ADAM  H
GLASS CAGE, THE  WILSON, COLIN  P
GLASS DRAGON, THE [DRAGON NIMBUS 1]  RADFORD, IRENE  C/P:2,P
GLASS DRAGONS [MOONWORLDS 2]  MCMULLEN, SEAN  P
GLASS EMPIRES [STAR TREK MIRROR]  POCKET BOOKS  LPA
GLASS HAMMER, THE  JETER, K. W.  C/P,P,H
GLASS HARMONICA, THE  MARLEY, LOUISE  C/P,LP
GLASS HOUSES [@MORGANVILLE VAMPIRES 1]  CAINE, RACHEL  P
GLASS HOUSES  MIXON, LAURA J.  P
GLASS MAN, THE [@AVENGER #34]  ROBESON, KENNETH  C/P,P
GLASS MOUNTAIN, THE [@AVENGER #8]  ROBESON, KENNETH  C/P:2,P
GLASS MOUNTAINS, THE  KADOHATA, CYNTHIA  P
GLASS OF DYSKORNIS, THE [GANDALARA 2]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
GLASS OF STARS, A  YOUNG, ROBERT F.  H
GLASS PRISON, THE [@FORGOTTEN REALMS]  COOK, MONTE  P
GLASS-SIDED ANTS’ NEST, THE  DICKINSON, PETER  C/P,P
GLASS SLIPPER, THE  FARJEON, ELEANOR  P
GLASS TEAT, THE  ELLISON, HARLAN  [C/RAN-P:2]
GLASS WARRIOR, THE [@DEMON CROWN #1]  VARDEMAN, ROBERT E.  P,SFWA-P
GLASSHOUSE  STROSS, CHARLES  H
GLASSWRIGHT’S APPRENTICE, THE [GLASSWRIGHT 1]  KLASKY, MINDY L.  C/P,P
GLASSWRIGHT’S JOURNEYMAN, THE [GLASSWRIGHT 3]  KLASKY, MINDY L.  P
GLASSWRIGHT’S PROGRESS, THE [GLASSWRIGHT 2]  KLASKY, MINDY L.  P
GLENRAVEN [GLENRAVEN 1]  BRADLEY, MARION ZIMMER/LISLE, HOLLY  C/P,P
GLIDE PATH  CLARKE, ARTHUR C.  SR-H,C/P:2,P
GLIMMERING  HAND, ELIZABETH  H,P
GLIMPSE OF THE PLANET ZINN, A  TERRY, MELVIN R.  H
GLIMPSES  SHINER, LEWIS  P
GLINDA OF OZ [OZ #14]  BAUM, L. FRANK  C/P,@P,H
GLITTERING PLAIN, THE  MORRIS, WILLIAM  H,LP
GLITTERSPIKE HALL  JEFFERIES, MIKE  P
GLOBAL BLACKOUT  VALLANCE, KARL  SR-P
GLOBAL GLOBULES AFFAIR, THE (BRITISH) [@GIRL FROM UNCLE #1]  LATTER, SIMON  SR-P
GLOBALHEAD  STERLING, BRUCE  C/P:2,P

GLOBE, THE [@SCIENCE OF DISCWORLD #2] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK

P

GLOBES OF LLARUM, THE  LANCOUR, GENE  H,SFWA
GLORIANA, OR THE UNFULFILL’D QUEEN  MOORCOCK, MICHAEL  H:2,LP
GLORIFYING TERRORISM  MENDLESOHN, FARAH  LPA
GLORIOUS POOL, THE  SMITH, THORNE  P
GLORIOUS TREASON [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 2]  RYAN, C. J.  P

The MITSFS Titledex  September 24, 2007  Page 148
TITLE  AUTHOR  CODE
GLORY [GOLDENWING 1]  COPPEL, ALFRED  H
GLORY BE TO MARS [AS IT IS ON MARS 3]  CRONIN, THOMAS W.  H
GLORY GAME, THE  LAUMER, KEITH  C/P,P,H,SFWA
GLORY LANE  FOSTER, ALAN DEAN  C/P:2,SFWA-P,P
GLORY PLANET  CHANDLER, A. BERTRAM  H:2
GLORY ROAD  HEINLEIN, ROBERT A.  C/P,P,H,SFWA
GLORY SEASON  BRIN, DAVID  C/P:2,P,H
GLORY SEASON (EXCERPT)  BRIN, DAVID  SR-LP
GLORY THAT WAS, THE  DE CAMP, L. SPRAGUE  C/P,P,H
GLORYHITS  STICKGOLD, BOB/NOBLE, MARK  C/P,P,H
GLORY’S END [TIME TOURS 2]  BARON, NICK  P
GLORY’S PEOPLE [GOLDENWING 3]  COPPEL, ALFRED  H,P
GLORY’S WAR [GOLDENWING 2]  COPPEL, ALFRED  P
GLOS PANA  LEM, STANISLAW  FP
GLOVE OF MAIDEN’S HAIR, THE  FRIEDMAN, MICHAEL JAN  P
GLOW, THE  STANWOOD, BROOKS  H,P
GLOW OF CANDLES AND OTHER STORIES, A  GRANT, CHARLES L.  P
GNARL!  RUCKER, RUDY  H,LP
GNOME MAN’S LAND [GNOME MAN’S LAND 1]  FRIESNER, ESTHER M.  C/P:2,SFWA-P:2,P
GNOME THERE WAS, A  PADGETT, LEWIS  H
GNOME’S ENGINE, THE [GOBLIN MOON 2]  EDGERTON, TERESA  C/P:2,SFWA-P:2,P
GNOMEWRENCH IN THE DWARFWORKS, THE [GNOMEWRENCH 1]  O’DONOHOE, NICK  C/P,P
GNOMEWRENCH IN THE PEOPLEWORKS, THE [GNOMEWRENCH 2]  O’DONOHOE, NICK  P
GNOMOBILE, THE  SINCLAIR, UPTON  P
GO ASK MALICE [@BUFFY]  LEVY, JOSEPH  H
GO FETCH! [@SABRINA THE TEENAGE WITCH #13]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
GO QUEST, YOUNG MAN  BOGEN, K. B.  P
GO QUEST, YOUNG MAN [@XENA]  EMERSON, RU  P
GO TELL THE SPARTANS [FALKENBERG’S LEGION]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
GOA [BLOOD OF THE GODDESS 1]  DALKEY, KARA  P
GOAT WITHOUT HORNS, THE  SWANN, THOMAS BURNETT  C/P:2,P
GOBLIN HERO [JIG THE GOBLIN 2]  HINES, JIM C.  P
GOBLIN MARKET [WARCHILD 2]  BOWES, RICHARD  P
GOBLIN MIRROR, THE  CHERRYH, C. J.  H:2
GOBLIN MOON [GOBLIN MOON 1]  EDGERTON, TERESA  C/P:2,SFWA-P,P
GOBLIN PLAIN WAR, THE  MILLER, CARL  C/P,P
GOBLIN RESERVATION, THE  SIMAK, CLIFFORD D.  C/P,P,H
GOBLIN TOWER, THE [RELUCTANT KING 1]  DE CAMP, L. SPRAGUE  C/P:2,P
GOBLINQUEST [JIG THE GOBLIN 1]  HINES, JIM C.  H
GOBLINS [X-FILES]  GRANT, CHARLES L.  C/P:2,P
GOBLINS [X-FILES]/WHIRLWIND [X-FILES]  GRANT, CHARLES L.  H,LP

GOBLINS, THE [DOC SAVAGE #125]/ SECRET OF THE SU, THE [DOC SAVAGE 126] ROBESON, KENNETH

@P

GOD AND GOLEM, INC.  WIENER, NORBERT  [C/RAN-H]
GOD BLESS YOU, MR. ROSEWATER, OR PEARLS BEFORE SWINE  VONNEGUT, KURT JR.  H:2,P
GOD BOX, THE  LONGYEAR, BARRY B.  C/P:2,P
GOD DRIVES A FLYING SAUCER  DIONE, R. L.  [C/RAN-P]
GOD EMPEROR OF DUNE [DUNE 4]  HERBERT, FRANK  C/P,P,H,SFWA
GOD GAME  GREELEY, ANDREW M.  H:2,SFWA,SFWA-P,P
GOD HEAD  CLINE, LEONARD  H
GOD IN THE IMAGE OF WOMAN  BERNARD, D. V.  LP
GOD IN THE MOON, THE [@AGE OF CONAN SOLDIER’S QUEST #1]  KNAAK, RICHARD A.  P
GOD KILLER, THE [HAWK & FISHER 3]  GREEN, SIMON R.  C/P:2,P
GOD MACHINE, THE  CAIDIN, MARTIN  C/P:2,SFWA-P,P
GOD MACHINE, THE  WATKINS, WILLIAM JOHN  H,SFWA
GOD OF DEATH [@CASCA #2]  SADLER, BARRY  P
GOD OF PLANET 607, THE  POHLMAN, EDWARD  LP
GOD OF TAROT [TAROT #1]  ANTHONY, PIERS  C/P:2,@P
GOD OF THE GOLDEN FLEECE [GODS 4]  SABERHAGEN, FRED  H
GOD PROJECT, THE  LEE, STANLEY R.  H
GOD PROJECT, THE  SAUL, JOHN  H:2
GOD STALK [JAMETHIEL DREAM-WEAVER 1]  HODGELL, P. C.  H:2,P
GOD: THE ULTIMATE AUTOBIOGRAPHY  PASCALL, JEREMY  H:2
GODA WAR, THE  BLAKENEY, JAY D.  C/P,P
GODBODY  STURGEON, THEODORE  H:2,P
GODBOND [SEA KING 3]  SPRINGER, NANCY  P,SFWA-P
GODDESS OF GANYMEDE, THE  RESNICK, MIKE  SR-H,P
GODDESS OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
GODDESS OF THE ICE REALM [LORD OF THE ISLES 5]  DRAKE, DAVID  H,P
GODDESS WORLDWEAVER, THE [@SEVEN CIRCLES #3]  NILES, DOUGLAS  LP
GODENGINE [DOCTOR WHO NEW ADVENTURES]  HINTON, CRAIG  P
GODFORSAKEN, THE  YARBRO, CHELSEA QUINN  C/P:2,P
GODHEADS  DEVENPORT, EMILY  C/P,P
GODLING, GO HOME!  SILVERBERG, ROBERT  P

The MITSFS Titledex  September 24, 2007  Page 149
TITLE  AUTHOR  CODE
GODMAKERS, THE  BRITAIN, DAN  C/P:2,P
GODMAKERS, THE  HERBERT, FRANK  C/P,P,H
GODMAKERS, THE  PENDLETON, DON  C/P:2,P
GODMAN!  BLOODSTONE, JOHN  SR-P
GODMOTHER, THE [GODMOTHER 1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
GODMOTHERS, THE  HALL, SANDI  LP
GODMOTHER’S APPRENTICE, THE [GODMOTHER 2]  SCARBOROUGH, ELIZABETH ANN  C/P,P
GODMOTHER’S WEB, THE [GODMOTHER 3]  SCARBOROUGH, ELIZABETH ANN  H
GODPLAYER  COOK, ROBIN  P
GODPLAYERS [PLAYERS IN THE CONTEST OF WORLDS 1]  BRODERICK, DAMIEN  LP
GODS ABIDE, THE  SWANN, THOMAS BURNETT  C/P:2,P
GODS ABOVE [@STAR TREK NEW FRONTIER 13]  DAVID, PETER  P
GODS AND ANDROIDS  NORTON, ANDRE  P
GODS AND GOLEMS  DEL REY, LESTER  C/P:2,P
GODS AND PAWNS [COMPANY 8]  BAKER, KAGE  H
GODS AWAKEN, THE [TIMURA 3]  COLE, ALLAN  P
GODS’ CONCUBINE [@TROY GAME #2]  DOUGLASS, SARA  P
GOD’S DICE  SWANN, S. ANDREW  P
GOD’S FIRES  ANTHONY, PATRICIA  C/P:2,P
GODS FOR TOMORROW  SANTESSON, HANS STEFAN  PA
GODS FROM OUTER SPACE  VON DANIKEN, ERICH  [C/RAN-H,C/RAN-P:2]
GODS HATE KANSAS, THE  MILLARD, JOSEPH  C/P:2,P
GODS IN A VORTEX  HOUSTON, DAVID  P
GODS IN ANGER, THE [@OMARAN #4]  COLE, ADRIAN  P
GODS IN WINTER, THE  MILES, PATRICIA  H
GODS LAUGHED, THE  ANDERSON, POUL  C/P:2,P
GODS, MEN AND GHOSTS  DUNSANY, LORD  H,LP
GODS OF AIR AND DARKNESS  MOONEY, RICHARD E.  [C/RAN-P]
GODS OF BAL-SAGOTH, THE  HOWARD, ROBERT E.  C/P,P
GODS OF CERUS MAJOR, THE  RUSE, GARY ALAN  H,SFWA
GODS OF DARKNESS  ZEIGFREID, KARL  SR-P
GODS OF FIRE AND THUNDER [GODS 5]  SABERHAGEN, FRED  H
GODS OF FOXCROFT, THE  LEVY, DAVID  C/P:2,P
GODS OF MARS, THE [BARSOOM #2]  BURROUGHS, EDGAR RICE  C/P,@P,H
GODS OF MARS, THE [BARSOOM 2]/WARLORD OF MARS, THE [BARSOOM 3] BURROUGHS, EDGAR RICE  H:2
GODS OF PEGANA, THE  DUNSANY, LORD  LP
GODS OF RIVERWORLD [RIVERWORLD 5]  FARMER, PHILIP JOSE  H:2,P
GODS OF THE GREATAWAY [SONG OF EARTH 2]  CONEY, MICHAEL G.  H:2
GODS OF THE GREATAWAY (PROOF) [SONG OF EARTH 2]  CONEY, MICHAEL G.  SR-LP
GODS OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 3] CHALKER, JACK L.  H:2,LP
GODS OF WAR, THE  BAEN BOOKS  PA,SFWA-P
GODS OF XUMA, THE  LAKE, DAVID J.  P
GODS OLD AND DARK [WORLD GATES 3]  LISLE, HOLLY  C/P,P
GODS OR DEMONS?  LIGHTNER, A. M.  H
GODS’ TEMPTRESS, THE  GREENFIELD, IRVING A.  C/P,P
GODS THEMSELVES, THE  ASIMOV, ISAAC  C/P:2,H:2,P,SFWA
GODSEND, THE  TAYLOR, BERNARD  P
GODSFIRE  FELICE, CYNTHIA  C/P:2,P
GODSHOME  SHECKLEY, ROBERT  H
GODSLAYER [@SUNDERING #2]  CAREY, JACQUELINE  H
GODSLAYER [BIFROST GUARDIANS 1]  REICHERT, MICKEY ZUCKER  C/P:2,P
GODSPEED  SHEFFIELD, CHARLES  C/P,P,H
GODWHALE, THE  BASS, T. J.  C/P:2,P
GODZILLA  MOLSTAD, STEPHEN  H,P
GODZILLA  THORNE, IAN  [C/RAN-H]
GODZILLA AT WORLD’S END  CERASINI, MARC  P
GOG  SINCLAIR, ANDREW  P
G.O.G. 666  TAINE, JOHN  SR-H
GOING FOR INFINITY  ANDERSON, POUL  H
GOING INTO SPACE  CLARKE, ARTHUR C.  [C/RAN-H]
GOING POSTAL [DISCWORLD 32]  PRATCHETT, TERRY  C/P,P,H
GOING TO SEE THE END OF THE SKY  WATKINS, WILLIAM JOHN  P
GOING TOWARD JASON  CLARK, VIRGINIA ROYCE  LP
GOJIRO  JACOBSON, MARK  P

GOLD AND IRON [@VANCE INTEGRAL EDITION #7]/ CLARGES [@VANCE INTEGRAL EDITION #7]/ LANGUAGES OF PAO, THE [@VANCE INTEGRAL EDITION #7] VANCE, JACK

SR-H

GOLD AT THE STARBOW’S END, THE  POHL, FREDERIK  C/P,P,H
GOLD BOMB, THE [AVENGERS 7]  LAUMER, KEITH  C/P,P
GOLD COAST, THE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  H,SFWA-P,P
GOLD FALCON, THE [@SILVER WYRM #1]  KERR, KATHARINE  H
GOLD OF THE GODS, THE  VON DANIKEN, ERICH  [C/RAN-H,C/RAN-P]
GOLD OGRE, THE [DOC SAVAGE #42]  ROBESON, KENNETH  C/P,@P
GOLD STAR  HUGHES, ZACH  C/P,P

The MITSFS Titledex  September 24, 2007  Page 150
TITLE  AUTHOR  CODE
GOLD: THE FINAL SCIENCE FICTION COLLECTION  ASIMOV, ISAAC  C/P,P,H
GOLD THE MAN  GREEN, JOSEPH  P
GOLD TOOTH, THE  TAINE, JOHN  SR-H
GOLD UNICORN [TANAQUIL 2]  LEE, TANITH  C/P,P,H
GOLDCAMP VAMPIRE, THE  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
GOLDCAMP VAMPIRE, THE (PROOF)  SCARBOROUGH, ELIZABETH ANN  SR-LP
GOLDEN, THE  SHEPARD, LUCIUS  P
GOLDEN AGE, THE  GRAHAME, KENNETH  LP
GOLDEN AGE, THE [GOLDEN AGE 1]  WRIGHT, JOHN C.  H:2
GOLDEN AGE OF SCIENCE FICTION, THE  CONKLIN, GROFF  HA
GOLDEN AGE SF  REYNOLDS, ERIC T.  HA
GOLDEN APPLE, THE [ILLUMINATUS! 2]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P:2,P
GOLDEN APPLES OF THE SUN, THE  BRADBURY, RAY  C/P:2,P,H
GOLDEN BAND OF EDDRIS, THE  MCKENZIE, ELLEN KINDT  H
GOLDEN BARGE, THE  MOORCOCK, MICHAEL  C/P,P
GOLDEN BLIGHT, THE  ENGLAND, GEORGE ALLAN  H
GOLDEN BLOOD  WILLIAMSON, JACK  C/P:2,P

GOLDEN BOATS OF TARADATA AFFAIR, THE (BRITISH)

[@GIRL FROM UNCLE #3] LATTER, SIMON

SR-P

GOLDEN BOOK OF ASTRONOMY  WYLER, ROSE/AMES, GERALD  [C/RAN-H]
GOLDEN CAT, THE [WILD ROAD 2]  KING, GABRIEL  P
GOLDEN CHALICE, THE  FANTHORPE, R. L.  SR-P
GOLDEN CIRCLE, THE [PHANTOM 5]  SHAWN, WILLIAM S.  P
GOLDEN CITIES, FAR  CARTER, LIN  SR-PA,C/PA
GOLDEN COMICS DIGEST #5  WESTERN BOOKS  [C/RAN-P]
GOLDEN COMPASS, THE [HIS DARK MATERIALS 1]  PULLMAN, PHILIP  C/P:2,P,H
GOLDEN DREAM: A FUZZY ODYSSEY [FUZZIES]  MAYHAR, ARDATH  C/P,LP
GOLDEN ENEMY, THE  KEY, ALEXANDER  H
GOLDEN FLEECE, THE  HAWTHORNE, JULIAN  H
GOLDEN FLEECE, THE  ODYSSEY PUBLICATIONS  SR-LPA
GOLDEN FLEECE  SAWYER, ROBERT J.  H,P
GOLDEN FOOL [TAWNY MAN 2]  HOBB, ROBIN  H,P
GOLDEN GIRL AND OTHER STORIES [@VANCE INTEGRAL EDITION #6]  VANCE, JACK  SR-H
GOLDEN GIRL OF HOCKEINBECK, THE  ROXBURY, KYLE  SR-L
GOLDEN GLOBE, THE  VARLEY, JOHN  H,P
GOLDEN GODDESS GAMBIT, THE [@AGENT OF TERRA #2]  MADDOCK, LARRY  C/P:2,P
GOLDEN GROVE, THE  KRESS, NANCY  C/P,SFWA,P
GOLDEN GRYPHON FEATHER, THE [KAPHTU 1]  PURTILL, RICHARD  C/P:2,P
GOLDEN HAWK OF ZANDRAYA, THE  SIROTA, MIKE  P
GOLDEN HELIX, THE  STURGEON, THEODORE  H:2,P
GOLDEN HORN, THE [@LAST VIKING #1]  ANDERSON, POUL  P
GOLDEN HORN, THE [HOUND AND FALCON #2]  TARR, JUDITH  C/P:2,@H,SFWA-P,P
GOLDEN KAZOO, THE  SCHNEIDER, JOHN G.  P
GOLDEN KEY, THE  MACDONALD, GEORGE  H

GOLDEN KEY, THE RAWN, MELANIE/ROBERSON, JENNIFER/

ELLIOTT, KATE

C/P:2,P

GOLDEN MAN, THE  DICK, PHILIP K.  C/P:2,P,H

GOLDEN MAN, THE [DOC SAVAGE #117]/ PERIL IN THE NORTH [DOC SAVAGE 118] ROBESON, KENNETH

@P

GOLDEN NAGINATA, THE [TOMOE GOZEN 2]  SALMONSON, JESSICA AMANDA  C/P:2,P
GOLDEN NINETIES, THE  MASON, LISA  C/P:2,P
GOLDEN OLIVE, THE  CRISTABEL  P
GOLDEN ONE, THE [@ALIEN CHRONICLES #1]  CHESTER, DEBORAH  C/P,P
GOLDEN ORB, THE [@DRAGONLANCE ICEWALL #2]  NILES, DOUGLAS  P
GOLDEN PEOPLE, THE  SABERHAGEN, FRED  C/P:2,DM103,P
GOLDEN PERIL, THE [DOC SAVAGE #55]  ROBESON, KENNETH  C/P,@P
GOLDEN QUEEN, THE [GOLDEN QUEEN 1]  WOLVERTON, DAVE  P
GOLDEN ROAD, THE  KNIGHT, DAMON  HA
GOLDEN ROOMS, THE  FISHER, VARDIS  H
GOLDEN SCORPIO [DRAY PRESCOT 18]  AKERS, ALAN BURT  P
GOLDEN SCORPION, THE  ROHMER, SAX  P
GOLDEN SHIELD OF IBF, THE  AHERN, JERRY/AHERN, SHARON  P
GOLDEN SKULL, THE [@RICK BRANT #10]  BLAINE, JOHN  H:2
GOLDEN SLAVE, THE  ANDERSON, POUL  C/P,P
GOLDEN SPACE, THE  SARGENT, PAMELA  H:2,SFWA,P
GOLDEN STEED, THE [@BLADE #13]  LORD, JEFFREY  P
GOLDEN SUNLANDS  ROWLEY, CHRISTOPHER  C/P,P
GOLDEN SWAN, THE [BOOKS OF THE ISLE 5]  SPRINGER, NANCY  C/P:2,P
GOLDEN SWORD, THE [SILISTRA 2]  MORRIS, JANET  C/P,P
GOLDEN SWORD, THE  PATTON, FIONA  P
GOLDEN, THE (PROOF)  SHEPARD, LUCIUS  SR-LP
GOLDEN THREAD, THE [VALENTINE MARSH 3]  CHARNAS, SUZY MCKEE  LP
GOLDEN TORC, THE [PLIOCENE EXILE 2]  MAY, JULIAN  C/P,P,H
GOLDEN TRANSCENDENCE, THE [GOLDEN AGE 3]  WRIGHT, JOHN C.  H

The MITSFS Titledex  September 24, 2007  Page 151
TITLE  AUTHOR  CODE
GOLDEN TRILLIUM [TRILLIUM 3]  NORTON, ANDRE  H:2,P
GOLDEN TROUBADOUR, THE  BELL, GORDON BENNETT  H
GOLDEN TROUBADOUR, THE (PROOF)  BELL, GORDON BENNETT  SR-LP
GOLDEN VANITY  POLLACK, RACHEL  C/P,P
GOLDEN VIRGIN, THE  DIPPER, ALAN  P
GOLDEN VOYAGE OF SINBAD, THE  HART, STEVE  P
GOLDEN WIND, THE  DE CAMP, L. SPRAGUE  H,P
GOLDEN WINGS AND OTHER STORIES  MORRIS, WILLIAM  LP
GOLDEN WITCHBREED [ORTHE 1]  GENTLE, MARY  C/P,P,H
GOLDEN WORLD, THE [AMERICA 2040 2]  INNES, EVAN  C/P,P
GOLDENGIRL  LEAR, PETER  P
GOLDFINGER [JAMES BOND 7]  FLEMING, IAN  C/P,P
GOLEM 100  BESTER, ALFRED  C/P,P,H
GOLEM IN THE GEARS [@XANTH #9]  ANTHONY, PIERS  C/P,P,H
GOLEM’S EYE, THE [@BARTIMAEUS #2]  STROUD, JONATHAN  H
GOLGOTHA  GARDNER, JOHN (BRITISH, 1926-)  P
GOLGOTHA FALLS  DE FELITTA, FRANK  H
GOLGOTHA RUN [@DARK FUTURE EDDIE KALISH 1]  STONE, DAVE  P
GOLIATH  ALTEN, STEVE  H
GOMORRAH  KARLINS, MARVIN/ANDREWS, LEWIS M.  H
GONE TO BE SNAKES NOW  BELL, NEAL  C/P,P
GONJI [GONJI]  RYPEL, T. C.  P
GOOD AMERICAN, THE  FLICKER, TED  LP
GOOD-BYE TO EARTH  KELLEY, LEO P.  SR-P
GOOD DAY TO DIE, A [@STAR TREK I.K.S. GORKON 1]  DECANDIDO, KEITH R. A.  P
GOOD FAIRIES OF NEW YORK, THE  MILLAR, MARTIN  LP
GOOD NEIGHBORS AND OTHER STRANGERS  PANGBORN, EDGAR  H:2,P
GOOD NEW STUFF, THE  DOZOIS, GARDNER  LPA
GOOD NEWS FROM OUTER SPACE  KESSEL, JOHN  C/P:2,P
GOOD OLD-FASHIONED FUTURE, A  STERLING, BRUCE  P
GOOD OLD STUFF, THE  DOZOIS, GARDNER  LPA
GOOD OMENS  GAIMAN, NEIL/PRATCHETT, TERRY  H:2,SFWA,LP
GOOD SWITCH, BAD SWITCH [@SABRINA THE TEENAGE WITCH #3]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
GOOD THAT MEN DO, THE [STAR TREK ENTERPRISE]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
GOOD, THE BAD, AND THE UNDEAD, THE [RACHEL MORGAN 2]  HARRISON, KIM  P
GOOSEFLESH!  GHIDALIA, VIC  PA
GORDON R. DICKSON’S SF BEST  DICKSON, GORDON R.  P
GORGON, THE  LEE, TANITH  P
GORGON AND OTHER BEASTLY TALES, THE  LEE, TANITH  H
GORGON CHILD [AUBRY KNIGHT 2]  BARNES, STEVEN  C/P:2,SFWA-P,P
GORGON FESTIVAL, THE  BOYD, JOHN  C/P,P,H
GORGON’S HEAD, THE  SERRAILLER, IAN  P
GORGONZOLA, WON’T YOU PLEASE COME HOME?  AMES, CLYDE  P
GORMENGHAST [GORMENGHAST 2]  PEAKE, MERVYN  C/P:2,P,H
GOSH! WOW! (SENSE OF WONDER)  ACKERMAN, FORREST J.  C/PA,PA
GOSPEL PROBE, THE  CURTIS, MYRON  LP
GOSSAMER AXE  BAUDINO, GAEL  C/P:2,P

GOSSAMER EYE, THE MCLAUGHLIN, MARK/GRAVES, RAIN/

WILSON, DAVID NIALL

LP

GOSSAMER PLAIN, THE [@FORGOTTEN REALMS EMPYREAN ODYSSEY #1]  REID, THOMAS M.  P
GOSSAMYR  HAUF, MICHELE  LP
GOTH OPERA [DOCTOR WHO MISSING ADVENTURES]  CORNELL, PAUL  C/P:2,P

GOTREK & FELIX: THE FIRST OMNIBUS [@WARHAMMER GOTREK & FELIX 1,2,3] KING, WILLIAM

H

GOTREK & FELIX: THE SECOND OMNIBUS [@WARHAMMER GOTREK & FELIX 4,5,6] KING, WILLIAM

H

GRACE  STEWART, MICHAEL  H
GRADISIL  ROBERTS, ADAM  LP
GRADUATED ROBOT, THE  ELWOOD, ROGER  SR-HA
GRAIL [PENDRAGON 5]  LAWHEAD, STEPHEN R.  H,P
GRAIL AND THE RING, THE [CELYDONN GWENLLIANT 2]  EDGERTON, TERESA  C/P,SFWA-P,P
GRAIL OF HEARTS, THE  SHWARTZ, SUSAN  C/P,P,H
GRAIL PRINCE  MCKENZIE, NANCY  LP
GRAIL WAR, THE [PARSIVAL 2]  MONACO, RICHARD  C/P,P
GRAILBLAZERS  HOLT, TOM  P

GRAILS: QUESTS OF THE DAWN GILLIAM, RICHARD/GREENBERG, MARTIN H./

KRAMER, EDWARD E.

HA,LPA,SFWA

GRAILS: VISITATIONS OF THE NIGHT GILLIAM, RICHARD/GREENBERG, MARTIN H./

KRAMER, EDWARD E.

LPA

GRAIN KINGS, THE  ROBERTS, KEITH  P

GRAND CONSPIRACY [WARS OF LIGHT AND SHADOW 4/ ALLIANCE OF LIGHT 2] WURTS, JANNY

H

GRAND CRUSADE, THE [DRAGONCROWN WAR 3]  STACKPOLE, MICHAEL A.  LP
GRAND CRUSADE, THE (PROOF) [DRAGONCROWN WAR 3]  STACKPOLE, MICHAEL A.  SR-LP

The MITSFS Titledex  September 24, 2007  Page 152
TITLE  AUTHOR  CODE
GRAND DESIGN, THE [TYRANTS AND KINGS 2]  MARCO, JOHN  C/P,LP
GRAND ELLIPSE, THE  VOLSKY, PAULA  P
GRAND ILLUSION, THE  STATTEN, VARGO  SR-P
GRAND MASTERS’ CHOICE  NORTON, ANDRE  PA
GRAND TOUR, THE  NYE, JODY LYNN  P
GRAND TOUR, THE [CECILIA & KATHERINE 2]  WREDE, PATRICIA C./STEVERMER, CAROLINE H
GRAND WHEEL, THE  BAYLEY, BARRINGTON J.  P
GRANDE DERIVE, LA  RANDA, PETER  FP
GRANGE, THE  CHETWYND-HAYES, R.  P,SFWA-P
GRANITE SHIELD, THE [LIVING FLAME 3]  PATTON, FIONA  P
GRANT COMES EAST [GETTYSBURG 2]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
GRANTVILLE GAZETTE, VOL. 1 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  PA
GRANTVILLE GAZETTE, VOL. 2 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  HA
GRANTVILLE GAZETTE, VOL. 3 [SIXTEEN THIRTY-TWO]  FLINT, ERIC  HA
GRAPPLE, THE [@SETTLING ACCOUNTS 3]  TURTLEDOVE, HARRY  H
GRASP THE STARS  WINGERT, JENNIFER  P
GRASS  TEPPER, SHERI S.  C/P,P,H,SFWA
GRASS FOR HIS PILLOW [@OTORI #2]  HEARN, LIAN  H:2
GRASS IS ALWAYS GREENER, THE  MALCOLM-SMITH, GEORGE  H
GRATITUDE OF KINGS, THE  BRADLEY, MARION ZIMMER  H:2
GRAUSTARK  MCCUTCHEON, GEORGE B.  H:2
GRAVE, THE  GRANT, CHARLES L.  P,SFWA-P
GRAVE COVENANT [@BATTLETECH #34/TWILIGHT OF THE CLANS 2]  STACKPOLE, MICHAEL A.  C/P,P
GRAVE MATTER, A  DAVIES, L. P.  H
GRAVE MATTER [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
GRAVE PERIL [@DRESDEN FILES #3]  BUTCHER, JIM  C/P,P
GRAVE SIGHT  HARRIS, CHARLAINE  H
GRAVE SURPRISE  HARRIS, CHARLAINE  H:2
GRAVELIGHT  BRADLEY, MARION ZIMMER  H:2
GRAVEN IMAGES  FERMAN, EDWARD L./MALZBERG, BARRY N. SR-HA,HA
GRAVEN IMAGES  KILPATRICK, NANCY/ROCHE, THOMAS S.  LPA
GRAVEYARD GAME, THE [COMPANY 4]  BAKER, KAGE  LP
GRAVEYARD HEART, THE  ZELAZNY, ROGER  TD24
GRAVEYARD READER, THE  CONKLIN, GROFF  PA,C/PA
GRAVITY DREAMS  MODESITT, L. E. JR.  H
GRAVITY WELLS  GARDNER, JAMES ALAN  LP
GRAVITY’S ANGELS  SWANWICK, MICHAEL  LP
GRAVITY’S RAINBOW  PYNCHON, THOMAS  C/P:2,LP
GRAY ALIENS, THE  HOLLY, JOAN HUNTER  H
GRAY LENSMAN [LENSMEN 4]  SMITH, E. E.  H:2,P
GRAY MAGIC  NORTON, ANDRE  LP
GRAY MATTER  BRAVER, GARY  H
GRAY MATTERS  HJORTSBERG, WILLIAM  H:2,P
GRAY PRINCE, THE  VANCE, JACK  H:2,P
GRAY VICTORY  SKIMIN, ROBERT  H
GRAY WOLF, THE  MACDONALD, GEORGE  P
GRAYHEART [WOLFWALKER 6]  HARPER, TARA K.  C/P:2,P
GRAYMANTLE [IRON ANGEL 2]  MORRESSY, JOHN  C/P:2,P
GRAYSPACE BEAST, THE  EKLUND, GORDON  H:2,SFWA,P
GRAZER CONSPIRACY, THE [MEN IN BLACK]  SMITH, DEAN WESLEY  P
GREAT AND SECRET SHOW, THE [ART 1]  BARKER, CLIVE  H,P
GREAT AWAKENING, THE  MERRILL, ALBERT ADAMS  SR-H
GREAT BOOK OF AMBER, THE [AMBER 1,2,3,4,5,6,7,8,9,10]  ZELAZNY, ROGER  LP
GREAT BRAIN ROBBERY, THE  FISHER, JAMES P.  C/P:2,P
GREAT DIVORCE, THE  LEWIS, C. S.  P
GREAT EXPLOSION, THE  RUSSELL, ERIC FRANK  H:2,P
GREAT FETISH, THE  DE CAMP, L. SPRAGUE  C/P,SFWA,P,H
GREAT GHOST STORIES  STERN, PHILIP VAN DOREN  PA
GREAT GOLD STEAL, THE  WHITE, TED  C/P,P
GREAT HORROR MOVIES  FRIEDMAN, FAVIUS  R/REF-P,C/REF-P
GREAT HUNT, THE [WHEEL OF TIME 2]  JORDAN, ROBERT  H:2,P,SFWA:2

GREAT INTERNATIONAL PAPER AIRPLANE BOOK, THE MANDER, JERRY/DIPPEL, GEORGE/

GOSSAGE, HOWARD

[C/RAN-H]

GREAT KEINPLATZ EXPERIMENT, THE  DOYLE, ARTHUR CONAN  SR-H
GREAT KINGS’ WAR [PARATIME POLICE]  GREEN, ROLAND J./CARR, JOHN F.  C/P:2,P
GREAT LOS ANGELES BLIZZARD, THE  RACINA, THOM  P
GREAT MISCHIEF  PINCKNEY, JOSEPHINE  H,P
GREAT NEWS: THE NEW TESTAMENT, THE (NEW INTERNATIONAL)  GOD  PA
GREAT ONES, THE  DEEGAN, JON J.  H
GREAT PRINCE SHAN, THE  OPPENHEIM, E. PHILLIPS  H
GREAT PYRAMID MYSTERY  MORGAN, DE WOLFE  P
GREAT QUILL, THE  GARSON, PAUL  H:2,SFWA
GREAT SCIENCE-FICTION  LICATA, TONY  C/PA,PA
GREAT SCIENCE FICTION ABOUT DOCTORS  CONKLIN, GROFF/FABRICANT, NOAH D.  PA,C/PA

The MITSFS Titledex  September 24, 2007  Page 153
TITLE  AUTHOR  CODE
GREAT SCIENCE FICTION ADVENTURES  SHAW, LARRY T.  PA,C/PA
GREAT SCIENCE FICTION BY SCIENTISTS  CONKLIN, GROFF  PA,C/PA
GREAT SCIENCE FICTION FROM THE MOVIES  EDELSON, EDWARD  R/REF-P

GREAT SCIENCE FICTION OF THE 20TH CENTURY SILVERBERG, ROBERT/

GREENBERG, MARTIN H.

HA

GREAT SCIENCE FICTION SERIES, THE POHL, FREDERIK/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

HA

GREAT SCIENCE FICTION STORIES  SMITH, CORDELIA TITCOMB  PA,C/PA

GREAT SCIENCE FICTION STORIES BY THE WORLD’S GREATEST SCIENTISTS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

GREAT SF STORIES 1, THE (1939)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 2, THE (1940)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA,C/PA
GREAT SF STORIES 3, THE (1941)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 4, THE (1942)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA,C/PA
GREAT SF STORIES 5, THE (1943)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 6, THE (1944)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 7, THE (1945)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 8, THE (1946)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 9, THE (1947)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 10, THE (1948)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 11, THE (1949)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 12, THE (1950)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 13, THE (1951)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 14, THE (1952)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 15, THE (1953)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 16, THE (1954)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 17, THE (1955)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 18, THE (1956)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  C/PA,PA
GREAT SF STORIES 19, THE (1957)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 20, THE (1958)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 21, THE (1959)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA
GREAT SF STORIES 23, THE (1961)  ASIMOV, ISAAC/GREENBERG, MARTIN H.  PA

GREAT SF STORIES 1964, THE SILVERBERG, ROBERT/

GREENBERG, MARTIN H.

HA

GREAT SHORT NOVELS OF SCIENCE FICTION  SILVERBERG, ROBERT  C/PA,PA
GREAT SKY RIVER [GALACTIC CENTER 3]  BENFORD, GREGORY  C/P,SFWA,P,H
GREAT SPACE BATTLES  COWLEY, STEWART/HERRIDGE, CHARLES  A
GREAT STARSHIP RACE, THE [STAR TREK]  CAREY, DIANE  C/P:3,P
GREAT STONE OF SARDIS, THE  STOCKTON, FRANK R.  H,P
GREAT STORIES FROM ROD SERLING’S THE TWILIGHT ZONE MAGAZINE  KLEIN, T. E. D.  VLPA
GREAT STORIES OF SPACE TRAVEL  CONKLIN, GROFF  PA,C/PA
GREAT SUPERMAN BOOK, THE  FLEISHER, MICHAEL L.  R/REF-H
GREAT TALES OF ACTION AND ADVENTURE  BENNETT, GEORGE  C/PA,PA

GREAT TALES OF CLASSIC SCIENCE FICTION ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA

GREAT TALES OF FANTASY AND IMAGINATION  STERN, PHILIP VAN DOREN  C/PA,PA
GREAT TALES OF HORROR  POE, EDGAR ALLAN  P
GREAT TALES OF JEWISH FANTASY AND THE OCCULT  NEUGROSCHEL, JOACHIM  HA,LPA
GREAT TALES OF TERROR AND THE SUPERNATURAL  WISE, HERBERT A./FRASER, PHYLLIS  HA
GREAT TIME MACHINE HOAX, THE  LAUMER, KEITH  C/P:2,P
GREAT TWENTY-FOUR HOUR "THING", THE  OFFUTT, ANDREW J.  SR-L
GREAT UNTOLD STORIES OF FANTASY AND HORROR  NORTON, ALDEN H./MOSKOWITZ, SAM  C/PA,PA
GREAT WAR OF 189–, THE  COLOMB, P.  SR-H
GREAT WHEEL, THE [BROKEN CITADEL 3]  GREGORIAN, JOYCE BALLOU  SFWA-P
GREAT WHITE SPACE, THE  COPPER, BASIL  C/P,P
GREAT WHITE WYRM, THE [@DRAGONLANCE CHAMPIONS]  ARCHER, PETER  P
GREAT WORK OF TIME  CROWLEY, JOHN  C/P,P
GREAT WORLD CIRCUS  KOTZWINKLE, WILLIAM  H
GREATER INFINITY, A  MCCOLLUM, MICHAEL  C/P:2,P
GREATER TRUMPS, THE  WILLIAMS, CHARLES  SR-P,C/P,H
GREATEST ADVENTURE, THE  TAINE, JOHN  C/P:2,P
GREATEST SHOW IN THE GALAXY, THE [DOCTOR WHO]  WYATT, STEPHEN  P
GREATHEART [BIRTHRIGHT 2]  MCKEONE, DIXIE LEE  P
GREATHEART SILVER  FARMER, PHILIP JOSE  C/P:2,P
GREEN AND THE GRAY, THE  ZAHN, TIMOTHY  H
GREEN BRAIN, THE  HERBERT, FRANK  C/P:2,P
GREEN CHILD, THE  READ, HERBERT  H
GREEN DEATH, THE [DOC SAVAGE #65]  ROBESON, KENNETH  C/P:2,@P
GREEN EAGLE, THE [DOC SAVAGE #24]  ROBESON, KENNETH  C/P,@P
GREEN ECHO  BOWEN, GARY  B
GREEN EYES  GRANT, MAXWELL  P
GREEN EYES  SHEPARD, LUCIUS  C/P:2,P
GREEN EYES OF BAST, THE  ROHMER, SAX  H,P
GREEN FIRE  TAINE, JOHN  SR-H

The MITSFS Titledex  September 24, 2007  Page 154
TITLE  AUTHOR  CODE
GREEN FLAMES OF ARIES, THE [@HORRORSCOPE #1]  LORY, ROBERT  P
GREEN FLASH, THE  AIKEN, JOAN  H,P
GREEN FUTURES OF TYCHO, THE  SLEATOR, WILLIAM  C/P,P
GREEN GENE, THE  DICKINSON, PETER  C/P,P,H
GREEN GIRL, THE  WILLIAMSON, JACK  SR-P
GREEN GOD, THE  DVORKIN, DAVID  P

GREEN GODS, THE HENNEBERG, NATHALIE/ HENNEBERG, CHARLES

C/P,P

GREEN HILLS OF EARTH, THE  HEINLEIN, ROBERT A.  C/P,P,H
GREEN HILLS OF EARTH, THE (DIMENSION X RADIO SHOW) (LP RECORD)  HEINLEIN, ROBERT A.  SR-MM
GREEN HORNET IN THE INFERNAL LIGHT, THE  FRIEND, ED  P
GREEN KILLER, THE [@AVENGER #20]  ROBESON, KENNETH  C/P,P
GREEN KINGDOM, THE  MADDUX, RACHEL  P
GREEN KNIGHT, THE [THREE DAMOSELS 1]  CHAPMAN, VERA  C/P,P
GREEN MAGIC  VANCE, JACK  SR-H,SFWA-P
GREEN MAN, THE  AMIS, KINGSLEY  H,P
GREEN MAN, THE  SHERMAN, HAROLD M.  SR-P
GREEN MAN, THE  TREECE, HENRY  C/P,P
GREEN MAN FROM SPACE, THE  ZAREM, LEWIS  H
GREEN MAN OF GRAYPEC, THE  PRAGNELL, FESTUS  H
GREEN MANSIONS  HUDSON, W. H.  H,P
GREEN MARS [MARS 2]  ROBINSON, KIM STANLEY  H:2,LP
GREEN MARS (NOVELLA) [MARS]  ROBINSON, KIM STANLEY  TD1
GREEN MILE, THE [GREEN MILE 1,2,3,4,5,6]  KING, STEPHEN  P
GREEN MILLENNIUM, THE  LEIBER, FRITZ  C/P,H,SR-P,D30300
GREEN MOUSE, THE  CHAMBERS, ROBERT W.  H
GREEN ODYSSEY, THE  FARMER, PHILIP JOSE  C/P,P,H
GREEN PEARL, THE [LYONESSE 2]  VANCE, JACK  H:2,SFWA-P,LP
GREEN PEARL, THE (VIE) [@VANCE INTEGRAL EDITION #37/LYONESSE 2]  VANCE, JACK  SR-H
GREEN PHOENIX  SWANN, THOMAS BURNETT  C/P:2,P
GREEN PLANET, THE  HOLLY, JOAN HUNTER  C/P:2,P
GREEN QUEEN, THE  SAINT CLAIR, MARGARET  DD176
GREEN RAIN, THE  TABORI, PAUL  C/P:2,P
GREEN RAY, THE/BLOCKADE RUNNERS, THE  VERNE, JULES  H
GREEN RIDER [GREEN RIDER 1]  BRITAIN, KRISTEN  H:2,P
GREEN SALIVA BLUES, THE [MEN IN BLACK]  SMITH, DEAN WESLEY  P
GREEN SHADOWS, WHITE WHALE  BRADBURY, RAY  P
GREEN TOAD, THE  MASTERMAN, WALTER S.  H
GREENBRIAR QUEEN  GILLULY, SHEILA  P
GREENCOMBER  TATE, PETER  H,SFWA
GREENDAUGHTER  LOGSTON, ANNE  P,SFWA-P:2
GREENER THAN YOU THINK  MOORE, WARD  H:2
GREENER THAN YOU THINK (ABRIDGED)  MOORE, WARD  C/P:2,P
GREENFIRE  TITCHENER, LOUISE  P
GREENHOUSE [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P:2,P
GREENHOUSE SUMMER  SPINRAD, NORMAN  H
GREENING OF MARS, THE  LOVELOCK, JAMES/ALLABY, MICHAEL  [C/RAN-P]
GREENMAGIC [CALINDOR 1]  KILIAN, CRAWFORD  P
GREENMANTLE  DE LINT, CHARLES  C/P:2,P
GREENSTONE GRAIL, THE  HEMINGWAY, AMANDA  H
GREENTHIEVES  FOSTER, ALAN DEAN  H:2,P
GREENWAR  GOULD, STEVEN C./MIXON, LAURA J.  P
GREENWITCH [DARK IS RISING 3]  COOPER, SUSAN  H:3,LP
GREETINGS  BISSON, TERRY  H
GREETINGS, CARBON-BASED BIPEDS!  CLARKE, ARTHUR C.  R/REF-H
GREETINGS FROM EARTH [@BATTLESTAR GALACTICA 1978 #8]  LARSON, GLEN A./GOULART, RON  P
GREKS BRING GIFTS, THE  LEINSTER, MURRAY  C/P:2,P
GREMLINS [GREMLINS]  GIPE, GEORGE  C/P,P
GREMLINS 2 [GREMLINS]  BISCHOFF, DAVID F.  P
GREMLINS GO HOME  BOVA, BEN/DICKSON, GORDON R.  C/P:2,P
GREMLINS GO HOME  DICKSON, GORDON R./BOVA, BEN  H
GRENDEL  GARDNER, JOHN (AMERICAN, 1933-1982)  C/P:2,P
GRENDEL: PAST PRIME  RUCKA, GREG  LP
GREY BEGINNING, THE  MICHAELS, BARBARA  SFWA-P:2
GREY HORSE, THE  MACAVOY, R. A.  C/P:2,P
GREY HUNTER [WARHAMMER 40K SPACE WOLF 3]  KING, WILLIAM  P
GREY KING, THE [DARK IS RISING 4]  COOPER, SUSAN  C/P,LP,H:2
GREY KNIGHTS [@WARHAMMER 40K GREY KNIGHTS 1]  COUNTER, BEN  P
GREY MAIDEN  SMITH, ARTHUR D. HOWDEN  C/P:2,P
GREY MANE OF MORNING, THE  CHANT, JOY  H:2,LP,P
GREY ONES, THE  LYMINGTON, JOHN  C/P,P
GREY ROOM, THE  PHILLPOTTS, EDEN  H
GREY STAR THE WIZARD [@LONE WOLF #1]  PAGE, IAN/DEVER, JOE  P
GREYBEARD  ALDISS, BRIAN W.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 155
TITLE  AUTHOR  CODE
GREYFAX GRIMWALD [@CIRCLE OF LIGHT #1]  HANCOCK, NIEL  C/P:2,P
GREYHAVEN  BRADLEY, MARION ZIMMER  PA,C/PA
GREYLAND  VAN ARNAM, DAVE  C/P:2,P
GREYLORN  LAUMER, KEITH  C/P:2,P
GREYSTONE BAY [GREYSTONE BAY 1]  GRANT, CHARLES L.  SFWA-P
GREYWALKER  RICHARDSON, KIT  LP
GRIDLINKED [POLITY GRIDLINKED 1]  ASHER, NEAL  C/P,P,H
GRIDLOCK AND OTHER STORIES  MCCOLLUM, MICHAEL  LP
GRIDRUNNER [STAR*DRIVE]  REID, THOMAS M.  C/P,P
GRIEVING TREE, THE [@EBERRON DRAGON BELOW #2]  BASSINGTHWAITE, DON  P
GRIFFIN’S EGG  SWANWICK, MICHAEL  H:2,LP
GRIM TUESDAY [@KEYS TO THE KINGDOM #2]  NIX, GARTH  LP

GRIMALKIN’S TALES WHITELAW, STELLA/GARDINER, JUDY/

RONSON, MARK

H

GRIMBOLD’S OTHER WORLD  GRAY, NICHOLAS STUART  P
GRIMM REALITY [DOCTOR WHO EIGHTH DOCTOR]  BUCHER-JONES, SIMON/HALE, KELLY  P
GRIMMER THAN HELL  DRAKE, DAVID  H,P
GRIMM’S WORLD  VINGE, VERNOR  C/P,P,H
GRIMUS  RUSHDIE, SALMAN  SR-P,SFWA
GRIP OF FEAR  RITCHIE, RITA  P
GRIPPING HAND, THE [MOTE IN GOD’S EYE 2]  NIVEN, LARRY/POURNELLE, JERRY  H:2,P
GROA’S OTHER EYE [@TWILIGHT OF THE GODS #2]  SCHMIDT, DENNIS  C/P:2,P
GROSSEN VON AGOPP, DIE [@PERRY RHODAN #138]  FRANCIS, H. G.  FPR
GROTTO OF THE FORMIGANS, THE  DA CRUZ, DANIEL  C/P:2,P
GROUND ZERO [X-FILES]  ANDERSON, KEVIN J.  P
GROUND ZERO [@DEATHLANDS]  AXLER, JAMES  P
GROUND ZERO MAN  SHAW, BOB  C/P:3,P
GROUNDED [STAR TREK TNG]  BISCHOFF, DAVID F.  C/P:2,P
GROUNDED! [NICOLE SHEA 2]  CLAREMONT, CHRIS  C/P:2,SFWA-P:2,P
GROUNDTIES  FANCHER, JANE S.  P
GROUP FEAST  SAXTON, JOSEPHINE  H:2,SFWA
GROUPMIND  CONNER, MIKE  P
GROWING UP IN TIER 3000  GOTSCHALK, FELIX C.  C/P,P
GROWING UP WEIGHTLESS  FORD, JOHN M.  C/P:2,LP
GRUDGE BEARER [@WARHAMMER]  THORPE, GAV  P

GRUMBLES FROM THE GRAVE HEINLEIN, ROBERT A.

R/REF-H,C/REF-H:3,

C/REF-P

GRUNE KOMET, DER  FRANKE, HERBERT W.  FH
GRUNTS!  GENTLE, MARY  C/P:2,P
GRYB, THE  VAN VOGT, A. E.  C/P,P
GRYPHON  KILIAN, CRAWFORD  C/P:2,P
GRYPHON IN GLORY [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE  H:2,P
GRYPHON KING, THE [OTHERWORLD]  DEITZ, TOM  C/P:2,P
GRYPHON’S EYRIE [WITCH WORLD HIGH HALLACK GRYPHON]  NORTON, ANDRE/CRISPIN, A. C.  H:2,SFWA-P:2,P
GUARD AGAINST DISHONOR [HAWK & FISHER 5]  GREEN, SIMON R.  C/P:2,P
GUARDIAN  HALDEMAN, JOE  H
GUARDIAN, THE  KONVITZ, JEFFREY  P
GUARDIAN  MONTELEONE, THOMAS F.  C/P,SFWA,P,H
GUARDIAN  SAUL, JOHN  H
GUARDIAN, THE  WELLS, ANGUS  P
GUARDIAN ANGEL  BEDWELL-GRIME, STEPHANIE  LP
GUARDIAN OF HONOR [LLADRANA 1]  OWENS, ROBIN D.  LP
GUARDIAN OF THE BALANCE [@MERLIN’S DESCENDANTS #1]  RADFORD, IRENE  H,P
GUARDIAN OF THE PROMISE [@MERLIN’S DESCENDANTS #4]  RADFORD, IRENE  H,P
GUARDIAN OF THE TRUST [@MERLIN’S DESCENDANTS #2]  RADFORD, IRENE  P
GUARDIAN OF THE VISION [@MERLIN’S DESCENDANTS #3]  RADFORD, IRENE  P

GUARDIAN:SAVIORS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 3] MCGOUGH, SCOTT

P

GUARDIANS, THE  ABBEY, LYNN  C/P,P
GUARDIANS, THE [@PERRY RHODAN #58]  ACE  C/PA:2,PA
GUARDIANS, THE [GUARDIANS 1]  AUSTIN, RICHARD  P
GUARDIANS, THE  CHRISTOPHER, JOHN  C/P,P,H
GUARDIAN’S KEY  LOGSTON, ANNE  P
GUARDIANS OF ALEXANDER [GOLDBANE 1]  WILSON, JOHN  LP
GUARDIANS OF THE CORAL THRONE [@BLADE #20]  LORD, JEFFREY  C/P,P
GUARDIANS OF THE FOREST [@WARHAMMER]  MCNEILL, GRAHAM  P
GUARDIANS OF THE GATE  TRIMBLE, LOUIS/TRIMBLE, JACQUELYN  C/P:2,P
GUARDIANS OF THE KEEP [BRIDGE OF D’ARNATH 2]  BERG, CAROL  P
GUARDIANS OF THE LOST [@SOVEREIGN STONE #2]  WEIS, MARGARET/HICKMAN, TRACY  H,P
GUARDIANS OF THE SINGREALE  MILLER, CALVIN  LP
GUARDIANS OF THE WEST [MALLOREON 1]  EDDINGS, DAVID  C/P,H,P,SFWA
GUARDIANS OF TIME  ANDERSON, POUL  C/P:2,P
GUARDS! GUARDS! [DISCWORLD 8/DISCWORLD WATCHMEN 1]  PRATCHETT, TERRY  H:2,P
GUARDSMAN, THE  BEESE, P. J./HAMILTON, TODD CAMERON  C/P,P

TITLE  AUTHOR  CODE
GUARDSMAN OF GOR [GOR 16]  NORMAN, JOHN  P
GUARDSMEN OF TOMORROW  GREENBERG, MARTIN H./SEGRIFF, LARRY  C/PA,PA
GUCKY UND DIE MORDWESPEN [@PERRY RHODAN #28]  DARLTON, CLARK  FPR
GUERNICA NIGHT  MALZBERG, BARRY N.  H:2,P
GUERRA FINALE  O’DONNELL, K. M.  FP
GUIDE TO BARSOOM, A  ROY, JOHN FLINT  R/REF-P,C/REF-P:2
GUIDE TO LARRY NIVEN’S RINGWORLD, THE  STEIN, KEVIN  R/REF-H
GUIDE TO MIDDLE-EARTH, A  FOSTER, ROBERT  R/REF-H,C/REF-P

GUIDE TO SCIENCE FICTION:

EXPLORING POSSIBILITIES AND ALTERNATIVES PAINE, DORIS M./MARTINEZ, DIANA

R/REF-P

GUIDE TO SUPERNATURAL FICTION, THE  BLEILER, EVERETT F.  R/IND
GUIDE TO THE STAR WARS UNIVERSE, A  VELASCO, RAYMOND L.  R/REF-P
GUIDED TOUR  DICKSON, GORDON R.  P,SFWA-P
GUIGNOIR AND OTHER FURIES  HATCH, GEORGE  LPA
GUILDPACT [MAGIC: THE GATHERING RAVNICA 2]  HERNDON, CORY J.  P
GUILT-EDGED IVORY [IVORY 3]  EGAN, DORIS  C/P:2,P
GUILTY PLEASURES [ANITA BLAKE 1]  HAMILTON, LAURELL K.  C/P:2,P
GUINEVERE [GUINEVERE 1]  NEWMAN, SHARAN  P
GUINEVERE EVERMORE [GUINEVERE 3]  NEWMAN, SHARAN  H
GUINNESS BOOK OF WORLD RECORDS, THE  MCWHIRTER, NORRIS/MCWHIRTER, ROSS  [C/RAN-P]
GUISES OF THE MIND [STAR TREK TNG]  NEASON, REBECCA  C/P:2,P
GUITAR  MORGAN, DAN  [C/RAN-P]
GULF RUN [@DEMONTECH 3]  SHERMAN, DAVID  P
GULLIBLE’S TRAVELS [ARCHONATE 1,2]  HUGHES, MATTHEW  H
GULLIVER OF MARS  ARNOLD, EDWIN LESTER  C/P:2,P
GULLIVER’S FUGITIVES [STAR TREK TNG]  SHAREE, KEITH  C/P:2,P
GULLIVER’S TRAVELS  SWIFT, JONATHAN  C/P:2,SR-L,P
GULLIVERZONE [WEB,THE]  BAXTER, STEPHEN  P
GULLY DWARVES, THE [@DRAGONLANCE LOST HISTORIES #5]  PARKINSON, DAN  P
GUMSHOE, THE WITCH & THE VIRTUAL CORPSE, THE  HARTMAN, KEITH  H,LP
GUN FOR DINOSAUR, A  DE CAMP, L. SPRAGUE  H:2,P
GUN, WITH OCCASIONAL MUSIC  LETHEM, JONATHAN  H:2,LP
GUNFIGHT AT LARAMIE  HOFFMAN, LEE  P
GUNFIGHTERS, THE [DOCTOR WHO]  COTTON, DONALD  SR-P
GUNNER CADE  JUDD, CYRIL  H,DD227
GUNNER CADE  KORNBLUTH, C. M./MERRIL, JUDITH  C/P:2,P
GUNPOWDER EMPIRE [CROSSTIME TRAFFIC 1]  TURTLEDOVE, HARRY  H
GUNPOWDER GOD [PARATIME POLICE]  PIPER, H. BEAM  P
GUNS, DRUGS, AND MONSTERS [CAL MCDONALD 2]  NILES, STEVE  P
GUNS OF AVALON, THE [AMBER 2]  ZELAZNY, ROGER  C/P:3,@P
GUNS OF DARKNESS [@THERE WILL BE WAR #6]  POURNELLE, JERRY  PA,C/PA,SFWA-P
GUNS OF EVERBLACK, THE [PERRY RHODAN 126]  SCHEER, K. H.  SR-LP
GUNS OF TANITH, THE [@WARHAMMER 40K GAUNT’S GHOSTS #5]  ABNETT, DAN  P
GUNS OF TERRA 10, THE  PENDLETON, DON  P
GUNS OF THE SOUTH, THE  TURTLEDOVE, HARRY  C/P,P,H
GUNSLINGER, THE [DARK TOWER 1]  KING, STEPHEN  C/P:2,SR-L,LP
GUNSLINGER, THE (REVISED) [DARK TOWER 1]  KING, STEPHEN  H
GUST FRONT [POSLEEN 2]  RINGO, JOHN  P
GUTBUCKET QUEST, THE  ANTHONY, PIERS/LEMING, RON  H
GUTHRIE METHOD, THE  GALLUN, RAYMOND Z.  B
GUTS  PREISS, BYRON  C/P:2,P
GUYS FROM SPACE  PINKWATER, DANIEL MANUS  SFWA-P
GWEN, IN GREEN  ZACHARY, HUGH  P
GYPSIES  WILSON, ROBERT CHARLES  H,SFWA-P,P
GYPSY  BRUST, STEVEN/LINDHOLM, MEGAN  H:2,P
GYPSY EARTH  HARPER, GEORGE W.  H,SFWA
GYPSY ENCHANTMENT, THE [@CHARMED]  JABLONSKI, CARLA  P
GYRATOR CONTROL  LANG, KING  SR-P

The MITSFS Titledex  September 24, 2007  Page 157
TITLE  AUTHOR  CODE
H-BOMB OVER AMERICA  SUTTON, JEFF  P
H. G. WELLS  WEST, ANTHONY  R/REF-H
H. G. WELLS: AUTHOR IN AGONY  BORRELLO, ALFRED  R/REF-H
H. G. WELLS: CRITIC OF PROGRESS  WILLIAMSON, JACK  R/REF-H
H. G. WELLS SCRAPBOOK, THE  WELLS, H. G.  R/REF-H,C/REF-H:2
H. RIDER HAGGARD: A VOICE FROM THE INFINITE  ELLIS, PETER BERRESFORD  R/REF-H
H. RIDER HAGGARD’S SHE  WARD, DON  SR-P
HAB THEORY, THE  ECKERT, ALLAN W.  C/P,P,H
HABITATS  SHWARTZ, SUSAN  C/PA,PA
HABU  JOHNSON, JAMES B.  P
HACKER, THE  DAY, CHET  SFWA-P
HACKER AND THE ANTS, THE  RUCKER, RUDY  C/P,P,H
HACKER CRACKDOWN, THE  STERLING, BRUCE  R/REF-P
HACKERS  BISCHOFF, DAVID F.  P
HACKERS  DANN, JACK/DOZOIS, GARDNER  PA
HADES’ DAUGHTER [TROY GAME #1]  DOUGLASS, SARA  H:2,@P
HADON OF ANCIENT OPAR  FARMER, PHILIP JOSE  C/P:2,P
HAG’S CONTRACT [BIRTHRIGHT 4]  BETANCOURT, JOHN GREGORY  P
HAGS, SIRENS, AND OTHER BAD GIRLS OF FANTASY  LITTLE, DENISE  PA
HAIL HIBBLER  GOULART, RON  C/P:2,P
HAIR OF THE DOG  DAVIS, BRETT  C/P,P
HAIRY HORROR TRICK, THE  CORBETT, SCOTT  P
HAL CLEMENT  HASSLER, DONALD M.  R/REF-P
HAL SPACEJOCK [@HAL SPACEJOCK 1]  HAYNES, SIMON  P
HAL SPACEJOCK - SECOND COURSE [@HAL SPACEJOCK 2]  HAYNES, SIMON  P
HALBERD, DREAM WARRIOR [DREAMQUEST 1]  SAINT ALCORN, LLOYD  P
HALCYON DRIFT, THE [GRAINGER 1]  STABLEFORD, BRIAN M.  C/P:2,P
HALF A LIFE  BULYCHEV, KIR  H,LP
HALF-HAUNTED SALOON, THE  SHATTUCK, RICHARD  H
HALF IN SHADOW  COUNSELMAN, MARY ELIZABETH  SR-H
HALF LIFE  CLEMENT, HAL  H:2,P
HALF-LIGHT  VITOLA, DENISE  P
HALF MAGIC  EAGER, EDWARD  H,LP
HALF PAST HUMAN  BASS, T. J.  C/P:2,P
HALF THE DAY IS NIGHT  MCHUGH, MAUREEN F.  C/P:2,P,H
HALFLIFE [DOCTOR WHO EIGHTH DOCTOR]  MICHALOWSKI, MARK  P
HALFLING AND OTHER STORIES, THE  BRACKETT, LEIGH  C/P:2,P
HALFLING’S GEM, THE [@FORGOTTEN REALMS ICEWIND DALE #3]  SALVATORE, R. A.  P
HALFLINGS, HOBBITS, WARROWS & WEEFOLK  SEARLES, BAIRD/THOMSEN, BRIAN M.  HA,LPA
HALFWAY ACROSS THE GALAXY AND TURN LEFT  KLEIN, ROBERT  P
HALFWAY HUMAN  GILMA, CAROLYN IVES  P
HALL OF THE MOUNTAIN KING, THE [@AVARYAN RISING #1]  TARR, JUDITH  H,SFWA,SFWA-P,P
HALL OF WHISPERS [HEIRS TO GNARLSMYRE]  JEFFERIES, MIKE  P
HALLOWED HUNT, THE [CHALION 3]  BUJOLD, LOIS MCMASTER  H:2
HALLOWED ISLE, THE [HALLOWED ISLE 1,2]  PAXSON, DIANA L.  P
HALLOWEEN HAVOC [@SABRINA THE TEENAGE WITCH #4]  GALLAGHER, DIANA G.  P
HALLOWEEN HORRORS  RYAN, ALAN  PA,SFWA
HALLOWEEN III: SEASON OF THE WITCH  MARTIN, JACK  P
HALLOWEEN RAIN [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  P
HALLOWEEN TREE, THE  BRADBURY, RAY  C/P,P,H
HALLOWING, THE  YARIV, FRAN POKRAS  P

HALLS OF STORMWEATHER, THE [@FORGOTTEN REALMS SEMBIA GATEWAY TO THE REALMS #1] ATHANS, PHILIP

PA

HALLS OF STORMWEATHER, THE [@FORGOTTEN REALMS SEMBIA] BYERS, RICHARD LEE/EMERY, CLAYTON/ GREENWOOD, ED/GROSS, DAVE/ KEMP, PAUL S./SMEDMAN, LISA/ WHITNEY-ROBINSON, VORONICA

P

HALO  COOK, PAUL  P
HALO  MADDOX, TOM  H:2,P
HALO: FIRST STRIKE [@HALO]  NYLUND, ERIC S.  P
HALO IN BRASS  EVANS, JOHN  P
HALO: THE FALL OF REACH [@HALO]  NYLUND, ERIC S.  P
HALO: THE FLOOD [@HALO]  DIETZ, WILLIAM C.  P
HAMELIN PLAGUE, THE  CHANDLER, A. BERTRAM  C/P:2,P
HAMMER, THE [@GENERAL #2]  STIRLING, S. M./DRAKE, DAVID  C/P:2,SFWA-P,P
HAMMER AND ANVIL [TIME OF TROUBLES 2]  TURTLEDOVE, HARRY  P
HAMMER AND AXE [@DRAGONLANCE DWARVEN NATIONS #2]  PARKINSON, DAN  P
HAMMER AND THE CROSS, THE [HAMMER AND THE CROSS 1]  HARRISON, HARRY  C/P:2,P
HAMMER AND THE HORN, THE  FRIEDMAN, MICHAEL JAN  C/P,P
HAMMER OF DARKNESS, THE  MODESITT, L. E. JR.  C/P,P
HAMMER OF GOD, THE  CLARKE, ARTHUR C.  C/P,P,H
HAMMER OF MARS [BUCK ROGERS MARTIAN WARS 2]  MURDOCK, M. S.  C/P,P
HAMMER OF THE SUN, THE [@WINTER OF THE WORLD #3]  ROHAN, MICHAEL SCOTT  H,P
HAMMERED [JENNY CASEY 1]  BEAR, ELIZABETH  C/P,P

The MITSFS Titledex  September 24, 2007  Page 158
TITLE  AUTHOR  CODE
HAMMERFALL [GENE WARS 1]  CHERRYH, C. J.  H:2,P
HAMMERJACK  GILLER, MARC D.  LP
HAMMERJACK  GILLESPIE, MARC D.  LP
HAMMERS OF ULRIC [WARHAMMER]  ABNETT, DAN/VINCENT, NIK/WALLIS, JAMES P
HAMMER’S SLAMMERS [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
HAMMER’S SLAMMERS (INCLUDES THE TANK LORDS)  DRAKE, DAVID  C/P:2,P
HAMPDENSHIRE WONDER, THE  BERESFORD, J. D.  H
HAN SOLO ADVENTURES, THE [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P
HAN SOLO AND THE LOST LEGACY [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HAN SOLO AT STARS’ END [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HAN SOLO’S REVENGE [STAR WARS HAN SOLO]  DALEY, BRIAN  C/P,P,H
HAND OF CHAOS, THE [@DEATH GATE #5]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
HAND OF CHAOS, THE (PROOF) [DEATH GATE 5]  WEIS, MARGARET/HICKMAN, TRACY  SR-LP
HAND OF DINOTOPIA, THE [DINOTOPIA]  FOSTER, ALAN DEAN  H
HAND OF DOOM  FANTHORPE, R. L.  H
HAND OF DRACULA, THE [@DRACULA #2]  LORY, ROBERT  C/P,P
HAND OF FIRE [@FORGOTTEN REALMS SHANDRIL #3]  GREENWOOD, ED  P
HAND OF FU-MANCHU, THE  ROHMER, SAX  H,P
HAND OF GANZ, THE  HAIBLUM, ISIDORE  P
HAND OF KANE, THE [SOLOMON KANE 2]  HOWARD, ROBERT E.  P
HAND OF OBERON, THE [AMBER 4]  ZELAZNY, ROGER  C/P,@P,H,SFWA
HAND OF PROPHECY  PARK, SEVERNA  C/P,P
HAND OF THE KING’S EVIL [@OUTREMER #5]  BRENCHLEY, CHAZ  P
HAND OF ZEI, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P,P,H,DF249
HANDBOOK FOR SPACE PIONEERS  WOLFE, L. STEPHEN/WYSACK, ROY L.  R/REF-H

HANDBOOK OF CHEMISTRY AND PHYSICS (11TH EDITION)

HODGMAN, CHARLES D./ LANGE, NORBERT A.

[C/RAN-P]

HANDFUL OF SILVER, A  SHORTER, PHILIP  P
HANDFUL OF STARS, A  STABENOW, DANA  P
HANDFUL OF TIME, A  BROWN, ROSEL GEORGE  C/P:2,P
HANDLING IT [PEMBROKE HALL 1,2]  FAUST, JOE CLIFFORD  H
HANDMAID’S TALE, THE  ATWOOD, MARGARET  C/P,P
HANDS OF GLORY  KANGILASKI, JAAN  C/P,P
HANDS OF LYR, THE  NORTON, ANDRE  C/P,P,H
HANGFIRE [@STARFIST 6]  SHERMAN, DAVID/CRAGG, DAN  C/P,P
HANGING MOUNTAINS [@CATACLYSM #3]  WILLIAMS, SEAN  H
HANGING STONES, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
HANGMAN, THE [@DARKMAN #1]  BOYLL, RANDALL  P
HANGMAN’S DOZEN, A  HITCHCOCK, ALFRED  PA
HANNES BOK ILLUSTRATION INDEX  BROOKS, NED/MARTIN, DON  R/IND,R/IND-F
HAPLOIDS, THE  SOHL, JERRY  SR-P,H
HAPPY BIRTHDAY, WANDA JUNE  VONNEGUT, KURT JR.  LP
HAPPY ENDINGS [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  P
HAPPY KILLERS, THE [@AVENGER #21]  ROBESON, KENNETH  C/P:2,P
HAPPY PLANET, THE  CLARKE, JOAN  H
HAPPY POLICEMAN  ANTHONY, PATRICIA  C/P,P
HARBINGER [STAR TREK VANGUARD]  MACK, DAVID  P
HARD BOILED  MILLER, FRANK/DARROW, GEOF  CX
HARD-BOILED WONDERLAND AND THE END OF THE WORLD  MURAKAMI, HARUKI  H
HARD CONTACT [@STAR WARS REPUBLIC COMMANDO 1]  TRAVISS, KAREN  C/P,P
HARD CRASH [UNREAL 1]  HUGHES, RYAN  P
HARD DRIVE  POGUE, DAVID  C/P,P
HARD LANDING  BUDRYS, ALGIS  P
HARD MERCHANDISE [@STAR WARS BOUNTY HUNTER WARS 3]  JETER, K. W.  C/P,P
HARD RAIN, A [STAR TREK TNG]  SMITH, DEAN WESLEY  C/P,P
HARD SELL  ANTHONY, PIERS  C/P:2,P
HARD SF RENAISSANCE, THE  HARTWELL, DAVID G./CRAMER, KATHRYN  HA
HARD TO BE A GOD  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P,SFWA,P,H
HARD WAY UP, THE [GRIMES]  CHANDLER, A. BERTRAM  D31755
HARD WORDS AND OTHER POEMS  LE GUIN, URSULA K.  SFWA
HARDBOILED CTHULHU  AMBUEHL, JAMES  LPA
HARDFOUGHT  BEAR, GREG  TD2,SFWA-TD2
HARDWIRED  WILLIAMS, WALTER JON  C/P,P,H
HAREM OF AMAN AKBAR, THE  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
HARILEK  GANPAT  H
HARKFAST: THE MAKING OF THE KING  RAE, HUGH C.  P
HARLAN ELLISON: A BIBLIOGRAPHICAL CHECKLIST  SWIGART, LESLIE KAY  R/IND
HARLAN ELLISON: UNREPENTANT HARLEQUIN  SLUSSER, GEORGE EDGAR  R/REF-P
HARLEQUIN, THE [ANITA BLAKE 15]  HAMILTON, LAURELL K.  H
HARLEQUIN [WARHAMMER INQUISITION WAR 2]  WATSON, IAN  P
HARLEQUIN’S DANCE [@OROKON #1]  ARDEN, TOM  P
HARLOT’S RUSE  FRIESNER, ESTHER M.  P
HARMONIES OF THE NET [GROUNDTIES 3]  FANCHER, JANE S.  C/P,P
HARMONY  KELLOGG, MARJORIE BRADLEY  P,SFWA-P:2

The MITSFS Titledex  September 24, 2007  Page 159
TITLE  AUTHOR  CODE
HARM’S WAY  GREENLAND, COLIN  C/P:2,SFWA-P,P
HARP AND THE BLADE, THE  MYERS, JOHN MYERS  P
HARP OF IMACH THYSSEL, THE [LYRA 5]  WREDE, PATRICIA C.  C/P:2,P
HARP OF THE GREY ROSE, THE  DE LINT, CHARLES  C/P:2,LP
HARP OF WINDS [AURIAN 2]  FUREY, MAGGIE  C/P:2,P
HARPER HALL OF PERN, THE [PERN HARPER HALL 1,2,3]  MCCAFFREY, ANNE  H:2
HARPIST IN THE WIND  MCKILLIP, PATRICIA A.  C/P,P,H
HARPY HIGH [GNOME MAN’S LAND 2]  FRIESNER, ESTHER M.  C/P,P
HARPY THYME [@XANTH #17]  ANTHONY, PIERS  C/P,P,H
HARPY’S FLIGHT [KI AND VANDIEN 1]  LINDHOLM, MEGAN  C/P,P
HARRAD EXPERIMENT, THE  RIMMER, ROBERT H.  C/P,P
HARRIET, THE HAUNTED [@PHENWICK #10]  KIMBROUGH, KATHERYN  P
HARRIGAN’S FILE  DERLETH, AUGUST  SR-H
HARROWING, THE  SKINNER, AINSLIE  P
HARROWING OF GWYNEDD, THE [DERYNI HEIRS 1]  KURTZ, KATHERINE  C/P,P,H
HARROWING THE DRAGON  MCKILLIP, PATRICIA A.  H
HARRY AND THE HENDERSONS  THOMPSON, JOYCE  P
HARRY KAPLAN’S ADVENTURES UNDERGROUND  STERN, STEVE  H

HARRY POTTER AND THE CHAMBER OF SECRETS (AMERICAN)

[HARRY POTTER 2] ROWLING, J. K.

H:2

HARRY POTTER AND THE DEATHLY HALLOWS [HARRY POTTER 7]  ROWLING, J. K.  H

HARRY POTTER AND THE GOBLET OF FIRE (AMERICAN)

[HARRY POTTER 4] ROWLING, J. K.

H:2,LP

HARRY POTTER AND THE GOBLET OF FIRE (BRITISH) [HARRY POTTER 4]  ROWLING, J. K.  H
HARRY POTTER AND THE HALF-BLOOD PRINCE [HARRY POTTER 6]  ROWLING, J. K.  H:2

HARRY POTTER AND THE ORDER OF THE PHOENIX (AMERICAN)

[HARRY POTTER 5] ROWLING, J. K.

H

HARRY POTTER AND THE PHILOSOPHER’S STONE (BRITISH)

[HARRY POTTER 1] ROWLING, J. K.

LP

HARRY POTTER AND THE PRISONER OF AZKABAN (AMERICAN)

[HARRY POTTER 3] ROWLING, J. K.

H

HARRY POTTER AND THE SORCERER’S STONE (AMERICAN)

[HARRY POTTER 1] ROWLING, J. K.

C/P,LP,H

HARRY POTTER UND DER STEIN DER WEISEN [HARRY POTTER 1]  ROWLING, J. K.  FH
HARRY SULLIVAN’S WAR [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
HARRY, THE FAT BEAR SPY  WILSON, GAHAN  LP
HARSH CRY OF THE HERON, THE [@OTORI #4]  HEARN, LIAN  H
HARSHINI [HYTHRUN 3]  FALLON, JENNIFER  H
HART’S HOPE  CARD, ORSON SCOTT  C/P:2,P
HARVEST, THE [BUFFY]  CUSICK, RICHIE TANKERSLEY  P
HARVEST, THE  WILSON, ROBERT CHARLES  LP
HARVEST BRIDE, THE  RICHARDS, TONY  SFWA-P
HARVEST MOON [@SABRINA THE TEENAGE WITCH #15]  ODOM, MEL  P
HARVEST OF BLOOD  COURTNEY, VINCENT  P
HARVEST OF CHANGELINGS  ROCHELLE, WARREN  H
HARVEST OF HOODWINKS, A  LORY, ROBERT  D52180
HARVEST OF STARS [HARVEST OF STARS 2]  ANDERSON, POUL  H:2,P
HARVEST, THE (PROOF)  WILSON, ROBERT CHARLES  SR-LP
HARVEST THE FIRE [HARVEST OF STARS 1]  ANDERSON, POUL  H:2,P
HASAN  ANTHONY, PIERS  C/P:2,SFWA-P,P
HASTUR CYCLE, THE (2ND EDITION)  PRICE, ROBERT M.  LPA
HAT FULL OF SKY, A [DISCWORLD 31/DISCWORLD TIFFANY ACHING 2]  PRATCHETT, TERRY  H
HATE GENIUS, THE [DOC SAVAGE #94]  ROBESON, KENNETH  @P
HATE MASTER, THE [@AVENGER #16]  ROBESON, KENNETH  C/P:2,P
HATRACK RIVER [ALVIN MAKER 1,2,3]  CARD, ORSON SCOTT  H
HAUNTED [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  P
HAUNTED [@ANGEL]  MARIOTTE, JEFF  P
HAUNTED AIR, THE [REPAIRMAN JACK]  WILSON, F. PAUL  P
HAUNTED AIRWAYS  BURTIS, THOMSON  H
HAUNTED BY DESIRE [@CHARMED]  DOKEY, CAMERON  P
HAUNTED DOLLHOUSE, THE [TWILIGHT #22]  BLAKE, SUSAN  P
HAUNTED EARTH, THE  KOONTZ, DEAN R.  C/P,P
HAUNTED HOLIDAYS  GREENBERG, MARTIN H./DAVIS, RUSSELL  PA
HAUNTED JESTER, THE  CORLEY, DONALD  H
HAUNTED MIND, THE  FODOR, NANDOR  [C/RAN-P]
HAUNTED OCEAN [DOC SAVAGE #51]  ROBESON, KENNETH  C/P,@P
HAUNTED STARS, THE  HAMILTON, EDMOND  H:2,P
HAUNTED STORIES  FURMAN, A. L.  PA
HAUNTED WIZARD, THE [WIZARD IN RHYME 6]  STASHEFF, CHRISTOPHER  P
HAUNTED WOMAN, THE  LINDSAY, DAVID  LP
HAUNTING OF ALAN MAIS, THE [GUARDIANS (SAXON) 3]  SAXON, PETER  C/P,P
HAUNTING OF HILL HOUSE, THE  JACKSON, SHIRLEY  C/P,P
HAUNTING OF TOBY JUGG, THE  WHEATLEY, DENNIS  C/P,P
HAUNTING THE DEAD [@WORLD OF DARKNESS ORPHEUS]  BOULLE, PHILIPPE  PA

The MITSFS Titledex  September 24, 2007  Page 160
TITLE  AUTHOR  CODE
HAUNTINGS [@SMALLVILLE]  HOLDER, NANCY  P
HAUNTINGS AND HORRORS  NORTON, ALDEN H.  PA
HAUNTINGS: TALES OF THE SUPERNATURAL  MAZZEO, HENRY  HA
HAUNTS AND ENIGMAS  HAWTHORNE, JULIAN  HA
HAUSER’S MEMORY  SIODMAK, CURT  H:2,P
HAVE ROBOT, WILL TRAVEL [@ASIMOV’S ROBOT MYSTERY]  IRVINE, ALEXANDER C.  H
HAVE SPACE SUIT — WILL TRAVEL  HEINLEIN, ROBERT A.  C/P,P,H:2
HAVE STARSHIP, WILL TRAVEL  WHEELER, RAMONA LOUISE  LP

HAVE TECH, WILL TRAVEL [@STAR TREK S.C.E. #1] DECANDIDO, KEITH R. A./DILMORE, KEVIN/ GOLDEN, CHRISTIE/SMITH, DEAN WESLEY/ WARD, DAYTON

P

HAVE TRENCHCOAT – WILL TRAVEL  SMITH, E. E.  SR-H
HAVE YOU SEEN THESE?  ASIMOV, ISAAC  SR-H
HAVEN [@PSI-MAN #6]  DAVID, PETER  P
HAVEN, THE [EMPIRE PRINCESS 0]  DIAMOND, GRAHAM  P
HAVEN OF DARKNESS [DUMAREST 16]  TUBB, E. C.  C/P:2,P
HAVOC IN ISLANDIA [ISLANDIA]  SAXTON, MARK  H
HAWAIIAN UFO ALIENS  GILDEN, MEL  C/P,P
HAWK & FISHER [HAWK & FISHER 1]  GREEN, SIMON R.  P
HAWK AMONG THE SPARROWS  MCLAUGHLIN, DEAN  H,SFWA
HAWK ETERNAL, THE [HAWK QUEEN 2]  GEMMELL, DAVID  P
HAWK IN SILVER, A  GENTLE, MARY  C/P,P,H
HAWK OF MAY [GWALCHMAI 1]  BRADSHAW, GILLIAN  H:2,P
HAWK OF MAY (PROOF) [GWALCHMAI 1]  BRADSHAW, GILLIAN  SR-LP
HAWK OF THE WILDERNESS [KIOGA 1]  CHESTER, WILLIAM L.  P
HAWKLINE MONSTER, THE  BRAUTIGAN, RICHARD  P
HAWKMISTRESS! [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
HAWKMOON [RUNESTAFF 1,2,3,4]  MOORCOCK, MICHAEL  H
HAWK’S FLIGHT  CHASE, CAROL  C/P,P
HAWK’S GRAY FEATHER, THE [KELTIAD ARTHUR 1]  KENNEALY-MORRISON, PATRICIA  C/P:2,P
HAWKS OF ARCTURUS, THE  SNYDER, CECIL  C/P,P
HAWKS OF FELLHEATH, THE [ASH STAFF 2]  FISHER, PAUL R.  C/P,P
HAWKSBILL STATION  SILVERBERG, ROBERT  SFWA-TD26
HAWKSBILL STATION (EXPANDED)  SILVERBERG, ROBERT  H:2,P
HAWKSHAW  GOULART, RON  H,SFWA,P
HAWKWOD’S VOYAGE [@MONARCHIES OF GOD #1]  KEARNEY, PAUL  P
HAYDN OF MARS [CATS OF MARS 1]  SARRANTONIO, AL  P
HAZARD  BROWNE, GERALD A.  P
HAZARD’S PRICE [CHRONICLES UNBINDING 1]  STONE, ROBERT S.  P
HAZEZON [@MAGIC: THE GATHERING LEGENDS 3]  EMERY, CLAYTON  P
HE COULD STOP THE WORLD [DOC SAVAGE #54]  ROBESON, KENNETH  C/P,@P
HE OWNED THE WORLD  MAINE, CHARLES ERIC  C/P,P,H
HE, SHE AND IT  PIERCY, MARGE  H,P
HE WALKED AMONG US  SPINRAD, NORMAN  LP
HE WHO SHAPES  ZELAZNY, ROGER  SFWA-TD12,TD12
HEAD GAMES [DOCTOR WHO NEW ADVENTURES]  LYONS, STEVE  C/P:2,P
HEADCRASH  BETHKE, BRUCE  H,P
HEADHUNTERS [@SHADOWRUN 27]  ODOM, MEL  P
HEADS [QUEEN OF ANGELS 2]  BEAR, GREG  C/P,P,H
HEADS (PROOF)  OSBORN, DAVID  SR-P
HEADS OF CERBERUS, THE  STEVENS, FRANCIS  SR-H,H
HEADS TO THE STORM  DRAKE, DAVID/MIESEL, SANDRA  PA,SFWA-P:2
HEALER, THE  BLUMLEIN, MICHAEL, M. D.  H
HEALER  JENSEN, KRIS  P
HEALER [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,P,H,SFWA
HEALER’S WAR, THE  SCARBOROUGH, ELIZABETH ANN  C/P,SFWA-P,P,H
HEALING OF CROSSROADS, THE [BJ VAUGHAN 3]  O’DONOHOE, NICK  P
HEALTHY DEAD, THE  ERIKSON, STEVEN  SR-H
HEART BEAST  LEE, TANITH  P
HEART OF A DOG, THE  BULGAKOV, MIKHAIL  LP
HEART OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 3]  THORPE, GAV  P
HEART OF GLASS [DARK ENCHANTMENT]  COOPER, LOUISE  P
HEART OF GOLD  SHINN, SHARON  C/P,LP,P
HEART OF MIDNIGHT [@RAVENLOFT]  KING, J. ROBERT  P
HEART OF MYRIAL, THE [SHADOWLEAGUE 1]  FUREY, MAGGIE  P
HEART OF RED IRON [BARRAZAN V 2]  GOTLIEB, PHYLLIS  H
HEART OF SPARRILL, THE [WHEEL 2]  FRANKOS, STEVEN  P
HEART OF STARS, THE [@RHIANNON’S RIDE #3]  FORSYTH, KATE  P
HEART OF STONE [@ASTROLOGER 1]  DEMARTINO, DENNY  P
HEART OF TARDIS [DOCTOR WHO PAST DOCTORS]  STONE, DAVE  P
HEART OF THE COMET  BENFORD, GREGORY/BRIN, DAVID  C/P,SFWA,P,H
HEART OF THE DRAGON  SHOWALTER, GENA  P
HEART OF THE SUN [STAR TREK]  SARGENT, PAMELA/ZEBROWSKI, GEORGE  C/P,P
HEART OF THE TIGER [@WING COMMANDER 4]  FORSTCHEN, WILLIAM R./KEITH, ANDREW  P

TITLE  AUTHOR  CODE
HEART OF THE WARRIOR, THE [STAR TREK DS9]  BETANCOURT, JOHN GREGORY  C/P:2,P
HEART OF THE WORLD  HAGGARD, H. RIDER  LP
HEART OF VALOR [NIGHT OF THE SOLSTICE 2]  SMITH, L. J.  H
HEART OF WHITENESSE  WALDROP, HOWARD  SR-H
HEART READERS  RUSCH, KRISTINE KATHRYN  C/P,P
HEART-SHAPED BOX  HILL, JOE  H
HEARTBEEPS  HILL, JOHN  P
HEARTFIRE [ALVIN MAKER 5]  CARD, ORSON SCOTT  H,P
HEARTLIGHT  BARRON, T. A.  P
HEARTLIGHT  BRADLEY, MARION ZIMMER  LP
HEARTS AND ARMOR [BRONWYN 3]  MILLER, RON  P
HEART’S BLOOD [PIT DRAGONS 2]  YOLEN, JANE  C/P,P
HEARTS OF CHAOS [@BATTLETECH 26]  MILAN, VICTOR  P
HEARTSEASE [CHANGES 2]  DICKINSON, PETER  H:2
HEARTSTONE & SABER  SINGLETON, JACQUI  LP
HEARTWOOD [TRICKSTER’S GAME 1]  CAMPBELL, BARBARA  P
HEAT  HERZOG, ARTHUR  P
HEAT [@BUFFY/ANGEL]  HOLDER, NANCY  P
HEAT STROKE [WEATHER WARDEN 2]  CAINE, RACHEL  P
HEAT WAVE [QUANTUM LEAP]  KENT, MELANIE  P
HEATHERN [DRYCO 3]  WOMACK, JACK  P,SFWA
HEAVEN  STEWART, IAN/COHEN, JACK  H
HEAVEN CAN WAIT  FLEISCHER, LEONORE  C/P,P
HEAVEN CENT [@XANTH #11]  ANTHONY, PIERS  H:2,P
HEAVEN CHRONICLES [HEAVEN 1,2]  VINGE, JOAN D.  C/P,P,H
HEAVEN KNOWS WHY  TAYLOR, SAMUEL W.  H
HEAVEN MAKERS, THE  HERBERT, FRANK  C/P:2,P
HEAVEN MAKERS, THE (REVISED)  HERBERT, FRANK  C/P,P
HEAVEN SENT  CROWTHER, PETER  PA
HEAVENLY BREAKFAST: AN ESSAY ON THE WINTER OF LOVE  DELANY, SAMUEL R.  R/REF-P
HEAVENLY HORSE FROM THE OUTERMOST WEST, THE  STANTON, MARY  C/P:2,P
HEAVENLY HOST, THE  ASIMOV, ISAAC  C/P,P
HEAVEN’S REACH [UPLIFT 6]  BRIN, DAVID  H:2,P
HEAVY PLANET  CLEMENT, HAL  LP
HEAVY TIME  CHERRYH, C. J.  C/P,P,H:2
HEAVY WEATHER  STERLING, BRUCE  C/P,P,H
HECATE’S CAULDRON  SHWARTZ, SUSAN  C/PA,PA
HECATE’S GLORY [ENEMY GLORY 2]  MICHALSON, KAREN  H
HECTOR SERVADAC  VERNE, JULES  FP
HEDERICK THE THEOCRAT [@DRAGONLANCE VILLAINS #4]  SEVERSON, ELLEN DODGE  C/P,P
HEDGE OF MIST, THE [KELTIAD ARTHUR 3]  KENNEALY-MORRISON, PATRICIA  H,P
HEECHEE RENDEZVOUS [HEECHEE 3]  POHL, FREDERIK  C/P,P,H
HEGIRA  BEAR, GREG  C/P,SFWA-P,P
HEIGHT OF THE SCREAM, THE  CAMPBELL, RAMSEY  SR-H
HEIGHTS OF HAVENREST, THE  SUBOND, VALERIE  P
HEILAND  SANDERS, FRANKLIN  P
HEINLEIN IN DIMENSION  PANSHIN, ALEXEI  R/REF-H,C/REF-H
HEINLEIN TRIO, A  HEINLEIN, ROBERT A.  H:2
HEIR APPARENT, THE [@GUARDIANS OF THE FLAME #4]  ROSENBERG, JOEL  C/P,P
HEIR OF AUTUMN [OHNDARIEN 1]  CARWYN, GILES/FAHNESTOCK, TODD  H
HEIR OF DARKNESS  RAHMAN, GLENN  P
HEIR OF RENGARTH [SWORD AND CIRCLET 2]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P
HEIR OF SEA AND FIRE  MCKILLIP, PATRICIA A.  C/P,P
HEIR OF STONE [@CLOUDMAGES #3]  FARRELL, S. L.  H
HEIR TO THE DRAGON [@BATTLETECH 28]  CHARRETTE, ROBERT N.  P
HEIR TO THE EMPIRE [STAR WARS]  ZAHN, TIMOTHY  H:2,P
HEIR TO THE KING [@SOONE 1]  SOMERS, JAMES  LP
HEIR TO THE SAND [HARAN OF OAS 2]  TAFF, JOHN F. D.  LP
HEIR TO THE SHADOWS [BLACK JEWELS 2]  BISHOP, ANNE  C/P,P
HEIRS OF BABYLON, THE  COOK, GLEN  C/P:2,P
HEIRS OF EARTH [ECHOES OF EARTH 3]  WILLIAMS, SEAN/DIX, SHANE  P
HEIRS OF EMPIRE [COLIN MACINTYRE 3]  WEBER, DAVID  C/P:2,P
HEIRS OF HAMMERFELL, THE [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P
HEIRS OF PROPHECY [@FORGOTTEN REALMS SEMBIA]  SMEDMAN, LISA  P
HEIRS OF THE FORCE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
HELDAN, THE  TALMADGE-BICKMORE, DEBORAH  P
HELENE [TROS OF SAMOTHRACE 4]  MUNDY, TALBOT  SR-P
HELIX  BROWN, ERIC  P
HELIX AND THE SWORD, THE  MCLOUGHLIN, JOHN C.  C/P,P,H
HELIX FILE, THE  BLANKENSHIP, WILLIAM D.  H
HELL  WILLIAMSON, CHET  P
HELL BELOW [DOC SAVAGE #99]/LOST GIANT, THE [DOC SAVAGE 100]  ROBESON, KENNETH  C/P,@P
HELL FRUIT, THE  ROSE, LAWRENCE F.  SR-P
HELL HATH NO FURY  WEBER, DAVID/EVANS, LINDA  H

TITLE  AUTHOR  CODE
HELL HEART [@VOR THE MAELSTROM]  VARDEMAN, ROBERT E.  P
HELL HOUSE  MATHESON, RICHARD  P
HELL ON EARTH [@DOOM #2]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P,P
HELL ON EARTH [@TRAVELER #10]  DRUMM, D. B.  P
HELL ON EARTH  REAVES, MICHAEL  H
HELL ON HIGH [SYMPATHY FOR THE DEVIL 3]  LISLE, HOLLY/NOLAN, TED  P
HELL RISING [@OUTLANDERS]  AXLER, JAMES  P
HELL-STORM [@WORLD OF DARKNESS WEREWOLF]  MOORE, JAMES A.  P
HELL TO PAY [NIGHTSIDE 7]  GREEN, SIMON R.  P
HELLBORN  BRANDNER, GARY  P
HELLBOUND [@CROW]  ATTANASIO, A. A.  H,LP
HELLBOY  NAVARRO, YVONNE  P
HELLBURNER  CHERRYH, C. J.  H:2,P
HELLER’S LEAP  WALLACE, IAN  C/P:2,P
HELLFIRE  SAUL, JOHN  P
HELLFIRE FILES OF JULES DE GRANDIN, THE  QUINN, SEABURY  P
HELLFIRE REBELLION, THE [TIMEWARS 10]  HAWKE, SIMON  C/P,P
HELLFLOWER [HELLFLOWER 1]  BES SHAHAR, ELUKI  C/P:2,SFWA-P:2,P
HELLFLOWER  SMITH, GEORGE O.  C/P,P
HELLHOUND MAGIC [@FAR SIDE OF FOREVER #2]  GREEN, SHARON  P
HELLHOUND PROJECT, THE  GOULART, RON  H,SFWA
HELLICONIA SPRING [HELLICONIA 1]  ALDISS, BRIAN W.  H:2,P,SFWA:2
HELLICONIA SUMMER [HELLICONIA 2]  ALDISS, BRIAN W.  H:2,P,SFWA
HELLICONIA WINTER [HELLICONIA 3]  ALDISS, BRIAN W.  C/P,SFWA-P,H:2,P,SFWA
HELLO AMERICA  BALLARD, J. G.  P
HELLO, GORGEOUS  DAVIDSON, MARYJANICE  LP
HELLO, LEMURIA HELLO  GOULART, RON  P
HELLO OUT THERE  MCDEVITT, JACK  LP
HELLO SUMMER, GOODBYE  CONEY, MICHAEL G.  H
HELLQUAD  GOULART, RON  C/P:2,P
HELL’S BELLES  KESSLER, JACKIE  LP
HELL’S CARTOGRAPHERS  ALDISS, BRIAN W./HARRISON, HARRY  SFWA,R/REF-H
HELL’S FAIRE (HARDCOVER AND CD) [POSLEEN 4]  RINGO, JOHN  H:2
HELL’S FAIRE (PROOF) [POSLEEN 4]  RINGO, JOHN  H,SR-LP
HELL’S FEAST [BLOODSHADOWS]  FARSHTEY, GREG  LPA
HELL’S GATE  KOONTZ, DEAN R.  P
HELL’S GATE [MULTIVERSE 1]  WEBER, DAVID/EVANS, LINDA  H
HELL’S PAVEMENT  KNIGHT, DAMON  SR-P,C/P:2
HELLSPARK  KAGAN, JANET  H:2,SFWA-P,P
HELLSTAR  REAVES, MICHAEL/PERRY, STEVE  C/P:2,P
HELLSTONE  SPRUILL, STEVEN G.  P
HELLSTROM’S HIVE  HERBERT, FRANK  C/P,P,H
HELLWALKER  VESSER, CAROLYN  C/P,SFWA-P,P
HELLWORLD  GREEN, SIMON R.  P
HELM  GOULD, STEVEN C.  H,P
HELMA [TROS OF SAMOTHRACE 2]  MUNDY, TALBOT  SR-P,C/P:2
HELMSMAN, THE [HELMSMAN 1]  BALDWIN, BILL  C/P:2,P
HELP, HELP, THE GLOBOLINKS!  MENOTTI, GIAN CARLO/DEAN, LEIGH  H
HELVERTI INVASION, THE [LIZARD WAR 2]  DALMAS, JOHN  P
HEMINGWAY HOAX, THE  HALDEMAN, JOE  H:2,P
HENCE  LEITHAUSER, BRAD  H
HENDRA’S BOOK [@PSI PATROL 2]  BENOIT, HENDRA  P
HENRY BROCKEN  DE LA MARE, WALTER J.  SR-H
HENRY MARTYN [HENRY MARTYN 1]  SMITH, L. NEIL  C/P,P
HEPHAESTUS PLAGUE, THE  PAGE, THOMAS  C/P,P,H
HER KLINGON SOUL [STAR TREK DAY OF HONOR 3/STAR TREK VOYAGER] FRIEDMAN, MICHAEL JAN  C/P:2,P
HER MAJESTY’S WIZARD [WIZARD IN RHYME 1]  STASHEFF, CHRISTOPHER  C/P,P,H
HER MONSTER  COLLIGNON, JEFF  P
HER PILGRIM SOUL  BRENNERT, ALAN  H,SFWA
HER SMOKE ROSE UP FOREVER  TIPTREE, JAMES JR.  H
HERALD, THE  SHAARA, MICHAEL  P

HERALDS OF THE STORM [@WORLD OF DARKNESS YEAR OF THE SCARAB #1] BATES, ANDREW

P

HERCULES TEXT, THE  MCDEVITT, JACK  C/P:2,P
HERDS  GOLDIN, STEPHEN  SR-P
HERE ABIDE MONSTERS  NORTON, ANDRE  C/P,P,H
HERE BE DAEMONS  COPPER, BASIL  H
HERE BE DEMONS [HERE BE DEMONS 1]  FRIESNER, ESTHER M.  C/P,P
HERE BE MONSTERS [BUFFY]  DOKEY, CAMERON  P
HERE BE MONSTERS [WARLOCK’S HEIRS 5]  STASHEFF, CHRISTOPHER  P
HERE COMES CIVILIZATION [COMPLETE TENN 2]  TENN, WILLIAM  H
HERE COMES THE SUN  HOLT, TOM  P
HERE I STAY  MICHAELS, BARBARA  SFWA-P
HERE, THERE & EVERYWHERE  ROBERSON, CHRIS  H

The MITSFS Titledex  September 24, 2007  Page 163
TITLE  AUTHOR  CODE
HERE, THERE BE DRAGONS  OWEN, JAMES A.  H
HERE THERE BE DRAGONS [STAR TREK TNG]  PEEL, JOHN  C/P:3,P
HERE TODAY  COATES, JOHN  H
HEREAFTER GANG, THE  BARRETT, NEAL JR.  H
HERESY [AQUASILVA 1]  AUDLEY, ANSELM  H
HERETIC KINGS, THE [@MONARCHIES OF GOD #2]  KEARNEY, PAUL  P

HERETIC SAINT [@WARHAMMER 40K DAEMONIFUGE] CAMPBELL, JIM/SPURRIER, SIMON/ RENNIE, GORDON/WALKER, KEV

SCX

HERETIC’S FAITH [@MECHWARRIOR DARK AGE/BATTLETECH]  BILLS, RANDALL N.  P
HERETICS OF DUNE [DUNE 5]  HERBERT, FRANK  C/P,LP,H
HERETICUS [WARHAMMER 40K EISENHORN 3]  ABNETT, DAN  P
HERIS SERRANO [HERIS SERRANO 1,2,3]  MOON, ELIZABETH  H
HERITAGE [EARTH FINAL CONFLICT]  DURGIN, DORANNA  H
HERITAGE [DOCTOR WHO PAST DOCTORS]  SMITH, DALE  P
HERITAGE OF FLIGHT  SHWARTZ, SUSAN  C/P:2,SFWA-P,P
HERITAGE OF HASTUR, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
HERITAGE OF STARS, A  SIMAK, CLIFFORD D.  H:2,P

HERITAGE OF THE LIZARD PEOPLE [@PERRY RHODAN #113]/ DEATH’S DEMAND [PERRY RHODAN 114] ACE

C/PA,PA

HERITAGE UNIVERSE, THE [HERITAGE UNIVERSE 1,2,3]  SHEFFIELD, CHARLES  H:2

HERITIC: BETRAYERS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 2] MCGOUGH, SCOTT

P

HERLAND  GILMAN, CHARLOTTE P.  H
HERMES FALL, THE  BAXTER, JOHN  C/P,P
HERMES STONE, THE  EILERS, ROBERT  P
HERMETECH  CONSTANTINE, STORM  C/P,P
HERMIT’S SWING  KOLUPAEV, VICTOR  H
HERO!  DUNCAN, DAVE  C/P:2,P
HERO  KERNS, DANIEL R.  P
HERO, THE  RINGO, JOHN/WILLIAMSON, MICHAEL Z.  H
HERO  ROSENBERG, JOEL  C/P,P
HERO AND THE CROWN, THE  MCKINLEY, ROBIN  H:2,P
HERO BORN, A [REALMS OF CHAOS 1]  STACKPOLE, MICHAEL A.  P
HERO FROM OTHERWHERE, THE  WILLIAMS, JAY  C/P:2,P
HERO IN THE SHADOWS [DRENAI]  GEMMELL, DAVID  P
HERO KING, THE [@VARAYAN #3]  SHELLEY, RICK  P
HERO OF DOWNWAYS, THE  CONEY, MICHAEL G.  C/P:2,P
HERO OF DREAMS [DREAMWORLD 1]  LUMLEY, BRIAN  LP
HERO OF VARAY, THE [@VARAYAN #2]  SHELLEY, RICK  P
HERO PULP INDEX, THE  WEINBERG, ROBERT/MCKINSTRY, LOHR  SRL/IND
HERO STRIKES BACK, THE [SHIELD AND SOURCE 2]  MOORE, MOIRA J.  P
HERO TALES FROM THE BRITISH ISLES  PICARD, BARBARA LEONIE  P
HEROD MEN, THE  KAMIN, NICK  D13805
HEROES, THE [GUARDIANS OF THE FLAME 4,5]  ROSENBERG, JOEL  H:2
HEROES [@LAWS OF THE BLOOD 5]  SIZEMORE, SUSAN  P
HEROES AND FOOLS [@DRAGONLANCE FIFTH AGE]  WEIS, MARGARET/HICKMAN, TRACY  C/PA,PA
HEROES AND HORRORS  LEIBER, FRITZ  P
HEROES AND VILLAINS  CARTER, ANGELA  C/P,P
HEROES DIE [HARI MICHAELSON 1]  STOVER, MATTHEW WOODRING  H,LP
HEROES IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA
HEROES INC.  CROCCO, KYLE  C/P,P
HEROES OF ASGARD, THE  KEARY, A./KEARY, E.  SR-H
HEROES OF ZARA KEEP  GREGORY, GUY  P
HEROES WANTED  CROCCO, KYLE  C/P,P
HEROIC FANTASY  PAGE, GERALD W./REINHARDT, HANK  C/PA,PA
HEROIC VISIONS  SALMONSON, JESSICA AMANDA  PA,C/PA
HEROIC VISIONS II  SALMONSON, JESSICA AMANDA  PA
HEROICS  EFFINGER, GEORGE ALEC  H,SFWA
HEROICS FOR BEGINNERS  MOORE, JOHN  C/P,P
HEROINE OF THE WORLD, A  LEE, TANITH  C/P,P,H
HEROINES  KELLY, JAMES PATRICK  SR-H
HEROING [JIANA 1]  AB HUGH, DAFYDD  C/P,P
HERO’S JUSTICE, A [@DRAGONLANCE ERGOTH #3]  THOMPSON, PAUL B./COOK, TONYA C.  P
HERO’S QUEST [@JUSTICE LEAGUE OF AMERICA GREEN LANTERN]  O’NEIL, DENNIS  P
HERO’S RETURN, THE [@WIZARD WAR #3]  COOK, HUGH  C/P:2,P
HERO’S SONG  PATTOU, EDITH  H
HERO’S TRIAL [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #1]  LUCENO, JAMES  P
HERO’S WALK  CRANE, ROBERT  C/P:2,P
HEROVIT’S WORLD  MALZBERG, BARRY N.  C/P:2,SFWA,P,H
HESTIA  CHERRYH, C. J.  C/P:2,P
HEX [DOC SAVAGE #37]  ROBESON, KENNETH  C/P:2,@P
HEX AND THE CITY [NIGHTSIDE 4]  GREEN, SIMON R.  P
HEX WITCH OF SELDOM, THE  SPRINGER, NANCY  C/P,P
HEXWOOD  JONES, DIANA WYNNE  H

TITLE  AUTHOR  CODE
HIDDEN CITY, THE [DUNGEON 5]  DE LINT, CHARLES  C/P:2,P
HIDDEN CITY, THE [TAMULI 3]  EDDINGS, DAVID  H:2
HIDDEN CITY, THE  KELLY, SHANE  SR-L
HIDDEN DEATH  GRANT, MAXWELL  P
HIDDEN DRAGON, THE [@KNIGHTS OF THE SILVER DRAGON #7]  TRUMBAUER, LISA TRUTKOFF  LP
HIDDEN ECHOES  JEFFERIES, MIKE  P
HIDDEN EMPIRE [SEVEN SUNS 1]  ANDERSON, KEVIN J.  C/P,P
HIDDEN FAMILY, THE [MERCHANT PRINCES 2]  STROSS, CHARLES  H
HIDDEN FIRES [NUALA 3]  KIMBRIEL, KATHERINE ELISKA  C/P:2,P
HIDDEN IN SIGHT [WEB SHIFTERS #3]  CZERNEDA, JULIE E.  P
HIDDEN KINGDOM, THE  BEEDING, FRANCIS  H
HIDDEN LAND, THE [SECRET COUNTRY 2]  DEAN, PAMELA C.  C/P,P
HIDDEN LAND, THE [DAUGHTERS OF BAST 1]  ISIDORE, SARAH  P
HIDDEN PLACE, A  WILSON, ROBERT CHARLES  C/P:2,SFWA-P,P
HIDDEN PLANET, THE  WOLLHEIM, DONALD A.  PA,C/PA
HIDDEN QUEEN, THE [ANGHARA 1]  ALEXANDER, ALMA  P
HIDDEN REALMS, THE [HIDDEN REALMS 1]  GREEN, SHARON  SFWA-P
HIDDEN SAINT, THE [@POLTERGEIST THE LEGACY]  HAUTALA, RICK  P
HIDDEN SIDE OF THE MOON, THE  RUSS, JOANNA  P
HIDDEN SPECTRE, THE  TRALINS, ROBERT  [C/RAN-P]
HIDDEN STARS, THE [RUNE OF UNMAKING 1]  HOWARD, MADELINE  H,LP
HIDDEN TALENTS  LUBAR, DAVID  H:2,LP
HIDDEN TEMPLE, THE [ELVEN KINGDOMS 3]  COOKE, CATHERINE  C/P,SFWA-P,P
HIDDEN TREASURE OF GLASTON, THE  JEWETT, ELEANORE M.  P
HIDDEN TRUTH [FIRST TRUTH 2]  COOK, DAWN  P
HIDDEN UNIVERSE, THE  FARLEY, RALPH M.  H
HIDDEN VARIABLES  SHEFFIELD, CHARLES  C/P,P
HIDDEN WAR, THE  ARMSTRONG, MICHAEL  P
HIDDEN WARRIOR [TOBIN OF SKALA 2]  FLEWELLING, LYNN  C/P,P
HIDDEN WORLD  COBLENTZ, STANTON A.  C/P,P
HIDDEN WORLD, THE [WATCHERS 2]  GORDON, STUART  P
HIDDEN WORLDS, THE  LANDON, KRISTIN  P
HIDDEN WORLDS OF ZANDRA, THE  ROTSLER, WILLIAM  H,SFWA
HIDE AND SEEK [@VENUS PRIME #3]  PREUSS, PAUL  C/P:2,P
HIER, L’AN 2000  SADOUL, JACQUES  A
HIERARCHIES  PHILLIFENT, JOHN T.  D53415
HIEROS GAMOS OF SAM AND AN SMITH, THE  SAXTON, JOSEPHINE  C/P,P
HIERO’S JOURNEY  LANIER, STERLING E.  C/P:2,P
HIGH AZTECH  HOGAN, ERNEST  P
HIGH COUCH OF SILISTRA [SILISTRA 1]  MORRIS, JANET  C/P:2,P
HIGH CRUSADE, THE  ANDERSON, POUL  H:2,SFWA-P,P
HIGH CRYSTAL [SIX MILLION DOLLAR MAN 3]  CAIDIN, MARTIN  H:2,@P
HIGH DERYNI [DERYNI CHRONICLES 3]  KURTZ, KATHERINE  C/P:2,SR-P,P
HIGH DESTINY, THE  MORGAN, DAN  C/P,P
HIGH EXPLOSIVE  PHILLIPS, GORDON  H
HIGH FANTASTIC  TEM, STEVE RASNIC  HA
HIGH FRONTIER  CLAREMONT, CHRIS  H
HIGH FRONTIER, THE  O’NEILL, GERARD K.  [C/RAN-P:2]
HIGH HEX, THE  JANIFER, LAURENCE M./TREIBICH, S. J.  D72400
HIGH HOUSE, THE [EVENMERE 1]  STODDARD, JAMES  H,P
HIGH HOUSE, THE (PROOF) [EVENMERE 1]  STODDARD, JAMES  SR-P
HIGH JUSTICE  POURNELLE, JERRY  C/P:2,P
HIGH KING, THE [PRYDAIN 5]  ALEXANDER, LLOYD  C/P:2,LP
HIGH KINGS, THE  CHANT, JOY  P
HIGH LORD, THE [BLACK MAGICIAN 3]  CANAVAN, TRUDI  C/P,P
HIGH MOON [@SPACE COPS 3]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
HIGH PLACE, THE  CABELL, JAMES BRANCH  C/P:2,P
HIGH QUEEN, THE [CHILD QUEEN 2]  MCKENZIE, NANCY  P
HIGH RHULAIN [@REDWALL]  JACQUES, BRIAN  P
HIGH-RISE  BALLARD, J. G.  H,P
HIGH SORCERY  NORTON, ANDRE  C/P:2,P
HIGH SPIRTS [CHARMED]  CIENCIN, SCOTT  P
HIGH STAKES [@PHOTON #2]  PETERS, DAVID  P
HIGH STEEL  HALDEMAN, JACK C. II/DANN, JACK  H,P
HIGH-TECH KNIGHT, THE [CONRAD STARGARD 2]  FRANKOWSKI, LEO  C/P:2,P
HIGH TENSION  ING, DEAN  P
HIGH THRUST  AUDEN, RENEE  SR-L
HIGH VACUUM  MAINE, CHARLES ERIC  C/P:2,P
HIGH WIZARDRY [YOUNG WIZARDS 3]  DUANE, DIANE  H:2,P
HIGHER EDUCATION  SHEFFIELD, CHARLES/POURNELLE, JERRY  C/P:2,P,H
HIGHER MYTHOLOGY [MYTHOLOGY 3]  NYE, JODY LYNN  P
HIGHER SPACE, THE  NASIR, JAMIL  P
HIGHER THINGS  HARRISON, MICHAEL  H
HIGHEST SCIENCE, THE [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 165
TITLE  AUTHOR  CODE
HIGHLANDER GAMBIT [@BATTLETECH 18]  PARDOE, BLAINE LEE  P
HIGHWAY OF ETERNITY  SIMAK, CLIFFORD D.  C/P,SFWA,P,H
HIGHWAYMAN, THE  SALVATORE, R. A.  P
HIGHWAYMEN: ROBBERS & ROGUES  ROBERSON, JENNIFER  C/PA,PA
HIGHWAYS IN HIDING  SMITH, GEORGE O.  H,P
HIGHWOOD  BARRETT, NEAL JR.  D33710
HIJACK  WELLEN, EDWARD  C/P:2,P
HILL OF DREAMS, THE  MACHEN, ARTHUR  H,SR-L
HILLS OF FARAWAY: A GUIDE TO FANTASY, THE  WAGGONER, DIANA  R/REF-H
HILLS WERE LIARS, THE  HUGHES, RILEY  P
HINTERLAND [GODSLAYER 2]  CLEMENS, JAMES  H
HIS CONQUERING SWORD [@SWORD OF HEAVEN #2]  ELLIOTT, KATE  C/P:2,SFWA-P,P
HIS FATHER’S SON  BENNETT, NIGEL/ELROD, P. N.  P
HIS LAST COMMAND [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  H
HIS MAJESTY’S DRAGON [TEMERAIRE 1]  NOVIK, NAOMI  P
HIS MASTER’S VOICE  LEM, STANISLAW  C/P:2,P
HIS NAME WAS DEATH  BROWN, FREDRIC  P
HIS SHARE OF GLORY  KORNBLUTH, C. M.  H:2
HISTORY 101 [DOCTOR WHO EIGHTH DOCTOR]  HALLIDAY, MAGS L.  P
HISTORY MAKER, A  GRAY, ALASDAIR  LP
HISTORY OF ALCHEMY, A  HUTIN, SERGE  [C/RAN-P]
HISTORY OF MR. POLLY, THE  WELLS, H. G.  H
HISTORY OF THE CALIPH WATHEK, THE  BECKFORD, WILLIAM  LP
HISTORY OF THE DC UNIVERSE  WOLFMAN, MARV  CX

HISTORY OF THE HUGO, NEBULA, AND INTERNATIONAL FANTASY AWARDS, A DE VORE, HOWARD

SRL/IND

HISTORY OF THE HUGO, NEBULA, AND INTERNATIONAL FANTASY AWARDS, A (1978)

FRANSON, DONALD/DEVORE, HOWARD

R/REF-P

HISTORY OF THE UNIVERSE, A [DOCTOR WHO]  PARKIN, LANCE  R/REF-P,C/REF-P
HISTORY OF UFO CRASHES, A  RANDLE, KEVIN D.  [C/RAN-P]
HIT OR MYTH [MYTH ADVENTURES 4]  ASPRIN, ROBERT LYNN  C/P:2,LP
HITCHHIKERS’ GUIDE TO THE GALAXY, THE [HITCHHIKERS 1]  ADAMS, DOUGLAS  C/P,H:2,P

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (5/6 CASSETTES)

[HITCHHIKERS 1] ADAMS, DOUGLAS

SR-MM

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (6 CASSETTES)

[HITCHHIKERS 1] ADAMS, DOUGLAS

SR-MM:2

HITCHHIKERS’ GUIDE TO THE GALAXY, THE (WITH CASABLANCA)

[HITCHHIKERS] ADAMS, DOUGLAS

SR-MM

HITLER VICTORIOUS BENFORD, GREGORY/

GREENBERG, MARTIN H.

HA,PA,SFWA-P

HITTING THE SKIDS IN PIXELTOWN CARD, ORSON SCOTT/OLEXA, KEITH/

O’TOOLE, CHRISTIAN

LPA

HIVE  CURRAN, TIM  LP
HOBBIT, THE  TOLKIEN, J. R. R.  H:2,SR-H,P

HOBBIT, THE TOLKIEN, J. R. R./DIXON, CHARLES/

DEMING, SEAN

CX

HOBBIT COMPANION, THE [LORD OF THE RINGS]  DAY, DAVID  R/REF-H
HOBBY [YOUNG MERLIN 2]  YOLEN, JANE  H
HOBGOBLIN  COYNE, JOHN  H,P
HOB’S BARGAIN, THE  BRIGGS, PATRICIA  P
HOCUS POCUS  VONNEGUT, KURT JR.  P
HOGAR, LORD OF THE ASYR  SHARPE, JOHN RUFUS III  P
HOGFATHER [DISCWORLD 20/DISCWORLD DEATH 4]  PRATCHETT, TERRY  H,P
HOGFOOT RIGHT AND BIRD-HANDS  KILWORTH, GARRY  LP
HOKA! [HOKA 1]  ANDERSON, POUL/DICKSON, GORDON R.  C/P,LP,H
HOKA! HOKA! HOKA! [HOKA 2]  ANDERSON, POUL/DICKSON, GORDON R.  P
HOKAS POKAS! [HOKA 3]  ANDERSON, POUL/DICKSON, GORDON R.  P
HOLD FAST TO LOVE  SIMON, JO ANN  P
HOLD ZERO!  GEORGE, JEAN CRAIGHEAD  H
HOLDER OF LIGHTNING [@CLOUDMAGES #1]  FARRELL, S. L.  H
HOLDER OF LIGHTNINGS (PROOF) [@CLOUDMAGES #1]  FARRELL, S. L.  SR-LP
HOLDING THE LINE [SPEC OPS SQUAD 1]  SHELLEY, RICK  P
HOLDING WONDER  HENDERSON, ZENNA  C/P,P,H
HOLDING YOUR EIGHT HANDS  LUCIE-SMITH, EDWARD  LPA
HOLE IN SPACE, A  NIVEN, LARRY  C/P:2,P
HOLE IN THE ZERO, THE  JOSEPH, M. K.  C/P,P,H
HOLLENTANZ DER MARIONETTEN [@PERRY RHODAN #47]  KNEIFEL, HANS  FPR

HOLLOW CHOCOLATE BUNNIES OF THE APOCALYPSE, THE [EDDIE BEAR 1] RANKIN, ROBERT

H

HOLLOW CROWN AFFAIR, THE [@MAN FROM UNCLE #17]  MCDANIEL, DAVID  C/P:2,P
HOLLOW EARTH, THE  BERNARD, RAYMOND  [C/RAN-P:2]
HOLLOW EARTH, THE  RUCKER, RUDY  H,SFWA
HOLLOW FACES, MERCILESS MOONS  HOME, WILLIAM SCOTT/FABIAN, STEPHEN E. VLP
HOLLOW HILLS, THE [ARTHURIAN 2]  STEWART, MARY  C/P,P,LP,H

The MITSFS Titledex  September 24, 2007  Page 166
TITLE  AUTHOR  CODE
HOLLOW LANDS, THE [DANCERS AT THE END OF TIME 2]  MOORCOCK, MICHAEL  H:2,SFWA,P
HOLLOW MAN, THE  SIMMONS, DAN  C/P,P,H
HOLLOW MEN [@STAR TREK DS9]  MCCORMACK, UNA  P
HOLLOW MEN, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH/DAY, MARTIN  C/P,P
HOLLOWING, THE [RYHOPE WOOD 2]  HOLDSTOCK, ROBERT P.  H,P
HOLLYWOOD NOIR [@ANGEL]  MARIOTTE, JEFF  P
HOLMES-DRACULA FILE, THE [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
HOLO MEN [CYBERPUNK 2.0.2.0.]  BILLIAS, STEPHEN  P
HOLOCAUST FOR HIRE  SILVA, JOSEPH  P
HOLOCAUST HORROR [@MUTANTS AMOK #4]  GRANT, MARK  C/P,P
HOLY BIBLE, THE (GIDEON)  GOD  HA
HOLY CITY, THE  PARIS, MATTHEW  LP
HOLY FIRE  STERLING, BRUCE  H:2,SR-P
HOLY TERROR, THE  WELLS, H. G.  SR-H
HOMBRE DEL BICENTENARIO, EL  ASIMOV, ISAAC  FP
HOME  FULLER, KIMBERLY  H,P
HOME FROM THE SHORE  DICKSON, GORDON R.  C/P:2,LP
HOME IS THE HANGMAN  ZELAZNY, ROGER  TD21,SFWA-TD21
HOME IS THE HUNTER [STAR TREK]  KRAMER-ROLLS, DANA  C/P:2,P
HOME SWEET HOME: 2010 A. D.  REYNOLDS, MACK/ING, DEAN  P
HOME — TO AVALON [CAMELOT 4]  LANDIS, ARTHUR H.  C/P,P
HOMEBODY  CARD, ORSON SCOTT  P
HOMEBREW  ANDERSON, POUL  SR-H
HOMECALLING AND OTHER STORIES  MERRIL, JUDITH  H
HOMECOMING  DALMAS, JOHN  P
HOMECOMING [@STAR TREK VOYAGER HOMECOMING 1]  GOLDEN, CHRISTIE  P
HOMECOMING  JACKSON, ANITA  SR-P
HOMECOMING, THE  LONGYEAR, BARRY B.  H,SFWA
HOMECOMING [@ROBOTECH #3]  MCKINNEY, JACK  P
HOMECOMING: EARTH [HOMECOMING 4,5]  CARD, ORSON SCOTT  H:2
HOMECOMING: HARMONY [HOMECOMING 1,2,3]  CARD, ORSON SCOTT  H:2
HOMEFALL [LAST LEGION 4]  BUNCH, CHRIS  C/P:2,P
HOMEGOING  POHL, FREDERIK  H:2,P
HOMELAND [@FORGOTTEN REALMS DARK ELF #1]  SALVATORE, R. A.  C/P:2,P
HOMEM DEMOLIDO, O  BESTER, ALFRED  FH
HOMESMIND  SARGENT, PAMELA  LP
HOMEWARD AND BEYOND  ANDERSON, POUL  C/P:2,SFWA,P
HOMEWARD BOUND [@DEATHLANDS]  AXLER, JAMES  P
HOMEWARD BOUND  TURTLEDOVE, HARRY  H
HOMEWARD BOUND  VERNE, JULES  H
HOMEWARD BOUNDERS, THE  JONES, DIANA WYNNE  H,P
HOMEWARD TO ITHAKA  WIBBERLEY, LEONARD  H
HOMEWORLD [TO THE STARS 1]  HARRISON, HARRY  C/P:2,P
HOMING, THE  CAMPBELL, JEFFREY  H
HOMING PIGEONS [@SCHRODINGER’S CAT #3]  WILSON, ROBERT ANTON  C/P,P
HOMINIDS [NEANDERTHAL PARALLAX 1]  SAWYER, ROBERT J.  H,P
HOMMUNCULUS, THE  KELLER, DAVID H., M. D.  H
HOMUNCULUS  BLAYLOCK, JAMES P.  P
HONEY FOR THE BEARS  BURGESS, ANTHONY  C/P,P
HONEYMOON IN HELL  BROWN, FREDRIC  C/P:2,P
HONG ON THE RANGE [JACK HONG]  WU, WILLIAM F.  H
HONOR AMONG ENEMIES [HONOR HARRINGTON 6]  WEBER, DAVID  H,P
HONOR AND FIDELITY [@FIFTH FOREIGN LEGION #2]  KEITH, ANDREW/KEITH, WILLIAM H. JR.  P
HONOR BLADE [STAR TREK RIHANNSU 4]  DUANE, DIANE  C/P,P
HONOR BOUND [@STAR TREK I.K.S. GORKON 2]  DECANDIDO, KEITH R. A.  P
HONOR OF THE GRAVE [@WARHAMMER ANGELIKA FLEISCHER 1]  LAWS, ROBIN D.  P
HONOR OF THE QUEEN, THE [HONOR HARRINGTON 2]  WEBER, DAVID  C/P:2,SFWA-P,P
HONOR OF THE REGIMENT [@BOLOS #1]  FAWCETT, BILL  SFWA-P,C/PA,PA
HONORABLE BARBARIAN, THE [RELUCTANT KING]  DE CAMP, L. SPRAGUE  H:2,P
HONORABLE DEFENSE, AN [CRISIS OF EMPIRE 1]  DRAKE, DAVID/THOMAS, THOMAS T.  C/P:2,SFWA-P,P

HONORED ENEMY [@LEGENDS OF THE RIFTWAR #1/RIFTWAR] FEIST, RAYMOND E./

FORSTCHEN, WILLIAM R.

H

HONORVERSE DISK, THE (CD) [HONOR HARRINGTON]  BAEN PUBLISHING ENTERPRISES  SR-MM
HONOUR GUARD [WARHAMMER 40K GAUNT’S GHOSTS 4]  ABNETT, DAN  P
HOOD [KING RAVEN 1]  LAWHEAD, STEPHEN R.  H
HOOK  BROOKS, TERRY  C/P,P,H
HOORAY FOR HELLYWOOD [HERE BE DEMONS 3]  FRIESNER, ESTHER M.  P,SFWA-P:2
HOPE [DOCTOR WHO EIGHTH DOCTOR]  CLAPHAM, MARK  P
HOPE OF EARTH [GEODYSSEY 3]  ANTHONY, PIERS  H
HOPE’S END  CHAMBERS, STEPHEN  C/P,P
HOPESHIP [USS SKIPJACK 3]  LANG, SIMON  P
HOPKINS MANUSCRIPT, THE  SHERRIFF, R. C.  H:2
HOPSCOTCH  ANDERSON, KEVIN J.  H
HORDE, THE  GREEN, JOSEPH  SR-P

The MITSFS Titledex  September 24, 2007  Page 167
TITLE  AUTHOR  CODE
HORIZON ALPHA  MASON, DOUGLAS R.  C/P:2,P
HORIZON STORMS [@SEVEN SUNS #3]  ANDERSON, KEVIN J.  H
HORIZONS  ROSENBLUM, MARY  H
HORMONE JUNGLE, THE  REED, ROBERT  H,P
HORN CROWN [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  H:2,P
HORN: GREEN [@PERRY RHODAN #96]  ACE  C/PA,PA
HORN OF TIME, THE  ANDERSON, POUL  C/P:2,P
HORNED WARRIOR, THE [@BERSERKER SWIFTAXE 3]  CARLSEN, CHRIS  P
HORNS OF ELFLAND, THE  VESS, CHARLES  VLP
HORREUR DU METRO, L  MONTELEONE, THOMAS F.  FP
HORRIBLE HUMES [RUNESWORD 4]  BILLIAS, STEPHEN  C/P,P
HORROR, THE [@PERRY RHODAN #66]  ACE  C/PA,PA
HORROR-7  BLOCH, ROBERT  P
HORROR AT FONTENAY  DUMAS, ALEXANDRE  SR-P
HORROR AT OAKDEENE AND OTHERS  LUMLEY, BRIAN  H
HORROR CHAMBERS OF JULES DE GRANDIN, THE  QUINN, SEABURY  P
HORROR HOUSE  WILLIAMSON, J. N.  P
HORROR HUNTERS  ELWOOD, ROGER/GHIDALIA, VIC  C/PA,PA
HORROR IN THE CINEMA  BUTLER, IVAN  R/REF-P
HORROR IN THE MUSEUM AND OTHER REVISIONS, THE  LOVECRAFT, H. P.  SR-H,C/P,P,H
HORROR LITERATURE  TYMN, MARSHALL B.  R/REF-H
HORROR ON THE BEACH, THE  FOSTER, ALAN DEAN  SR-P
HORROR SHOW  KIHN, GREG  H
HORROR TALES  ELWOOD, ROGER  SR-VLHA
HORROR TIMES TEN  NORTON, ALDEN H.  PA
HORROR WEARS BLUE  CARTER, LIN  SFWA
HORRORS  GRANT, CHARLES L.  PA
HORRORS AND UNPLEASANTRIES (ARKHAM HOUSE BIBLIOGRAPHY)  JAFFERY, SHELDON  R/REF-H
HORRORS BEYOND  JONES, WILLIAM  LPA
HORRORS IN HIDING  MOSKOWITZ, SAM/NORTON, ALDEN H.  PA
HORRORS OF THE DANCING GODS [DANCING GODS 5]  CHALKER, JACK L.  C/P:2,P
HORRORS UNKNOWN  MOSKOWITZ, SAM  HA,PA
HORRORS UNSEEN  MOSKOWITZ, SAM  C/PA,PA
HORSE AND HIS BOY, THE [NARNIA 5]  LEWIS, C. S.  C/P:2,P

HORSE FANTASTIC GREENBERG, MARTIN H./

GREENBERG, ROSALIND M.

PA,C/PA

HORSE IN THE GRAY FLANNEL SUIT, THE  HATCH, ERIC  P
HORSE KING, THE [DRAGONREALM 7]  KNAAK, RICHARD A.  P
HORSE LORD, THE [BOOK OF YEARS 1]  MORWOOD, PETER  P
HORSE OF FLAME, THE  SHERMAN, JOSEPHA  P
HORSECLANS ODYSSEY [HORSECLANS 7]  ADAMS, ROBERT  C/P:2,P
HORSEGIRL, THE [HORSEGIRL 1]  ASH, CONSTANCE  C/P,P
HORSELORDS [FORGOTTEN REALMS EMPIRES 1]  COOK, DAVID  C/P:2,P
HORSEMASTER  SINGER, MARILYN  H
HORSEMEN FROM NOWHERE  ABRAMOV, A./ABRAMOV, S.  H
HORSES!  DANN, JACK/DOZOIS, GARDNER  PA
HORSES’ ASTEROID  FRITCH, CHARLES E.  P
HORSES OF HEAVEN  BRADSHAW, GILLIAN  C/P,P,H
HORSES OF THE NORTH [HORSECLANS 13]  ADAMS, ROBERT  P
HORSEWOMAN IN GODSLAND, A  EDWARDS, CLAUDIA J.  C/P,P
HORUS RISING [@WARHAMMER 40K HORUS HERESY 1]  ABNETT, DAN  P
HOSPITAL HORROR, THE  BINDER, OTTO  P
HOSPITAL STATION [SECTOR GENERAL 1]  WHITE, JAMES  C/P:2,P
HOST, THE  EMSHWILLER, PETER R.  C/P,P
HOST OF DRAGONS, A [DRAGON DELASANGRE 4]  TROOP, ALAN F.  P
HOSTAGE FOR HINTERLAND, A  DARNAY, ARSEN  C/P,P
HOSTAGE OF ZIR, THE [KRISHNA]  DE CAMP, L. SPRAGUE  H:2,P
HOSTAGE OF ZIR, THE (PROOF) [KRISHNA]  DE CAMP, L. SPRAGUE  SR-LP
HOSTILE TAKEOVER [@BLACK HOLE TRAVEL AGENCY #4]  MCKINNEY, JACK  C/P:2,P
HOSTILE WORLDS  HUNT, GILL  SR-P
HOSTS [REPAIRMAN JACK]  WILSON, F. PAUL  P

HOT BLOODED FEEHAN, CHRISTINE/SHAYNE, MAGGIE/ HOLLY, EMMA/KNIGHT, ANGELA

P

HOT RAIN  PORTNOY, HOWARD N.  P
HOT ROCK, THE  WESTLAKE, DONALD E.  P
HOT SKY AT MIDNIGHT  SILVERBERG, ROBERT  H:2,P
HOT SLEEP: THE WORTHING CHRONICLE [WORTHING]  CARD, ORSON SCOTT  C/P:2,P
HOT SPOT, THE [LAND OF THE GIANTS 2]  LEINSTER, MURRAY  P
HOT TIME IN OLD TOWN [MATHEW SWAIN 1]  MCQUAY, MIKE  P
HOT-WIRED DODO, THE [@WONDERLAND GAMBIT #3]  CHALKER, JACK L.  H:2,P
HOT ZONE, THE  PRESTON, RICHARD  R/REF-P
HOT ZONE [HORN]  SLOANE, BEN  P
HOTEL ANDROMEDA  CHALKER, JACK L.  C/PA,PA,SFWA-P
HOTEL COSMOS  BURKE, JONATHAN  SR-P

The MITSFS Titledex  September 24, 2007  Page 168
TITLE  AUTHOR  CODE
HOTEL TRANSYLVANIA [SAINT-GERMAIN 1]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
HOTHOUSE  ALDISS, BRIAN W.  C/P:2,P
HOUDINI DIRECTIVE, THE  CONNER, MIKE  P
HOUDINI: THE UNTOLD STORY  CHRISTOPHER, MILBOURNE  [C/RAN-P]
HOUND AND THE FALCON, THE [HOUND AND FALCON 1,2,3]  TARR, JUDITH  H:2
HOUND OF THE BASKERVILLES, THE  DOYLE, ARTHUR CONAN  P
HOUNDS OF DRACULA  JOHNSON, KEN  P
HOUNDS OF GOD, THE [HOUND AND FALCON #3]  TARR, JUDITH  P,@H,SFWA-P
HOUNDS OF SKAITH, THE [SKAITH 2]  BRACKETT, LEIGH  C/P:2,P
HOUNDS OF THE MORRIGAN, THE  O’SHEA, PAT  C/P,P
HOUNDS OF TINDALOS, THE  LONG, FRANK BELKNAP  C/P,P,H
HOUNDS OF TINDALOS: THE EARLY LONG, THE  LONG, FRANK BELKNAP  P
HOUNGAN, THE  WILLIAMSON, J. N.  P
HOUR 25 SHOW, THE (5 CASSETTES)  HODEL, MIKE/GILDEN, MEL  SR-MM
HOUR OF JUDGEMENT [JUDICIARY 3]  MATTHEWS, SUSAN R.  C/P,P
HOUR OF THE DRAGON, THE [BERKLEY CONAN 1]  HOWARD, ROBERT E.  H,P
HOUR OF THE GATE, THE [SPELLSINGER 2]  FOSTER, ALAN DEAN  C/P:2,P
HOUR OF THE GREMLINS  DICKSON, GORDON R./BOVA, BEN  P
HOUR OF THE HORDE  DICKSON, GORDON R.  C/P,P,H
HOUR OF THE OCTOPUS [D’SHAI 2]  ROSENBERG, JOEL  P
HOUR OF THE OXRUN DEAD, THE  GRANT, CHARLES L.  H,SFWA-P,P
HOUR OF THE PHOENIX, THE  RICHARDS, HENRY  H
HOUR OF THE SCORPION [@TIME WARRIOR #2]  COSTELLO, MATTHEW J.  P
HOUR OF THE WOLF [TERMINATOR 2]  TIEDEMANN, MARK W.  LP
HOUSE BETWEEN THE WORLDS, THE  BRADLEY, MARION ZIMMER  H:2,P,SFWA
HOUSE-BOAT ON THE STYX, A  BANGS, JOHN KENDRICK  H
HOUSE HAUNTED  SARRANTONIO, AL  P
HOUSE IN NOVEMBER, THE  LAUMER, KEITH  H:2,P
HOUSE IN THE HIGH WOOD, THE [WESTERN LIGHTS 2]  BARLOUGH, JEFFREY E.  LP
HOUSE LIKE A LOTUS, A  L’ENGLE, MADELEINE  P
HOUSE NEXT DOOR, THE  SIDDONS, ANNE RIVERS  P
HOUSE OF CAINE, THE  EULO, KEN  SFWA-P
HOUSE OF CARDS [@STAR TREK NEW FRONTIER #1]  DAVID, PETER  C/P,P
HOUSE OF CARDS [@FARSCAPE]  DECANDIDO, KEITH R. A.  C/P,P
HOUSE OF CHAINS [MALAZAN 4]  ERIKSON, STEVEN  H
HOUSE OF CTHULHU, THE [PRIMAL LAND 1]  LUMLEY, BRIAN  SR-VLP
HOUSE OF DARK SHADOWS  ROSS, MARILYN  P
HOUSE OF DEATH [@AVENGER #15]  ROBESON, KENNETH  C/P,P
HOUSE OF DOOM, HOUSE OF DESIRE  WAGNER, SHARON  P
HOUSE OF DOORS, THE  LUMLEY, BRIAN  C/P,SFWA-P,P
HOUSE OF ELNORA GARLAND, THE  LUTTRELL, WANDA  P
HOUSE OF ENTROPY  SHELDON, ROY  SR-P,H
HOUSE OF FULFILMENT: THE ROMANCE OF A SOUL, THE  BECK, L. ADAMS  H
HOUSE OF GAIAN [TIR ALAINN 3]  BISHOP, ANNE  P
HOUSE OF HADES  JACKSON, STEVE  P
HOUSE OF LEAVES  DANIELEWSKI, MARK Z.  LP
HOUSE OF MANY WORLDS, THE  MERWIN, SAM JR.  GN2.12,C/P,P,H
HOUSE OF MOONS [HOUSE OF MOONS 2]  WENTWORTH, K. D.  P
HOUSE OF REEDS [TIME OF THE SIXTH SUN 2]  HARLAN, THOMAS  H
HOUSE OF SCORPIONS [@CHILL #6]  SHERMAN, JORY  P
HOUSE OF SECRETS [WORLD OF DARKNESS]  MOORE, JAMES A./MURPHY, KEVIN  P
HOUSE OF SHADOWS  NORTON, ANDRE/MILLER, PHYLLIS  C/P:2,SFWA-P,P
HOUSE OF SHARDS [@CHARMED]  OSTOW, MICOL  P
HOUSE OF SHARDS [DRAKE MAIJSTRAL 2]  WILLIAMS, WALTER JON  P,SFWA-P
HOUSE OF STAIRS  SLEATOR, WILLIAM  C/P:2,P
HOUSE OF THE DEADLY NIGHTSHADE  LONG, LYDA BELKNAP  P
HOUSE OF THE HUNTER  TAYLOR, FRANK  SR-P
HOUSE OF THE LIONS, THE  STUART, L. T.  P
HOUSE OF THE SCORPION, THE  FARMER, NANCY  H
HOUSE OF THE SECRET, THE  FARRERE, CLAUDE  SR-H
HOUSE OF THE SEVEN GABLES, THE  HAWTHORNE, NATHANIEL  C/P,P
HOUSE OF THE SUN [@SHADOWRUN 17]  FINDLEY, NIGEL D.  P
HOUSE OF THE WOLF [PHOENIX LEGACY 3]  WREN, M. K.  C/P:2,P
HOUSE OF THE WOLFINGS, THE  MORRIS, WILLIAM  LP
HOUSE OF THE WORM, THE  MYERS, GARY  H
HOUSE OF WAR  TARR, JUDITH  LP
HOUSE OF ZEOR [SIME/GEN 1]  LICHTENBERG, JACQUELINE  C/P,P,H,SFWA
HOUSE ON FIRE  OBOLER, ARCH  C/P,P

HOUSE ON LILY STREET, THE [@VANCE INTEGRAL EDITION #11]/ VIEW FROM CHICKWEE’S WINDOW, THE [@VANCE INTEGRAL EDITION #11] VANCE, JACK

SR-H

HOUSE ON THE BORDERLAND, THE  HODGSON, WILLIAM HOPE  C/P,LP,H
HOUSE THAT JACK BUILT, THE [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
HOUSE THAT STOOD STILL, THE  VAN VOGT, A. E.  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 169
TITLE  AUTHOR  CODE
HOUSE WITH A CLOCK IN ITS WALLS, THE  BELLAIRS, JOHN  LP
HOUSEHOLD GODS  TARR, JUDITH/TURTLEDOVE, HARRY  C/P,P
HOUSES OF ISZM, THE  VANCE, JACK  D77525,DF265

HOUSES OF ISZM AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #8] VANCE, JACK

SR-H

HOUSES OF STONE  MICHAELS, BARBARA  P

HOUSES OF THE KZINTI [MAN-KZIN WARS] POURNELLE, JERRY/STIRLING, S. M./

ING, DEAN

LP

HOUSES WITHOUT DOORS  STRAUB, PETER  P
HOUSTON, HOUSTON, DO YOU READ?  TIPTREE, JAMES JR.  SFWA-TD11,TD11
HOW AND WHY WONDER BOOK OF STARS  HOSS, NORMAN  [C/RAN-H]
HOW ARE THE MIGHTY FALLEN  SWANN, THOMAS BURNETT  C/P:2,P
HOW DO YOU GO TO THE BATHROOM IN SPACE?  POGUE, WILLIAM R.  R/REF-P
HOW FEW REMAIN [CONFEDERATE STATES]  TURTLEDOVE, HARRY  C/P:2,P
HOW I SURVIVED MY SUMMER VACATION, VOL. 1 [BUFFY]  TWENTIETH CENTURY FOX  C/PA,PA
HOW LIKE A GOD [PEARL OF IMMORTALITY 1]  CLOUGH, BRENDA W.  H,P
HOW MUCH FOR JUST THE PLANET? [STAR TREK]  FORD, JOHN M.  C/P,P,H
HOW PRECIOUS WAS THAT WHILE  ANTHONY, PIERS  R/REF-H,C/REF-H
HOW THE GODS WOVE IN KYRANNON  MAYHAR, ARDATH  C/P,SFWA,P
HOW THE OTHER HALF LIVES  LOVEGROVE, JAMES  MB001
HOW TO CONTACT SPACE PEOPLE  OWENS, TED  [C/RAN-H]
HOW TO SAVE THE WORLD  SHEFFIELD, CHARLES  SFWA,LP
HOW TO SELL THROUGH MAIL ORDER  GRAHAM, IRVING  [C/RAN-H]
HOW TO WRITE SCIENCE FICTION AND FANTASY  CARD, ORSON SCOTT  R/REF-H
HOWARD COLLECTOR, THE  HOWARD, ROBERT E.  P
HOWARD PHILLIPS LOVECRAFT: DREAMER ON THE NIGHTSIDE  LONG, FRANK BELKNAP  R/REF-H
HOWARD THE DUCK  WEINER, ELLIS  C/P,P
HOWARD WHO?  WALDROP, HOWARD  H,SFWA
HOWLIDAY INN [BUNNICULA 2]  HOWE, JAMES  LP
HOWLING, THE [@HOWLING #1]  BRANDNER, GARY  P
HOWLING DELVE, THE [@FORGOTTEN REALMS DUNGEONS]  JOHNSON, JALEIGH  P
HOWLING II, THE [@HOWLING #2]  BRANDNER, GARY  P
HOWLING III, THE [@HOWLING #3]  BRANDNER, GARY  P
HOWLING MAD  DAVID, PETER  C/P,P
HOWLING STONES, THE [HUMANX]  FOSTER, ALAN DEAN  H:2,P
HOWL’S MOVING CASTLE [CASTLE 1]  JONES, DIANA WYNNE  P
H.P. LOVECRAFT: A LIFE  JOSHI, S. T.  R/REF-H
H.P. LOVECRAFT INSTITUTE, THE  BISCHOFF, DAVID F.  H
HROLF KRAKI’S SAGA  ANDERSON, POUL  C/P:2,SFWA-P,P

HUB: DANGEROUS TERRITORY, THE [TELZEY] SCHMITZ, JAMES H./FLINT, ERIC/

GORDON, GUY

C/P,P

HUBSCHMANN EFFECT, THE  MCMAHON, MICK  P
HUGO, NEBULA AND WORLD FANTASY AWARDS, THE  DEVORE, HOWARD  R/REF-P
HUGO WINNERS 1, THE  ASIMOV, ISAAC  PA,C/PA
HUGO WINNERS 1 AND 2, THE  ASIMOV, ISAAC  HA
HUGO WINNERS 2, THE  ASIMOV, ISAAC  HA,SFWA
HUGO WINNERS 3, THE  ASIMOV, ISAAC  SFWA,HA
HUGO WINNERS 3 BOOK 1, THE  ASIMOV, ISAAC  C/PA,PA
HUGO WINNERS 3 BOOK 2, THE  ASIMOV, ISAAC  C/PA,PA
HUGO WINNERS 4, THE  ASIMOV, ISAAC  HA,SFWA
HUGO WINNERS 5, THE  ASIMOV, ISAAC  HA
HULK [HULK]  DAVID, PETER  C/P,P
HUMAN?  MERRIL, JUDITH  SR-PA,C/PA
HUMAN AND INHUMAN STORIES  SAYERS, DOROTHY L.  PA
HUMAN AND OTHER BEINGS  DEGRAEFF, ALLEN  PA
HUMAN ANGLE, THE  TENN, WILLIAM  C/P:2,P
HUMAN APES, THE  CARLSON, DALE  H
HUMAN BAT, THE  HOME-GALL, EDWARD R.  SR-P
HUMAN BAT VS. THE ROBOT GANGSTER, THE  HOME-GALL, EDWARD R.  SR-P
HUMAN, BEWARE! [CALIFORNIA DRAGONS 2]  GUNNARSSON, THORARINN  P
HUMAN BODY, THE  ASIMOV, ISAAC  [C/RAN-P]
HUMAN CHORD, THE  BLACKWOOD, ALGERNON  SR-H
HUMAN EDGE, THE  DICKSON, GORDON R.  P
HUMAN EQUATION, THE  NOLAN, WILLIAM F.  HA
HUMAN ERROR  PREUSS, PAUL  H:2
HUMAN-MACHINES  SCORTIA, THOMAS N./ZEBROWSKI, GEORGE PA,C/PA
HUMAN NATURE [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
HUMAN RESOURCE [INCONSTANT MOON 1]  ASKEGREN, PIERCE  P
HUMAN TERMITES, THE  KELLER, DAVID H.  SR-LP
HUMAN TO HUMAN [BECOMING ALIEN 3]  ORE, REBECCA  H:2,SFWA-P,P
HUMAN ZERO, THE  GARDNER, ERLE STANLEY  H
HUMAN ZERO, THE  MOSKOWITZ, SAM/ELWOOD, ROGER  PA,C/PA
HUMANITY [ROBOTS AND ALIENS 6]  OLTION, JERRY  P
HUMANITY PRIME  MCALLISTER, BRUCE  C/P:2,P

TITLE  AUTHOR  CODE
HUMANOID TOUCH, THE  WILLIAMSON, JACK  C/P,P,H
HUMANOIDS, THE  BOWEN, CHARLES  [C/RAN-P]
HUMANOIDS, THE  WILLIAMSON, JACK  C/P,GN4.21,H,SR-P,P
HUMANOIDS, THE (EXPANDED)  WILLIAMSON, JACK  C/P,P,H
HUMANS [NEANDERTHAL PARALLAX 2]  SAWYER, ROBERT J.  C/P,P
HUMANS  WESTLAKE, DONALD E.  P
HUMANS KEEP OUT! [PERRY RHODAN 135]  VOLTZ, WILLIAM  SR-LP
HUMANS OF ZIAX II, THE  MORRESSY, JOHN  H
HUMILITY GARDEN [HUMILITY GARDEN 1]  SAVAGE, FELICITY  C/P,P
HUMOR HORROR AND THE SUPERNATURAL  SAKI  P
HUMPTY DUMPTY: AN OVAL  KNIGHT, DAMON  H,LP
100TH MILLENNIUM, THE  BRUNNER, JOHN  DD362
HUNG OUT [STAR OF THE GUARDIANS MAG FORCE 7]  WEIS, MARGARET/PERRIN, DON  P
HUNGER, THE  STRIEBER, WHITLEY  C/P,P,H
HUNGER AND OTHER STORIES, THE  BEAUMONT, CHARLES  C/P,P
HUNGER IN THE SOUL, A  RESNICK, MIKE  LP
HUNGRE LIKE NO OTHER, A [IMMORTALS AFTER DARK 1]  COLE, KRESLEY  H
HUNGRY MOON, THE  CAMPBELL, RAMSEY  SFWA-P
HUNT, THE  KUNSTLER, JAMES HOWARD  SFWA-P
HUNT, THE [@LAWS OF THE BLOOD 1]  SIZEMORE, SUSAN  C/P,P
HUNT FOR DEFAY, THE [WARHAMMER INQUISITOR ASCENDANT]  ABNETT, DAN  CX
HUNT FOR RED OCTOBER, THE  CLANCY, TOM  C/P:2,P
HUNT FOR THE METEOR, THE  VERNE, JULES  H,P
HUNT THE BEAST DOWN [@SPECTROS #2]  WINTERS, LOGAN  P
HUNT THE HEAVENS [@SHADOW WARRIOR #2]  BUNCH, CHRIS  P
HUNTED, THE [VAMPIRE HUNTRESS 3]  BANKS, L. A.  LP
HUNTED [LEAGUE OF PEOPLES 3]  GARDNER, JAMES ALAN  H:2,P
HUNTED, THE  PTACEK, KATHRYN  SFWA
HUNTER, THE [HUNTER 1]  ESTES, ROSE  C/P,P
HUNTER OF THE LIGHT  ARATYR, RISA  P
HUNTER OF WORLDS [HANAN REBELLION 2]  CHERRYH, C. J.  C/P,P,H
HUNTER ON ARENA, THE [HUNTER 2]  ESTES, ROSE  P
HUNTER OUT OF TIME, THE  FOX, GARDNER F.  C/P:2,P
HUNTER RETURNS, THE  DRAKE, DAVID/KJELGAARD, JIM  C/P:2,SFWA-P,P
HUNTER/VICTIM  SHECKLEY, ROBERT  P
HUNTER VICTORIOUS, THE [HUNTER 3]  ESTES, ROSE  P
HUNTERS, THE [@BATTLETECH #35/TWILIGHT OF THE CLANS 3]  GRESSMAN, THOMAS S.  C/P,P
HUNTERS, THE  LOVEJOY, JACK  C/P,P
HUNTERS, THE  WETANSON, BURT/HOOBLER, THOMAS  H,SFWA,P
HUNTER’S DEATH [AVERALAAN 2]  WEST, MICHELLE  C/P,P
HUNTER’S FIRE, A  DALE, FLOYD D.  P
HUNTERS’ HAUNT, THE [OMAR 2]  DUNCAN, DAVE  C/P:2,P
HUNTER’S OATH [AVERALAAN 1]  WEST, MICHELLE  C/P,P
HUNTERS OF DUNE [DUNE 7]  HERBERT, BRIAN/ANDERSON, KEVIN J.  H
HUNTERS OF GOR [GOR 8]  NORMAN, JOHN  C/P:2,P
HUNTERS OF JUNDAGAI, THE [KEYS TO THE DIMENSIONS 6]  BULMER, KENNETH  D68310
HUNTERS OF THE DARK SEA  ODOM, MEL  H
HUNTERS OF THE DEEP [@MECHWARRIOR DARK AGE/BATTLETECH]  BILLS, RANDALL N.  P
HUNTERS OF THE RED MOON  BRADLEY, MARION ZIMMER  C/P:2,P
HUNTER’S PLANET [@ALIENS VS. PREDATOR 2]  BISCHOFF, DAVID F.  C/P,P
HUNTING GROUND [@BLACK OAK 4]  GRANT, CHARLES L.  P
HUNTING ON KUNDERER  BARTON, WILLIAM  D48245
HUNTING PARTY [HERIS SERRANO 1]  MOON, ELIZABETH  C/P:2,SFWA-P,P
HUNTING THE CORRIGAN’S BLOOD  LISLE, HOLLY  P
HUNTING THE GHOST DANCER  ATTANASIO, A. A.  H,P
HUNTRESS AND THE SPHINX, THE [@XENA]  EMERSON, RU  P
HUNTSMAN, THE [HUNTSMAN 1]  HILL, DOUGLAS  H
HUON OF THE HORN  NORTON, ANDRE  C/P:2,P
HUROK OF THE STONE AGE [ERIC CARSTAIRS 3]  CARTER, LIN  P
HURRICANE HEX [@CHARMED]  GALLAGHER, DIANA G.  P
HURRICANE MOON  LATNER, ALEXIS GLYNN  LP
HURRICANE’S CHILDREN, THE  CARMER, CARL  H
HUSBAND FOR KUTANI, A  OWEN, FRANK  SR-H
HUTT GAMBIT, THE [@STAR WARS HAN SOLO 2]  CRISPIN, A. C.  P
HUYSMAN’S PETS  WILHELM, KATE  H,P
HYACINTHS  YARBRO, CHELSEA QUINN  H,SFWA
HYADES CONTACT, THE  WHITEFORD, WYNNE  C/P,P
HYBRID, THE  JAKES, JOHN  C/P:2,P
HYBRID, THE  VANCE, STEVE  P
HYBRIDS [NEANDERTHAL PARALLAX 3]  SAWYER, ROBERT J.  H
HYDRA MONSTER, THE [PHANTOM 8]  SHAWN, WILLIAM S.  P
HYDRABYSS RED [@TIMEQUEST #2]  TEDFORD, WILLIAM  P
HYDROGEN STEEL  BEDFORD, K. A.  LP
HYDRONAUTS, THE  BIEMILLER, CARL L.  H

TITLE  AUTHOR  CODE
HYDROSPHERE  MERAK, A. J.  H
HYMN BEFORE BATTLE, A [POSLEEN 1]  RINGO, JOHN  C/P,P
HYPERBOREA  SMITH, CLARK ASHTON  SR-P,C/P:2
HYPERION [HYPERION 1]  SIMMONS, DAN  C/P:2,P
HYPERION CANTOS [HYPERION 1,2]  SIMMONS, DAN  H:3
HYPERSEUCHE, DIE [@PERRY RHODAN SILBERBANDE #69]  PABEL MOEWIG KG  FPR
HYPERSPACE  FANTHORPE, R. L.  SR-P
HYPERTHOUGHT  BUCKNER, M. M.  P

TITLE  AUTHOR  CODE
I, ALEPPO  SOHL, JERRY  SR-P
I — ALIEN  REAVES, MICHAEL  C/P:2,P
I, ALIEN  RESNICK, MIKE  PA
I AM A BARBARIAN  BURROUGHS, EDGAR RICE  C/P,P,H
I AM A DALEK [DOCTOR WHO]  ROBERTS, GARETH  LP
I AM ALIVE AND YOU ARE DEAD  CARRERE, EMMANUEL  R/REF-H
I AM FRANKENSTEIN  ANDERSSON, C. DEAN  P
I AM JADE FALCON [@BATTLETECH 17]  THURSTON, ROBERT  P
I AM LEGEND  MATHESON, RICHARD  H:2,SR-P
I AM MORDRED  SPRINGER, NANCY  P
"I AM NOT A CROOK"  BUCHWALD, ART  [C/RAN-P]
I AM NOT SPOCK  NIMOY, LEONARD  R/REF-P,C/REF-P
I AM NOT THE OTHER HOUDINI  CONNER, MIKE  H,SFWA,P
I AM SPOCK  NIMOY, LEONARD  R/REF-H,C/REF-H:2
I AM THINKING OF MY DARLING  MCHUGH, VINCENT  SR-P,H
I AM VIDOCQ  MCCONNOR, VINCENT  SFWA-P
I. ASIMOV  ASIMOV, ISAAC  R/REF-H
I BEEN THERE BEFORE  CARKEET, DAVID  H
I DARE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  LP
I DIED HERE  SHEA, GEORGE  SR-P
I DREAM OF JEANNIE  BREWSTER, DENNIS  LP
I FIGHT FOR MARS  TUBB, E. C.  LP
I FOUND CLEOPATRA  KELLEY, THOMAS P.  SR-LP
I HAVE NO MOUTH AND I MUST SCREAM  ELLISON, HARLAN  C/P:2,P
I HOPE I SHALL ARRIVE SOON  DICK, PHILIP K.  H,SFWA,P
I INSIDE, THE  FOSTER, ALAN DEAN  C/P:2,P
I, JEDI [@STAR WARS]  STACKPOLE, MICHAEL A.  C/P,P,H
I LEFT MY SNEAKERS IN DIMENSION X [ROD ALLBRIGHT 2]  COVILLE, BRUCE  LP
I LIVE WITH YOU  EMSHWILLER, CAROL  LP
I LOVE GALESBURG IN THE SPRINGTIME  FINNEY, JACK  SR-P
I, MARTHA ADAMS  WINSLOW, PAULINE GLEN  C/P,P
I, Q [STAR TREK TNG]  DE LANCIE, JOHN/DAVID, PETER  P
I, ROBOT  ASIMOV, ISAAC  C/P,SR-P,P,H:3
I, ROBOT: THE ILLUSTRATED SCREENPLAY  ELLISON, HARLAN/ASIMOV, ISAAC  LP
I, SAID THE FLY  SHEA, MICHAEL  LP
I SHUDDER AT YOUR TOUCH  SLUNG, MICHELE  C/PA,PA
I SING THE BODY ELECTRIC!  BRADBURY, RAY  C/P:2,P,H
I SPEAK FOR EARTH  WOODCOTT, KEITH  DD497
I SPY  STATTEN, VARGO  SR-P
I, STRAHD [@RAVENLOFT]  ELROD, P. N.  C/P:2,P
I, THE MACHINE  FAIRMAN, PAUL W.  C/P:2,P
I, THE SUN [HATTI EMPIRE]  MORRIS, JANET  P
I, VAMPIRE  ROMKEY, MICHAEL  P
I, VAMPIRE  SCOTT, JODY  C/P:2,P
I WANT THE STARS  PURDOM, TOM  DF289
I, WEAPON  RUNYON, CHARLES W.  C/P,SFWA,P,H
I WHO HAVE NEVER KNOWN MEN  HARPMAN, JACQUELINE  P
I WILL FEAR NO EVIL  HEINLEIN, ROBERT A.  H:2,P
I, ZOMBIE  SELBY, CURT  C/P,P
IAPETUS  KIRBY, WILLIAM S.  P
IBIS  STEELE, LINDA  C/P,P
ICARUS DESCENDING  HAND, ELIZABETH  C/P,P
ICARUS DESCENDING (PROOF)  HAND, ELIZABETH  SR-LP
ICARUS HUNT, THE  ZAHN, TIMOTHY  H,P
ICE!  FEDERBUSH, ARNOLD  P
ICE  KAVAN, ANNA  H,P
ICE AND FIRE [@DEATHLANDS]  AXLER, JAMES  P
ICE AND IRON  TUCKER, WILSON  H:2,P,SFWA
ICE BEAST, THE  JAVOR, FRANK A.  C/P,P
ICE CROWN  NORTON, ANDRE  H:2,P
ICE DRAGON, THE [@DRAGON KNIGHTS #3]  D’ARC, BIANCA  LP
ICE DRAGON [@BLADE #10]  LORD, JEFFREY  C/P,P
ICE DRAGON, THE  MARTIN, GEORGE R. R.  H
ICE FROM SPACE [TALES OF TOMORROW 4]  HOUSTON, DAVID  P
ICE IS COMING, THE  WRIGHTSON, PATRICIA  C/P:2,P
ICE KING, THE  ROHAN, MICHAEL SCOTT/SCOTT, ALLAN  P
ICE MAGE [TIGUAFAYA 1]  GRAY, JULIA  P
ICE PEOPLE, THE  BARJAVEL, RENE  H:2,P
ICE PEOPLE, THE  GEE, MAGGIE  LP
ICE PRISON  SKY, KATHLEEN  SR-P
ICE PROPHET [ICE 1]  FORSTCHEN, WILLIAM R.  C/P:2,P
ICE SCHOONER, THE  MOORCOCK, MICHAEL  H:2,SFWA,P
ICE STATION ZEBRA  MACLEAN, ALISTAIR  P
ICE TRAP [STAR TREK]  GRAF, L. A.  C/P:2,P

TITLE  AUTHOR  CODE
ICEBERG [DOCTOR WHO NEW ADVENTURES]  BANKS, DAVID  C/P,P
ICEBERG EXPRESS, THE  CORY, DAVID  H
ICEBLOOD [@OUTLANDERS]  AXLER, JAMES  P
ICEBONES [MAMMOTHS]  BAXTER, STEPHEN  H,LP
ICEBORN  BENFORD, GREGORY/CARTER, PAUL A.  TD14,SFWA-TD14
ICEBREAKER [JAMES BOND 17]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
ICED ON ARAN [DREAMLAND 4]  LUMLEY, BRIAN  P
ICEDRAGON [DRAGONREALM 2]  KNAAK, RICHARD A.  C/P,P
ICEFALCON’S QUEST [DARWATH 5]  HAMBLY, BARBARA  C/P:2,P,H
ICEHENGE  ROBINSON, KIM STANLEY  C/P:2,SFWA-P,LP
ICEMAN  FELICE, CYNTHIA  C/P:2,P
ICEQUAKE  KILIAN, CRAWFORD  P
ICEQUAKE  SPENCER, JOHN R.  P
ICERIGGER [HUMANX ICERIGGER 1]  FOSTER, ALAN DEAN  C/P,P,H
ICETOWERS  MCGEARY, DUNCAN  P
ICEWORLD  CLEMENT, HAL  SR-H,C/P:2,P
ICEWORLD CONNECTION, THE [@SPACEWAYS #11]  CLEVE, JOHN  P
IDEAL COMMONWEALTHS  MORLEY, HENRY  SR-HA
IDEAL WAR [@BATTLETECH 9]  KUBASIK, CHRIS  C/P,P
IDENTITY MATRIX, THE  CHALKER, JACK L.  C/P:2,P
IDENTITY PLUNDERERS, THE  HAIBLUM, ISIDORE  P
IDENTITY SEVEN  LORY, ROBERT  C/P,P
IDES OF TOMORROW, THE  CARR, TERRY  SR-HA,HA
IDIC EPIDEMIC, THE [STAR TREK]  LORRAH, JEAN  C/P:2,P
IDLE PLEASURES  EFFINGER, GEORGE ALEC  P
IDLEWILD [IDLEWILD 1]  SAGAN, NICK  H
IDLEWILD (PROOF) [IDLEWILD 1]  SAGAN, NICK  SR-LP
IDOL FROM PASSA, THE [@PERRY RHODAN #98]  ACE  PA
IDOL OF THE FLIES, THE  RICE, JANE  SR-H
IDOLIN  BUDZ, MARK  P
IDORU  GIBSON, WILLIAM  H:2,P
IDUNA’S UNIVERSE [DUMAREST 21]  TUBB, E. C.  C/P,P
IDYLL OF THE WHITE LOTUS, THE  COLLINS, MABEL  P
IDYLLS OF THE QUEEN, THE  KARR, PHYLLIS ANN  C/P,P
IF ALL ELSE FAILS...  STRETE, CRAIG  SFWA
IF ALL THE REBELS DIE  SOUTHWELL, SAMUEL  P
IF AT FAUST YOU DON’T SUCCEED [AZZIE 2]  ZELAZNY, ROGER/SHECKLEY, ROBERT  P,LP
IF AT FAUST YOU DON’T SUCCEED (PROOF) [AZZIE 2]  ZELAZNY, ROGER/SHECKLEY, ROBERT  SR-LP
IF I NEVER GET BACK [SAMUEL CLEMENS FOWLER 1]  BROCK, DARRYL  H
IF I PAY THEE NOT IN GOLD  ANTHONY, PIERS/LACKEY, MERCEDES  H,SFWA,P
IF I WERE AN EVIL OVERLORD  GREENBERG, MARTIN H./DAVIS, RUSSELL  PA
IF READER OF SCIENCE FICTION, THE  POHL, FREDERIK  C/PA,PA
SECOND IF READER OF SCIENCE FICTION, THE  POHL, FREDERIK  C/PA,PA
IF THE SOUTH HAD WON THE CIVIL WAR  KANTOR, MACKINLAY  P
IF THE STARS ARE GODS  BENFORD, GREGORY/EKLUND, GORDON  C/P:2,P,H
IF THIS GOES ON  NUETZEL, CHARLES  SR-PA,C/PA
IF WHISPERS CALL [DARK MATTER 2]  BASSINGTHWAITE, DON  P
IF WISHES WERE HORSES  MCCAFFREY, ANNE  H
IF YOU BELIEVE THE SOLDIERS  CORDELL, ALEXANDER  H,SFWA
IF YOU COULD SEE ME NOW  STRAUB, PETER  C/P,P
IGNITING THE REACHES [REACHES 1]  DRAKE, DAVID  C/P,P,H
IGNITION  ANDERSON, KEVIN J./BEASON, DOUG  H,P
IGNORANT ARMIES [WARHAMMER]  PRINGLE, DAVID  SR-LP
IK, ROBOT  ASIMOV, ISAAC  FP
ILE DES DAUPHINS, L’  CLARKE, ARTHUR C.  FH
ILENE, THE SUPERSTITIOUS [@PHENWICK #14]  KIMBROUGH, KATHERYN  P
ILIUM [ILIUM 1]  SIMMONS, DAN  H
ILL FATE MARSHALLING, AN [DREAD EMPIRE 4/RAGNARSON 6]  COOK, GLEN  C/P,SFWA-P,P
ILL-MADE KNIGHT, THE  WHITE, T. H.  H
ILL-MADE MUTE, THE [BITTERBYNDE 1]  DART-THORNTON, CECILIA  C/P:2,P
ILL MET BY MOONLIGHT [WILLIAM SHAKESPEARE 1]  HOYT, SARAH A.  P
ILL MET BY MOONLIGHT [SCEPTER’D ISLE 2]  LACKEY, MERCEDES/GELLIS, ROBERTA  H
ILL MET BY MOONLIGHT (PROOF) [SCEPTER’D ISLE 2]  LACKEY, MERCEDES/GELLIS, ROBERTA  SR-LP
ILL MET IN LANKHMAR [FAFHRD 1,2]  LEIBER, FRITZ  H,SFWA-TD19,TD19
ILL MET IN LANKHMAR (AND OTHER STORIES) [FAFHRD]  LEIBER, FRITZ  P
ILL MET IN LANKHMAR (SHORT STORY) [FAFHRD]  LEIBER, FRITZ  H
ILL WIND  ANDERSON, KEVIN J./BEASON, DOUG  H:2,P
ILL WIND [WEATHER WARDEN 1]  CAINE, RACHEL  P
I’LL ZAP MANHATTAN [@SABRINA THE TEENAGE WITCH #18]  ODOM, MEL  P
ILLEARTH WAR, THE [THOMAS COVENANT 2]  DONALDSON, STEPHEN R.  C/P,H,LP
ILLEGAL ALIEN  LUCENO, JAMES  C/P:2,P
ILLEGAL ALIEN  SAWYER, ROBERT J.  C/P,P
ILLEGAL ALIEN [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE/PERRY, ROBERT  C/P:2,P
ILLEGAL ALIENS  POLLOTTA, NICK/FOGLIO, PHIL  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 175
TITLE  AUTHOR  CODE
ILLEGAL REBIRTH OF BILLY THE KID, THE  ORE, REBECCA  H,P
ILLUMINATI PAPERS, THE  WILSON, ROBERT ANTON  LP
ILLUMINATION [ILLUMINATION 1]  MCGARRY, TERRY  H
ILLUMINATUS! TRILOGY, THE [ILLUMINATUS! 1,2,3]  SHEA, ROBERT/WILSON, ROBERT ANTON  H:3,LP
ILLUSION  VOLSKY, PAULA  C/P,P,H
ILLUSIONISTS, THE  MILLER, FAREN  P
ILLUSIONS OF VICTORY [@BATTLETECH 47]  COLEMAN, LOREN L.  P
ILLUSTRATED BOOK OF SCIENCE FICTION LISTS, THE  ASHLEY, MIKE  R/REF-H
ILLUSTRATED DUNE, THE [DUNE 1]  HERBERT, FRANK  LP
ILLUSTRATED HARLAN ELLISON, THE  ELLISON, HARLAN  CX
ILLUSTRATED HITCHHIKER’S GUIDE TO THE GALAXY, THE [HITCHHIKERS] ADAMS, DOUGLAS  A
ILLUSTRATED MAN, THE  BRADBURY, RAY  C/P,H,SR-P
ILLUSTRATED ROGER ZELAZNY, THE  ZELAZNY, ROGER/MORROW, GRAY  CX
ILLYRIAN ADVENTURE, THE  ALEXANDER, LLOYD  P
ILLYWHACKER  CAREY, PETER  H
ILSE WITCH [@VOYAGE OF THE JERLE SHANNARA #1/SHANNARA JERLE]  BROOKS, TERRY  H
ILUSTRE FAMILIA ANDROIDE, LA  DESLEAL, ALVARO MENEN  FP
IM MAHLSTROM DER STERNE [@PERRY RHODAN SILBERBANDE #77]  PABEL MOEWIG KG  FPR
IMAGE [@ANGEL]  ODOM, MEL  P
IMAGE OF THE BEAST, THE [HERALD CHILDE 1]  FARMER, PHILIP JOSE  C/P,SR-L
IMAGE OF THE BEAST [HERALD CHILDE 1,2]/BLOWN [HERALD CHILDE 1,2] FARMER, PHILIP JOSE  C/P:2,P
IMAGE OF VOICES, AN [@WINDHOVER TAPES #1]  NORWOOD, WARREN  C/P,P
IMAGES OF TOMORROW  NUETZEL, CHARLES  SR-P
IMAGINARY LANDS  MCKINLEY, ROBIN  C/PA,PA
IMAGINARY MAGNITUDE  LEM, STANISLAW  H,P
IMAGINARY WORLDS  CARTER, LIN  R/REF-P,C/REF-P
IMAGINATION FULLY DILATED: SCIENCE FICTION  KRUGER, ROBERT/SWENSON, PATRICK  LPA
IMAGINATION UNLIMITED  BLEILER, EVERETT F./DIKTY, T. E.  C/PA,PA
IMAGINATIVE SEX  NORMAN, JOHN  [C/RAN-P:2]
IMAGINED SLIGHTS  LOVEGROVE, JAMES  LP
IMAGININGS  DECANDIDO, KEITH R. A.  LPA
IMAGO [XENOGENESIS 3]  BUTLER, OCTAVIA E.  C/P,P,H
IMAGO SEQUENCE, THE  BARRON, LAIRD  H
IMAJICA  BARKER, CLIVE  H:2,LP
IMARO [IMARO 1]  SAUNDERS, CHARLES R.  C/P,P
IMBALANCE [STAR TREK TNG]  MITCHELL, V. E.  C/P:2,P
IMMACULATE, THE  MORRIS, MARK  P
IMMENSE JOURNEY, THE  EISELEY, LOREN  [C/RAN-P:2]
IMMINENT CRISIS [@MECHWARRIOR]  BILLS, RANDALL N.  P
IMMODEST PROPOSALS [COMPLETE TENN 1]  TENN, WILLIAM  H
IMMORTAL  DANN, JACK  HA,SFWA
IMMORTAL [BUFFY]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
IMMORTAL, THE  GUNN, JAMES E.  P

IMMORTAL BAD BOYS YORK, REBECCA/LAUREY, ROSEMARY/

THOMAS-SUNDSTROM, LINDA

LP

IMMORTAL COIL [STAR TREK TNG]  LANG, JEFFREY  P
IMMORTAL GAME [QUEST TRIAD 3]  NILES, DOUGLAS  P
IMMORTAL GIRL, THE  RUCK, BERTA  H
IMMORTAL IN DEATH [EVE DALLAS]  ROBB, J. D.  P
IMMORTAL OF WORLD’S END, THE [GONDWANE 3]  CARTER, LIN  C/P,P
IMMORTAL STORM, THE  MOSKOWITZ, SAM  R/REF-H
IMMORTAL UNKNOWN, THE [@PERRY RHODAN #13]  ACE  PA,C/PA
IMMORTALIS [DEMONWARS II 3]  SALVATORE, R. A.  H
IMMORTALITY, INC.  SHECKLEY, ROBERT  C/P:2,P
IMMORTALITY OPTION, THE  HOGAN, JAMES P.  C/P,P,H
IMMORTALITY: THE SCIENTIFIC EVIDENCE  SMITH, ALSON J.  [C/RAN-P]
IMMORTALS, THE  BARJAVEL, RENE  C/P,P
IMMORTALS, THE  BRETT, LEO  SR-P
IMMORTALS  DANN, JACK/DOZOIS, GARDNER  PA
IMMORTALS, THE  GARNER, ROLF  SR-P
IMMORTALS, THE  GUNN, JAMES E.  C/P:2,P
IMMORTALS, THE  HICKMAN, TRACY  H,P
IMMORTALS, THE [IMMORTALS (PIERCE) 1,2,3,4]  PIERCE, TAMORA  H
IMPACT!  FODOR, R. V./TAYLOR, G. J.  P
IMPACT-20  NOLAN, WILLIAM F.  C/P:2,P
IMPACT PARAMETER  LANDIS, GEOFFREY A.  H:2
IMPAKTO  CALDER, RICHARD  P
IMPERATOR PLOT, THE  SPRUILL, STEVEN G.  P,SFWA,SFWA-P
IMPERIAL BOUNTY [SAM MCCADE 2]  DIETZ, WILLIAM C.  P
IMPERIAL EARTH  CLARKE, ARTHUR C.  C/P,P,H,SR-H
IMPERIAL LADY  NORTON, ANDRE/SHWARTZ, SUSAN  C/P,SFWA,SFWA-P,P
IMPERIAL MOON [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  C/P,P
IMPERIAL STARS, THE [DOMINIC FLANDRY]  ANDERSON, POUL  H
IMPERIAL STARS [FAMILY D’ALEMBERT #1]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P

TITLE  AUTHOR  CODE
IMPERIO DOS MUTANTES, O  WUL, STEFAN  FP
IMPERIUM AFIRE [@PANGAEA #2]  MASON, LISA  P
IMPERIUM GAME, THE  WENTWORTH, K. D.  P
IMPERIUM WITHOUT END [@PANGAEA #1]  MASON, LISA  P
IMPETUS OF WAR [@BATTLETECH 30]  PARDOE, BLAINE LEE  P
IMPLOSION  JONES, D. F.  H:2,P
IMPLOSION EFFECT, THE  PAULSEN, GARY  P
IMPOSSIBLE?  JANIFER, LAURENCE M.  C/P:2,P
IMPOSSIBLE BIRD, THE  O’LEARY, PATRICK  LP
IMPOSSIBLE MAN, THE  BALLARD, J. G.  C/P:2,P
IMPOSSIBLE ODDS [KING’S BLADES]  DUNCAN, DAVE  H,P
IMPOSSIBLE PLACES  FOSTER, ALAN DEAN  P
IMPOSSIBLE POSSIBILITIES  PAUWELS, LOUIS/BERGIER, JACQUES  [C/RAN-P]
IMPOSSIBLE THINGS  WILLIS, CONNIE  C/P,P,H
IMPOSSIBLE WORLD, THE  BINDER, EANDO  C/P:2,P
IMPOSSIBLES, THE  PHILLIPS, MARK  C/P:2,P
IMPOSTOR  FREIREICH, VALERIE J.  C/P,P
IMPRESSIONS [@ANGEL]  DURGIN, DORANNA  P
IMRAM  BARNWELL, WILLIAM  P
IMZADI [STAR TREK TNG]  DAVID, PETER  H:2,P
IMZADI FOREVER [STAR TREK TNG]  DAVID, PETER  LP
IN A DARK DREAM  GRANT, CHARLES L.  P,SFWA-P
...IN A WORLD NOT OF HIS OWN MAKING  PAPSON, STEPHEN  LP
IN ALIEN FLESH  BENFORD, GREGORY  C/P,SFWA-P,P,H
IN ALIEN HANDS [@QUEST FOR TOMORROW 2]  SHATNER, WILLIAM  P
IN AND OUT OF QUANDRY  HOFFMAN, LEE  SR-DNESFA2
IN ANOTHER COUNTRY [VINTAGE SEASON 2]  SILVERBERG, ROBERT  SFWA-TD18,TD18
IN BAD COMPANY [COMPANY 3,4]  BAKER, KAGE  H
IN BETWEEN DRAGONS  KANDEL, MICHAEL  C/P:2,P
IN CAMELOT’S SHADOW  ZETTEL, SARAH  LP
IN CONQUEST BORN  FRIEDMAN, C. S.  C/P:2,P
IN CONQUEST BORN (REVISED)  FRIEDMAN, C. S.  P
IN DARK PLACES  RUSSELL, JOHN  H
IN DEATH GROUND [STARFIRE 3]  WEBER, DAVID/WHITE, STEVE  C/P,P
IN DEEP  KNIGHT, DAMON  C/P:2,P
IN DEFENCE OF FANTASY  SWINFEN, ANN  R/REF-P
IN DREAMS AWAKE  FIEDLER, LESLIE A.  PA
IN ENDLESS TWILIGHT [FOREVER HERO 3]  MODESITT, L. E. JR.  C/P,P
IN ENEMY HANDS [HONOR HARRINGTON 7]  WEBER, DAVID  H:2,P
IN FLUID SILENCE [DARK MATTER 3]  TIRPA, G. W.  P
IN FURY BORN  WEBER, DAVID  H
IN GHOSTLY JAPAN  HEARN, LAFCADIO  P
IN GREEN’S JUNGLES [BOOK OF THE SHORT SUN 2]  WOLFE, GENE  LP
IN HIGH PLACES  HAILEY, ARTHUR  H
IN HIGH PLACES [CROSSTIME TRAFFIC 3]  TURTLEDOVE, HARRY  H
IN HIS IMAGE [@CHRIST CLONE #1]  BEAUSEIGNEUR, JAMES  P
IN HOLLOW HOUSES [DARK MATTER 1]  BRAUNBECK, GARY A.  P
IN HOT BLOOD  POPESCU, PETRU  P
IN IRON YEARS  DICKSON, GORDON R.  C/P,P,H,SFWA
IN JOY STILL FELT (AUTOBIOGRAPHY 1954–1978)  ASIMOV, ISAAC  R/REF-H,C/REF-H
IN LANDS THAT NEVER WERE  VAN GELDER, GORDON  LPA
IN LEGEND BORN  RESNICK, LAURA  H:2
IN LUNACY [@MUTANT CHRONICLES APOSTLE OF INSANITY 1]  WU, WILLIAM F.  P
IN MAYAN SPLENDOR  LONG, FRANK BELKNAP  SR-H
IN MEMORY OF SARAH BAILEY  COOPER, LOUISE  SR-P
IN MEMORY YET GREEN (AUTOBIOGRAPHY 1920–1954)  ASIMOV, ISAAC  R/REF-H
IN MY PARENTS’ BEDROOM  MALZBERG, BARRY N.  SR-P
IN NORTHERN TWILIGHT [@EXALTED #3]  HARTLEY, JESS  P
IN OTHER WORLDS [RADIX 2]  ATTANASIO, A. A.  C/P,P
IN OUR HANDS, THE STARS  HARRISON, HARRY  H
IN QUEST OF QALARA [@SPACEWAYS #9]  CLEVE, JOHN  P
IN REALMS UNKNOWN  BELL, ROBERT  H
IN SEARCH OF ANCIENT MYSTERIES  LANDSBURG, ALAN/LANDSBURG, SALLY  [C/RAN-P]
IN SEARCH OF DRACULA  MCNALLY, RAYMOND T./FLORESCU, RADU  R/REF-P,C/REF-P
IN SEARCH OF FRANKENSTEIN  FLORESCU, RADU  R/REF-P
IN SEARCH OF MOM [@PHOTON #3]  PETERS, DAVID  P
IN SEARCH OF THE UNKNOWN  CHAMBERS, ROBERT W.  LP
IN SEARCH OF WONDER  KNIGHT, DAMON  R/REF-H,C/REF-H:2
IN SILENCE SEALED  PTACEK, KATHRYN  SFWA-P
IN SOLITARY  KILWORTH, GARRY  P
IN STONE’S GRASP [KEVLA OF ARUKAN 2]  GOLDEN, CHRISTIE  LP
IN SYLVAN SHADOWS [@FORGOTTEN REALMS CLERIC #2]  SALVATORE, R. A.  P
IN THE BEGINNING [@BABYLON 5]  DAVID, PETER  C/P,P
IN THE BEGINNING  DOUGLAS, NORMAN  H

TITLE  AUTHOR  CODE
IN THE BEGINNING  MULLER, JOHN E.  SR-P
IN THE BEGINNING  SILVERBERG, ROBERT  SR-HA
IN THE BLOOD  COLLINS, NANCY A.  P
IN THE BONE  DICKSON, GORDON R.  P,SFWA-P
IN THE CAVES OF EXILE [NEDAO 2]  EMERSON, RU  P,SFWA-P
IN THE CENTER OF THE GALAXY (NOVEL) [PERRY RHODAN]  ACE  PA
IN THE CHINKS OF THE WORLD MACHINE  LEFANU, SARAH  R/REF-P
IN THE CIRCLE OF TIME  ANDERSON, MARGARET J.  P
IN THE CLAWS OF THE TIGER [@EBERRON WAR-TORN #3]  WYATT, JAMES  P
IN THE COMPANY OF MIND  PIZIKS, STEVEN  C/P:2,P
IN THE COMPANY OF OGRES  MARTINEZ, A. LEE  LP
IN THE COMPANY OF OTHERS  CZERNEDA, JULIE E.  H,P
IN THE COUNTRY OF THE BLIND  FLYNN, MICHAEL  C/P:2,SFWA-P,P
IN THE COUNTRY OF THE BLIND (REVISED)  FLYNN, MICHAEL  H,P
IN THE CUBE  SMITH, DAVID ALEXANDER  C/P,P,H
IN THE DAYS OF DREAD  MARKS, ANTON  LP
IN THE DAYS OF THE COMET  WELLS, H. G.  C/P,P,H
IN THE DRIFT  SWANWICK, MICHAEL  C/P:2,P
IN THE EMPIRE OF SHADOW [THREE WORLDS 2]  WATT-EVANS, LAWRENCE  LP
IN THE ENCLOSURE  MALZBERG, BARRY N.  P
IN THE EYE OF HEAVEN  KECK, DAVID  H
IN THE EYE OF TERROR [WARHAMMER BLOODQUEST 2]  RENNIE, GORDON/MACNEIL, COLIN  CX
IN THE FACE OF DEATH [@DARKMAN #4]  BOYLL, RANDALL  P
IN THE FACE OF MY ENEMY  DELANEY, JOSEPH H.  C/P:2,P
IN THE FIELD OF FIRE  DANN, JEANNE VAN BUREN/DANN, JACK  PA,SFWA-P,SFWA
IN THE FLESH  BARKER, CLIVE  H:2,P
IN THE FOREST OF FORGETTING  GOSS, THEODORA  H
IN THE FORESTS OF SERRE  MCKILLIP, PATRICIA A.  H
IN THE GARDEN OF IDEN [COMPANY 1]  BAKER, KAGE  C/P:2,P
IN THE GREEN STAR’S GLOW [GREEN STAR 5]  CARTER, LIN  P
IN THE GRIP OF TERROR  CONKLIN, GROFF  PA
IN THE HALL OF THE DRAGON KING [@DRAGON KING #1]  LAWHEAD, STEPHEN R.  C/P,P
IN THE HALL OF THE MARTIAN KING [JAK JINNAKA 3]  BARNES, JOHN  C/P,P
IN THE HAND OF THE GODDESS [SONG OF THE LIONESS 2]  PIERCE, TAMORA  P
IN THE HANDS OF GLORY  EISENSTEIN, PHYLLIS  C/P,P
IN THE HEART OF DARKNESS [BELISARIUS 2]  FLINT, ERIC/DRAKE, DAVID  C/P,P
IN THE HEART OF THE VALLEY OF LOVE  KADOHATA, CYNTHIA  H
IN THE ICE KING’S PALACE [WORLD IN AMBER 2]  ORR, A.  H,SFWA
IN THE KEEP OF TIME  ANDERSON, MARGARET J.  P
IN THE KINGDOM OF THE BEASTS [DIES IRAE 2]  STABLEFORD, BRIAN M.  C/P:2,P
IN THE KING’S SERVICE [@DERYNI]  KURTZ, KATHERINE  P
IN THE LIGHT OF SIGMA DRACONIS [@WOMAN OF THE IRON PEOPLE #1]  ARNASON, ELEANOR  P
IN THE MIST AND OTHER UNCANNY ENCOUNTERS  WALTER, ELIZABETH  H
IN THE MOOD [MILLENNIUM QUARTET 2]  GRANT, CHARLES L.  H
IN THE MOONS OF BOREA  LUMLEY, BRIAN  C/P,P
IN THE MORNING OF TIME  KING, CYNTHIA  H,P
IN THE MOTHER’S LAND  VONARBURG, ELISABETH  P
IN THE NAME OF HONOR [STAR TREK]  WARD, DAYTON  P
IN THE NET OF DREAMS [DREAMLAND 1]  SIMMONS, WM. MARK  P
IN THE NIGHT GARDEN [@ORPHAN’S TALES #1]  VALENTE, CATHERYNNE M  LP
IN THE NIGHT ROOM  STRAUB, PETER  H
IN THE OCEAN OF NIGHT [GALACTIC CENTER 1]  BENFORD, GREGORY  C/P,P,H
IN THE POCKET AND OTHER S-F STORIES  O’DONNELL, K. M.  D27415
IN THE PRESENCE OF MINE ENEMIES  TURTLEDOVE, HARRY  H
IN THE PROBLEM PIT  POHL, FREDERIK  P
IN THE REALM OF THE WOLF [DRENAI WAYLANDER 2]  GEMMELL, DAVID  P
IN THE RED LORD’S REACH  EISENSTEIN, PHYLLIS  C/P,SFWA-P,P
IN THE RIFT [GLENRAVEN 2]  BRADLEY, MARION ZIMMER/LISLE, HOLLY  H:2,P
IN THE RUINS [CROWN OF STARS 6]  ELLIOTT, KATE  H
IN THE SEA NYMPH’S LAIR [SPELL WARS 2]  HALE, F. J.  P
IN THE SERVICE OF DRAGONS [KEEPER MARTIN 5]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS II [KEEPER MARTIN 6]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS III [KEEPER MARTIN 7]  STANEK, ROBERT  LP
IN THE SERVICE OF DRAGONS IV [KEEPER MARTIN 8]  STANEK, ROBERT  LP
IN THE SHADOW OF EVIL  GREENBERG, MARTIN H./HELFERS, JOHN  PA
IN THE SHADOW OF EVIL: THE HANDKERCHIEF [@BLACKSTONE #4]  SAUL, JOHN  C/P,P
IN THE SHADOW OF OMEN [MATTHEWS 4]  BURGAUER, STEVEN  LP
IN THE SHADOW OF OMIZANTRIM [WAR OF POWERS 5]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P,P
IN THE SHADOW OF THE GARGOYLE  KILPATRICK, NANCY/ROCHE, THOMAS S.  PA
IN THE SHADOW OF THE OAK KING  JONES, COURTNEY  H
IN THE TOMB OF THE BISHOP [@TALES OF THE WEREWOLF CLAN #1]  MUNN, H. WARNER  SR-H
IN THE UPPER ROOM AND OTHER LIKELY STORIES  BISSON, TERRY  H
IN THE WET  SHUTE, NEVIL  H,P
IN THE WRONG HANDS  GIBSON, EDWARD  P

The MITSFS Titledex  September 24, 2007  Page 178
TITLE  AUTHOR  CODE
IN THIS AGE OF STONE  SOSNA, SHARON  C/P,P
IN WAR TIMES  GOONAN, KATHLEEN ANN  H
IN WINTER’S SHADOW  BRADSHAW, GILLIAN  H:2,P
IN-WORLD, THE  ROBERTS, LIONEL  SR-P
IN YOUR DREAMS [PAUL CARPENTER 2]  HOLT, TOM  H
INADEQUATE ADEPT, THE [MARVIN BREWSTER 2]  HAWKE, SIMON  C/P:2,P
INCANDESCENT ONES, THE  HOYLE, FRED/HOYLE, GEOFFREY  P
INCARNATIONS  BARKER, CLIVE  H,LP
INCIDENT AT ARBUK [STAR TREK VOYAGER]  BETANCOURT, JOHN GREGORY  P
INCIDENT AT EXETER  FULLER, JOHN G.  [C/RAN-P]
INCIDENT ON ATH [DUMAREST 18]  TUBB, E. C.  C/P:2,P
INCOMER, THE  ELPHINSTONE, MARGARET  LP
INCOMPETENCE  GRANT, ROB  H
INCOMPLEAT NIFFT, THE  SHEA, MICHAEL  P
INCOMPLETE ENCHANTER, THE [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  SR-H,C/P:2,P
INCONSTANT MOON (BRITISH)  NIVEN, LARRY  SR-P,C/P
INCONSTANT STAR [MAN-KZIN WARS]  ANDERSON, POUL  C/P:2,P

INCORPORATED KNIGHT, THE DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

H,SFWA-P

INCREDIBLE ADVENTURES #1  MARSHAL, GENE/WAEDT, CARL F.  SR-LPA
INCREDIBLE ADVENTURES OF DENNIS DORGAN, THE  HOWARD, ROBERT E.  SR-H,P
INCREDIBLE PLANET, THE  CAMPBELL, JOHN W. JR.  SR-H
INCREDIBLE RADIO EXPLOITS OF DOC SAVAGE 1, THE  ROBESON, KENNETH  VLP
INCREDIBLE SHRINKING MAN, THE  MATHESON, RICHARD  P
INCREDIBLE SHRINKING MORK, THE [MORK AND MINDY 2]  WAGNER, ROBIN S.  C/P,P
INCREDIBLE TALES  SAKI  P
INCREDIBLE TIDE, THE  KEY, ALEXANDER  H
INCREDIBLE UMBRELLA, THE  KAYE, MARVIN  P,SFWA
INCUBUS  RUSSELL, RAY  P
INCUBUS DREAMS [ANITA BLAKE 12]  HAMILTON, LAURELL K.  H
INDA  SMITH, SHERWOOD  H
INDECENT RELATIONS  MCKAY, KENNETH  P
INDEPENDENCE [QUANTUM LEAP]  PEEL, JOHN  P
INDEPENDENCE DAY [DOCTOR WHO PAST DOCTORS]  DARVILL-EVANS, PETER  C/P,P

INDEPENDENCE DAY [@ID4] DEVLIN, DEAN/EMMERICH, ROLAND/

MOLSTAD, STEPHEN

P

INDEPENDENT COMMAND, THE [FLIGHT ENGINEER 3]  DOOHAN, JAMES/STIRLING, S. M.  H,P
INDESTRUCTIBLE, THE  GARNER, ROLF  SR-P
INDESTRUCTIBLE MAN, THE [@DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  P
INDEX AND SHORT HISTORY OF UNKNOWN, AN  METZGER, ARTHUR  R/IND
INDEX J-81 ARBEITET FUR MR. FAUST  VERLAG VOLK UND WELT  FH
INDEX OF POSSIBILITIES, AN  CATALOGUE, THE  SFWA-VLP
INDEX OF SCIENCE FICTION MAGAZINES, 1951–1965, THE  METCALF, NORMAN  R/IND
INDEX TILL SVENSKA GALAXY  NILSSON, INGEMAR  R/IND
INDEX TILL...HAPNA!  ANDERSSON, LEIF  R/IND
INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 1  STONE, GRAHAM  R/IND
INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 2  STONE, GRAHAM  R/IND
INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 4  STONE, GRAHAM  R/IND
INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 5  STONE, GRAHAM  R/IND
INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 6  STONE, GRAHAM  R/IND

INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 VOLUME 2 (AUTHORS)

STONE, GRAHAM

R/IND

INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 VOLUME 3 (TITLES)

STONE, GRAHAM

R/IND

INDEX TO FANTASY AND SCIENCE FICTION IN MUNSEY PUBLICATIONS  ANONYMOUS  R/IND
INDEX TO FICTION IN RADIO NEWS AND OTHER MAGAZINES  COCKCROFT, T. G. L.  R/IND-F
INDEX TO MISCELLANEOUS FANTASY MAGAZINES  JOHNSON, KENNETH R.  R/IND
INDEX TO PERRY RHODAN — US EDITION 1–25  NESFA  R/IND-F
INDEX TO PERRY RHODAN — US EDITION 26–50  NESFA  R/IND-F
INDEX TO SCIENCE FICTION ANTHOLOGIES AND COLLECTIONS  CONTENTO, WILLIAM G.  R/IND
INDEX TO SCIENCE FICTION ANTHOLOGIES AND COLLECTIONS 1977–1983  CONTENTO, WILLIAM G.  R/IND
INDEX TO SHORT SCIENCE FICTION 1987  NESFA  R/IND
INDEX TO SHORT SCIENCE FICTION FOR 1988  NESFA  R/IND
INDEX TO SHORT SCIENCE FICTION FOR 1989  NESFA  R/IND
INDEX TO THE S-F MAGAZINES 1966, THE  STRAUSS, ERWIN S.  R/IND:2
INDEX TO THE SCIENCE FICTION MAGAZINES, 1926–1950  DAY, DONALD B.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1961  LEWIS, AL  SRL/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1966  NESFA  SRL/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1966–1970  NESFA  R/IND:2
INDEX TO THE SCIENCE FICTION MAGAZINES 1967  NESFA  SRL/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1968  NESFA  SRL/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1969  NESFA  SRL/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1977  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1978  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1979  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1980  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND

TITLE  AUTHOR  CODE
INDEX TO THE SCIENCE FICTION MAGAZINES 1981  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1982  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1983  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SCIENCE FICTION MAGAZINES 1984  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1971–1972 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1973 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1974 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1975 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1976 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1977–1978 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1979–1980 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1981 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1982 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1983 NESFA

R/IND

INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL ANTHOLOGIES 1986 NESFA

R/IND

INDEX TO THE SELECTED LETTERS OF H. P. LOVECRAFT, AN  JOSHI, S. T.  R/IND

INDEX TO THE SELECTED LETTERS OF H. P. LOVECRAFT (SECOND EDITION, KEEP WITH THE SELECTED LETTERS), AN JOSHI, S. T. =LOVECRAFT, H. P.

R/REF-H

INDEX TO THE SEMI-PROFESSIONAL FANTASY MAGAZINES 1982  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SEMI-PROFESSIONAL FANTASY MAGAZINES 1983  BOYAJIAN, JERRY/JOHNSON, KENNETH R.  R/IND
INDEX TO THE SF MAGAZINES, 1951–1965  STRAUSS, ERWIN S.  R/IND:2,SRL/IND
INDEX TO THE WEIRD FICTION MAGAZINES — AUTHOR  COCKCROFT, T. G. L.  SRL/IND
INDEX TO THE WEIRD FICTION MAGAZINES — TITLE  COCKCROFT, T. G. L.  SRL/IND
INDIAN IN THE CUPBOARD, THE [INDIAN IN THE CUPBOARD 1]  BANKS, LYNNE REID  C/P:2,P
INDIANA JONES AND THE DANCE OF THE GIANTS [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE GENESIS DELUGE [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE INTERIOR WORLD [INDIANA JONES]  MACGREGOR, ROB  C/P,P
INDIANA JONES AND THE LAST CRUSADE [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE SEVEN VEILS [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE SKY PIRATES [INDIANA JONES]  CAIDIN, MARTIN  C/P,P
INDIANA JONES AND THE UNICORN’S LEGACY [INDIANA JONES]  MACGREGOR, ROB  P
INDIANA JONES AND THE WHITE WITCH [INDIANA JONES]  CAIDIN, MARTIN  P
INDIANS WON, THE  SMITH, MARTIN CRUZ  P
INDIA’S STORY  STARBUCK, KATHLYN S.  C/P:2,P
INDIGO  JOYCE, GRAHAM  P
INDISTINGUISHABLE FROM MAGIC  FORWARD, ROBERT L.  C/P:2,P

INDIVIDUATED HOBBIT:

JUNG, TOLKIEN AND THE ARCHETYPES OF MIDDLE-EARTH, THE O’NEILL, TIMOTHY R.

R/REF-H

INDOCTRINAIRE  PRIEST, CHRISTOPHER  C/P,P,H
INDOMITABLE THUNDER [@WORLD OF ADEN THUNDERSCAPE]  ACRES, MARK  P

INDUSTRIAL MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 2] ARMSTRONG, KELLEY

C/P,P

INEXPERIENCED GHOST AND NINE OTHER STORIES, THE  WELLS, H. G.  C/P,P
INFANTA [@INDIGO #3]  COOPER, LOUISE  C/P,SFWA-P,P
INFECTION [@STAR TREK TNG DOUBLE HELIX 1]  BETANCOURT, JOHN GREGORY  P
INFECTRESS  COOL, TOM  P
INFERNAL [REPAIRMAN JACK]  WILSON, F. PAUL  H
INFERNAL DESIRE MACHINES OF DOCTOR HOFFMAN, THE  CARTER, ANGELA  LP
INFERNAL DEVICES  JETER, K. W.  H
INFERNAL DEVICES [HUNGRY CITY 3]  REEVE, PHILIP  H
INFERNAL DEVICES (PROOF) [@HUNGRY CITY 3]  REEVE, PHILIP  SR-LP
INFERNAL SKY [@DOOM #3]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P:2,P
INFERNO [CALIBAN 2]  ALLEN, ROGER MACBRIDE  P
INFERNO [@INDIGO #2]  COOPER, LOUISE  C/P:2,SFWA-P,P
INFERNO, THE  CREASEY, JOHN  P
INFERNO [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
INFERNO, THE  HOYLE, FRED/HOYLE, GEOFFREY  H:2
INFERNO  NIVEN, LARRY/POURNELLE, JERRY  C/P:2,P

INFERNO [@STAR TREK DS9 MILLENIUM #3] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

INFERNO  RESNICK, MIKE  H
INFERNO!  STATTEN, VARGO  SR-P
INFILTRATOR [STAR TREK TNG]  THOMPSON, W. R.  C/P,P
INFINITE ARENA, THE  CARR, TERRY  HA
INFINITE BATTLE, THE [STARHOUNDS 1]  BISCHOFF, DAVID F.  C/P:2,P
INFINITE CAGE, THE  LAUMER, KEITH  H:2,P

The MITSFS Titledex  September 24, 2007  Page 180
TITLE  AUTHOR  CODE
INFINITE DREAMS  HALDEMAN, JOE  H:2,P
INFINITE JESTS: THE LIGHTER SIDE OF SCIENCE FICTION  SILVERBERG, ROBERT  HA
INFINITE KINGDOMS, THE  RUTHERFORD, MICHAEL  LP
INFINITE LOOP  CONSTANTINE, LARRY  HA
INFINITE MAN, THE  GALOUYE, DANIEL F.  C/P:2,P
INFINITE MOMENT, THE  WYNDHAM, JOHN  C/P:2,P
INFINITE REQUIEM [DOCTOR WHO NEW ADVENTURES]  BLYTHE, DANIEL  C/P:2,P
INFINITE SEA, THE [CHAOS CHRONICLES 3]  CARVER, JEFFREY A.  H,P
INFINITE SUMMER, AN  PRIEST, CHRISTOPHER  C/P,P,H
INFINITE WEB, THE  SILVERBERG, ROBERT  HA
INFINITIES  CROWTHER, PETER  HA
INFINITIVE OF GO, THE  BRUNNER, JOHN  C/P:2,P
INFINITY ONE  HOSKINS, ROBERT  C/PA,PA
INFINITY TWO  HOSKINS, ROBERT  PA,C/PA
INFINITY THREE  HOSKINS, ROBERT  HA,PA
INFINITY FOUR  HOSKINS, ROBERT  PA,C/PA
INFINITY FIVE  HOSKINS, ROBERT  PA
INFINITY BEACH  MCDEVITT, JACK  H:2

INFINITY BOX, THE WILHELM, KATE SFWA-TD12,C/P,SFWA,

TD12,P

INFINITY CONCERTO, THE [MICHAEL PERRIN 1]  BEAR, GREG  C/P:2,P
INFINITY DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
INFINITY FLIGHT [@PERRY RHODAN #24]  ACE  PA,C/PA
INFINITY HOLD  LONGYEAR, BARRY B.  C/P:2,P
INFINITY LINK, THE  CARVER, JEFFREY A.  C/P,SFWA-P:2,P,H:2
INFINITY MACHINE  MULLER, JOHN E.  SR-P
INFINITY PLAGUE, THE [@BIOWARRIORS 1]  VARDEMAN, ROBERT E.  P
INFINITY RACE, THE [DOCTOR WHO EIGHTH DOCTOR]  MESSINGHAM, SIMON  P
INFINITY WELCOMES CAREFUL DRIVERS [@RED DWARF 1]  NAYLOR, GRANT  C/P:2,SFWA-P,P
INFINITY’S SHORE [UPLIFT 5]  BRIN, DAVID  H:2,P
INFINITY’S WEB  FINCH, SHEILA  C/P,P
INFLATABLE VOLUNTEER, THE  AYLETT, STEVE  LP
INFOQUAKE [JUMP 225 1]  EDELMAN, DAVID LOUIS  LP
INFORMAL BIOGRAPHY OF SCROOGE MCDUCK, AN  CHALKER, JACK L.  R/REF-P
INFORMED SOURCES  KAMARCK, LAWRENCE  P
INGATHERING  HENDERSON, ZENNA  H:2
INGULF THE MAD [DARK BORDER]  ZIMMER, PAUL EDWIN  C/P,SFWA-P,P
INHERIT THE EARTH  NUNES, CLAUDE  DG580
INHERIT THE EARTH [EMORTALITY]  STABLEFORD, BRIAN M.  H
INHERIT THE EARTH  WALKER, IRMA  H
INHERIT THE EARTH [@HUNTER THE RECKONING]  WIECK, STEWART  PA
INHERIT THE STARS [GIANTS 1]  HOGAN, JAMES P.  C/P:2,SR-P
INHERIT THE WITCH [@CHARMED]  BURNS, LAURA J.  P
INHERITANCE, THE [@DRAGONLANCE CLASSICS]  BERBERICK, NANCY VARIAN  P
INHERITANCE  BROOKES, OWEN  P
INHERITANCE [@KEYS OF POWER #1]  BROWN, SIMON  P
INHERITOR, THE [INHERITOR 1]  BRADLEY, MARION ZIMMER  C/P:2,LP
INHERITOR [ATEVI 3]  CHERRYH, C. J.  H:2,P
INHERITORS, THE [GRIMES]  CHANDLER, A. BERTRAM  D37062
INHERITORS, THE [GRIMES]/GATEWAY TO NEVER [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
INHERITORS, THE  GOLDING, WILLIAM  P
INHERITORS OF EARTH  EKLUND, GORDON/ANDERSON, POUL  C/P,P,H
INHUMAN BEINGS  CARROLL, JERRY JAY  C/P,LP
INHUMAN CONDITION, THE  BARKER, CLIVE  H:2,P
INITIATE, THE [TIME MASTER 1]  COOPER, LOUISE  C/P:2,SFWA-P,P
INITIATE BROTHER, THE [INITIATE BROTHER 1]  RUSSELL, SEAN  C/P:2,P
INITIATION [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  SR-H
INITIATION, THE [TWILIGHT #3]  BRUNN, ROBERT  P
INITIATION  HUGHES, MARIAN  P
INITIATION, THE [SECRET CIRCLE 1]  SMITH, L. J.  P
INITIATION TO WAR [@MECHWARRIOR/BATTLETECH]  CHARRETTE, ROBERT N.  P
INK [BOOK OF ALL HOURS 2]  DUNCAN, HAL  LP
INKARRI  MILLER, RYAN S. J.  LP
INKLINGS, THE  CARPENTER, HUMPHREY  R/REF-H
INNAMORATI, THE  SNYDER, MIDORI  H
INNER CIRCLE  PETERKIEWICZ, JERZY  H
INNER COSMOS, THE  STATTEN, VARGO  SR-P
INNER ECLIPSE  RUSSO, RICHARD PAUL  P,SFWA-P
INNER LANDSCAPE, THE  PAPERBACK LIBRARY  C/PA,PA
INNER WHEEL, THE  ROBERTS, KEITH  H,P
INNERVERSE  DE CHANCIE, JOHN  P
INNERVISIONS  GRIBBIN, JOHN  P
INNKEEPER’S SONG, THE  BEAGLE, PETER S.  H,LP
INNOCENT, THE [@OUTER LIMITS]  PEEL, JOHN  LP

The MITSFS Titledex  September 24, 2007  Page 181
TITLE  AUTHOR  CODE
INNOCENTS ABOARD  WOLFE, GENE  H
INQUISITION [@AQUASILVA #2]  AUDLEY, ANSELM  P
INQUISITION WAR, THE [@WARHAMMER INQUISITION WAR 1,2,3]  WATSON, IAN  LP
INQUISITOR, THE [NETWORK-CONSORTIUM 3]  FRANKLIN, CHERYL J.  C/P:2,P
INSANE CITY, THE  BULMER, KENNETH  C/P:2,P
INSANITY, ILLINOIS [@NEWS FROM THE EDGE #2]  SUMNER, MARK  P
INSECT INVASION, THE  CUMMINGS, RAY  H
INSECT WARRIORS, THE  LEVIE, REX DEAN  C/P:2,P
INSELN DER BEWAHRUNG, DIE [@PERRY RHODAN #130]  KNEIFEL, HANS  FPR
INSIDE  MORGAN, DAN  C/P:2,P
INSIDE JOB [@STAR PRECINCT #3]  RANDLE, KEVIN/DRISCOLL, RICHARD  C/P,P
INSIDE JOB  WILLIS, CONNIE  H
INSIDE OUTER SPACE  JARVIS, SHARON  R/REF-H
INSIDE OUTSIDE  FARMER, PHILIP JOSE  H:2,P
INSIDE THE FUNHOUSE  RESNICK, MIKE  PA,SFWA-P
INSIDIOUS DR. FU MANCHU, THE  ROHMER, SAX  C/P,P,H
INSOMNIA  KING, STEPHEN  P
INSTANT GOLD  O’ROURKE, FRANK  H:2
INSTEAD OF THREE WISHES  TURNER, MEGAN WHALEN  LP
INSTITUTE, THE  PETYO, ROBERT  C/P,P
INSTRUMENT OF FATE  GOLDEN, CHRISTIE  P
INSTRUMENTALITY OF MANKIND, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P
INSTRUMENTS OF DARKNESS [DOCTOR WHO]  RUSSELL, GARY  P
INSULATORS, THE  CREASEY, JOHN  P
INSURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #2]  REID, THOMAS M.  H
INSURRECTION [STARFIRE 1]  WEBER, DAVID/WHITE, STEVE  C/P:2,SFWA-P,P
INTEGRAL TREES, THE [INTEGRAL TREES 1]  NIVEN, LARRY  H:2,P,SFWA
INTEGRATED MAN, THE  BERLYN, MICHAEL  C/P,P
INTELLIVORE [STAR TREK TNG]  DUANE, DIANE  C/P:2,P
INTER ICE AGE 4  ABE, KOBO  H:2,P
INTERESTING TIMES [DISCWORLD 17/DISCWORLD RINCEWIND 5]  PRATCHETT, TERRY  C/P,P,H
INTERFACE [INTERFACE 1]  ADLARD, MARK  C/P,P
INTERFACE  BURY, STEPHEN  C/P,LP
INTERFACE MASQUE  LEWITT, SHARIANN  H
INTERFACES  LE GUIN, URSULA K./KIDD, VIRGINIA  HA,PA

INTERFERENCE BOOK ONE: SHOCK TACTIC [DOCTOR WHO PAST DOCTORS] MILES, LAWRENCE

P

INTERFERENCE BOOK TWO: THE HOUR OF THE GEEK [DOCTOR WHO EIGHTH DOCTOR] MILES, LAWRENCE

P

INTERFICTIONS  SHERMAN, DELIA/GOSS, THEODORA  LPA
INTERGALAC AGENT  HASSLER, KENNETH W.  H

INTERGALACTIC EMPIRES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

INTERGALACTIC MERCENARIES  WILLIAMS, SHEILA/MANSON, CYNTHIA  PA
INTERGALACTIC TRAVEL (CHINESE)  RODDENBERRY, GENE  FP
INTERIM: HANNES BOK ILLUSTRATION INDEX (SUMMER 1967)  BROOKS, NED  SRL/IND
INTERIM: VAUGHN F. BODE ILLUSTRATION INDEX (SPRING 1971)  BROOKS, NED/BEAHM, GEORGE W.  R/IND-F
INTERIOR LIFE, THE  BLAKE, KATHERINE  C/P:2,P
INTERLOPERS  FOSTER, ALAN DEAN  C/P,P
INTERLUDE ON SILIKO 5 [@PERRY RHODAN #59]  ACE  C/PA,PA
INTERMIND  LUTHER, RAY  P
INTERN, THE  X, DOCTOR  [C/RAN-P]
INTERNATIONAL INCIDENTS [BERKLEY SIX MILLION DOLLAR MAN 2]  JAHN, MICHAEL  P
INTERNATIONAL SCIENCE FICTION YEARBOOK 1979, THE  LESTER, COLIN  R/REF-P,C/REF-H
INTERPLANETARY ADVENTURERS, THE  BERTIN, JACK  H
INTERPLANETARY HUNTER  BARNES, ARTHUR K.  H:2,P
INTERRUPTED JOURNEY, THE  FULLER, JOHN G.  [C/RAN-H,C/RAN-P]
INTERSECTION  WORLD SCIENCE FICTION SOCIETY  R/REF-H

INTERSECTIONS: THE SYCAMORE HILL ANTHOLOGY KESSEL, JOHN/VAN NAME, MARK L./

BUTNER, RICHARD

LPA

INTERSTELLAR EMPIRE  BRUNNER, JOHN  C/P:2,P
INTERSTELLAR ESPIONAGE  DEL MARTIA, ASTRON  SR-P
INTERSTELLAR INTRIGUE (PROOF) [@OUTREACH INVESTIGATIONS #1]  FOOTE, KEITH D.  SR-P
INTERSTELLAR PATROL [INTERSTELLAR PATROL]  ANVIL, CHRISTOPHER  P

INTERSTELLAR PATROL II: THE FEDERATION OF HUMANITY [INTERSTELLAR PATROL] ANVIL, CHRISTOPHER

H

INTERSTELLAR PIG  SLEATOR, WILLIAM  H,P
INTERSTELLAR TWO FIVE  RANKINE, JOHN  P
INTERVENTION [INTERVENTION 1,2]  MAY, JULIAN  H:2,SFWA
INTERVIEW WITH THE VAMPIRE [LESTAT 1]  RICE, ANNE  C/P:2,P
INTERWORLD  HAIBLUM, ISIDORE  C/P,P

INTERZONE: THE FIRST ANTHOLOGY [INTERZONE 1] CLUTE, JOHN/GREENLAND, COLIN/

PRINGLE, DAVID

HA

INTERZONE: THE SECOND ANTHOLOGY [INTERZONE 2] CLUTE, JOHN/PRINGLE, DAVID/

OUNSLEY, SIMON PA

The MITSFS Titledex  September 24, 2007  Page 182
TITLE  AUTHOR  CODE
INTIMATIONS OF MORTALITY [FOREVER KNIGHT]  GARRETT, SUSAN M.  P
INTO A DARK REALM [DARKWAR 2]  FEIST, RAYMOND E.  H
INTO DEEPEST SPACE  HOYLE, FRED/HOYLE, GEOFFREY  H:2,P
INTO PLUTONIAN DEPTHS  COBLENTZ, STANTON A.  SR-P
INTO THE ABYSS  VERNE, JULES  H
INTO THE AETHER  LUPOFF, RICHARD A.  C/P:2,P
INTO THE ALTERNATE UNIVERSE [GRIMES]  CHANDLER, A. BERTRAM  DM107

INTO THE ALTERNATE UNIVERSE [GRIMES]/ CONTRABAND FROM OTHERSPACE [GRIMES] CHANDLER, A. BERTRAM

C/P,P

INTO THE DARK LANDS [SUNDERED 1]  SAGARA, MICHELLE  P,LP
INTO THE DARKNESS [@FAMILY #2]  MCCARTHY, KEVIN/SILVA, DAVID  P
INTO THE DARKNESS [INTO THE DARKNESS 1]  TURTLEDOVE, HARRY  C/P,P,H
INTO THE DREAM  SLEATOR, WILLIAM  C/P,P
INTO THE EIGHTH DECADE  WILLIAMSON, JACK  SR-H
INTO THE FIRE [HEL’S CRUCIBLE 2/MITHGAR 4]  MCKIERNAN, DENNIS L.  P
INTO THE FORGE [HEL’S CRUCIBLE 1/MITHGAR 3]  MCKIERNAN, DENNIS L.  P
INTO THE FRINGE  TURNER, KARLA  SFWA-P
INTO THE GREEN  DE LINT, CHARLES  C/P,P,H
INTO THE LABYRINTH [@DEATH GATE #6]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
INTO THE LAND OF THE UNICORNS  COVILLE, BRUCE  LP
INTO THE LOOKING GLASS [LOOKING GLASS 1]  RINGO, JOHN  H
INTO THE MAELSTROM [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
INTO THE NEBULA [STAR TREK TNG]  DEWEESE, GENE  C/P:2,P
INTO THE NIGER BEND  VERNE, JULES  C/P,P,H
INTO THE OUT OF  FOSTER, ALAN DEAN  H:2
INTO THE SEA OF STARS  FORSTCHEN, WILLIAM R.  C/P,P
INTO THE SHADOWS [SHADOWRUN 7]  WEISMAN, JOSEPH K.  C/PA,PA
INTO THE SLAVE NEBULA  BRUNNER, JOHN  C/P:2,P
INTO THE SUN AND OTHER STORIES  MILNE, ROBERT DUNCAN  SR-H
INTO THE TENTH MILLENNIUM  CAPON, PAUL  SR-P
INTO THE THINKING KINGDOMS [JOURNEYS OF THE CATECHIST 2]  FOSTER, ALAN DEAN  P
INTO THE UNKNOWN  CARR, TERRY  HA
INTO THE UNKNOWN  PHILMUS, ROBERT M.  R/REF-H
INTO THE VOID [@STAR TREK NEW FRONTIER #2]  DAVID, PETER  C/P,P
INTO THE VOID [@SPELLJAMMER CLOAKMASTER #2]  FINDLEY, NIGEL D.  C/P,P
INTRIGUES [@BLENDING ENTHRONED #1]  GREEN, SHARON  C/P,P
INTRODUCCION A LA CIENCIA FICCION  HURTADO, OSCAR  FP
INTRODUCTION TO THE MOON, AN  SALISBURY, JOHN W.  [C/RAN-H]
INTRUDER, THE  LEIMAS, BROOKE  P
INTRUDER [ROBOTS AND ALIENS 3]  THURSTON, ROBERT  C/P,SFWA-P:2,P
INTRUDERS, THE  FANE, BRON  SR-P
INTUIT  CLEMENT, HAL  H
INU-YASHA  TAKAHASHI, RUMIKO  CX
INUIT  BELL, M. SHAYNE  P
INVADER [ATEVI 2]  CHERRYH, C. J.  H:2
INVADER  HILL, ALBERT FAY/HILL, DAVID CAMPBELL  C/P,P
INVADER [@ROBOTS IN TIME 6]  WU, WILLIAM F.  P
INVADER ON MY BACK  HIGH, PHILIP E.  DH085
INVADERS, THE  CARR, ROBERT SPENCER  B
INVADERS!  DANN, JACK/DOZOIS, GARDNER  PA
INVADERS!  DICKSON, GORDON R.  C/P,P
INVADERS, THE [INVADERS 1]  LAUMER, KEITH  C/P:2,P
INVADERS, THE [@OUTER LIMITS]  PEEL, JOHN  LP
INVADERS: ALIEN MISSILE THREAT, THE  NEWMAN, PAUL S.  C/P,BLB
INVADERS ARE COMING, THE  NOURSE, ALAN E./MEYER, J. A.  C/P:2,P
INVADERS AT GROUND ZERO [TALES OF TOMORROW 1]  HOUSTON, DAVID  P
INVADERS FROM EARTH  SILVERBERG, ROBERT  DD286,P,C/P:2
INVADERS FROM EARTH/TO WORLDS BEYOND  SILVERBERG, ROBERT  P
INVADERS FROM MARS  GARTON, RAY  P
INVADERS FROM RIGEL  PRATT, FLETCHER  C/P:2,P
INVADERS FROM SPACE  SILVERBERG, ROBERT  HA
INVADERS FROM THE CENTRE [ASGARD 2]  STABLEFORD, BRIAN M.  P
INVADERS FROM THE INFINITE  CAMPBELL, JOHN W. JR.  H:2,P
INVADERS OF EARTH  CONKLIN, GROFF  HA
INVADERS OF EARTH (15 STORIES)  CONKLIN, GROFF  PA,C/PA
INVADERS OF EARTH (17 STORIES)  CONKLIN, GROFF  PA,C/PA
INVADERS OF SPACE  LEINSTER, MURRAY  C/P:2,P
INVADERS ON THE MOON  NEVILLE, KRIS  C/P:2,P
INVADERS PLAN, THE [@MISSION EARTH #1]  HUBBARD, L. RON  SFWA-P,H,P
INVADING ASTEROID, THE  WELLMAN, MANLY WADE  B
INVASION!  CHAMBERS, WHITMAN  H
INVASION  COOK, ROBIN  P
INVASION! [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
INVASION [@MAGIC: THE GATHERING INVASION 1]  KING, J. ROBERT  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 183
TITLE  AUTHOR  CODE
INVASION  MACKSEY, KENNETH  H
INVASION, THE [DOCTOR WHO]  MARTER, IAN  SR-P,P
INVASION [@WAR WORLD #4]  POURNELLE, JERRY  PA
INVASION  VAN LOON, HENDRIK WILLEM  H
INVASION  WOLFE, AARON  SR-P
INVASION: 2200 A.D.  GOLDSMITH, HOWARD  H
INVASION AMERICA [INVASION AMERICA 1]  GOLDEN, CHRISTIE  P
INVASION: EARTH  HARRISON, HARRY  C/P,LP
INVASION FROM 2500  EDWARDS, NORMAN  C/P,P
INVASION FROM MARS, THE  CANTRIL, HADLEY  R/REF-P
INVASION FROM MARS: INTERPLANETARY STORIES  WELLES, ORSON  SR-PA,C/PA
INVASION FROM SPACE [@PERRY RHODAN #4]  ACE  PA,C/PA
INVASION FROM SPACE  BRADFORD, MATTHEW C.  SR-P
INVASION OF MARS  SERVISS, GARRETT P.  SR-P
INVASION OF THE BLUE LIGHTS  GLICK, RUTH  P
INVASION OF THE BODY SNATCHERS  FINNEY, JACK  C/P:2,LP
INVASION OF THE CAT-PEOPLE [DOCTOR WHO MISSING ADVENTURES]  RUSSELL, GARY  C/P:2,P
INVASION OF THE ORMAZOIDS [DOCTOR WHO]  MARTIN, PHILIP  [C/RAN-P]
INVASION OF THE ROBOTS  ELWOOD, ROGER  PA,C/PA
INVASIONE DELLA GALASSIA, L  HAMILTON, EDMOND  FP

INVASIONS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

INVASIONS [@GAMBLER’S STAR #3]  HOLDER, NANCY  P
INVASORES INVADIDOS, LOS  PARRISH, GLENN  FP
INVENTING MEMORY  HARRIS, ANNE  LP
INVERSIA  RICHARD-BESSIERE, F.  FP
INVERSIONS  BANKS, IAIN M.  H:2
INVERTED WORLD, THE  PRIEST, CHRISTOPHER  H:2,SFWA,P
INVERTED WORLD/FUGUE FOR A DARKENING ISLAND  PRIEST, CHRISTOPHER  P
INVESTIGATION, THE  LEM, STANISLAW  C/P:2,P
INVID INVASION [@ROBOTECH #10]  MCKINNEY, JACK  C/P,P
INVINCIBLE, THE  LEM, STANISLAW  C/P:3,SFWA,P
INVINCIBLE DOC SAVAGE, THE  GRUSSKIN, EDWARD  B
INVISIBILITY AFFAIR, THE [@MAN FROM UNCLE #11]  STRATTON, THOMAS  C/P,P
INVISIBLE BARRIERS  OSBORNE, DAVID  H:2
INVISIBLE COMPANY, THE  SANDERS, SCOTT RUSSELL  P,SFWA-P
INVISIBLE COUNTRY, THE  MCAULEY, PAUL J.  P
INVISIBLE DEATH [ZARKON 2]  CARTER, LIN  H:2,SFWA,P
INVISIBLE EMPIRE, THE [OPERATOR 5 #2]  STEELE, CURTIS  SR-P,C/P
INVISIBLE FIRE  GRAVERSEN, PAT  P
INVISIBLE FORCE, THE [@NEW TOM SWIFT #10]  APPLETON, VICTOR  LP
INVISIBLE MAN, THE  JAHN, MICHAEL  P
INVISIBLE MAN, THE  WELLS, H. G.  SR-P,C/P:2
INVISIBLE MAN, THE/WAR OF THE WORLDS, THE  WELLS, H. G.  C/P:2,P
INVISIBLE MEN  DAVENPORT, BASIL  PA,C/PA
INVISIBLE PYRAMID, THE  EISELEY, LOREN  [C/RAN-H]
INVISIBLE RESIDENTS  SANDERSON, IVAN T.  [C/RAN-P]
INVISIBLE RING, THE [BLACK JEWELS]  BISHOP, ANNE  C/P,P
INVISIBLES, THE [MARK HOOD 9]  DARK, JAMES  P
INVISIBLES, THE  HURWOOD, BERNHARDT J.  P
INVITATION, THE [@BOOKS OF MAGIC 1]  JABLONSKI, CARLA  P
INVITATION TO CAMELOT  GODWIN, PARKE  C/PA,PA
INVITATION TO THE GAME  HUGHES, MONICA  LP
INVOKING DARKNESS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #3]  CAVELOS, JEANNE  C/P,P
INVOLUTION OCEAN  STERLING, BRUCE  C/P,P
ION WAR, THE  KAPP, COLIN  C/P:2,P
IOWA BASEBALL CONFEDERACY, THE  KINSELLA, W. P.  H
IPOMOEA  RACKHAM, JOHN  D37250
I.Q. MERCHANT, THE  BOYD, JOHN  H
IRDA, THE [DRAGONLANCE LOST HISTORIES 2]  BAKER, LINDA P.  C/P,P
IRETA ADVENTURE, THE [DINOSAUR PLANET 1,2]  MCCAFFREY, ANNE  H
IRIS  BARTON, WILLIAM/CAPOBIANCO, MICHAEL C/P,SFWA-P,P,H
IRIS, THE BEWITCHED [@PHENWICK #39]  KIMBROUGH, KATHERYN  P
IRON ARMY, THE  ADDISCOMBE, JOHN  B
IRON BUTTERFLIES  NORTON, ANDRE  P
IRON CAGE  NORTON, ANDRE  H:2,P
IRON COUNCIL [NEW CROBUZON]  MIEVILLE, CHINA  H:2
IRON DAWN [BARRA THE PICT 1]  STOVER, MATTHEW WOODRING  C/P,P
IRON DRAGON’S DAUGHTER, THE  SWANWICK, MICHAEL  H:2,P
IRON DREAM, THE  SPINRAD, NORMAN  C/P:2,P
IRON FIST [@STAR WARS X-WING 6]  ALLSTON, AARON  C/P,P
IRON GOD, THE  WILLIAMSON, JACK  GRD18
IRON GRAIL, THE [MERLIN CODEX 2]  HOLDSTOCK, ROBERT P.  P
IRON HANDS [@WARHAMMER 40K]  GREEN, JONATHAN  P

The MITSFS Titledex  September 24, 2007  Page 184
TITLE  AUTHOR  CODE
IRON HEEL, THE  LONDON, JACK  H,LP
IRON HORSE [@AFRIKORPS #2]  DOLAN, BILL  P
IRON LANCE, THE [CELTIC CRUSADES 1]  LAWHEAD, STEPHEN R.  P
IRON LORDS, THE [WAR OF THE GODS ON EARTH 1]  OFFUTT, ANDREW J.  C/P:2,P
IRON MAIDEN, THE [BIO OF A SPACE TYRANT 6]  ANTHONY, PIERS  LP
IRON MAN, THE  HOWARD, ROBERT E.  P
IRON MAN: THE ARMOR TRAP  COX, GREG  C/P,P
IRON MASTER [AMTRAK WARS 3]  TILLEY, PATRICK  C/P,P
IRON RAIN, THE  MALCOLM, DONALD  SR-P
IRON SHADOWS  BARNES, STEVEN  H
IRON SKULL, THE [@AVENGER #35]  ROBESON, KENNETH  P
IRON STAR, THE  TAINE, JOHN  SR-H,LP
IRON SUNRISE [SINGULARITY SKY 2]  STROSS, CHARLES  H
IRON THANE, THE  HENDERSON, JASON  P,SFWA-P
IRON THRONE, THE [BIRTHRIGHT 1]  HAWKE, SIMON  P
IRON TONGUE [CENOTAPH ROAD 4]  VARDEMAN, ROBERT E.  C/P,P
IRON TREE, THE [CROWTHISTLE 1]  DART-THORNTON, CECILIA  H,P
IRON WARS, THE [@MONARCHIES OF GOD #3]  KEARNEY, PAUL  P
IRONBRAND [IRON ANGEL 1]  MORRESSY, JOHN  C/P,P
IRONCASTLE  ROSNY, J. H.  C/P,P
IRONCROWN MOON [BOREAL MOON 2]  MAY, JULIAN  H
IRONHAND’S DAUGHTER [HAWK QUEEN]  GEMMELL, DAVID  P
IRONHELM [@FORGOTTEN REALMS MAZTICA #1]  NILES, DOUGLAS  P
IROSHI [IROSHI 1]  OSBORNE, CARY  C/P,P
IRRATIONAL FEARS  SPENCER, WILLIAM BROWNING  H
IRRATIONAL NUMBERS  EFFINGER, GEORGE ALEC  H,SFWA
IRRATIONALS, THE  LESSER, MILTON  B
IRRESISTIBLE FORCES  ASARO, CATHERINE  LPA
IRSUD [DIADEM 3]  CLAYTON, JO  C/P:2,P
IS ANOTHER WORLD WATCHING?  HEARD, GERALD  [C/RAN-P]
IS ANYONE THERE?  ASIMOV, ISAAC  [C/RAN-P:2]
IS ELVIS ALIVE?  BREWER-GIORGIO, GAIL  [C/RAN-P]
IS SOMETHING UP THERE?  WHITE, DALE  [C/RAN-P]
IS THAT YOU SIMON?  PIOMBO, AKBAR DEL  SR-P
IS THIS APOCALYPSE NECESSARY?  BRITTAIN, C. DALE  LP
ISAAC ASIMOV  GREEN, SCOTT E.  R/REF-H

ISAAC ASIMOV OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

R/REF-H

ISAAC ASIMOV: A CHECKLIST...  MILLER, MARJORIE M.  R/IND

ISAAC ASIMOV PRESENTS:

THE BEST SCIENCE FICTION OF THE 19TH CENTURY ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

ISAAC ASIMOV PRESENTS THE GOLDEN YEARS OF SCIENCE FICTION  ASIMOV, ISAAC/GREENBERG, MARTIN H.  HA:3
ISAAC ASIMOV’S ADVENTURES OF SCIENCE FICTION  SCITHERS, GEORGE  HA
ISAAC ASIMOV’S ALIENS  DOZOIS, GARDNER  C/PA,PA
ISAAC ASIMOV’S ALIENS AND OUTWORLDERS  MCCARTHY, SHAWNA  C/PA,LPA
ISAAC ASIMOV’S CAMELOT  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S CHRISTMAS  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S CYBERDREAMS  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S DETECTIVES  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S EARTH  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S FANTASY!  MCCARTHY, SHAWNA  HA,SFWA
ISAAC ASIMOV’S FATHER’S DAY  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S GHOSTS  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S HALLOWEEN  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S HISTORY OF I-BOTICS  CHAMBERS, JAMES  HA
ISAAC ASIMOV’S MARS  DOZOIS, GARDNER  C/PA,PA
ISAAC ASIMOV’S MASTERS OF SCIENCE FICTION  SCITHERS, GEORGE  HA
ISAAC ASIMOV’S MOONS  DOZOIS, GARDNER/WILLIAMS, SHEILA  C/PA,PA
ISAAC ASIMOV’S MOTHER’S DAY  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S ROBOTS  DOZOIS, GARDNER/WILLIAMS, SHEILA  C/PA,PA
ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 1  SCITHERS, GEORGE  LPA
ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 2  SCITHERS, GEORGE  PA
ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 3  SCITHERS, GEORGE  PA
ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 4  SCITHERS, GEORGE  C/PA,PA
ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 5  SCITHERS, GEORGE  PA
ISAAC ASIMOVS SCIENCE FICTION MAGAZIN 25  WAHREN, FRIEDEL  FP
ISAAC ASIMOV’S SCIENCE FICTION SERIE 7 (NORWEGIAN)  ASIMOV, ISAAC  FP
ISAAC ASIMOV’S SKIN DEEP  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S SOLAR SYSTEM  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S SPACE OF HER OWN  MCCARTHY, SHAWNA  SFWA
ISAAC ASIMOV’S SPACE OF HER OWN (ABRIDGED)  MCCARTHY, SHAWNA  PA
ISAAC ASIMOV’S TOMORROW’S VOICES  DAVIS PUBLICATIONS  HA,SFWA
ISAAC ASIMOV’S UTOPIAS  DOZOIS, GARDNER/WILLIAMS, SHEILA  C/PA,PA
ISAAC ASIMOV’S VALENTINES  DOZOIS, GARDNER/WILLIAMS, SHEILA  C/PA,PA

The MITSFS Titledex  September 24, 2007  Page 185
TITLE  AUTHOR  CODE
ISAAC ASIMOV’S VAMPIRES  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISAAC ASIMOV’S WEREWOLVES  DOZOIS, GARDNER/WILLIAMS, SHEILA  PA
ISABELLE, THE FRANTIC [@PHENWICK #22]  KIMBROUGH, KATHERYN  P
ISARD’S REVENGE [@STAR WARS X-WING 8]  STACKPOLE, MICHAEL A.  C/P,P
ISEULT  MEANEY, DEE MORRISON  P
ISHMAEL [STAR TREK]  HAMBLY, BARBARA  C/P:3,P
ISHMAEL  QUINN, DANIEL  H
ISIS RISING [ISIS 2]  STEWART, JEAN  LP
ISKIIR  EASTON, M. COLEMAN  C/P,P
ISLAND  HUXLEY, ALDOUS  P
ISLAND, THE  WRIGHT, T. M.  SFWA-P
ISLAND AND THE RING, THE  STEVENSON, LAURA C.  P
ISLAND AT THE TOP OF THE WORLD  CAMERON, IAN  P
ISLAND CALLED MOREAU, AN  ALDISS, BRIAN W.  C/P,P,H
ISLAND IN THE SEA OF TIME [NANTUCKET 1]  STIRLING, S. M.  C/P,P
ISLAND IN THE SKY  WELLMAN, MANLY WADE  H
ISLAND OF CAPTAIN SPARROW, THE  WRIGHT, S. FOWLER  H
ISLAND OF DEATH [DOCTOR WHO PAST DOCTORS]  LETTS, BARRY  P

ISLAND OF DOCTOR DEATH AND OTHER STORIES AND OTHER STORIES, THE WOLFE, GENE

C/P:3,P

ISLAND OF DOCTOR MOREAU  SILVA, JOSEPH  P
ISLAND OF DOCTOR MOREAU, THE  WELLS, H. G.  C/P:2,P
ISLAND OF FEAR AND OTHER SCIENCE FICTION STORIES  SAMBROT, WILLIAM  C/P:2,P
ISLAND OF FU MANCHU, THE  ROHMER, SAX  C/P:2,P
ISLAND OF THE DAY BEFORE, THE  ECO, UMBERTO  H,LP
ISLAND OF THE GREAT MOTHER, THE  HAUPTMANN, GERHART  H
ISLAND OF THE MIGHTY, THE [MABINOGI 4]  WALTON, EVANGELINE  C/P:2,P
ISLAND OF THE SEQUINED LOVE NUN  MOORE, CHRISTOPHER  LP
ISLAND OF THE VOODOO DOLLS  DANGERFIELD, PAUL  SR-L
ISLAND PEOPLE, THE  COBLENTZ, STANTON A.  P
ISLAND SNATCHERS, THE [ARWIN 5]  SMITH, GEORGE H.  C/P,P
ISLAND STALLION RACES, THE  FARLEY, WALTER  H
ISLAND UNDER THE EARTH, THE  DAVIDSON, AVRAM  C/P:2,P
ISLAND WORLDS, THE [ISLAND WORLDS 2]  KOTANI, ERIC/ROBERTS, JOHN MADDOX  C/P,P
ISLANDER, THE [STORMLANDS 1]  ROBERTS, JOHN MADDOX  C/P:2,SFWA-P,P
ISLANDIA  WRIGHT, AUSTIN TAPPAN  SR-H,C/P,P
ISLANDS  RANDALL, MARTA  C/P:2,P
ISLANDS IN THE NET  STERLING, BRUCE  C/P:2,SFWA-P,P
ISLANDS IN THE SKY  CLARKE, ARTHUR C.  C/P,P,H
ISLANDS OF SPACE  CAMPBELL, JOHN W. JR.  SR-H,C/P:2,P
ISLANDS OF TOMORROW  BUSBY, F. M.  P
ISLANDS OUT OF TIME  THOMPSON, WILLIAM IRWIN  H,SFWA
ISLAR, THE  SAXTON, MARK  C/P,P
ISLE BEYOND TIME, THE  DOUGLAS, L. WARREN  P
ISLE OF BATTLE, THE [@SWANS’ WAR #2]  RUSSELL, SEAN  H
ISLE OF DEAD SHIPS, THE  MARRIOTT, CRITTENDEN  H
ISLE OF DESTINY  FLINT, KENNETH C.  C/P,P
ISLE OF GLASS, THE [HOUND AND FALCON #1]  TARR, JUDITH  C/P,@H,SFWA-P,P
ISLE OF THE DEAD  ZELAZNY, ROGER  C/P:2,SFWA-P,P
ISLE OF THE SHAPESHIFTERS  COONTZ, OTTO  P
ISLE OF VIEW [@XANTH #13]  ANTHONY, PIERS  H:2,SFWA,P
ISLE OF WOMAN [GEODYSSEY 1]  ANTHONY, PIERS  H:2,P
ISLES OF THE BLEST, THE  LLYWELYN, MORGAN  C/P,SFWA-P,P
ISOLA NEL CIELO, L  CLARKE, ARTHUR C.  FP
ISOTOPE MAN, THE  MAINE, CHARLES ERIC  H:2
ISSOLA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P,P,H
ISSUE AT HAND, THE  ATHELING, WILLIAM JR.  R/REF-H,C/REF-H
IT  KING, STEPHEN  C/P,P,H
IT CAME FROM ON HIGH  HARMAN, ANDREW  P
IT CAME FROM SCHENECTADY  LONGYEAR, BARRY B.  H:2,SFWA,P

IT CAME FROM THE DRIVE-IN PARTRIDGE, NORMAN/

GREENBERG, MARTIN H.

PA

IT CAN’T HAPPEN HERE  LEWIS, SINCLAIR  C/P,P
IT LIVES AGAIN!  DIXON, JAMES  P
IT LOOKS ALIVE TO ME!  BAUM, THOMAS  LP
IT WAS THE DAY OF THE ROBOT  LONG, FRANK BELKNAP  C/P:2,P
ITERATIONS  SAWYER, ROBERT J.  LP
ITHANALIN’S RESTORATION [ETHSHAR]  WATT-EVANS, LAWRENCE  H
ITINERANT EXORCIST, THE [DEMON WARS 2]  MCCONNELL, ASHLEY  P
IT’S A MAD, MAD, MAD GALAXY  LAUMER, KEITH  C/P:2,P
IT’S ALIVE!  WOODLEY, RICHARD  P
IT’S ALL IN YOUR MIND  BLOCH, ROBERT  SR-P
IT’S ONLY TEMPORARY  SHAPIRO, ERIC  LP
IVANHOE  SCOTT, WALTER  H
IVANHOE GAMBIT, THE [TIMEWARS 1]  HAWKE, SIMON  C/P:2,P

TITLE  AUTHOR  CODE
IVORY  RESNICK, MIKE  C/P:2,SFWA-P,P
IVORY AND THE HORN, THE [NEWFORD]  DE LINT, CHARLES  C/P,P,H
IVORY DISC, THE  BREBNER, PERCY JAMES  H
IVORY STAR, THE  HEERMANN, TRAVIS  P

The MITSFS Titledex  September 24, 2007  Page 187
TITLE  AUTHOR  CODE
J. R. R. TOLKIEN  EVANS, ROBLEY  R/REF-P,C/REF-P
J. R. R. TOLKIEN: ARCHITECT OF MIDDLE EARTH  GROTTA-KURSKA, DANIEL  R/REF-P,C/REF-P
J. R. R. TOLKIEN CALENDAR, THE (1973)  CALENDAR  CAL
J. R. R. TOLKIEN CALENDAR, THE (1975)  CALENDAR  CAL
J. R. R. TOLKIEN CALENDAR, THE (1976)  CALENDAR  CAL
J. R. R. TOLKIEN CALENDAR, THE (1979)  CALENDAR  CAL
J. R. R. TOLKIEN CALENDAR, THE (1980)  CALENDAR  CAL
J. R. R. TOLKIEN CALENDAR, THE (1981)  CALENDAR  CAL
JABBA’S PALACE POP-UP BOOK [STAR WARS]  ANDERSON, KEVIN J./MOESTA, REBECCA  SR-MM
JACK FAUST  SWANWICK, MICHAEL  H:2,LP
JACK LONDON, STAR WARRIOR  BISCHOFF, DAVID F.  LP
JACK OF EAGLES  BLISH, JAMES  C/P,GN3.19,P,H
JACK OF KINROWAN [JACKY ROWAN 1,2]  DE LINT, CHARLES  P
JACK OF SHADOWS  ZELAZNY, ROGER  C/P,P,H,SFWA
JACK OF SWORDS [DUMAREST 14]  TUBB, E. C.  C/P:2,P
JACK THE BODILESS [GALACTIC MILIEU 1]  MAY, JULIAN  H:2,P
JACK, THE GIANT-KILLER [JACKY ROWAN 1]  DE LINT, CHARLES  C/P,SFWA-P,P,H
JACK THE GIANT KILLER (PROOF) [JACKY ROWAN 1]  DE LINT, CHARLES  SR-LP
JACK VANCE  UNDERWOOD, TIM/MILLER, CHUCK  R/REF-H
JACK VANCE TREASURY, THE  VANCE, JACK  H
JACK WILLIAMSON: A PRIMARY AND SECONDARY BIBLIOGRAPHY  MYERS, ROBERT E.  R/IND
JACKAL HELIX, THE [GLOBAL 2000 2]  BASILE, GLORIA VITANZA  P
JACKAL OF NAR, THE [TYRANTS AND KINGS 1]  MARCO, JOHN  C/P:2,LP
JACKAL’S HEAD, THE  PETERS, ELIZABETH  SFWA-P
JACKEEN J. O’MALLEY FOR CONGRESS [@BARNABY #3]  JOHNSON, CROCKETT  SCX
JACOB [@NIGHTWALKERS]  FRANK, JACQUELYN  P
JADE DARCY AND THE AFFAIR OF HONOR [@JADE DARCY #1]  GOLDIN, STEPHEN/MASON, MARY  C/P,P
JADE DARCY AND THE ZEN PIRATES [@JADE DARCY #2]  GOLDIN, STEPHEN/MASON, MARY  C/P:2,P
JADE ENCHANTRESS, THE  PRICE, E. HOFFMANN  C/P:2,P
JADE OGRE, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P:2,P
JADE UNICORN, THE  HALPERN, JAY  P
JADE WARRIOR, THE [@BLADE #2]  LORD, JEFFREY  C/P,P
JAGGED ORBIT, THE  BRUNNER, JOHN  C/P:2,P,H
JAGGED STEELE [STEELE 4]  MASTERS, J. D.  P
JAGUAR  RANSOM, BILL  C/P:2,SFWA-P:2,P
JAGUAR AND THE WOLF, THE  CUTTER, LEAH R.  P
JAGUAR HUNTER, THE  SHEPARD, LUCIUS  H,LP
JAGUAR KNIGHTS, THE [KING’S BLADES]  DUNCAN, DAVE  H
JAGUAR PRINCESS, THE  BELL, CLARE  P
JAMAIS VU PAPERS, THE (PROOF)  COLEMAN, WIM/PERRIN, PAT  SFWA
JAMES AND THE GIANT PEACH  DAHL, ROALD  P
JAMES BOND AND MOONRAKER  WOOD, CHRISTOPHER  P
JAMES H. SCHMITZ: A BIBLIOGRAPHY  OWINGS, MARK  R/IND

JAMES TIPTREE AWARD ANTHOLOGY 1, THE FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

LPA

JAMES TIPTREE AWARD ANTHOLOGY 2, THE FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

LPA

JAMES TIPTREE AWARD ANTHOLOGY 3, THE FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

LPA

JAMIE THE RED  DICKSON, GORDON R./GREEN, ROLAND  C/P:2,P
JAN IN INDIA  KLINE, OTIS ADELBERT  SR-LP
JAN OF THE JUNGLE  KLINE, OTIS ADELBERT  P
JANDAR OF CALLISTO [JANDAR 1]  CARTER, LIN  C/P:2,P
JANDRAX  LOGSDON, SYD  P
JANE, THE COURAGEOUS [@PHENWICK #2]  KIMBROUGH, KATHERYN  P
JANISSARIES [@JANISSARIES 1]  POURNELLE, JERRY  C/P,LP,H
JANUARY PLATOON, THE [JEFFERSON’S WAR 4]  RANDLE, KEVIN  C/P,P
JANUS CONJUNCTION, THE [DOCTOR WHO EIGHTH DOCTOR]  BAXENDALE, TREVOR  P
JANUS EQUATION, THE  SPRUILL, STEVEN G.  BS4
JANUS GATE, THE [STAR TREK JANUS GATE 1,2,3]  GRAF, L. A.  H
JANUS MASK, THE  KNAAK, RICHARD A.  P
JANUS SYNDROME, THE  MCDONALD, STEVEN E.  C/P,P
JAPAN SINKS  KOMATSU, SAKYO  H
JAPANESE FAIRY TALES  OZAKI, YEI THEODORA  HA
JARAN [JARAN 1]  ELLIOTT, KATE  C/P:2,SFWA-P,P
JARGOON PARD, THE [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P,P,H
JARKA RUUS [HIGH DRUID OF SHANNARA 1]  BROOKS, TERRY  H
JARKA RUUS (PROOF) [HIGH DRUID OF SHANNARA 1]  BROOKS, TERRY  SR-LP
JASON COSMO [JASON COSMO 1]  MCGIRT, DAN  C/P,P
JASON LIVES  HAWKE, SIMON  P
JASPER FOREST, THE [@GUARDIAN #2]  GRAY, JULIA  P
JAUFRY THE KNIGHT AND THE FAIR BRUNISSENDE  ANONYMOUS  LP
JAWS  BENCHLEY, PETER  P
JAWS OF DARKNESS [INTO THE DARKNESS 5]  TURTLEDOVE, HARRY  P

TITLE  AUTHOR  CODE
JAWS OF MENX, THE  MAXWELL, ANN  C/P:2,P
JAWS THAT BITE, THE CLAWS THAT CATCH, THE  CONEY, MICHAEL G.  P
JAYDIUM  WHEELER, DEBORAH  C/P,P
JAZZ, THE  SCOTT, MELISSA  H
JED THE DEAD  FOSTER, ALAN DEAN  C/P,P
JEDI BOUNTY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
JEDI ECLIPSE [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #2]  LUCENO, JAMES  P
JEDI MASTER’S QUIZBOOK, THE  MILLER, RUSTY  [C/RAN-P]
JEDI SEARCH [@STAR WARS JEDI ACADEMY 1]  ANDERSON, KEVIN J.  C/P:2,P
JEDI TRIAL [@STAR WARS CLONE WARS]  SHERMAN, DAVID/CRAGG, DAN  H
JEDI UNDER SIEGE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
JEDIT [@MAGIC: THE GATHERING LEGENDS 2]  EMERY, CLAYTON  P
JEHAD  YERMAKOV, NICHOLAS  C/P:2,P
JEHOVAH CONTRACT, THE  KOMAN, VICTOR  H,SFWA,P
JEKYLL LEGACY, THE  BLOCH, ROBERT/NORTON, ANDRE  P
JEM  POHL, FREDERIK  C/P,P,H
JENNA STARBORN  SHINN, SHARON  H,LP
JENNIFER GOVERNMENT  BARRY, MAX  H:2
JENNIFER MORGUE, THE [ATROCITY ARCHIVES 2]  STROSS, CHARLES  H:2
JENNIFER MURDLEY’S TOAD [MAGIC SHOP 1]  COVILLE, BRUCE  LP
JEREMY CASE  DEWEESE, GENE  SR-P,C/P
JEREMY THATCHER, DRAGON HATCHER [MAGIC SHOP 2]  COVILLE, BRUCE  LP
JERICHO ITERATION, THE  STEELE, ALLEN  P
JERICHO MOON [BARRA THE PICT 2]  STOVER, MATTHEW WOODRING  LP
JERLAYNE  ABBEY, LYNN  C/P,P
JERNIGAN’S EGG  MUELLER, RICHARD  LP
JERRY BROWN: THE MAN ON THE WHITE HORSE  LORENZ, J. D.  [C/RAN-H]
JERUSALEM FIRE  MELUCH, R. M.  C/P,P
JERUSALEM MAN, THE [STONES OF POWER 3]  GEMMELL, DAVID  P,SFWA-P
JESTER AT SCAR, THE [DUMAREST 5]  TUBB, E. C.  C/P,D81610
JESUS CHRISTS  LANGGUTH, A. J.  C/P,P
JESUS FACTOR, THE  CORLEY, EDWIN  P
JESUS INCIDENT, THE [SHIP 2]  HERBERT, FRANK/RANSOM, BILL  C/P,P,H,SFWA
JESUS ON MARS  FARMER, PHILIP JOSE  C/P:2,P
JESUS VIDEO, DAS  ESCHBACH, ANDREAS  FP
JETMAN MEETS THE MAD MADAM  SAVAGE, HARDLEY  SR-L
JEWEL-HINGED JAW, THE  DELANY, SAMUEL R.  R/REF-P,C/REF-H:2
JEWEL IN THE SKULL, THE [RUNESTAFF 1]  MOORCOCK, MICHAEL  C/P:2,P
JEWEL OF ARWEN, THE  BRADLEY, MARION ZIMMER  B
JEWEL OF BAS, THE  BRACKETT, LEIGH  TD22,SFWA-TD22
JEWEL OF EQUILIBRANT, THE [WHEEL 1]  FRANKOS, STEVEN  P
JEWEL OF FIRE, THE [WESTRIA 7]  PAXSON, DIANA L.  C/P,P
JEWEL OF JARHEN [@CAP KENNEDY #5]  KERN, GREGORY  P
JEWEL OF SEVEN STARS, THE  STOKER, BRAM  C/P,VLP
JEWEL OF THARN [@BLADE #3]  LORD, JEFFREY  C/P:2,P
JEWEL OF THE MOON  KOTZWINKLE, WILLIAM  H
JEWEL OF TURMISH, THE [@FORGOTTEN REALMS CITIES]  ODOM, MEL  P
JEWELS OF APTOR, THE  DELANY, SAMUEL R.  DF173
JEWELS OF APTOR, THE (RESTORED)  DELANY, SAMUEL R.  C/P:2,P
JEWELS OF ELSEWHEN, THE  WHITE, TED  C/P,P
JEWELS OF ELVISH, THE [JEWELS OF ELVISH 1]  BERBERICK, NANCY VARIAN  P
JEWELS OF THE DRAGON  WOLD, ALLEN L.  P
JHEREG [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,P
JIGSAW WOMAN, THE  ANTIEAU, KIM  LP
JIMGRIM  MUNDY, TALBOT  C/P,SR-P,H
JINGO [DISCWORLD 21/DISCWORLD WATCHMEN 4]  PRATCHETT, TERRY  H:2,P
JINIAN FOOTSEER [TRUE GAME 4]  TEPPER, SHERI S.  C/P,SFWA-P,P
JINIAN STAR-EYE [TRUE GAME 6]  TEPPER, SHERI S.  C/P,SFWA-P,P
JINX HIGH [DIANA TREGARDE]  LACKEY, MERCEDES  C/P:2,P
JINX ON A TERRAN INHERITANCE [HOBART FLOYT 2]  DALEY, BRIAN  C/P:2,P
JIREL OF JOIRY  MOORE, C. L.  H:2,P
JITTERBUG  MCQUAY, MIKE  P
JITTERBUG PERFUME  ROBBINS, TOM  H,P
JIU SAN [DOC SAVAGE #107]/BLACK, BLACK WITCH, THE [DOC SAVAGE 108] ROBESON, KENNETH  @P
JIZZLE  WYNDHAM, JOHN  SR-P
JOAN-OF-ARC REPLAY, THE  BARBET, PIERRE  C/P:2,P
JOANNE, THE UNPREDICTABLE [@PHENWICK #8]  KIMBROUGH, KATHERYN  P
JOB: A COMEDY OF JUSTICE  HEINLEIN, ROBERT A.  C/P,P,H
JOE GOSH  DE HAVEN, TOM  H
JOE HALDEMAN  GORDON, JOAN  R/REF-P
JOE MAUSER, MERCENARY FROM TOMORROW  REYNOLDS, MACK/BANKS, MICHAEL  C/P:2,P
JOG RUMMAGE  WRIGHT, GRAHAME  P
JOHAN [@MAGIC: THE GATHERING LEGENDS 1]  EMERY, CLAYTON  P
JOHN BIRCH COLORING BOOK, THE  COHEN, MICHAEL  CX

TITLE  AUTHOR  CODE
JOHN CARSTAIRS: SPACE DETECTIVE  LONG, FRANK BELKNAP  SR-P,H
JOHN CARTER OF MARS [BARSOOM #11]  BURROUGHS, EDGAR RICE  C/P:2,@P
JOHN CARTER OF MARS (ORIGINAL) [BARSOOM 11]  BURROUGHS, EDGAR RICE  SR-L
JOHN COLLIER READER, THE  COLLIER, JOHN  H:2
JOHN DOUGH AND THE CHERUB  BAUM, L. FRANK  LP
JOHN GAYTHER’S GARDEN AND THE STORIES TOLD THEREIN  STOCKTON, FRANK R.  H
JOHN GRIMES: LIEUTENANT OF THE SURVEY SERVICE [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: RESERVE COMMODORE [GRIMES]  CHANDLER, A. BERTRAM  H:2
JOHN GRIMES: RIM COMMANDER [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: RIM RUNNER [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: SURVEY CAPTAIN [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN GRIMES: TRAMP CAPTAIN [GRIMES]  CHANDLER, A. BERTRAM  H
JOHN SCHOENHERR SF CHECKLIST, A  STEVER, DAVID/BOYAJIAN, JERRY  R/IND:2
JOHN THE BALLADEER [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  P,SFWA-P
JOHN VARLEY READER, THE  VARLEY, JOHN  H
JOHN W. CAMPBELL AWARDS 5, THE  MARTIN, GEORGE R. R.  HA,LPA,SFWA
JOHN W. CAMPBELL LETTERS 1, THE  CAMPBELL, JOHN W. JR.  R/REF-P
JOHNNY AND THE DEAD [JOHNNY MAXWELL 2]  PRATCHETT, TERRY  H
JOHNNY MAXWELL TRILOGY, THE [JOHNNY MAXWELL 1,2,3]  PRATCHETT, TERRY  H
JOHNNY MNEMONIC  BISSON, TERRY  P
JOHNNY ZED  BETANCOURT, JOHN GREGORY  C/P,P
JOINING OF THE STONE, THE [YNELL 5]  MURPHY, SHIRLEY ROUSSEAU  H,P
JOKER, THE  MALAQUAIS, JEAN  P
JOKERS WILD [@WILD CARDS #3]  MARTIN, GEORGE R. R.  C/PA,PA
JOKERTOWN SHUFFLE [@WILD CARDS #9]  MARTIN, GEORGE R. R.  C/PA,PA
JONAH KIT, THE  WATSON, IAN  H,P
JONAS MCFEE, A.T.P.  SARGENT, SARAH  SFWA
JONATHAN LIVINGSTON SEAGULL  BACH, RICHARD  P
JONATHAN SEGAL CHICKEN  WEINSTEIN, SOL/ALBRECHT, HOWARD  [C/RAN-P]
JONATHAN STRANGE & MR. NORRELL  CLARKE, SUSANNA  H
JONDELLE [DUMAREST 10]  TUBB, E. C.  C/P,P
JONGOR FIGHTS BACK  WILLIAMS, ROBERT MOORE  P
JONGOR OF LOST LAND  WILLIAMS, ROBERT MOORE  C/P,P
JONUTA RISING! [@SPACEWAYS #13]  CLEVE, JOHN  P
JORIS OF THE ROCK [NEUSTRIAN CYCLE 2]  BARRINGER, LESLIE  LP
JOSHUA SON OF NONE  FREEDMAN, NANCY  C/P,P,H
JOURNAL OF NICHOLAS THE AMERICAN, THE  KENNEDY, LEIGH  C/P:2,LP,H
JOURNALS OF THE PLAGUE YEARS  SPINRAD, NORMAN  LP
JOURNEY  RANDALL, MARTA  C/P:2,P
JOURNEY [@FIVE WORLDS #2]  SARRANTONIO, AL  P
JOURNEY ACROSS THREE WORLDS  MIR PUBLISHERS  HA
JOURNEY BEHIND THE WIND  WRIGHTSON, PATRICIA  C/P,P
JOURNEY BETWEEN WORLDS  ENGDAHL, SYLVIA LOUISE  H
JOURNEY BEYOND THE BEYOND [MAGIC LANDS]  STANEK, ROBERT  LP
JOURNEY BEYOND TOMORROW  SHECKLEY, ROBERT  C/P:2,P
JOURNEY FROM FLESH  YERMAKOV, NICHOLAS  C/P,P
JOURNEY INTO DARKNESS  LONG, FRANK BELKNAP  C/P:2,P
JOURNEY INTO SPACE  CHILTON, CHARLES  SR-P,C/P
JOURNEY INTO TERROR [@PLANET OF THE APES #3]  EFFINGER, GEORGE ALEC  C/P,P
JOURNEY INTO THE GATEWAY OF DIMENSIONS [TAVAN 1]  BURTON, JOSH MICHAEL  LP
JOURNEY INTO THE VOID [SOVEREIGN STONE 3]  WEIS, MARGARET/HICKMAN, TRACY  H
JOURNEY INTO THE VOID (PROOF) [SOVEREIGN STONE 3]  WEIS, MARGARET/HICKMAN, TRACY  SR-LP
JOURNEY OF NIELS KLIM TO THE WORLD UNDERGROUND  HOLBERG, LUDVIG  LP
JOURNEY OF THE OCEANAUTS  WOLFE, LOUIS  C/P,P
JOURNEY OUTSIDE  STEELE, MARY Q.  H
JOURNEY TO ANOTHER STAR  ELWOOD, ROGER  SR-HA
JOURNEY TO APRILIOTH  KERNAGHAN, EILEEN  C/P:2,P
JOURNEY TO FUSANG  SANDERS, WILLIAM  P
JOURNEY TO IXTLAN  CASTANEDA, CARLOS  [C/RAN-P]
JOURNEY TO JUPITER [@DIG ALLEN #3]  GREENE, JOSEPH  H
JOURNEY TO JUPITER  WALTERS, HUGH  H
JOURNEY TO MARS  POPE, GUSTAVUS W.  LP
JOURNEY TO MEMBLIAR [NAPHAR 2]  BAKER, SHARON  C/P,P
JOURNEY TO MESHARRA [@RO-LAN #3]  SIROTA, MIKE  P
JOURNEY TO RILLA [DR. BONES]  WYLDE, THOMAS  P
JOURNEY TO THE CENTER [ASGARD 1]  STABLEFORD, BRIAN M.  H,P
JOURNEY TO THE CENTER OF THE EARTH, A (BRADLEY TRANSLATION)  VERNE, JULES  C/P,P
JOURNEY TO THE CENTRE OF THE EARTH, A (EVANS TRANSLATION)  VERNE, JULES  H
JOURNEY TO THE CENTRE OF THE EARTH, A (UNCREDITED TRANSLATION) VERNE, JULES  H:2,P
JOURNEY TO THE EARTH’S INTERIOR, A  GARDNER, MARSHALL B.  H
JOURNEY TO THE FLAME  MONACO, RICHARD  P
JOURNEY TO THE STARS  JASTROW, ROBERT  [C/RAN-H]
JOURNEY TO THE STARS  RAYER, FRANCIS G.  H
JOURNEY TO THE UNDERGROUND WORLD [ERIC CARSTAIRS 1]  CARTER, LIN  P

TITLE  AUTHOR  CODE
JOURNEY TO YESTERDAY [STACY GORDON 1]  SHIPLETT, JUNE LUND  P
JOURNEY UNDER THE SEA  MONTGOMERY, R. A.  C/P,P
JOURNEYS  ALLEN, VERGIE  P
JOURNEYS OF FRODO  STRACHEY, BARBARA  R/REF-P
JOURNEYS TO THE TWILIGHT ZONE  SERLING, CAROL  HA,SFWA-P
JOUST [DRAGON JOUSTERS 1]  LACKEY, MERCEDES  H:2
JOVAH’S ANGEL [SAMARIA 2]  SHINN, SHARON  H,P
JOY DEVICE, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  P
JOY MACHINE, THE [STAR TREK]  GUNN, JAMES E.  C/P,P
JOY MAKERS, THE  GUNN, JAMES E.  C/P:2,P
JOY WAGON, THE  HADLEY, ARTHUR T.  C/P,P
JOYCE OF THE SECRET SQUADRON  WINTERBOTHAM, RUSS  H
JOYCE, THE BELOVED [@PHENWICK #27]  KIMBROUGH, KATHERYN  P
JOYLEG  MOORE, WARD/DAVIDSON, AVRAM  C/P,P,H
J.R R. TOLKIEN’S THE FELLOWSHIP OF THE RING  MORRISON, LOUISE D.  R/REF-P
J.R.R. TOLKIEN CALENDAR 1998  CALENDAR  CAL
J.R.R. TOLKIEN UNIVERSITY  BISCHOFF, DAVID F.  LP
JUBILEE  DANN, JACK  H
JUDAS GLASS, THE  CADNUM, MICHAEL  H
JUDAS GOAT, THE  EDGLEY, LESLIE  SR-DD013
JUDAS MANDALA, THE  BRODERICK, DAMIEN  C/P:2,P
JUDAS MANDATE, THE  EGLETON, CLIVE  P
JUDAS ROSE, THE [NATIVE TONGUE 2]  ELGIN, SUZETTE HADEN  C/P:2,P
JUDAS UNCHAINED [COMMONWEALTH 2]  HAMILTON, PETER F.  H
JUDGE [@WORLD OF DARKNESS PREDATOR & PREY #2]  FLEMING, GHERBOD  P
JUDGE DREDD [JUDGE DREDD]  BARRETT, NEAL JR.  P
JUDGES OF HADES, THE  HOCH, EDWARD D.  SR-P,C/P
JUDGMENT [@MAGIC: THE GATHERING ODYSSEY 3]  MCDERMOTT, WILL  P
JUDGMENT DAY  JENSEN, JANE  P
JUDGMENT DAY  SHARMAN, NICK  P
JUDGMENT DAY  WINSLOW, PAULINE GLEN  P
JUDGMENT IN DEATH [EVE DALLAS]  ROBB, J. D.  P
JUDGMENT NIGHT  DOYLE, DEBRA/MACDONALD, JAMES D.  P
JUDGMENT NIGHT  MOORE, C. L.  C/P,P
JUDGMENT NIGHT [@BUREAU 13]  POLLOTTA, NICK  LP
JUDGMENT NIGHT (COLLECTIONS)  MOORE, C. L.  C/P:2,P
JUDGMENT OF DRAGONS, A  GOTLIEB, PHYLLIS  C/P:3,P
JUDGMENT OF EVE, THE  PANGBORN, EDGAR  C/P:2,P
JUDGMENT ON JANUS [JANUS 1/DIPPLE 2]  NORTON, ANDRE  H:2,P
JUDSON’S EDEN  LAUMER, KEITH  C/P:2,P
JULES VERNE: A BIOGRAPHY  JULES-VERNE, JEAN  R/REF-H
JULES VERNE COMPANION, THE  HAINING, PETER  R/REF-H
JULES VERNE: INVENTOR OF SCIENCE FICTION  COSTELLO, PETER  SFWA
JULES VERNE REDISCOVERED  EVANS, ARTHUR B.  R/REF-H
JULES VERNE: THE MAN WHO INVENTED THE FUTURE  BORN, FRANZ  R/REF-P
JULIUS CAESAR IS ALIVE AND WELL  GREENFIELD, IRVING A.  P
JUMBEE AND OTHER UNCANNY TALES  WHITEHEAD, HENRY S.  H
JUMP PAY [LUCKY 13TH]  SHELLEY, RICK  P
JUMPER [JUMPER 1]  GOULD, STEVEN C.  H:2,P
JUMPING OFF THE PLANET [DINGILLIAD 1]  GERROLD, DAVID  H,P
JUNCTION  DANN, JACK  P
JUNGLE, THE  DRAKE, DAVID  H:2,P
JUNGLE ASSAULT [STAR COMMANDOS 8]  GRIFFIN, P. M.  P
JUNGLE JUNGLE  NUETZEL, CHARLES  SR-P
JUNGLE OF HORRORS, THE [LONE WOLF 8]  DEVER, JOE  P
JUNGLE OF STARS, A  CHALKER, JACK L.  C/P:2,P
JUNGLE TALES OF TARZAN [TARZAN #6]  BURROUGHS, EDGAR RICE  H:2,@P
JUNIPER, GENTIAN, AMD ROSEMARY  DEAN, PAMELA  LP
JUNIPER TIME  WILHELM, KATE  H:2,SFWA,P
JUNKTION [@NECROMUNDA]  FARRER, MATTHEW  P
JUPITER  BOVA, BEN  H,P
JUPITER  POHL, CAROL/POHL, FREDERIK  C/PA,PA
JUPITER EXPERIMENT, THE  MOON, MARGARET/MOON, MAURINE  P
JUPITER LEGACY, THE  HARRISON, HARRY  C/P:2,P
JUPITER MAGNIFIED  ROBERTS, ADAM  SR-LP
JUPITER PLAGUE, THE  HARRISON, HARRY  P,SFWA-P
JUPITER PROJECT  BENFORD, GREGORY  C/P:2,P
JUPITER THEFT, THE  MOFFITT, DONALD  C/P:2,P
JUPITER WAR, THE [@WAR YEARS #3]  FAWCETT, BILL  PA
JUPITER’S DAUGHTER  HYMAN, TOM  H,SFWA-P
JURASSIC PARK [JURASSIC PARK 1]  CRICHTON, MICHAEL  H:2,P
JURGEN  CABELL, JAMES BRANCH  H:2,SR-P
JURY [@WORLD OF DARKNESS PREDATOR & PREY #4]  FLEMING, GHERBOD  P
JUST COMPENSATION [SHADOWRUN 19]  CHARRETTE, ROBERT N.  C/P:2,P

TITLE  AUTHOR  CODE
JUST DETERMINATION, A [PAUL SINCLAIR 1]  HEMRY, JOHN G.  P
JUST FOR YOU: THE WORLD’S FIRST SCIENCE FICTION BIRTHDAY CARD  HALLMARK  WRD
JUST WAR [DOCTOR WHO NEW ADVENTURES]  PARKIN, LANCE  C/P:2,P
JUSTICE AND HER BROTHERS [JUSTICE 1]  HAMILTON, VIRGINIA  C/P,P
JUSTICE INC. [@AVENGER #1]  ROBESON, KENNETH  C/P:2,P
JUXTAPOSITION [APPRENTICE ADEPT #3]  ANTHONY, PIERS  C/P,@P,H

TITLE  AUTHOR  CODE
K-9 AND COMPANY [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P,C/P
K-9 CORPS [K-9 CORPS 1]  VON GUNDEN, KENNETH  P
K-MACHINES [PLAYERS IN THE CONTEST OF WORLDS 2]  BRODERICK, DAMIEN  LP
K-PAX [K-PAX 1]  BREWER, GENE  P
KA OF GIFFORD HILLARY, THE  WHEATLEY, DENNIS  C/P,P
KADUNA MEMORIES  MCKINNEY, JACK  C/P:2,P
KAETI & COMPANY  ROBERTS, KEITH  LP
KAGONESTI, THE [@DRAGONLANCE LOST HISTORIES #1]  NILES, DOUGLAS  P
KAHLESS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  P
KAI LUNG UNROLLS HIS MAT [KAI LUNG]  BRAMAH, ERNEST  C/P,P
KAI LUNG’S GOLDEN HOURS [KAI LUNG]  BRAMAH, ERNEST  H:2,P
KAJIRA OF GOR [GOR 19]  NORMAN, JOHN  P
KAL JERICO [WARHAMMER]  RENNIE, GORDON/KOPINSKI, KARL  CX
KALEIDOSCOPE  TURTLEDOVE, HARRY  C/P,P
KALEIDOSCOPE CENTURY  BARNES, JOHN  H,P
KALEVIDE, THE  GOBLE, LOU  C/P:2,P
KALGAN THE GOLDEN  TUBB, E. C.  LP
KALIF’S WAR, THE [REGIMENT 3]  DALMAS, JOHN  C/P,SFWA-P,P
KALIMANTAN  SHEPARD, LUCIUS  LP
KALIN [DUMAREST 4]  TUBB, E. C.  C/P,D42800,P
KALKI  VIDAL, GORE  H,P
KALPA IMPERIAL  GORODISCHER, ANGELICA  LP
KAMIKAZE L’AMOUR  KADREY, RICHARD  H
KAMONGO  SMITH, HOMER W.  H
KAMPF IM UTOPIA  BERNING, FRANK W.  FP
KAMPUS  GUNN, JAMES E.  P
KANDAR  BULMER, KENNETH  C/P,P
KANE’S ODYSSEY  CLINTON, JEFF  SR-P
KAR-CHEE REIGN, THE  DAVIDSON, AVRAM  DG574
KAR-CHEE REIGN, THE/ROGUE DRAGON  DAVIDSON, AVRAM  C/P,P
KAR KABALLA [ARWIN 3]  SMITH, GEORGE H.  D42900
KAR KALIM  CHRISTIAN, DEBORAH  C/P,P
KARAVANS [KARAVANS 1]  ROBERSON, JENNIFER  H
KARL GRIER  TRACY, LOUIS  H
KARMA  DARNAY, ARSEN  H,P
KARMA AFFAIR, THE  DARNAY, ARSEN  P
KARMA CORPS, THE  BARRETT, NEAL JR.  C/P:2,P
KARMA MACHINE, THE  DAVIDSON, MICHAEL  C/P,P
KASPAR’S BOX [THREE KINGS 3]  CHALKER, JACK L.  P
KASPAR’S BOX (PROOF) [THREE KINGS 3]  CHALKER, JACK L.  SR-LP
KATE, THE CURIOUS [@PHENWICK #13]  KIMBROUGH, KATHERYN  P
KATHERINE, THE RETURNED [@PHENWICK #30]  KIMBROUGH, KATHERYN  P
KAVIN’S WORLD  MASON, DAVID  C/P,P
KAZ THE MINOTAUR [@DRAGONLANCE HEROES II #1]  KNAAK, RICHARD A.  C/P:2,P
KECKSIES AND OTHER TWILIGHT TALES  BOWEN, MARJORIE  H
KEDRIGERN AND THE CHARMING COUPLE [KEDRIGERN]  MORRESSY, JOHN  P
KEDRIGERN IN WONDERLAND [KEDRIGERN 3]  MORRESSY, JOHN  P
KEEP, THE [ADVERSARY 1]  WILSON, F. PAUL  H,P
KEEP OF FIRE, THE [LAST RUNE 2]  ANTHONY, MARK  P
KEEP ON THE BORDERLANDS [@GREYHAWK]  EMERSON, RU  P
KEEP THE GIRAFFE BURNING  SLADEK, JOHN  SR-P
KEEPER  HOLLY, JOAN HUNTER  SR-P
KEEPER MARTIN’S TALE [RUIN MIST 1]  STANEK, ROBERT  LP
KEEPER OF CATS [SKYLA 3]  BOYER, ELIZABETH H.  C/P,P
KEEPER OF THE CHILDREN  HALLAHAN, WILLIAM H.  P
KEEPER OF THE CITY [@GUARDIANS OF THE THREE 2]  MORWOOD, PETER/DUANE, DIANE  C/P,P
KEEPER OF THE ISIS LIGHT, THE [ISIS 1]  HUGHES, MONICA  H,P
KEEPER OF THE KEYS [CYCLE OF FIRE 2]  WURTS, JANNY  C/P:2,P
KEEPER OF THE KING  BENNETT, NIGEL/ELROD, P. N.  C/P,P
KEEPER OF THE UNIVERSE  LAWRENCE, LOUISE  H
KEEPERS OF EDANVANT [SWORD AND CIRCLET 1]  DOUGLAS, CAROLE NELSON  H:2,SFWA-P,P
KEEPERS OF THE FLAME [@WARHAMMER STEFAN KUMANSKY 3]  MCINTOSH, NEIL  P
KEEPERS OF THE GATE  SPRUILL, STEVEN G.  C/P,SFWA,P,H
KEEPERS OF THE SUN [@DEATHLANDS]  AXLER, JAMES  P
KEEPER’S PRICE, THE [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
KEEPING IT REAL [@QUANTUM GRAVITY #1]  ROBSON, JUSTINA  LP
KEEPING TIME  BEAR, DAVID  H,P
KELLORY THE WARLOCK  CARTER, LIN  SFWA
KELLY COUNTRY  CHANDLER, A. BERTRAM  C/P,P
KELWIN  BARRETT, NEAL JR.  C/P,P
KEMLO AND THE CRAZY PLANET  ELIOT, E. C.  H
KEMLO AND THE STAR MEN  ELIOT, E. C.  SR-P
KEMLO AND THE ZONES OF SILENCE  ELIOT, E. C.  H
KENDER, GULLY DWARVES, AND GNOMES [@DRAGONLANCE TALES #2]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA

TITLE  AUTHOR  CODE
KENDERMORE [@DRAGONLANCE PRELUDES 2]  KIRCHOFF, MARY  C/P:2,P
KENSHO  SCHMIDT, DENNIS  C/P:2,P
KENT MONTANA AND THE ONCE AND FUTURE THING [KENT MONTANA 3] FENN, LIONEL  C/P,P

KENT MONTANA AND THE REALLY UGLY THINGS FROM MARS [KENT MONTANA 1] FENN, LIONEL

C/P:2,SFWA-P:2,P

KENT MONTANA AND THE REASONABLY INVISIBLE MAN [KENT MONTANA 2] FENN, LIONEL

C/P,P

KESRICK [TERRA MAGICA 1]  CARTER, LIN  C/P:2,P
KESTREL, THE [WESTMARK 2]  ALEXANDER, LLOYD  P
KEVIN BROWNE’S NIGHTALES  TIGGES, JOHN  LP
KEY FOR THE NONESUCH, A [FADING WORLDS 1]  GRAVEL, GEARY  C/P:2,P
KEY OF ICE AND STEEL [KEYS TO PARADISE 3]  VARDEMAN, ROBERT E.  P,SFWA-P
KEY OF THE KEPLAN, THE [WITCH WORLD SECRETS 1]  NORTON, ANDRE/MCCONCHIE, LYN  P
KEY OUT OF TIME [TIME TRADERS 4]  NORTON, ANDRE  C/P:2,P
KEY TO CHROMA [CHROMAGIC 2]  ANTHONY, PIERS  H
KEY TO CONFLICT  GRYPHON, TALIA  P
KEY TO DESTINY [CHROMAGIC 3]  ANTHONY, PIERS  H
KEY TO HAVOC [CHROMAGIC 1]  ANTHONY, PIERS  LP
KEY TO IRUNIUM, THE [KEYS TO THE DIMENSIONS 2]  BULMER, KENNETH  DH020
KEY TO LIBERTY [CHROMAGIC 4]  ANTHONY, PIERS  H
KEY TO THE GREAT GATE, THE  GOTTLIEB, HINKO  H
KEY TO THE GRIFFON’S LAIR [@KNIGHTS OF THE SILVER DRAGON #9]  RANSOM, CANDICE  LP
KEY TO VENUDINE, THE [KEYS TO THE DIMENSIONS 3]  BULMER, KENNETH  DH065
KEYS TO D’ESPERANCE, THE  BRENCHLEY, CHAZ  B
KHAI OF ANCIENT KHEM  LUMLEY, BRIAN  P
KHAI OF KHEM  LUMLEY, BRIAN  LP
KHALED  CRAWFORD, F. MARION  C/P:2,P
KHAN’S PERSUASION, THE  FELICE, CYNTHIA  C/P:2,P
KHARE — CITYPORT OF TRAPS [@SORCERY! #2]  JACKSON, STEVE  [C/RAN-P:2]
KHI TO FREEDOM  MAYHAR, ARDATH  C/P,P
KHYBER CONNECTION, THE [TIMEWARS 6]  HAWKE, SIMON  P
KHYREN  KAPLAN, ALINE BOUCHER  C/P:2,SFWA-P,P
KID FROM MARS, THE  FRIEND, OSCAR J.  H
KIDNAPPER, THE  BLOCH, ROBERT  P,SFWA-P
KIF STRIKE BACK, THE [CHANUR 3]  CHERRYH, C. J.  C/P,P,H
KILL CITY [@ENFORCER #3]  SUGAR, ANDREW  P
KILL DEADLINE [@ENFORCER #4]  SUGAR, ANDREW  P
KILL RATIO  MORRIS, JANET/DRAKE, DAVID  P
KILL RIFF, THE  SCHOW, DAVID J.  SR-LP,SFWA-P
KILL STATION [@SPACE COPS 1]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
KILL TEAM [WARHAMMER 40K LAST CHANCERS 2]  THORPE, GAV  P
KILL THE DEAD  LEE, TANITH  C/P,P
KILL THE EDITOR  ROBINSON, SPIDER  SR-LP
KILL THE POOR  ROSE, JOEL  SFWA
KILLASHANDRA [CRYSTAL SINGER 2]  MCCAFFREY, ANNE  C/P,P,H,SFWA
KILLBIRD  HUGHES, ZACH  P
KILLDOZER! [COMPLETE STORIES OF THEODORE STURGEON 3]  STURGEON, THEODORE  H
KILLER  DRAKE, DAVID/WAGNER, KARL EDWARD  C/P:2,SFWA-P,P
KILLER CRABS  SMITH, GUY N.  P
KILLER PINE [MATTHEW DILKE 2]  GUTTERIDGE, LINDSAY  H,P
KILLER PLANTS, THE  ELWOOD, ROGER  SR-HA
KILLER PLANTS OF BINAARK [@BLADE #33]  LORD, JEFFREY  P
KILLER SPORES [MAN FROM ATLANTIS 3]  WOODLEY, RICHARD  P
KILLER STATION  CAIDIN, MARTIN  P
KILLER THING, THE  WILHELM, KATE  H:2,P
KILLER TO COME  MERWIN, SAM JR.  H:2,GN4.22
KILLERBOWL  WOLF, GARY K.  SFWA
KILLERBOWL T-SHIRT  DOUBLEDAY BOOKS  SFWA-VLP
KILLERS FROM HYPERSPACE [PERRY RHODAN 120]  VOLTZ, WILLIAM  SR-LP
KILLER’S KEEP [@WARLORD #6]  FROST, JASON  P
KILLER’S TOWN [PHANTOM 9]  FALK, LEE  P
KILLING BLOW [@STAR TREK ERRAND OF VENGEANCE #2]  RYAN, KEVIN  P
KILLING BONE, THE [GUARDIANS (SAXON) 1]  SAXON, PETER  P
KILLING DANCE, THE [ANITA BLAKE 6]  HAMILTON, LAURELL K.  C/P:2,P
KILLING FIELDS, THE [@BATTLETECH #45/CAPELLAN SOLUTION 2]  COLEMAN, LOREN L.  P
KILLING FIELDS [WARKEEP 2030]  KASNER, MICHAEL  P
KILLING GROUND [DOCTOR WHO MISSING ADVENTURES]  LYONS, STEVE  C/P:2,P
KILLING GROUND  POWE, BRUCE  P
KILLING MACHINE, THE [DEMON PRINCES 2]  VANCE, JACK  C/P:2,P

KILLING MACHINE (VIE), THE [@VANCE INTEGRAL EDITION #23/ DEMON PRINCES 2] VANCE, JACK

SR-H

KILLING ME SOFTLY  DOZOIS, GARDNER  PA,LPA
KILLING OF WORLDS, THE [SUCCESSION 2]  WESTERFELD, SCOTT  P
KILLING SONG, THE [@EBERRON DRAGON BELOW #3]  BASSINGTHWAITE, DON  P

The MITSFS Titledex  September 24, 2007  Page 195
TITLE  AUTHOR  CODE

KILLING STAR, THE PELLEGRINO, CHARLES/ ZEBROWSKI, GEORGE

C/P:2,P

KILLING TIME (ORIGINAL) [STAR TREK]  VAN HISE, DELLA  C/P:2,SR-P
KILLING TIME (REVISED) [STAR TREK]  VAN HISE, DELLA  C/P:3,P
KILLOBYTE  ANTHONY, PIERS  C/P,SFWA-P:2,P
KILN, THE [FIRE-US]  ARMSTRONG, JENNIFER/BUTCHER, NANCY  H
KILN PEOPLE  BRIN, DAVID  H:2,P
KILTERNAN LEGACY, THE  MCCAFFREY, ANNE  C/P,P
KIN OF ATA ARE WAITING FOR YOU, THE  BRYANT, DOROTHY  SFWA
KINDA [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
KINDLING  FARREN, MICK  H
KINDLY ONES, THE [@SANDMAN #9]  GAIMAN, NEIL  H,CX
KINDLY ONES, THE  SCOTT, MELISSA  H:2,SFWA-P,P
KINDRED  BUTLER, OCTAVIA E.  C/P,P,H
KINDRED SPIRITS [DRAGONLANCE MEETINGS 1]  ANTHONY, MARK/PORATH, ELLEN  C/P:2,P
KINDRED SPIRITS  BRENNERT, ALAN  P
KINDRED SPIRITS  ELLIOT, JEFFREY M.  LP
KING, THE  BARTHELME, DONALD  H
KING, THE [RODRIGO OF CALEDON 2]  FEINTUCH, DAVID  H
KING, THE [@TELNARIAN 3]  NORMAN, JOHN  P
KING & RAVEN  JAMES, CARY  SFWA
KING AND EMPEROR [HAMMER AND THE CROSS 3]  HARRISON, HARRY/HOLM, JOHN  C/P,P,H
KING AND QUEEN OF SWORDS, THE [@OROKON #2]  ARDEN, TOM  P
KING AND RAVEN  JAMES, CARY  P
KING ARTHUR’S DAUGHTER [THREE DAMOSELS 3]  CHAPMAN, VERA  C/P:2,P
KING BENEATH THE MOUNTAIN, A [AWAKENED KING 2]  CHARRETTE, ROBERT N.  P
KING BETRAYED, THE [SWORD,RING AND CHALICE PHERESA 3]  CHESTER, DEBORAH  P
KING BEYOND THE GATE, THE [DRENAI 2]  GEMMELL, DAVID  P
KING CHONDOS’ RIDE [@DARK BORDER #2]  ZIMMER, PAUL EDWIN  C/P:2,P
KING DAVID’S SPACESHIP  POURNELLE, JERRY  C/P,P,H
KING DRAGON  OFFUTT, ANDREW J.  C/P,P
KING IMPERILLED, THE [SWORD,RING AND CHALICE PHERESA 4]  CHESTER, DEBORAH  P
KING IN YELLOW, THE  CHAMBERS, ROBERT W.  SR-H,C/P:3,P
KING IS A FINK, THE  PARKER, BRANT/HART, JOHNNY  [C/RAN-P]
KING JAVAN’S YEAR [DERYNI HEIRS 2]  KURTZ, KATHERINE  H
KING KELSON’S BRIDE [DERYNI KELSON 4]  KURTZ, KATHERINE  H
KING KOBOLD [WARLOCK 2]  STASHEFF, CHRISTOPHER  C/P:2,P
KING KOBOLD REVIVED [WARLOCK 2]  STASHEFF, CHRISTOPHER  C/P:2,P
KING KONG  LOVELACE, DELOS W.  C/P:2,P
KING KONG  WALLACE, EDGAR/COOPER, MERIAN C.  P
KING KULL [KULL]  HOWARD, ROBERT E./CARTER, LIN  C/P,P
KING MAKER, THE [DOC SAVAGE #80]  ROBESON, KENNETH  C/P:2,@P
KING MUST DIE, THE  RENAULT, MARY  P
KING MUST DIE, THE/BULL FROM THE SEA, THE  RENAULT, MARY  H
KING OF ARGENT  PHILLIFENT, JOHN T.  C/P:2,P
KING OF CATS, THE [@SABRINA THE TEENAGE WITCH SALEM #4]  DUBOWSKI, MARK  P
KING OF DREAMS, THE [MAJIPOOR PRESTIMION 3]  SILVERBERG, ROBERT  H,P
KING OF ELFLAND’S DAUGHTER, THE  DUNSANY, LORD  C/P:2,P
KING OF EOLIM, THE  JONES, RAYMOND F.  SR-P
KING OF FOXES [CONCLAVE OF SHADOWS 2]  FEIST, RAYMOND E.  H
KING OF FOXES (PROOF) [CONCLAVE OF SHADOWS 2]  FEIST, RAYMOND E.  SR-LP
KING OF INFINITE SPACE, A [NEAR SPACE]  STEELE, ALLEN  H:2,P
KING OF LIGHT AND SHADOWS, THE [@OMARAN #3]  COLE, ADRIAN  P
KING OF MORNING, QUEEN OF DAY  MCDONALD, IAN  C/P,P
KING OF SATAN’S EYES, THE [LINCOLN BLACKTHORNE 1]  MARSH, GEOFFREY  H,SFWA,SFWA-P
KING OF SHADOWS [SWORD IN EXILE 2]  MARSTON, ANN  P
KING OF SLEEP, THE [@WATCHERS #2]  MOR, CAISEAL  P
KING OF TERROR, THE [DOCTOR WHO PAST DOCTORS]  TOPPING, KEITH  C/P,P
KING OF THE DEAD [@RAVENLOFT]  DEWEESE, GENE  P
KING OF THE DEAD [@BUFFY LOST SLAYER #3]  GOLDEN, CHRISTOPHER  C/P:2
KING OF THE DEAD [LENS OF THE WORLD 2]  MACAVOY, R. A.  H:2
KING OF THE FOURTH PLANET  WILLIAMS, ROBERT MOORE  DF149
KING OF THE GREY  KNAAK, RICHARD A.  C/P,P
KING — OF THE KHYBER RIFLES  MUNDY, TALBOT  H,P
KING OF THE MURGOS [MALLOREON 2]  EDDINGS, DAVID  H:2,P
KING OF THE NORTH  TURTLEDOVE, HARRY  P
KING OF THE SCEPTER’D ISLE  CONEY, MICHAEL G.  H:2,P
KING OF THE SEA  BICKERTON, DEREK  P
KING OF THE SLAVERS [@SPACEWAYS #19]  CLEVE, JOHN  C/P,P
KING OF THE STARS  KELLEY, LEO P.  SR-P
KING OF THE SUN, THE  MCFEDRIES, A. A.  P
KING OF THE SWORDS, THE [CORUM 3]  MOORCOCK, MICHAEL  C/P:2,P
KING OF THE VAGABONDS [BAROQUE CYCLE 1B]  STEPHENSON, NEAL  P
KING OF THE WOOD  ROBERTS, JOHN MADDOX  H,SFWA,SFWA-P,P

TITLE  AUTHOR  CODE
KING OF THE WORLD’S EDGE  MUNN, H. WARNER  C/P:2,P
KING OF YS, THE, VOLUME I [KING OF YS 1,2]  ANDERSON, POUL/ANDERSON, KAREN  H
KING OF YS, THE, VOLUME II [KING OF YS 3,4]  ANDERSON, POUL/ANDERSON, KAREN  H
KING OF ZUNGA [@BLADE #12]  LORD, JEFFREY  P
KING PINCH [@FORGOTTEN REALMS NOBLES #1]  COOK, DAVID  P
KING RAT  MIEVILLE, CHINA  LP
KING SOLOMON’S MINES [ALLAN QUATERMAIN]  HAGGARD, H. RIDER  C/P,P,H
KING SOLOMON’S MINES  WEBB, JEAN FRANCIS  SR-P
KINGDOM ALLIANCE [RUIN MIST 2]  STANEK, ROBERT  LP
KINGDOM COME [@TRAVELER #2]  DRUMM, D. B.  C/P,P
KINGDOM COME  MAGGIN, ELLIOT S.  C/P,P
KINGDOM COME  WAID, MARK/ROSS, ALEX  CX
KINGDOM OF CAGES  ZETTEL, SARAH  H,P
KINGDOM OF KEVIN MALONE, THE  CHARNAS, SUZY MCKEE  H
KINGDOM OF ROYTH [@BLADE #9]  LORD, JEFFREY  P
KINGDOM OF SHADOW, THE [@DIABLO #3]  KNAAK, RICHARD A.  C/P,P
KINGDOM OF SORROW [@KING’S QUEST 2]  MORR, KENYON  P
KINGDOM OF SUMMER [GWALCHMAI 2]  BRADSHAW, GILLIAN  H:2,P
KINGDOM OF THE BLIND [@JUDGE DREDD]  BISHOP, DAVID  P
KINGDOM OF THE CATS, THE  GOTLIEB, PHYLLIS  C/P:2,P
KINGDOM OF THE GRAIL  TARR, JUDITH  LP
KINGDOM OF THE SPIDERS  HURWOOD, BERNHARDT J.  P
KINGDOM RIVER [SNOWFALL 2]  SMITH, MITCHELL  P
KINGDOMS AND ELVES OF THE REACHES, THE [KEEPER MARTIN 1]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES II, THE [KEEPER MARTIN 2]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES III, THE [KEEPER MARTIN 3]  STANEK, ROBERT  LP
KINGDOMS AND ELVES OF THE REACHES IV, THE [KEEPER MARTIN 4]  STANEK, ROBERT  LP
KINGDOM’S FURY [@STARFIST 8]  SHERMAN, DAVID/CRAGG, DAN  P
KINGDOMS OF ELFIN  WARNER, SYLVIA TOWNSEND  H:2,LP
KINGDOMS OF LIGHT  FOSTER, ALAN DEAN  H
KINGDOMS OF SORCERY  CARTER, LIN  HA,SFWA
KINGDOMS OF TERROR, THE [LONE WOLF 6]  DEVER, JOE  P
KINGDOMS OF THE NIGHT [ANTEROS]  COLE, ALLAN/BUNCH, CHRIS  C/P,P
KINGDOMS OF THE WALL  SILVERBERG, ROBERT  H:2,P
KINGDOMS OF THE WALL (PROOF)  SILVERBERG, ROBERT  SR-LP
KINGDOM’S SWORDS [@STARFIST 7]  SHERMAN, DAVID/CRAGG, DAN  C/P,P
KINGLESS LAND, THE [BAND OF FOUR]  GREENWOOD, ED  P
KINGMAKER’S SWORD [RUNE BLADE 1]  MARSTON, ANN  P
KING’S BLOOD  TARR, JUDITH  LP
KING’S BLOOD FOUR [TRUE GAME 1]  TEPPER, SHERI S.  C/P:2,P
KING’S BUCCANEER, THE [RIFTWAR]  FEIST, RAYMOND E.  H:2,P
KING’S CURE [FANUILH 5]  HOOD, DANIEL  P
KING’S DAMOSEL, THE [THREE DAMOSELS 2]  CHAPMAN, VERA  C/P,P
KING’S DAUGHTER [ATLAN]  GASKELL, JANE  C/P,P
KING’S DEMONS, THE [DOCTOR WHO]  DUDLEY, TERRENCE  SR-P
KING’S DRAGON [CROWN OF STARS 1]  ELLIOTT, KATE  C/P,P,H
KINGS IN HELL [HEROES IN HELL]  CHERRYH, C. J./MORRIS, JANET  C/P,P
KING’S JUSTICE, THE [DERYNI KELSON 2]  KURTZ, KATHERINE  H,P,SFWA
KING’S MAN & THIEF  GOLDEN, CHRISTIE  P
KING’S MATE [GIRL FACTORY 2]  MURPHY, ROBERT FRANKLIN  P
KING’S NAME, THE [TIR TANAGIRI 2]  WALTON, JO  H
KINGS OF INFINITE SPACE  BALCHIN, NIGEL  SR-P,H
KINGS OF THE HIGH FRONTIER [HIGH PILGRIMAGE 1]  KOMAN, VICTOR  H
KING’S OWN, THE [@BORDERLANDS (IVERSTERRE)]  FREEMAN, LORNA  P
KING’S PEACE, THE [TIR TANAGIRI 1]  WALTON, JO  H
KING’S SACRIFICE [@STAR OF THE GUARDIANS #3]  WEIS, MARGARET  C/P:2,P
KING’S SON, MAGIC’S SON  SHERMAN, JOSEPHA  P
KING’S TEST [@STAR OF THE GUARDIANS #2]  WEIS, MARGARET  C/P:2,P
KINGSBANE [IRON ANGEL 3]  MORRESSY, JOHN  P
KINGSLAYER  JAMES, L. DEAN  P
KINSHIP WITH THE STARS  ANDERSON, POUL  C/P:2,P
KINSLAYER WARS, THE [@DRAGONLANCE ELVEN NATIONS #2]  NILES, DOUGLAS  P
KINSMAN  BOVA, BEN  H:2,P
KINSMAN SAGA, THE  BOVA, BEN  C/P,SFWA-P,P,H
KIOGA OF THE UNKNOWN LAND [KIOGA 4]  CHESTER, WILLIAM L.  C/P,P
KIOGA OF THE WILDERNESS [KIOGA 2]  CHESTER, WILLIAM L.  P
KIPLING’S SCIENCE FICTION  KIPLING, RUDYARD  H
KIPPER’S GAME  EHRENREICH, BARBARA  H
KIRINYA [CHAGA 2]  MCDONALD, IAN  P
KIRINYAGA (NOVEL)  RESNICK, MIKE  H
KIRINYAGA (SHORT STORY)  RESNICK, MIKE  B
KIRLIAN QUEST [CLUSTER #3]  ANTHONY, PIERS  C/P:2,@P
KISS BEFORE LOVING, A  REYNOLDS, MACK  P
KISS KISS  DAHL, ROALD  SR-P

The MITSFS Titledex  September 24, 2007  Page 197
TITLE  AUTHOR  CODE
KISS OF DARKNESS [@CHARMED]  ALEXANDER, BRANDON  P
KISS OF GLORY, THE  BOYLAN, GRACE DUFFIE  H
KISS OF SHADOWS, A [MEREDITH GENTRY 1]  HAMILTON, LAURELL K.  H
KISSING SIN [RILEY JENSON 2]  ARTHUR, KERI  P
KISSING THE BEEHIVE  CARROLL, JONATHAN  LP
KITEMAN  REYNOLDS, ALFRED  P
KITEMAN OF KARANGA  REYNOLDS, ALFRED  H
KITEWORLD  ROBERTS, KEITH  C/P,SFWA,P,H
KIT’S BOOK [WINTER WORLD 3]  MILLS, C. J.  C/P,P
KITSUNE [@TIME HUNTER 4]  CATTON, JOHN PAUL  LP
KITTATINNY  RUSS, JOANNA  VLP
KITTY AND THE MIDNIGHT HOUR [KITTY NORVILLE 1]  VAUGHN, CARRIE  P
KITTY GOES TO WASHINGTON [KITTY NORVILLE 2]  VAUGHN, CARRIE  P
KITTY TAKES A HOLIDAY [KITTY NORVILLE 3]  VAUGHN, CARRIE  P
KLEINE HOBBIT, DER  TOLKIEN, J. R. R.  FH
KLEINE HOBBNIX, DER  ROBERTS, A. R. R. R.  FP
KLINGON DICTIONARY, THE [STAR TREK]  OKRAND, MARC  R/REF-P
KLINGON DICTIONARY, THE (WITH 1992 ADDENDUM) [STAR TREK]  OKRAND, MARC  R/REF-P,C/REF-H
KLINGON GAMBIT, THE [STAR TREK]  VARDEMAN, ROBERT E.  C/P:3,P
KNAVE AND THE GAME [GERALD KNAVE]  JANIFER, LAURENCE M.  H,SFWA
KNAVE IN HAND [GERALD KNAVE 2]  JANIFER, LAURENCE M.  C/P,P
KNAVE OF DREAMS  NORTON, ANDRE  C/P,P,H
KNEE-DEEP IN THE DEAD [@DOOM #1]  AB HUGH, DAFYDD/LINAWEAVER, BRAD  C/P,P
KNEE DEEP IN THE THUNDER  MOON, SHELIA  LP
KNEES UP MOTHER EARTH [BRENTFORD 7/WITCHES OF CHISWICK 2]  RANKIN, ROBERT  H
KNIFE OF DREAMS [WHEEL OF TIME 11]  JORDAN, ROBERT  H:2
KNIGHT, THE [WIZARD KNIGHT 1]  WOLFE, GENE  H
KNIGHT AMONG KNAVES, A [AWAKENED KING 3]  CHARRETTE, ROBERT N.  P
KNIGHT AND KNAVE OF SWORDS, THE [FAFHRD]  LEIBER, FRITZ  C/P,SFWA-P:2,P,H
KNIGHT FANTASTIC  GREENBERG, MARTIN H./HELFERS, JOHN  PA
KNIGHT LIFE [ARTHUR PENN 1]  DAVID, PETER  P,SFWA-P
KNIGHT MOVES [@PALADINS #2]  ROSENBERG, JOEL  H
KNIGHT MOVES  WILLIAMS, WALTER JON  C/P,SFWA-P,P
KNIGHT OF DARKNESS [LORDS OF AVALON 1]  MACGREGOR, KINLEY  H
KNIGHT OF DELUSIONS  LAUMER, KEITH  C/P:2,SFWA-P,P
KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY]  ANDERSON, POUL  H:2,P
KNIGHT OF GHOSTS AND SHADOWS [ERIC BANYON 1]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
KNIGHT OF ILLUSION [ENDLESS QUEST 33]  KIRCHOFF, MARY  C/P,P
KNIGHT OF SHADOWS [AMBER 9]  ZELAZNY, ROGER  C/P,H,@P,SFWA
KNIGHT OF THE BLACK ROSE [RAVENLOFT]  LOWDER, JAMES  P
KNIGHT OF THE DEMON QUEEN [WINTERLANDS 3]  HAMBLY, BARBARA  P
KNIGHT OF THE SWORDS, THE [CORUM 1]  MOORCOCK, MICHAEL  C/P:2,P
KNIGHT OF THE WORD, A [VOID & WORD 2]  BROOKS, TERRY  H,P
KNIGHT OR KNAVE [OAK,YEW,ASH,AND ROWAN 2]  NORTON, ANDRE/MILLER, SASHA  H
KNIGHT RIDER  LARSON, GLEN A./HILL, ROGER  P
KNIGHT, THE HARP AND THE MAIDEN, THE  BUSH, ANNE KELLEHER  C/P,P
KNIGHTHOOD OF THE DRAGON [@DRAGONMASTER #2]  BUNCH, CHRIS  LP
KNIGHT’S CASTLE  EAGER, EDWARD  LP
KNIGHT’S DAWN [@RED PAVILIONS #1]  HUNTER, KIM  LP
KNIGHTS OF CAWDOR, THE [ELUNDIUM 5]  JEFFERIES, MIKE  P
KNIGHTS OF DARK RENOWN  GEMMELL, DAVID  P
KNIGHTS OF MADNESS  HAINING, PETER  PA

KNIGHTS OF THE BLACK EARTH, THE [STAR OF THE GUARDIANS MAG FORCE 7] WEIS, MARGARET/PERRIN, DON

C/P:2,P

KNIGHTS OF THE CROWN [@DRAGONLANCE WARRIORS #1]  GREEN, ROLAND J.  P
KNIGHTS OF THE MORNINGSTAR [QUANTUM LEAP]  RAWN, MELANIE  C/P:2,P
KNIGHTS OF THE ROSE [@DRAGONLANCE WARRIORS #5]  GREEN, ROLAND J.  P
KNIGHTS OF THE SWORD [@DRAGONLANCE WARRIORS #3]  GREEN, ROLAND J.  P
KNOT IN THE GRAIN, A  MCKINLEY, ROBIN  P
KNOWN AND THE UNKNOWN, THE  CARTIER, EDD  A
KOBAYASHI MARU, THE [STAR TREK]  ECKLAR, JULIA  C/P:2,P
KODEZEICHEN GROSSER BAR  SCHEER, K. H.  FH
KOIEC COROLLARY, THE  GHOLSTON, HOMER N.  P
KOMARR [MILES VORKOSIGAN 8]  BUJOLD, LOIS MCMASTER  H:2,P
KONETS VECHNOSTI  ASIMOV, ISAAC  FH
KONGA  OWEN, DEAN  SR-P
KONRAD [WARHAMMER KONRAD 1]  FERRING, DAVID  C/P,LP
KONRAD SAGA, THE [@WARHAMMER KONRAD 1,2,3]  FERRING, DAVID  H
KONTAKTE MIT DER EWIGKEIT [@PERRY RHODAN SILBERBANDE #72]  PABEL MOEWIG KG  FPR
KONZIL DER SIEBEN [@PERRY RHODAN SILBERBANDE #74]  PABEL MOEWIG KG  FPR
KOP [JUNO 1]  HAMMOND, WARREN  H
KOREN  LUKEMAN, TIM  H:2,P
KORPS DER VERZWEIFELTEN  SCHEER, K. H.  FP
KOTHAR AND THE CONJURER’S CURSE [KOTHAR 4]  FOX, GARDNER F.  C/P,P

TITLE  AUTHOR  CODE
KOTHAR AND THE DEMON QUEEN [KOTHAR 3]  FOX, GARDNER F.  P
KOTHAR AND THE WIZARD SLAYER [KOTHAR 5]  FOX, GARDNER F.  C/P,P
KOTHAR — BARBARIAN SWORDSMAN [KOTHAR 1]  FOX, GARDNER F.  P
KOTHAR OF THE MAGIC SWORD [KOTHAR 2]  FOX, GARDNER F.  P
KRAKATIT  CAPEK, KAREL  H
KRAKEN WAKES, THE  WYNDHAM, JOHN  SR-P,C/P:2
KRAZY KAT  CANTOR, JAY  H:2
KRILOV CONTINUUM, THE [GUARDIANS 1]  LOVEGROVE, J. M. H.  LP
KRIS LONGKNIFE: DEFIANT [KRIS LONGKNIFE 3]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: DESERTER [KRIS LONGKNIFE 2]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: MUTINEER [KRIS LONGKNIFE 1]  SHEPHERD, MIKE  P
KRIS LONGKNIFE: RESOLUTE [KRIS LONGKNIFE 4]  SHEPHERD, MIKE  P
KRISPOS OF VIDESSOS [KRISPOS 2]  TURTLEDOVE, HARRY  C/P:2,P
KRISPOS RISING [KRISPOS 1]  TURTLEDOVE, HARRY  C/P:2,P
KRISPOS THE EMPEROR [KRISPOS 3]  TURTLEDOVE, HARRY  P
KROKODIL TEARS [DARK FUTURE]  YEOVIL, JACK  LP
KRONDOR: TEAR OF THE GODS [RIFTWAR LEGACY 3]  FEIST, RAYMOND E.  P
KRONDOR: THE ASSASSINS [RIFTWAR LEGACY 2]  FEIST, RAYMOND E.  P
KRONDOR: THE BETRAYAL [RIFTWAR LEGACY 1]  FEIST, RAYMOND E.  H:2,P
KRONE EXPERIMENT, THE  WHEELER, J. CRAIG  P
KRONK  COOPER, EDMUND  P
KRONO  HARNESS, CHARLES L.  C/P,P
KRONOS CONDITION, THE  DEVENPORT, EMILY  P
KROTONS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
KROZAIR OF KREGEN [DRAY PRESCOT 14]  AKERS, ALAN BURT  P
KRULL  FOSTER, ALAN DEAN  C/P,P
KRUTON INTERFACE, THE [U.S.S. REPULSE 1]  DE CHANCIE, JOHN  C/P,P
KRYTOS TRAP, THE [@STAR WARS X-WING 3]  STACKPOLE, MICHAEL A.  C/P,P
KUBLA KHAN CAPER, THE  PRATHER, RICHARD S.  SFWA-P
KULDESAK  COWPER, RICHARD  SR-P,H,SFWA
KULL [KULL]  HOWARD, ROBERT E.  C/P,P
KULL: EXILE OF ATLANTIS [KULL]  HOWARD, ROBERT E.  H
KUNDALINI EQUATION, THE  BARNES, STEVEN  C/P,SFWA-P,P
KURSAAL [DOCTOR WHO EIGHTH DOCTOR]  ANGHELIDES, PETER  C/P,P
KURT SINGER’S GHOST OMNIBUS  SINGER, KURT  SR-PA
KURT VONNEGUT: FANTASIST OF FIRE AND ICE  GOLDSMITH, DAVID H.  R/REF-P
KURT VONNEGUT, JR.  REED, PETER J.  R/REF-P

KURT VONNEGUT:

THE GOSPEL FROM OUTER SPACE OR YES WE HAVE NO NIRVANAS MAYO, CLARK

R/REF-P,C/REF-P

KUSHIEL’S AVATAR [KUSHIEL 3]  CAREY, JACQUELINE  H,P
KUSHIEL’S CHOSEN [KUSHIEL 2]  CAREY, JACQUELINE  H,P
KUSHIEL’S DART [KUSHIEL 1]  CAREY, JACQUELINE  H,P
KUSHIEL’S JUSTICE [KUSHIEL 5]  CAREY, JACQUELINE  H
KUSHIEL’S SCION [KUSHIEL 4]  CAREY, JACQUELINE  H
KYRIK AND THE LOST QUEEN [KYRIK 4]  FOX, GARDNER F.  P
KYRIK AND THE WIZARD’S SWORD [KYRIK 3]  FOX, GARDNER F.  P
KYRIK FIGHTS THE DEMON WORLD [KYRIK 2]  FOX, GARDNER F.  P
KYRIK: WARLOCK WARRIOR [KYRIK 1]  FOX, GARDNER F.  P

TITLE  AUTHOR  CODE
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE  BUDRYS, ALGIS  C/PA,PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 2  BUDRYS, ALGIS  C/PA,PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 3  BUDRYS, ALGIS  C/PA,PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 4  BUDRYS, ALGIS  C/PA,PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 5  BUDRYS, ALGIS  PA,SFWA-P
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 8  BUDRYS, ALGIS  C/PA,PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 9  WOLVERTON, DAVE  PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 11  WOLVERTON, DAVE  PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 13  WOLVERTON, DAVE  PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 14  WOLVERTON, DAVE  PA
L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 18  BUDRYS, ALGIS  PA
LA-BAS  HUYSMANS, J. K.  LP
LA CIVILTA DEI SOLARI  SPINRAD, NORMAN  FP
LABORATORY X  SHAW, DAVID  SR-P
LABYRINTH [@ALIENS 7]  PERRY, S. D.  P
LABYRINTH [QUESTIONER 1]  SCHMIDT, DENNIS  P
LABYRINTH  SMITH, A. C. H.  P
LABYRINTH GATE, THE  RASMUSSEN, ALIS A.  C/P:2,SFWA-P,P
LABYRINTH KEY, THE  HENDRIX, HOWARD V.  LP
LABYRINTH OF DREAMS, THE [@G.O.D. INC. #1]  CHALKER, JACK L.  C/P:2,SFWA-P,P
LABYRINTH OF EVIL [@STAR WARS]  LUCENO, JAMES  H
LABYRINTH OF LIES  BERRY, ADRIAN  H
LABYRINTH OF NIGHT  STEELE, ALLEN  C/P:2,SFWA-P,P
LABYRINTH OF WORLDS [@STAR REQUIEM #4]  COLE, ADRIAN  C/P,SFWA-P,P
LABYRINTHS  BORGES, JORGE LUIS  LP
LACEY AND HIS FRIENDS  DRAKE, DAVID  C/P:2,P
LAD AND THE LION, THE  BURROUGHS, EDGAR RICE  C/P:2,P
LAD OF THE GAD, THE  GARNER, ALAN  H
LADDER IN THE SKY, THE  WOODCOTT, KEITH  DF141
LADIES’ DAY  BLOCH, ROBERT  SR-DB0080
LADIES OF MANDRIGYN, THE [SUNWOLF AND STARHAWK 1]  HAMBLY, BARBARA  C/P:3,P
LADY, THE  MCCAFFREY, ANNE  H,P
LADY AND THE TIGER, THE [TAYLOR’S ARK 3]  NYE, JODY LYNN  C/P,P
LADY BLADE, LORD FIGHTER  GREEN, SHARON  C/P,P
LADY COTTINGTON’S PRESSED FAIRY BOOK  TURNER PUBLISHING  A,H
LADY COTTINGTON’S PRESSED FAIRY CALENDAR 1997  COTTINGTON, LADY  CAL
LADY CRYMSYN [@VAMPIRE FILES 9]  ELROD, P. N.  H
LADY EL  STARLIN, JIM/GRAZIUNAS, DIANA  C/P,SFWA-P,P
LADY FROM VENUS, THE  RADCLIFFE, GARNETT  H
LADY IN GIL [LADY IN GIL 1]  BRADLEY, REBECCA  C/P,P
LADY IN THE LOCH, THE  SCARBOROUGH, ELIZABETH ANN  P
LADY INTO FOX  GARNETT, DAVID  H
LADY KNIGHT [@PROTECTOR OF THE SMALL 4]  PIERCE, TAMORA  H
LADY OF AVALON  BRADLEY, MARION ZIMMER  H
LADY OF DARKNESS [WESTRIA 2]  PAXSON, DIANA L.  C/P,P
LADY OF HAN-GILEN, THE [@AVARYAN RISING #2]  TARR, JUDITH  H,SFWA-P,P
LADY OF HORSES [HORSE GODDESS PEOPLE]  TARR, JUDITH  LP
LADY OF LIGHT [WESTRIA 1]  PAXSON, DIANA L.  P
LADY OF MAZES  SCHROEDER, KARL  H:2
LADY OF MERCY [SUNDERED 3]  SAGARA, MICHELLE  P
LADY OF MERCY [SUNDERED 3]  WEST, MICHELLE  LP
LADY OF POISON [@FORGOTTEN REALMS]  CORDELL, BRUCE R.  P
LADY OF SHERWOOD  ROBERSON, JENNIFER  LP
LADY OF THE BEES  SWANN, THOMAS BURNETT  C/P:2,P
LADY OF THE FOREST  ROBERSON, JENNIFER  P
LADY OF THE HAVEN [EMPIRE PRINCESS 1]  DIAMOND, GRAHAM  P
LADY OF THE SERPENTS, THE [@VAMPYRICON 2]  CLEGG, DOUGLAS  H
LADY OF THE SHROUD, THE  STOKER, BRAM  P
LADY OF THE SNOWMIST, THE [WAR OF THE GODS ON EARTH 3]  OFFUTT, ANDREW J.  C/P:2,P
LADY OF THE SORROWS, THE [BITTERBYNDE 2]  DART-THORNTON, CECILIA  C/P,P
LADY OF THE TRILLIUM [TRILLIUM 4]  BRADLEY, MARION ZIMMER  C/P,P,H
LADY OR THE TIGER AND OTHER STORIES, THE  STOCKTON, FRANK R.  H,P
LADY PAIN [LADY IN GIL 3]  BRADLEY, REBECCA  P
LADY SATIVA  LAURIA, FRANK  P
LADY SLINGS THE BOOZE [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P,P,H
LADY WITH THE X-RAY EYES, THE  MINKOV, SVETOSLAV  H
LADYHAWKE  VINGE, JOAN D.  C/P:2,P
LADYLORD  MILLER, SASHA  C/P,P,H
LAERTIAN GAMBLE, THE [STAR TREK DS9]  SHECKLEY, ROBERT  P
LAGRANGE FIVE  REYNOLDS, MACK  C/P:2,P
LAGRANGISTS, THE  REYNOLDS, MACK  P
LAID IN THE FUTURE  GRAY, ROD  SR-L
LAIR  HERBERT, JAMES  P
LAIR OF BONES, THE [RUNELORDS 4]  FARLAND, DAVID  H

The MITSFS Titledex  September 24, 2007  Page 200
TITLE  AUTHOR  CODE
LAIR OF THE CYCLOPS  WOLD, ALLEN L.  P
LAKE, THE  COOKE, JOHN PAYTON  P
LAKE, THE  JENSEN, RUBY JEAN  SFWA-P
LAKE OF DEATH, THE [@DRAGONLANCE AGE OF MORTALS]  RABE, JEAN  P
LAKE OF FIRE, THE [DUNGEON 4]  BAILEY, ROBIN WAYNE  C/P:2,SFWA-P,P
LAKE OF THE LONG SUN [LONG SUN 2]  WOLFE, GENE  H,P
LAKE OF THE SUN  WHITEFORD, WYNNE  C/P,P
LALLIA [DUMAREST 6]  TUBB, E. C.  D71082
LAMARCHOS [DIADEM 2]  CLAYTON, JO  C/P:2,P
LAMBDA I AND OTHER STORIES  CARNELL, JOHN  PA,C/PA
LAMIA  GARDINE, MICHAEL  P
LAMMAS FIELD, THE  FISHER, CATHERINE  H,LP
LAMMAS NIGHT  KURTZ, KATHERINE  C/P:2,P
LAMMAS NIGHT  SHERMAN, JOSEPHA  PA,C/PA
LAMP FOR MEDUSA, A  TENN, WILLIAM  SR-DB0077
LAMP FROM THE WARLOCK’S TOMB, THE  BELLAIRS, JOHN  P
LANAGUE CHRONICLES, THE  WILSON, F. PAUL  P,SFWA-P
LANARK  GRAY, ALASDAIR  H,LP
LANCE THROWER, THE [CAMULOD]  WHYTE, JACK  P
LANCELOT DU LETHE  KING, J. ROBERT  P
LANCES OF NENGESDUL  TAYLOR, KEITH  SR-P
LAND, THE  VIGLIANTE, MARY  P
LAND AUS FEUER UND WASSER  DOMINIK, HANS  FP
LAND BEYOND THE GATE, THE [LAND BEYOND THE GATE 1]  ESHBACH, LLOYD ARTHUR  C/P,P
LAND BEYOND THE MAP [KEYS TO THE DIMENSIONS 1]  BULMER, KENNETH  DM111
LAND LEVIATHAN, THE [OSWALD BASTABLE 2]  MOORCOCK, MICHAEL  C/P,SFWA,P,H
LAND OF ALWAYS-NIGHT [DOC SAVAGE #13]  ROBESON, KENNETH  C/P:2,@P
LAND OF DREAMS  BLAYLOCK, JAMES P.  C/P,SFWA,P,H
LAND OF ESA, THE  CHARLES, NEIL  SR-P
LAND OF FEAR, THE [DOC SAVAGE #75]  ROBESON, KENNETH  C/P:2,@P
LAND OF FOAM, THE  YEFREMOV, IVAN  H
LAND OF HIDDEN MEN, THE  BURROUGHS, EDGAR RICE  C/P:2,P
LAND OF LAUGHS, THE  CARROLL, JONATHAN  C/P,P
LAND OF LONG JUJU [DOC SAVAGE #47]  ROBESON, KENNETH  C/P,@P
LAND OF MIST AND SNOW  DOYLE, DEBRA/MACDONALD, JAMES D.  P
LAND OF NO SHADOW, THE  CLAUDY, CARL H.  H:2
LAND OF OZ, THE [OZ #2]  BAUM, L. FRANK  H,@P,@C/P:2
LAND OF TERROR [PELLUCIDAR #6]  BURROUGHS, EDGAR RICE  C/P:2,@P
LAND OF TERROR, THE [DOC SAVAGE #8]  ROBESON, KENNETH  C/P,@P
LAND OF THE DEAD [@WORLD OF DARKNESS YEAR OF THE SCARAB #3]  BATES, ANDREW  P
LAND OF THE GIANTS [LAND OF THE GIANTS 1]  LEINSTER, MURRAY  C/P:2,P
LAND OF THE MINOTAURS [@DRAGONLANCE LOST HISTORIES #4]  KNAAK, RICHARD A.  P
LAND OF THE POSSIBLE, THE  WHITE, MARY ALICE  P
LAND OF UNREASON  PRATT, FLETCHER/DE CAMP, L. SPRAGUE  C/P:2,P
LAND THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 1]  BURROUGHS, EDGAR RICE  C/P:2,P
LAND THAT TIME FORGOT, THE/MOON MAID, THE  BURROUGHS, EDGAR RICE  LP

LAND THAT TIME FORGOT, THE (OMNIBUS EDITION)

[LAND THAT TIME FORGOT 1,2,3] BURROUGHS, EDGAR RICE

H:2

LAND UNDER ENGLAND  O’NEILL, JOSEPH  SR-P,H:2
LANDO CALRISSIAN ADVENTURES, THE [STAR WARS]  SMITH, L. NEIL  P
LANDO CALRISSIAN AND THE FLAMEWIND OF OSEON [STAR WARS]  SMITH, L. NEIL  P
LANDO CALRISSIAN AND THE MINDHARP OF SHARU [STAR WARS]  SMITH, L. NEIL  C/P,P
LANDO CALRISSIAN AND THE STARCAVE OF THONBOKA [STAR WARS]  SMITH, L. NEIL  P
LAND’S END  POHL, FREDERIK/WILLIAMSON, JACK  H,SFWA-P,P
LANDSCAPE OF DARKNESS, A  BLAIR, JOHN  C/P,P
LANDSLAYER’S LAW [OTHERWORLD]  DEITZ, TOM  P
LANDSLIDE, THE  GILBERT, STEPHEN  H
LANGENSCHEIDT’S GERMAN-ENGLISH ENGLISH-GERMAN DICTIONARY  KLATTE, E./GOLZE, G.  DICT
LANGUAGE OF THE NIGHT, THE  LE GUIN, URSULA K.  SFWA,R/REF-H
LANGUAGES OF MIDDLE EARTH, THE  NOEL, RUTH S.  R/REF-P,C/REF-P
LANGUAGES OF PAO, THE  VANCE, JACK  C/P:3,SFWA-P,P
LANGUAGES OF TOLKIEN’S MIDDLE-EARTH, THE  NOEL, RUTH S.  R/REF-P
LANI PEOPLE, THE  BONE, J. F.  C/P,P
LANKAR OF CALLISTO [JANDAR 6]  CARTER, LIN  C/P,P
LANKHMAR [FAFHRD]  LEIBER, FRITZ  P
LANTERN OF GOD, THE  DALMAS, JOHN  C/P,SFWA-P,P
LANTERN ROAD  ARGO, JOHN  LP
LANTERNS OF GOD, THE [@SCAR #3]  SARTI, RON  P
LAPTOP OF THE GODS  CHIPPINDALE, PETER  LP
LARA CROFT TOMB RAIDER  STERN, DAVE  P
LAREN, DIE [@PERRY RHODAN SILBERBANDE #75]  PABEL MOEWIG KG  FPR
LARISSA  DEVENPORT, EMILY  C/P,P
LARK AND THE WREN, THE [BARDIC VOICES 1]  LACKEY, MERCEDES  C/P:2,P,H
LARK AND THE WREN, THE (EXCERPT) [BARDIC VOICES 1]  LACKEY, MERCEDES  SR-LP

TITLE  AUTHOR  CODE
LARKLIGHT  REEVE, PHILIP  H
LAROUSSE’S FRENCH-ENGLISH ENGLISH-FRENCH DICTIONARY  DUBOIS, MARGUERITE-MARIE, ET AL.  DICT
LARQUE ON THE WING  SPRINGER, NANCY  P
LASERTOWN BLUES [@SAND WARS #2]  INGRID, CHARLES  C/P,P
LASGUN WEDDING [@NECROMUNDA]  MCDERMOTT, WILL  P
LASHER [MAYFAIR WITCHES 2]  RICE, ANNE  C/P,P,H
LASOMBRA [@WORLD OF DARKNESS DARK AGES CLAN #5]  WILSON, DAVID NIALL  P
LAST ADVENTURER: THE LIFE OF TALBOT MUNDY, THE  ELLIS, PETER BERRESFORD  R/REF-H
LAST AERIE, THE [NECROSCOPE 7/VAMPIRE WORLD 2]  LUMLEY, BRIAN  H
LAST ALIEN, THE [@PRIVATE SCHOOL #6]  CHARLES, STEVEN  P
LAST AMAZON, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
LAST AMERICAN, THE [MATTHEWS 2]  BURGAUER, STEVEN  H,LP
LAST AMERICAN, THE  MITCHELL, J. A.  H
LAST AMERICAN, THE [@DOOMSDAY WARRIOR #3]  STACY, RYDER  P
LAST AND FIRST MEN  STAPLEDON, OLAF  SR-H,LP
LAST AND FIRST MEN/LAST MEN IN LONDON  STAPLEDON, OLAF  SR-P,C/P
LAST AND FIRST MEN/STAR MAKER  STAPLEDON, OLAF  H:2,LP
LAST ARABIAN NIGHT, THE [OTHER SINBAD 3]  GARDNER, CRAIG SHAW  C/P,SFWA-P,P
LAST ASTRONAUT, THE  TORRO, PEL  SR-P,C/P
LAST BATTLE, THE [@DRAGONMASTER #3]  BUNCH, CHRIS  LP
LAST BATTLE, THE [NARNIA 7]  LEWIS, C. S.  C/P:2,P
LAST BATTLEMAGE, THE [DRAGON NIMBUS HISTORY 2]  RADFORD, IRENE  P
LAST BREATH, THE  SHAFFER, EUGENE CARL  SR-P
LAST CALL  POWERS, TIM  H:2,P
LAST CALL (PROOF)  POWERS, TIM  SR-P
LAST CALL FOR THE STARS  NUETZEL, CHARLES  H
LAST CALL OF MOURNING, THE  GRANT, CHARLES L.  H,SFWA-P
LAST CASTLE, THE  VANCE, JACK  TD15,DH021, SFWA-TD15
LAST CELT: A BIO-BIBLIOGRAPHY OF ROBERT ERVIN HOWARD, THE  LORD, GLENN  R/IND
LAST CHANCERS, THE [@WARHAMMER 40K]  THORPE, GAV  LP
LAST COIN, THE  BLAYLOCK, JAMES P.  C/P,SFWA-P,P,H
LAST COLONY, THE [OLD MAN’S WAR 3]  SCALZI, JOHN  H
LAST COME THE CHILDREN  HAGBERG, DAVID  SFWA-P
LAST COMING, THE  MOORE, ROBIN/GRIFFIN, GERALD G.  P
LAST COMMAND, THE [STAR WARS]  ZAHN, TIMOTHY  H:2,P
LAST COMMUNION  YERMAKOV, NICHOLAS  C/P:2,P
LAST CONTINENT, THE  COOPER, EDMUND  C/P,P
LAST CONTINENT, THE [DISCWORLD 22/DISCWORLD RINCEWIND 6]  PRATCHETT, TERRY  C/P,P
LAST CRIME, THE  DOMATILLA, JOHN  H,SFWA
LAST DANCER, THE [CONTINUING TIME 3]  MORAN, DANIEL KEYS  C/P,P
LAST DAY, THE  KLEIER, GLENN  P
LAST DAY IN LIMBO [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
LAST DAY OF CREATION, THE  JESCHKE, WOLFGANG  H
LAST DAYS OF ATLANTIS, THE [@PERRY RHODAN #62]  ACE  PA
LAST DAYS OF CHRIST THE VAMPIRE  ECCARIUS, J. G.  P
LAST DAYS OF THE EDGE OF THE WORLD, THE  STABLEFORD, BRIAN M.  P
LAST DAYS OF THE LATE, GREAT STATE OF CALIFORNIA, THE  GENTRY, CURT  C/P,P
LAST DEATHSHIP OFF ANTARES, THE  WATKINS, WILLIAM JOHN  C/P,P
LAST DEFENDER OF CAMELOT, THE  ZELAZNY, ROGER  C/P,P,H
LAST DOOR TO AIYA  GINSBURG, MIRRA  HA
LAST DRAGONLORD, THE [DRAGONLORD LINDEN 1]  BERTIN, JOANNE  H,P
LAST DREAM, THE  DICKSON, GORDON R.  C/P,P
LAST ECHO, THE [INVASION: EARTH]  CAVE, PETER  LP
LAST ENCHANTMENT, THE [ARTHURIAN 3]  STEWART, MARY  H:2,LP,P
LAST ENEMY, THE (DRAFT? MANUSCRIPT)  LONGYEAR, BARRY B.  SR-L
LAST FALL  STOLBOV, BRUCE  H,SFWA
LAST FATHOM, THE  CAIDIN, MARTIN  P
LAST FOURTEEN, THE  BARR, TYRONE C.  SR-P,C/P
LAST FULL MEASURE [STAR TREK ENTERPRISE]  MARTIN, MICHAEL A./MANGELS, ANDY  P
LAST GASP, THE  HOYLE, TREVOR  H,P
LAST GENE, THE  LONGO, CHRIS  P
LAST GREEN TREE, THE  GRIMSLEY, JIM  H
LAST GUARDIAN, THE [@STONES OF POWER #4]  GEMMELL, DAVID  P
LAST GUARDIAN, THE [@WARCRAFT #3]  GRUBB, JEFF  P
LAST HARBOR, THE  FOY, GEORGE  P
LAST HAWK, THE [SKOLIAN EMPIRE 3]  ASARO, CATHERINE  C/P,P,H
LAST HERALD MAGE, THE [LAST HERALD MAGE 1,2,3]  LACKEY, MERCEDES  H:2
LAST HERO, THE [DISCWORLD]  PRATCHETT, TERRY  VLP
LAST HOPE OF EARTH, THE  WRIGHT, LAN  C/P:2,P
LAST HOT TIME, THE  FORD, JOHN M.  H
LAST HUMAN, THE [@KING’S TRAMP #3]  DE HAVEN, TOM  C/P,LP
LAST HUMAN, THE (PROOF) [KING’S TRAMP 3]  DE HAVEN, TOM  SR-LP
LAST HURRAH OF THE GOLDEN HORDE, THE  SPINRAD, NORMAN  H:2,P

The MITSFS Titledex  September 24, 2007  Page 202
TITLE  AUTHOR  CODE
LAST IMMORTAL, THE  JEPPSON, J. O.  C/P,P,H
LAST INCANTATION, THE  SMITH, CLARK ASHTON  C/P,P
LAST KNIGHT OF ALBION, THE  HANRATTY, PETER  LP
LAST LEGENDS OF EARTH, THE [RADIX 4]  ATTANASIO, A. A.  C/P:2,SFWA,LP
LAST LEGION, THE [LAST LEGION 1]  BUNCH, CHRIS  C/P,P
LAST LIGHT OF THE SUN, THE  KAY, GUY GAVRIEL  H:2,LP
LAST MAGICIAN, THE  KELLER, DAVID H.  SR-LP
LAST MAGICIANS, THE  JAKES, JOHN  C/P,P
LAST MAN ALIVE, THE  NEILL, A. S.  H,P

LAST MAN ON EARTH, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

LAST MAN ON EARTH  FANE, BRON  SR-P
LAST MAN RUNNING [DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  C/P:2,P
LAST MARTIAN, THE  STATTEN, VARGO  SR-P
LAST MASTER, THE  DICKSON, GORDON R.  C/P:2,P
LAST MEN IN LONDON  STAPLEDON, OLAF  P
LAST MORTAL MAN, THE [DEATHLESS 1]  MITCHELL, SYNE  P
LAST MOUNTAIN  FLEET, ROBERT C.  P,SFWA-P:2
LAST NIGHT AT THE BRAIN THIEVES BALL  SPENCER, SCOTT  C/P,P
LAST OF LAZARUS, THE  GOLDSTON, ROBERT  P
LAST OF THE DANU’S CHILDREN, THE  RUSH, ALISON  P
LAST OF THE DRAGONS AND SOME OTHERS, THE  NESBIT, E.  P
LAST OF THE GADERENE [DOCTOR WHO PAST DOCTORS]  GATISS, MARK  C/P,P
LAST OF THE RENSHAI, THE [RENSHAI 1]  REICHERT, MICKEY ZUCKER  C/P:2,P
LAST OF THE WILDS [AGE OF THE FIVE 2]  CANAVAN, TRUDI  P
LAST ORDERS  ALDISS, BRIAN W.  C/P:2,P
LAST PLANET, THE [STAR SOLDIERS 2]  NORTON, ANDRE  DD096,C/P:2,P
LAST POST FOR A PARTISAN  EGLETON, CLIVE  P
LAST PRESIDENT, THE  KURLAND, MICHAEL/BARTON, S. W.  P
LAST RAINBOW, THE  GODWIN, PARKE  C/P:2,LP
LAST RECALL, THE [MARKED MAN 2]  INGRID, CHARLES  P
LAST RECKONING, THE [@DESTINY MAKERS #5]  SHUPP, MIKE  C/P,P
LAST REFUGE  SCARBOROUGH, ELIZABETH ANN  C/P,P,H
LAST RESORT, THE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  P
LAST RIDE, THE  HOWARD, ROBERT E.  P
LAST RITES  SMITH, PERRY MICHAEL  H:2
LAST ROUNDUP, THE [@STAR TREK]  GOLDEN, CHRISTIE  P
LAST SHERLOCK HOLMES STORY, THE  DIBDIN, MICHAEL  P
LAST SHIP, THE  BRINKLEY, WILLIAM  H
LAST SPACESHIP, THE  LEINSTER, MURRAY  GN4.25,H,SR-P
LAST STAND [@BOLOS #4]  FAWCETT, BILL  C/PA,PA
LAST STAND, THE [STAR TREK TNG]  FERGUSON, BRAD  C/P:3,P
LAST STAND OF THE DNA COWBOYS, THE  FARREN, MICK  C/P,P
LAST STAR AT DAWN, THE [@LIGHTBRINGER #3]  JOHNSON, OLIVER  H,P
LAST STARFIGHTER, THE  FOSTER, ALAN DEAN  C/P:2,P
LAST STARSHIP FROM EARTH, THE  BOYD, JOHN  H:2,P
LAST SWORD OF POWER [@STONES OF POWER #2]  GEMMELL, DAVID  P
LAST THANE, THE [@DRAGONLANCE CHAOS WAR 2]  NILES, DOUGLAS  P
LAST TRAIN TO LIMBO  PLAYBOY BOOKS  C/PA,PA
LAST TRANSACTION, THE  MALZBERG, BARRY N.  P
LAST UNICORN, THE  BEAGLE, PETER S.  C/P:2,P
LAST WAR, THE  GODFREY, MARTYN  P
LAST WAR [SKYRIDER 3]  MICHAELS, MELISA C.  P,SFWA-P
LAST WARRIOR QUEEN, THE  MACKEY, MARY  C/P,P,H
LAST WEEK’S APOCALYPSE  LAIN, DOUGLAS  LP
LAST WIZARD, THE [WYRDRUNE 8]  HAWKE, SIMON  C/P,P
LAST WIZARD, THE  HUFF, TANYA  C/P,P
LAST WORLD WAR, THE  WARD, DAYTON  P
LAST YGGDRASIL, THE  YOUNG, ROBERT F.  C/P,P
LASTBORN OF ELVINWOOD, THE  HALDEMAN, LINDA  P
LATE GREAT CREATURE, THE  BROWER, BROCK  P
LATE GREAT FUTURE, THE  FITZGERALD, GREGORY/DILLON, JOHN  PA
LATE GREAT PLANET EARTH, THE  LINDSEY, HAL  [C/RAN-P:2]
LATE KNIGHT EDITION  KNIGHT, DAMON  R/REF-H
LATHE OF HEAVEN, THE  LE GUIN, URSULA K.  C/P,SFWA,P,H
LATITUDE ZERO [@DEATHLANDS]  AXLER, JAMES  P
LATTER DAY SCIENCE FICTION 2  URRUTIA, BENJAMIN  SR-LPA
LATTER DAY SCIENCE FICTION 3  URRUTIA, BENJAMIN  SR-LPA

LAUGH OF DEATH, THE [DOC SAVAGE #119]/ KING OF TERROR, THE [DOC SAVAGE 120] ROBESON, KENNETH

@P

LAUGH WAS ON LAZARUS, THE [@AVENGERS 2]  GARFORTH, JOHN  P
LAUGHING CORPSE, THE [ANITA BLAKE 2]  HAMILTON, LAURELL K.  C/P:2,P
LAUGHING GAS  WODEHOUSE, P. G.  P
LAUGHTER IN SPACE  STATTEN, VARGO  SR-P

The MITSFS Titledex  September 24, 2007  Page 203
TITLE  AUTHOR  CODE
LAUGHTER OF CARTHAGE, THE [PYAT 2]  MOORCOCK, MICHAEL  H
LAUGHTER OF DARK GODS, THE [WARHAMMER]  PRINGLE, DAVID  PA
LAURA, THE IMPERILED [@PHENWICK #38]  KIMBROUGH, KATHERYN  P
LAVA [@JOURNEYS OF MCGILL FEIGHAN #3]  O’DONNELL, KEVIN JR.  C/P:2,P
LAVALITE WORLD, THE [WORLD OF TIERS 5]  FARMER, PHILIP JOSE  C/P:2,P
LAVENDER-GREEN MAGIC  NORTON, ANDRE  C/P,P,H
LAVONDYSS  HOLDSTOCK, ROBERT P.  H:2,P
LAW FOR THE STARS, A  MORRESSY, JOHN  SR-P,C/P
LAW OF BECOMING, THE [@JARAN #4]  ELLIOTT, KATE  C/P,P
LAW OF LOVE, THE  ESQUIVEL, LAURA  H
LAW OF SURVIVAL [JANI KILIAN 3]  SMITH, KRISTINE  C/P:2,P
LAW OF THE JUNGLE [@X-MEN]  SMEDS, DAVE  P
LAW OF WAR, THE [BENTON HAWKES 2]  SHATNER, WILLIAM  H,P

LAY DOWN WITH LIONS [@WORLD OF DARKNESS YEAR OF THE SCARAB #2] BATES, ANDREW

P

LAYING THE MUSIC TO REST  SMITH, DEAN WESLEY  C/P,SFWA-P,P
LAYS OF BELERIAND, THE [HISTORY OF MIDDLE-EARTH 3]  TOLKIEN, J. R. R.  H:2,P
LAZARUS  HARTENFELS, JEROME  P
LAZARUS EFFECT, THE [SHIP 3]  HERBERT, FRANK/RANSOM, BILL  C/P,P,H
LAZARUS HEART, THE [@CROW]  BRITE, POPPY Z.  LP
LAZARUS RISING [@STARFIST 9]  SHERMAN, DAVID/CRAGG, DAN  H,P
LAZERWARZ  SHEPHERD, MARK  P
LE MORTE D’ARTHUR (BAINES RENDITION)  MALORY, SIR THOMAS  P
LEADER OF THE RESISTANCE  VERNE, JULES  H
LEAGUE OF GREY-EYED WOMEN, THE  FAST, JULIUS  H
LEAGUE OF THE CRIMSON CRESCENT, THE [WIZARDRY]  REAGEN, JAMES E.  P
LEAKY ESTABLISHMENT, THE  LANGFORD, DAVID  LP
LEAP IN THE DARK  GORDON, DONALD  P
LEAP POINT  KENYON, KAY  P
LEAPING TO THE STARS [DINGILLIAD 3]  GERROLD, DAVID  H
LEARNING FEAR [FEAR 3]  CHEPAITIS, B. A.  P
LEARNING MAZE, THE  ELWOOD, ROGER  HA
LEARNING THE WORLD  MACLEOD, KEN  H

LEAR’S DAUGHTERS KELLOGG, MARJORIE BRADLEY/

ROSSOW, WILLIAM B.

H

LEATHER WINGS [EARTH 2]  VORNHOLT, JOHN  P
LEAVES OF OCTOBER, THE  SAKERS, DON  C/P,SFWA-P,P
LEAVES OF TIME, THE  BARRETT, NEAL JR.  C/P:2,P
LECHEROUS LIMERICKS  ASIMOV, ISAAC  [C/RAN-P]
LEESHORE, THE  REED, ROBERT  H:2,P
LEFT BEHIND [@LEFT BEHIND #1]  LAHAYE, TIM/JENKINS, JERRY B.  H:2,LP
LEFT HAND OF DARKNESS, THE  LE GUIN, URSULA K.  H:3,P
LEFT HAND OF DEATH, THE [@EBERRON LANTERNLIGHT #1]  DEWOLF, PARKER  P
LEFT HAND OF DESTINY BOOK ONE, THE [@STAR TREK DS9]  HERTZLER, J. G./LANG, JEFFREY  P
LEFT HAND OF DESTINY BOOK TWO, THE [@STAR TREK DS9]  HERTZLER, J. G./LANG, JEFFREY  P
LEFT HAND OF THE ELECTRON, THE  ASIMOV, ISAAC  [C/RAN-P]
LEFT-HANDED HUMMINGBIRD, THE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P:2,P
LEGACIES [COREAN 1]  MODESITT, L. E. JR.  H
LEGACIES  SINCLAIR, ALISON  C/P,P
LEGACIES [REPAIRMAN JACK]  WILSON, F. PAUL  P
LEGACIES (PROOF) [COREAN 1]  MODESITT, L. E. JR.  SR-LP
LEGACIES AND LIES [@GAMBLER’S STAR #2]  HOLDER, NANCY  P
LEGACY [EON 0]  BEAR, GREG  H:2,P
LEGACY  BONE, J. F.  SR-P
LEGACY [@SHARING KNIFE 2]  BUJOLD, LOIS MCMASTER  H
LEGACY [@WARHAMMER 40K SHIRA CALPURNIA 2]  FARRER, MATTHEW  P
LEGACY [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:3,P
LEGACY [DOCTOR WHO NEW ADVENTURES]  RUSSELL, GARY  C/P:2,P
LEGACY, THE [FORGOTTEN REALMS LEGACY OF THE DROW 1]  SALVATORE, R. A.  C/P,P,H
LEGACY  SCHMITZ, JAMES H.  C/P:2,P
LEGACY [EARTH FINAL CONFLICT]  SIXBURY, GLENN R.  LP
LEGACY [HEAVEN 1]  VINGE, JOAN D.  BS4
LEGACY [DISINHERITED 2]  WHITE, STEVE  P
LEGACY OF BLOOD [@DIABLO #1]  KNAAK, RICHARD A.  P
LEGACY OF EARTH [@CHILDREN OF THE STARS #3]  COULSON, JUANITA  C/P,P

LEGACY OF HEOROT, THE [HEOROT 1] NIVEN, LARRY/POURNELLE, JERRY/

BARNES, STEVEN

H:2,P

LEGACY OF LEHR, THE  KURTZ, KATHERINE  C/P,P,H
LEGACY OF MOREVI [@ARATHELLEAN WARS #1]  MORRIS, TEE  LP

LEGACY OF OLAF STAPLEDON, THE MCCARTHY, PATRICK A./ELKINS, CHARLES/

GREENBERG, MARTIN H.

R/REF-H

LEGACY OF THE ANCIENTS [@SCAR #2]  SARTI, RON  P
LEGACY OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  PEEL, JOHN  C/P,P
LEGACY OF THE DARK SWORD [DARKSWORD]  WEIS, MARGARET/HICKMAN, TRACY  H

The MITSFS Titledex  September 24, 2007  Page 204
TITLE  AUTHOR  CODE
LEGACY OF THE STARS  GREGORY, JOHN  P
LEGACY OF THE SWORD [CHEYSULI 3]  ROBERSON, JENNIFER  C/P:2,P
LEGACY OF THE WOLF [@CHEYSULI #3,4]  ROBERSON, JENNIFER  P
LEGACY OF VASHNA, THE [LONE WOLF 16]  DEVER, JOE  P
LEGACY OF WOLVES [@EBERRON INQUISITIVES]  ROCKWELL, MARSHEILA  P
LEGEND [DRENAI 1 (REVISED)]  GEMMELL, DAVID  C/P:2,P
LEGEND OF BLACKJACK SAM, THE  HOFFMAN, LEE  P
LEGEND OF DEATHWALKER  GEMMELL, DAVID  P
LEGEND OF HOLDERLY HALL, THE [HOLDERLY HALL 1]  CAMERON, KATE  P
LEGEND OF HUMA, THE [@DRAGONLANCE HEROES #1]  KNAAK, RICHARD A.  C/P:2,P
LEGEND OF LOST EARTH  WALLIS, G. MCDONALD  DF187
LEGEND OF LUKE, THE [REDWALL]  JACQUES, BRIAN  LP
LEGEND OF MIAREE, THE  HUGHES, ZACH  C/P,P
LEGEND OF NIGHTFALL, THE [NIGHTFALL 1]  REICHERT, MICKEY ZUCKER  C/P,P
LEGEND OF THE DUELIST  ETHERIDGE, RUTLEDGE  P,SFWA-P:2
LEGEND OF THE FIREFISH, THE [TROPHY CHASE 1]  POLIVKA, GEORGE BRYAN  LP
LEGEND REBORN, A [ELIXIR OF LIFE 3]  FRANKOS, STEVEN  P
LEGEND THAT WAS EARTH, THE  HOGAN, JAMES P.  H
LEGENDS [ETERNAL ARCHIVES #1]  WEIS, MARGARET  PA
LEGENDS (I) (VOL. 2)  SILVERBERG, ROBERT  C/PA,PA
LEGENDS (I) (VOL.S 1-3)  SILVERBERG, ROBERT  HA
LEGENDS FROM THE END OF TIME [DANCERS AT THE END OF TIME 4]  MOORCOCK, MICHAEL  H,SFWA,P
LEGENDS II  SILVERBERG, ROBERT  HA
LEGENDS OF THE FERENGI [STAR TREK DS9]  QUARK  LP
LEGENDS REBORN  FLINT, KENNETH C.  C/P,P
LEGENDS WALKING [ATHANOR 2]  LINDSKOLD, JANE  P
LEGION [EXORCIST 2]  BLATTY, WILLIAM PETER  H:2
LEGION  GRANT, CHARLES L.  P
LEGION FROM THE SHADOWS [BRAN MAK MORN 2]  WAGNER, KARL EDWARD  C/P:2,SFWA-P,P

LEGION OF SPACE, THE WILLIAMSON, JACK C/P,SFWA-P,SR-H,SR-P,P,

H,GN1.02

LEGION OF THE DAMNED [LEGION OF THE DAMNED 1]  DIETZ, WILLIAM C.  C/P:2,SFWA-P,P
LEGION OF THE DEAD  CAVE, HUGH B.  C/P,P
LEGION OF THE LOST  TORRO, PEL  SR-P
LEGION OF THUNDER [ORCS FIRST BLOOD 2]  NICHOLLS, STAN  LP
LEGION OF TIME, THE  WILLIAMSON, JACK  C/P,SR-H,H
LEGION OF VIDESSOS, THE [VIDESSOS 3]  TURTLEDOVE, HARRY  C/P:2,P
LEGIONNAIRE, THE [@CASCA #11]  SADLER, BARRY  P
LEGIONS [@MAGIC: THE GATHERING ONSLAUGHT 2]  KING, J. ROBERT  P
LEGIONS OF ANTARES [DRAY PRESCOT 25]  AKERS, ALAN BURT  P
LEGIONS OF HELL [HEROES IN HELL]  CHERRYH, C. J.  C/P,P
LEGIONS OF SPACE  LAUMER, KEITH  P
LEGIONS OF STARVATION  STEELE, CURTIS  SR-LPA
LEMOYNE HERITAGE, THE  LONG, LYDA BELKNAP  P
LEMURIAN STONE, THE  HICKMAN, STEPHEN F.  C/P,P
LENINGRAD NIGHTS  JOYCE, GRAHAM  MB001
LENS OF THE WORLD [LENS OF THE WORLD #1]  MACAVOY, R. A.  H:2,@P
LENSMAN FROM RIGEL [LENSMEN 8]  KYLE, DAVID A.  C/P:2,P
LEONARD NIMOY SCIENCE FICTION (DEMO CD)  NIMOY, LEONARD  SR-MM
LEOPARD IN EXILE [CAROLUS REX 2]  NORTON, ANDRE/EDGHILL, ROSEMARY  H,P
LEOPARD LORD  MORLAND, ALANNA  P
LEOPARD’S DAUGHTER  KILLOUGH, LEE  C/P:2,P
LEOPARD’S TOOTH, A  KOTZWINKLE, WILLIAM  P
LERIOS MECCA, THE [DIRSHAN 1]  LANCOUR, GENE  H,SFWA
LERONI OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
LES DOSSIERS D’INTERPOL  BELLEMARE, PIERRE  FP
LESS THAN HUMAN  CLARKE, ROBERT  C/P:2,P
LESSER EVIL [STAR TREK DS9 MISSION GAMMA #4]  SIMPSON, ROBERT  C/P,P
LESSER KINDRED, THE [LANEN KAELAR 2]  KERNER, ELIZABETH  P
LEST DARKNESS FALL  DE CAMP, L. SPRAGUE  GN4.24,P,H:2
LEST DARKNESS FALL/TO BRING THE LIGHT  DE CAMP, L. SPRAGUE/DRAKE, DAVID  P
LEST EARTH BE CONQUERED  LONG, FRANK BELKNAP  C/P:2,P
LEST WE FORGET THEE, EARTH  KNOX, CALVIN M.  DD291
LET OUT THE BEAST  FISCHER, LEONARD  SR-P
LET THE FIRE FALL  WILHELM, KATE  C/P,P,H
LET THE GALAXY BURN [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  LPA
LET THE SPACEMEN BEWARE!  ANDERSON, POUL  DF209
LETHAL EXPOSURE [CRAIG KREIDENT 3]  ANDERSON, KEVIN J./BEASON, DOUG  C/P,P
LETHAL HERITAGE [@BATTLETECH #20/BLOOD OF KERENSKY 1]  STACKPOLE, MICHAEL A.  P,SFWA-P
LETHAL INTERFACE  ODOM, MEL  P
LETHE  SULLIVAN, TRICIA  C/P,P
LETITIA, THE DREAMER [@PHENWICK #35]  KIMBROUGH, KATHERYN  P
LETTER, THE WITCH, AND THE RING, THE  BELLAIRS, JOHN  H,LP
LETTERS FROM ATLANTIS  SILVERBERG, ROBERT  P,SFWA

The MITSFS Titledex  September 24, 2007  Page 205
TITLE  AUTHOR  CODE
LETTERS OF JAMES BRANCH CABELL, THE  CABELL, JAMES BRANCH  R/REF-H
LETTERS OF J.R.R. TOLKIEN, THE  CARPENTER, HUMPHREY  R/REF-H,C/REF-H
LETTERS TO STAR TREK  SACKETT, SUSAN  R/REF-P,C/REF-P:2
LETTERS TO THE AIR FORCE ON UFO’S  ADLER, BILL  [C/RAN-P:2]
LEVEL 7  ROSHWALD, MORDECAI  C/P,P,H
LEVEN IN HET HEELAL  JACKSON, FRANCIS/MOORE, PATRICK  FP
LEVIATHAN [ILLUMINATUS! 3]  SHEA, ROBERT/WILSON, ROBERT ANTON  C/P:2,P
LEVIATHAN 3  AGUIRRE, FORREST/VANDERMEER, JEFF  LPA
LEVIATHAN’S DEEP  CARR, JAYGE  H:2,SFWA,P
LEXICON DER SCIENCE FICTION LITERATUR  ALPERS, FUCHS/HAHN/JESCHKE  R/REF-P
LIADEN UNIVERSE COMPANION: VOLUME I [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  LP
LIAFAIL [TROS OF SAMOTHRACE 3]  MUNDY, TALBOT  SR-P,C/P:2
LIAR’S HOUSE  SHEPARD, LUCIUS  SR-H
LIAR’S OATH [GIRD 2]  MOON, ELIZABETH  C/P:2,SFWA-P,P
LIAR’S PEAK [@WARHAMMER ANGELIKA FLEISCHER 3]  LAWS, ROBIN D.  P
LIAVEK [LIAVEK 1]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: FESTIVAL WEEK [LIAVEK 5]  SHETTERLY, WILL/BULL, EMMA  PA
LIAVEK: PLAYERS OF LUCK, THE [LIAVEK 2]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: SPELLS OF BINDING [LIAVEK 4]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIAVEK: WIZARD’S ROW [LIAVEK 3]  SHETTERLY, WILL/BULL, EMMA  C/PA,PA
LIBERATED, THE [@STAR TREK DS9 REBELS #3]  AB HUGH, DAFYDD  P
LIBERATED FUTURE, THE  HOSKINS, ROBERT  PA,C/PA
LIBERATOR OF JEDD [@BLADE #5]  LORD, JEFFREY  C/P:2,P
LIBERTY’S CRUSADE [@STARCRAFT #1]  GRUBB, JEFF  P
LIBERTY’S WORLD  KILLOUGH, LEE  C/P,P
LIBRO DE ARENA, EL  BORGES, JORGE LUIS  FP
LICENCE DENIED [DOCTOR WHO]  CORNELL, PAUL  R/REF-P
LICENSE INVOKED  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  C/P,P
LICENSE RENEWED [JAMES BOND 15]  GARDNER, JOHN (BRITISH, 1926-)  P
LIE DESTROYER, THE  STATTEN, VARGO  SR-P
LIE TEN NIGHTS AWAKE  VAN THAL, HERBERT  PA
LIEGE-KILLER [PARATWA 1]  HINZ, CHRISTOPHER  C/P:2,P,H
LIES, INC.  DICK, PHILIP K.  LP
LIES OF LOCKE LAMORA, THE  LYNCH, SCOTT  H
LIEUTENANT [DIRIGENT MERCENARY CORPS 2]  SHELLEY, RICK  C/P,P
LIEUTENANT COLONEL [DIRIGENT MERCENARY CORPS 5]  SHELLEY, RICK  C/P,P
LT. LEARY, COMMANDING [RCN 2]  DRAKE, DAVID  C/P,P,H
LIFE  JONES, GWYNETH  LP
LIFE & ADVENTURES OF PETER WILKINS, THE  PALTOCK, ROBERT  H
LIFE AND ADVENTURES OF SANTA CLAUS, THE  BAUM, L. FRANK  C/P,P,H
LIFE DURING WARTIME [BERNICE SUMMERFIELD]  CORNELL, PAUL  HA
LIFE DURING WARTIME  SHEPARD, LUCIUS  C/P,P
LIFE EVERLASTING  KELLER, DAVID H.  SR-H,LP
LIFE EXPECTANCY  KOONTZ, DEAN R.  H
LIFE EXTENSION  PEARSON, DURK/SHAW, SANDY  [C/RAN-H:3]
LIFE FOR KREGEN, A [DRAY PRESCOT 19]  AKERS, ALAN BURT  P
LIFE FOR THE STARS, A  BLISH, JAMES  C/P:2,P
LIFE FORCE  HUGHES, ZACH  C/P,P
LIFE FORM  FOSTER, ALAN DEAN  C/P:2,P
LIFE HUNT [@PERRY RHODAN #43]  ACE  PA,C/PA
LIFE IN DARWIN’S UNIVERSE  BYLINSKY, GENE  SFWA

LIFE IN PIECES, A [BERNICE SUMMERFIELD] STONE, DAVE/SUTTON, PAUL/

LIDSTER, JOSEPH

H

LIFE IN THE DAY OF...AND OTHER SHORT STORIES, A  ROBINSON, FRANK M.  P
LIFE IN THE FUTURE, A [BUCK ROGERS]  CAIDIN, MARTIN  H
LIFE IN THE UNIVERSE  OVENDEN, MICHAEL W.  [C/RAN-P]
LIFE OF SURPRISES, A [BERNICE SUMMERFIELD]  CORNELL, PAUL  HA
LIFE OF THE WORLD TO COME, THE [COMPANY 5]  BAKER, KAGE  H
LIFE ON OTHER WORLDS  JONES, H. SPENCER  [C/RAN-P:2]
LIFE ON THE BORDER [BORDERLANDS 3]  WINDLING, TERRI  PA,C/PA
LIFE PROBE [LIFE PROBE 1]  MCCOLLUM, MICHAEL  C/P:2,P
LIFE SIGNS [STAR TREK]  JENKINS, SUSAN/JENKINS, ROBERT  R/REF-H
LIFE, THE UNIVERSE AND EVERYTHING [HITCHHIKERS 3]  ADAMS, DOUGLAS  C/P,P,H:2
LIFE TRANSFORM MIRROR (CHINESE)  CHEUNG, H. K.  FP
LIFE WITH LANCELOT  PHILLIFENT, JOHN T.  D48245
LIFE WORTH LIVING, A [BERNICE SUMMERFIELD]  GUERRIER, SIMON  HA
LIFEBLOOD [@VAMPIRE FILES #2]  ELROD, P. N.  C/P:2,P
LIFEBOAT  WHITE, JAMES  C/P:2,P
LIFEBOAT EARTH  SCHMIDT, STANLEY  C/P,P
LIFEBURST  WILLIAMSON, JACK  C/P,P,H
LIFEDEATH [DOCTOR WHO]  COWE-SPIGAI, HERETH/NEIGHLY, PATRICK  SR-PA
LIFEHOUSE  ROBINSON, SPIDER  C/P:2,P
LIFEKEEPER  MCQUAY, MIKE  C/P:2,P
LIFELINE  ANDERSON, KEVIN J./BEASON, DOUG  C/P:2,P

TITLE  AUTHOR  CODE
LIFESHIP, THE  DICKSON, GORDON R./HARRISON, HARRY  C/P:2
LIFESHIP, THE  HARRISON, HARRY/DICKSON, GORDON R.  P,SFWA
LIFTED VEIL, THE  WILLIAMS, A. SUSAN  LPA
LIFTER  KILIAN, CRAWFORD  P
LIGHT [LIGHT 1]  HARRISON, M. JOHN  H
LIGHT A LAST CANDLE  KING, VINCENT  C/P,P,H
LIGHT AGES, THE  MACLEOD, IAN R.  H
LIGHT AT THE END, THE  SKIPP, JOHN/SPECTOR, CRAIG  C/P,P
LIGHT AT THE END OF THE UNIVERSE, THE  CARR, TERRY  C/P,P
LIGHT BEARER, THE  NICHOLSON, SAM  P
LIGHT FANTASTIC, THE  HARRISON, HARRY  SFWA,LPA
LIGHT FANTASTIC, THE [DISCWORLD 2/DISCWORLD RINCEWIND 2]  PRATCHETT, TERRY  C/P:2,P,H
LIGHT FROM A LONE STAR  VANCE, JACK  SR-H
LIGHT IN EXILE, THE [NETWORK-CONSORTIUM 1]  FRANKLIN, CHERYL J.  C/P,P
LIGHT MUSIC [NANOTECH]  GOONAN, KATHLEEN ANN  H
LIGHT MY FIRE [AISLING GREY 3]  MACALISTER, KATIE  C/P,P
LIGHT OF EIDON [GUARDIAN-KING 1]  HANCOCK, KAREN  LP
LIGHT OF LILITH, THE  WALLIS, G. MCDONALD  DF108
LIGHT OF MARS  CHARKIN, PAUL  SR-P
LIGHT OF OTHER DAYS, THE  CLARKE, ARTHUR C./BAXTER, STEPHEN  C/P,P,H
LIGHT OF THE WORLD [@CHARMED]  CIENCIN, SCOTT  P
LIGHT ON DARK CORNERS  JEFFERIS, B. G./NICHOLS, J. L.  [C/RAN-H]
LIGHT ON QUESTS MOUNTAIN  KIRCHOFF, MARY/WARD, JAMES M.  C/P,P
LIGHT ON THE SOUND [INQUESTOR 1]  SUCHARITKUL, SOMTOW  C/P,P
LIGHT RAID  WILLIS, CONNIE/FELICE, CYNTHIA  C/P,SFWA-P:2,H,P,SFWA
LIGHT THAT NEVER WAS, THE  BIGGLE, LLOYD JR.  C/P:2,SFWA,P,H
LIGHT-YEARS BENEATH MY FEET, THE  FOSTER, ALAN DEAN  H
LIGHTER SIDE, THE  LAUMER, KEITH  P
LIGHTHOUSE AT THE END OF THE WORLD, THE  VERNE, JULES  H
LIGHTLESS DOME, THE [APOTHEOSIS 1]  HILL, DOUGLAS  P
LIGHTLESS KINGDOM, THE [@UNBALANCED EARTH #2]  WYLIE, JONATHAN  C/P,P
LIGHTNING  KOONTZ, DEAN R.  C/P,P

LIGHTNING FROM A CLEAR SKY:

TOLKIEN, THE TRILOGY, AND THE SILMARILLION MATHEWS, RICHARD

R/REF-P

LIGHTNING THIEF, THE [@PERCY JACKSON & THE OLYMPIANS #1]  RIORDAN, RICK  H
LIGHTNING WORLD  THORPE, TREBOR  SR-P
LIGHTNING’S DAUGHTER [DARK HORSE 2]  HERBERT, MARY H.  P
LIGHTPATHS  HENDRIX, HOWARD V.  P
LIGHTS IN THE SKY ARE STARS, THE  BROWN, FREDRIC  C/P:2,P,H
LIGHTS OF BARBRIN, THE  BURGO, JOSEPH  P
LIGHTS OF ZETAR, THE [BALZAN 3]  MOORE, WALLACE  P
LIGHTSABERS [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P:2,P
LIGHTSTONE, THE [LIGHTSTONE 1]  ZINDELL, DAVID  H
LIGHTSTORM [WEB,THE]  HAMILTON, PETER F.  P
LIGHTWING  HARPER, TARA K.  C/P:2,P
LIKELY STORY, A  WESTLAKE, DONALD E.  SFWA-P
LIKENESS TO VOICES, A  SAVAGE, MARY  P
LILIOM  MOLNAR, FRANZ  H
LILITH  MACDONALD, GEORGE  C/P:2,P
LILITH: A SNAKE IN THE GRASS [@FOUR LORDS OF THE DIAMOND #1]  CHALKER, JACK L.  C/P:2,P
LILLIPUT LEGION, THE [TIMEWARS 9]  HAWKE, SIMON  C/P,P
LILY  WALTERS, R. R.  SFWA-P
LIMBO  WOLFE, BERNARD  C/P,P,H
LIMBO ’90  WOLFE, BERNARD  SR-P
LIMBO OF THE LOST  SPENCER, JOHN  [C/RAN-P]
LIMBO SEARCH  GODWIN, PARKE  C/P,P
LIMBO SYSTEM  COOK, RICK  C/P:2,SFWA-P,P
LIMBRETH GATE, THE [KI AND VANDIEN 3]  LINDHOLM, MEGAN  C/P,P
LIMIERS DE L’INFINI, LES  BARBET, PIERRE  FP
LIMIT OF VISION  NAGATA, LINDA  H
LIMITS  NIVEN, LARRY  H:2,P
LIMITS OF INFINITY: THE AMERICAN SCIENCE FICTION FILM, THE  SOBCHACK, VIVIAN CAROL  R/REF-H
LIN CARTER’S ANTON ZARNAK SUPERNATURAL SLEUTH  PRICE, ROBERT M.  LPA
LINCOLN HUNTERS, THE  TUCKER, WILSON  H,P
LINCOLN’S DREAMS  WILLIS, CONNIE  H:2,P
LINDA LOVELACE FOR PRESIDENT  MARGOLIS, JACK S.  SR-P
LINE BETWEEN, THE  BEAGLE, PETER S.  LP
LINE OF FIRE [@STAR TREK TNG STARFLEET ACADEMY #2]  DAVID, PETER  H,LP
LINE OF LOVE, THE  CABELL, JAMES BRANCH  LP
LINE OF POLITY, THE [POLITY GRIDLINKED 2]  ASHER, NEAL  C/P,P
LINE TO TOMORROW  PADGETT, LEWIS  C/P:2,P
LINERS OF TIME  FEARN, JOHN RUSSELL  H:2
LINEUP, THE  REED, ALISON  SR-L
LINKS  PANATI, CHARLES  P

The MITSFS Titledex  September 24, 2007  Page 207
TITLE  AUTHOR  CODE
LION, THE [FIVE RINGS CLAN WAR 7]  SULLIVAN, STEPHEN D.  P
LION GAME, THE [TELZEY 3]  SCHMITZ, JAMES H.  C/P:2,P
LION IN THE VALLEY  PETERS, ELIZABETH  SFWA-P
LION MEN OF MONGO [FLASH GORDON 1]  STEFFANSON, CON  P
LION MOUNTAIN [@AFRIKORPS #5]  DOLAN, BILL  P
LION OF BOAZ-JACHIN AND JACHIN-BOAZ, THE  HOBAN, RUSSELL  C/P,P
LION OF COMARRE, THE/AGAINST THE FALL OF NIGHT  CLARKE, ARTHUR C.  H:2,P
LION OF FARSIDE, THE [CURTIS MACURDY 1]  DALMAS, JOHN  C/P,P
LION OF JUDAH IN NEVER-NEVER LAND, THE  LINDSKOOG, KATHRYN  R/REF-P
LION OF MACEDON [MACEDON 1]  GEMMELL, DAVID  H,LP
LION OF SENET, THE [SECOND SONS 1]  FALLON, JENNIFER  P
LION ON THARTHEE, A  CALLIN, GRANT  C/P:2,SFWA-P,P
LION RETURNS, THE [CURTIS MACURDY 3]  DALMAS, JOHN  C/P,P
LION, THE WITCH AND THE WARDROBE, THE [NARNIA 1]  LEWIS, C. S.  C/P:2,P
LION THRONE, THE [@CHEYSULI #7,8]  ROBERSON, JENNIFER  P
LION TIME IN TIMBUCTOO  SILVERBERG, ROBERT  SR-H
LIONESS, THE [@DRAGONLANCE AGE OF MORTALS]  BERBERICK, NANCY VARIAN  P
LIONESS, THE [REALMS OF ARKANIA THALIONMEL 1]  KRAMER, INA  P
LIONESS RAMPANT [SONG OF THE LIONESS 4]  PIERCE, TAMORA  H,P
LIONORS  JOHNSON, BARBARA FERRY  P
LION’S BLOOD [LION’S BLOOD 1]  BARNES, STEVEN  C/P,P
LION’S EYE, THE [@ILARIO 1]  GENTLE, MARY  LP
LION’S HEART [FIFTH MILLENNIUM]  WEHRSTEIN, KAREN  C/P,P
LIONS OF AL-RASSAN, THE  KAY, GUY GAVRIEL  H,P
LION’S SOUL [FIFTH MILLENNIUM]  WEHRSTEIN, KAREN  C/P:2,P
LIQUID DIET  TEDFORD, WILLIAM  P
LIQUID MAN, THE  GILFORD, C. B.  P
LIQUID SKY  CARLISLE, ANNE  LP
LIST OF 7, THE  FROST, MARK  H
LISTEN, LISTEN  WILHELM, KATE  C/P,P
LISTEN, MOON!  CLINE, LEONARD  H
LISTEN! THE STARS!  BRUNNER, JOHN  DF215
LISTENERS, THE  GUNN, JAMES E.  H:2,P,SFWA
LISTENERS, THE  PIKE, CHRISTOPHER  P

LITANY OF SH’REEV, THE WATKINS, WILLIAM JOHN/ SNYDER, EUGENE V.

H,SFWA

LITANY OF THE LONG SUN [LONG SUN 1,2]  WOLFE, GENE  H,LP
LITERARY SWORDSMEN AND SORCERERS  DE CAMP, L. SPRAGUE  R/REF-H
LITERARY WEREWOLF, THE  OTTEN, CHARLOTTE F.  LPA
LITERATURE OF TERROR, THE  PUNTER, DAVID  R/REF-H
LITERATURE OF THE SUPERNATURAL  BECK, ROBERT E.  LP
LITTLE, BIG  CROWLEY, JOHN  C/P,LP,H,SFWA
LITTLE BLACK BOOK OF STORIES  BYATT, A. S.  LP
LITTLE BROTHER  MCNEIL, JOHN  H,P
LITTLE COUNTRY, THE  DE LINT, CHARLES  H:2,LP
LITTLE DEATHS  DATLOW, ELLEN  PA
LITTLE FEAR, A  WRIGHTSON, PATRICIA  SFWA
LITTLE FUZZY [FUZZIES 1]  PIPER, H. BEAM  C/P:2,P
LITTLE GODS  PRATT, TIM  LP
LITTLE GREEN MEN  BUCKLEY, CHRISTOPHER  H
LITTLE GREEN MEN, THE [@FLYING NUN #4]  JOHNSTON, WILLIAM  P
LITTLE HELLIAD, THE [HEROES IN HELL]  MORRIS, JANET/MORRIS, CHRIS  C/P,SFWA-P,P
LITTLE HEROES  SPINRAD, NORMAN  H:2,P
LITTLE KNOWLEDGE, A  BISHOP, MICHAEL  H,P
LITTLE LAME PRINCE, THE/ADVENTURES OF A BROWNIE, THE/POOR PRIN MULOCK, MISS  SR-H
LITTLE MONSTERS, THE  ELWOOD, ROGER/GHIDALIA, VIC  C/PA,PA
LITTLE MYTH MARKER [MYTH ADVENTURES 6]  ASPRIN, ROBERT LYNN  C/P,SFWA-P,P,H
LITTLE NIGHT MUSIC, A  HUNT, BARBARA  H
LITTLE PEOPLE, THE  CHRISTOPHER, JOHN  H:2,P
LITTLE PEOPLE!  DANN, JACK/DOZOIS, GARDNER  PA
LITTLE PEOPLE  HOLT, TOM  H
LITTLE PRINCE, THE  SAINT EXUPERY, ANTOINE DE  P
LITTLE RED RIDING HOOD IN THE BIG BAD CITY  GREENBERG, MARTIN H./HELFERS, JOHN  PA
LITTLE THINGS [@BUFFY]  MOESTA, REBECCA  P
LITTLE WHITE HORSE, THE  GOUDGE, ELIZABETH  P
LITTLE WITCH  BENNETT, ANNA ELIZABETH  C/P,P
LITTLEST REBELS, THE [@FLYING NUN #2]  JOHNSTON, WILLIAM  C/P,P
LIVE AND LET DIE [JAMES BOND 2]  FLEMING, IAN  C/P,P
LIVE! FROM PLANET EARTH  EFFINGER, GEORGE ALEC  H
LIVE GIRLS  GARTON, RAY  SR-L
LIVE ROBOTS [COBB ANDERSON 1,2]  RUCKER, RUDY  P
LIVE WITHOUT A NET  ANDERS, LOU  LPA
LIVERY OF EVE, THE  BAIN, F. W.  H

The MITSFS Titledex  September 24, 2007  Page 208
TITLE  AUTHOR  CODE

LIVES AND TIMES OF JERRY CORNELIUS, THE (1979 EDITION)

[CORNELIUS 5] MOORCOCK, MICHAEL

P

LIVES AND TIMES OF JERRY CORNELIUS, THE (2003 EDITION)

[CORNELIUS 5] MOORCOCK, MICHAEL

LP

LIVES OF CHRISTOPHER CHANT, THE [CHRESTOMANCI 4]  JONES, DIANA WYNNE  LP
LIVES OF DAX, THE [@STAR TREK]  PALMIERI, MARCO  LPA
LIVES OF THE MONSTER DOGS  BAKIS, KIRSTEN  H
LIVING BOMB, THE  AVALLONE, MICHAEL  P
LIVING DEAD, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
LIVING DEAD IN DALLAS [SOUTHERN VAMPIRES 2]  HARRIS, CHARLAINE  C/P:2,P
LIVING END, THE  ELKIN, STANLEY  P
LIVING FIRE MENACE, THE [DOC SAVAGE #61]  ROBESON, KENNETH  C/P:2,@P
LIVING GOD, THE [HANDFUL OF MEN 4]  DUNCAN, DAVE  C/P,P,H
LIVING IN ETHER  GEARY, PATRICIA  P
LIVING LEGEND, THE [@BATTLESTAR GALACTICA 1978 #6]  LARSON, GLEN A./YERMAKOV, NICHOLAS  P
LIVING NEXT DOOR TO THE GOD OF LOVE  ROBSON, JUSTINA  LP
LIVING ONE, THE  GANNETT, LEWIS  SFWA
LIVING ONE, THE  HAWKINS, JIM  C/P,P
LIVING REAL  BASSETT, JAMES C.  C/P,P
LIVING SHADOW, THE  GRANT, MAXWELL  C/P,P
LIVING STONE, THE  ROLAND, HOWELL  P
LIVING THINGS  RUSSO, JOHN  P
LIVING WAY OUT  GUIN, WYMAN  C/P:2,P
LIVING WITH ALIENS  DE CHANCIE, JOHN  C/P,P
LIVING WITH THE REPTILES  DISILVESTRO, ROGER L.  H
LIZARD MUSIC  PINKWATER, DANIEL MANUS  P
LIZARD WAR, THE [LIZARD WAR 1]  DALMAS, JOHN  C/P:2,P
LLANA OF GATHOL [BARSOOM #10]  BURROUGHS, EDGAR RICE  C/P:2,@P
LLANA OF GATHOL [BARSOOM 10]/JOHN CARTER OF MARS [BARSOOM 11] BURROUGHS, EDGAR RICE  H:2
LLEGAN LOS DRAGONES  SOUTO, MARCIAL  FP
LO!  FORT, CHARLES  [C/RAN-H,C/RAN-P:2]
LOADING LAS VEGAS  POTTS, CHARLES  LP
LOAFERS OF REFUGE, THE  GREEN, JOSEPH  C/P:2,P
LOCAL CUSTOM [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  P
LOCH, THE  CAIRD, JANET  P
LOCK  LOCK, MASTER  WRD
LOCKDOWN [GARRISON 1]  SCHARTZ, VIJAYA  LP
LOCKED IN TIME  DUNCAN, LOIS  P
LOCUS AWARDS, THE  BROWN, CHARLES N./STRAHAN, JONATHAN LPA
LODESTAR [STARS 3] FLYNN, MICHAEL H,P

LODGE OF THE LYNX, THE [@ADEPT #2] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P,P

LOG OF A MOON EXPEDITION  PESEK, LUDEK  H

LOGAN [LOGAN 1,2,3] NOLAN, WILLIAM F./

JOHNSON, GEORGE CLAYTON

H,SFWA,LP

LOGAN’S RUN [LOGAN 1] NOLAN, WILLIAM F./

JOHNSON, GEORGE CLAYTON

C/P:2,P

LOGAN’S SEARCH [LOGAN 3]  NOLAN, WILLIAM F.  C/P,P
LOGAN’S WORLD [LOGAN 2]  NOLAN, WILLIAM F.  C/P,P
LOGICAL MAGICIAN, A [LOGICAL MAGICIAN 1]  WEINBERG, ROBERT  C/P:2,P
LOGOPOLIS [DOCTOR WHO]  BIDMEAD, CHRISTOPHER H.  SR-P,C/P:2
LOGOS RUN [JAK REBO 2]  DIETZ, WILLIAM C.  H
LOIS & CLARK [LOIS & CLARK]  CHERRYH, C. J.  H
"LOMOKOME" PAPERS, THE  WOUK, HERMAN  C/P,P
LONE DROW, THE [@FORGOTTEN REALMS HUNTER’S BLADES #2]  SALVATORE, R. A.  P
LONE SENTINEL, THE  DERESKE, JO  SFWA
LONE STAR [STELLAR RANGER 2]  PERRY, STEVE  P
LONE STAR UNIVERSE  PROCTOR, GEO. W./UTLEY, STEVEN  SR-HA,HA
LONE WOLF [@SHADOWRUN 12]  FINDLEY, NIGEL D.  P
LONELIEST MAGICIAN, THE [DRAGON NIMBUS 3]  RADFORD, IRENE  P
LONELY ASTRONOMER, THE  GRIDBAN, VOLSTED  SR-P
LONELY PLANET, THE  LEINSTER, MURRAY  B
LONELY VIGILS  WELLMAN, MANLY WADE  SR-H
LONG AFTER MIDNIGHT  BRADBURY, RAY  C/P,P,H
LONG AFTERNOON OF EARTH, THE  ALDISS, BRIAN W.  C/P:2,P
LONG ARM OF GIL HAMILTON, THE [KNOWN SPACE]  NIVEN, LARRY  SR-P,C/P:2
LONG DARK TEA-TIME OF THE SOUL, THE [DIRK GENTLY 2]  ADAMS, DOUGLAS  C/P,P,H
LONG FORGETTING, THE  BYERS, EDWARD A.  C/P:2,P
LONG HOT SUMMONING [KEEPER’S CHRONICLES 3]  HUFF, TANYA  P
LONG HUNT, THE [MAGEWORLDS 5]  DOYLE, DEBRA/MACDONALD, JAMES D.  P
LONG LOST, THE  CAMPBELL, RAMSEY  P
LONG LOUD SILENCE, THE  TUCKER, WILSON  H:2,P
LONG MYND, THE  HUGHES, EDWARD P.  C/P,P
LONG NIGHT, THE  ANDERSON, POUL  C/P,P

The MITSFS Titledex  September 24, 2007  Page 209
TITLE  AUTHOR  CODE

LONG NIGHT, THE [STAR TREK DS9] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P:2,P

LONG NIGHT OF CENTAURI PRIME, THE [@BABYLON 5 LEGIONS OF FIRE #1] DAVID, PETER

P

LONG NIGHT OF THE GRAVE, THE  GRANT, CHARLES L.  H
LONG NIGHT OF WAITING  ELWOOD, ROGER  SR-HA
LONG ORBIT, THE  FARREN, MICK  P
LONG PATROL, THE [REDWALL]  JACQUES, BRIAN  P
LONG PATROL, THE [@BATTLESTAR GALACTICA 1978 #10]  LARSON, GLEN A./GOULART, RON  P
LONG RESULT, THE  BRUNNER, JOHN  C/P:2,P
LONG RUN, THE [CONTINUING TIME 2]  MORAN, DANIEL KEYS  C/P,P
LONG SHOT FOR ROSINANTE [ROSINANTE 2]  GILLILAND, ALEXIS A.  C/P:2,P
LONG SLEEP, THE  HILL, JOHN  P
LONG TOMORROW, THE  BRACKETT, LEIGH  H:2,P
LONG TWILIGHT, THE  LAUMER, KEITH  C/P,P,H
LONG VACATION, A  VERNE, JULES  H
LONG VIEW, THE  BUSBY, F. M.  C/P,P,H
LONG WAY HOME, THE  ANDERSON, POUL  C/P:2,P
LONG WAY HOME [@BUFFY ANGEL UNSEEN 3]  HOLDER, NANCY/MARIOTTE, JEFF  P
LONG WINTER, THE  CHRISTOPHER, JOHN  C/P:2,P
LONGEST DAY [DOCTOR WHO EIGHTH DOCTOR]  COLLIER, MICHAEL  C/P,P
LONGEST NIGHT 1, THE [@ANGEL]  TWENTIETH CENTURY FOX  PA
LONGEST VOYAGE, THE  ANDERSON, POUL  SFWA-TD30,TD30
LONGEST WAY HOME, THE  SILVERBERG, ROBERT  H
LONGING RING, THE [@EARTHDAWN]  KUBASIK, CHRISTOPHER  P
LONGTUSK [MAMMOTHS 2]  BAXTER, STEPHEN  C/P,P
LOOK BACK TO EARTH  PFEIL, DONALD J.  P
LOOK BEHIND YOU!  BURKS, ARTHUR J.  SR-H
LOOK INTO THE SUN [MESSENGER 2]  KELLY, JAMES PATRICK  C/P,SFWA-P,P,H
LOOK TO WINDWARD [CULTURE]  BANKS, IAIN M.  H
LOOKING BACKWARD 2000–1887  BELLAMY, EDWARD  H,SR-H,SR-L,P
LOOKING BACKWARD FROM THE YEAR 2000  REYNOLDS, MACK  C/P:2,P
LOOKING FOR BLUCHER  WODHAMS, JACK  SR-P
LOOKING FOR JAKE  MIEVILLE, CHINA  LP
LOOKING FOR THE GENERAL  MILLER, WARREN  P
LOOKING FOR THE MAHDI  WOOD, N. LEE  C/P,LP
LOOKING FORWARD  LESSER, MILTON  HA
LOOKING GLASS WARS, THE  BEDDOR, FRANK  H
LOOSE CANNON [LIADEN UNIVERSE 0/LIADEN ADVENTURES 7]  LEE, SHARON/MILLER, STEVE  SR-LP
LOOTERS OF THARN [@BLADE #19]  LORD, JEFFREY  P
L’ORBE ET LA ROUE  JEURY, MICHEL  FP
LORD BROCKTREE [REDWALL]  JACQUES, BRIAN  H
LORD CONRAD’S LADY [CONRAD STARGARD 5]  FRANKOWSKI, LEO  C/P:2,P
LORD DARCY [LORD DARCY 1,2,3]  GARRETT, RANDALL  H
LORD DARCY (BAEN) [LORD DARCY 1,2,3,4]  GARRETT, RANDALL  LP
LORD DARCY INVESTIGATES [LORD DARCY 3]  GARRETT, RANDALL  C/P:2,P
LORD DEMON  ZELAZNY, ROGER/LINDSKOLD, JANE  H,P
LORD FOUL’S BANE [THOMAS COVENANT 1]  DONALDSON, STEPHEN R.  C/P,H,LP
LORD HALIFAX’S GHOST BOOK  HALIFAX, LORD  [C/RAN-H,C/RAN-P]
LORD KALVAN OF OTHERWHEN [PARATIME POLICE]  PIPER, H. BEAM  C/P,P,H
LORD KELVIN’S MACHINE  BLAYLOCK, JAMES P.  P,SFWA-P:2
LORD OF BLOOD  VAN ARNAM, DAVE  C/P,P
LORD OF CASTLE BLACK, THE [DRAGAERA ADRILANKHA 2]  BRUST, STEVEN  H
LORD OF CHAOS, THE [WIZARD’S WAR 4]  BOYER, ELIZABETH H.  C/P:2,P
LORD OF CHAOS [WHEEL OF TIME 6]  JORDAN, ROBERT  H:2,P
LORD OF CRAGSCLAW [GUARDIANS OF THE THREE 1]  FAWCETT, BILL/RANDALL, NEIL  C/P:2,P
LORD OF DARKNESS  SILVERBERG, ROBERT  H
LORD OF DEATH, THE/QUEEN OF LIFE, THE  FLINT, HOMER EON  C/P,P
LORD OF EMPERORS [SARANTINE MOSAIC 2]  KAY, GUY GAVRIEL  C/P,P,H
LORD OF HORSES, THE [WODAN’S CHILDREN 3]  PAXSON, DIANA L.  P
LORD OF LANCES [JANISSARIES]  ACRES, MARK  P
LORD OF LIGHT  ZELAZNY, ROGER  C/P,P,H
LORD OF NIGHTMARES [@DREAM LORDS #2]  COLE, ADRIAN  P
LORD OF SHADOW KEEP, THE  JOHNSON, OLIVER  P
LORD OF SNOW AND SHADOWS [TEARS OF ARTAMON 1]  ASH, SARAH  C/P,P
LORD OF SNOW AND SHADOWS (PROOF) [TEARS OF ARTAMON 1]  ASH, SARAH  SR-LP
LORD OF STORMWEATHER [@FORGOTTEN REALMS SEMBIA]  GROSS, DAVE  P
LORD OF THE BLACK RIVER [CONAN]  CARPENTER, LEONARD  P
LORD OF THE CLANS [@WARCRAFT #2]  GOLDEN, CHRISTIE  P
LORD OF THE CROOKED PATHS [TITANS 1]  ADKINS, PATRICK H.  P
LORD OF THE DEAD  HOLLAND, TOM  H,LP
LORD OF THE FANTASTIC  GREENBERG, MARTIN H.  HA,LPA
LORD OF THE FIRE LANDS [KING’S BLADES]  DUNCAN, DAVE  P
LORD OF THE FLIES  GOLDING, WILLIAM  C/P:2,P

TITLE  AUTHOR  CODE
LORD OF THE GREEN PLANET [GREEN PLANET 1]  PETAJA, EMIL  DH022
LORD OF THE ISLES [LORD OF THE ISLES 1]  DRAKE, DAVID  C/P:2,P,H
LORD OF THE LIBRARIES [ROVER 3]  ODOM, MEL  P
LORD OF THE NECROPOLIS [RAVENLOFT]  DEWEESE, GENE  P
LORD OF THE NIGHT [@WARHAMMER 40K]  SPURRIER, SIMON  P
LORD OF THE RED SUN  SILENT, WILLIAM T.  H:2
LORD OF THE RINGS, THE (12 CASSETTES, BOXED) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
LORD OF THE RINGS, THE (13 CASSETTES, BOXED) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
LORD OF THE RINGS, THE (COLLECTORS EDITION) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-VLH
LORD OF THE RINGS, THE (MOVIE PROMOTION)  FANTASY FILMS  A
LORD OF THE ROSE [@DRAGONLANCE RISE OF SOLAMNIA #1]  NILES, DOUGLAS  P
LORD OF THE SEA  SHIEL, M. P.  H,LP
LORD OF THE SHADOWS [SECOND SONS 3]  FALLON, JENNIFER  P

LORD OF THE SILENT KINGDOM [@INSTRUMENTALITIES OF THE NIGHT #2] COOK, GLEN

H

LORD OF THE SILVER BOW [@TROY 1]  GEMMELL, DAVID  LP
LORD OF THE SPIDERS, THE [MICHAEL KANE 2]  MOORCOCK, MICHAEL  C/P,P
LORD OF THE STARS  SUTTON, JEAN/SUTTON, JEFF  H
LORD OF THE TREES  FARMER, PHILIP JOSE  D51375,P
LORD OF THE TREES/MAD GOBLIN, THE  FARMER, PHILIP JOSE  C/P,P
LORD OF THE TROLL-BATS  GILLILAND, ALEXIS A.  P
LORD OF THE TWO LANDS  TARR, JUDITH  H
LORD OF THUNDER [HOSTEEM STORM 2]  NORTON, ANDRE  C/P:2,P
LORD OF TRANERICA  COBLENTZ, STANTON A.  LP
LORD PRESTIMION [MAJIPOOR PRESTIMION 2]  SILVERBERG, ROBERT  H:2,P
LORD SOTH [@DRAGONLANCE WARRIORS #6]  VAN BELKOM, EDO  P
LORD TEDRIC [LORD TEDRIC #1]  EKLUND, GORDON  H,@P,@C/P:2
LORD TOEDE [DRAGONLANCE VILLAINS 5]  GRUBB, JEFF  P
LORD TYGER  FARMER, PHILIP JOSE  C/P:2,P
LORD VALENTINE’S CASTLE [MAJIPOOR 1]  SILVERBERG, ROBERT  H:2,SFWA,P
LORD VALENTINE’S CASTLE (PROOF) [MAJIPOOR 1]  SILVERBERG, ROBERT  SR-LP
LORDS AND LADIES [DISCWORLD 14/DISCWORLD WITCHES 4]  PRATCHETT, TERRY  C/P:2,LP,P
LORDS AND LADIES (PROOF) [DISCWORLD 14/DISCWORLD WITCHES 4]  PRATCHETT, TERRY  SR-LP
LORDS OF ATLANTIS  WEST, WALLACE  C/P:2,P
LORDS OF CREATION  BINDER, EANDO  C/P:2,P
LORDS OF CREATION  SULLIVAN, TIM  C/P,P
LORDS OF DARKNESS, THE [@FLAT EARTH 3,4,5]  LEE, TANITH  H:2
LORDS OF DESTRUCTION [DEATH DEALER 2]  SILKE, JAMES  P,SFWA-P
LORDS OF DUS, THE [LORDS OF DUS 1,2,3,4]  WATT-EVANS, LAWRENCE  H
LORDS OF SWORDS  BLACKSTON, DANIEL E.  LPA
LORDS OF THE CRIMSON RIVER, THE [@BLADE #35]  LORD, JEFFREY  P
LORDS OF THE MIDDLE DARK [@RINGS OF THE MASTER #1]  CHALKER, JACK L.  C/P:2,P
LORDS OF THE PSYCHON  GALOUYE, DANIEL F.  C/P:2,P
LORDS OF THE SEVENTH SWARM [GOLDEN QUEEN 3]  WOLVERTON, DAVE  H,P
LORDS OF THE SHADOWS [@RAVEN #4]  KIRK, RICHARD  SR-P
LORDS OF THE SKY [LORDS OF THE SKY 1]  WELLS, ANGUS  P
LORDS OF THE STARSHIP  GESTON, MARK S.  C/P:2,P
LORDS OF THE STORM [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
LORDS OF THE SWORD [@WIZARD WAR]  COOK, HUGH  P,SFWA-P:2
LORDS OF THE TRIPLE MOONS  MAYHAR, ARDATH  C/P:2,SFWA,P,H
LORDS OF THLASSA MEY [THLASSA MEY 3]  MCCARTY, DENNIS  C/P,P
LORDS OF VALOUR [WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
LORD’S PINK OCEAN, THE  WALKER, DAVID  C/P,P,H
LORDS TEMPORAL  DELANEY, JOSEPH H.  C/P,P
LORE OF THE WITCH WORLD [WITCH WORLD]  NORTON, ANDRE  C/P:2,P
LOREMASTERS, THE  GADALLAH, LESLIE  C/P:2,P
LORI  BLOCH, ROBERT  SFWA-P
LOS ANGELES: A.D. 2017  WYLIE, PHILIP  C/P:2,P
LOSERS, THE  EDDINGS, DAVID  P
LOSERS, THE & HIGH HUNT  EDDINGS, DAVID  H
LOST, THE  AYCLIFFE, JONATHAN  H,P
LOST, THE [@OUTER LIMITS]  PEEL, JOHN  LP
LOST: A MOON  CAPON, PAUL  H
LOST AEONS  CAMERON, BERL  SR-P
LOST AND FOUND [@GAMMA QUEST #1]  COX, GREG  P
LOST AND FOUND [TAKEN 1]  FOSTER, ALAN DEAN  H
LOST AND THE LURKING, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  H:2,SFWA,P
LOST BOYS  CARD, ORSON SCOTT  H:2,P
LOST BOYS, THE  GARDNER, CRAIG SHAW  P
LOST BURGUNDY [@ASH #4]  GENTLE, MARY  C/P,P
LOST CITIES AND VANISHED CIVILIZATIONS  SILVERBERG, ROBERT  [C/RAN-P]
LOST CITY, THE [@RICK BRANT #2]  BLAINE, JOHN  H:2
LOST CITY OF THE DAMNED  RIVERE, ALEC  SR-L
LOST CITY OF ZORK, THE [ZORK]  BAILEY, ROBIN WAYNE  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 211
TITLE  AUTHOR  CODE
LOST COLONY, THE [JEFFERSON’S WAR 3]  RANDLE, KEVIN  C/P,P
LOST CONTINENT, THE  BURROUGHS, EDGAR RICE  C/P:2,P
LOST CONTINENT, THE  HYNE, C. J. CUTCLIFFE  C/P:2,P
LOST CONTINENT OF MU, THE  CHURCHWARD, JAMES  [C/RAN-P]
LOST CONTINENTS  DE CAMP, L. SPRAGUE  [C/RAN-P]

LOST CONTINENTS & THE HOLLOW EARTH CHILDRESS, DAVID HATCHER/

SHAVER, RICHARD

[C/RAN-H]

LOST DAYS [@LOST MILLENNIUM 3]  MOSCOE, MIKE  C/P,P
LOST DESTINY [@BATTLETECH #22/BLOOD OF KERENSKY 3]  STACKPOLE, MICHAEL A.  P
LOST DORSAI [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,LP,H
LOST DRAGONS OF BARAKHAI, THE [@BARAKHAI #2]  REICHERT, MICKEY ZUCKER  H,P
LOST FACE, THE  NESVADBA, JOSEF  H
LOST FANTASIES #4  WEINBERG, ROBERT  SR-LPA
LOST FANTASIES #5  WEINBERG, ROBERT  SR-LPA
LOST FANTASIES #6  WEINBERG, ROBERT  SR-LPA
LOST FANTASIES #8: THE LAKE OF LIFE  WEINBERG, ROBERT  SR-LPA
LOST: FIFTY SUNS  VAN VOGT, A. E.  P
LOST GUARDIAN, THE [@ETERNAL GUARDIANS #2]  CROSS, RONALD ANTHONY  H,P
LOST HORIZON  HILTON, JAMES  C/P,SR-P,H
LOST IN A GOOD BOOK [THURSDAY NEXT 2]  FFORDE, JASPER  LP
LOST IN SPACE  SMITH, GEORGE O.  DD431
LOST IN SPACE [LOST IN SPACE]  VAN ARNAM, DAVE/ARCHER, RON  P
LOST IN SPACE [LOST IN SPACE]  VINGE, JOAN D.  P
LOST IN SPACE SCRAPBOOK VOL. TWO, THE [LOST IN SPACE]  ANCHORS, WILLIAM E. JR./MITCHELL, FLINT VLP

LOST IN THE RENTHARPIAN HILLS:

SPANNING THE DECADES WITH CARL JACOBI SMITH, R. DIXON

R/REF-P

LOST IN TIME AND SPACE WITH LEFTY FEEP  BLOCH, ROBERT  H:2,LP
LOST IN TRANSLATION  BALL, MARGARET  C/P:2,P
LOST IN TRANSMISSION [QUEENDOM OF SOL 3]  MCCARTHY, WIL  C/P,P
LOST KAER [@EARTHDAWN]  FINDLEY, NIGEL D.  P
LOST KING, THE [@STAR OF THE GUARDIANS #1]  WEIS, MARGARET  C/P:2,P
LOST LAND, THE  MARSHALL, EDISON  P
LOST LANDS OF THE WITCH WORLD [WITCH WORLD ESTCARP 1,2,3]  NORTON, ANDRE  H

LOST LIBRARY OF CORMANTHYR, THE [@FORGOTTEN REALMS LOST EMPIRES] ODOM, MEL

P

LOST LOVE  CARTLAND, BARBARA  P
LOST MILLENNIUM, THE  RICHMOND, WALT/RICHMOND, LEIGH  DH029
LOST MYTHOLOGIES  ZACKEY, CHRISTOPHER A.  SR-LP
LOST OASIS, THE [DOC SAVAGE #6]  ROBESON, KENNETH  C/P:2,@P
LOST ON THE MOON  ROCKWOOD, ROY  H
LOST ON VENUS [VENUS #2]  BURROUGHS, EDGAR RICE  C/P:2,@P
LOST ONES, THE [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P:2,P
LOST PAGES  DI FILIPPO, PAUL  LP
LOST PRINCE, THE [WOLFKINGS]  WOOD, BRIDGET  P
LOST PRINCE, THE [@DARK BORDER #1]  ZIMMER, PAUL EDWIN  C/P:2,P
LOST PRINCESS, THE  MACDONALD, GEORGE  P
LOST PRINCESS OF OZ, THE [OZ #11]  BAUM, L. FRANK  H,@P
LOST RACE OF MARS  SILVERBERG, ROBERT  C/P:2,P
LOST ROAD AND OTHER WRITINGS, THE [HISTORY OF MIDDLE-EARTH 5]  TOLKIEN, J. R. R.  H:2,P
LOST RUNES, THE [RUNESPELL 2]  WELCH, JANE  P
LOST SLAYER, THE [@BUFFY LOST SLAYER 1,2,3,4]  GOLDEN, CHRISTOPHER  P
LOST SOULS  BRITE, POPPY Z.  C/P,P
LOST STAR, THE  HOOVER, H. M.  C/P,P
LOST TALE, A  ESTEY, DALE  P
LOST TRAVELLER, THE  WILSON, STEVE  C/P,P,H
LOST TRUTH [FIRST TRUTH 4]  COOK, DAWN  P
LOST VALLEY OF ISKANDER, THE  HOWARD, ROBERT E.  SR-H,P
LOST VILLAGE, THE  HALL, MARK EDWARD  H
LOST VIOL, THE  SHIEL, M. P.  H
LOST WORLD, THE [JURASSIC PARK 2]  CRICHTON, MICHAEL  H:2,P
LOST WORLD, THE  DOYLE, ARTHUR CONAN  C/P:2,P

LOST WORLD, THE [PROF. CHALLENGER 1]/ POISON BELT, THE [PROF. CHALLENGER 1] DOYLE, ARTHUR CONAN

SR-H

LOST WORLD OF TIME  CARTER, LIN  P
LOST WORLDS  CARTER, LIN  P
LOST WORLDS  SMITH, CLARK ASHTON  H
LOST WORLDS I  SMITH, CLARK ASHTON  SR-P
LOST WORLDS II  SMITH, CLARK ASHTON  SR-P
LOST WORLDS OF 2001, THE  CLARKE, ARTHUR C.  C/P:2,P
LOST WORLDS OF CRONUS, THE [CAGEWORLD 2]  KAPP, COLIN  C/P,P
LOST WORLDS, UNKNOWN HORIZONS  SILVERBERG, ROBERT  HA
LOST YEARS, THE [STAR TREK]  DILLARD, J. M.  C/P,P,H:2
LOST YEARS OF MERLIN, THE [LOST YEARS OF MERLIN 1]  BARRON, T. A.  P
LOT & LOT’S DAUGHTER  MOORE, WARD  LP

The MITSFS Titledex  September 24, 2007  Page 212
TITLE  AUTHOR  CODE
LOTSA LUCK [@SABRINA THE TEENAGE WITCH #10]  GALLAGHER, DIANA G.  P
LOTUS AND THE ROSE, THE  CIENCIN, SCOTT  P
LOTUS CAVES, THE  CHRISTOPHER, JOHN  P,SFWA-P
LOTUS VELLUM, THE  HUNT, CHARLOTTE  P

LOVE 3000 WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA

LOVE AIN’T NOTHING BUT SEX MISSPELLED  ELLISON, HARLAN  C/P:2,P
LOVE AMERICAN STYLE  FAIRMAN, PAUL W.  P
LOVE AMERICAN STYLE #2  FAIRMAN, PAUL W.  P
LOVE AND DEATH [@ANGEL]  MARIOTTE, JEFF  P
LOVE AND NAPALM: EXPORT U.S.A.  BALLARD, J. G.  H
LOVE AND SLEEP [AEGYPT 2]  CROWLEY, JOHN  LP
LOVE AND WAR [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
LOVE AND WAR [@DRAGONLANCE TALES #3]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA
LOVE BITE  GOTTLIEB, SHERRY  P
LOVE BOMB, THE  SACKERMAN, HENRY  P
LOVE CHILD  NEIDERMAN, ANDREW  P,SFWA-P
LOVE CONQUERS ALL  SABERHAGEN, FRED  C/P:2,P
LOVE IN VEIN  BRITE, POPPY Z.  C/PA,LPA
LOVE IN VEIN II  BRITE, POPPY Z.  LPA
LOVE IS FOREVER — WE ARE FOR TONIGHT  WILLIAMS, ROBERT MOORE  R/REF-P
LOVE MACHINE, THE [@THESE LAWLESS WORLDS #1]  COMSTOCK, JARROD  P
LOVE MACHINERY, THE  FLINDERS, KARL  SR-L
LOVE NOT HUMAN  DICKSON, GORDON R.  C/P:2,SFWA-P,P
LOVE ONCE AGAIN  SIMON, JO ANN  P
LOVE ONCE IN PASSING  SIMON, JO ANN  P
LOVE OUT OF TIME  ASHLEY, STEVEN  P
LOVECRAFT: A BIOGRAPHY  DE CAMP, L. SPRAGUE  R/REF-H
LOVECRAFT: A LOOK BEHIND THE CTHULHU MYTHOS  CARTER, LIN  R/REF-P,C/REF-P
LOVECRAFT AT LAST  LOVECRAFT, H. P./CONOVER, WILLIS  R/REF-H
LOVEDEATH  SIMMONS, DAN  C/P,P
LOVEDEATH (PROOF)  SIMMONS, DAN  SR-LP
LOVELOCK [MAYFLOWER 1]  CARD, ORSON SCOTT/KIDD, KATHRYN H.  H,P
LOVERS, THE  FARMER, PHILIP JOSE  C/P,P,H
LOVERS AND OTHER MONSTERS  KAYE, MARVIN  HA
LOVE’S BODY, DANCING IN TIME  DUCHAMP, L. TIMMEL  LP
LOVING THE ALIEN [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE/PERRY, ROBERT  P
LOW-FLYING AIRCRAFT  BALLARD, J. G.  SR-LP
LOW PORT  LEE, SHARON/MILLER, STEVE  HA
LOW RED MOON  KIERNAN, CAITLIN R.  LP
LOWLAND RIDER  WILLIAMSON, CHET  SFWA-P
LOXFINGER  WEINSTEIN, SOL  LP
LOYALTIES [@STAR TREK TNG STARFLEET ACADEMY #10]  BARNES-SVARNEY, PATRICIA  LP
LUANA  FOSTER, ALAN DEAN  P
LUCHSHII IZ MIROV  MITROKHINA, S.  FH
LUCIFER COMET, THE  WALLACE, IAN  C/P,P
LUCIFER CRUSADE, THE [WINGMAN 3]  MALONEY, MACK  P
LUCIFER DECK, THE [SHADOWRUN 23]  SMEDMAN, LISA  P
LUCIFER JONES [LUCIFER JONES 2]  RESNICK, MIKE  C/P,P
LUCIFER RISING [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY/MORTIMORE, JIM  C/P:2,P
LUCIFER SOCIETY, THE  HAINING, PETER  PA
LUCIFER’S DRAGON  GRIMWOOD, JON COURTENAY  LP
LUCIFER’S HAMMER  NIVEN, LARRY/POURNELLE, JERRY  H:2,P
LUCK BE A LADY [@CHARMED]  CIENCIN, SCOTT  P
LUCK IN THE SHADOWS [NIGHTRUNNER 1]  FLEWELLING, LYNN  C/P,P
LUCK MACHINE, THE  TUBB, E. C.  LP
LUCK OF BRIN’S FIVE, THE [TORIN 1]  WILDER, CHERRY  H,P
LUCK OF RELIAN KRU, THE  VOLSKY, PAULA  C/P,SFWA-P,P
LUCK OF THE WHEELS [KI AND VANDIEN 4]  LINDHOLM, MEGAN  C/P,P
LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 4]  ASIMOV, ISAAC  C/P:2,P
LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5]  ASIMOV, ISAAC  C/P:2,P

LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5,6]/ LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 5,6] ASIMOV, ISAAC

P

LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3]  ASIMOV, ISAAC  C/P,P,H

LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3,4]/ LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 3,4] ASIMOV, ISAAC

P

LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  ASIMOV, ISAAC  C/P:2,P
LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 6]  ASIMOV, ISAAC  C/P:2,P
LUCY’S BLADE  LAMBSHEAD, JOHN  H
LUD-IN-THE-MIST  MIRRLEES, HOPE  H,P
LUDLOW’S MILL  WALTERS, R. R.  P
LUKAN WAR  COLLINS, MICHAEL  C/P,P
LUKE SUTTON: OUTRIDER  KELLEY, LEO P.  SFWA
LUMINOUS  EGAN, GREG  H,P

TITLE  AUTHOR  CODE
LUNA DE MIEL EL INFIERNO (SPANISH)  BROWN, FREDRIC  FH
LUNA MARINE [@HERITAGE #2]  DOUGLAS, IAN  C/P,P
LUNAR ACTIVITY  MOON, ELIZABETH  C/P:2,SFWA-P,P
LUNAR ATTACK [SPACE: 1999 5]  RANKINE, JOHN  C/P,P
LUNAR DESCENT  STEELE, ALLEN  C/P:2,SFWA-P:2,P
LUNAR EYE, THE  WILLIAMS, ROBERT MOORE  DF261
LUNAR FLIGHT  FEAR, W. H.  SR-P
LUNAR JUSTICE  HARNESS, CHARLES L.  P
LUNAR RESOURCES AND THEIR UTILIZATION  PHINNEY, WILLIAM C., ET AL.  [C/RAN-H:2]
LUNATIC CAFE, THE [ANITA BLAKE 4]  HAMILTON, LAURELL K.  C/P:2,P
LUNATICS  DENTON, BRADLEY  LP
LUNGBARROW [DOCTOR WHO NEW ADVENTURES]  PLATT, MARC  P
L’UOMO ETERNO (ITALIAN)  DICKSON, GORDON R.  FP
LUPE  THOMPSON, GENE  P
LURE OF THE BASILISK, THE [LORDS OF DUS 1]  WATT-EVANS, LAWRENCE  C/P:2,P
LURID DREAMS  HARNESS, CHARLES L.  C/P,P
LURKER  PETRUCHA, STEFAN/PENDLETON, THOMAS  LP
LURKER AT THE THRESHOLD, THE  LOVECRAFT, H. P./DERLETH, AUGUST  SR-H,C/P:2,P
LURKING FEAR AND OTHER STORIES, THE (AVON)  LOVECRAFT, H. P.  SR-P,C/P
LURKING FEAR AND OTHER STORIES, THE (BALLANTINE)  LOVECRAFT, H. P.  C/P:2,P
LUST FOR A VAMPIRE  HUGHES, WILLIAM  P
LUST LIZARD OF MELANCHOLY COVE, THE  MOORE, CHRISTOPHER  H,LP
LUTA INTERGALACTICA  HAMILTON, EDMOND  FP
LUTE AND THE GLOVE, THE  EYRE, KATHERINE WIGMORE  P
LUTHIEN’S GAMBLE [CRIMSON SHADOW 2]  SALVATORE, R. A.  P
LYCANTHIA, OR, THE CHILDREN OF WOLVES  LEE, TANITH  C/P,P
LYON’S PRIDE  MCCAFFREY, ANNE  C/P,P,H
LYREC  FROST, GREGORY  C/P:2,P
LYTHANDE [THIEVES’ WORLD]  BRADLEY, MARION ZIMMER  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 215
TITLE  AUTHOR  CODE
M IS FOR MAGIC  GAIMAN, NEIL  H
M IS FOR MAGIC (PROOF)  GAIMAN, NEIL  SR-LP
M33 IN ANDROMEDA  VAN VOGT, A. E.  C/P:2,P
MA QUI AND OTHER PHANTOMS  BRENNERT, ALAN  SR-LP
MABINOGI, THE  FORD, PATRICK K.  LP
MABINOGION, THE  ANONYMOUS  P
MACABRE ONES, THE  FANE, BRON  SR-P
MACABRE READER, THE  WOLLHEIM, DONALD A.  PA,C/PA
MACABRE TALES  O’BRIEN, FITZ-JAMES  SFWA
MACAH [ANITA BLAKE 13]  HAMILTON, LAURELL K.  P
MACCHINA DEL TEMPO, LA  DE CAMP, L. SPRAGUE  FP
MACE OF SOULS, THE [SIX KINGDOMS]  FERGUSSON, BRUCE  C/P,P,H
MACH 1  ADLER, ALLEN  H:2
MACHIAVELLI INTERFACE, THE [MATADOR 3]  PERRY, STEVE  C/P:2,P
MACHINATIONS [@ULTIMA TECHNOCRAT WAR #1]  ANDREWS, AUSTEN  P
MACHINATIONS  ANDREWS, CHRIS  LPA
MACHINE-GOD LAUGHS, THE  GRIFFIN PUBLISHING  HA
MACHINE IN SHAFT TEN, THE  HARRISON, M. JOHN  SR-P
MACHINE IN WARD ELEVEN, THE  WILLEFORD, CHARLES  C/P:2,P
MACHINE SEX  DORSEY, CANDAS JANE  LP
MACHINE THAT THOUGHT, THE  CALLAHAN, WILLIAM  B
MACHINE TO KILL, THE  LEROUX, GASTON  H
MACHINERIES OF JOY, THE  BRADBURY, RAY  C/P:3,P
MACHINES AND MEN  ROBERTS, KEITH  SR-P
MACHINE’S CHILD, THE [COMPANY 7]  BAKER, KAGE  H

MACHINES THAT KILL SABERHAGEN, FRED/

GREENBERG, MARTIN H.

C/PA,PA

MACRA TERROR, THE [DOCTOR WHO]  BLACK, IAN STUART  SR-P
MACRO  DAVIDSON, FRANK P./COX, JOHN STUART  [C/RAN-H]
MACROCOSMIC CONFLICT, THE [STARHOUNDS 3]  BISCHOFF, DAVID F.  C/P,P
MACROLIFE  ZEBROWSKI, GEORGE  C/P,SFWA,P
MACROSCOPE  ANTHONY, PIERS  C/P:2,P
MAD AMOS  FOSTER, ALAN DEAN  C/P,P
MAD DOGS AND ENGLISHMEN [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL  P
MAD EMPRESS OF CALLISTO [JANDAR 4]  CARTER, LIN  C/P,P
MAD EYES [DOC SAVAGE #34]  ROBESON, KENNETH  C/P,@P
MAD GOBLIN, THE  FARMER, PHILIP JOSE  C/P,D51375
MAD GOD’S AMULET, THE [RUNESTAFF 2]  MOORCOCK, MICHAEL  C/P:2,P
MAD KING, THE  BURROUGHS, EDGAR RICE  C/P,P
MAD MAUDLIN [ERIC BANYON 5]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H:2
MAD MAX BEYOND THUNDERDOME  VINGE, JOAN D.  P
MAD MERLIN  KING, J. ROBERT  P
MAD MESA [DOC SAVAGE #66]  ROBESON, KENNETH  C/P:2,@P
MAD METROPOLIS, THE  HIGH, PHILIP E.  P,DM135
MAD MOON OF DREAMS [DREAMLAND 3]  LUMLEY, BRIAN  P
MAD PROFESSOR  RUCKER, RUDY  LP
MAD ROY’S LIGHT  KING, PAULA  P,SFWA-P
MAD SCIENTIST AFFAIR, THE [@MAN FROM UNCLE #5]  PHILLIFENT, JOHN T.  C/P:2,P
MAD SCIENTISTS  SCHIFF, STUART DAVID  HA
MAD SCIENTISTS’ CLUB, THE  BRINLEY, BERTRAND R.  C/P:2,P
MAD SHIP [LIVESHIP TRADERS 2]  HOBB, ROBIN  C/P,P,H
MAD THRONE, THE  MUNSON, BRAD  P
MADAGASCAR MANIFESTO, THE [MADAGASCAR MANIFESTO 1,2,3]  BERLINER, JANET/GUTHRIDGE, GEORGE  LP
MADBOND [SEA KING 1]  SPRINGER, NANCY  P,SFWA-P
MADE IN GOATSWOOD  ANIOLOWSKI, SCOTT DAVID  LP
MADLANDS  JETER, K. W.  H
MADMAN’S ARMY [HORSECLANS 17]  ADAMS, ROBERT  P

MADNESS OF PRIESTS, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #2] BOULLE, PHILIPPE

P

MADNESS SEASON, THE  FRIEDMAN, C. S.  C/P:2,P
MADOUC [LYONESSE 3]  VANCE, JACK  C/P,SFWA-P:2,LP,H
MADOUC (VIE) [@VANCE INTEGRAL EDITION #38/LYONESSE 3]  VANCE, JACK  SR-H
MADWAND [POL DETSON 2]  ZELAZNY, ROGER  C/P,H,LP
MAELSTROM [@ULTIMA TECHNOCRAT WAR #3]  ANDREWS, AUSTEN  P

MAELSTROM [TWINS OF PETAYBEE 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

MAELSTROM [@VENUS PRIME #2]  PREUSS, PAUL  C/P:2,P
MAELSTROM [RIFTERS 2]  WATTS, PETER  H,P
MAELSTROM’S EYE, THE [@SPELLJAMMER CLOAKMASTER #3]  MOORE, ROGER E.  P
MAEVE [DIADEM 4]  CLAYTON, JO  C/P:2,P
MAGAZINE BEACH  GANNETT, LEWIS  H,LP

MAGAZINE OF FANTASY AND SCIENCE FICTION:

A 30-YEAR RETROSPECTIVE, THE FERMAN, EDWARD L.

HA,SFWA

MAGE [@WORLD OF DARKNESS PREDATOR & PREY #5]  BOWEN, CARL  P

The MITSFS Titledex  September 24, 2007  Page 216
TITLE  AUTHOR  CODE
MAGE-BORN CHILD, THE [SERVANTS OF ARK 3]  WYLIE, JONATHAN  C/P:2,P
MAGE HEART [DION HOLYHANDS 2]  ROUTLEY, JANE  C/P,P
MAGE IN THE IRON MASK, THE [@FORGOTTEN REALMS NOBLES #4]  THOMSEN, BRIAN M.  P
MAGE OF CLOUDS [@CLOUDMAGES #2]  FARRELL, S. L.  P
MAGE QUEST [WIZARD OF YURT 3]  BRITTAIN, C. DALE  C/P,SFWA-P,P
MAGEBORN TRAITOR, THE [@EXILES #2]  RAWN, MELANIE  C/P,P,H
MAGEHOUND, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #1]  CUNNINGHAM, ELAINE  P
MAGELLAN  ANDERSON, COLIN  H,P
MAGES DE DEREB, LES  RICHARD-BESSIERE, F.  FP
MAGESTONE, THE [WITCH WORLD SECRETS]  NORTON, ANDRE/SCHAUB, MARY H.  C/P,P
MAGESTORM [@WARHAMMER]  GREEN, JONATHAN  P
MAGIA NEGRA: AMULETOS Y TALISMANES, LA  ANONYMOUS  FP
MAGIC & MALICE [MAIRELON AND KIM 1,2]  WREDE, PATRICIA C.  H
MAGIC AND THE HEALING, THE [BJ VAUGHAN 1]  O’DONOHOE, NICK  P,SFWA-P
MAGIC BITES  ANDREWS, ILONA  P
MAGIC BOOKS, THE  NORTON, ANDRE  P
MAGIC CARPET AND THE CEMENT WALL, THE  VIXEN, RICHARD M.  H,VLP
MAGIC CARPET MAGAZINE, THE  ODYSSEY PUBLICATIONS  SR-LPA
MAGIC CASEMENT [MAN OF HIS WORD 1]  DUNCAN, DAVE  C/P,P
MAGIC COTTAGE, THE  HERBERT, JAMES  H
MAGIC CUP, THE  GREELEY, ANDREW M.  C/P:2,P
MAGIC DEAD, THE [@CHANGELING #3]  GARRISON, PETER  P
MAGIC ENGINEER, THE [RECLUCE 3]  MODESITT, L. E. JR.  P
MAGIC FOR BEGINNERS  LINK, KELLY  LP
MAGIC FOR SALE  DAVIDSON, AVRAM  C/PA,PA
MAGIC GOES AWAY, THE [MAGIC 1]  NIVEN, LARRY  H:2,LP
MAGIC IN ITHKAR 1  NORTON, ANDRE/ADAMS, ROBERT  C/PA,LPA,SFWA
MAGIC IN ITHKAR 2  NORTON, ANDRE/ADAMS, ROBERT  SFWA,SFWA-P,LP
MAGIC IN ITHKAR 3  NORTON, ANDRE/ADAMS, ROBERT  PA,SFWA-P
MAGIC IN ITHKAR 4  NORTON, ANDRE/ADAMS, ROBERT  C/PA,PA,SFWA-P
MAGIC ISLAND, THE [DOC SAVAGE #89]  ROBESON, KENNETH  C/P,@P
MAGIC KINGDOM FOR SALE — SOLD! [LANDOVER 1]  BROOKS, TERRY  C/P,H:2,P
MAGIC LABYRINTH, THE [RIVERWORLD 4]  FARMER, PHILIP JOSE  C/P,P,H,SFWA
MAGIC LOST, TROUBLE FOUND  SHEARIN, LISA  P
MAGIC-LOVER’S TREASURY OF THE FANTASTIC, A  WEIS, MARGARET  LPA
MAGIC MAN, THE  BEAUMONT, CHARLES  C/P:2,P
MAGIC MAY RETURN, THE [MAGIC 2]  NIVEN, LARRY  C/PA,LPA
MAGIC MOSCOW, THE  PINKWATER, DANIEL  SFWA-P
MAGIC OF ATLANTIS, THE  CARTER, LIN  C/PA,PA

MAGIC OF CHRISTMAS, THE SILBERSACK, JOHN/

SCHELLING, CHRISTOPHER

PA

MAGIC OF KRYNN, THE [@DRAGONLANCE TALES #1]  WEIS, MARGARET/HICKMAN, TRACY  PA,C/PA
MAGIC OF OZ, THE [OZ #13]  BAUM, L. FRANK  H,@P
MAGIC OF RECLUCE, THE [RECLUCE 1]  MODESITT, L. E. JR.  H:2,P
MAGIC OF XANTH, THE [XANTH 1,2,3]  ANTHONY, PIERS  H
MAGIC OR NOT? [MAGIC OR NOT? 1]  EAGER, EDWARD  H,LP
MAGIC SHOP, THE  DOLBIER, MAURICE  H
MAGIC SHOP, THE  LITTLE, DENISE  PA
MAGIC SHOW [@DARK FORCES 2]  BRIDGES, LAURIE/ALEXANDER, PAUL  P
MAGIC STONE, THE  KOOIKER, LEONIE  SFWA
MAGIC STREET  CARD, ORSON SCOTT  H
MAGIC STUDY [YELINA 2]  SNYDER, MARIA V.  H
MAGIC TAILS  GREENBERG, MARTIN H./PACK, JANET  PA
MAGIC: THE FINAL FANTASY COLLECTION  ASIMOV, ISAAC  P
MAGIC THREE OF SOLATIA, THE  YOLEN, JANE  P
MAGIC TIME  REED, KIT  H,SFWA,P
MAGIC TIME [MAGIC TIME]  ZICREE, MARC SCOTT/HAMBLY, BARBARA  H

MAGIC TIME: ANGELFIRE [MAGIC TIME] ZICREE, MARC SCOTT/

BOHNHOFF, MAYA KAATHRYN

H

MAGIC TO DO  MCKENZIE, MELINDA  P
MAGIC TOUCH, THE  NYE, JODY LYNN  P
MAGIC TOYBOX, THE  LITTLE, DENISE  PA
MAGIC TOYSHOP, THE  CARTER, ANGELA  P
MAGIC TUNNEL, THE  EMERSON, CAROLINE D.  C/P:2,LP
MAGIC WARS, THE [@WILD MAGIC #3]  CLAYTON, JO  C/P,P
MAGICAL BEGINNINGS  SILVER, STEVEN H./GREENBERG, MARTIN H. PA
MAGICAL COACH, THE  ADLER, C. S.  P
MAGICAL MIND ADVENTURE OF HANNAH AND COLDY COLDY, THE  ROBERTS, THOM  H
MAGICAL MONARCH OF MO, THE  BAUM, L. FRANK  H,LP

MAGICAL WISHES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

MAGICATS!  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
MAGICATS 2  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
MAGICIAN [RIFTWAR 1,2]  FEIST, RAYMOND E.  H,SFWA:2

The MITSFS Titledex  September 24, 2007  Page 217
TITLE  AUTHOR  CODE
MAGICIAN: APPRENTICE [RIFTWAR 1]  FEIST, RAYMOND E.  C/P:3,P
MAGICIAN: MASTER [RIFTWAR 2]  FEIST, RAYMOND E.  C/P:2,P
MAGICIAN OF MARS, THE  HAMILTON, EDMOND  C/P:2,P
MAGICIANS, THE  GUNN, JAMES E.  C/P,SFWA,P,H
MAGICIAN’S GAMBIT [BELGARIAD 3]  EDDINGS, DAVID  C/P:2,P
MAGICIANS’ GUILD, THE [BLACK MAGICIAN 1]  CANAVAN, TRUDI  C/P,P
MAGICIAN’S NEPHEW, THE [NARNIA 6]  LEWIS, C. S.  C/P:2,P
MAGICIANS OF CAPRONA, THE [CHRESTOMANCI 3]  JONES, DIANA WYNNE  C/P,P
MAGICIANS OF GOR [GOR 25]  NORMAN, JOHN  C/P:2,P
MAGICIANS OF NIGHT, THE [SUN-CROSS 2]  HAMBLY, BARBARA  C/P:2,P
MAGICIANS OF QUALITY [CECILIA & KATHERINE 1,2]  WREDE, PATRICIA C./STEVERMER, CAROLINE H
MAGICIAN’S SLEEVE, THE  CONAWAY, J. C.  P
MAGICIAN’S WARD [MAIRELON AND KIM 2]  WREDE, PATRICIA C.  H
MAGICK OF CAMELOT, THE [CAMELOT 3]  LANDIS, ARTHUR H.  C/P:2,P
MAGICKERS, THE [MAGICKERS 1]  DRAKE, EMILY  P
MAGICKS AND CEREMONYES  JEANYNE  [C/RAN-P]
MAGICNET  DE CHANCIE, JOHN  C/P,P,H
MAGIC’S PAWN [LAST HERALD MAGE 1]  LACKEY, MERCEDES  C/P:2,P
MAGIC’S PRICE [LAST HERALD MAGE 3]  LACKEY, MERCEDES  C/P:2,P
MAGIC’S PROMISE [LAST HERALD MAGE 2]  LACKEY, MERCEDES  C/P:2,P
MAGIC’S SILKEN SNARE [SILKEN MAGIC 1]  GILLIGAN, ELIZABETH  P
MAGI’I OF CYADOR [RECLUCE 10]  MODESITT, L. E. JR.  H:2,P
MAGNETIC BRAIN  GRIDBAN, VOLSTED  SR-P
MAGNETIC MAN, THE [AVENGERS #8]  DANIELS, NORMAN  C/P:2,P
MAGNIFI-CAT  SHEEHAN, CAROLYN/SHEEHAN, EDMUND  SFWA
MAGNIFICAT [GALACTIC MILIEU 3]  MAY, JULIAN  C/P,P,H
MAGNIFICENT GALLERY, THE  MONTELEONE, THOMAS F.  SFWA-P

MAGNIFICENT SHOWBOATS OF THE LOWER VISSEL RIVER, LUNE XXIII SOUTH, BIG PLANET, THE [@VANCE INTEGRAL EDITION #19] VANCE, JACK

SR-H

MAGNIFICENT WILF, THE  DICKSON, GORDON R.  H,P
MAGUS, THE  FOWLES, JOHN  P
MAGUS REX  LOVEJOY, JACK  C/P:2,P
MAGYK [SEPTIMUS HEAP 1]  SAGE, ANGIE  H
MAHABHARATA  ANONYMOUS  H
MAHARAJAH AND OTHER STORIES, THE  WHITE, T. H.  C/P:2,P
MAHARS OF PELLUCIDAR  HOLMES, JOHN ERIC  P
MAHOGANY TRINROSE [SIME/GEN 4]  LICHTENBERG, JACQUELINE  C/P:2,SFWA,P
MAIA  ADAMS, RICHARD  C/P,P,H
MAIA’S VEIL [CLOUDSHIPS OF ORION 2]  MCALLISTER, P. K.  P
MAID OF THURO  LANE, JOHN  H
MAIDEN FLIGHT  VINICOFF, ERIC  C/P:2,SFWA-P,P
MAIDEN, MATRON, CRONE  GREENBERG, MARTIN H./HUGHES, KERRIE  C/PA,PA
MAIDEN OF PAIN [@FORGOTTEN REALMS PRIESTS]  FRANKLIN, KAMERON M.  P
MAIN EVENT [@BATTLETECH 10]  LONG, JIM  C/P,P
MAIN EXPERIMENT, THE  HODDER-WILLIAMS, CHRISTOPHER  H:2,P
MAIN STREET D.O.A. [@PSI-MAN #3]  DAVID, PETER  P
MAINLINE  CHRISTIAN, DEBORAH  P
MAINSPRING  LAKE, JAY  H
MAIRELON THE MAGICIAN [MAIRELON AND KIM 1]  WREDE, PATRICIA C.  H,P
MAITRE DU MONDE  VERNE, JULES  FP
MAJESTIC  STRIEBER, WHITLEY  H
MAJESTRUM [ARCHONATE/HENGHIS HAPTHORN]  HUGHES, MATTHEW  H
MAJII, THE [DOC SAVAGE #60]  ROBESON, KENNETH  C/P,@P
MAJIPOOR CHRONICLES [MAJIPOOR 2]  SILVERBERG, ROBERT  H:2,P
MAJOR [DIRIGENT MERCENARY CORPS 4]  SHELLEY, RICK  C/P,P
MAJOR INGREDIENTS  RUSSELL, ERIC FRANK  H
MAJOR MATT MASON: MOON MISSION  ELRICK, GEORGE S.  BLB
MAJOR OPERATION [SECTOR GENERAL 3]  WHITE, JAMES  C/P:2,P
MAJYK BY ACCIDENT [KENDAR GANGLE 1]  FRIESNER, ESTHER M.  C/P,SFWA-P,P
MAJYK BY DESIGN [KENDAR GANGLE]  FRIESNER, ESTHER M.  C/P,P
MAJYK BY HOOK OR CROOK [KENDAR GANGLE 2]  FRIESNER, ESTHER M.  P,SFWA-P
MAKE IT SO [STAR TREK TNG]  ROBERTS, WESS/ROSS, BILL  H
MAKE ROOM! MAKE ROOM!  HARRISON, HARRY  C/P,P,H
MAKE WAY FOR DRAGONS! [CALIFORNIA DRAGONS 1]  GUNNARSSON, THORARINN  C/P,P
MAKER [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
MAKER OF DUNE, THE  O’REILLY, TIMOTHY  SFWA
MAKER OF GARGOYLES, THE  SMITH, CLARK ASHTON  LP
MAKER OF MOONS, THE  CHAMBERS, ROBERT W.  H
MAKER OF UNIVERSES, THE [WORLD OF TIERS 1]  FARMER, PHILIP JOSE  C/P:2,P
MAKESHIFT GOD, THE  GRIFFIN, RUSSELL M.  C/P:2,P
MAKESHIFT ROCKET, THE  ANDERSON, POUL  DF139
MAKING BOOK  NIELSEN HAYDEN, TERESA  SR-LP
MAKING HISTORY  FRY, STEPHEN  H,LP
MAKING HISTORY  MCAULEY, PAUL J.  GZD07306

The MITSFS Titledex  September 24, 2007  Page 218
TITLE  AUTHOR  CODE
MAKING OF A MONSTER, THE  PETERSEN, GAIL  P
MAKING OF A MOON, THE  CLARKE, ARTHUR C.  R/REF-H
MAKING OF DOCTOR WHO, THE  DICKS, TERRANCE/HULKE, MALCOLM  R/REF-P,C/REF-P:2
MAKING OF INDEPENDENCE DAY, THE  ABERLY, RACHEL/ENGEL, VOLKER  R/REF-H
MAKING OF KUBRICK’S 2001, THE  AGEL, JEROME  R/REF-P,C/REF-P:2
MAKING OF QUANTUM LEAP, THE [QUANTUM LEAP]  SCHUSTER, HAL  R/REF-P
MAKING OF SPACE: 1999, THE  HEALD, TIM  R/REF-P,C/REF-P

MAKING OF STAR TREK, THE [STAR TREK] WHITFIELD, STEPHEN E./ RODDENBERRY, GENE

R/REF-P,C/REF-P:2

MAKING OF STAR TREK DEEP SPACE NINE, THE [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

R/REF-H

MAKING OF STAR TREK II: THE WRATH OF KHAN, THE [STAR TREK]  ASHERMAN, ALLAN  R/REF-P
MAKING OF STAR TREK: THE MOTION PICTURE, THE [STAR TREK]  SACKETT, SUSAN/RODDENBERRY, GENE  R/REF-P,C/REF-H:2
MAKING OF STAR WARS: THE RETURN OF THE JEDI, THE  PEECHER, JOHN PHILLIP  R/REF-P

MAKING OF THE REPRESENTATIVE FOR PLANET 8, THE [CANOPUS IN ARGOS 4] LESSING, DORIS

H,LP

MAKING OF THE TREK CONVENTIONS, OR HOW TO THROW A PARTY FOR 12,000 OF YOUR MOST INTIMATE FRIENDS, THE WINSTON, JOAN

SFWA

MAKRA CHORIA  MAYHAR, ARDATH  SFWA
MALACIA TAPESTRY, THE  ALDISS, BRIAN W.  C/P,SFWA,P,H
MALADY OF MAGICKS, A [EBENEZUM 1]  GARDNER, CRAIG SHAW  C/P:2,P
MALAFRENA  LE GUIN, URSULA K.  H:2,SFWA,P
MALE RESPONSE, THE  ALDISS, BRIAN W.  SR-P
MALEDICTE  ROBINS, LANE  LP
MALEVIL  MERLE, ROBERT  C/P,P,H
MALICIOUS INTENT [@BATTLETECH 24]  STACKPOLE, MICHAEL A.  P
MALIFESTRO QUEST, THE [ZORK 2]  MERETZKY, S. ERIC  P
MALL, THE  KAHN, STEVE  P
MALL, MAYHEM AND MAGIC  LISLE, HOLLY/GUIN, CHRIS  C/P:2,SFWA-P,P
MALL PURCHASE NIGHT  COOK, RICK  C/P:2,SFWA-P,P
MALLEUS [WARHAMMER 40K EISENHORN 2]  ABNETT, DAN  P
MALLWORLD [MALLWORLD]  SUCHARITKUL, SOMTOW  H:2,P
MALZBERG AT LARGE  MALZBERG, BARRY N.  P
MAMMALIA  LANE, JOHN  SR-P
MAMMOTH  VARLEY, JOHN  H
MAMMOTH BOOK OF ARTHURIAN LEGENDS, THE  ASHLEY, MIKE  LPA
MAMMOTH BOOK OF AWESOME COMIC FANTASY, THE  ASHLEY, MIKE  LPA
MAMMOTH BOOK OF BEST NEW HORROR, THE  JONES, STEPHEN  LPA
MAMMOTH BOOK OF COMIC FANTASY, THE  ASHLEY, MIKE  HA,LPA
MAMMOTH BOOK OF COMIC FANTASY II, THE  ASHLEY, MIKE  HA
MAMMOTH BOOK OF EXTREME SCIENCE FICTION, THE  ASHLEY, MIKE  LPA

MAMMOTH BOOK OF FANTASTIC SCIENCE FICTION, THE [@MAMMOTH SF 5] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA,LPA

MAMMOTH BOOK OF FANTASY, THE  ASHLEY, MIKE  LPA
MAMMOTH BOOK OF FRANKENSTEIN, THE  JONES, STEPHEN  PA
MAMMOTH BOOK OF FUTURE COPS, THE  JAKUBOWSKI, MAXIM/CHRISTIAN, M.  LPA
MAMMOTH BOOK OF GHOST STORIES, THE  DALBY, RICHARD  LPA

MAMMOTH BOOK OF GOLDEN AGE SCIENCE FICTION, THE ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

MAMMOTH BOOK OF MODERN SCIENCE FICTION, THE [@MAMMOTH SF 6] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

MAMMOTH BOOK OF NEW COMIC FANTASY, THE  ASHLEY, MIKE  HA
MAMMOTH BOOK OF NEW JULES VERNE ADVENTURES, THE  ASHLEY, MIKE/BROWN, ERIC  LPA
MAMMOTH BOOK OF SCIENCE FICTION, THE  ASHLEY, MIKE  LPA
MAMMOTH BOOK OF SORCERERS’ TALES, THE  ASHLEY, MIKE  LPA
MAMMOTH BOOK OF 20TH CENTURY GHOST STORIES, THE  HAINING, PETER  LPA
MAMMOTH BOOK OF VAMPIRE STORIES BY WOMEN, THE  JONES, STEPHEN  HA
MAMMOTH BOOK OF VAMPIRES, THE  JONES, STEPHEN  LPA

MAMMOTH BOOK OF VINTAGE SCIENCE FICTION, THE [@MAMMOTH SF 3] WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

MAMMOTH HUNTERS, THE [EARTH’S CHILDREN 3]  AUEL, JEAN M.  C/P,P,H
MAN AGAINST TOMORROW  NOLAN, WILLIAM F.  PA,C/PA
MAN AND MONSTER [@PERRY RHODAN #36]  ACE  PA,C/PA
MAN AND THE COMPUTER  KEMENY, JOHN G.  [C/RAN-P]
MAN BETRAYED, A [BOOK OF WORDS 2]  JONES, J. V.  C/P,LP,H
MAN CALLED DESTINY, A  WRIGHT, LAN  DD311
MAN CALLED MILO MORAI, A [HORSECLANS 14]  ADAMS, ROBERT  C/P,P
MAN-EATER, THE  BURROUGHS, EDGAR RICE  SR-LP
MAN-EATERS OF CASCALON, THE [DIRSHAN 4]  LANCOUR, GENE  SFWA
MAN FROM ATLANTIS, THE  MCWILLIAMS, JUDITH  P
MAN FROM ATLANTIS, THE [@AVENGER #25]  ROBESON, KENNETH  C/P:2,P
MAN FROM ATLANTIS [MAN FROM ATLANTIS 1]  WOODLEY, RICHARD  C/P,P
MAN FROM BEYOND, THE  MULLER, JOHN E.  SR-P
MAN FROM EARTH, THE  DICKSON, GORDON R.  C/P:2,P
MAN FROM HARDIN, THE  VELON, ERNEST  LP

The MITSFS Titledex  September 24, 2007  Page 219
TITLE  AUTHOR  CODE
MAN FROM MUNDANIA [@XANTH #12]  ANTHONY, PIERS  C/P:2,P
MAN FROM P.I.G., THE  HARRISON, HARRY  C/P,P
MAN FROM TOMORROW, THE  STATTEN, VARGO  SR-P
MAN FROM TOMORROW  TUCKER, WILSON  C/P:2,P
MAN FROM U.N.C.L.E., THE [@MAN FROM UNCLE #1]  AVALLONE, MICHAEL  C/P:2,P
MAN IN A CAGE  STABLEFORD, BRIAN M.  H
MAN IN DUPLICATE  STATTEN, VARGO  SR-P
MAN IN HIS TIME  ALDISS, BRIAN W.  C/P,P,H
MAN IN QUESTION, THE (COVER FROM LAUMER’S NIGHT OF DELUSIONS)  GODEY, JOHN  SR-L
MAN IN THE BIRD CAGE, THE  MACLEAN, KATHERINE  P

MAN IN THE CAGE, THE [@VANCE INTEGRAL EDITION #14]/ DEADLY ISLES, THE [@VANCE INTEGRAL EDITION #14] VANCE, JACK

SR-H

MAN IN THE DARKSUIT, THE  CARO, DENNIS R.  C/P,P
MAN IN THE HIGH CASTLE, THE  DICK, PHILIP K.  C/P,P,H:2
MAN IN THE MAZE, THE  SILVERBERG, ROBERT  C/P:2,P
MAN IN THE MOON  FOLEY, JOHN  SR-P
MAN IN THE MOON MUST DIE, THE  BREDENBERG, JEFF  C/P,SFWA-P,P
MAN IN THE MOONE AND OTHER LUNAR FANTASIES, THE  PIZOR, FAITH K./COMP, T. ALLAN  HA
MAN IN THE RUBBER MASK, THE [RED DWARF]  LLEWELLYN, ROBERT  R/REF-P
MAN IN THE TREE, THE  KNIGHT, DAMON  P
MAN IN THE VELVET MASK, THE [DOCTOR WHO MISSING ADVENTURES]  O’MAHONY, DANIEL  C/P,P
MAN-KZIN WARS, THE [@MAN-KZIN WARS #1]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 2 [@MAN-KZIN WARS #2]  NIVEN, LARRY  PA,C/PA,SFWA-P:2
MAN-KZIN WARS 3 [@MAN-KZIN WARS #3]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 4 [@MAN-KZIN WARS #4]  NIVEN, LARRY  PA,C/PA
MAN-KZIN WARS 5 [@MAN-KZIN WARS #5]  NIVEN, LARRY  PA,C/PA,SFWA-P
MAN-KZIN WARS 6 [@MAN-KZIN WARS #6]  NIVEN, LARRY  C/PA,PA
MAN-KZIN WARS 7 [@MAN-KZIN WARS #7]  NIVEN, LARRY  PA,C/PA
MAN-KZIN WARS 9 [@MAN-KZIN WARS #9]  NIVEN, LARRY  C/PA,PA
MAN-KZIN WARS 11 [@MAN-KZIN WARS #11]  NIVEN, LARRY  HA
MAN-MADE SATELLITES  LEY, WILLY  [C/RAN-H]
MAN-MADE WOMAN, THE [GIRL FACTORY 3]  MURPHY, ROBERT FRANKLIN  P
MAN O’ WAR [BENTON HAWKES 1]  SHATNER, WILLIAM  H,P
MAN OBSESSED, A  NOURSE, ALAN E.  DD096
MAN OF BRONZE, THE [@LARA CROFT TOMB RAIDER]  GARDNER, JAMES ALAN  P
MAN OF BRONZE, THE [DOC SAVAGE #1]  ROBESON, KENNETH  C/P:2,@P
MAN OF DOUBLE DEED, A  DAVENTRY, LEONARD  C/P,P,H
MAN OF EARTH  BUDRYS, ALGIS  C/P:2,P
MAN OF GOLD, THE [TEKUMEL 1]  BARKER, M. A. R.  C/P,P
MAN OF MANY MINDS  EVANS, E. EVERETT  C/P:2,P
MAN OF METAL  TORRO, PEL  H
MAN OF TWO WORLDS  HERBERT, FRANK/HERBERT, BRIAN  C/P,SFWA-P,P,H
MAN OF TWO WORLDS  JONES, RAYMOND F.  C/P:2,P
MAN OF TWO WORLDS  SCHWARTZ, JULIUS/THOMSEN, BRIAN M.  R/REF-P
MAN OF TWO WORLDS  STATTEN, VARGO  SR-P
MAN ON ANOTHER WORLD  EHRENSVARD, GOSTA  [C/RAN-H]
MAN ON THE MOUNTAIN  CARROLL, GLADYS HASTY  P
MAN PLUS [MAN PLUS 1]  POHL, FREDERIK  C/P,P,H,SFWA
MAN RESPONSIBLE, THE  ROBINETT, STEPHEN  C/P:2,P
MAN RIDES THROUGH, A [MORDANT’S NEED 2]  DONALDSON, STEPHEN R.  H:2,P
MAN THE FUGITIVE [@PLANET OF THE APES #1]  EFFINGER, GEORGE ALEC  C/P:2,P
MAN, THE HUNTED ANIMAL [@RETURN TO THE PLANET OF THE APES #3]  ARROW, WILLIAM  P
MAN THE WORLDS REJECTED, THE  DICKSON, GORDON R.  C/P:2,SFWA-P,P
MAN UNWEPT  WHALEY, STEPHAN V./COOK, STANLEY J.  LP
MAN WHO ATE THE WORLD, THE  POHL, FREDERIK  C/P:2,P
MAN WHO AWOKE, THE  MANNING, LAURENCE  C/P:2,P
MAN WHO CAME BACK, THE  THANET, NEIL  SR-P
MAN WHO CONQUERED TIME, THE  MULLER, JOHN E.  SR-P
MAN WHO CORRUPTED EARTH, THE  EDMONDSON, G. C.  C/P,P
MAN WHO COULD CHEAT DEATH, THE  LYNDON, BARRE/SANGSTER  P
MAN WHO COULDN’T SLEEP, THE  MAINE, CHARLES ERIC  H:2
MAN WHO COUNTS, THE  ANDERSON, POUL  C/P:2,P
MAN WHO FELL TO EARTH, THE  TEVIS, WALTER  C/P:2,LP,P
MAN WHO FOLDED HIMSELF, THE  GERROLD, DAVID  H:2,P,SFWA
MAN WHO FOLDED HIMSELF, THE (REVISED)  GERROLD, DAVID  LP
MAN WHO HAD NO IDEA, THE  DISCH, THOMAS M.  C/P:2,P
MAN WHO JAPED, THE  DICK, PHILIP K.  DD193
MAN WHO KILLED HIS BROTHER, THE  STEPHENS, REED  C/P,P
MAN WHO LIKED WOMEN, THE  BRANDEL, MARC  P
MAN WHO LIVED FOREVER, THE  MILLER, R. DEWITT/HUNGER, ANNA  DD162
MAN WHO LIVED IN INNER SPACE, THE  FEDERBUSH, ARNOLD  C/P,P

MAN WHO LOST THE SEA, THE [COMPLETE STORIES OF THEODORE STURGEON 10] STURGEON, THEODORE

H

MAN WHO LOVED MARS, THE [MAN WHO LOVED MARS 1]  CARTER, LIN  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 220
TITLE  AUTHOR  CODE
MAN WHO LOVED THE MIDNIGHT LADY, THE  MALZBERG, BARRY N.  SFWA
MAN WHO MASTERED TIME, THE  CUMMINGS, RAY  H,DD173
MAN WHO MELTED, THE  DANN, JACK  H,P
MAN WHO MET HIMSELF AND OTHER STORIES, THE  CRAWSHAY-WILLIAMS, ELIOT  H
MAN WHO NEVER MISSED, THE [MATADOR 1]  PERRY, STEVE  C/P,P
MAN WHO PULLED DOWN THE SKY, THE  BARNES, JOHN  C/P:2,P
MAN WHO SAW TOMORROW, THE  SUTTON, JEFF  DH095
MAN WHO SHOOK THE EARTH, THE [DOC SAVAGE #43]  ROBESON, KENNETH  C/P:2,@P
MAN WHO SOLD THE MOON, THE  HEINLEIN, ROBERT A.  C/P:2,B,P

MAN WHO STOLE TOMORROW, THE [AVENGERS (MARVEL COMICS CHARACTERS)] MICHELINIE, DAVID

C/P,P

MAN WHO TURNED INTO HIMSELF, THE  AMBROSE, DAVID  H
MAN WHO UPSET THE UNIVERSE, THE  ASIMOV, ISAAC  C/P:2,P
MAN WHO USED THE UNIVERSE, THE  FOSTER, ALAN DEAN  H:2,P
MAN WHO WANTED STARS, THE  MCLAUGHLIN, DEAN  C/P:2,P
MAN WHO WAS THURSDAY, THE  CHESTERTON, G. K.  C/P,P
MAN WHO WOULD NOT DIE, THE  PAGE, THOMAS  P
MAN WHOSE NAME WOULDN’T FIT, THE  TYLER, THEODORE  H:2,P
MAN WITH A THOUSAND NAMES, THE  VAN VOGT, A. E.  C/P:2,P
MAN WITH ABSOLUTE MOTION, THE  WATER, SILAS  SR-P
MAN WITH NINE LIVES, THE  ELLISON, HARLAN  DD413
MAN WITH ONLY ONE HEAD, THE  BARR, DENSIL NEVE  SR-P
MAN WITH THE GOLDEN GUN, THE [JAMES BOND 13]  FLEMING, IAN  C/P,P
MAN WITH THE GOLDEN TORC, THE  GREEN, SIMON R.  H
MAN WITH THREE EYES, THE  ARCH, E. L.  H
MAN WITH TWO FACES, THE [@PERRY RHODAN #104]  ACE  C/PA,PA
MAN WITHOUT A PLANET, THE [GREAT IMPERIUM]  CARTER, LIN  DG606
MAN WITHOUT A PLANET, THE  DEL REY, LESTER  P
MANAGRA [DOCTOR WHO MISSING ADVENTURES]  MARLEY, STEPHEN  C/P:2,P
MANALACOR OF VELTAKIN  CRISTABEL  P
MANALONE  KAPP, COLIN  SR-P
MANCHURIAN CANDIDATE, THE  CONDON, RICHARD  P
MANDALA  BISCHOFF, DAVID F.  C/P:2,P
MANDALORIAN ARMOR, THE [@STAR WARS BOUNTY HUNTER WARS 1]  JETER, K. W.  C/P,P
MANDEN FRA I GAR/OJNE I VERDENSRUMMET  TUCKER, WILSON/CRANE, ROBERT  FP
MANDRAKE  COOPER, SUSAN  SR-P,C/P
MANDRAKE THE MAGICIAN  ASHMAN, HOWARD  C/P,P
MANDRICARDO [TERRA MAGICA 3]  CARTER, LIN  P
MANDRILL  GARDNER, RICHARD  P
MANFAC  CAIDIN, MARTIN  C/P,P
MANHATTAN TRANSFER  STITH, JOHN E.  H,P
MANHOUNDS OF ANTARES [DRAY PRESCOT 6]  AKERS, ALAN BURT  C/P,P
MANHUNTRESS, THE [@SPACEWAYS #7]  CLEVE, JOHN  P
MANIFEST DESTINY  LONGYEAR, BARRY B.  C/P:2,P
MANIFOLD: ORIGIN [REID MALENFANT 3]  BAXTER, STEPHEN  C/P,P,H
MANIFOLD: SPACE [REID MALENFANT 2]  BAXTER, STEPHEN  C/P,P,H:2
MANIFOLD: TIME [REID MALENFANT 1]  BAXTER, STEPHEN  C/P,P,H
MANITOU, THE  MASTERTON, GRAHAM  P,SFWA-P
MANJINN MOON [TY MERRICK 3]  VITOLA, DENISE  P
MANKIND ON THE RUN  DICKSON, GORDON R.  DD164
MANKIND UNDER THE LEASH  DISCH, THOMAS M.  DG597
MANKIND WITCH, A [SHADOW OF THE LION 3]  FREER, DAVE  H
MANLIB  CLEVE, JOHN  SR-L
MANN AUS DEM ANDEREN JAHRTAUSEND, DER  GROSS, RICHARD  FH
MANNA  CORREY, LEE  C/P:2,P
MANNA ENZYME, THE  HOYT, RICHARD  SFWA-P
MANNA FROM HEAVEN  ZELAZNY, ROGER  H
MANNED SPACECRAFT  GATLAND, KENNETH W.  [C/RAN-H]
MANNER FUR LACERTAE [@PERRY RHODAN #64]  DARLTON, CLARK  FPR
MAN’S MORTALITY  ARLEN, MICHAEL  H
MANSE, THE [MANSE 1]  CANTRELL, LISA W.  SFWA-P
MANSEED  WILLIAMSON, JACK  H:2,P
MANSHAPE  BRUNNER, JOHN  C/P:2,P
MANSIONS OF DARKNESS [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  LP
MANSIONS OF SPACE, THE  MORRESSY, JOHN  P
MANTA’S GIFT  ZAHN, TIMOTHY  H
MANTLE OF KENDIS-DAI, THE [@STARSHIELD 1]  WEIS, MARGARET/HICKMAN, TRACY  P
MANTON’S WORLD  FEARN, JOHN RUSSELL  LP
MANY-COLORED LAND, THE [PLIOCENE EXILE 1]  MAY, JULIAN  C/P,P,H

MANY-COLORED LAND, THE [PLIOCENE EXILE 1,2]/ GOLDEN TORC, THE [PLIOCENE EXILE 1,2] MAY, JULIAN

H:2

MANY DIMENSIONS  WILLIAMS, CHARLES  H,LP
MANY FACES OF VAN HELSING, THE  CAVELOS, JEANNE  LPA
MANY FUTURES, MANY WORLDS  CLARESON, THOMAS D.  R/REF-H

TITLE  AUTHOR  CODE
MANY WATERS  L’ENGLE, MADELEINE  C/P,P,H
MANY WORLDS OF ANDRE NORTON, THE  NORTON, ANDRE  H
MANY WORLDS OF BARRY MALZBERG, THE  MALZBERG, BARRY N.  C/P,P
MANY WORLDS OF LARRY NIVEN, THE  GUPTILL, PAUL/DRUMM, CHRIS  R/REF-P
MANY WORLDS OF MAGNUS RIDOLPH, THE  VANCE, JACK  DM141
MANY WORLDS OF MAGNUS RIDOLPH, THE (EXPANDED)  VANCE, JACK  C/P:2,P
MANY WORLDS OF SCIENCE FICTION, THE  BOVA, BEN  SR-HA,HA
MAORI  FOSTER, ALAN DEAN  C/P,SFWA-P,P
MAP OF DREAMS  RICKERT, M.  H
MAPPA MUNDI  ROBSON, JUSTINA  LP
MAPS  SLADEK, JOHN  LP
MAPS IN A MIRROR  CARD, ORSON SCOTT  H:2,SFWA
MAQUESTA KAR-THON [@DRAGONLANCE WARRIORS #2]  DANIELL, TINA  P
MAQUINAS DA ALEGRIA, AS  BRADBURY, RAY  FP
MAQUISARDE, THE  MARLEY, LOUISE  P
MARACOT DEEP, THE  DOYLE, ARTHUR CONAN  SR-H,C/P,P
MARATHON [MARATHON 1]  SMITH, DAVID ALEXANDER  P
MARATHON PHOTOGRAPH, THE  SIMAK, CLIFFORD D.  P
MARAUDER [@ROBOTS IN TIME 2]  WU, WILLIAM F.  P,SFWA-P
MARAUDERS OF GOR [GOR 9]  NORMAN, JOHN  C/P:2,P
MARBLEHEART  CALLANDER, DON  P
MARCH HARE NETWORK, THE [@WONDERLAND GAMBIT #2]  CHALKER, JACK L.  H:2,P
MARCH INTO DARKNESS, A [DESTINIES OF BLOOD AND STONE 2]  NEWCOMB, ROBERT  H
MARCH OR DIE [@FIFTH FOREIGN LEGION #1]  KEITH, ANDREW/KEITH, WILLIAM H. JR.  P
MARCH TO THE SEA [MARCH UPCOUNTRY 2]  WEBER, DAVID/RINGO, JOHN  H,P
MARCH TO THE STARS [MARCH UPCOUNTRY 3]  WEBER, DAVID/RINGO, JOHN  H,P
MARCH UPCOUNTRY [MARCH UPCOUNTRY 1]  WEBER, DAVID/RINGO, JOHN  H
MARCHERS OF VALHALLA  HOWARD, ROBERT E.  C/P:2,P
MARCHING MORONS, THE  KORNBLUTH, C. M.  C/P:2,P
MARCHING THROUGH GEORGIA [DRAKA 1]  STIRLING, S. M.  C/P:2,SFWA-P,P
MARCHING THROUGH PEACHTREE [WAR BETWEEN THE PROVINCES 2]  TURTLEDOVE, HARRY  H
MARCIA, THE INNOCENT [@PHENWICK #12]  KIMBROUGH, KATHERYN  P
MARCO POLO [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P
MARCO POLO AND THE SLEEPING BEAUTY  DAVIDSON, AVRAM/DAVIS, GRANIA  P,SFWA-P
MARCO’S MILLIONS [MARCO 1]  SLEATOR, WILLIAM  H,LP
MARGARET AND I  WILHELM, KATE  C/P,P
MARGARET, THE FAITHFUL [@PHENWICK #3]  KIMBROUGH, KATHERYN  P
MARGARETS, THE  TEPPER, SHERI S.  H
MARIA LOONEY AND THE COSMIC CIRCUS  BEATTY, JEROME JR.  LP
MARIA LOONEY AND THE REMARKABLE ROBOT  BEATTY, JEROME JR.  LP
MARIA LOONEY ON THE RED PLANET  BEATTY, JEROME JR.  LP
MARIANNE, THE MADAME, AND THE MOMENTARY GODS [MARIANNE 2]  TEPPER, SHERI S.  C/P:2,P
MARIANNE, THE MAGUS, AND THE MANTICORE [MARIANNE 1]  TEPPER, SHERI S.  C/P:2,P
MARIANNE, THE MATCHBOX, AND THE MALACHITE MOUSE [MARIANNE 3] TEPPER, SHERI S.  C/P,P
MARIEL OF REDWALL [REDWALL]  JACQUES, BRIAN  C/P,P,H
MARINER-MARS 1964/FINAL PROJECT REPORT  NASA  [C/RAN-H]
MARINERS OF SPACE  COLLINS, ERROL  H
MARION ISLE  HAGGARD, H. RIDER  H
MARION ZIMMER BRADLEY’S DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
MARION’S WALL  FINNEY, JACK  C/P,P
MARK OF CHAOS [@WARHAMMER MARK OF CHAOS]  REYNOLDS, ANTHONY  P
MARK OF DAMNATION [@WARHAMMER MARK OF CHAOS]  WALLIS, JAMES  P
MARK OF HERESY [@WARHAMMER]  WALLIS, JAMES  P
MARK OF MERLIN, THE  MCCAFFREY, ANNE  C/P,P
MARK OF RAN, THE [@SEA BEGGARS 1]  KEARNEY, PAUL  LP
MARK OF THE BEAST  MULLER, JOHN E.  SR-P
MARK OF THE CAT, THE  NORTON, ANDRE  C/P,SFWA,SFWA-P,P
MARK OF THE DRAGON [RUIN MIST 4]  STANEK, ROBERT  LP
MARK OF THE MANDRAGORA, THE [DOCTOR WHO]  MARVEL COMICS  CX
MARK OF THE MODERATELY VICIOUS VAMPIRE, THE [KENT MONTANA 4]  FENN, LIONEL  P
MARK OF THE RANI, THE [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  P
MARK OF THE VULTURE, THE [@G-8 #8]  HOGAN, ROBERT J.  SR-P
MARK OF THE YUAN-TI [@KNIGHTS OF THE SILVER DRAGON #12]  ROBERTS, KERRY DANIELS  LP
MARKED CARDS [@WILD CARDS NEW CYCLE 2]  MARTIN, GEORGE R. R.  PA
MARKED FOR DEATH [@EBERRON LOST MARK #1]  FORBECK, MATT  P
MARKED MAN, THE [MARKED MAN 1]  INGRID, CHARLES  C/P:2,P
MARKET FORCES  MORGAN, RICHARD  H
MARKET FORCES (PROOF)  MORGAN, RICHARD  SR-LP
MARKS OF OUR BROTHERS  LINDSKOLD, JANE  P
MARLBOROUGH STREET  BOWKER, RICHARD  C/P:2,SFWA,P
MARLFOX [REDWALL]  JACQUES, BRIAN  P
MAROONED [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  C/P,P
MAROONED (ORIGINAL)  CAIDIN, MARTIN  SR-P
MAROONED (REVISED)  CAIDIN, MARTIN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 222
TITLE  AUTHOR  CODE
MAROONED IN ORBIT  BALLOU, ARTHUR W.  H
MAROONED IN REALTIME  VINGE, VERNOR  C/P,P,H

MAROONED ON EDEN [ROCHEWORLD 3] FORWARD, ROBERT L./

FORWARD, MARGARET DODSON

C/P,SFWA-P,P

MAROONED ON MARS  DEL REY, LESTER  H:2,P
MARQUE AND REPRISAL [VATTA’S WAR 2]  MOON, ELIZABETH  H,P
MARRIAGE OF STICKS, THE  CARROLL, JONATHAN  LP
MARRIAGE OF THE LIVING DARK, THE [@CHUNG KUO 8]  WINGROVE, DAVID  P

MARRIAGES BETWEEN ZONES THREE, FOUR, AND FIVE, THE [CANOPUS IN ARGOS 2] LESSING, DORIS

H:2,LP

MARROW [MARROW 1]  REED, ROBERT  H:2,P
MARROW EATERS, THE  MOORE, HARRIS  P
MARS [JAMIE WATERMAN 1]  BOVA, BEN  C/P:2,P,H
MARS ARENA, THE [@DEATHLANDS]  AXLER, JAMES  C/P,P
MARS ATTACKS! [@MARS ATTACKS]  GEMS, JONATHAN  P
MARS CROSSING  LANDIS, GEOFFREY A.  H,P
MARS IS MY DESTINATION  LONG, FRANK BELKNAP  C/P:2,P
MARS MONOPOLY, THE  SOHL, JERRY  DD162
MARS ONE CREW MANUAL, THE  JOELS, KERRY MARK  R/REF-H
MARS PLUS [MAN PLUS 2]  POHL, FREDERIK/THOMAS, THOMAS T.  H,P
MARS PRIME  DIETZ, WILLIAM C.  P
MARS PROBES  CROWTHER, PETER  PA
MARS — THE RED PLANET  FARREN, MICK  P
MARS UNDERGROUND  HARTMAN, WILLIAM K.  P
MARS, WE LOVE YOU  HIPOLITO, JANE/MCNELLY, WILLIS E.  HA,PA,SFWA
MARTIAN, THE  DU MAURIER, GEORGE  H
MARTIAN CHILD, THE  GERROLD, DAVID  LP
MARTIAN CHRONICLES, THE  BRADBURY, RAY  H:2,SFWA,SR-P
MARTIAN DEATHTRAP [@MARS ATTACKS]  ARCHER, NATHAN  C/P,P,H
MARTIAN INCA, THE  WATSON, IAN  C/P,P,H
MARTIAN KNIGHTLIFE  HOGAN, JAMES P.  H,P
MARTIAN MARTYRS  COLERIDGE, JOHN  B
MARTIAN MISSILE, THE  GRINNELL, DAVID  P,DD465
MARTIAN ODYSSEY, A  WEINBAUM, STANLEY G.  SR-H,C/P:2,P
MARTIAN QUEST  BRACKETT, LEIGH  H
MARTIAN RACE, THE  BENFORD, GREGORY  H
MARTIAN RAINBOW  FORWARD, ROBERT L.  C/P,P,H
MARTIAN SEXPOT  O’NEILL, SCOTT  SR-L
MARTIAN SPHINX, THE  WOODCOTT, KEITH  C/P:2,P
MARTIAN SPRING  WILLIAMS, MICHAEL LINDSAY  C/P,P
MARTIAN TIME-SLIP  DICK, PHILIP K.  C/P:2,P
MARTIAN VIKING, THE  SULLIVAN, TIM  C/P,P
MARTIAN VISITORS, THE  LONG, FRANK BELKNAP  H
MARTIAN WAY, THE  ASIMOV, ISAAC  SR-H
MARTIAN WAY AND OTHER STORIES, THE  ASIMOV, ISAAC  H:2,P
MARTIAN WAY, THE (AUTOGRAPHED)  ASIMOV, ISAAC  SR-P
MARTIANS, THE [MARS]  ROBINSON, KIM STANLEY  H
MARTIANS AND MADNESS  BROWN, FREDRIC  H:2
MARTIANS, GO HOME  BROWN, FREDRIC  C/P,P,H
MARTIN THE WARRIOR [REDWALL]  JACQUES, BRIAN  P
MARTYR [@STAR TREK NEW FRONTIER #5]  DAVID, PETER  P
MARTYR  UTLEY, BRIAN W.  C/P,P
MARUNE: ALASTOR 933 [ALASTOR 2]  VANCE, JACK  C/P:2,P

MARUNE: ALASTOR 933 (VIE) [@VANCE INTEGRAL EDITION #30/ ALASTOR 2] VANCE, JACK

SR-H

MARVEL SUPERHEROES, THE  WEIN, LEN/WOLFMAN, MARV  PA
MARVELOUS LAND OF OZ, THE [OZ 2]  BAUM, L. FRANK  H:2,LP
MARY AND THE GIANT  DICK, PHILIP K.  H:2
MARY POPPINS OPENS THE DOOR  TRAVERS, P. L.  H
MARY REILLY  MARTIN, VALERIE  P

MARY-SUE EXTRUSION, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] STONE, DAVE

P

MARY’S MONSTER  LATHAM, WILLIAM  P
MASK AND THE SORCERESS, THE [PANDRAGORE 2]  JONES, DENNIS  H
MASK FOR THE GENERAL, A  GOLDSTEIN, LISA  H,P
MASK OF CHAOS  JAKES, JOHN  D78400
MASK OF CIRCE, THE  KUTTNER, HENRY  C/P:2,P
MASK OF CTHULHU, THE  DERLETH, AUGUST  C/P:2,P
MASK OF FU MANCHU, THE  ROHMER, SAX  C/P,P
MASK OF JON CULON, THE  LUDWIG, EDWARD W.  H
MASK OF LOKI, THE  ZELAZNY, ROGER/THOMAS, THOMAS T.  C/P:2,SFWA-P,P
MASK OF THE PHANTASM [@BATMAN]  GRAVEL, GEARY  P
MASK OF THE SORCERER, THE  SCHWEITZER, DARRELL  H,P
MASK OF THE SUN, THE  SABERHAGEN, FRED  C/P:2,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 223
TITLE  AUTHOR  CODE
MASK OF THE WIZARD [ELVEN KINGDOMS 1]  COOKE, CATHERINE  C/P,SFWA-P,P
MASKE: THAERY  VANCE, JACK  H:2,P
MASKE: THAERY (VIE) [@VANCE INTEGRAL EDITION #33]  VANCE, JACK  SR-H
MASKED INVASION, THE [OPERATOR 5 #1]  STEELE, CURTIS  SR-P
MASKERADE [DISCWORLD 18/DISCWORLD WITCHES 5]  PRATCHETT, TERRY  C/P,P,H
MASKS  PLAYBOY BOOKS  C/PA:2,PA
MASKS [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
MASKS OF SCORPIO [DRAY PRESCOT 31]  AKERS, ALAN BURT  C/P,P
MASKS OF THE ILLUMINATI  WILSON, ROBERT ANTON  H:2,P
MASKS OF THE MARTYRS [@RINGS OF THE MASTER #4]  CHALKER, JACK L.  C/P:2,P
MASKS OF THE OUTCASTS [DIPPLE 1,3]  NORTON, ANDRE  H
MASKS OF TIME, THE  SILVERBERG, ROBERT  C/P:2,P
MASKS OF TIME, THE/BORN WITH THE DEAD/DYING INSIDE  SILVERBERG, ROBERT  P
MASQUE  WILSON, F. PAUL/COSTELLO, MATTHEW J.  P
MASQUE DE L’OUBLI, LE  KOONTZ, DEAN R.  FP
MASQUE OF MANANA, THE  SHECKLEY, ROBERT  H
MASQUE OF SATAN [LUCIFER COVE 4]  COFFMAN, VIRGINIA  P
MASQUE WORLD [ANTHONY VILLIERS 3]  PANSHIN, ALEXEI  C/P:2,P
MASQUERADE [@ULTIMA TECHNOCRAT WAR #2]  ANDREWS, AUSTEN  P
MASQUERADES [FORGOTTEN REALMS HARPERS 10]  NOVAK, KATE/GRUBB, JEFF  P
MASQUES  BRIGGS, PATRICIA  C/P,P
MASQUES [@BEAUTY AND THE BEAST]  EMERSON, RU  P
MASQUES  PRONZINI, BILL  P
MASQUES  WILLIAMSON, J. N.  SR-HA,SFWA
MASQUES 2  WILLIAMSON, J. N.  SR-HA
MASQUES 3  WILLIAMSON, J. N.  SR-HA
MASQUES 4  WILLIAMSON, J. N.  SR-HA
MASS MEDIA AND MASS CULTURE  STANTON, FRANK  [C/RAN-H]
MASSACRE, THE [DOCTOR WHO]  LUCAROTTI, JOHN  SR-P
MASSACRE IN THE 22ND CENTURY [@FLASH GORDON #1]  ANONYMOUS  C/P:2,P
MASTER, THE [TIME MASTER 3]  COOPER, LOUISE  C/P:2,SFWA-P,P
MASTER, THE  WHITE, T. H.  C/P,P
MASTER AND FOOL [BOOK OF WORDS 3]  JONES, J. V.  C/P:2,LP
MASTER AND MARGARITA, THE  BULGAKOV, MIKHAIL  P
MASTER CURE, THE  BUCK, CHARLES H.  P
MASTER GOES HOME, THE [@TALES OF THE WEREWOLF CLAN #2]  MUNN, H. WARNER  SR-H
MASTER KEY: AN ELECTRICAL FAIRY TALE, THE  BAUM, L. FRANK  H,LP
MASTER-MIND MENACE  LUIGI, BELLI  SR-P
MASTER OF BORANGA [@RO-LAN #1]  SIROTA, MIKE  P
MASTER OF CHAINS [@FORGOTTEN REALMS FIGHTERS]  LEBOW, JESS  P
MASTER OF CHAOS, THE [SENTIENCE 2]  ADAMS, TERRY A.  C/P:2,P
MASTER OF DRAGONS [DRAGONVARLD 3]  WEIS, MARGARET  H

MASTER OF EARTH AND WATER [FIONN MAC CUMHAL 1] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

H,P

MASTER OF EVIL, THE [FALL OF THE FIRST WORLD 1]  SMITH, DAVID C.  P
MASTER OF HAWKS  BUSHYAGER, LINDA E.  C/P,P
MASTER OF LIES  MASTERTON, GRAHAM  H
MASTER OF LIFE AND DEATH  SILVERBERG, ROBERT  SFWA-P,DD237,P,C/P:2
MASTER OF MANY TREASURES [UNEXPECTED DRAGON 2]  BROWN, MARY  P
MASTER OF MEN! [@SPIDER #2]  STOCKBRIDGE, GRANT  P

MASTER OF MIDDLE-EARTH: THE FICTION OF J. R. R. TOLKIEN KOCHER, PAUL H.

R/REF-P,C/REF-H,

C/REF-P

MASTER OF MISFIT [@SPACEWAYS #5]  CLEVE, JOHN  P
MASTER OF NONE  WOOD, N. LEE  LP
MASTER OF NORRIYA, THE [ERTHRING 3]  DREW, WAYLAND  P
MASTER OF PAXWAX [GARDNER 1]  MANN, PHILLIP  H
MASTER OF SPACE  CLARKE, ARTHUR C.  C/P,P
MASTER OF SPACE AND TIME  RUCKER, RUDY  C/P:2,SFWA,P,H
MASTER OF THE CAULDRON [LORD OF THE ISLES 6]  DRAKE, DAVID  P
MASTER OF THE DARK GATE  JAKES, JOHN  P
MASTER OF THE DAY OF JUDGMENT, THE  PERUTZ, LEO  SR-P
MASTER OF THE ETRAX  LORY, ROBERT  C/P,P
MASTER OF THE FEARFUL DEPTHS [TITANS 2]  ADKINS, PATRICK H.  P
MASTER OF THE FIVE MAGICS  HARDY, LYNDON  C/P:2,P
MASTER OF THE HASHOMI [@BLADE #27]  LORD, JEFFREY  P
MASTER OF THE MICROBE, THE  SERVICE, ROBERT W.  H
MASTER OF THE NIGHT [MAGEVERSE MASTERS 1]  KNIGHT, ANGELA  P
MASTER OF THE SIDHE [SIDHE 3]  FLINT, KENNETH C.  C/P:2,P
MASTER OF THE STARS  HOSKINS, ROBERT  SR-P
MASTER OF THE UNDEAD  PAUL, HUGO  P
MASTER OF THE WORLD  VERNE, JULES  C/P:2,P,H
MASTER OF WHITE STORM, THE  WURTS, JANNY  C/P:2,SFWA-P,P
MASTER WALK [ADVOCACY 1]  LEE, SHARON/MILLER, STEVE  SR-LP
MASTER WOLF [@GREYHAWK #3]  ESTES, ROSE  C/P,P

The MITSFS Titledex  September 24, 2007  Page 224
TITLE  AUTHOR  CODE
MASTERHARPER OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
MASTERLESS MAN, THE  VERNE, JULES  H
MASTERMIND OF MARS, THE [BARSOOM #6]  BURROUGHS, EDGAR RICE  C/P,@P,H

MASTERMIND OF MARS, THE [BARSOOM 6]/ FIGHTING MAN OF MARS, A [BARSOOM 7] BURROUGHS, EDGAR RICE

H:2

MASTERMINDS OF FALKENSTEIN [@CASTLE FALKENSTEIN]  DE CHANCIE, JOHN  P
MASTERPIECE  UMPATHY, RANJAN  H
MASTERPIECES  CARD, ORSON SCOTT  HA
MASTERPIECES OF FANTASY AND ENCHANTMENT  HARTWELL, DAVID G.  HA
MASTERPIECES OF FANTASY AND WONDER  HARTWELL, DAVID G.  HA
MASTERPIECES OF SCIENCE FICTION  MOSKOWITZ, SAM  LPA
MASTERPIECES OF TERROR AND THE SUPERNATURAL  KAYE, MARVIN  SFWA
MASTERPLAY  WU, WILLIAM F.  C/P:2,P
MASTERS’ GAMBIT, THE [@ROBOTECH #20]  MCKINNEY, JACK  C/P,P
MASTERS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA
MASTERS OF DARKNESS, THE [LONE WOLF 12]  DEVER, JOE  C/P,P
MASTERS OF DARKNESS  ETCHISON, DENNIS  SFWA-P
MASTERS OF DARKNESS 2  ETCHISON, DENNIS  SFWA-P
MASTERS OF EVERON  DICKSON, GORDON R.  C/P,H,LP
MASTERS OF EVOLUTION  KNIGHT, DAMON  DD375
MASTERS OF FLUX AND ANCHOR [@SOUL RIDER #3]  CHALKER, JACK L.  C/P:2,SFWA-P,P
MASTERS OF GLASS  EASTON, M. COLEMAN  C/P,P
MASTERS OF HORROR  NORTON, ALDEN H.  PA
MASTERS OF SCIENCE FICTION  BELMONT BOOKS  PA,C/PA
MASTERS OF SOLITUDE, THE  KAYE, MARVIN/GODWIN, PARKE  H:2,P,SFWA
MASTERS OF SPACE  SMITH, E. E.  C/P:2,P
MASTERS OF TERROR I  HODGSON, WILLIAM HOPE  SR-P
MASTERS OF THE FIST  HUGHES, EDWARD P.  C/P,SFWA-P,P
MASTERS OF THE LAMP  LORY, ROBERT  D52180
MASTERS OF THE MAZE  DAVIDSON, AVRAM  C/P:2,P
MASTERS OF THE PIT OR BARBARIANS OF MARS [MICHAEL KANE 3]  MOORCOCK, MICHAEL  C/P:2,P
MASTERS OF THE VORTEX  SMITH, E. E.  C/P:2,P
MASTERS OF TIME  VAN VOGT, A. E.  C/P,P,H
MASTERS OF WAR [@MECHWARRIOR DARK AGE]  STACKPOLE, MICHAEL A.  P
MASTODONIA  SIMAK, CLIFFORD D.  H:2,P
MATADORA [MATADOR 2]  PERRY, STEVE  C/P:2,P
MATCH OF THE DAY [@DOCTOR WHO PAST DOCTORS]  BOUCHER, CHRIS  P
MATE  BURKA, LAUREN P.  SR-L
MATHEMAGICS [CHICKS IN CHAINMAIL 4]  BALL, MARGARET  C/P,P
MATHEMATICAL MAGPIE, THE  FADIMAN, CLIFTON  [C/RAN-H]
MATHEMATICIANS IN LOVE  RUCKER, RUDY  H
MATHEMATICS IN FUN AND IN EARNEST  COURT, NATHAN  [C/RAN-P]
MATHENAUTS  RUCKER, RUDY  HA,SFWA
MATHIAS SANDORF TOME II  VERNE, JULES  FP
MATILDA  DAHL, ROALD  H
MATILDA’S STEPCHILDREN [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
MATRIARCH [WESS’HAR WARS 4]  TRAVISS, KAREN  P
MATRIX  MASON, DOUGLAS R.  C/P,P
MATRIX [DOCTOR WHO PAST DOCTORS]  PERRY, ROBERT/TUCKER, MIKE  C/P,P
MATRIX CUBED [BUCK ROGERS INNER PLANETS 3]  BLOOM, BRITTON  C/P,P
MATRIX MAN  DIETZ, WILLIAM C.  C/P:2,P
MATTER FOR MEN, A [WAR AGAINST THE CHTORR 1]  GERROLD, DAVID  C/P,P

MATTER FOR MEN, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 1] GERROLD, DAVID

C/P,LP

MATTER OF HONOR, A [@STARGATE SG-1 #3]  MALCOLM, SALLY  P
MATTER OF HONOR, A [TWO-EDGED SWORD 2]  MARTIN, THOMAS K.  P
MATTER OF METALAW, A  CORREY, LEE  C/P,P
MATTER OF OATHS, A  WRIGHT, HELEN S.  C/P,P
MATTER OF TASTE, A [DRACULA]  SABERHAGEN, FRED  C/P,P,H
MATTER OF TIME, A  COOK, GLEN  C/P,P
MATTER’S END  BENFORD, GREGORY  C/P:2,P
MATTERS OF FORM  WHEELER, SCOTT  P
MATTHEW LOONEY AND THE SPACE PIRATES  BEATTY, JEROME JR.  C/P,LP
MATTHEW LOONEY IN THE OUTBACK  BEATTY, JEROME JR.  LP
MATTHEW LOONEY’S INVASION OF THE EARTH  BEATTY, JEROME JR.  C/P,LP
MATTHEW LOONEY’S VOYAGE TO THE EARTH  BEATTY, JEROME JR.  H,LP
MATTHEW’S DRAGON  COOPER, SUSAN  SFWA
MATTIMEO [REDWALL]  JACQUES, BRIAN  P
MAUGIS, YE SORCERER  GILHOOLEY, LORD  SR-H
MAURAI AND KITH  ANDERSON, POUL  C/P:2,P
MAUREEN BIRNBAUM, BARBARIAN SWORDSPERSON  EFFINGER, GEORGE ALEC  H:2
MAVERICK [ROBOTS AND ALIENS 5]  BETHKE, BRUCE  P,SFWA-P:2
MAWDRYN UNDEAD [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P:2
MAX AND ME AND THE TIME MACHINE  GREER, GERY/RUDDICK, BOB  P

The MITSFS Titledex  September 24, 2007  Page 225
TITLE  AUTHOR  CODE
MAX HEADROOM: 20 MINUTES INTO THE FUTURE  ROBERTS, STEVE  VLP
MAX ON EARTH [@OUT OF THIS WORLD #1]  KAYE, MARILYN  P
MAX SMART AND THE GHASTLY GHOST AFFAIR [GET SMART 9]  JOHNSTON, WILLIAM  P
MAX SMART AND THE PERILOUS PELLETS [GET SMART 4]  JOHNSTON, WILLIAM  C/P,P
MAX SMART LOSES CONTROL [GET SMART 8]  JOHNSTON, WILLIAM  P
MAX SMART — THE SPY WHO WENT OUT TO THE COLD [GET SMART 7]  JOHNSTON, WILLIAM  P
MAXIE’S DEMON  ROHAN, MICHAEL SCOTT  P
MAXIMUM EFFORT  EDMONDSON, G. C./KOTLAN, C. M.  P
MAXIMUM ICE [MAXIMUM ICE 1]  KENYON, KAY  P
MAXIMUM LIGHT  KRESS, NANCY  P
MAX’S BOOK [@PSI PATROL 3]  HURLEY, MAXWELL  P
MAXWELL’S DEMON  SHERWOOD, MARTIN  SR-P
MAXWELL’S DEMONS  BOVA, BEN  C/P,P,H
MAYA THE BEE  BONSELLS, WALDEMAR  H
MAYDAY  BLOCK, THOMAS H.  P
MAYDAY ORBIT  ANDERSON, POUL  P,DF104
MAYENNE [DUMAREST 9]  TUBB, E. C.  C/P:2,P
MAYFLIES  O’DONNELL, KEVIN JR.  C/P:2,P
MAYHEM IN MANHATTAN [SPIDER-MAN]  WEIN, LEN/WOLFMAN, MARV  P
MAZA OF THE MOON  KLINE, OTIS ADELBERT  C/P,P,H
MAZE IN THE HEART OF THE CASTLE, THE  GILMAN, DOROTHY  P
MAZE IN THE MIRROR, THE [@G.O.D. INC. #3]  CHALKER, JACK L.  C/P:2,SFWA-P,P
MAZE MAKER, THE  AYRTON, MICHAEL  P
MAZE OF DEATH, A  DICK, PHILIP K.  C/P,P,H
MAZE OF MOONLIGHT [STRANDS OF STARLIGHT 2]  BAUDINO, GAEL  C/P,P
MAZE OF PERIL, THE  HOLMES, JOHN ERIC  LP
MAZE OF STARS, A  BRUNNER, JOHN  C/P:2,P,H
MAZE OF WORLDS  LUMLEY, BRIAN  P
MAZES OF SCORPIO [DRAY PRESCOT 27]  AKERS, ALAN BURT  P
MAZEWAY  WILLIAMSON, JACK  H:2,P

MAZIRIAN THE MAGICIAN [@VANCE INTEGRAL EDITION #1/ DYING EARTH 1] VANCE, JACK

SR-H

MCANDREW CHRONICLES, THE  SHEFFIELD, CHARLES  C/P:2,P
MCATRIX DERIDED, THE  BROTHERS, ROBERTSKI  H
MCCADE ON THE RUN [SAM MCCADE 3,4]  DIETZ, WILLIAM C.  LP
MCCADE’S BOUNTY [SAM MCCADE 4]  DIETZ, WILLIAM C.  P

MCKAY’S MODERN ENGLISH-SWEDISH SWEDISH-ENGLISH DICTIONARY NOJD, RUBEN/TORNBERG, ASTRID/

ANGSTROM, MARGARITA

DICT

MCLENDON’S SYNDROME  FREZZA, ROBERT  C/P,P,H
MCSWEENEY’S ENCHANTED CHAMBER OF ATONISHING STORIES  CHABON, MICHAEL  LPA
MCSWEENEY’S MAMMOTH TREASURY OF THRILLING TALES  CHABON, MICHAEL  LPA
M.D., THE  DISCH, THOMAS M.  C/P,P
ME  THOMAS, THOMAS T.  C/P:2,P
MEAN STREETS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  DICKS, TERRANCE  P
MEANING IN STAR TREK  BLAIR, KARIN  R/REF-P,C/REF-P:2
MEANING OF LIFE, THE  SCHWEITZER, DARRELL  B
MEANING OF LIFF, THE  ADAMS, DOUGLAS/LLOYD, JOHN  [C/RAN-P]
MEANWHILE  HANDLEY, MAX  LP

MEASURE AND THE TRUTH, THE [@DRAGONLANCE RISE OF SOLAMNIA #3] NILES, DOUGLAS

P

MEASURE OF A HERO [@BATTLETECH 48]  PARDOE, BLAINE LEE  P
MEASURE OF A MAN [@HIGHLANDER]  HOLDER, NANCY  P
MEASURE OF THE UNIVERSE, THE  ASIMOV, ISAAC  R/REF-H
MEASURING A MERIDIAN  VERNE, JULES  H
MECHANICAL MONARCH, THE  TUBB, E. C.  DD266
MECHANISMO  HARRISON, HARRY  A
MECHASM  SLADEK, JOHN  C/P:2,P
MED SERIES, THE [MED 1,2,3]  LEINSTER, MURRAY  C/P:2,P
MED SHIP [MED]  LEINSTER, MURRAY  C/P,P
MEDALON [HYTHRUN 1]  FALLON, JENNIFER  H
MEDDLERS, THE  BONE, J. F.  SR-P
MEDEA: HARLAN’S WORLD  ELLISON, HARLAN  HA,LPA
MEDICINE FOR MELANCHOLY, A  BRADBURY, RAY  C/P,P
MEDICINE SHOW [TAYLOR’S ARK 2]  NYE, JODY LYNN  P
MEDIEVAL LEGENDS  JENNINGS, PHILIP S.  HA
MEDSTAR [@STAR WARS CLONE WARS MEDSTAR 1,2]  REAVES, MICHAEL/PERRY, STEVE  H
MEDUSA  INNES, HAMMOND  SFWA-P
MEDUSA  MORRIS, JANET/MORRIS, CHRIS  C/P:2,P
MEDUSA: A TIGER BY THE TAIL [@FOUR LORDS OF THE DIAMOND #4]  CHALKER, JACK L.  C/P:2,P
MEDUSA EFFECT, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  RICHARDS, JUSTIN  P,PA
MEDUSA ENCOUNTER, THE [@VENUS PRIME #4]  PREUSS, PAUL  C/P:3,P
MEDUSA FREQUENCY, THE  HOBAN, RUSSELL  LP
MEDUSA IN THE SHIELD, THE [DARK DESCENT 2]  HARTWELL, DAVID G.  PA
MEDUSA PLAGUE, THE [@DRAGONLANCE DEFENDERS OF MAGIC #2]  KIRCHOFF, MARY  P

The MITSFS Titledex  September 24, 2007  Page 226
TITLE  AUTHOR  CODE
MEDUSA WAR, THE [@ABC WARRIORS]  MILLS, PAT/MITCHELL, ALAN  P
MEDUSA’S CHILDREN  SHAW, BOB  SFWA-P,H:2,P,SFWA
MEEK, THE  MACKAY, SCOTT  P
MEET ME AT INFINITY  TIPTREE, JAMES JR.  H
MEET ME IN THE MOON ROOM  VUKCEVICH, RAY  LP
MEETING AT CORVALLIS, A [EMBERVERSE 3]  STIRLING, S. M.  H
MEETING AT INFINITY  BRUNNER, JOHN  C/P:2,DD507,P
MEETING OF THE WATERS, THE [@WATCHERS #1]  MOR, CAISEAL  P
MEETING THE BEAR: JOURNAL OF THE BLACK WARS  ZIMPEL, LLOYD  C/P,P
MEETING WITH MEDUSA, A  CLARKE, ARTHUR C.  TD1
MEGALOMANIA  WALLACE, IAN  P
MEGATOKYO 2  GALLAGHER, FRED  SCX
MEGLOS [DOCTOR WHO]  DICKS, TERRANCE  SR-P
MELANCHOLY ELEPHANTS  ROBINSON, SPIDER  C/P:2,SFWA-P,P
MELCHIOR’S FIRE [THREE KINGS 2]  CHALKER, JACK L.  H
MELMOTH THE WANDERER  MATURIN, CHARLES ROBERT  LP
MELOME [DUMAREST 28]  TUBB, E. C.  C/P,P
MELTDOWN [@TWENTY NINETY-NINE #5]  PEEL, JOHN  LP
MELUSINE  MONETTE, SARAH  H
MEMENTO MORI  LEWITT, SHARIANN  LP
MEMINI  PEARLMAN, DANIEL  H
MEMNOCH THE DEVIL [LESTAT 5]  RICE, ANNE  C/P,P,H
MEMOIRS FOUND IN A BATHTUB  LEM, STANISLAW  C/P,SFWA,P,H
MEMOIRS OF A SPACE TRAVELER  LEM, STANISLAW  H:2,P
MEMOIRS OF A SPACEWOMAN  MITCHISON, NAOMI  C/P:2,P
MEMOIRS OF A SURVIVOR, THE  LESSING, DORIS  P
MEMOIRS OF ALCHERINGIA, THE [ERTHRING 1]  DREW, WAYLAND  C/P:2,P
MEMOIRS OF AN INVISIBLE MAN  SAINT, H. F.  H,SFWA
MEMOIRS OF SCHLOCK HOMES, THE [BAGEL STREET DOZEN 2]  FISH, ROBERT L.  H
MEMORANDA [WELL-BUILT CITY 2]  FORD, JEFFREY  P
MEMORIES  MCQUAY, MIKE  C/P:2,P
MEMORIES AND VISIONS  STURGIS, SUSANNA J.  LPA
MEMORIES END [@WEB WARRIORS #1]  LUCENO, JAMES  P
MEMORIES OF ICE [MALAZAN 3]  ERIKSON, STEVEN  H,LP
MEMORIES OF MILO MORAI, THE [HORSECLANS 15]  ADAMS, ROBERT  C/P,P
MEMORIES OF THE FUTURE  KNOX, RICHARD A.  SR-H
MEMORIES OF THE SPACE AGE  BALLARD, J. G.  H
MEMORY [MILES VORKOSIGAN 7]  BUJOLD, LOIS MCMASTER  H:2,P
MEMORY  NAGATA, LINDA  H
MEMORY [@SCAVENGER #3]  PARKER, K. J.  LP
MEMORY AND DREAM [NEWFORD]  DE LINT, CHARLES  H,P
MEMORY BANK, THE  WEST, WALLACE  C/P,P,H
MEMORY BLANK  STITH, JOHN E.  C/P:2,P
MEMORY OF DRAGONS, A  JOHNSON, ANNABEL/JOHNSON, EDGAR  H
MEMORY OF EARTH, THE [HOMECOMING 1]  CARD, ORSON SCOTT  C/P,P,H
MEMORY OF FIRE, THE  FOY, GEORGE  P
MEMORY OF FIRE [WORLD GATES 1]  LISLE, HOLLY  C/P,P
MEMORY OF WHITENESS, THE  ROBINSON, KIM STANLEY  SFWA-P,H:2,P,SFWA

MEMORY PRIME [STAR TREK] REEVES-STEVENS, GARFIELD/ REEVES-STEVENS, JUDITH

C/P:3,P

MEMORY WIRE  WILSON, ROBERT CHARLES  C/P:2,P
MEMORYMAKERS  HERBERT, BRIAN/LANDIS, MARIE  P
MEMOS FROM PURGATORY  ELLISON, HARLAN  [C/RAN-P:2]
MEN AGAINST THE STARS  GREENBERG, MARTIN  PA,C/PA
MEN AND MACHINES  SILVERBERG, ROBERT  C/PA,PA
MEN AND THE MIRROR, THE  ROCKLYNNE, ROSS  C/P,P
MEN AND VOLTS AT WAR  MILLER, JOHN A.  [C/RAN-P]
MEN AT ARMS [DISCWORLD 15/DISCWORLD WATCHMEN 2]  PRATCHETT, TERRY  C/P,P,H
MEN FROM ARCTURUS, THE  WINTERBOTHAM, RUSS  H
MEN FROM ARIEL, THE  WOLLHEIM, DONALD A.  SR-H
MEN FROM P.I.G. AND R.O.B.O.T., THE  HARRISON, HARRY  H:2
MEN HUNTING THINGS [STARHUNTERS 1]  DRAKE, DAVID  SFWA-P,C/PA,PA
MEN IN BLACK [MEN IN BLACK]  PERRY, STEVE/SOLOMON, ED  C/P:2,P
MEN IN BLACK II [MEN IN BLACK]  FRIESNER, ESTHER M.  P
MEN IN THE JUNGLE, THE  SPINRAD, NORMAN  C/P:2,P
MEN INSIDE, THE  MALZBERG, BARRY N.  P
MEN INTO SPACE  LEINSTER, MURRAY  C/P:2,P
MEN LIKE GODS  WELLS, H. G.  H,P
MEN LIKE RATS  CHILSON, ROBERT  P
MEN, MARTIANS AND MACHINES  RUSSELL, ERIC FRANK  C/P,P,H
MEN OF WAR [@LOST REGIMENT #8]  FORSTCHEN, WILLIAM R.  C/P,P
MEN OF WAR [@THERE WILL BE WAR #2]  POURNELLE, JERRY  C/PA,PA
MEN ON THE MOON  WOLLHEIM, DONALD A.  DD277,PA,C/PA
MEN WHO DIE TWICE [MIND BROTHERS 3]  HEATH, PETER  P

TITLE  AUTHOR  CODE
MEN WHO SMILED NO MORE, THE [DOC SAVAGE #45]  ROBESON, KENNETH  C/P:2,@P
MEN WITHOUT BONES  KERSH, GERALD  C/P:2,P
MEN WRITING SCIENCE FICTION AS WOMEN  RESNICK, MIKE  PA
MENACE FROM EARTH, THE  HEINLEIN, ROBERT A.  C/P:2,P
MENACE FROM MERCURY  BADGER BOOKS  SR-PA
MENACE OF THE MUTANT MASTER [@PERRY RHODAN #18]  ACE  C/PA,PA
MENACE OF THE SAUCERS  BINDER, EANDO  C/P,P
MENACE UNDER MARSWOOD  LANIER, STERLING E.  C/P,P
MENACING SLEEP, THE  SHELDON, ROY  SR-P
MENAGERIE, THE [DOCTOR WHO MISSING ADVENTURES]  DAY, MARTIN  C/P:2,P
MENAGERIE, THE [STAR TREK]  RODDENBERRY, GENE  SR-S
MENDELOV CONSPIRACY, THE  CAIDIN, MARTIN  P
MENDOZA IN HOLLYWOOD [COMPANY 3]  BAKER, KAGE  P
MENNYMS IN THE WILDERNESS [MENNYMS]  WAUGH, SYLVIA  P
MENS IN HET UITDIJENDE HEELAL 1, DE  DE VRIES, TJOMME  FP
MENS IN HET UITDIJENDE HEELAL 2, DE  DE VRIES, TJOMME  FP
MENSCHHEIT AM SCHEIDEWEG [@PERRY RHODAN SILBERBANDE #80]  PABEL MOEWIG KG  FPR
MENTAL WIZARD, THE [DOC SAVAGE #53]  ROBESON, KENNETH  C/P,@P
MENTION MY NAME IN ATLANTIS  JAKES, JOHN  C/P:2,P
MENTOR  JENSEN, KRIS  P
MEPHISTO WALTZ, THE  STEWART, FRED MUSTARD  C/P:2,P
MEQ, THE  CASH, STEVE  LP
MERCEDES NIGHTS  WEAVER, MICHAEL D.  H,P
MERCENARIES, THE [HELMSMAN 4]  BALDWIN, BILL  C/P:2,P
MERCENARIES OF GOR [GOR 21]  NORMAN, JOHN  C/P,P
MERCENARY [@BIO OF A SPACE TYRANT #2]  ANTHONY, PIERS  C/P:2,P
MERCENARY, THE  POURNELLE, JERRY  C/P:2,P
MERCENARY FROM TOMORROW  REYNOLDS, MACK  DH065
MERCENARY’S STAR [@BATTLETECH 7]  KEITH, WILLIAM H. JR.  C/P:2,P
MERCHANT PRINCE, THE [MERCHANT PRINCE 1]  SHIMERMAN, ARMIN/SCOTT, MICHAEL  P
MERCHANTER’S LUCK [UNION/ALLIANCE]  CHERRYH, C. J.  C/P,P,H
MERCHANTS OF DISASTER [DOC SAVAGE #41]  ROBESON, KENNETH  C/P:2,@P
MERCHANTS OF SOULS, THE [THOUSAND CULTURES 3]  BARNES, JOHN  H,P
MERCHANTS’ WAR, THE [SPACE MERCHANTS 2]  POHL, FREDERIK  C/P,P,H
MERCURY  BOVA, BEN  H
MERCY MEN, THE  NOURSE, ALAN E.  C/P,P,H
MERCYCLE  ANTHONY, PIERS  C/P:2,SFWA-P:2,P
MERI, THE [MERI 1]  BOHNHOFF, MAYA KAATHRYN  C/P,SFWA-P,P
MERLIN  GREENBERG, MARTIN H.  PA
MERLIN [@PENDRAGON #2]  LAWHEAD, STEPHEN R.  P
MERLIN  NYE, ROBERT  C/P,P
MERLIN CHRONICLES, THE  ASHLEY, MIKE  PA
MERLIN CONSPIRACY, THE  JONES, DIANA WYNNE  H
MERLIN CONSPIRACY, THE (PROOF)  JONES, DIANA WYNNE  H,SR-LP
MERLIN EFFECT, THE  BARRON, T. A.  C/P,P
MERLIN OF ST. GILLES’ WELL, THE [JOAN OF ARC 1]  CHAMBERLIN, ANN  LP
MERLIN OF THE OAK WOOD, THE  CHAMBERLIN, ANN  P
MERLIN’S BONES  SABERHAGEN, FRED  C/P,P,H
MERLIN’S GIFT [MORDRED’S CURSE 2]  MCDOWELL, IAN  P
MERLIN’S GODSON  MUNN, H. WARNER  C/P:2,P
MERLIN’S HARP  CROMPTON, ANNE ELIOT  P
MERLIN’S KEEP  BRENT, MADELEINE  P
MERLIN’S MIRROR  NORTON, ANDRE  C/P:2,P
MERLIN’S MISTAKE  NEWMAN, ROBERT  LP
MERLIN’S RING  MUNN, H. WARNER  SR-P,C/P:2
MERMAID IN THE SWIMMING POOL, THE  WALLOP, DOUGLASS  H
MERMAIDS!  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
MERMAID’S SONG  VAN GORES, ALIDA  P
MERMAN’S CHILDREN, THE  ANDERSON, POUL  C/P,P,H,SFWA
MERRICK [MAYFAIR WITCHES/LESTAT]  RICE, ANNE  H:2
MERRO TREE, THE  WAITMAN, KATIE  P
MERRY ANIMAL TALES  BIGHAM, MADGE A.  H
MERRY MEN, THE/STRANGE CASE OF DR. JEKYLL AND MR. HYDE, THE  STEVENSON, ROBERT LOUIS  H
MESSAGE FROM EARTH, A  HASSLER, KENNETH W.  H
MESSAGE FROM THE EOCENE  SAINT CLAIR, MARGARET  DM105
MESSAGES FOUND IN AN OXYGEN BOTTLE  SHAW, BOB  SR-DNESFA1
MESSENGER, THE  DICKENS, MONICA  P
MESSENGER, THE [@DRAGONLANCE ICEWALL #1]  NILES, DOUGLAS  P
MESSENGER, THE  WILLIAMS, MONA  P
MESSENGER OF ZHUVASTOU  OFFUTT, ANDREW J.  C/P:2,P
MESSENGERS OF DARKNESS [MAGIRA 3]  WALKER, HUGH  P
MESSIAH  VIDAL, GORE  C/P:2,P
MESSIAH AT THE END OF TIME, A [DANCERS AT THE END OF TIME 5]  MOORCOCK, MICHAEL  C/P:2,P
MESSIAH CHOICE, THE  CHALKER, JACK L.  H:2,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 228
TITLE  AUTHOR  CODE
MESSIAH NODE [ARCHANGEL 3]  MOREHOUSE, LYDA  P
MESSIAH OF THE CYLINDER, THE  ROUSSEAU, VICTOR  H,LP
MESSIAH STONE, THE [MESSIAH STONE 1]  CAIDIN, MARTIN  C/P,P
METACONCERT, THE [@INTERVENTION #2]  MAY, JULIAN  C/P:2,P
METAL ANGEL  SPRINGER, NANCY  P,SFWA-P
METAL EATER, THE  SHELDON, ROY  SR-P
METAL FATIGUE  WILLIAMS, SEAN  H
METAL FIRE [@ROBOTECH #8]  MCKINNEY, JACK  C/P,P

METAL MAN AND OTHERS, THE [@COLLECTED STORIES JACK WILLIAMSON #1] WILLIAMSON, JACK

H

METAL MASTER, THE [DOC SAVAGE #72]  ROBESON, KENNETH  C/P:2,@P
METAL MONSTER, THE  MERRITT, A.  SR-P,C/P:2,SR-LP
METAL OF NIGHT [SECANTIS 2]  TIEDEMANN, MARK W.  LP
METAL SMILE, THE  KNIGHT, DAMON  PA,C/PA
METALLIC LOVE [SILVER METAL LOVER 2]  LEE, TANITH  P
METALLIC LOVE (PROOF) [SILVER METAL LOVER 2]  LEE, TANITH  SR-P
METALLIC MUSE, THE  BIGGLE, LLOYD JR.  H:2,SFWA,P
METALSTORM [@PHOENIX #4]  ALEXANDER, DAVID  P
METAMORPHOSES OF SCIENCE FICTION  SUVIN, DARKO  R/REF-H
METAMORPHOSIS [STAR TREK TNG]  LORRAH, JEAN  C/P:2,P
METAMORPHOSIS [@ROBOTECH #11]  MCKINNEY, JACK  C/P,P
METAMORPHOSIS [COCOON 2]  SAPERSTEIN, DAVID  P
METAPHASE [STARFARERS 3]  MCINTYRE, VONDA N.  C/P:2,P
METAPLANETARY [METAPLANETARY 1]  DANIEL, TONY  H
METEOR  NORTH, EDMUND H./COEN, FRANKLIN  C/P:2,P
METEOR MENACE [DOC SAVAGE #3]  ROBESON, KENNETH  C/P:2,@P
METHINKS THE LADY...  ENDORE, GUY  H
METHUSELAH ENZYME, THE  STEWART, FRED MUSTARD  C/P,P
METHUSELAH’S CHILDREN  HEINLEIN, ROBERT A.  C/P:2,P,H
METRO POUR L’INCONNU  LIMAT, MAURICE  FP
METROGNOME, THE  FOSTER, ALAN DEAN  C/P:2,P
METROPHAGE  KADREY, RICHARD  C/P,SFWA-P,P
METROPOLIS  VON HARBOU, THEA  FP,C/P:2,P
METROPOLIS (BRITISH)  VON HARBOU, THEA  SR-P
METROPOLITAN [AIAH 1]  WILLIAMS, WALTER JON  H,P
MEZENTIAN GATE, THE [ZIMIAMVIA 3]  EDDISON, E. R.  C/P:2,P
MICHAEL AND THE MAGIC MAN  SIDNEY, KATHLEEN M.  C/P,P
MICHAEL AND THE SECRET WAR  GOLDS, CASSANDRA  SFWA
MICHAEL: THE MASTER  LAMBERT, W. III  P
MICHAEL WHELAN’S YEAR OF WONDER (1989)  CALENDAR  CAL
MICHAELMAS  BUDRYS, ALGIS  H:2,P
MICKEY MOUSE ADVENTURE IN OUTER SPACE  DAVIE, GEORGE E.  BLB
MICRO INFINITY  MULLER, JOHN E.  SR-P
MICRO MEN, THE  STATTEN, VARGO  SR-P
MICRO-TECHS, THE [@PERRY RHODAN #55]  ACE  C/PA,PA
MICROBE DETECTIVE, LE  CLEMENT, HAL  FP
MICROBE HUNTERS  DE KRUIF, PAUL  [C/RAN-P]
MICROBOTS, THE [@TOM SWIFT (ARCHWAY) #8]  APPLETON, VICTOR  P
MICROCOLONY, THE [MICRONAUTS 2]  WILLIAMS, GORDON  P
MICROCOSMIC GOD  MOSKOWITZ, SAM  PA
MICROCOSMIC GOD [@COMPLETE STORIES OF THEODORE STURGEON 2]  STURGEON, THEODORE  LP

MICROCOSMIC TALES ASIMOV, ISAAC/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

HA

MICROCOSMS  BENFORD, GREGORY  PA
MICRONAUTS [TIME TRAVELER 3]  LYONS, STEVE  P
MICRONAUTS, THE [MICRONAUTS 1]  WILLIAMS, GORDON  P
MICROVERSE, THE  PREISS, BYRON  SR-VLHA
MICROWAVE CAPER, THE  TABLER, JOSEPH  P
MICROWORLDS  LEM, STANISLAW  R/REF-H
MID-FLINX [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P,P,H
MID-WAKE [@SURVIVALIST]  AHERN, JERRY  P
MIDAS MAN, THE [DOC SAVAGE #46]  ROBESON, KENNETH  C/P:2,@P
MIDAS WORLD  POHL, FREDERIK  C/P,SFWA-P,H:2,P
MIDDLE-EARTH ALBUM, A  WYATT, JOAN  A
MIDDLE KINGDOM, THE [@CHUNG KUO #1]  WINGROVE, DAVID  C/P,P,H
MIDDLE OF NOWHERE, THE [STAR WOLF 2]  GERROLD, DAVID  C/P,P
MIDDLE OF NOWHERE, THE [@DRAGONLANCE CROSSROADS]  THOMPSON, PAUL B.  P
MIDNIGHT  GRANT, CHARLES L.  PA
MIDNIGHT  KOONTZ, DEAN R.  H
MIDNIGHT  RUSSO, JOHN  P
MIDNIGHT AT THE WELL OF SOULS [WELL WORLD 1]  CHALKER, JACK L.  C/P:3,P
MIDNIGHT BLUE [SONJA BLUE 1,2,3]  COLLINS, NANCY A.  H
MIDNIGHT CAFE, THE [ANITA BLAKE 4,5,6]  HAMILTON, LAURELL K.  H
MIDNIGHT DANCERS, THE  CONWAY, GERARD F.  C/P:2,P

TITLE  AUTHOR  CODE
MIDNIGHT FALCON [RIGANTE 2]  GEMMELL, DAVID  P
MIDNIGHT FOLK, THE  MASEFIELD, JOHN  LP
MIDNIGHT HARVEST [SAINT-GERMAIN]  YARBRO, CHELSEA QUINN  H
MIDNIGHT JUSTICE [SPIDER-MAN]  DELRIO, MARTIN  C/P,P
MIDNIGHT MASS  WILSON, F. PAUL  H
MIDNIGHT MURDER [@AVENGER #24]  ROBESON, KENNETH  C/P:2,P
MIDNIGHT PEOPLE, THE  HAINING, PETER  PA
MIDNIGHT PLEASURES  BLOCH, ROBERT  H,SFWA
MIDNIGHT READER  STERN, PHILIP VAN DOREN  SR-PA
MIDNIGHT ROBBER  HOPKINSON, NALO  H,LP
MIDNIGHT SUN  WAGNER, KARL EDWARD  H
MIDNIGHT TIDES [MALAZAN 5]  ERIKSON, STEVEN  H
MIDNIGHT’S MASK [@FORGOTTEN REALMS EREVIS CALE #3]  KEMP, PAUL S.  P
MIDSHIPMAN’S HOPE [NICHOLAS SEAFORT 1]  FEINTUCH, DAVID  C/P:2,P
MIDSUMMER CENTURY  BLISH, JAMES  H:2,P,SFWA
MIDSUMMER TEMPEST, A  ANDERSON, POUL  C/P,SFWA-P,H,P,SFWA
MIDSUMMER’S NIGHTMARE, A  KILWORTH, GARRY  P
MIDWAY BETWEEN [DOUBLE SPIRAL WAR 1]  NORWOOD, WARREN  C/P:2,P
MIDWICH CUCKOOS, THE  WYNDHAM, JOHN  H:2,SR-H,P
MIDWORLD [HUMANX]  FOSTER, ALAN DEAN  H:2,P
MIGHTIEST MACHINE, THE  CAMPBELL, JOHN W. JR.  SR-H,C/P,P,H
MIGHTY ATOM, THE  CORELLI, MARIE  P
MIGHTY BARBARIANS, THE  SANTESSON, HANS STEFAN  PA
MIGHTY GOOD ROAD  SCOTT, MELISSA  C/P:2,SFWA-P,P
MIGHTY MANSLAYER, THE  LAMB, HAROLD  P
MIGHTY SWORDSMEN, THE  SANTESSON, HANS STEFAN  PA
MIGRATION [@SPECIES IMPERATIVE #2]  CZERNEDA, JULIE E.  H
MIKE MARS AND THE MYSTERY SATELLITE [@MIKE MARS #7]  WOLLHEIM, DONALD A.  P
MIKE MARS AROUND THE MOON [@MIKE MARS #8]  WOLLHEIM, DONALD A.  P
MIKE MARS ASTRONAUT [@MIKE MARS #1]  WOLLHEIM, DONALD A.  P
MIKE MARS AT CAPE KENNEDY [@MIKE MARS #3]  WOLLHEIM, DONALD A.  P
MIKE MARS FLIES THE DYNA-SOAR [@MIKE MARS #5]  WOLLHEIM, DONALD A.  P
MIKE MARS FLIES THE X-15 [@MIKE MARS #2]  WOLLHEIM, DONALD A.  P
MIKE MARS IN ORBIT [MIKE MARS #4]  WOLLHEIM, DONALD A.  H,@P
MIKE MARS, SOUTH POLE SPACEMAN [@MIKE MARS #6]  WOLLHEIM, DONALD A.  P
MILE BEYOND THE MOON, A  KORNBLUTH, C. M.  H:2,P
MILE LONG SPACESHIP, THE  WILHELM, KATE  C/P:2,P
MILICE DES MUTANTS, LA (FRENCH) [@PERRY RHODAN #3]  SCHEER, K. H.  FPR
MILITARY DIMENSION, THE  DRAKE, DAVID  C/P:2,P
MILITARY DIMENSION: MARK II, THE  DRAKE, DAVID  P
MILLENIUM  ANDERSON, JACK  LP
MILLENNIAL RITES [DOCTOR WHO MISSING ADVENTURES]  HINTON, CRAIG  C/P:2,P
MILLENNIAL WOMEN  KIDD, VIRGINIA  HA,PA
MILLENNIUM  ANDERSON, JACK  P
MILLENNIUM  BOVA, BEN  H:2,SFWA,P
MILLENNIUM  VARLEY, JOHN  H:2,LP
MILLENNIUM 3001  GREENBERG, MARTIN H./DAVIS, RUSSELL  PA
MILLENNIUM EVE  LOVE, KENNY  H
MILLENNIUM RISING  JENSEN, JANE  H
MILLENNIUM SHOCK [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN  P
MILLIJOY, THE DETERMINED [@PHENWICK #15]  KIMBROUGH, KATHERYN  P
MILLION CITIES, THE  MCINTOSH, J. T.  C/P:2,P
MILLION OPEN DOORS, A [THOUSAND CULTURES 1]  BARNES, JOHN  C/P,P,H
MILLION YEAR HUNT, THE  BULMER, KENNETH  DF285
MILWAUKEE THE BEAUTIFUL  JAMES, DAKOTA  H
MIMICS OF DEPHENE [@CAP KENNEDY #15]  KERN, GREGORY  C/P,P
MINA  KIRALY, MARIE  H:2
MIND ANGEL, THE  ELWOOD, ROGER  SR-HA
MIND AT BAY  EVANS, CHRISTOPHER  SR-PA
MIND BEHIND THE EYE, THE  GREEN, JOSEPH  C/P:2,P
MIND BENDERS, THE  KENNAWAY, JAMES  C/P,P
MIND-BREAKS OF SPACE [@SPACE: 1999 YEAR 2 #2]  BUTTERWORTH, MICHAEL  C/P,P
MIND BROTHERS, THE [MIND BROTHERS 1]  HEATH, PETER  C/P:2,P
MIND CAGE, THE  VAN VOGT, A. E.  H:2,P
MIND-CALL  BELDEN, WILANNE SCHNEIDER  H
MIND CALL  SHAPIRO, NEIL  P
MIND CHANGER [SECTOR GENERAL 11]  WHITE, JAMES  H
MIND FIELDS  YERKA, JACK  A
MIND FOR TRADE, A [SOLAR QUEEN 6]  NORTON, ANDRE/SMITH, SHERWOOD  P
MIND FORCE  BRETT, LEO  H
MIND FORCE WARRIOR [@PSI-MAN #1]  DAVID, PETER  P
MIND GAME [GHOSTWALKERS 2]  FEEHAN, CHRISTINE  P
MIND GAME, THE  SPINRAD, NORMAN  P
MIND GAMES [@TOM SWIFT (ARCHWAY) #10]  APPLETON, VICTOR  P

TITLE  AUTHOR  CODE
MIND GUEST [DIANA SANTEE 1]  GREEN, SHARON  P
MIND-HOLD  BELDEN, WILANNE, SCHNEIDER  H:2
MIND IN CHAINS  EVANS, CHRISTOPHER  SR-PA
MIND IN MOTION  WARRICK, PATRICIA S.  R/REF-H
MIND MASTER, THE  GUNN, JAMES E.  P
MIND MASTER, THE  HURWOOD, BERNHARDT J.  P
MIND MASTERS, THE [@MIND MASTERS #1]  ROSSMANN, JOHN F.  P
MIND MELD [STAR TREK]  VORNHOLT, JOHN  C/P,P
MIND MONSTERS, THE  CORY, HOWARD L.  DG602
MIND NET, THE  FRANKE, HERBERT W.  P
MIND OF EVIL, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,P
MIND OF MR. SOAMES, THE  MAINE, CHARLES ERIC  SR-P
MIND OF MY MIND  BUTLER, OCTAVIA E.  C/P,SFWA,P,H
MIND OF THE MAGIC [FAIA RISSEDOTTE 3]  LISLE, HOLLY  C/P:2,P
MIND ONE  DOLINSKY, MIKE  P
MIND OUT OF TIME  TONKS, ANGELA  H
MIND PARASITES, THE  WILSON, COLIN  H:2,LP
MIND PARTNER  GOLD, HORACE L.  PA,C/PA
MIND POOL, THE  SHEFFIELD, CHARLES  C/P:2,SFWA-P,P
MIND-RIDERS, THE  STABLEFORD, BRIAN M.  P
MIND SLAVER [STAR COMMANDOS 5]  GRIFFIN, P. M.  P
MIND SNARE  GREENO, GAYLE  C/P,P
MIND-SPEAKERS’ CALL [@GHATTI’S TALE #2]  GREENO, GAYLE  C/P,SFWA-P,P
MIND SPIDER AND OTHER STORIES, THE  LEIBER, FRITZ  C/P:2,DD491,P
MIND-SURFER  LEON, MARK  P
MIND SWITCH  KNIGHT, DAMON  C/P:2,P
MIND THING, THE  BROWN, FREDRIC  C/P:2,P
MIND TO MIND  SILVERBERG, ROBERT  HA,PA
MIND TRADERS, THE  HOLLY, JOAN HUNTER  C/P,P,H
MIND TRANSFER  ASIMOV, JANET  C/P,SFWA-P,P,H
MIND TRAP  MORGAN, DAN  C/P:2,P
MIND TWISTER AFFAIR, THE [@MAN FROM UNCLE #12]  STRATTON, THOMAS  C/P:2,P
MIND WAR  SNYDER, GENE  P
MIND WIZARDS OF CALLISTO [JANDAR 5]  CARTER, LIN  C/P,P
MINDBEND  COOK, ROBIN  P
MINDBENDER  COHEN, JAMES/RUBIE, PETER  C/P,P
MINDBLAST [@SPACE COPS 2]  DUANE, DIANE/MORWOOD, PETER  C/P,P
MINDBLOCKED MAN, THE  SUTTON, JEFF  C/P:2,P
MINDBOND [SEA KING 2]  SPRINGER, NANCY  P
MINDBRIDGE  HALDEMAN, JOE  H:2,P
MINDFLIGHT  GOLDIN, STEPHEN  C/P:2,P
MINDFOGGER  ROGERS, MICHAEL  C/P,P,H
MINDHOPPER  JOHNSON, JAMES B.  C/P,P
MINDKILLER [MINDKILLER 1]  ROBINSON, SPIDER  C/P,P,H
MINDLIGHT  DAVIS, MARGARET  C/P,P
MINDMIX  KELLEY, LEO P.  C/P,P
MINDPLAYERS  CADIGAN, PAT  C/P:2,P
MINDPLAYERS (PROOF)  CADIGAN, PAT  SR-LP
MINDS APART  DAVIS, MARGARET  P
MINDS, MACHINES AND EVOLUTION  HOGAN, JAMES P.  C/P:2,P
MINDS UNLEASHED  CONKLIN, GROFF  C/PA
MINDSCAN  SAWYER, ROBERT J.  H
MINDSCAPE  HAIRSTON, ANDREA  LP
MINDSHADOW [STAR TREK]  DILLARD, J. M.  C/P:2,P
MINDSHIP  CONWAY, GERARD F.  C/P:2,P
MINDSLAYER [@STAR PRECINCT #2]  RANDLE, KEVIN/DRUSCOLL, RICHARD  C/P,P
MINDSONG  COX, JOAN  C/P:2,P
MINDSPAN  DICKSON, GORDON R.  C/P:2,P
MINDSTAR RISING [GREG MANDEL 1]  HAMILTON, PETER F.  C/P:2,P
MINDSWAP  SHECKLEY, ROBERT  H:2,P
MINDSWORD’S STORY [@LOST SWORDS #6]  SABERHAGEN, FRED  C/P,SFWA,P,H
MINDWARP [DOCTOR WHO]  MARTIN, PHILIP  P
MINDWARPERS, THE  RUSSELL, ERIC FRANK  C/P:2,P
MINDWIPE!  HAHN, STEVE  SR-P
MINDWORLDS [FLESH AND GOLD 3]  GOTLIEB, PHYLLIS  H
MINDWORM, THE  TANDEM BOOKS  SR-PA
MINERS IN THE SKY  LEINSTER, MURRAY  C/P:2,P
MINERVA WAKES  LISLE, HOLLY  C/P:2,SFWA-P,P
MINERVAN EXPERIMENT, THE [GIANTS 1,2,3]  HOGAN, JAMES P.  H:2
MINES OF BEHEMOTH, THE [NIFFT 2]  SHEA, MICHAEL  P
MINES OF MORIA  MATHEWS, SUSAN/RUEMMLER, J. D.  P
MINESWEEPER  LOVISI, GARY  GRD8
MINIKINS OF YAM, THE  SWANN, THOMAS BURNETT  C/P:2,P
MINING THE OORT  POHL, FREDERIK  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 231
TITLE  AUTHOR  CODE
MINION [VAMPIRE HUNTRESS 1]  BANKS, L. A.  LP
MINION (SPECIAL HUNTRESS EDITION) [VAMPIRE HUNTRESS]  BANKS, L. A.  P
MINIONS OF THE MOON  BOWES, RICHARD  H
MINISTERO DELLA FELICITA, IL  RAMBELLI, ROBERTA  FP
MINOR ARCANA  JONES, DIANA WYNNE  P
MINORITY REPORT, THE [COLLECTED STORIES OF PHILIP K. DICK 4]  DICK, PHILIP K.  LP
MINOS OF SARDANES  STILSON, CHARLES B.  H
MINOTAUR MAZE  SHECKLEY, ROBERT  SR-H
MINUTEGIRLS, THE  PHILLIES, GEORGE  LP
MIOCENE ARROW, THE  MCMULLEN, SEAN  LP
MIR [VIRTUAL REALITY 2]  BESHER, ALEXANDER  H
MIRABILE  KAGAN, JANET  H
MIRACLE AND OTHER CHRISTMAS STORIES  WILLIS, CONNIE  C/P,P
MIRACLE AT SAN TANCO [@FLYING NUN #1]  JOHNSTON, WILLIAM  P
MIRACLE OF RARE DESIGN, A  RESNICK, MIKE  H:2,P
MIRACLE OF THE AGES: THE GREAT PYRAMID OF GIZEH  SMITH, WORTH  [C/RAN-P]
MIRACLE ON 34TH STREET  DAVIES, VALENTINE  P
MIRACLE SEASON, THE  CLINE, LINDA  P
MIRACLE VISITORS  WATSON, IAN  C/P,P

MIRACLE WORKERS [@STAR TREK S.C.E. #2] DECANDIDO, KEITH R. A./DILMORE, KEVIN/

MACK, DAVID/WARD, DAYTON

P

MIRACLE WORKERS, THE  STEARN, JESS  [C/RAN-P]

MIRACLE WORKERS, THE [@VANCE INTEGRAL EDITION #9]/ DRAGON MASTERS, THE [@VANCE INTEGRAL EDITION #9]/ LAST CASTLE, THE [@VANCE INTEGRAL EDITION #9] VANCE, JACK

SR-H

MIRAGE  COOPER, LOUISE  C/P:2,SFWA-P,P
MIRAGE [@ASIMOV’S ROBOT MYSTERY]  TIEDEMANN, MARK W.  LP
MIRKHEIM  ANDERSON, POUL  H:2,P
MIROIRS DE SANG  KOONTZ, DEAN R.  FP
MIRROR  MASTERTON, GRAHAM  SFWA-P
MIRROR, THE  MILLHISER, MARLYS  H
MIRROR DANCE [MILES VORKOSIGAN 5]  BUJOLD, LOIS MCMASTER  H,SFWA,P
MIRROR FOR OBSERVERS, A  PANGBORN, EDGAR  H:2,LP,SFWA
MIRROR FOR PRINCES, A  DE HAAN, TOM  H
MIRROR FRIEND, MIRROR FOE  ASPRIN, ROBERT LYNN/TAKEI, GEORGE  C/P:2,P
MIRROR FRIEND, MIRROR FOE  TAKEI, GEORGE/ASPRIN, ROBERT LYNN  SR-P,C/P
MIRROR IMAGE  CONEY, MICHAEL G.  C/P,P
MIRROR IMAGE  DUNCAN, BRUCE  SR-DB0081
MIRROR IMAGE [@CHARMED]  MARIOTTE, JEFF  P
MIRROR IN THE SKY  GARNETT, DAV  P
MIRROR MAZE, THE  HOGAN, JAMES P.  C/P:2,P
MIRROR, MIRROR [STAR TREK]  BIXBY, JEROME  SR-S
MIRROR OF DESTINY  NORTON, ANDRE  P
MIRROR OF DREAMS  GANPAT  H
MIRROR OF HELEN, THE [KAPHTU 3]  PURTILL, RICHARD  P
MIRROR OF HER DREAMS, THE [MORDANT’S NEED 1]  DONALDSON, STEPHEN R.  H:2,P,SFWA
MIRROR OF INFINITY, THE  SILVERBERG, ROBERT  PA
MIRROR OF MERLIN, THE [LOST YEARS OF MERLIN 4]  BARRON, T. A.  P
MIRROR TO THE SKY  GESTON, MARK S.  H:2,SFWA-P
MIRROR WORLD  WILLIAMS, TAD  H
MIRROR’S EDGE [@QUANTUM LEAP]  DAVIS, CAROL/REESE, ESTHER D.  P
MIRRORSHADES  STERLING, BRUCE  PA,SFWA
MIRTH OF YORE  KING, JAMES B.  SCX
MISADVENTURES & MERFOLK  RENO, KELLY  LP
MISBEGOTTEN KING, THE  BUSH, ANNE KELLEHER  C/P,P

MISCAST BARBARIAN: A BIOGRAPHY OF ROBERT E. HOWARD, THE (1906–1936)

DE CAMP, L. SPRAGUE

R/REF-H

MISCAST GENTLEMAN, THE  EASTON, EDWARD  SR-P
MISENCHANTED SWORD, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
MISERY  KING, STEPHEN  C/P,P,H
MISHMASH AND THE ROBOT  CONE, MOLLY  P
MISLAID CHARM, THE  PHILLIPS, ALEXANDER M.  SR-H
MISLAID MAGICIAN, THE [CECILIA & KATHERINE 3]  WREDE, PATRICIA C./STEVERMER, CAROLINE H:2
MISPLACED LEGION, THE [VIDESSOS 1]  TURTLEDOVE, HARRY  C/P:2,P
MISPLACED PERSONS  HARDING, LEE  P
MISS FINNEY KILLS NOW AND THEN  DEMPSEY, AL  SFWA-P
MISS OSBORNE-THE-MOP  GAGE, WILSON  P
MISS PICKERELL AND THE WEATHER SATELLITE [@MISS PICKERELL #2]  MACGREGOR, ELLEN/PANTELL, DORA  P
MISS PICKERELL GOES TO MARS [@MISS PICKERELL]  MACGREGOR, ELLEN  C/P:2,LP
MISS PICKERELL MEETS MR. H.U.M. [@MISS PICKERELL #3]  MACGREGOR, ELLEN/PANTELL, DORA  P
MISS PICKERELL ON THE MOON [@MISS PICKERELL #1]  MACGREGOR, ELLEN/PANTELL, DORA  H:2,P
MISS PICKERELL TAKES THE BULL BY THE HORNS [@MISS PICKERELL #4]  MACGREGOR, ELLEN/PANTELL, DORA  P
MISS PIGGY COVER GIRL FANTASY CALENDAR (1981)  CALENDAR  CAL
MISSED IT BY THAT MUCH! [GET SMART 5]  JOHNSTON, WILLIAM  P

The MITSFS Titledex  September 24, 2007  Page 232
TITLE  AUTHOR  CODE
MISSILE GAP  STROSS, CHARLES  H
MISSILE LORDS, THE  SUTTON, JEFF  P
MISSILES, MOON PROBES, AND MEGAPARSECS  LEY, WILLY  [C/RAN-P]
MISSING IN ACTION [STAR TREK NEW FRONTIER 16]  DAVID, PETER  H
MISSING MAN  MACLEAN, KATHERINE  H:2,P
MISSING MATTER, THE [NEXT WAVE 3]  MCDONOUGH, THOMAS R.  C/P,P
MISSING MEN OF SATURN  LATHAM, PHILIP  H:2
MISSING PERSONS LEAGUE, THE  BONHAM, FRANK  P
MISSING PIECES [DOCTOR WHO]  PHIPPEN, MARK/LYON, SHAUN  VLPA
MISSING WORLD, THE  ELWOOD, ROGER  SR-HA
MISSION  TILLEY, PATRICK  LP
MISSION ACCOMPLISHED  WALKER, JERRY  H
MISSION CHILD  MCHUGH, MAUREEN F.  H:2,P
MISSION FROM MARS  CONROY, RICK  SR-P
MISSION IMPOSSIBLE  TIGER, JOHN  P
MISSION: IMPRACTICAL [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P,P
MISSION: INTERPLANETARY  VAN VOGT, A. E.  C/P:2,P
MISSION: MANSTOP  NEVILLE, KRIS  SR-P
MISSION OF GRAVITY  CLEMENT, HAL  C/P,P,H
MISSION OF MAGIC [CAITHAN CRUSADE 2]  SMITH, JULIE DEAN  P
MISSION TO HORATIUS [STAR TREK]  REYNOLDS, MACK  H:2
MISSION TO MAGNUS [DOCTOR WHO]  MARTIN, PHILIP  P
MISSION TO MINERVA [GIANTS 5]  HOGAN, JAMES P.  H
MISSION TO MOULOKIN [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  H:2,P
MISSION TO MOULOKIN (PROOF) [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  SR-LP
MISSION TO THE HEART STARS  BLISH, JAMES  C/P:2,P
MISSION TO THE STARS  VAN VOGT, A. E.  C/P:2,P
MISSION TO UNIVERSE  DICKSON, GORDON R.  C/P:2,SFWA-P,P
MISSION TO UNIVERSE (REVISED)  DICKSON, GORDON R.  C/P:2,P
MISSION TO VENUS [DOCTOR WHO]  EMMS, WILLIAM  [C/RAN-P]
MISSION: TORI  BOLTON, JOHANNA M.  P
MISSION UNDERGROUND [STAR COMMANDOS 3]  GRIFFIN, P. M.  C/P:2,P
MISSIONARIES, THE  COMPTON, D. G.  C/P:2,P
MISSISSIPPI BLUES [NANOTECH VERITY 2]  GOONAN, KATHLEEN ANN  LP
MISSPENT YOUTH  HAMILTON, PETER F.  H

MIST, THE [STAR TREK CAPTAIN’S TABLE 3/STAR TREK DS9] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P:2,P

MIST AND STONE [@CHARMED]  GALLAGHER, DIANA G.  P
MIST LIZARD, THE  ALEXANDER, MARC  P
MISTBORN [MISTBORN 1]  SANDERSON, BRANDON  H
MISTED CLIFFS, THE [LOST CONTINENT 1]  ASARO, CATHERINE  LP
MR. ADAM  FRANK, PAT  H:2,P
MR. ANTIPHILOS, SATYR  DE GOURMONT, REMY  H
MR. BASS’S PLANETOID  CAMERON, ELEANOR  H
MR. CHANG’S CRIME RAY  APPLE, A. E.  H
MISTER DA V.  REED, KIT  C/P:2,P
MR. FAIRLIE’S FINAL JOURNEY  DERLETH, AUGUST  SR-H
MR. GEORGE AND OTHER ODD PERSONS  GRENDON, STEPHEN  SR-H
MR. GODLY BESIDE HIMSELF  BULLETT, GERALD  H
MISTER JUSTICE  PISERCHIA, DORIS  H,D53415
MR. LIMPET  PRATT, THEODORE  H
MISTER MERLIN 1  ROTSLER, WILLIAM  C/P:2,LP
MISTER MERLIN 2  ROTSLER, WILLIAM  C/P:2,LP
MR. MIRAKEL  OPPENHEIM, E. PHILLIPS  H
MISTER MONDAY [@KEYS TO THE KINGDOM #1]  NIX, GARTH  LP
MR. MOON  KNOBEL, PHILIP  P
MR. MUNCHAUSEN  BANGS, JOHN KENDRICK  SR-H
MR. NOAH AND THE SECOND FLOOD  BURNFORD, SHEILA  P
MR. O’MALLEY AND THE HAUNTED HOUSE [@BARNABY #2]  JOHNSON, CROCKETT  SCX
MR. O’MALLEY GOES FOR THE GOLD [@BARNABY #4]  JOHNSON, CROCKETT  SCX
MR. O’MALLEY, WIZARD OF WALL STREET [@BARNABY #5]  JOHNSON, CROCKETT  SCX
MR. PYE  PEAKE, MERVYN  SR-P
MR. TOMORROW  SELLERS, CON  P
MR. TOMPKINS IN PAPERBACK  GAMOW, GEORGE  LP
MISTER TOUCH  BOSSE, MALCOLM  H
MR. WILMER  LAWSON, ROBERT  H
MISTRAL’S KISS [MEREDITH GENTRY 5]  HAMILTON, LAURELL K.  H
MISTRESS MASHAM’S REPOSE  WHITE, T. H.  C/P,P,H
MISTRESS OF AMBIGUITIES [SILVERGLASS 4]  RIVKIN, J. F.  P
MISTRESS OF DEATH  ANTHONY, PIERS/FUENTES, ROBERTO  P
MISTRESS OF DRAGONS [DRAGONVARLD 1]  WEIS, MARGARET  H
MISTRESS OF MISTRESSES [ZIMIAMVIA 1]  EDDISON, E. R.  SR-H,C/P:2,P
MISTRESS OF THE CATACOMBS [LORD OF THE ISLES 4]  DRAKE, DAVID  C/P:2,P,H
MISTRESS OF THE EMPIRE [MARA 3]  FEIST, RAYMOND E./WURTS, JANNY  H:2,P

The MITSFS Titledex  September 24, 2007  Page 233
TITLE  AUTHOR  CODE
MISTRESS OF THE JEWELS, THE [WESTRIA 1,2]  PAXSON, DIANA L.  C/P,P
MISTRESS OF THE NIGHT [@FORGOTTEN REALMS PRIESTS]  BASSINGTHWAITE, DON/GROSS, DAVE  P
MISTRESS OF WINTER [OHNDARIEN 2]  CARWYN, GILES/FAHNESTOCK, TODD  H

MISTS FROM BEYOND, THE WEINBERG, ROBERT/

DZIEMIANOWICZ, STEFAN R./ GREENBERG, MARTIN H.

SFWA,SFWA-P

MISTS OF AVALON, THE [MISTS OF AVALON 1]  BRADLEY, MARION ZIMMER  H:2,LP
MISTS OF DOOM, THE [CORMAC 1]  OFFUTT, ANDREW J.  C/P,P
MISTS OF EVERNESS [WAR OF THE DREAMING 2]  WRIGHT, JOHN C.  H
MISTS OF THE AGES  GREEN, SHARON  P
MISTS OF TIME, THE  ANDERSON, MARGARET J.  H
MISTWALKER  HEALD, DENISE LOPES  P
MISTWORLD  GREEN, SIMON R.  C/P:2,P
MITHRANDIR  MILLER, STEPHEN O.  R/REF-P
MIXED DOUBLES  DA CRUZ, DANIEL  C/P,P
MIXED FEELINGS  EFFINGER, GEORGE ALEC  H:2
MIXED MAGICS [CHRESTOMANCI]  JONES, DIANA WYNNE  LP
MIXED MEN, THE  VAN VOGT, A. E.  H:2
M’LADY WITCH [WARLOCK’S HEIRS 2]  STASHEFF, CHRISTOPHER  C/P,P
MNOGOYARUSNEE MIR  FARMER, PHILIP JOSE  FH

MOB MAGIC THOMSEN, BRIAN M./

GREENBERG, MARTIN H.

PA

MOBY DICK (ABRIDGED)  MELVILLE, HERMAN  P
MOCKING PROGRAM, THE  FOSTER, ALAN DEAN  C/P,P,H
MOCKINGBIRD  TEVIS, WALTER  H,P
MOCKYMEN  WATSON, IAN  H
MODEL, THE  AICKMAN, ROBERT  H
MODERAN  BUNCH, DAVID R.  C/P:2,P
MODERN CLASSIC SHORT NOVELS OF SCIENCE FICTION  DOZOIS, GARDNER  HA
MODERN CLASSICS OF FANTASY  DOZOIS, GARDNER  LPA
MODERN CLASSICS OF SCIENCE FICTION  DOZOIS, GARDNER  HA
MODERN FANTASY  PRINGLE, DAVID  R/REF-P
MODERN GREATS OF SCIENCE FICTION  STRAHAN, JONATHAN  LPA
MODERN MASTERPIECES OF SCIENCE FICTION  MOSKOWITZ, SAM  H,LPA
MODERN MASTERS OF HORROR  COFFEY, FRANK  HA
MODERN SCIENCE FICTION  SPINRAD, NORMAN  C/PA,PA
MODERN SCIENCE FICTION: ITS MEANING AND ITS FUTURE  BRETNOR, REGINALD  R/REF-H
MODERN UTOPIA, A  WELLS, H. G.  LP
MODERNE NATUURWETENSCHAPPEN 1, DE  ASIMOV, ISAAC  FP
MODERNE NATUURWETENSCHAPPEN 2, DE  ASIMOV, ISAAC  FP
MODERNE THEORIEEN OVER HET HEELAL  COLEMAN, JAMES A.  FP
MODULAR MAN, THE  ALLEN, ROGER MACBRIDE  C/P,P,H
MOJAVE WELLS  JAMES, L. DEAN  P
MOJO AND THE PICKLE JAR  BELL, DOUGLAS  C/P,P
MOJO: CONJURE STORIES (PROOF)  HOPKINSON, NALO  SR-LPA
MOLDIES & MEATBOPS [WARE 1,2,3]  RUCKER, RUDY  H
MOLECULAR CAFE, THE  MIR PUBLISHERS  SR-PA
MOLECULE MEN, THE  HOYLE, FRED/HOYLE, GEOFFREY  H:2
MOLESWORTH’S GUIDE TO THE ATOMMIC AGE  WILLANS, GEOFFREY/SEARLE, RONALD  [C/RAN-H]
MOLLY DEAR  FINE, STEPHEN  H
MOLLY ZERO  ROBERTS, KEITH  SR-P
MOLT BROTHER [KREN 1]  LICHTENBERG, JACQUELINE  C/P,P
MOMENT OF ECLIPSE  ALDISS, BRIAN W.  H,SFWA
MOMENT OF NO RETURN  KING-HALL, STEPHEN  P
MOMENT OF THE MAGICIAN, THE [SPELLSINGER 4]  FOSTER, ALAN DEAN  C/P,P,H
MOMENT OUT OF TIME  SHELDON, ROY  SR-P
MOMO  ENDE, MICHAEL  SFWA,LP
MONA LISA AWAKENING [MONERE]  SUNNY  LP
MONA LISA OVERDRIVE [NEUROMANCER 3]  GIBSON, WILLIAM  C/P,SFWA-P,H:2,P
MONAD 2  KNIGHT, DAMON  R/REF-P
MONADIC UNIVERSE, THE  ZEBROWSKI, GEORGE  C/P,P
MONARCHS OF SOL, THE [QUEENDOM OF SOL 1,2]  MCCARTHY, WIL  H

MONDADORI’S POCKET ITALIAN-ENGLISH ENGLISH-ITALIAN DICTIONARY TEDESCHI, ALBERTO/

FANTONETTI, CARLO ROSSI

DICT

MONDAY BEGINS ON SATURDAY  STRUGATSKY, ARKADY/STRUGATSKY, BORIS P
MONDO DI GRIMM, IL (ITALIAN)  VINGE, VERNOR  FP
MONEY PLAYS  BECKNER, MORTON  H
MONGOL, THE [@CASCA #22]  SADLER, BARRY  P
MONITOR FOUND IN ORBIT  CONEY, MICHAEL G.  P
MONITOR, THE MINERS AND THE SHREE, THE  KILLOUGH, LEE  C/P:2,P
MONITORS, THE  LAUMER, KEITH  C/P:2,P
MONKEY SHINES  STEWART, MICHAEL  H,P
MONKEY SONATAS  CARD, ORSON SCOTT  P
MONKEY STATION  MAYHAR, ARDATH/FORTIER, RON  P

The MITSFS Titledex  September 24, 2007  Page 234
TITLE  AUTHOR  CODE
MONKEY VISITS A BLACK HOLE, THE (CHINESE)  XI TSUN TSAI  FP
MONK’S MAGIC  DE CONEAU, ALEXANDER  H
MONODYNE CATASTROPHE, THE  RENARD, JOSEPH  P
MONOLITH [@ANGEL]  PASSARELLA, JOHN  C/P,P
MONSTER, THE  BOND, NELSON  B
MONSTER  CROWCROFT, PETER  P
MONSTER  KONVITZ, JEFFREY  P

MONSTER BRIGADE 3000 GREENBERG, MARTIN H./ WAUGH, CHARLES G.

PA

MONSTER EXPRESS  MOOSER, STEPHEN  P
MONSTER FLY, THE  PIDDOCK, CHARLES  P
MONSTER FROM EARTH’S END, THE  LEINSTER, MURRAY  C/P:2,P
MONSTER FROM OUT OF TIME  LONG, FRANK BELKNAP  P

MONSTER ISLAND [@BUFFY/ANGEL] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

MONSTER MACHINE [@TOM SWIFT (ARCHWAY) #5]  APPLETON, VICTOR  P
MONSTER MAKERS, INC.  YEP, LAURENCE  P,SFWA
MONSTER MEN, THE  BURROUGHS, EDGAR RICE  C/P:2,P
MONSTER MIX  ARTHUR, ROBERT  PA
MONSTER OF METELAZE [@CAP KENNEDY #3]  KERN, GREGORY  P
MONSTER OF MINNESOTA, THE [@NEWS FROM THE EDGE 1]  SUMNER, MARK  P
MONSTER OF THE MAZE [@BLADE #6]  LORD, JEFFREY  C/P,P
MONSTER OF THE PROPHECY, THE  SMITH, CLARK ASHTON  P
MONSTER TALES  ELWOOD, ROGER  SR-VLHA
MONSTER WHEEL AFFAIR, THE [@MAN FROM UNCLE #8]  MCDANIEL, DAVID  C/P:2,P

MONSTERS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

MONSTERS, THE [DOC SAVAGE #7]  ROBESON, KENNETH  C/P:2,@P
MONSTERS  VAN VOGT, A. E.  C/P:2,P
MONSTERS AMONG US  LEE, JOHN/MOORE, B.  [C/RAN-P]
MONSTERS AND ALIENS FROM GEORGE LUCAS  CARRAU, BOB  H
MONSTERS AND MAGICIANS [STAIRWAY TO FOREVER 2]  ADAMS, ROBERT  C/P:2,P
MONSTERS AND MEDICS  WHITE, JAMES  C/P:2,P
MONSTERS AND SUCH  LEINSTER, MURRAY  C/P:2,P
MONSTERS AND THE CRITICS AND OTHER ESSAYS  TOLKIEN, J. R. R.  R/REF-H
MONSTERS GALORE  HURWOOD, BERNHARDT J.  PA
MONSTERS IN ORBIT  VANCE, JACK  DM125
MONSTERS INSIDE, THE [DOCTOR WHO NINTH DOCTOR]  COLE, STEPHEN  H
MONSTERS OF MAGIC, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  PA
MONSTERS OF STAR TREK, THE  COHEN, DANIEL  R/REF-P
MONSTER’S RING, THE [MAGIC SHOP 3]  COVILLE, BRUCE  C/P,LP
MONSTROCITY  THOMAS, JEFFREY  LP
MONSTROUS REGIMENT [DISCWORLD 30]  PRATCHETT, TERRY  H:2,P
MONT CANT GOLD  FISHER, PAUL R.  C/P,P
MONTE CRISTO #99  JAKES, JOHN  P
MONTEZUMA STRIP  FOSTER, ALAN DEAN  C/P,P
MONTEZUMA’S DAUGHTER  HAGGARD, H. RIDER  H
MONTEZUMA’S PEARL  JONES, DAVID LEE  P
MONTH FOR MANKIND, A  WILLIAMS, T. OWEN  H
MONTHLY TERRORS  PARNELL, FRANK H./ASHLEY, MIKE  R/IND
MONTY PYTHON AND THE HOLY GRAIL BOOK  MONTY PYTHON  VLP
MONUMENT  BIGGLE, LLOYD JR.  H:2,P,SFWA
MONUMENT  GRAHAM, IAN  C/P,P,H
MOON AHEAD  GREENER, LESLIE  H
MOON AHEAD (CONDENSED)  GREENER, LESLIE/HUTCHINSON, JOHN  SR-P
MOON AND THE FACE, THE  MCKILLIP, PATRICIA A.  C/P,P,H
MOON AND THE SUN, THE  MCINTYRE, VONDA N.  H,P
MOON AND THE THORN, THE [CELYDONN GWENLLIANT 3]  EDGERTON, TERESA  C/P,P
MOON BASE  NEPHEW, WILLIAM/CHESTER, MICHAEL  [C/RAN-P]
MOON BASE  TUBB, E. C.  C/P:2,P
MOON CALLED [MERCY THOMPSON 1]  BRIGGS, PATRICIA  P
MOON CALLED  NORTON, ANDRE  C/P,LP,P
MOON CHILDREN, THE  WILLIAMSON, JACK  C/P,P,H
MOON COLONY, THE  BELL, WILLIAM DIXON  H
MOON CONQUERORS, THE  ROMANS, R. H.  SR-P
MOON DOGS  SWANWICK, MICHAEL  SR-H,H
MOON DREAMS  STRICKLAND, BRAD  C/P:2,P
MOON ERA, THE  MOSKOWITZ, SAM  PA
MOON EXPRESS [AVENGERS #9]  DANIELS, NORMAN  C/P,P
MOON FACE  LONDON, JACK  H:2
MOON FATE [@DEATHLANDS]  AXLER, JAMES  P
MOON-FLASH  MCKILLIP, PATRICIA A.  C/P,P,H
MOON GODDESS AND THE SON, THE  KINGSBURY, DONALD  H:2,P
MOON IN HIDING, THE [CELYDONN GREEN LION 2]  EDGERTON, TERESA  C/P:2,SFWA-P,P

TITLE  AUTHOR  CODE
MOON IN SCIENCE & FICTION, THE  ALICKE, WALTER  R/REF-H
MOON IN THE GROUND  ANTILL, KEITH  H
MOON IS A HARSH MISTRESS, THE  HEINLEIN, ROBERT A.  C/P:2,P
MOON IS HELL!, THE (INCLUDES THE ELDER GODS)  CAMPBELL, JOHN W. JR.  SR-H,C/P,SR-LP,H
MOON IS RED, THE  ROHMER, SAX  H
MOON LAMP, THE  SMITH, MARK  P
MOON MAID, THE  BURROUGHS, EDGAR RICE  C/P:2,P
MOON MAID, THE  ROBERTSON, R. GARCIA  H
MOON MEN, THE  BURROUGHS, EDGAR RICE  C/P:2,P
MOON METAL, THE  SERVISS, GARRETT P.  H
MOON MIRROR  NORTON, ANDRE  H:2,SFWA-P,P
MOON MISSING  SOREL, EDWARD  LP
MOON MONSTERS, THE  MARQUIS, ROY  SR-P
MOON MOTH AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #17]  VANCE, JACK  SR-H
MOON ODYSSEY [SPACE: 1999 2]  RANKINE, JOHN  C/P:2,P
MOON OF GOMRATH, THE [ALDERLY 2]  GARNER, ALAN  C/P:2,P
MOON OF ICE  LINAWEAVER, BRAD  H
MOON OF MUTINY  DEL REY, LESTER  H:2,P
MOON OF SKULLS, THE [SOLOMON KANE 1]  HOWARD, ROBERT E.  C/P,P
MOON OF THE SPIDER [@DIABLO]  KNAAK, RICHARD A.  P
MOON OF THE WOLF  WHITTEN, LESLIE H.  P
MOON OF THREE RINGS [MOON SINGER 1]  NORTON, ANDRE  C/P:2,P
MOON ON AN IRON MEADOW  TATE, PETER  H,SFWA
MOON PEOPLE, THE  COBLENTZ, STANTON A.  C/P,P
MOON POOL, THE  MERRITT, A.  C/P,SR-LP,P,H:2
MOON RAIDERS, THE  BOUNDS, SYDNEY J.  SR-P
MOON RIDERS, THE  TOMLINSON, THERESA  H
MOON ROCKET  MULLER, JOHN E.  H
MOON SHOTS  CROWTHER, PETER  C/PA,PA
MOON TRIP  NEPHEW, WILLIAM/CHESTER, MICHAEL  [C/RAN-P]
MOON WAR  HUGHES, DENIS  SR-P
MOON ZERO TWO  BURKE, JOHN  C/P,P
MOONBALL, THE  WILLIAMS, URSULA MORAY  P
MOONBANE  SARRANTONIO, AL  C/P,SFWA-P,P
MOONBANE MAGE, THE  MARKS, LAURIE J.  P
MOONBASE ONE  JONES, RAYMOND F.  H
MOONBEAST  VAN VOGT, A. E.  P
MOONBIRD  DAVIS, GRANIA  H
MOONCHILD  CROWLEY, ALEISTER  C/P,P
MOONCHILD, THE  MCKENNEY, KENNETH  P
MOONCROW  MASSA, JACK  P
MOONDUST  SWANN, THOMAS BURNETT  C/P:2,P
MOONDUST AND MADNESS  TAYLOR, JANELLE  P
MOONFALL  MCDEVITT, JACK  C/P,P
MOONFERNS AND STARSONGS  SILVERBERG, ROBERT  C/P:2,P
MOONGATHER [DUEL OF SORCERY 1]  CLAYTON, JO  C/P:2,P
MOONHARE, THE  HAMPTON, KIRK  LP
MOONHEART [TAMSON HOUSE 1]  DE LINT, CHARLES  C/P:2,P
MOONLIGHT AND VINES [NEWFORD]  DE LINT, CHARLES  C/P,P,H

MOONLIGHT TRAVELER:

GREAT TALES OF FANTASY AND IMAGINATION, THE STERN, PHILIP VAN DOREN

HA

MOONLOVERS, THE  LORNQUEST, OLAF  C/P,P
MOONRAKER [JAMES BOND 3]  FLEMING, IAN  C/P:2,P
MOONRAPERS, THE  BANKS, RAYMOND E.  SR-L
MOONRISE [MOONBASE 1]  BOVA, BEN  C/P,P
MOONRISE [SNOWFALL 3]  SMITH, MITCHELL  P
MOON’S FIRE-EATING DAUGHTER, THE  MYERS, JOHN MYERS  H,LP
MOONS FOR SALE  GRIDBAN, VOLSTED  SR-P
MOONS OF MIRRODIN, THE [@MAGIC: THE GATHERING MIRRODIN 1]  MCDERMOTT, WILL  P
MOONS OF TRIOPUS  RANKINE, JOHN  C/P,P
MOON’S SHADOW, THE [SKOLIAN EMPIRE 8]  ASARO, CATHERINE  H
MOON’S WIFE, THE  ATTANASIO, A. A.  H
MOONSCATTER [DUEL OF SORCERY 2]  CLAYTON, JO  C/P:2,P
MOONSEED  BAXTER, STEPHEN  H,P
MOONSHINE [CAL AND NIKO 2]  THURMAN, ROB  P
MOONSINGER [MOON SINGER 1,2]  NORTON, ANDRE  H
MOONSINGER’S FRIENDS  SHWARTZ, SUSAN  HA,PA,SFWA-P
MOONSPEAKER [HOUSE OF MOONS 1]  WENTWORTH, K. D.  P
MOONSPIN  CARPENTER, ELMER J.  P
MOONSTALKER [@TOM SWIFT (ARCHWAY) #7]  APPLETON, VICTOR  P
MOONSTAR ODYSSEY  GERROLD, DAVID  C/P:2,P
MOONSTRUCK  LERNER, EDWARD M.  H
MOONWALK  FYFE, H. B.  B
MOONWAR [MOONBASE 2]  BOVA, BEN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 236
TITLE  AUTHOR  CODE
MOONWIND  LAWRENCE, LOUISE  H
MOONWISE  GILMAN, GREER ILENE  C/P,P
MORAL RE-ARMAMENT — WHAT IS IT?  ENTWISTLE, BASIL/ROOTS, JOHN MCCOOK  [C/RAN-P]
MORBID TASTE FOR BONES, A  PETERS, ELLIS  P
MORDENHEIM [RAVENLOFT]  WILLIAMSON, CHET  P

MORDRED [BUCK ROGERS (NIVEN & POURNELLE’S) 2] HOLMES, JOHN ERIC/NIVEN, LARRY/

POURNELLE, JERRY

C/P,P

MORDRED, BASTARD SON  CLEGG, DOUGLAS  H
MORDRED’S CURSE [MORDRED’S CURSE 1]  MCDOWELL, IAN  P
MORE ADVENTURES IN TIME AND SPACE  HEALY, RAYMOND J./MCCOMAS, J. FRANCIS PA,C/PA
MORE ADVENTURES OF SAMURAI CAT [SAMURAI CAT 2]  ROGERS, MARK E.  SFWA
MORE ADVENTURES ON OTHER PLANETS  WOLLHEIM, DONALD A.  PA,C/PA
MORE AMAZING STORIES  MOHAN, KIM  HA
MORE GHOST STORIES  FURMAN, A. L.  PA
MORE ISSUES AT HAND  ATHELING, WILLIAM JR.  R/REF-H
MORE LIMEHOUSE NIGHTS  BURKE, THOMAS  H
MORE LITTLE MONSTERS  ELWOOD, ROGER/GHIDALIA, VIC  PA
MORE MACABRE  WOLLHEIM, DONALD A.  PA
MORE MAGIC [MAGIC 3]  NIVEN, LARRY  LP
MORE NIGHTMARES  BLOCH, ROBERT  P
MORE PERFECT UNION, A  STAPP, ROBERT  H,P
MORE SCIENCE FICTION TALES  ELWOOD, ROGER  HA,SR-VLHA
MORE SHORT & SHIVERY  SAN SOUCCI, ROBERT D.  H
MORE SHORT TRIPS [DOCTOR WHO]  COLE, STEPHEN  PA
MORE SOVIET SCIENCE FICTION  COLLIER BOOKS  PA,C/PA
MORE STORIES FROM THE HUGO WINNERS II  ASIMOV, ISAAC  PA,C/PA
MORE STORIES FROM THE TWILIGHT ZONE  SERLING, ROD  C/P:2,P
MORE TALES FROM THE FORBIDDEN PLANET  KAVENY, ROZ  HA
MORE TALES OF PIRX THE PILOT  LEM, STANISLAW  C/P,P,H
MORE TALES OF THE BLACK WIDOWERS  ASIMOV, ISAAC  C/P:2,P
MORE TALES TO TREMBLE BY  SUTTON, STEPHEN P.  HA
MORE THAN FIRE [WORLD OF TIERS 6]  FARMER, PHILIP JOSE  H
MORE THAN HONOR [HONOR HARRINGTON/WORLDS OF HONOR 1]  WEBER, DAVID  C/PA,PA
MORE THAN HUMAN  STURGEON, THEODORE  H:2,SR-P

MORE THAN HUMAN: THE GRAPHIC STORY VERSION STURGEON, THEODORE/NINO, ALEX/

MOENCH, DOUG

CX

MORE THAN MORTAL [VICTOR RENQUIST 2]  FARREN, MICK  H
MORE THAN ONE UNIVERSE  CLARKE, ARTHUR C.  C/P,P
MORE THAN SUPERHUMAN  VAN VOGT, A. E.  SR-P,C/P
MORE THINGS IN HEAVEN  BRUNNER, JOHN  C/P:2,P
’MORE THINGS IN HEAVEN...’  OWEN, WALTER  H
MORE TOMORROW & OTHER STORIES  SMITH, MICHAEL MARSHALL  H
MORE WANDERING STARS  DANN, JACK  SFWA
MORE WHATDUNITS  RESNICK, MIKE  PA
MORE WOMEN OF WONDER  SARGENT, PAMELA  C/PA,PA
MOREAU FACTOR, THE  CHALKER, JACK L.  P
MORETA: DRAGONLADY OF PERN [PERN]  MCCAFFREY, ANNE  H:2,P
MOREVI [RAFE AND ASKANA]  LEE, LISA/MORRIS, TEE  LP
MORGAWR [@VOYAGE OF THE JERLE SHANNARA 3/SHANNARA JERLE]  BROOKS, TERRY  H
MORGOTH’S RING [HISTORY OF MIDDLE-EARTH 10]  TOLKIEN, J. R. R.  H
MORIGU: THE DEAD [MORIGU 1]  PERRY, MARK C.  P
MORIGU: THE DESECRATION [MORIGU 2]  PERRY, MARK C.  P
MORK AND MINDY  CHURCH, RALPH  P
MORLAC: THE QUEST OF THE GREEN MAGICIAN  RUSE, GARY ALAN  P
MORLOCK NIGHT  JETER, K. W.  P
MORNING CHILD AND OTHER STORIES  DOZOIS, GARDNER  LP
MORNING OF CREATION [@DESTINY MAKERS #2]  SHUPP, MIKE  C/P:2,P
MORNING OF THE MAGICIANS, THE  PAUWELS, LOUIS/BERGIER, JACQUES  [C/RAN-P]
MORNING STAR  HAGGARD, H. RIDER  C/P,P
MORNINGSTAR  GEMMELL, DAVID  P
MORPHEUS FACTOR, THE [@STARGATE SG-1 (MCCONNELL) 4]  MCCONNELL, ASHLEY  P
MORPHODITE, THE [MORPHODITE 1]  FOSTER, M. A.  C/P:2,P
MORRIGAN’S CROSS [CIRCLE 1]  ROBERTS, NORA  H
MORT [DISCWORLD 4/DISCWORLD DEATH 1]  PRATCHETT, TERRY  C/P,H,P
MORTAL BANE, A  GELLIS, ROBERTA  P

MORTAL CONSEQUENCES [@FORGOTTEN REALMS ARCANE AGE NETHERIL #3] EMERY, CLAYTON

P

MORTAL ENGINES  LEM, STANISLAW  C/P,P
MORTAL ENGINES [@HUNGRY CITY 1]  REEVE, PHILIP  H
MORTAL ENGINES (PROOF) [@HUNGRY CITY 1]  REEVE, PHILIP  SR-LP
MORTAL FEAR [@BUFFY]  CIENCIN, SCOTT/CIENCIN, DENISE  LP
MORTAL FEAR  COOK, ROBIN  C/P,P
MORTAL GLAMOUR, A  YARBRO, CHELSEA QUINN  C/P,P
MORTAL GODS  FAST, JONATHAN  P

The MITSFS Titledex  September 24, 2007  Page 237
TITLE  AUTHOR  CODE
MORTAL IMMORTAL, THE  SHELLEY, MARY  SR-LP
MORTAL IMMORTALS, THE  CRISTABEL  H,SFWA
MORTAL INSTRUMENTS, THE  BETHANCOURT, T. ERNESTO  P
MORTAL KOMBAT [MORTAL KOMBAT 1]  DELRIO, MARTIN  P
MORTAL KOMBAT [MORTAL KOMBAT 0]  ROVIN, JEFF  P
MORTAL MASK [CHIA BLACK DRAGON 2]  MARLEY, STEPHEN  LP
MORTALIS [DEMONWARS 4]  SALVATORE, R. A.  H
MORTALS AND MONSTERS  DEL REY, LESTER  C/P:2,P
MORTALS OF RENI, THE  GRUEN, VON  SR-P,H
MORTE DELLA TERRA  CUSTER, FRED  FP
MOSAIC [@STAR TREK VOYAGER]  TAYLOR, JERI  C/P,P

MOSCOW 1979 VON KUEHNELT-LEDDIHN, ERIK R. CHRISTIANE

SR-H

MOSCOW 2042  VOINOVICH, VLADIMIR  H
MOSSFLOWER [REDWALL]  JACQUES, BRIAN  C/P,P
MOST AMAZING MAN WHO EVER LIVED, THE [CORNELIUS MURPHY 3]  RANKIN, ROBERT  P
MOST ANCIENT SONG [GODS OF IRELAND 1]  FLYNN, CASEY  C/P,P
MOSTLY HARMLESS [HITCHHIKERS 5]  ADAMS, DOUGLAS  H:2
MOSUTHA’S MAGIC [@ORON #3]  SMITH, DAVID C.  P
MOTE IN GOD’S EYE, THE [MOTE IN GOD’S EYE 1]  NIVEN, LARRY/POURNELLE, JERRY  C/P,P,H:2,SFWA
MOTE IN TIME’S EYE, THE  KLEIN, GERARD  C/P:2,P
MOTHER EARTH [MORGAN OF ALB’NY 2]  MCQUAY, MIKE  P
MOTHER EARTH FATHER SKY  HARRISON, SUE  H
MOTHER GRIMM  WELLS, CATHERINE  P
MOTHER LODE  HUGHES, ZACH  P
MOTHER LONDON  MOORCOCK, MICHAEL  LP
MOTHER NIGHT  VONNEGUT, KURT JR.  C/P:3,P
MOTHER OCEAN, DAUGHTER SEA [BRIERLY 1]  MARCELLAS, DIANA  P
MOTHER OF DEMONS  FLINT, ERIC  C/P:2,P
MOTHER OF INVENTION [@FLYING NUN #3]  JOHNSTON, WILLIAM  P
MOTHER OF KINGS  ANDERSON, POUL  H
MOTHER OF LIES [CHILDREN OF CHAOS 2]  DUNCAN, DAVE  H
MOTHER OF PLENTY [PLENTY 3]  GREENLAND, COLIN  P
MOTHER OF STORMS  BARNES, JOHN  C/P,P
MOTHER OF STORMS [@STAR REQUIEM #1]  COLE, ADRIAN  P,SFWA-P
MOTHER OF WINTER [DARWATH 4]  HAMBLY, BARBARA  H:2,P
MOTHER WAS A LOVELY BEAST  FARMER, PHILIP JOSE  C/PA,PA
MOTHERLINES [HOLDFAST 2]  CHARNAS, SUZY MCKEE  H,P
MOTION MENACE [DOC SAVAGE #64]  ROBESON, KENNETH  C/P,@P
MOTION OF LIGHT IN WATER, THE  DELANY, SAMUEL R.  R/REF-H
MOULOKIN MISSION, DIE [HUMANX ICERIGGER 2]  FOSTER, ALAN DEAN  FP
MOUNT, THE  EMSHWILLER, CAROL  LP
MOUNTAIN MADE OF LIGHT, THE [MOUNTAIN MADE OF LIGHT 1]  MYERS, EDWARD  P

MOUNTAIN MAGIC DRAKE, DAVID/FLINT, ERIC/SPOOR, RYK E./

KUTTNER, HENRY

C/PA,PA

MOUNTAIN MONSTER, THE [DOC SAVAGE #84]  ROBESON, KENNETH  C/P:2,@P
MOUNTAIN OF BLACK GLASS [@OTHERLAND #3]  WILLIAMS, TAD  C/P,P,H
MOUNTAIN OF MIRRORS  ESTES, ROSE  [C/RAN-P:2]
MOUNTAIN OF STOLEN DREAMS [@PLANET BUILDERS #1]  TALLIS, ROBYN  C/P,P
MOUNTAIN WALKS, THE [PEACE COMPANY]  GREEN, ROLAND J.  P
MOUNTAINS AND MADNESS [@IRON DRAGONS]  ESTES, ROSE  P,SFWA-P
MOUNTAINS AT THE BOTTOM OF THE WORLD, THE  CAMERON, IAN  P
MOUNTAIN’S CALL, THE [AURELIAN EMPIRE 1]  BRENNAN, CAITLIN  LP
MOUNTAINS OF BREGA, THE [@BLADE #17]  LORD, JEFFREY  P
MOUNTAINS OF CHANNADRAN, THE [WINTER KING’S WAR 3]  DEXTER, SUSAN  C/P,P
MOUNTAINS OF MAJIPOOR, THE  SILVERBERG, ROBERT  H:2
MOUNTAINS OF THE SUN, THE  LEOURIER, CHRISTIAN  C/P:2,P
MOUSE AND HIS CHILD, THE  HOBAN, RUSSELL  P
MOUSE ON THE MOON, THE  WIBBERLEY, LEONARD  C/P:2,P
MOUSE ON WALL STREET, THE  WIBBERLEY, LEONARD  P
MOUSE THAT ROARED, THE  WIBBERLEY, LEONARD  C/P,P,H
MOUSE THAT SAVED THE WEST, THE  WIBBERLEY, LEONARD  H
MOUTHFUL OF TONGUES, A  DI FILIPPO, PAUL  H
MOVE UNDER GROUND  MAMATAS, NICK  LP
MOVEMENT OF MOUNTAINS, THE  BLUMLEIN, MICHAEL  H,P
MOVIE MONSTERS  ORMSBY, ALAN  R/REF-P
MOVING FINGER, THE  LIVINGSTON, BERKELEY  B
MOVING MARS [QUEEN OF ANGELS 3]  BEAR, GREG  H:2,P
MOVING PICTURES [DISCWORLD 10]  PRATCHETT, TERRY  C/P:3,P,H
MR. BRITLING SEES IT THROUGH  WELLS, H. G.  H
MR. FOX AND OTHER FERAL TALES  PARTRIDGE, NORMAN  SR-LP
MR. LARS, SOGNATORE D’ARMI  DICK, PHILIP K.  FP
MRS. CALIBAN  INGALLS, RACHEL  LP
MRS. COVERLET’S MAGICIANS  NASH, MARY  P

TITLE  AUTHOR  CODE
MRS. DEMMING AND THE MYTHICAL BEAST  SULLIVAN, FAITH  P,SFWA-P
MRS. FRISBY AND THE RATS OF NIMH [NIMH 1]  O’BRIEN, ROBERT C.  P
MU REVEALED  EARLL, TONY  [C/RAN-P]
MUCH OF MAGIC AND MIRACLES  HARDY, NAOMI  P
MUCKER, THE  BURROUGHS, EDGAR RICE  H
MUDD IN YOUR EYE [STAR TREK]  OLTION, JERRY  C/P:2,P
MUDDLE EARTH  BRUNNER, JOHN  C/P:2,P
MUDD’S ANGELS [STAR TREK]  LAWRENCE, J. A.  C/P:2,P
MUDD’S WOMEN [STAR TREK]  RODDENBERRY, GENE/KANDEL, S.  SR-S
MULENGRO  DE LINT, CHARLES  C/P,P
MULLER-FOKKER EFFECT, THE  SLADEK, JOHN  C/P:2,P,H
MULTIFACE [INTERFACE 3]  ADLARD, MARK  P
MULTIPLE MAN, THE  BOVA, BEN  C/P,SFWA-P,P,H
MULTIPLEX MAN, THE  HOGAN, JAMES P.  H:2,P
MULTIPLEX MAN, THE (PROOF)  HOGAN, JAMES P.  SR-LP
MULTITUDE OF MONSTERS, A [EBENEZUM 2]  GARDNER, CRAIG SHAW  C/P:2,P
MULTITUDE OF VISIONS, A  CHAUVIN, CY  R/REF-P
MUMMY, THE  DREADSTONE, CARL  P
MUMMY, THE  RICE, ANNE  C/P,P,H
MUMMY, THE WILL, AND THE CRYPT, THE  BELLAIRS, JOHN  LP
MUMMY WALKS AMONG US, THE  GHIDALIA, VIC  PA
MUNDO DE CRISTAL, O  BALLARD, J. G.  FP
MUNDOS CREADOS I  CLARESON, THOMAS D.  FP
MUNDOS CREADOS II  CLARESON, THOMAS D.  FP
MUNITIONS MASTER, THE [DOC SAVAGE #58]  ROBESON, KENNETH  C/P:2,@P
MUNSTERS, THE  COOPER, MORTON  P
MUNSTERS AND THE GREAT CAMERA CAPER, THE  JOHNSTON, WILLIAM  H
MURAL MASTER, THE  JONES, ADRIENNE  H
MURASAKI  SILVERBERG, ROBERT  SR-HA,C/PA,PA
MURDER AND MAGIC [LORD DARCY 1]  GARRETT, RANDALL  C/P:2,P
MURDER AT THE ABA  ASIMOV, ISAAC  H:2,P
MURDER AT THE GALACTIC WRITERS’ SOCIETY  ASIMOV, JANET  C/P,P
MURDER BY MAGIC: TWENTY TALES OF CRIME AND THE SUPERNATURAL  EDGHILL, ROSEMARY  LPA
MURDER BY TELECOPTER  HUGHES, DENIS  SR-P
MURDER GAME, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
MURDER IN CORMYR [@FORGOTTEN REALMS]  WILLIAMSON, CHET  P
MURDER IN MARIENBURG, A [@WARHAMMER]  BISHOP, DAVID  P
MURDER IN MILLENNIUM VI  GRAY, CURME  H:2
MURDER IN SPACE  REED, DAVID V.  C/P:2,GN4.23
MURDER IN SPACE  TUBB, E. C.  LP
MURDER IN TARSIS [@DRAGONLANCE CLASSICS]  ROBERTS, JOHN MADDOX  P
MURDER IN THE SOLID STATE  MCCARTHY, WIL  C/P:2,P
MURDER IN THE SOLID STATE (PROOF)  MCCARTHY, WIL  SR-LP
MURDER MELODY [DOC SAVAGE #15]  ROBESON, KENNETH  C/P:2,@P
MURDER MIRAGE [DOC SAVAGE #71]  ROBESON, KENNETH  C/P:2,@P
MURDER OF A MISTRESS  KUTTNER, HENRY  P
MURDER OF THE U.S.A., THE  JENKINS, WILL F.  SR-P
MURDER ON USHER’S PLANET  NOEL, ATANIELLE ANNYN  P
MURDER ON WHEELS [@AVENGER #13]  ROBESON, KENNETH  C/P:2,P
MURDERCON  PURTILL, RICHARD  SFWA
MURDERED SUN, THE [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  P
MURDERER INVISIBLE, THE  WYLIE, PHILIP  C/P,P
MURDERERS, THE  BROWN, FREDRIC  P
MURDERMOON  KUPPERBERG, PAUL  P
MURDOCK  SIMMONS, GEOFFREY  H
MURGUNSTRUMM  CAVE, HUGH B.  SR-H
MURMURS OF EARTH  SAGAN, CARL  SFWA
MURPHY’S GAMBIT  MITCHELL, SYNE  C/P,P
MURPHY’S LAW  BLOCH, ARTHUR  [C/RAN-P]
MUS OF KERBRIDGE  KIDD, PAUL  P
MUSASHI FLEX, THE  PERRY, STEVE  P
MUSE OF ART [GEODYSSEY 4]  ANTHONY, PIERS  H
MUSIC FROM THE PAST [HOLDERLY HALL 5]  CAMERON, KATE  P
MUSIC OF RAZORS, THE  ROGERS, CAMERON  LP
MUSIC OF THE SPEARS [@ALIENS 8]  NAVARRO, YVONNE  P
MUSIC TO MY SORROW [ERIC BANYON 6]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H
MUST LOVE DRAGONS  ROWE, STEPHANIE  P
MUSTAPHA AND HIS WISE DOG [TWELVE KINGDOMS 1]  FRIESNER, ESTHER M.  C/P:2,P
MUTAGENESIS  COLLINS, HELEN  C/P:2,P
MUTANT  KUTTNER, HENRY  C/P:2,P
MUTANT  PADGETT, LEWIS  H
MUTANT 59: THE PLASTIC-EATERS  PEDLER, KIT/DAVIS, GERRY  H:2,P
MUTANT BEACH [@TOM SWIFT (ARCHWAY) #11]  APPLETON, VICTOR  P
MUTANT FILES, THE  GREENBERG, MARTIN H./HELFERS, JOHN  PA

The MITSFS Titledex  September 24, 2007  Page 239
TITLE  AUTHOR  CODE
MUTANT HELL [@MUTANTS AMOK #2]  GRANT, MARK  C/P,P
MUTANT LEGACY [MUTANT SEASON 4]  HABER, KAREN  P
MUTANT PRIME, THE [MUTANT SEASON 2]  HABER, KAREN  C/P:2,LP
MUTANT SEASON, THE [MUTANT SEASON 1]  HABER, KAREN/SILVERBERG, ROBERT  C/P,LP,P
MUTANT STAR [MUTANT SEASON 3]  HABER, KAREN  P
MUTANT WEAPON, THE [MED 1]  LEINSTER, MURRAY  D66525,DD403

MUTANTS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA

MUTANTS  DICKSON, GORDON R.  C/P:2,P
MUTANTS, THE  NEVILLE, KRIS  C/P,P
MUTANTS  SILVERBERG, ROBERT  HA
MUTANTS AMOK [@MUTANTS AMOK #1]  GRANT, MARK  C/P,P
MUTANTS ARE COMING, THE  HAIBLUM, ISIDORE  C/P,SFWA,P
MUTANTS REBEL, THE  TUBB, E. C.  SR-P
MUTANTS VS. MUTANTS [@PERRY RHODAN #19]  ACE  PA
MUTATION  COOK, ROBIN  H,P
MUTE  ANTHONY, PIERS  C/P:2,P
MUTINEERS’ MOON [COLIN MACINTYRE 1]  WEBER, DAVID  C/P:2,SFWA-P,P
MUTINOUS WIND: A SORCERER’S TALE, THE  REYNARD, ELIZABETH  H
MUTINY IN SPACE  DAVIDSON, AVRAM  C/P:2,P
MUTINY ON THE ENTERPRISE [STAR TREK]  VARDEMAN, ROBERT E.  C/P:3,P
MY BEST SCIENCE FICTION STORY  MARGULIES, LEO/FRIEND, OSCAR J.  C/PA,PA
MY BIG FAT SUPERNATURAL WEDDING  ELROD, P. N.  LPA
MY BROTHER’S KEEPER  SHEFFIELD, CHARLES  C/P:2,P
MY CROWD  ADDAMS, CHARLES  CX
MY DIRTY LITTLE BOOK OF STOLEN TIMES  JENSEN, LIZ  H
MY ENEMY, MY ALLY [STAR TREK RIHANNSU 1]  DUANE, DIANE  C/P:3,P
MY FATHER IMMORTAL  WEAVER, MICHAEL D.  H,P
MY FAVORITE FANTASY STORY  GREENBERG, MARTIN H.  PA
MY FAVORITE SCIENCE FICTION STORY  GREENBERG, MARTIN H.  PA

MY FIRST TWO THOUSAND YEARS (ABRIDGED)

VIERECK, GEORGE SYLVESTER/ ELDRIDGE, PAUL

C/P,P

MY LADY EVIL  COOPER, PARLEY J.  P
MY LORD BARBARIAN  OFFUTT, ANDREW J.  C/P,P
MY NAME IS DEATH  BIRKIN, CHARLES  P
MY NAME IS LEGION  ZELAZNY, ROGER  C/P:2,P
MY NAME IS VLADIMIR SLOIFOISKI  ALPER, GERALD A.  P
MY PETITION FOR MORE SPACE  HERSEY, JOHN  C/P,H
MY ROBOT BUDDY  SLOTE, ALFRED  C/P,LP
MY SCIENCE PROJECT  MCQUAY, MIKE  C/P,P
MY SON, THE WIZARD [WIZARD IN RHYME 5]  STASHEFF, CHRISTOPHER  C/P,LP
MY SOUL TO KEEP  DAVIS, ELIZABETH  P
MY SOUL TO KEEP  DUE, TANANARIVE  LP
MY TEACHER FLUNKED THE PLANET [MY TEACHERS 4]  COVILLE, BRUCE  LP
MY TEACHER FRIED MY BRAINS [MY TEACHERS 2]  COVILLE, BRUCE  P
MY TEACHER GLOWS IN THE DARK [MY TEACHERS 3]  COVILLE, BRUCE  P
MY TEACHER IS AN ALIEN [MY TEACHERS 1]  COVILLE, BRUCE  P
MY TRIP TO ALPHA 1  SLOTE, ALFRED  LP
MYRIAD, THE [@TOUR OF THE MERRIMACK #1]  MELUCH, R. M.  H
MYRREN’S GIFT [QUICKENING 1]  MCINTOSH, FIONA  H,LP

MYST: THE BOOK OF ATRUS [@MYST 1] MILLER, RAND/MILLER, ROBYN/

WINGROVE, DAVID

C/P,P

MYST: THE BOOK OF D’NI [@MYST 2]  MILLER, RAND/WINGROVE, DAVID  P
MYST: THE BOOK OF TI’ANA [@MYST 3]  MILLER, RAND/WINGROVE, DAVID  H,P
MYSTERE DES SUPS, LE  BRUSS, B. R.  FP
MYSTERE JAUNE, LE  DE LA HIRE, JEAN  FP
MYSTERIES OF THE SKIES: UFO’S IN PERSPECTIVE  LORE, GORDON I. R./DENEAULT, HAROLD H. [C/RAN-H]
MYSTERIES OF THE WORM  BLOCH, ROBERT  C/P,P
MYSTERIOUS CIRCUS, THE (PROOF)  LANGTON, JANE  SR-LP
MYSTERIOUS CURE, THE  JEPPSON, J. O.  SFWA
MYSTERIOUS DOCUMENT, THE  VERNE, JULES  H
MYSTERIOUS FIRES AND LIGHTS  GADDIS, VINCENT H.  [C/RAN-P]
MYSTERIOUS ISLAND, THE (ABRIDGED, BAIR TRANSLATION)  VERNE, JULES  C/P,P
MYSTERIOUS ISLAND, THE (KINGSTON TRANSLATION)  VERNE, JULES  H,P
MYSTERIOUS PLANET, THE  DEL REY, LESTER  C/P,P
MYSTERIOUS PLANET, THE [DOCTOR WHO]  DICKS, TERRANCE  P
MYSTERIOUS PLANET, THE  WRIGHT, KENNETH  H:2

MYSTERIOUS VISIONS WAUGH, CHARLES G./

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

HA

MYSTERIUM  WILSON, ROBERT CHARLES  H,LP
MYSTERY FOR MR. BASS, A  CAMERON, ELEANOR  H
MYSTERY MEN OF MARS, THE  CLAUDY, CARL H.  H:2
MYSTERY OF ARTHUR GORDON PYM, THE  POE, EDGAR ALLAN/VERNE, JULES  H

TITLE  AUTHOR  CODE
MYSTERY OF COLLINWOOD, THE  ROSS, MARILYN  P
MYSTERY OF SATELLITE 7  COOMBS, CHARLES  C/P,P
MYSTERY OF THE ANTI, THE [@PERRY RHODAN #88]  ACE  C/PA,PA
MYSTERY OF THE BURNING OCEAN, THE  LYLE, MEL  H
MYSTERY OF THE DRAGON’S SHADOW, THE  WEINBERG, ROBERT  SR-LP
MYSTERY OF THE LOST MOON, THE  CHESTER, MICHAEL  H

MYSTERY OF THE MISSING CREW [@STAR TREK TNG STARFLEET ACADEMY #6] FRIEDMAN, MICHAEL JAN

H,LP

MYSTERY OF THE SEA HORSE, THE [PHANTOM 7]  SHAWN, WILLIAM S.  P
MYSTERY OF THE THIRD MINE  LOWNDES, ROBERT A. W.  H:2

MYSTERY OF THE WIZARD’S TOMB [@KNIGHTS OF THE SILVER DRAGON #11] PLUMMER, RACHEL

LP

MYSTERY OF THE YELLOW ROOM, THE  LEROUX, GASTON  H
MYSTERY ON HAPPY BONES, THE [DOC SAVAGE #96]  ROBESON, KENNETH  @P
MYSTERY ON THE SNOW, THE [DOC SAVAGE #69]  ROBESON, KENNETH  C/P:2,@P
MYSTERY PLANET, THE  KENNEDY, EDGAR REES  SR-P
MYSTERY UNDER THE SEA [DOC SAVAGE #27]  ROBESON, KENNETH  C/P,@P
MYSTERY WALK  MCCAMMON, ROBERT R.  C/P,P
MYSTIC AND RIDER [TWELVE HOUSES]  SHINN, SHARON  H
MYSTIC EMPIRE [@BRONZE CANTICLES #3]  HICKMAN, TRACY/HICKMAN, LAURA  H
MYSTIC KNOLL [@CHARMED]  GALLAGHER, DIANA G.  P
MYSTIC MULLAH, THE [DOC SAVAGE #9]  ROBESON, KENNETH  @P
MYSTIC QUEST [@BRONZE CANTICLES #2]  HICKMAN, TRACY/HICKMAN, LAURA  H:2
MYSTIC ROSE, THE [@CELTIC CRUSADES 3]  LAWHEAD, STEPHEN R.  P
MYSTIC WARRIOR [@BRONZE CANTICLES #1]  HICKMAN, TRACY/HICKMAN, LAURA  P
MYTH ADVENTURES [MYTH ADVENTURES 1,2,3,4]  ASPRIN, ROBERT LYNN  H:2
MYTH ALLIANCES (OMNIBUS EDITION) [MYTH ADVENTURES 5,6,7]  ASPRIN, ROBERT LYNN  H:2
MYTH ALLIANCES (SINGLE NOVEL) [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  H,P
MYTH CONCEPTIONS [MYTH ADVENTURES 2]  ASPRIN, ROBERT LYNN  C/P,P,H
MYTH DIRECTIONS [MYTH ADVENTURES 3]  ASPRIN, ROBERT LYNN  C/P,LP,H

MYTH DIRECTIONS [MYTH ADVENTURES 3,4]/ HIT OR MYTH [MYTH ADVENTURES 3,4] ASPRIN, ROBERT LYNN

P

MYTH-GOTTEN GAINS [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  LP
MYTH HUNTERS, THE [VEIL 1]  GOLDEN, CHRISTOPHER  LP
M.Y.T.H. INC. IN ACTION [MYTH INC 3]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P,H
M.Y.T.H. INC. LINK [MYTH ADVENTURES 7/MYTH INC 1]  ASPRIN, ROBERT LYNN  C/P,SFWA-P,H,LP
MYTH-ING PERSONS [MYTH ADVENTURES 5]  ASPRIN, ROBERT LYNN  C/P,P,H
MYTH-ION IMPROBABLE [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P,P
MYTH MAKERS, THE [DOCTOR WHO]  COTTON, DONALD  C/P,P
MYTH-NOMERS AND IM-PERVECTIONS [MYTH INC 2]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P,P,H
MYTH, SYMBOL AND RELIGION IN THE LORD OF THE RINGS  MIESEL, SANDRA  R/REF-P
MYTH-TAKEN IDENTITY [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  LP
MYTH-TOLD TALES [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN/NYE, JODY LYNN  P
MYTHAGO WOOD [RYHOPE WOOD 1]  HOLDSTOCK, ROBERT P.  H:2,SFWA,P
MYTHIC PERSPECTIVES FICTION 1998  GNAWING IDEAS  LPA

MYTHICAL BEASTIES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

MYTHMASTER  KELLEY, LEO P.  C/P,P
MYTHOLOGY 101 [MYTHOLOGY 1]  NYE, JODY LYNN  P
MYTHOLOGY ABROAD [MYTHOLOGY 2]  NYE, JODY LYNN  C/P,P
MYTHOPOEIKON  WOODROFFE, PATRICK  A
MYTHS FOR THE MODERN AGE  ECKERT, WIN SCOTT  LPA
MYTHS, LEGENDS, AND TRUE HISTORY  LANDIS, GEOFFREY A.  SR-H
MYTHS OF MAGIC ANTHOLOGY, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA

TITLE  AUTHOR  CODE
N-SPACE  NIVEN, LARRY  H:2,P
NADA THE LILY  HAGGARD, H. RIDER  LP
NADYA  MURPHY, PAT  P
NAGASAKI VECTOR, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P

NAIL AND THE ORACLE, THE [COMPLETE STORIES OF THEODORE STURGEON 11] STURGEON, THEODORE

H

NAIL DOWN THE STARS  MORRESSY, JOHN  H,SFWA
NAKED CAME THE ROBOT  LONGYEAR, BARRY B.  C/P:2,P
NAKED CAME THE SASQUATCH  BOSTON, JOHN  P
NAKED EMPIRE [SWORD OF TRUTH 8]  GOODKIND, TERRY  H
NAKED FLESH OF FEELING, THE  WILLIAMSON, J. N.  SR-H
NAKED GOD, THE [NIGHT’S DAWN 3]  HAMILTON, PETER F.  H
NAKED STORM, THE  EISNER, SIMON  P
NAKED SUN, THE [LIJE BALEY 2]  ASIMOV, ISAAC  C/P:3,P,H
NAKED TO THE STARS  DICKSON, GORDON R.  C/P:2,SFWA-TD31,P
NAME FOR EVIL, A  LYTLE, ANDREW  P
NAME OF A SHADOW  MAXWELL, ANN  C/P:2,P
NAME OF THE GAME, THE [@GREYHAWK #6]  ESTES, ROSE  C/P:2,P
NAME OF THE ROSE, THE  ECO, UMBERTO  H,P
NAME OF THE SUN, THE [AVERIDAN 3]  CLOUGH, BRENDA W.  C/P:2,P
NAME OF THE WIND, THE [KINGKILLER 1]  ROTHFUSS, PATRICK  H
NAME TO CONJURE WITH, A  AAMODT, DONALD  C/P:2,P
NAMELESS, THE  CAMPBELL, RAMSEY  H,P
NAMELESS DAY, THE [CRUCIBLE 1]  DOUGLASS, SARA  P
NAMELESS MAGERY [MENNENKALTENEI 1]  TURNER, DELIA MARSHALL  P
NAMELESS PLACES  PAGE, GERALD W.  SR-HA,HA
NAMELESS WITCH, A  MARTINEZ, A. LEE  LP
NAMING OF KINZEL, THE  LEE, SHARON/MILLER, STEVE  SR-LP
NANCY, THE DARING [@PHENWICK #11]  KIMBROUGH, KATHERYN  P
NANNY AND THE PROFESSOR [NANNY AND THE PROFESSOR 1]  JOHNSTON, WILLIAM  P
NANO FLOWER, THE [GREG MANDEL 3]  HAMILTON, PETER F.  H
NANODREAMS  ELLIOTT, ELTON  PA,SFWA-P
NANOTECH  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
NANOTECH CHRONICLES, THE  FLYNN, MICHAEL  C/P,SFWA-P,P
NANOTECH WAR, THE [STAR TREK VOYAGER]  PIZIKS, STEVEN  P
NAPOLEON DISENTIMED [MACNAIR OF MACNAIR 1]  PEIRCE, HAYFORD  C/P:2,SFWA-P,P
NAPOLEON WAGER, THE [GAMESTER WARS 3]  FORSTCHEN, WILLIAM R.  P
NAPOLEONS OF ERIDANUS, THE  BARBET, PIERRE  C/P,P
NARABEDLA, LTD.  POHL, FREDERIK  C/P,P,H
NARCISSUS IN CHAINS [ANITA BLAKE 10]  HAMILTON, LAURELL K.  C/P,P,H
NARGUN AND THE STARS, THE  WRIGHTSON, PATRICIA  LP
NARROW LAND, THE  VANCE, JACK  P
NARROWS, THE  IRVINE, ALEXANDER C.  LP
NASA-UNIVERSITY CONFERENCE 1  NASA  [C/RAN-H]
NASA-UNIVERSITY CONFERENCE 2  NASA  [C/RAN-H]
NASTY STORIES  MCNAUGHTON, BRIAN  LP
NATCHEZ [@SPECTROS #3]  WINTERS, LOGAN  P
NATHANIEL  SAUL, JOHN  C/P,P
NATIONAL LAMPOON’S DOON  WEINER, ELLIS  C/P:2,P
NATIONS OF THE NIGHT, THE [@LIGHTBRINGER #2]  JOHNSON, OLIVER  H,P
NATIVE TONGUE [NATIVE TONGUE 1]  ELGIN, SUZETTE HADEN  C/P:2,P
NATIVES OF SPACE  CLEMENT, HAL  C/P:2,P
NATURAL HISTORY  ROBSON, JUSTINA  LP
NATURAL SELECTION [@BATTLETECH #5]  STACKPOLE, MICHAEL A.  C/P,P
NATURE OF THE BEAST  FAWKES, RICHARD  P
NATURE’S END  STRIEBER, WHITLEY/KUNETKA, JAMES  H,P
NATURE’S GOD [@HISTORICAL ILLUMINATUS #3]  WILSON, ROBERT ANTON  P

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 1 [@NAUSICAA PERFECT #1] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 2 [@NAUSICAA PERFECT #2] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 3 [@NAUSICAA PERFECT #3] MIYAZAKI, HAYAO

CX

NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 4 [@NAUSICAA PERFECT #4] MIYAZAKI, HAYAO

CX

NAUTILUS [STARFARERS 4]  MCINTYRE, VONDA N.  C/P,P
NAUTILUS SANCTION, THE [TIMEWARS 5]  HAWKE, SIMON  C/P:2,P
NAUTIPUSS  ALLISON, CLYDE  SR-L
NAVIGATING THE GOLDEN COMPASS  YEFFETH, GLENN  R/REF-H
NAVIGATOR OF RHADA, THE [RHADA 3]  GILMAN, ROBERT CHAM  C/P,P,H
NAVIGATOR’S SINDROME [RABELAIS 1]  CARR, JAYGE  H,SFWA
NAVOHAR  BELL, HILARI  P
NEANDERTHAL  DARNTON, JOHN  C/P,P,H
NEANDERTHAL PLANET  ALDISS, BRIAN W.  H:2,P

The MITSFS Titledex  September 24, 2007  Page 242
TITLE  AUTHOR  CODE

NEANDERTHALS SILVERBERG, ROBERT/

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

C/PA,PA

NEAREST FIRE, THE [TORIN 2]  WILDER, CHERRY  H,P
NEBULA ALERT  CHANDLER, A. BERTRAM  DG632
NEBULA AWARD STORIES  KNIGHT, DAMON  PA,C/PA
NEBULA AWARD STORIES 2  ALDISS, BRIAN W./HARRISON, HARRY  HA,PA
NEBULA AWARD STORIES 3  ZELAZNY, ROGER  PA,C/PA
NEBULA AWARD STORIES 4  ANDERSON, POUL  C/PA,PA
NEBULA AWARD STORIES 5  BLISH, JAMES  C/PA,PA
NEBULA AWARD STORIES 6  SIMAK, CLIFFORD D.  C/PA,PA,SFWA
NEBULA AWARD STORIES 7  BIGGLE, LLOYD JR.  C/PA,PA
NEBULA AWARD STORIES 8  ASIMOV, ISAAC  C/PA,PA
NEBULA AWARD STORIES 9  WILHELM, KATE  C/PA,PA
NEBULA AWARD STORIES 10  GUNN, JAMES E.  C/PA,PA
NEBULA AWARD STORIES 11  LE GUIN, URSULA K.  HA,PA
NEBULA AWARD STORIES 16  POURNELLE, JERRY  C/PA,PA
NEBULA AWARD STORIES 17  HALDEMAN, JOE  C/PA,PA
NEBULA AWARD STORIES 1965  KNIGHT, DAMON  HA
NEBULA AWARD-WINNING NOVELLAS  GREENBERG, MARTIN H.  LPA
NEBULA AWARDS 18, THE  SILVERBERG, ROBERT  C/PA,PA
NEBULA AWARDS 19, THE  RANDALL, MARTA  HA
NEBULA AWARDS 20  ZEBROWSKI, GEORGE  HA,LPA
NEBULA AWARDS 21  ZEBROWSKI, GEORGE  LPA
NEBULA AWARDS 22  ZEBROWSKI, GEORGE  LPA
NEBULA AWARDS 23  BISHOP, MICHAEL  HA
NEBULA AWARDS 24  BISHOP, MICHAEL  SFWA,LPA
NEBULA AWARDS 25  BISHOP, MICHAEL  LPA
NEBULA AWARDS 26  MORROW, JAMES  HA,LPA,SFWA
NEBULA AWARDS 27  MORROW, JAMES  HA
NEBULA AWARDS 28  MORROW, JAMES  HA,LPA
NEBULA AWARDS 29  SARGENT, PAMELA  LPA
NEBULA AWARDS 30  SARGENT, PAMELA  LPA
NEBULA AWARDS 31  SARGENT, PAMELA  LPA
NEBULA AWARDS 32  DANN, JACK  LPA
NEBULA AWARDS 33  WILLIS, CONNIE  LPA
NEBULA AWARDS SHOWCASE 2000  BENFORD, GREGORY  LPA
NEBULA AWARDS SHOWCASE 2001  SILVERBERG, ROBERT  LPA
NEBULA AWARDS SHOWCASE 2002  ROBINSON, KIM STANLEY  LPA
NEBULA AWARDS SHOWCASE 2003  KRESS, NANCY  LPA
NEBULA AWARDS SHOWCASE 2004  MCINTYRE, VONDA N.  LPA
NEBULA AWARDS SHOWCASE 2005, THE  DANN, JACK  LPA
NEBULA AWARDS SHOWCASE 2006  DOZOIS, GARDNER  LPA
NEBULA AWARDS SHOWCASE 2007  RESNICK, MIKE  LPA
NEBULA EXPRESS  SUNBORD, TERRY  LP
NEBULA MAKER/FOUR ENCOUNTERS  STAPLEDON, OLAF  LP
NEBULA WINNERS 12  DICKSON, GORDON R.  C/PA,PA,SFWA
NEBULA WINNERS 13  DELANY, SAMUEL R.  C/PA,PA,SFWA
NEBULA WINNERS 14  POHL, FREDERIK  C/PA,PA
NEBULA WINNERS 15  HERBERT, FRANK  C/PA,PA,SFWA
NEBULA-X  STATTEN, VARGO  SR-P
NEBULON HORROR, THE  CAVE, HUGH B.  P
NECESSARY BEGGAR, THE  PALWICK, SUSAN  P
NECKLACE OF RAINDROPS, A  AIKEN, JOAN  P
NECROM  FARREN, MICK  C/P,P
NECROMANCER [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H
NECROMANCER [@WARHAMMER]  GREEN, JONATHAN  P
NECROMANCER  HOLDSTOCK, ROBERT P.  P
NECROMANCER NINE [TRUE GAME 2]  TEPPER, SHERI S.  C/P:2,P
NECROMANCERS, THE  BENSON, ROBERT HUGH  LP
NECROMANCERS, THE  HAINING, PETER  HA

NECROMANCIES AND NETHERWORLDS SCHWEITZER, DARRELL/ VAN HOLLANDER, JASON

SR-H

NECRONOMICON  TYSON, DONALD  LP
NECROPOLIS [WARHAMMER 40K GAUNT’S GHOSTS 3]  ABNETT, DAN  P
NECROPOLIS  COPPER, BASIL  H
NECROSCOPE [NECROSCOPE 1]  LUMLEY, BRIAN  SFWA-P:2
NECROSCOPE: AVENGERS [NECROSCOPE 13/E-BRANCH 3]  LUMLEY, BRIAN  H
NECROSCOPE: INVADERS [NECROSCOPE 11/E-BRANCH 1]  LUMLEY, BRIAN  P
NECROSCOPE: RESURGENCE [NECROSCOPE 10/LOST YEARS 2]  LUMLEY, BRIAN  P
NECROSCOPE: THE LOST YEARS [NECROSCOPE 9/LOST YEARS 1]  LUMLEY, BRIAN  P
NECTAR OF HEAVEN [DUMAREST 24]  TUBB, E. C.  C/P:2,P
NEEDFUL THINGS  KING, STEPHEN  C/P,P,H
NEEDING GHOSTS  CAMPBELL, RAMSEY  H,LP
NEEDLE  CLEMENT, HAL  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 243
TITLE  AUTHOR  CODE
NEEDLE IN A TIMESTACK  SILVERBERG, ROBERT  C/P:2,P
NEEDLE ON FULL, THE  FORBES, CAROLINE  LP
NEENA GATHERING  COLANDER, VALERIE N.  C/P:2,P
NEGATIVE MINUS  FANTHORPE, R. L.  SR-P
NEGATIVE ONES, THE  MULLER, JOHN E.  SR-P
NEGATIVE ZONE, THE [@TOM SWIFT (ARCHWAY) #2]  APPLETON, VICTOR  P

NEGLECTED VISIONS MALZBERG, BARRY N./

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

HA

NEIL GAIMAN’S MIDNIGHT DAYS  GAIMAN, NEIL/WAGNER, MATT  H,CX
NEKROPOLIS  MCHUGH, MAUREEN F.  H
NEL NOME DELL’UOMO  MONTANARI, GIANNI  FP
NELLIE, THE OBVIOUS [@PHENWICK #21]  KIMBROUGH, KATHERYN  P
NEMESIS  ASIMOV, ISAAC  C/P,P,H
NEMESIS [@ANGEL]  CIENCIN, SCOTT/CIENCIN, DENISE  P
NEMESIS [@INDIGO #1]  COOPER, LOUISE  C/P,SFWA-P,P
NEMESIS  FANE, BRON  SR-P
NEMESIS [@RESIDENT EVIL #5]  PERRY, S. D.  P
NEMESIS [@MAGIC: THE GATHERING MASQUERADE 2]  THOMPSON, PAUL B.  P
NEMESIS [STAR TREK TNG]  VORNHOLT, JOHN  LP
NEMESIS FROM TERRA, THE  BRACKETT, LEIGH  DF123,P,SFWA-TD8,TD8
NEMESIS MISSION, THE  ING, DEAN  H:2
NEMESIS OF EVIL, THE [ZARKON 1]  CARTER, LIN  P,SFWA
NEMO  GOULART, RON  C/P,P
NEMYDIA DEEP [@TIMEQUEST #3]  TEDFORD, WILLIAM  P
NEON LOTUS  LAIDLAW, MARC  C/P,P
NEPTUNE CROSSING [CHAOS CHRONICLES 1]  CARVER, JEFFREY A.  C/P:2,P,H
NEPTUNE’S CAULDRON  CONEY, MICHAEL G.  P
NERILKA’S STORY [PERN]  MCCAFFREY, ANNE  H:2,P
NERVE  GOLDBERG, MARSHALL  P
NERVES  DEL REY, LESTER  C/P:2,P
NESFA HYMNAL, THE, VOLUME 1  NESFA  [C/RAN-H]
NESFA HYMNAL, THE, VOLUME 1, 2ND EDITION  NESFA  [C/RAN-H:2]
NESFA HYMNAL, THE, VOLUME 2  NESFA  [C/RAN-H:2]
NET, THE  MACGREGOR, LOREN J.  P,SFWA-P
NETHER SCROLL, THE [@FORGOTTEN REALMS LOST EMPIRES]  ABBEY, LYNN  P
NETHERWORLD [WORLD OF DARKNESS VAMPIRE]  BYERS, RICHARD LEE  P
NETS OF SPACE, THE  PETAJA, EMIL  C/P,P
NEUN-PLANETEN-SPIEL, DAS  SHOLS, W. W.  FP
NEUROLINK  BUCKNER, M. M.  P
NEUROMANCER [NEUROMANCER 1]  GIBSON, WILLIAM  H:2,P
NEURON WORLD  FANTHORPE, R. L.  SR-P
NEUTRAL STARS, THE  MORGAN, DAN/KIPPAX, JOHN  C/P:2,P
NEUTRINO DRAG  DI FILIPPO, PAUL  H,LP
NEUTRON SOLSTICE [@DEATHLANDS]  AXLER, JAMES  P
NEUTRON STAR [KNOWN SPACE]  NIVEN, LARRY  C/P:2,SR-P
NEUTRON STARS  FITZGERALD, GREGORY  PA
NEUTRON TWO IS CRITICAL  DUNNING, LAWRENCE  P
NEUTRONIUM ALCHEMIST [NIGHT’S DAWN 2]  HAMILTON, PETER F.  H
NEUTRONIUM ALCHEMIST 1: CONSOLIDATION [NIGHT’S DAWN 2]  HAMILTON, PETER F.  C/P,P
NEUTRONIUM ALCHEMIST 2: CONFLICT [NIGHT’S DAWN 2]  HAMILTON, PETER F.  P
NEVER AFTER  LICKISS, REBECCA  P
NEVER CALL RETREAT [GETTYSBURG 3]  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H
NEVER DEAL WITH A DRAGON [SHADOWRUN 1/SECRETS OF POWER 1]  CHARRETTE, ROBERT N.  C/P:2,P
NEVER IN THIS WORLD  STONE, IDELLA PURNELL  C/PA,PA
NEVER LET ME GO  ISHIGURO, KAZUO  H
NEVER SOUND RETREAT [@LOST REGIMENT #6]  FORSTCHEN, WILLIAM R.  P
NEVER THE TWAIN  MITCHELL, KIRK  C/P,P
NEVER TRUST A DEAD MAN  VELDE, VIVIAN VANDE  H
NEVER TRUST A MARTIAN  FAIRMAN, PAUL W.  B
NEVER TRUST AN ELF [SHADOWRUN 6]  CHARRETTE, ROBERT N.  P
NEVERENDING STORY, THE  ENDE, MICHAEL  H:2,LP
NEVERMORE AFFAIR, THE  WILHELM, KATE  C/P,P
NEVERNESS [REQUIEM FOR HOMO SAPIENS 1]  ZINDELL, DAVID  C/P:2,P,H
NEVERNEVER [BORDERLANDS 6]  SHETTERLY, WILL  C/P,P,H
NEVERWHERE  CORBEN, RICHARD  CX
NEVERWHERE  GAIMAN, NEIL  C/P:2,P
NEVERYONA  DELANY, SAMUEL R.  H:2,LP
NEVLO [@AVENGER #17]  ROBESON, KENNETH  C/P,P
NEW ADAM, THE  WEINBAUM, STANLEY G.  C/P:2,P
NEW ADVENTURES OF THE MAD SCIENTISTS’ CLUB, THE  BRINLEY, BERTRAND R.  P
NEW AMAZONS  WEIS, MARGARET  C/PA,PA
NEW AMERICA [RUSTUM]  ANDERSON, POUL  C/P:2,P
NEW AMERICAN REVIEW 15  SOLOTAROFF, THEODORE  PA

The MITSFS Titledex  September 24, 2007  Page 244
TITLE  AUTHOR  CODE
NEW AMSTERDAM  BEAR, ELIZABETH  H
NEW ARRIVALS, OLD ENCOUNTERS  ALDISS, BRIAN W.  H,SFWA,P
NEW ATLANTEANS, THE  ARETON, EMIL COHEN  LP
NEW ATLANTIS, THE  LE GUIN, URSULA K.  SFWA-TD13,TD13
NEW ATLANTIS, THE  SILVERBERG, ROBERT  C/PA,PA
NEW ATOMS’ BOMBSHELL, THE  BROWNE, ROBERT  C/P,P

NEW AWARENESS: RELIGION THROUGH SCIENCE FICTION, THE WARRICK, PATRICIA S./

GREENBERG, MARTIN H.

HA

NEW BARBARIANS [PROCURATOR 2]  MITCHELL, KIRK  P
NEW CONSTELLATIONS  DISCH, THOMAS M./NAYLOR, CHARLES  HA,SFWA
NEW DAWN, A  CAMPBELL, JOHN W. JR.  H
NEW DESTINIES 1  BAEN, JAMES  PA,C/PA
NEW DESTINIES 2  BAEN, JAMES  PA,C/PA
NEW DESTINIES 3  BAEN, JAMES  C/PA,PA,SFWA-P
NEW DESTINIES 4  BAEN, JAMES  PA,C/PA,SFWA-P
NEW DESTINIES 6  BAEN, JAMES  PA,C/PA,SFWA-P
NEW DESTINIES 7  BAEN, JAMES  PA,C/PA,SFWA-P
NEW DESTINIES 8  BAEN, JAMES  C/PA,PA,SFWA-P:2
NEW DESTINIES 9  BAEN, JAMES  C/PA,PA
NEW DIMENSIONS 1  SILVERBERG, ROBERT  C/PA,PA
NEW DIMENSIONS 2  SILVERBERG, ROBERT  C/PA,PA,SFWA
NEW DIMENSIONS 3  SILVERBERG, ROBERT  C/PA,PA
NEW DIMENSIONS 4  SILVERBERG, ROBERT  C/PA,PA
NEW DIMENSIONS 5  SILVERBERG, ROBERT  HA,SFWA
NEW DIMENSIONS 6  SILVERBERG, ROBERT  HA,SFWA
NEW DIMENSIONS 7  SILVERBERG, ROBERT  HA,SFWA
NEW DIMENSIONS 8  SILVERBERG, ROBERT  HA,SFWA
NEW DIMENSIONS 9  SILVERBERG, ROBERT  HA,SFWA
NEW DIMENSIONS 9 (PROOF)  SILVERBERG, ROBERT  SR-LPA
NEW DIMENSIONS 10  SILVERBERG, ROBERT  HA
NEW DIMENSIONS 11  SILVERBERG, ROBERT/RANDALL, MARTA  C/PA,PA
NEW DIMENSIONS 12  RANDALL, MARTA/SILVERBERG, ROBERT  PA,C/PA
NEW DINOSAURS, THE  DIXON, DOUGAL  A
NEW DREAMS FOR OLD  RESNICK, MIKE  LP
NEW DREAMS THIS MORNING  BLISH, JAMES  PA,C/PA
NEW ENCYCLOPEDIA OF SCIENCE FICTION, THE  GUNN, JAMES E.  R/REF-H
NEW ENGLAND RESISTANCE, THE [V]  SULLIVAN, TIM  C/P,P
NEW ENGLAND’S MARVELOUS MONSTERS  CAHILL, ROBERT ELLIS  [C/RAN-H]
NEW ENGLAND’S THINGS THAT GO BUMP IN THE NIGHT  CAHILL, ROBERT ELLIS  [C/RAN-H]
NEW ENGLAND’S WITCHES AND WIZARDS  CAHILL, ROBERT ELLIS  [C/RAN-H]
NEW FRONTIERS OF THE MIND  RHINE, J. B.  [C/RAN-H]
NEW GULLIVER, THE  DODDERIDGE, ESME  LP

NEW HANDBOOK OF THE HEAVENS BERNHARD, HUBERT J./

BENNETT, DOROTHY A./RICE, HUGH S.

[C/RAN-P:2]

NEW HOPE, A [STAR WARS MOVIE 4]  LUCAS, GEORGE  P
NEW HUGO WINNERS, THE  ASIMOV, ISAAC  HA,PA
NEW HUGO WINNERS 2, THE  ASIMOV, ISAAC  HA,PA,SFWA-P
NEW HUGO WINNERS 3, THE  WILLIS, CONNIE  PA
NEW HUGO WINNERS 4, THE  GREENBERG, MARTIN H.  C/PA,PA
NEW IMPROVED SUN, THE  DISCH, THOMAS M.  SFWA
NEW LIFE, THE  COLERIDGE, JOHN  B
NEW LIFE FOR OLD [RACK 5]  JAMES, LAURENCE  SR-P
NEW LOVECRAFT CIRCLE, THE  PRICE, ROBERT M.  LPA
NEW MAPS OF HELL  AMIS, KINGSLEY  R/REF-P,C/REF-H:2
NEW MERRIAM-WEBSTER POCKET DICTIONARY, THE  MERRIAM-WEBSTER  DICT
NEW MIND, THE [FRONTIERS 2]  ELWOOD, ROGER  C/PA,PA
NEW MINDS, THE  MORGAN, DAN  C/P,P
NEW MOON  RELLING, WILLIAM JR.  P
NEW MOON [QUEEN’S QUARTER 1]  SNYDER, MIDORI  C/P:2,P
NEW MORNING DRAGON  HENSTELL, DIANA  P

NEW NATURE OF THE CATASTROPHE, THE (OMNIBUS EDITION)

[CORNELIUS] JONES, LANGDON/MOORCOCK, MICHAEL

PA

NEW NEIGHBORS, THE  PLANTE, EDMUND  P
NEW REBELLION, THE [@STAR WARS]  RUSCH, KRISTINE KATHRYN  C/P,P
NEW SATELLITE, THE  STATTEN, VARGO  SR-P
NEW SCIENCE FICTION BOOKS PUBLISHED IN GREAT BRITAIN 1968, 1969  BISHOP, GERALD  R/IND-F
NEW SETTLEMENT OF OLD SCORES, A  BRUNNER, JOHN  VLP
NEW SKIES  NIELSEN HAYDEN, PATRICK  PA
NEW SOVIET SCIENCE FICTION  COLLIER BOOKS  HA,LPA
NEW SOVIET SCIENCE FICTION  MACMILLAN PUBLISHING  HA
NEW SPACE HANDBOOK  BUCHEIM, ROBERT  [C/RAN-P]
NEW SPACE OPERA, THE  DOZOIS, GARDNER/STRAHAN, JONATHAN  SR-LPA
NEW SPRING [WHEEL OF TIME 0]  JORDAN, ROBERT  H
NEW SPRINGTIME, THE [WINTER 2]  SILVERBERG, ROBERT  C/P,P,H:2
NEW STORIES FROM THE TWILIGHT ZONE  SERLING, ROD  C/P:2,P

TITLE  AUTHOR  CODE
NEW TALES OF SPACE AND TIME  HEALY, RAYMOND J.  C/PA,SR-PA
NEW TALES OF THE CTHULHU MYTHOS  CAMPBELL, RAMSEY  SR-HA,HA
NEW TERRORS 1  CAMPBELL, RAMSEY  SR-PA
NEW TERRORS 2  CAMPBELL, RAMSEY  SR-PA
NEW TESTAMENT, THE (KING JAMES)  GOD  PA
NEW TESTAMENT, THE (REVISED STANDARD)  GOD  HA
NEW TOLKIEN COMPANION, THE  TYLER, J. E. A.  R/REF-H
NEW TOMORROWS, THE  SPINRAD, NORMAN  PA
NEW VISIONS: A COLLECTION OF MODERN SCIENCE FICTION ART, THE  DOUBLEDAY BOOKS  SFWA
NEW VOICES 1  MARTIN, GEORGE R. R.  C/PA,PA
NEW VOICES 2  MARTIN, GEORGE R. R.  C/PA,PA,SFWA-P
NEW VOICES 3  MARTIN, GEORGE R. R.  C/PA,PA
NEW VOICES 4  MARTIN, GEORGE R. R.  C/PA,PA
NEW VOICES IN SCIENCE FICTION  MARTIN, GEORGE R. R.  HA
NEW VOICES IN SCIENCE FICTION  RESNICK, MIKE  PA
NEW VOYAGE TO THE COUNTRY OF THE HOUYHNHNMS, A  HODGART, MATTHEW  H
NEW WOMEN OF WONDER, THE  SARGENT, PAMELA  C/PA,PA
NEW WORLD WRITING 9  MENTOR BOOKS  PA
NEW WORLDS 1  GARNETT, DAVID  HA,SR-LPA
NEW WORLDS 1  MOORCOCK, MICHAEL  SR-PA
NEW WORLDS 2  MOORCOCK, MICHAEL  SR-PA
NEW WORLDS 3  MOORCOCK, MICHAEL  SR-PA
NEW WORLDS 5  MOORCOCK, MICHAEL  SR-PA
NEW WORLDS #5  MOORCOCK, MICHAEL/PLATT, CHARLES  LPA
NEW WORLDS #6  PLATT, CHARLES/BAILEY, HILARY  LPA
NEW WORLDS 9  BAILEY, HILARY  SR-PA
NEW WORLDS 10  BAILEY, HILARY  SR-PA
NEW WORLDS FOR OLD  CARTER, LIN  SR-PA
NEW WORLDS FOR OLD  KETTERER, DAVID  R/REF-P,C/REF-P
NEW WORLDS OF FANTASY  CARR, TERRY  C/PA,PA
NEW WORLDS OF FANTASY #2  CARR, TERRY  PA,C/PA
NEW WORLDS OF FANTASY #3  CARR, TERRY  PA
NEW WORLDS QUARTERLY #1  MOORCOCK, MICHAEL  C/PA,PA
NEW WORLDS QUARTERLY #2  MOORCOCK, MICHAEL  C/PA,PA
NEW WORLDS QUARTERLY #3  MOORCOCK, MICHAEL  PA,C/PA
NEW WORLDS QUARTERLY #4  MOORCOCK, MICHAEL  C/PA,PA
NEW WRITINGS IN SF 1  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 2  CARNELL, JOHN  SR-PA,C/PA
NEW WRITINGS IN SF 3  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 4  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 5  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 6  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 7  CARNELL, JOHN  SR-HA,PA,C/PA
NEW WRITINGS IN SF 8  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 9  CARNELL, JOHN  PA,C/PA
NEW WRITINGS IN SF 10  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 12  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 13  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 14  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 15  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 16  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 17  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 18  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 19  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 20  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 21  CARNELL, JOHN  SR-PA
NEW WRITINGS IN SF 22  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 23  BULMER, KENNETH  SR-PA,C/PA
NEW WRITINGS IN SF 24  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 25  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 26  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 27  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 28  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 29  BULMER, KENNETH  SR-PA
NEW WRITINGS IN SF 30  BULMER, KENNETH  SR-PA
NEW YORK BY KNIGHT  FRIESNER, ESTHER M.  C/P,P
NEWER YORK  WATT-EVANS, LAWRENCE  PA
NEWMAN FACTOR, THE  PRAGER, J. SIMON  P
NEWS FOR HEAVEN  DELL, JEFFREY  SR-P
NEWS FROM ELSEWHERE  COOPER, EDMUND  C/P:2,P
NEWTON AND THE QUASI-APPLE  SCHMIDT, STANLEY  C/P,P,H
NEWTON’S CANNON [AGE OF UNREASON 1]  KEYES, J. GREGORY  H:2,LP
NEWTON’S WAKE  MACLEOD, KEN  H
NEXT, THE  RANDALL, BOB  P
NEXT, AFTER LUCIFER  RHODES, DANIEL  SFWA-P

The MITSFS Titledex  September 24, 2007  Page 246
TITLE  AUTHOR  CODE
NEXT DOOR TO THE SUN  COBLENTZ, STANTON A.  LP
NEXT ENCOUNTER, THE  THOMPSON, DONALD  C/P,P
NEXT OF KIN  RUSSELL, ERIC FRANK  C/P:2,P
NEXT STOP — MARS!  EDWARDS, DAVID  H
NEXT STOP THE STARS  SILVERBERG, ROBERT  C/P,DF145,SFWA-P
NEXT TEN THOUSAND YEARS, THE  BERRY, ADRIAN  P
NEXUS, THE  MCQUAY, MIKE  C/P,P
NICK OF TIME, THE  EFFINGER, GEORGE ALEC  H,SFWA
NICOJI  BELL, M. SHAYNE  C/P,SFWA-P,P
NICOLAE [@LEFT BEHIND #3]  LAHAYE, TIM/JENKINS, JERRY B.  LP
NIFFT THE LEAN [NIFFT 1]  SHEA, MICHAEL  C/P,P
NIGHT  RODGERS, ALAN  P
NIGHT ARRANT [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
NIGHT BOAT, THE  MCCAMMON, ROBERT R.  P
NIGHT CALLS  KIMBRIEL, KATHERINE ELISKA  P
NIGHT CHILD  BAKER, SCOTT  H,SFWA,P
NIGHT CHILLS  KOONTZ, DEAN R.  H,P
NIGHT CHILLS  MCCAULEY, KIRBY  C/PA,PA
NIGHT CHURCH, THE  STRIEBER, WHITLEY  H
NIGHT FACE, THE  ANDERSON, POUL  C/P:2,P
NIGHT FANTASTIC, THE  ANDERSON, POUL/ANDERSON, KAREN  PA
NIGHT FEAR  LONG, FRANK BELKNAP  C/P,P
NIGHT GALLERY  SERLING, ROD  C/P:2,P
NIGHT GALLERY 2  SERLING, ROD  P
NIGHT GAMES  HARRIS, MARILYN  SFWA-P
NIGHT GIFT, THE  MCKILLIP, PATRICIA A.  H,LP
NIGHT HAS A THOUSAND EYES  IRISH, WILLIAM  SR-P
NIGHT HAWK, THE  BREARLEY, JOHN  B
NIGHT HORSEMAN, THE  BRAND, MAX  P
NIGHT HUNTER  REAVES, MICHAEL  SFWA
NIGHT IN FUNLAND  BRONDFIELD, JEROME  PA
NIGHT IN THE LONESOME OCTOBER, A  ZELAZNY, ROGER  C/P,SFWA,P,H
NIGHT IN THE NETHERHELLS, A [EBENEZUM 3]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
NIGHT LAMP  VANCE, JACK  H
NIGHT LAMP (VIE) [@VANCE INTEGRAL EDITION #42]  VANCE, JACK  SR-H
NIGHT LAND 1, THE  HODGSON, WILLIAM HOPE  P
NIGHT LAND 2, THE  HODGSON, WILLIAM HOPE  C/P,P
NIGHT LAND AND OTHER PERILOUS ROMANCES, THE  HODGSON, WILLIAM HOPE  H
NIGHT LAUNCH  GARN, JAKE/COHEN, STEPHEN PAUL  P
NIGHT LIFE OF THE GODS, THE  SMITH, THORNE  C/P,P,H
NIGHT MAGIC  TRYON, TOM  H
NIGHT MAGIC  TYRON, TOM  H
NIGHT MARE [@XANTH #6]  ANTHONY, PIERS  C/P,P,H
NIGHT MARE  GOLDING, MORTON J.  C/P:2,P
NIGHT MASKS [@FORGOTTEN REALMS CLERIC #3]  SALVATORE, R. A.  P
NIGHT MAYOR, THE  NEWMAN, KIM  P
NIGHT MONSTERS  LEIBER, FRITZ  D30300
NIGHT OF BLOOD [@DRAGONLANCE MINOTAUR WARS #1]  KNAAK, RICHARD A.  P
NIGHT OF DELUSIONS  LAUMER, KEITH  H:2,P
NIGHT OF DELUSIONS (COVER FROM GODEY’S THE MAN IN QUESTION)  LAUMER, KEITH  SR-L
NIGHT OF DRAGONS  SALSITZ, R. A. V.  C/P,P
NIGHT OF FOUR HUNDRED RABBITS, THE  PETERS, ELIZABETH  SFWA-P
NIGHT OF GHOSTS AND LIGHTNING [@PLANET BUILDERS #2]  TALLIS, ROBYN  C/P,P
NIGHT OF GLORY [@ELVEN WAYS #3]  CIENCIN, SCOTT  P
NIGHT OF KADAR, THE  KILWORTH, GARRY  C/P,P
NIGHT OF LIGHT  FARMER, PHILIP JOSE  C/P:2,P
NIGHT OF MADNESS [ETHSHAR]  WATT-EVANS, LAWRENCE  H
NIGHT OF MASKS [DIPPLE 3]  NORTON, ANDRE  C/P,P,H
NIGHT OF MORNINGSTAR, THE [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
NIGHT OF POWER  ROBINSON, SPIDER  C/P,P,H
NIGHT OF TEARS  KIMBRO, JOHN M.  P
NIGHT OF THE APOCALYPSE  EASTERMAN, DANIEL  H
NIGHT OF THE BIG FIRE  MULLER, JOHN E.  SR-P
NIGHT OF THE BIG HEAT  LYMINGTON, JOHN  C/P,P
NIGHT OF THE BLACK HORROR  NORWOOD, VICTOR  SR-P
NIGHT OF THE COOTERS  WALDROP, HOWARD  H,SFWA-P
NIGHT OF THE CRABS [CRABS]  SMITH, GUY N.  P

NIGHT OF THE DRAGONSTAR [DRAGONSTAR 2] BISCHOFF, DAVID F./

MONTELEONE, THOMAS F.

C/P:2,P

NIGHT OF THE EYE [@DRAGONLANCE DEFENDERS OF MAGIC #1]  KIRCHOFF, MARY  C/P,P
NIGHT OF THE LIVING DEAD  RUSSO, JOHN  P
NIGHT OF THE LIVING GATOR! [@MELVINGE 3]  LUPOFF, RICHARD  P
NIGHT OF THE LIVING RAT [MELVINGE 2]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,P
NIGHT OF THE LIVING RERUN [BUFFY]  COVER, ARTHUR BYRON  P

The MITSFS Titledex  September 24, 2007  Page 247
TITLE  AUTHOR  CODE
NIGHT OF THE LIVING SHARK [MELVINGE 1]  BISCHOFF, DAVID F.  P
NIGHT OF THE LONG SHADOWS [@EBERRON INQUISITIVES]  CRILLEY, PAUL  P
NIGHT OF THE PHOENIX [GUARDIANS 4]  AUSTIN, RICHARD  P
NIGHT OF THE RIPPER, THE  BLOCH, ROBERT  H,SFWA
NIGHT OF THE SAUCERS  BINDER, EANDO  C/P:2,P
NIGHT OF THE SPHINX  ELWOOD, ROGER  SR-HA
NIGHT OF THE TRIFFIDS, THE [TRIFFIDS 2]  CLARK, SIMON  P
NIGHT OF THE VAMPIRE  GILES, RAYMOND  P
NIGHT OF THE WEREWOLF  PACKARD, EDWARD  LP
NIGHT OF THE WOLF [WOLF RICE 1]  BORCHARDT, ALICE  P
NIGHT OF THE WOLF  CALLAHAN, JAY  P
NIGHT OF THE WOLF, THE  LEIBER, FRITZ  C/P,P
NIGHT OF THE WOLF, THE  LONG, FRANK BELKNAP  P
NIGHT OF TWO NEW MOONS [@PLANET BUILDERS #6]  TALLIS, ROBYN  P
NIGHT PARADE, THE [FORGOTTEN REALMS HARPERS #4]  CIENCIN, SCOTT  C/P,P
NIGHT PLAGUE  MASTERTON, GRAHAM  SFWA-P
NIGHT PROBE!  CUSSLER, CLIVE  P
NIGHT RELICS  BLAYLOCK, JAMES P.  SFWA
NIGHT RIDE AND OTHER JOURNEYS  BEAUMONT, CHARLES  P
NIGHT SHAPES, THE  BLISH, JAMES  C/P:2,P
NIGHT SHIFT  KING, STEPHEN  C/P,P,H
NIGHT SHOW  LAYMON, RICHARD  SFWA-P
NIGHT-SIDE  OATES, JOYCE CAROL  P
NIGHT SKY MINE  SCOTT, MELISSA  H,LP
NIGHT SLAVES  SOHL, JERRY  C/P:2,P
NIGHT SONGS  GRANT, CHARLES L.  C/P,P
NIGHT SPIDERS, THE  LYMINGTON, JOHN  H,P
NIGHT STALKER, THE  RICE, JEFF  P
NIGHT STRANGLER, THE  RICE, JEFF  C/P,P
NIGHT THINGS  MONTELEONE, THOMAS F.  P
NIGHT TIDE, A  INOUYE, JON  SR-P
NIGHT-TRAIN  MONTELEONE, THOMAS F.  P
NIGHT TRAIN TO RIGEL  ZAHN, TIMOTHY  H:2
NIGHT VISIONS 1  RYAN, ALAN  SR-HA
NIGHT VISIONS 2  GRANT, CHARLES L.  SR-HA
NIGHT VISIONS 3  MARTIN, GEORGE R. R.  SR-HA
NIGHT VISIONS 4  DARK HARVEST  SR-HA
NIGHT VISIONS 5  WINTER, DOUGLAS E.  SR-HA
NIGHT VISIONS 6  DARK HARVEST  SR-HA
NIGHT VISIONS 7  WIATER, STANLEY  SR-HA
NIGHT VISIONS 8  DARK HARVEST  SR-HA
NIGHT VISIONS 9  DARK HARVEST  SR-HA
NIGHT WALK  SHAW, BOB  SR-P,C/P:2
NIGHT WARRIORS  MASTERTON, GRAHAM  SFWA-P
NIGHT WATCH [NIGHTWATCH 1]  LUKYANENKO, SERGEI  H,LP
NIGHT WATCH [DISCWORLD 28/DISCWORLD WATCHMEN 6]  PRATCHETT, TERRY  H:2
NIGHT WATCH, THE [RESURRECTION MAN 3]  STEWART, SEAN  C/P:2,P
NIGHT WE BURIED ROAD DOG, THE  CADY, JACK  H
NIGHT WHISTLERS, THE [@NIGHT WHISTLERS #1]  TREVOR, DAN  P
NIGHT WINDS  WAGNER, KARL EDWARD  C/P:2,P
NIGHT-WORLD  BLOCH, ROBERT  P,SFWA-P
NIGHTBRINGER [WARHAMMER 40K ULTRAMARINES 1]  MCNEILL, GRAHAM  P
NIGHTCHILD [CHRONICLES OF THE RAVEN 3]  BARCLAY, JAMES  LP
NIGHTDREAMERS [DOCTOR WHO]  ARDEN, TOM  SR-H
NIGHTEYES  REEVES-STEVENS, GARFIELD  H:2,SR-LP
NIGHTFALL  ASIMOV, ISAAC/SILVERBERG, ROBERT  C/P:2,P,H
NIGHTFALL AND OTHER STORIES  ASIMOV, ISAAC  H:2,P
NIGHTFALL AT ALGEMRON [STAR*DRIVE HARBINGER 3]  DUANE, DIANE  P

NIGHTFEEDER [GALEN SWORD 2] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

P

NIGHTFLYERS  MARTIN, GEORGE R. R.  BS5,SFWA-P,LP,C/P:2
NIGHTINGALE, THE  DALKEY, KARA  H,SFWA,P
NIGHTINGALE’S LAMENT [NIGHTSIDE 3]  GREEN, SIMON R.  P
NIGHTLIFE [CAL AND NIKO 1]  THURMAN, ROB  P
NIGHTMARE  BRETT, LEO  SR-P
NIGHTMARE  CHILTON, I. M.  P
NIGHTMARE [SILENT EMPIRE 2]  HARPER, STEVEN  P
NIGHTMARE [DOCTOR STRANGE]  ROTSLER, WILLIAM  C/P,P
NIGHTMARE AGE  POHL, FREDERIK  C/PA,PA
NIGHTMARE ASYLUM [@ALIENS #2]  PERRY, STEVE  P
NIGHTMARE BABY  DUBREUIL, LINDA  P
NIGHTMARE BEGINS, THE [@SURVIVALIST #2]  AHERN, JERRY  P

NIGHTMARE BLUE DOZOIS, GARDNER/

EFFINGER, GEORGE ALEC

C/P,P

NIGHTMARE DREAM, THE [@TORG POSSIBILITY WARS #3]  CASPIAN, JONATHA ARIADNE  C/P,P

The MITSFS Titledex  September 24, 2007  Page 248
TITLE  AUTHOR  CODE
NIGHTMARE EXPRESS  HAIBLUM, ISIDORE  C/P,P
NIGHTMARE FACTOR, THE  SCORTIA, THOMAS N./ROBINSON, FRANK M. P
NIGHTMARE FAIR, THE [DOCTOR WHO]  WILLIAMS, GRAHAM  SR-P,P
NIGHTMARE GARDEN  GHIDALIA, VIC  PA
NIGHTMARE IN RED  MARTEN, JACQUELINE  P
NIGHTMARE JOURNEY  KOONTZ, DEAN R.  H,P
NIGHTMARE MACHINE, THE [@BATTLESTAR GALACTICA 1978 #11]  LARSON, GLEN A./THURSTON, ROBERT  P
NIGHTMARE MACHINE, THE [@STAR WARS GALAXY OF FEAR #4]  WHITMAN, JOHN  P
NIGHTMARE OF BLACK ISLAND, THE [@DOCTOR WHO]  TUCKER, MIKE  H
NIGHTMARE OF GOD, THE [ZORACHUS 2]  ROGERS, MARK E.  P
NIGHTMARE ON VEGA 3  HUNTINGTON, CHARLES  C/P:2,P
NIGHTMARE PASSAGE [@DEATHLANDS]  AXLER, JAMES  P
NIGHTMARE PEOPLE, THE  WATT-EVANS, LAWRENCE  LP
NIGHTMARE READER, THE  HAINING, PETER  HA,SFWA
NIGHTMARE SEASONS  GRANT, CHARLES L.  P,SFWA
NIGHTMARE SESSION [@PRIVATE SCHOOL #1]  CHARLES, STEVEN  P
NIGHTMARE WORLD [DR. BONES]  STERN, DAVID  P
NIGHTMARES  BLOCH, ROBERT  P
NIGHTMARES  GRANT, CHARLES L.  PA
NIGHTMARES AND DAYDREAMS  BOND, NELSON  SR-H
NIGHTMARES AND DREAMSCAPES  KING, STEPHEN  C/P,P,H
NIGHTMARES AND GEEZENSTACKS  BROWN, FREDRIC  P
NIGHTMARE’S DISCIPLE  PULVER, JOSEPH S., SR.  LP
NIGHTMARES OF EMINENT PERSONS  RUSSELL, BERTRAND  H
NIGHTMARES ON ELM STREET, THE [ELM STREET 1,2,3]  COOPER, JEFFREY  P
NIGHTPOOL  MURPHY, SHIRLEY ROUSSEAU  H
NIGHTREAVER [BLOODFANG 2]  WEAVER, MICHAEL D.  P
NIGHTRUNNERS, THE  LANSDALE, JOE R.  SFWA-P
NIGHT’S ANGEL  BAILEY, ROBIN WAYNE  LP
NIGHTS AT THE CIRCUS  CARTER, ANGELA  H
NIGHT’S BLACK AGENTS (ABRIDGED)  LEIBER, FRITZ  C/P,P
NIGHT’S BLACK AGENTS (REVISED)  LEIBER, FRITZ  C/P:2,P
NIGHT’S DAUGHTER  BRADLEY, MARION ZIMMER  C/P,P
NIGHT’S DAUGHTER [@FLAT EARTH 1,2]  LEE, TANITH  H:2
NIGHT’S MASTER [@FLAT EARTH #1]  LEE, TANITH  C/P,P
NIGHT’S PAWN [@SHADOWRUN 10]  DOWD, TOM  C/P,P
NIGHT’S SORCERIES [@FLAT EARTH #5]  LEE, TANITH  P
NIGHT’S YAWNING PEAL  DERLETH, AUGUST  PA
NIGHTSEER  HAMILTON, LAURELL K.  C/P:2,P
NIGHTSHADE  BUTLER, JACK  SFWA
NIGHTSHADE [DOCTOR WHO NEW ADVENTURES]  GATISS, MARK  C/P:2,P
NIGHTSHADE [STAR TREK TNG]  HAMILTON, LAURELL K.  C/P:2,P
NIGHTSHADE [WEIRD HEROES 4]  KING, TAPPAN/MEACHAM, BETH  P
NIGHTSHADE AND DAMNATIONS  KERSH, GERALD  C/P:2,P
NIGHTSHADES  LEE, TANITH  P
NIGHTSIDE CITY  WATT-EVANS, LAWRENCE  P
NIGHTSIDE THE LONG SUN [LONG SUN 1]  WOLFE, GENE  H,P
NIGHTSWORD [@STARSHIELD 3]  WEIS, MARGARET/HICKMAN, TRACY  H

NIGHTTOUCH: JOURNEYING INTO THE REALMS OF NIGHTMARE GOLDBERG, GERRY/STOROSCHUK, STEPHEN/

CORBETT, FRED

VLPA

NIGHTWALKER, THE  TESSIER, THOMAS  P
NIGHTWATCH [GREYHAWK]  BAILEY, ROBIN WAYNE  C/P,P
NIGHTWATCH  STEPHENSON, ANDREW M.  C/P,P
NIGHTWING  SMITH, MARTIN CRUZ  C/P:2,P

NIGHTWINGS SILVERBERG, ROBERT C/P:2,TD15,SFWA-TD15,

P

NIGHTWITCH DEVIL, THE [@AVENGER #29]  ROBESON, KENNETH  C/P:2,P
NIGHTWORLD  BISCHOFF, DAVID F.  C/P:2,P
NIGHTWORLD [ADVERSARY 6]  WILSON, F. PAUL  P
NILE GOLD  KNITTEL, JOHN  H

NIMBLE MAN, THE [@MENAGERIE] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

NIMBUS  JABLOKOV, ALEXANDER  H:2,SFWA,P
NIMISHA’S SHIP  MCCAFFREY, ANNE  C/P:2,P,H
NIMROD HUNT, THE  SHEFFIELD, CHARLES  C/P:2,P
NINE BILLION NAMES OF GOD, THE  CLARKE, ARTHUR C.  C/P:2,P
NINE BY LAUMER  LAUMER, KEITH  C/P:2,P
NINE GATES, THE  BRUGALETTE, PHILLIP  P
NINE GEES  PHILLIES, GEORGE  LP
NINE HORRORS AND A DREAM  BRENNAN, JOSEPH PAYNE  P
NINE HUNDRED GRANDMOTHERS  LAFFERTY, R. A.  C/P:2,P
NINE LAYERS OF SKY  WILLIAMS, LIZ  P
NINE LAYERS OF SKY (PROOF)  WILLIAMS, LIZ  SR-P
NINE LIVES  FAYLE, HELEN/COX, SUE  SR-LPA

The MITSFS Titledex  September 24, 2007  Page 249
TITLE  AUTHOR  CODE
NINE LIVES  LE GUIN, URSULA K.  B
NINE LIVES OF CATSEYE GOMEZ, THE [WYRDRUNE]  HAWKE, SIMON  C/P,P
998  HYAMS, EDWARD  H
NINE PLANETS  NOURSE, ALAN E.  [C/RAN-P]
NINE PRINCES IN AMBER [AMBER 1]  ZELAZNY, ROGER  C/P:2,@P
NINE QUESTIONS, THE  FENTON, EDWARD  H
NINE RITTENHOUSE SQUARE  WYLIE, PHILIP  P
NINE STRANGE STORIES  OWEN, BETTY M.  C/PA,PA
9 TALES OF SPACE AND TIME  HEALY, RAYMOND J.  SR-HA
NINE TIMES NINE  BOUCHER, ANTHONY  P
NINE TOMORROWS  ASIMOV, ISAAC  H:2,P
NINE UNKNOWN, THE  MUNDY, TALBOT  SR-P,C/P,H
1985  BURGESS, ANTHONY  H
1985  DALOS, GYORGY  H:2,LP
NINETEEN EIGHTY-FOUR  ORWELL, GEORGE  C/P,P,H
1984: SPRING — A CHOICE OF FUTURES  CLARKE, ARTHUR C.  [C/RAN-P]
NINETEEN EIGHTY-NINE: POPULATION DOOMSDAY  PENDLETON, DON  P
1953 CHECKDEX, THE  RIDDLE, CHARLES LEE  SRL/IND
1945  GINGRICH, NEWT/FORSTCHEN, WILLIAM R. H,SFWA,P
1999: THE WORLD OF TOMORROW  CORNISH, EDWARD  [C/RAN-H]
1901  CONROY, ROBERT  P
1968  HALDEMAN, JOE  P
19 TALES OF TERROR  BURNETT, WHIT/BURNETT, HALLIE  PA
98.4  HODDER-WILLIAMS, CHRISTOPHER  SR-P
97TH STEP, THE [MATADOR 0]  PERRY, STEVE  C/P:2,SFWA-P,P
NINETY TRILLION FAUSTS, THE [QUINTARA 3]  CHALKER, JACK L.  H:2,SFWA-P,P
NINTH TALISMAN, THE [CHOSEN 2]  WATT-EVANS, LAWRENCE  H
NITPICKER’S GUIDE FOR CLASSIC TREKKERS, THE [STAR TREK]  FARRAND, PHIL  R/REF-H

NITPICKER’S GUIDE FOR NEXT GENERATION TREKKERS, THE [STAR TREK TNG] FARRAND, PHIL

R/REF-H

NITROGEN FIX, THE  CLEMENT, HAL  C/P,P,H
NIVEN’S LAWS  NIVEN, LARRY  SR-H
NO BLADE OF GRASS  CHRISTOPHER, JOHN  H:2,LP
NO BLOOD SPILLED [DON SEBASTIAN 5]  DANIELS, LES  P,SFWA-P
NO BOUNDARIES  KUTTNER, HENRY/MOORE, C. L.  SR-H,C/P:2,P
NO BROTHER, NO FRIEND [ERIC MATHERS 2]  MEREDITH, RICHARD C.  C/P,P,H
NO CLOCK IN THE FOREST  WILLIS, PAUL J.  P,SFWA-P
NO DAWN AND NO HORIZON  MERAK, A. J.  H
NO DEALS, MR. BOND [JAMES BOND 20]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
NO DIRECTION HOME  SPINRAD, NORMAN  C/P:2,P
NO DOMINION  HUSTON, CHARLIE  LP
NO DOORS, NO WINDOWS  ELLISON, HARLAN  P
NO EARTHLY SHORE [@AREIA DARENGA #3]  MEZO, FRANCINE  C/P,P
NO EARTHLY SUNNE  BALL, MARGARET  C/P,P
NO ENEMY BUT TIME  BISHOP, MICHAEL  SFWA-P,H:2,P,SFWA
NO FLYING IN THE HOUSE  BROCK, BETTY  P
NO-FRILLS BOOK: SCIENCE FICTION  ANONYMOUS  C/P,P
NO FUTURE [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
NO FUTURE FOR DRAGONS [DRAGON WARS 3]  GENTILE, GARY  P
NO FUTURE IN IT  BRUNNER, JOHN  C/P,SR-P,H
NO HUMANS INVOLVED [WOMEN OF THE OTHERWORLD]  ARMSTRONG, KELLEY  H
NO LIMITS [@STAR TREK NEW FRONTIER]  DAVID, PETER  LPA
NO LIMITS  FERMAN, JOSEPH W.  PA,C/PA
NO LIMITS [@VIRTUAL WORLD]  FINDLEY, NIGEL D.  P
NO MAN ON EARTH  MOUDY, WALTER  C/P:2,P
NO MAN’S LAND [STAR TREK GATEWAYS #5/STAR TREK VOYAGER]  GOLDEN, CHRISTIE  P
NO MAN’S WORLD  BULMER, KENNETH  DF104
NO MAN’S WORLD  CAIDIN, MARTIN  H
NO MAS DUENDES  DICKSON, GORDON R./BOVA, BEN  FP
NO MIND OF MAN  HAWTHORN BOOKS  HA
NO MIND OF MAN  SILVERBERG, ROBERT  C/PA,PA
NO NIGHT WITHOUT STARS  NORTON, ANDRE  H:2,P
NO ONE GOES THERE NOW  WALLING, WILLIAM H.  H,SFWA
NO ONE NOTICED THE CAT  MCCAFFREY, ANNE  H:2
NO OTHER MAN  NOYES, ALFRED  H
NO PHULE LIKE AN OLD PHULE [PHULE 5]  ASPRIN, ROBERT LYNN/HECK, PETER J.  P
NO PLACE LIKE EARTH  WYNDHAM, JOHN  SR-H
NO PLACE ON EARTH  CHARBONNEAU, LOUIS  H:2,P
NO QUARTER [QUARTER SINGERS 3]  HUFF, TANYA  C/P:2,P
NO REFUGE  BOLAND, JOHN  SR-P
NO RENT TO PAY  JACKSON, ANITA  SR-P
NO REST FOR THE WICKED [IMMORTALS AFTER DARK 2]  COLE, KRESLEY  H
NO ROOM FOR MAN [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
NO SAFE PLACE  MOROZ, ANNE  C/P,P

The MITSFS Titledex  September 24, 2007  Page 250
TITLE  AUTHOR  CODE
NO STARS FOR US  RAY, ROBERT  SR-P

NO SURRENDER [@STAR TREK S.C.E. #4] COLLINS, MIKE/

EDGINTON, IAN/GREENBARGER, ROBERT/ HAUMAN, GLENN/MARIOTTE, JEFF

P

NO TIME FOR HEROES  LUNDWALL, SAM J.  D58880
NO TIME LIKE THE FUTURE  BOND, NELSON  C/P:2,P
NO TIME LIKE TOMORROW  ALDISS, BRIAN W.  C/P:2,P
NO TIME LIKE TOMORROW  WHITE, TED  H
NO TRUCE WITH KINGS  ANDERSON, POUL  TD5
NO TRUCE WITH TERRA  HIGH, PHILIP E.  DF275
NO WAY BACK  ZEIGFREID, KARL  SR-P
NO WORLD OF THEIR OWN  ANDERSON, POUL  P,DD110
NOAH II  DIXON, ROGER  C/P,P,H
NOBLEST EXPERIMENT IN THE GALAXY, THE  TRIMBLE, LOUIS  D11560
NOBODY GETS THE GIRL  MAXEY, JAMES  LP
NOBODY LIVES FOREVER [JAMES BOND 19]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
NOBODY’S SON  STEWART, SEAN  P
NOCTULPA NO. 4  HATCH, GEORGE  LPA
NOCTURNE [@INDIGO #4]  COOPER, LOUISE  C/P,SFWA-P,P
NOCTURNE (NOVEL EXCERPTS)  HARLEQUIN BOOKS S. A.  PA
NOCTURNE FOR A DANGEROUS MAN  MATZ, MARC  P
NOIR  JETER, K. W.  P
NOISE  CLEMENT, HAL  H
NOMAD  ALEXANDER, DAVID  P
NOMAD, THE [@DARK SUN TRIBE OF ONE #3]  HAWKE, SIMON  P
NOMAD OF TIME, THE  MOORCOCK, MICHAEL  H:2
NOMADS  YARBRO, CHELSEA QUINN  C/P,P
NOMADS OF GOR [GOR 4]  NORMAN, JOHN  C/P:2,P
NOMADS OF THE SKY [BUCK ROGERS INVADERS OF CHARON 2]  KEITH, WILLIAM H. JR.  P
NON-STATISTICAL MAN, THE  JONES, RAYMOND F.  C/P:2,P
NON-STOP  ALDISS, BRIAN W.  H:2,P
NONBORN KING, THE [PLIOCENE EXILE 3]  MAY, JULIAN  C/P:2,SFWA,P

NONBORN KING, THE [PLIOCENE EXILE 3,4]/ ADVERSARY, THE [PLIOCENE EXILE 3,4] MAY, JULIAN

H:2

NONE BUT MAN  DICKSON, GORDON R.  C/P,P,H
NONE BUT MAN (WITH ADDITIONAL STORY)  DICKSON, GORDON R.  P
NONE SO BLIND  HALDEMAN, JOE  H
NONINTERFERENCE  TURTLEDOVE, HARRY  C/P,P
NOON: 22ND CENTURY  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H,LP
NOONAN  FISHER, LEONARD EVERETT  H
NOONBLAZE  CHIBA, MILAN  P
NOONSHADE [CHRONICLES OF THE RAVEN 2]  BARCLAY, JAMES  LP
NOOSE REPORT  HITCHCOCK, ALFRED  PA
NOR CRYSTAL TEARS [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
NOR OF HUMAN...  MALONEY, GEOFFREY  LPA
NORBY AND THE COURT JESTER [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  P
NORBY AND THE OLDEST DRAGON [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  SFWA-P
NORBY AND YOBO’S GREAT ADVENTURE [NORBY 4]  ASIMOV, JANET/ASIMOV, ISAAC  C/P,SFWA-P:2,P
NORBY CHRONICLES, THE [NORBY 1]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:2,P
NORBY DOWN TO EARTH [NORBY 5]  ASIMOV, JANET/ASIMOV, ISAAC  P,SFWA-P:2
NORBY: ROBOT FOR HIRE [NORBY 2]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:3,P
NORBY THROUGH TIME AND SPACE [NORBY]  ASIMOV, JANET/ASIMOV, ISAAC  C/P:2,P
...NORDPOL RUFT ATLANTIS...  BERNING, FRANK W.  FP
NORMAN CONQUEST 2066  MCINTOSH, J. T.  SR-P
NORSTRILIA [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P

NORSTRILIA (COMPLETE AND CORRECTED)

[INSTRUMENTALITY OF MANKIND] SMITH, CORDWAINER

SR-H

NORTH CAPE  POYER, JOE  H
NORTH DIMENSION  ROGERSOHN, WILLIAM  SR-P
NORTH WIND [ALEUTIAN 2]  JONES, GWYNETH  H,LP
NORTHERN GIRL, THE [TORNOR 3]  LYNN, ELIZABETH A.  C/P:2,SFWA,P
NORTHERN STARS  HARTWELL, DAVID G./GRANT, GLENN  HA
NORTHERN STORM [@ALDABRESHIN COMPASS #2]  MCKENNA, JULIET E.  P
NORTHERN SUNS  HARTWELL, DAVID G./GRANT, GLENN  LPA
NORTHLIGHT  WHEELER, DEBORAH  P
NORTHSHORE [@AWAKENERS #1]  TEPPER, SHERI S.  C/P,SFWA,SFWA-P,P
NORTHSTAR RISING [DEATHLANDS]  AXLER, JAMES  P
NORTHWEST SMITH  MOORE, C. L.  C/P,P
NORTHWORLD [NORTHWORLD 1]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
NORTHWORLD JUSTICE [NORTHWORLD 3]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
NORTHWORLD TRILOGY [NORTHWORLD 1,2,3]  DRAKE, DAVID  C/P,P
NORTHWORLD VENGEANCE [NORTHWORLD 2]  DRAKE, DAVID  C/P,SFWA-P:2,P
NOSFERATU [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  FLEMING, GHERBOD  P
NOSFERATU [SHADOWRUN 14]  SARGENT, CARL/GASCOIGNE, MARC  C/P,P

TITLE  AUTHOR  CODE
NOSFERATU THE VAMPYRE  MONETTE, PAUL  P
NOSTRADAMUS ATE MY HAMSTER  RANKIN, ROBERT  P

NOT EXACTLY THE THREE MUSKETEERS [GUARDIANS OF THE FLAME CASTLE CULLINANE 1] ROSENBERG, JOEL

H

NOT FOR GLORY  ROSENBERG, JOEL  H:2,SFWA-P,P
NOT FORGOTTEN [@ANGEL]  HOLDER, NANCY  P
NOT IN OUR STARS  SWYCAFFER, JEFFERSON P.  C/P,P
NOT IN SOLITUDE  GANTZ, KENNETH F.  H:2,P
NOT LONG FOR THIS WORLD  DERLETH, AUGUST  H,P
NOT-MEN, THE  WILLIAMSON, JACK  P
NOT OF WOMAN BORN  ASH, CONSTANCE  PA

NOT QUITE SCARAMOUCHE [GUARDIANS OF THE FLAME CASTLE CULLINANE 2] ROSENBERG, JOEL

H

NOT REALLY THE PRISONER OF ZENDA [GUARDIANS OF THE FLAME CASTLE CULLINANE 3] ROSENBERG, JOEL

H

NOT THIS AUGUST  KORNBLUTH, C. M.  H:2,P
NOT THIS AUGUST (REVISED)  KORNBLUTH, C. M.  P,SFWA-P
NOT TO MENTION CAMELS  LAFFERTY, R. A.  H
NOT TOO NARROW...NOT TOO DEEP  SALE, RICHARD  SR-P
NOT WANTED ON THE VOYAGE  FINDLEY, TIMOTHY  H
NOT WITH A BANG  PINCHER, CHAPMAN  H:3,P
NOT WITHOUT SORCERY  STURGEON, THEODORE  C/P:2,P
NOT-WORLD, THE  SWANN, THOMAS BURNETT  C/P:2,P
NOTES FROM THE FUTURE  AMOSOFF, N.  H
NOTES TO A SCIENCE FICTION WRITER  BOVA, BEN  R/REF-H
NOTHING BURNS IN HELL  FARMER, PHILIP JOSE  P
NOTHING BUT BLUE SKIES  HOLT, TOM  P
NOTHING HUMAN  KRESS, NANCY  H
NOTHING LIKE THE SUN  BURGESS, ANTHONY  C/P,P
NOTHING SACRED  SCARBOROUGH, ELIZABETH ANN  H:2,LP,P
NOTIONS: UNLIMITED  SHECKLEY, ROBERT  C/P:2,P
NOVA [@STARCRAFT GHOST]  DECANDIDO, KEITH R. A.  P
NOVA  DELANY, SAMUEL R.  H:2,P
NOVA 1  HARRISON, HARRY  C/PA,PA
NOVA 2  HARRISON, HARRY  HA,PA
NOVA 3  HARRISON, HARRY  HA,SFWA
NOVA 4  HARRISON, HARRY  C/PA,PA,SFWA
NOVA COMMAND [@STAR TREK TNG STARFLEET ACADEMY #9]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
NOVA EXPRESS  BURROUGHS, WILLIAM S.  H
NOVA SWING [LIGHT 2]  HARRISON, M. JOHN  H
NOVEL IDEAS: FANTASY  THOMSEN, BRIAN M.  PA
NOVEL SAMPLER (CD)  RUGGIERO, TONY  SR-MM:2
NOVELETS OF SCIENCE FICTION  HOWARD, IVAN  PA,C/PA
NOVELTIES & SOUVENIRS  CROWLEY, JOHN  LP
NOVELTY  CROWLEY, JOHN  SR-LP
NOVICE, THE [BLACK MAGICIAN 2]  CANAVAN, TRUDI  C/P,P
NOW AND BEYOND  BELMONT BOOKS  PA,C/PA
NOW BEGINS TOMORROW  KNIGHT, DAMON  C/PA,PA
NOW COMES TOMORROW  WILLIAMS, ROBERT MOORE  P
NOW, THE GODS  FRENCH, ROBERT  P
NOW THEN!  BRUNNER, JOHN  C/P:2,P
NOW WAIT FOR LAST YEAR  DICK, PHILIP K.  H:2,P

NOW YOU SEE HER, NOW YOU DON’T [@SABRINA THE TEENAGE WITCH #16] GALLAGHER, DIANA G.

P

NOW YOU SEE IT/HIM/THEM...  DEWEESE, GENE/COULSON, ROBERT  H,SFWA
NOWHERE HUNT, THE [DIADEM 6]  CLAYTON, JO  C/P:2,P
NOWHERE ON EARTH  ELDER, MICHAEL  C/P,P
NOWHERE PLACE, THE  LYMINGTON, JOHN  H:2,SFWA
NTH DOCTOR, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
NUKES  MACLAY, JOHN  SFWA
NULL-A THREE  VAN VOGT, A. E.  C/P:2,P
NULL-A WORLDS OF A. E. VAN VOGT, THE  DRAKE, H. L.  R/REF-P
NULL-FREQUENCY IMPULSER, THE  COLEMAN, JAMES NELSON  C/P:2,P
NUL’S QUEST  STRICKLAND, BRAD  P
NUMBER OF THE BEAST, THE  HEINLEIN, ROBERT A.  C/P,H,LP
NUMBERLAND  WEINBERG, GEORGE  H
NUMBERS DON’T LIE  BISSON, TERRY  LP
NUNQUAM  DURRELL, LAWRENCE  C/P,P
NUTRO 29  NORRIS, FRANK  H
NUTZENBOLTS AND MORE TROUBLES WITH MACHINES  GOULART, RON  H:2,SFWA
NYARLATHOTEP CYCLE, THE  PRICE, ROBERT M.  LPA
NYLON ANGEL [PARRISH PLESSIS 1]  DE PIERRES, MARIANNE  P
NYMPHOMATION  NOON, JEFF  P
NYPD 2025  STRYKER, HAL  C/P,P

The MITSFS Titledex  September 24, 2007  Page 253
TITLE  AUTHOR  CODE
O GREENEST BRANCH! [@WATER! #1]  BAUDINO, GAEL  P
O MASTER CALIBAN! [BARRAZAN V 1]  GOTLIEB, PHYLLIS  C/P,SFWA,P,H
O PIONEER!  POHL, FREDERIK  H
O-ZONE  THEROUX, PAUL  H,P
OAK ABOVE THE KINGS, THE [KELTIAD ARTHUR 2]  KENNEALY-MORRISON, PATRICIA  C/P,SFWA,P
OAK AND THE RAM, THE [CORUM 5]  MOORCOCK, MICHAEL  C/P,P,H
OAKDALE AFFAIR, THE  BURROUGHS, EDGAR RICE  C/P:2,P
OAKEN THRONE, THE [@DEPTFORD HISTORIES #2]  JARVIS, ROBIN  LP
OASIS  HODGE, BRIAN  SFWA-P
OATH AND THE MEASURE, THE [@DRAGONLANCE MEETINGS 4]  WILLIAMS, MICHAEL  P
OATH OF FEALTY  NIVEN, LARRY/POURNELLE, JERRY  C/P,P,H:2
OATH OF GOLD [PAKSENARRION 3]  MOON, ELIZABETH  C/P:2,P
OATH OF NERULL [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
OATH OF STONEKEEP, THE  DENNING, TROY  P
OATH OF SWORDS [HRADANI 2]  WEBER, DAVID  C/P,P
OATH OF THE RENUNCIATES [DARKOVER]  BRADLEY, MARION ZIMMER  H
OATH TO MIDA, AN [JALAV 2]  GREEN, SHARON  C/P,P
OATHBLOOD [VOWS AND HONOR 3]  LACKEY, MERCEDES  C/P:2,P
OATHBOUND, THE [VOWS AND HONOR 1]  LACKEY, MERCEDES  C/P:2,P
OATHBOUND WIZARD, THE [WIZARD IN RHYME 2]  STASHEFF, CHRISTOPHER  C/P,P,H
OATHBREAKERS [VOWS AND HONOR 2]  LACKEY, MERCEDES  C/P:2,P
OATHS AND MIRACLES  KRESS, NANCY  C/P,P
OBELISK  JONES, JUDITH KAYE  P
OBERNEWTYN [OBERNEWTYN 1]  CARMODY, ISOBELLE  P
OBERNEWTYN & THE FARSEEKERS [OBERNEWTYN 1,2]  CARMODY, ISOBELLE  H
OBJECTIF TAMAX  RANDA, PETER  FP
OBJECTIVE: BAJOR [STAR TREK DS9]  PEEL, JOHN  C/P,P
OBLIQUE APPROACH, AN [BELISARIUS 1]  FLINT, ERIC/DRAKE, DAVID  C/P,P
OBLIVION  BONANSINGA, JAY  SR-H
OBLIVION [@STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
OBLIVION [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  STONE, DAVE  C/P,P
OBLIVION TAPES, THE  MURARI, TIMERI  P
OBSERVERS, THE [CV 2]  KNIGHT, DAMON  H,SFWA-P
OBSESSION  CAMPBELL, RAMSEY  P
OBSESSION OF SALLY WING, THE  MARTIN, RUSS  SFWA-P
OBSESSIONS [@QUANTUM LEAP]  DAVIS, CAROL  P
OBSIDIAN ALLIANCES [STAR TREK MIRROR]  POCKET BOOKS  LPA
OBSIDIAN APE, THE  LEATH, ROBERT NEAL  VLP
OBSIDIAN BUTTERFLY [ANITA BLAKE 9]  HAMILTON, LAURELL K.  C/P,P,H
OBSIDIAN FATE [@BUFFY]  GALLAGHER, DIANA G.  C/P,P
OBSIDIAN KEY, THE [ASAHIEL 2]  THOMPSON, ELDON  H
OBSIDIAN ORACLE, THE [@DARK SUN PRISM #4]  DENNING, TROY  P
OBSIDIAN TOWER, THE [JEWELFIRE 3]  WARRINGTON, FREDA  P
OCCAM’S RAZOR  DUNCAN, DAVID  C/P,P,H
OCCULT DETECTIVES OF C.J. HENDERSON, THE  HENDERSON, C. J.  LP
OCCULT LOVECRAFT, THE  LOVECRAFT, H. P./RAVEN, ANTHONY  R/REF-P
OCEAN IN ALL ITS DEVICES, THE  SPENCER, WILLIAM BROWNING  SR-H
OCEAN OF YEARS, THE [SOLACE 2]  ALLEN, ROGER MACBRIDE  P
OCEAN ON TOP  CLEMENT, HAL  C/P:2,P

OCEAN UNDER THE ICE [ROCHEWORLD 4] FORWARD, ROBERT L./

FORWARD, MARTHA DODSON

C/P,P

OCEANS DU CIEL, LES  STEINER, KURT  FP

OCEANS OF MAGIC THOMSEN, BRIAN M./

GREENBERG, MARTIN H.

C/PA,PA

OCEANS OF SPACE THOMSEN, BRIAN M./

GREENBERG, MARTIN H.

PA

OCEANSPACE  STEELE, ALLEN  H:2,P
OCTAGON  SABERHAGEN, FRED  C/P:2,P
OCTAGON MAGIC  NORTON, ANDRE  H:2
OCTAGONAL RAVEN, THE  MODESITT, L. E. JR.  H,P
OCTOBER COUNTRY, THE  BRADBURY, RAY  C/P:2,P
OCTOBER DREAMS  CHIZMAR, RICHARD T./MORRISH, ROBERT  LPA
OCTOBER SCENARIO, THE  RANDLE, KEVIN  [C/RAN-P]
OCTOBER THE FIRST IS TOO LATE  HOYLE, FRED  H:2,P
OCTOBERLAND [@DOMINIONS OF IRTH #3]  LEE, ADAM  H,P
OCTOBER’S BABY [DREAD EMPIRE 2/RAGNARSON 4]  COOK, GLEN  C/P:2,P
OCTOBER’S GHOST  PEARSON, RYNE DOUGLAS  P
OCTOPUS AGENDA, THE [@SPIDERMAN]  DUANE, DIANE  H
OD MAGIC  MCKILLIP, PATRICIA A.  H:2
ODALISQUE [PERCHERON 1]  MCINTOSH, FIONA  LP
ODD JOB #101  GOULART, RON  H,SFWA
ODD JOHN  STAPLEDON, OLAF  C/P:2,P,GN2.08
ODD JOHN AND SIRIUS  STAPLEDON, OLAF  H,LP
ODD SCIENCE FICTION  LONG, FRANK BELKNAP  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 254
TITLE  AUTHOR  CODE
ODD THOMAS  KOONTZ, DEAN R.  H
ODD WARLOCK OUT [WARLOCK 7,8,9]  STASHEFF, CHRISTOPHER  H
ODDBALLS, THE  GHIDALIA, VIC  PA
ODDS AND GODS  HOLT, TOM  P
ODIOUS ONES, THE  SOHL, JERRY  H
ODYSSEA NA MARSU  WEINBAUM, STANLEY G.  FP
ODYSSEUS SOLUTION, THE  BANKS, MICHAEL/LAMBE, DEAN R.  C/P,P
ODYSSEY, THE  HOMER  P
ODYSSEY [ROBOT CITY 1]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
ODYSSEY  LAUMER, KEITH  P
ODYSSEY  MCDEVITT, JACK  H
ODYSSEY [@MAGIC: THE GATHERING ODYSSEY 1]  MOORE, VANCE  P
ODYSSEY [@QUANTUM LEAP]  WALTON, BARBARA E.  P
ODYSSEY FILE, THE  CLARKE, ARTHUR C./HYAMS, PETER  R/REF-P
ODYSSEY IN SPACE  BRACK, VEKTIS  SR-P
ODYSSEY OF THE DAMNED  BHATT, JACK  P
ODYSSEY PROJECT, THE  DE MARINO, LAWRENCE  P
ODYSSEY TO EARTHDEATH  KELLEY, LEO P.  P
OEDIPUS CYCLE, THE  SOPHOCLES  LP
OENIPS D’ORLON, L  LE MAY, J./LE MAY, D.  FP
OF A FIRE ON THE MOON  MAILER, NORMAN  [C/RAN-P]
OF AGED ANGELS [DARK MATTER 4]  COOK, MONTE  P
OF ALIEN BONDAGE [@SPACEWAYS #1]  CLEVE, JOHN  P
OF ALL POSSIBLE WORLDS  TENN, WILLIAM  C/P,P,H
OF BEGINNINGS AND ENDINGS [@CASTAWAYS IN TIME #6]  ADAMS, ROBERT  P
OF CHIEFS AND CHAMPIONS [@CASTAWAYS IN TIME #4]  ADAMS, ROBERT  P
OF DARKNESS, LIGHT, AND FIRE (OMNIBUS) [HUFF 3,4]  HUFF, TANYA  P
OF GODLIKE POWER  REYNOLDS, MACK  C/P:2,P
OF MASQUES AND MARTYRS [SHADOWS 3]  GOLDEN, CHRISTOPHER  P
OF MEN AND MONSTERS  TENN, WILLIAM  C/P:2,P
OF MYTHS AND MONSTERS [@CASTAWAYS IN TIME #5]  ADAMS, ROBERT  P
OF OTHER WORLDS  LEWIS, C. S.  LP
OF QUESTS AND KINGS [@CASTAWAYS IN TIME #3]  ADAMS, ROBERT  C/P:2,P
OF SAINTS AND SHADOWS [SHADOWS 1]  GOLDEN, CHRISTOPHER  C/P,P
OF SUCH AS THESE  COX, IRVING JR.  B
OF SWORDS AND SPELLS [MENNENKALTENEI 2]  TURNER, DELIA MARSHALL  P
OF TANGIBLE GHOSTS [JOHAN ESBACH 1]  MODESITT, L. E. JR.  P
OF THE CITY OF THE SAVED... [@FACTION PARADOX 2]  PURSER-HALLARD, PHILIP  LP
OF THE FALL  MCAULEY, PAUL J.  P
OF TIME AND SPACE AND OTHER THINGS  ASIMOV, ISAAC  [C/RAN-P:2]
OF WAR AND HONOR [@HARRIERS #1]  DICKSON, GORDON R.  PA,C/PA,SFWA-P
OF WORLDS BEYOND: THE SCIENCE OF SCIENCE FICTION WRITING  ESHBACH, LLOYD ARTHUR  R/REF-H
OFF ARMAGEDDON REEF  WEBER, DAVID  H
OFF CENTER  KNIGHT, DAMON  C/P,DM113,P
OFF LIMITS  DATLOW, ELLEN  C/PA,PA
OFF-LINE  GOLDSTONE, LAWRENCE  H
OFF ON A COMET (ACE BOOKS, 1957, ABRIDGED)  VERNE, JULES  C/P:2,P
OFF SEASON, THE  CADY, JACK  H
OFF THE BEATEN ORBIT  MERRIL, JUDITH  PA,C/PA
OFF THE MAIN SEQUENCE  HEINLEIN, ROBERT A.  H
OFF THE WALL AT CALLAHAN’S [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  LP
OFF-WORLDERS, THE  BAXTER, JOHN  DG588
OFFICER-CADET [DIRIGENT MERCENARY CORPS 1]  SHELLEY, RICK  C/P,P

OFFICIAL BATTLESTAR GALACTICA SCRAPBOOK, THE [BATTLESTAR GALACTICA 1978] NEYLAND, JAMES

R/REF-H

OFFICIAL DOCTOR WHO AND THE DALEKS BOOK, THE  PEEL, JOHN/NATION, TERRY  R/REF-H
OFFICIAL GUIDE TO FANTASTIC LITERATURE, THE  RESNICK, MIKE  R/REF-P
OFFICIAL PRICE GUIDE: STAR TREK AND STAR WARS COLLECTIBLES, THE  CORNWELL, SUE/KOTT, MIKE  R/REF-P
OFFICIAL STAR TREK QUIZ BOOK, THE [STAR TREK]  MAGLIO, MITCHELL  [C/RAN-P]
OFFSPRING [SILENT EMPIRE 4]  HARPER, STEVEN  P
OFFWORLD  DANIELS, MAX  P
OGDEN ENIGMA, THE  SNYDER, GENE  C/P,P
OGILVIE, TALLANT & MOON  YARBRO, CHELSEA QUINN  H
OGRE CASTLE [SPELL WARS 1]  HALE, F. J.  P
OGRE DOWNSTAIRS, THE  JONES, DIANA WYNNE  H
OGRE, OGRE [@XANTH #5]  ANTHONY, PIERS  C/P:2,P
OGRE’S PACT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #1]  DENNING, TROY  P
OH, GOD!  CORMAN, AVERY  C/P,P
OH NO IT ISN’T! [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  CORNELL, PAUL  C/P,P
OH SUSANNAH!  WILHELM, KATE  H
OI, ROBOT  FERMAN, EDWARD L.  [C/RAN-P]
OISEAUX DE VEGA, LES  LIMAT, MAURICE  FP
OKLA HANNALI  LAFFERTY, R. A.  P
OLAF STAPLEDON, A MAN DIVIDED  FIEDLER, LESLIE A.  R/REF-H

The MITSFS Titledex  September 24, 2007  Page 255
TITLE  AUTHOR  CODE
OLD FRIEND OF THE FAMILY, AN [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
OLD FUNNY STUFF, THE [@AUTHOR’S CHOICE MONTHLY #1]  EFFINGER, GEORGE ALEC  SR-H
OLD GODS WAKEN, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  C/P,SFWA,P,H
OLD GUARD [@BOLOS #5]  FAWCETT, BILL  PA
OLD MAN’S WAR [OLD MAN’S WAR 1]  SCALZI, JOHN  H:2
OLD NATHAN  DRAKE, DAVID  C/P:2,SFWA-P,P
OLD SOLDIERS [BOLOS]  WEBER, DAVID  H
OLD TIN SORROWS [GARRETT 4]  COOK, GLEN  C/P,P
OLD TWENTIETH  HALDEMAN, JOE  H
OLD WOUNDS [@STAR TREK VOYAGER SPIRIT WALK 1]  GOLDEN, CHRISTIE  P
OLDER KIND OF MAGIC, AN  WRIGHTSON, PATRICIA  P
OLDEST CONFESSION, THE  CONDON, RICHARD  P
OLDEST GOD, THE  MCKENNA, STEPHEN  H
OLE DOC METHUSELAH  HUBBARD, L. RON  C/P:2,P
OLGA ROMANOFF  GRIFFITH, GEORGE  LP
OLGA, THE DISILLUSIONED [@PHENWICK #32]  KIMBROUGH, KATHERYN  P
OLIVIA, THE TORMENTED [@PHENWICK #9]  KIMBROUGH, KATHERYN  P
OLYMPOS [ILIUM 2]  SIMMONS, DAN  H

OLYMPUS GREENBERG, MARTIN H./ ARTHURS, BRUCE D.

PA

OM: THE SECRET OF AHBOR VALLEY  MUNDY, TALBOT  C/P,P
OM: THE SECRET OF AHBOR VALLEY (COVER MISSPELLED)  MUNDY, TALBOT  LP
OMBRIA IN SHADOW  MCKILLIP, PATRICIA A.  H
OMEGA  ELWOOD, ROGER  C/PA,PA,SFWA
OMEGA  FARRAR, STEWART  H
OMEGA  FLAMMARION, CAMILLE  SR-H
OMEGA [PRISCILLA HUTCHINS 4]  MCDEVITT, JACK  H
OMEGA CAGE, THE [MATADOR 4]  PERRY, STEVE/REAVES, MICHAEL  C/P,SFWA-P,P
OMEGA EXPEDITION, THE [EMORTALITY]  STABLEFORD, BRIAN M.  H
OMEGA OPERATION, THE [HUNTER 1]  CONWAY, NORMAN  P
OMEGA POINT, THE  ZEBROWSKI, GEORGE  C/P:2,P
OMEGA POINT TRILOGY, THE  ZEBROWSKI, GEORGE  P
OMEGA SUB [@OMEGA SUB #1]  CAMERON, J. D.  P
OMEGA THREAT, THE  WASHBURN, MARK  P
OMEN, THE [OMEN 1]  SELTZER, DAVID  P
OMENS OF KREGEN [DRAY PRESCOT 36]  AKERS, ALAN BURT  P
OMHA ABIDES  MACAPP, C. C.  C/P,P
OMICRON INVASION, THE [FAMILY D’ALEMBERT #9]  SMITH, E. E./GOLDIN, STEPHEN  @P
OMNI BEST SCIENCE FICTION TWO  DATLOW, ELLEN  LPA
OMNI STRAIN, THE  PATTON, CLIFF  P
OMNI VISIONS TWO  DATLOW, ELLEN  LPA
OMNIBUS OF SCIENCE FICTION  CONKLIN, GROFF  HA
OMNIBUS OF TIME, THE  FARLEY, RALPH M.  H
OMNIFICENCE FACTOR, THE  HUGHES, ZACH  C/P,P
OMNIFIX  MACKAY, SCOTT  P
OMNI’S SCREEN FLIGHTS/SCREEN FANTASIES  PEARY, DANNY  SFWA
OMNIUMGATHUM  BACON, JONATHAN/TROYANOVICH, STEVE  SR-VLPA
OMNIVORE [OF MAN AND MANTA 1]  ANTHONY, PIERS  C/P,P,H
ON  ROBERTS, ADAM  H,LP
ON A BEAM OF LIGHT [K-PAX 2]  BREWER, GENE  P
ON A DARKLING PLAIN  BYERS, RICHARD LEE  P
ON A LARK TO THE PLANETS  MONTGOMERY, FRANCES T.  H
ON A PALE HORSE [INCARNATIONS OF IMMORTALITY 1]  ANTHONY, PIERS  H:2,P
ON A PLANET ALIEN  MALZBERG, BARRY N.  C/P:2,P
ON A TORN-AWAY WORLD  ROCKWOOD, ROY  H
ON ALIEN WINGS [@VAMPIRELLA #2]  GOULART, RON  C/P,P
ON AN ODD NOTE  KERSH, GERALD  C/P,P
ON BASILISK STATION [HONOR HARRINGTON 1]  WEBER, DAVID  C/P:2,SFWA-P,P
ON BLUE’S WATERS [BOOK OF THE SHORT SUN 1]  WOLFE, GENE  H,LP
ON BOARD THE FLYING SAUCERS  GIBBONS, GAVIN  [C/RAN-P]
ON BORROWED TIME  WATKIN, LAWRENCE EDWARD  H
ON COMPANY TIME [COMPANY 1,2]  BAKER, KAGE  H
ON CRUSADE: MORE TALES OF THE KNIGHTS TEMPLAR  KURTZ, KATHERINE  LPA
ON FIRE’S WINGS [KEVLA OF ARUKAN 1]  GOLDEN, CHRISTIE  LP
ON HER MAJESTY’S SECRET SERVICE [JAMES BOND 11]  FLEMING, IAN  C/P,P
ON MY WAY TO PARADISE  WOLVERTON, DAVE  C/P:2,SFWA-P,P
ON OUR WAY TO THE FUTURE  CARR, TERRY  PA,C/PA
ON SPEC: THE FIRST FIVE YEARS  ON SPEC EDITORIAL COLLECTIVE  PA
ON STRANGER TIDES  POWERS, TIM  H:2,SFWA-P,P
ON THE BEACH  SHUTE, NEVIL  C/P,P,H
ON THE BOUNDARIES OF DARKNESS [@WILDERNESS OF FOUR #3]  HANCOCK, NIEL  P
ON THE EIGHTH DAY  OKUN, LAWRENCE  P
ON THE GOOD SHIP ENTERPRISE  TRIMBLE, BJO  R/REF-P
ON THE LOST CONTINENT OF MONGO [FLASH GORDON]  WILLIAMSON, AL  SCX

The MITSFS Titledex  September 24, 2007  Page 256
TITLE  AUTHOR  CODE
ON THE NECESSITY OF BESTIALIZING THE HUMAN FEMALE  SIMS, MARGOT  [C/RAN-P]
ON THE OCEANS OF ETERNITY [NANTUCKET 3]  STIRLING, S. M.  C/P:2,P
ON THE RED WORLD  KELLEY, LEO P.  SR-P
ON THE RIM OF THE MANDALA  COOK, PAUL  P
ON THE RUN  DICKSON, GORDON R.  C/P:2,SFWA-P,P
ON THE RUN [INVASION AMERICA 2]  GOLDEN, CHRISTIE  P
ON THE SEAS OF DESTINY  EMERSON, RU  P
ON THE SECRET SERVICE OF HIS MAJESTY, THE QUEEN  WEINSTEIN, SOL  LP
ON THE SHOULDERS OF GIANTS [DREAMQUEST 2]  SAINT ALCORN, LLOYD  P
ON THE SYMB-SOCKET CIRCUIT  BULMER, KENNETH  C/P:2,P
ON THE TRAIL OF THE SPACE PIRATES [TOM CORBETT 3]  ROCKWELL, CAREY  H:2
ON THE VERGE [STAR*DRIVE 1]  GREEN, ROLAND J.  P
ON WHEELS  JAKES, JOHN  C/P,P

ON WINGS OF MAGIC [WITCH WORLD TURNING 3] NORTON, ANDRE/MATHEWS, PATRICIA/

MILLER, SASHA

H:2

ON WINGS OF SONG  DISCH, THOMAS M.  C/P,P,H
ONCE...  HERBERT, JAMES  H
ONCE A HERO [HERIS SERRANO 4]  MOON, ELIZABETH  C/P:2,P
ONCE A HERO  STACKPOLE, MICHAEL A.  C/P:2,P
ONCE AND FUTURE KING, THE  WHITE, T. H.  C/P,P,H
ONCE AROUND THE BLOCH  BLOCH, ROBERT  R/REF-H
ONCE AROUND THE REALMS [@FORGOTTEN REALMS]  THOMSEN, BRIAN M.  P

ONCE BURNED [STAR TREK CAPTAIN’S TABLE 5/ STAR TREK NEW FRONTIER] DAVID, PETER

C/P,P

ONCE DEPARTED  REYNOLDS, MACK  C/P,P
ONCE ON A TIME  MILNE, A. A.  C/P,P
ONCE THERE WAS A GIANT  LAUMER, KEITH  SR-H,H:2,SFWA
ONCE THERE WAS A GIANT/NO SHIP BOOTS IN FAIRYLAND  LAUMER, KEITH  C/P,P
ONCE UPON A GALAXY  ARNOLD, ALAN  R/REF-P,C/REF-P

ONCE UPON A GALAXY MCCARTHY, WIL/GREENBERG, MARTIN H./

HELFERS, JOHN

PA

ONCE UPON A GALAXY  SHERMAN, JOSEPHA  R/REF-P
ONCE UPON A SPACE  CAMPBELL, H. J.  SR-P
ONCE UPON A TIME  DEL REY, LESTER/KESSLER, RISA  SR-HA,HA
ONCE UPON A TIME IN THE EAST [DIEGO 1]  FENN, LIONEL  P
ONCE UPON A WINTER’S NIGHT  MCKIERNAN, DENNIS L.  P
ONCE UPON STILETTOS [KATIE CHANDLER 2]  SWENDSON, SHANNA  LP
ONE  BACH, RICHARD  H
ONE  KARP, DAVID  SR-P,H
ONE AGAINST A WILDERNESS [KIOGA 3]  CHESTER, WILLIAM L.  P
ONE AGAINST ETERNITY  VAN VOGT, A. E.  DD094
ONE AGAINST HERCULUM  SOHL, JERRY  DD381
ONE AGAINST THE LEGION  WILLIAMSON, JACK  C/P:2,P
ONE AGAINST TIME  DEL MARTIA, ASTRON  C/P:2,P
ONE-ARMED QUEEN, THE [ALTA 3]  YOLEN, JANE  C/P,P
ONE DOOR AWAY FROM HEAVEN  KOONTZ, DEAN R.  P
ONE-EYE  GORDON, STUART  P
ONE-EYED JACKS [@WILD CARDS #8]  MARTIN, GEORGE R. R.  PA,C/PA

ONE-EYED MYSTIC [DOC SAVAGE #111]/ MAN WHO FELL UP, THE [DOC SAVAGE 112] ROBESON, KENNETH

@P

ONE FOOT IN THE GRAVE [HALFLIFE 1]  SIMMONS, WM. MARK  C/P:2,P
ONE FOR SORROW, TWO FOR JOY  WOODALL, CLIVE  H
ONE FOR THE MORNING GLORY  BARNES, JOHN  C/P,P,H
ONE FOR THE ROAD  BROWN, FREDRIC  P
ONE GIFT, THE  SHEEHAN, PERLEY POORE  SR-LP
ONE GOOD KNIGHT [FIVE HUNDRED KINGDOMS]  LACKEY, MERCEDES  H
ONE HALF OF THE WORLD  BARLOW, JAMES  H
ONE HUMAN MINUTE  LEM, STANISLAW  H,P
ONE HUNDRED AND TWO H-BOMBS  DISCH, THOMAS M.  C/P:2,P

100 ASTOUNDING LITTLE ALIEN STORIES DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

HA

100 CREEPY LITTLE CREATURE STORIES DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

HA

100 FATHOMS UNDER [@RICK BRANT #4]  BLAINE, JOHN  H

100 GHASTLY LITTLE GHOST STORIES DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

HA

100 GREAT FANTASY SHORT SHORT STORIES ASIMOV, ISAAC/CARR, TERRY/ GREENBERG, MARTIN H.

PA,SFWA

100 GREAT SCIENCE FICTION SHORT SHORT STORIES ASIMOV, ISAAC/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

PA,SFWA

100 HAIR-RAISING LITTLE HORROR STORIES  SARRANTONIO, AL/GREENBERG, MARTIN H. HA

The MITSFS Titledex  September 24, 2007  Page 257
TITLE  AUTHOR  CODE
ONE HUNDRED PAGES FOR THE FUTURE  PECCEI, AURELIO  [C/RAN-P]

100 VICIOUS LITTLE VAMPIRE STORIES WEINBERG, ROBERT/

DZIEMIANOWICZ, STEFAN R./ GREENBERG, MARTIN H.

HA

100 WICKED LITTLE WITCH STORIES DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

HA

100 WILD LITTLE WEIRD TALES DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

HA,LPA

ONE HUNDRED YEARS OF SCIENCE FICTION ILLUSTRATION  FREWIN, ANTHONY  A
ONE IN THREE HUNDRED  MCINTOSH, J. T.  DD113,H:2
ONE JUMP AHEAD  VAN NAME, MARK L.  H
ONE KING, ONE SOLDIER  IRVINE, ALEXANDER C.  LP
ONE KINGDOM, THE [@SWANS’ WAR #1]  RUSSELL, SEAN  H
ONE KING’S WAY [HAMMER AND THE CROSS 2]  HARRISON, HARRY/HOLM, JOHN  C/P,P
ONE KNIGHT ONLY [ARTHUR PENN 2]  DAVID, PETER  P
ONE LAMP  VAN GELDER, GORDON  LPA
ONE LAND, ONE DUKE [@NIGHT-THREADS #3]  EMERSON, RU  C/P,P
ONE-MAN SHOW  THAYER, TIFFANY  H
ONE MAN’S UNIVERSE  SHEFFIELD, CHARLES  LP
ONE MILLION A.D.  DOZOIS, GARDNER  HA
ONE MILLION CENTURIES  LUPOFF, RICHARD A.  C/P:2,P
ONE MILLION TOMORROWS  SHAW, BOB  C/P:2,P
ONE MIND’S EYE  TYERS, KATHY  P
ONE MORE FOR THE ROAD  BRADBURY, RAY  H
ONE OF OUR ASTEROIDS IS MISSING  KNOX, CALVIN M.  DF253
ONE OF OUR H BOMBS IS MISSING  BRENNAN, FREDERICK H.  P
ONE OF US  SMITH, MICHAEL MARSHALL  H,LP
ONE ON ME  HUNTLEY, TIM  C/P,P
ONE QUEST, HOLD THE DRAGONS [CUPS AND SORCERY 2]  COSTIKYAN, GREG  P
ONE SIDE LAUGHING  KNIGHT, DAMON  H
ONE SMALL STEP [STAR TREK GATEWAYS #1]  WRIGHT, SUSAN  P
ONE STEP FROM EARTH  HARRISON, HARRY  H:2,SFWA-P,P
1000-YEAR VOYAGE  STATTEN, VARGO  SR-P
ONE TREE, THE [THOMAS COVENANT 5]  DONALDSON, STEPHEN R.  C/P,P,H
ONE, TWO, THREE...INFINITY  GAMOW, GEORGE  [C/RAN-P:2]
ONE WHO WALKED ALONE: ROBERT E. HOWARD  ELLIS, NOVALYNE PRICE  R/REF-H
ONION GIRL, THE [NEWFORD]  DE LINT, CHARLES  H,LP
ONLY APPARENTLY REAL  WILLIAMS, PAUL  SFWA,R/REF-P
ONLY BEGOTTEN DAUGHTER  MORROW, JAMES  SFWA-P,H:2,P
ONLY FORWARD  SMITH, MICHAEL MARSHALL  C/P,P,H
ONLY HUMAN  HOLT, TOM  P
ONLY HUMAN [DOCTOR WHO NINTH DOCTOR]  ROBERTS, GARETH  H
ONLY LOVERS LEFT ALIVE  WALLIS, DAVE  C/P,P
ONLY YOU CAN SAVE MANKIND [JOHNNY MAXWELL 1]  PRATCHETT, TERRY  H
ONSLAUGHT [@MAGIC: THE GATHERING ONSLAUGHT 1]  KING, J. ROBERT  P
ONSLAUGHT [@DEMONTECH 1]  SHERMAN, DAVID  P
ONSLAUGHT [@STAR WARS NEW JEDI ORDER DARK TIDE #1]  STACKPOLE, MICHAEL A.  P
OPAL-EYED FAN, THE  NORTON, ANDRE  H:2,P
OPALITE MOON [TY MERRICK 2]  VITOLA, DENISE  P
OPEN SPACE VOL. 1, NO. 1  KALISH, CAROL/BUSIEK, KURT  CX
OPENER OF THE WAY, THE  BLOCH, ROBERT  SR-H,H
OPERATION ARES  WOLFE, GENE  C/P:2,P
OPERATION "ASTREE" (FRENCH) [@PERRY RHODAN #1]  SCHEER, K. H.  FPR
OPERATION AUDACITY [BATTLETECH]  PARDOE, BLAINE LEE  P
OPERATION BOUDOIR  O’SHEA, SEAN  P
OPERATION CHAOS [OPERATION CHAOS 1]  ANDERSON, POUL  C/P:2,SFWA,P
OPERATION DAMOCLES  FELLOWS, OSCAR L.  P
OPERATION DOOMSDAY [BARONESS 5]  KENYON, PAUL  P
OPERATION EXCALIBUR [@BATTLETECH 27]  KEITH, WILLIAM H. JR.  P
OPERATION EXILE  PRICE, E. HOFFMANN  P
OPERATION FUTURE  CONKLIN, GROFF  PA,C/PA
OPERATION INTERSTELLAR  SMITH, GEORGE O.  SR-P
OPERATION IRON FIST [@WARBOTS #7]  STINE, G. HARRY  P
OPERATION ISIS  PRICE, E. HOFFMANN  P
OPERATION LIGHTNING STRIKE  POPE, NICK  P
OPERATION LONGLIFE  PRICE, E. HOFFMANN  C/P:2,P
OPERATION LUNA [OPERATION CHAOS 2]  ANDERSON, POUL  H,P
OPERATION MALACCA  POYER, JOE  H,P
OPERATION MISFIT  PRICE, E. HOFFMANN  C/P:2,P
OPERATION NUKE [@CYBORG #2]  CAIDIN, MARTIN  P
OPERATION OCTOPUS [MARK HOOD 8]  DARK, JAMES  P
OPERATION OMINA  STARR, ROLAND  H

The MITSFS Titledex  September 24, 2007  Page 258
TITLE  AUTHOR  CODE
OPERATION: OUTER SPACE  LEINSTER, MURRAY  C/P:2,SR-LP,LP
OPERATION: PHANTASY  WOLLHEIM, DONALD A.  SR-H
OPERATION ROSWELL  RANDLE, KEVIN D.  H
OPERATION SATELLITE  FEAR, W. H.  SR-P
OPERATION SHATTERHAND  PAGE, JAKE  P
OPERATION SIERRA-75 [@VOR THE MAELSTROM]  GRESSMAN, THOMAS S.  P
OPERATION STEEL BAND [@WARBOTS #2]  STINE, G. HARRY  C/P,P
OPERATION SURVIVAL EARTH  DENAERDE, STEFAN  P
OPERATION TERROR  LEINSTER, MURRAY  C/P:2,P
OPERATION TIME SEARCH  NORTON, ANDRE  C/P:2,P
OPERATION UMANAQ  RANKINE, JOHN  C/P:2,P
OPERATION VENUS  FEARN, JOHN RUSSELL  SR-P
OPHELIA, THE ANXIOUS [@PHENWICK #18]  KIMBROUGH, KATHERYN  P
OPHIDIAN CONSPIRACY, THE  CARR, JOHN F.  P
OPHIUCHI HOTLINE, THE  VARLEY, JOHN  C/P,P,H
OPIUM GENERAL, THE  MOORCOCK, MICHAEL  H
OPOPONAX INVASION, THE  BROSNAN, JOHN  LP
OPOWIADANIA WYBRANE  LEM, STANISLAW  FP

OPPORTUNITY FOR PROFIT, AN [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #5] GROSS, DAVE

P

OPTIMAN  STABLEFORD, BRIAN M.  C/P:2,P
OPTIMISM ONE  ESFANDIARY, S. M.  [C/RAN-P]
OPTION LOCK [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  C/P:2,P
OPTIONS  SHECKLEY, ROBERT  C/P:2,P
OPUS 100  ASIMOV, ISAAC  [C/RAN-H,C/RAN-P:2]
OPUS 200  ASIMOV, ISAAC  [C/RAN-P:2]
OPUS 300  ASIMOV, ISAAC  [C/RAN-H]
OPUS 300 (PROOF)  ASIMOV, ISAAC  SR-LP
OR ALL THE SEAS WITH OYSTERS  DAVIDSON, AVRAM  C/P:2,P
ORACLE, THE [@WIZARD WAR #4]  COOK, HUGH  P
ORACLE  O’DONNELL, KEVIN JR.  C/P:2,P
ORACLE  RESNICK, MIKE  C/P,SFWA-P:2,P
ORACLE  WATSON, IAN  P
ORACLE BETRAYED, THE [ORACLE 1]  FISHER, CATHERINE  P
ORACLES OF NOSTRADAMUS  NOSTRADAMUS  [C/RAN-H]
ORACLE’S QUEEN, THE [TOBIN OF SKALA 3]  FLEWELLING, LYNN  C/P,P
ORANGE R, THE  CLAGETT, JOHN  C/P,P
ORB OF XORIAT, THE [@EBERRON WAR-TORN #2]  BOLME, EDWARD  P
ORBIS  MACKAY, SCOTT  C/P,P
ORBIT  BLOCK, THOMAS H.  H,P
ORBIT 1  KNIGHT, DAMON  C/PA,PA
ORBIT 2  KNIGHT, DAMON  PA,C/PA
ORBIT 3  KNIGHT, DAMON  PA,C/PA
ORBIT 4  KNIGHT, DAMON  C/PA,PA
ORBIT 5  KNIGHT, DAMON  C/PA,PA
ORBIT 6  KNIGHT, DAMON  C/PA,PA
ORBIT 7  KNIGHT, DAMON  C/PA,PA
ORBIT 8  KNIGHT, DAMON  C/PA,PA
ORBIT 9  KNIGHT, DAMON  C/PA,PA
ORBIT 10  KNIGHT, DAMON  C/PA,PA
ORBIT 11  KNIGHT, DAMON  HA,PA
ORBIT 12  KNIGHT, DAMON  C/PA,PA
ORBIT 13  KNIGHT, DAMON  C/PA,PA
ORBIT 14  KNIGHT, DAMON  HA,SFWA
ORBIT 15  KNIGHT, DAMON  HA,SFWA
ORBIT 16  KNIGHT, DAMON  HA,SFWA:2
ORBIT 17  KNIGHT, DAMON  SFWA
ORBIT 18  KNIGHT, DAMON  HA,SFWA
ORBIT 19  KNIGHT, DAMON  HA,SFWA
ORBIT 20  KNIGHT, DAMON  HA,SFWA
ORBIT 21  KNIGHT, DAMON  HA,SFWA
ORBIT ONE  JAY, MEL  P
ORBIT ONE  MULLER, JOHN E.  SR-P
ORBIT UNLIMITED [RUSTUM]  ANDERSON, POUL  C/P:2,P
ORBITAL DECAY  STEELE, ALLEN  SR-P,C/P:2,SFWA-P,P
ORBITAL RESONANCE  BARNES, JOHN  C/P,SFWA,P,H
ORBITER  CZERNEDA, JULIE E.  LPA
ORBITSVILLE  SHAW, BOB  C/P:2,SFWA-P,P
ORBITSVILLE DEPARTURE  SHAW, BOB  C/P:2,SFWA-P,P
ORC WARS, THE [@YNGLING #1,2]  DALMAS, JOHN  C/P:2,SFWA-P,P
ORCA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
ORCHARD, THE  GRANT, CHARLES L.  SFWA-P
ORCHID CAGE, THE  FRANKE, HERBERT W.  C/P,P
ORCHIDEENKAFIG, DER  FRANKE, HERBERT W.  FH

TITLE  AUTHOR  CODE
ORC’S OPAL [KELVIN HACKLEBERRY 4]  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:2,P
ORDA DI CRISTALLO, L  TAINE, JOHN  FP
ORDEAL, THE [@SURVIVALIST #17]  AHERN, JERRY  P
ORDEAL  SHUTE, NEVIL  P
ORDEAL IN OTHERWHERE [FORERUNNER 2]  NORTON, ANDRE  C/P:2,P
ORDER OF THE DAY, THE  SOUZA, MARCIO  P
ORDER WAR, THE [RECLUCE 4]  MODESITT, L. E. JR.  H,P
ORDERMASTER [RECLUCE 13]  MODESITT, L. E. JR.  H
ORDINARY PEOPLE  ARNASON, ELEANOR  LP
OREGON INVASION, THE [V]  TANNEHILL, JAYNE  C/P,SFWA-P,P
ORGAN BANK FARM, THE  BOYD, JOHN  C/P,P,H
ORGANIC DESTINY  ROMILUS, ARN  SR-P
ORIENTAL STORIES  ODYSSEY PUBLICATIONS  SR-LPA
ORIENTAL TALES  YOURCENAR, MARGUERITE  H
ORIGIN IN DEATH [EVE DALLAS]  ROBB, J. D.  H
ORIGIN OF THE CRABS, THE [CRABS 1]  SMITH, GUY N.  P
ORIGINAL HITCHHIKER RADIO SCRIPTS, THE [HITCHHIKERS]  ADAMS, DOUGLAS  H,LP
ORIGINAL MR. ED, THE  BROOKS, WALTER  P
ORIGINAL SIN [DOCTOR WHO NEW ADVENTURES]  LANE, ANDY  C/P:2,P
ORIGINAL SINS [@BUFFY LOST SLAYER #4]  GOLDEN, CHRISTOPHER  C/P
ORIGINS OF MARVEL COMICS [MARVEL COMICS]  LEE, STAN  CX
ORION [ORION 1]  BOVA, BEN  C/P,SFWA-P:2,P,H
ORION AMONG THE STARS [ORION 4]  BOVA, BEN  C/P,SFWA,P,H
ORION AND THE CONQUEROR [ORION 3]  BOVA, BEN  C/P,P
ORION ARM [RAMPART WORLDS 2]  MAY, JULIAN  C/P,P
ORION IN THE DYING TIME [ORION 2]  BOVA, BEN  C/P:2,P,H
ORION SHALL RISE  ANDERSON, POUL  H:2,LP
ORION’S DAGGER [CLOUDSHIPS OF ORION 3]  MCALLISTER, P. K.  P
ORION’S HOUNDS [STAR TREK TITAN]  BENNETT, CHRISTOPHER L.  P
ORION’S SWORD [@FUTURE AT WAR #3]  BRETNOR, REGINALD  PA,C/PA,SFWA-P
ORIZZONTE PERDUTO  HILTON, JAMES  FP
ORLANDO  WOOLF, VIRGINIA  P
ORN [OF MAN AND MANTA 2]  ANTHONY, PIERS  C/P,P,H
ORNAMENT TO HIS PROFESSION, AN  HARNESS, CHARLES L.  H
ORON [@ORON #1]  SMITH, DAVID C.  C/P,P
ORPHAN, THE  STALLMAN, ROBERT  C/P,P
ORPHAN OF CREATION  ALLEN, ROGER MACBRIDE  P,SFWA-P
ORPHAN STAR [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
ORPHANAGE [JASON WANDER 1]  BUETTNER, ROBERT  C/P,P
ORPHAN’S DESTINY [JASON WANDER 2]  BUETTNER, ROBERT  P
ORPHANS OF CHAOS [ORPHANS OF CHAOS 1]  WRIGHT, JOHN C.  H:2
ORPHANS OF EARTH [ECHOES OF EARTH 2]  WILLIAMS, SEAN/DIX, SHANE  P
ORPHANS OF THE SKY  HEINLEIN, ROBERT A.  H:2,P
ORPHEUS MACHINE, THE [@EMANCIPATOR #3]  ALDRIDGE, RAY  C/P,P
ORPHEUS PROCESS, THE  GOWER, DANIEL H.  P
ORSINIAN TALES  LE GUIN, URSULA K.  H:2,P
ORSON WELLES’ ...WAR OF THE WORLDS (CASSETTE)  KOCH, HOWARD  SR-MM
ORYX AND CRAKE  ATWOOD, MARGARET  H
OSCILLATING UNIVERSE, THE  OPIK, ERNST J.  [C/RAN-P]
OSSIAN’S RIDE  HOYLE, FRED  H:2,P
OTHER [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H
OTHER, THE  TRYON, THOMAS  C/P:2,P
OTHER AMERICAS  SPINRAD, NORMAN  C/P,P
OTHER DAYS, OTHER EYES  SHAW, BOB  C/P:2,P
OTHER DIMENSION, THE  ROSNER, SARA  PA
OTHER DIMENSIONS  SILVERBERG, ROBERT  HA,PA
OTHER DIMENSIONS  SMITH, CLARK ASHTON  SR-H
OTHER END OF TIME, THE [ESCHATON 1]  POHL, FREDERIK  C/P:2,P
OTHER FOOT, THE  KNIGHT, DAMON  C/P,P
OTHER GLASS TEAT, THE  ELLISON, HARLAN  [C/RAN-P:2]
OTHER HUMAN RACE, THE [FUZZIES 2]  PIPER, H. BEAM  C/P:2,P
OTHER LOG OF PHILEAS FOGG, THE  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
OTHER NATURE  SMITH, STEPHANIE A.  LP
OTHER NINETEENTH CENTURY, THE  DAVIDSON, AVRAM  LP
OTHER PASSENGER, THE  CROSS, JOHN KEIR  C/P,P
OTHER PEOPLE, THE  BRISCO, PAT  P
OTHER PLACE, THE  PRIESTLEY, J. B.  H
OTHER SENSES, OTHER WORLDS  JONAS, DORIS/JONAS, DAVID  R/REF-H
OTHER SIDE, THE  CHETWYND-HAYES, R.  SFWA-P
OTHER SIDE, THE  KUBILIUS, WALTER  B
OTHER SIDE OF HERE, THE  LEINSTER, MURRAY  DD094
OTHER SIDE OF NIGHT, THE  CHARKIN, PAUL  SR-P
OTHER SIDE OF NOWHERE, THE  LEINSTER, MURRAY  C/P:2,P
OTHER SIDE OF THE CLOCK, THE  STERN, PHILIP VAN DOREN  PA

The MITSFS Titledex  September 24, 2007  Page 260
TITLE  AUTHOR  CODE
OTHER SIDE OF THE MIRROR, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
OTHER SIDE OF THE MOON, THE  DERLETH, AUGUST  PA,C/PA
OTHER SIDE OF THE SKY, THE  CLARKE, ARTHUR C.  C/P,P,H
OTHER SIDE OF THE UNIVERSE, THE  DREIFUSS, KURT  H
OTHER SIDE OF TIME, THE [IMPERIUM 2]  LAUMER, KEITH  SR-H,C/P:2,P
OTHER SIDE OF TOMORROW, THE  ELWOOD, ROGER  C/PA,PA
OTHER SINBAD, THE [OTHER SINBAD 1]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
OTHER STORIES AND THE ATTACK OF THE GIANT BABY  REED, KIT  P
OTHER TIME, THE  REYNOLDS, MACK/ING, DEAN  C/P,P
OTHER TIMES, OTHER WORLDS  MACDONALD, JOHN D.  C/P:2,P
OTHER TIMES THAN PEACE  DRAKE, DAVID  H
OTHER WAYS [SECANTIS]  TIEDEMANN, MARK W.  SR-LP
OTHER WIND, THE [EARTHSEA]  LE GUIN, URSULA K.  LP
OTHER WORLD, THE [DOC SAVAGE #29]  ROBESON, KENNETH  @P
OTHER WORLD, THE (BRITISH)  BOND, J. HARVEY  SR-P,C/P
OTHER WORLDS  COLLINS, PAUL  SR-HA
OTHER WORLDS  SAGAN, CARL  [C/RAN-P]
OTHER WORLDS  STONG, PHIL  HA
OTHER WORLDS 1  TORGESON, ROY  PA
OTHER WORLDS 2  TORGESON, ROY  PA
OTHER WORLDS OF CLIFFORD SIMAK  SIMAK, CLIFFORD D.  C/P:2,P
OTHER WORLDS, OTHER GODS  MOHS, MAYO  C/PA,PA

OTHER WORLDS, OTHER SEAS:

SCIENCE FICTION STORIES FROM SOCIALIST COUNTRIES SUVIN, DARKO

HA,PA

OTHER WORLDS, OTHER TIMES  MOSKOWITZ, SAM/ELWOOD, ROGER  C/PA,PA
OTHER WORLDS, OTHER UNIVERSES  STEIGER, BRAD/WHITE, JOHN  SFWA
OTHER WORLDS: THE FANTASY GENRE  TIMMERMAN, JOHN H.  R/REF-P
OTHERNESS  BRIN, DAVID  C/P:2,P
OTHERS, THE [OTHERS (BONANNO) 1]  BONANNO, MARGARET WANDER  H,SFWA
OTHERS, THE  CARR, TERRY  PA,C/PA
OTHERS, THE  GREENFIELD, IRVING A.  C/P:2,P

OTHERWERE GILMAN, LAURA ANNE/

DECANDIDO, KEITH R. A.

PA

OTHERWHERE [OTHERS (BONANNO) 2]  BONANNO, MARGARET WANDER  H
OTHERWISE [OTHERS (BONANNO) 3]  BONANNO, MARGARET WANDER  H
OTHERWORLD  FLINT, KENNETH C.  C/P,P

OTHERWORLD, THE LACKEY, MERCEDES/SHEPHERD, MARK/

LISLE, HOLLY

P

OU LU KHEN AND THE BEAUTIFUL MADWOMAN  SALMONSON, JESSICA AMANDA  C/P:2,P
OUR BEST  POHL, FREDERIK/KORNBLUTH, C. M.  C/P,P
OUR CHILDREN’S CHILDREN  SIMAK, CLIFFORD D.  C/P,P,H
OUR FRIEND: PUBLIC NUISANCE #1 [@SAMANTHA SLADE #3]  SMITH, SUSAN  P
OUR FRIENDS FROM FROLIX 8  DICK, PHILIP K.  H:2,P
OUR HAUNTED PLANET  KEEL, JOHN A.  [C/RAN-P:2]
OUR LADY OF DARKNESS  LEIBER, FRITZ  C/P,TD36,P,H
OUR MAN IN SPACE  RONALD, BRUCE W.  DM117
OUR MOTHER’S HOUSE  GLOAG, JULIAN  P

OUR NEXT PRESIDENT:

THE INCREDIBLE STORY OF WHAT HAPPENED IN THE 1968 ELECTIONS BAKER, RUSSELL

P

OUR PLACE IN TIME  BANCHI, CHARLOTTE A  LP
OUR SPACE CRAFT IS MISSING!  RICHARDS, PAUL  P
OUR SUN AND THE WORLDS AROUND IT  LYON, JENE  [C/RAN-H]
OUR TOPIC TODAY IS D&D (RADIO SERMON CASSETTE)  LARSON, BOB  SR-MM
OUR WORLD IN SPACE ATLAS  HAMMOND AND COMPANY  [C/RAN-H]
OUT FROM GANYMEDE  MALZBERG, BARRY N.  P
OUT OF A FOREST CLEARING  PLATT, RANDALL BETH  LP
OUT OF AVALON  ROBERSON, JENNIFER  PA
OUT OF BOUNDS  MERRIL, JUDITH  C/P:2,P
OUT OF MY HEAD  BLOCH, ROBERT  SR-H
OUT OF MY MIND  BRUNNER, JOHN  C/P:2,P
OUT OF PHAZE [APPRENTICE ADEPT #4]  ANTHONY, PIERS  C/P:2,@P
OUT OF SIGHT, OUT OF MIND  AARON, CHESTER  C/P,P
OUT OF SPACE AND TIME I  SMITH, CLARK ASHTON  SR-P
OUT OF SPACE AND TIME II  SMITH, CLARK ASHTON  SR-P
OUT OF SYNC  HAIBLUM, ISIDORE  P
OUT OF THE ASHES [ASHES 1]  JOHNSTONE, WILLIAM W.  C/P,P
OUT OF THE BODY EXPERIENCES  GREEN, CELIA  [C/RAN-P]
OUT OF THE DARK WORLD  CHETWIN, GRACE  P
OUT OF THE DARKNESS [@BABYLON 5 LEGIONS OF FIRE #3]  DAVID, PETER  C/P,P
OUT OF THE DARKNESS  FANTHORPE, R. L.  SR-P
OUT OF THE DARKNESS [INTO THE DARKNESS 6]  TURTLEDOVE, HARRY  P
OUT OF THE DEAD CITY [FALL OF THE TOWERS 1]  DELANY, SAMUEL R.  P
OUT OF THE DEEPS  WYNDHAM, JOHN  C/P:2,P
OUT OF THE EVERYWHERE  ASIMOV, ISAAC  R/REF-H
OUT OF THE EVERYWHERE  TIPTREE, JAMES JR.  C/P,P

The MITSFS Titledex  September 24, 2007  Page 261
TITLE  AUTHOR  CODE
OUT OF THE HOUSE OF LIFE [SAINT-GERMAIN 7]  YARBRO, CHELSEA QUINN  H
OUT OF THE MADHOUSE [@BUFFY GATEKEEPER 1]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  C/P,P
OUT OF THE MOUTH OF THE DRAGON  GESTON, MARK S.  C/P:2,P
OUT OF THE NIGHT  MULLER, JOHN E.  SR-P
OUT OF THE SILENCE  COX, ERLE  H
OUT OF THE SILENT NORTH  DRAGO, HENRY  H
OUT OF THE SILENT PLACES  KENT, BRAD  SR-P
OUT OF THE SILENT PLANET [PERELANDRA 1]  LEWIS, C. S.  C/P,P,H
OUT OF THE STORM  HODGSON, WILLIAM HOPE  H
OUT OF THE SUN  BOVA, BEN  SR-H,C/P,P
OUT OF THE UNKNOWN  VAN VOGT, A. E./HULL, E. MAYNE  SR-H:2,SR-P
OUT OF THE VOID  STONE, LESLIE F.  H
OUT OF THEIR MINDS  SIMAK, CLIFFORD D.  C/P:2,P
OUT OF THIS WORLD  FAST, JULIUS  SR-PA,C/PA
OUT OF THIS WORLD  MACKLIN, JOHN  [C/RAN-P]

OUT OF THIS WORLD [EVE DALLAS] ROBB, J. D./HAMILTON, LAURELL K./ KRINARD, SUSAN/SHAYNE, MAGGIE

C/P,P

OUT OF THIS WORLD [THREE WORLDS 1]  WATT-EVANS, LAWRENCE  H:2,P
OUT OF TIME [EMMA MERRIGAN 1]  ABBEY, LYNN  C/P,P
OUT OF TIME  COONEY, CAROLINE B.  P
OUT OF TIME  HOGAN, JAMES P.  P
OUT OF TIME’S ABYSS [LAND THAT TIME FORGOT 3]  BURROUGHS, EDGAR RICE  C/P:2,P
OUT ON BLUE SIX  MCDONALD, IAN  P
OUT THERE WHERE THE BIG SHIPS GO  COWPER, RICHARD  C/P,P
OUTBACK STARS, THE  MCDONALD, SANDRA  H
OUTBANKER  MADDEN, TIMOTHY A.  C/P,P
OUTBOUND  MCDEVITT, JACK  H
OUTBOUND FLIGHT [@STAR WARS]  ZAHN, TIMOTHY  H
OUTBREAK  COOK, ROBIN  H,P
OUTBREAK  RUUTH, MARIANNE  P
OUTCAST, THE [TIME MASTER 2]  COOPER, LOUISE  C/P:2,SFWA-P,P
OUTCAST, THE [@DARK SUN TRIBE OF ONE #1]  HAWKE, SIMON  C/P,P
OUTCAST OF REDWALL, THE [REDWALL]  JACQUES, BRIAN  P
OUTCASTS [RUNESWORD 1]  EMERY, CLAYTON  C/P,P
OUTCASTS OF HEAVEN BELT, THE [HEAVEN 2]  VINGE, JOAN D.  C/P:2,P
OUTDATED MAN, THE  HARRISON, HARRY  C/PA,PA
OUTER FLEET, THE  MATZKIN, M.  C/P,P
OUTER LIMITS: THE OFFICIAL COMPANION, THE  SCHOW, DAVID J./FRENTZEN, JEFFREY  R/REF-H
OUTER LIMITS VOLUME ONE, THE  NOTKIN, DEBBIE/STEWART, ROGER  LP
OUTER LIMITS VOLUME TWO, THE  NOTKIN, DEBBIE  LP
OUTER PERIMETER  GODDARD, KEN  P
OUTER REACHES, THE  DERLETH, AUGUST  HA,PA
OUTER SPACE STORIES  FURMAN, A. L.  C/PA,PA
OUTERSPACE SEX ORGY  FABER, ARTHUR  SR-L
OUTERWORLD  HAIBLUM, ISIDORE  BS3
OUTLAND  FOSTER, ALAN DEAN  C/P:2,P
OUTLAND STRIP [HORN]  SLOANE, BEN  P
OUTLAND: THE MOVIE NOVEL  ANOBILE, RICHARD J.  VLP
OUTLANDER [OUTLANDER (GABALDON) 1]  GABALDON, DIANA  H:2,P
OUTLANDER [@NECROMUNDA]  KEEFE, MATT  P
OUTLANDERS, THE [LON TOBYN 2]  COE, DAVID B.  C/P:2,P

OUTLAW: CHAMPIONS OF KAMIGAWA [@MAGIC: THE GATHERING KAMIGAWA 1] MCGOUGH, SCOTT

C/P,P

OUTLAW OF GOR [GOR 2]  NORMAN, JOHN  C/P:2,P
OUTLAW OF TORN, THE  BURROUGHS, EDGAR RICE  C/P,P,H
OUTLAW SCHOOL  ORE, REBECCA  H,LP
OUTLAW SWORDS [DYLAN MATHESON 2]  LEE, J. ARDIAN  LP
OUTLAW WORLD  HAMILTON, EDMOND  C/P:2,P
OUTLAWS OF MARS, THE  KLINE, OTIS ADELBERT  C/P,P,H
OUTLAWS OF SHERWOOD, THE  MCKINLEY, ROBIN  C/P:2,SFWA-P,P
OUTLAWS OF THE MOON  HAMILTON, EDMOND  C/P:2,P
OUTLINE OF HISTORY, THE: VOLUME 1 [HISTORY 1]  WELLS, H. G.  [C/RAN-H]
OUTLINE OF HISTORY, THE: VOLUME 2 [HISTORY 2]  WELLS, H. G.  [C/RAN-H]
OUTNUMBERING THE DEAD  POHL, FREDERIK  H
OUTPASSAGE  MORRIS, JANET/MORRIS, CHRIS  C/P:2,P
OUTPOST  MACKAY, SCOTT  LP
OUTPOST, THE  RESNICK, MIKE  H
OUTPOST MARS  JUDD, CYRIL  C/P:2,P
OUTPOST OF JUPITER  DEL REY, LESTER  C/P,P,H
OUTPOSTER, THE  DICKSON, GORDON R.  C/P,SFWA-P:2,P,H
OUTPOSTS IN SPACE  WEST, WALLACE  H

OUTRAGEOUS FORTUNE [@MERCHANT PRINCE #2] SHIMERMAN, ARMIN/

YARBRO, CHELSEA QUINN

P

OUTREACH [DUSHAU 3]  LICHTENBERG, JACQUELINE  C/P,P

TITLE  AUTHOR  CODE
OUTRIDER, THE [@OUTRIDER #1]  HARDING, RICHARD  P
OUTRIDERS [BIRTHRIGHT PROJECT 1]  MACKEL, KATHRYN  LP
OUTSIDE  NORTON, ANDRE  C/P,P,H
OUTSIDE THE BOX  ANDERS, LOU  LPA
OUTSIDE THE DOG MUSEUM  CARROLL, JONATHAN  C/P,P
OUTSIDE THE GATES  GLOSS, MOLLY  H
OUTSIDE THE UNIVERSE  HAMILTON, EDMOND  C/P:2,P
OUTSIDERS  TENN, WILLIAM  C/PA,PA
OUTSKIRTER’S SECRET, THE [STEERSWOMAN 2]  KIRSTEIN, ROSEMARY  C/P,P
OUTSTRETCHED SHADOW, THE [OBSIDIAN 1]  LACKEY, MERCEDES/MALLORY, JAMES  H
OUTWARD BOUND [@CHILDREN OF THE STARS #2]  COULSON, JUANITA  C/P,P,H
OUTWARD BOUND  HOGAN, JAMES P.  P
OUTWARD URGE, THE  WYNDHAM, JOHN/PARKES, LUCAS  C/P:2,P
OUTWORLD CATS  LOVEJOY, JACK  P
OUTWORLDER [GREAT IMPERIUM]  CARTER, LIN  C/P:2,P
OVER MY SHOULDER  ESHBACH, LLOYD ARTHUR  R/REF-H
OVER SEA, UNDER STONE [DARK IS RISING 1]  COOPER, SUSAN  C/P,SR-P,H:2
OVER THE BORDER  ROBERTSON, MORGAN  H
OVER THE EDGE  DERLETH, AUGUST  SR-PA
OVER THE EDGE  ELLISON, HARLAN  C/P,P
OVER THE RAINBOW  OCTOPUS BOOKS  HA
OVER THE RIVER AND THROUGH THE WOODS  SIMAK, CLIFFORD D.  H
OVERCLOCKED  DOCTOROW, CORY  LP
OVERLAY  MALZBERG, BARRY N.  C/P:2,P
OVERLORD [@SURVIVALIST #15]  AHERN, JERRY  P
OVERLORD NEW YORK  ELLIOT, LEE  SR-P
OVERLORDS FROM SPACE  KELLEAM, JOSEPH E.  DD173
OVERLORDS OF ANDROMEDA  FALKNER, JOHN  SR-P
OVERLORDS OF WAR, THE  KLEIN, GERARD  H:2,P,SFWA
OVERMAN CULTURE, THE  COOPER, EDMUND  H:2,P
OVERSHOOT  CLEE, MONA  P
OVERTIME  HOLT, TOM  P
OWL SERVICE, THE  GARNER, ALAN  C/P,P,H
OWL TIME  FOSTER, M. A.  C/P:2,P
OWLFLIGHT [DARIAN 1]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
OWLKNIGHT [DARIAN 3]  LACKEY, MERCEDES/DIXON, LARRY  C/P:2,P,H
OWLS HOOT IN THE DAYTIME  WELLMAN, MANLY WADE  H
OWLS’ WATCH  SAUL, GEORGE BRANDON  C/PA,PA
OWLSIGHT [DARIAN 2]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
OWLSTONE CROWN, THE  KENNEDY, X. J.  P
OX [OF MAN AND MANTA 3]  ANTHONY, PIERS  C/P,P,H
OXFORD BOOK OF FANTASY STORIES, THE  SHIPPEY, TOM  PA
OXFORD BOOK OF SCIENCE FICTION STORIES, THE  SHIPPEY, TOM  HA:2
OXYGEN BARONS, THE  FEELEY, GREGORY  C/P,P
OZ ENCOUNTER, THE [WEIRD HEROES 5]  WOLFMAN, MARV/WHITE, TED  C/P,P
OZ: INTO THE WILD [@BUFFY]  GOLDEN, CHRISTOPHER  P
OZARK TRILOGY, THE [OZARK 1,2,3]  ELGIN, SUZETTE HADEN  H:2
OZINE CONQUEST, THE  GILBERT, C. M.  P
OZMA OF OZ [OZ #3]  BAUM, L. FRANK  H:2,@P,@C/P
OZYMANDIAS  MONTELEONE, THOMAS F.  H

The MITSFS Titledex  September 24, 2007  Page 263
TITLE  AUTHOR  CODE
PACIFIC ADVANCE  MELDE, G. R.  SR-P
PACIFIC BOOK OF SCIENCE FICTION, THE  BAXTER, JOHN  HA
PACIFIC EDGE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  H,P

PACIFICA BETANCOURT, JOHN GREGORY/

BUSHYAGER, LINDA E.

LP

PACK, THE  FULLER, JOHN G.  P
PACKAGE HOLIDAY  VERNE, JULES  H
PACT WITH SATAN, A  HOLTON, LEONARD  P
PAGAN PASSIONS  GARRETT, RANDALL/HARRIS, LARRY M.  SR-P
PAGE [@PROTECTOR OF THE SMALL 2]  PIERCE, TAMORA  H
PAIL OF AIR, A  LEIBER, FRITZ  C/P,P
PAIN GAIN  CARR, JOHN F.  P
PAINFUL FIELD, THE [STARCRUISER SHENANDOAH 5]  GREEN, ROLAND J.  P
PAINGOD  ELLISON, HARLAN  C/P:2,P
PAINTED BIRD, THE (REVISED)  KOSINSKI, JERZY  P
PAINTED DEVILS  AICKMAN, ROBERT  H,SFWA
PAINTER KNIGHT, THE [LIVING FLAME 2]  PATTON, FIONA  P
PAIR FROM SPACE: GIANTS IN THE EARTH, A  BLISH, JAMES  SR-DB0612
PAIR FROM SPACE: WE, THE MARAUDERS, A  SILVERBERG, ROBERT  SR-DB0612
PALACE [PINCH 1]  KERR, KATHARINE/KREIGHBAUM, MARK  P
PALACE, THE [SAINT-GERMAIN 2]  YARBRO, CHELSEA QUINN  H:2,SFWA-P,P
PALACE OF ETERNITY, THE  SHAW, BOB  C/P:2,P
PALACE OF KINGS [ELUNDIUM 2]  JEFFERIES, MIKE  P
PALACE OF LOVE, THE [DEMON PRINCES 3]  VANCE, JACK  C/P:2,P

PALACE OF LOVE (VIE), THE [@VANCE INTEGRAL EDITION #24/ DEMON PRINCES 3] VANCE, JACK

SR-H

PALACE OF THE RED SUN [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  P
PALACES AND PRISONS  MILLER, RON  C/P,P
PALACES AND PRISONS (REVISED) [BRONWYN 1]  MILLER, RON  LP
PALADIN, THE  CHERRYH, C. J.  H:2,SFWA-P,P
PALADIN OF SOULS [CHALION 2]  BUJOLD, LOIS MCMASTER  H
PALADIN OF THE NIGHT, THE [@ROSE OF THE PROPHET #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
PALADINS [PALADINS 1]  ROSENBERG, JOEL  P
PALADINS, THE [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #2]  WARD, JAMES S./WISE, DAVID  P
PALE COMPANION, THE  VON GUNDEN, KENNETH  C/P,P

PALE COUNTRY PURSUIT [@ATLAN #3]/ ROBOT THREAT: NEW YORK (SPECIAL) [PERRY RHODAN] ACE

PA

PALEO [@BUFFY]  NAVARRO, YVONNE  C/P,P
PALIMPSESTS  SCHOLZ, CARTER/HARCOURT, GLENN  P
PALINDOR [THREE LANDS 1]  EVANS, D. R.  LP
PALLAS  SMITH, L. NEIL  P
PALOMA [RETRIEVAL ARTIST 5]  RUSCH, KRISTINE KATHRYN  P
PAN  DIBBLE, BIRNEY  P
PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  PA
THIRD PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  C/PA,PA
FOURTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
FIFTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
SIXTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
EIGHTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
ELEVENTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
THIRTEENTH PAN BOOK OF HORROR STORIES, THE  VAN THAL, HERBERT  SR-PA
PAN SAGITTARIUS  WALLACE, IAN  C/P,P,H
PAN SATYRUS  WORMSER, RICHARD  C/P:2,P
PANCHRONICON PLOT, THE  GOULART, RON  C/P,P
PANDA RAY  KANDEL, MICHAEL  H:2
PANDEMIC  SIMMONS, GEOFFREY  P
PANDORA [NEW TALES 1]  RICE, ANNE  C/P,P,H
PANDORA BY HOLLY HOLLANDER  WOLFE, GENE  H
PANDORA DIRECTIVE, THE [@TEX MURPHY]  CONNERS, AARON  P
PANDORA EFFECT, THE  WILLIAMSON, JACK  C/P,P
PANDORA PRINCIPLE, THE [STAR TREK]  CLOWES, CAROLYN  C/P:3,P
PANDORA STONE, THE  GREENLEAF, WILLIAM  P
PANDORA’S BOX  TUBB, E. C.  LP
PANDORA’S CHILDREN [PANDORA 2]  LANCE, KATHRYN  P
PANDORA’S GENES [PANDORA 1]  LANCE, KATHRYN  C/P,P
PANDORA’S LEGIONS  ANVIL, CHRISTOPHER  P
PANDORA’S PLANET  ANVIL, CHRISTOPHER  C/P,SFWA,P,H
PANDORA’S STAR [COMMONWEALTH 1]  HAMILTON, PETER F.  H
PANGLOR  CARVER, JEFFREY A.  P
PANIC 7  SHAFFER, EUGENE CARL  P
PANIC BROADCAST, THE  KOCH, HOWARD  R/REF-P,C/REF-P
PANORAMA EGG, THE  SILAS, A. E.  P
PANTHER BOOK OF HORROR, THE  RAMPTON, ANTHONY  SR-PA
PANTHER’S HOARD, THE  BERBERICK, NANCY VARIAN  P,SFWA-P:2

TITLE  AUTHOR  CODE
PANZER SOLDIER [@CASCA #4]  SADLER, BARRY  C/P,P
PANZER SPIRIT  TOWNSEND, TOM  C/P,P
PAPER DOLLS, THE  DAVIES, L. P.  H,P
PAPER DRAGONS  BLAYLOCK, JAMES P.  B
PAPER GRAIL, THE  BLAYLOCK, JAMES P.  C/P:2,SFWA,P,H
PAPER MAGE  CUTTER, LEAH R.  C/P,P
PARA  ABRAMS, R. VAUGHN  H
PARA-ROBOT  CHARLES, NEIL  SR-P
PARABLE OF THE SOWER  BUTLER, OCTAVIA E.  C/P,P,H
PARABLE OF THE TALENTS  BUTLER, OCTAVIA E.  C/P,LP,H
PARADIGM OF EARTH, A  DORSEY, CANDAS JANE  LP
PARADIGM RED  KING, HAROLD  H
PARADIS [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  P
PARADISE  HENDERSON, DAN  P
PARADISE  RESNICK, MIKE  C/P,SFWA-P,P,H
PARADISE DRIFT [ANDROMEDA (GENE RODDENBERRY’S)]  SMITH, SHERWOOD  H
PARADISE GAME, THE [GRAINGER 4]  STABLEFORD, BRIAN M.  C/P,P
PARADISE IS NOT ENOUGH  ELDER, MICHAEL  C/P,P
PARADISE 1  HARRINGTON, ALAN  H
PARADISE PLOT, THE  NAHA, ED  P
PARADISE SNARE, THE [@STAR WARS HAN SOLO 1]  CRISPIN, A. C.  C/P,P
PARADISE TOWERS [DOCTOR WHO]  WYATT, STEPHEN  P
PARADISE TREE, THE  PAXSON, DIANA L.  P
PARADISE WAR, THE [SONG OF ALBION #1]  LAWHEAD, STEPHEN R.  H,@P
PARADOX  ERICKSON, LYNN  P
PARADOX [NULAPEIRON 1]  MEANEY, JOHN  P
PARADOX ALLEY [SKYWAY 3]  DE CHANCIE, JOHN  C/P,P
PARADOX GATE [GATES OF TIME 3]  PARKINSON, DAN  P
PARADOX LOST  BROWN, FREDRIC  H:2,SFWA,P
PARADOX MEN, THE  HARNESS, CHARLES L.  DD118
PARADOX MEN, THE (REVISED)  HARNESS, CHARLES L.  H
PARADOX OF THE SETS, THE [DAEDALUS 6]  STABLEFORD, BRIAN M.  C/P:2,P
PARADOX PLANET, THE  SPRUILL, STEVEN G.  C/P,P,H
PARADOXICON, THE  FALLETTA, NICHOLAS  SFWA
PARAFAITH WAR, THE  MODESITT, L. E. JR.  P
PARAGAEA  ROBERSON, CHRIS  LP
PARAGON LOST [@KING’S BLADES]  DUNCAN, DAVE  P
PARAGONS  WILSON, ROBIN  HA
PARALLAX RED [@OUTLANDERS]  AXLER, JAMES  P
PARALLEL 59 [DOCTOR WHO EIGHTH DOCTOR]  DALLAIRE, NATALIE/COLE, STEPHEN  P
PARALLEL MAN, THE  PURTILL, RICHARD  C/P:2,P
PARALLELITIES  FOSTER, ALAN DEAN  C/P,P,H
PARASAURIANS, THE  WELLS, ROBERT  P
PARASITE, THE  ASHER, NEAL  LP
PARASITE, THE  CAMPBELL, RAMSEY  H,SFWA-P,P
PARASITE [DOCTOR WHO NEW ADVENTURES]  MORTIMORE, JIM  C/P:2,P
PARASITE PLANET  WEINBAUM, STANLEY G.  B
PARASITE WAR, THE  SULLIVAN, TIM  C/P,P
PARATIME [PARATIME POLICE]  PIPER, H. BEAM  C/P:2,P
PARATWA, THE [PARATWA 3]  HINZ, CHRISTOPHER  SFWA
PARCHED SEA, THE [@FORGOTTEN REALMS HARPERS #1]  DENNING, TROY  C/P,P
PARIS IN THE TWENTIETH CENTURY  VERNE, JULES  H,LP
PARLIAMENT OF BLOOD  CIENCIN, SCOTT  P
PARSECS AND PARABLES  SILVERBERG, ROBERT  H
PARSIVAL OR A KNIGHT’S TALE [PARSIVAL 1]  MONACO, RICHARD  C/P,P,H
PARTICLE THEORY  BRYANT, EDWARD  C/P,P
PARTICOLORED UNICORN, THE  DE CLES, JON  P
PARTING OF ARWEN, THE  BRADLEY, MARION ZIMMER  B
PARTISAN [@HOSTILE TAKEOVER #2]  SWANN, S. ANDREW  P
PARTNERS [@LAWS OF THE BLOOD 2]  SIZEMORE, SUSAN  C/P,P
PARTNERS IN NECESSITY [LIADEN UNIVERSE 1,2,3]  LEE, SHARON/MILLER, STEVE  H
PARTNERS IN WONDER  ELLISON, HARLAN  C/P:2,P
PARTNERSHIP [SHELLPEOPLE 2]  MCCAFFREY, ANNE/BALL, MARGARET  C/P,SFWA-P,P,H
PARTNERSHIP (EXCERPT) [SHELLPEOPLE 2]  MCCAFFREY, ANNE/BALL, MARGARET  SR-LP
PAS DE GONIA POUR LES GHARKANDES  RICHARD-BESSIERE, F.  FP
PASHAZADE [EL ISKANDRYIA 1]  GRIMWOOD, JON COURTENAY  H
PASIPHAE  SMETHURST, WILLIAM  H
PASQUALE’S ANGEL  MCAULEY, PAUL J.  P
PASSAGE, THE  WARTOFSKY, VICTOR  C/P,P
PASSAGE  WILLIS, CONNIE  H,P
PASSAGE AT ARMS [STARFISHERS]  COOK, GLEN  H,P
PASSAGE OF STARS, A [@HIGHROAD #1]  RASMUSSEN, ALIS A.  C/P:2,P
PASSAGE OF THE LIGHT, THE  MALZBERG, BARRY N.  LP
PASSAGE TO DAWN [FORGOTTEN REALMS LEGACY OF THE DROW 4]  SALVATORE, R. A.  P

TITLE  AUTHOR  CODE
PASSAGE TO PLUTO  WALTERS, HUGH  H
PASSENGERS, THE [LOST IN SPACE NEW JOURNEYS]  KRULIK, NANCY  LP
PASSING FANCY [ALIEN NATION 6]  SPENCER, DAVID  C/P,P
PASSING FOR HUMAN  SCOTT, JODY  C/P,P
PASSING OF GLORIA MUNDAY, THE [@AVENGERS 3]  GARFORTH, JOHN  C/P,P
PASSING OF THE DRAGONS, THE  ROBERTS, KEITH  C/P,P
PASSING OF THE GODS, THE [FALL OF THE FIRST WORLD 3]  SMITH, DAVID C.  P
PASSING THROUGH THE FLAME  SPINRAD, NORMAN  P
PASSION FOR DEAD LEAVES [ENEMIES OF SOCIETY 3]  DAVID, JOHN  LP
PASSION OF MOLLY T., THE  SANDERS, LAWRENCE  P
PASSION OF NEW EVE, THE  CARTER, ANGELA  H,LP
PASSION PLAY  STEWART, SEAN  C/P:2,SFWA-P,P
PASSIONATE INVADERS, THE  CLARE, JOHN  H
PASSIONATE WITCH, THE  SMITH, THORNE  C/P,P
PASSIONELLA AND OTHER STORIES  FEIFFER, JULES  SCX
PASSPORT TO ETERNITY  BALLARD, J. G.  C/P:2,P
PAST IMPERATIVE [@GREAT GAME #1]  DUNCAN, DAVE  H,P
PAST IMPERFECT  GREENBERG, MARTIN H./SEGRIFF, LARRY  PA
PAST LIVES, PRESENT TENSE  SCARBOROUGH, ELIZABETH ANN  LPA
PAST MASTER  LAFFERTY, R. A.  C/P:2,P
PAST OF FOREVER, THE [@CHILDREN OF THE STARS #4]  COULSON, JUANITA  C/P,P
PAST, PRESENT AND FUTURE PERFECT  WOLF, JACK C./FITZ GERALD, GREGORY  PA
PAST PROLOGUE [@STAR TREK JANUS GATE #3]  GRAF, L. A.  P
PAST THE SIZE OF DREAMING [MATILDA BLACK 2]  HOFFMAN, NINA KIRIKI  P
PAST THROUGH TOMORROW, THE  HEINLEIN, ROBERT A.  C/P,P,H
PAST TIMES  ANDERSON, POUL  C/P:2,SFWA-P,P
PASTEL CITY, THE [VIRICONIUM 1]  HARRISON, M. JOHN  H:2,SFWA,P
PASTWATCH  CARD, ORSON SCOTT  C/P,P,H
PATCH OF THE ODIN SOLDIER, THE [LINCOLN BLACKTHORNE 3]  MARSH, GEOFFREY  H,SFWA,SFWA-P
PATCHWORK GIRL, THE [KNOWN SPACE]  NIVEN, LARRY  C/P,LP,H
PATCHWORK GIRL OF OZ, THE [OZ #7]  BAUM, L. FRANK  H,@P
PATH, THE [@HIGHLANDER]  NEASON, REBECCA  P
PATH BEYOND THE STARS, THE  PETAJA, EMIL  C/P:2,P
PATH INTO THE UNKNOWN: THE BEST OF SOVIET SCIENCE FICTION  DELACORTE PRESS  HA
PATH INTO THE UNKNOWN: THE BEST OF SOVIET SCIENCE FICTION  DELL BOOKS  PA,C/PA
PATH OF BLOOD [REISIL 3]  FRANCIS, DIANA PHARAOH  C/P,P
PATH OF DAGGERS, THE [WHEEL OF TIME 8]  JORDAN, ROBERT  H:2,P
PATH OF DESTRUCTION [@STAR WARS DARTH BANE]  KARPYSHYN, DREW  H
PATH OF EXOTERRA, THE  MCBAIN, GORDON  P
PATH OF FATE [REISIL 1]  FRANCIS, DIANA PHARAOH  C/P,P
PATH OF FIRE [@PATTERNS OF CHAOS #2]  INGRID, CHARLES  P
PATH OF GLORY [@BATTLETECH 49]  BILLS, RANDALL N.  P
PATH OF HONOR [REISIL 2]  FRANCIS, DIANA PHARAOH  C/P,P
PATH OF THE BOLD [SILVER AGE SENTINELS]  LOWDER, JAMES  LPA
PATH OF THE ECLIPSE [SAINT-GERMAIN 4]  YARBRO, CHELSEA QUINN  H,SFWA,SFWA-P,P
PATH OF THE FURY  WEBER, DAVID  C/P:2,SFWA-P,P
PATH OF THE HERO  WOLVERTON, DAVE  P
PATH OF THE JUST [SILVER AGE SENTINELS]  LOWDER, JAMES  LPA
PATH OF UNREASON, THE  SMITH, GEORGE O.  C/P,P,H
PATH TO CONQUEST [V]  WEINSTEIN, HOWARD  C/P,SFWA-P,P
PATH TO SAVAGERY  ALTER, ROBERT EDMOND  C/P,P
PATHFINDERS, THE [AUTUMN WORLD MOSAIC 2]  GRAVEL, GEARY  P
PATHLESS TRAIL, THE  FRIEL, ARTHUR O.  C/P,P
PATHLESS WAY, THE  NICHOLS, ADAM  H
PATHS NOT TAKEN [NIGHTSIDE 5]  GREEN, SIMON R.  P
PATHS OF EVIL [@GAUNTLET DARK LEGACY 1]  WHITE, RICHARD C.  P
PATHS OF THE DEAD, THE [DRAGAERA ADRILANKHA 1]  BRUST, STEVEN  H:2,P
PATHS OF THE PERAMBULATOR, THE [SPELLSINGER 5]  FOSTER, ALAN DEAN  C/P,P,H
PATHS TO OTHERWHERE  HOGAN, JAMES P.  H
PATHS TO OTHERWHERE (PROOF)  HOGAN, JAMES P.  SR-LP
PATHWAYS [@STAR TREK VOYAGER]  TAYLOR, JERI  C/P,P
PATRIAN TRANSGRESSION, THE [STAR TREK]  HAWKE, SIMON  C/P:2,P
PATRIARCH’S HOPE [NICHOLAS SEAFORT 6]  FEINTUCH, DAVID  C/P,P,H
PATRICIA, THE BEAUTIFUL [@PHENWICK #4]  KIMBROUGH, KATHERYN  P
PATRICK SON OF IRELAND  LAWHEAD, STEPHEN R.  H
PATRIMONY, THE [HORSECLANS 6]  ADAMS, ROBERT  C/P:2,P
PATRIOTS  DRAKE, DAVID  C/P:2,P
PATRIOTS AND TYRANTS [@BATTLETECH 52]  COLEMAN, LOREN L.  P
PATRIOTS IN ARMS [SCOTT ST. ANDREW 3]  WEAVER, BEN  P
PATRIOT’S STAND [@MECHWARRIOR DARK AGE/BATTLETECH]  MOSCOE, MIKE  P
PATRON OF THE ARTS  ROTSLER, WILLIAM  P
PATRON SAINT OF PLAGUES, THE  ANDERSON, BARTH  LP
PATTERN [@SCAVENGER #2]  PARKER, K. J.  LP
PATTERN FOR CONQUEST  SMITH, GEORGE O.  H

The MITSFS Titledex  September 24, 2007  Page 266
TITLE  AUTHOR  CODE
PATTERN RECOGNITION  GIBSON, WILLIAM  C/P,LP,H
PATTERNMASTER  BUTLER, OCTAVIA E.  C/P,SFWA,P,H
PATTERNS  CADIGAN, PAT  LP
PATTERNS OF CHAOS  KAPP, COLIN  C/P:2,P
PATTERNS OF CHAOS, VOLUME ONE, THE [PATTERNS OF CHAOS 1,2]  INGRID, CHARLES  P
PATTERNS OF CHAOS, VOLUME TWO, THE [PATTERNS OF CHAOS 3,4]  INGRID, CHARLES  P
PATTON’S SPACESHIP [@TIMELINE WARS #1]  BARNES, JOHN  C/P:2,P
PAUSE TO WONDER  FISCHER, MARJORIE/HUMPHRIES, ROLFE  HA
PAVANE  ROBERTS, KEITH  C/P:2,P
PAWN OF CHAOS [ETERNAL CHAMPION]  MOORCOCK, MICHAEL  LPA
PAWN OF PROPHECY [BELGARIAD 1]  EDDINGS, DAVID  C/P:2,P
PAWN OF THE OMPHALOS  TUBB, E. C.  C/P,P
PAWN OF TIME  CARSON, ROBIN  H
PAWN TO INFINITY  SABERHAGEN, FRED/SABERHAGEN, JOAN  C/PA,PA
PAWNS AND SYMBOLS [STAR TREK]  LARSON, MAJLISS  C/P:2,P
PAWN’S DREAM  NYLUND, ERIC S.  P
PAWNS OF CHAOS [WARHAMMER 40K]  CRAIG, BRIAN  P
PAYCHECK  DICK, PHILIP K.  LP
PAYING THE PIPER [HAMMER’S SLAMMERS]  DRAKE, DAVID  P
PEACE  WOLFE, GENE  C/P,SFWA,P,H
PEACE & MEMORY [SECANTIS 3]  TIEDEMANN, MARK W.  LP
PEACE COMPANY [PEACE COMPANY]  GREEN, ROLAND J.  P
PEACE ON EARTH  LEM, STANISLAW  H
PEACE WAR, THE  VINGE, VERNOR  C/P:2,SFWA,P,H:3
PEACEKEEPERS  BOVA, BEN  C/P:2,SFWA-P,P
PEACEKEEPERS, THE [STAR TREK TNG]  DEWEESE, GENE  C/P:3,P
PEACEMAKERS, THE  CASEWIT, CURTIS W.  C/P,P
PEARL OF PATMOS [@BLADE #7]  LORD, JEFFREY  C/P,P
PEARL OF THE SOUL OF THE WORLD, THE [DARKANGEL 3]  PIERCE, MEREDITH ANN  P
PEARLS FROM PEORIA  FARMER, PHILIP JOSE  H
PEARLS OF LUTRA [REDWALL]  JACQUES, BRIAN  P
PEBBLE IN THE SKY [GALACTIC EMPIRE 3]  ASIMOV, ISAAC  C/P,P,H,GN3.14
PECULIAR LIVES [@TIME HUNTER 7]  PURSER-HALLARD, PHILIP  LP
PECULIAR MISS PICKETT, THE  JULIAN, NANCY R.  LP
PEDESTAL, THE  LANNING, GEORGE  P
PEEPS  WESTERFELD, SCOTT  LP
PEGASUS IN FLIGHT [PEGASUS 2]  MCCAFFREY, ANNE  H:2,SR-H,P
PEGASUS IN SPACE [PEGASUS 3]  MCCAFFREY, ANNE  H:2,P
PEGGY, THE CONCERNED [@PHENWICK #31]  KIMBROUGH, KATHERYN  P
PELLUCIDAR [PELLUCIDAR #2]  BURROUGHS, EDGAR RICE  C/P,@P,H
PELOS CAMINHOS DO ESPACO  RICHARD-BESSIERE, F.  FP
PENAL COLONY, THE  HERLEY, RICHARD  P
PENDRAGON, THE  CHRISTIAN, CATHERINE  P
PENDRAGON [PENDRAGON 4]  LAWHEAD, STEPHEN R.  C/P,P
PENDRAGON CHRONICLES, THE  ASHLEY, MIKE  HA
PENDULUM  CHRISTOPHER, JOHN  H:2,P
PENDULUM  VAN VOGT, A. E.  C/P:2,P
PENELOPE  ANDERSON, WILLIAM C.  C/P,P
PENGUIN BOOK OF VAMPIRE STORIES, THE  RYAN, ALAN  LPA
PENGUIN BOOK OF WITCHES AND WARLOCKS, THE  KAYE, MARVIN  HA,LPA
PENGUIN ISLAND  FRANCE, ANATOLE  H
PENGUIN SCIENCE FICTION  ALDISS, BRIAN W.  SR-PA
PENGUIN SCIENCE FICTION OMNIBUS, THE  ALDISS, BRIAN W.  C/PA,PA
PENNANT FOR THE KREMLIN, A  MOLLOY, PAUL  P
PENNTERRA  MOFFETT, JUDITH  H,P
PENULTIMATE TRUTH, THE  DICK, PHILIP K.  C/P:2,P
PENUMBRA  BROWN, ERIC  P
PEOPLE BEYOND THE WALL, THE  TALL, STEPHEN  C/P,P
PEOPLE EXCHANGE, THE  BAYLUS, ROBERT F.  SR-P,SFWA-P
PEOPLE MACHINES  WILLIAMSON, JACK  C/P,P
PEOPLE MAKER, THE  KNIGHT, DAMON  SR-P
PEOPLE MINUS X  GALLUN, RAYMOND Z.  DD291
PEOPLE: NO DIFFERENT FLESH, THE  HENDERSON, ZENNA  C/P:2,P
PEOPLE OF ASA  ASHTON, MARVIN  SR-P
PEOPLE OF THE AX, THE  WILLIAMS, JAY  C/P,P
PEOPLE OF THE BLACK CIRCLE, THE [BERKLEY CONAN 3]  HOWARD, ROBERT E.  P
PEOPLE OF THE COMET  HALL, AUSTIN  SR-H,H
PEOPLE OF THE DARK  HOWARD, ROBERT E.  H
PEOPLE OF THE DARK, THE  WRIGHT, T. M.  SFWA-P:2

PEOPLE OF THE EARTH [PEOPLE OF THE WOLF 1] GEAR, W. MICHAEL/

GEAR, KATHLEEN O’NEAL

C/P,P

PEOPLE OF THE FIRE [PEOPLE OF THE WOLF 2] GEAR, W. MICHAEL/

GEAR, KATHLEEN O’NEAL

SFWA-P

PEOPLE OF THE MIST, THE  HAGGARD, H. RIDER  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 267
TITLE  AUTHOR  CODE
PEOPLE OF THE RUINS, THE  SHANKS, EDWARD  H

PEOPLE OF THE SEA GEAR, W. MICHAEL/

GEAR, KATHLEEN O’NEAL

H

PEOPLE OF THE SKY [RATHA 5]  BELL, CLARE  C/P,SFWA-P,P,H
PEOPLE OF THE SOUTH POLE  THOMSON, K. GRAHAM  H
PEOPLE OF THE TALISMAN  BRACKETT, LEIGH  DM101
PEOPLE OF THE TWILIGHT  KANER, H.  H:2
PEOPLE OF THE WIND, THE  ANDERSON, POUL  C/P:2,SFWA-P,P
PEOPLE OF THE WIND, THE/DAY OF THEIR RETURN, THE  ANDERSON, POUL  P
PEOPLE THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 2]  BURROUGHS, EDGAR RICE  C/P:2,P
PEOPLE TRAP, THE  SHECKLEY, ROBERT  C/P:2,P
PEOPLE TRAP, THE/MINDSWAP  SHECKLEY, ROBERT  C/P,P
PEOPLES OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 12]  TOLKIEN, J. R. R.  H
PERCHANCE  KURLAND, MICHAEL  P
PERCHANCE TO DREAM  KNIGHT, DAMON  C/PA,PA,SFWA
PERCHANCE TO DREAM  LITTLE, DENISE  PA
PERCHANCE TO DREAM  REYNOLDS, MACK  C/P:2,P
PERCHANCE TO DREAM  STACKPOLE, MICHAEL A.  H
PERCHANCE TO DREAM [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:2,P
PERCIVAL’S ANGEL  CROMPTON, ANNE ELIOT  P
PERDIDO NO ESPACO 1  CAIDIN, MARTIN  FP
PERDIDO NO ESPACO 2  CAIDIN, MARTIN  FP
PERDIDO STREET STATION  MIEVILLE, CHINA  C/P,LP,H
PEREGRINE, THE  ANDERSON, POUL  C/P:2,P
PEREGRINE: PRIMUS  DAVIDSON, AVRAM  H:2,SFWA,P
PEREGRINE: SECUNDUS  DAVIDSON, AVRAM  P
PERELANDRA [PERELANDRA 2]  LEWIS, C. S.  C/P:2,P,H
PERFECT CIRCLE  STEWART, SEAN  H
PERFECT HOST, THE [COMPLETE STORIES OF THEODORE STURGEON 5]  STURGEON, THEODORE  H
PERFECT LOVER, THE  PRIEST, CHRISTOPHER  C/P,P,H
PERFECT PEOPLE  LIEBERMAN, ROBERT  P
PERFECT PLANET, THE  SMITH, EVELYN E.  C/P,P,H
PERFECT PRINCESS, THE [DRAGON NIMBUS 2]  RADFORD, IRENE  C/P:2,P
PERFECT TIMING [DOCTOR WHO]  PHIPPEN, MARK/FAYLE, HELEN  SR-LPA
PERFECT TIMING 2 [DOCTOR WHO]  FAYLE, HELEN/EALES, JULIAN  SR-LPA
PERFECT VACUUM, A  LEM, STANISLAW  LP
PERFECTION OF YOGA, THE  PRABHUPADA, SWAMI  [C/RAN-P]
PERFUME  SUSKIND, PATRICK  P
PERFUMERY TECHNOLOGY  BILLOT, MARCEL/WELLS, F. V.  [C/RAN-H]
PERIHELION [@ROBOT CITY #6]  WU, WILLIAM F.  C/P:2,SFWA-P,P
PERIL OF BARNABAS COLLINS, THE  ROSS, MARILYN  P
PERIL OF THE STARMEN  NEVILLE, KRIS  SR-DB0759
PERIL ON ICE PLANET [@PERRY RHODAN #23]  ACE  PA,C/PA
PERIL UNLIMITED [PERRY RHODAN 130]  BRAND, KURT  SR-LP
PERILOUS DREAMS  NORTON, ANDRE  C/P:2,P
PERILOUS GALAXY  MULLER, JOHN E.  SR-P
PERILOUS GARD, THE  POPE, ELIZABETH MARIE  C/P,P
PERILOUS PLANETS  ALDISS, BRIAN W.  C/PA,PA
PERILOUS SEAS [MAN OF HIS WORD 3]  DUNCAN, DAVE  C/P,P
PERIL’S GATE [WARS OF LIGHT AND SHADOW 5/ALLIANCE OF LIGHT 3]  WURTS, JANNY  H
PERMANENCE  SCHROEDER, KARL  H
PERMANENT IMPLOSION, THE  CAMPBELL, JOHN W. JR.  C/PA,PA
PERMUTATION CITY  EGAN, GREG  C/P:2,P
PERPETUAL LIGHT  RYAN, ALAN  HA,PA
PERPETUITY BLUES AND OTHER STORIES  BARRETT, NEAL JR.  H

PERRY RHODAN: DIE GROESSTE SCIENCE FICTION-SERIE DER WELT (CD)

[PERRY RHODAN] PABEL MOEWIG KG

SR-MM

PERRY’S PLANET [STAR TREK]  HALDEMAN, JACK C. II  C/P,P
PERSEA [IROSHI 3]  OSBORNE, CARY  P
PERSECUCION COSMICA  CLEMENT, HAL  FP
PERSEIDS AND OTHER STORIES, THE  WILSON, ROBERT CHARLES  H
PERSEUS BREED, THE  EGAN, KEVIN  C/P,P
PERSEUS SPUR [RAMPART WORLDS 1]  MAY, JULIAN  C/P,P
PERSIAN, THE [@CASCA #6]  SADLER, BARRY  P
PERSIMMION SEQUENCE, THE  FRETLAND, DON J.  SR-P,C/P
PERSISTENCE OF MEMORY, THE [@SLOW WORLD #1]  RIPLEY, KAREN  P
PERSISTENCE OF VISION, THE  VARLEY, JOHN  H:2,P
PERSONAL AGENDAS [@BABYLON 5 #8]  SARRANTONIO, AL  C/P,P
PERSONAL DARKNESS [BLOOD OPERA 2]  LEE, TANITH  H

PERSONAL DEMON, A BISCHOFF, DAVID F./BROWN, RICH/

RICHARDSON, LINDA

C/P:2,P

PERSONALITY OF ANIMALS, THE  FOX, H. MUNRO  [C/RAN-P]
PERSONIFIED PROJECT, THE  BARTLETT, R. E.  LP
PERTURBED SPIRITS  BULL, R. C.  PA

The MITSFS Titledex  September 24, 2007  Page 268
TITLE  AUTHOR  CODE
PESCATONS, THE [@DOCTOR WHO]  PEMBERTON, VICTOR  P
PESHAWAR LANCERS, THE  STIRLING, S. M.  H,P
PESTIS 18  WEBB, SHARON  C/P,P
PET, THE  GRANT, CHARLES L.  P,SFWA-P
PET PEEVE [@XANTH #29]  ANTHONY, PIERS  P
PET SEMATARY  KING, STEPHEN  C/P,SFWA,P,H
PETER AND THE SHADOW THIEVES [PETER 2]  BARRY, DAVE/PEARSON, RIDLEY  H
PETER AND THE STARCATCHERS [PETER 1]  BARRY, DAVE/PEARSON, RIDLEY  H
PETER DAVISON’S BOOK OF ALIEN PLANETS  DAVISON, PETER  SR-PA
PETER PAN  BARRIE, JAMES M.  P
PETER S. BEAGLE’S IMMORTAL UNICORN  BEAGLE, PETER S./BERLINER, JANET  HA
PETER S. BEAGLE’S IMMORTAL UNICORN, VOLUME 1  BEAGLE, PETER S./BERLINER, JANET  PA
PETER S. BEAGLE’S IMMORTAL UNICORN, VOLUME 2  BEAGLE, PETER S./BERLINER, JANET  PA
PETRIFIED PLANET, THE  STATTEN, VARGO  SR-P
PETRIFIED PLANET, THE  TWAYNE BOOKS  HA
PETROGYPSIES  HARPER, RORY  P,SFWA-P
PETTY PEWTER GODS [GARRETT 8]  COOK, GLEN  C/P:2,P
PFLANZEN DES TODES, DIE [@PERRY RHODAN #43]  KNEIFEL, HANS  FPR
PHAEDRA  BRINDEL, JUNE RACHUY  H
PHAETON CONDITION, THE  MASON, DOUGLAS R.  H,P
PHAID THE GAMBLER [PHAID 1]  FARREN, MICK  C/P:2,P
PHANTASIES AND LILITH  MACDONALD, GEORGE  LP
PHANTASMAGORIA  MOBLEY, JANE  PA
PHANTASTES  MACDONALD, GEORGE  C/P:2,P
PHANTOM [SWORD OF TRUTH 11]  GOODKIND, TERRY  H
PHANTOM  KAY, SUSAN  P
PHANTOM AND BARNABAS COLLINS, THE  ROSS, MARILYN  P
PHANTOM AND THE SCORPIA MENACE, THE [PHANTOM 3]  COPPER, BASIL  P
PHANTOM BANJO [@SONGKILLER #1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
PHANTOM CITY, THE [DOC SAVAGE #10]  ROBESON, KENNETH  C/P,@P
PHANTOM FLEET [@PERRY RHODAN #97]  ACE  PA

PHANTOM HORDE [PERRY RHODAN 137]/ STAR MAN #1 THROUGH #5 [PERRY RHODAN 137] MASTER PUBLICATIONS

SR-LP

PHANTOM MENACE, THE [STAR WARS MOVIE 1]  BROOKS, TERRY  H:2,P
PHANTOM OF THE OPERA, THE  LEROUX, GASTON  H,P
PHANTOM ONES, THE  TORRO, PEL  SR-P
PHANTOM REGIMENTS  ADAMS, ROBERT  C/PA,PA,SFWA-P
PHANTOM RICKSHAW AND OTHER TALES, THE  KIPLING, RUDYARD  SR-H
PHANTOM SHARK, THE [@RICK BRANT #6]  BLAINE, JOHN  H:2
PHANTOM-STATION [@PERRY RHODAN #16]  EWERS, H. G.  FPR
PHANTOM TOLLBOOTH, THE  JUSTER, NORTON  H:2,LP
PHANTOME UNTER UNS  BROLL, W. W.  FP

PHANTOMS GREENBERG, MARTIN H./

GREENBERG, ROSALIND M.

PA

PHANTOMS AND FANCIES  DE CAMP, L. SPRAGUE  SR-H
PHANTOMS ON THE WIND [@DEMON CROWN #2]  VARDEMAN, ROBERT E.  P,SFWA-P
PHARAOH CONTRACT, THE [@EMANCIPATOR #1]  ALDRIDGE, RAY  C/P:2,P

PHAROAH FANTASTIC GREENBERG, MARTIN H./ KOREN, BRITTANY A.

PA

PHAROAH’S GHOST, THE [DOC SAVAGE #101]/ TIME TERROR, THE [DOC SAVAGE 102] ROBESON, KENNETH

@P

PHASE IV  MALZBERG, BARRY N.  P
PHASE SPACE [REID MALENFANT]  BAXTER, STEPHEN  H
PHASE TWO  RICHMOND, WALT/RICHMOND, LEIGH  C/P,P
PHASES  MOON, ELIZABETH  C/P:2,P
PHASES IN CHAOS [@ISAAC’S UNIVERSE #2]  GREENBERG, MARTIN H.  PA
PHASES OF GRAVITY  SIMMONS, DAN  C/P:2,P
PHAZE DOUBT [APPRENTICE ADEPT #7]  ANTHONY, PIERS  C/P,@P,SFWA-P:2,H
PHENDRIC [THREE LANDS 3]  EVANS, D. R.  LP
PHENOMENA X  MULLER, JOHN E.  SR-P
PHENOMENON: FORTY YEARS OF FLYING SAUCERS  SPENCER, JOHN/EVANS, HILARY  [C/RAN-P]

PHILIP WALLER, MARK AERIAL/CHONG, HEMAN/

COSTINAS, COSMIN/HEATHER, ROSEMARY/ MCKEE, FRANCIS/REINFURT, DAVID/ RUSHTON, STEVE/TANGEN, LEIF MAGNE

SR-P,C/P

PHILIP JOSE FARMER  BRIZZI, MARY T.  R/REF-P
PHILIP K. DICK  FBI  R/REF-H

PHILIP K. DICK OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

R/REF-H

PHILIP K. DICK: ELECTRIC SHEPHERD  GILLESPIE, BRUCE  R/REF-P
PHILIP K. DICK HIGH  BISCHOFF, DAVID F.  LP
PHILIP K. DICK IS DEAD, ALAS  BISHOP, MICHAEL  LP
PHILOSOPHER’S STONE, THE  LARSEN, J. ANKER  H
PHILOSOPHER’S STONE, THE  LITTLE, JANE  H

TITLE  AUTHOR  CODE
PHILOSOPHER’S STONE, THE  WILSON, COLIN  C/P,P,H
PHILOSOPHICAL CORPS, THE  COLE, E. B.  H:2
PHILOSOPHICAL INVESTIGATION, A  KERR, PHILIP  SFWA
PHILOSOPHICAL STRANGLER, THE [JOE’S WORLD 1]  FLINT, ERIC  H,P
PHILOSOPHY AND LIVING 2  STAPLEDON, OLAF  [C/RAN-P]
PHOBOS  DRAGO, TY  H
PHOEBE WHO? [@CHARMED]  HARRISON, EMMA  P
PHOENIX [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:2,SFWA-P:2,P
PHOENIX  COWPER, RICHARD  H:2,P
PHOENIX, THE [@CASCA #14]  SADLER, BARRY  P
PHOENIX, THE [FIVE RINGS CLAN WAR 4]  SULLIVAN, STEPHEN D.  P
PHOENIX AND ASHES [ELEMENTAL MASTERS 3]  LACKEY, MERCEDES  H:2
PHOENIX AND THE CARPET, THE  NESBIT, E.  H
PHOENIX AND THE MIRROR, THE  DAVIDSON, AVRAM  C/P,P,H
PHOENIX BELLS, THE [TEN THOUSAND WILLOWS 1]  GRANT, KATHRYN  P,SFWA-P
PHOENIX CAFE [ALEUTIAN 3]  JONES, GWYNETH  LP
PHOENIX CODE, THE  ASARO, CATHERINE  P
PHOENIX EXULTANT, THE [GOLDEN AGE 2]  WRIGHT, JOHN C.  H
PHOENIX FIRE  FORREST, ELIZABETH  P,SFWA-P
PHOENIX GUARDS, THE [DRAGAERA KHAAVREN 1]  BRUST, STEVEN  C/P,P,H
PHOENIX IN FLIGHT, THE [EXORDIUM 1]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
PHOENIX IN OBSIDIAN [ETERNAL CHAMPION 2]  MOORCOCK, MICHAEL  SR-P
PHOENIX IN THE ASHES  VINGE, JOAN D.  H,SFWA-P,P
PHOENIX MAN, THE [@CHILL #5]  SHERMAN, JORY  P
PHOENIX OF MEGARON [SPACE: 1999 10]  RANKINE, JOHN  P
PHOENIX PRIME  WHITE, TED  C/P:2,P
PHOENIX SEDUCTION, THE  PARLIER, CAP  P
PHOENIX SHIP  RICHMOND, WALT/RICHMOND, LEIGH  D66160
PHOENIX TREE, THE  BOYER, ROBERT H./ZAHORSKI, KENNETH J. C/PA,PA
PHOENIX WITHOUT ASHES  BRYANT, EDWARD/ELLISON, HARLAN  C/P:2,P
PHOTOMESIS  CAMERON, BERL  SR-P
PHTHOR [CHTHON #2]  ANTHONY, PIERS  C/P:2,@P
PHULE AND HIS MONEY, A [PHULE 3]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P,P
PHULE ME TWICE [PHULE 4]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P:2,P
PHULE’S COMPANY [PHULE 1]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P
PHULE’S ERRAND [PHULE 6]  ASPRIN, ROBERT LYNN/HECK, PETER J.  C/P,P
PHULE’S PARADISE [PHULE 2]  ASPRIN, ROBERT LYNN  C/P:2,SFWA-P:2,P
PHYLOGENESIS [HUMANX FOUNDING COMMONWEALTH 1]  FOSTER, ALAN DEAN  H
PHYLUM MONSTERS  PEIRCE, HAYFORD  C/P,SFWA-P,P
PHYSICS OF STAR TREK, THE [STAR TREK]  KRAUSS, LAWRENCE M.  R/REF-P,C/REF-H
PHYSIOGNOMY, THE [WELL-BUILT CITY 1]  FORD, JEFFREY  C/P:2,LP
PHYTOSPHERE  MACKAY, SCOTT  P
PIANETA NELL’OCCHIO DEL TEMPO, IL  EARNSHAW, BRIAN  FP
PICKING THE BALLAD’S BONES [@SONGKILLER #2]  SCARBOROUGH, ELIZABETH ANN  C/P,P
PICKUP ARTIST, THE  BISSON, TERRY  H
PICNIC ON NEARSIDE  VARLEY, JOHN  C/P,P
PICNIC ON PARADISE  RUSS, JOANNA  C/P:2,P
PICOVERSE  METZGER, ROBERT A.  H
PICTORIAL HISTORY OF HORROR MOVIES, A  GIFFORD, DENIS  R/REF-H
PICTORIAL HISTORY OF SCIENCE FICTION FILMS, A  ROVIN, JEFF  R/REF-H
PICTURE OF DORIAN GRAY, THE  WILDE, OSCAR  C/P,P,H
PICTURE PERFECT [@CHARMED]  DOKEY, CAMERON  P
PICTURES AT 11  SPINRAD, NORMAN  C/P,P
PICTURES AT AN EXHIBITION  WATSON, IAN  SR-PA
PICTURES FROM AN EXPEDITION  IRVINE, ALEXANDER C.  LP
PICTURES OF DEATH [@AVENGER #19]  ROBESON, KENNETH  C/P:2,P
PICTURES OF PAVANNE, THE  WRIGHT, LAN  DH048
PIECE OF MARTIN CANN, A  JANIFER, LAURENCE M.  C/P,P
PIECE OF RESISTANCE, A  EGLETON, CLIVE  P
PIECE OF THE ACTION, A [STAR TREK]  HARMON, DAVID P./COON, GENE L.  SR-S
PIECES OF EIGHT [TIME TRAP 3]  DALTON, SEAN  P
PIED PIPER [@CHARMED]  VIGUIE, DEBBIE  P
PIERCING, THE  COYNE, JOHN  P
PIERS ANTHONY’S VISUAL GUIDE TO XANTH [XANTH]  ANTHONY, PIERS/NYE, JODY LYNN  R/REF-P,C/REF-H:2
PIG PLANTAGENET, THE  ANDREWS, ALLEN  C/P,P
PIG, THE PRINCE, AND THE UNICORN, THE  BRUSH, KAREN A.  P
PIG WORLD  RUNYON, CHARLES W.  H,SFWA,P
PIGEONS FROM HELL  HOWARD, ROBERT E.  C/P:2,P
PIGS DON’T FLY [UNEXPECTED DRAGON 1]  BROWN, MARY  C/P:2,P
PILGERMANN  HOBAN, RUSSELL  LP
PILGRIM [WAYFARER REDEMPTION 5]  DOUGLASS, SARA  H
PILGRIM [PILGRIM 1,2]  SABERHAGEN, FRED  P
PILGRIM PROJECT, THE  SEARLS, HANK  C/P:2,P
PILGRIM STARS [@WING COMMANDER]  TELEP, PETER  P

The MITSFS Titledex  September 24, 2007  Page 270
TITLE  AUTHOR  CODE
PILGRIMAGE  HENDERSON, ZENNA  C/P:2,P
PILGRIMAGE  MENDELSON, DREW  C/P,P
PILGRIMAGE TO EARTH  SHECKLEY, ROBERT  C/P:2,P
PILGRIMAGE TO HELL [@DEATHLANDS]  ADRIAN, JACK  C/P,P
PILGRIM’S REGRESS, THE  LEWIS, C. S.  P
PILGRIMS THROUGH SPACE AND TIME  BAILEY, J. O.  R/REF-H,C/REF-H
PILLAR OF FIRE AND OTHER PLAYS  BRADBURY, RAY  P

PILLAR OF FLAME, THE [@DRAGONLANCE NEW ADVENTURES ELEMENTS #1] SOESBEE, REE

LP

PILLAR OF NIGHT [CENOTAPH ROAD 6]  VARDEMAN, ROBERT E.  C/P:2,P
PILLAR OF THE SKY  HOLLAND, CECELIA  H
PILLAR OF THE SUN [@EXALTED #4]  BOWEN, CARL  P
PILLARS OF CREATION, THE [SWORD OF TRUTH 7]  GOODKIND, TERRY  H:2
PILLARS OF ETERNITY, THE  BAYLEY, BARRINGTON J.  P
PILLARS OF PENTEGARN  ESTES, ROSE  [C/RAN-P]
PILLARS OF SALT  PAUL, BARBARA  C/P,P
PILLARS OF THE WORLD, THE [TIR ALAINN 1]  BISHOP, ANNE  C/P,P
PILLOW FRIEND, THE  TUTTLE, LISA  H
PILOT ERROR [@SIX MILLION DOLLAR MAN #4]  BARBREE, JAY  P
PILOTS CHOICE [LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  H,LP
PIMPERNEL PLOT, THE [TIMEWARS 3]  HAWKE, SIMON  C/P:2,P
PINDHAREE  RICHARDS, JOEL  P,SFWA-P
PINHOE EGG, THE [CHRESTOMANCI]  JONES, DIANA WYNNE  H
PINOCCHIO  COLLODI, CARLO  H
PIONEER, 1990  STATTEN, VARGO  SR-P
PIONEERING THE SPACE FRONTIER  NATIONAL COMMISSION ON SPACE  R/REF-H
PIONEERS  ARGO, JOHN  LP
PIPER AT THE GATE  STANTON, MARY  C/P,SFWA-P,P
PIPES OF ORPHEUS, THE  LINDSKOLD, JANE  P
PIRATE, THE [@CASCA #15]  SADLER, BARRY  P
PIRATE ISLE [DOC SAVAGE #115]/SPEAKING STONE, THE [DOC SAVAGE 116] ROBESON, KENNETH  @P
PIRATE OF THE PACIFIC [DOC SAVAGE #19]  ROBESON, KENNETH  C/P,@P
PIRATE OF WORLD’S END, THE [GONDWANE 5]  CARTER, LIN  P
PIRATE PRINCE [SKYRIDER 4]  MICHAELS, MELISA C.  P,SFWA-P
PIRATE’S GHOST, THE [DOC SAVAGE #62]  ROBESON, KENNETH  C/P:2,@P
PIRATES OF CERBERUS  MISTRAL, BENGO  SR-P
PIRATES OF GOHAR [@BLADE #32]  LORD, JEFFREY  P
PIRATES OF ROSINANTE, THE [ROSINANTE 3]  GILLILAND, ALEXIS A.  C/P:2,P
PIRATES OF SHAN, THE [@RICK BRANT #14]  BLAINE, JOHN  H
PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  ASIMOV, ISAAC  C/P,P
PIRATES OF THE THUNDER [@RINGS OF THE MASTER #2]  CHALKER, JACK L.  C/P:2,P
PIRATES OF THE UNIVERSE  BISSON, TERRY  H,LP
PIRATES OF VENUS [VENUS #1]  BURROUGHS, EDGAR RICE  C/P:2,@P
PIRATES OF ZAN, THE  LEINSTER, MURRAY  D66525,DD403
PIT, THE [DOCTOR WHO NEW ADVENTURES]  PENSWICK, NEIL  C/P:2,P
PIT DRAGON TRILOGY, THE [PIT DRAGONS 1,2,3]  YOLEN, JANE  H
PIT PLANET  DVORKIN, DAVID  H
PITCH BLACK  LAURIA, FRANK  P
PITFALL [ALIEN SPEEDWAY 2]  WYLDE, THOMAS  C/P:2,P
PITFALL (PROOF) [ALIEN SPEEDWAY 2]  WYLDE, THOMAS  SR-LP
PITY ABOUT EARTH  HILL, ERNEST  DH056
PIXEL EYE, THE [PHIL D’AMATO 3]  LEVINSON, PAUL  H
PIXEL JUICE  NOON, JEFF  P

PIXILATED PEERESS, THE DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

C/P,P,H

PKD: A PHILIP K. DICK BIBLIOGRAPHY  LEVACK, DANIEL J. H.  R/IND
PLACE, THE  WRIGHT, T. M.  SFWA-P
PLACE AMONG THE FALLEN, A [@OMARAN #1]  COLE, ADRIAN  H,SFWA,P
PLACE BEYOND MAN, A  NEEPER, CARY  C/P,SFWA,P,H
PLACE OF THE LION, THE  WILLIAMS, CHARLES  LP
PLACE SO FOREIGN, A  DOCTOROW, CORY  H,LP
PLACEBO EFFECT [DOCTOR WHO EIGHTH DOCTOR]  RUSSELL, GARY  C/P:2,P
PLAGUE  MASTERTON, GRAHAM  P
PLAGUE DAEMON [WARHAMMER ORFEO 2]  CRAIG, BRIAN  LP
PLAGUE DOGS, THE  ADAMS, RICHARD  H,P
PLAGUE FROM SPACE  HARRISON, HARRY  C/P,P,H
PLAGUE IN PARADISE, A [@MASTERS OF SPACE #3]  VARDEMAN, ROBERT E.  P
PLAGUE LORDS OF RUEL, THE [LONE WOLF 13]  DEVER, JOE  P
PLAGUE OF ALL COWARDS, A  BARTON, WILLIAM  C/P:2,P
PLAGUE OF ANGELS, A  TEPPER, SHERI S.  C/P:2,P,H
PLAGUE OF CHANGE, A  DOUGLAS, L. WARREN  P
PLAGUE OF DEMONS, A  LAUMER, KEITH  C/P:2,P
PLAGUE OF ICE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
PLAGUE OF KNIVES [DEATH DEALER 4]  SILKE, JAMES  C/P,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 271
TITLE  AUTHOR  CODE
PLAGUE OF MEMORY [@STARDOC]  VIEHL, S. L.  P
PLAGUE OF NIGHTMARES, A [@DREAM LORDS #1]  COLE, ADRIAN  C/P:2,P
PLAGUE OF OBLIVION, THE [@PERRY RHODAN #28]  ACE  PA,C/PA
PLAGUE OF PYTHONS, A  POHL, FREDERIK  C/P:2,P
PLAGUE OF SILENCE, THE  CREASEY, JOHN  P
PLAGUE OF SOUND, THE [FLASH GORDON 2]  STEFFANSON, CON  P
PLAGUE SHIP [SOLAR QUEEN 2]  NORTH, ANDREW  DD345
PLAGUE SHIP [SOLAR QUEEN 2]  NORTON, ANDRE  C/P,P,H
PLAGUE TALES, THE  BENSON, ANN  P
PLAGUE YEARS [GUARDIANS 12]  AUSTIN, RICHARD  P
PLAINS OF PASSAGE, THE [EARTH’S CHILDREN 4]  AUEL, JEAN M.  P
PLAINS OF THE SEA, THE [@WILDERNESS OF FOUR #2]  HANCOCK, NIEL  C/P,P
PLAN B [LIADEN UNIVERSE 4]  LEE, SHARON/MILLER, STEVE  H:2,LP
PLAN FOR CONQUEST  GLYNN, A. A.  SR-P,H
PLAN S.03  RAYJEAN, MAX-ANDRE  FP
PLAN(E)T ENGINEERING  WOLFE, GENE  SR-H
PLANAR CHAOS [@MAGIC: THE GATHERING TIME SPIRAL #2]  MCGOUGH, SCOTT/SANDERS, TIMOTHY  P
PLANESHIFT [@MAGIC: THE GATHERING INVASION 2]  KING, J. ROBERT  C/P,P
PLANESWALKER [MAGIC: THE GATHERING ARTIFACTS 2]  ABBEY, LYNN  P
PLANET AMERICA [@STARHAWK 2]  MALONEY, MACK  P
PLANET BIG ZERO  HADLEY, FRANKLIN  C/P:2,P
PLANET BUYER, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P:2,P
PLANET CALLED TREASON, A  CARD, ORSON SCOTT  H:2,P
PLANET CALLED UTOPIA, A  MCINTOSH, J. T.  P
PLANET DREAMS  CARLOCK, MICHAELA  LP
PLANET DWELLER, THE [PLANET DWELLER 1]  PALMER, JANE  P
PLANET EXPLORER, THE  LEINSTER, MURRAY  P
PLANET FEDERATION  SHAW, DAVID  SR-P
PLANET FINDERS  DERMOTT, VERN  C/P,P,H
PLANET FOR TEXANS, A  PIPER, H. BEAM/MCGUIRE, JOHN J.  DD299
PLANET IN ARMS, A  BARR, DONALD  C/P:2,P
PLANET IN PERIL  CHRISTOPHER, JOHN  C/P:2,P
PLANET IN PERIL [@CYBORG COMMANDO #1]  O’NEILL, PAMELA/MOHAN, KIM  C/P:2,P
PLANET KILLERS, THE  SILVERBERG, ROBERT  DD407
PLANET MASTERS, THE  WOLD, ALLEN L.  H
PLANET MASTERS, THE (PROOF)  WOLD, ALLEN L.  SR-LP
PLANET MURDERER, THE [@SPACEWAYS #16]  CLEVE, JOHN  P
PLANET OF ADVENTURE  VANCE, JACK  H
PLANET OF DREAD [@CAP KENNEDY #10]  KERN, GREGORY  C/P,P
PLANET OF EXILE [HAINISH 2]  LE GUIN, URSULA K.  C/P:2,P,SFWA,DG597
PLANET OF FEAR  DETZER, DIANE  H
PLANET OF FIRE [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P
PLANET OF FLOWERS [@WINDHOVER TAPES #4]  NORWOOD, WARREN  C/P:2,P
PLANET OF JUDGMENT [STAR TREK]  HALDEMAN, JOE  C/P:2,P
PLANET OF LIGHT  JONES, RAYMOND F.  H:2
PLANET OF NIGHTMARES [@NEW TOM SWIFT #11]  APPLETON, VICTOR  LP
PLANET OF NO RETURN  ANDERSON, POUL  DD199
PLANET OF NO RETURN  HARRISON, HARRY  C/P,LP,H
PLANET OF PERIL  KLINE, OTIS ADELBERT  C/P,P,H
PLANET OF TEARS, THE  REINIUS, TRISH  C/P:2,P
PLANET OF THE APES  BOULLE, PIERRE  C/P,P,H
PLANET OF THE APES  QUICK, W. T.  H
PLANET OF THE BLIND, THE  COREY, PAUL  C/P,P
PLANET OF THE BLIND [@RACK #4]  JAMES, LAURENCE  P
PLANET OF THE DAMNED  HARRISON, HARRY  C/P:2,SFWA-P,P
PLANET OF THE DEAD  WISMER, DON  P,SFWA-P
PLANET OF THE DOUBLE SUN, THE [PROFESSOR JAMESON 1]  JONES, NEIL R.  C/P:2,P
PLANET OF THE DREAMERS  MACDONALD, JOHN D.  C/P:2,P
PLANET OF THE DYING SUN, THE [@PERRY RHODAN #11]  ACE  PA,C/PA
PLANET OF THE GAWFS  VANCE, STEVE  C/P,P
PLANET OF THE GODS [@PERRY RHODAN #27]  ACE  PA,C/PA
PLANET OF THE VOLES  PLATT, CHARLES  H,P
PLANET OF THE WARLORD [LAST LEGIONARY 4]  HILL, DOUGLAS  H
PLANET OF TREACHERY [FAMILY D’ALEMBERT #7]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
PLANET OF TWILIGHT [@STAR WARS]  HAMBLY, BARBARA  P
PLANET OF WHISPERS [MESSENGER 1]  KELLY, JAMES PATRICK  C/P,SFWA,SFWA-P,P,H
PLANET OF YOUR OWN, A  BRUNNER, JOHN  DG592
PLANET OF YOUTH, THE  COBLENTZ, STANTON A.  H
PLANET ON THE TABLE, THE  ROBINSON, KIM STANLEY  H,SFWA-P,P
PLANET PATROL  DORMAN, SONYA  H

PLANET PIRATES, THE [PLANET PIRATES 1,2,3] MCCAFFREY, ANNE/MOON, ELIZABETH/

NYE, JODY LYNN

LP

PLANET PLAGUE [@STAR WARS GALAXY OF FEAR #3]  WHITMAN, JOHN  P
PLANET PROBABILITY  BALL, BRIAN N.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 272
TITLE  AUTHOR  CODE
PLANET RUN  LAUMER, KEITH/DICKSON, GORDON R.  C/P:2,P
PLANET RUN AND TWO STORIES  LAUMER, KEITH/DICKSON, GORDON R.  C/P,P
PLANET SAVERS, THE [DARKOVER]  BRADLEY, MARION ZIMMER  DF153

PLANET SAVERS, THE [DARKOVER]/ SWORD OF ALDONES, THE [DARKOVER] BRADLEY, MARION ZIMMER

C/P,P

PLANET SAVERS, THE (INCLUDES THE WATERFALL) [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P
PLANET SEEKERS, THE  BARTON, ERLE  SR-P
PLANET STRAPPERS, THE  GALLUN, RAYMOND Z.  C/P,P
PLANET THA  CHARLES, NEIL  SR-P
PLANET THAT WASN’T, THE  ASIMOV, ISAAC  SFWA
PLANET TOPIDE, PLEASE REPLY! [@PERRY RHODAN #75]  ACE  C/PA,PA
PLANET WAR  FYSH  SR-P
PLANET WITHOUT A NAME  SHAPIRO, NEIL  P
PLANET WIZARD, THE  JAKES, JOHN  C/P:2,P
PLANET X [@STAR TREK TNG/X-MEN]  FRIEDMAN, MICHAEL JAN  C/P,P
PLANET X  HUNT, GILL  SR-P
PLANET X  KIEFER, MILDRED S.  H
PLANETA DE LAS BRUJAS, EL  NORTON, ANDRE  FP
PLANETA ROJO, EL  CAMPBELL, H. J.  FP
PLANETARY AGENT X  REYNOLDS, MACK  DM131
PLANETE EN PERIL  GUIEU, JIMMY  FP
PLANETEERS, THE  CAMPBELL, JOHN W. JR.  DG585
PLANETENKONIG, DER [@PERRY RHODAN #61]  MAHR, KURT  FPR
PLANETFALL [PLANETFALL 1]  COVER, ARTHUR BYRON  C/P:2,P
PLANETOID DISPOSALS LTD.  GRIDBAN, VOLSTED  SR-P
PLANETOID PERIL, THE  BROWN, GEORGE SHELDON  SR-P
PLANETOIDE 13  LIMAT, MAURICE  FP
PLANETS  BINDER, OTTO  [C/RAN-H]
PLANETS, THE  PREISS, BYRON  SR-VLHA
PLANETS AND DIMENSIONS  SMITH, CLARK ASHTON  R/REF-H
PLANETS FOR SALE  VAN VOGT, A. E./HULL, E. MAYNE  SR-P,C/P:2
PLANETS IN PERIL  HAMILTON, EDMOND  C/P,P
PLANETS OF ADVENTURE  LEINSTER, MURRAY  P
PLANETS OF ADVENTURE  WELLS, BASIL  SR-LP
PLANETS OF DEATH, THE  COLLINS, MICHAEL  C/P:2,P
PLANETS OF PERIL [@SPACE: 1999 YEAR 2 #1]  BUTTERWORTH, MICHAEL  C/P,P
PLANETS OF WONDER  CARR, TERRY  HA
PLANETS THREE  POHL, FREDERIK  C/P,P
PLANIVERSE, THE  DEWDNEY, A. K.  LP
PLANT PEOPLE, THE  CARLSON, DALE  P
PLANTS, THE  MCKENNEY, KENNETH  P
PLASM [WORLDS OF CHTHON]  PLATT, CHARLES  C/P,P
PLASMA MONSTER, THE [@PERRY RHODAN #95]  ACE  C/PA,PA
PLASTIC MAN, THE  BRENT, JEREMY  P
PLATFORMS  MAXIM, JOHN  P,SFWA-P
PLATINUM POHL  POHL, FREDERIK  H:2
PLATO PAPERS, THE  ACKROYD, PETER  LP
PLATYPUS OF DOOM AND OTHER NIHILISTS, THE  COVER, ARTHUR BYRON  P
PLAY TO LIVE [TWILIGHT 7]  VELEY, CHARLES  P
PLAYBOY BOOK OF HORROR AND THE SUPERNATURAL, THE  PLAYBOY BOOKS  C/PA,PA
PLAYBOY BOOK OF SCIENCE FICTION, THE  TURNER, ALICE K.  HA,LPA
PLAYBOY BOOK OF SCIENCE FICTION AND FANTASY, THE  PLAYBOY BOOKS  PA,C/PA
PLAYBOY’S GAHAN WILSON  WILSON, GAHAN  CX
PLAYBOY’S LITTLE ANNIE FANNY  KURTZMAN, HARVEY/ELDER, WILL  CX
PLAYBOY’S STORIES OF THE SINISTER AND STRANGE  PLAYBOY BOOKS  PA
PLAYER OF GAMES, THE [CULTURE]  BANKS, IAIN M.  C/P:3,P,H
PLAYER PIANO  VONNEGUT, KURT JR.  H:2,P
PLAYERS [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  C/P,P
PLAYERS AT THE GAME OF PEOPLE  BRUNNER, JOHN  H:2,P
PLAYERS OF GILEAN, THE [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
PLAYERS OF GOR [GOR 20]  NORMAN, JOHN  C/P:2,P
PLAYERS OF HELL, THE  VAN ARNAM, DAVE  SR-DB0077
PLAYERS OF NULL-A, THE  VAN VOGT, A. E.  C/P:2,P
PLAYGROUND, THE  WRIGHT, T. M.  SFWA-P
PLAYGROUNDS OF THE MIND  NIVEN, LARRY  H,P
PLAYING GOD  ZETTEL, SARAH  H,P
PLAYING IN WONDERLAND  BETANCOURT, JOHN GREGORY  LP
PLAYMASTERS, THE  DALMAS, JOHN/MARTIN, ROD  C/P,P
PLAYS OF GODS AND MEN  DUNSANY, LORD  H
PLEASANT DREAMS  BLOCH, ROBERT  P
PLEASURE TUBE, THE  ONOPA, ROBERT  C/P:2,P
PLEASURING OF RORY MALONE, THE  PANATI, CHARLES  P

PLIOCENE COMPANION, A MAY, JULIAN R/REF-H,C/REF-H:2,

C/REF-P

PLOT AGAINST EARTH, THE  KNOX, CALVIN M.  DD358

The MITSFS Titledex  September 24, 2007  Page 273
TITLE  AUTHOR  CODE
PLOT TO SAVE SOCRATES, THE  LEVINSON, PAUL  H
PLOTTERS, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
PLUCKER, THE  BROM  H
PLUNDER  GOULART, RON  C/P:2,P
PLURIBUS  KURLAND, MICHAEL  C/P,SFWA,P,H
PLUS  MCELROY, JOSEPH  LP
PLUTONIA: AN ADVENTURE THROUGH PREHISTORY  OBRUCHEV, V. A.  H
PLUTONIUM BLONDE, THE [ZACHARY NIXON JOHNSON 1]  ZAKOUR, JOHN/GANEM, LAWRENCE  C/P,P
PNUME, THE  VANCE, JACK  C/P:2,P
POCKET AND THE PENDANT, THE  JEFFREY, MARK  LP
POCKET BOOK OF SCIENCE FICTION, THE  WOLLHEIM, DONALD A.  SR-PA,C/PA
POCKETFUL OF STARS, A  KNIGHT, DAMON  HA,PA,SFWA
POCKETS OF RESISTANCE  SUNDOWN, WILL  P
PODKAYNE OF MARS  HEINLEIN, ROBERT A.  C/P,P,H
PODKAYNE OF MARS (WITH BOTH ENDINGS)  HEINLEIN, ROBERT A.  H,P
POE MUST DIE  OLDEN, MARC  P
POETIC FANTASTIC, THE  MURPHY, PATRICK D./HYLES, VERNON  R/REF-H
POHLSTARS  POHL, FREDERIK  C/P:2,P
POINT COUNTER POINT  HUXLEY, ALDOUS  P
POINT MAN, THE  ENGLEHART, STEPHEN  C/P,P
POINT OF DREAMS [NICOLAS RATHE 2]  SCOTT, MELISSA/BARNETT, LISA A.  H
POINT OF HONOR, A  HEYDT, DOROTHY J.  C/P:2,P
POINT OF HOPES [NICOLAS RATHE 1]  SCOTT, MELISSA/BARNETT, LISA A.  H
POINT ULTIMATE  SOHL, JERRY  H:2,P
POINTBLANK [@STARFIST FORCE RECON 2]  SHERMAN, DAVID/CRAGG, DAN  P
POINTS OF DEPARTURE  MURPHY, PAT  C/P,P
POISON AGENDAS [SHADOWRUN]  KENSON, STEPHEN  P
POISON BELT, THE  DOYLE, ARTHUR CONAN  C/P:2,P
POISON FRUIT [CHARLIE MOON 3]  YARBRO, CHELSEA QUINN  P
POISON ISLAND [DOC SAVAGE #57]  ROBESON, KENNETH  C/P,@P
POISON MASTER, THE  WILLIAMS, LIZ  C/P,P
POISON ORACLE, THE  DICKINSON, PETER  SFWA
POISON STUDY [YELINA 1]  SNYDER, MARIA V.  H
POISONED LANDS, THE [STORMLANDS 3]  ROBERTS, JOHN MADDOX  C/P,P
POISONED MEMORIES [@EARTHDAWN]  KUBASIK, CHRISTOPHER  P
POLAR CITY BLUES [POLAR CITY 1]  KERR, KATHARINE  C/P,P
POLAR CITY NIGHTMARE [POLAR CITY 2]  KERR, KATHARINE/DANIEL, KATE  LP
POLAR FLEET [DOUBLE SPIRAL WAR 2]  NORWOOD, WARREN  C/P,P
POLAR TREASURE, THE [DOC SAVAGE #4]  ROBESON, KENNETH  C/P,@P
POLARIS [ALEX BENEDICT 2]  MCDEVITT, JACK  H
POLARIS  PERKINS, SHELDON  P
POLARIS AND THE IMMORTALS  STILSON, CHARLES B.  H
POLARIS OF THE SNOWS  STILSON, CHARLES B.  H
POLGARA THE SORCERESS [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  H:3
POLICE PATROL: 2000 A.D.  REYNOLDS, MACK  C/P:2,P
POLICE YOUR PLANET  DEL REY, LESTER/VAN LHIN, ERIK  C/P:2,P

POLITICAL SCIENCE FICTION GREENBERG, MARTIN H./ WARRICK, PATRICIA S.

LPA

POLITICIAN [@BIO OF A SPACE TYRANT #3]  ANTHONY, PIERS  C/P:2,P
POLITY AGENT [POLITY GRIDLINKED 4]  ASHER, NEAL  H
POLLEN  NOON, JEFF  LP
POLLINATORS OF EDEN, THE  BOYD, JOHN  H:2,P
POLLY, THE WORRIED [@PHENWICK #25]  KIMBROUGH, KATHERYN  P
POLTERGEIST  KAHN, JAMES  H,P
POLTERGEIST II: THE OTHER SIDE  KAHN, JAMES  P
POLYMATH  BRUNNER, JOHN  C/P:2,P
POLYMORPH  WESTERFELD, SCOTT  P
POLYSTOM  ROBERTS, ADAM  H
POMEGRANATES FULL AND FINE [@WORLD OF DARKNESS]  BASSINGTHWAITE, DON  P
PONY SOLDIERS [@DEATHLANDS]  AXLER, JAMES  P
POOL OF FIRE, THE [TRIPODS 3]  CHRISTOPHER, JOHN  C/P:2,P
POOL OF RADIANCE [@FORGOTTEN REALMS RUINS OF MYTH DRANNOR]  BEBRIS, CARRIE A.  P
POOL OF RADIANCE [@FORGOTTEN REALMS]  WARD, JAMES M./HONG, JANE COOPER  C/P:2,P
POOL OF TWILIGHT [@FORGOTTEN REALMS]  WARD, JAMES M./BROWN, ANNE K.  C/P,P
POOL OF TWO MOONS, THE [WITCHES OF EILEANAN 2]  FORSYTH, KATE  C/P:2,P
POOLS OF DARKNESS [@FORGOTTEN REALMS]  WARD, JAMES M./BROWN, ANNE K.  C/P,P
POPES AND PHANTOMS  WHITBOURN, JOHN  H
POPEYE IN GHOST SHIP TO TREASURE ISLAND  NEWMAN, PAUL S.  BLB
PORNUCOPIA  ANTHONY, PIERS  SR-L
PORT ETERNITY  CHERRYH, C. J.  H:2,P
PORT OF PERIL, THE  KLINE, OTIS ADELBERT  C/P,P,H
PORTABLE DOOR, THE [PAUL CARPENTER 1]  HOLT, TOM  H
PORTABLE EDGAR ALLAN POE, THE  POE, EDGAR ALLAN  P
PORTABLE NOVELS OF SCIENCE, THE  WOLLHEIM, DONALD A.  HA

The MITSFS Titledex  September 24, 2007  Page 274
TITLE  AUTHOR  CODE
PORTABLE VOLTAIRE, THE  VOLTAIRE  P
PORTAL  SWIGART, ROB  H
PORTAL THROUGH TIME [@BUFFY]  HENDERSON, ALICE  P
PORTAL TO E’EWERE  WALKER, IRMA  H
PORTALS IN A NORTHERN SKY  HAYES, CHARLES DOUGLAS  H
PORTALS OF TOMORROW  DERLETH, AUGUST  HA
PORTENT  HERBERT, JAMES  H,LP
PORTERO METHOD, THE [@SIMS #2]  WILSON, F. PAUL  SR-H
PORTRAIT OF JENNIE  NATHAN, ROBERT  P
PORTRAITS IN MOONLIGHT  JACOBI, CARL  SR-H
PORTRAITS OF HIS CHILDREN  MARTIN, GEORGE R. R.  H,P
PORTRAITS OF THE PAST [HOLDERLY HALL 6]  CAMERON, KATE  P
PORTS OF CALL [MYRON TANY 1]  VANCE, JACK  H
PORTS OF CALL & LURULU [MYRON TANY 1,2]  VANCE, JACK  H
PORTS OF CALL (VIE) [@VANCE INTEGRAL EDITION #43/MYRON TANY 1]  VANCE, JACK  SR-H
POSEIDONIS  SMITH, CLARK ASHTON  SR-P,C/P
POSITIVE CHARGE  RICHMOND, WALT/RICHMOND, LEIGH  D27235
POSITRONIC MAN, THE  ASIMOV, ISAAC/SILVERBERG, ROBERT  C/P,P,H:2
POSLEEN DISK, THE (CD) [POSLEEN]  BAEN PUBLISHING ENTERPRISES  SR-MM
POSSESS AND CONQUER  BROWN, WENZELL  P
POSSESSION [STAR TREK TNG]  DILLARD, J. M./O’MALLEY, KATHLEEN  C/P:2,P
POSSESSION OF IMMANUEL WOLF, THE  KAYE, MARVIN  H,SFWA
POSSESSION OF TAMARA, THE  LAWSON, SYBIL  SR-L
POSSESSORS, THE  CHRISTOPHER, JOHN  H:2,P
POSSIBLE WORLDS OF SCIENCE FICTION  CONKLIN, GROFF  PA,C/PA
POST READER OF FANTASY AND SCIENCE FICTION, THE  SATURDAY EVENING POST  C/PA,PA
POSTMAN, THE  BRIN, DAVID  C/P:2,SFWA,P
POSTMARKED THE STARS [SOLAR QUEEN 3]  NORTON, ANDRE  C/P,P,H
POWER, THE [TWILIGHT 6]  HAYNES, BETSY  P
POWER  JANIFER, LAURENCE M.  C/P,P
POWER, THE  ROBINSON, FRANK M.  H:2,P
POWER  STIRLING, S. M.  C/PA,PA,SFWA-P
POWER AND THE PROPHET, THE [PELMEN THE POWERSHAPER 3]  HUGHES, ROBERT DON  C/P:2,P
POWER CUBE AFFAIR, THE [@MAN FROM UNCLE #19]  PHILLIFENT, JOHN T.  P
POWER HUNGRY [STAR TREK TNG]  WEINSTEIN, HOWARD  C/P:3,P
POWER KEY [@PERRY RHODAN #78]  ACE  PA
POWER KLINGON (CASSETTE) [STAR TREK]  LEVINE, BARRY/OKRAND, MARC  SR-MM

POWER LINES [PETAYBEE 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

POWER OF BLACKNESS, THE  WILLIAMSON, JACK  H,P
POWER OF PERSUASION [BUFFY]  MASSIE, ELIZABETH  P
POWER OF STARS, THE  LAWRENCE, LOUISE  H
POWER OF THE DALEKS, THE [DOCTOR WHO]  PEEL, JOHN  C/P,P
POWER OF THE GODS, THE  VERNAM, GLENN R.  P
POWER OF THE SERPENT, THE  TIMLETT, PETER VALENTINE  C/P,P
POWER OF THREE  JONES, DIANA WYNNE  C/P,P
POWER OF THREE, THE [@CHARMED]  WILLARD, ELIZA  P
POWER OF TIME, THE  SAXTON, JOSEPHINE  LP
POWER OF X, THE  SELLINGS, ARTHUR  C/P:3,P

POWER PLAY [PETAYBEE 3] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

POWER SPHERE  BRETT, LEO  SR-P
POWER THAT PRESERVES, THE [THOMAS COVENANT 3]  DONALDSON, STEPHEN R.  C/P,LP,H:2
POWERS OF DETECTION  STABENOW, DANA  LPA
POWERS OF TWO  POWERS, TIM  H:2
POWER’S PRICE [@PERRY RHODAN #89]  ACE  C/PA,PA

POWERS THAT BE [PETAYBEE 1] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H:2,P

POWERS THAT BE  NICHOLS, BEVERLY  [C/RAN-P:2]
POWERSAT  BOVA, BEN  H
PRACTICAL DEMONKEEPING  MOORE, CHRISTOPHER  H,P
PRACTICAL MAGIC  HOFFMAN, ALICE  P
PRACTICE EFFECT, THE  BRIN, DAVID  C/P,SR-P,H
PRADOR MOON [POLITY]  ASHER, NEAL  LP
PRAESIDIUM OF ARCHIVE, THE  SWYCAFFER, JEFFERSON P.  P
PRAISE ALL THE MOONS OF MORNING  STONE, JOSEPHINE RECTOR  H
PRAXIS, THE [@DREAD EMPIRE’S FALL 1]  WILLIAMS, WALTER JON  P
PRAXIS, THE (PROOF) [@DREAD EMPIRE’S FALL 1]  WILLIAMS, WALTER JON  SR-P
PRAYER MACHINE, THE  HODDER-WILLIAMS, CHRISTOPHER  H
PRAYERS FOR THE ASSASSIN  FERRIGNO, ROBERT  H
PRAYERS TO BROKEN STONES  SIMMONS, DAN  C/P:2,P
PRE-EMPT  VORHIES, JOHN  P
PRE-GARGANTUA  CHARLES, NEIL  SR-P
PRECIOUS CARGO [ANGEL’S LUCK 2]  FAUST, JOE CLIFFORD  H,P

The MITSFS Titledex  September 24, 2007  Page 275
TITLE  AUTHOR  CODE
PRECIOUS DRAGON [DETECTIVE INSPECTOR CHEN]  WILLIAMS, LIZ  H
PRECIPICE, THE [ASTEROID WARS 1]  BOVA, BEN  H,P
PRECOG  DUIGON, LEE  P
PRECURSOR [ATEVI 4]  CHERRYH, C. J.  H:2,P
PREDATOR [PREDATOR]  MONETTE, PAUL  C/P,P
PREDATOR [@ROBOTS IN TIME 1]  WU, WILLIAM F.  P,SFWA-P
PREDATOR 2 [PREDATOR]  HAWKE, SIMON  P
PREDATOR AND PREY [@PROWLERS 3]  GOLDEN, CHRISTOPHER  P
PREDATORS  GORMAN, ED/GREENBERG, MARTIN H.  SFWA
PREDATORS  GRAZIUNAS, DAINA/STARLIN, JIM  P
PREDATOR’S GOLD [HUNGRY CITY 2]  REEVE, PHILIP  H
PREDATOR’S SMILE [@DAREDEVIL]  GOLDEN, CHRISTOPHER  P
PREFERRED RISK  MCCANN, EDSON  H:2,P
PREFERRED RISK  POHL, FREDERIK/DEL REY, LESTER  C/P,P
PRELUDE [QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,P
PRELUDE TO CHAOS  LLEWELLYN, EDWARD  C/P:2,P
PRELUDE TO FOUNDATION [FOUNDATION 0]  ASIMOV, ISAAC  C/P,P,H
PRELUDE TO FOUNDATION PAPER BLOCK  DOUBLEDAY BOOKS  SFWA-VLP
PRELUDE TO MARS  CLARKE, ARTHUR C.  H:2,SR-H
PRELUDE TO SPACE  CLARKE, ARTHUR C.  C/P:2,GN1.03,SR-P,P
PRELUDES & NOCTURNES [@SANDMAN #1]  GAIMAN, NEIL  H,CX
PREMONITION  WILLIAMSON, J. N.  P
PRENTICE ALVIN [ALVIN MAKER 3]  CARD, ORSON SCOTT  C/P,SFWA-P,P,H
PRESENCE, THE  CLEMENS, RODGERS  P
PRESENT TENSE [@GREAT GAME #2]  DUNCAN, DAVE  H
PRESENT TENSE [@STAR TREK JANUS GATE #1]  GRAF, L. A.  P
PRESERVER [MORPHODITE 3]  FOSTER, M. A.  C/P:2,P

PRESERVER [STAR TREK MIRROR 3] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

PRESERVING MACHINE, THE  DICK, PHILIP K.  C/P,P,H
PRESIDENT FU MANCHU  ROHMER, SAX  C/P,P
PRESIDENT IS BORN, A  HURST, FANNIE  H
PRESIDENT KISSINGER  ROSENTHAL, MONROE/MUNSON, DONALD  P
PRESIDENT MCGOVERN’S FIRST TERM  MAX, NICHOLAS  H
PRESIDENTIAL PLOT, THE  JOHNSON, STANLEY  H
PRESIDENTIAL YEAR  POHL, FREDERIK/KORNBLUTH, C. M.  P
PRESIDENT’S DAUGHTER, THE  WHITE, ELLEN EMERSON  P
PRESIDENT’S PLANE IS MISSING, THE  SERLING, ROBERT  P
PRESS ENTER  VARLEY, JOHN  H,SFWA-TD26
PRESSURE MAN  HUGHES, ZACH  P
PRESTIGE, THE  PRIEST, CHRISTOPHER  H,LP
PRETENDER  ANTHONY, PIERS/HALL, FRANCES  C/P:2,SFWA-P,LP
PRETENDER [ATEVI 8]  CHERRYH, C. J.  H
PRETENDER, THE [CHAOS GATE 2]  COOPER, LOUISE  P
PRETERNATURAL [PRETERNATURAL 1]  BONANNO, MARGARET WANDER  C/P,P
PRETERNATURAL 3 [PRETERNATURAL 3]  BONANNO, MARGARET WANDER  H:2
PRETERNATURAL TOO: GYRE [PRETERNATURAL 2]  BONANNO, MARGARET WANDER  H,LP
PRETTY MAIDS ALL IN A ROW [BUFFY SPIKE & DRU]  GOLDEN, CHRISTOPHER  H,P
PREY  CRICHTON, MICHAEL  H:2
PREY, THE [@TRAVELER #12]  DRUMM, D. B.  P
PREY [@ALIENS VS. PREDATOR 1]  PERRY, STEVE/PERRY, STEPHANI  P
PREY, THE  SMITH, ROBERT ARTHUR  P
PREYING FOR KEEPS [@SHADOWRUN 21]  ODOM, MEL  P
PRIAM’S LENS  CHALKER, JACK L.  P
PRICE OF BLOOD AND HONOR, THE [LANDUC 3]  WILLEY, ELIZABETH  C/P:2,P
PRICE OF COMMAND, THE [JEFFERSON’S WAR 2]  RANDLE, KEVIN  C/P,P
PRICE OF FEAR, THE [@DARKMAN #2]  BOYLL, RANDALL  P
PRICE OF FREEDOM, THE [@WING COMMANDER 5]  FORSTCHEN, WILLIAM R./OHLANDER, BEN  P
PRICE OF GLORY, THE [@BATTLETECH 8]  KEITH, WILLIAM H. JR.  C/P:2,P
PRICE OF ORANGES, THE  KRESS, NANCY  B
PRICE OF PARADISE, THE [DOCTOR WHO TENTH DOCTOR]  BRAKE, COLIN  H
PRICE OF PEACE, THE [SOCIETY OF HUMANITY/UNITY PARTY 2]  MOSCOE, MIKE  C/P,P
PRICE OF POWER [@GREYHAWK #4]  ESTES, ROSE  C/P:2,P
PRICE OF RANSOM, THE [@HIGHROAD #3]  RASMUSSEN, ALIS A.  C/P:2,P
PRICE OF THE PHOENIX, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
PRICE OF THE STARS, THE [MAGEWORLDS 1]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P:2,P
PRICE OF VICTORY, THE [THIEVES’ WORLD 10,11,12]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  H:2
PRICE YOU PAY, THE [@STARGATE SG-1 (MCCONNELL) 2]  MCCONNELL, ASHLEY  P
PRIDE OF CHANUR, THE [CHANUR 1]  CHERRYH, C. J.  C/P,P,H
PRIDE OF KINGS  TARR, JUDITH  LP
PRIDE OF MONSTERS, A  SCHMITZ, JAMES H.  H:2,P
PRIDE OF PRINCES, A [CHEYSULI 5]  ROBERSON, JENNIFER  C/P:2,P
PRIEST-KINGS OF GOR [GOR 3]  NORMAN, JOHN  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 276
TITLE  AUTHOR  CODE
PRIESTESS, THE  LAURIA, FRANK  P

PRIESTESS OF AVALON [MISTS OF AVALON 3] BRADLEY, MARION ZIMMER/

PAXSON, DIANA L.

H

PRIESTESS OF THE DAMNED [LUCIFER COVE 2]  COFFMAN, VIRGINIA  P
PRIESTESS OF THE WHITE [AGE OF THE FIVE 1]  CANAVAN, TRUDI  P
PRIESTS OF FERRIS, THE  GEE, MAURICE  SR-P
PRIMAL URGE, THE  ALDISS, BRIAN W.  C/P:2,P
PRIMARY INVERSION [SKOLIAN EMPIRE 1]  ASARO, CATHERINE  C/P,P
PRIMAVERA  BLOCK, FRANCESCA LIA  P

PRIME DIRECTIVE [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

H:2,P

PRIME EVIL [@BUFFY]  GALLAGHER, DIANA G.  C/P,P
PRIME EVIL  WINTER, DOUGLAS E.  HA,PA
PRIME NUMBER  HARRISON, HARRY  C/P:2,P
PRIME SQUARED [BUCK ROGERS INNER PLANETS 2]  MURDOCK, M. S.  C/P:2,P
PRIME TIME [DOCTOR WHO PAST DOCTORS]  TUCKER, MIKE  P
PRIMEROS HOMBRES EN LA LUNA, LOS  WELLS, H. G.  FP
PRIMITIVE, THE  TUBB, E. C.  SR-P,C/P
PRINCE AMONG MEN, A  CHARRETTE, ROBERT N.  P
PRINCE CASPIAN [NARNIA 2]  LEWIS, C. S.  C/P:2,P
PRINCE COMMANDS, THE  NORTON, ANDRE  P
PRINCE IN CAMELOT, A [DRAGON’S HEIRS 3]  JONES, COURTWAY  P
PRINCE IN WAITING, THE [LUKE 1]  CHRISTOPHER, JOHN  H:2,P
PRINCE OF ANNWN [MABINOGI 1]  WALTON, EVANGELINE  C/P:2,P
PRINCE OF AYODHYA [@RAMAYANA #1]  BANKER, ASHOK K.  P
PRINCE OF AYODHYA (PROOF) [RAMAYANA 1]  BANKER, ASHOK K.  SR-LP
PRINCE OF CHAOS [AMBER 10]  ZELAZNY, ROGER  H:2,@P
PRINCE OF DARKNESS  RUSSELL, RAY  SR-P
PRINCE OF DEMONS [RENSHAI CHRONICLES 2]  REICHERT, MICKEY ZUCKER  P
PRINCE OF DOGS [CROWN OF STARS 2]  ELLIOTT, KATE  C/P,P,H
PRINCE OF DREAMS, THE [SEVEN BROTHERS 2]  BENJAMIN, CURT  P
PRINCE OF FIRE AND ASHES [TIELMARAN 3]  REIMANN, KATYA  P
PRINCE OF GRAUSTARK, THE  MCCUTCHEON, GEORGE B.  H
PRINCE OF HAVOC [@BATTLETECH #39/TWILIGHT OF THE CLANS 7]  STACKPOLE, MICHAEL A.  P
PRINCE OF ILL LUCK, THE [WARHORSE OF ESDRAGON 1]  DEXTER, SUSAN  P
PRINCE OF LIES [FORGOTTEN REALMS AVATAR 4]  LOWDER, JAMES  P
PRINCE OF MERCENARIES [FALKENBERG’S LEGION]  POURNELLE, JERRY  C/P:2,SFWA-P,P
PRINCE OF MORNING BELLS, THE  KRESS, NANCY  C/P,P
PRINCE OF PERIL  KLINE, OTIS ADELBERT  C/P,P,H
PRINCE OF SCORPIO [DRAY PRESCOT 5]  AKERS, ALAN BURT  C/P,P
PRINCE OF SHADOW, THE [SEVEN BROTHERS 1]  BENJAMIN, CURT  C/P,P
PRINCE OF SPACE, THE  WILLIAMSON, JACK  GRD17
PRINCE OF SPARTA [FALKENBERG’S LEGION]  POURNELLE, JERRY/STIRLING, S. M.  C/P:2,SFWA-P,P
PRINCE OF SUNSET [BASIL CASTELLAN 1]  WHITE, STEVE  C/P,P
PRINCE OF THE BLOOD [RIFTWAR]  FEIST, RAYMOND E.  C/P,P,H,SFWA
PRINCE OF THE CITY [WORLD OF DARKNESS VAMPIRE]  HERBER, KEITH  P
PRINCE OF THE GODBORN [SEVEN CITADELS 1]  HARRIS, GERALDINE  P
PRINCE OF THE NORTH [GERIN THE FOX 2]  TURTLEDOVE, HARRY  C/P,P
PRINCE OF WHALES, THE  FISHER, R. L.  P,SFWA-P
PRINCE OMBRA  MACLEISH, RODERICK  C/P,P
PRINCE ZALESKI/CUMMINGS KING MONK  SHIEL, M. P.  SR-H
PRINCES OF EARTH, THE  KURLAND, MICHAEL  H
PRINCES OF THE AIR, THE  FORD, JOHN M.  C/P:2,SFWA-P,P
PRINCESS AND CURDIE, THE  MACDONALD, GEORGE  SR-P,C/P
PRINCESS AND THE GOBLIN, THE  MACDONALD, GEORGE  SR-P,C/P:2
PRINCESS AND THE GOBLIN, THE/PRINCESS AND CURDIE, THE  MACDONALD, GEORGE  P
PRINCESS AND THE HOUND, THE  HARRISON, METTE IVIE  H
PRINCESS AND THE THORN, THE [ASH STAFF 3]  FISHER, PAUL R.  P
PRINCESS AT SEA [DECOY PRINCESS 2]  COOK, DAWN  C/P,P
PRINCESS BRIDE, THE  GOLDMAN, WILLIAM  SR-P,C/P:2
PRINCESS HYNCHATTI AND SOME OTHER SURPRISES  LEE, TANITH  H
PRINCESS NEVERMORE  REGAN, DIAN CURTIS  P
PRINCESS OF ALL LANDS, THE  KIRK, RUSSELL  H
PRINCESS OF DHAGABAD, THE [SPIRITS OF ANCIENT SANDS 1]  KASHINA, ANNA  H
PRINCESS OF FLAMES, THE  EMERSON, RU  P
PRINCESS OF MARS, A [BARSOOM #1]  BURROUGHS, EDGAR RICE  H:3,@P,SR-H

PRINCESS OF MARS, A [BARSOOM 1]/ FIGHTING MAN OF MARS, A [BARSOOM 7] BURROUGHS, EDGAR RICE

LP

PRINCESS OF ROUMANIA, A [ROUMANIA 1]  PARK, PAUL  P
PRINCESS OF THE AERIE, A [JAK JINNAKA 2]  BARNES, JOHN  P
PRINCESS OF THE ATOM, THE  CUMMINGS, RAY  C/P,P
PRINCESS OF THE CHAMELN, A [RULERS OF HYLOR 1]  WILDER, CHERRY  H,P
PRINCESS OF THE STARS [SEIKAI CREST OF THE STARS #1]  MORIOKA, HIROYUKI  LP
PRINCESS OF WANDS  RINGO, JOHN  H

The MITSFS Titledex  September 24, 2007  Page 277
TITLE  AUTHOR  CODE
PRINCIPIA DISCORDIA  MALACLYPSE THE YOUNGER  [C/RAN-H,C/RAN-P]
PRINCIPLES OF DESOLATION [@MECHWARRIOR DARK AGE/BATTLETECH]  HARDY, JASON M./BILLS, RANDALL N.  P
PRINTER’S DEVIL, THE  KIDD, CHICO  P
PRISM  BAY, AUSTIN  P
PRISM, THE  PETAJA, EMIL  DH051
PRISON OF ICE  AXTON, DAVID  P
PRISON OF NIGHT [DUMAREST 17]  TUBB, E. C.  C/P:2,P
PRISON OF SOULS [BARD’S TALE 3]  LACKEY, MERCEDES/SHEPHERD, MARK  C/P,SFWA-P,P
PRISON PLANET  DIETZ, WILLIAM C.  C/P,P
PRISON PLANET  HEYWOOD, VICTOR D.  SR-P
PRISON SATELLITE  KELLEY, LEO P.  SR-P
PRISON SHIP  CAIDIN, MARTIN  C/P:2,SFWA-P,P
PRISON UNSOUGHT, A [EXORDIUM 3]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
PRISONER, THE [PRISONER]  DISCH, THOMAS M.  C/P,P
PRISONER #2, THE [PRISONER]  MCDANIEL, DAVID  C/P,P
PRISONER #3, THE [PRISONER]  STINE, HANK  C/P:2,P
PRISONER IN FAIRYLAND, A  BLACKWOOD, ALGERNON  SR-H
PRISONER IN THE SKULL  DYE, CHARLES  H,P
PRISONER OF BLACKWOOD CASTLE, THE  GOULART, RON  C/P:2,P
PRISONER OF CABIN 13 [@SABRINA THE TEENAGE WITCH #11]  VORNHOLT, JOHN  P
PRISONER OF CONSCIENCE [JUDICIARY 2]  MATTHEWS, SUSAN R.  C/P,P
PRISONER OF DREAMS  RIPLEY, KAREN  C/P,P
PRISONER OF FIRE  COOPER, EDMUND  SR-P
PRISONER OF HAVEN [@DRAGONLANCE AGE OF MORTALS]  BERBERICK, NANCY VARIAN  P
PRISONER OF PSI  JOHNSON, ANNABEL/JOHNSON, EDGAR  H
PRISONER OF REGLATHIUM, THE [@DANNUS #1]  SIROTA, MIKE  P
PRISONER OF THE HORNED HELMET [DEATH DEALER 1]  SILKE, JAMES  C/P:2,SFWA-P,P
PRISONER OF THE IRON TOWER [TEARS OF ARTAMON 2]  ASH, SARAH  H
PRISONER OF TIME [@PERRY RHODAN #56]  ACE  C/PA:2,PA
PRISONER OF ZENDA, THE [ZENDA 1]  HOPE, ANTHONY  SR-H
PRISONER OF ZHAMANAK, THE [KRISHNA]  DE CAMP, L. SPRAGUE  H,P
PRISONER WITHIN, THE [LANNAT 2]  MCQUINN, DONALD E.  P
PRISONERS AND PAWNS [V]  WEINSTEIN, HOWARD  C/P,SFWA-P,P
PRISONER’S DILEMMA, THE [PRISONER]  BLUM, JONATHAN/BOOTH, RUPERT  LP
PRISONER’S HOPE [NICHOLAS SEAFORT 3]  FEINTUCH, DAVID  C/P:2,P
PRISONERS OF ARIONN  HERBERT, BRIAN  C/P,SFWA,P,H
PRISONERS OF PARADISE  CROSS, RONALD ANTHONY  H
PRISONERS OF PEACE [@STAR TREK DS9]  PEEL, JOHN  H,LP
PRISONERS OF POWER  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P,P,H
PRISONERS OF SPACE  DEL REY, LESTER  H
PRISONERS OF THE SKY  MACAPP, C. C.  C/P,P
PRISONERS OF TIME, THE [LONE WOLF 11]  DEVER, JOE  P
PRISONLAND [@WARLORD #4]  FROST, JASON  P
PRITCHER MASS, THE  DICKSON, GORDON R.  C/P,SFWA-P,P,H,SFWA
PRIVATE COSMOS, A [WORLD OF TIERS 3]  FARMER, PHILIP JOSE  C/P:2,P
PRIVATE LIFE OF HENRY PERKINS, THE  REYNOLDS, JOHN MURRAY  H
PRIVATE LIFE OF SHERLOCK HOLMES, THE  HARDWICK, MICHAEL/HARDWICK, MOLLIE P
PRIVATEER, THE [FLIGHT ENGINEER 2]  DOOHAN, JAMES/STIRLING, S. M.  C/P,P
PRIVATEERS [PRIVATEERS 1]  BOVA, BEN  C/P:2,SFWA-P,P,H
PRIVILEGE OF THE SWORD, THE [SWORDSPOINT]  KUSHNER, ELLEN  H,LP
PRIZE IN THE GAME, THE [TIR TANAGIRI 3]  WALTON, JO  H,P
PRIZE IN THE GAME, THE (PROOF) [TIR TANAGIRI 3]  WALTON, JO  SR-LP
PRO  DICKSON, GORDON R.  C/P:2,P
PROBABILITY BROACH, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P

PROBABILITY BROACH, THE (ORIGINAL)

[NORTH AMERICAN CONFEDERACY] SMITH, L. NEIL

H,LP

PROBABILITY CORNER, THE  RICHMOND, WALT/RICHMOND, LEIGH  C/P:2,P
PROBABILITY MAN, THE  BALL, BRIAN N.  C/P:2,P
PROBABILITY MOON [FALLERS 1]  KRESS, NANCY  H:2,P
PROBABILITY PAD, THE  WATERS, T. A.  P
PROBABILITY SPACE [FALLERS 3]  KRESS, NANCY  H
PROBABILITY SUN [FALLERS 2]  KRESS, NANCY  H
PROBE [STAR TREK]  BONANNO, MARGARET WANDER  C/P,P,H

PROBE [PROBE 1] DOUGLAS, CAROLE NELSON C/P,SFWA-P:2,H,LP,

SFWA

PROBE  LERNER, EDWARD M.  C/P:2,P
PROBLEM OF PAIN, THE  LEWIS, C. S.  [C/RAN-P]
PROBLEMS AND PROGRESS IN BANANA DISEASE RESEARCH  UNITED FRUIT COMPANY  [C/RAN-H]
PROCEEDINGS; CHICON III, THE  KEMP, EARL  R/REF-P
PROCEEDINGS; DISCON, THE  ENEY, DICK  R/REF-P
PROCEEDINGS OF THE INSTITUTE FOR TWENTY-FIRST CENTURY STUDIES  COGSWELL, THEODORE R.  R/REF-H
PROCESS OF AGEING, THE  COMFORT, ALEX  [C/RAN-P]
PROCESSION OF THE DAMNED  TUCKER, WILSON  P
PROCURATOR [PROCURATOR 1]  MITCHELL, KIRK  P

TITLE  AUTHOR  CODE
PROCYON’S PROMISE [LIFE PROBE 2]  MCCOLLUM, MICHAEL  C/P:2,P
PRODIGAL  GILLER, MARK D.  LP
PRODIGAL SORCERER, THE [MAGIC: THE GATHERING]  SUMNER, MARK  P
PRODIGAL SUN, THE  HIGH, PHILIP E.  C/P:2,P
PRODIGAL TROLL, THE  FINLAY, CHARLES COLEMAN  LP
PRODIGY  CLARK, IAN  P
PRODIGY [ROBOT CITY 4]  COVER, ARTHUR BYRON  C/P:2,SFWA-P,P
PRODUCTIONS OF TIME, THE  BRUNNER, JOHN  C/P:2,P

PROFESOR BERNICE SUMMERFIELD AND THE DEAD MEN DIARIES [BERNICE SUMMERFIELD] CORNELL, PAUL

PA

PROFESOR BERNICE SUMMERFIELD AND THE GODS OF THE UNDERWORLD [BERNICE SUMMERFIELD 2] COLE, STEPHEN

C/P,P

PROFESSOR BERNICE SUMMERFIELD AND THE DOOMSDAY MANUSCRIPT [BERNICE SUMMERFIELD 1] RICHARDS, JUSTIN

P

PROFESSOR BERNICE SUMMERFIELD AND THE GLASS PRISON [BERNICE SUMMERFIELD 5] RAYNER, JACQUELINE

P

PROFESSOR BERNICE SUMMERFIELD AND THE INFERNAL NEXUS [BERNICE SUMMERFIELD 4] STONE, DAVE

P

PROFESSOR BERNICE SUMMERFIELD AND THE SQUIRE’S CRYSTAL [BERNICE SUMMERFIELD 3] RAYNER, JACQUELINE

P

PROFESSOR DOWELL’S HEAD  BELIAEV, ALEXANDER  H:2,LP
PROFILES OF THE FUTURE  CLARKE, ARTHUR C.  [C/RAN-P:2]
PROFITEER [@HOSTILE TAKEOVER #1]  SWANN, S. ANDREW  P
PROFUNDIS  COWPER, RICHARD  P
PROGENITOR [@STAR TREK TNG STARGAZER #2]  FRIEDMAN, MICHAEL JAN  P
PROJECT AURA [PSYCHIC WARRIOR]  DOHERTY, ROBERT  P
PROJECT AVALON [BLAKE’S 7]  HOYLE, TREVOR  C/P,P
PROJECT: EARTHSAVE [@PERRY RHODAN #38]  ACE  PA,C/PA
PROJECT FARCRY  ASHWELL, PAULINE  LP
PROJECT FARCRY (PROOF)  ASHWELL, PAULINE  SR-LP
PROJECT: GENIUS  HAYES, WILLIAM D.  LP
PROJECT GOD  THOMPSON, V. M.  P
PROJECT ICARUS  KLEIMAN, LOUIS A.  [C/RAN-H]
PROJECT JOVE  GLASBY, JOHN  D68310
PROJECT MALDON  ATACK, CHRIS  P
PROJECT MEGALON  HODES, RICHARD  LP
PROJECT: MILLENNIUM  HOFFMANN, CURTIS H.  C/P,P
PROJECT PENDULUM  SILVERBERG, ROBERT  C/P,SFWA-P,P,H
PROJECT POPE  SIMAK, CLIFFORD D.  H:2,P
PROJECT SATELLITE  GATLAND, KENNETH W.  [C/RAN-H]
PROJECT SOLAR SAIL  CLARKE, ARTHUR C.  C/PA,PA
PROJECT WEB  ROGERS, BARBARA  H
PROJECTILE WAR  LANG, KING  SR-P
PROJECTION INFINITY  ZEIGFREID, KARL  SR-P,C/P
PROJECTIONS  ROBINETT, STEPHEN  C/P,P
PROJECTS: SPACE  VIORST, JUDITH  [C/RAN-P]
PROJET "KOZNA"  RAYJEAN, MAX-ANDRE  FP
PROLOGUE TO ANALOG  CAMPBELL, JOHN W. JR.  HA
PROM NIGHT  SPRINGER, NANCY  PA,C/PA
PROM TIME [@SABRINA THE TEENAGE WITCH #21]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
PROMETHEAN BALANCE, THE [@APPLESEED #4]  SHIROW, MASAMUNE  CX
PROMETHEAN CHALLENGE, THE [@APPLESEED #1]  SHIROW, MASAMUNE  CX
PROMETHEANS  BOVA, BEN  C/P:2,P
PROMETHEUS [JUSTIN BELL 3]  FORSTCHEN, WILLIAM R.  C/P,P
PROMETHEUS CRISIS, THE  SCORTIA, THOMAS N./ROBINSON, FRANK M. H:2,P
PROMETHEUS DESIGN, THE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
PROMETHEUS MAN, THE  NELSON, RAY FARADAY  LP
PROMETHEUS PROJECT, THE  WHITE, STEVE  H
PROMETHEUS PROJECT, THE (PROOF)  WHITE, STEVE  SR-LP
PROMETHEUS ROAD  BALFOUR, BRUCE  P
PROMETHEUS UNBOUND [@APPLESEED #2]  SHIROW, MASAMUNE  CX
PROMISE, THE [SANDWRITER 2]  HUGHES, MONICA  H
PROMISE, THE  LINDSEY, HAL  [C/RAN-P]
PROMISE OF SPACE, THE  CLARKE, ARTHUR C.  [C/RAN-P]
PROMISE OF THE WITCHKING [@FORGOTTEN REALMS SELLSWORDS #2]  SALVATORE, R. A.  P
PROMISED LAND [LOST IN SPACE]  CADIGAN, PAT  P
PROMISED LAND [GRAINGER 3]  STABLEFORD, BRIAN M.  H:2,P
PROMISED LAND  WILLIS, CONNIE/FELICE, CYNTHIA  C/P,P,H
PROMISING PLANET, A  STRIKE, JEREMY  D24100
PROOF HOUSE, THE [FENCER 3]  PARKER, K. J.  LP
PROPELLER ISLAND  VERNE, JULES  H
PROPHECIES [@BUFFY LOST SLAYER #1]  GOLDEN, CHRISTOPHER  C/P
PROPHECIES OF NOSTRADAMUS, THE  CHEETHAM, ERIKA  [C/RAN-P]
PROPHECY [@EARTHDAWN]  GORDEN, GREG  P

TITLE  AUTHOR  CODE
PROPHECY [@BLENDING #5]  GREEN, SHARON  C/P,P
PROPHECY [SYMPHONY OF AGES 2]  HAYDON, ELIZABETH  H,P
PROPHECY [@MAGIC: THE GATHERING MASQUERADE 3]  MOORE, VANCE  P
PROPHECY  SELTZER, DAVID  P
PROPHECY AND CHANGE [STAR TREK DS9]  PALMIERI, MARCO  LPA
PROPHECY MACHINE, THE  BARRETT, NEAL JR.  C/P,P
PROPHECY OF DARKNESS [@XENA]  HOWARD, STELLA  P

PROPHECY OF THE DRAGONS [KNIGHTS OF THE DRAGONS REVELATIONS 1] FORBECK, MATT

LP

PROPHET, THE [@SURVIVALIST #7]  AHERN, JERRY  P
PROPHET  RESNICK, MIKE  C/P,SFWA-P:2,P
PROPHET OF AKHRAN, THE [@ROSE OF THE PROPHET #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
PROPHET OF FIRE, THE  CREASEY, JOHN  H
PROPHET OF LAMATH, THE [PELMEN THE POWERSHAPER 1]  HUGHES, ROBERT DON  C/P:2,P
PROPHET OF MOONSHAE [@FORGOTTEN REALMS DRUIDHOME #1]  NILES, DOUGLAS  C/P,P
PROPHETS FOR THE END OF TIME  DONNELLY, MARCOS  C/P,P
PROPHETS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA,SFWA-P
PROPHET’S POWER [UNREAL 2]  SMITH, DEAN WESLEY  P
PROSE BOWL  PRONZINI, BILL/MALZBERG, BARRY N.  H
PROSPERO’S CHILDREN [FERN CHAPEL 1]  SIEGEL, JAN  P
PROSTHO PLUS  ANTHONY, PIERS  C/P:2,SFWA-P,P
PROTECTOR [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
PROTECTOR OF THE FLIGHT [LLADRANA 3]  OWENS, ROBIN D.  LP
PROTECTOR OF THE REALM [@SUPREME CONSTELLATIONS #1]  BROOKE, GUN  LP
PROTECTORATE  FARREN, MICK  P
PROTECTOR’S WAR, THE [EMBERVERSE 2]  STIRLING, S. M.  H
PROTEKTOR  PLATT, CHARLES  C/P,P
PROTEUS  MCENROE, RICHARD S.  C/PA,PA
PROTEUS COMBINED [PROTEUS]  SHEFFIELD, CHARLES  P
PROTEUS IN THE UNDERWORLD [PROTEUS 3]  SHEFFIELD, CHARLES  C/P,P,H
PROTEUS MANIFEST [PROTEUS 1,2]  SHEFFIELD, CHARLES  H
PROTEUS OPERATION, THE  HOGAN, JAMES P.  H:2,P,SFWA
PROTEUS UNBOUND [PROTEUS 2]  SHEFFIELD, CHARLES  C/P,P
PROTOSTARS  GERROLD, DAVID  PA,C/PA
PROUD ENEMY, THE [DEMU 2]  BUSBY, F. M.  C/P:2,P
PROUD HELIOS [STAR TREK DS9]  SCOTT, MELISSA  C/P:3,P
PROVEN GUILTY [DRESDEN FILES 8]  BUTCHER, JIM  H
PROVENANCE OF SHADOWS [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
PROXIES  MIXON, LAURA J.  C/P,P,H
PROXIMA PROJECT, THE  RACKHAM, JOHN  DH091
PROXY INTELLIGENCE AND OTHER MIND BENDERS, THE  VAN VOGT, A. E.  C/P:2,P
PRYDAIN CHRONICLES, THE [PRYDAIN 1,2,3,4,5,6]  ALEXANDER, LLOYD  H
PSALMS OF HEROD, THE [BECCA 1]  FRIESNER, ESTHER M.  P
PSEUDO-CITY  WILSON, D. HARLAN  LP
PSEUDO ONE, THE [@PERRY RHODAN #44]  ACE  PA,C/PA
PSEUDO-PEOPLE, THE  NOLAN, WILLIAM F.  HA,PA
PSHRINKS ANONYMOUS  ASIMOV, JANET  C/P,P
PSI-ENCE FICTION [DOCTOR WHO]  BOUCHER, CHRIS  P
PSI HIGH AND OTHERS  NOURSE, ALAN E.  C/P:2,P
PSI HUNT  KURLAND, MICHAEL  C/P:2,P
PSI/NET  WILLIAMS, BILLY DEE/MACGREGOR, ROB  P
PSION [CAT 1]  VINGE, JOAN D.  C/P:2,P
PSIONIC MENACE, THE  WOODCOTT, KEITH  DF199
PSTALEMATE  DEL REY, LESTER  C/P:2,P,H
PSYCHAMOK [PSYCHOMECH 3]  LUMLEY, BRIAN  C/P,P
PSYCHEDELIC-40  CHARBONNEAU, LOUIS  C/P:2,P
PSYCHIC  HURKOS, PETER  [C/RAN-P]
PSYCHIC KITTY [@SABRINA THE TEENAGE WITCH SALEM #6]  DUBOWSKI, CATHY EAST  P
PSYCHIC WORLD OF BISHOP PIKE, THE  HOLZER, HANS  [C/RAN-P]
PSYCHLONE  BEAR, GREG  C/P:2,SFWA-P,P
PSYCHO [PSYCHO 1]  BLOCH, ROBERT  SFWA-P
PSYCHO-CYBERNETICS  MALTZ, MAXWELL, M. D., F. I. C. S.  [C/RAN-P]
PSYCHO HOUSE [PSYCHO 3]  BLOCH, ROBERT  SFWA-P
PSYCHO II [PSYCHO 2]  BLOCH, ROBERT  SFWA-P
PSYCHODROME [PSYCHODROME 1]  HAWKE, SIMON  C/P,SFWA-P,P
PSYCHOGEIST  DAVIES, L. P.  H:2,P
PSYCHOHISTORICAL CRISIS  KINGSBURY, DONALD  H:2,P
PSYCHOMECH [PSYCHOMECH 1]  LUMLEY, BRIAN  C/P,P
PSYCHOMORPH, THE [@SPACE: 1999 YEAR 2 #4]  BUTTERWORTH, MICHAEL  C/P,P
PSYCHOPATH PLAGUE, THE  SPRUILL, STEVEN G.  C/P,SFWA,SFWA-P,P
PSYCHOPATHS  BLOCH, ROBERT  SFWA
PSYCHOSHOP  BESTER, ALFRED/ZELAZNY, ROGER  H:2,LP
PSYCHOSPHERE [PSYCHOMECH 2]  LUMLEY, BRIAN  C/P,P
PSYCHOTECHNIC LEAGUE, THE  ANDERSON, POUL  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 280
TITLE  AUTHOR  CODE
PSYCHOTROPE [@SHADOWRUN 33]  SMEDMAN, LISA  P
PSYCOP PARTNERS  PRICE, JORDAN CASTILLO  LP
PSYKOSIS  BAIRD, WILHELMINA  P
PTOLEMY’S GATE [@BARTIMAEUS #3]  STROUD, JONATHAN  H
PTOLIA [PTOLIA 1]  BAKER, FRED  P
PUCK OF POOK’S HILL  KIPLING, RUDYARD  SR-H,C/P,SR-P
PUCKY’S GREATEST HOUR [@PERRY RHODAN #81]  ACE  PA
PUERTA DE IVREL, LA [MORGAINE 1]  CHERRYH, C. J.  FP
PUGNACIOUS PEACEMAKER, THE  TURTLEDOVE, HARRY  TD20,SFWA-TD20
PUISSANCE: FACTEUR 3  RAYJEAN, MAX-ANDRE  FP
PULITZER [QUANTUM LEAP]  STORM, L. ELIZABETH  C/P,P
PULLING THROUGH  ING, DEAN  C/P:2,P
PULPHOUSE 1  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 2  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 3  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 4  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 5  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 6  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 7  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 8  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 9  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 10  RUSCH, KRISTINE KATHRYN  SR-HA
PULPHOUSE 11  RUSCH, KRISTINE KATHRYN  SR-HA
PULPS, THE  GOODSTONE, TONY  R/REF-H
PULSAR 1  HAY, GEORGE  SR-PA
PUMPKINS OF TIME, THE  GILDEN, MEL  P
PUNISH THE SINNERS  SAUL, JOHN  P
PUNISHER, THE  GRANT, STEVEN/DUFFY, JO  CX
PUNKTOWN (REVISED) [PUNKTOWN]  THOMAS, JEFFREY  H
PUPPET KING, THE [@DRAGONLANCE CHAOS WAR 4]  NILES, DOUGLAS  P
PUPPET MASTER, THE  DALMAS, JOHN  P
PUPPET MASTERS, THE  HEINLEIN, ROBERT A.  C/P:3,P
PUPPET MASTERS, THE (UNABRIDGED)  HEINLEIN, ROBERT A.  P
PURCELL PAPERS, THE  LE FANU, J. SHERIDAN  SR-H
PURCHASE OF THE NORTH POLE, THE  VERNE, JULES  C/P,P,H
PURE BLOOD [MORGAN OF ALB’NY 1]  MCQUAY, MIKE  P
PURE COLD LIGHT, THE  FROST, GREGORY  P,SFWA-P
PURE PRODUCT, THE  KESSEL, JOHN  LP
PURGATORIO, THE (CIARDI TRANSLATION)  ALIGHIERI, DANTE  P
PURGATORY  RESNICK, MIKE  C/P:2,P
PURGATORY ZONE, THE  DARNAY, ARSEN  C/P,P
PURITY IN DEATH [EVE DALLAS]  ROBB, J. D.  P
PURITY PLOT, THE [FAMILY D’ALEMBERT #6]  SMITH, E. E./EKLUND, GORDON  C/P,@P
PURLOINED PLANET, THE  CARTER, LIN  SR-DB1010
PURLOINED PRINCE, THE  WALLACE, IAN  H
PURPLE-6  BRINTON, HENRY  H,P
PURPLE ACES [@G-8 #2]  HOGAN, ROBERT J.  P
PURPLE BOOK, THE  FARMER, PHILIP JOSE  C/P:2,P
PURPLE CLOUD, THE  SHIEL, M. P.  C/P,P,H
PURPLE DRAGON, THE [DOC SAVAGE #91]  ROBESON, KENNETH  @P
PURPLE ISLANDS  CARTER, DEE  SR-P
PURPLE PIRATE, THE  MUNDY, TALBOT  SR-P
PURPLE PRINCE OF OZ, THE  THOMPSON, RUTH PLUMLY  H
PURPLE PTERODACTYLS, THE  DE CAMP, L. SPRAGUE  SR-H,C/P:2,P
PURPLE SAPPHIRE, THE  TAINE, JOHN  SR-H
PURPLE WIZARD, THE  GRIDBAN, VOLSTED  SR-P
PURPLE ZOMBIE, THE [@AVENGER #27]  ROBESON, KENNETH  C/P:2,P

PURRFECT CRIME [PURRFECT SERIES] WAUGH, CAROL-LYNN ROSSEL/

GREENBERG, MARTIN/ASIMOV, ISAAC

PA

PURRFECT PLUNDER [@SPACEWAYS #6]  CLEVE, JOHN  P
PURSUIT [@SURVIVALIST #13]  AHERN, JERRY  P
PURSUIT OF DIANA, THE [V]  WOLD, ALLEN L.  C/P,SFWA-P,P
PURSUIT OF THE PHOENIX, THE  PRICE, JOHN-ALLEN  P
PURSUIT OF THE SCREAMER [KANTMORIE 1]  DIBELL, ANSEN  C/P:2,P
PURSUIT ON GANYMEDE  RESNICK, MIKE  P
PUSHCART WAR, THE  MERRILL, JEAN  LP
PUSHING ICE  REYNOLDS, ALASTAIR  LP
PUSS IN BOOTS  JONES, DIANA WYNNE  LP
PUTESHESTVIYA V PROSHLOE I BUDUSHCHEE  OBRUCHEV, V. A.  FH
PUTTING UP ROOTS  SHEFFIELD, CHARLES  C/P,P,H
PUZZLE [EARTH 2]  DALTON, SEAN  C/P,P
PUZZLE OF THE SPACE PYRAMIDS  BINDER, EANDO  P
PUZZLE PLANET, THE  LOWNDES, ROBERT A. W.  DD485
PYRAMID SCHEME  FREER, DAVE/FLINT, ERIC  H

TITLE  AUTHOR  CODE
PYRAMIDS [DISCWORLD 7]  PRATCHETT, TERRY  H:2,P
PYRAMIDS [PILGRIM 1]  SABERHAGEN, FRED  C/P:2,P
PYRAMIDS FROM SPACE, THE  BERTIN, JACK  C/P,P
PYROMANCER [MANCER 1]  CALLANDER, DON  P
PYTHON ISLE [DOC SAVAGE]  ROBESON, KENNETH  C/P:2,P
PYTHON MEN OF LOST CITY  HAWKS, CHESTER  SR-LP
PZYCHE  HEMINGWAY, AMANDA  H

TITLE  AUTHOR  CODE
Q COLONY  THURSTON, ROBERT  P
Q-IN-LAW [STAR TREK TNG]  DAVID, PETER  C/P:2,P
Q-SPACE [STAR TREK TNG]  COX, GREG  C/P,P
Q-SQUARED [STAR TREK TNG]  DAVID, PETER  H:2,P
Q-STRIKE [@STAR TREK TNG]  COX, GREG  C/P,P
Q-ZONE [STAR TREK TNG]  COX, GREG  C/P:2,P
QE2 IS MISSING, THE  HARRISON, HARRY  P
QINMEARTHA AND THE GIRL CHILD LOCHI/TOMB OF THE OLD ONES  GRANT, JOHN/WILSON, COLIN  LP
QUAD WORLD  METZGER, ROBERT A.  LP
QUADRIPHOBIA  RYAN, ALAN  H,SFWA
QUAG KEEP [QUAG KEEP 1]  NORTON, ANDRE  H:2,P
QUAKING LANDS, THE [@JADE DEMONS #1]  VARDEMAN, ROBERT E.  P
QUALINESTI, THE [@DRAGONLANCE ELVEN NATIONS #3]  THOMPSON, PAUL B./CARTER, TONYA R.  P
QUALITY OF MERCY, THE  COMPTON, D. G.  P
QUANTUM ARCHANGEL, THE [DOCTOR WHO PAST DOCTORS]  HINTON, CRAIG  P
QUANTUM CONNECTION, THE  TAYLOR, TRAVIS S.  H
QUANTUM GATE  HAWKINS, JANE E.  C/P,P
QUANTUM LEAP [QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,SFWA-P,P
QUANTUM MOON [TY MERRICK 1]  VITOLA, DENISE  P
QUANTUM MURDER, A [GREG MANDEL 2]  HAMILTON, PETER F.  C/P,P,H
QUANTUM ROSE, THE [SKOLIAN EMPIRE 6]  ASARO, CATHERINE  H,P
QUARANTINE [@STAR TREK VOYAGER STARFLEET ACADEMY #3]  BARNES-SVARNEY, PATRICIA  LP
QUARANTINE  EGAN, GREG  C/P,P,H
QUARANTINE [@STAR TREK TNG DOUBLE HELIX 4]  VORNHOLT, JOHN  P
QUARANTINE WORLD  LEINSTER, MURRAY  P
QUARK 1  DELANY, SAMUEL R./HACKER, MARILYN  PA,C/PA
QUARK 2  DELANY, SAMUEL R./HACKER, MARILYN  C/PA,PA
QUARK 3  DELANY, SAMUEL R./HACKER, MARILYN  C/PA,PA
QUARK 4  DELANY, SAMUEL R./HACKER, MARILYN  PA,C/PA
QUARREL WITH THE MOON  CONWAY, J. C.  SFWA-P
QUARRELING, THEY MET THE DRAGON [NAPHAR 1]  BAKER, SHARON  P
QUARTERED SEA, THE [QUARTER SINGERS 4]  HUFF, TANYA  C/P,P
QUARTET  MARTIN, GEORGE R. R.  SR-H
QUAS STARBRITE  BERRY, JAMES R.  C/P,P
QUASAR  NASIR, JAMIL  C/P,P
QUATERMASS  KNEALE, NIGEL  SR-P
QUATERMASS AND THE PIT  KNEALE, NIGEL  SR-P
QUATERMASS EXPERIMENT, THE  KNEALE, NIGEL  SR-P
QUATERMASS II  KNEALE, NIGEL  SR-P
QUEEN CITY JAZZ [NANOTECH VERITY 1]  GOONAN, KATHLEEN ANN  C/P,P
QUEEN CLEOPATRA  MUNDY, TALBOT  SR-P
QUEEN IN DANGER  HALL, ADAM  P
QUEEN OF AIR AND DARKNESS AND OTHER STORIES, THE  ANDERSON, POUL  C/P:2,P
QUEEN OF ANGELS [QUEEN OF ANGELS 1]  BEAR, GREG  C/P:2,P
QUEEN OF ASHES, THE [CALEDON 2]  HARRIS, DEBORAH TURNER  C/P,P
QUEEN OF BLADES [@STARCRAFT]  ROSENBERG, AARON  P
QUEEN OF BLOOD  NUETZEL, CHARLES  SR-L
QUEEN OF COINS  HALL, STEPHANIE  P
QUEEN OF DARKNESS, THE  CONNER, MIGUEL  P
QUEEN OF DEATH, THE [@EBERRON LOST MARK #3]  FORBECK, MATT  P
QUEEN OF DEMONS [LORD OF THE ISLES 2]  DRAKE, DAVID  C/P,P,H:2
QUEEN OF DENIAL [DREWCILA QWAH 1]  ROSEN, SELINA  LP
QUEEN OF KNIGHTS  BARRIE, MONICA  P
QUEEN OF MARS [CATS OF MARS 3]  SARRANTONIO, AL  P
QUEEN OF SORCERY [BELGARIAD 2]  EDDINGS, DAVID  C/P:2,P
QUEEN OF SPRINGTIME, THE [WINTER 2]  SILVERBERG, ROBERT  LP
QUEEN OF SWORDS  KOTZWINKLE, WILLIAM  H,P
QUEEN OF SWORDS  TARR, JUDITH  H
QUEEN OF THE AMAZONS  TARR, JUDITH  H:2,LP
QUEEN OF THE DAMNED, THE [LESTAT 3]  RICE, ANNE  H:2,P
QUEEN OF THE DARKNESS [BLACK JEWELS 3]  BISHOP, ANNE  C/P,P
QUEEN OF THE DAWN  HAGGARD, H. RIDER  H
QUEEN OF THE DEMONWEB PITS [@GREYHAWK]  KIDD, PAUL  P
QUEEN OF THE DEPTHS [@FORGOTTEN REALMS PRIESTS]  BYERS, RICHARD LEE  P
QUEEN OF THE LEGION, THE  WILLIAMSON, JACK  C/P,P
QUEEN OF THE SLAYERS [@BUFFY]  HOLDER, NANCY  H,LP
QUEEN OF THE SUMMER STARS  WOOLLEY, PERSIA  H
QUEEN OF THE SWORDS, THE [CORUM 2]  MOORCOCK, MICHAEL  C/P:2,P
QUEEN OF ZAMBA, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P:2,P
QUEEN SHEBA’S RING  HAGGARD, H. RIDER  SR-L
QUEEN VICTORIA’S REVENGE  HARRISON, HARRY  SFWA-P
QUEEN ZIXI OF IX; OR, THE STORY OF THE MAGIC CLOAK  BAUM, L. FRANK  H,LP
QUEENMAGIC, KINGMAGIC  WATSON, IAN  H,P

The MITSFS Titledex  September 24, 2007  Page 284
TITLE  AUTHOR  CODE
QUEEN’S CURSE, THE [@CHARMED]  HARRISON, EMMA  P
QUEEN’S GAMBIT, THE [SWORD,RING AND CHALICE PHERESA 1]  CHESTER, DEBORAH  C/P,P
QUEEN’S GAMBIT, THE  TEVIS, WALTER  LP
QUEEN’S GAMBIT DECLINED  SNODGRASS, MELINDA M.  P,SFWA-P
QUEEN’S KNIGHT, THE [SWORD,RING AND CHALICE PHERESA 2]  CHESTER, DEBORAH  P
QUEEN’S NECKLACE, THE  EDGERTON, TERESA  LP
QUEENS OF DELIRIA  BUTTERWORTH, MICHAEL  SR-P
QUEENS OF LAND AND SEA [STORMLANDS 5]  ROBERTS, JOHN MADDOX  P
QUEEN’S OWN [HERALDS OF VALDEMAR 1,2,3]  LACKEY, MERCEDES  H
QUEEN’S SQUADRON, THE  MELUCH, R. M.  SFWA-P
QUEENS WALK IN THE DUSK  SWANN, THOMAS BURNETT  SR-H
QUEENSBLADE  SHWARTZ, SUSAN  C/P,P
QUEER DESTINIES  BOWEN, GARY  SR-L
QUEST, THE [@SURVIVALIST #3]  AHERN, JERRY  P
QUEST BEGINS, THE [COMPLETE ELFQUEST]  PINI, WENDY/PINI, RICHARD  LP
QUEST BEYOND THE STARS  HAMILTON, EDMOND  P
QUEST CROSSTIME [CROSSTIME]  NORTON, ANDRE  H:2,P
QUEST FOR APOLLO  LAHEY, MICHAEL  P
QUEST FOR ATLANTIS  LEONARD, R. CEDRIC  [C/RAN-P]
QUEST FOR CUSH, THE [IMARO 2]  SAUNDERS, CHARLES R.  C/P:2,P
QUEST FOR FIRE, THE  ROSNY, J. H.  H
QUEST FOR LOST HEROES [DRENAI 3]  GEMMELL, DAVID  P
QUEST FOR SAINT CAMBER, THE [DERYNI KELSON 3]  KURTZ, KATHERINE  C/P,H,P
QUEST FOR SIMBILIS, A [DYING EARTH]  SHEA, MICHAEL  C/P,P
QUEST FOR TANELORN, THE [CASTLE BRASS 3]  MOORCOCK, MICHAEL  C/P:2,P
QUEST FOR TANELORN, THE (BRITISH) [CASTLE BRASS 3]  MOORCOCK, MICHAEL  SR-P

QUEST FOR THE FALLEN STAR ANTHONY, PIERS/RICHEY, JAMES/

RIGGS, ALAN

H

QUEST FOR THE FARADAWN [FARADAWN 1]  FORD, RICHARD  H,P
QUEST FOR THE FUTURE  VAN VOGT, A. E.  H:2,P
QUEST FOR THE 36, THE  BILLIAS, STEPHEN  C/P,P
QUEST FOR THE TRILOGY, THE [ROVER 4]  ODOM, MEL  H
QUEST FOR THE WELL OF SOULS [WELL WORLD 3]  CHALKER, JACK L.  C/P:3,P
QUEST FOR THE WHITE WITCH  LEE, TANITH  P
QUEST OF EXCALIBUR, THE  WIBBERLEY, LEONARD  H,LP
QUEST OF KADJI, THE  CARTER, LIN  C/P:2,P
QUEST OF QUI [DOC SAVAGE #12]  ROBESON, KENNETH  C/P:2,@P
QUEST OF THE DARK LADY  READE, QUINN  P
QUEST OF THE DAWN MAN  ROSNY, J. H.  P
QUEST OF THE DAWNSTAR  MCBAIN, GORDON  P
QUEST OF THE GYPSY [GYPSY 1]  GOULART, RON  C/P,P
QUEST OF THE SPIDER [DOC SAVAGE #68]  ROBESON, KENNETH  C/P:2,@P
QUEST OF THE THREE WORLDS  SMITH, CORDWAINER  C/P:2,P
QUEST THROUGH SPACE AND TIME [@PERRY RHODAN #9]  ACE  PA,C/PA
QUEST TO RIVERWORLD [RIVERWORLD]  FARMER, PHILIP JOSE  PA
QUESTER’S ENDGAME [DIADEM 9]  CLAYTON, JO  C/P:2,P
QUESTING HERO, THE [@WIZARD WAR #2]  COOK, HUGH  C/P:2,P
QUESTING OF KEDRIGERN, THE [KEDRIGERN 2]  MORRESSY, JOHN  C/P,P
QUESTION AND ANSWER  ANDERSON, POUL  C/P:2,P
QUESTION OF DESTINY, A  SERVICE, PAMELA F.  H
QUESTION OF TIME, A [DRACULA]  SABERHAGEN, FRED  C/P,P,H
QUESTION QUEST [@XANTH #14]  ANTHONY, PIERS  C/P:2,P
QUESTOR TAPES, THE  FONTANA, D. C.  C/P:2,P
QUEST’S END [@COMPLETE ELFQUEST #4]  PINI, WENDY/PINI, RICHARD  CX
QUETE COSMIQUE, LA (FRENCH) [@PERRY RHODAN #11]  SCHEER, K. H./DARLTON, CLARK  FPR
QUI PARLE DE CONQUETE  WRIGHT, LAN  FP
QUICK, THE  COLE, BURT  H
QUICK AND EASY MATH  ASIMOV, ISAAC  R/REF-H
QUICKENING, THE  DE ANGELES, LY  LP
QUICKER THAN THE EYE  BRADBURY, RAY  P
QUICKSAND  BRUNNER, JOHN  H:2,P
QUICKSILVER [BAROQUE CYCLE 1]  STEPHENSON, NEAL  H:3,LP
QUICKSILVER SCREEN, THE  DEBRANDT, DON H.  P
QUICKSILVER’S KNIGHT [WARLOCK’S HEIRS 3]  STASHEFF, CHRISTOPHER  P
QUIET INVASION, THE  ZETTEL, SARAH  H:2,P
QUIET KNIVES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 9]  LEE, SHARON/MILLER, STEVE  SR-LP
QUIET MAGIC  LEE, SHARON/MILLER, STEVE  SR-LP
QUIET NIGHT OF FEAR, A  GRANT, CHARLES L.  C/P,P
QUIET OF STONE, A  LEIGH, STEPHEN  C/P:2,P
QUIET PLACE, THE [@STAR TREK NEW FRONTIER #7]  DAVID, PETER  P
QUIET PLACE, THE  MAYNARD, RICHARD  P
QUIET POOLS, THE  KUBE-MCDOWELL, MICHAEL P.  H:2,SFWA:2,SFWA-P:2,P
QUILLIAN SECTOR, THE [DUMAREST 19]  TUBB, E. C.  C/P,P

TITLE  AUTHOR  CODE
QUINCEY MORRIS, VAMPIRE  ELROD, P. N.  C/P:2,P
QUINCUNX OF TIME, THE  BLISH, JAMES  C/P:2,P
QUINTESSENTIAL WORLD OF DARKNESS, THE [WORLD OF DARKNESS]  WIECK, STEWART/BRANSCOME, ANNA  LPA
QUOTABLE STAR WARS, THE  SANSWEET, STEPEHN J.  R/REF-P
QUOTATIONS FROM CHAIRMAN MAO TSE-TUNG  MAO TSE-TUNG  [C/RAN-P]
QUOZL  FOSTER, ALAN DEAN  C/P:2,SFWA-P,P
QUY EFFECT, THE  SELLINGS, ARTHUR  C/P:2,P

TITLE  AUTHOR  CODE
R. CRUMB’S CARLOAD O’ COMICS  CRUMB, R.  CX
R IS FOR ROCKET  BRADBURY, RAY  C/P:2,P
R-MASTER, THE  DICKSON, GORDON R.  H:2,P
R-T, MARGARET, AND THE RATS OF NIMH [NIMH 3]  CONLY, JANE LESLIE  H
R. U. R.  CAPEK, KAREL  P
RABBI’S CAT, THE  SFAR, JOANN  CX
RABELAISIAN REPRISE [RABELAIS 3]  CARR, JAYGE  H,SFWA
RACE ACROSS THE STARS [@SPACEWAYS #18]  CLEVE, JOHN  P
RACE AGAINST TIME  ANTHONY, PIERS  C/P:2,SFWA-P,P
RACE AGAINST TIME [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  [C/RAN-P]
RACE FOR GOD, THE  HERBERT, BRIAN  C/P,SFWA-P,P
RACE TO THE STARS  MARGULIES, LEO/FRIEND, OSCAR J.  PA
RACHEL IN LOVE  MURPHY, PAT  B
RACHEL, THE POSESSESED [@PHENWICK #5]  KIMBROUGH, KATHERYN  P
RADAR ALERT  ZEIGFREID, KARL  SR-P,C/P
RADIANCE  SCHOLZ, CARTER  H,LP
RADIANT [LEAGUE OF PEOPLES 6]  GARDNER, JAMES ALAN  H
RADIANT DAWN [RADIANT DAWN 1]  GOODFELLOW, CODY  LP
RADIANT DOME, THE [@PERRY RHODAN #2]  ACE  PA,C/PA
RADIANT DRAGON, THE [SPELLJAMMER CLOAKMASTER 4]  CUNNINGHAM, ELAINE  C/P,P
RADIANT SEAS, THE [SKOLIAN EMPIRE 4]  ASARO, CATHERINE  C/P,P,H
RADIANT WARRIOR, THE [CONRAD STARGARD 3]  FRANKOWSKI, LEO  C/P:2,P
RADICAL UTOPIAS  BOOK-OF-THE-MONTH CLUB  HA
RADIO BEASTS, THE  FARLEY, RALPH M.  C/P:2,P
RADIO FREE ALBEMUTH  DICK, PHILIP K.  H:2,P
RADIO MAN, THE  FARLEY, RALPH M.  SR-LP
RADIO PLANET, THE  FARLEY, RALPH M.  C/P:2,P
RADIOACTIVE CAMEL AFFAIR, THE [@MAN FROM UNCLE #7]  LESLIE, PETER  C/P:2,P
RADIOACTIVE REDHEAD, THE [ZACHARY NIXON JOHNSON 3]  ZAKOUR, JOHN/GANEM, LAWRENCE  P
RADIUM POOL, THE  REPP, ED EARL  SR-H,H
RADIUM TERRORS, THE  DORRINGTON, ALBERT  SR-H
RADIUS OF DOUBT [@PATTERNS OF CHAOS #1]  INGRID, CHARLES  P
RADIX [RADIX 1]  ATTANASIO, A. A.  C/P:2,P
RADON FILE, THE [TY MERRICK 5]  VITOLA, DENISE  P
RAFT [XEELEE]  BAXTER, STEPHEN  C/P:2,SFWA-P,P
RAG, A BONE, AND A HANK OF HAIR, A  FISK, NICHOLAS  H
RAGA SIX  LAURIA, FRANK  P
RAGAMUFFIN  BUCKELL, TOBIAS S.  H
RAGE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #1]  BYERS, RICHARD LEE  P
RAGE FOR REVENGE, A [WAR AGAINST THE CHTORR 3]  GERROLD, DAVID  C/P:2,P
RAGE IN HEAVEN, A (COMPLETE) [@YAMATO #1]  KATO, KEN  H
RAGE IN HEAVEN, A (PART 1) [YAMATO]  KATO, KEN  P
RAGE IN HEAVEN, A (PART 2) [@YAMATO #1]  KATO, KEN  C/P,P
RAGE OF A DEMON KING [SERPENTWAR 3]  FEIST, RAYMOND E.  C/P,P,H
RAGGED ASTRONAUTS, THE [TOLLER MARAQUINE 1]  SHAW, BOB  H:2,SFWA,P
RAGGED EDGE, THE  CHRISTOPHER, JOHN  H:2,P
RAGGED WORLD, THE  MOFFETT, JUDITH  H,P
RAGNAROCK [@SHADOWRUN 38]  KENSON, STEPHEN  P
RAGNAROK [STAR TREK VOYAGER]  ARCHER, NATHAN  P
RAGNAR’S CLAW [WARHAMMER 40K SPACE WOLF 2]  KING, WILLIAM  P
RAGS [DOCTOR WHO PAST DOCTORS]  LEWIS, MICK  P
RAGS AND OLD IRON  SHANNON, LORELEI  H
RAGWITCH, THE  NIX, GARTH  P
RAHNE [LIVING PLANET 1]  COON, SUSAN  P
RAHXEPHON, VOL. 1 [RAHXEPHON 1]  IZUBUCHI, YUTAKA/BONES  SCX
RAHXEPHON, VOL. 2 [RAHXEPHON 2]  IZUBUCHI, YUTAKA/BONES  SCX
RAID OF "LE VENGEUR", THE  GRIFFITH, GEORGE  SR-LP
RAIDERS FROM THE RINGS  NOURSE, ALAN E.  C/P:2,P
RAIDERS OF GOR [GOR 6]  NORMAN, JOHN  C/P:2,P
RAIDERS OF NOOMAS  NUETZEL, CHARLES  SR-P
RAIDERS OF THE LOST ARK  BLACK, CAMPBELL  C/P,P
RAIDERS OF THE LOST CAR PARK [CORNELIUS MURPHY 2]  RANKIN, ROBERT  P
RAIDERS OF THE REVOLUTION [@FREEDOM’S RANGERS 2]  ANDREWS, KEITH WILLIAM  P
RAIN IN THE DOORWAY  SMITH, THORNE  C/P:2,P
RAINBOW ABYSS, THE [SUN-CROSS 1]  HAMBLY, BARBARA  C/P:2,P
RAINBOW AFFAIR, THE [@MAN FROM UNCLE #13]  MCDANIEL, DAVID  C/P,P
RAINBOW ANNALS, THE  DAVIS, GRANIA  P
RAINBOW CADENZA, THE  SCHULMAN, J. NEIL  C/P,SFWA,P,H
RAINBOW MAN  ENGH, M. J.  LP
RAINBOW MARS  NIVEN, LARRY  H:2,P
RAINBOW SWORD, THE  MARTINE-BARNES, ADRIENNE  C/P:2,P
RAINBOWS END  VINGE, VERNOR  H
RAINS OF ERIDAN, THE  HOOVER, H. M.  C/P,P,H
RAISE THE TITANIC!  CUSSLER, CLIVE  C/P,P

The MITSFS Titledex  September 24, 2007  Page 288
TITLE  AUTHOR  CODE
RAISING THE PAST  ROBINSON, JEREMY  LP
RAISING THE STONES  TEPPER, SHERI S.  H:2,P
RAJAN  LUKEMAN, TIM  H,SFWA,P
RAKEHELLS OF HEAVEN, THE  BOYD, JOHN  C/P,P,H
RAKKETY TAM [REDWALL]  JACQUES, BRIAN  LP
RALESTONE LUCK  NORTON, ANDRE  C/P,SFWA-P,P
RALLY CRY [@LOST REGIMENT 1]  FORSTCHEN, WILLIAM R.  H,P
RALLY POINT [@DEMONTECH 2]  SHERMAN, DAVID  P
RALPH 124C41+  GERNSBACK, HUGO  C/P:2,P,H
RAM SONG [EARTH SONG 3]  WEBB, SHARON  C/P,SFWA,P,H
RAMA II [RAMA 2]  CLARKE, ARTHUR C./LEE, GENTRY  C/P,P,H
RAMA REVEALED [RAMA 4]  CLARKE, ARTHUR C./LEE, GENTRY  H:2,P
RAMPAGE [@DOOM’S DAY 1]  FINGEROTH, DANNY/FEIN ERIC  P
RAMSGATE PARADOX, THE  TALL, STEPHEN  C/P,P
RANDOM ACCESS MESSAGES OF THE COMPUTER AGE  MONTELEONE, THOMAS F.  SR-LPA
RANDOM ACTS OF SENSELESS VIOLENCE [DRYCO 5]  WOMACK, JACK  H,LP
RANDOM FACTOR  SHERMAN, JOEL HENRY  P
RANDOM HOUSE BOOK OF FANTASY STORIES, THE  ASHLEY, MIKE  LPA
RANDOM HOUSE BOOK OF SCIENCE FICTION STORIES, THE  ASHLEY, MIKE  LPA
RANDOM MEASURES [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,P
RANGER TO THE MOON  LEY, WILLY  [C/RAN-P]
RANIA  RUDHYAR, DANE  P
RANKIN: ENEMY OF THE STATE  OSIER, JOHN  P
RANKS OF BRONZE [RANKS OF BRONZE]  DRAKE, DAVID  C/P:2,P
RANSOM OF BLACK STEALTH ONE, THE  ING, DEAN  P,SFWA-P
RAPE OF THE SUN, THE  WALLACE, IAN  C/P,P
RAPHAEL [DAMIANO 3]  MACAVOY, R. A.  C/P:2,P

RAPPAREE, THE [@VANCE INTEGRAL EDITION #4]/ BIG PLANET [@VANCE INTEGRAL EDITION #4]/ VANDALS OF THE VOID [@VANCE INTEGRAL EDITION #4] VANCE, JACK

SR-H

RAPTOR RED  BAKKER, ROBERT T.  C/P,P
RAPTURE EFFECT, THE  CARVER, JEFFREY A.  H:2,SFWA-P,P
RAPTURE IN DEATH [EVE DALLAS]  ROBB, J. D.  P
RARE BREED, THE  STURGEON, THEODORE  P
RARE SCIENCE FICTION  HOWARD, IVAN  PA,C/PA
RASHANYN DARK [@TIMEQUEST #1]  TEDFORD, WILLIAM  C/P,P
RAT KING [@DEATHLANDS]  AXLER, JAMES  P
RAT RACE, THE  FRANKLIN, JAY  C/P,GN2.10,SR-LP
RAT TRAP, THE [GALACTIC MI 2]  RANDLE, KEVIN  C/P,P
RATH AND STORM [@MAGIC: THE GATHERING]  ARCHER, PETER  PA
RATHA’S CREATURE [RATHA 1]  BELL, CLARE  C/P,SFWA,P
RATNER’S STAR  DE LILLO, DON  H,P
RATS AND GARGOYLES  GENTLE, MARY  C/P,SFWA,P
RATS, BATS, & VATS [RATS & BATS 1]  FREER, DAVE/FLINT, ERIC  H
RATS, THE BATS & THE UGLY, THE [RATS & BATS 2]  FLINT, ERIC/FREER, DAVE  H
RAUM  SHERRELL, CARL  C/P,P
RAUMSCHIFF ERDE [@PERRY RHODAN SILBERBANDE #76]  PABEL MOEWIG KG  FPR
RAVEN [CLEVELAND UNDEAD 1]  SWINIARSKI, S. A.  P
RAVEN GAME, THE [@WATCHERS #3]  MOR, CAISEAL  P
RAVEN RING, THE [LYRA 2]  WREDE, PATRICIA C.  P
RAVEN RISING [@OMEGA SUB #6]  CAMERON, J. D.  P
RAVEN WARRIOR, THE [TALES OF GUINEVERE 2]  BORCHARDT, ALICE  H
RAVENGERS, THE [CYBERPUNK 2.0.2.0.]  BILLIAS, STEPHEN  P
RAVENHEART [RIGANTE 3]  GEMMELL, DAVID  H,P
RAVENNA AND HER GHOSTS  LEE, VERNON  SR-P
RAVENOR [@WARHAMMER 40K RAVENOR 1]  ABNETT, DAN  P
RAVENOR RETURNED [@WARHAMMER 40K RAVENOR 2]  ABNETT, DAN  H
RAVENOR ROGUE [@WARHAMMER 40K RAVENOR 3]  ABNETT, DAN  H
RAVENOUS DUSK [RADIANT DAWN 2]  GOODFELLOW, CODY  LP
RAVENS’ GATHERING [@BARD #4]  TAYLOR, KEITH  P
RAVENS OF THE MOON, THE  GRANT, CHARLES L.  H,SFWA
RAVEN’S SHADOW [SERAPH 1]  BRIGGS, PATRICIA  P
RAVEN’S STRIKE [SERAPH 2]  BRIGGS, PATRICIA  P
RAVNICA: CITY OF GUILDS [MAGIC: THE GATHERING RAVNICA 1]  HERNDON, CORY J.  P
RAVNOS [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  RYAN, KATHLEEN  P
RAVYN’S FLIGHT [JARVED 9 1]  O’SHEA, PATTI  P
RAW PAIN MAX  ANDERSSON, C. DEAN  P
RAX  CONEY, MICHAEL G.  P

RAY BRADBURY OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

R/REF-H

RAYMOND E. FEIST’S WORLD OF KRONDOR (CD) [RIFTWAR]  SIERRA STUDIOS  SR-MM
RAZOR’S EDGE [SHOLAN ALLIANCE 4]  NORMAN, LISANNE  P
RE-ANIMATOR  ROVIN, JEFF  P
RE-BIRTH  WYNDHAM, JOHN  C/P:2,P

TITLE  AUTHOR  CODE
RE: COLONISED PLANET 5, SHIKASTA [CANOPUS IN ARGOS 1]  LESSING, DORIS  H,LP
RE-ENTER FU MANCHU  ROHMER, SAX  C/P:2,P
RE-ENTRY  PREUSS, PAUL  C/P:2,P
REACH  GIBSON, EDWARD  H,P
REACH FOR TOMORROW  CLARKE, ARTHUR C.  SR-H,C/P:2,P
REACTOR XK9  MULLER, JOHN E.  SR-P
READERS’ FAVORITES FROM ANALOG 1980 TO 1990  SCHMIDT, STANLEY  B

READER’S GUIDE TO FANTASY, A SEARLES, BAIRD/MEACHAM, BETH/

FRANKLIN, MICHAEL

R/REF-P,C/REF-P

READER’S GUIDE TO SCIENCE FICTION, A SEARLES, BAIRD/LAST, MARTIN/

MEACHAM, BETH/FRANKLIN, MICHAEL

R/REF-P,C/REF-P:2

READER’S GUIDE TO THE SILMARILLION, A  KOCHER, PAUL H.  R/REF-H,C/REF-H
READINGS IN FOLK-LORE  SKINNER, HUBERT M.  [C/RAN-H]
REAL STORY, THE [GAP 1]  DONALDSON, STEPHEN R.  H:2,P
REAL UFO INVASION, THE  PALMER, RAYMOND A.  [C/RAN-P]
REALITY DUST  BAXTER, STEPHEN  GZD07306
REALITY DYSFUNCTION [NIGHT’S DAWN 1]  HAMILTON, PETER F.  H
REALITY DYSFUNCTION 1: EMERGENCE [NIGHT’S DAWN 1]  HAMILTON, PETER F.  C/P:2,P
REALITY DYSFUNCTION 2: EXPANSION [NIGHT’S DAWN 1]  HAMILTON, PETER F.  C/P,P
REALITY FORBIDDEN  HIGH, PHILIP E.  P,DG609
REALITY MATRIX, THE  DALMAS, JOHN  C/P:2,P
REALITY TRIP AND OTHER IMPLAUSIBILITIES, THE  SILVERBERG, ROBERT  C/P:2,P
REALM BENEATH, THE [AVERIDAN 2]  CLOUGH, BRENDA W.  C/P:2,P
REALM OF CHAOS [WARHAMMER]  GASCOIGNE, MARC/JONES, ANDY  PA
REALM OF LIGHT [REIGN OF SHADOWS 3]  CHESTER, DEBORAH  P
REALM OF THE GODS [ARRIS 2]  COOKE, CATHERINE  P
REALM OF THE TRI-PLANETS [@PERRY RHODAN #31]  ACE  PA,C/PA
REALMS OF INFAMY [FORGOTTEN REALMS]  LOWDER, JAMES  PA
REALMS OF MAGIC [@FORGOTTEN REALMS]  THOMSEN, BRIAN M./KING, J. ROBERT  PA
REALMS OF MYSTERY [@FORGOTTEN REALMS]  ATHANS, PHILIP  PA

REALMS OF SHADOW [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS] BALDWIN, LIZZ

PA

REALMS OF TARTARUS, THE  STABLEFORD, BRIAN M.  C/P,P
REALMS OF THE ARCANE [@FORGOTTEN REALMS]  THOMSEN, BRIAN M.  PA
REALMS OF THE DEEP [@FORGOTTEN REALMS THREAT FROM THE SEA]  ATHANS, PHILIP  PA

REALMS OF THE DRAGONS [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS] ATHANS, PHILIP

PA

REALMS OF THE DRAGONS II [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS] ATHANS, PHILIP

PA

REALMS OF THE ELVES [@FORGOTTEN REALMS LAST MYTHAL]  ATHANS, PHILIP  PA
REALMS OF THE GODS, THE [@IMMORTALS (PIERCE) #4]  PIERCE, TAMORA  P
REALMS OF THE UNDERDARK [@FORGOTTEN REALMS UNDERDARK]  KING, J. ROBERT  PA
REALMS OF TOLKIEN [LORD OF THE RINGS]  HARPERPRISM  A
REALMS OF VALOR [FORGOTTEN REALMS]  LOWDER, JAMES  PA
REALMS OF WIZARDRY  CARTER, LIN  HA,SFWA
REALTIME  TIEDEMANN, MARK W.  LP
REALTIME INTERRUPT  HOGAN, JAMES P.  C/P,P,H
REALTY CHECK  ANTHONY, PIERS  LP
REALWARE [WARE 4]  RUCKER, RUDY  LP
REAP THE EAST WIND [DREAD EMPIRE 5/RAGNARSON 7]  COOK, GLEN  C/P:2,SFWA-P,P
REAP THE WHIRLWIND [SWORD OF KNOWLEDGE 3]  CHERRYH, C. J./LACKEY, MERCEDES  C/P:2,SFWA-P:2,P
REAP THE WHIRLWIND [@STAR TREK VANGUARD]  MACK, DAVID  P
REAPER MAN [DISCWORLD 11/DISCWORLD DEATH 2]  PRATCHETT, TERRY  H:2,SFWA-P,P
REAPER OF SOULS [@WARHAMMER MALUS DARKBLADE]  ABNETT, DAN/LEE, MIKE  P
REAPING, THE  TAYLOR, BERNARD  P
REASONABLE WORLD, A [CV 3]  KNIGHT, DAMON  H:2,SFWA,P
REASSEMBLED MAN, THE  KASTLE, HERBERT D.  C/P:2,P
REAVE THE JUST AND OTHER TALES  DONALDSON, STEPHEN R.  H
REAVER ROAD, THE  DUNCAN, DAVE  H,P
REAVERS OF SKAITH, THE [SKAITH 3]  BRACKETT, LEIGH  C/P:2,P
REAVERS OF THE BLOOD SEA [@DRAGONLANCE CHAOS WAR 5]  KNAAK, RICHARD A.  P
REBECCA, THE MYSTERIOUS [@PHENWICK #7]  KIMBROUGH, KATHERYN  P
REBEL AGENT [@STAR WARS DARK FORCES 2]  DIETZ, WILLIAM C.  H
REBEL ATTACK [@MUTANTS AMOK #3]  GRANT, MARK  P
REBEL DAWN [@STAR WARS HAN SOLO 3]  CRISPIN, A. C.  P
REBEL DREAM [@STAR WARS NEW JEDI ORDER ENEMY LINES 1]  ALLSTON, AARON  P
REBEL DYNASTY 2, THE [HULZEIN DYNASTY]  BUSBY, F. M.  P
REBEL FAY [NOBLE DEAD SERIES 1 5]  HENDEE, BARB/HENDEE, J. C.  H:2
REBEL FROM ALPHORION [@PLANET BUILDERS #3]  TALLIS, ROBYN  P
REBEL ICE [@STARDOC 5]  VIEHL, S. L.  P
REBEL IN TIME, A  HARRISON, HARRY  C/P:2,P
REBEL MOON  BETHKE, BRUCE/DAY, VOX  C/P,P
REBEL OF ANTARES [DRAY PRESCOT 24]  AKERS, ALAN BURT  P
REBEL OF RHADA, THE [RHADA 1]  GILMAN, ROBERT CHAM  H:2,P

The MITSFS Titledex  September 24, 2007  Page 290
TITLE  AUTHOR  CODE
REBEL PRINCE, THE  GREEN, SHARON  C/P,P
REBEL STAND [@STAR WARS NEW JEDI ORDER ENEMY LINES 2]  ALLSTON, AARON  P
REBEL SUTRA  LEWITT, SHARIANN  H,P
REBEL WINTER [WARHAMMER 40K]  PARKER, STEVE  P
REBEL WITCH, THE  LOVEJOY, JACK  H
REBEL WORLDS, THE [DOMINIC FLANDRY]  ANDERSON, POUL  C/P:2,P

REBEL WORLDS, THE [DOMINIC FLANDRY]/ KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY] ANDERSON, POUL

P

REBELLERS, THE  ROBERTS, JANE  DF215
REBELLION, THE [@SURVIVALIST #12]  AHERN, JERRY  P
REBELLION [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/RODGERS, ALAN  H
REBELLION [STARGATE]  MCCAY, BILL  C/P:2,P
REBELLION 2456 [BUCK ROGERS MARTIAN WARS 1]  MURDOCK, M. S.  C/P,P
REBELLIOUS STARS, THE  ASIMOV, ISAAC  DD084
REBELS AND TYRANTS [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
REBEL’S CAGE [@ELITA #4]  JACOBY, KATE  LP
REBELS IN ARMS [SCOTT ST. ANDREW 2]  WEAVER, BEN  P
REBELS IN HELL [HEROES IN HELL]  MORRIS, JANET  PA,C/PA
REBELS OF MERKA  FUNNELL, AUGUSTINE  SR-P
REBELS OF THE RED PLANET  FONTENAY, CHARLES L.  P,DF113
REBELS OF TUGLAN, THE [@PERRY RHODAN #12]  ACE  PA,C/PA
REBEL’S QUEST [@SUPREME CONSTELLATIONS #2]  BROOKE, GUN  LP
REBEL’S QUEST [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
REBELS’ SEED [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
REBIRTH [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #2]  KEYES, J. GREGORY  P
REBIRTH OF WONDER, THE  WATT-EVANS, LAWRENCE  C/P,P
REBIRTH — WHEN EVERYONE FORGOT  MCCLARY, THOMAS CALVERT  SR-P
REBORN  SIMON, LEONARD  C/P,P
REBORN [ADVERSARY 4]  WILSON, F. PAUL  P
RECALL NOT EARTH  MACAPP, C. C.  C/P:2,P
RECALLED TO LIFE  SILVERBERG, ROBERT  C/P:2,SFWA,P
RECEIVE THE GIFT [CANTRIX 3]  MARLEY, LOUISE  P
RECKLESS ENGINEERING [DOCTOR WHO EIGHTH DOCTOR]  WALTERS, NICK  P
RECKLESS SLEEP  LEVY, ROGER  P
RECKONING INFINITY  STITH, JOHN E.  P
RECLAMATION  ZETTEL, SARAH  C/P,P,H
RECOIL  NUNES, CLAUDE/NUNES, RHODA  D71082
RECONINGS [@BOOKS OF MAGIC #6]  JABLONSKI, CARLA  P
RECONNAISSANCE [STARGATE]  MCCAY, BILL  P
RECOVERY [STAR TREK]  DILLARD, J. M.  C/P,P
RECRUIT FOR ANDROMEDA  LESSER, MILTON  DD358,P
RECRUITS FOR ARKON [@PERRY RHODAN #76]  ACE  C/PA,PA
RECUSRION [RECURSION 1]  BALLANTYNE, TONY  P
RECYCLED [DREWCILA QWAH 2]  ROSEN, SELINA  LP
RECYCLED SOULS [MIND MASTERS 5]  ROSS, IAN  P
RED ALERT  BRYANT, PETER  C/P:2,P
RED AMERICA [@DOOMSDAY WARRIOR #2]  STACY, RYDER  P
RED ARMY  PETERS, RALPH  H
RED AS BLOOD, OR, TALES FROM THE SISTERS GRIMMER  LEE, TANITH  H:2,P
RED BRANCH  LLYWELYN, MORGAN  P
RED BRIDE  FOWLER, CHRISTOPHER  SFWA
RED CHINDVIT CONSPIRACY, THE  HOLZER, HANS  P
RED DEATH [JONATHAN BARRETT 1]  ELROD, P. N.  P,SFWA-P:2
RED DICE [@LAST VAMPIRE #3]  PIKE, CHRISTOPHER  P
RED DRAGON  HARRIS, THOMAS  [C/RAN-P]
RED DUST [TALES OF TOMORROW 2]  HOUSTON, DAVID  P
RED DUST  MCAULEY, PAUL J.  C/P:2,P
RED DWARF [RED DWARF 1,2]  NAYLOR, GRANT  H
RED DWARF LOG NO. 1996 [RED DWARF]  ALEXANDER, PAUL  [C/RAN-H]
RED EVE  HAGGARD, H. RIDER  H
RED EYE OF BETELGEUSE [@PERRY RHODAN #40]  ACE  PA,C/PA
RED FLAME BURNING [HARRY BORG 1]  HAWKINS, WARD  C/P,P
RED FRIDAY  TURNER, GEORGE KIBBE  SR-H
RED GENESIS [NEXT WAVE 1]  SYKES, S. C.  C/P:2,P
RED HART MAGIC  NORTON, ANDRE  C/P,P,H
RED HEART OF MEMORIES, A [MATILDA BLACK]  HOFFMAN, NINA KIRIKI  LP
RED HEROIN  POURNELLE, JERRY  P
RED HOLOCAUST [@DEATHLANDS]  AXLER, JAMES  P
RED INSECTS, THE  STATTEN, VARGO  SR-P
RED IRON NIGHTS [GARRETT 6]  COOK, GLEN  P,SFWA-P:2
RED KING, THE [STAR TREK TITAN]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
RED LAMP, THE  RINEHART, MARY ROBERTS  H,P
RED LIGHTNING [RED THUNDER 2]  VARLEY, JOHN  H
RED LIMIT FREEWAY [SKYWAY 2]  DE CHANCIE, JOHN  C/P:3,P

The MITSFS Titledex  September 24, 2007  Page 291
TITLE  AUTHOR  CODE
RED MAGIC [@FORGOTTEN REALMS HARPERS #3]  RABE, JEAN  P
RED MAGICIAN, THE  GOLDSTEIN, LISA  C/P:2,SFWA,SFWA-P,LP
RED MARS [MARS 1]  ROBINSON, KIM STANLEY  H:2,P
RED MEN OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
RED MOON [@AVENGER #26]  ROBESON, KENNETH  C/P:2,P
RED MOON AND BLACK MOUNTAIN  CHANT, JOY  H:2,P
RED NAILS [BERKLEY CONAN 2]  HOWARD, ROBERT E.  P
RED NAPOLEON, THE  GIBBONS, FLOYD  H,P
RED ORC’S RAGE  FARMER, PHILIP JOSE  H,P
RED PERI, THE  WEINBAUM, STANLEY G.  SR-H,SR-LP
RED PLANET, THE  CHILTON, CHARLES  SR-P
RED PLANET  HEINLEIN, ROBERT A.  C/P:2,SR-LP,P
RED PLANET  TELEP, PETER  P
RED PLANET, THE  WINTERBOTHAM, RUSS  C/P:2,P
RED PLANET (UNABRIDGED)  HEINLEIN, ROBERT A.  C/P,P
RED PLANET RUN  STABENOW, DANA  P
RED PROPHET [ALVIN MAKER 2]  CARD, ORSON SCOTT  H:2,SFWA-P,P
RED QUEEN, THE  DRAULANS, DIRK  H
RED RAIN [BATMAN]  MOENCH, DOUG  CX
RED ROSE RAGES (BLEEDING), THE  DUCHAMP, L. TIMMEL  LP
RED SANDS  THOMPSON, PAUL B./CARTER, TONYA R.  P
RED SECTOR [@STAR TREK TNG DOUBLE HELIX 3]  CAREY, DIANE  P
RED SHADOWS  NAVARRO, YVONNE  P
RED SHIFT  GARNER, ALAN  H:2,P
RED SKELTON’S FAVORITE GHOST STORIES  SKELTON, RED  PA
RED SKULL, THE [DOC SAVAGE #17]  ROBESON, KENNETH  C/P:2,@P
RED SKY FILE, THE [TY MERRICK 4]  VITOLA, DENISE  P
RED SLIPPERS [@MITHGAR]  MCKIERNAN, DENNIS L.  H,P
RED SNOW [DOC SAVAGE #38]  ROBESON, KENNETH  C/P,@P
RED SPIDER, THE [DOC SAVAGE #95]  ROBESON, KENNETH  @P
RED STAR, THE  COVER, ARTHUR BYRON  LP
RED STAR: THE FIRST SOVIET UTOPIA  BOGDANOV, ALEXANDER  LP
RED STARS  MCGUIRE, PATRICK L.  R/REF-H
RED STORM  STORM, BRIAN  SR-P
RED STORM RISING  CLANCY, TOM  C/P:2,P
RED SUN OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA

RED TALONS [WORLD OF DARKNESS WEREWOLF 3]/ FIANNA [WORLD OF DARKNESS WEREWOLF 3] BOULLE, PHILIPPE/GRIFFIN, ERIC

PA

RED TAPE WAR, THE CHALKER, JACK L./RESNICK, MIKE/

EFFINGER, GEORGE ALEC

C/P,P,H,SFWA

RED TERRORS, THE [DOC SAVAGE #83]  ROBESON, KENNETH  C/P,@P
RED THIRST [WARHAMMER]  PRINGLE, DAVID  LP
RED THUNDER [RED THUNDER 1]  VARLEY, JOHN  H
RED TIDE  CHAPMAN, D. D./TARZAN, DELORIS LEHMAN P
RED TWILIGHT/WORLD’S END  STARMONT HOUSE, INC.  SR-LPA
RED UNICORN [TANAQUIL 3]  LEE, TANITH  H,P
RED WORLD OF POLARIS  SMITH, CLARK ASHTON  H
RED WYVERN, THE [DEVERRY 9/DRAGON MAGE 1]  KERR, KATHARINE  C/P,LP
REDBEARD  RESNICK, MIKE  C/P:2,P
REDEEMING THE LOST  KERNER, ELIZABETH  P
REDEMOLISHED  BESTER, ALFRED  P
REDEMPTION [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/LINAWEAVER, BRAD  H
REDEMPTION [@ANGEL]  ODOM, MEL  P
REDEMPTION [@DRAGONLANCE DHAMON #3]  RABE, JEAN  H
REDEMPTION ARK [REVELATION SPACE 2]  REYNOLDS, ALASTAIR  H:2
REDEMPTION OF ALTHALUS, THE  EDDINGS, DAVID/EDDINGS, LEIGH  H:2,P
REDEMPTION OF LIGHT [GAMANT 3]  O’NEAL, KATHLEEN M.  P

REDISCOVERY [DARKOVER] BRADLEY, MARION ZIMMER/ LACKEY, MERCEDES

H:2,P

REDISCOVERY OF MAN, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  SR-H
REDLINE THE STARS  NORTON, ANDRE/GRIFFIN, P. M.  H
REDLINERS  DRAKE, DAVID  C/P:2,P
REDMAGIC [CALINDOR 2]  KILIAN, CRAWFORD  P
REDROBE  GRIMWOOD, JON COURTENAY  H
REDSHIFT  SARRANTONIO, AL  HA
REDSHIFT RENDEZVOUS  STITH, JOHN E.  C/P,P
REDUCTION IN ARMS  PURDOM, TOM  C/P:2,P
REDWALL [REDWALL]  JACQUES, BRIAN  C/P,P
REDWARD EDWARD PAPERS, THE  DAVIDSON, AVRAM  H,SFWA
REDWORK  BEDARD, MICHAEL  SFWA-P
REDWORLD  HARNESS, CHARLES L.  C/P,P
REEFS [@JOURNEYS OF MCGILL FEIGHAN #2]  O’DONNELL, KEVIN JR.  C/P,P
REEFS OF EARTH, THE  LAFFERTY, R. A.  C/P,P
REEFS OF SPACE, THE  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 292
TITLE  AUTHOR  CODE
REEFSONG  SEVERANCE, CAROL  C/P,P
REEL  JANIFER, LAURENCE M.  H:2,SFWA
REENCUENTRO  KLEIN, GERARD  FP
REFERENCE GUIDE TO FANTASTIC FILMS 1 (A–F)  LEE, WALT  R/IND
REFERENCE GUIDE TO FANTASTIC FILMS 2 (G–O)  LEE, WALT  R/IND
REFERENCE GUIDE TO FANTASTIC FILMS 3 (P–Z)  LEE, WALT  R/IND
REFLECTIONS IN A MIRAGE  DAVENTRY, LEONARD  C/P,P
REFLECTIONS IN A MIRAGE/TICKING IS IN YOUR HEAD, THE  DAVENTRY, LEONARD  H
REFLECTIONS OF A. E. VAN VOGT  VAN VOGT, A. E.  R/REF-P
REFLEX [JUMPER 2]  GOULD, STEVEN C.  H
REFORMER, THE [GENERAL]  STIRLING, S. M./DRAKE, DAVID  H,P
REFUGE [ROBOT CITY 5]  CHILSON, ROBERT  C/P:2,SFWA-P,P
REFUGE FOR TONIGHT  WILLIAMS, ROBERT MOORE  B
REFUGEE [@BIO OF A SPACE TYRANT #1]  ANTHONY, PIERS  C/P:2,P
REFUGEE [@STAR WARS NEW JEDI ORDER FORCE HERETIC #2]  WILLIAMS, SEAN/DIX, SHANE  P
REFUTATION OF DARWINISM, THE  O’NEILL, T. WARREN  R/REF-H
REGAN’S PLANET  SILVERBERG, ROBERT  C/P,P
REGENERATION [@SPECIES IMPERATIVE #3]  CZERNEDA, JULIE E.  H
REGENERATION [@X-FILES]  OWENS, EVERETT  H
REGENESIS  ECKLAR, JULIA  C/P:2,P
REGIMENT, THE [REGIMENT 1]  DALMAS, JOHN  C/P:2,P
REGIMENT OF WOMEN  BERGER, THOMAS  P
REGIMENTS OF NIGHT, THE  BALL, BRIAN N.  C/P:2,P
REGIMENT’S WAR, THE [REGIMENT 4]  DALMAS, JOHN  C/P,SFWA-P,P
REGINA’S SONG  EDDINGS, DAVID/EDDINGS, LEIGH  H
REGRESO AL PASADO  KLEIN, GERARD  FP
REGULATORS, THE  BACHMAN, RICHARD  H
REGULATORS, THE  KING, STEPHEN  P
REHEARSAL FOR A RENAISSANCE  CLARK, DOUGLAS W.  C/P,P
REHEARSALS FOR OBLIVION: ACT I  WORTHY, PETER A.  LPA

REIGN OF FIRE [LEAR’S DAUGHTERS 2] KELLOGG, MARJORIE BRADLEY/

ROSSOW, WILLIAM B.

P

REIGN OF ISTAR, THE [@DRAGONLANCE TALES II #1]  TSR  C/PA,PA
REIGN OF SHADOWS [REIGN OF SHADOWS 1]  CHESTER, DEBORAH  C/P,P
REIGN OF TERROR, THE [DOCTOR WHO]  MARTER, IAN  SR-P
REIGN OF TERROR  PARRY, MICHEL  PA
REIGN OF THE BROWN MAGICIAN, THE [THREE WORLDS 3]  WATT-EVANS, LAWRENCE  P
REIGN OF WIZARDRY, THE  WILLIAMSON, JACK  C/P:2,P
REINCARNATION IN VENICE  EHRLICH, MAX  P
REINCARNATION OF PETER PROUD, THE  EHRLICH, MAX  H
REINCARNATION OF REECE ERIKSON, THE  OBSTFELD, RAYMOND  P,SFWA-P
REINCARNATION — THE WHOLE STARTLING STORY  MILLER, R. DEWITT  [C/RAN-P]
REINDEER MOON  THOMAS, ELIZABETH MARSHALL  H,P
REINDEER PEOPLE, THE  LINDHOLM, MEGAN  C/P,SFWA-P,P
REISE NACH ANTARIA, DIE  VON KORFF, FREYA GRAFIN  FH
RELATIVE DEMENTIAS [DOCTOR WHO PAST DOCTORS]  MICHALOWSKI, MARK  P
RELATIVE MAGIC  HUFF, TANYA  H,LP
RELATIVES  EFFINGER, GEORGE ALEC  H:2,SFWA,P
RELATIVITY  SAWYER, ROBERT J.  H
RELATIVITY FOR THE LAYMAN  COLEMAN, JAMES A.  [C/RAN-P]
RELIC [RELIC 1]  PRESTON, DOUGLAS/CHILD, LINCOLN  P
RELIC OF EMPIRE [@FORBIDDEN BORDERS #2]  GEAR, W. MICHAEL  C/P:2,SFWA-P,P
RELIC OF THE DAWN [@EXALTED #2]  WILSON, DAVID NIALL  P
RELICS [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P,P
RELICS AND OMENS [@DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  PA
RELIGION, THE  CONDE, NICHOLAS  P
RELIQUARY [RELIC 2]  PRESTON, DOUGLAS/CHILD, LINCOLN  H
RELIQUARY [STARGATE ATLANTIS #2]  WELLS, MARTHA  P
RELUCTANT GOD, THE  SERVICE, PAMELA F.  C/P,P
RELUCTANT KING, THE [RELUCTANT KING 1,2,3]  DE CAMP, L. SPRAGUE  H:2
RELUCTANT SHAMAN AND OTHER FANTASTIC TALES, THE  DE CAMP, L. SPRAGUE  C/P:2,P
RELUCTANT SORCERER, THE [MARVIN BREWSTER 1]  HAWKE, SIMON  C/P,P
RELUCTANT SPY, THE  BLACKBURN, JOHN  P
RELUCTANT SWORDSMAN, THE [SEVENTH SWORD 1]  DUNCAN, DAVE  C/P:2,P
RELUCTANT VOYAGERS  VONARBURG, ELISABETH  P
RELUCTANT WIZARD, THE  NEWELL, NEIL K.  P
REMAKE  WILLIS, CONNIE  H
REMAKING HISTORY  ROBINSON, KIM STANLEY  H,LP
REMAKING OF SIGMUND FREUD, THE  MALZBERG, BARRY N.  C/P,P
REMARKABLE BIRTH OF PLANET EARTH, THE  MORRIS, HENRY M.  [C/RAN-P]
REMARKABLE EXPLOITS OF LANCELOT BIGGS: SPACEMAN, THE  BOND, NELSON  H,LP
REMARKABLES, THE  REED, ROBERT  C/P,P
REMEMBER GETTYSBURG! [REMEMBER! 2]  RANDLE, KEVIN/CORNETT, ROBERT  P
REMEMBER THE ALAMO! [REMEMBER! 1]  RANDLE, KEVIN/CORNETT, ROBERT  P

The MITSFS Titledex  September 24, 2007  Page 293
TITLE  AUTHOR  CODE
REMEMBER THE LITTLE BIGHORN! [REMEMBER! 3]  RANDLE, KEVIN/CORNETT, ROBERT  P
REMEMBRANCE FOR KEDRIGERN, A [KEDRIGERN]  MORRESSY, JOHN  P
REMEMBRANCE OF THE DALEKS [DOCTOR WHO]  AARONOVITCH, BEN  C/P,P
REMEMORY  BETANCOURT, JOHN GREGORY  C/P,P
REMIX  GRIMWOOD, JON COURTENAY  LP
REMNANT [@STAR WARS NEW JEDI ORDER FORCE HERETIC #1]  WILLIAMS, SEAN/DIX, SHANE  P
REMNANT POPULATION  MOON, ELIZABETH  C/P,P,H
REMO  MURPHY, WARREN/SAPIR, RICHARD BEN  P
REMSCELA  FROST, GREGORY  C/P,P
RENAISSANCE  JONES, RAYMOND F.  H
RENAISSANCE  VAN VOGT, A. E.  C/P:2,P
RENAISSANCE FAIRE  NORTON, ANDRE/RABE, JEAN  PA
RENDEZVOUS [MARATHON 2]  SMITH, DAVID ALEXANDER  P
RENDEZVOUS ON A LOST WORLD  CHANDLER, A. BERTRAM  DF117
RENDEZVOUS WITH RAMA [RAMA 1]  CLARKE, ARTHUR C.  H:2,P
RENDING OF FALCONS, A [@MECHWARRIOR DARK AGE/BATTLETECH]  MILAN, VICTOR  P
RENEGADE [STAR TREK]  DEWEESE, GENE  C/P:2,P
RENEGADE [@MARQ’SSAN #2]  DUCHAMP, L. TIMMEL  LP
RENEGADE [ROBOTS AND ALIENS 2]  SCOTTEN, CORDELL  C/P:3,SFWA-P,P
RENEGADE DRAGON, THE [DRAGON NIMBUS HISTORY 3]  RADFORD, IRENE  P
RENEGADE OF CALLISTO [JANDAR 8]  CARTER, LIN  P
RENEGADE OF KREGEN [DRAY PRESCOT 13]  AKERS, ALAN BURT  P
RENEGADE STAR, THE  STATTEN, VARGO  SR-P
RENEGADE’S HONOR  KEITH, WILLIAM H. JR.  P
RENEGADES OF GOR [GOR 23]  NORMAN, JOHN  P
RENEGADES OF PERN, THE [PERN]  MCCAFFREY, ANNE  H:2,P
RENEGADES OF THE FUTURE [@PERRY RHODAN #65]  ACE  C/PA,PA
RENEGADES OF TIME  JONES, RAYMOND F.  SR-P,C/P:2
RENSIME [SIME/GEN 6]  LICHTENBERG, JACQUELINE  C/P,SFWA,P,H
RENUNCIATES OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
REPAIRMEN OF CYCLOPS, THE  BRUNNER, JOHN  C/P:2,DM115

REPENTANT, THE THOMSEN, BRIAN M./

GREENBERG, MARTIN H.

PA

REPLAY  GRIMWOOD, KEN  C/P,SFWA,P,H
REPLICA  BOWKER, RICHARD  C/P,P
REPLICA  SABEN, LIONEL  P
REPLICANT NIGHT [@BLADE RUNNER]  JETER, K. W.  P
REPLICATOR RUN  REY, RAINER  P
REPORT FROM GROUP 17, A  O’BRIEN, ROBERT C.  H,P

REPORT OF THE PRESIDENTIAL COMMISSION ON THE SPACE SHUTTLE CHALLENGER ACCIDENT PRESIDENTIAL COMMISSION

R/REF-H

REPORT ON PLANET THREE AND OTHER SPECULATIONS  CLARKE, ARTHUR C.  [C/RAN-H,C/RAN-P:2]
REPORT ON PROBABILITY A  ALDISS, BRIAN W.  C/P,P
REPORT ON UNIDENTIFIED FLYING OBJECTS, THE  RUPPELT, EDWARD J.  [C/RAN-H:3,C/RAN-P]
REPORT TO THE MEN’S CLUB  EMSHWILLER, CAROL  LP
REPOSSESSED  JENKINS, A. M.  SR-LP
REPRIEVE FROM PARADISE  ELLIOTT, H. CHANDLER  H
REPRISAL, THE [@SURVIVALIST #11]  AHERN, JERRY  P
REPRISAL [ADVERSARY 5]  WILSON, F. PAUL  P
REPRODUCTIVE SYSTEM, THE  SLADEK, JOHN  LP
REPUBLIC [@STAR TREK MY BROTHER’S KEEPER #1]  FRIEDMAN, MICHAEL JAN  P
REPUBLIC AND EMPIRE [@IMPERIAL STARS #2]  POURNELLE, JERRY  C/PA,PA,SFWA-P
REQUIEM [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN/RYAN, KEVIN  C/P:3,P
REQUIEM  HEINLEIN, ROBERT A.  C/P,P,H
REQUIEM  JOYCE, GRAHAM  LP
REQUIEM FOR A RULER OF WORLDS [HOBART FLOYT 1]  DALEY, BRIAN  C/P:2,P
REQUIEM FOR ANTHI [ANTHI]  BLAKENEY, JAY D.  P
REQUIEM FOR ASTOUNDING, A  ROGERS, ALVA  R/REF-H,C/REF-H
REQUIEM FOR BOONE [@EARTH FINAL CONFLICT]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,LP
REQUIEM FOR THE CONQUEROR [@FORBIDDEN BORDERS #1]  GEAR, W. MICHAEL  C/P:2,SFWA-P:2,P
REQUIEM FOR THE SUN [SYMPHONY OF AGES 4]  HAYDON, ELIZABETH  H
REQUIEM FOR THE SUN (PROOF) [SYMPHONY OF AGES 4]  HAYDON, ELIZABETH  SR-LP
REQUIEM OF STARS [SONGS OF THE STELLAR WIND 1]  HICKMAN, TRACY  P
RESCUE, THE [DOCTOR WHO]  MARTER, IAN  SR-P,C/P
RESCUE MISSION, THE [@NEW TOM SWIFT #6]  APPLETON, VICTOR  C/P,LP
RESCUE OF ATHENA ONE, THE [@SIX MILLION DOLLAR MAN #5]  JAHN, MICHAEL  C/P,P
RESCUE RUN [PERN]  MCCAFFREY, ANNE  H:2

RESCUED FROM PARADISE [ROCHEWORLD 5] FORWARD, ROBERT L./

FULLER, JULIE FORWARD

C/P,P

RESEARCH GUIDE TO SCIENCE FICTION STUDIES, A TYMN, MARSHALL B./SCHLOBIN, ROGER C./

CURREY, L. W.

R/IND

RESEARCH MAGNIFICENT, THE  WELLS, H. G.  H:2
RESENTING THE HERO [SHIELD AND SOURCE 1]  MOORE, MOIRA J.  P
RESISTANCE [STARGATE]  MCCAY, BILL  C/P,P

The MITSFS Titledex  September 24, 2007  Page 294
TITLE  AUTHOR  CODE
RESISTANCE, THE [@FEY 4]  RUSCH, KRISTINE KATHRYN  C/P,P
RESOLUTION [NULAPEIRON 3]  MEANEY, JOHN  H
RESONANCE  DOLLEY, CHRIS  H
RESONANCE OF BLOOD, THE  VARDEMAN, ROBERT E.  P
REST IN AGONY  FAIRMAN, PAUL W.  P
REST IN AGONY  JORGENSEN, IVAR  P
REST IN AGONY  YOUNGER, JACK  P
REST MUST DIE, THE  FOSTER, RICHARD  P
REST OF THE ROBOTS, THE  ASIMOV, ISAAC  H:2
RESTAURANT AT THE END OF THE UNIVERSE, THE [HITCHHIKERS 2]  ADAMS, DOUGLAS  C/P,P,H

RESTAURANT AT THE END OF THE UNIVERSE, THE (1 CASSETTE)

[HITCHHIKERS 2] ADAMS, DOUGLAS

SR-MM

RESTORATION [SEYONNE 3]  BERG, CAROL  P
RESTORATION [NOEL KEDRAN 1]  DALTON, SEAN  P
RESTOREE  MCCAFFREY, ANNE  C/P:2,P
RESULTS OF THE 1ST US MANNED ORBITAL SPACE FLIGHT  NASA  [C/RAN-H]
RESULTS OF THE 2ND US MANNED SUBORBITAL SPACE FLIGHT  NASA  [C/RAN-H]
RESUME WITH MONSTERS  SPENCER, WILLIAM BROWNING  P
RESURGENCE [HERITAGE UNIVERSE 5]  SHEFFIELD, CHARLES  H:2
RESURGENT DUST  GARNER, ROLF  SR-P
RESURRECTED MAN, THE  WILLIAMS, SEAN  H
RESURRECTING RAVANA [BUFFY]  GARTON, RAY  C/P,P
RESURRECTION [@WARTORN 1]  ASPRIN, ROBERT LYNN/DEL CARLO, ERIC  P
RESURRECTION [WASTEWORLD 2]  BARTON, JAMES  P
RESURRECTION  DAYTON, ARWEN ELYS  P
RESURRECTION, THE [WAR OF THE WORLDS 1]  DILLARD, J. M.  C/P,P
RESURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #6]  KEMP, PAUL S.  H
RESURRECTION  KERR, KATHARINE  C/P,P
RESURRECTION [@SPACE: 1999]  LATHAM, WILLIAM  P
RESURRECTION [VAMPIRE LEGACY]  TAYLOR, KAREN E.  C/P,P
RESURRECTION 2027  ECCARIUS, J. G.  P
RESURRECTION CASKET, THE [DOCTOR WHO TENTH DOCTOR]  RICHARDS, JUSTIN  H
RESURRECTION DAY  DUBOIS, BRENDAN  H,P
RESURRECTION DAY [DOC SAVAGE #36]  ROBESON, KENNETH  C/P:2,@P
RESURRECTION DAYS  TUCKER, WILSON  C/P:2,P
RESURRECTION, INC.  ANDERSON, KEVIN J.  C/P,P
RESURRECTION MAN [RESURRECTION MAN 1]  STEWART, SEAN  C/P,LP
RESURRECTION MAN’S LEGACY, THE  BAILEY, DALE  H
RESURRECTION OF ROGER DIMENT, THE  MASON, DOUGLAS R.  C/P,P
RESURRECTIONIST, THE  WOLF, GARY K.  H,SFWA
RETALIATION [STARGATE]  MCCAY, BILL  C/P,P
RETALIATION  ROHMER, RICHARD  H,P
RETIEF! [RETIEF]  LAUMER, KEITH  P
RETIEF: AMBASSADOR TO SPACE [RETIEF]  LAUMER, KEITH  H,P
RETIEF AND THE PANGALACTIC PAGEANT OF PULCHRITUDE [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF AND THE RASCALS [RETIEF]  LAUMER, KEITH  P,SFWA-P
RETIEF AND THE WARLORDS [RETIEF]  LAUMER, KEITH  C/P,P,H
RETIEF AT LARGE [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: DIPLOMAT AT ARMS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: EMISSARY TO THE STARS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF: EMISSARY TO THE STARS (EXPANDED COLLECTION) [RETIEF]  LAUMER, KEITH  P

RETIEF: EMISSARY TO THE STARS (MORE EXPANDED COLLECTION)

[RETIEF] LAUMER, KEITH

C/P,P

RETIEF: ENVOY TO NEW WORLDS [RETIEF]  LAUMER, KEITH  P
RETIEF IN THE RUINS [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF OF THE CDT [RETIEF]  LAUMER, KEITH  C/P,P,H,SR-H,SFWA
RETIEF TO THE RESCUE [RETIEF]  LAUMER, KEITH  C/P,P,H
RETIEF UNBOUND [RETIEF]  LAUMER, KEITH  C/P:2,P
RETIEF’S PEACE [RETIEF]  KEITH, WILLIAM H. JR.  P
RETIEF’S RANSOM [RETIEF]  LAUMER, KEITH  C/P,SR-H,P,H
RETIEF’S WAR [RETIEF]  LAUMER, KEITH  C/P:2,P
RETREAD SHOP  KING, T. JACKSON  P
RETRIBUTION [STARGATE]  MCCAY, BILL  P
RETRIBUTION [@WARHAMMER VON CARSTEIN 3]  SAVILE, STEVEN  P
RETRO LIVES  GRIMES, LEE  P,SFWA-P
RETURN, THE  ALDRIN, BUZZ/BARNES, JOHN  H:2
RETURN, THE  DE LA MARE, WALTER J.  LP
RETURN, THE  HAIBLUM, ISIDORE  P
RETURN, THE [AMERICA 2040 4]  INNES, EVAN  P
RETURN, THE [@LAST ON EARTH #3]  KAYE, MARILYN  P
RETURN, THE  MAYNARD, RICHARD  P
RETURN [@FIVE WORLDS #3]  SARRANTONIO, AL  P

The MITSFS Titledex  September 24, 2007  Page 295
TITLE  AUTHOR  CODE

RETURN, THE [STAR TREK ODYSSEY 2] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

P

RETURN, THE  TORRO, PEL  SR-P
RETURN, THE [@DUMAREST #32]  TUBB, E. C.  LP
RETURN ENGAGEMENT [SETTLING ACCOUNTS 1]  TURTLEDOVE, HARRY  H
RETURN FIRE [@SAND WARS #5]  INGRID, CHARLES  C/P,P
RETURN FROM LUNA  HALACY, D. S.  H
RETURN FROM THE DEAD [@STARSHIP ORPHEUS #1]  JADE, SYMON  P
RETURN FROM THE STARS  LEM, STANISLAW  H:2,P
RETURN FROM THE VOID [@PERRY RHODAN #51]  ACE  PA,C/PA
RETURN FROM WITCH MOUNTAIN  KEY, ALEXANDER  C/P,P
RETURN OF CAPTAIN CONQUER, THE  GILDEN, MEL  SFWA
RETURN OF COUNT ELECTRIC, THE  SPENCER, WILLIAM BROWNING  P
RETURN OF DR. FU-MANCHU, THE  ROHMER, SAX  C/P,P
RETURN OF JONGOR, THE  WILLIAMS, ROBERT MOORE  C/P,P
RETURN OF KAVIN, THE  MASON, DAVID  C/P:2,P
RETURN OF MERLIN, THE  CHOPRA, DEPAK  H
RETURN OF NATHAN BRAZIL, THE [WELL WORLD 4]  CHALKER, JACK L.  C/P:2,P
RETURN OF RETIEF, THE [RETIEF]  LAUMER, KEITH  C/P,P
RETURN OF SANTIAGO, THE [SANTIAGO 2]  RESNICK, MIKE  H
RETURN OF SHERLOCK HOLMES, THE  DOYLE, ARTHUR CONAN  H
RETURN OF SKULL-FACE, THE  HOWARD, ROBERT E./LUPOFF, RICHARD A.  SR-H
RETURN OF SUMURU  ROHMER, SAX  P
RETURN OF SWAMP THING, THE  DAVID, PETER  P
RETURN OF TARZAN, THE [TARZAN #2]  BURROUGHS, EDGAR RICE  H:2,@P
RETURN OF THE BREAKNECK BOYS [FADING WORLDS 2]  GRAVEL, GEARY  P
RETURN OF THE DAMNED [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
RETURN OF THE DARK SIDE [@STAR WARS LAST OF THE JEDI #6]  WATSON, JUDE  LP
RETURN OF THE DINOSAURS  RESNICK, MIKE/GREENBERG, MARTIN H.  C/PA,PA
RETURN OF THE EAGLE  DORER, NANCY/DORER, FRANCES  P
RETURN OF THE EMPEROR, THE [STEN 6]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
RETURN OF THE EXILE [@DRAGONLANCE LINSHA #3]  HERBERT, MARY H.  P
RETURN OF THE GUARDIAN-KING [GUARDIAN-KING 4]  HANCOCK, KAREN  LP
RETURN OF THE JEDI [STAR WARS MOVIE 6]  KAHN, JAMES  C/P,P,H
RETURN OF THE JEDI PORTFOLIO  MCQUARRIE, RALPH  A
RETURN OF THE JEDI SOUVENIR BOOK  LUCASFILM LTD.  A
RETURN OF THE KING, THE [@LORD OF THE RINGS #3]  TOLKIEN, J. R. R.  SR-H,H:2,P
RETURN OF THE KING, THE (ACE) [@LORD OF THE RINGS #3]  TOLKIEN, J. R. R.  SR-P
RETURN OF THE LIVING DAD [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
RETURN OF THE SHADOW  GIBSON, WALTER B.  P
RETURN OF THE SHADOW, THE [HISTORY OF MIDDLE-EARTH 6]  TOLKIEN, J. R. R.  H

RETURN OF THE SORCERESS [@DRAGONLANCE NEW ADVENTURES #4/ DRAGONLANCE SPELLBINDER QUARTET 4] WAGGONER, TIM

LP

RETURN OF THE STRANGER  ROBERTS, DOROTHY JAMES  H
RETURN OF THE TIME MACHINE, THE  FRIEDELL, EGON  C/P:2,P
RETURN OF ZEUS  MULLER, JOHN E.  SR-P
RETURN TO AVALON  ROBERSON, JENNIFER  C/PA,PA
RETURN TO BROKEN CROSSING  HOFFMAN, LEE  P
RETURN TO BROOKMERE  ESTES, ROSE  [C/RAN-P]
RETURN TO CAMEREIN [SECOND COMMONWEALTH 3]  SHELLEY, RICK  P
RETURN TO CENTAUR [XANTH]  PINI, RICHARD  CX
RETURN TO CHAOS [@BUFFY]  GARDNER, CRAIG SHAW  C/P:2,P
RETURN TO DOOMSTAR [DOOM STAR]  MEYERS, RICHARD S.  C/P,P
RETURN TO EARTH  BERRY, BRYAN  SR-P
RETURN TO EARTH  HOOVER, H. M.  C/P,P
RETURN TO EDDARTA [GANDALARA 6]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
RETURN TO EDEN [EDEN 3]  HARRISON, HARRY  H:2,P
RETURN TO ELYSIUM  GRANT, JOAN  P
RETURN TO FANGLITH  DALMAS, JOHN  C/P:2,P
RETURN TO KALDAK [@BLADE #36]  LORD, JEFFREY  P
RETURN TO LOVECRAFT COUNTRY  ANIOLOWSKI, SCOTT DAVID  LPA
RETURN TO MARS [JAMIE WATERMAN 2]  BOVA, BEN  C/P,P,H
RETURN TO ORD MANTELL [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
RETURN TO OTHERNESS  KUTTNER, HENRY  C/P:2,P
RETURN TO OZ [OZ]  VINGE, JOAN D.  C/P,P
RETURN TO QUAG KEEP [QUAG KEEP 2]  NORTON, ANDRE/RABE, JEAN  H:2

RETURN TO ROCHEWORLD [ROCHEWORLD 2] FORWARD, ROBERT L./

FULLER, JULIE FORWARD

C/P:2,SFWA-P,P

RETURN TO SAC PRAIRIE  DERLETH, AUGUST  LP

RETURN TO THE FRACTURED PLANET [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] STONE, DAVE

P

RETURN TO THE STARS  HAMILTON, EDMOND  C/P:2,P
RETURN TO THE TWILIGHT ZONE  SERLING, CAROL  PA

The MITSFS Titledex  September 24, 2007  Page 296
TITLE  AUTHOR  CODE
RETURN TO THE WHORL [BOOK OF THE SHORT SUN 3]  WOLFE, GENE  LP
RETURN TO TOMORROW  HUBBARD, L. RON  C/P:2,P
RETURN TO WAR [STAR COMMANDOS 6]  GRIFFIN, P. M.  C/P,P
RETURN TO YESTERDAY [STACY GORDON 2]  SHIPLETT, JUNE LUND  P
RETURNING CREATION [SILISTRA (REVISED) 1]  MORRIS, JANET  C/P,P
REUNION [@WATER #2]  DALKEY, KARA  LP
REUNION [HUMANX FLINX]  FOSTER, ALAN DEAN  H
REUNION [STAR TREK TNG]  FRIEDMAN, MICHAEL JAN  C/P:2,P
REUNION  GRIBBIN, JOHN/CHOWN, MARCUS  H
REUNION [@STAR WARS NEW JEDI ORDER FORCE HERETIC #3]  WILLIAMS, SEAN/DIX, SHANE  P
REUNION ON NEVEREND  STITH, JOHN E.  P
REVELATION [SEYONNE 2]  BERG, CAROL  P
REVELATION  HARBINSON, W. A.  P
REVELATION [@MASS EFFECT]  KARPYSHYN, DREW  P
REVELATION SPACE [REVELATION SPACE 1]  REYNOLDS, ALASTAIR  C/P,P,H
REVELATIONS  MALZBERG, BARRY N.  P
REVELATIONS IN BLACK  JACOBI, CARL  P
REVELATIONS OF DOCTOR MODESTO, THE  HARRINGTON, ALAN  P
REVELRY IN HELL  WEINBERG, ROBERT  SR-LPA
REVENANT [@INDIGO #7]  COOPER, LOUISE  C/P,P
REVENANT, THE  LOBEL, BRANA  P
REVENANT [BUFFY]  ODOM, MEL  C/P,P
REVENANT, THE  TEM, MELANIE  P
REVENANTS, THE  TEPPER, SHERI S.  C/P:2,P
REVENGE OF INCREASE SEWALL, THE  GRAAT, HEINRICH  P
REVENGE OF TAURUS, THE [@HORRORSCOPE #2]  LORY, ROBERT  P
REVENGE OF THE DAMNED [STEN 5]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
REVENGE OF THE FLUFFY BUNNIES [CINEVERSE 3]  GARDNER, CRAIG SHAW  C/P,SFWA-P:2,P
REVENGE OF THE HORSECLANS [HORSECLANS 3]  ADAMS, ROBERT  C/P:2,P
REVENGE OF THE MANITOU  MASTERTON, GRAHAM  P
REVENGE OF THE RAINBOW DRAGONS  ESTES, ROSE  [C/RAN-P:2]
REVENGE OF THE ROSE, THE [ELRIC]  MOORCOCK, MICHAEL  H:2,SFWA:2,P
REVENGE OF THE SENIOR CITIZENS ** PLUS, THE  REED, KIT  H
REVENGE OF THE SITH [@STAR WARS]  STOVER, MATTHEW WOODRING  H:2
REVENGE OF THE VALKYRIE [SONG OF THE DWARVES 2]  GUNNARSSON, THORARINN  C/P,P
REVEREND MAMA  COOPER, PARLEY J.  P
REVES DE SABLE, CHATEAUX DE SANG  FREMION, YVES  FP
REVISED H. P. LOVECRAFT BIBLIOGRAPHY, THE  OWINGS, MARK/CHALKER, JACK L.  R/IND
REVISIONS  CZERNEDA, JULIE E./SZPINDEL, ISAAC  C/PA:2,PA
REVOLT  MCMASTERS, WILLIAM H.  H
REVOLT!  PENDLETON, DON  SR-P,C/P
REVOLT AND REBIRTH [@CONCORDAT #3]  SWYCAFFER, JEFFERSON P.  P
REVOLT IN 2100  HEINLEIN, ROBERT A.  C/P,P,H
REVOLT IN 2100/METHUSELAH’S CHILDREN  HEINLEIN, ROBERT A.  P
REVOLT OF THE ANGELS, THE  FRANCE, ANATOLE  H
REVOLT OF THE DWARVES  ESTES, ROSE  [C/RAN-P]
REVOLT OF THE GALAXY [FAMILY D’ALEMBERT #10]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
REVOLT OF THE HUMANS  BURKE, JONATHAN  SR-P
REVOLT OF THE MICRONAUTS [MICRONAUTS 3]  WILLIAMS, GORDON  P
REVOLT OF THE TRIFFIDS  WYNDHAM, JOHN  SR-P,C/P
REVOLT ON ALPHA C  SILVERBERG, ROBERT  C/P,P,H
REVOLT ON JUPITER  MARTIN, JOHN  P
REVOLT ON MAJIPOOR  COSTELLO, MATTHEW J.  P,SFWA-P
REVOLT ON VENUS, THE [TOM CORBETT 5]  ROCKWELL, CAREY  H:2
REVOLTA NA LUA (VOLUME 1 OF 2)  HEINLEIN, ROBERT A.  FP
REVOLUTION FROM ROSINANTE, THE [ROSINANTE 1]  GILLILAND, ALEXIS A.  P
REVOLUTION MAN [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
REVOLUTIONARY [@HOSTILE TAKEOVER #3]  SWANN, S. ANDREW  C/P,P
REVOLUTION’S SHORE [@HIGHROAD #2]  RASMUSSEN, ALIS A.  P
REVOLVING BOY, THE  FRIEDBERG, GERTRUDE  C/P,P,H
REWARD FOR RETIEF [RETIEF]  LAUMER, KEITH  C/P:2,SFWA-P,P
REWARDS AND FAIRIES [@WORKS OF KIPLING #25]  KIPLING, RUDYARD  SR-H,H
REWIND  ENGLAND, TERRY  C/P,P
REX COLE JR. AND THE GRINNING GHOST  CHAPMAN, GORDON  H
RHAPSODY [SYMPHONY OF AGES 1]  HAYDON, ELIZABETH  H,P
RHAPSODY FOR A UNICORN  CAPPELLI, OSCAR  LP
RHAPSODY IN BLACK [GRAINGER 2]  STABLEFORD, BRIAN M.  C/P:2,P
RHIALTO THE MARVELLOUS [DYING EARTH 4]  VANCE, JACK  H,P

RHIALTO THE MARVELLOUS (VIE) [@VANCE INTEGRAL EDITION #34/ DYING EARTH 4] VANCE, JACK

SR-H

RHIANA  HAUF, MICHELE  LP
RHINEGOLD  GRUNDY, STEPHAN  H,P

RHINOCEROS WHO QUOTED NIETZCHE AND OTHER ODD ACQUAINTANCES, THE BEAGLE, PETER S.

LP

The MITSFS Titledex  September 24, 2007  Page 297
TITLE  AUTHOR  CODE
RHYTHM RIDES THE ROCKET  OLSEN, BOB  B
RIBOFUNK  DI FILIPPO, PAUL  P
RICHTER 10  CLARKE, ARTHUR C./MCQUAY, MIKE  C/P:2,P
RIDDLE AND THE RUNE, THE [TALES OF GOM]  CHETWIN, GRACE  P
RIDDLE IN THE STONE [@KNIGHTS OF THE SILVER DRAGON #2]  SOESBEE, REE  LP
RIDDLE-MASTER OF HED, THE  MCKILLIP, PATRICIA A.  C/P,P,H
RIDDLE OF STARS  MCKILLIP, PATRICIA A.  H:2
RIDDLE OF THE GRIFFON [@CRIMSON CRYSTAL 1]  LAWSON, SUSAN  P
RIDDLE OF THE SEVEN REALMS  HARDY, LYNDON  C/P:2,P
RIDDLE OF THE WREN, THE  DE LINT, CHARLES  C/P,P
RIDDLED MAN, THE [@BLOOD OF THE LAMB #3]  ROGERS, MARK E.  P
RIDDLED NIGHT, THE [EVERIEN 2]  LEITH, VALERY  C/P,LP
RIDDLEY WALKER  HOBAN, RUSSELL  H:2,LP
RIDE THE GRAY PLANET  SAVAGE, BLAKE  C/P,P
RIDE THE GREEN DRAGON  NORTON, ANDRE/MILLER, PHYLLIS  H
RIDE THE LIGHTNING [ALO NUDGER]  LUTZ, JOHN  SFWA-P
RIDE THE NIGHTMARE  MATHESON, RICHARD  P
RIDER, THE  BURROUGHS, EDGAR RICE  C/P:2,P
RIDER AT THE GATE  CHERRYH, C. J.  H:2,P
RIDERS IN THE SKY  GRANT, CHARLES L.  H
RIDERS OF LEVIATHAN [TYPHON 2]  ANZETTI, TONI  P
RIDERS OF THE DEAD [@WARHAMMER]  ABNETT, DAN  H
RIDERS OF THE DRAGON  DANE, CHRISTOPHER  P
RIDERS OF THE PURPLE WAGE  FARMER, PHILIP JOSE  C/P,P
RIDERS OF THE SIDHE, THE [SIDHE 1]  FLINT, KENNETH C.  C/P:2,P
RIDERS OF THE WINDS [CHANGEWINDS 2]  CHALKER, JACK L.  C/P:2,SFWA-P,P
RIDERS TO THE STARS  SIODMAK, CURT  C/P:2,P
RIDGE, THE  CANTRELL, LISA W.  SFWA-P
RIDING THE ROCK [XEELEE]  BAXTER, STEPHEN  SR-LP
RIDING THE TORCH  SPINRAD, NORMAN  TD23,BS1
RIEN QU’UNE ETOILE  LIMAT, MAURICE  FP
RIFKIND’S CHALLENGE  ABBEY, LYNN  H
RIFT, THE [STAR TREK]  DAVID, PETER  C/P:3,P
RIFT  KENYON, KAY  P
RIFT, THE  WILLIAMS, WALTER JON  P
RIFTER’S COVENANT, THE [EXORDIUM 4]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
RIGHT HAND OF DEXTRA, THE  LAKE, DAVID J.  C/P,P
RIGHT-HANDED WILDERNESS  WELLS, ROBERT  C/P:2,P
RIGHT OFF THE MAP  MONTAGUE, C. E.  H:2
RIHANNSU [STAR TREK RIHANNSU 1,2,3,4]  DUANE, DIANE  H
RIM [VIRTUAL REALITY 1]  BESHER, ALEXANDER  P
RIM GODS, THE [GRIMES]  CHANDLER, A. BERTRAM  D72400
RIM OF SPACE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,DF133,P,H
RIM OF SPACE, THE [GRIMES]/SHIP FROM OUTSIDE, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
RIM OF THE UNKNOWN, THE  LONG, FRANK BELKNAP  SR-H,P
RIM-WORLD LEGACY, THE  JAVOR, FRANK A.  C/P:2,P
RIM-WORLD LEGACY AND BEYOND, THE  JAVOR, FRANK A.  P
RIME ISLE  LEIBER, FRITZ  SR-H
RIMRUNNERS [UNION/ALLIANCE]  CHERRYH, C. J.  C/P,P,H

RINCEWIND THE WIZARD [DISCWORLD 1,2,5,9/ DISCWORLD RINCEWIND 1,2,3,4] PRATCHETT, TERRY

H

RING, THE  ANTHONY, PIERS/MARGROFF, ROBERT E.  C/P:2,SFWA-P,P
RING [XEELEE]  BAXTER, STEPHEN  C/P:2,P
RING, THE [@SWORD,RING AND CHALICE 2]  CHESTER, DEBORAH  P
RING, THE  MORAN, DANIEL KEYS  H
RING-A-DING U F O’S, THE [MISS FROM SIS]  TRALINS, ROBERT  P
RING AROUND THE SUN  SIMAK, CLIFFORD D.  H:2,P,SR-DD061
RING OF ALLAIRE, THE [WINTER KING’S WAR 1]  DEXTER, SUSAN  C/P:2,P
RING OF ANGELICA, THE [@ORLANDO FURIOSO 1]  ARIOSTO, LUDOVICO  C/P,P
RING OF CHARON, THE [HUNTED EARTH 1]  ALLEN, ROGER MACBRIDE  C/P,SFWA-P,P,H
RING OF DESTINY [DANCE OF THE RINGS 3]  FANCHER, JANE S.  P
RING OF ENDLESS LIGHT, A  L’ENGLE, MADELEINE  P
RING OF FEAR  MCCAFFREY, ANNE  C/P,P
RING OF FIRE  FLINT, ERIC  PA
RING OF FIRE, THE [YNELL 1]  MURPHY, SHIRLEY ROUSSEAU  H,P
RING OF FIRE (PROOF) [SIXTEEN THIRTY-TWO]  FLINT, ERIC  SR-LPA
RING OF FIVE DRAGONS, THE [PEARL 1]  LUSTBADER, ERIC VAN  H:2,P
RING OF FIVE DRAGONS, THE (PROOF) [PEARL 1]  LUSTBADER, ERIC VAN  SR-LP
RING OF IKRIBU, THE [@RED SONJA #1]  SMITH, DAVID C./TIERNEY, RICHARD L.  C/P,P
RING OF INTRIGUE [DANCE OF THE RINGS 2]  FANCHER, JANE S.  C/P,P
RING OF LIGHTNING [DANCE OF THE RINGS 1]  FANCHER, JANE S.  P
RING OF RITORNEL, THE  HARNESS, CHARLES L.  C/P:2,P
RING OF SWORDS  ARNASON, ELEANOR  H
RING OF TRUTH, THE  LAKE, DAVID J.  SR-P,C/P

The MITSFS Titledex  September 24, 2007  Page 298
TITLE  AUTHOR  CODE
RING OF VIOLENCE  MASON, DOUGLAS R.  C/P,P
RING OF WINTER, THE [FORGOTTEN REALMS HARPERS #5]  LOWDER, JAMES  P
RINGING CHANGES  LAFFERTY, R. A.  C/P,P
RINGS  HARNESS, CHARLES L.  H
RINGS OF ICE  ANTHONY, PIERS  C/P:2,P
RINGS OF TANTALUS, THE [EXPENDABLES 2]  AVERY, RICHARD  C/P:2,P

RINGS OF TAUTEE, THE [STAR TREK] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

RINGSTONES  SARBAN  P
RINGWORLD (ORIGINAL) [KNOWN SPACE RINGWORLD 1]  NIVEN, LARRY  SR-P
RINGWORLD (REVISED) [KNOWN SPACE RINGWORLD 1]  NIVEN, LARRY  C/P,P
RINGWORLD ENGINEERS, THE [KNOWN SPACE RINGWORLD 2]  NIVEN, LARRY  H:2,P
RINGWORLD THRONE, THE [KNOWN SPACE RINGWORLD 3]  NIVEN, LARRY  C/P:2,P
RINGWORLD’S CHILDREN [KNOWN SPACE RINGWORLD 4]  NIVEN, LARRY  H:2
RINKITINK IN OZ [OZ #10]  BAUM, L. FRANK  C/P,@P,H
RINN’S STAR  DOWNING, PAULA E.  C/P,P
RIP TIDE [DOCTOR WHO]  COOPER, LOUISE  SR-H
RIPPER!  DOZOIS, GARDNER/CASPER, SUSAN  SFWA-P
RIPPING TIME [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
RISE AND FALL OF A DRAGON KING, THE [@DARK SUN ATHAS #5]  ABBEY, LYNN  P
RISE OF A HERO [@FARSALA #2]  BELL, HILARI  H
RISE OF A MERCHANT PRINCE [SERPENTWAR 2]  FEIST, RAYMOND E.  H:2,P
RISE OF A MERCHANT PRINCE (PROOF) [SERPENTWAR 2]  FEIST, RAYMOND E.  SR-LP
RISE OF ENDYMION, THE [HYPERION 4]  SIMMONS, DAN  H:2,P
RISE OF THE HORDE [WORLD OF WARCRAFT]  GOLDEN, CHRISTIE  P
RISEN, THE [WORLD OF DARKNESS WRAITH]  DITCHBURN, ELIZABETH/GROVE, HEATHER  SR-VLH
RISEN EMPIRE, THE [SUCCESSION 1]  WESTERFELD, SCOTT  P
RISING, THE [FLIGHT ENGINEER 1]  DOOHAN, JAMES/STIRLING, S. M.  P
RISING [STARGATE ATLANTIS #1]  MALCOLM, SALLY  P
RISING OF THE MOON, THE  CONNOLLY, FLYNN  P
RISING SON [@STAR TREK DS9]  PERRY, S. D.  P
RISING STARS: BORN IN FIRE [RISING STARS]  STRACZYNSKI, J. MICHAEL  CX
RISING STARS: FIRE AND ASH [@RISING STARS #3]  STRACZYNSKI, J. MICHAEL  CX
RISING STARS: POWER [RISING STARS]  STRACZYNSKI, J. MICHAEL  CX
RISING TIDE [@FORGOTTEN REALMS THREAT FROM THE SEA #1]  ODOM, MEL  P
RISSA AND TREGARE [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
RISSA KERGUELEN [HULZEIN DYNASTY]  BUSBY, F. M.  C/P,P,H
RITE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #2]  BYERS, RICHARD LEE  P
RITE  WILLIAMS, TAD  H
RITE OF CONQUEST  TARR, JUDITH  P
RITE OF PASSAGE  PANSHIN, ALEXEI  C/P:2,P
RITES OF OHE, THE  BRUNNER, JOHN  DF242
RITES OF THE DEMON [@DEATHWALKER #1]  CASTEVANO, ROMAN  P
RITHIAN TERROR, THE  KNIGHT, DAMON  C/P,DM113,P
RITNYM’S DAUGHTER  GILLULY, SHEILA  C/P:2,P
RITUAL, THE  WALTERS, R. R.  P
RITUALS OF INFINITY, THE  MOORCOCK, MICHAEL  C/P,P
RIVAL, THE [@FEY 3]  RUSCH, KRISTINE KATHRYN  P
RIVAL RIGELIANS, THE  REYNOLDS, MACK  DG632
RIVAL RIGELIANS, THE/PLANETARY AGENT X  REYNOLDS, MACK  P
RIVALS OF FRANKENSTEIN  PARRY, MICHEL  PA
RIVAL’S SON [KYDAN 2]  BROWN, SIMON  P
RIVAN CODEX, THE [BELGARIAD/MALLOREON]  EDDINGS, DAVID/EDDINGS, LEIGH  P
RIVEN SHIELD, THE [@SUN SWORD #5]  WEST, MICHELLE  P
RIVER  CAMERON, SKYLA DAWN  LP
RIVER AND THE DREAM, THE  JONES, RAYMOND F.  SR-P
RIVER AT GREEN KNOWE, THE [GREEN KNOWE 3]  BOSTON, L. M.  H,LP
RIVER JOURNEY, THE  NATHAN, ROBERT  H
RIVER OF BLOOD [@STAR TREK ERRAND OF VENGEANCE #3]  RYAN, KEVIN  P
RIVER OF BLUE FIRE [@OTHERLAND #2]  WILLIAMS, TAD  H,P
RIVER OF DANCING GODS, THE [DANCING GODS 1]  CHALKER, JACK L.  C/P:2,P
RIVER OF DUST  JABLOKOV, ALEXANDER  P
RIVER OF ETERNITY  FARMER, PHILIP JOSE  H
RIVER OF GOD, THE [SHATTERZONE 3]  FARSHTEY, GREG  P
RIVER OF GODS  MCDONALD, IAN  LP
RIVER OF ICE [@AVENGER #11]  ROBESON, KENNETH  C/P:2,P
RIVER OF OUR DESTINY, THE  RODGERS, ALAN  LP
RIVER OF TIME, THE  BRIN, DAVID  C/P:2,P,H
RIVER OF TIME  WEST, WALLACE  H
RIVER WALL, THE [GANDALARA 7]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
RIVERRUN [RIVERRUN 1]  SOMTOW, S. P.  P
RIVERRUN TRILOGY [RIVERRUN 1,2,3]  SOMTOW, S. P.  LP
RIVERS OF TIME  DE CAMP, L. SPRAGUE  P
RIVERS OF WAR, THE [RIVERS OF WAR 1]  FLINT, ERIC  H

The MITSFS Titledex  September 24, 2007  Page 299
TITLE  AUTHOR  CODE
RIVERWIND THE PLAINSMAN [DRAGONLANCE PRELUDES II 1]  THOMPSON, PAUL B./CARTER, TONYA R.  C/P,P
RIVERWORLD AND OTHER STORIES [RIVERWORLD]  FARMER, PHILIP JOSE  C/P:2,P
RIVET IN GRANDFATHER’S NECK, THE  ELLIOTT, BRUCE  C/P:2,P
RIVETS!!!  KELLER, MARK M./ANDERSON, SUE  LP
ROAD, THE  MCCARTHY, CORMAC  H
ROAD GHOST, THE [@TRAVELER #7]  DRUMM, D. B.  P
ROAD GOES EVER ON, THE  TOLKIEN, J. R. R./SWANN, DONALD  R/REF-H,C/REF-H
ROAD HOME, THE [GUARDIANS OF THE FLAME]  ROSENBERG, JOEL  P
ROAD OF AZRAEL, THE  HOWARD, ROBERT E.  P
ROAD OF KINGS, THE [BANTAM CONAN 4]  WAGNER, KARL EDWARD  P
ROAD OF THE PATRIARCH [@FORGOTTEN REALMS SELLSWORDS #3]  SALVATORE, R. A.  H
ROAD OF THE SEA HORSE, THE [@LAST VIKING #2]  ANDERSON, POUL  C/P:2,P
ROAD TO CORLAY, THE [BIRD OF KINSHIP 1]  COWPER, RICHARD  C/P,P,H
ROAD TO DAMASCUS, THE [BOLOS]  RINGO, JOHN/EVANS, LINDA  H
ROAD TO DAMASCUS, THE (PROOF) [BOLOS]  RINGO, JOHN/EVANS, LINDA  SR-LP
ROAD TO DEATH, THE [@EBERRON LOST MARK #2]  FORBECK, MATT  P

ROAD TO DUNE, THE [DUNE] HERBERT, FRANK/HERBERT, BRIAN/

ANDERSON, KEVIN J.

R/REF-H

ROAD TO EHVENOR, THE [GUARDIANS OF THE FLAME 6]  ROSENBERG, JOEL  H,P
ROAD TO HELL, THE [@HORIZON WAR #1/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P
ROAD TO MARS, THE  IDLE, ERIC  H
ROAD TO OZ, THE [OZ #5]  BAUM, L. FRANK  C/P,@P,@C/P,H
ROAD TO SCIENCE FICTION #1: FROM GILGAMESH TO WELLS, THE  GUNN, JAMES E.  C/PA,PA
ROAD TO SCIENCE FICTION #2: FROM WELLS TO HEINLEIN, THE  GUNN, JAMES E.  C/PA,PA
ROAD TO SCIENCE FICTION #3: FROM HEINLEIN TO HERE, THE  GUNN, JAMES E.  C/PA,PA
ROAD TO SCIENCE FICTION #4: FROM HERE TO FOREVER, THE  GUNN, JAMES E.  C/PA,PA
ROAD TO SCIENCE FICTION #5: THE BRITISH WAY, THE  GUNN, JAMES E.  LPA
ROAD TO SCIENCE FICTION #6: AROUND THE WORLD, THE  GUNN, JAMES E.  LPA
ROAD TO THE MIDDLE ISLANDS, THE [@WILDERNESS OF FOUR #4]  HANCOCK, NIEL  P
ROAD TO THE RIM, THE [GRIMES]  CHANDLER, A. BERTRAM  DH029
ROAD TO THE RIM, THE [GRIMES]/HARD WAY UP, THE [GRIMES]  CHANDLER, A. BERTRAM  P
ROAD TO UNDERFALL, THE [ELUNDIUM 1]  JEFFERIES, MIKE  C/P:2,P
ROAD WAR [@TRAVELER #5]  DRUMM, D. B.  P
ROAD WARS [@DEATHLANDS]  AXLER, JAMES  P
ROAD WEST, THE  WRIGHT, GARY  P
ROADKILL OF MIDDLE EARTH, THE  CARNELL, JOHN  LP
ROADMARKS  ZELAZNY, ROGER  C/P:2,P
ROADS BETWEEN THE WORLDS, THE  MOORCOCK, MICHAEL  LP
ROADS NOT TAKEN  DOZOIS, GARDNER/SCHMIDT, STANLEY  PA
ROADS OF HEAVEN, THE [SILENCE LEIGH 1,2,3]  SCOTT, MELISSA  H:2
ROADSIDE PICNIC  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P
ROADSIDE PICNIC/TALE OF THE TROIKA  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H:2,P
ROAR DEVIL, THE [DOC SAVAGE #88]  ROBESON, KENNETH  C/P,@P
ROAR OF HONOR [@MECHWARRIOR]  PARDOE, BLAINE LEE  P
ROBERT A. HEINLEIN: AMERICA AS SCIENCE FICTION  FRANKLIN, H. BRUCE  R/REF-H
ROBERT A. HEINLEIN: STRANGER IN HIS OWN LAND  SLUSSER, GEORGE EDGAR  R/REF-P

ROBERT ADAMS’ BOOK OF ALTERNATE WORLDS ADAMS, ROBERT/GREENBERG, MARTIN H./

ADAMS, PAMELA CRIPPEN

C/PA,PA

ROBERT ADAMS’ BOOK OF SOLDIERS ADAMS, ROBERT/GREENBERG, MARTIN H./

ADAMS, PAMELA CRIPPEN

C/PA,PA

ROBERT BLOCH: A BIO-BIBLIOGRAPHY  FLANAGAN, GRAEME  R/IND
ROBERT BLOCH: APPRECIATIONS OF THE MASTER  MATHESON, RICHARD/MAINHARDT, RICIA  R/REF-P
ROBERT E. HOWARD’S WORLD OF HEROES  HOWARD, ROBERT E.  P
ROBERT SILVERBERG OMNIBUS, A  SILVERBERG, ROBERT  H,SFWA
ROBERT SILVERBERG’S WORLDS OF WONDER  SILVERBERG, ROBERT  LPA
ROBIN AND THE KESTREL, THE [BARDIC VOICES 2]  LACKEY, MERCEDES  C/P,@H,P
ROBIN AND THE KING [SHERWOOD 2]  GODWIN, PARKE  C/P,P
ROBIN HOOD AMBUSH, THE [TIME TOURS 1]  WU, WILLIAM F.  P
ROBINSHEUGH  DUNLOP, EILEEN  P
ROBINSON CRUSOE 1,000,000 A.D.  SUNBORD, TERRY  LP

ROBINSON CRUSOE SCHOOL, THE/STAR OF THE SOUTH, THE/ PURCHASE OF THE NORTH POLE VERNE, JULES

SR-H

ROBOCOP  NAHA, ED  C/P,P
ROBOPATHS  YABLONSKY, LEWIS  [C/RAN-H,C/RAN-P]
ROBOT ADEPT [APPRENTICE ADEPT #5]  ANTHONY, PIERS  C/P:2,@P,SFWA-P
ROBOT AND THE MAN, THE  GREENBERG, MARTIN  HA
ROBOT BLUES [STAR OF THE GUARDIANS MAG FORCE 7]  WEIS, MARGARET/PERRIN, DON  C/P,P
ROBOT BRAINS, THE  BOUNDS, SYDNEY J.  C/P,P
ROBOT DREAMS  ASIMOV, ISAAC  C/P,SFWA-P:2,P
ROBOT IN THE CLOSET, THE  GOULART, RON  P
ROBOT INVITATION, THE [PERRY RHODAN 136]  SCHEER, K. H.  SR-LP
ROBOT JOX  THURSTON, ROBERT  C/P,P
ROBOT NOVELS, THE  ASIMOV, ISAAC  H:3
ROBOT ROCKET, THE [TOM CORBETT 8]  ROCKWELL, CAREY  H:2

The MITSFS Titledex  September 24, 2007  Page 300
TITLE  AUTHOR  CODE
ROBOT VISIONS  ASIMOV, ISAAC  C/P,SFWA-P:2,P,H
ROBOT VISIONS (PROOF)  ASIMOV, ISAAC  SR-LP

ROBOT WARRIORS WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA

ROBOT WHO LOOKED LIKE ME, THE  SHECKLEY, ROBERT  C/P:2,P
ROBOT WHO LOOKED LIKE ME, THE (BRITISH)  SHECKLEY, ROBERT  SR-P
ROBOT WORLD [INTERPLANETARY SPY 3]  MCEVOY, SETH  SCX
ROBOTECH [@ROBOTECH #1,2,3]  MCKINNEY, JACK  P
ROBOTECH ART 2  REYNOLDS, KAY  A
ROBOTECH ART 3: THE SENTINELS  MACEK, CARL  A
ROBOTICS  MINSKY, MARVIN  SFWA
ROBOTS AND CHANGELINGS  DEL REY, LESTER  C/P:2,P
ROBOTS AND EMPIRE  ASIMOV, ISAAC  C/P:2,P,H,SFWA
ROBOTS, BOMBS AND MUTANTS [PERRY RHODAN 125]  VOLTZ, WILLIAM  SR-LP
ROBOTS FROM ASIMOV’S  WILLIAMS, SHEILA  B
ROBOTS HAVE NO TAILS  KUTTNER, HENRY  C/P:2,P
ROBOTS, LES  ASIMOV, ISAAC  FP
ROBOTS OF DAWN, THE [LIJE BALEY 3]  ASIMOV, ISAAC  C/P:2,P,H
ROBOTS OF SATURN [@DIG ALLEN #5]  GREENE, JOSEPH  H
ROBUR LE CONQUERANT  VERNE, JULES  FP
ROC AND A HARD PLACE [@XANTH #19]  ANTHONY, PIERS  C/P:2,P
ROCANNON’S WORLD [HAINISH 1]  LE GUIN, URSULA K.  C/P:2,SFWA,DG574,P
ROCHEWORLD [ROCHEWORLD 1]  FORWARD, ROBERT L.  C/P:2,SFWA-P,P
ROCK, THE  DOHERTY, ROBERT  P
ROCK AND A HARD PLACE, A [STAR TREK TNG]  DAVID, PETER  C/P:2,P
ROCK ’N’ ROLL BABES FROM OUTER SPACE  JAIVIN, LINDA  H
ROCK OF AGES [DRAKE MAIJSTRAL 3]  WILLIAMS, WALTER JON  H,SFWA
ROCK RATS, THE [ASTEROID WARS 2]  BOVA, BEN  P
ROCK THE BIG ROCK  PETERSON, JOHN VICTOR  P
ROCKET INVASION  LANG, KING  SR-P
ROCKET JOCKEY  DEL REY, LESTER  C/P:2,P
ROCKET JOCKEY  SAINT JOHN, PHILIP  H
ROCKET JUMPER [@RICK BRANT #21]  BLAINE, JOHN  H
ROCKET MANUAL FOR AMATEURS  BRINLEY, B. R., CAPT.  [C/RAN-P]
ROCKET MOUSE  PRIESTLEY, LEE  P
ROCKET PILOT  SAINT JOHN, PHILIP  H
ROCKET RIDERS OVER THE DESERT  GARIS, HOWARD R.  H
ROCKET SCIENCE  LAKE, JAY  LP
ROCKET SHIP GALILEO  HEINLEIN, ROBERT A.  C/P,P,H
ROCKET TO LIMBO  NOURSE, ALAN E.  DD385
ROCKET TO LUNA  MARSTEN, RICHARD  H
ROCKET TO THE MORGUE  BOUCHER, ANTHONY  SR-P,C/P:2
ROCKETEER, THE  DAVID, PETER  P
ROCKETS IN URSA MAJOR  HOYLE, FRED/HOYLE, GEOFFREY  H:2,P
ROCKETS, JETS, GUIDED MISSILES, AND SPACE SHIPS  COGGINS, JACK/PRATT, FLETCHER  [C/RAN-H]
ROCKETS, REDHEADS & REVOLUTION  HOGAN, JAMES P.  C/P:2,P
ROCKET’S SHADOW, THE [@RICK BRANT #1]  BLAINE, JOHN  H:2
ROCKETS: THE FUTURE OF TRAVEL BEYOND THE STRATOSPHERE  LEY, WILLY  R/REF-H
ROCKETS TO NOWHERE  SAINT JOHN, PHILIP  H:2
ROD SERLING (BIOGRAPHY)  ENGEL, JOEL  R/REF-H
ROD SERLING’S DEVILS AND DEMONS  SERLING, ROD  C/PA,PA
ROD SERLING’S OTHER WORLDS  SERLING, ROD  PA
ROD SERLING’S THE TWILIGHT ZONE  GIBSON, WALTER B.  H,P
ROD SERLING’S TRIPLE W  SERLING, ROD  PA,C/PA
ROD SERLING’S TWILIGHT ZONE REVISITED  GIBSON, WALTER B.  H,P
RODENT MUTATION  FANE, BRON  SR-P
RODERICK [RODERICK THE ROBOT 1]  SLADEK, JOHN  C/P:2,SFWA-P,P
RODERICK AT RANDOM [RODERICK THE ROBOT 2]  SLADEK, JOHN  P
ROGER ZELAZNY  YOKE, CARL B.  R/REF-P
ROGER ZELAZNY’S VISUAL GUIDE TO CASTLE AMBER  ZELAZNY, ROGER/RANDALL, NEIL  R/REF-P,C/REF-H:3
ROGERS’ RANGERS [BUCK ROGERS (NIVEN & POURNELLE’S) 5]  SILBERSACK, JOHN  P
ROGUE [@STAR TREK TNG SECTION 31]  MANGELS, ANDY/MARTIN, MICHAEL A.  P
ROGUE [@ALIENS 6]  SCHOFIELD, SANDY  C/P,P
ROGUE BERSERKER [BERSERKER]  SABERHAGEN, FRED  H
ROGUE BOLO [BOLOS]  LAUMER, KEITH  C/P:2,P
ROGUE CLONE [CLONE REPUBLIC 2]  KENT, STEVEN L.  P
ROGUE DRAGON  DAVIDSON, AVRAM  C/P:2,P
ROGUE EMPEROR [CHRONOPLANE WARS 2]  KILIAN, CRAWFORD  C/P,P
ROGUE GOLEM  KENYON, ERNEST M.  P
ROGUE IN SPACE  BROWN, FREDRIC  C/P:2,P
ROGUE MOON  BUDRYS, ALGIS  C/P,SR-P,H
ROGUE OF GOR [GOR 15]  NORMAN, JOHN  C/P:2,P
ROGUE PIRATE  BETANCOURT, JOHN GREGORY  P
ROGUE PLANET [@STAR WARS]  BEAR, GREG  P

TITLE  AUTHOR  CODE
ROGUE PLANET [SPACE: 1999 9]  TUBB, E. C.  C/P,P
ROGUE POWERS [TORCH OF HONOR 2]  ALLEN, ROGER MACBRIDE  C/P,P
ROGUE QUEEN  DE CAMP, L. SPRAGUE  H:2,P
ROGUE SAUCER [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
ROGUE SHIP  VAN VOGT, A. E.  H:2,P
ROGUE SQUADRON [@STAR WARS X-WING 1]  STACKPOLE, MICHAEL A.  C/P,P
ROGUE STAR [STARS 2]  FLYNN, MICHAEL  P
ROGUE STAR [@WARHAMMER 40K]  HOARE, ANDY  P
ROGUE STAR  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
ROGUE SWORD  ANDERSON, POUL  P
ROIL OF STARS, A  WISMER, DON  P,SFWA-P
ROLE OF HONOR [JAMES BOND 18]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
ROLIND OF MERU  LYLE, PETER  C/P,P
ROLLBACK  SAWYER, ROBERT J.  H:2
ROLLER COASTER WORLD  BULMER, KENNETH  P
ROLLERBALL  HARRISON, WILLIAM  C/P,P
ROLLING HOT [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,SFWA-P:2,P
ROLLING STONES, THE  HEINLEIN, ROBERT A.  C/P:2,P
ROLLTOWN  REYNOLDS, MACK  C/P:2,P
ROMA ETERNA  SILVERBERG, ROBERT  H
ROMA MATER [@KING OF YS 1]  ANDERSON, POUL/ANDERSON, KAREN  C/P:2,P
ROMAN  JONES, DOUGLAS C.  SFWA-P
ROMANCE OF ATLANTIS, THE  CALDWELL, TAYLOR/STEARN, JESS  P
ROMANCE OF CRIME, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
ROMANCE OF THE EQUATOR, A  ALDISS, BRIAN W.  H:2
ROMANCE OF TWO WORLDS, A  CORELLI, MARIE  H
ROMANOV’S RUSSIAN-ENGLISH/ENGLISH-RUSSIAN DICTIONARY  WEDEL, E./ROMANOV, A. S.  DICT
ROMANTIC GOTHIC TALES 1790-1840  THOMPSON, G. RICHARD  PA
ROMBELLA SHUTTLE, THE  CONVERTITO, BILL  C/P,P
ROMULAN PRIZE, THE [STAR TREK TNG]  HAWKE, SIMON  C/P:2,P
ROMULAN STRATAGEM, THE [STAR TREK TNG]  GREENBERGER, ROBERT  C/P:2,P
ROMULAN WAY, THE [STAR TREK RIHANNSU 2]  DUANE, DIANE/MORWOOD, PETER  C/P:2,P
RONIA, THE ROBBER’S DAUGHTER  LINDGREN, ASTRID  LP
ROOFWORLD  FOWLER, CHRISTOPHER  P
ROOM WITH NO DOORS, THE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
ROOMS OF PARADISE  HARDING, LEE  SR-HA,HA
ROOTS OF HORROR IN THE FICTION OF H. P. LOVECRAFT, THE  SAINT ARMAND, BARTON LEVI  R/REF-H
ROOTS OF THE MOUNTAINS, THE  MORRIS, WILLIAM  LP
RORK!  DAVIDSON, AVRAM  C/P:2,P
ROSE, THE  HARNESS, CHARLES L.  C/P,P
ROSE AND THE KEY, THE  LE FANU, J. SHERIDAN  LP
ROSE AND THE RING, THE  THACKERAY, W. M.  SR-P
ROSE DAUGHTER  MCKINLEY, ROBIN  P
ROSE FOR ARMAGEDDON, A  SCHENCK, HILBERT  P,SFWA-P
ROSE FOR ECCLESIASTES, A  ZELAZNY, ROGER  P
ROSE MADDER  KING, STEPHEN  C/P,P
ROSE OF THE WORLD, THE [FOOL’S GOLD 3]  FISHER, JUDE  H
ROSE-RED CITY, A  DUNCAN, DAVE  P
ROSE SEA, THE  STIRLING, S. M./LISLE, HOLLY  P
ROSEMARY’S BABY  LEVIN, IRA  C/P,P
ROSEMARY’S BRAIN  SOUKUP, MARTHA  P
ROSETTA [STAR TREK ENTERPRISE]  STERN, DAVE  P
ROSTMA LURE, THE [@STARHAWKS #4]  DALTON, SEAN  C/P,P
ROSWELL [@STARGATE SG-1 #9]  WHITELAW, SONNY/FALLON, JENNIFER  P
ROTTEN RELATIONS  LITTLE, DENISE  PA
ROUES DES TENEBRES, LES  JAKES, JOHN  FP
ROUGH TRAILS [@STAR TREK NEW EARTH #3]  GRAF, L. A.  C/P,P
ROUND THE MOON  VERNE, JULES  H
ROUND TRIP TO THE YEAR 2000, A  COOK, WILLIAM W.  LP
ROUNDED WITH SLEEP  CHILSON, ROBERT  P
ROUNDHEADS, THE [DOCTOR WHO PAST DOCTORS]  GATISS, MARK  C/P:2,P
ROUSE A SLEEPING CAT [NIMNESTL 1]  CRAWFORD, DAN  P
ROUTE 666 [DARK FUTURE]  PRINGLE, DAVID  LP
ROUTE 666 [DARK FUTURE DEMON DOWNLOAD 2]  YEOVIL, JACK  P
ROVER, THE [ROVER 1]  ODOM, MEL  P
ROVERANDOM  TOLKIEN, J. R. R.  H
ROWAN, THE [ROWAN 1]  MCCAFFREY, ANNE  C/P,SFWA-P,H:2,P,SFWA
ROYAL ASSASSIN [FARSEER 2]  HOBB, ROBIN  C/P:2,P
ROYAL CHAOS  MCGIRT, DAN  C/P,P
ROYAL FOUR, THE [ONE LAND 2]  STOUT, AMY  P
RUBICON [@SENTINELS #5/ROBOTECH 17]  MCKINNEY, JACK  C/P,P
RUBY  STEWART, KERRY  P
RUBY FLAMES [SHATTERED LIGHT 1]  KNAAK, RICHARD A.  P
RUBY GUARDIAN, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #2]  REID, THOMAS M.  P

The MITSFS Titledex  September 24, 2007  Page 302
TITLE  AUTHOR  CODE
RUBY KNIGHT, THE [ELENIUM 2]  EDDINGS, DAVID  H:2,P
RUBY RAY MYSTERY, THE [@RICK BRANT #19]  BLAINE, JOHN  H
RUBY SLIPPERS, GOLDEN TEARS [COMPLEMENTARY FAIRY COLORS 3]  DATLOW, ELLEN/WINDLING, TERRI  HA,PA
RUDE ASTRONAUTS  STEELE, ALLEN  C/P,P
RUDE MECHANICALS [COMPANY]  BAKER, KAGE  SR-H
RUEDA DE CIELO, LA  LE GUIN, URSULA K.  FP
RUIMTEVAART  GAIL, O. W./PETRI, W.  FP
RUIN, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #3]  BYERS, RICHARD LEE  P
RUIN [@STAR WARS NEW JEDI ORDER DARK TIDE #2]  STACKPOLE, MICHAEL A.  P
RUINS [X-FILES]  ANDERSON, KEVIN J.  H,P
RUINS OF AMBRAI, THE [@EXILES #1]  RAWN, MELANIE  C/P,P,H

RUINS OF DANTOOINE [@STAR WARS GALAXIES] WHITNEY-ROBINSON, VORONICA/

BLACKMAN, HADEN

P

RUINS OF EARTH, THE  DISCH, THOMAS M.  PA,C/PA
RUINS OF ISIS, THE  BRADLEY, MARION ZIMMER  C/P,P,H
RUINS OF KALDAK, THE [@BLADE #34]  LORD, JEFFREY  P
RUINS OF POWER, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  VARDEMAN, ROBERT E.  P
RULE GOLDEN  KNIGHT, DAMON  TD34
RULE GOLDEN AND OTHER STORIES  KNIGHT, DAMON  P
RULE OF EVIDENCE [PAUL SINCLAIR 3]  HEMRY, JOHN G.  P
RULE OF THE DOOR AND OTHER FANCIFUL REGULATIONS, THE  BIGGLE, LLOYD JR.  C/P,P,H
RULE OF THE PAGBEASTS, THE  MCINTOSH, J. T.  SR-P,C/P
RULED BRITANNIA  TURTLEDOVE, HARRY  H
RULER OF NAUGHT [EXORDIUM 2]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
RULER OF THE WORLD  MCINTOSH, J. T.  SR-P,C/P:2
RULERS OF MEN  SANTESSON, HANS STEFAN  PA,C/PA
RULERS OF THE DARKNESS [INTO THE DARKNESS 4]  TURTLEDOVE, HARRY  H
RULES OF ASCENSION [WINDS OF THE FORELANDS 1]  COE, DAVID B.  H
RULES OF CONFLICT [JANI KILIAN 2]  SMITH, KRISTINE  C/P,P
RULES OF ENGAGEMENT [HERIS SERRANO 5]  MOON, ELIZABETH  C/P,P,H
RULES OF ENGAGEMENT [STAR TREK]  MORWOOD, PETER  C/P:2,P
RUMOR OF ANGELS, A  KELLOGG, MARJORIE BRADLEY  C/P,P
RUMORS OF SPRING  GRANT, RICHARD  C/P,SFWA,P,H
RUMOUR IN THE FOREST, THE  COUPPEY, MADELEINE  H
RUN, COME SEE JERUSALEM!  MEREDITH, RICHARD C.  C/P:2,P
RUN FOR THE STARS  ELLISON, HARLAN  TD32
RUN HARD, DIE FAST [@SHADOWRUN 35]  ODOM, MEL  P
RUN TO CHAOS KEEP, THE [QUINTARA 2]  CHALKER, JACK L.  SFWA-P:2,H:2,P,SFWA:2

RUN TO STARLIGHT GREENBERG, MARTIN H./ OLANDER, JOSEPH D./ WARRICK, PATRICIA S.

HA

RUN TO THE STARS  ROHAN, MICHAEL SCOTT  P
RUNAWAY ROBOT, THE  DEL REY, LESTER  LP
RUNCIBLE TALES [POLITY]  ASHER, NEAL  SR-LP
RUNELORDS, THE [RUNELORDS 1]  FARLAND, DAVID  C/P:2,P
RUNES  MONACO, RICHARD  C/P:2,P
RUNES & RIDDLES [EMPIRE OF ELWOLF]  CARRADINE, F. J.  LP
RUNES OF AUTUMN  ELMORE, LARRY/ELMORE, ROBERT  P
RUNES OF SORCERY, THE [RUNESPELL 3]  WELCH, JANE  P
RUNES OF THE EARTH, THE [LAST THOMAS COVENANT 1]  DONALDSON, STEPHEN R.  H
RUNES OF THE LYRE  MAYHAR, ARDATH  P
RUNES OF THE LYRE, THE  MAYHAR, ARDATH  H
RUNES OF WAR, THE [RUNESPELL 1]  WELCH, JANE  P
RUNESPEAR  MILAN, VICTOR/SNODGRASS, MELINDA  C/P,P
RUNESTAFF, THE [RUNESTAFF 4]  MOORCOCK, MICHAEL  C/P:2,P
RUNNER [JAK REBO 1]  DIETZ, WILLIAM C.  P
RUNNING FROM THE DEITY [HUMANX FLINX]  FOSTER, ALAN DEAN  H
RUNNING MAN, THE  BACHMAN, RICHARD  SR-L
RUNNING MAN, THE  HOLLY, JOAN HUNTER  C/P:2,P
RUNNING MAN, THE  KING, STEPHEN  P
RUNNING OUT OF TIME  HADDIX, MARGARET PETERSON  LP
RUNNING WITH THE DEMON [VOID & WORD 1]  BROOKS, TERRY  H,P
RUNTS OF 61 CYGNI C  GRAZIER, JAMES  C/P,P
RUPERT OF HENTZAU [ZENDA 2]  HOPE, ANTHONY  SR-H
RUSALKA  CHERRYH, C. J.  C/P,P,H
RUSSIAN SCIENCE FICTION  MAGIDOFF, ROBERT  HA
RUSSIAN SCIENCE FICTION 1956–1974: A BIBLIOGRAPHY  SUVIN, DARKO  R/IND
RUSSIAN SCIENCE FICTION: 1968  MAGIDOFF, ROBERT  HA
RUSSIAN SCIENCE FICTION: 1969  MAGIDOFF, ROBERT  HA
RUSSIAN SPRING  SPINRAD, NORMAN  C/P,P,H
RUST  CALIF, R. C.  P
RUSTLE IN THE GRASS, A  HAWDON, ROBIN  P,SFWA,SFWA-P
RUTH, THE UNSUSPECTING [@PHENWICK #17]  KIMBROUGH, KATHERYN  P
RYN  WODHAMS, JACK  SR-P

TITLE  AUTHOR  CODE
RYNOSSEROS  DOWLING, TERRY  H:2

The MITSFS Titledex  September 24, 2007  Page 304
TITLE  AUTHOR  CODE
S-F 2  MEYERS, RICHARD S.  R/REF-H
S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY  MERRIL, JUDITH  C/PA,PA
S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 2  MERRIL, JUDITH  PA,C/PA
S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 3  MERRIL, JUDITH  PA,C/PA
S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 4  MERRIL, JUDITH  PA,C/PA
S IS FOR SPACE  BRADBURY, RAY  C/P,P,H
S. O. S. LUA  CLARKE, ARTHUR C.  FP
SABAZEL  CARL, LILLIAN STEWART  P
SABBAT MARTYR [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
SABELLA, OR, THE BLOOD STONE  LEE, TANITH  C/P,P
SABER AND SHADOW [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,P
SABERHAGEN: MY BEST  SABERHAGEN, FRED  C/P,P
SABLE MOON, THE [BOOKS OF THE ISLE 3]  SPRINGER, NANCY  C/P:2,P
SABLE, SHADOW, AND ICE  FRANKLIN, CHERYL J.  C/P,P
SABOTAGE [@DOOM’S DAY #2]  ASKEGREN, PIERCE/FINGEROTH, DANNY  C/P,P
SABOTAGE IN SPACE [TOM CORBETT 7]  ROCKWELL, CAREY  H

SABOTEURS IN A-1 [@PERRY RHODAN #115]/ PSYCHO DUEL, THE [PERRY RHODAN 116] ACE

C/PA,PA

SABRETOOTH UNLEASHED [X-MEN]  KAMIDA, VICKI  [C/RAN-P]
SABRINA GOES TO ROME [@SABRINA THE TEENAGE WITCH]  ODOM, MEL  P
SABRINA THE TEENAGE WITCH [@SABRINA THE TEENAGE WITCH #1]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
SACRAMENT  BARKER, CLIVE  H
SACRAMENT OF NIGHT  COOPER, LOUISE  H
SACRED FIRE [@DRAGONLANCE KINGPRIEST #3]  PIERSON, CHRIS  P
SACRED FLESH [@WARHAMMER ANGELIKA FLEISCHER 2]  LAWS, ROBIN D.  P
SACRED GIRAFFE, THE  MADARIAGA, SALVADOR DE  H
SACRED GROUND  LACKEY, MERCEDES  C/P,P,H
SACRED LOCOMOTIVE FLIES  LUPOFF, RICHARD A.  C/P,P
SACRED POOL, THE [SORCERER’S TALE 1]  DOUGLAS, L. WARREN  H
SACRED SCIMITER, THE  BELL, WILLIAM DIXON  H
SACRED SEVEN, THE  STOUT, AMY  P
SACRED SYMBOLS OF MU, THE  CHURCHWARD, JAMES  [C/RAN-P:2]
SACRED VISIONS  GREELEY, ANDREW M./CASSUTT, MICHAEL  HA
SACRIFICE, THE [REALMS OF ARKANIA THALIONMEL 2]  KRAMER, INA  P
SACRIFICE, THE [@FEY 1]  RUSCH, KRISTINE KATHRYN  C/P,P
SACRIFICE, THE  SUTTON, HENRY  P
SACRIFICE [STAR WARS LEGACY OF THE FORCE 5]  TRAVISS, KAREN  H
SACRIFICE MOON [@STARGATE SG-1 #2]  FORTUNE, JULIE  P

...SACRIFICE OF ANGELS [@STAR TREK DOMINION WAR #4/ STAR TREK DS9] CAREY, DIANE

P

SACRIFICE OF THE WIDOW [@FORGOTTEN REALMS LADY PENITENT #1]  SMEDMAN, LISA  C/P,P
SADAR’S KEEP [QUEEN’S QUARTER 2]  SNYDER, MIDORI  C/P:2,SFWA-P,P
SAFARI OF THE MIND, A  RESNICK, MIKE  LP
SAFFRON AND BRIMSTONE  HAND, ELIZABETH  LP
SAGA OF A STAR WORLD [@BATTLESTAR GALACTICA 1978]  LARSON, GLEN A./THURSTON, ROBERT  P
SAGA OF CUCKOO, THE  POHL, FREDERIK/WILLIAMSON, JACK  H:2
SAGA OF FILSTER STEIN, THE  LOVISI, GARY  B
SAGA OF GRITTEL SUNDOTHA, THE  MAYHAR, ARDATH  H
SAGA OF LOST EARTHS [KALEVALA 1]  PETAJA, EMIL  C/P:2,P
SAGA OF OLD CITY [@GREYHAWK #1]  GYGAX, GARY  C/P:2,P
SAGA OF THE REINDEER PEOPLE, A  LINDHOLM, MEGAN  H
SAGA OF THE RENUNCIATES, THE [DARKOVER]  BRADLEY, MARION ZIMMER  P
SAGAN DIARY, THE [OLD MAN’S WAR]  SCALZI, JOHN  H
SAGE, THE [@STAR STONE #2]  STASHEFF, CHRISTOPHER  H,P
SAGE OF SARE [CAITHAN CRUSADE 3]  SMITH, JULIE DEAN  C/P,P
SAGES & SWORDS  BLACKSTON, DANIEL E.  LPA
SAGITTARIUS  RUSSELL, RAY  C/P:2,P
SAGITTARIUS IS BLEEDING [@BATTLESTAR GALACTICA 2003]  DAVID, PETER  LP
SAGITTARIUS WHORL [RAMPART WORLDS 3]  MAY, JULIAN  C/P,P
SAGRAMANDA  FOSTER, ALAN DEAN  H
SAIL WITH PIRATES [@TIME MACHINE #4]  GASPERINI, JIM  P
SAILING BRIGHT ETERNITY [GALACTIC CENTER 6]  BENFORD, GREGORY  C/P:2,P
SAILING TO BYZANTIUM  SILVERBERG, ROBERT  SFWA-TD10,TD10
SAILING TO BYZANTIUM (COLLECTION)  SILVERBERG, ROBERT  LP
SAILING TO SARANTIUM [SARANTINE MOSAIC 1]  KAY, GUY GAVRIEL  H
SAILING TO SARANTIUM (BRITISH) [SARANTINE MOSAIC 1]  KAY, GUY GAVRIEL  C/P,P,H
SAILING TO UTOPIA (OMNIBUS EDITION)  MOORCOCK, MICHAEL  P
SAILOR ON THE SEAS OF FATE, THE [ELRIC 2]  MOORCOCK, MICHAEL  C/P:2,P
SAILS OF TAU CETI, THE  MCCOLLUM, MICHAEL  C/P:2,P
SAINT AND THE HUNCHBACK, THE  STAUFFER, DONALD A.  H
SAINT CAMBER [DERYNI CAMBER 2]  KURTZ, KATHERINE  C/P,P,H
SAINT ERRANT  CHARTERIS, LESLIE  P
SAINT FIRE  LEE, TANITH  H
SAINT-GERMAIN CHRONICLES, THE [SAINT-GERMAIN 6]  YARBRO, CHELSEA QUINN  H,P

The MITSFS Titledex  September 24, 2007  Page 305
TITLE  AUTHOR  CODE
SAINT LEIBOWITZ AND THE WILD HORSE WOMAN [LEIBOWITZ 2]  MILLER, WALTER M. JR.  H:2,LP
SAINT OF THE DRAGONS, THE  HIGHTMAN, JASON  H
ST. PATRICK’S GARGOYLE  KURTZ, KATHERINE  H
SAINT PETER’S WOLF  CADNUM, MICHAEL  H
SAINT STEPS IN, THE  CHARTERIS, LESLIE  H
SAINT VIDICON TO THE RESCUE  STASHEFF, CHRISTOPHER  P
SAINTS  CARD, ORSON SCOTT  SFWA-P
SAINTS OF THE SWORD, THE [TYRANTS AND KINGS 3]  MARCO, JOHN  H,LP
SALAMANDASTRON [REDWALL]  JACQUES, BRIAN  C/P,P
SALAMANDER WAR  CARR, CHARLES  SR-P
SALEM ON TRIAL [@SABRINA THE TEENAGE WITCH #8]  WEISS, DAVID CODY/WEISS, BOBBI JG  P
’SALEM’S LOT  KING, STEPHEN  C/P,P,H
SALIVA TREE, THE  ALDISS, BRIAN W.  SR-P,TD3,SFWA-TD3
SALMON OF DOUBT, THE  ADAMS, DOUGLAS  H
SALON FANTASTIQUE  DATLOW, ELLEN/WINDLING, TERRI  LPA
SAL’S BOOK [@PSI PATROL 1]  LIQUORI, SAL  P
SALT  ROBERTS, ADAM  H,LP
SALTFLOWER  VAN SCYOC, SYDNEY J.  C/P:2,P
SALUKAN GAMBIT [@STARHAWKS #6]  DALTON, SEAN  P
SALUTE TO BAZARADA AND OTHER STORIES  ROHMER, SAX  H
SALVAGE AND DESTROY  LLEWELLYN, EDWARD  C/P:2,P
SALVAGE FROM THE CYNTHIA  VERNE, JULES  H
SALVATION [@NECROMUNDA]  GOTO, C. S.  P
SALVATION [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
SALVATION BEHIND BARS  ELWOOD, ROGER  [C/RAN-P]
SALYUT IN ORBIT  VASILIEV, M. P., ET AL.  FH
SAM GUNN FOREVER [SAM GUNN]  BOVA, BEN  P
SAM GUNN OMNIBUS, THE [SAM GUNN]  BOVA, BEN  H
SAM GUNN, UNLIMITED [SAM GUNN]  BOVA, BEN  P
SAM WESKIT ON THE PLANET FRAMINGHAM  JOHNSTON, WILLIAM  C/P:2,P
SAMANTHA SLADE: MONSTER-SITTER [@SAMANTHA SLADE #1]  SMITH, SUSAN  P
SAMARKAND  DIAMOND, GRAHAM  P
SAMARKAND DAWN  DIAMOND, GRAHAM  P
SAMARKAND SOLUTION, THE [@DANGEROUS JOURNEYS 2]  GYGAX, GARY  C/P,P
SAMAX, THE GLADIATOR  FARADAY, ROBERT  LP
SAME LIE TWICE, THE  GOULART, RON  P
SAMURAI, THE [@CASCA #19]  SADLER, BARRY  P
SAMURAI CAT GOES TO HELL [SAMURAI CAT 5]  ROGERS, MARK E.  LP
SAMURAI CAT IN THE REAL WORLD [SAMURAI CAT 3]  ROGERS, MARK E.  VLP
SAMURAI WIZARD, THE [WYRDRUNE 5]  HAWKE, SIMON  C/P,P
SAN DIEGO LIGHTFOOT SUE  REAMY, TOM  C/P:2,P
SANCTION, THE [V]  TAGGERT, BRIAN  SR-S
SANCTUARY [THIEVES’ WORLD]  ABBEY, LYNN  H
SANCTUARY  ASPRIN, ROBERT LYNN  HA
SANCTUARY [DRAGON JOUSTERS 3]  LACKEY, MERCEDES  H:2
SANCTUARY [@ANGEL]  MARIOTTE, JEFF  P
SANCTUARY [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
SANCTUARY [@DRAGONLANCE ELVEN EXILES #1]  THOMPSON, PAUL B./COOK, TONYA C.  P
SANCTUARY [STAR TREK]  VORNHOLT, JOHN  C/P:3,P
SANCTUARY IN THE SKY  BRUNNER, JOHN  DD471
SAND AND STARS [STAR TREK]  POCKET BOOKS  LPA
SAND OF SATURN  FRENCH, ROBERT  P
SANDCATS OF RHYL, THE  VARDEMAN, ROBERT E.  P
SANDEAGOZU  JENNER, JANANN V.  H
SANDKINGS  MARTIN, GEORGE R. R.  C/P:2,P,H
SANDMAN BOOK OF DREAMS, THE [SANDMAN]  GAIMAN, NEIL/KRAMER, ED  LPA
SANDRA EFFECT, THE  LAMPP, JAMES  P
SANDRY’S BOOK [@CIRCLE OF MAGIC #1]  PIERCE, TAMORA  P
SANDS OF KALAVEN, THE [@SHUNLAR 2]  HELLER, CAROL  P
SANDS OF KARAKORUM, THE  ULLMAN, JAMES R.  P
SANDS OF MARS, THE  CLARKE, ARTHUR C.  H:2,SR-H,P
SANDS OF THE SOUL [@FORGOTTEN REALMS SEMBIA]  WHITNEY-ROBINSON, VORONICA  P
SANDS OF TIME  DAGMAR, PETER  H
SANDS OF TIME, THE [DOCTOR WHO MISSING ADVENTURES]  RICHARDS, JUSTIN  C/P,P
SANDWORLD  LUPOFF, RICHARD A.  C/P,P
SANNIKOV LAND  OBRUCHEV, VLADIMIR  SR-H
SANTA CLAUS: THE MOVIE  VINGE, JOAN D.  P
SANTA STEPS OUT  DEVEREAUX, ROBERT  P
SANTANA MORNING  DOLAN, MIKE  SR-P,C/P
SANTAROGA BARRIER, THE  HERBERT, FRANK  C/P:2,P
SANTA’S LITTLE HELPER [@SABRINA THE TEENAGE WITCH #5]  DUBOWSKI, CATHY EAST  P
SANTIAGO [SANTIAGO 1]  RESNICK, MIKE  C/P:2,SFWA-P,P

SAPPHIRE CRESCENT, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #1] REID, THOMAS M.

P

The MITSFS Titledex  September 24, 2007  Page 306
TITLE  AUTHOR  CODE
SAPPHIRE ROAD  WHITEFORD, WYNNE  SR-P,C/P
SAPPHIRE ROSE, THE [ELENIUM 3]  EDDINGS, DAVID  C/P,P,H
SAPPHIRE THRONE, THE [JEWELFIRE 2]  WARRINGTON, FREDA  P
SAR  RUSSELL, JOHN ROBERT  C/P:2,P
SARABAND OF LOST TIME  GRANT, RICHARD  C/P,P
SARAH [WOMEN OF GENESIS 1]  CARD, ORSON SCOTT  H
SARAH HALL’S SEA GOD  DUBOIS, THEODORA  P
SARATOGA [STAR TREK DS9]  FRIEDMAN, MICHAEL JAN  P
SARDONICUS AND OTHER STORIES  RUSSELL, RAY  P
SARDONYX NET, THE  LYNN, ELIZABETH A.  C/P,P,H
SAREK [STAR TREK]  CRISPIN, A. C.  H:2,P
SARGASSO  CORLEY, EDWIN  C/P,P
SARGASSO OF SPACE [SOLAR QUEEN 1]  NORTH, ANDREW  DD249
SARGASSO OF SPACE [SOLAR QUEEN 1]  NORTON, ANDRE  C/P:2,P
SARGASSO OGRE, THE [DOC SAVAGE #18]  ROBESON, KENNETH  H,@P
SARSEN WITCH, THE  KERNAGHAN, EILEEN  P
SASQUATCH: MONSTER OF THE NORTHWEST WOODS  KNERR, M. E.  P
SASSINAK [PLANET PIRATES 1]  MCCAFFREY, ANNE/MOON, ELIZABETH  C/P:2,SFWA-P,P
SATAN BLACK [DOC SAVAGE #97]/CARGO UNKNOWN [DOC SAVAGE 98]  ROBESON, KENNETH  @P
SATAN BUG, THE  MACLEAN, ALISTAIR  P
SATAN BUG, THE  STUART, IAN  H

SATAN: HIS PSYCHOTHERAPY AND CURE BY THE UNFORTUNATE DR.

KASSLER, J.S.P.S.

LEVEN, JEREMY

H

SATAN IN GORAY  SINGER, ISAAC BASHEVIS  LP
SATAN, SORCERY AND SEX  LAURENCE, THEODOR  [C/RAN-H]
SATANA ENSLAVED [@SPACEWAYS #4]  CLEVE, JOHN  P
SATANIC MILL, THE  PREUSSLER, OTFRIED  SFWA-P
SATANIC VERSES, THE  RUSHDIE, SALMAN  H
SATANIST, THE  WHEATLEY, DENNIS  P
SATANISTS, THE  HAINING, PETER  PA
SATAN’S CHANCE  SHRADER, ALAN ROSS  P
SATAN’S LOVE CHILD  MCNAUGHTON, BRIAN  SR-P
SATAN’S MISTRESS  MCNAUGHTON, BRIAN  P
SATAN’S PETS  GHIDALIA, VIC  PA
SATAN’S SEDUCTRESS  MCNAUGHTON, BRIAN  P
SATAN’S SEED [@CHILL #1]  SHERMAN, JORY  P
SATAN’S SISTER  JENSEN, RUBY JEAN  P
SATANS UNIVERSUM [@PERRY RHODAN #85]  VLCEK, ERNST  FPR
SATAN’S WORLD  ANDERSON, POUL  SR-H,H:2,P
SATELLITE!  BERGAUST, ERIK/BELLER, WILLIAM  [C/RAN-H,C/RAN-P:2]
SATELLITE 54-ZERO  MASON, DOUGLAS R.  C/P,P
SATELLITE B. C.  LE PAGE, RAND  SR-P
SATELLITE CITY  REYNOLDS, MACK  C/P:2,P
SATELLITE E ONE  CASTLE, JEFFERY LLOYD  H:2,P
SATELLITE IN SPACE  LOW, PROF. A. M.  H
SATELLITE NIGHT FEVER [SATELLITE NIGHT 3]  HOPKINS, JACK  C/P,P
SATELLITE NIGHT NEWS [SATELLITE NIGHT 1]  HOPKINS, JACK  C/P,P
SATELLITE NIGHT SPECIAL [SATELLITE NIGHT 2]  HOPKINS, JACK  C/P,P
SATELLITES, ROCKETS AND OUTER SPACE  LEY, WILLY  [C/RAN-P]
SATI  PIKE, CHRISTOPHER  P
SATORI  SCHMIDT, DENNIS  C/P,P
SATURDAY EVENING POST FANTASY STORIES, THE  FLES, BARTHOLD  SR-PA
SATURDAY EVENING POST STORIES 1953, THE  SATURDAY EVENING POST  HA
SATURN  BOVA, BEN  H
SATURN GAME, THE  ANDERSON, POUL  H,TD14,SFWA-TD14
SATURN PATROL  TUBB, E. C.  LP
SATURN RETURNS  WILLIAMS, SEAN  P
SATURN RUKH  FORWARD, ROBERT L.  C/P,P,H
SATURNALIA  CALLIN, GRANT  C/P:2,P

SATURN’S CHILD NICHOLS, NICHELLE/

BONANNO, MARGARET WANDER

H

SATURN’S RACE  NIVEN, LARRY/BARNES, STEVEN  H:2,P
SATYR TREK  KAINEN, RAY  SR-L
SATYRDAY  BAUER, STEVEN  H,P
SATYRICON  PETRONIUS ARBITER, TITUS  P
SAUCER  COONTS, STEPHEN  P
SAUCER HILL  ADLER, PAUL  P

SAUCER OF LONELINESS, A [COMPLETE STORIES OF THEODORE STURGEON 7] STURGEON, THEODORE

H

SAUCER WISDOM  RUCKER, RUDY  LP
SAURON DEFEATED [HISTORY OF MIDDLE-EARTH 9]  TOLKIEN, J. R. R.  H
SAURON DOMINION [@WAR WORLD #3]  POURNELLE, JERRY  SFWA-P,C/PA,PA
SAVAGE AMUSEMENT, THE [JUDGE DREDD]  BISHOP, DAVID  P
SAVAGE CAVES, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P

The MITSFS Titledex  September 24, 2007  Page 307
TITLE  AUTHOR  CODE
SAVAGE CITY [@WARHAMMER]  EARL, ROBERT  P
SAVAGE EMPIRE [SAVAGE EMPIRE 1]  LORRAH, JEAN  C/P:2,P
SAVAGE HEROES  PARRY, MICHEL  HA
SAVAGE HORDE, THE [@SURVIVALIST #6]  AHERN, JERRY  P
SAVAGE MESSIAH [DESTINIES OF BLOOD AND STONE 1]  NEWCOMB, ROBERT  H
SAVAGE MOUNTAINS, THE [HORSECLANS 5]  ADAMS, ROBERT  C/P:2,P
SAVAGE PELLUCIDAR [PELLUCIDAR #7]  BURROUGHS, EDGAR RICE  C/P:2,@P
SAVAGE REPORT: 1994, THE  RHEINGOLD, HOWARD  SR-P,C/P
SAVAGE SCORPIO [DRAY PRESCOT 16]  AKERS, ALAN BURT  C/P:2,P
SAVAGE STARS, THE  REINSMITH, RICHARD  P
SAVAGE TALES OF SOLOMON KANE, THE [SOLOMON KANE]  HOWARD, ROBERT E.  H
SAVAGE TOMORROW  DONOHUE, TREVOR  SR-P,C/P
SAVAGES, THE [DOCTOR WHO]  BLACK, IAN STUART  SR-P
SAVAGES OF GOR [GOR 17]  NORMAN, JOHN  C/P,P
SAVING SOLACE [@DRAGONLANCE CHAMPIONS]  CLARK, DOUGLAS W.  P
SAVING WORLDS  ELWOOD, ROGER/KIDD, VIRGINIA  HA

SAVIOR OF THE EMPIRE [@PERRY RHODAN #117]/ SHADOWS ATTACK, THE [PERRY RHODAN 118] ACE

PA

SAVOY BOOK, THE  BRITTON, DAVID/BUTTERWORTH, MICHAEL SR-PA
SAW  KATZ, STEVE  H
SAXON SHORE, THE [CAMULOD 4]  WHYTE, JACK  P
SCALEHUNTER’S BEAUTIFUL DAUGHTER, THE  SHEPARD, LUCIUS  H
SCALES OF INJUSTICE, THE [DOCTOR WHO MISSING ADVENTURES]  RUSSELL, GARY  C/P:2,P
SCALES OF JUSTICE [@THESE LAWLESS WORLDS #2]  COMSTOCK, JARROD  P
SCALES OF JUSTICE [FANUILH 4]  HOOD, DANIEL  P
SCALES OF PROMETHEUS, THE [@APPLESEED #3]  SHIROW, MASAMUNE  CX
SCALES OF THE SERPENT [@DIABLO SIN WAR #2]  KNAAK, RICHARD A.  P
SCALLION STONE, THE  SMITH, CANON BASIL A.  H
SCANNER DARKLY, A  DICK, PHILIP K.  C/P,H:3,LP,SFWA
SCANNERS  WHITESON, LEON  P
SCANNERS II: THE NEW ORDER  KIMBALL, PROFESSOR JANUS  P
SCAPESCOPE  STITH, JOHN E.  P
SCAR, THE  MIEVILLE, CHINA  H,LP
SCAR NIGHT [DEEPGATE 1]  CAMPBELL, ALAN  H
SCARABIAN NIGHTS [@SABRINA THE TEENAGE WITCH #24]  HOLDER, NANCY  P
SCARDOWN [JENNY CASEY 2]  BEAR, ELIZABETH  P
SCARE CARE  MASTERTON, GRAHAM  C/PA,PA,SFWA-P
SCARE TACTICS  FARRIS, JOHN  SFWA-P
SCARECROW OF OZ, THE [OZ #9]  BAUM, L. FRANK  H:2,@P
SCARED STIFF  CAMPBELL, RAMSEY  LP
SCARED STIFF (PROOF)  CAMPBELL, RAMSEY  SR-LP:2
SCARF, THE  BLOCH, ROBERT  SR-P
SCARFACE  NORTON, ANDRE  P
SCARLET ARENA 30303  MOORE, SILAS  H
SCARLET EMPIRE, THE  PARRY, DAVID M.  H
SCARLET EMPRESS, THE [DOCTOR WHO EIGHTH DOCTOR]  MAGRS, PAUL  C/P,P
SCARLET FEATHER  GRANT, JOAN  P
SCARLET LAKE MYSTERY, THE [@RICK BRANT #13]  BLAINE, JOHN  H
SCARLET SKULL, THE  OWSTON, C. E.  C/P,P
SCARS AND OTHER DISTINGUISHING MARKS  MATHESON, RICHARD CHRISTIAN  SFWA-P
SCARS OF DRACULA, THE  HALL, ANGUS  P
SCATHACH AND MAEVE’S DAUGHTERS, THE  WALKER, MARY ALEXANDER  SFWA
SCATTER OF STARDUST, A  TUBB, E. C.  D79975,LP
SCATTERBRAIN  NIVEN, LARRY  H
SCATTERED SUNS [SEVEN SUNS 4]  ANDERSON, KEVIN J.  H
SCATTERING OF JADES, A  IRVINE, ALEXANDER C.  H,P
SCAVENGER, THE [@TIMECOP 2]  PARKINSON, DAN  P
SCAVENGER HUNT  GOLDIN, STEPHEN  SR-P
SCAVENGERS  SKAL, DAVID J.  P
SCAVENGERS IN SPACE  NOURSE, ALAN E.  C/P:2,P
SCENT OF MAGIC  NORTON, ANDRE  C/P,P,H
SCENT OF NEW-MOWN HAY, A  BLACKBURN, JOHN  H:2
SCENT OF SHADOWS, THE [ZODIAC 1]  PETTERSON, VICKI  P
SCEPTERS [COREAN 3]  MODESITT, L. E. JR.  H
SCEPTER’S RETURN, THE [@SCEPTER OF MERCY #3]  CHERNENKO, DAN  LP
SCHACH DER FINSTERNIS [@PERRY RHODAN SILBERBANDE #73]  PABEL MOEWIG KG  FPR
SCHATTEN DES KRISTALLENEN TODES, DIE [@PERRY RHODAN #32]  KNEIFEL, HANS  FPR
SCHATZKAMMER DER STERNE [@PERRY RHODAN #3]  BRAND, KURT  FPR
SCHEME OF THINGS, THE  DEL REY, LESTER  P
SCHEMES OF DRAGONS, THE [TANAGARAN 2]  SMEDS, DAVE  P
SCHILD’S LADDER  EGAN, GREG  H:2
SCHIMMELHORN FILE: MEMOIRS OF A DIRTY OLD GENIUS, THE  BRETNOR, REGINALD  C/P:2,P
SCHIMMELHORN’S GOLD  BRETNOR, REGINALD  C/P:2,SFWA-P,P
SCHISM [@TRIAD #1]  ASARO, CATHERINE  H

The MITSFS Titledex  September 24, 2007  Page 308
TITLE  AUTHOR  CODE
SCHISMATRIX  STERLING, BRUCE  C/P,SFWA,P,H
SCHIZOGENIC MAN, THE  HARRIS, RAYMOND  P
SCHOLAR OF DECAY [@RAVENLOFT]  HUFF, TANYA  P
SCHOLAR OF MAGICS, A [COLLEGE OF MAGICS 2]  STEVERMER, CAROLINE  H
SCHOLARS AND SOLDIERS  GENTLE, MARY  P
SCHOLARS OF NIGHT, THE  FORD, JOHN M.  P,SFWA-P
SCHOOL FOR CRUSOES, A  VERNE, JULES  H
SCHOOL OF DARKNESS, THE [JOHN THUNSTONE]  WELLMAN, MANLY WADE  H,SFWA
SCHOOL OF FIRE [@STARFIST 2]  SHERMAN, DAVID/CRAGG, DAN  P
SCHOOL OF LIGHT  NYE, JODY LYNN  P
SCHRODINGER’S CAT TRILOGY [SCHRODINGER’S CAT 1,2,3]  WILSON, ROBERT ANTON  H,LP
SCHRODINGER’S KITTEN  EFFINGER, GEORGE ALEC  B
SCHWERTER UND EISZAUBER  LEIBER, FRITZ  FP
SCI-FI NOW  FRANK, ALAN  R/REF-H

SCI-FI PRIVATE EYE WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA

SCI FI TV: FROM THE TWILIGHT ZONE TO DEEP SPACE NINE  VAN HISE, JAMES  R/REF-P
SCIENCE AGAINST MAN  CHEETHAM, ANTHONY  PA,C/PA
SCIENCE AND SORCERY  FORD, GARRET  PA
SCIENCE BOOK OF SPACE TRAVEL, THE  GOODWIN, HAROLD LELAND  [C/RAN-P:2]

SCIENCE FICTION ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA

SCIENCE FICTION BRODKIN, SYLVIA Z./

PEARSON, ELIZABETH J.

HA

SCIENCE FICTION  ROSE, MARK  R/REF-P
SCIENCE FICTION: A HISTORICAL ANTHOLOGY  RABKIN, ERIC S.  LPA
SCIENCE FICTION ADVENTURES IN DIMENSION  CONKLIN, GROFF  PA,C/PA
SCIENCE FICTION ADVENTURES IN MUTATION  CONKLIN, GROFF  C/PA,PA
SCIENCE FICTION: AN ILLUSTRATED HISTORY  LUNDWALL, SAM J.  R/REF-H
SCIENCE FICTION: AN INTRODUCTION  ALLEN, L. DAVID  R/REF-H,C/REF-H

SCIENCE FICTION AND FANTASY LITERATURE 1:

INDEXES TO THE LITERATURE REGINALD, R.

R/IND

SCIENCE FICTION AND FANTASY LITERATURE 2:

CONTEMPORARY SCIENCE FICTION AUTHORS REGINALD, R.

R/IND

SCIENCE FICTION AND FANTASY LITERATURE, 1975–1991  REGINALD, ROBERT  R/IND
SCIENCE FICTION AND FANTASY PSEUDONYMS  MCGHAN, BARRY  R/IND
SCIENCE FICTION AND HEROIC FANTASY AUTHOR INDEX, THE  WELLS, STUART W. III  R/IND
SCIENCE FICTION AND THE NEW DARK AGE  BERGER, HAROLD L.  R/REF-P
SCIENCE FICTION ARGOSY, A  KNIGHT, DAMON  HA
SCIENCE FICTION ART  FOSS, CHRIS  A
SCIENCE FICTION AT LARGE  NICHOLLS, PETER  R/REF-H,C/REF-H
SCIENCE FICTION AUDIENCES  TULLOCH, JOHN/JENKINS, HENRY  R/REF-H
SCIENCE FICTION AUTHORS (PLAYING CARDS)  WHITEHALL LINE  WRD
SCIENCE FICTION: AUTORES SELECIONADOS  VON SCHMIDT, CARLOS  FH
SCIENCE FICTION BEFORE 1900  ALKON, PAUL K.  R/REF-H
SCIENCE FICTION BESTIARY, THE  SILVERBERG, ROBERT  HA,PA
SCIENCE FICTION BOOK, THE  ROTTENSTEINER, FRANZ  R/REF-H
SCIENCE FICTION BOOK REVIEW INDEX 1971  HALL, HAL W.  R/IND
SCIENCE FICTION BOOK REVIEW INDEX 1972 VOLUME 3  HALL, HAL W.  R/IND
SCIENCE FICTION BOOK REVIEW INDEX 1980 VOLUME 11  HALL, HAL W.  R/IND
SCIENCE FICTION BOOKS PUBLISHED IN 1967  BURGER, JOANNE  R/IND
SCIENCE FICTION BY ASIMOV  ASIMOV, ISAAC  SR-P

SCIENCE FICTION BY GASLIGHT: A HISTORY AND ANTHOLOGY OF SCIENCE FICTION IN THE POPULAR MAGAZINES, 1891–1911 MOSKOWITZ, SAM

HA

SCIENCE FICTION BY THE RIVALS OF H. G. WELLS  RUSSELL, ALAN  HA
SCIENCE FICTION CALENDAR 1980  DOUBLEDAY BOOKS  SFWA-VLP
SCIENCE-FICTION CARNIVAL  BROWN, FREDRIC/REYNOLDS, MACK  PA,C/PA
SCIENCE FICTION CENTURY, THE  HARTWELL, DAVID G.  HA
SCIENCE-FICTION CLASSICS  ACKERMAN, FORREST J.  LPA
SCIENCE FICTION CLASSICS [BARSOOM 4,5,6/PELLUCIDAR 2,3]  BURROUGHS, EDGAR RICE  H

SCIENCE FICTION COLLECTIONS IN THE GEORGE ARENTS RESEARCH LIBRARY AT SYRACUSE UNIVERSITY MOONEY, PHILIP F.

R/REF-H

SCIENCE FICTION: CONTEMPORARY MYTHOLOGY WARRICK, PATRICIA S./

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

HA,SFWA

SCIENCE FICTION CRITICISM: AN ANNOTATED CHECKLIST  CLARESON, THOMAS D.  R/IND
SCIENCE FICTION DIALOGUES  WOLFE, GARY  R/REF-P
SCIENCE FICTION DISCOVERIES  POHL, CAROL/POHL, FREDERIK  PA,C/PA
SCIENCE FICTION EMPHASIS #1  GERROLD, DAVID  PA,C/PA
SCIENCE FICTION ENCYCLOPEDIA, THE  NICHOLLS, PETER  R/REF-H,C/REF-H

SCIENCE FICTION, FANTASY, & WEIRD FICTION MAGAZINE INDEX (1890–2003) (CD)

MILLER, STEPHEN T./ CONTENTO, WILLIAM G.

SR-MM

SCIENCE FICTION, FANTASY, AND HORROR: 1984 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION, FANTASY, AND HORROR: 1986 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

The MITSFS Titledex  September 24, 2007  Page 309
TITLE  AUTHOR  CODE

SCIENCE FICTION, FANTASY, AND HORROR: 1987 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION, FANTASY, AND HORROR: 1988 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION, FANTASY, AND HORROR: 1989 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION, FANTASY, AND HORROR: 1990 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION, FANTASY, AND HORROR: 1991 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION FILM  GIFFORD, DENIS  R/REF-P
SCIENCE FICTION FIRST EDITIONS  LOCKE, GEORGE  R/IND
SCIENCE FICTION FOR PEOPLE WHO HATE SCIENCE FICTION  CARR, TERRY  HA,PA
SCIENCE FICTION FROM CHINA  WU, DINGBO/MURPHY, PATRICK D.  HA
SCIENCE FICTION GALAXY, THE  CONKLIN, GROFF  SR-PA
SCIENCE FICTION HALL OF FAME  SILVERBERG, ROBERT  C/PA,PA
SCIENCE FICTION HALL OF FAME 2A, THE  BOVA, BEN  SFWA,PA,HA
SCIENCE FICTION HALL OF FAME 2B, THE  BOVA, BEN  C/PA,PA,SFWA
SCIENCE FICTION HALL OF FAME 3, THE  CLARKE, ARTHUR C./PROCTOR, GEORGE W. HA
SCIENCE FICTION HALL OF FAME 4, THE CARR, TERRY HA,PA

SCIENCE FICTION HANDBOOK (REVISED)

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK

R/REF-H,C/REF-H

SCIENCE FICTION HANDBOOK FOR READERS AND WRITERS  ELRICK, GEORGE S.  R/REF-H
SCIENCE FICTION: HISTORY — SCIENCE — VISION  SCHOLES, ROBERT/RABKIN, ERIC S.  R/REF-P,C/REF-H

SCIENCE FICTION IN OLD SAN FRANCISCO:

HISTORY OF THE MOVEMENT 1854–1890 MOSKOWITZ, SAM

R/REF-H

SCIENCE FICTION IN PRINT: 1985 BROWN, CHARLES N./

CONTENTO, WILLIAM G.

R/IND

SCIENCE FICTION IN THE CINEMA  BAXTER, JOHN  [C/RAN-P]
SCIENCE FICTION INVENTIONS  KNIGHT, DAMON  C/PA,PA
SCIENCE FICTION MAGAZINE CHECKLIST 1  JOHNSON, KENNETH R.  R/IND
SCIENCE FICTION MAGAZINE CHECKLIST 2  JOHNSON, KENNETH R.  R/IND
SCIENCE-FICTION MAGAZINE CHECKLIST 1961–1972, THE  DESMOND, WILLIAM H.  R/IND:3
SCIENCE FICTION MASTERPIECES  ASIMOV, ISAAC  HA
SCIENCE FICTION NOVEL, THE  ADVENT BOOKS  R/REF-H
SCIENCE FICTION ODDITIES  CONKLIN, GROFF  PA,C/PA
SCIENCE FICTION OF EDGAR ALLAN POE, THE  POE, EDGAR ALLAN  C/P,P
SCIENCE FICTION OF ISAAC ASIMOV, THE  PATROUCH, JOSEPH F. JR.  R/REF-H,C/REF-H
SCIENCE FICTION OF MARK CLIFTON, THE  CLIFTON, MARK  H

SCIENCE FICTION OF THE 40’S POHL, FREDERIK/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

LPA

SCIENCE FICTION OF THE 50’S GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

HA,LPA

SCIENCE FICTION OF THE THIRTIES  KNIGHT, DAMON  HA
SCIENCE FICTION OMNIBUS  CONKLIN, GROFF  SR-PA,C/PA
SCIENCE FICTION ORIGINS  NOLAN, WILLIAM F./GREENBERG, MARTIN H. PA
SCIENCE FICTION QUIZBOOK, THE  LAST, MARTIN/SEARLES, BAIRD  [C/RAN-P:2]
SCIENCE FICTION READER, A  HARRISON, HARRY/PUGNER, CAROL  LPA
SCIENCE FICTION READER’S GUIDE  ALLEN, L. DAVID  R/REF-P,C/REF-P:2
SCIENCE FICTION ROLL OF HONOR, THE  POHL, FREDERIK  HA,SFWA
SCIENCE FICTION SHOWCASE  KORNBLUTH, MARY  HA,PA
SCIENCE FICTION STORIES  BLISHEN, EDWARD  LPA
SCIENCE FICTION STORIES  BOARDMAN, TOM JR.  HA
SCIENCE FICTION STORIES  ELAM, RICHARD M. JR.  C/P:2,LP
SCIENCE FICTION STORIES  YOUTH PUBLICATIONS  SR-VLPA
SCIENCE FICTION STORIES 4  SPIEGL, WALTER  FP
SCIENCE FICTION STORIES 23  SPIEGL, WALTER  FP
SCIENCE FICTION STORY-READER 4 (GERMAN)  FRANKE, HERBERT W.  FP
SCIENCE FICTION STUDIES IN FILM  POHL, FREDERIK/POHL, FREDERIK IV  R/REF-H
SCIENCE-FICTION SUBTREASURY, THE  TUCKER, WILSON  H
SCIENCE FICTION TALES  ELWOOD, ROGER  SR-VLHA
SCIENCE FICTION TERROR TALES  CONKLIN, GROFF  C/PA,PA
SCIENCE FICTION THE BEST OF 2001  SILVERBERG, ROBERT/HABER, KAREN  C/PA,PA
SCIENCE FICTION THE BEST OF 2002  SILVERBERG, ROBERT/HABER, KAREN  PA
SCIENCE FICTION: THE BEST OF THE YEAR 2007  HORTON, RICH  LPA
SCIENCE FICTION: THE CLASSROOM IN ORBIT  FRIEND, BEVERLY  R/REF-P
SCIENCE-FICTION THE EARLY YEARS  BLEILER, EVERETT F.  R/IND
SCIENCE FICTION: THE FUTURE  ALLEN, DICK  LP
SCIENCE-FICTION THE GERNSBACK YEARS  BLEILER, EVERETT F.  R/REF-H
SCIENCE FICTION: THE GREAT YEARS  POHL, CAROL/POHL, FREDERIK  C/PA,PA
SCIENCE FICTION: THE GREAT YEARS II  POHL, CAROL/POHL, FREDERIK  C/PA,PA
SCIENCE FICTION: THE ILLUSTRATED ENCYCLOPEDIA  CLUTE, JOHN  R/REF-H

SCIENCE FICTION:

THE SCIENCE FICTION RESEARCH ASSOCIATION ANTHOLOGY WARRICK, PATRICIA S./WAUGH, CHARLES G./

GREENBERG, MARTIN H.

LPA

SCIENCE FICTION THE BEST OF 2003  HABER, KAREN/STRAHAN, JONATHAN  PA

The MITSFS Titledex  September 24, 2007  Page 310
TITLE  AUTHOR  CODE
SCIENCE FICTION THE BEST OF 2004  HABER, KAREN/STRAHAN, JONATHAN  PA
SCIENCE FICTION: THE VERY BEST OF 2005  STRAHAN, JONATHAN  LPA
SCIENCE FICTION THINKING MACHINES  CONKLIN, GROFF  PA,C/PA
SCIENCE FICTION THROUGH THE AGES 2  EVANS, I. O.  SR-PA
SCIENCE FICTION TITLE CHANGES  VIGGIANO, MICHAEL/FRANSON, DONALD  R/IND
SCIENCE FICTION, TODAY AND TOMORROW  BRETNOR, REGINALD  SFWA,R/REF-P,C/REF-H
SCIENCE FICTION VOICES #4  ELLIOT, JEFFREY M.  R/REF-P

SCIENCE FICTION WEIGHT LOSS BOOK, THE ASIMOV, ISAAC/MARTIN, GEORGE R. R./

GREENBERG, MARTIN H.

HA

SCIENCE FICTION: WHAT IT’S ALL ABOUT  LUNDWALL, SAM J.  R/REF-P,C/REF-P:2
SCIENCE FICTION WRITER’S WORKSHOP-1  LONGYEAR, BARRY B.  R/REF-P
SCIENCE FICTION YEARBOOK, THE  POURNELLE, JERRY  HA

SCIENCE FICTIONAL DINOSAUR, THE SILVERBERG, ROBERT/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA

SCIENCE FICTIONAL OLYMPICS, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

SCIENCE FICTIONAL SOLAR SYSTEM, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,SFWA

SCIENCE MADE STUPID  WELLER, TOM  VLP
SCIENCE METROPOLIS  STATTEN, VARGO  SR-P
SCIENCE, NUMBERS, AND I  ASIMOV, ISAAC  [C/RAN-P:2]

SCIENCE OF DISCWORLD, THE (REVISED) [SCIENCE OF DISCWORLD 1/ DISCWORLD] PRATCHETT, TERRY/STEWART, IAN/

COHEN, JACK

P

SCIENCE OF LIFE, THE  WELLS, H. G./HUXLEY, JULIAN/WELLS, G. P. [C/RAN-H]
SCIENCE OF POWER, THE [@NIGHT-THREADS 6] EMERSON, RU C/P,P

SCIENCE OF THE HITCHHIKER’S GUIDE TO THE GALAXY, THE [HITCHHIKERS] HANLON, MICHAEL

R/REF-H

SCIENCE PAST — SCIENCE FUTURE  ASIMOV, ISAAC  SFWA
SCIENCE POLITIQUE ET SCIENCE-FICTION (FRENCH)  HOTTERBEEX, MARCEL  R/REF-H
SCIENCE PROJECTS HANDBOOK  MOORE, SHIRLEY  [C/RAN-P]
SCIENCE SERVICE SCIENCE PROGRAM (4 BOOKLETS, BOXED)  NELSON DOUBLEDAY, INC.  [C/RAN-H]

SCIENTIFIC RESEARCH ABSTRACTS IN REPUBLIC OF CHINA I: ABSTRACTS SCIENCE AND TECHNOLOGY INFORMATION CENTER, NATIONAL SCIENCE COUNCIL, REPUBLIC OF CHINA

[C/RAN-H]

SCIENTIFIC RESEARCH ABSTRACTS IN REPUBLIC OF CHINA II: INDEXES SCIENCE AND TECHNOLOGY INFORMATION CENTER, NATIONAL SCIENCE COUNCIL, REPUBLIC OF CHINA

[C/RAN-H]

SCIENTIFIC ROMANCE, A  WRIGHT, RONALD  H,LP
SCIENTIFIC STUDY OF UNIDENTIFIED FLYING OBJECTS  CONDON, EDWARD U.  [C/RAN-P]
SCIENTIST SPECULATES, THE  GOOD, I. J.  [C/RAN-H]
SCIMITAR [HIGHLANDER]  MCCONNELL, ASHLEY  P
SCION OF CYADOR [RECLUCE 11]  MODESITT, L. E. JR.  H:2
SCION OF THE SERPENT [@AGE OF CONAN ANOK #1]  YORK, J. STEVEN  P
SCION’S LADY [LADY IN GIL 2]  BRADLEY, REBECCA  P
SCIONS OF SHANNARA, THE [SHANNARA HERITAGE 1]  BROOKS, TERRY  H:2,P
SCISSORS CUT PAPER WRAP STONE  MCDONALD, IAN  C/P,P
SCOP  MALZBERG, BARRY N.  C/P,P
SCOR-STING  JAVOR, FRANK A.  P
SCORPIANNE  DEVENPORT, EMILY  C/P,P
SCORPIO [SCORPIO 1]  MCDONOUGH, ALEX  C/P:2,P
SCORPIO ATTACK [BLAKE’S 7]  HOYLE, TREVOR  P
SCORPIO DESCENDING [SCORPIO 3]  MCDONOUGH, ALEX  C/P,P
SCORPIO RISING [SCORPIO 2]  MCDONOUGH, ALEX  C/P,P
SCORPION  LINAKER, MICHAEL R.  P
SCORPION, THE [FIVE RINGS CLAN WAR 1]  SULLIVAN, STEPHEN D.  P
SCORPION JAR, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  HARDY, JASON M.  P
SCORPION SHARDS [STAR SHARDS 1]  SHUSTERMAN, NEAL  P
SCORPIUS [JAMES BOND 21]  GARDNER, JOHN (BRITISH, 1926-)  P
SCOT, THE WITCH AND THE WARDROBE, THE  BLAIR, ANNETTE  P
SCOTLAND THE BRAVE [@HIGHLANDER]  ROBERSON, JENNIFER  P
SCOUNDREL WORLDS, THE [STAR RISK,LTD. 2]  BUNCH, CHRIS  P
SCOURGE [@MAGIC: THE GATHERING ONSLAUGHT 3]  KING, J. ROBERT  P
SCOURGE OF SCREAMERS, A  GALOUYE, DANIEL F.  C/P:2,P
SCOURGE OF THE STEEL MASK [@G-8 #1]  HOGAN, ROBERT J.  P
SCOUT’S PROGRESS [@LIADEN UNIVERSE]  LEE, SHARON/MILLER, STEVE  P
SCRATCH MONKEY  STROSS, CHARLES  H
SCREAM FOR JEEVES  CANNON, P. H.  H,SR-LP
SCREAM OF STONE [@FORGOTTEN REALMS WATERCOURSE #3]  ATHANS, PHILIP  P
SCREAM OF THE SHALKA [@DOCTOR WHO NINTH (ANIMATED) DOCTOR]  CORNELL, PAUL  P
SCREAMING ACROSS THE SKY, A [@GAMMALAW 2]  DALEY, BRIAN  C/P,P
SCREAMING FACE, THE  LYMINGTON, JOHN  P
SCREAMING SCIENCE FICTION  LUMLEY, BRIAN  H
SCREEN  MALZBERG, BARRY N.  P
SCREWTAPE LETTERS, THE  LEWIS, C. S.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 311
TITLE  AUTHOR  CODE
SCREWTOP  MCINTYRE, VONDA N.  TD7,SFWA-TD7
SCRIBBLINGS  DE CAMP, L. SPRAGUE  SR-H,SR-L
SCROLL OF LUCIFER, THE [LAND BEYOND THE GATE 4]  ESHBACH, LLOYD ARTHUR  C/P:2,P
SCROLL OF MAN, THE  DALMAS, JOHN  P,SFWA-P
SCROLLS OF THE ANCIENTS, THE [BLOOD AND STONE 3]  NEWCOMB, ROBERT  H,P
SCRYER  GRAY, LINDA CROCKETT  P,SFWA-P
SCRYING GAME, THE  HARMAN, ANDREW  P
’SCUSE ME WHILE I KISS THIS GUY  EDWARDS, GAVIN  [C/RAN-P]
SEA ANGEL, THE [DOC SAVAGE #49]  ROBESON, KENNETH  C/P:2,@P
SEA BEASTS, THE  CHANDLER, A. BERTRAM  P
SEA CHANGE  HUNT, BARBARA  H

SEA DEVIL’S EYE, THE [@FORGOTTEN REALMS THREAT FROM THE SEA #3] ODOM, MEL

P

SEA DRAGON HEIR [MAGRAVANDIAS 1]  CONSTANTINE, STORM  H:2
SEA GOLD [@RICK BRANT #3]  BLAINE, JOHN  H
SEA HAG, THE [CRYSTAL WALLS 1]  DRAKE, DAVID  C/P,SFWA-P,P
SEA HORSE IN THE SKY  COOPER, EDMUND  H:2,P
SEA IS BOILING HOT, THE  BAMBER, GEORGE  C/P:2,P
SEA IS FULL OF STARS, THE [WELL WORLD 6]  CHALKER, JACK L.  C/P,P
SEA-KING’S DAUGHTER, THE  MICHAELS, BARBARA  P
SEA LARK’S SONG, THE [BRIERLY 2]  MARCELLAS, DIANA  P
SEA MAGIC  HARRIS, ROSEMARY  H
SEA MAGICIAN, THE [DOC SAVAGE #44]  ROBESON, KENNETH  C/P:2,@P
SEA OF DEATH [GORD THE ROGUE]  GYGAX, GARY  C/P:2,P
SEA OF GLASS  LONGYEAR, BARRY B.  H:2,SFWA,P
SEA OF MIST, THE [MIGHT AND MAGIC]  ODOM, MEL  P
SEA OF PLAGUE [@OUTLANDERS HEART OF THE WORLD 2]  AXLER, JAMES  P
SEA OF SILENCE, THE [WINDAMEIR CIRCLE 2]  HANCOCK, NIEL  P
SEA OF SILVER LIGHT [@OTHERLAND #4]  WILLIAMS, TAD  H,P
SEA OF SORROWS [@SUN SWORD #4]  WEST, MICHELLE  P
SEA OF SPACE, A  NOLAN, WILLIAM F.  PA,C/PA
SEA OF SWORDS [@FORGOTTEN REALMS]  SALVATORE, R. A.  P
SEA OF TIME [CRETACEOUS SEA 2]  HUBBELL, WILL  P
SEA PEOPLE, THE  LUKENS, ADAM  H
SEA SERPENT, THE  VERNE, JULES  H
SEA SIEGE  NORTON, ANDRE  DF147,P,C/P:2
SEA STALLION [@AFRIKORPS #4]  DOLAN, BILL  P
SEA STAR, THE [WESTRIA 5]  PAXSON, DIANA L.  P,SFWA-P
SEA SWORD, THE  MARTINE-BARNES, ADRIENNE  C/P,P
SEA WITHOUT A SHORE [MOONTIDE AND MAGIC RISE 2]  RUSSELL, SEAN  P
SEADEMONS  YEP, LAURENCE  C/P,SFWA,P,H
SEADRAGON’S DAUGHTER, THE [DRAGON DELASANGRE 3]  TROOP, ALAN F.  P
SEAFORT’S HOPE [NICHOLAS SEAFORT 1,2]  FEINTUCH, DAVID  H
SEAGULLS UNDER GLASS  TATE, PETER  H
SEAHARP HOTEL, THE [GREYSTONE BAY 3]  GRANT, CHARLES L.  SFWA-P
SEANCE FOR A VAMPIRE [DRACULA]  SABERHAGEN, FRED  H,P
SEANCE FOR SUSAN  CLARK, LYDIA BENSON  P
SEAQUEST DSV: THE NOVEL [@SEAQUEST DSV 1]  DUANE, DIANE/MORWOOD, PETER  P
SEARCH, THE [STAR TREK DS9]  CAREY, DIANE  C/P,P
SEARCH  WEVERKA, ROBERT  P
SEARCH AND DESTROY [@FREEDOM’S RANGERS #3]  ANDREWS, KEITH WILLIAM  P
SEARCH AND RESCUE [QUANTUM LEAP]  CRANDALL, MELISSA  C/P,P
SEARCH FOR A DEAD NYMPHO  FAIRMAN, PAUL W.  P
SEARCH FOR BRIDEY MURPHY, THE  BERNSTEIN, MOREY  [C/RAN-P]
SEARCH FOR DINOSAURS [@TIME MACHINE #2]  BISCHOFF, DAVID F.  P
SEARCH FOR FIERRA, THE [@EMPYRION #1]  LAWHEAD, STEPHEN R.  LP
SEARCH FOR JOSEPH TULLY, THE  HALLAHAN, WILLIAM H.  P
SEARCH FOR KA, THE [GANDALARA 5]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P,P
SEARCH FOR LIFE ON MARS, THE  COOPER, HENRY S. F. JR.  [C/RAN-H]
SEARCH FOR MAGIC, THE [DRAGONLANCE WAR OF SOULS]  WEIS, MARGARET/HICKMAN, TRACY  PA

SEARCH FOR MORAG, THE CAMPBELL, ELIZABETH MONTGOMERY/

SOLOMON, DAVID

SFWA

SEARCH FOR PLANET X, THE  SIMON, TOBY  [C/RAN-P]
SEARCH FOR POWER, THE [@DRAGONLANCE]  WEIS, MARGARET  PA
SEARCH FOR THE DOCTOR [DOCTOR WHO]  MARTIN, DAVID  [C/RAN-P]
SEARCH FOR THE STARBLADE [DANANS 3]  TAYLOR, KEITH  P
SEARCH FOR THE SUN [CAGEWORLD 1]  KAPP, COLIN  C/P:2,P
SEARCH FOR ZEI, THE [KRISHNA]  DE CAMP, L. SPRAGUE  DF249,H
SEARCH OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 3]  TEPPER, SHERI S.  C/P:2,P
SEARCH THE SKY  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
SEARCH THE SKY (REVISED)  POHL, FREDERIK/KORNBLUTH, C. M.  C/P,P
SEARCHER [@FRONTIER EARTH 2]  BOXLEITNER, BRUCE  P
SEARCHING FOR DRAGONS [ENCHANTED FOREST 2]  WREDE, PATRICIA C.  P
SEARING, THE  COYNE, JOHN  H

The MITSFS Titledex  September 24, 2007  Page 312
TITLE  AUTHOR  CODE
SEAROAD  LE GUIN, URSULA K.  H:2,LP
SEAS OF ERNATHE  CARVER, JEFFREY A.  SR-P
SEASERPENTS!  DANN, JACK/DOZOIS, GARDNER  PA
SEASON FOR SLAUGHTER, A [WAR AGAINST THE CHTORR 4]  GERROLD, DAVID  C/P,P
SEASON OF MISTS [@SANDMAN #4]  GAIMAN, NEIL  H,CX
SEASON OF PASSAGE, THE  PIKE, CHRISTOPHER  H
SEASON OF SACRIFICE  KLASKY, MINDY L.  P
SEASON OF SHADOWS [@SUMMERLANDS #1]  FOXXE, ELLEN  P
SEASON OF STORMS [@SUMMERLANDS #2]  FOXXE, ELLEN  P
SEASON OF THE SPELLSONG [SPELLSINGER 1,2,3]  FOSTER, ALAN DEAN  H:2
SEASONS IN FLIGHT  ALDISS, BRIAN W.  H
SEASONS OF PLENTY [PLENTY 2]  GREENLAND, COLIN  C/P,P
SEAWARD  COOPER, SUSAN  C/P,SFWA,P
SEBASTIAN [EPHEMERA]  BISHOP, ANNE  H
SEBASTIAN OF MARS [CATS OF MARS 2]  SARRANTONIO, AL  P
SECOND ANGEL, THE  KERR, PHILIP  H
SECOND ATLANTIS, THE  WILLIAMS, ROBERT MOORE  C/P:2,P
SECOND BODY  PAYER, SUE  P
SECOND BOOK OF FRITZ LEIBER, THE  LEIBER, FRITZ  C/P,P
SECOND BOOK OF ROBERT E. HOWARD, THE  HOWARD, ROBERT E.  C/P,P
SECOND BOOK OF SWORDS, THE [@SWORDS #2]  SABERHAGEN, FRED  C/P,P,LP,H
SECOND BOOK OF THE COSMIC FORCES OF MU, THE  CHURCHWARD, JAMES  [C/RAN-P]
SECOND CHANCE AT EDEN, A  HAMILTON, PETER F.  P
SECOND CHANCES  SHWARTZ, SUSAN  H
SECOND CHILD  SAUL, JOHN  P
2ND COMING, THE  MCFADDEN, EDWARD J. III  LPA
SECOND COMING OF LUCAS BROKAW, THE  BRAUN, MATTHEW  P
SECOND CONTACT  AUSTIN, J. D.  P
SECOND CONTACT  RESNICK, MIKE  C/P:2,SFWA-P,P
SECOND CONTACT [@STARSEA INVADERS 2]  STINE, G. HARRY  C/P,P
SECOND DEATH  ZUROY, MICHAEL  SFWA
SECOND DELUGE, THE  SERVISS, GARRETT P.  LP

SECOND DOCTOR, THE [DOCTOR WHO] HOWE, DAVID J./STAMMERS, MARK/

WALKER, STEPHEN JAMES

R/REF-P

SECOND EARTH, THE  WOODROFFE, PATRICK  A
SECOND EMPIRE, THE [@MONARCHIES OF GOD #4]  KEARNEY, PAUL  P
SECOND ENDING  WHITE, JAMES  DF173
SECOND EXPERIMENT, THE  JEPPSON, J. O.  P
SECOND FIRE [@LOST MILLENNIUM 2]  MOSCOE, MIKE  C/P,P
SECOND FOUNDATION [FOUNDATION 3]  ASIMOV, ISAAC  C/P:2,P,H:3,SR-P,SR-H
SECOND GAME  DE VET, CHARLES V./MACLEAN, KATHERINE C/P,P
SECOND GENERATION, THE [DRAGONLANCE]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H
SECOND GENESIS [GENESIS 2]  MOFFITT, DONALD  C/P:2,P
SECOND GHOST BOOK, THE  ASQUITH, CYNTHIA  SR-PA
SECOND INFINITY, A  WEAVER, MICHAEL D.  P
SECOND KINGDOM, THE [@VISION OF BEASTS #2]  LOVEJOY, JACK  C/P,SFWA-P,P
SECOND LIFE OF CECILY PRIDE  DALTON, CLAIRE  P
SECOND NATURE  WILDER, CHERRY  P
SECOND OMNI BOOK OF SCIENCE FICTION, THE  DATLOW, ELLEN  PA
SECOND PACIFIC BOOK OF SCIENCE FICTION, THE  BAXTER, JOHN  HA
SECOND SATELLITE  RICHARDSON, ROBERT  H
SECOND SIGHT  WILLIAMS, DAVID  P
SECOND SON, THE  SAILOR, CHARLES  H
SECOND STAGE LENSMEN [LENSMEN 5]  SMITH, E. E.  C/P,P,H
SECOND STAR  STABENOW, DANA  P
SECOND SUMMONING, THE [KEEPER’S CHRONICLES 2]  HUFF, TANYA  C/P:2,P
SECOND TIME AROUND [@TIMESHARE 2]  DANN, JOSHUA  P
SECOND TIME AROUND [TIME LOTTERY 2]  MOSER, NANCY  LP
SECOND TOM HOLT OMNIBUS, THE  HOLT, TOM  H
SECOND TRIAD, THE [@LOST SWORDS #4,5,6]  SABERHAGEN, FRED  H:2
SECOND TRIP, THE  SILVERBERG, ROBERT  H:2,P
SECOND VARIETY  DICK, PHILIP K.  B
SECOND WAR OF THE WORLDS, THE [ARWIN 4]  SMITH, GEORGE H.  P

SECOND WAVE [ACORNA’S CHILDREN 2] MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

H

SECOND WORLD OF IF, THE  QUINN, JAMES/WULFF, EVE  SR-PA
SECOND YEAR ASHORE  VERNE, JULES  H
SECONDE FONDATION [FOUNDATION 3]  ASIMOV, ISAAC  FP
SECONDS  ELY, DAVID  P
SECRET, THE  MALONE, ADRIAN  H
SECRET AGENT OF TERRA  BRUNNER, JOHN  DF133
SECRET ASCENSION, THE  BISHOP, MICHAEL  C/P,SFWA-P,P,H
SECRET ATLAS, A [AGE OF DISCOVERY 1]  STACKPOLE, MICHAEL A.  H
SECRET BARRIER X [@PERRY RHODAN #16]  ACE  C/PA,PA

The MITSFS Titledex  September 24, 2007  Page 313
TITLE  AUTHOR  CODE
SECRET BOOKS OF PARADYS I & II, THE [PARADYS 1,2]  LEE, TANITH  H:2
SECRET BOOKS OF PARADYS III & IV, THE [PARADYS 3,4]  LEE, TANITH  H
SECRET BOOKS OF VENUS I & II, THE [VENUS 1,2]  LEE, TANITH  H
SECRET BOOKS OF VENUS III & IV, THE [VENUS 3,4]  LEE, TANITH  H
SECRET CITY, THE  EMSHWILLER, CAROL  LP
SECRET COUNTRY, THE [SECRET COUNTRY 1]  DEAN, PAMELA C.  P
SECRET DES QUASARS, LE  BARBET, PIERRE  FP
SECRET GALACTICS, THE  VAN VOGT, A. E.  H,LP
SECRET HISTORY, A [@ASH #1]  GENTLE, MARY  C/P:2,P
SECRET HISTORY OF TIME TO COME, A  MACAULEY, ROBIE  H:2
SECRET IN THE SKY, THE [DOC SAVAGE #20]  ROBESON, KENNETH  C/P:2,@P
SECRET KINGDOM, THE  ROBERTS, MURRAY  B
SECRET LIFE  VANDERMEER, JEFF  H
SECRET LIFE OF LASZLO, COUNT DRACULA, THE  ANSCOMBE, RODERICK  H
SECRET LIFE OF PLANTS, THE  TOMPKINS, PETER/BIRD, CHRISTOPHER  [C/RAN-P]
SECRET MARTIANS, THE  SHARKEY, JACK  DD471
SECRET MASTERS, THE  KERSH, GERALD  H:2,P
SECRET MISSION: MOLUK [@PERRY RHODAN #84]  ACE  PA
SECRET OF BARNABAS COLLINS, THE  ROSS, MARILYN  P
SECRET OF BIGFOOT PASS, THE [BERKLEY SIX MILLION DOLLAR MAN 1]  JAHN, MICHAEL  P
SECRET OF DRAGONHOME, THE  PEEL, JOHN  P
SECRET OF HOLM PEEL, THE  ROHMER, SAX  P
SECRET OF LIFE, THE  MCAULEY, PAUL J.  H:2
SECRET OF LIFE, THE  RUCKER, RUDY  H
SECRET OF PLATFORM 13, THE  IBBOTSON, EVA  LP
SECRET OF SAREK, THE  LEBLANC, MAURICE  H
SECRET OF SATURN’S RINGS, THE  WOLLHEIM, DONALD A.  SR-H,H,P
SECRET OF SINHARAT, THE  BRACKETT, LEIGH  DM101
SECRET OF SPRING, THE  ANTHONY, PIERS/TAEUSCH, JO ANNE  H:2,P
SECRET OF THE BLACK PLANET  LESSER, MILTON  C/P:2,P
SECRET OF THE CRATER, THE  OSBORNE, DUFFIELD  H
SECRET OF THE ISLAND, THE  VERNE, JULES  H
SECRET OF THE KNIGHTS [@TIME MACHINE #1]  GASPERINI, JIM  C/P,P
SECRET OF THE LONA, THE [DR. BONES 1]  LEIGH, STEPHEN  C/P,P
SECRET OF THE LOST MESA, THE  STRIKER, FRAN  H
SECRET OF THE LOST RACE  NORTON, ANDRE  C/P:2,DD381,P
SECRET OF THE MARAUDER SATELLITE  WHITE, TED  H
SECRET OF THE MARTIAN MOONS, THE  WOLLHEIM, DONALD A.  H,SR-H,P
SECRET OF THE NINTH PLANET, THE  WOLLHEIM, DONALD A.  P,H,SR-H
SECRET OF THE RED SPOT  BINDER, EANDO  P
SECRET OF THE ROUND BEAST, THE [BESTIARY 2]  FORRESTER, JOHN  P
SECRET OF THE RUNESTAFF, THE [RUNESTAFF 4]  MOORCOCK, MICHAEL  P
SECRET OF THE SIXTH MAGIC  HARDY, LYNDON  C/P:2,P

SECRET OF THE SPIRITKEEPER, THE [@KNIGHTS OF THE SILVER DRAGON #1] FORBECK, MATT

LP

SECRET OF THE SUNLESS WORLD  CAPPS, CARROLL M.  C/P,P
SECRET OF THE TIME VAULT, THE [@PERRY RHODAN #6]  ACE  PA,C/PA
SECRET OF TWO-EDGE, THE [@COMPLETE ELFQUEST #6]  PINI, WENDY/PINI, RICHARD  CX
SECRET OF WILHELM STORITZ, THE  VERNE, JULES  H
SECRET OF ZI, THE  BULMER, KENNETH  DD331
SECRET PASSAGES  PREUSS, PAUL  H,P
SECRET PEOPLE, THE  HARRIS, JOHN BEYNON  C/P:2,P
SECRET PEOPLE, THE  JONES, RAYMOND F.  H
SECRET PEOPLE, THE  WYNDHAM, JOHN  C/P,P
SECRET PROPHECIES OF NOSTRADAMUS, THE  STERNAU, CYNTHIA/GREENBERG, MARTIN H. C/PA,PA
SECRET REALMS  COOL, TOM  H,P
SECRET SCORPIO [DRAY PRESCOT 15]  AKERS, ALAN BURT  P
SECRET SEA, THE  MONTELEONE, THOMAS F.  P
SECRET SHARER, THE  SILVERBERG, ROBERT  H
SECRET SONGS, THE  LEIBER, FRITZ  SR-P
SECRET SORCERESS  LOWERY, LINDA  P
SECRET UNDER THE SEA  DICKSON, GORDON R.  LP
SECRET VISITORS, THE  WHITE, JAMES  DD237,P,C/P:2
SECRET WARS [MARVEL’S FINEST]  SHOOTER, JIM  CX
SECRETS OF DOC SAVAGE  MURRAY, WILL  R/REF-P,C/REF-H
SECRETS OF DOCTOR TAVERNER, THE  FORTUNE, DION  H
SECRETS OF MAGIC, THE [@MAGIC: THE GATHERING]  LEBOW, JESS  PA
SECRETS OF STARDEEP  JAKES, JOHN  P
SECRETS OF SYNCHRONICITY, THE  FAST, JONATHAN  C/P:2,P
SECRETS OF THE BERMUDA TRIANGLE  LANDSBURG, ALAN  [C/RAN-P]
SECRETS OF THE DEEP  DICKSON, GORDON R.  P
SECTION G: UNITED PLANETS  REYNOLDS, MACK  C/P:2,P
SECTOR GENERAL [SECTOR GENERAL 5]  WHITE, JAMES  C/P:2,P
SECULAR WIZARD, THE [WIZARD IN RHYME 4]  STASHEFF, CHRISTOPHER  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 314
TITLE  AUTHOR  CODE
SEDICI MAPPE DEL NOSTRO FUTURO  GALASSIA  FP
SEDUCED BY MOONLIGHT [MEREDITH GENTRY 3]  HAMILTON, LAURELL K.  H
SEDUCTIONS  GARTON, RAY  SR-L
SEE NO WEEVIL [@KING’S QUEST 3]  MORR, KENYON  P
SEED, THE  THOMAS, DAN  C/P:2,P
SEED OF EARTH, THE  SILVERBERG, ROBERT  P,DF145
SEED OF LIGHT  COOPER, EDMUND  C/P:3,P
SEED OF STARS  MORGAN, DAN/KIPPAX, JOHN  C/P:2,P
SEED OF THE DREAMERS  PETAJA, EMIL  D06707
SEED OF THE GODS  HUGHES, ZACH  C/P:2,P
SEED UPON THE WIND [TALISWOMAN 2]  DOUGLAS, CAROLE NELSON  H:2
SEEDBEARERS, THE  TIMLETT, PETER VALENTINE  P
SEEDING, THE  SHOBIN, DAVID  H
SEEDING THE CLOUDS (CHINESE)  XI TSUN TSAI  FP
SEEDLING STARS, THE  BLISH, JAMES  C/P:2,P,H
SEEDS OF BETRAYAL [WINDS OF THE FORELANDS 2]  COE, DAVID B.  P
SEEDS OF CHANGE  MONTELEONE, THOMAS F.  SR-P,C/P:2
SEEDS OF DEATH, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
SEEDS OF DESTINY [ORGANIC FUTURE]  EASTON, THOMAS A.  P
SEEDS OF LIFE  TAINE, JOHN  SR-H,C/P:2,GN3.13
SEEDS OF LIFE/WHITE LILY  TAINE, JOHN  H:2,LP
SEEDS OF RAGE [@STAR TREK ERRAND OF FURY #1]  RYAN, KEVIN  P

SEEDS OF RUIN [@PERRY RHODAN #111]/ PLANET MECHANICA [PERRY RHODAN 112] ACE

C/PA,PA

SEEDS OF TIME, THE  KENYON, KAY  C/P,P
SEEDS OF TIME, THE  WYNDHAM, JOHN  SR-P,C/P
SEEDS OF WAR [SEEDS OF WAR 1]  RANDLE, KEVIN/CORNETT, ROBERT  P

SEEING, THE MCGIVERN, WILLIAM P./ MCGIVERN, MAUREEN

P

SEEING I [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P:2,P
SEEING RED  SCHOW, DAVID J.  SFWA-P

SEEKER, THE BISCHOFF, DAVID F./

LAMPTON, CHRISTOPHER

SR-P,C/P

SEEKER, THE [@DARK SUN TRIBE OF ONE #2]  HAWKE, SIMON  P
SEEKER [ALEX BENEDICT 3]  MCDEVITT, JACK  H
SEEKER FROM THE STARS  COLEMAN, JAMES NELSON  C/P:2,P
SEEKERS, THE  DVORKIN, DAVID  H:2
SEEKERS AND THE SWORD, THE  FRIEDMAN, MICHAEL JAN  C/P:2,P
SEEKER’S MASK [JAMETHIEL DREAM-WEAVER 3]  HODGELL, P. C.  SR-H,H:2
SEEKERS OF SHAR-NUHN, THE  MAYHAR, ARDATH  H,SFWA,P
SEEKERS OF TOMORROW  MOSKOWITZ, SAM  R/REF-P,C/REF-P:2
SEEKING SWORD, THE  KANGILASKI, JAAN  C/P:2,P
SEEKING THE DREAM BROTHER  BENNETT, MARCIA J.  C/P:2,P
SEEKING THE MYTHICAL FUTURE [Q 1]  HOYLE, TREVOR  C/P,P
SEEKLIGHT  JETER, K. W.  SR-P
SEER, THE [ANTRIAN MESSENGER 2]  WISLER, G. CLIFTON  H
SEER AND THE SWORD, THE  HANLEY, VICTORIA  P
SEER KING, THE [DAMASTES 1]  BUNCH, CHRIS  C/P,P
SEERESS OF KELL [MALLOREON 5]  EDDINGS, DAVID  C/P,H:2,P
SEER’S BLOOD  DURGIN, DORANNA  P
SEETEE ALERT! [@CAP KENNEDY #6]  KERN, GREGORY  C/P,P
SEETEE SHIP  STEWART, WILL  SR-H
SEETEE SHIP  WILLIAMSON, JACK  P
SEETEE SHIP/SEETEE SHOCK  WILLIAMSON, JACK  C/P:2,P
SEETEE SHOCK  STEWART, WILL  H
SEETEE SHOCK  WILLIAMSON, JACK  P
SEG THE BOWMAN [DRAY PRESCOT 32]  AKERS, ALAN BURT  C/P,P
SELECTED LETTERS I: 1911–1924  LOVECRAFT, H. P.  R/REF-H
SELECTED LETTERS II: 1925–1929  LOVECRAFT, H. P.  R/REF-H
SELECTED LETTERS III: 1929–1931  LOVECRAFT, H. P.  R/REF-H
SELECTED LETTERS IV: 1932–1934  LOVECRAFT, H. P.  R/REF-H
SELECTED LETTERS V: 1934–1937  LOVECRAFT, H. P.  R/REF-H
SELECTED POEMS  SMITH, CLARK ASHTON  SR-H
SELECTED SCIENCE FICTION AND FANTASY STORIES  LONDON, JACK  H
SELECTED STORIES  STURGEON, THEODORE  LP
SELECTED TALES OF GUY DE MAUPASSANT  DE MAUPASSANT, GUY  SR-VLH
SELF-DISCOVERY  SAVCHENKO, VLADIMIR  LP
SELKIE, THE  SHEFFIELD, CHARLES/BISCHOFF, DAVID F.  H,P
SEMINATORI DI ABISSI, I (ITALIAN)  BRUSSOLO, SERGE  FP

SEMIOTEXT[E] SF 14 RUCKER, RUDY/WILSON, PETER LAMBORN/

WILSON, ROBERT ANTON

HA,SR-LPA

SEMPER MARS [@HERITAGE #1]  DOUGLAS, IAN  C/P,P
SEND HIM VICTORIOUS  HURD, DOUGLAS/OSMOND, ANDREW  P
SEND IN THE SWORDS [ENEMIES OF SOCIETY 4]  DAVID, JOHN  LP

The MITSFS Titledex  September 24, 2007  Page 315
TITLE  AUTHOR  CODE
SENDING OF DRAGONS, A [PIT DRAGONS 3]  YOLEN, JANE  H,P
SENSE OF SHADOW, A  WILHELM, KATE  C/P,P,H
SENSE OF WONDER, A  MOSKOWITZ, SAM  PA

SENSITIVE DEPENDENCE ON INITIAL CONDITIONS, A [@AUTHOR’S CHOICE MONTHLY #20] ROBINSON, KIM STANLEY

SR-LP

SENSITIVES, THE  BURKHOLZ, HERBERT  P
SENSITIVES, THE  CHARBONNEAU, LOUIS  P
SENSORITES, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P
SENSUOUS DIRTY OLD MAN, THE  "A", DOCTOR  [C/RAN-H,C/RAN-P]
SENSUOUS SCIENCE FICTION FROM THE WEIRD AND SPICY PULPS  JAFFERY, SHELDON  SR-LPA
SENTENCED TO PRISM [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
SENTIENCE [SENTIENCE 1]  ADAMS, TERRY A.  C/P:2,P
SENTINEL, THE  CLARKE, ARTHUR C.  H:2,P
SENTINEL, THE [@WORLD OF ADEN THUNDERSCAPE]  MCKEONE, DIXIE LEE  P
SENTINEL, THE [@CASCA #9]  SADLER, BARRY  P
SENTINEL STARS, THE  CHARBONNEAU, LOUIS  C/P:2,P
SENTINELS, THE [@SENTINELS #1,2,3/ROBOTECH 13,14,15]  MCKINNEY, JACK  P
SENTINELS [@STARSHIELD 2]  WEIS, MARGARET/HICKMAN, TRACY  H:2
SENTINELS FROM SPACE  RUSSELL, ERIC FRANK  C/P:2,P
SENTINELS OF SOLITUDE [PERRY RHODAN 127]  DARLTON, CLARK  SR-LP
SENTINELS OF SPACE  RUSSELL, ERIC FRANK  SR-DD044
SENTRY PEAK [WAR BETWEEN THE PROVINCES 1]  TURTLEDOVE, HARRY  H,P
SEPARATE REALITY, A  CASTANEDA, CARLOS  [C/RAN-P]
SEPARATE STAR, A  DRAKE, DAVID/MIESEL, SANDRA  C/PA,PA,SFWA-P
SEPARATE WAR AND OTHER STORIES, A  HALDEMAN, JOE  H
SEPOLTURA, LA  MONTANARI, GIANNI  FP
SEPT DIALOGUES DE BETES  WILLY, COLETTE  FP
SEPTIEME CONTINENT, LE  RAYJEAN, MAX-ANDRE  FP
SERAPHIM  HAUF, MICHELE  LP
SERAPHIM RISING, THE  DEVOS, ELISABETH  P
SEREN CENACLES, THE  NORWOOD, WARREN/MYLIUS, RALPH  P
SERENITY  DECANDIDO, KEITH R. A.  P
SERENITY [SERENITY]  WHEDON, JOSS  R/REF-H

SERENITY 1 [SERENITY 1] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY 2 [SERENITY 2] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY 3 [SERENITY 3] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

CX

SERENITY: THOSE LEFT BEHIND [SERENITY 1,2,3] WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA

H

SERLING  SANDER, GORDON F.  R/REF-H
SERPENT, THE [ATLAN]  GASKELL, JANE  C/P:2,P
SERPENT AND THE GRAIL, THE  ATTANASIO, A. A.  P
SERPENT AND THE ROSE, THE [WAR OF THE ROSE 1]  BRYAN, KATHLEEN  LP
SERPENT BRIDE, THE [DARKGLASS MOUNTAIN 1]  DOUGLASS, SARA  H
SERPENT BRIDE, THE (PROOF) [DARKGLASS MOUNTAIN 1]  DOUGLASS, SARA  SR-LP
SERPENT CATCH  WOLVERTON, DAVE  C/P:2,P
SERPENT MAGE, THE [MICHAEL PERRIN 2]  BEAR, GREG  C/P:2,P
SERPENT MAGE [@DEATH GATE #4]  WEIS, MARGARET/HICKMAN, TRACY  C/P,P,H
SERPENT WALTZ [@DANCER #2]  CLAYTON, JO  C/P:2,P
SERPENTS AMONG THE RUINS [@STAR TREK LOST ERA (2311 AD)]  GEORGE, DAVID R. III  P
SERPENT’S CIRCLE, THE  HARPUR, PATRICK  H
SERPENT’S EGG, THE [SERPENT’S EGG 1]  MCCURDY, J. FITSGERALD  LP
SERPENT’S EGG, THE  STEVERMER, CAROLINE  P
SERPENT’S GIFT [@STARBRIDGE #4]  CRISPIN, A. C./MARSHALL, DEBORAH A.  C/P,SFWA-P:2,P
SERPENT’S REACH  CHERRYH, C. J.  C/P,P,H
SERPENT’S SHADOW, THE [ELEMENTAL MASTERS 1]  LACKEY, MERCEDES  H
SERPENT’S SILVER [KELVIN HACKLEBERRY 2]  ANTHONY, PIERS/MARGROFF, ROBERT E.  H:2,P
SERPENT’S TOOTH, THE  PAXSON, DIANA L.  H,P
SERUM 2000  FREDD, STEPHEN  P
SERVANT OF THE DRAGON [LORD OF THE ISLES 3]  DRAKE, DAVID  C/P:3,P,H
SERVANT OF THE EMPIRE [MARA 2]  FEIST, RAYMOND E./WURTS, JANNY  H:2,P
SERVANT OF THE SHARD [@FORGOTTEN REALMS SELLSWORDS #1]  SALVATORE, R. A.  P
SERVANTS OF THE WANKH  VANCE, JACK  C/P:2,P

SERVICE FOR THE DEAD [@MECHWARRIOR DARK AGE PROVING GROUNDS 3/BATTLETECH] DELRIO, MARTIN

P

SERVICE OF THE SWORD, THE [HONOR HARRINGTON/ WORLDS OF HONOR 4] WEBER, DAVID

HA

SERVING IN TIME  EKLUND, GORDON  SR-P,C/P
SET OF WHEELS  THURSTON, ROBERT  P
SET PIECE [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P,P
SETH PAPERS, THE  LAURIA, FRANK  P
SETHRA LAVODE [DRAGAERA ADRILANKHA 3]  BRUST, STEVEN  H

The MITSFS Titledex  September 24, 2007  Page 316
TITLE  AUTHOR  CODE
SETI  FICHMAN, FREDERICK  C/P:2,P
SETITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  RYAN, KATHLEEN  P
SEVEN AGATE DEVILS, THE [DOC SAVAGE #73]  ROBESON, KENNETH  C/P:2,@P
SEVEN ALTARS OF DUSARRA, THE [LORDS OF DUS 2]  WATT-EVANS, LAWRENCE  C/P:2,P
SEVEN AMERICAN NIGHTS  WOLFE, GENE  SFWA-TD10,TD10

7 CARDINAL VIRTUES OF SCIENCE FICTION, THE ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

C/PA,PA

SEVEN COME INFINITY  CONKLIN, GROFF  PA,C/PA
SEVEN CONQUESTS  ANDERSON, POUL  C/P,SR-H,H,P
SEVEN CROWN [@BUFFY/ANGEL]  VORNHOLT, JOHN  C/P,P
SEVEN DAYS IN MAY  KNEBEL, FLETCHER/BAILEY, CHARLES W. II C/P:2,P
SEVEN DAYS IN NEW CRETE GRAVES, ROBERT C/P,P

7 DEADLY SINS AND CARDINAL VIRTUES OF SCIENCE FICTION, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

HA

7 DEADLY SINS OF SCIENCE FICTION, THE ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

HA,PA

SEVEN FAMOUS NOVELS  WELLS, H. G.  H
SEVEN FOOTPRINTS TO SATAN  MERRITT, A.  SR-P,C/P:2
SEVEN FOOTPRINTS TO SATAN/BURN WITCH BURN  MERRITT, A.  H
SEVEN FROM THE STARS  BRADLEY, MARION ZIMMER  C/P,DF127,P
SEVEN FROZEN SAILORS  FENN, GEORGE MANVILLE  H
SEVEN GOTHIC TALES  DINESEN, ISAK  P
SEVEN LAST YEARS, THE  BALIZET, CAROL  P
SEVEN MAGICAL JEWELS OF IRELAND, THE [@CASTAWAYS IN TIME #2]  ADAMS, ROBERT  C/P,P
SEVEN MASTERPIECES OF GOTHIC HORROR  SPECTOR, ROBERT DONALD  PA
SEVEN NO-TRUMP [AMBER]  RANDALL, NEIL  SFWA-P
SEVEN OF NINE [STAR TREK VOYAGER]  GOLDEN, CHRISTIE  C/P:2,P
SEVEN OF SWORDS [SWORD AND CIRCLET 3]  DOUGLAS, CAROLE NELSON  C/P,SFWA-P,P,H
SEVEN OUT OF TIME  ZAGAT, ARTHUR LEO  SR-H
SEVEN-PER-CENT SOLUTION, THE  MEYER, NICHOLAS  H,P
SEVEN SCIENCE FICTION NOVELS OF H.G. WELLS  WELLS, H. G.  H:2
SEVEN SERPENTS, THE [@SORCERY! #3]  JACKSON, STEVE  [C/RAN-P]
SEVEN SEXES, THE  TENN, WILLIAM  C/P:2,P
SEVEN SINS  ROHMER, SAX  H
SEVEN SONGS OF MERLIN, THE [LOST YEARS OF MERLIN 2]  BARRON, T. A.  P
SEVEN SPEARS OF THE W’DCH’CK, THE  FENN, LIONEL  C/P:2,SFWA-P,P
SEVEN SPELLS TO SUNDAY  NORTON, ANDRE/MILLER, PHYLLIS  H:2,P
SEVEN STEPS TO THE ARBITER  HUBBARD, L. RON  P
SEVEN STEPS TO THE SUN  HOYLE, FRED/HOYLE, GEOFFREY  C/P:2,P
SEVEN TICKETS TO HELL  WILLIAMS, ROBERT MOORE  C/P,P
SEVEN TOWERS, THE  WREDE, PATRICIA C.  C/P:2,P
SEVEN TRIPS THROUGH TIME AND SPACE  CONKLIN, GROFF  PA,C/PA
SEVEN WORLDS  CARAKER, MARY  C/P:2,P
17 X INFINITY  CONKLIN, GROFF  PA,C/PA
SEVENTH DAY, THE  KIRST, HANS HELLMUT  C/P,P
SEVENTH DIMENSION, THE  LA SALLE, VICTOR  SR-P
SEVENTH GATE, THE [SEVEN CITADELS 4]  HARRIS, GERALDINE  C/P,P
SEVENTH GATE, THE [@DEATH GATE #7]  WEIS, MARGARET/HICKMAN, TRACY  H:2,P
SEVENTH HEART, THE  FITCH, MARINA  P
SEVENTH MOUNTAIN, THE [MAGI 1]  CURTIS, GENE  LP
SEVENTH POWER, THE  MILLS, JAMES  P
SEVENTH SENSE  THOMPSON, HARRISON R.  P
SEVENTH SENTINEL, THE [@DRAGONLANCE DEFENDERS OF MAGIC #3]  KIRCHOFF, MARY  P
SEVENTH SON [ALVIN MAKER 1]  CARD, ORSON SCOTT  SFWA-P,H:2,P
SEVENTH SWAN, THE  GRAY, NICHOLAS STUART  P
SEVENTY-FIVE  WILLIAMSON, JACK  H
SEVERAL MINDS, THE  MORGAN, DAN  SR-P,C/P
SEVERED MAN, THE [@TIME HUNTER 5]  MANN, GEORGE  LP
SEWER, GAS & ELECTRIC  RUFF, MATT  P
SEX AND THE BLACK MACHINE  ALDISS, BRIAN W.  B
SEX AND THE HIGH COMMAND  BOYD, JOHN  C/P:2,P,H
SEX AND THE SINGLE VAMPIRE  MACALISTER, KATIE  P
SEX, LIES, AND VAMPIRES  MACALISTER, KATIE  P
SEX LIFE OF THE GODS, THE  KNERR, MICHAEL  SR-L
SEX LIFE OF THE IMMORTALS  YOUNG, RED  SR-L
SEX MAGICIANS, THE  WILSON, ROBERT ANTON  SR-L
SEX MAGICK  TAN, CECILIA  SR-L
SEX SPHERE, THE  RUCKER, RUDY  P
SEX WAR, THE (ORIGINALLY THE WHITE WIDOWS)  MERWIN, SAM JR.  C/P
SEXMAX  COOPER, HUGHES  P
SEXORCIST, THE  CLEVE, JOHN  SR-L
SEXORCIST, THE  FEDAHK, FATIMA  SR-L
SEXTOPIA  TAN, CECILIA  SR-L
SEXUAL CHEMISTRY  STABLEFORD, BRIAN M.  P
SEXUALIS ’95  STERNBERG, JACQUES  P

The MITSFS Titledex  September 24, 2007  Page 317
TITLE  AUTHOR  CODE
SF: ’57: THE YEAR’S GREATEST SCIENCE FICTION AND FANTASY  MERRIL, JUDITH  HA
SF: ’59: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY  MERRIL, JUDITH  HA
SF: AUTHORS’ CHOICE  HARRISON, HARRY  PA,C/PA
SF: AUTHORS’ CHOICE 2  HARRISON, HARRY  PA,C/PA
SF: AUTHORS’ CHOICE 3  HARRISON, HARRY  C/PA,PA
SF: AUTHORS’ CHOICE 4  HARRISON, HARRY  C/PA,PA
SF BIBLIOGRAPHIES  BRINEY, ROBERT/WOOD, EDWARD  R/IND
SF BOOK OF LISTS, THE  JAKUBOWSKI, MAXIM/EDWARDS, MALCOLM R/REF-P
SF IN DIMENSION  PANSHIN, ALEXEI/PANSHIN, CORY  R/REF-H
SF PUBLISHED IN 1968  BURGER, JOANNE  R/IND
SF PUBLISHED IN 1969  BURGER, JOANNE  R/IND:2
SF PUBLISHED IN 1970  BURGER, JOANNE  R/IND
SF PUBLISHED IN 1971  BURGER, JOANNE  R/IND
SF PUBLISHED IN 1972  BURGER, JOANNE  R/IND
SF PUBLISHED IN 1973  BURGER, JOANNE  R/IND
SF REPRISE 4  BONFIGLIOLI, KYRIL  PA
SF REPRISE 5  MOORCOCK, MICHAEL  PA
SF REPRISE 6  BONFIGLIOLI, KYRIL  PA
SF: THE BEST OF THE BEST  MERRIL, JUDITH  PA,C/PA
SF: THE NEW WRITERS 1973 & 1974  WHYTE, ANDREW ADAMS  R/IND
SF: THE OTHER SIDE OF REALISM  CLARESON, THOMAS D.  R/REF-H,C/REF-H
SF VOICES  SCHWEITZER, DARRELL  R/REF-P
SFWA EUROPEAN HALL OF FAME, THE  MORROW, JAMES/MORROW, KATHRYN  HA
SFWA GRAND MASTERS 1, THE  POHL, FREDERIK  LPA
SFWA GRAND MASTERS 2, THE  POHL, FREDERIK  HA,LPA
SFWA GRAND MASTERS 3, THE  POHL, FREDERIK  HA
SHACKLE AND SWORD  MORLAND, ALANNA  P
SHADA [DOCTOR WHO]  ADAMS, DOUGLAS  SR-P
SHADE  DEVENPORT, EMILY  P
SHADE AND SHADOW  WOODBURY, FRANCINE G.  P
SHADE OF LO MAN GONG, THE [JACK HONG 1]  WU, WILLIAM F.  B
SHADE OF THE TREE  ANTHONY, PIERS  C/P,SFWA-P,P,H
SHADE’S CHILDREN  NIX, GARTH  P
SHADES OF EVIL  WAGNER, SHARON  P
SHADES OF NIGHT, FALLING [@TWILIGHT ZONE 1]  MILLER, JOHN J.  LP
SHADOW  DUNCAN, DAVE  P
SHADOW [THREE LANDS 2]  EVANS, D. R.  LP
SHADOW [SHADOW 1]  LOGSTON, ANNE  C/P:2,SFWA-P:2,P
SHADOW, THE  LUCENO, JAMES  P
SHADOW [@SCAVENGER #1]  PARKER, K. J.  LP

SHADOW [@STAR TREK VOYAGER SECTION 31] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

SHADOW  SWEENEY, JOYCE  H
SHADOW ACADEMY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
SHADOW AND THE NIGHT, THE [@LAMB AMONG THE STARS #1]  WALLEY, CHRIS  H
SHADOW BLACK  ARDEN, TOM  LP
SHADOW CASTLE  COCKRELL, MARIAN  P
SHADOW CHASER, THE  BIRTOLO, DYLAN  LP
SHADOW CLIMBER [BIFROST GUARDIANS 2]  REICHERT, MICKEY ZUCKER  C/P,P
SHADOW DANCE [SHADOW 3]  LOGSTON, ANNE  P
SHADOW DANCERS  CARL, LILLIAN STEWART  C/P,SFWA-P,P
SHADOW DANCERS, THE [@G.O.D. INC. #2]  CHALKER, JACK L.  C/P:2,SFWA-P,P
SHADOW DAWN [SHADOW WAR 2]  CLAREMONT, CHRIS/LUCAS, GEORGE  H:2
SHADOW: DESTINATION MOON, THE  GRANT, MAXWELL  P
SHADOW EATER, THE [@DOMINIONS OF IRTH #2]  LEE, ADAM  P
SHADOW FLOW  KAO, TERESA  LP
SHADOW GAME [GHOSTWALKERS 1]  FEEHAN, CHRISTINE  H
SHADOW GAMES [BLACK COMPANY 4/BLACK COMPANY SOUTH 1]  COOK, GLEN  C/P,SFWA-P,P
SHADOW GATE, THE  BALL, MARGARET  C/P,P
SHADOW GIRL, THE  CUMMINGS, RAY  C/P:2,P
SHADOW — GO MAD!  GRANT, MAXWELL  P
SHADOW HAWK  NORTON, ANDRE  C/P,P,H
SHADOW HUNT [SHADOW 2]  LOGSTON, ANNE  P
SHADOW HUNTER  BAKER, WILL  P
SHADOW HUNTER, THE  MURPHY, PAT  LP
SHADOW IN SUMMER, A [LONG PRICE 1]  ABRAHAM, DANIEL  H
SHADOW IN THE DEEP [@BINDING OF THE BLADE #3]  GRAHAM, L. B.  LP
SHADOW IN THE GLASS, THE [DOCTOR WHO PAST DOCTORS]  RICHARDS, JUSTIN/COLE, STEPHEN  P
SHADOW KINGDOMS  HOWARD, ROBERT E.  H
SHADOW LEADER [WOLFWALKER 2]  HARPER, TARA K.  C/P,P
SHADOW LORD [STAR TREK]  YEP, LAURENCE  C/P:3,P
SHADOW MAGIC [LYRA 3]  WREDE, PATRICIA C.  C/P:2,P
SHADOW MAN  SCOTT, MELISSA  H,SFWA,LP
SHADOW MATRIX, THE [DARKOVER]  BRADLEY, MARION ZIMMER  H:2,P

The MITSFS Titledex  September 24, 2007  Page 318
TITLE  AUTHOR  CODE
SHADOW-MAZE  SMITH, MARK/SMITH, JULIA  P
SHADOW MOON [SHADOW WAR 1]  LUCAS, GEORGE/CLAREMONT, CHRIS  C/P,P,H
SHADOW OF A BROKEN MAN [MONGO FREDERICKSON 1]  CHESBRO, GEORGE C.  H,P
SHADOW OF A DARK QUEEN [SERPENTWAR 1]  FEIST, RAYMOND E.  C/P,P,H
SHADOW OF ALBION, THE [CAROLUS REX 1]  NORTON, ANDRE/EDGHILL, ROSEMARY  H:2,P
SHADOW OF ALL NIGHT FALLING, A [DREAD EMPIRE 1/RAGNARSON 3]  COOK, GLEN  C/P:2,P
SHADOW OF ALPHA, THE  GRANT, CHARLES L.  C/P:2,P
SHADOW OF ARARAT, THE [OATH OF EMPIRE 1]  HARLAN, THOMAS  C/P,P
SHADOW OF ARARAT, THE (PROOF) [OATH OF EMPIRE 1]  HARLAN, THOMAS  SR-LP
SHADOW OF EARTH  EISENSTEIN, PHYLLIS  C/P:2,P
SHADOW OF FU MANCHU, THE  ROHMER, SAX  C/P:2,P
SHADOW OF HEAVEN [@STAR TREK VOYAGER DARK MATTERS #3]  GOLDEN, CHRISTIE  P
SHADOW OF HEAVEN  SHAW, BOB  C/P:2,P
SHADOW OF HIS WINGS, THE [SIX KINGDOMS]  FERGUSSON, BRUCE  H,SFWA,P
SHADOW OF OBSESSION [@HIGHLANDER]  NEASON, REBECCA  P
SHADOW OF SAGANAMI, THE [HONOR HARRINGTON SAGANAMI 1]  WEBER, DAVID  H
SHADOW OF SORCERY, THE [THIEVES’ WORLD SHADOWSPAWN]  OFFUTT, ANDREW J.  P,SFWA-P:2
SHADOW OF THE CROWN  MILLS, CRAIG  P
SHADOW OF THE FLAME [@DRAGONLANCE TALADAS #3]  PIERSON, CHRIS  P
SHADOW OF THE GIANT [ENDER 1.4]  CARD, ORSON SCOTT  H:2,P
SHADOW OF THE HEGEMON [ENDER 1.2]  CARD, ORSON SCOTT  C/P,P,H

SHADOW OF THE LION, THE [SHADOW OF THE LION 1] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H:2,P

SHADOW OF THE MUTANT MASTER [@PERRY RHODAN #47]  ACE  PA,C/PA
SHADOW OF THE RED MOON  MYERS, WALTER DEAN  P
SHADOW OF THE SEVENTH MOON  BERBERICK, NANCY VARIAN  C/P,SFWA-P:2,P
SHADOW OF THE SHIP, THE  FRANSON, ROBERT WILFRED  C/P:2,P
SHADOW OF THE SPHINX [@CHARMED]  JABLONSKI, CARLA  P
SHADOW OF THE STORM [FALLEN CLOUD 3]  GIAMBASTIANI, KURT R. A.  P
SHADOW OF THE SWAN [PHOENIX LEGACY 2]  WREN, M. K.  C/P:2,P
SHADOW OF THE TORTURER, THE [NEW SUN 1]  WOLFE, GENE  C/P,P,H
SHADOW OF THE WARMASTER  CLAYTON, JO  C/P:2,P

SHADOW OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 2] CHALKER, JACK L.

H:2,LP

SHADOW OF THE WOLF [@BERSERKER SWIFTAXE 1]  CARLSEN, CHRIS  P
SHADOW OF THE XEL’NAGA [@STARCRAFT #2]  MESTA, GABRIEL  P
SHADOW OF TOMORROW  POHL, FREDERIK  SR-PA,C/PA
SHADOW OF WENG-CHIANG, THE [DOCTOR WHO MISSING ADVENTURES]  MCINTEE, DAVID A.  C/P,P
SHADOW ON THE GLASS, A [@VIEW FROM THE MIRROR #1]  IRVINE, IAN  C/P:2,P
SHADOW ON THE HEARTH  MERRIL, JUDITH  H
SHADOW ON THE SAND [LONE WOLF 5]  DEVER, JOE  P
SHADOW ON THE STARS  MARCUS, ROBERT B. JR.  SR-P
SHADOW ON THE STONES [SACRED STONES 3]  CALDECOTT, MOYRA  P
SHADOW OVER INNSMOUTH AND OTHER STORIES OF HORROR, THE  LOVECRAFT, H. P.  P
SHADOW OVER KIRIATH [GUARDIAN-KING 3]  HANCOCK, KAREN  LP
SHADOW OVER THE EARTH  WILDING, PHILIP  H
SHADOW PEOPLE, THE  SAINT CLAIR, MARGARET  C/P:2,P
SHADOW PLANET [QUEST FOR TOMORROW 5]  SHATNER, WILLIAM  H
SHADOW PLAY  BEAUMONT, CHARLES  SR-P
SHADOW POINT [@WARHAMMER 40K]  RENNIE, GORDON  P
SHADOW PUPPETS [ENDER 1.3]  CARD, ORSON SCOTT  H:2,P
SHADOW RISING, THE [WHEEL OF TIME 4]  JORDAN, ROBERT  C/P:2,P
SHADOW ROADS, THE [SWANS’ WAR 3]  RUSSELL, SEAN  H
SHADOW SHAIA, THE [WIZENBEAK 2]  GILLILAND, ALEXIS A.  P
SHADOW SINGER  BENNETT, MARCIA J.  C/P:2,P
SHADOW SORCERESS, THE [SPELLSONG 4]  MODESITT, L. E. JR.  C/P,P,H
SHADOW STAR [SHADOW WAR 3]  CLAREMONT, CHRIS/LUCAS, GEORGE  P
SHADOW STEED [DRAGONREALM 4]  KNAAK, RICHARD A.  P
SHADOW STONE [@FORGOTTEN REALMS]  BAKER, RICHARD  P
SHADOW WAR [REIGN OF SHADOWS 2]  CHESTER, DEBORAH  P
SHADOW WITHIN, THE [@BABYLON 5 #7]  CAVELOS, JEANNE  C/P,P
SHADOW WITHIN, THE [GUARDIAN-KING 2]  HANCOCK, KAREN  LP
SHADOW WORLD [@DEATHLANDS]  AXLER, JAMES  P
SHADOW WORLD [@STARBRIDGE #3]  CRISPIN, A. C./ELLIOTT, JANNEAN  C/P:2,SFWA-P:2,P
SHADOWBORN [@RAVENLOFT]  CONNORS, WILLIAM W./BEBRIS, CARRIE A.  P
SHADOWBOXER [SHADOWRUN 25]  POLLOTTA, NICK  P
SHADOWBRED [@FORGOTTEN REALMS TWILIGHT WAR #1]  KEMP, PAUL  P
SHADOWBREED [WARHAMMER KONRAD 2]  FERRING, DAVID  LP
SHADOWDALE [FORGOTTEN REALMS AVATAR 1]  AWLINSON, RICHARD  C/P:2,P
SHADOWDALE [@FORGOTTEN REALMS AVATAR #1]  CIENCIN, SCOTT  P
SHADOWDANCE  BAILEY, ROBIN WAYNE  H,P
SHADOWED BY WINGS [@DRAGON TEMPLE #2]  CROSS, JANINE  LP
SHADOWEYES  PTACEK, KATHRYN  P
SHADOWFALL [@DEATHLANDS]  AXLER, JAMES  P

The MITSFS Titledex  September 24, 2007  Page 319
TITLE  AUTHOR  CODE
SHADOWFALL [GODSLAYER 1]  CLEMENS, JAMES  H:2
SHADOWFANE [CYCLE OF FIRE 3]  WURTS, JANNY  C/P:2,P
SHADOWFIRES  NICHOLS, LEIGH  P
SHADOWHEART [LEGENDS OF THE RAVEN 2]  BARCLAY, JAMES  LP
SHADOWKEEP  FOSTER, ALAN DEAN  C/P:2,P
SHADOWKILL [SHADITH’S QUEST 3/DIADEM 12]  CLAYTON, JO  C/P,P
SHADOWKINGS [@SHADOWKINGS #1]  COBLEY, MICHAEL  LP
SHADOWLAND  STRAUB, PETER  H:2,P
SHADOWLIGHT  HYMAN, JACKIE  C/P,P
SHADOWLIGHT [ELUNDIUM 3]  JEFFERIES, MIKE  P
SHADOWLINE [STARFISHERS 1]  COOK, GLEN  P
SHADOWMAN  PROCTOR, GEO. W.  P
SHADOWMANCER  TAYLOR, G. P.  H
SHADOWMARCH [SHADOWMARCH 1]  WILLIAMS, TAD  H

SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #2 (EXCERPT, CHAPTERS 3 AND 4) [SHADOWMARCH 1] WILLIAMS, TAD

SR-LP

SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #3 (EXCERPT, CHAPTERS 5 AND 6) [SHADOWMARCH 1] WILLIAMS, TAD

SR-LP

SHADOWMIND [DOCTOR WHO NEW ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
SHADOWPLAY [SHADITH’S QUEST 1/DIADEM 10]  CLAYTON, JO  C/P:2,P
SHADOWPLAY [@SHADOWRUN 9]  FINDLEY, NIGEL D.  C/P,P
SHADOWPLAY [SHADOWMARCH 2]  WILLIAMS, TAD  H
SHADOWS [@SMALLVILLE]  GALLAGHER, DIANA G.  P
SHADOWS  GRANT, CHARLES L.  HA,PA,SFWA
SHADOWS  SAUL, JOHN  P
SHADOWS [@CHILL #7]  SHERMAN, JORY  C/P,P
SHADOWS 2  GRANT, CHARLES L.  HA,PA
SHADOWS 3  GRANT, CHARLES L.  C/PA,PA,SFWA
SHADOWS 4  GRANT, CHARLES L.  HA,PA,SFWA
SHADOWS 5  GRANT, CHARLES L.  HA,PA,SFWA
SHADOWS 6  GRANT, CHARLES L.  SFWA
SHADOWS 7  GRANT, CHARLES L.  HA,PA,SFWA
SHADOWS 8  GRANT, CHARLES L.  HA,PA,SFWA
SHADOWS 9  GRANT, CHARLES L.  HA,SFWA
SHADOWS 10  GRANT, CHARLES L.  HA,SFWA
SHADOWS (PROOF) [@SMALLVILLE]  GALLAGHER, DIANA G.  SR-P,C/P
SHADOWS AFTER DARK  CROZIER, OUIDA  LP
SHADOWS AND LIGHT [TIR ALAINN 2]  BISHOP, ANNE  C/P,P
SHADOWS AND SHADES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 8]  LEE, SHARON/MILLER, STEVE  SR-LP
SHADOWS BEND  BARBOUR, DAVID/RALEIGH, RICHARD  LP
SHADOWS BITE [MAGE AND TAKUMO 2]  DEDMAN, STEPHEN  H,LP
SHADOW’S DAUGHTER [FIFTH MILLENNIUM]  MEIER, SHIRLEY  C/P:2,P
SHADOW’S END  TEPPER, SHERI S.  H:2,P
SHADOWS FALL  GREEN, SIMON R.  C/P,P
SHADOWS IN THE DARKNESS [CHANGELING DETECTIVE 1]  CUNNINGHAM, ELAINE  H
SHADOWS IN THE STARLIGHT [CHANGELING DETECTIVE 2]  CUNNINGHAM, ELAINE  H
SHADOWS IN THE SUN  OLIVER, CHAD  H:2,P
SHADOWS IN THE WATCHGATE  JEFFERIES, MIKE  P
SHADOWS, KITH AND KIN, THE  LANSDALE, JOE R.  SR-H
SHADOWS LINGER [BLACK COMPANY 2]  COOK, GLEN  SR-P,C/P:2
SHADOWS OF AGGAR [AGGAR 1]  WOLFE, CHRIS ANNE  LP
SHADOWS OF AMBER [@ROGER ZELAZNY’S DAWN OF AMBER 4]  BETANCOURT, JOHN GREGORY  H
SHADOWS OF AMN [@FORGOTTEN REALMS BALDUR’S GATE II]  ATHANS, PHILIP  P
SHADOWS OF AVALON, THE [DOCTOR WHO]  CORNELL, PAUL  P
SHADOWS OF DOOM [@FORGOTTEN REALMS SHADOW OF AVATAR #1]  GREENWOOD, ED  P
SHADOWS OF DOOM [@IRON TOWER #2/MITHGAR 8]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
SHADOWS OF ECSTASY  WILLIAMS, CHARLES  H,LP
SHADOWS OF GOD, THE [AGE OF UNREASON 4]  KEYES, J. GREGORY  H,LP
SHADOWS OF ICE [@WARCRAFT SUNWELL 2]  KNAAK, RICHARD A.  SCX
SHADOWS OF IMAGINATION  HILLEGAS, MARK R.  R/REF-H
SHADOWS OF MEDUSA  ENKE, BRIAN  H,LP
SHADOWS OF MYTH  LEE, RACHEL  LP
SHADOWS OF SANCTUARY [@THIEVES’ WORLD #3]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
SHADOWS OF SANCTUARY (BRITISH) [THIEVES’ WORLD 3]  ASPRIN, ROBERT LYNN  PA
SHADOWS OF THE EMPIRE [@STAR WARS]  PERRY, STEVE  H:2,P
SHADOWS OF THE PAST [HOLDERLY HALL 2]  CAMERON, KATE  P
SHADOWS OF THE PAST [@BATMAN ANIMATED SERIES 1]  GRAVEL, GEARY  P
SHADOWS OF THE WHITE SUN  HARRIS, RAYMOND  H:2
SHADOWS OF WAR [@BATTLETECH #38/TWILIGHT OF THE CLANS 6]  GRESSMAN, THOMAS S.  P
SHADOWS ON THE HILL [@IMMORTAL EYES #2]  CASSADA, JACKIE  P
SHADOWS ON THE MOON [WHISPERING HILLS 3]  CAMERON, KATE  P
SHADOWS ON THE STARS [GREAT TREE OF AVALON 2]  BARRON, T. A.  P
SHADOWS ON THE SUN [STAR TREK]  FRIEDMAN, MICHAEL JAN  C/P:2,P
SHADOWS OUT OF HELL [WAR OF THE GODS ON EARTH 2]  OFFUTT, ANDREW J.  C/P,P

The MITSFS Titledex  September 24, 2007  Page 320
TITLE  AUTHOR  CODE
SHADOWS OVER BAKER STREET  REAVES, MICHAEL/PELAN, JOHN  HA
SHADOWS OVER LYRA [LYRA 1,2,3]  WREDE, PATRICIA C.  H,LP
SHADOW’S REALM [BIFROST GUARDIANS 4]  REICHERT, MICKEY ZUCKER  P
SHADOW’S REVENGE, THE  GRANT, MAXWELL  P

SHADOW’S SON [FIFTH MILLENNIUM] MEIER, SHIRLEY/STIRLING, S. M./

WEHRSTEIN, KAREN

C/P,SFWA-P,P

SHADOWS WITH EYES  LEIBER, FRITZ  C/P,P
SHADOW’S WITNESS [@FORGOTTEN REALMS SEMBIA]  KEMP, PAUL S.  P
SHADOWSINGER [SPELLSONG 5]  MODESITT, L. E. JR.  H
SHADOWSMITH, THE [@MIGHT AND MAGIC #2]  GRAVEL, GEARY  C/P,P
SHADOWSPAWN [THIEVES’ WORLD SHADOWSPAWN]  OFFUTT, ANDREW J.  H:2,P
SHADOWSPEER [SHADITH’S QUEST 2/DIADEM 11]  CLAYTON, JO  C/P:2,P
SHADRACH IN THE FURNACE  SILVERBERG, ROBERT  C/P,P,H
SHAGGY D.A., THE  CRUME, VIC  P
SHAGGY DOG, THE  GRIFFEN, ELIZABETH L.  P
SHAGGY PLANET  GOULART, RON  C/P:2,P
SHAI’S DESTINY  WALTHER, DANIEL  P
SHAKEDOWN [@ANGEL]  DEBRANDT, DON  P
SHAKEDOWN [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P:2,P
SHAKESPEARE’S PLANET  SIMAK, CLIFFORD D.  H:2,P
SHAKING EARTH [@DEATHLANDS]  AXLER, JAMES  P
SHAKWA, THE  MARCH, GENE  P
SHALLOWS OF NIGHT [SUNSET WARRIOR 2]  LUSTBADER, ERIC VAN  C/P,SFWA,P,H
SHAMAN, THE  COFFEY, FRANK  H,P
SHAMAN  MIESEL, SANDRA  C/P,SFWA-P,P
SHAMAN, THE [@STAR STONE #1]  STASHEFF, CHRISTOPHER  P
SHAMAN’S CROSSING [SOLDIER SON 1]  HOBB, ROBIN  H
SHAMBALLAH [@MIND MASTERS #2]  ROSSMANN, JOHN F.  P
SHAMBLEAU  MOORE, C. L.  GN5.31,C/P,SR-P
SHAMBLEAU AND OTHERS  MOORE, C. L.  SR-H,H
SHAME OF MAN [GEODESSEY 2]  ANTHONY, PIERS  H,P
SHAMER’S SIGNET, THE [SHAMER 2]  KAABERSOL, LENE  H
SHAMROCK SHENANIGANS [@SABRINA THE TEENAGE WITCH #19]  GALLAGHER, DIANA G.  P
SHAMUTANTI HILLS, THE [@SORCERY! #1]  JACKSON, STEVE  [C/RAN-P:2]
SHANEE, THE  MYERS, BARRY  P
SHANEE WOMAN  MYERS, BARRY  P
SHANJI [SHANJI INTERSTELLAR EMPIRE 1]  GLASS, JAMES C.  P
SHAPE CHANGER, THE [LAFAYETTE O’LEARY 3]  LAUMER, KEITH  C/P,P,H
SHAPE-CHANGER’S WIFE, THE  SHINN, SHARON  C/P,LP
SHAPE OF FURTHER THINGS, THE  ALDISS, BRIAN W.  SFWA,R/REF-H
SHAPE OF SEX TO COME, THE  HILL, DOUGLAS  SR-PA
SHAPE OF SPACE, THE  NIVEN, LARRY  C/P:2,P

SHAPE OF TERROR, THE [DOC SAVAGE #109]/ DEATH HAD YELLOW EYES [DOC SAVAGE 110] ROBESON, KENNETH

@P

SHAPE OF THINGS, THE  KNIGHT, DAMON  PA,C/PA
SHAPE OF TOMORROW, THE  SOULE, GEORGE  [C/RAN-P:2]
SHAPECHANGER SCENARIO, THE [PSYCHODROME 2]  HAWKE, SIMON  C/P,SFWA-P,P
SHAPECHANGERS [CHEYSULI 1]  ROBERSON, JENNIFER  C/P:2,P
SHAPECHANGER’S SONG (INCLUDES "KINSPIRIT") [@CHEYSULI #1,2,3]  ROBERSON, JENNIFER  P
SHAPERS  CHASE, ROBERT R.  P
SHAPER’S LEGACY [SHAPER EXILE 2]  FINCH, SHEILA  C/P:2,P
SHAPERS OF DARKNESS [WINDS OF THE FORELANDS 4]  COE, DAVID B.  H
SHAPES OF MIDNIGHT, THE  BRENNAN, JOSEPH PAYNE  P
SHAPES OF THE SUPERNATURAL  MANLEY, SEON/LEWIS, GOGO  HA
SHAPES OF THEIR HEARTS, THE  SCOTT, MELISSA  H
SHAPING OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 4]  TOLKIEN, J. R. R.  H,P
SHAPING THE DAWN [SHAPER EXILE 3]  FINCH, SHEILA  C/P:2,SFWA-P,P
SHARDIK  ADAMS, RICHARD  H:2,P
SHARDS OF A BROKEN CROWN [SERPENTWAR 4]  FEIST, RAYMOND E.  H:2,P
SHARDS OF ALDERAAN [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  C/P,P
SHARDS OF DARKNESS  BERGLUND, E. P.  LP
SHARDS OF EMPIRE  SHWARTZ, SUSAN  H,P
SHARDS OF HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1]  BUJOLD, LOIS MCMASTER  C/P:2,P,H
SHARDS OF SPACE  SHECKLEY, ROBERT  C/P:2,P
SHARED TOMORROWS: SCIENCE FICTION IN COLLABORATION  PRONZINI, BILL/MALZBERG, BARRY N.  HA
SHAREWORLD  HERSHMAN, MORRIS  H,SFWA
SHARP END, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,SFWA,P,H
SHARP PRACTICE  FARRIS, JOHN  SFWA-P
SHARPER THAN A SERPENT’S TOOTH [NIGHTSIDE 6]  GREEN, SIMON R.  P
SHARPEST EDGE, THE [FIFTH MILLENNIUM]  STIRLING, S. M./MEIER, SHIRLEY  C/P,P
SHARPNESS ON THE NECK, A [DRACULA]  SABERHAGEN, FRED  H,P
SHARRA’S EXILE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
SHAST WAR, THE [@ARNA 2]  ROWLEY, CHRISTOPHER  P
SHATTERDAY  ELLISON, HARLAN  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 321
TITLE  AUTHOR  CODE
SHATTERED ALLIANCE, THE [@MAGIC: THE GATHERING ICE AGE 3]  GRUBB, JEFF  P
SHATTERED AND OTHER STORIES [SHATTERZONE]  FARSHTEY, GREG  LPA
SHATTERED CHAIN, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
SHATTERED CHAINS [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
SHATTERED DANCE [AURELIAN EMPIRE 3]  BRENNAN, CAITLIN  LP
SHATTERED GLASS  BERGSTROM, ELAINE  C/P,P
SHATTERED GODDESS, THE  SCHWEITZER, DARRELL  H,LP
SHATTERED HORSE, THE  SOMTOW, S. P.  H,SFWA-P,P
SHATTERED LAND, THE [@EBERRON DREAMING DARK #2]  BAKER, KEITH  P
SHATTERED MASK, THE [@FORGOTTEN REALMS SEMBIA]  BYERS, RICHARD LEE  P
SHATTERED OATH, THE [PRINCE OF THE SIDHE 1]  SHERMAN, JOSEPHA  P
SHATTERED PEOPLE, THE  HOSKINS, ROBERT  H,SFWA
SHATTERED RING, THE  ROSE, LOIS/ROSE, STEPHEN  R/REF-H
SHATTERED SKY [STAR SHARDS 3]  SHUSTERMAN, NEAL  P
SHATTERED SPHERE, THE [HUNTED EARTH 2]  ALLEN, ROGER MACBRIDE  H,P
SHATTERED SPHERE, THE  ASPRIN, ROBERT LYNN/ABBEY, LYNN  HA
SHATTERED STARS, THE [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  C/P:2,P
SHATTERED TWILIGHT [@BUFFY WICKED WILLOW #2]  NAVARRO, YVONNE  P
SHATTERED WORLD, THE  REAVES, MICHAEL  C/P,LP,H
SHATTERPOINT [@STAR WARS CLONE WARS]  STOVER, MATTHEW WOODRING  P
SHAUNESSY FONG [JACK HONG 2]  WU, WILLIAM F.  B
SHAVING OF SHAGPAT, THE  MEREDITH, GEORGE  P
SHE [AYESHA 1]  HAGGARD, H. RIDER  H:2,P
SHE AND ALLAN [AYESHA 4/ALLAN QUATERMAIN]  HAGGARD, H. RIDER  LP
SHE-BEAST, THE [@MAN FROM PLANET X #1]  ADAMS, HUNTER  SR-P
SHE DEVIL, THE  HOWARD, ROBERT E.  P

SHE IS THE DARKNESS [@BLACK COMPANY GLITTERING STONE 2/ BLACK COMPANY 7] COOK, GLEN

H,P

SHE WAITS  CLEMENT, HENRY  P
SHEEP LOOK UP, THE  BRUNNER, JOHN  H:2,P
SHEEPFARMER’S DAUGHTER [PAKSENARRION 1]  MOON, ELIZABETH  C/P:3,SFWA-P:2,P
SHELL GAME [STAR TREK]  CRANDALL, MELISSA  C/P:2,P
SHELL SHOCK [DOCTOR WHO]  FORWARD, SIMON A.  SR-H
SHELLBREAK  GROVES, J. W.  C/P:2,P
SHELTER  ASHER, MARTY  H,SFWA,LP
SHELTER, THE  KITTREDGE, MARY/O’DONNELL, KEVIN JR.  SFWA-P
SHELTER  LJOKA, DAN  P
SHELTER  PALWICK, SUSAN  LP
SHELTERED LIVES  OBERNDORF, CHARLES  C/P,P
SHEN OF THE SEA: CHINESE STORIES FOR CHILDREN  CHRISMAN, ARTHUR BOWIE  H
SHEPHERD  HOLLY, JOAN HUNTER  SR-P
SHEPHERD IS MY LORD, THE  GAT, DIMITRI V.  H,SFWA
SHEPHERD KINGS, THE  TARR, JUDITH  H
SHERIFF OF PURGATORY, THE  MORRIS, JIM  SFWA
SHERLOCK HOLMES IN ORBIT  RESNICK, MIKE/GREENBERG, MARTIN H.  C/PA,PA

SHERLOCK HOLMES THROUGH TIME AND SPACE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

LPA

SHERLOCK HOLMES VS. DRACULA  ESTLEMAN, LOREN  P
SHERLOCK HOLMES’S WAR OF THE WORLDS  WELLMAN, MANLY WADE/WELLMAN, WADE C/P,P
SHERWOOD [SHERWOOD 1]  GODWIN, PARKE  P,SFWA-P
SHERWOOD GAME, THE  FRIESNER, ESTHER M.  P
SHERWOOD RING, THE  POPE, ELIZABETH MARIE  P
SHE’S NO FAERIE PRINCESS [OTHERS (WARREN) 2]  WARREN, CHRISTINE  P
SHIELD  ANDERSON, POUL  C/P:2,P

SHIELD BETWEEN THE WORLDS, THE [FIONN MAC CUMHAL 2] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

H,P

SHIELD OF LIES [STAR WARS BLACK FLEET CRISIS 2]  KUBE-MCDOWELL, MICHAEL P.  C/P:2,P
SHIELD OF THE SKY  KRINARD, SUSAN  LP
SHIELD OF THUNDER [@TROY 2]  GEMMELL, DAVID  H
SHIELD OF TIME, THE  ANDERSON, POUL  H:2,P
SHIELDBREAKER’S STORY [@LOST SWORDS #8]  SABERHAGEN, FRED  C/P,P,H
SHIFT, THE  FOY, GEORGE  P
SHIFT KEY, THE  BRUNNER, JOHN  P

SHIFTER [@GALEN SWORD #1] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P:2,P

SHIMMERING DOOR, THE  KERR, KATHARINE  LPA
SHINING, THE  KING, STEPHEN  H,P
SHINING CITY, THE [@RHIANNON’S RIDE #2]  FORSYTH, KATE  P
SHINING COURT, THE [@SUN SWORD #3]  WEST, MICHELLE  C/P,P
SHINING FALCON, THE  SHERMAN, JOSEPHA  C/P:2,P
SHINING ISLE  DE ANGELES, LY  LP
SHINING ONES, THE [TAMULI 2]  EDDINGS, DAVID  H:2,P
SHINING ONES, THE [@VENUS PRIME #6]  PREUSS, PAUL  C/P,P,H
SHINING POOL, THE  CARLSON, DALE/CARLSON, DANNY  H

The MITSFS Titledex  September 24, 2007  Page 322
TITLE  AUTHOR  CODE
SHINING STEEL  WATT-EVANS, LAWRENCE  C/P,P
SHIP AVENGED, THE [SHELLPEOPLE 7]  STIRLING, S. M.  C/P,P,H
SHIP ERRANT, THE [SHELLPEOPLE 6]  NYE, JODY LYNN  C/P,P
SHIP FROM ATLANTIS, THE  MUNN, H. WARNER  DG618
SHIP FROM NOWHERE, THE (DOUBLE WITH THE MOON MIRAGE)  PATZER, SIDNEY  B
SHIP FROM OUTSIDE, THE [GRIMES]  CHANDLER, A. BERTRAM  DF237
SHIP OF DESTINY [LIVESHIP TRADERS 3]  HOBB, ROBIN  H
SHIP OF DESTINY  SLATER, HENRY J.  H
SHIP OF DREAMS [DREAMWORLD 2]  LUMLEY, BRIAN  H
SHIP OF FOOLS  RUSSO, RICHARD PAUL  C/P,LP,H
SHIP OF FOOLS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  STONE, DAVE  C/P,P
SHIP OF GHOSTS [@FARSCAPE]  BISCHOFF, DAVID F.  C/P,P
SHIP OF ISHTAR, THE  MERRITT, A.  SFWA-P,H:2,SR-LP,P
SHIP OF MAGIC [LIVESHIP TRADERS 1]  HOBB, ROBIN  H:2,P
SHIP OF SHADOWS  LEIBER, FRITZ  TD5
SHIP OF STRANGERS  SHAW, BOB  C/P,P
SHIP OF THE DAMNED  DAVID, JAMES F.  P
SHIP OF THE LINE [STAR TREK TNG]  CAREY, DIANE  P
SHIP THAT SAILED THE TIME STREAM, THE [TIME STREAM 1]  EDMONDSON, G. C.  DM109
SHIP THAT SAILED THE TIME STREAM, THE (REVISED) [TIME STREAM 1]  EDMONDSON, G. C.  C/P:3,P
SHIP WHO SANG, THE [SHELLPEOPLE 1]  MCCAFFREY, ANNE  C/P,P,H
SHIP WHO SEARCHED, THE [SHELLPEOPLE 3]  MCCAFFREY, ANNE/LACKEY, MERCEDES  C/P,P,H
SHIP WHO WON, THE [SHELLPEOPLE 5]  MCCAFFREY, ANNE/NYE, JODY LYNN  C/P,P,H
SHIPBUILDING  TAVERNA PRESS  LPA
SHIPS FROM THE WEST [@MONARCHIES OF GOD #5]  KEARNEY, PAUL  C/P,P
SHIPS OF AIR, THE [FALL OF ILE-RIEN 2]  WELLS, MARTHA  H
SHIPS OF DUROSTORUM, THE [KEYS TO THE DIMENSIONS 5]  BULMER, KENNETH  D76096
SHIPS OF EARTH, THE [HOMECOMING 3]  CARD, ORSON SCOTT  H:2,P

SHIPS OF MERIOR (COMPLETE) [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 1,2] WURTS, JANNY

H

SHIPS OF MERIOR (PARTIAL) [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 1] WURTS, JANNY

C/P:2,P

SHIPS OF VERO  SHAW, BRIAN  SR-P,H
SHIPS TO THE STARS  LEIBER, FRITZ  DF285
SHIPWRECK  LOGAN, CHARLES  P
SHIVA ACCUSED  BRENNAN, J. H.  SFWA
SHIVA DESCENDING  BENFORD, GREGORY/ROTSLER, WILLIAM  C/P:2,P
SHIVA IN STEEL [BERSERKER 8]  SABERHAGEN, FRED  C/P,P,H
SHIVA OPTION, THE [STARFIRE 4]  WEBER, DAVID/WHITE, STEVE  C/P,P
SHIVERING WORLD  TYERS, KATHY  C/P:2,P
SHOCK!  MATHESON, RICHARD  C/P:2,P
SHOCK II  MATHESON, RICHARD  C/P:2,P
SHOCK III  MATHESON, RICHARD  C/P:2,P
SHOCK WAVE  RICHMOND, WALT/RICHMOND, LEIGH  DG614
SHOCK WAVES  MATHESON, RICHARD  C/P,P
SHOCKBALL [STARDOC 4]  VIEHL, S. L.  C/P,P
SHOCKING THING, A  KNIGHT, DAMON  PA
SHOCKWAVE [STAR TREK ENTERPRISE]  RUDITIS, PAUL  H,@P
SHOCKWAVE [@FIREBRATS #4]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
SHOCKWAVE RIDER, THE  BRUNNER, JOHN  H:2,P
SHOELESS JOE  KINSELLA, W. P.  P
SHOES OF THE FISHERMAN, THE  WEST, MORRIS L.  P
SHOGGOTH ON THE ROOF, A: THE DOCUMENTARY (DVD)  BRANNEY, SEAN/LEMAN, ANDREW  SR-MM
SHOGGOTH ON THE ROOF, A: THE HPLHS CAST ALBUM  HPLHS INC.  SR-MM
SHOGGOTH ON THE ROOF, A: THE MUSICAL  BRANNEY, SEAN/LEMAN, ANDREW  SR-VLP
SHOOT, THE  TREVOR, ELLESTON  H
SHOOT AT THE MOON  TEMPLE, WILLIAM F.  C/P:2,P,H:2
SHOOTING OF THE GREEN, THE  POYER, JOE  P
SHOOTING SCRIPTS, THE [DOCTOR WHO TENTH DOCTOR]  BBC BOOKS  R/REF-H
SHORE LEAVE [STAR TREK]  STURGEON, THEODORE  SR-S
SHORE OF WOMEN, THE  SARGENT, PAMELA  C/P,SFWA,P,H:2
SHORES BENEATH, THE  SALLIS, JAMES  PA,C/PA
SHORES OF ANOTHER SEA, THE  OLIVER, CHAD  C/P,P
SHORES OF DEATH, THE  MOORCOCK, MICHAEL  C/P,P
SHORES OF KANSAS, THE  CHILSON, ROBERT  C/P,P
SHORES OF SPACE, THE  MATHESON, RICHARD  C/P:2,P
SHORES OF TOMORROW, THE [SOLACE 3]  ALLEN, ROGER MACBRIDE  P
SHORES OF TOMORROW, THE  MASON, DAVID  C/P:2,P
SHORES OF TOMORROW, THE  SILVERBERG, ROBERT  H
SHORT BIOGRAPHY OF ROBERT E. HOWARD, A  BURKE, RUSTY  R/REF-H
SHORT BLADE  EMSHWILLER, PETER R.  P
SHORT DICTIONARY OF MYTHOLOGY  WOODCOCK, P. G.  [C/RAN-H]
SHORT HISTORY OF THE FUTURE, A (SECOND ED.)  WAGAR, W. WARREN  LP
SHORT REIGN OF PIPPIN IV, THE  STEINBECK, JOHN  H,P

The MITSFS Titledex  September 24, 2007  Page 323
TITLE  AUTHOR  CODE
SHORT, SHARP SHOCK, A  ROBINSON, KIM STANLEY  H,SFWA-TD28,P
SHORT STORIES OF H. G. WELLS, THE  WELLS, H. G.  H:2
SHORT TRIPS [DOCTOR WHO]  COLE, STEPHEN  C/PA,PA
SHORT TRIPS: A CHRISTMAS TREASURY [DOCTOR WHO]  CORNELL, PAUL  HA
SHORT TRIPS: A DAY IN THE LIFE [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: A UNIVERSE OF TERRORS [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS AND SIDE STEPS [DOCTOR WHO]  COLE, STEPHEN/RAYNER, JACQUELINE  PA
SHORT TRIPS: COMPANIONS [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SHORT TRIPS: DALEK EMPIRE [DOCTOR WHO]  BRIGGS, NICHOLAS/GUERRIER, SIMON  HA
SHORT TRIPS: LIFE SCIENCES [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: MONTSERS [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: PAST TENSE [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: REPERCUSSIONS [DOCTOR WHO]  RUSSELL, GARY  HA
SHORT TRIPS: SEVEN DEADLY SINS [DOCTOR WHO]  BAILEY, DAVID  HA
SHORT TRIPS: STEEL SKIES [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: THE CENTENARIAN [DOCTOR WHO]  FARRINGTON, IAN  HA
SHORT TRIPS: THE MUSES [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SHORT TRIPS: THE SOLAR SYSTEM [DOCTOR WHO]  RUSSELL, GARY  HA
SHORT TRIPS: 2040 [DOCTOR WHO]  BINNS, JOHN  HA
SHORT TRIPS: ZODIAC [DOCTOR WHO]  RAYNER, JACQUELINE  HA
SHORT VICTORIOUS WAR, THE [HONOR HARRINGTON 3]  WEBER, DAVID  C/P,SFWA-P,P
SHORT VISIT TO ERGON  OSBORN, E. M.  LP
SHORTCUT IN TIME, A  DICKINSON, CHARLES  LP
SHOT IN THE DARK  MERRIL, JUDITH  SR-PA
SHOTLAND PUZZLE, THE (CHINESE)  CHEUNG, H. K.  FP

SHOW BUSINESS IS MURDER WAUGH, CAROL-LYNN ROSSEL/

GREENBERG, MARTIN H./ASIMOV, ISAAC

PA

SHOW ON ICE, THE  VERNE, JULES  H
SHOWBOAT WORLD  VANCE, JACK  C/P:2,SFWA-P,P
SHOWCASE  ELWOOD, ROGER  SR-HA
SHOWCASE OF FANTASY ART  PETAJA, EMIL  A
SHOWDOWN [TIME TRAP 2]  DALTON, SEAN  P
SHOWDOWN AT CENTERPOINT [STAR WARS CORELLIAN 3]  ALLEN, ROGER MACBRIDE  C/P:2,P
SHOWDOWN AT THE MALL [@SABRINA THE TEENAGE WITCH #2]  GALLAGHER, DIANA G.  P
SHRIEK: AN AFTERWORD [AMBERGRIS]  VANDERMEER, JEFF  LP
SHRINE OF LIGHT [DAUGHTERS OF BAST 2]  ISIDORE, SARAH  P
SHRINE OF STARS [@CONFLUENCE #3]  MCAULEY, PAUL J.  H
SHRINE OF THE DESERT MAGE [PARSINA 1]  GOLDIN, STEPHEN  P
SHRINKING MAN, THE  MATHESON, RICHARD  C/P:2,P
SHROUD, THE  COYNE, JOHN  P
SHROUD OF SHADOW [STRANDS OF STARLIGHT 3]  BAUDINO, GAEL  C/P:2,P
SHROUDED PLANET, THE  RANDALL, ROBERT  SR-P,C/P:2,H
SHROUDED REALM, THE [DRAGONREALM ORIGINS 1]  KNAAK, RICHARD A.  P
SHROUDED WALKS OF BORANGA, THE [@RO-LAN #2]  SIROTA, MIKE  P
SHUDDER PULPS, THE  JONES, ROBERT KENNETH  R/REF-H
SHUDDERCHILD  NORWOOD, WARREN  P
SHUTEYE FOR THE TIMEBROKER  DI FILIPPO, PAUL  LP
SHUTTERED ROOM AND OTHER TALES OF HORROR, THE  LOVECRAFT, H. P./DERLETH, AUGUST  C/P:2,P
SHUTTLE  ONLEY, DAVID C.  P
SHUTTLE  POWERS, ROBERT M.  [C/RAN-P]
SHUTTLE DOWN  CORREY, LEE  C/P:2,P
SHUTTLE INTO SPACE  STINE, G. HARRY  [C/RAN-H]
SHUTTLE PEOPLE, THE  BISHOP, GEORGE  P
SHY LEOPARDESS [NEUSTRIAN CYCLE 3]  BARRINGER, LESLIE  LP
SHY STEGOSAURUS OF CRICKET CREEK, THE  LAMPMAN, EVELYN SIBLEY  C/P,LP,H
SHY STEGOSAURUS OF INDIAN SPRINGS, THE  LAMPMAN, EVELYN SIBLEY  H
SIBLING, THE  HALL, ADAM  P
SIBYL SUE BLUE  BROWN, ROSEL GEORGE  H:2
SIDE SHOW [LUCKY 13TH]  SHELLEY, RICK  P
SIDESHOW [GALACTIC MIDWAY 1]  RESNICK, MIKE  C/P,P
SIDESHOW  TEPPER, SHERI S.  C/P,P,H
SIDESHOW  THOMPSON, W. R.  C/P:2,SFWA-P,P
SIDESLIP  WHITE, TED/VAN ARNAM, DAVE  C/P:2,P
SIDEWISE IN TIME  LEINSTER, MURRAY  SR-H
SIDHE DEVIL [DOC SIDHE 2]  ALLSTON, AARON  P
SIDNEY’S COMET  HERBERT, BRIAN  C/P:2,P
SIDONIA THE SORCERESS AND THE AMBER WITCH I  MEINHOLD, WILLIAM  H
SIDONIA THE SORCERESS AND THE AMBER WITCH II  MEINHOLD, WILLIAM  H
SIDURI’S NET [CLOUDSHIPS OF ORION 1]  MCALLISTER, P. K.  P
SIE KAMEN VON KENDARKOS [@PERRY RHODAN #142]  DARLTON, CLARK  FPR
SIEGE, THE [HELMSMAN 6]  BALDWIN, BILL  C/P,P
SIEGE  CORLEY, EDWIN  P
SIEGE, THE [STAR TREK DS9]  DAVID, PETER  C/P:2,P
SIEGE, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #2]  DENNING, TROY  P

The MITSFS Titledex  September 24, 2007  Page 324
TITLE  AUTHOR  CODE
SIEGE  HOYT, RICHARD  SFWA-P
SIEGE AT BLUE MOUNTAIN [@COMPLETE ELFQUEST #5]  PINI, WENDY/PINI, RICHARD  CX
SIEGE OF ARISTA, THE [@WAR YEARS #2]  FAWCETT, BILL  C/PA,PA
SIEGE OF ARRANDIN, THE [@ARRANDIN #1]  HERNIMAN, MARCUS  P
SIEGE OF DARKNESS [FORGOTTEN REALMS LEGACY OF THE DROW 3]  SALVATORE, R. A.  P
SIEGE OF DOME, THE [@EMPYRION #2]  LAWHEAD, STEPHEN R.  LP
SIEGE OF EARTH  FAUCETTE, JOHN M.  C/P:2,P
SIEGE OF ETERNITY, THE [ESCHATON 2]  POHL, FREDERIK  C/P,P,H
SIEGE OF FALTARA, THE  DARNAY, ARSEN  P
SIEGE OF HARLEM, THE  MILLER, WARREN  P
SIEGE OF MT. NEVERMIND, THE [@DRAGONLANCE CHAOS WAR 6]  RYAN, FERGUS  P
SIEGE OF ORBITOR  NEWMAN, RICHARD LOUIS  P
SIEGE OF SHADOWS  ABBEY, LYNN  C/P,P
SIEGE OF STONE [@SEARCHERS #3]  WILLIAMSON, CHET  C/P:2,P
SIEGE OF THE UNSEEN  VAN VOGT, A. E.  DD391
SIEGE OF WONDER, THE  GESTON, MARK S.  H,SFWA,P
SIEGE PERILOUS  DEL REY, LESTER  C/P:2,P
SIERRA MADRE [@WARBOTS #4]  STINE, G. HARRY  P
SIGHT OF PROTEUS [PROTEUS 1]  SHEFFIELD, CHARLES  C/P:2,P
SIGHTBLINDER’S STORY [@LOST SWORDS #2]  SABERHAGEN, FRED  H:2,P
SIGMA CURVE, THE  BARNWELL, WILLIAM  P
SIGN AT SIX, THE  WHITE, STEWART EDWARD  H
SIGN OF CHAOS [AMBER 8]  ZELAZNY, ROGER  C/P,H,@P
SIGN OF THE LABRYS  SAINT CLAIR, MARGARET  C/P:2,P
SIGN OF THE MOONBOW, THE [CORMAC 6]  OFFUTT, ANDREW J.  P
SIGN OF THE MUTE MEDUSA, THE  WALLACE, IAN  P
SIGN OF THE RAVEN, THE [@LAST VIKING #3]  ANDERSON, POUL  C/P,P
SIGN OF THE SHAPESHIFTER [@KNIGHTS OF THE SILVER DRAGON #3]  DONOVAN, DALE/JOHNS, LINDA  LP
SIGN OF THE UNICORN [AMBER 3]  ZELAZNY, ROGER  C/P:2,@P
SIGNAL SHATTERED, A [SIGNAL TO NOISE 2]  NYLUND, ERIC S.  H,P
SIGNAL TO NOISE [SIGNAL TO NOISE 1]  NYLUND, ERIC S.  C/P,P
SIGNALS  DEUTCHMAN, DEBORAH  H
SIGNALS  DEUTSCHMAN, DEBORAH  P
SIGNALS [@EXPLORATION #1]  RANDLE, KEVIN  C/P,P
SIGNS AND PORTENTS  YARBRO, CHELSEA QUINN  H,P
SIGNS OF LIFE  WILDER, CHERRY  P
SIGNS OF LOVE AND GLORY  HOLZER, HANS  [C/RAN-P]
SILENCE [@SMALLVILLE]  HOLDER, NANCY  P
SILENCE DESCENDS  CASE, GEORGE  LP
SILENCE IN SOLITUDE [SILENCE LEIGH 2]  SCOTT, MELISSA  C/P:2,P

SILENCE IN THE HEAVENS, A [@MECHWARRIOR DARK AGE PROVING GROUNDS 1/BATTLETECH] DELRIO, MARTIN

P

SILENCE IS DEADLY [JAN DARZEK]  BIGGLE, LLOYD JR.  H:2,SFWA
SILENCE OF GOM, THE [@PERRY RHODAN #39]  ACE  PA,C/PA
SILENCE OF THE LANGFORD, THE  LANGFORD, DAVID  SR-LP
SILENT BLADE, THE [FORGOTTEN REALMS PATHS OF DARKNESS 1]  SALVATORE, R. A.  H
SILENT CITY, THE  VONARBURG, ELISABETH  C/P,P
SILENT DANCES [@STARBRIDGE #2]  CRISPIN, A. C./O’MALLEY, KATHLEEN  C/P:2,SFWA-P:2,P
SILENT DEATH, THE  GRANT, MAXWELL  P
SILENT ENEMY, THE  JAHN, ERNEST T.  P
SILENT ENEMY, THE [@AGE OF CONAN SOLDIER’S QUEST #3]  KNAAK, RICHARD A.  P
SILENT GALAXY  TEDFORD, WILLIAM  P
SILENT GONDOLIERS, THE  GOLDMAN, WILLIAM  C/P,P
SILENT GONDOLIERS, THE (PROOF)  MORGENSTERN, S.  SR-LP
SILENT INTERVIEWS  DELANY, SAMUEL R.  R/REF-H
SILENT INVADERS, THE  SILVERBERG, ROBERT  C/P,DF195,SFWA-P,P
SILENT MOON  RELLING, WILLIAM JR.  SFWA-P
SILENT MULTITUDE, THE  COMPTON, D. G.  C/P:2,P
SILENT RUNNING  THOMPSON, HARLAN  SR-P
SILENT SKY, THE  BIGGLE, LLOYD JR.  P
SILENT SONGS [@STARBRIDGE #5]  CRISPIN, A. C./O’MALLEY, KATHLEEN  P,SFWA-P
SILENT STARS GO BY, THE  WHITE, JAMES  C/P:2,P
SILENT STRENGTH OF STONES, THE  HOFFMAN, NINA KIRIKI  P

SILENT STRIDERS [WORLD OF DARKNESS WEREOLF 2]/ BLACK FURIES [WORLD OF DARKNESS WEREOLF 2] BOWEN, CARL/FLEMING, GHERBOD

PA

SILENT THUNDER  ING, DEAN  TD31
SILENT TOWER, THE [WINDROSE 1]  HAMBLY, BARBARA  C/P:2,P
SILENT WAR, THE [ASTEROID WARS 3]  BOVA, BEN  P
SILENT WARRIOR, THE [FOREVER HERO 2]  MODESITT, L. E. JR.  C/P,P
SILENT ZONE [@ID4]  MOLSTAD, STEPHEN  H,P
SILICON DAGGER, THE  WILLIAMSON, JACK  H
SILICON DREAMS  GREENBERG, MARTIN H./SEGRIFF, LARRY  PA
SILICON MAGE, THE [WINDROSE 2]  HAMBLY, BARBARA  C/P:2,P
SILICON MAN, THE  PLATT, CHARLES  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 325
TITLE  AUTHOR  CODE
SILK  KIERNAN, CAITLIN R.  C/P,LP
SILK AND STEEL [BRONWYN 2]  MILLER, RON  C/P,P
SILK CODE, THE [PHIL D’AMATO 1]  LEVINSON, PAUL  H:2,P
SILK ROADS AND SHADOWS  SHWARTZ, SUSAN  C/P:2,SFWA-P,P
SILKEN SHROUD, THE [SILKEN MAGIC 2]  GILLIGAN, ELIZABETH  P
SILKIE, THE  VAN VOGT, A. E.  C/P:2,P
SILMARILLION, THE  TOLKIEN, J. R. R.  C/P,P,H
SILMARILLION, VOL. 1, THE (4 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
SILMARILLION, VOL. 2, THE (4 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
SILMARILLION, VOL. 3, THE (5 CDS) [LORD OF THE RINGS]  TOLKIEN, J. R. R.  SR-MM
SILVABAMBA  D’AMELIO, DAN  SR-P
SILVER BIRCH, BLOOD MOON [COMPLEMENTARY FAIRY COLORS 5]  DATLOW, ELLEN/WINDLING, TERRI  HA,LPA
SILVER BOUGH, THE  TUTTLE, LISA  H
SILVER BRANCH, THE [KELTIAD 3]  KENNEALY-MORRISON, PATRICIA  P
SILVER CALL, THE (OMNIBUS) [MITHGAR 10,11]  MCKIERNAN, DENNIS L.  P
SILVER CANYON, THE [@SPECTROS #4]  WINTERS, LOGAN  P
SILVER CHAIR, THE [NARNIA 4]  LEWIS, C. S.  C/P:2,P
SILVER CROWN, THE [WORLD OF DARKNESS RAGE]  BRIDGES, WILLIAM  C/P,P
SILVER CROWN, THE  O’BRIEN, ROBERT C.  C/P,LP
SILVER CROWN, THE [@GUARDIANS OF THE FLAME #3]  ROSENBERG, JOEL  C/P:2,P
SILVER EGGHEADS, THE  LEIBER, FRITZ  C/P:2,P
SILVER GLOVE, THE [VALENTINE MARSH 2]  CHARNAS, SUZY MCKEE  H
SILVER GRYPHON, THE [MAGE WARS 3]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
SILVER GRYPHON, THE  TURNER, GARY/HALPERN, MARTY  HA
SILVER HAND, THE [@SONG OF ALBION #2]  LAWHEAD, STEPHEN R.  P
SILVER HORSE, THE  LYNN, ELIZABETH A.  SFWA
SILVER LAKE, THE [@WARRIORS OF ESTAVIA #1]  PATTON, FIONA  H
SILVER LOCUSTS, THE  BRADBURY, RAY  SR-P
SILVER MAY TARNISH [WITCH WORLD]  NORTON, ANDRE/MCCONCHIE, LYN  H
SILVER METAL LOVER, THE [SILVER METAL LOVER 1]  LEE, TANITH  H:2,P
SILVER MISTRESS, THE [MODESTY BLAISE]  O’DONNELL, PETER  SFWA-P
SILVER MOONS, BLACK STEEL [WOLFWALKER 5]  HARPER, TARA K.  P
SILVER MOUNTAIN, THE  FRIESNER, ESTHER M.  C/P,P
SILVER NAILS [WARHAMMER GENEVIEVE 4]  YEOVIL, JACK  P
SILVER ON THE TREE [DARK IS RISING 5]  COOPER, SUSAN  C/P,P,H:2
SILVER PRINCESS, GOLDEN KNIGHT [ALEXIA AND TIRAN 1]  GREEN, SHARON  P,SFWA-P
SILVER SCREAM  SCHOW, DAVID J.  SFWA-P
SILVER SCREEN  ROBSON, JUSTINA  P
SILVER SHADOWS [@FORGOTTEN REALMS HARPERS #13]  CUNNINGHAM, ELAINE  P
SILVER SHIP AND THE SEA, THE  COOPER, BRENDA  H
SILVER SKULL, THE [DON SEBASTIAN 2]  DANIELS, LES  P
SILVER SPELL, THE [@KNIGHTS OF THE SILVER DRAGON #8]  BANERJEE, ANJALI  LP
SILVER SPIKE, THE [BLACK COMPANY]  COOK, GLEN  C/P:2,SFWA-P,P
SILVER STAIR, THE [@DRAGONLANCE BRIDGES OF TIME]  RABE, JEAN  P
SILVER STALLION, THE  CABELL, JAMES BRANCH  C/P,P,H
SILVER STONE, THE [@KEEPERS OF THE HIDDEN WAYS #2]  ROSENBERG, JOEL  C/P,P
SILVER SUN, THE [BOOKS OF THE ISLE 2]  SPRINGER, NANCY  C/P:2,P
SILVER SWORD, THE [LIGHTSTONE 2]  ZINDELL, DAVID  H
SILVER TOWER  BROWN, DALE  H,P
SILVER WARRIORS, THE [EREKOSE 2]  MOORCOCK, MICHAEL  C/P:2,P
SILVER WINGS AND LEATHER JACKETS [EAGLEHEART 1]  WESTCOTT, C. T.  P
SILVER WOLF, THE [WOLF RICE 2]  BORCHARDT, ALICE  C/P:2,P
SILVER WOLF, BLACK FALCON [MITHGAR 13]  MCKIERNAN, DENNIS L.  H
SILVERADO [@SPECTROS #1]  WINTERS, LOGAN  P
SILVERFALL [@FORGOTTEN REALMS]  GREENWOOD, ED  P
SILVERGLASS [SILVERGLASS 1]  RIVKIN, J. F.  C/P:2,P
SILVERHAIR [MAMMOTHS 1]  BAXTER, STEPHEN  C/P,P
SILVERHAIR THE WANDERER [WESTRIA 3]  PAXSON, DIANA L.  P
SILVERHAND [@ARCANA #1]  LLYWELYN, MORGAN/SCOTT, MICHAEL  P

SILVERHEART MOORCOCK, MICHAEL/ CONSTANTINE, STORM

H

SILVERLEAF SYNDROME, THE  ROBINSON, ELEANOR  P
SILVERLIGHT [@ARCANA #2]  LLYWELYN, MORGAN/SCOTT, MICHAEL  P
SILVERLOCK  MYERS, JOHN MYERS  SR-P,C/P:2
SILVERLOCK COMPANION, A  LERNER, FRED  R/REF-H
SILVER’S BANE [SIDHE AND SHADOWLANDS 2]  KELLEHER, ANNE  LP
SILVER’S LURE  KELLEHER, ANNE  LP
SILVERTHORN [RIFTWAR 3]  FEIST, RAYMOND E.  C/P,P,H,SFWA
SIMBUL’S GIFT, THE [@FORGOTTEN REALMS NOBLES #6]  ABBEY, LYNN  P
SIMPLY HUMAN  DOUGLAS, L. WARREN  P
SIMS  WILSON, F. PAUL  H
SIMULACRA, THE  DICK, PHILIP K.  C/P:2,P
SIMULACRON-3  GALOUYE, DANIEL F.  C/P:2,P
SIMULATIONS  JACOBSON, KARIE  HA,LPA

The MITSFS Titledex  September 24, 2007  Page 326
TITLE  AUTHOR  CODE
SIMULTANEOUS MAN, THE  BLUM, RALPH  C/P:2,P
SIN EATER, THE  WALTER, ELIZABETH  H
SIN EATERS, THE  WEINBERG, ROBERT  SR-LPA
SIN IN SPACE  JUDD, CYRIL  SR-P,C/P
SIN OF ORIGIN  BARNES, JOHN  P
SINAI TAPESTRY  WHITTEMORE, EDWARD  P
SINBAD AND THE EYE OF THE TIGER  HALL, JOHN RYDER  P
SINDBAD: THE THIRTEENTH VOYAGE  LAFFERTY, R. A.  H:2,LP
SINFUL ONES, THE  LEIBER, FRITZ  C/P:2,P
SING A SONG OF TREKKIN’  ROGOW, ROBERTA  [C/RAN-H]
SING THE FOUR QUARTERS [QUARTER SINGERS 1]  HUFF, TANYA  C/P:2,P
SING THE LIGHT [CANTRIX 1]  MARLEY, LOUISE  P
SING THE WARMTH [CANTRIX 2]  MARLEY, LOUISE  P
SINGER, THE  MILLER, CALVIN  LP
SINGER AND THE SEA, THE [WINTER OF THE WORLD]  ROHAN, MICHAEL SCOTT  P
SINGER ENIGMA, THE  MAXWELL, ANN  C/P:2,P
SINGER FROM THE SEA  TEPPER, SHERI S.  H
SINGER IN THE STONE, THE  WILLETT, JOHN  H
SINGER’S CROWN, THE [DURHYS FAMILY 1]  ISAAK, ELAINE  LP
SINGERS OF TIME, THE  POHL, FREDERIK/WILLIAMSON, JACK  C/P,LP,H
SINGING CITADEL, THE  MOORCOCK, MICHAEL  P
SINGING SANDS, THE  TEY, JOSEPHINE  P
SINGING STONES, THE  COULSON, JUANITA  DH077
SINGLE COMBAT [SYSTEMIC SHOCK 2]  ING, DEAN  C/P:2,P

SINGLE WHITE VAMPIRE SEEKS SAME GREENBERG, MARTIN H./ KOREN, BRITTANY A.

PA

SINGULAR TRAVELS, CAMPAIGNS AND ADVENTURES OF BARON MUNCHAUSEN RASPE, R. E.

LP

SINGULARITIES  QUICK, W. T.  P
SINGULARITY  SLEATOR, WILLIAM  C/P:2,P
SINGULARITY PROJECT, THE  BUSBY, F. M.  C/P,P
SINGULARITY SKY [SINGULARITY SKY 1]  STROSS, CHARLES  C/P,P,H
SINGULARITY STATION  BALL, BRIAN N.  C/P,P
SINISTER BARRIER  RUSSELL, ERIC FRANK  C/P:2,P,GN1.01
SINISTER FORCES  WESTWOOD, ALVIN  SR-P
SINISTER POWER, THE [PERRY RHODAN 124]  BRAND, KURT  SR-LP
SINISTER RESEARCHES OF C. P. RANSOM, THE  NEARING, H. JR.  P
SINISTRE  HATCH, GEORGE  LPA
SINK THE ARMADA! [@FREEDOM’S RANGERS #5]  ANDREWS, KEITH WILLIAM  P
SINNER [WAYFARER REDEMPTION 4]  DOUGLASS, SARA  P
SINNERS OF HWANG, THE  DEXTER, JOHN  SR-L
SIN’S DOORWAY  WELLMAN, MANLY WADE  H
SINS OF COMMISSION [STAR TREK TNG]  WRIGHT, SUSAN  C/P:2,P
SINS OF OMISSION  YARBRO, CHELSEA QUINN  C/P:2,P
SINS OF THE BLOOD  RUSCH, KRISTINE KATHRYN  P
SINS OF THE FATHER [BUFFY]  GOLDEN, CHRISTOPHER  C/P,P
SINS OF THE FATHERS [WORLD OF DARKNESS WRAITH]  CHUPP, SAM  P
SINS OF THE FATHERS [@GABRIEL KNIGHT]  JENSEN, JANE  P
SINS OF THE FATHERS, THE  SCHMIDT, STANLEY  P
SINS OF THE FLESH  DAVIS, DON/DAVIS, JAY  SFWA-P
SINS OF THE NIGHT [@DARK-HUNTER]  KENYON, SHERRILYN  P
SIOUX SPACEMAN, THE  NORTON, ANDRE  C/P:2,P,DD437
SIR APROPOS OF NOTHING [SIR APROPOS OF NOTHING 1]  DAVID, PETER  C/P,P
SIR GAWAIN AND THE GREEN KNIGHT/PEARL/SIR ORFEO  TOLKIEN, J. R. R.  HA,PA
SIR HAROLD AND THE GNOME KING [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE  LP
SIR HENRY  NATHAN, ROBERT  H,LP
SIR STALWART [KING’S DAGGERS 1]  DUNCAN, DAVE  P
SIR THURSDAY [@KEYS TO THE KINGDOM #4]  NIX, GARTH  LP
SIREN SONG [@STARGATE SG-1 #6]  SCOTT, HOLLY/DUNCAN, JAMES  P
SIREN STARS, THE  CARRIGAN, RICHARD/CARRIGAN, NANCY  C/P,P
SIRENS  DATLOW, ELLEN/WINDLING, TERRI  PA
SIRENS OF TITAN, THE  VONNEGUT, KURT JR.  C/P:2,P
SIRIAN EXPERIMENTS, THE [CANOPUS IN ARGOS 3]  LESSING, DORIS  H:2,LP
SIRIUS  STAPLEDON, OLAF  SR-P,C/P

SIRIUS THE DOG STAR GREENBERG, MARTIN H./ POTTER, ALEXANDER

PA

SISTER ALICE  REED, ROBERT  H
SISTER BLOOD  HABER, KAREN  P
SISTER EARTH  BREDE, ARNOLD  SR-P
SISTER EMILY’S LIGHTSHIP AND OTHER STORIES  YOLEN, JANE  H
SISTER LIGHT, SISTER DARK [ALTA 1]  YOLEN, JANE  P,SFWA-P
SISTER OF THE DEAD [NOBLE DEAD SERIES 1 3]  HENDEE, BARB/HENDEE, J. C.  P
SISTER OF THE SKY  ALLAN, MARGARET  P
SISTER OF THE SWORD [@DRAGONLANCE BARBARIANS #3]  THOMPSON, PAUL B./COOK, TONYA C.  P

TITLE  AUTHOR  CODE
SISTER TO THE RAIN  MICHAELS, MELISA C.  P
SISTERHOOD OF THE BLUE STORM [@OROKON #4]  ARDEN, TOM  H,P
SISTERS  BEAR, GREG  B
SISTERS IN FANTASY  SHWARTZ, SUSAN/GREENBERG, MARTIN H.  PA
SISTERS IN FANTASY 2  SHWARTZ, SUSAN/GREENBERG, MARTIN H.  PA
SISTERS OF THE NIGHT  HAMBLY, BARBARA/GREENBERG, MARTIN H. C/PA,LPA
SISTERS OF THE RAVEN [SISTERS OF THE RAVEN 1]  HAMBLY, BARBARA  H:2,LP
SITUATION RED: THE UFO SIEGE  STRINGFIELD, LEONARD H.  SFWA
SIVA!  RICHMOND, WALT/RICHMOND, LEIGH  C/P,P
6 AND THE SILENT SCREAM  HOWARD, IVAN  C/PA,PA
6 X H  HEINLEIN, ROBERT A.  C/P:2,P
SIX DECADES  DAVIS PUBLICATIONS  SR-LPA
SIX FAMILIES, THE [@GAMBLER’S STAR #1]  HOLDER, NANCY  P
6 FINGERS OF TIME, THE  MACFADDEN BOOKS  PA,C/PA
6 FROM WORLDS BEYOND  DIKTY, T. E.  PA,C/PA
SIX GATES FROM LIMBO  MCINTOSH, J. T.  C/P:2,P
6 GREAT SHORT NOVELS OF SCIENCE FICTION  CONKLIN, GROFF  PA,C/PA
SIX GREAT SHORT SCIENCE FICTION NOVELS  CONKLIN, GROFF  PA,C/PA
SIX-GUN PLANET  JAKES, JOHN  C/P,P
SIX-GUN SOLUTION, THE [TIMEWARS 12]  HAWKE, SIMON  C/P,P
668: THE NEIGHBOR OF THE BEAST [KENT MONTANA 5]  FENN, LIONEL  C/P,P
SIX MILLION DOLLAR MAN AND THE BIONIC WOMAN, THE  COHEN, JOEL H.  [C/RAN-P]
SIX MOON DANCE  TEPPER, SHERI S.  C/P,P,H:2
SIX NOVELS OF THE SUPERNATURAL  WAGENKNECHT, EDWARD  HA
SIX OF SWORDS  DOUGLAS, CAROLE NELSON  C/P:2,P
SIX SCIENCE FICTION PLAYS  ELWOOD, ROGER  C/PA,PA
SIX WORLDS YONDER  RUSSELL, ERIC FRANK  DD315,D77785
SIXERS  KAVANAGH, JOHN PATRICK  H
1635: THE CANNON LAW [SIXTEEN THIRTY-TWO]  FLINT, ERIC/DENNIS, ANDREW  H
1634: THE GALILEO AFFAIR [SIXTEEN THIRTY-TWO 3]  FLINT, ERIC/DENNIS, ANDREW  H:2
1633 [SIXTEEN THIRTY-TWO 2]  WEBER, DAVID/FLINT, ERIC  H
1632 [SIXTEEN THIRTY-TWO 1]  FLINT, ERIC  H
1634: THE RAM REBELLION [SIXTEEN THIRTY-TWO]  FLINT, ERIC/DEMARCE, VIRGINIA  H
SIXTH COLUMN  HEINLEIN, ROBERT A.  C/P:2,SFWA-P,P
SIXTH DIRECTORATE, THE  HONE, JOSEPH  P
SIXTH WINTER, THE  ORGILL, DOUGLAS/GRIBBIN, JOHN  H
SIXTY DAYS AND COUNTING [FORTY SIGNS OF RAIN 3]  ROBINSON, KIM STANLEY  H
’68  SCAEVOLA, PETER  H
SKAVENSLAYER [WARHAMMER GOTREK & FELIX 2]  KING, WILLIAM  P
SKEEN’S LEAP [SKEEN 1]  CLAYTON, JO  C/P:2,P
SKEEN’S RETURN [SKEEN 2]  CLAYTON, JO  C/P:2,P
SKEEN’S SEARCH [SKEEN 3]  CLAYTON, JO  C/P:2,P
SKELETON CLOSET OF JULES DE GRANDIN, THE  QUINN, SEABURY  C/P:2,P
SKELETON CREW  KING, STEPHEN  C/P,P,H:2
SKELETON KEY [@PRIVATE SCHOOL #4]  CHARLES, STEVEN  P
SKELETON LORD’S KEY, THE [KEYS TO PARADISE 2]  VARDEMAN, ROBERT E.  C/P,SFWA-P,P
SKEWED THRONE, THE [VARIS 1]  PALMATIER, JOSHUA  H
SKIES DISCROWNED, THE  POWERS, TIM  SR-P
SKIES OF PERN, THE [PERN]  MCCAFFREY, ANNE  H,P
SKIN [@X-FILES]  MEZRICH, BEN  H
SKIN AND BONES  SMITH, THORNE  H:2,P
SKIN DEEP [@PHOTON #6]  PETERS, DAVID  P
SKIN FOLK  HOPKINSON, NALO  LP
SKIN GAME [@DARK ANGEL]  COLLINS, MAX ALLAN  C/P,P
SKIN TRADE, THE  WINTER, DOUGLAS E.  PA
SKINNER, THE [POLITY SKINNER 1]  ASHER, NEAL  H:2,LP
SKINNER [FAR STARS AND FUTURE TIMES]  MCENROE, RICHARD S.  C/P:2,P
SKINNER  RAE, HUGH C.  P
SKINNY DIPPING IN THE LAKE OF THE DEAD  DENIRO, ALAN  LP
SKIRMISH [SKYRIDER 1]  MICHAELS, MELISA C.  C/P,SFWA-P:2,P
SKIRMISH  SIMAK, CLIFFORD D.  C/P,P,H
SKOOK, THE  MILLER, J. P.  H,P
SKRAELINGS  SHERRELL, CARL  P
SKRAYLING TREE, THE [ELRIC/VON BEK]  MOORCOCK, MICHAEL  P
SKRYLING’S BLADE [RUNESWORD 2]  ESTES, ROSE/WHAM, TOM  C/P,SFWA-P:2,P
SKULL-FACE  HOWARD, ROBERT E.  P
SKULL-FACE AND OTHERS  HOWARD, ROBERT E.  SR-H
SKULL GATE [FROST 2]  BAILEY, ROBIN WAYNE  C/P,SFWA-P,P
SKULL OF THE WORLD, THE [WITCHES OF EILEANAN 5]  FORSYTH, KATE  C/P,P
SKULLS IN THE STARS [SOLOMON KANE 4]  HOWARD, ROBERT E.  P
SKY BLOCK, THE  FRAZEE, STEVE  C/P,P
SKY COYOTE [COMPANY 2]  BAKER, KAGE  P
SKY IS FALLING, THE  DEL REY, LESTER  D76960
SKY IS FALLING, THE/BADGE OF INFAMY  DEL REY, LESTER  SR-P,C/P:2

The MITSFS Titledex  September 24, 2007  Page 328
TITLE  AUTHOR  CODE
SKY IS FILLED WITH SHIPS, THE  MEREDITH, RICHARD C.  P
SKY ISLAND  BAUM, L. FRANK  H
SKY LORDS, THE [SKY LORDS 1]  BROSNAN, JOHN  SFWA
SKY OF SWORDS [KING’S BLADES]  DUNCAN, DAVE  H
SKY PEOPLE, THE  STIRLING, S. M.  H
SKY PIRATES! [DOCTOR WHO NEW ADVENTURES]  STONE, DAVE  C/P:2,P
SKY PIRATES OF CALLISTO [JANDAR 3]  CARTER, LIN  C/P,P
SKY ROAD, THE [FALL REVOLUTION]  MACLEOD, KEN  H
SKY ROAD  ZEDDIES, ANN TONSOR  P
SKY SO BIG AND BLACK, THE  BARNES, JOHN  H:2
SKY TRILLIUM [TRILLIUM]  MAY, JULIAN  H,P
SKY WALKER, THE [@AVENGER #3]  ROBESON, KENNETH  C/P:2,P
SKYBORN  KRUCHTEN, MARCIA H.  P
SKYBOWL [DRAGON STAR 3]  RAWN, MELANIE  C/P,SFWA-P,H,P,SFWA
SKYBREAKER [MATT CRUSE 2]  OPPEL, KENNETH  H
SKYCLIMBER  GALLUN, RAYMOND Z.  P
SKYDANCER  ARCHER, GEOFFREY  P,SFWA-P
SKYDARK SPAWN [@DEATHLANDS]  AXLER, JAMES  P
SKYFALL [SKOLIAN EMPIRE 9]  ASARO, CATHERINE  P
SKYFALL  HARRISON, HARRY  C/P,SFWA-P,P,H
SKYLARK DUQUESNE [SKYLARK 4]  SMITH, E. E.  C/P:2,P
SKYLARK OF SPACE, THE [SKYLARK 1]  SMITH, E. E.  C/P:2,P,H
SKYLARK OF VALERON [SKYLARK 3]  SMITH, E. E.  C/P,P,H:2
SKYLARK THREE [SKYLARK 2]  SMITH, E. E.  C/P:2,P,H
SKYLOCK  KOZERSKI, PAUL  P
SKYNAPPERS, THE  BRUNNER, JOHN  DD457
SKYPORT  SIODMAK, CURT  C/P,P,H
SKYRIPPER  DRAKE, DAVID  C/P,P
SKYROCKET, THE  HADER, BERTA/HADER, ELMER  H
SKYROCKET STEELE  GOULART, RON  P

SKYROCKET STEELE CONQUERS THE UNIVERSE [@AUTHOR’S CHOICE MONTHLY #11] GOULART, RON

SR-H

SKYSCRAPER DOOM, THE  ZIEROLD, NORMAN  SFWA
SKYSTONE, THE [CAMULOD 1]  WHYTE, JACK  P
SKYWALKING  POLLOCK, DALE  R/REF-P,C/REF-P
SLAG LIKE ME [@ALIEN NATION #5]  LONGYEAR, BARRY B.  C/P,P
SLAMMERS DOWN! [HAMMER’S SLAMMERS]  JOHNSON, TODD  P,SFWA-P
SLAN  VAN VOGT, A. E.  C/P,P,H
SLAN HUNTER  VAN VOGT, A. E./ANDERSON, KEVIN J.  H
SLANT [QUEEN OF ANGELS 4]  BEAR, GREG  C/P:2,P
SLAPSTICK  VONNEGUT, KURT JR.  P
SLATER’S PLANET  MOORE, HARRIS  C/P:2,P
SLAUGHTER OF THE INNOCENT  RUESCH, HANS  [C/RAN-P]
SLAUGHTERHOUSE-FIVE, OR THE CHILDREN’S CRUSADE  VONNEGUT, KURT JR.  C/P:2,P
SLAVE AND THE FREE, THE [HOLDFAST 1,2]  CHARNAS, SUZY MCKEE  LP
SLAVE GIRL OF GOR [GOR 11]  NORMAN, JOHN  C/P,P
SLAVE MASTERS [SLAVE TRADE 2]  WRIGHT, SUSAN  P
SLAVE OF FRANKENSTEIN, THE  MYERS, ROBERT J.  P
SLAVE OF SARMA [@BLADE #4]  LORD, JEFFREY  C/P:2,P
SLAVE PLANET  JANIFER, LAURENCE M.  C/P:2,P
SLAVE SHIP [@STAR WARS BOUNTY HUNTER WARS 2]  JETER, K. W.  C/P,P
SLAVE SHIP  POHL, FREDERIK  C/P:2,P
SLAVE SHIP FROM SERGAN [@CAP KENNEDY #2]  KERN, GREGORY  C/P,P
SLAVE TRADE [SLAVE TRADE 1]  WRIGHT, SUSAN  P
SLAVE TRADERS OF THE SKY  FUTURE, STEVE  SR-P
SLAVERS OF SPACE  BRUNNER, JOHN  DD421
SLAVES OF HEAVEN, THE  COOPER, EDMUND  H:2,SFWA,P
SLAVES OF IJAX  FEARN, JOHN RUSSELL  SR-P
SLAVES OF LOMOORO, THE  AUGUSTUS, ALBERT JR.  SR-P
SLAVES OF REGLATHIUM, THE [@DANNUS #5]  SIROTA, MIKE  P
SLAVES OF SLEEP  HUBBARD, L. RON  C/P:2,P
SLAVES OF SPACE [YOLANDA 2]  VERSEAU, DOMINIQUE  SR-L
SLAVES OF THE BLOODWOLVES  WEINBERG, ROBERT  SR-LPA
SLAVES OF THE KLAU  VANCE, JACK  DD295
SLAVES OF THE VOLCANO GOD [CINEVERSE 1]  GARDNER, CRAIG SHAW  C/P:2,SFWA-P,P
SLAVES UNCHAINED [SLAVE TRADE 3]  WRIGHT, SUSAN  P
SLAY AND RESCUE  MOORE, JOHN  C/P:2,SFWA-P,P
SLAY RIDE  HITCHCOCK, ALFRED  C/PA,PA
SLAYER OF SOULS, THE  CHAMBERS, ROBERT W.  H
SLEEP, THE  CREASEY, JOHN  P
SLEEP EATERS, THE  LYMINGTON, JOHN  P
SLEEP NO MORE  DERLETH, AUGUST  HA
SLEEP OF REASON, THE [@DOCTOR WHO EIGHTH DOCTOR]  DAY, MARTIN  P
SLEEP OF STONE, THE  COOPER, LOUISE  C/P,SFWA,P

The MITSFS Titledex  September 24, 2007  Page 329
TITLE  AUTHOR  CODE
SLEEP TWO, THREE, FOUR!  NEUFELD, JOHN  C/P,P
SLEEPERS, THE [@PERRY RHODAN #79]  ACE  PA
SLEEPERS OF MARS  WYNDHAM, JOHN  SR-P,C/P
SLEEPING CITY, THE [MALKAR 2]  TUBB, E. C.  LP
SLEEPING DRAGON, THE [@GUARDIANS OF THE FLAME #1]  ROSENBERG, JOEL  C/P:2,P
SLEEPING IN FLAME  CARROLL, JONATHAN  P,LP
SLEEPING PLANET  BURKETT, WILLIAM R. JR.  C/P,P,H
SLEEPING SORCERESS, THE [ELRIC]  MOORCOCK, MICHAEL  C/P:2,P
SLEEPING WITH THE FISHES  DAVIDSON, MARYJANICE  P
SLEEPLESS NIGHTS IN THE PROCRUSTEAN BED  ELLISON, HARLAN  R/REF-P
SLEEPWALK  SAUL, JOHN  P
SLEEPWALKER’S WORLD  DICKSON, GORDON R.  C/P,P,H
SLEEPY [DOCTOR WHO NEW ADVENTURES]  ORMAN, KATE  C/P:2,P
SLEIPNIR  EVANS, LINDA  P,SFWA-P
SLIDERS  LINAWEAVER, BRAD  P
SLIDING SCALES [HUMANX FLINX]  FOSTER, ALAN DEAN  P
SLIGHT DETOUR, A [@STARSHIP TROUPERS #3]  STASHEFF, CHRISTOPHER  C/P:2,P
SLIPBACK [DOCTOR WHO]  SAWARD, ERIC  SR-P,C/P
SLIPPAGE  ELLISON, HARLAN  LP
SLIPPERY MAP, THE (PROOF)  BODE, N. E.  SR-LP
SLIPSTREAMS  GREENBERG, MARTIN H./HELFERS, JOHN  PA
SLIPT  FOSTER, ALAN DEAN  C/P:2,P
SLOB  MILLER, REX  P
SLOW BURNER  HAGGARD, WILLIAM  H
SLOW EMPIRE, THE [DOCTOR WHO EIGHTH DOCTOR]  STONE, DAVE  P
SLOW FALL TO DAWN  LEIGH, STEPHEN  C/P:2,P
SLOW FREIGHT  BUSBY, F. M.  C/P:2,P
SLOW FUNERAL  ORE, REBECCA  H,P

SLOW LIGHTNING POPKES, STEVEN TD51170,

SFWA-TD51170

SLOW RIVER  GRIFFITH, NICOLA  H,LP
SMALL ARMAGEDDON, A  ROSHWALD, MORDECAI  P
SMALL CHANGES  CLEMENT, HAL  H
SMALL COLONIAL WAR, A [VERESHCHAGIN 1]  FREZZA, ROBERT  C/P:2,P
SMALL GODS [DISCWORLD 13]  PRATCHETT, TERRY  H:2,P
SMALL GODS AND MR. BARNUM, THE  TRELL, MAX  H
SMALL SHADOWS CREEP  NORTON, ANDRE  HA
SMALL WORLD  KING, TABITHA  H,P

SMART DRAGONS, FOOLISH ELVES FOSTER, ALAN DEAN/

GREENBERG, MARTIN H.

C/PA,PA,SFWA-P

SMASHED WORLD, THE  SLATER, HENRY J.  H
SMILE ON THE VOID  GORDON, STUART  C/P,SFWA,P,H
SMILING DOGS, THE [@AVENGER #10]  ROBESON, KENNETH  C/P:2,P
SMILLA’S SENSE OF SNOW  HOEG, PETER  C/P,P
SMITH OF WOOTTON MAJOR  TOLKIEN, J. R. R.  H
SMITH OF WOOTTON MAJOR/FARMER GILES OF HAM  TOLKIEN, J. R. R.  C/P,P,H
SMITHSONIAN INSTITUTION, THE  VIDAL, GORE  H
SMOG, THE  CREASEY, JOHN  P
SMOKE  WESTLAKE, DONALD E.  P
SMOKE AND ASHES [TONY FOSTER 3]  HUFF, TANYA  H:2
SMOKE AND MIRRORS  GAIMAN, NEIL  H
SMOKE AND MIRRORS [TONY FOSTER 2]  HUFF, TANYA  H
SMOKE AND MIRRORS  LINDSKOLD, JANE  P
SMOKE AND SHADOWS [TONY FOSTER 1]  HUFF, TANYA  H
SMOKE FROM THE ASHES [ASHES 7]  JOHNSTONE, WILLIAM W.  P
SMOKE ON THE WATER [@GAMMALAW 1]  DALEY, BRIAN  C/P,P
SMOKE RING, THE [INTEGRAL TREES 2]  NIVEN, LARRY  H:2,P,SFWA
SMUGGLED ATOM BOMB, THE  WYLIE, PHILIP  C/P,P
SMUGGLERS, THE [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P
SMUGGLER’S GOLD [@MEROVINGEN NIGHTS #4]  CHERRYH, C. J.  PA,C/PA
SMUGGLER’S REEF [@RICK BRANT #7]  BLAINE, JOHN  H
SNAIL  MILLER, RICHARD  LP
SNAIL ON THE SLOPE, THE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS C/P,P
SNAKE, THE  POWELL, F. INGLIS  H
SNAKE OIL WARS, THE [BARION AND COYUL 2]  GODWIN, PARKE  H
SNAKEDANCE [DOCTOR WHO]  DICKS, TERRANCE  SR-P
SNARE  KERR, KATHARINE  H,P
SNARES OF IBEX, THE [DIADEM 8]  CLAYTON, JO  C/P:2,P
SNARKOUT BOYS AND THE AVOCADO OF DEATH, THE  PINKWATER, DANIEL MANUS  P
SNEAK PREVIEW  BLOCH, ROBERT  C/P:2,P
SNEAK PREVIEW WINTER 2000  WIZARDS OF THE COAST  SR-P
SNEEZE ON SUNDAY  NORTON, ANDRE/HOGARTH, GRACE ALLEN H
SNOW, THE  ROBERTS, ADAM  LP
SNOW CRASH  STEPHENSON, NEAL  C/P:4,P,H

TITLE  AUTHOR  CODE
SNOW-EYES [SNOW-EYES 1]  SMITH, STEPHANIE A.  C/P,P,H
SNOW FURY  HOLDEN, RICHARD CORT  H:2,P
SNOW KILL [@FREEDOM’S RANGERS #6]  ANDREWS, KEITH WILLIAM  P
SNOW QUEEN, THE [SNOW QUEEN 1]  VINGE, JOAN D.  C/P,SFWA-P,H,P,SFWA
SNOW SHADOW  NORTON, ANDRE  P
SNOW WHITE AND ROSE RED  WREDE, PATRICIA C.  H:2,P
SNOW WHITE AND THE GIANTS  MCINTOSH, J. T.  C/P,P
SNOW WHITE AND THE SEVEN SAMURAI  HOLT, TOM  C/P,P
SNOW WHITE, BLOOD RED [COMPLEMENTARY FAIRY COLORS 1]  DATLOW, ELLEN/WINDLING, TERRI  SR-HA,PA,C/PA,SFWA
SNOWBROTHER [FIFTH MILLENNIUM]  STIRLING, S. M.  C/P:2,P
SNOWCASTLES  MCGEARY, DUNCAN  P
SNOWFALL [SNOWFALL 1]  SMITH, MITCHELL  P
SNOWMAN  BOGNER, NORMAN  P
SNOWMAN IN FLAMES [@PERRY RHODAN #25]  ACE  PA,C/PA
SNOWS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
SNOWS OF GANYMEDE, THE  ANDERSON, POUL  DD303
SNOWS OF JASPRE, THE  CARAKER, MARY  H,P
SNUFF FICTION  RANKIN, ROBERT  P
SO BEAUTIFUL AND SO DANGEROUS  MCKIE, ANGUS  CX
SO BRIGHT THE VISION  SIMAK, CLIFFORD D.  C/P,DH095
SO CLOSE TO HOME  BLISH, JAMES  C/P:2,P
SO DARK A HERITAGE  LONG, FRANK BELKNAP  P
SO LONG, AND THANKS FOR ALL THE FISH [HITCHHIKERS 4]  ADAMS, DOUGLAS  C/P,P,H
SO LONG BEEN DREAMING  HOPKINSON, NALO/MEHAN, UPPINDER  LPA
SO MOSES WAS BORN  GRANT, JOAN  P
SO THIS IS HOW IT ENDS [@AVATARS #1]  SUTHERLAND, TUI T.  H
SO VILE A SIN [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN/ORMAN, KATE  C/P,P
SO YOU WANT TO BE A DOCTOR  NOURSE, ALAN E.  [C/RAN-P]
SO YOU WANT TO BE A WIZARD [YOUNG WIZARDS 1]  DUANE, DIANE  C/P,P,H
SOARER’S CHOICE [COREAN 6]  MODESITT, L. E. JR.  H
SOCHINENIE (COLLECTED WORKS) I  YEFREMOV, IVAN  FH
SOCHINENIE (COLLECTED WORKS) II  YEFREMOV, IVAN  FH
SOCHINENIE (COLLECTED WORKS) III BOOK 1  YEFREMOV, IVAN  FH
SOCHINENIE (COLLECTED WORKS) III BOOK 2  YEFREMOV, IVAN  FH
SOCIETA DEL TEMPO, LA  BRUNNER, JOHN  FP
SOCIETY, THE  SAINTCROW, LILITH  LP
SOCIETY OF THE MIND  HARRY, ERIC L.  P
SODOM AND GOMORRAH BUSINESS, THE  MALZBERG, BARRY N.  P
SOFT AND OTHERS  WILSON, F. PAUL  H,P
SOFT COME THE DRAGONS  KOONTZ, DEAN R.  D13793
SOFT KILL, THE  FREE, COLIN  C/P,P
SOFT TARGETS  ING, DEAN  C/P,P,H
SOFT WHISPER OF THE DEAD, THE  GRANT, CHARLES L.  SR-H
SOFTLY BY MOONLIGHT  FANE, BRON  SR-P
SOFTLY WALKS THE BEAST  HUNTER, THOMAS O’D.  P
SOFTWAR  BRETON, THIERRY/BENEICH, DENIS  H,SR-LP
SOFTWARE [WARE 1]  RUCKER, RUDY  C/P:2,P
SOJAN  MOORCOCK, MICHAEL  SR-LP
SOJARR OF TITAN  WELLMAN, MANLY WADE  SR-P,C/P
SOJOURN [@FORGOTTEN REALMS DARK ELF 3]  SALVATORE, R. A.  C/P,P
SOLAR ASSASSINS [@PERRY RHODAN #49]  ACE  PA,C/PA
SOLAR INVASION, THE  WELLMAN, MANLY WADE  SR-P,C/P:2
SOLAR KILL [@SAND WARS #1]  INGRID, CHARLES  C/P:2,P
SOLAR LOTTERY  DICK, PHILIP K.  C/P,DD103,SFWA-P,P

SOLAR POWER FROM SATELLITES COMMITTEE ON AERONAUTICAL AND SPACE

SCIENCES

[C/RAN-H]

SOLAR SYSTEM AND BACK, THE  ASIMOV, ISAAC  [C/RAN-P]
SOLARIANS, THE  SPINRAD, NORMAN  C/P:2,P
SOLARIS  LEM, STANISLAW  H:2,P
SOLARIS BOOK OF NEW SCIENCE FICTION, THE  MANN, GEORGE  PA
SOLARSTATION & KELWITTS STERN  ESCHBACH, ANDREAS  FP
SOLAY  HARRIS, MARK JONATHAN  SFWA
SOLDADO ANT, THE  JENKINS, WILL F.  B
SOLDIER, ASK NOT [CHILDE CYCLE]  DICKSON, GORDON R.  C/P:2,P
SOLDIER BOY  SHAARA, MICHAEL  C/P,P
SOLDIER FOR THE EMPIRE [@STAR WARS DARK FORCES 1]  DIETZ, WILLIAM C.  H
SOLDIER OF ANOTHER FORTUNE [@DESTINY MAKERS #3]  SHUPP, MIKE  P
SOLDIER OF ARETE [SOLDIER OF THE MIST 2]  WOLFE, GENE  H:2,SFWA-P,P
SOLDIER OF FORTUNE [@CASCA #8]  SADLER, BARRY  P
SOLDIER OF GIDEON [@CASCA #20]  SADLER, BARRY  P
SOLDIER OF LIGHT  DE LANCIE, JOHN/COOL, TOM  H,P
SOLDIER OF SIDON  WOLFE, GENE  H
SOLDIER OF THE MIST [SOLDIER OF THE MIST 1]  WOLFE, GENE  H:2,P
SOLDIERS  DALMAS, JOHN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 331
TITLE  AUTHOR  CODE

SOLDIERS LIVE [@BLACK COMPANY GLITTERING STONE 4/ BLACK COMPANY 9] COOK, GLEN

H:2

SOLDIERS OF FEAR, THE [STAR TREK INVASION! 2/STAR TREK TNG] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

SOLDIERS OF ICE [@FORGOTTEN REALMS HARPERS #7]  COOK, DAVID  P
SOLDIERS OF PARADISE [STARBRIDGE CHRONICLES 1]  PARK, PAUL  C/P,SFWA,P,H
SOLE SURVIVOR [SHATTERZONE 2]  HENSLEY, SHANE LACY  P
SOLE SURVIVOR  KOONTZ, DEAN R.  P
SOLEIL DE GLACE, LE  LIMAT, MAURICE  FP
SOLID GOLD KIDNAPPING [@SIX MILLION DOLLAR MAN #2]  RICHARDS, EVAN  P
SOLIP:SYSTEM  WILLIAMS, WALTER JON  SR-H
SOLIS  ATTANASIO, A. A.  H,P
SOLITAIRE  ESKRIDGE, KELLEY  H
SOLITARY MAN [@ANGEL]  MARIOTTE, JEFF  P
SOLO COMMAND [@STAR WARS X-WING 7]  ALLSTON, AARON  P
SOLO FLIGHTS THROUGH SHARED WORLDS  RESNICK, MIKE  LP
SOLO KILL  BOULT, S. KYE  C/P,P
SOLOMON KANE [SOLOMON KANE 3]  HOWARD, ROBERT E.  P
SOLOMON’S KNIFE  KOMAN, VICTOR  H
SOL’S CHILDREN  RABE, JEAN/GREENBERG, MARTIN H.  PA
SOLSTICE  HEWSON, DAVID  H
SOLUTION T-25  DUBOIS, THEODORA  SR-P,C/P:2
SOLUTION THREE  MITCHISON, NAOMI  H,P

SOME ASSEMBLY REQUIRED [@STAR TREK S.C.E. #3] BRODEUR, GREG/

CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ GALANTER, DAVE/JOLLEY, DAN/ ROSENBERG, AARON

P

SOME DREAMS ARE NIGHTMARES  GUNN, JAMES E.  H,SFWA
SOME GOLDEN HARBOR [RCN 5]  DRAKE, DAVID  H
SOME OF YOUR BLOOD  STURGEON, THEODORE  C/P:2,P
SOME SUMMER LANDS [ATLAN]  GASKELL, JANE  C/P,P,H
SOME THINGS DARK AND DANGEROUS  KAHN, JOAN  HA
SOME THINGS STRANGE AND SINISTER  KAHN, JOAN  HA
SOME THINGS WEIRD AND WICKED  KAHN, JOAN  PA
SOME TIME NEVER  DAHL, ROALD  H
SOME WILL NOT DIE  BUDRYS, ALGIS  C/P,P
SOME WILL NOT DIE (REVISED)  BUDRYS, ALGIS  LP
SOMEBODY COME AND PLAY  MCNALLY, CLARE  SFWA-P
SOMEONE COMES TO TOWN, SOMEONE LEAVES TOWN  DOCTOROW, CORY  H
SOMEONE IN THE DARK  DERLETH, AUGUST  C/P,P
SOMEONE TO WATCH OVER ME  SULLIVAN, TRICIA  P
SOMEPLACE TO BE FLYING [NEWFORD]  DE LINT, CHARLES  C/P:2,P
SOMERSET DREAMS AND OTHER FICTIONS  WILHELM, KATE  H,SFWA
SOMETHING ABOUT EVE  CABELL, JAMES BRANCH  C/P,P
SOMETHING FROM THE NIGHTSIDE [NIGHTSIDE 1]  GREEN, SIMON R.  C/P,P
SOMETHING MORE  CORNELL, PAUL  LP
SOMETHING M.Y.T.H. INC. [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P,P
SOMETHING NEAR  DERLETH, AUGUST  SR-H
SOMETHING RICH AND STRANGE  MCKILLIP, PATRICIA A.  LP
SOMETHING ROTTEN [THURSDAY NEXT 4]  FFORDE, JASPER  H
SOMETHING WICCAN THIS WAY COMES [@CHARMED]  HARRISON, EMMA  P
SOMETHING WICKED THIS WAY COMES  BRADBURY, RAY  C/P,P,H
SOMETHING’S CALLING ME HOME  O’CALLAGHAN, MAXINE  P
SOMETIME AFTER THE EQUINOX  BENT, JORJ  P
SOMETIME, NEVER  BALLANTINE BOOKS  SR-PA,C/PA
SOMETIME NEVER... [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
SOMETIMES, AFTER SUNSET  LEE, TANITH  H
SOMETIMES THE MAGIC WORKS  BROOKS, TERRY  R/REF-H
SOMEWHERE A VOICE  RUSSELL, ERIC FRANK  C/P:2,P
SOMEWHERE EAST OF LIFE  ALDISS, BRIAN W.  H
SOMEWHERE OUT THERE  FANE, BRON  SR-P
SON OF A WITCH  MAGUIRE, GREGORY  H
SON OF AVONAR [BRIDGE OF D’ARNATH 1]  BERG, CAROL  P
SON OF DARKNESS  SHERMAN, JOSEPHA  C/P,P,H
SON OF FLUBBER  FULLER, ROGER  P
SON OF MAN  SILVERBERG, ROBERT  C/P:2,P
SON OF ORIGINS OF MARVEL COMICS [MARVEL COMICS]  LEE, STAN  CX
SON OF SPELLSINGER [SPELLSINGER 7]  FOSTER, ALAN DEAN  C/P:2,SFWA-P,P
SON OF SUMMER STARS, THE [FIREBRINGER 3]  PIERCE, MEREDITH ANN  P
SON OF TARZAN, THE [TARZAN #4]  BURROUGHS, EDGAR RICE  C/P,@P,H
SON OF THE HERO [@VARAYAN #1]  SHELLEY, RICK  P
SON OF THE MORNING AND OTHER STORIES  GOTLIEB, PHYLLIS  P
SON OF THE SHADOWS [SEVENWATERS 2]  MARILLIER, JULIET  H
SON OF THE STARS  JONES, RAYMOND F.  H:2

The MITSFS Titledex  September 24, 2007  Page 332
TITLE  AUTHOR  CODE
SON OF THE SWORD [DYLAN MATHESON 1]  LEE, J. ARDIAN  LP
SON OF THE TREE  VANCE, JACK  D77525,DF265
SON OF THE TREE AND OTHER STORIES [@VANCE INTEGRAL EDITION #5]  VANCE, JACK  SR-H
SON OF THUNDER [@FORGOTTEN REALMS FIGHTERS]  LEEDER, MURRAY J. D.  P
SONG FOR ARBONNE, A  KAY, GUY GAVRIEL  C/P:2,SFWA-P,P
SONG FOR LYA AND OTHER STORIES, A  MARTIN, GEORGE R. R.  C/P:2,P
SONG IN THE DARK [@VAMPIRE FILES 11]  ELROD, P. N.  H
SONG IN THE SILENCE [LANEN KAELAR 1]  KERNER, ELIZABETH  C/P,P
SONG OF CTHULHU  RAINEY, STEPHEN MARK  LPA
SONG OF HOMANA, THE [CHEYSULI 2]  ROBERSON, JENNIFER  C/P:2,P
SONG OF KALI  SIMMONS, DAN  C/P,SFWA-P,P,H
SONG OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 1]  TEPPER, SHERI S.  C/P:2,P
SONG OF ORPHEUS [BEAUTY AND THE BEAST]  HAMBLY, BARBARA  C/P:3,P
SONG OF RHIANNON, THE [MABINOGI 3]  WALTON, EVANGELINE  C/P,SFWA,P
SONG OF ROLAND, THE  LUQUIENS, FREDERICK BLISS  LP
SONG OF SORCERY [ARGONIA 1]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
SONG OF STONE, A  BANKS, IAIN M.  P
SONG OF SUSANNAH [@DARK TOWER #6]  KING, STEPHEN  H
SONG OF THE AXE, THE [PELBAR 6]  WILLIAMS, PAUL O.  C/P:2,P
SONG OF THE BEAST  BERG, CAROL  P
SONG OF THE DWARVES [SONG OF THE DWARVES 1]  GUNNARSSON, THORARINN  P
SONG OF THE KINGDOM  STONE, ANDY  P
SONG OF THE PEARL  NICHOLS, RUTH  P
SONG OF THE SAURIALS [FORGOTTEN REALMS FINDER’S STONE 3]  NOVAK, KATE/GRUBB, JEFF  P
SONG OF THE SWAN, THE  D’ALEMBERT, ARTHUR  LP
SONG OF THE SWAN, PART 2  D’ALEMBERT, ARTHUR  LP
SONG OF THE VOLSUNGS AND THE NIBELUNGS  ANONYMOUS  P
SONG OF TIME [MAGIC: THE GATHERING]  MCLAREN, TERI  P
SONG OF UNMAKING [AURELIAN EMPIRE 2]  BRENNAN, CAITLIN  LP
SONGBIRDS OF PAIN, THE  KILWORTH, GARRY  P
SONGMASTER  CARD, ORSON SCOTT  C/P,SFWA,P,H
SONGS FROM THE SEASHELL ARCHIVES [ARGONIA 1,2]  SCARBOROUGH, ELIZABETH ANN  P
SONGS FROM THE STARS  SPINRAD, NORMAN  H:2,P
SONGS FROM THE STARS (PROOF)  SPINRAD, NORMAN  SR-LP
SONGS OF CHAOS  LEWITT, SHARIANN  C/P,SFWA-P:2,P
SONGS OF DISTANT EARTH, THE  CLARKE, ARTHUR C.  C/P,P,H,SFWA
SONGS OF EARTH & POWER [MICHAEL PERRIN (REVISED) 1,2]  BEAR, GREG  P
SONGS OF STARS AND SHADOWS  MARTIN, GEORGE R. R.  C/P,P
SONGS OF THE DANCING GODS [DANCING GODS 4]  CHALKER, JACK L.  C/P:2,P
SONGS OF VICTORY [@AGE OF CONAN KERN #3]  COLEMAN, LOREN L.  P
SONGSMITH [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE/CRISPIN, A. C.  H:2,P
SONJA BLUE COLLECTION, THE [SONJA BLUE 1,2,3]  COLLINS, NANCY A.  LP
SONS OF ENTROPY [@BUFFY GATEKEEPER 3]  GOLDEN, CHRISTOPHER/HOLDER, NANCY  P
SONS OF FENRIS [@WARHAMMER 40K]  LIGHTNER, LEE  P
SONS OF HEAVEN, THE [COMPANY 9]  BAKER, KAGE  H
SONS OF SOLOMON  REID, C. LESTOCK  H
SONS OF THE BEAR-GOD  PAGE, NORVELL W.  P
SONS OF THE OAK [RUNELORDS 5]  FARLAND, DAVID  H
SONS OF THE OCEAN DEEPS  WALTON, BRYCE  H:2
SONS OF THE TITANS [TITANS 3]  ADKINS, PATRICK H.  P
SONS OF THE WOLF  LUKENS, ADAM  H
SOOTERKIN, THE  GILLING, TOM  LP
SOOTHSAYER  RESNICK, MIKE  C/P:2,SFWA-P:2,P
SOPHIE’S LEGACY  MORGAN, DAREION  LP
SOPRANO SORCERESS, THE [SPELLSONG 1]  MODESITT, L. E. JR.  H,P
SORCERER, THE  BATO, JOSEPH  H

SORCERER, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #3] DENNING, TROY

P

SORCERER, THE  ERICSON, ERIC  H
SORCERER AND A GENTLEMAN, A [LANDUC 2]  WILLEY, ELIZABETH  C/P:2,SFWA,P
SORCERER AND A GENTLEMAN, A (PROOF) [LANDUC 2]  WILLEY, ELIZABETH  SR-LP
SORCERER: METAMORPHOSIS, THE [CAMULOD 6]  WHYTE, JACK  P
SORCERERS!  DANN, JACK/DOZOIS, GARDNER  C/PA:2,PA
SORCERER’S ACADEMY, THE  LITTLE, DENISE  PA
SORCERER’S AMULET [RUNESTAFF 2]  MOORCOCK, MICHAEL  P
SORCERER’S APPENDIX, THE  HARMAN, ANDREW  P
SORCERER’S APPRENTICE, THE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
SORCERER’S APPRENTICE, A  WILLIAMS, MICHAEL  P
SORCERER’S BLOOD [@WARRIOR OF VENGEANCE #1]  COE, ROSS ANTON  P
SORCERER’S CURSE, THE  VOLSKY, PAULA  C/P,P
SORCERER’S GUN, THE [@CHANGELING #2]  GARRISON, PETER  C/P,LP
SORCERER’S HEIR, THE  VOLSKY, PAULA  C/P,P
SORCERER’S LEGACY  WURTS, JANNY  C/P:2,P
SORCERERS OF MAJIPOOR [MAJIPOOR PRESTIMION 1]  SILVERBERG, ROBERT  H:2,P

TITLE  AUTHOR  CODE
SORCERERS OF THE FROZEN ISLES [SAVAGE EMPIRE 5]  LORRAH, JEAN  C/P,P
SORCERER’S SACRED ISLE, THE [DANANS 1]  TAYLOR, KEITH  C/P,P
SORCERER’S SHADOW, THE [@ORON #2]  SMITH, DAVID C.  P
SORCERER’S SHIP, THE  BOK, HANNES  C/P:2,P
SORCERER’S SKULL, THE  MASON, DAVID  P
SORCERER’S SKULL, THE [CENOTAPH ROAD 2]  VARDEMAN, ROBERT E.  C/P:2,P
SORCERER’S SON [CRAY ORMORU 1]  EISENSTEIN, PHYLLIS  C/P:2,P
SORCERER’S SON (PROOF) [CRAY ORMORU 1]  EISENSTEIN, PHYLLIS  SR-LP
SORCERER’S STONE  JAMES, L. DEAN  C/P,P
SORCERER’S TREASON, A [ISAVALTA 1]  ZETTEL, SARAH  H
SORCERER’S WORLD  BRODERICK, DAMIEN  C/P:2,P
SORCERESS [WEB,THE]  FUREY, MAGGIE  P
SORCERESS AND THE CYGNET, THE [CYGNET 1]  MCKILLIP, PATRICIA A.  C/P:2,SFWA,SFWA-P:2,P
SORCERESS OF AMBERMERE, THE [AMBERMERE 2]  PIERCE, J. CALVIN  P,SFWA-P
SORCERESS OF DARSHIVA [MALLOREON 4]  EDDINGS, DAVID  H:2,P
SORCERESS OF FAITH [LLADRANA 2]  OWENS, ROBIN D.  LP
SORCERESS OF QAR, THE  WHITE, TED  C/P:2,P
SORCERESS OF SCATH, THE [LAND BEYOND THE GATE 3]  ESHBACH, LLOYD ARTHUR  P
SORCERESS OF THE WITCH WORLD [WITCH WORLD ESTCARP 3]  NORTON, ANDRE  C/P:2,P
SORCEROR’S APPRENTICE, THE  JOHNSON, CHARLES  H
SORCEROR’S LADY, THE  VOLSKY, PAULA  P
SORCEROUS SEA [ISLAND WARRIOR 3]  SEVARANCE, CAROL  P
SORCERY AND CECELIA [CECILIA & KATHERINE 1]  WREDE, PATRICIA C./STEVERMER, CAROLINE C/P,P
SORCERY IN SHAD [PRIMAL LAND 3]  LUMLEY, BRIAN  H
SORCERY RISING [FOOL’S GOLD 1]  FISHER, JUDE  C/P,P
SORCERY SPELL BOOK, THE [SORCERY!]  JACKSON, STEVE  [C/RAN-P]
SORCERY WITHIN, THE [TANAGARAN 1]  SMEDS, DAVE  C/P,P
SORCIER DE LINN, LE  VAN VOGT, A. E.  FP
SORROWING VENGEANCE [FALL OF THE FIRST WORLD 2]  SMITH, DAVID C.  P
SORROWS OF SAP’ED, THE  ROCHE, JAMES JEFFREY  H
SORROWS OF SATAN, THE  CORELLI, MARIE  H
SORRY, CHIEF... [GET SMART 2]  JOHNSTON, WILLIAM  P
SOS BOBOMOBILE!  WAHL, JAN  LP
S.O.S. FROM THREE WORLDS [MED 2]  LEINSTER, MURRAY  C/P:2,P
!S.O.S. PLUTON!  THELS, H. S.  FP
SOS: SPACESHIP TITAN! [@PERRY RHODAN #34]  ACE  PA,C/PA
SOS THE ROPE [BATTLE CIRCLE 1]  ANTHONY, PIERS  C/P,@P,H
SOUL, THE  GORTON, RON  C/P,P
SOUL CATCHER  HERBERT, FRANK  C/P:2,P
SOUL DRINKER, THE [SOUL DRINKER 1,2,3]  CLAYTON, JO  H:2
SOUL DRINKER [@WARHAMMER 40K SOUL DRINKERS 1]  COUNTER, BEN  P
SOUL EATER, THE  RESNICK, MIKE  C/P,P
SOUL FIRE [@PATTERNS OF CHAOS #4]  INGRID, CHARLES  P
SOUL HARVEST [@LEFT BEHIND #4]  LAHAYE, TIM/JENKINS, JERRY B.  LP
SOUL MUSIC [DISCWORLD 16/DISCWORLD DEATH 3]  PRATCHETT, TERRY  C/P:2,P
SOUL OF A NEW MACHINE, THE  KIDDER, TRACY  [C/RAN-P]
SOUL OF ANNA KLANE, THE  MIEDANER, TERREL  C/P,P
SOUL OF THE BRIDE [@CHARMED]  LENHARD, ELIZABETH  P
SOUL OF THE CITY [@THIEVES’ WORLD #8]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
SOUL OF THE FIRE [SWORD OF TRUTH 5]  GOODKIND, TERRY  H:2,P
SOUL OF THE ROBOT  BAYLEY, BARRINGTON J.  C/P,SFWA,P,H
SOUL-SINGER OF TYRNOS  MAYHAR, ARDATH  C/P,P,H
SOUL STEALERS, THE  HUNTINGTON, CHARLES  C/P,P
SOUL TRADE [@ANGEL]  SNIEGOSKI, THOMAS E.  P
SOUL WEAVER, THE [BRIDGE OF D’ARNATH 3]  BERG, CAROL  P
SOULFORGE, THE  PHILLIPS, TERRY  P
SOULFORGE, THE [@DRAGONLANCE RAISTLIN 1]  WEIS, MARGARET  P
SOULMATE  RUNYON, CHARLES W.  P
SOULS  RUSS, JOANNA  SFWA-TD11,TD11
SOULS IN METAL  ASHLEY, MIKE  HA,PA
SOULS IN PAWN  HATCH, GEORGE  LPA
SOULS IN THE GREAT MACHINE [GREAT MACHINE 1]  MCMULLEN, SEAN  LP
SOULSMITH [SOULSMITH 1]  DEITZ, TOM  P
SOULSTRING  SNYDER, MIDORI  C/P:2,P
SOULSUCKER, THE  SABINE, TED  C/P:2,P
SOUND AND THE FURRY, THE [HOKA 2,3]  ANDERSON, POUL/DICKSON, GORDON R.  H
SOUND OF HIS HORN, THE  SARBAN  C/P:2,P
SOUND OF MIDNIGHT, THE  GRANT, CHARLES L.  H,SFWA-P
SOUND OF WINGS, THE  DUNMORE, SPENCER  C/P,SFWA-P,P
SOUND OF WINTER, THE  COVER, ARTHUR BYRON  P
SOUNDING  SEARLS, HANK  LP
SOUNDING STILLNESS, THE  VON GUNDEN, KENNETH  C/P,P
SOURCE, THE [NECROSCOPE 3]  LUMLEY, BRIAN  C/P,SFWA-P:2,P
SOURCE OF EVIL  VIGLIANTE, MARY  P

The MITSFS Titledex  September 24, 2007  Page 334
TITLE  AUTHOR  CODE
SOURCE OF MAGIC, THE [@XANTH #2]  ANTHONY, PIERS  C/P:2,P
SOURCERY [DISCWORLD 5/DISCWORLD RINCEWIND 3]  PRATCHETT, TERRY  C/P,P,H
SOUTH POLE TERROR, THE [DOC SAVAGE #77]  ROBESON, KENNETH  C/P,@P
SOUTHERN CROSS [@ROBOTECH #7]  MCKINNEY, JACK  C/P,P
SOUTHERN FIRE [ALDABRESHIN COMPASS 1]  MCKENNA, JULIET E.  H
SOUTHERN FRIED SCI-FI AND JAMBALAYA GENRES  COCHRAN, LIN NIELSEN/HANCOCK, SCOTT  SR-L,LPA
SOUTHERN STAR MYSTERY, THE  VERNE, JULES  H
SOUTHSHORE [@AWAKENERS #2]  TEPPER, SHERI S.  C/P:2,SFWA-P,P,H
SOUVENIR OF MONIQUE  BRADLEY, MARION ZIMMER  C/P,P
SOVEREIGN [@KEYS OF POWER #3]  BROWN, SIMON  P
SOVEREIGN  MELUCH, R. M.  C/P,P
SOVIET SCIENCE FICTION  COLLIER BOOKS  PA,C/PA
SOVREMENNAYA ZARUBEZHNAYA FANTASTIKA  MITROKHINA, S./VOSKRESENSKAYA, E.  FH
SOWERS OF THE THUNDER, THE  HOWARD, ROBERT E.  SR-H,P
SOWN IN THE DARKNESS A.D. 2,000  TWIFORD, WILLIAM RICHARD  H
SPACE  MICHENER, JAMES A.  C/P,P,H
SPACE: 1999 YEAR TWO [SPACE: 1999 YEAR 2]  BUTTERWORTH, MICHAEL  SR-VLH
SPACE (1999)  POMEGRANATE  CAL
SPACE: ABOVE AND BEYOND [SPACE: ABOVE AND BEYOND 1]  TELEP, PETER  P
SPACE ADVENTURES  FURMAN, A. L.  PA
SPACE AGE, THE [DOCTOR WHO EIGHTH DOCTOR]  LYONS, STEVE  P
SPACE AND BEYOND  MONTGOMERY, R. A.  C/P:2,P
SPACE AND SCIENCE FICTION PLAYS FOR YOUNG PEOPLE  KAMERMAN, SYLVIA E.  HA
SPACE ANGEL [SPACE ANGEL 1]  ROBERTS, JOHN MADDOX  C/P:2,P
SPACE APPRENTICE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H:2
SPACE ARK  HUBSCHMAN, THOMAS  P
SPACE ART  MILLER, RON  A
SPACE BALTIC  HOLLO, ANSELM  OVD-2
SPACE BARBARIANS, THE  GODWIN, TOM  C/P:2,P
SPACE BARBARIANS, THE  REYNOLDS, MACK  D77710
SPACE BEAM  ROBB, JOHN  SR-P
SPACE BEYOND, THE  CAMPBELL, JOHN W. JR.  C/P:2,P
SPACE-BORN, THE  TUBB, E. C.  DD193
SPACE-BORNE  BADGER BOOKS  SR-PA
SPACE BY THE TALE  BIXBY, JEROME  C/P:2,P
SPACE CADET  HEINLEIN, ROBERT A.  C/P,P,H
SPACE CAMP [STAR TREK DS9]  PEDERSEN, TED  LP
SPACE CAPTAIN  LEINSTER, MURRAY  DM135
SPACE CHANTEY  LAFFERTY, R. A.  DH056
SPACE CHILD’S MOTHER GOOSE, THE  WINSOR, FREDERICK/PARRY, MARIAN  H,LP
SPACE CIRCUS, THE [FLASH GORDON 3]  STEFFANSON, CON  C/P,P
SPACE DICTIONARY  ASIMOV, ISAAC  [C/RAN-P]
SPACE DOCTOR  CORREY, LEE  C/P:2,P
SPACE DOGFIGHTS  DICKSON, GORDON R.  C/PA,PA

SPACE DREADNOUGHTS DRAKE, DAVID/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA

SPACE DREAMERS, THE  CLARKE, ARTHUR C.  P
SPACE EAGLE: OPERATION DOOMSDAY, THE  PEARL, JACK  H
SPACE EAGLE: OPERATION STAR VOYAGE, THE  PEARL, JACK  H
SPACE EATER, THE  LANGFORD, DAVID  C/P,P
SPACE EGG, THE  WINTERBOTHAM, RUSS  C/P:2,P
SPACE ENTERPRISE, THE  STINE, G. HARRY  [C/RAN-P]
SPACE FLIGHT  DEL REY, LESTER  [C/RAN-H]
SPACE FLIGHT  HUNT, GILL  SR-P
SPACE FOLK  ANDERSON, POUL  C/P:2,SFWA-P,P
SPACE FOR HIRE [SAM SPACE]  NOLAN, WILLIAM F.  C/P:2,P
SPACE FRONTIERS, THE  VERNON, ROGER LEE  C/P:2,P
SPACE FURY  FANTHORPE, R. L.  SR-P,C/P:2
SPACE GHOST/SORCERESS OF CYBA-3, THE  CHRISTENSEN, DON R.  BLB

SPACE GLADIATORS DRAKE, DAVID/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

C/PA,PA

SPACE GUARDIAN, THE  DANIELS, MAX  P
SPACE GUARDIANS, THE [SPACE: 1999 3]  BALL, BRIAN N.  C/P:2,P
SPACE GYPSIES  LEINSTER, MURRAY  C/P:2,P
SPACE HAWK  GILMORE, ANTHONY  H
SPACE HAWKS [@STARHAWKS #1]  DALTON, SEAN  C/P:2,P
SPACE HOSTAGES  FISK, NICHOLAS  SR-P
SPACE INC.  CZERNEDA, JULIE E.  C/PA,PA

SPACE INFANTRY DRAKE, DAVID/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA

SPACE-JACKERS, THE [@SPACE: 1999 YEAR 2 #3]  BUTTERWORTH, MICHAEL  P
SPACE JOKES, COSMIC CARTOONS, AND MARTIAN LAUGHS  BREDESON, LENORE/NELSON, DICK  [C/RAN-P]
SPACE LASH  CLEMENT, HAL  C/P:2,P
SPACE LAWYER  SCHACHNER, NAT  H

The MITSFS Titledex  September 24, 2007  Page 335
TITLE  AUTHOR  CODE
SPACE LINE  LANG, KING  SR-P
SPACE LORDS  SMITH, CORDWAINER  C/P:2,P
SPACE MACHINE, THE  DANIEL, JERRY C.  H
SPACE MACHINE, THE  PRIEST, CHRISTOPHER  H,SFWA,P
SPACE MACHINE, THE/DREAM OF WESSEX, A  PRIEST, CHRISTOPHER  P
SPACE MAGICIANS, THE  NORTON, ALDEN H./MOSKOWITZ, SAM  C/PA,PA

SPACE MAIL ASIMOV, ISAAC/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

PA,C/PA

SPACE MAIL II ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

SPACE MAVERICKS, THE [SPACE MAVERICKS 1]  KRING, MICHAEL  P
SPACE MEN  SHAW, DAVID  SR-P
SPACE MERCENARIES  CHANDLER, A. BERTRAM  DM133
SPACE MERCHANTS, THE [SPACE MERCHANTS 1]  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
SPACE NO BARRIER  TORRO, PEL  SR-P
SPACE ODYSSEY  GALLERY BOOKS  HA
SPACE ODYSSEYS  ALDISS, BRIAN W.  C/PA,PA,SFWA
SPACE ODYSSEYS OF ARTHUR C. CLARKE, THE  SLUSSER, GEORGE EDGAR  R/REF-P
SPACE OLYMPICS, THE  LIGHTNER, A. M.  P
SPACE ON MY HANDS  BROWN, FREDRIC  SR-P,C/P:2
SPACE OPERA  ALDISS, BRIAN W.  C/PA,PA

SPACE OPERA MCCAFFREY, ANNE/

SCARBOROUGH, ELIZABETH ANN

C/PA,PA

SPACE OPERA  VANCE, JACK  C/P:2,P
SPACE OPERA (VIE) [@VANCE INTEGRAL EDITION #18]  VANCE, JACK  SR-H
SPACE OPERA RENAISSANCE, THE  HARTWELL, DAVID G./CRAMER, KATHRYN  HA
SPACE PIONEERS  NORTON, ANDRE  HA
SPACE PIONEERS, THE [TOM CORBETT 4]  ROCKWELL, CAREY  H
SPACE PIRATES  DEL MARTIA, ASTRON  SR-P
SPACE PIRATES, THE [DOCTOR WHO]  DICKS, TERRANCE  P
SPACE PIRATES [@LORD TEDRIC #2]  EKLUND, GORDON  P
SPACE PLAGUE  SMITH, GEORGE O.  C/P:2,P
SPACE PLATFORM  LEINSTER, MURRAY  C/P,P,H
SPACE POLICE (TEACHER’S GUIDE)  FEARON-PITMAN  R/REF-P
SPACE POWER  STINE, G. HARRY  [C/RAN-P]
SPACE PRISON  GODWIN, TOM  C/P,P
SPACE PRODIGAL, THE  SHERRELL, CARL  C/P:2,P
SPACE RANGER [LUCKY STARR 1]  ASIMOV, ISAAC  P
SPACE RELATIONS  BARR, DONALD  C/P,P,H
SPACE SCAVENGERS, THE  CARTMILL, CLEVE  P
SPACE SEARCH  REYNOLDS, MACK  C/P:2,P
SPACE SEED [STAR TREK]  COON, GENE L./RODDENBERRY, GENE  SR-S
SPACE SERVICE  NORTON, ANDRE  HA
SPACE SHIP UNDER THE APPLE TREE, THE  SLOBODKIN, LOUIS  C/P,P
SPACE SHUTTLE AT WORK, THE  ALLAWAY, HOWARD  [C/RAN-H]

SPACE SHUTTLES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

SPACE SKIMMER  GERROLD, DAVID  C/P:2,P
SPACE SLAVES  RUSK, JAMES JR.  C/P,P
SPACE SOLDIERS  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
SPACE, SPACE, SPACE  SLOANE, WILLIAM  HA
SPACE STATION #1  LONG, FRANK BELKNAP  DD242,P,C/P
SPACE STATION EIGHT  ALLEN, CARLTON C.  H
SPACE STATION ICE-3  COVILLE, BRUCE  C/P,P
SPACE STATIONS  GREENBERG, MARTIN H./HELFERS, JOHN  PA
SPACE SUITS  SHEFFIELD, CHARLES  LP
SPACE SWAP 1984–8612  PIPER, WALTER  SR-L
SPACE SWIMMERS, THE  DICKSON, GORDON R.  C/P:2,P
SPACE, THE FIDO FRONTIER [@STAR WRECK #7]  REWOLINSKI, LEAH  P
SPACE, TIME, AND CRIME  DEFORD, MIRIAM ALLEN  PA,C/PA
SPACE, TIME AND NATHANIEL  ALDISS, BRIAN W.  P
SPACE-TIME JOURNAL, THE  MERRIL, JUDITH  PA
SPACE-TIME JUGGLER, THE  BRUNNER, JOHN  DF227
SPACE-TIME TASK FORCE  YORKE, PRESTON  SR-P
SPACE TRAP  BELL, THORNTON  SR-P,H
SPACE TRAP  COULSON, JUANITA  SR-P
SPACE TRAVEL  BOVA, BEN/LEWIS, ANTHONY R.  R/REF-H
SPACE TREASON  CLARKE, A. V./BULMER, H. K.  SR-P

SPACE TREK GLENN, JEROME CLAYTON/ ROBINSON, GEORGE S.

[C/RAN-P]

SPACE TUG  LEINSTER, MURRAY  C/P:2,P
SPACE VAMPIRES, THE  WILSON, COLIN  C/P,P,H,SFWA
SPACE VECTORS [@BIOWARRIORS #3]  VARDEMAN, ROBERT E.  C/P,P
SPACE VIKING  PIPER, H. BEAM  C/P:2,P
SPACE VISITOR  REYNOLDS, MACK  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 336
TITLE  AUTHOR  CODE
SPACE VOID  JOHNS, MARSTON  H
SPACE WAR [PROFESSOR JAMESON 3]  JONES, NEIL R.  C/P:2,P
SPACE WAR BLUES  LUPOFF, RICHARD A.  P
SPACE WARP, THE  STATTEN, VARGO  SR-P

SPACE WARS WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,SFWA-P

SPACE WARS: WORLDS AND WEAPONS  EISLER, STEVEN  A
SPACE WILLIES, THE  RUSSELL, ERIC FRANK  DD315,D77785
SPACE WINNERS  DICKSON, GORDON R.  C/P,SFWA-P,P,H
SPACE WOLF [WARHAMMER 40K SPACE WOLF 1]  KING, WILLIAM  P
SPACE WOLF OMNIBUS, THE [@WARHAMMER 40K SPACE WOLF 1,2,3]  KING, WILLIAM  H
SPACEACHE  WILSON, SNOO  C/P,P
SPACEBREAD  SENN, STEVE  H
SPACECAMP  CLARO, JOE  P
SPACECRAFT  HAGGERTY, JAMES J. JR.  [C/RAN-P:2]
SPACECRAFT 2000 TO 2100 AD  COWLEY, STEWART  A
SPACECRAFT IN FACT AND FICTION  HARRISON, HARRY/EDWARDS, MALCOLM  A

SPACEFIGHTERS HALDEMAN, JOE/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA

SPACEFLIGHT — VENUS  WILDING, PHILIP  H
SPACEHAWK  WOLVERTON, BASIL  CX
SPACEHAWK, INC.  GOULART, RON  C/P:2,P
SPACEHIVE  SUTTON, JEFF  C/P:2,P
SPACEHOUNDS OF IPC  SMITH, E. E.  C/P,P,H
SPACEJACKS, THE  WELLS, ROBERT  P
SPACELAND  RUCKER, RUDY  H
SPACELING  PISERCHIA, DORIS  H:2,P
SPACEMEN, GO HOME  LESSER, MILTON  H:2
SPACEPAW  DICKSON, GORDON R.  C/P,SFWA-P,P,H
SPACEPOWER  COX, DONALD/STOIKO, MICHAEL  [C/RAN-H]
SPACER DREAMS  SEGRIFF, LARRY  P
SPACER: WINDOW OF THE MIND [SPACE ANGEL 2]  ROBERTS, JOHN MADDOX  P
SPACESHIP BUILT OF STONE, A  TUTTLE, LISA  P
SPACESHIP FOR THE KING, A  POURNELLE, JERRY  C/P,P
SPACESHIP MEDIC  HARRISON, HARRY  SR-P
SPACESHIP OF ANCESTORS [@PERRY RHODAN #73]  ACE  PA,C/PA
SPACESHIP TITANIC  DUPREY, RICHARD/O’LEARY, BRIAN  H
SPACESHIP TO SATURN  WALTERS, HUGH  H
SPACESHIPS OF EZEKIEL, THE  BLUMRICH, JOSEF  [C/RAN-P:2]
SPACESHIPS OF THE MIND  CALDER, NIGEL  R/REF-H
SPACETIME DONUTS  RUCKER, RUDY  P
SPACEWALKS  ROBERTS, BURGERT  H
SPACEWARD BOUND  BROWN, SLATER  H
SPACEWAYS  MAINE, CHARLES ERIC  SR-P
SPACEWRECK  COWLEY, STEWART  A
SPACIAL DELIVERY  DICKSON, GORDON R.  DF119,C/P,SFWA-P
SPACING DUTCHMAN  VINICOFF, ERIC/MARTIN, MARCIA  SR-LP
SPADEFUL OF SPACETIME, A  SABERHAGEN, FRED  C/PA,PA
SPANISH CAVE, THE  HOUSEHOLD, GEOFFREY  P
SPARES  SMITH, MICHAEL MARSHALL  P
SPARK AND BURN [@BUFFY]  GALLAGHER, DIANA G.  P
SPARROW, THE  RUSSELL, MARY DORIA  H:2,LP
SPARROWHAWK [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P,P
SPARTACUS [STAR TREK TNG]  MANCOUR, T. L.  C/P:3,P

SPARTACUS FILE, THE WATT-EVANS, LAWRENCE/ PARLAGRECO, CARL

LP

SPARTAN PLANET [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
SPATIAL RAY  HUNT, GILL  SR-P
SPAWN  GLUT, DONALD F.  SR-P
SPAWN OF CTHULHU, THE  CARTER, LIN  PA,C/PA
SPAWN OF LABAN [@CAP KENNEDY #11]  KERN, GREGORY  C/P,P
SPAWN OF LOKI, THE  HENDERSON, JASON  P
SPAWN OF SATAN  BIRKIN, CHARLES  P
SPAWN OF SPACE  HARKON, FRANZ  SR-P
SPAWN OF THE DEATH MACHINE, THE  WHITE, TED  C/P:2,P
SPAWN OF THE VORTEX  GAYLE, HENRY K.  H
SPAWN OF THE WINDS  LUMLEY, BRIAN  C/P,P
SPAWNING, THE  RAVENSWOOD, FRITZEN  P
SPEAK OF THE DEVIL  NORTH, STERLING/BOUTELL, C. B.  HA
SPEAKER FOR THE DEAD [ENDER 2]  CARD, ORSON SCOTT  C/P:2,SFWA,SFWA-P,P
SPEAKER TO HEAVEN  NOEL, ATANIELLE ANNYN  H,SFWA-P
SPEAKING DREAMS  PARK, SEVERNA  P
SPEAKING IN TONGUES  MCDONALD, IAN  P
SPEAKING OF SCIENCE FICTION: THE PAUL WALKER INTERVIEWS  WALKER, PAUL  R/REF-H,C/REF-H
SPEAKING SCIENCE FICTION  SAWYER, ANDY/SEED, DAVID  R/REF-P

The MITSFS Titledex  September 24, 2007  Page 337
TITLE  AUTHOR  CODE
SPEAKING STONES  LEIGH, STEPHEN  P
SPEAR, THE  HERBERT, JAMES  P
SPEAR OF FIRE  RUSOFF, GARRY  P
SPEAR OF HEAVEN [@AVARYAN RISING #5]  TARR, JUDITH  P
SPEAR OF MARS, THE [@FUTURE AT WAR #2]  BRETNOR, REGINALD  PA,C/PA,SFWA-P
SPEARMEN OF ARN  DOW DELL, DEL  P
SPEARS OF GOD  HENDRIX, HOWARD V.  LP

SPECIAL COLLECTIONS SCIENCE/FICTION COLLECTIONS:

FANTASY, SUPERNATURAL AND WEIRD TALES HAWORTH PRESS

R/REF-P

SPECIAL DELIVERANCE  SIMAK, CLIFFORD D.  C/P,P,H:2
SPECIAL DELIVERY  NEVILLE, KRIS  SR-DB0788
SPECIAL EDITIONS SAMPLER 1992  BANTAM SPECTRA  B
SPECIAL EFFECTS [@FAMILY #1]  MCCARTHY, KEVIN/SILVA, DAVID  P
SPECIAL FEATURE  DE VET, CHARLES V.  C/P,P
SPECIAL MISSION  MULLER, JOHN E.  SR-P
SPECIAL WONDER 1  MCCOMAS, J. FRANCIS  PA
SPECIAL WONDER 2  MCCOMAS, J. FRANCIS  PA,C/PA

SPECIAL WONDER: THE ANTHONY BOUCHER MEMORIAL ANTHOLOGY OF FANTASY AND SCIENCE FICTION MCCOMAS, J. FRANCIS

HA

SPECIALIST, THE  WHITEFORD, WYNNE  C/P,P
SPECIES  NAVARRO, YVONNE  P
SPECIMEN DAYS  CUNNINGHAM, MICHAEL  LP
SPECIMENS  SABERHAGEN, FRED  C/P:2,P
SPECTER BOOKS (TEACHER’S GUIDE)  FEARON-PITMAN  R/REF-P
SPECTER IS HAUNTING TEXAS, A  LEIBER, FRITZ  H:2,P,SFWA
SPECTER OF THE PAST [STAR WARS]  ZAHN, TIMOTHY  P
SPECTERS OF THE DAWN [MOREAU 3]  SWANN, S. ANDREW  C/P:2,P
SPECTERWORLD  HAIBLUM, ISIDORE  C/P,P
SPECTRA SAMPLER 1991  BANTAM SPECTRA  B
SPECTRA TO THE STARS!  BANTAM SPECTRA  B
SPECTRA TO THE STARS: SPECTRA SAMPLER 1993  BANTAM SPECTRA  B
SPECTRE  LAWS, STEPHEN  SFWA-P

SPECTRE [STAR TREK MIRROR 1] SHATNER, WILLIAM/

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD

C/P,P

SPECTRE OF DARKNESS  MULLER, JOHN E.  SR-P

SPECTRE OF THE BLACK ROSE [RAVENLOFT] LOWDER, JAMES/

WHITNEY-ROBINSON, VORONICA

C/P,P

SPECTRUM  AMIS, KINGSLEY/CONQUEST, ROBERT  C/PA,PA
SPECTRUM 2  AMIS, KINGSLEY/CONQUEST, ROBERT  C/PA,PA
SPECTRUM 3  AMIS, KINGSLEY/CONQUEST, ROBERT  PA,C/PA
SPECTRUM 4  AMIS, KINGSLEY/CONQUEST, ROBERT  C/PA,PA
SPECTRUM 5  AMIS, KINGSLEY/CONQUEST, ROBERT  C/PA,PA
SPECTRUM (SELECTIONS)  AMIS, KINGSLEY/CONQUEST, ROBERT  C/PA,PA
SPECTRUM OF A FORGOTTEN SUN [DUMAREST 15]  TUBB, E. C.  C/P:2,P
SPECTRUM OF WORLDS, A  CLARESON, THOMAS D.  HA
SPECULATIONS  ASIMOV, ISAAC/LAURANCE, ALICE  SR-HA

SPECULATIONS CLYMER, THEODORE/RUTH, LEO/ EVANECHKO, PETER/HIGGS, JULIA

LP

SPEED OF DARK, THE  MOON, ELIZABETH  H
SPEED OF DARKNESS [@STARCRAFT #3]  HICKMAN, TRACY  P
SPEED OF FLIGHT [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P,P
SPELL AT SCOGGIN’S CROSSING, A  RILEY, LOUISE  H
SPELL BOUND  EMERSON, RU  P,SFWA-P:2
SPELL-BOUND SCHOLAR, THE [WARLOCK’S HEIRS]  STASHEFF, CHRISTOPHER  P
SPELL FANTASTIC  GREENBERG, MARTIN H./SEGRIFF, LARRY  C/PA,PA
SPELL FOR CHAMELEON, A [@XANTH #1]  ANTHONY, PIERS  C/P:2,P
SPELL FOR CHAMELEON, A (ABRIDGED; CASSETTE)  ANTHONY, PIERS  SR-MM
SPELL FOR OLD BONES, A  LINKLATER, ERIC  H
SPELL IS CAST, A  CAMERON, ELEANOR  P
SPELL OF APOCALYPSE [DANCE OF GODS 4]  BRENNER, MAYER ALAN  P
SPELL OF DECEIT, A  GOODMAN, LAURIE  P
SPELL OF FATE [DANCE OF GODS 3]  BRENNER, MAYER ALAN  P
SPELL OF INTRIGUE [DANCE OF GODS 2]  BRENNER, MAYER ALAN  P
SPELL OF SEVEN, THE  DE CAMP, L. SPRAGUE  PA,C/PA
SPELL OF THE BLACK DAGGER, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
SPELL OF THE HIGHLANDER  MONING, KAREN MARIE  H
SPELL OF THE WINTER WIZARD  LOWERY, LINDA  P
SPELL OF THE WITCH WORLD [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
SPELL SINGERS  NEWCOMER, ALAN BARD  PA,C/PA
SPELL SWORD, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
SPELLBINDER  RAWN, MELANIE  H
SPELLBOUND  BEEDING, FRANCIS  H
SPELLBOUND  VELDE, VIVIAN VANDE  H
SPELLCOATS, THE [DALEMARK 3]  JONES, DIANA WYNNE  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 338
TITLE  AUTHOR  CODE
SPELLFIRE [FORGOTTEN REALMS SHANDRIL 1]  GREENWOOD, ED  C/P:2,P
SPELLKEY TRILOGY, THE [SPELLKEY (REVISED) 1,2,3]  DOWNER, ANN  P

SPELLS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

SPELLS OF MORTAL WEAVING [TWELVE KINGDOMS 2]  FRIESNER, ESTHER M.  C/P,P
SPELLS OF WONDER  BRADLEY, MARION ZIMMER  C/PA,PA
SPELLSINGER [SPELLSINGER 1]  FOSTER, ALAN DEAN  C/P:2,P
SPELLSINGER’S SCHERZO [SPELLSINGER 4,5,6]  FOSTER, ALAN DEAN  H
SPELLSONG WAR, THE [SPELLSONG 2]  MODESITT, L. E. JR.  H,P
SPELLSTONE OF SHALTUS, THE  BUSHYAGER, LINDA E.  C/P:2,P
SPHERE  CRICHTON, MICHAEL  H:2,P
SPHERELAND  BURGER, DIONYS  H,LP
SPHERES OF HEAVEN, THE  SHEFFIELD, CHARLES  H:2,P
SPHERICAL HARMONIC [SKOLIAN EMPIRE 7]  ASARO, CATHERINE  H
SPHERO NOVA  CAMERON, BERL  SR-P
SPHYNXES WILD  FRIESNER, ESTHER M.  P

SPIDER DESERT [@ATLAN #1]/ WASP MEN ATTACK, THE (SPECIAL) [PERRY RHODAN] ACE

PA

SPIDER ISLAND [@COLLECTED STORIES JACK WILLIAMSON #4]  WILLIAMSON, JACK  H
SPIDER KISS  ELLISON, HARLAN  C/P:2,P
SPIDER LEGS  ANTHONY, PIERS/PICKOVER, CLIFFORD A.  H,P
SPIDER-MAN [SPIDER-MAN]  DAVID, PETER  P
SPIDER-MAN 2  DAVID, PETER  H
SPIDER-MAN: CARNAGE IN NEW YORK [SPIDER-MAN]  MICHELINIE, DAVID/SMITH, DEAN WESLEY  P
SPIDER-MAN: THE VENOM FACTOR [SPIDER-MAN]  DUANE, DIANE  P
SPIDER PLAY [JANNA BRILL 2]  KILLOUGH, LEE  P
SPIDERS, THE  LEWIS, RICHARD  P
SPIDER’S TEST, THE [@BIRTHRIGHT 5]  MCKEONE, DIXIE LEE  P
SPIDERS’ WAR  WRIGHT, S. FOWLER  H
SPIKE VS. DRACULA [BUFFY SPIKE]  DAVID, PETER  CX
SPIN  WILSON, ROBERT CHARLES  H
SPIN CONTROL  MORIARTY, CHRIS  LP
SPIN STATE  MORIARTY, CHRIS  H,LP
SPIN STATE (PROOF)  MORIARTY, CHRIS  SR-LP
SPINDLE’S END  MCKINLEY, ROBIN  H:2
SPINDOC  PERRY, STEVE  C/P:2,SFWA-P,P
SPINDRIFT  STEELE, ALLEN  H
SPINE OF THE WORLD, THE [@FORGOTTEN REALMS]  SALVATORE, R. A.  P
SPINNER, THE  PISERCHIA, DORIS  H:2,P
SPINNER OF SILVER AND THISTLE  BOK, HANNES  SR-H
SPINNERET  ZAHN, TIMOTHY  C/P,P,H
SPIRAL DANCE, THE  GARCIA Y ROBERTSON, R.  H,P
SPIRAL OF FIRE [MAGES OF GARILLON 3]  HARRIS, DEBORAH TURNER  C/P,P
SPIRAL SCRATCH [DOCTOR WHO PAST DOCTORS]  RUSSELL, GARY  P
SPIRES OF SPIRIT [STRANDS OF STARLIGHT]  BAUDINO, GAEL  P
SPIRIT CROSSINGS  PECK, CLAUDIA  P

SPIRIT FOX REICHERT, MICKEY ZUCKER/

WINGERT, JENNIFER

C/P,P

SPIRIT GATE, THE  BOHNHOFF, MAYA KAATHRYN  P
SPIRIT GATE [CROSSROADS 1]  ELLIOTT, KATE  H
SPIRIT OF BAMBATSE, THE  HAGGARD, H. RIDER  LP
SPIRIT OF DORSAI, THE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,LP,H
SPIRIT OF THE HAWK [LOST LANDS 3]  ESTES, ROSE  P
SPIRIT OF THE STONE, THE [SHADOWLEAGUE 2]  FUREY, MAGGIE  P
SPIRIT OF THE WOLF [@CHARMED]  GALLAGHER, DIANA G.  P
SPIRIT OF THUNDER, THE [FALLEN CLOUD 2]  GIAMBASTIANI, KURT R. A.  H,P
SPIRIT RING, THE  BUJOLD, LOIS MCMASTER  H:2,P
SPIRIT STONE, THE [@SILVER WYRM #2]  KERR, KATHARINE  H
SPIRIT WAS WILLING, THE  LUBAN, MILTON  H
SPIRITRIDE  SHEPHERD, MARK  P
SPIRITS IN THE WIRES  DE LINT, CHARLES  H
SPIRITS OF CAVERN AND HEARTH  EASTON, M. COLEMAN  H
SPIRITS OF CHRISTMAS  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
SPIRITS OF FLUX AND ANCHOR [@SOUL RIDER #1]  CHALKER, JACK L.  C/P:2,P
SPIRITS THAT WALK IN SHADOW  HOFFMAN, NINA KIRIKI  H
SPIRITS WHITE AS LIGHTNING [ERIC BANYON 4]  LACKEY, MERCEDES/EDGHILL, ROSEMARY  H
SPIRITUAL TRAILS TO HAPPINESS  FOSTER, RABBI SOLOMON  [C/RAN-H]
SPIRITWALK [TAMSON HOUSE 2]  DE LINT, CHARLES  C/P,P,H
SPLENDID CHAOS, A  SHIRLEY, JOHN  H
SPLENDID FREEDOM, THE  DARNAY, ARSEN  C/P,P
SPLENDID YOU  BAILEY, LYNNE  P
SPLENDOUR FALLS, THE  KELLY, ERIN E.  PA
SPLINTER OF THE MIND’S EYE [STAR WARS]  FOSTER, ALAN DEAN  H:2,P
SPLINTERED LIGHT  FLIEGER, VERLYN  R/REF-P

The MITSFS Titledex  September 24, 2007  Page 339
TITLE  AUTHOR  CODE
SPLINTERED SUNGLASSES AFFAIR, THE [@MAN FROM UNCLE #16]  LESLIE, PETER  C/P,P
SPLINTERS  HAMILTON, ALEX  SR-PA

SPLIT HEIRS WATT-EVANS, LAWRENCE/ FRIESNER, ESTHER M.

C/P,P,H

SPLIT INFINITY [APPRENTICE ADEPT #1]  ANTHONY, PIERS  C/P:2,@P
SPLIT SECOND  KILWORTH, GARRY  P
SPLIT WORLDS  BARR, TYRONE C.  SR-P
SPLITTING INFINITY  WSFS  VLP

SPOCK, MESSIAH! [STAR TREK] COGSWELL, THEODORE R./ SPANO, CHARLES A.

C/P:3,P

SPOCK MUST DIE! [STAR TREK]  BLISH, JAMES  C/P:2,P
SPOCK’S WORLD [STAR TREK]  DUANE, DIANE  H:2,P
SPOILS OF WAR, THE [@DAMNED #3]  FOSTER, ALAN DEAN  C/P:2,P
SPOOK HOLE [DOC SAVAGE #70]  ROBESON, KENNETH  C/P:2,@P
SPOOK LEGION, THE [DOC SAVAGE #16]  ROBESON, KENNETH  C/P:2,@P
SPOOKS, SPOOKS, SPOOKS  HOKE, HELEN  HA
SPOOKY SOUTHWEST  SCHLOSSER, S. E.  LP
SPOOR OF THE ANTIS [@PERRY RHODAN #102]  ACE  C/PA,PA
SPORE 7  CARLILE, CLANCY  P
SPORTING BLOOD [@UFO #2]  MIALL, ROBERT  P
SPORTING CHANCE [HERIS SERRANO 2]  MOON, ELIZABETH  C/P:2,P
SPOT OF LIFE, THE  HALL, AUSTIN  C/P,P
SPOTTED MEN, THE [DOC SAVAGE #87]  ROBESON, KENNETH  @P
SPREAD, THE  MALZBERG, BARRY N.  SR-P,C/P
SPRIGGAN MIRROR, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  LP
SPRINGWAR [ERON 2]  DEITZ, TOM  LP
SPROUT MASK REPLICA [TRILOGY THAT DARE NOT SPEAK ITS NAME 1]  RANKIN, ROBERT  P
SPROUTS OF WRATH, THE [BRENTFORD 4]  RANKIN, ROBERT  P
SPURIOUS SUN  BORODIN, GEORGE  H
SPURLOCK: SHERIFF OF PURGATORY  MORRIS, JIM  P
SPY AND THE ATOM GUN, THE  SETH, RONALD  H
SPY IN ISENGARD, A  AMTHOR, TERRY K.  P
SPY IN SPACE  MOORE, PATRICK  P
SPY WHO LOVED ME, THE [JAMES BOND 10]  FLEMING, IAN  C/P:2,P
SPY WHO SPOKE PORPOISE, THE  WYLIE, PHILIP  C/P,P
SPYBOT! [@PERRY RHODAN #53]  ACE  PA,C/PA
SPYING EYES [@SABRINA THE TEENAGE WITCH #14]  HOLDER, NANCY  P
SQUADRON ALERT [STARCRUISER SHENANDOAH 1]  GREEN, ROLAND J.  C/P:2,P
SQUARE DEAL, THE [@CAR WARRIORS #1]  DRAKE, DAVID  LP
SQUARE ROOT OF MAN, THE  TENN, WILLIAM  C/P:2,P
SQUARES OF THE CITY, THE  BRUNNER, JOHN  C/P:2,SFWA-P,P
SQUARING THE CIRCLE [@CIRCLE OF LIGHT #4]  HANCOCK, NIEL  C/P:2,P
SQUAT  ROME, DAVID  SR-P
SQUEAKING GOBLIN, THE [DOC SAVAGE #35]  ROBESON, KENNETH  C/P:2,@P
SQUIRE [@PROTECTOR OF THE SMALL 3]  PIERCE, TAMORA  H:2
SQUIRE [SQUIRE 1]  TELEP, PETER  C/P,P
SQUIRE’S BLOOD [SQUIRE 2]  TELEP, PETER  P
SQUIRE’S HONOR [SQUIRE 3]  TELEP, PETER  P
SRDCE TECHNOPOPU  OLSA, JAROSLAV JR.  FP
SS-GB  DEIGHTON, LEN  H,SFWA,P
ST ANTHONY’S FIRE [DOCTOR WHO NEW ADVENTURES]  GATISS, MARK  C/P:2,P
STABLE STRATEGIES AND OTHERS  GUNN, EILEEN  LP
STADIUM BEYOND THE STARS  LESSER, MILTON  H
STAGE FRIGHT  MCNALLY, CLARE  SFWA-P
STAINLESS  GRIMSON, TODD  C/P,SFWA,LP
STAINLESS STEEL RAT, THE [STAINLESS STEEL RAT 1]  HARRISON, HARRY  C/P,P
STAINLESS STEEL RAT FOR PRESIDENT, THE [STAINLESS STEEL RAT 5]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT GETS DRAFTED, THE [STAINLESS STEEL RAT 6]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT GOES TO HELL, THE [STAINLESS STEEL RAT 8]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT IS BORN, A [STAINLESS STEEL RAT 0]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT JOINS THE CIRCUS, THE [STAINLESS STEEL RAT 9]  HARRISON, HARRY  H

STAINLESS STEEL RAT JOINS THE CIRCUS, THE (PROOF)

[STAINLESS STEEL RAT 9] HARRISON, HARRY

SR-LP

STAINLESS STEEL RAT SAVES THE WORLD, THE [STAINLESS STEEL RAT 2]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT SINGS THE BLUES, THE [STAINLESS STEEL RAT 7]  HARRISON, HARRY  H:2,P
STAINLESS STEEL RAT WANTS YOU!, THE [STAINLESS STEEL RAT 4]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL RAT’S REVENGE, THE [STAINLESS STEEL RAT 3]  HARRISON, HARRY  C/P,P,H
STAINLESS STEEL VISIONS [STAINLESS STEEL RAT]  HARRISON, HARRY  C/P,P,H
STAIRWAY TO DANGER [@RICK BRANT #9]  BLAINE, JOHN  H:2
STAIRWAY TO FOREVER [STAIRWAY TO FOREVER 1]  ADAMS, ROBERT  C/P,SFWA-P,P
STALK, THE  MORRIS, JANET/MORRIS, CHRIS  P
STALKER [@PSI-MAN #5]  DAVID, PETER  P
STALKER ANALOG  ODOM, MEL  P
STALKER FROM THE STARS [HULK]  WEIN, LEN/WOLFMAN, MARV/SILVA, JOSEPH C/P,P

TITLE  AUTHOR  CODE
STALKER LORD  SPIELMAN, BRENDA GATES  P
STALKERS, THE [@TRAVELER #3]  DRUMM, D. B.  P
STALKERS  GORMAN, ED/GREENBERG, MARTIN H.  PA
STALKING DARKNESS [NIGHTRUNNER 2]  FLEWELLING, LYNN  C/P,P
STALKING HORSE, THE [HORSEGIRL 2]  ASH, CONSTANCE  C/P,P
STALKING THE NIGHTMARE  ELLISON, HARLAN  H,P
STALKING THE UNICORN  RESNICK, MIKE  C/P:2,SFWA-P,P
STALKING THE WILD RESNICK  RESNICK, MIKE  SR-H
STALKING TIME, THE [@TRAVELER #9]  DRUMM, D. B.  P
STALLION QUEEN, THE  ASH, CONSTANCE  C/P,SFWA-P,P
STAN LEE’S RIFTWORLD ODYSSEY  LEE, STAN/MCCAY, BILL  P
STAND, THE  KING, STEPHEN  C/P:2,P
STAND & DELIVER  NORTON, ANDRE  P
STAND BY FOR MARS! [TOM CORBETT 1]  ROCKWELL, CAREY  H
STAND ON ZANZIBAR  BRUNNER, JOHN  H:2,P
STAND, THE (UNABRIDGED)  KING, STEPHEN  H
STAND TO HORSE  NORTON, ANDRE  P
STANDARD CANDLES  MCDEVITT, JACK  H
STANDING DEAD, THE [STONE DANCE OF THE CHAMELEON 2]  PINTO, RICARDO  H
STANDING JOY, THE  GUIN, WYMAN  C/P,P
STANDING WAVE  HENDRIX, HOWARD V.  P
STANISAW LEM READER, A  SWIASKI, PETER  R/REF-H
STANLEY KUBRICK: A FILM ODYSSEY  PHILLIPS, GENE D.  R/REF-P
STANLEY KUBRICK’S A CLOCKWORK ORANGE  KUBRICK, STANLEY  LP
STAR (PSI CASSIOPEIA)  DEFONTENAY, C. I.  P
STAR ABOVE IT, A  OLIVER, CHAD  H
STAR-ANCHORED, STAR-ANGERED [COYOTE JONES 4]  ELGIN, SUZETTE HADEN  H:2,SFWA,P
STAR AND THE STAR-KINGS  HAMILTON, EDMOND/BRACKETT, LEIGH  H
STAR ASCENDANT [TIME MASTER 0]  COOPER, LOUISE  C/P,P,H
STAR AXE  MCGEARY, DUNCAN  P
STAR BARBARIAN  VAN ARNAM, DAVE  C/P,P
STAR BEAST, THE  HEINLEIN, ROBERT A.  C/P:2,P
STAR-BEGOTTEN  WELLS, H. G.  C/P,P,H
STAR BEGOTTEN (TWO VINYL RECORDS)  RODDEN, MARK  SR-MM
STAR BORN [PAX]  NORTON, ANDRE  C/P:2,P,DD299
STAR BRIDGE  WILLIAMSON, JACK/GUNN, JAMES E.  C/P,P,H
STAR BRIGHT  CAIDIN, MARTIN  C/P:2,SFWA-P:2,P
STAR BROTHERS [@VOYAGERS #3]  BOVA, BEN  C/P:2,SFWA-P,P,H
STAR BY STAR [@STAR WARS NEW JEDI ORDER]  DENNING, TROY  P
STAR CHILD  HOGAN, JAMES P.  C/P:2,P
STAR CHILD  STEWART, FRED MUSTARD  H,P
STAR CIRCUS  LIGHTNER, A. M.  H
STAR CITY [@HOOK #3]  ZETFORD, TULLY  P
STAR COLONIES  GREENBERG, MARTIN H./HELFERS, JOHN  C/PA,PA
STAR COLONY  LAUMER, KEITH  C/P,P,H
STAR COMMANDOS [STAR COMMANDOS 1]  GRIFFIN, P. M.  C/P:2,P
STAR CONTROL INTERBELLUM [@STAR CONTROL]  QUICK, W. T.  P
STAR CORPS [@LEGACY #1]  DOUGLAS, IAN  P
STAR COUNTRY, THE  CASSUTT, MICHAEL  H,SFWA
STAR COURIER [GRIMES]  CHANDLER, A. BERTRAM  P
STAR-CROWNED KINGS, THE  CHILSON, ROBERT  C/P:2,P
STAR CRUSADER  BALFOUR, BRUCE  C/P,P
STAR DANCERS, THE [STARDANCE 1,2]  ROBINSON, SPIDER/ROBINSON, JEANNE  P
STAR DESTROYER  BRAMSTRUP, M. A.  SR-P
STAR DIARIES, THE  LEM, STANISLAW  C/P:2,P,H
STAR DRAGON  BROTHERTON, MIKE  H
STAR DRIFTER  AYCOCK, DALE  P
STAR DRIVER  CORREY, LEE  C/P:2,P
STAR DWELLERS, THE  BLISH, JAMES  C/P:2,P
STAR FALL  BISCHOFF, DAVID F.  C/P:2,P
STAR FIGHTERS  MILLS, ROBERT E.  P
STAR FIRE  SWANN, INGO  C/P,P
STAR FLEET MEDICAL REFERENCE MANUAL  PALESTINE, EILEEN  R/REF-H,C/REF-H
STAR FLEET TECHNICAL MANUAL  JOSEPH, FRANZ  R/REF-H,C/REF-H:3
S.T.A.R. FLIGHT  TUBB, E. C.  C/P:2,P
STAR FORCE  MILLS, ROBERT E.  P
STAR FOURTEEN  POHL, FREDERIK  PA
STAR FOX, THE  ANDERSON, POUL  SR-H,H:2,P
STAR FRACTION, THE [FALL REVOLUTION]  MACLEOD, KEN  H,LP
STAR GATE  NORTON, ANDRE  C/P:2,P
STAR GHOST, THE [@STAR TREK DS9]  STRICKLAND, BRAD  LP
STAR GIANT  SKINKLE, DOROTHY  P
STAR GIRL  WINTERFELD, HENRY  C/P,LP
STAR GLADIATOR  VAN ARNAM, DAVE  SR-DB0788

The MITSFS Titledex  September 24, 2007  Page 341
TITLE  AUTHOR  CODE
STAR GOD  WOLD, ALLEN L.  H,P
STAR GOD (PROOF)  WOLD, ALLEN L.  SR-LP
STAR GODS  LOVEJOY, JACK  P
STAR GOLD  KELLEY, LEO P.  SR-P
STAR GRIFFIN  KURLAND, MICHAEL  H,SFWA
STAR GUARD [STAR SOLDIERS 1]  NORTON, ANDRE  C/P,H:2,P,DD199
STAR HEALER [SECTOR GENERAL 6]  WHITE, JAMES  C/P:2,P
STAR HUNTER  NORTON, ANDRE  DD509
STAR HUNTER AND VOODOO PLANET  NORTON, ANDRE  C/P:2,P
STAR HUNTERS [DIADEM 5]  CLAYTON, JO  C/P:2,P
STAR KA’AT  NORTON, ANDRE/MADLEE, DOROTHY  C/P,P,H
STAR KA’AT WORLD  NORTON, ANDRE/MADLEE, DOROTHY  C/P,P,H
STAR KA’ATS AND THE PLANT PEOPLE  NORTON, ANDRE/MADLEE, DOROTHY  C/P,P,H
STAR KING, THE [DEMON PRINCES 1]  VANCE, JACK  C/P:2,P

STAR KING (VIE), THE [@VANCE INTEGRAL EDITION #22/ DEMON PRINCES 1] VANCE, JACK

SR-H

STAR KINGS, THE  HAMILTON, EDMOND  C/P,P,H
STAR LIGHT  CLEMENT, HAL  C/P:2,P
STAR LIGHT, STAR BRIGHT  BESTER, ALFRED  H
STAR LOOT [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
STAR LORD  LAWRENCE, LOUISE  P
STAR LORD [@BATTLETECH 23]  PHILLIPS, DONALD G.  P

STAR LOST, THE [STAR TREK,THE NEXT GENERATION] FRIEDMAN, MICHAEL JAN/KRAUSE, PETER/

MARCO, PABLO

CX

STAR MAGICIANS, THE  CARTER, LIN  DG588
STAR MAIDENS  EVANS, IAN  SR-P,C/P
STAR MAKER  STAPLEDON, OLAF  C/P,P,LP,H
STAR MAN’S SON  NORTON, ANDRE  C/P:2,P
STAR MARINES [@LEGACY #3]  DOUGLAS, IAN  P
STAR MEN, THE  FRIEND, OSCAR J.  H
STAR MILL, THE [KALEVALA 2]  PETAJA, EMIL  C/P:2,P
STAR MOTHER  VAN SCYOC, SYDNEY J.  H:2,P
STAR OF CURSRAH [FORGOTTEN REALMS LOST EMPIRES]  EMERY, CLAYTON  P
STAR OF DANGER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
STAR OF DOOM [@RED SONJA #6]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
STAR OF ERENGRAD [@WARHAMMER]  MCINTOSH, NEIL  P
STAR OF GYPSIES  SILVERBERG, ROBERT  H:2,P
STAR OF ILL-OMEN  WHEATLEY, DENNIS  SR-P
STAR OF LIFE, THE  HAMILTON, EDMOND  H:2,P
STAR OF STARS  POHL, FREDERIK  C/PA,PA
STAR OF THE MORNING [NINE KINGDOMS]  KURLAND, LYNN  LP
STAR OF THE UNBORN  WERFEL, FRANZ  C/P:2,P
STAR PEACE: ASSURED SURVIVAL  BOVA, BEN  SFWA
STAR PIT, THE  DELANY, SAMUEL R.  SFWA-TD4,TD4
STAR PRECINCT [@STAR PRECINCT #1]  RANDLE, KEVIN/DRISCOLL, RICHARD  P
STAR PRINCE CHARLIE  ANDERSON, POUL/DICKSON, GORDON R.  C/P:2,P
STAR PROBE  GREEN, JOSEPH  P
STAR QUEST  CALDWELL, STEVEN  A
STAR QUEST  KOONTZ, DEAN R.  DH070
STAR QUEST  MILLS, ROBERT E.  P
STAR RAIDERS, THE  SUDDABY, DONALD  SR-P
STAR-RAKER  GORDON, DONALD  H
STAR RANGERS [STAR SOLDIERS 2]  NORTON, ANDRE  C/P,P,H
STAR REBEL [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
STAR RIDER  PISERCHIA, DORIS  C/P:2,P
STAR RIGGER’S WAY  CARVER, JEFFREY A.  C/P,SR-P,H:2
STAR RISK, LTD. [STAR RISK,LTD. 1]  BUNCH, CHRIS  P
STAR ROAD, THE  DICKSON, GORDON R.  H:2,SFWA,P
STAR ROGUE [GREAT IMPERIUM]  CARTER, LIN  C/P:2,P
STAR ROVER, THE  LONDON, JACK  H:2
STAR SAILORS, THE  BENNETT, GARY L.  H:2
STAR SCIENCE FICTION STORIES  POHL, FREDERIK  SR-PA,C/PA
STAR SCIENCE FICTION STORIES 2  POHL, FREDERIK  SR-PA,C/PA
STAR SCIENCE FICTION STORIES 3  POHL, FREDERIK  SR-PA,C/PA
STAR SCIENCE FICTION STORIES 4  POHL, FREDERIK  SR-PA,C/PA
STAR SCIENCE FICTION STORIES 5  POHL, FREDERIK  SR-PA,C/PA
STAR SCIENCE FICTION STORIES 6  POHL, FREDERIK  SR-PA,C/PA
STAR SCROLL, THE [DRAGON PRINCE 2]  RAWN, MELANIE  C/P:2,P
STAR-SEARCH  KAPP, COLIN  C/P:2,P
STAR SEEKERS, THE  LESSER, MILTON  H:2
STAR SEEKERS, THE  RAYER, FRANCIS G.  SR-P
STAR SHINE  BROWN, FREDRIC  C/P:2,P
STAR SHORT NOVELS  POHL, FREDERIK  SR-PA,C/PA
STAR SISTER  COULSON, JUANITA  C/P,P

The MITSFS Titledex  September 24, 2007  Page 342
TITLE  AUTHOR  CODE
STAR SLAVER [@SPACEWAYS #12]  CLEVE, JOHN  P
STAR SMASHERS OF THE GALAXY RANGERS  HARRISON, HARRY  C/P,P,H
STAR SOLDIERS [STAR SOLDIERS 1,2]  NORTON, ANDRE  P
STAR SONGS OF AN OLD PRIMATE  TIPTREE, JAMES JR.  C/P:2,P
STAR-SPANGLED FUTURE, THE  SPINRAD, NORMAN  C/P:2,P
STAR SPLIT  LASKY, KATHRYN  H
STAR SPRING  BISCHOFF, DAVID F.  C/P:2,P
STAR SURGEON  NOURSE, ALAN E.  C/P,P,H
STAR SURGEON [SECTOR GENERAL 2]  WHITE, JAMES  C/P:2,P
STAR TO EVERY WANDERING, THE [STAR TREK CRUCIBLE]  GEORGE, DAVID R. III  P
STAR TREASURE, THE  LAUMER, KEITH  SR-H,C/P:2,P
STAR TREK [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 2 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 3 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 4 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 5 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 6 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 7 [STAR TREK]  BLISH, JAMES  C/P:3,P
STAR TREK 8 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 9 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 10 [STAR TREK]  BLISH, JAMES  C/P:2,P
STAR TREK 11 [STAR TREK]  BLISH, JAMES  C/P,P,H
STAR TREK 12 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  C/P:2,P
STAR TREK (STARDATE) CALENDAR (1976)  CALENDAR  CAL
STAR TREK (VIEWMASTER STEREO PICTURES) [STAR TREK]  ANONYMOUS  SR-MM

STAR TREK ALIENS & ARTIFACTS [STAR TREK] WESTMORE, MICHAEL/SIMS, ALAN/

LOOK, BRADLEY M./BIRNES, WILLIAM J.

R/REF-P

STAR TREK — AN ANALYSIS OF A PHENOMENON IN SCIENCE FICTION  S. C. ENTERPRISES  R/REF-H
STAR TREK BLUEPRINTS  JOSEPH, FRANZ  A
STAR TREK CALENDAR 1991  CALENDAR  CAL
STAR TREK CATALOG [STAR TREK]  TURNBULL, GERRY  [C/RAN-H]
STAR TREK COMPENDIUM, THE (REVISED 1986) [STAR TREK]  ASHERMAN, ALLEN  R/REF-H
STAR TREK COMPENDIUM, THE (REVISED 1993) [STAR TREK]  ASHERMAN, ALLEN  R/REF-H
STAR TREK CONCORDANCE, THE  TRIMBLE, BJO  R/REF-H
STAR TREK CREATOR  ALEXANDER, DAVID  SFWA
STAR TREK DEEP SPACE NINE CALENDAR (1994) [STAR TREK DS9]  CALENDAR  CAL

STAR TREK ENCYCLOPEDIA, THE OKUDA, MICHAEL/OKUDA, DENISE/

MIREK, DEBBIE

R/REF-H,C/REF-H

STAR TREK FEDERATION PASSPORT [STAR TREK]  DILLARD, J. M.  B
STAR TREK FEDERATION TRAVEL GUIDE [STAR TREK]  FRIEDMAN, MICHAEL JAN  B
STAR TREK: FIRST CONTACT [STAR TREK TNG]  DILLARD, J. M.  H:2
STAR TREK GENERATIONS [STAR TREK]  DILLARD, J. M.  H
STAR TREK GUIDE, THE [STAR TREK]  RODDENBERRY, GENE  SR-S
STAR TREK II SHORT STORIES [STAR TREK]  ROTSLER, WILLIAM  P
STAR TREK II: THE WRATH OF KHAN [STAR TREK]  MCINTYRE, VONDA N.  C/P,P,H
STAR TREK III SHORT STORIES [STAR TREK]  ROTSLER, WILLIAM  P
STAR TREK III: THE SEARCH FOR SPOCK [STAR TREK]  MCINTYRE, VONDA N.  H:2,P
STAR TREK INSURRECTION [STAR TREK TNG]  DILLARD, J. M.  H
STAR TREK INTERVIEW BOOK, THE  ASHERMAN, ALLAN  R/REF-P
STAR TREK IV: THE VOYAGE HOME [STAR TREK]  MCINTYRE, VONDA N.  C/P:2,P,H

STAR TREK IV: THE VOYAGE HOME [STAR TREK] MEERSON, STEVE/KRIKES, PETER/

BENNETT, HARVE/MEYER, NICHOLAS

SR-S

STAR TREK IV, THE VOYAGE HOME  PARAMOUNT  A
STAR TREK IV: THE VOYAGE HOME (FOTONOVEL) [STAR TREK]  LERANGIS, PETER  LP

STAR TREK KLINGON [@STAR TREK TNG] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

STAR TREK LIVES!

LICHTENBERG, JACQUELINE/ MARSHAK, SONDRA/WINSTON, JOAN

R/REF-P,C/REF-P:2

STAR TREK LOG 1 [STAR TREK]  FOSTER, ALAN DEAN  C/P:3,P
STAR TREK LOG ONE LOG TWO LOG THREE [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 2 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 3 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 4 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG FOUR LOG FIVE LOG SIX [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 5 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 6 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 7 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG SEVEN LOG EIGHT LOG NINE [STAR TREK]  FOSTER, ALAN DEAN  P
STAR TREK LOG 8 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 9 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK LOG 10 [STAR TREK]  FOSTER, ALAN DEAN  C/P:2,P
STAR TREK MAPS  MAYNARD, JEFF  A
STAR TREK MEMORIES  SHATNER, WILLIAM/KRESKI, CHRIS  R/REF-H
STAR TREK MOVIE MEMORIES  SHATNER, WILLIAM  R/REF-H,C/REF-P

The MITSFS Titledex  September 24, 2007  Page 343
TITLE  AUTHOR  CODE
STAR TREK: NEMESIS [STAR TREK TNG]  DILLARD, J. M.  H
STAR TREK PUZZLE MANUAL  RAZZI, JAMES  [C/RAN-P:2]
STAR TREK QUIZ BOOK  ANDREWS, BART/DUNNING, BRAD  [C/RAN-P:2]
STAR TREK QUOTE JOURNAL [STAR TREK]  PARAMOUNT PICTURES  [C/RAN-H]
STAR TREK READER, THE [STAR TREK]  BLISH, JAMES  H:2
STAR TREK READER II, THE [STAR TREK]  BLISH, JAMES  H
STAR TREK READER III, THE [STAR TREK]  BLISH, JAMES  H
STAR TREK READER IV, THE [STAR TREK]  BLISH, JAMES  H
STAR TREK SPACEFLIGHT CHRONOLOGY [STAR TREK]  GOLDSTEIN, STAN/GOLDSTEIN, FRED  VLP

STAR TREK SPEAKS SACKETT, SUSAN/GOLSTEIN, FRED/

GOLDSTEIN, STAN

[C/RAN-H]

STAR TREK: STRANGE NEW WORLDS [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

HA:2,LPA

STAR TREK: STRANGE NEW WORLDS II [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

C/PA,PA

STAR TREK: STRANGE NEW WORLDS III [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS IV [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS V [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS VI [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS VII [STAR TREK] SMITH, DEAN WESLEY/ORDOVER, JOHN J./ KASSIN, ELISA J./BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 8 [STAR TREK] SMITH, DEAN WESLEY/KASSIN, ELISA J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 9 [STAR TREK] SMITH, DEAN WESLEY/KASSIN, ELISA J./

BLOCK, PAULA M.

LPA

STAR TREK: STRANGE NEW WORLDS 10 [STAR TREK]  SMITH, DEAN WESLEY/BLOCK, PAULA M.  LPA
STAR TREK: THE CLASSIC EPISODES 1 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  P
STAR TREK: THE CLASSIC EPISODES 2 [STAR TREK]  BLISH, JAMES  P
STAR TREK: THE CLASSIC EPISODES 3 [STAR TREK]  BLISH, JAMES/LAWRENCE, J. A.  P

STAR TREK: THE KEY COLLECTION [STAR TREK KEY COLLECTION 1] RODDENBERRY, GENE/ZACCARA, NEVIO/

GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION: VOLUME 2 [STAR TREK KEY COLLECTION 2] RODDENBERRY, GENE/GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION VOLUME 3 [STAR TREK KEY COLLECTION 3] GOLD KEY COMICS

CX

STAR TREK: THE KEY COLLECTION: VOLUME 4 [STAR TREK KEY COLLECTION 4] RODDENBERRY, GENE/ZACCARA, NEVIO/

GIOLITTI, ALBERTO

CX

STAR TREK: THE KEY COLLECTION: VOLUME 5 [STAR TREK KEY COLLECTION 5] RODDENBERRY, GENE/MCWILLIAMS, AL/ ZACCARA, NEVIO/GIOLITTI, ALBERTO

CX

STAR TREK: THE MAGNA [STAR TREK]  SHINSEI, SHINSEI  SCX
STAR TREK: THE MOTION PICTURE [STAR TREK]  RODDENBERRY, GENE  C/P,P,H
STAR TREK: THE MOTION PICTURE [STAR TREK]  RODDENBERRY, GENE/LIVINGSTON, HAROLD SR-S
STAR TREK: THE NEW VOYAGES [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  SR-PA,C/PA:2
STAR TREK: THE NEW VOYAGES 2 [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  SR-PA,C/PA:2

STAR TREK THE NEXT GENERATION TECHNICAL MANUAL [STAR TREK TNG] STERNBACH, RICK/OKUDA, MICHAEL

R/REF-H,C/REF-H

STAR TREK THE NEXT GENERATION TECHNICAL MANUAL (FIFTH SEASON EDITION) [STAR TREK TNG] STERNBACH, RICK/OKUDA, MICHAEL

R/REF-H

STAR TREK V: THE FINAL FRONTIER [STAR TREK]  DILLARD, J. M.  C/P:2,P
STAR TREK VI: THE UNDISCOVERED COUNTRY [STAR TREK]  DILLARD, J. M.  C/P:2,P
STAR TREK: VOYAGE TO ADVENTURE  DODGE, MICHAEL J.  P
STAR TREK’S GREATEST GUEST STARS  MCDONNELL, DAVID  R/REF-P,C/REF-P
STAR TRIAL  GREENFIELD, IRVING A.  P
STAR VENTURERS, THE  BULMER, KENNETH  D22600
STAR VIRUS, THE  BAYLEY, BARRINGTON J.  D78400
STAR VOYAGER ACADEMY [JUSTIN BELL 1]  FORSTCHEN, WILLIAM R.  C/P,P
STAR WARPED  ROBERTS, A3R  H
STAR WARS [STAR WARS]  LUCAS, GEORGE  C/P,H,SR-P
STAR WARS [STAR WARS]  THOMAS, ROY/CHAYKIN, HOWARD  SCX
STAR WARS (RADIO SHOW) (7 CASSETTES) [STAR WARS]  LUCAS, GEORGE  SR-MM
STAR WARS BLUEPRINTS [STAR WARS]  BALLANTINE BOOKS  A
STAR WARS BOOK ONE [STAR WARS]  DAVIDS, PAUL/DAVIDS, HOLLACE  H
STAR WARS CALENDAR (1978)  CALENDAR  CAL
STAR WARS EPISODE 1 JOURNAL: DARTH MAUL [STAR WARS]  WATSIN, JUDE  LP
STAR WARS EPISODE I: INCREDIBLE CROSS-SECTIONS  REYNOLDS, DAVID WEST  [C/RAN-H]
STAR WARS EPISODE I: THE VISUAL DICTIONARY  REYNOLDS, DAVID WEST  [C/RAN-H]
STAR WARS INTERGALACTIC PASSPORT, THE  BJORKLUND, CHARLES R.  R/REF-P

STAR WARS: NATIONAL PUBLIC RADIO DRAMATIZATION, THE [STAR WARS] DALEY, BRIAN

LP

STAR WARS POSTER ART CALENDAR (1980)  CALENDAR  CAL

The MITSFS Titledex  September 24, 2007  Page 344
TITLE  AUTHOR  CODE
STAR WARS RETURN OF THE JEDI (1984)  CALENDAR  CAL
STAR WARS SKETCHBOOK, THE [STAR WARS]  JOHNSTON, JOE  A
STAR WARS THE EMPIRE STRIKES BACK CALENDAR (1981)  CALENDAR  CAL
STAR WARS THE GREATEST BATTLES (3-D)  GOLDEN BOOKS  A

STAR WARS: THE SCRIPTS [STAR WARS] LUCAS, GEORGE/BRACKETT, LEIGH/

KASDAN, LAWRENCE

SR-S

STAR WARS TRILOGY, THE [STAR WARS]  BALLANTINE BOOKS  LP
STAR WASPS, THE  WILLIAMS, ROBERT MOORE  DF177
STAR WATCHMAN  BOVA, BEN  C/P:2,P
STAR WAYS  ANDERSON, POUL  C/P:2,P,DD255
STAR WEB  COX, JOAN  C/P,P
STAR WEB, THE  ZEBROWSKI, GEORGE  SR-P
STAR WELL [ANTHONY VILLIERS 1]  PANSHIN, ALEXEI  SR-P,C/P:2,P
STAR WIND  WOOLVERTON, LINDA  SFWA
STAR WINDS  BAYLEY, BARRINGTON J.  C/P,P
STAR WITCHES, THE  LYMINGTON, JOHN  P
STAR WOLF!  WHITE, TED  C/P:2,P
STARBLOOD  KOONTZ, DEAN R.  C/P,P
STARBORNE  SILVERBERG, ROBERT  C/P,P,H
STARBRAT  MORRESSY, JOHN  C/P,P,H
STARBREED  CLOW, MARTHA DEMEY  C/P:2,P
STARBRIDGE [@STARBRIDGE #1]  CRISPIN, A. C.  C/P:2,SFWA-P:2,P
STARBRIGHT  CASTLE, DAMON  P
STARBURST  BESTER, ALFRED  C/P:2,P
STARBURST  POHL, FREDERIK  C/P,P,H
STARCHILD  POHL, FREDERIK/WILLIAMSON, JACK  C/P,P,H
STARCHILD TRILOGY, THE  POHL, FREDERIK/WILLIAMSON, JACK  H:2,P
STARCROSSED, THE  BOVA, BEN  C/P:2,SFWA-P,P
STARDANCE [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  H:2,P
STARDANCE (NOVELLA) [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  H
STARDANCE (PROOF) [STARDANCE 1]  ROBINSON, SPIDER/ROBINSON, JEANNE  SR-LP
STARDEATH  TUBB, E. C.  C/P,P
STARDOC [STARDOC 1]  VIEHL, S. L.  C/P,P
STARDREAMER  SMITH, CORDWAINER  C/P:2,P
STARDRIFT  MORRESSY, JOHN  C/P,P
STARDRIFT  PETAJA, EMIL  H
STARDROPPERS, THE  BRUNNER, JOHN  C/P:2,P
STARDUST  CZERNEDA, JULIE E.  LPA
STARDUST  GAIMAN, NEIL  C/P,P,H
STARDUST VOYAGES, THE  TALL, STEPHEN  C/P:2,P
STARFALL [STAR*DRIVE]  GREENBERG, MARTIN H.  PA
STARFALL [@STAR TREK TNG STARFLEET ACADEMY #8]  STRICKLAND, BRAD/STRICKLAND, BARBARA LP
STARFARERS  ANDERSON, POUL  H:2
STARFARERS [STARFARERS 1]  MCINTYRE, VONDA N.  C/P:2,SFWA-P,P
STARFAWN  PREISS, BYRON/FABIAN, STEPHEN  C/P:2,SCX
STARFIGHTERS OF ADUMAR [@STAR WARS X-WING 9]  ALLSTON, AARON  C/P,P
STARFINDER  YOUNG, ROBERT F.  C/P:2,P
STARFIRE  BUCKNER, ROBERT  C/P:2,P
STARFIRE  PREUSS, PAUL  C/P:2,SFWA-P,P,H
STARFIRE  SHEFFIELD, CHARLES  C/P,LP,H
STARFIRE DOWN [CONEK HAYDEN]  MCKEONE, LEE  P
STARFISH [RIFTERS 1]  WATTS, PETER  H,P
STARFISHERS [STARFISHERS 2]  COOK, GLEN  C/P,P
STARFLEET ACADEMY [STAR TREK]  CAREY, DIANE  C/P,P
STARFLEET YEAR ONE [STAR TREK]  FRIEDMAN, MICHAEL JAN  P
STARFLIGHT 3000  MACKELWORTH, R. W.  C/P,P
STARFLIGHT TO FAROUL  DEAREN, PATRICK  P
STARFOLLOWERS OF CORAMONDE, THE [CORAMONDE 2]  DALEY, BRIAN  C/P:2,P
STARGATE [STARGATE]  DEVLIN, DEAN/EMMERICH, ROLAND  P
STARGATE  GEDGE, PAULINE  C/P,P,H
STARGATE  ROBINETT, STEPHEN  C/P,P
STARGATE (ADAPTED FOR YOUNGER READERS) [STARGATE]  BLACK, SHEILA  LP
STARGATE SG-1 [@STARGATE SG-1 (MCCONNELL) 1]  MCCONNELL, ASHLEY  C/P,P
STARGAZER (LARGE PAPERBACK AND CASSETTE) [STARGAZER 1]  HAUSMAN, GERALD  SR-MM
STARHAMMER  ROWLEY, CHRISTOPHER  C/P,P
STARHAVEN  JORGENSEN, IVAR  DD351
STARHAWK [@STARHAWK 1]  MALONEY, MACK  P
STARHIKER  DANN, JACK  H:2,SFWA
STARHUNT [STAR WOLF 0]  GERROLD, DAVID  P
STARJACKED!  GREENLEAF, WILLIAM  P
STARKADDER [HATHER LAMBISSON 1]  KING, BERNARD  H
STARKAHN OF RHADA, THE [RHADA 4]  GILMAN, ROBERT CHAM  C/P,P,H
STARK’S COMMAND [ETHAN STARK 2]  HEMRY, JOHN G.  P
STARK’S CRUSADE [ETHAN STARK 3]  HEMRY, JOHN G.  P

The MITSFS Titledex  September 24, 2007  Page 345
TITLE  AUTHOR  CODE
STARK’S WAR [ETHAN STARK 1]  HEMRY, JOHN G.  C/P,P
STARLESS NIGHT [FORGOTTEN REALMS LEGACY OF THE DROW 2]  SALVATORE, R. A.  H,P
STARLESS REALM, THE [@PERRY RHODAN #87]  ACE  PA
STARLESS WORLD, THE [STAR TREK]  EKLUND, GORDON  C/P:2,LP
STARLIGHT  BESTER, ALFRED  H:3,P
STARLIGHT 1  NIELSEN HAYDEN, PATRICK  LPA
STARLIGHT 2  NIELSEN HAYDEN, PATRICK  HA,LPA
STARLIGHT 3  NIELSEN HAYDEN, PATRICK  HA
STARLIGHT CRYSTAL, THE  PIKE, CHRISTOPHER  P
STARLINER  DRAKE, DAVID  C/P,P

STARLOG: SCIENCE FICTION TRIVIA MCDONNELL, DAVID/SAYERS, JOHN/

SMITH, TIM L.

[C/RAN-P]

STARLUCK  WISMER, DONALD  P
STARMAN [WAYFARER REDEMPTION 3]  DOUGLASS, SARA  H:2
STARMAN  FOSTER, ALAN DEAN  C/P:2,P
STARMAN JONES  HEINLEIN, ROBERT A.  C/P:2,P
STARMAN’S QUEST  SILVERBERG, ROBERT  H
STARMASTERS’ GAMBIT  KLEIN, GERARD  C/P:2,P
STARMEN, THE  BRACKETT, LEIGH  H
STARMEN OF LLYRDIS, THE  BRACKETT, LEIGH  C/P:2,P
STARMIND [STARDANCE 3]  ROBINSON, SPIDER/ROBINSON, JEANNE  C/P,P,H:2
STARMIND  VAN ARNAM, DAVE  C/P:2,P
STAROAMER’S FATE  ROTHMAN, CHUCK  C/P:2,P
STARPILOT’S GRAVE [MAGEWORLDS 2]  DOYLE, DEBRA/MACDONALD, JAMES D.  C/P,P
STARPIRATE’S BRAIN [EXCHAMELEON 2]  GOULART, RON  C/P:2,P
STARPLEX  SAWYER, ROBERT J.  C/P,P,H
STARQUAKE  FORWARD, ROBERT L.  C/P,P,H,SFWA
STARRIGGER [SKYWAY 1]  DE CHANCIE, JOHN  C/P:2,P
STARRISE AT CORRIVALE [STAR*DRIVE HARBINGER 1]  DUANE, DIANE  P
STARRY MESSENGER: THE BEST OF GALILEO  RYAN, CHARLES C.  HA
STARRY RIFT, THE  TIPTREE, JAMES JR.  H:2,SFWA-P
STARS  IAN, JANICE/RESNICK, MIKE  HA
STARS  ZIM, HERBERT/BAKER, ROBERT H.  [C/RAN-P]
STARS & STRIPES FOREVER [STARS & STRIPES 1]  HARRISON, HARRY  C/P,P,H
STARS & STRIPES IN PERIL [STARS & STRIPES 2]  HARRISON, HARRY  H,P
STARS & STRIPES TRIUMPHANT [STARS & STRIPES 3]  HARRISON, HARRY  P
STARS ARE ALSO FIRE, THE [HARVEST OF STARS 3]  ANDERSON, POUL  H,P
STARS ARE OURS, THE  BULMER, H. K.  H
STARS ARE OURS!, THE [PAX]  NORTON, ANDRE  C/P,H,DD121,P
STARS ARE THE STYX, THE  STURGEON, THEODORE  C/P,P,H
STARS ARE TOO HIGH, THE  BAHNSON, AGNEW H. JR.  C/P:2,P
STARS AROUND US, THE  HOSKINS, ROBERT  C/PA,PA
STARS ASUNDER, THE [MAGEWORLDS 6]  DOYLE, DEBRA/MACDONALD, JAMES D.  H:2
STARS AT WAR, THE [@IMPERIAL STARS #1]  POURNELLE, JERRY  C/PA,PA
STARS CAME DOWN, THE  SAXON, RICHARD  H
STARS COMPEL, THE [TOMMASO 2]  ROESSNER, MICHAELA  P
STARS DISPOSE, THE [TOMMASO 1]  ROESSNER, MICHAELA  C/P,P
STARS’ END [STARFISHERS 3]  COOK, GLEN  C/P,P
STARS IN MY POCKET LIKE GRAINS OF SAND  DELANY, SAMUEL R.  C/P,P,H
STARS IN SHROUD, THE  BENFORD, GREGORY  H:3,P
STARS IN THEIR COURSES, THE  ASIMOV, ISAAC  [C/RAN-P]
STARS, LIKE DUST, THE [GALACTIC EMPIRE 1]  ASIMOV, ISAAC  C/P:2,P,H
STARS MUST WAIT, THE [BOLOS]  LAUMER, KEITH  C/P:2,SFWA-P,P
STARS MY DESTINATION, THE  BESTER, ALFRED  C/P,H:2,P,SFWA
STARS OVER STARS  WENTWORTH, K. D.  P
STARS THROUGH MAGIC CASEMENTS, THE  WILLIAMSON, JULIA  [C/RAN-H]
STARS WILL JUDGE, THE  GREENFIELD, IRVING A.  P
STARS WILL SPEAK, THE [SUNSPACE 2]  ZEBROWSKI, GEORGE  H
STARS WILL WAIT, THE  HASSE, HENRY L.  H
STARSEED [STARDANCE 2]  ROBINSON, SPIDER/ROBINSON, JEANNE  C/P,SFWA-P:2,P,H
STARSEEKERS, THE  GARNETT, DAV  C/P:2,P
STARSHADOWS  SARGENT, PAMELA  C/P:2,P
STARSHINE  STURGEON, THEODORE  C/P:2,P
STARSHIP  ALDISS, BRIAN W.  C/P:2,P
STARSHIP  ANDERSON, POUL  C/P:2,P
STARSHIP [@EXPLORATION #2]  RANDLE, KEVIN  C/P,P
STARSHIP AND HAIKU  SOMTOW, S. P.  C/P:2,P
STARSHIP AND HAIKU  SUCHARITKUL, SOMTOW  C/P
STARSHIP AND THE CANOE, THE  BROWER, KENNETH  [C/RAN-P]
STARSHIP DEATH  GARRETT, RANDALL  C/P,P
STARSHIP INVINCIBLE  KELLY, FRANK K.  LP
STARSHIP INVINCIBLE (PROOF)  KELLY, FRANK K.  SR-VLP
STARSHIP: MUTINY [STARSHIP THEODORE ROOSEVELT 1]  RESNICK, MIKE  H
STARSHIP:PIRATE [STARSHIP 2]  RESNICK, MIKE  H

The MITSFS Titledex  September 24, 2007  Page 346
TITLE  AUTHOR  CODE
STARSHIP SAPPHIRE [@SPACEWAYS #15]  CLEVE, JOHN  P
STARSHIP TRAP, THE [STAR TREK]  GILDEN, MEL  C/P:2,P
STARSHIP TROOPERS  HEINLEIN, ROBERT A.  C/P,P,H

STARSHIPS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

STARSILK  VAN SCYOC, SYDNEY J.  P
STARSKY AND HUTCH #6: THE PSYCHIC  FRANKLIN, MAX  P
STARSONG  PARKINSON, DAN  C/P:2,P
STARSONGS AND UNICORNS  NORDEN, ERIC  C/P,P
STARSPAWN  VON GUNDEN, KENNETH  P
STARSPINNER  AYCOCK, DALE  P
STARSTONE, THE [TALES OF GOM]  CHETWIN, GRACE  P
STARSTREAK  OWEN, BETTY M.  PA
STARSTRIKE  GEAR, W. MICHAEL  C/P:2,P
STARSWARM  ALDISS, BRIAN W.  C/P:2,P
STARSWARM  POURNELLE, JERRY  H,P
STARTIDE RISING [UPLIFT 2]  BRIN, DAVID  C/P:2,P
STARTIDE RISING (REVISED) [UPLIFT 2]  BRIN, DAVID  C/P,P,H
STARTLING STORIES  MINES, SAMUEL  HA
STARTLING WORLDS OF HENRY KUTTNER, THE  KUTTNER, HENRY  P
STARWATER STRAINS  WOLFE, GENE  H
STARWINGS  PROCTOR, GEO. W.  P
STARWOLF [STARWOLF 1,2,3]  HAMILTON, EDMOND  C/P:2,P
STARWOLVES, THE [STARWOLVES 1]  GUNNARSSON, THORARINN  C/P,P
STARWOLVES DREADNOUGHT [STARWOLVES 4]  GUNNARSSON, THORARINN  P
STARWORLD [TO THE STARS 3]  HARRISON, HARRY  C/P:2,P
STATE OF CHANGE [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P:2,P
STATE OF DISOBEDIENCE, A (PROOF)  KRATMAN, TOM  SR-LP
STATE OF GRACE  WILHELM, KATE  SR-H
STATE OF THE ART, THE [CULTURE]  BANKS, IAIN M.  H,P
STATESMAN [@BIO OF A SPACE TYRANT #5]  ANTHONY, PIERS  C/P:2,P
STATION 7  HUNT, GILL  SR-P
STATION GEHENNA  WEINER, ANDREW  P
STATION IN SPACE  GUNN, JAMES E.  C/P:2,P
STATION OF THE INVISIBLES [PERRY RHODAN 133]  MAHR, KURT  SR-LP
STATION RAGE [STAR TREK DS9]  CAREY, DIANE  C/P,P
STATION X  WINSOR, G. MCLEOD  H
STATION ZERO-ZERO  SULLIVAN, MIKE  P
STATIONFALL [PLANETFALL 2]  COVER, ARTHUR BYRON  C/P,P
STATIONS OF THE NIGHTMARE  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
STATIONS OF THE TIDE  SWANWICK, MICHAEL  C/P,P,H
STATUS CIVILIZATION, THE  SHECKLEY, ROBERT  C/P:2,P
STATUS CIVILIZATION, THE/NOTIONS: UNLIMITED  SHECKLEY, ROBERT  C/P:2,P
STATUS: DEADZONE [WARHAMMER 40K]  GASCOIGNE, MARC/JONES, ANDY  PA
STATUS QUOTIENT: THE CARRIER  SPERRY, RALPH A.  C/P:2,P
STAY OUT OF THE BASEMENT [GOOSEBUMPS]  STINE, R. L.  LP
STAYING DEAD [RETRIEVERS]  GILMAN, LAURA ANNE  H,LP
STEAL THE DRAGON  BRIGGS, PATRICIA  P
STEALER OF DREAMS, THE [DOCTOR WHO NINTH DOCTOR]  LYONS, STEVE  H
STEALER OF SOULS, THE [ELRIC]  MOORCOCK, MICHAEL  SR-H,C/P:2
STEALER’S SKY [@THIEVES’ WORLD #12]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA,SFWA-P
STEALING MAGIC  HUFF, TANYA  H,LP
STEALING THE ELF-KING’S ROSES  DUANE, DIANE  C/P,P
STEALING THUNDER [@DESCENT 2]  TELEP, PETER  P
STEALTH PLANET  BLACKHALL, RAY  H,LP
STEALTH PLANET (PROOF)  BLACKHALL, RAY  H,SR-LP
STEAM BIRD  SCHENCK, HILBERT  P,SFWA-P
STEAM-DRIVEN BOY, THE  SLADEK, JOHN  SR-P,C/P
STEAMPUNK TRILOGY, THE  DI FILIPPO, PAUL  H
STEEL, THE [@GENERAL #4]  STIRLING, S. M./DRAKE, DAVID  P,SFWA-P
STEEL AND STONE [@DRAGONLANCE MEETINGS 5]  PORATH, ELLEN  C/P,P
STEEL BEACH  VARLEY, JOHN  C/P:2,P,H
STEEL BREAKFAST ERA, THE  MELLICK, CARLTON III  ERD3
STEEL BROTHER  DICKSON, GORDON R.  C/P:2,SFWA-P,P
STEEL CROCODILE, THE  COMPTON, D. G.  C/P:2,P
STEEL GAUNTLET [@STARFIST 3]  SHERMAN, DAVID/CRAGG, DAN  P
STEEL HELIX  ZEDDIES, ANN TONSOR  P
STEEL KINGS, THE [STORMLANDS 4]  ROBERTS, JOHN MADDOX  P
STEEL MAGIC  NORTON, ANDRE  P
STEEL OF RAITHSKAR, THE [GANDALARA 1]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
STEEL RAIN [@SHADOWRUN 24]  SMITH, NYX  P
STEEL ROSE  DALKEY, KARA  C/P,P
STEEL, THE MIST, AND THE BLAZING SUN, THE  ANVIL, CHRISTOPHER  P
STEEL THRONE, THE [FIVE RINGS FOUR WINDS 0]  BOLME, EDWARD  P

TITLE  AUTHOR  CODE
STEEL TRAP [@ENFORCER #6]  SUGAR, ANDREW  P
STEEL TSAR, THE [OSWALD BASTABLE 3]  MOORCOCK, MICHAEL  C/P,P
STEELDRIVER  DEBRANDT, DON H.  P
STEELE [STEELE 1]  MASTERS, J. D.  P
STEELEYE: WATERSPACE [STEELEYE]  DUNN, SAUL  SR-P
STEELHEART  DIETZ, WILLIAM C.  P
STEERSWOMAN, THE [STEERSWOMAN 1]  KIRSTEIN, ROSEMARY  C/P:2,P
STELLA RADIUM DISCHARGE  LUNA, KRIS  SR-P,H
STELLAR #1  DEL REY, JUDY-LYNN  PA,C/PA
STELLAR #2  DEL REY, JUDY-LYNN  PA,C/PA
STELLAR #3  DEL REY, JUDY-LYNN  C/PA,PA
STELLAR #4  DEL REY, JUDY-LYNN  C/PA,PA
STELLAR #5  DEL REY, JUDY-LYNN  PA,C/PA,SFWA-P
STELLAR #6  DEL REY, JUDY-LYNN  PA
STELLAR #7  DEL REY, JUDY-LYNN  PA,C/PA
STELLAR DEATH PLAN, THE [@MASTERS OF SPACE #1]  VARDEMAN, ROBERT E.  C/P,P
STELLAR FIST  PROCTOR, GEO. W.  P
STELLAR MISSILES, THE  REPP, ED EARL  H,SR-LP
STELLAR RANGER [STELLAR RANGER 1]  PERRY, STEVE  C/P,P
STELLAR SHORT NOVELS  DEL REY, JUDY-LYNN  PA,C/PA
STEN [STEN 1]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
STEP FARTHER OUT, A  POURNELLE, JERRY  [C/RAN-P:2]
STEP INTO CHAOS [@QUEST FOR TOMORROW 3]  SHATNER, WILLIAM  P
STEP TO THE STARS  DEL REY, LESTER  H:2,P
STEPFATHER BANK  POYER, D. C.  H:2,P
STEPFORD WIVES, THE  LEVIN, IRA  P

STEPHEN KING’S THE DARK TOWER A CONCORDANCE VOLUME I [DARK TOWER] FURTH, ROBIN

R/REF-P

STEPPE  ANTHONY, PIERS  H:2,P
STEPPENWOLF  HESSE, HERMAN  C/P,P
STEPPING FROM THE SHADOWS  MCKILLIP, PATRICIA A.  C/P,P
STEPS OF THE SUN, THE  TEVIS, WALTER  C/P,SFWA,SFWA-P,P,H
STEPS UP THE CHIMNEY, THE [MAGICIAN’S HOUSE 1]  CORLETT, WILLIAM  P
STEPSONS OF TERRA  SILVERBERG, ROBERT  C/P,DD311
STEPWATER [ARBITER]  DOUGLAS, L. WARREN  P
STERKARM HANDSHAKE, THE  PRICE, SUSAN  H
STERNEN KAMPFER  VOLTZ, WILLIAM CH.  FH
STEVEN BRUST’S JHEREG [DRAGAERA VLAD TALTOS]  ZELENETZ, ALAN/PIERARD, JOHN  CX
STICKMAN  PFEFFERLE, SETH  SFWA-P
STILL, THE [RODRIGO OF CALEDON 1]  FEINTUCH, DAVID  C/P:2,LP
STILL FORMS ON FOXFIELD  SLONCZEWSKI, JOAN  C/P,P
STILL I PERSIST IN WONDERING  PANGBORN, EDGAR  C/P:2,P
STILL LIFE  HORLAK, E. E.  C/P,P
STILL RIVER  CLEMENT, HAL  H:2,P
STILL, SMALL VOICE OF TRUMPETS, THE [CULTURAL SURVEY 1]  BIGGLE, LLOYD JR.  C/P,P,H:2
STING OF THE SCORPION, THE [GLOBAL 2000 3]  BASILE, GLORIA VITANZA  P
STINGER  KRESS, NANCY  C/P,P
STINGER  MCCAMMON, ROBERT R.  P
STINGRAY  THEYDON, JOHN  SR-P
STINKER FROM SPACE [STINKER 1]  SERVICE, PAMELA F.  P
STINKER’S RETURN [STINKER 2]  SERVICE, PAMELA F.  SFWA
STIR OF ECHOES, A  MATHESON, RICHARD  C/P,P
STITCH IN SNOW  MCCAFFREY, ANNE  P
STITCH IN TIME, A [STAR TREK DS9]  ROBINSON, ANDREW J.  C/P,P
STOCHASTIC MAN, THE  SILVERBERG, ROBERT  C/P:2,P,H
STOCKHOLDERS IN DEATH [@AVENGER #7]  ROBESON, KENNETH  C/P:2,P
STOLEN [WOMEN OF THE OTHERWORLD 2]  ARMSTRONG, KELLEY  P
STOLEN CHILD, THE  DONOHUE, KEITH  LP
STOLEN FACES  BISHOP, MICHAEL  P
STOLEN GODDESS, THE [KAPHTU 2]  PURTILL, RICHARD  C/P:2,P
STOLEN RIVER, THE  WILLIS, PAUL J.  P

STOLEN SPACEFLEET, THE [@PERRY RHODAN #109]/ SGT. ROBOT [PERRY RHODAN 110] ACE

PA

STOLEN SUN, THE [KALEVALA 3]  PETAJA, EMIL  DG618
STOLEN THRONE, THE [TIME OF TROUBLES 1]  TURTLEDOVE, HARRY  P
STONE  ROBERTS, ADAM  C/P,P,H
STONE & SEA [STONE 2]  EDWARDS, GRAHAM  P
STONE & SKY [STONE 1]  EDWARDS, GRAHAM  P
STONE & SUN [STONE 3]  EDWARDS, GRAHAM  P
STONE 588  BROWNE, GERALD A.  P
STONE AND ANVIL [@STAR TREK NEW FRONTIER 14]  DAVID, PETER  P
STONE AND THE FLUTE, THE  BEMMANN, HANS  H,LP
STONE AND THE MAIDEN, THE [PANDRAGORE 1]  JONES, DENNIS  C/P,P,H
STONE ANGELS  JEFFERIES, MIKE  P

The MITSFS Titledex  September 24, 2007  Page 348
TITLE  AUTHOR  CODE
STONE CANAL, THE [FALL REVOLUTION]  MACLEOD, KEN  C/P,LP
STONE CARNATION, THE  HINTZE, NAOMI A.  P

STONE-COLD DEAD IN THE MARKET AFFAIR, THE [@MAN FROM UNCLE #22] ORAM, JOHN

C/P,P

STONE-COLD DEAD IN THE MARKET AFFAIR, THE (BRITISH)

[@MAN FROM UNCLE #4] ORAM, JOHN

SR-P

STONE DOGS, THE [DRAKA 3]  STIRLING, S. M.  C/P,P
STONE GARDEN, THE  ROSENBLUM, MARY  C/P,P
STONE GIANT, THE  BLAYLOCK, JAMES P.  SFWA-P
STONE GOD AWAKENS, THE  FARMER, PHILIP JOSE  C/P:2,P
STONE IN HEAVEN, A [DOMINIC FLANDRY]  ANDERSON, POUL  H:2,P
STONE MAN, THE [DOC SAVAGE #81]  ROBESON, KENNETH  @P
STONE MOVERS, THE  MULLEN, PATRICIA  P
STONE OF CHASTITY, THE  SHARP, MARGERY  H
STONE OF FAREWELL [MEMORY,SORROW AND THORN 2]  WILLIAMS, TAD  H,P
STONE OF STARS, THE [@DRAGON THRONE #1]  BAIRD, ALISON  LP
STONE OF TEARS [SWORD OF TRUTH 2]  GOODKIND, TERRY  C/P:2,P
STONE OF TIME, THE [RUNESWORD 6]  ESTES, ROSE/WHAM, TOM  P
STONE PRINCE, THE [LIVING FLAME 1]  PATTON, FIONA  C/P:2,P
STONE ROSE, THE [@DOCTOR WHO]  RAYNER, JACQUELINE  H
STONE THAT NEVER CAME DOWN, THE  BRUNNER, JOHN  C/P,SFWA,P,H:2
STONE WAR, THE  ROBINS, MADELEINE E.  C/P,P
STONE WITHIN, THE [CHUNG KUO 4]  WINGROVE, DAVID  C/P:2,LP
STONE WIZARD, THE  STOREY, MARGARET  P
STONECLIFF  NATHAN, ROBERT  P
STONECUTTER’S STORY [@LOST SWORDS #3]  SABERHAGEN, FRED  H:2,SFWA-P,P
STONEFACE [@DEATHLANDS]  AXLER, JAMES  P
STONEHENGE  HARRISON, HARRY/STOVER, LEON E.  C/P,SFWA,P
STONEHENGE GATE, THE  WILLIAMSON, JACK  H

STONES OF NOMURU, THE DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

C/P:2,SFWA-P,P

STONES OF STIGA, THE [@SHUNLAR 3]  HELLER, CAROL  P
STONESKIN’S REVENGE [OTHERWORLD 5]  DEITZ, TOM  C/P,P
STOP THAT WITCH!  CLARK, MARY  P
STOPPING AT SLOWYEAR  POHL, FREDERIK  C/P:2,P
STOPPING FOR A SPELL  JONES, DIANA WYNNE  LP
STOPWATCH  HAY, GEORGE  SR-PA
STORE OF INFINITY  SHECKLEY, ROBERT  C/P:2,P

STOREHOUSES OF THE SNOW WOODARD, EDWIN/

BISCHOFF, HEATHER WOODARD

P

STOREYS FROM THE OLD HOTEL  WOLFE, GENE  H
STORIES  YEFREMOV, IVAN  H

STORIES BY MAMA LANSDALE’S YOUNGEST BOY [@AUTHOR’S CHOICE MONTHLY #18] LANSDALE, JOE R.

SR-H

STORIES FOR AN ENCHANTED AFTERNOON  RUSCH, KRISTINE KATHRYN  H
STORIES FOR THE DEAD OF NIGHT  CONGDON, DON  PA
STORIES FROM THE HUGO WINNERS II  ASIMOV, ISAAC  C/PA,PA
STORIES FROM THE TWILIGHT ZONE  SERLING, ROD  C/P:2,P
STORIES OF DARKNESS AND DREAD  BRENNAN, JOSEPH PAYNE  SR-H
STORIES OF OTHER WORLDS & HONEYMOON IN SPACE, A  GRIFFITH, GEORGE  LP
STORIES OF RAY BRADBURY, THE  BRADBURY, RAY  H
STORIES OF SUSPENSE  SCHOLASTIC BOOKS  PA
STORIES OF THE SUPERNATURAL  OWEN, BETTY M.  PA
STORIES OF YOUR LIFE AND OTHERS  CHIANG, TED  H
STORK FACTOR, THE  HUGHES, ZACH  C/P:2,P
STORK NAKED [@XANTH #30]  ANTHONY, PIERS  H
STORM AT ELDALA [STAR*DRIVE HARBINGER 2]  DUANE, DIANE  P
STORM BREAKING [MAGE STORMS 3]  LACKEY, MERCEDES  C/P,P,H:2
STORM CALLER [ISLAND WARRIOR 2]  SEVERANCE, CAROL  P
STORM CHILD [TWILIGHT 17]  NETTER, SUSAN  P
STORM FORCE [LAST LEGION 3]  BUNCH, CHRIS  C/P,P
STORM FRONT [@DRESDEN FILES #1]  BUTCHER, JIM  C/P,P
STORM HARVEST [DOCTOR WHO PAST DOCTORS]  PERRY, ROBERT/TUCKER, MIKE  C/P,P
STORM KNIGHTS [@TORG POSSIBILITY WARS #1]  SLAVICSEK, BILL/TRAMONTANA, C. J.  P
STORM LORD, THE [VIS 1]  LEE, TANITH  C/P,P
STORM OF DUST  RANDALL, NEIL  P,SFWA-P
STORM OF HEAVEN, THE [OATH OF EMPIRE 3]  HARLAN, THOMAS  H
STORM OF IRON [WARHAMMER 40K]  MCNEILL, GRAHAM  P
STORM OF SWORDS, A [SONG OF ICE AND FIRE 3]  MARTIN, GEORGE R. R.  H:2,P
STORM OF SWORDS, A (BRITISH) [SONG OF ICE AND FIRE 3]  MARTIN, GEORGE R. R.  H
STORM OF WINGS [@DRAGONMASTER #1]  BUNCH, CHRIS  LP
STORM OF WINGS, A [VIRICONIUM 2]  HARRISON, M. JOHN  C/P,SFWA,P
STORM OF WINGS, A (BRITISH) [VIRICONIUM 2]  HARRISON, M. JOHN  SR-P
STORM OVER SATURN [@STARHAWK 5]  MALONEY, MACK  P

TITLE  AUTHOR  CODE
STORM OVER VALLIA [DRAY PRESCOT 35]  AKERS, ALAN BURT  P
STORM OVER WARLOCK [FORERUNNER 1]  NORTON, ANDRE  C/P,P,H
STORM RISING [MAGE STORMS 2]  LACKEY, MERCEDES  H:2,SFWA,P
STORM RUNNER [WOLFWALKER 3]  HARPER, TARA K.  C/P,P
STORM SEASON [@THIEVES’ WORLD #4]  ASPRIN, ROBERT LYNN  PA,C/PA
STORM SEED [THIEVES’ WORLD TEMPUS 7]  MORRIS, JANET/MORRIS, CHRIS  P
STORM UPON ULSTER, A  FLINT, KENNETH C.  C/P:2,P
STORM WARNING [MAGE STORMS 1]  LACKEY, MERCEDES  H:2,SFWA-P,P
STORM WARRIORS [WARHAMMER ORFEO 3]  CRAIG, BRIAN  H,P
STORM WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  SAINTCROW, LILITH  LP
STORMBLADE [DRAGONLANCE HEROES 2]  BERBERICK, NANCY VARIAN  C/P:2,P
STORMBRINGER [ELRIC]  MOORCOCK, MICHAEL  SR-H,C/P:2,P
STORMBRINGER (REVISED) [ELRIC 6]  MOORCOCK, MICHAEL  C/P,P
STORMING INTREPID  HARRISON, PAYNE  C/P:2,P
STORMLIGHT [@FORGOTTEN REALMS HARPERS #14]  GREENWOOD, ED  P
STORMQUEEN! [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
STORMRIDER [RIGANTE 4]  GEMMELL, DAVID  P
STORMRIDERS [@SHADOW WORLD 4]  HAMMELL, IAN  P
STORMS OF DESTINY [EXILES OF BOQ’URAIN 1]  CRISPIN, A. C.  P
STORMS OF FATE [@BATTLETECH]  COLEMAN, LOREN L.  C/P,P
STORMS OF VICTORY [WITCH WORLD TURNING 1]  NORTON, ANDRE/GRIFFIN, P. M.  C/P,SFWA,P
STORMS OF VICTORY [@JANISSARIES 3]  POURNELLE, JERRY/GREEN, ROLAND  C/P,SFWA,P,H
STORMTRACK  SUTHERLAND, JAMES  C/P:2,P
STORMWARDEN [CYCLE OF FIRE 1]  WURTS, JANNY  C/P:2,SFWA-P,P
STORY BEHIND THE STORY, THE  MERRITT, A.  [C/RAN-H]
STORY INDEX #1 1955–1965  VEJCHAR, ALFRED  R/IND
STORY OF ADAMSVILLE, THE  RIEDMANN, AGNES  LP
STORY OF "F", THE  HUNTER, ALAN  SR-L
STORY OF THE AMULET, THE  NESBIT, E.  C/P,SR-P,H
STORY OF THE STONE, THE [MASTER LI 2]  HUGHART, BARRY  C/P,P,H
STORY OF THE VEILED LADY, THE [PHANTOM 4]  SHAWN, WILLIAM S.  P
STORYTELLER  THOMSON, AMY  LP
STORYTELLER  YOLEN, JANE  H
STORYTELLER AND THE JAHN, THE [PARSINA 2]  GOLDIN, STEPHEN  P
STORYTELLING STONE, THE  FELDMANN, SUSAN  PA
STOWAWAY TO MARS  BEYNON, JOHN  SR-P,C/P:2
STOWAWAY TO MARS  WYNDHAM, JOHN  C/P:2,P
STOWAWAY TO THE MOON  SHELTON, WILLIAM R.  H
STOWAWAY TO THE MUSHROOM PLANET  CAMERON, ELEANOR  P
STOWAWAYS [STAR TREK DS9]  STRICKLAND, BRAD  H,LP
STR8 BOLT  WHITMAN, J. T.  LP
STRACZYNSKI UNPLUGGED  STRACZYNSKI, J. MICHAEL  LP
STRAIGHT SILVER [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
STRAKEN [HIGH DRUID OF SHANNARA 3]  BROOKS, TERRY  H
STRANDED [@TIME POLICE #3]  NORWOOD, WARREN/ODOM, MEL  C/P,P
STRANDS OF STARLIGHT [STRANDS OF STARLIGHT 1]  BAUDINO, GAEL  C/P:2,P
STRANDS OF SUNLIGHT [STRANDS OF STARLIGHT 4]  BAUDINO, GAEL  C/P,P
STRANGE ADVENTURES OF RANGERGIRL, THE  PRATT, TIM  LP
STRANGE AND ANCIENT NAME, A  SHERMAN, JOSEPHA  C/P,P,H
STRANGE AND UNCANNY, THE  MACKLIN, JOHN  [C/RAN-P]
STRANGE ATTRACTORS [CHAOS CHRONICLES 2]  CARVER, JEFFREY A.  C/P,P,H
STRANGE ATTRACTORS  HENIGHAN, TOM  P
STRANGE BEASTS AND UNNATURAL MONSTERS  STERN, PHILIP VAN DOREN  PA
STRANGE BEDFELLOWS  SCORTIA, THOMAS N.  C/PA,PA
STRANGE BUT NOT A STRANGER  KELLY, JAMES PATRICK  H
STRANGE CANDY  HAMILTON, LAURELL K.  H
STRANGE CARGO [WESTERN LIGHTS 3]  BARLOUGH, JEFFREY E.  LP
STRANGE CASE OF DR. JEKYLL AND MR. HYDE AND OTHER STORIES, THE STEVENSON, ROBERT LOUIS  P
STRANGE CITY [WORLD OF DARKNESS]  KRAUSE, STALEY/WIECK, STEWART  C/PA,PA
STRANGE CONFLICT  WHEATLEY, DENNIS  SR-P,C/P:2
STRANGE DAYS  DOZOIS, GARDNER  H
STRANGE DEATH OF ADOLF HITLER, THE  ANONYMOUS  H
STRANGE DELIVERANCE  BROWN, MARY  P
STRANGE DESTINIES  MACKLIN, JOHN  [C/RAN-P]
STRANGE DEVICES OF THE SUN AND MOON  GOLDSTEIN, LISA  H
STRANGE DOINGS  LAFFERTY, R. A.  H:2,SFWA,P
STRANGE DREAMS  DONALDSON, STEPHEN R.  HA,LPA
STRANGE ECSTASIES  PARRY, MICHEL  PA
STRANGE ENCOUNTERS WITH GHOSTS  STEIGER, BRAD  [C/RAN-P]
STRANGE ENGLAND [DOCTOR WHO NEW ADVENTURES]  MESSINGHAM, SIMON  C/P:2,P
STRANGE EONS  BLOCH, ROBERT  P
STRANGE EVIL  GASKELL, JANE  P
STRANGE FORCES  LUGONES, LEOPOLDO  LP
STRANGE GATEWAYS  PRICE, E. HOFFMANN  SR-H

The MITSFS Titledex  September 24, 2007  Page 350
TITLE  AUTHOR  CODE
STRANGE GIFTS  SILVERBERG, ROBERT  HA
STRANGE GODS  ELWOOD, ROGER  PA
STRANGE HAPPENINGS  BANNISTER, PAUL  [C/RAN-H]
STRANGE HAPPENINGS  VITARELLI, ROBERT  SR-PA
STRANGE HARVEST  WANDREI, DONALD  SR-H
STRANGE HIGHWAYS  KOONTZ, DEAN R.  P
STRANGE HORIZONS  MOSKOWITZ, SAM  SFWA,R/REF-H
STRANGE INVADERS, THE  LLEWELLYN, ALUN  SR-P
STRANGE INVASION  KANDEL, MICHAEL  C/P,P
STRANGE ITINERARIES  POWERS, TIM  LP

STRANGE MAINE WAUGH, CHARLES G./

GREENBERG, MARTIN H./ MCSHERRY, FRANK D. JR.

LP

STRANGE MANUSCRIPT FOUND IN A COPPER CYLINDER, A  DE MILLE, JAMES  H
STRANGE MONSTERS OF THE RECENT PAST  WALDROP, HOWARD  C/P,P
STRANGE NEWS FROM HEAVEN  GRIFFITHS, ALAN  H
STRANGE ONES, THE  TORRO, PEL  SR-P
STRANGE ORBIT  SIMPSON, MARGARET  LP
STRANGE PORTS OF CALL  DERLETH, AUGUST  C/PA,PA
STRANGE POWERS OF HEALING  STEIGER, BRAD  [C/RAN-P]
STRANGE POWERS OF UNUSUAL PEOPLE  HOUSE, BRANT  [C/RAN-P]
STRANGE RELATIONS  FARMER, PHILIP JOSE  C/P,P,H
STRANGE SEAS AND SHORES  DAVIDSON, AVRAM  H:2,P
STRANGE SEED  WRIGHT, T. M.  P,SFWA-P
STRANGE SHADOWS  SMITH, CLARK ASHTON  SR-H
STRANGE SIGNPOSTS  ELWOOD, ROGER/MOSKOWITZ, SAM  HA
STRANGE STARS & ALIEN SHADOWS  SCHWADER, ANN K.  H
STRANGE STORY OF OUR EARTH, THE  VERRILL, A. HYATT  [C/RAN-P]
STRANGE TALENTS  HURWOOD, BERNHARDT J.  [C/RAN-P]
STRANGE TALES  ODYSSEY PUBLICATIONS  SR-LPA
STRANGE TALES FROM CBS RADIO MYSTERY THEATER  BROWN, HIMAN  PA
STRANGE TALES FROM THE NILE EMPIRE [TORG POSSIBILITY WARS]  WEST END BOOKS  PA
STRANGE TOMORROWS  HOSKINS, ROBERT  C/PA,PA
STRANGE TOYS  GEARY, PATRICIA  C/P,P
STRANGE TRADES  DI FILIPPO, PAUL  H
STRANGE TRAVELERS  WOLFE, GENE  LP
STRANGE VISITORS [@SMALLVILLE]  STERN, ROGER  P
STRANGE VOYAGE OF KIAN THE MARINER, THE  BAKST, HAROLD  P
STRANGE WAYS OF MAN, THE  PIKE, E. ROYSTON  [C/RAN-P]
STRANGE WINE  ELLISON, HARLAN  C/P,SFWA,P,H
STRANGE WORLD  EDWARDS, FRANK  [C/RAN-P:2]
STRANGE WORLD OF SCIENCE FICTION  NEW ENGLISH LIBRARY  SR-VLHA
STRANGELOVE GAMBIT, THE [@NIKOLAI DANTE]  BISHOP, DAVID  P
STRANGENESS  DISCH, THOMAS M./NAYLOR, CHARLES  C/PA,PA
STRANGER, THE  DICKSON, GORDON R.  C/P,SFWA-P,P
STRANGER, THE [JENNY SIXA 2]  FULLILOVE, ERIC JAMES  P
STRANGER AT GREEN KNOWE, A [GREEN KNOWE 4]  BOSTON, L. M.  H,LP
STRANGER AT THE CROSSROADS  HARD, ANNETTE  P
STRANGER AT THE WEDDING  HAMBLY, BARBARA  C/P:2,P
STRANGER BY NIGHT [HOT BLOOD]  GELB, JEFF/GARRETT, MICHAEL  PA
STRANGER FROM THE DEPTHS  TURNER, GERRY  C/P:2,P
STRANGER FROM THE STARS  ETCHEMENDY, NANCY  LP
STRANGER IN A STRANGE LAND  HEINLEIN, ROBERT A.  C/P,P
STRANGER IN A STRANGE LAND (UNCUT)  HEINLEIN, ROBERT A.  H:2,SFWA:2,LP
STRANGER INSIDE  KRANE, STEVEN  P
STRANGER SOULS [SHADOWRUN 26/DRAGON HEART 1]  KOKE, JAK  P
STRANGER SUNS  ZEBROWSKI, GEORGE  C/P,P
STRANGER THAN LIFE  MILLER, R. DEWITT  [C/RAN-P]
STRANGER THAN SCIENCE  EDWARDS, FRANK  [C/RAN-P]
STRANGER THAN YOU THINK  EDMONDSON, G. C.  DM109
STRANGER THINGS HAPPEN  LINK, KELLY  LP
STRANGER TO THE SUN [@ANGEL]  MARIOTTE, JEFF  P
STRANGERS  DOZOIS, GARDNER  C/P,P,H
STRANGERS  KOONTZ, DEAN R.  H
STRANGERS AT COLLINS HOUSE  ROSS, MARILYN  C/P,P
STRANGERS FROM EARTH  ANDERSON, POUL  SR-P,C/P:2
STRANGERS FROM THE SKIES  STEIGER, BRAD  [C/RAN-P]
STRANGERS FROM THE SKY [STAR TREK]  BONANNO, MARGARET WANDER  C/P:2,P
STRANGERS IN PARADISE  ANVIL, CHRISTOPHER  C/P:2,P
STRANGERS IN THE UNIVERSE  SIMAK, CLIFFORD D.  H:2,P
STRANGERS ON THE HEIGHTS  WELLMAN, MANLY WADE  H
STRANGEST OF ALL  EDWARDS, FRANK  [C/RAN-P:2]
STRANGEWOOD  GOLDEN, CHRISTOPHER  LP
STRANGLER’S MOON [FAMILY D’ALEMBERT #2]  SMITH, E. E./GOLDIN, STEPHEN  C/P:2,@P
STRATA  PRATCHETT, TERRY  H:2

The MITSFS Titledex  September 24, 2007  Page 351
TITLE  AUTHOR  CODE
STRATEGY IN POKER, BUSINESS, AND WAR  MCDONALD, JOHN  [C/RAN-P]
STRATEGY OF TECHNOLOGY, THE  POSSONY, STEFAN T./POURNELLE, JERRY  R/REF-H
STRAY  VINCENT, RACHEL  P
STRAY LAMB, THE  SMITH, THORNE  C/P,P,H
STRAYED SHEEP OF CHARUN, THE  ROBERTS, JOHN MADDOX  H,SFWA
STREAMS OF SILVER [@FORGOTTEN REALMS ICEWIND DALE #2]  SALVATORE, R. A.  C/P:2,P
STREET MAGIC  REAVES, MICHAEL  H
STREETLETHAL [AUBRY KNIGHT 1]  BARNES, STEVEN  C/P:2,SFWA-P,P
STREETS OF BLOOD [@SHADOWRUN 8]  SARGENT, CARL/GASCOIGNE, MARC  C/P,P
STRENGTH OF STONES  BEAR, GREG  C/P,P
STRESS ANALYSIS OF A STRAPLESS EVENING GOWN, A  BAKER, ROBERT A.  [C/RAN-H]
STRESS OF HER REGARD, THE  POWERS, TIM  H,SFWA,SFWA-P
STRESS PATTERN  BARRETT, NEAL JR.  C/P,P
STRICKEN FIELD, THE [HANDFUL OF MEN 3]  DUNCAN, DAVE  C/P,P,H
STRICKLAND DEMON, THE  SHACKLEFORD, JACK D.  SR-P
STRIKE EAGLES  COYNE, JAMES P.  LP
STRIKE ZONE [STAR TREK TNG]  DAVID, PETER  C/P:3,P
STRINGS  DUNCAN, DAVE  C/P,P
STRIPED HOLES  BRODERICK, DAMIEN  P
STRIPER ASSASSIN [SHADOWRUN 11]  SMITH, NYX  C/P,P
STROKE OF MIDNIGHT, A [MEREDITH GENTRY 4]  HAMILTON, LAURELL K.  H

STROKE OF MIDNIGHT KENYON, SHERRILYN/ASHLEY, AMANDA/ BANKS, L. A./HANDELAND, LORI

P

STRONG SPIRITS  DECARLO, ELISA  P
STRONGHOLD [DRAGON STAR 1]  RAWN, MELANIE  C/P,P,H
STRONGHOLD RISING [SHOLAN ALLIANCE 6]  NORMAN, LISANNE  P
STRUCTURAL FABULATION  SCHOLES, ROBERT  R/REF-P
STRUM AGAIN? [@SONGKILLER #3]  SCARBOROUGH, ELIZABETH ANN  P
STUDY IN SCARLET, A/SIGN OF FOUR, THE  DOYLE, ARTHUR CONAN  P
STUDY IN SORCERY, A [LORD DARCY 6]  KURLAND, MICHAEL  C/P:2,P

STUDY OF THE PERFORMANCE OF AN ASTRONAUT DURING INGRESS AND EGRESS MANEUVERS THROUGH AIRLOCKS AND PASSAGEWAYS, A NASA

[C/RAN-H]

STUDY WAR NO MORE  HALDEMAN, JOE  C/PA,PA
STUNTS  GRANT, CHARLES L.  SFWA
STURGEON IN ORBIT  STURGEON, THEODORE  C/P:2,P
STURGEON IS ALIVE AND WELL...  STURGEON, THEODORE  SR-H,H:2,P
SUB-ZERO  WALKER, ROBERT W.  P
SUBB  MACAPP, C. C.  C/P,P
SUBMARINE BOYS’ TRIAL TRIP  DURHAM, VICTOR  H
SUBMARINE MYSTERY, THE [DOC SAVAGE #63]  ROBESON, KENNETH  C/P:2,@P
SUBSPACE ENCOUNTER  SMITH, E. E.  C/P:2,P
SUBSPACE EXPLORERS  SMITH, E. E.  C/P,P,H
SUBSTANCE X [TALES OF TOMORROW 3]  HOUSTON, DAVID  P
SUBSTITUTE, THE  SUMNER, M. C.  P
SUBTERRANEAN GALLERY  RUSSO, RICHARD PAUL  P,SFWA-P
SUBTLE KNIFE, THE [HIS DARK MATERIALS 2]  PULLMAN, PHILIP  C/P,P,H
SUBURBAN BOOK OF THE DEAD, THE [ARMAGEDDON 3]  RANKIN, ROBERT  P
SUCCESSION [SUCCESSION 1,2]  WESTERFELD, SCOTT  H
SUCCUBUS BLUES  MEAD, RICHELLE  LP
SUCH A GOOD BABY  JENSEN, RUBY JEAN  SFWA-P
SUCH A PRETTY FACE  MARTINDALE, LEE  HA,LPA
SUCH PAIN [@WORLD OF DARKNESS MAGE 1]  BASSINGTHWAITE, DON  C/P,P
SUCH STUFF AS SCREAMS ARE MADE OF  BLOCH, ROBERT  C/P,P
SUCHE NACH DER ERDE [@PERRY RHODAN SILBERBANDE #78]  PABEL MOEWIG KG  FPR
SUCKER PUNCH [@SPEC OPS SQUAD 3]  SHELLEY, RICK  P
SUDANNA, SUDANNA  HERBERT, BRIAN  C/P:2,P,H
SUDDEN STAR, THE  SARGENT, PAMELA  C/P,P
SUDDEN WILD MAGIC, A  JONES, DIANA WYNNE  P,SFWA
SUDDENLY  KARP, MARVIN ALLEN/SETTEL, IRVING  C/PA,PA
SUFFER THE CHILDREN  SAUL, JOHN  P
SUGAR FESTIVAL, THE [STARBRIDGE CHRONICLES 1,2]  PARK, PAUL  H:2
SUGAR RAIN [STARBRIDGE CHRONICLES 2]  PARK, PAUL  C/P,P,H
SUICIDE ACADEMY  STERN, DANIEL  P
SUICIDE, INC.  GOULART, RON  C/P:2,P
SUICIDE PLAGUE, THE  NAHA, ED  C/P,P
SUICIDERS, THE  MCINTOSH, J. T.  C/P,P
SUISAN  AGINS, PHYLLIS CAROL  P,SFWA-P
SUITING, THE  WILDE, KELLEY  SFWA-P
SUITORS DUEL [QUEST TRIAD 2]  NILES, DOUGLAS  P
SULDRUN’S GARDEN [LYONESSE 1]  VANCE, JACK  C/P,LP,H
SULDRUN’S GARDEN (VIE) [@VANCE INTEGRAL EDITION #36/LYONESSE 1]  VANCE, JACK  SR-H
SULTAN OF THE MOON AND STARS [@OROKON #3]  ARDEN, TOM  P
SUM OF THINGS, THE [@STARCRUISER SHENANDOAH #3]  GREEN, ROLAND J.  P
SUM VII  HARD, T. W.  H

The MITSFS Titledex  September 24, 2007  Page 352
TITLE  AUTHOR  CODE
SUM VII (PROOF)  HARD, T. W.  SR-LP
SUMMER COUNTRY, THE [MAUREEN PIERCE 1]  HETLEY, JAMES A.  LP
SUMMER IN THREE THOUSAND  MARTIN, PETER  H
SUMMER KING, WINTER FOOL  GOLDSTEIN, LISA  P
SUMMER KNIGHT [@DRESDEN FILES #4]  BUTCHER, JIM  C/P,P
SUMMER LIGHT  GUTTENBERG, ELYSE  P
SUMMER OF LOVE  MASON, LISA  LP
SUMMER OF NIGHT  SIMMONS, DAN  C/P,P
SUMMER OF THE APOCALYPSE  VAN PELT, JAMES  LP
SUMMER QUEEN, THE [SNOW QUEEN 3]  VINGE, JOAN D.  H:2,P
SUMMER TREE, THE [FIONAVAR TAPESTRY 1]  KAY, GUY GAVRIEL  H:2,SFWA,P
SUMMER VISITORS, THE  LEIMAS, BROOKE  P
SUMMERBLOOD [ERON 3]  DEITZ, TOM  P
SUMMERFAIR [KANTMORIE 3]  DIBELL, ANSEN  P
SUMMERHILL HOUNDS  KING, J. ROBERT  P
SUMMERLAND  JAMES, L. DEAN  P
SUMMERS AT CASTLE AUBURN  SHINN, SHARON  H,LP
SUMMER’S KING, THE [RULERS OF HYLOR 3]  WILDER, CHERRY  H:2,P
SUMMERTIDE [HERITAGE UNIVERSE 1]  SHEFFIELD, CHARLES  C/P,P,H
SUMMIT, THE  MYERS, EDWARD  P
SUMMON THE KEEPER [KEEPER’S CHRONICLES 1]  HUFF, TANYA  C/P:2,P
SUMMON THE THUNDER [STAR TREK VANGUARD]  WARD, DAYTON/DILMORE, KEVIN  P
SUMMONED, THE [@ANGEL]  DOKEY, CAMERON  P
SUMMONED TO TOURNEY [ERIC BANYON 2]  LACKEY, MERCEDES/GUON, ELLEN  C/P:2,SFWA-P,P
SUMMONER, THE [@NECROMANCER #1]  MARTIN, GAIL Z.  P

SUMMONING, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #1] DENNING, TROY

P

SUMMONING LIGHT [@BABYLON 5 PASSING OF THE TECHNO-MAGES #2]  CAVELOS, JEANNE  C/P,P
SUMURU  ROHMER, SAX  P
SUN AND ITS FAMILY, THE  ADLER, IRVING  [C/RAN-P]
SUN-CROSS [SUN-CROSS 1,2]  HAMBLY, BARBARA  H:2
SUN DESTROYERS, THE  ROCKLYNNE, ROSS  D93900
SUN DOGS  MCGARRY, MARK J.  C/P,P
SUN GROWS COLD, THE  BERK, HOWARD  H,P
SUN IN ECLIPSE, THE  VERNE, JULES  H
SUN IN GLORY [VALDEMAR]  LACKEY, MERCEDES  PA
SUN LORD [@OUTLANDERS]  AXLER, JAMES  P
SUN-MAKERS, THE  STATTEN, VARGO  SR-P
SUN OF SUNS [VIRGA 1]  SCHROEDER, KARL  H
SUN QUEEN, THE  KANER, H.  H:2
SUN-RUNNER’S FIRE [DRAGON PRINCE 3]  RAWN, MELANIE  C/P:2,P
SUN SABOTEURS, THE  KNIGHT, DAMON  DF108
SUN SHINES BRIGHT, THE  ASIMOV, ISAAC  [C/RAN-P]
SUN SMASHER, THE  HAMILTON, EDMOND  DD351,P
SUN SWORD, THE [@SUN SWORD #6]  WEST, MICHELLE  C/P,P
SUN, THE MOON, AND THE STARS, THE  BRUST, STEVEN  H:2,SFWA,P
SUNBIRD, THE  SMITH, WILBUR  P
SUNBORN, THE  BENFORD, GREGORY  H
SUNBOUND, THE  FELICE, CYNTHIA  C/P,P
SUNBURST  GOTLIEB, PHYLLIS  C/P:2,P
SUNDANCE AND OTHER SCIENCE FICTION STORIES  SILVERBERG, ROBERT  H
SUNDER, ECLIPSE & SEED  GUTTENBERG, ELYSE  C/P,P
SUNDERED, THE [@STAR TREK LOST ERA (2298 AD)]  MARTIN, MICHAEL A./MANGELS, ANDY  P
SUNDERED ARMS, THE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
SUNDERED REALM, THE [WAR OF POWERS 1]  VARDEMAN, ROBERT E./MILAN, VICTOR  C/P:2,P
SUNDERED WORLDS, THE  MOORCOCK, MICHAEL  C/P:2,P
SUNDERING, THE [@WARCRAFT WAR OF THE ANCIENTS #3]  KNAAK, RICHARD A.  P
SUNDERING, THE [@DREAD EMPIRE’S FALL 2]  WILLIAMS, WALTER JON  P
SUNDERING FLOOD, THE  MORRIS, WILLIAM  C/P,P
SUNDERLIES SEEKING [@GHATTENS’ GAMBIT #1]  GREENO, GAYLE  P
SUNDIAL, THE  JACKSON, SHIRLEY  P
SUNDIPPER  THOMPSON, PAUL B.  P
SUNDIVER [UPLIFT 1]  BRIN, DAVID  C/P:2,P
SUNDOG  BALL, BRIAN N.  C/P:2,P
SUNDOWNER [NICOLE SHEA 3]  CLAREMONT, CHRIS  P
SUNDRINKER  HUGHES, ZACH  P
SUNFALL  CHERRYH, C. J.  C/P,P,H
SUNG IN BLOOD  COOK, GLEN  SR-H,H
SUNG IN SHADOW  LEE, TANITH  C/P:2,P
SUNGLASSES AFTER DARK [SONJA BLUE 1]  COLLINS, NANCY A.  C/P,P
SUNKEN CITY, THE  MCNEILL, JAMES  LP
SUNKEN GARDEN, THE  CRANE, NATHALIA  H
SUNKEN WORLD, THE  COBLENTZ, STANTON A.  SR-LP
SUNLESS WORLD, THE [PROFESSOR JAMESON 2]  JONES, NEIL R.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 353
TITLE  AUTHOR  CODE
SUNRISE ALLEY [CHARON 1]  ASARO, CATHERINE  H
SUNRISE OF THE MEDITERRANEAN  FRANK, SUZANNE  H
SUNRISE ON MERCURY  SILVERBERG, ROBERT  H
SUNRISE WEST  CARLSON, WILLIAM K.  H
SUN’S END  LUPOFF, RICHARD A.  C/P:2,P
SUNS IN DUO  LA SALLE, VICTOR  SR-P
SUNS OF CARESH, THE [DOCTOR WHO PAST DOCTORS]  SAINT, PAUL  P
SUNS OF SCORPIO, THE [DRAY PRESCOT 2]  AKERS, ALAN BURT  C/P,P
SUNSET [@WARRIORS NEW PROPHECY #6]  HUNTER, ERIN  H
SUNSET WARRIOR, THE [SUNSET WARRIOR 1]  LUSTBADER, ERIC VAN  C/P:2,SFWA,P
SUNSET WITH A BEARD  MELLICK, CARLTON III  LP
SUNSHAKER’S WAR [OTHERWORLD 4]  DEITZ, TOM  C/P:2,P
SUNSHINE  MCKINLEY, ROBIN  H
SUNSHINE (PROOF)  MCKINLEY, ROBIN  SR-LP
SUNSHINE PATRIOTS  CAMPBELL, BILL  LP
SUNSMOKE  KILLUS, JAMES  C/P,P
SUNSPACER [SUNSPACE 1]  ZEBROWSKI, GEORGE  LP
SUNSTOP 8  FISHER, LOU  P
SUNSTORM [TIME ODYSSEY 2]  CLARKE, ARTHUR C./BAXTER, STEPHEN  H
SUNSTRIKE  CARPOZI, GEORGE JR.  C/P,P
SUNSTROKE  KAGAN, DAVID  C/P,P
SUNWAIFS  VAN SCYOC, SYDNEY J.  H,P
SUNWATCH  DORN, FRANK  P
SUNWORLD  KELLEY, LEO P.  SR-P
SUPER BARBARIANS, THE  BRUNNER, JOHN  C/P:2,P
SUPER-CANNES  BALLARD, J. G.  H
SUPER HUGOS, THE  ASIMOV, ISAAC  PA
SUPER SCIENCE  WHITE, MICHAEL  [C/RAN-P:2]
SUPER SCIENCE STORIES  ELAM, RICHARD M. JR.  LP
SUPER SEX CAPTAIN  PECK, VERNON  SR-L
SUPER-STATE  ALDISS, BRIAN W.  H
SUPERBABY  MENDELSOHN, FELIX JR.  C/P,P
SUPERGIRL  MAZER, NORMA FOX  C/P,P
SUPERHEROES  PARRY, MICHEL  SR-PA
SUPERHEROES  VARLEY, JOHN/MAINHARDT, RICIA  HA,LPA
SUPERHORROR  CAMPBELL, RAMSEY  SR-HA
SUPERIOR BEINGS [DOCTOR WHO PAST DOCTORS]  WALTERS, NICK  P
SUPERJET  ZAREM, LEWIS/CANTWELL, RAY  H
SUPERLUMINAL [METAPLANETARY 2]  DANIEL, TONY  H
SUPERLUMINAL  MCINTYRE, VONDA N.  C/P:2,SFWA,P,H
SUPERMAN: DOOMSDAY & BEYOND [DC COMICS]  SIMONSON, LOUISE  LP
SUPERMAN III [DC COMICS]  KOTZWINKLE, WILLIAM  C/P,P
SUPERMAN: LAST SON OF KRYPTON [DC COMICS]  MAGGIN, ELLIOT S.  C/P:2,P
SUPERMAN: MIRACLE MONDAY [DC COMICS]  MAGGIN, ELLIOT S.  P
SUPERMAN ON RADIO (4 CASSETTES, BOXED)  RADIO SPIRITS, INC.  SR-MM
SUPERMAN: THE MAN OF STEEL [DC COMICS]  HELFER, ANDREW  P

SUPERMEN ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

SUPERMEN  DOZOIS, GARDNER  LPA
SUPERMIND  PHILLIPS, MARK  C/P:2,P
SUPERMIND  VAN VOGT, A. E.  C/P:2,P
SUPERMONSTERS  COHEN, DANIEL  R/REF-P
SUPERNATURAL, THE  HILL, DOUGLAS/WILLIAMS, PAT  [C/RAN-H]
SUPERNATURAL  HIRSCH, PHIL  [C/RAN-P]
SUPERNATURAL CATS  NECKER, CLAIRE  PA,SFWA
SUPERNATURAL OMNIBUS: BEST GHOST STORIES, THE  SUMMERS, MONTAGUE  HA
SUPERNATURAL READER, THE  CONKLIN, GROFF  HA,PA
SUPERNATURAL STORIES #36  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #41  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #43  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #47  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #49  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #51  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #53  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #55  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #57  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #59  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #61  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #63  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #65  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #67  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #69  BADGER BOOKS  SR-PA,C/PA
SUPERNATURAL STORIES #71  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #73  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #75  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #77  BADGER BOOKS  SR-PA

The MITSFS Titledex  September 24, 2007  Page 354
TITLE  AUTHOR  CODE
SUPERNATURAL STORIES #79  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #81  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #83  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #85  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #87  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #89  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #91  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #93  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #95  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #97  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #99  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #103  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #105  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #107  BADGER BOOKS  SR-PA
SUPERNATURAL STORIES #109  BADGER BOOKS  SR-PA
SUPERNOVA  ALLEN, ROGER MACBRIDE/KOTANI, ERIC  C/P,P
SUPERNOVA [@PERRY RHODAN #66]  MAHR, KURT  FPR

SUPERTANKS HALDEMAN, JOE/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA,C/PA,SFWA-P

SUPERTOYS LAST ALL SUMMER LONG  ALDISS, BRIAN W.  LP
SUPPORT YOUR LOCAL WIZARD [YOUNG WIZARDS 1,2,3]  DUANE, DIANE  H
SURAK’S SOUL [STAR TREK ENTERPRISE]  DILLARD, J. M.  P
SURE DEATH OF A MOUSE, THE [NIMNESTL 2]  CRAWFORD, DAN  C/P,P
SURELY YOU’RE JOKING, MR. FEYNMAN  FEYNMAN, RICHARD P.  [C/RAN-H]
SURFACE ACTION  DRAKE, DAVID  C/P:2,SFWA-P:2,P
SURFING SAMURAI ROBOTS  GILDEN, MEL  C/P:2,SFWA-P:2,P
SURPRISE GUESTS, THE [LAND OF THE LOST]  GRAHAM, KENNON  H
SURRENDER NONE: THE LEGACY OF GIRD [GIRD 1]  MOON, ELIZABETH  C/P:2,SFWA-P,P
SURRENDER THE GALACTICA! [@BATTLESTAR GALACTICA 1978 #14]  LARSON, GLEN A./THURSTON, ROBERT  P
SURRENDER YOUR DREAMS [@MECHWARRIOR DARK AGE/BATTLETECH]  PARDOE, BLAINE LEE  P
SURROGATE, THE  SHARMAN, NICK  P
SURVEILLANCE, THE [@INTERVENTION #1]  MAY, JULIAN  C/P:2,P
SURVEY SHIP  BRADLEY, MARION ZIMMER  C/P:2,LP
SURVIVAL [@SPACE: 1999]  BALL, BRIAN N.  P
SURVIVAL [@SPECIES IMPERATIVE #1]  CZERNEDA, JULIE E.  H
SURVIVAL [@STAR TREK TNG STARFLEET ACADEMY #3]  DAVID, PETER  H,LP
SURVIVAL!  DICKSON, GORDON R.  C/P:2,P
SURVIVAL [DOCTOR WHO]  MUNRO, RONA  SR-P
SURVIVAL  SHERWOOD, MARTIN  P
SURVIVAL FROM INFINITY  ELWOOD, ROGER  SR-HA
SURVIVAL GAME, THE  KAPP, COLIN  C/P,P
SURVIVAL IN SPACE  LEBEDEV, VLADIMIR/GAGARIN, YURI  [C/RAN-P]
SURVIVAL INSTINCT [@NECROMUNDA]  CHAMBERS, ANDY  P
SURVIVAL MARGIN  MAINE, CHARLES ERIC  C/P:2,P
SURVIVAL OF FREEDOM, THE  POURNELLE, JERRY/CARR, TERRY  C/PA,PA
SURVIVAL OF THE FITTEST [@STARGATE SG-1 #7]  BAUER, SABINE C.  P
SURVIVAL OF THE FITTEST [@CHARMED]  MARIOTTE, JEFF  P

SURVIVAL PRINTOUT ALLISON, LEONARD/JENKIN, LEONARD/

PERRAULT, ROBERT

C/PA,PA

SURVIVAL PROJECT  MULLER, JOHN E.  SR-P
SURVIVAL SHIP AND OTHER STORIES  MERRIL, JUDITH  SR-P,C/P
SURVIVAL WORLD  LONG, FRANK BELKNAP  C/P,P
SURVIVANTS DE KOR, LES  RANDA, PETER  FP
SURVIVING ANTARCTICA: REALITY TV 2083  WHITE, ANDREA  H
SURVIVOR  BUTLER, OCTAVIA E.  C/P,SFWA,P,H
SURVIVOR, THE  HERBERT, JAMES  P
SURVIVOR [GERALD KNAVE 1]  JANIFER, LAURENCE M.  C/P:2,P
SURVIVOR, THE  PARRY, DENNIS  H
SURVIVOR AND OTHERS, THE  LOVECRAFT, H. P./DERLETH, AUGUST  C/P:2,P
SURVIVOR IN DEATH [EVE DALLAS]  ROBB, J. D.  H

SURVIVORS, THE BRADLEY, MARION ZIMMER/ ZIMMER, PAUL EDWIN

C/P:2,P

SURVIVORS, THE  GODWIN, TOM  H
SURVIVORS [STAR TREK TNG]  LORRAH, JEAN  C/P:2,P
SURVIVORS, THE [DINOSAUR PLANET 2]  MCCAFFREY, ANNE  P
SURVIVORS [@FIREBRATS #2]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
SURVIVOR’S QUEST [STAR WARS]  ZAHN, TIMOTHY  H
SUSANNAH, THE RIGHTEOUS [@PHENWICK #6]  KIMBROUGH, KATHERYN  P
SUSPENSION  FANE, BRON  SR-P
SUSPICION [ROBOT CITY #2]  MCQUAY, MIKE  C/P:2,P
SVAHA  DE LINT, CHARLES  C/P:2,SFWA-P,P
SWAMP THING  HOUSTON, DAVID/WEIN, LEN  P
SWAN SISTER  DATLOW, ELLEN/WINDLING, TERRI  LPA
SWAN SONG  MCCAMMON, ROBERT R.  C/P,P
SWAN SONG [GRAINGER 6]  STABLEFORD, BRIAN M.  P
SWAN SONGS [GRAINGER 1,2,3,4,5,6]  STABLEFORD, BRIAN M.  LP

The MITSFS Titledex  September 24, 2007  Page 355
TITLE  AUTHOR  CODE
SWAN’S WING  SYNGE, URSULA  C/P:2,P
SWARM, THE  HERZOG, ARTHUR  P
SWASHBUCKLING EDITOR STORIES  BETANCOURT, JOHN GREGORY  LP
SWATTING AT THE COSMOS  MORROW, JAMES  SR-H
SWEENEY’S ISLAND  CHRISTOPHER, JOHN  P
SWEET DREAMS, SWEET PRINCES  REYNOLDS, MACK/BANKS, MICHAEL  C/P:2,P
SWEET HEART  JAMES, PETER  SFWA
SWEET MYTH-TERY OF LIFE [MYTH ADVENTURES]  ASPRIN, ROBERT LYNN  C/P:2,P,H
SWEET RIDE, THE  PRATHER, RICHARD S.  SFWA-P
SWEET SILVER BLUES [GARRETT 1]  COOK, GLEN  C/P:2,P
SWEET SIXTEEN [@BUFFY]  CIENCIN, SCOTT  P
SWEET STARFIRE  KRENTZ, JAYNE ANN  P
SWEET, SWEET SUMMER, A  GASKELL, JANE  H
SWEET TALKIN’ DEMON [@CHARMED]  BURNS, LAURA J.  P
SWEETWATER  YEP, LAURENCE  H,LP
SWELL FOOP [XANTH #25]  ANTHONY, PIERS  H,P
SWIFTLY TILTING PLANET, A [MEG MURRY 3]  L’ENGLE, MADELEINE  C/P,P,H
SWIM THE MOON  BRANDON, PAUL  LP
SWIMMERS BENEATH THE BRIGHT  EASTON, M. COLEMAN  C/P,P
SWITCH BITCH  DAHL, ROALD  P
SWITCH ON THE NIGHT  BRADBURY, RAY  H
SWITCHING WELL  GRIFFIN, PENI R.  SFWA
SWORD, THE [@SWORD,RING AND CHALICE 1]  CHESTER, DEBORAH  P
SWORD, THE [@GENERAL #5]  STIRLING, S. M./DRAKE, DAVID  C/P:2,P
SWORD ABOVE THE NIGHT, A  LYMINGTON, JOHN  P
SWORD AND FIRE [@BATTLETECH #37/TWILIGHT OF THE CLANS 5]  GRESSMAN, THOMAS S.  C/P,P
SWORD AND SHADOW [SWORD IN EXILE 3]  MARSTON, ANN  P
SWORD AND SORCERESS  BRADLEY, MARION ZIMMER  PA,C/PA
SWORD AND SORCERESS II  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS III  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS IV  BRADLEY, MARION ZIMMER  PA,C/PA
SWORD AND SORCERESS V  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS VI  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS VII  BRADLEY, MARION ZIMMER  PA,C/PA
SWORD AND SORCERESS VIII  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS IX  BRADLEY, MARION ZIMMER  C/PA,PA,SFWA-P
SWORD AND SORCERESS X  BRADLEY, MARION ZIMMER  PA
SWORD AND SORCERESS XI  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS XII  BRADLEY, MARION ZIMMER  PA,C/PA
SWORD AND SORCERESS XIII  BRADLEY, MARION ZIMMER  PA,C/PA

SWORD AND SORCERESS XIV BRADLEY, MARION ZIMMER/ HOLMAN, RACHEL E.

PA,C/PA

SWORD AND SORCERESS XV  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS XVI  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS XVII  BRADLEY, MARION ZIMMER  C/PA,PA
SWORD AND SORCERESS XVIII  BRADLEY, MARION ZIMMER  PA
SWORD AND SORCERESS XIX  BRADLEY, MARION ZIMMER  PA
SWORD AND SORCERESS XX  BRADLEY, MARION ZIMMER  PA
SWORD AND SORCERESS XXI  PAXSON, DIANA L.  PA
SWORD AND THE CHAIN, THE [@GUARDIANS OF THE FLAME #2]  ROSENBERG, JOEL  C/P:2,P
SWORD AND THE DAGGER, THE [@BATTLETECH]  MAYHAR, ARDATH  P
SWORD AND THE EYE, THE [HOUSE OF EIGIN 1]  LEIBER, JUSTIN  C/P,SFWA-P,P
SWORD AND THE FLAME, THE [@DRAGON KING #3]  LAWHEAD, STEPHEN R.  C/P:2,SFWA-P,P
SWORD AND THE LION, THE  CRAY, ROBERTA  P,SFWA-P:2
SWORD AND THE SATCHEL, THE [ALFAR 1]  BOYER, ELIZABETH H.  C/P:2,P
SWORD AND THE SORCERER, THE  WINSKI, NORMAN  C/P,P
SWORD AND THE STALLION, THE [CORUM 6]  MOORCOCK, MICHAEL  C/P,P,H
SWORD AND THE TOWER, THE [HOUSE OF EIGIN 2]  LEIBER, JUSTIN  P,SFWA-P
SWORD AT SUNSET  SUTCLIFF, ROSEMARY  SFWA-P
SWORD-BORN [TIGER 5]  ROBERSON, JENNIFER  P
SWORD-BREAKER [TIGER 4]  ROBERSON, JENNIFER  C/P:2,SFWA-P:2,P
SWORD-DANCER [TIGER 1]  ROBERSON, JENNIFER  C/P:2,P
SWORD FOR A DRAGON, A [BAZIL BROKETAIL 2]  ROWLEY, CHRISTOPHER  C/P:2,SFWA-P,P
SWORD FOR KREGEN, A [DRAY PRESCOT 20]  AKERS, ALAN BURT  P
SWORD FOR THE EMPIRE [DIRSHAN 3]  LANCOUR, GENE  H,SFWA
SWORD IN THE STORM [RIGANTE 1]  GEMMELL, DAVID  P
SWORD IS DRAWN, THE  NORTON, ANDRE  H
SWORD IS FORGED, THE  WALTON, EVANGELINE  P
SWORD-MAKER [TIGER 3]  ROBERSON, JENNIFER  C/P:2,P
SWORD OF ALDONES, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,DF153,P
SWORD OF ALLAH, THE  ELLIOTT, RICHARD  P
SWORD OF ANGELS, THE [EYES OF GOD 3]  MARCO, JOHN  H
SWORD OF BEDWYR, THE [CRIMSON SHADOW 1]  SALVATORE, R. A.  P
SWORD OF BHELEU, THE [LORDS OF DUS 3]  WATT-EVANS, LAWRENCE  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 356
TITLE  AUTHOR  CODE
SWORD OF CALANDRA, THE [WINTER KING’S WAR 2]  DEXTER, SUSAN  C/P,P
SWORD OF CHAOS [DARKOVER]  BRADLEY, MARION ZIMMER  PA,C/PA
SWORD OF CULANN, THE  LEVIN, BETTY  H
SWORD OF DARKNESS [LORDS OF AVALON 2]  MACGREGOR, KINLEY  H
SWORD OF FIRE [HARRY BORG 2]  HAWKINS, WARD  P

SWORD OF FIRE AND SHADOW [FIONN MAC CUMHAL 3] PAXSON, DIANA L./

MARTINE-BARNES, ADRIENNE

P

SWORD OF FLAME [AURIAN 3]  FUREY, MAGGIE  C/P:2,P
SWORD OF FORBEARANCE, THE [PELBAR 7]  WILLIAMS, PAUL O.  P

SWORD OF FOREVER, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] MORTIMORE, JIM

C/P,P

SWORD OF GENGHIS KHAN, THE [MARK HOOD 7]  DARK, JAMES  P
SWORD OF ICE [VALDEMAR]  LACKEY, MERCEDES  C/PA,PA
SWORD OF KING JAMES [DYLAN MATHESON 3]  LEE, J. ARDIAN  LP

SWORD OF KNOWLEDGE, THE [SWORD OF KNOWLEDGE 1,2,3] CHERRYH, C. J./LACKEY, MERCEDES/

ASIRE, NANCY/FISH, LESLIE

LP

SWORD OF LANKOR, THE  CORY, HOWARD L.  C/P:2,P
SWORD OF MAIDEN’S TEARS, THE [TWELVE TREASURES 1]  EDGHILL, ROSEMARY  C/P,P
SWORD OF MARY, THE [BECCA 2]  FRIESNER, ESTHER M.  P
SWORD OF MORNING STAR, THE  MEADE, RICHARD  C/P:2,P
SWORD OF ORION [@BENEATH STRANGE SKIES #1]  LEE, SHARON/MILLER, STEVE  LP
SWORD OF ORLEY, THE  FARRAR, STEWART  H
SWORD OF POYANA [THORGRIM 2]  BAILEY, GERALD EARL  P
SWORD OF RHIANNON, THE  BRACKETT, LEIGH  C/P:2,SR-DD036,P
SWORD OF SAGAMORE, THE [SAGAMORE 2]  DALKEY, KARA  C/P,P
SWORD OF SAMURAI CAT, THE [SAMURAI CAT 4]  ROGERS, MARK E.  LP
SWORD OF SEDITION [@MECHWARRIOR DARK AGE/BATTLETECH]  COLEMAN, LOREN L.  P
SWORD OF SHANDAR, THE  BESAW, VICTOR  C/P,P
SWORD OF SHANNARA, THE [SHANNARA 1]  BROOKS, TERRY  C/P,LP,H,SFWA:2
SWORD OF SKELOS, THE [BANTAM CONAN 3]  OFFUTT, ANDREW J.  C/P,P
SWORD OF STRAW, THE [SANGREAL 2]  HEMINGWAY, AMANDA  LP
SWORD OF THE DAWN [RUNESTAFF 3]  MOORCOCK, MICHAEL  C/P:2,P
SWORD OF THE DECEIVER [ISAVALTA 4]  ZETTEL, SARAH  H
SWORD OF THE DEMON  LUPOFF, RICHARD A.  C/P,SFWA,P,H
SWORD OF THE GAEL [CORMAC 4]  OFFUTT, ANDREW J.  C/P:2,P
SWORD OF THE GOLEM, THE  ROTHBERG, ABRAHAM  C/P,P,H
SWORD OF THE LAMB [PHOENIX LEGACY 1]  WREN, M. K.  C/P:2,P
SWORD OF THE LAND, THE [RILSIN 1]  BRENNAN, NOEL-ANNE  P
SWORD OF THE LICTOR, THE [NEW SUN 3]  WOLFE, GENE  C/P,SFWA-P,P,H
SWORD OF THE NURLINGAS [THORGRIM 1]  BAILEY, GERALD EARL  P
SWORD OF THE PROPHETS  GARLAND, MARK A.  P
SWORD OF THE SAMURAI  REAVES, MICHAEL/PERRY, STEVE  P
SWORD OF THE SPIRITS, THE [LUKE 3]  CHRISTOPHER, JOHN  C/P,P,H
SWORD OF THE WHITE ROSE  LEE, J. ARDIAN  LP
SWORD OF WINTER, THE  RANDALL, MARTA  H:2,P
SWORD OF ZAGAN, THE  SMITH, CLARK ASHTON  LP
SWORD PLAY [@FORGOTTEN REALMS ARCANE AGE NETHERIL #1]  EMERY, CLAYTON  P
SWORD-SINGER [TIGER 2]  ROBERSON, JENNIFER  C/P:2,P
SWORD SMITH, THE  ARNASON, ELEANOR  P
SWORD SWALLOWER, THE  GOULART, RON  C/P,P,H
SWORD-SWORN [TIGER 6]  ROBERSON, JENNIFER  H,P
SWORD, THE JEWEL AND THE MIRROR, THE [CINGULUM 3]  ROBERTS, JOHN MADDOX  P,SFWA-P
SWORD WOMAN, THE  HOWARD, ROBERT E.  C/P,P
SWORDBEARER, THE  COOK, GLEN  C/P:2,SFWA-P,P
SWORDFISH ISLAND COVENANT, THE  GRAVES, STEPHAN  LP
SWORDHUNT [STAR TREK RIHANNSU 3]  DUANE, DIANE  C/P:2,P
SWORDS AGAINST DARKNESS I  OFFUTT, ANDREW J.  C/PA,PA
SWORDS AGAINST DARKNESS II  OFFUTT, ANDREW J.  C/PA,PA
SWORDS AGAINST DARKNESS III  OFFUTT, ANDREW J.  C/PA,PA
SWORDS AGAINST DARKNESS IV  OFFUTT, ANDREW J.  C/PA,PA
SWORDS AGAINST DARKNESS V  OFFUTT, ANDREW J.  PA
SWORDS AGAINST DEATH [FAFHRD 2]  LEIBER, FRITZ  C/P:2,P
SWORDS AGAINST THE SHADOWLAND [FAFHRD]  BAILEY, ROBIN WAYNE  H,LP
SWORDS AGAINST TOMORROW  HOSKINS, ROBERT  PA,C/PA
SWORDS AGAINST WIZARDRY [FAFHRD 4]  LEIBER, FRITZ  C/P:2,P
SWORDS AND DEVILTRY [FAFHRD 1]  LEIBER, FRITZ  C/P:2,P
SWORDS AND ICE MAGIC [FAFHRD 6]  LEIBER, FRITZ  C/P:2,P
SWORDS AND SORCERY  DE CAMP, L. SPRAGUE  PA,C/PA
SWORDS IN THE MIST [FAFHRD 3]  LEIBER, FRITZ  C/P:2,P
SWORDS’ MASTERS [FAFHRD 4,5,6]  LEIBER, FRITZ  H:2

SWORDS OF EVENINGSTAR [@FORGOTTEN REALMS KNIGHTS OF MYTH DRANNOR #1] GREENWOOD, ED

H

SWORDS OF HAVEN [@HAWK & FISHER 1,2,3]  GREEN, SIMON R.  P
SWORDS OF LANKHMAR, THE [FAFHRD 5]  LEIBER, FRITZ  SR-P,C/P:2

The MITSFS Titledex  September 24, 2007  Page 357
TITLE  AUTHOR  CODE
SWORDS OF MARS [BARSOOM #8]  BURROUGHS, EDGAR RICE  C/P:2,@P
SWORDS OF MARS [BARSOOM 8]/SYNTHETIC MEN OF MARS [BARSOOM 9] BURROUGHS, EDGAR RICE  H:2
SWORDS OF NIGHT AND DAY, THE [SKILGANNON THE DAMNED]  GEMMELL, DAVID  H
SWORDS OF SHAHRAZAR  HOWARD, ROBERT E.  C/P,P
SWORDS OF THE BARBARIANS  BULMER, KENNETH  C/P,P
SWORDS OF THE EMPIRE [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
SWORDS OF THE HORSECLANS [HORSECLANS 2]  ADAMS, ROBERT  C/P:2,P
SWORDS OF THE LEGION [VIDESSOS 4]  TURTLEDOVE, HARRY  C/P:2,P
SWORDS OF THE RAINBOW  GARBER, ERIC/GOMEZ, JEWELLE  LPA

SWORDS OF ZINJABAN, THE [KRISHNA] DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

P

SWORDS TRILOGY, THE [SWORDS 1,2,3]  MOORCOCK, MICHAEL  C/P,P,H
SWORDSHEATH SCROLL, THE [@DRAGONLANCE DWARVEN NATIONS #3]  PARKINSON, DAN  P
SWORDSHIPS OF SCORPIO [DRAY PRESCOT 4]  AKERS, ALAN BURT  C/P,P
SWORDSMAN OF MARS, THE  KLINE, OTIS ADELBERT  C/P:2,P
SWORDSMAN’S OATH, THE [EINARINN 2]  MCKENNA, JULIET E.  H,P
SWORDSMEN AND SUPERMEN  CENTAUR PRESS  PA
SWORDSMEN IN THE SKY  WOLLHEIM, DONALD A.  PA,C/PA
SWORDSMEN OF VISTAR  NUETZEL, CHARLES  SR-P
SWORDSMISTRESS OF CHAOS [@RAVEN #1]  KIRK, RICHARD  SR-P
SWORDSPOINT [SWORDSPOINT 1]  KUSHNER, ELLEN  C/P,SFWA-P,P,H
SWORDSWOMAN, THE  SALMONSON, JESSICA AMANDA  C/P,SFWA-P,H:2,P
SWORN ALLIES [@FLEET #4]  DRAKE, DAVID/FAWCETT, BILL  PA,SFWA-P:2
SYKAOS PAPERS, THE  THOMPSON, E. P.  H
SYLVA  VERCORS  FP
SYMBIOSE  CLEMENT, HAL  FH
SYMBIOTE’S CROWN  BAKER, SCOTT  P
SYMBOL OF TERRA [DUMAREST 30]  TUBB, E. C.  C/P,P
SYMPATHY FOR THE DEVIL [SYMPATHY FOR THE DEVIL 1]  LISLE, HOLLY  C/P,P
SYMPHONY [MILLENNIUM QUARTET 1]  GRANT, CHARLES L.  P
SYMPHONY OF LIGHT [@ROBOTECH #12]  MCKINNEY, JACK  P
SYMPHONY OF STORMS, A [@DEMON CROWN #3]  VARDEMAN, ROBERT E.  C/P:2,SFWA-P,P
SYMPHONY OF TERROR [V]  SUCHARITKUL, SOMTOW  SFWA-P
SYN  JONES, RAYMOND F.  C/P:2,P
SYNDIC, THE  KORNBLUTH, C. M.  H:2,P
SYNERGY 1  ZEBROWSKI, GEORGE  PA
SYNERGY 2  ZEBROWSKI, GEORGE  PA
SYNERGY 3  ZEBROWSKI, GEORGE  PA
SYNERGY 4  ZEBROWSKI, GEORGE  PA
SYNNERS  CADIGAN, PAT  C/P:2,P
SYNTHAJOY  COMPTON, D. G.  C/P:2,P
SYNTHESIS & OTHER VIRTUAL REALITIES  ROSENBLUM, MARY  H
SYNTHESPIANS[TM] [@DOCTOR WHO PAST DOCTORS]  HINTON, CRAIG  P
SYNTHETIC MAN, THE  STURGEON, THEODORE  C/P:2,P
SYNTHETIC MEN OF MARS [BARSOOM #9]  BURROUGHS, EDGAR RICE  C/P:2,@P
SYSTEM 2000 — 3. PLANET  BERNING, FRANK W.  FP
SYSTEM OF THE WORLD, THE [BAROQUE CYCLE 3]  STEPHENSON, NEAL  H
SYSTEM SHOCK [DOCTOR WHO MISSING ADVENTURES]  RICHARDS, JUSTIN  C/P:2,P
SYSTEMIC SHOCK [SYSTEMIC SHOCK 1]  ING, DEAN  C/P:2,P
SYSTEMS  QUICK, W. T.  C/P:2,P
SYZYGY  CONEY, MICHAEL G.  C/P:2,P
SYZYGY  POHL, FREDERIK  C/P:2,P

TITLE  AUTHOR  CODE
T ZERO  CALVINO, ITALO  C/P,P,H
T2: INFILTRATOR [TERMINATOR T 2 1]  STIRLING, S. M.  H
T2: RISING STORM [TERMINATOR T 2 2]  STIRLING, S. M.  P
T2: THE FUTURE WAR [TERMINATOR T 2 3]  STIRLING, S. M.  P
TA  RUSSELL, JOHN ROBERT  C/P,P
TABOO, MAGIC, SPIRITS  BURRISS, ELI EDWARD  [C/RAN-H]
TACHYON WEB, THE  PIKE, CHRISTOPHER  P
TACTICAL ERROR [STARWOLVES 3]  GUNNARSSON, THORARINN  P
TACTICS OF CONQUEST  MALZBERG, BARRY N.  C/P:2,P
TACTICS OF DUTY [@BATTLETECH 19]  KEITH, WILLIAM H. JR.  P
TACTICS OF MISTAKE [CHILDE CYCLE]  DICKSON, GORDON R.  C/P,P,H,SFWA
TAFLAK LYSANDRA [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  P
TAGGERUNG [REDWALL]  JACQUES, BRIAN  P
TAHARA BOY MYSTIC OF INDIA  SHERMAN, HAROLD M.  H
TAHARA IN THE LAND OF YUCATAN  SHERMAN, HAROLD M.  H
TAIL OF THE ARABIAN, KNIGHT, THE [LINCOLN BLACKTHORNE 2]  MARSH, GEOFFREY  P,SFWA,SFWA-P
TAILCHASER’S SONG  WILLIAMS, TAD  C/P,P,H
TAILS YOU LOSE [@SHADOWRUN 39]  SMEDMAN, LISA  P
TAIN  FROST, GREGORY  C/P,P
TAINT, THE [DOCTOR WHO EIGHTH DOCTOR]  COLLIER, MICHAEL  C/P,P
TAINT, THE  WALLACE, PATRICIA  P
TAINT OF EVIL [@WARHAMMER STEFAN KUMANSKY 2]  MCINTOSH, NEIL  P
TAINT OF LOVECRAFT, THE  SARGENT, STANLEY C.  LP
TAINTED, THE [@ISLES OF GLORY #3]  LARKE, GLENDA  P
TAINTED SWORD, THE [@DUNGEONS & DRAGONS PENHALIGON #1]  HEINRICH, D. J.  P
TAINTED TRAIL [UKIAH OREGON 2]  SPENCER, WEN  C/P,P
TAJI’S SYNDROME  YARBRO, CHELSEA QUINN  C/P:2,P
TAKE A THIEF [VALDEMAR]  LACKEY, MERCEDES  H:2
TAKE AWAY THE FLOWERS  CHILTON, IRMA  H
TAKE BACK PLENTY [PLENTY 1]  GREENLAND, COLIN  C/P:2,P
TAKEOFF!  GARRETT, RANDALL  LP
TAKEOFF  KORNBLUTH, C. M.  C/P,P,H
TAKEOFF TOO!  GARRETT, RANDALL  LP
TAKEOVER, THE  EDMONDSON, G. C./KOTLAN, C. M.  P
TAKERS, THE  AHERN, JERRY  C/P,P
TAKERS, THE [@OZ CHRONICLES #1]  RIDLEY, R. W.  LP
TAKING, THE  KOONTZ, DEAN R.  H
TAKING FLIGHT [ETHSHAR]  WATT-EVANS, LAWRENCE  P
TAKING OF PLANET 5, THE [DOCTOR WHO EIGHTH DOCTOR]  BUCHER-JONES, SIMON/CLAPHAM, MARK  P
TAKING OF SATCON STATION, THE  COHEN, BARNEY/BAEN, JAMES  P
TAKING TIME [EMMA MERRIGAN 3]  ABBEY, LYNN  P
TAKING WING [@STAR TREK TITAN 1]  MARTIN, MICHAEL A./MANGELS, ANDY  P
TALBOT MUNDY: MESSENGER OF DESTINY  GRANT, DONALD M.  R/REF-H
TALBOTT AGREEMENT, THE  GARVIN, RICHARD M./ADDEO, EDMOND G.  H,P
TALE AND ITS MASTER, THE  RUTHERFORD, MICHAEL  SFWA
TALE OF THE BODY THIEF, THE [LESTAT 4]  RICE, ANNE  H:2,P
TALE OF THE CHILDREN OF HURIN, THE [LORD OF THE RINGS]  TOLKIEN, J. R. R.  H:2
TALE OF THE COMET  GREEN, ROLAND J.  P
TALE OF THE THUNDERBOLT [@VAMPIRE EARTH #3]  KNIGHT, E. E.  P
TALE OF THREE LIONS, A  HAGGARD, H. RIDER  H
TALE OF TIME CITY, A  JONES, DIANA WYNNE  LP
TALE OF TWO CLOCKS, A  SCHMITZ, JAMES H.  H:2,P
TALE OF TWO PIPERS, A [@CHARMED]  HARRISON, EMMA  P
TALE SPIN  CLADINGBOEL, NEIL C.  SR-P
TALE THAT WAGS THE GOD, THE  BLISH, JAMES  R/REF-H
TALENT FOR THE INVISIBLE, A  GOULART, RON  C/P:2,P
TALENT FOR WAR, A [ALEX BENEDICT 1]  MCDEVITT, JACK  H,SFWA-P,P
TALENTS, INCORPORATED  LEINSTER, MURRAY  C/P:2,P
TALES BEFORE TOLKIEN  ANDERSON, DOUGLAS A.  HA
TALES BY MOONLIGHT  SALMONSON, JESSICA AMANDA  SR-HA,PA,SFWA-P
TALES BY MOONLIGHT II  SALMONSON, JESSICA AMANDA  PA,SFWA-P
TALES FROM EARTHSEA [EARTHSEA]  LE GUIN, URSULA K.  H
TALES FROM EARTHSEA [EARTHSEA]/OTHER WIND, THE [EARTHSEA]  LE GUIN, URSULA K.  H
TALES FROM GAVAGAN’S BAR (EXPANDED)  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  C/P,SR-H,P,H
TALES FROM JABBA’S PALACE [STAR WARS]  ANDERSON, KEVIN J.  PA,C/PA
TALES FROM PLANET EARTH  CLARKE, ARTHUR C.  H:2,LP,P
TALES FROM TETHEDRIL  SIEGEL, SCOTT  PA
TALES FROM THE ARABIAN NIGHTS (BURTON TRANSLATION)  ANONYMOUS  HA
TALES FROM THE CAPTAIN’S TABLE [STAR TREK]  DECANDIDO, KEITH R. A.  LPA
TALES FROM THE CRYPT  OLECK, JACK  C/P,P
TALES FROM THE DARK MILLENNIUM [@WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
TALES FROM THE DARKSIDE 1  GALIN, MITCHELL/ALLEN, TOM  PA
TALES FROM THE EMPIRE [@STAR WARS]  SCHWEIGHOFER, PETER  PA
TALES FROM THE FORBIDDEN PLANET  KAVENY, ROZ  SR-LPA

TITLE  AUTHOR  CODE
TALES FROM THE HOUSE OF MYSTERY #1  OLECK, JACK  C/P,P
TALES FROM THE HOUSE OF MYSTERY #2  OLECK, JACK  C/P,P
TALES FROM THE MOS EISLEY CANTINA [STAR WARS]  ANDERSON, KEVIN J.  C/PA,PA
TALES FROM THE NEW REPUBLIC [@STAR WARS]  SCHWEIGHOFER, PETER/CAREY, CRAIG  PA
TALES FROM THE NEW TWILIGHT ZONE  STRACZYNSKI, J. MICHAEL  P,SFWA-P
TALES FROM THE NIGHTSIDE  GRANT, CHARLES L.  SR-H,SFWA-P
TALES FROM THE PLANET EARTH  POHL, FREDERIK/HULL, ELIZABETH ANNE  SR-HA,C/PA,PA
TALES FROM THE SPACEPORT BAR  SCITHERS, GEORGE/SCHWEITZER, DARRELL PA,C/PA
TALES FROM THE TEN-TAILED CAT [WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  CX
TALES FROM THE UNKNOWN  SINGER, KURT  PA
TALES FROM THE VULGAR UNICORN [@THIEVES’ WORLD #2]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
TALES FROM THE "WHITE HART"  CLARKE, ARTHUR C.  H:2,P
TALES FROM WATERSHIP DOWN [WATERSHIP DOWN 2]  ADAMS, RICHARD  C/P,P
TALES IN A JUGULAR VEIN  BLOCH, ROBERT  P
TALES IN SPACE  CROWTHER, PETER  HA,LPA
TALES OF FANTASY  JENKS, TUDOR  HA
TALES OF GOOSEFLESH AND LAUGHTER  WYNDHAM, JOHN  C/P:2,P
TALES OF HEARSAY  CONRAD, JOSEPH  H
TALES OF HORROR 1  MACHEN, ARTHUR  C/P:2,P
TALES OF HORROR AND THE SUPERNATURAL 2  MACHEN, ARTHUR  P
TALES OF INSPECTOR LEGRASSE, THE  LOVECRAFT, H. P./HENDERSON, C. J.  LP
TALES OF KNOWN SPACE [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P
TALES OF LOVE AND HORROR  CONGDON, DON  PA
TALES OF MITHGAR [MITHGAR 6]  MCKIERNAN, DENNIS L.  LP
TALES OF MYSTERY  POE, EDGAR ALLAN  P
TALES OF NEVERYON  DELANY, SAMUEL R.  C/P:2,P
TALES OF OLD EARTH  SWANWICK, MICHAEL  H
TALES OF OUTER SPACE  WOLLHEIM, DONALD A.  DD073
TALES OF PIRX THE PILOT  LEM, STANISLAW  H:2,P
TALES OF POWER  CASTANEDA, CARLOS  [C/RAN-P]
TALES OF RAVENLOFT [RAVENLOFT]  THOMSEN, BRIAN M.  C/PA:2,PA
TALES OF RIVERWORLD [RIVERWORLD]  FARMER, PHILIP JOSE  C/PA,PA
TALES OF ROBIN HOOD  EMERY, CLAYTON  P,SFWA-P
TALES OF SCIENCE AND SORCERY  SMITH, CLARK ASHTON  SR-P
TALES OF SCIENCE FICTION  BALL, BRIAN N.  SR-PA,C/PA
TALES OF TEN WORLDS  CLARKE, ARTHUR C.  C/P:2,P
TALES OF TERROR  FRENCH, JOSEPH L.  HA
TALES OF TERROR  KARLOFF, BORIS  HA
TALES OF TERROR AND THE SUPERNATURAL  COLLINS, WILKIE  LP
TALES OF THE BLACK WIDOWERS  ASIMOV, ISAAC  C/P:2,P
TALES OF THE BOUNTY HUNTERS [STAR WARS]  ANDERSON, KEVIN J.  PA
TALES OF THE CTHULHU MYTHOS  TURNER, JIM  HA
TALES OF THE CTHULHU MYTHOS 1  DERLETH, AUGUST  PA,C/PA
TALES OF THE CTHULHU MYTHOS 2  DERLETH, AUGUST  PA,C/PA
TALES OF THE DOMINION WAR [@STAR TREK]  DECANDIDO, KEITH R. A.  LPA
TALES OF THE FLYING MOUNTAINS  ANDERSON, POUL  C/P,P,H
TALES OF THE FRIGHTENED  AVALLONE, MICHAEL  C/P,P
TALES OF THE GRAND TOUR  BOVA, BEN  H
TALES OF THE HORSECLANS [HORSECLANS 1,2,3]  ADAMS, ROBERT  H
TALES OF THE INCREDIBLE  BALLANTINE BOOKS  SCX
TALES OF THE KNIGHTS TEMPLAR  KURTZ, KATHERINE  C/PA,PA
TALES OF THE LAST WAR, THE [@EBERRON]  SEHESTEDT, MARK  PA
TALES OF THE LOVECRAFT MYTHOS  PRICE, ROBERT M.  HA
TALES OF THE MYSTERIOUS AND MACABRE  BLACKWOOD, ALGERNON  H
TALES OF THE OCCULT  WOLF, JACK C./WOLF, BARBARA  PA
TALES OF THE OLD WORLD [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  LPA
TALES OF THE QUINTANA ROO  TIPTREE, JAMES JR.  H
TALES OF THE RIGANTE [RIGANTE 1,2]  GEMMELL, DAVID  H
TALES OF THE SLAYER, VOL. 1 [BUFFY]  POCKET PULSE  LP
TALES OF THE SLAYER, VOL. 4 [@BUFFY]  SIMON SPOTLIGHT ENTERTAINMENT  LPA
TALES OF THE SLAYER: VOLUMES 2 & 3 [BUFFY]  SFBC  HA
TALES OF THE STAR WOLF [STAR WOLF 1,2,3]  GERROLD, DAVID  H
TALES OF THE UNCANNY AND SUPERNATURAL  BLACKWOOD, ALGERNON  H
TALES OF THE WHITE WOLF [ELRIC]  KRAMER, EDWARD E.  PA
TALES OF THE WITCH WORLD [WITCH WORLD]  NORTON, ANDRE  SR-HA,HA
TALES OF THE WITCH WORLD 2 [WITCH WORLD]  NORTON, ANDRE  PA,SFWA-P
TALES OF THE WITCH WORLD 3 [WITCH WORLD]  NORTON, ANDRE  C/PA,PA
TALES OF THREE HEMISPHERES  DUNSANY, LORD  SR-H
TALES OF UNCLE TRAPSPRINGER [@DRAGONLANCE]  MCKEONE, DIXIE LEE  P
TALES OF UNEASE  BURKE, JOHN  PA
TALES OF WITCHCRAFT  DALBY, RICHARD  HA
TALES OF WONDER  WELLS, H. G.  H
TALES OF ZOTHIQUE [ZOTHIQUE]  SMITH, CLARK ASHTON  LP
TALES OUT OF DUNWICH  PRICE, ROBERT M.  LPA

The MITSFS Titledex  September 24, 2007  Page 361
TITLE  AUTHOR  CODE
TALES OUT OF INNSMOUTH  PRICE, ROBERT M.  LPA
TALES TO BE TOLD IN THE DARK  DAVENPORT, BASIL  PA
TALES TO TREMBLE BY  SUTTON, STEPHEN P.  HA
TALIESIN [@PENDRAGON #1]  LAWHEAD, STEPHEN R.  C/P:2,P
TALION REVENANT  STACKPOLE, MICHAEL A.  P
TALISKER [@LAST CLANSMAN #1]  LAU, MILLER  P
TALISMAN, THE  KING, STEPHEN/STRAUB, PETER  C/P,P,H
TALISMAN [@EARTHDAWN]  LEWIS, SAM  PA
TALISMAN OF DEATH  JACKSON, STEVE/LIVINGSTONE, IAN  P
TALISMANS OF SHANNARA, THE [SHANNARA HERITAGE 4]  BROOKS, TERRY  H:2
TALKERS, THE  CHAMBERS, ROBERT W.  H
TALKING BACK  DUCHAMP, L. TIMMEL  LPA
TALKING COFFINS OF CRYO-CITY, THE  PARENTEAU, SHIRLEY  H

TALKING DEVIL, THE [DOC SAVAGE #113]/ TEN TON SNAKES, THE [DOC SAVAGE 114] ROBESON, KENNETH

@P

TALKING MAN  BISSON, TERRY  H,SFWA,P
TALKING TO DRAGONS [ENCHANTED FOREST 4]  WREDE, PATRICIA C.  P
TALL STONES, THE [SACRED STONES 1]  CALDECOTT, MOYRA  P
TALLEY’S TRUTH  ROSS, PHILIP  SFWA
TALON AND FANG [OUTLANDERS HEART OF THE WORLD 1]  AXLER, JAMES  P
TALON AND THE DRAGONS OF CRINNELIA  METZ, DIANA  P
TALON OF THE SILVER HAWK [CONCLAVE OF SHADOWS 1]  FEIST, RAYMOND E.  P
TALONS [@WITCHBLADE]  DE CHANCIE, JOHN  P
TALONS [@WITCHBLADE]/TERRIBLE BEAUTY, A [@WITCHBLADE]  DE CHANCIE, JOHN/MILLER, JOHN J.  P
TALONS OF SCORPIO [DRAY PRESCOT 30]  AKERS, ALAN BURT  P
TALONS OF TIME  TWITCHELL, PAUL  P
TALTOS [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
TALTOS [MAYFAIR WITCHES 3]  RICE, ANNE  H:2,P
TALYN [KORRE]  LISLE, HOLLY  H,P
TAM LIN  DEAN, PAMELA C.  H,P
TAM, SON OF THE TIGER  KLINE, OTIS ADELBERT  H
TAMA OF THE LIGHT COUNTRY  CUMMINGS, RAY  C/P:2,P
TAMA, PRINCESS OF MERCURY  CUMMINGS, RAY  C/P,P
TAMASTARA  LEE, TANITH  C/P,P
TAMBU  ASPRIN, ROBERT LYNN  H:2,LP
TAMING THE FOREST KING  EDWARDS, CLAUDIA J.  C/P,P
TAMINY [MERI 2]  BOHNHOFF, MAYA KAATHRYN  P,SFWA-P
TAMSIN  BEAGLE, PETER S.  H:2
TAN MING  STORMONT, LAN  H
TANAKH  JEWISH PUBLICATION SOCIETY  PA
TANAR OF PELLUCIDAR [PELLUCIDAR #3]  BURROUGHS, EDGAR RICE  C/P:2,@P
TANELORN ARCHIVES, THE  BILYEU, RICHARD  R/IND:2
TANEQUIL [HIGH DRUID OF SHANNARA 2]  BROOKS, TERRY  H
TANGENTS  BEAR, GREG  H:2,SFWA,P
TANGERINE  GRAY, LINDA CROCKETT  SFWA-P
TANGLE BOX, THE [LANDOVER 4]  BROOKS, TERRY  C/P,P,H
TANGLED LANDS, THE  SHETTERLY, WILL  C/P,P
TANGLED UP IN BLUE [SNOW QUEEN 4]  VINGE, JOAN D.  H:2
TANGLED WEBS [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 2]  CUNNINGHAM, ELAINE  P
TANGLED WEBS  MUDD, STEVE  P
TANGLEWOOD TALES  HAWTHORNE, NATHANIEL  P
TANGO CHARLIE AND FOXTROT ROMEO  VARLEY, JOHN  SFWA-TD4,TD4
TANIS, THE SHADOW YEARS [DRAGONLANCE PRELUDES II 3]  SIEGEL, BARBARA/SIEGEL, SCOTT  C/P:2,P
TANITH  SHACKLEFORD, JACK D.  SR-P
TANK LORDS, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P
TANTRAS [FORGOTTEN REALMS AVATAR 2]  AWLINSON, RICHARD  C/P:2,P
TANTRAS [@FORGOTTEN REALMS AVATAR #2]  CIENCIN, SCOTT  P
TAPESTRIES [MAGIC: THE GATHERING]  ICE, KATHY  LP
TAPESTRY OF DARK SOULS [@RAVENLOFT]  BERGSTROM, ELAINE  P
TAPESTRY OF LIONS, A [CHEYSULI 8]  ROBERSON, JENNIFER  C/P,P
TAPESTRY OF MAGICS, A  DALEY, BRIAN  C/P,P
TAPESTRY OF TIME, A [BIRD OF KINSHIP 3]  COWPER, RICHARD  P
TAPESTRY WARRIORS, THE [TORIN 3]  WILDER, CHERRY  H
TAPPING THE DREAM TREE [NEWFORD]  DE LINT, CHARLES  H
TAR-AIYM KRANG, THE [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
TARA OF THE TWILIGHT  CARTER, LIN  SR-P
TARAN WANDERER [PRYDAIN 4]  ALEXANDER, LLOYD  C/P:3,LP
TARGET  MORRIS, JANET/DRAKE, DAVID  C/P,SFWA-P,P
TARGET EARTH [@PRIMORTALS]  PERRY, STEVE  P
TARGET MOON  WARSHOFSKY, FRED  [C/RAN-H]
TARGET OF OPPORTUNITY [@MECHWARRIOR DARK AGE/BATTLETECH]  PARDOE, BLAINE LEE  P
TARGET STAR, THE [@PERRY RHODAN #92]  ACE  C/PA:2,PA
TARGET: TERRA  JANIFER, LAURENCE M./TREIBICH, S. J.  DH091
TARK AND THE GOLDEN TIDE  MACCONNELL, COLUM  P

The MITSFS Titledex  September 24, 2007  Page 362
TITLE  AUTHOR  CODE
TARNISHED UTOPIA  JAMESON, MALCOM  GN5.27
TARNSMAN OF GOR [GOR 1]  NORMAN, JOHN  C/P:2,P
TAROT [TAROT 1,2,3]  ANTHONY, PIERS  LP

TAROT FANTASTIC GREENBERG, MARTIN H./ SCHIMEL, LAWRENCE

C/PA,PA

TAROT TALES  POLLACK, RACHEL/MATTHEWS, CAITLIN  PA
TARRA KHASH: HROSSAK! [PRIMAL LAND 2]  LUMLEY, BRIAN  H
TARRANO THE CONQUEROR  CUMMINGS, RAY  H
TARTARUS INCIDENT, THE  GREENLEAF, WILLIAM  P
TARZAN ALIVE  FARMER, PHILIP JOSE  C/P,SFWA,P,H
TARZAN AND THE ANT-MEN [TARZAN #10]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN AND THE CASTAWAYS [TARZAN #24]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN AND THE CITY OF GOLD [TARZAN #16]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN AND THE FORBIDDEN CITY [TARZAN #20]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AND "THE FOREIGN LEGION" [TARZAN #22]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND THE GOLDEN LION [TARZAN #9]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN AND THE JEWELS OF OPAR [TARZAN #5]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN AND THE LEOPARD MEN [TARZAN #18]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND THE LION MAN [TARZAN #17]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN AND THE LOST EMPIRE [TARZAN #12]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN AND THE MADMAN [TARZAN #23]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN AND THE SILVER GLOBE [TARZAN]  WERPER, BARTON  SR-P
TARZAN AND THE SNAKE PEOPLE [TARZAN]  WERPER, BARTON  SR-P
TARZAN AND THE VALLEY OF GOLD [TARZAN]  LEIBER, FRITZ  P
TARZAN AND THE WELL OF SLAVES [TARZAN]  NILES, DOUGLAS  P
TARZAN AT THE EARTH’S CORE [TARZAN #13/PELLUCIDAR 4]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN, LORD OF THE JUNGLE [TARZAN #11]  BURROUGHS, EDGAR RICE  C/P,@P,H:2
TARZAN OF THE APES [TARZAN #1]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN: THE EPIC ADVENTURES [TARZAN]  SALVATORE, R. A.  P
TARZAN THE INVINCIBLE [TARZAN #14]  BURROUGHS, EDGAR RICE  C/P:2,@P
TARZAN THE MAGNIFICENT [TARZAN #21]  BURROUGHS, EDGAR RICE  C/P,@P
TARZAN THE TERRIBLE [TARZAN #8]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN THE UNTAMED [TARZAN #7]  BURROUGHS, EDGAR RICE  H:2,@P
TARZAN TRIUMPHANT [TARZAN #15]  BURROUGHS, EDGAR RICE  C/P,@P,H
TARZAN TWINS, THE [TARZAN]  BURROUGHS, EDGAR RICE  H
TARZAN’S QUEST [TARZAN #19]  BURROUGHS, EDGAR RICE  @P
TASHKENT CRISIS, THE  CRAIG, WILLIAM  P
TASK FLIGHT  LANG, KING  SR-P
TASK FORCE MARS [@STARSTRIKE]  DOCKERY, KEVIN/NILES, DOUGLAS  P
TASTE FOR BLOOD, A  GREENBERG, MARTIN H.  LPA
TASTE OF MAGIC, A  NORTON, ANDRE/RABE, JEAN  H
TASTE OF NIGHT, THE [ZODIAC 2]  PETTERSON, VICKI  P
TASTE OF THE FUTURE, A: FREE SAMPLER  ROC BOOKS  PA
TASTEN IN DE TOEKOMST  CLARKE, ARTHUR C.  FP
TATHAM MOUND  ANTHONY, PIERS  H,SFWA-P,P
TATJA GRIMM’S WORLD  VINGE, VERNOR  C/P:2,P
TATOOINE GHOST [@STAR WARS]  DENNING, TROY  P
TAU ZERO  ANDERSON, POUL  H:2,P
TAURUS FOUR  VALE, RENA  P
TAUSENDJAHRES PLAN, DER  ASIMOV, ISAAC  FP
TAYLOR’S ARK [TAYLOR’S ARK 1]  NYE, JODY LYNN  P
TEA FROM AN EMPTY CUP [KONSTANTIN 1]  CADIGAN, PAT  H,P
TEA WITH THE BLACK DRAGON  MACAVOY, R. A.  C/P:2,P
TEACHER’S PET [@SABRINA THE TEENAGE WITCH SALEM #2]  BARNES-SVARNEY, PATRICIA  P
TEACHINGS OF DON JUAN: A YAQUI WAY OF KNOWLEDGE, THE  CASTANEDA, CARLOS  [C/RAN-P:2]
TEACHINGS OF THE COMPASSIONATE BUDDHA, THE  BUDDHA, THE  PA

TEARS OF THE FURIES [@MENAGERIE] GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

P

TEARS OF THE ORACLE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] RICHARDS, JUSTIN

P

TEARS OF THE SINGERS, THE [STAR TREK]  SNODGRASS, MELINDA M.  C/P:3,P
TEARS OF TIME  ASIRE, NANCY  P,SFWA-P
TECH BATTLEGROUND [@C.A.D.S. #2]  SIEVERT, JOHN  P
TECH COMMANDO [@C.A.D.S. #3]  SIEVERT, JOHN  P
TECH-HEAVEN  NAGATA, LINDA  P
TECH SATAN [@C.A.D.S. #5]  SIEVERT, JOHN  P
TECH STRIKE FORCE [@C.A.D.S. #4]  SIEVERT, JOHN  P
TECHNICOLOR TIME MACHINE, THE  HARRISON, HARRY  C/P:2,SFWA-P,P
TECHNOBABEL [@SHADOWRUN 31]  KENSON, STEPHEN  P
TECHNOGENESIS  MITCHELL, SYNE  P
TECHNOKILL [@STARFIST 5]  SHERMAN, DAVID/CRAGG, DAN  P
TECHNOS [DUMAREST 7]  TUBB, E. C.  C/P,D79975
TECHNOSEX  TAN, CECILIA  SR-L
TECKLA [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P

The MITSFS Titledex  September 24, 2007  Page 363
TITLE  AUTHOR  CODE
TEEN-AGE GHOST STORIES  FURMAN, A. L.  HA
TEEN-AGE OUTER SPACE STORIES  FURMAN, A. L.  HA
TEEN TERMINATORS  LAGOWSKI, BARBARA/MUMMA, RICK  P
TEENOCRACY  SHIRLEY, ROBERT  P
TEETH AND TONGUE LANDSCAPE  MELLICK, CARLTON III  ERD2
TEGNE: SOUL WARRIOR  LA PLANTE, RICHARD  P
TEG’S 1994  THEOBALD, ROBERT/SCOTT, J. M.  C/P,P
TEHANU [EARTHSEA 4]  LE GUIN, URSULA K.  H:2,P
TEK KILL [TEK 8]  SHATNER, WILLIAM  P
TEK MONEY [TEK 7]  SHATNER, WILLIAM  H,P
TEK NET [TEK 9]  SHATNER, WILLIAM  H,P
TEK POWER [TEK 6]  SHATNER, WILLIAM  P
TEK SECRET [TEK 5]  SHATNER, WILLIAM  P
TEK VENGEANCE [TEK 4]  SHATNER, WILLIAM  H,SFWA-P:2,P
TEKLAB [TEK 3]  SHATNER, WILLIAM  C/P,SFWA-P,P
TEKLORDS [TEK 2]  SHATNER, WILLIAM  C/P:2,SFWA-P:2,P
TEKWAR [TEK 1]  SHATNER, WILLIAM  C/P,SFWA-P:2,P,H
TELEMANN TOUCH, THE  HAGGARD, WILLIAM  H
TELEMPATH  ROBINSON, SPIDER  C/P:2,SFWA-P,P
TELEPATH  SELLINGS, ARTHUR  C/P:2,P
TELEPATHS DON’T NEED SAFEWORDS  TAN, CECILIA  SR-L
TELEPOWER  HOFFMAN, LEE  SR-DB0779
TELESPHERES  ESFANDIARY, S. M.  [C/RAN-P]
TELLING, THE [HAINISH]  LE GUIN, URSULA K.  H:2
TELLING, THE (PROOF) [HAINISH]  LE GUIN, URSULA K.  SR-LP
TELOS  VERSLUIS, ARTHUR  P
TELZEY AMBERDON [TELZEY]  SCHMITZ, JAMES H./FLINT, ERIC  P
TELZEY TOY, THE [TELZEY 2]  SCHMITZ, JAMES H.  C/P:2,P
TEMPER OF WISDOM, THE [ULTIMA]  ABBEY, LYNN  C/P,P
TEMPEST [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  BULIS, CHRISTOPHER  C/P,P
TEMPEST [@STAR WARS LEGACY OF THE FORCE 3]  DENNING, TROY  P
TEMPEST [@WARHAMMER 40K DAWN OF WAR 3]  GOTO, C. S.  P
TEMPEST, THE [STAR TREK DS9]  WRIGHT, SUSAN  C/P,P
TEMPING FATE  FRIESNER, ESTHER M.  H

TEMPLAR TREASURE, THE [@ADEPT #3] KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

C/P:2,SFWA-P,P

TEMPLE AND THE CROWN, THE KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

P

TEMPLE AND THE STONE, THE KURTZ, KATHERINE/

HARRIS, DEBORAH TURNER

H

TEMPLE AT LANDFALL, THE [CELAENO]  FLETCHER, JANE  LP
TEMPLE HILL [FORGOTTEN REALMS CITIES]  KARPYSHYN, DREW  P
TEMPLE OF DEATH  TUBB, E. C.  LP
TEMPLE OF ELEMENTAL EVIL, THE [@GREYHAWK]  REID, THOMAS M.  P
TEMPLE OF HASHEM, THE  BECKER, HYAM YONA  C/P,P
TEMPLE OF NIGHT [@CROW]  SOMTOW, S. P.  P

TEMPLE OF THE DRAGONSLAYER [@DRAGONLANCE NEW ADVENTURES #1/ DRAGONLANCE SPELLBINDER QUARTET 1] WAGGONER, TIM

LP

TEMPLE OF THE PAST, THE  WUL, STEFAN  H,SFWA
TEMPLE OF THE SUN, THE [SACRED STONES 2]  CALDECOTT, MOYRA  H,P
TEMPLE OF THE TEN, THE  BEDFORD-JONES, H./ROBERTSON, W. C.  SR-H
TEMPLE OF THE WINDS [SWORD OF TRUTH 4]  GOODKIND, TERRY  H:2,P
TEMPLE OF TRUTH, THE [DUMAREST 31]  TUBB, E. C.  C/P,P
TEMPLE SERVANT AND OTHER STORIES, THE  MORROUGH, E. R.  H
TEMPLES OF AYOCAN, THE [@BLADE #14]  LORD, JEFFREY  C/P,P
TEMPO DI STREGHE  GUNN, JAMES E.  FP
TEMPORARY AGENCY [LIVING WORLD 2]  POLLACK, RACHEL  H:2,LP
TEMPS [@TEMPS 1]  GAIMAN, NEIL/STEWART, ALEX  SR-PA
TEMPTATION OF ELMINSTER, THE [@FORGOTTEN REALMS]  GREENWOOD, ED  P
TEMPTED CHAMPIONS [@BUFFY]  NAVARRO, YVONNE  C/P,P
TEMPTING DANGER [LILY YU 1]  WILKS, EILEEN  P
TEMPTING EVIL [RILEY JENSON 3]  ARTHUR, KERI  P
TEMPTING FATE [SAINT-GERMAIN 5]  YARBRO, CHELSEA QUINN  H:2
TEMPTING OF THE WITCH KING, THE  BLACKFORD, RUSSELL  SR-P
TEMPUS [THIEVES’ WORLD TEMPUS 4]  MORRIS, JANET  C/P:2,P
TEMPUS UNBOUND [THIEVES’ WORLD TEMPUS 6]  MORRIS, JANET/MORRIS, CHRIS  C/P:2,P
TEN DRAGON TAILS  TUTT, C. T.  H
TEN EVER-LOVIN’ BLUE-EYED YEARS WITH POGO  KELLY, WALT  CX
TEN FROM INFINITY  JORGENSEN, IVAR  C/P,P
TEN LITTLE ALIENS [DOCTOR WHO PAST DOCTORS]  COLE, STEPHEN  P
TEN LITTLE WIZARDS [LORD DARCY 5]  KURLAND, MICHAEL  C/P:2,SFWA-P,P
TEN MILE TREASURE  NORTON, ANDRE  P
TEN MILLION YEARS TO FRIDAY  LYMINGTON, JOHN  H:2,P

The MITSFS Titledex  September 24, 2007  Page 364
TITLE  AUTHOR  CODE
TEN SCI-FI SHORTS  ABANTE, ABRAHAM  H,LP
TEN TALES TALL AND TRUE  GRAY, ALASDAIR  H
TEN THOUSAND LIGHT-YEARS FROM HOME  TIPTREE, JAMES JR.  C/P:2,P
TEN THOUSAND YEARS IN A BLOCK OF ICE  PARET, JOHN  H
TEN TOMORROWS  ELWOOD, ROGER  C/PA,PA
TEN YEARS TO DOOMSDAY  ANDERSON, CHESTER/KURLAND, MICHAEL C/P:2,P
TENANT, THE  TOPOR, ROLAND  P
TENANT, THE  WEST, CHARLES  H
TENDELEO’S STORY [CHAGA]  MCDONALD, IAN  GZD07305,SR-LP
TENEBREA RISING [@TENEBREA #3]  DAWSON, ROXANN/GRAHAM, DANIEL  P
TENEBREA’S HOPE [@TENEBREA #2]  DAWSON, ROXANN/GRAHAM, DANIEL  P
TENTACLES OF DAWN, THE  WILSON, ROBERT  P
TENTH CLASS, THE  RIPLEY, KAREN  C/P:2,P
10TH KINGDOM, THE  WESLEY, KATHRYN  C/P,P
TENTH PLANET, THE  COOPER, EDMUND  C/P,P,H
TENTH PLANET, THE  MELAMED, LEO  LP

TENTH PLANET, THE [@TENTH PLANET 1] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

TENTH PLANET, THE  STERLING, BRETT  P

TENTH PLANET: FINAL ASSAULT, THE [@TENTH PLANET 3] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

TENTH PLANET: OBLIVION, THE [@TENTH PLANET 2] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

10TH VICTIM, THE  SHECKLEY, ROBERT  C/P,P
TEOT’S WAR [NAGA TEOT 1]  GLADNEY, HEATHER  P,SFWA-P
TEQUILA MOCKINGBIRD  POLLOTTA, NICK  LP
TERA BEYOND, THE  MACCLOUD, MALCOLM  H:2
TERANESIA  EGAN, GREG  C/P,P,H
TERMINAL BEACH  BALLARD, J. G.  C/P,P
TERMINAL BEACH, THE  BALLARD, J. G.  SR-P
TERMINAL CAFE  MCDONALD, IAN  P
TERMINAL EXPERIMENT, THE  SAWYER, ROBERT J.  H:2,P
TERMINAL FREEDOM  MORAN, DANIEL KEYS/MORAN, JODI  SR-H
TERMINAL GAMES  PERRIMAN, COLE  P
TERMINAL ISLAND [@WARLORD #5]  FROST, JASON  P
TERMINAL MAN, THE  CRICHTON, MICHAEL  C/P,P,H
TERMINAL ROAD [@TRAVELER #8]  DRUMM, D. B.  P
TERMINATE [@NIGHT WHISTLERS #3]  TREVOR, DAN  P
TERMINATION [NOEL KEDRAN 3]  DALTON, SEAN  C/P,P
TERMINATOR, THE [TERMINATOR]  FRAKES, RANDALL/WISHER, BILL  P
TERMINATOR 2 [TERMINATOR]  FRAKES, RANDALL  C/P,P
TERMINATOR 3: RISE OF THE MACHINES [TERMINATOR]  HAGBERG, DAVID  P
TERMINATOR DREAMS [TERMINATOR 3]  ALLSTON, AARON  P
TERMINATOR: THE BURNING EARTH [TERMINATOR]  FORTIER, RON  CX
TERMINUS  EDWARDS, PETER  H
TERMINUS [DOCTOR WHO]  LYDECKER, JOHN  SR-P,C/P:2
TERMINUS EXPERIMENT, THE [@SHADOWRUN 34]  BOND, JONATHAN E./KOKE, JAK  P
TERRA!  BENNI, STEFANO  H,LP
TERRA!  LANG, KING  SR-P
TERRA DATA, THE [DUMAREST 22]  TUBB, E. C.  C/P,P
TERRA SF  NOLANE, RICHARD D.  PA
TERRA SF 2  NOLANE, RICHARD D.  C/PA,PA
TERRAFORMING EARTH  WILLIAMSON, JACK  H
TERRAN GIRLS MAKE WONDERFUL WIVES  REASONER, JAMES M.  GRD8
TERRAPIN OR  MANADAY, TILPER  H:2,LP
TERRAPLANE [DRYCO 2]  WOMACK, JACK  C/P,SFWA-P,P,H
TERRARIUM  SANDERS, SCOTT RUSSELL  C/P,SFWA-P,P
TERRE A PEUR, LA (FRENCH) [@PERRY RHODAN #2]  SCHEER, K. H.  FPR
TERRE N’EST PAS RONDE, LA  LIMAT, MAURICE  FP
TERRESTRIAL INDEX, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
TERRIBLE AWAKENING, THE  DESMOND, HUGH  H
TERRIBLE BEAUTY, A [@WITCHBLADE]  MILLER, JOHN J.  P
TERRIBLE SWIFT SWORD [@LOST REGIMENT 3]  FORSTCHEN, WILLIAM R.  C/P,P
TERRIDAE, THE [DUMAREST 25]  TUBB, E. C.  C/P,P
TERRIER [@BEKA COOPER #1]  PIERCE, TAMORA  H
TERROR, THE [@SURVIVALIST #14]  AHERN, JERRY  P
TERROR, THE  CREASEY, JOHN  C/P,P
TERROR!  HAINING, PETER  A
TERROR, THE  MACHEN, ARTHUR  VLP
TERROR  POHL, FREDERIK  C/P:2,P
TERROR!  SHAW, LARRY T.  C/PA,PA
TERROR, THE  SIMMONS, DAN  H
TERROR AT PLAY [NOT QUITE HUMAN 5]  MCEVOY, SETH  P
TERROR BY NIGHT  HURWOOD, BERNHARDT J.  [C/RAN-P]

The MITSFS Titledex  September 24, 2007  Page 365
TITLE  AUTHOR  CODE
TERROR IN THE MODERN VEIN  WOLLHEIM, DONALD A.  HA
TERROR IN THE NAVY, THE [DOC SAVAGE #33]  ROBESON, KENNETH  C/P:2,@P
TERROR OF SAINT TRINIAN’S, THE  SHY, TIMOTHY/SEARLE, RONALD  H
TERROR OF THE VERVOIDS [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  SR-P,C/P
TERROR ON PLANET IONUS  ADLER, ALLEN  C/P:2,P
TERROR ON THE MOONS OF JUPITER [@NEW TOM SWIFT #2]  APPLETON, VICTOR  LP
TERROR-TRAP!  LAZENBY, NORMAN  B
TERRORISTS OF IRUSTAN, THE  MARLEY, LOUISE  C/P:2,LP
TERRORS OF ROCK AND ROLL, THE [@SAMANTHA SLADE #4]  SMITH, SUSAN  P
TERRY CARR’S BEST SCIENCE FICTION OF THE YEAR (#14)  CARR, TERRY  C/PA,PA,SFWA-P
TERRY CARR’S BEST SCIENCE FICTION OF THE YEAR #15  CARR, TERRY  PA,SFWA-P
TERRY CARR’S BEST SCIENCE FICTION AND FANTASY OF THE YEAR #16  CARR, TERRY  HA
TERRY NATION’S BLAKE’S 7 THE PROGRAMME GUIDE [BLAKE’S 7]  ATTWOOD, TONY  R/REF-P

TERRY PRATCHETT 2000 DISCWORLD COLLECTOR’S EDITION CALENDAR, THE CALENDAR

CAL

TERRY’S UNIVERSE  MEACHAM, BETH  HA,SFWA-P
TERY, THE  WILSON, F. PAUL  BS2
TERY, THE (COMPLETE)  WILSON, F. PAUL  C/P,SFWA-P,P
TERZO DAL SOLE (ITALIAN)  MATHESON, RICHARD  FP
TESSERACT  ADDISON, JOSEPH  C/P:2,P
TESSERACTS  MERRIL, JUDITH  LPA
TESSERACTS 2  GOTLIEB, PHYLLIS/BARBOUR, DOUGLAS  LPA
TESSERACTS 3  DORSEY, CANDAS JANE/TRUSCOTT, GERRY  C/PA,PA
TESSERACTS 4  TOOLIS, LORNA/SKEET, MICHAEL  PA
TESSERACTS 5  RUNTE, ROBERT/MEYNARD, YVES  C/PA,PA
TESSERACTS 6  SAWYER, ROBERT J./CLINK, CLAYTON  LPA
TESSERACTS 7  JOHANSON, PAULA/TRUDEL, JEAN-LOUIS  LPA
TESSERACTS 8  CLUTE, JOHN/DORSEY, CANDAS JANE  LPA
TESSERACTS NINE [TESSERACTS 9]  HOPKINSON, NALO/RYMAN, GEOFF  LPA
TESSERACTS Q  VONARBURG, ELISABETH/BRIERLEY, JANE  C/PA,PA
TEST OF FIRE  BOVA, BEN  C/P,P
TEST OF THE TWINS [@DRAGONLANCE LEGENDS #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
TEST OF VENGEANCE [@BATTLETECH 51]  NYSTUL, BRYAN  P
TEST YOUR ESP  EBON, MARTIN  [C/RAN-P]
TESTAMENT [HARMONY OF WORLDS]  FREIREICH, VALERIE J.  C/P,P
TESTAMENT OF ANDROS, THE  BLISH, JAMES  P
TESTAMENT OF THE DRAGON  WEIS, MARGARET/BALDWIN, DAVID  HA
TESTAMENT XXI  SNYDER, GUY  P
TESTING  OBERNDORF, CHARLES  P
TESTING (PROOF)  OBERNDORF, CHARLES  SR-LP
TETRARCH  COMFORT, ALEX  H
TETRARCH (PROOF)  COMFORT, ALEX  SR-LP
TETRASOMY TWO  ROSSITER, OSCAR  C/P,SFWA,P,H
TEX AND MOLLY IN THE AFTERLIFE  GRANT, RICHARD  P
TEXAR THE SOUTHERNER  VERNE, JULES  H
TEXAS-ISRAELI WAR: 1999, THE  SAUNDERS, JAKE/WALDROP, HOWARD  C/P:2,P
TEXAS ON THE ROCKS [TEXAS 2]  DA CRUZ, DANIEL  C/P,P
TEXAS RUN, THE [V]  PROCTOR, GEO. W.  C/P,P
TEXAS TRIUMPHANT [TEXAS 3]  DA CRUZ, DANIEL  C/P,P
TEXTS OF FESTIVAL, THE  FARREN, MICK  C/P,P
TEXTUAL POACHERS  JENKINS, HENRY  R/REF-H

THACKERY T. LAMBSHEAD POCKET GUIDE TO ECCENTRIC & DISCREDITED DISEASES, THE VANDERMEER, JEFF/ROBERTS, MARK

LPA

THANATOS SYNDROME, THE  PERCY, WALDER  H,P
THAT DARN SQUID GOD  POLLOTTA, NICK/CLAY, JAMES  H
THAT GIRL  FAIRMAN, PAUL W.  P
THAT HIDEOUS STRENGTH [PERELANDRA 3]  LEWIS, C. S.  C/P:2,P
THAT MAN ON BETA [@BUCK ROGERS #2]  STEELE, ADDISON E.  C/P,P
THAT WAY LIES CAMELOT  WURTS, JANNY  H,P
THEATER OF TIMESMITHS  KILWORTH, GARRY  P
THEATRE OF WAR [DOCTOR WHO NEW ADVENTURES]  RICHARDS, JUSTIN  C/P:2,P
THEBES OF THE HUNDRED GATES  SILVERBERG, ROBERT  C/P,P
THEIR MAJESTIES’ BUCKETEERS  SMITH, L. NEIL  P
THEIR MAN IN THE WHITE HOUSE  ARDIES, TOM  SFWA
THEIR MASTER’S WAR  FARREN, MICK  P
T.H.E.M.  EDMONDSON, G. C.  H
THEM  FRENCH, ROBERT  P
THEM BONES  WALDROP, HOWARD  C/P,P
THEN BEGGARS COULD RIDE  NELSON, RAY FARADAY  SR-P,C/P
"THEN IS FINISHED THE MYSTERY OF GOD"  WATCHTOWER  [C/RAN-H]
THENDARA HOUSE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
THEODORE STURGEON  MENGER, LUCY  R/REF-P
THERE AND BACK AGAIN [MAX MERRIWELL 1]  MURPHY, PAT  P
THERE ARE DOORS  WOLFE, GENE  H,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 366
TITLE  AUTHOR  CODE
THERE IS NO DARKNESS  HALDEMAN, JOE/HALDEMAN, JACK C. II  C/P:2,P
THERE SHALL BE DARKNESS  MOORE, C. L.  B
THERE WAS A LITTLE MAN  JONES, GUY/JONES, CONSTANCE  H
THERE WILL BE DRAGONS (HARDCOVER AND CD) [DRAGONS 1]  RINGO, JOHN  H:2
THERE WILL BE DRAGONS (PROOF) [DRAGONS 1]  RINGO, JOHN  SR-LP
THERE WILL BE TIME  ANDERSON, POUL  H:2,P
THERE WILL BE WAR [@THERE WILL BE WAR #1]  POURNELLE, JERRY  PA,C/PA
THERE WON’T BE WAR  HARRISON, HARRY/MCALLISTER, BRUCE  C/PA,PA
THERE’S A NEW WORLD COMING  LINDSEY, HAL  [C/RAN-P:2]
THERION CHRONICLES, THE  CALVO, FRANCESCO  LP
THERMONUCLEAR WARFARE  ANDERSON, POUL  [C/RAN-P]
THEROS IRONFELD [@DRAGONLANCE WARRIORS #4]  PERRIN, DON  P
THESE GREEN FOREIGN HILLS [PEACE COMPANY]  GREEN, ROLAND J.  C/P,P
THESE OUR ACTORS [@BUFFY]  MCCONNELL, ASHLEY/KOOGLER, DORI  C/P,P
THESE OUR REVELS [@FOREVER KNIGHT]  HATHAWAY-NAYNE, ANNE  P
THESE SAVAGE FUTURIANS  HIGH, PHILIP E.  DG623
THESE WILL CHILL YOU  WRIGHT, LEE/SHEEHAN, RICHARD G.  PA
THESSALONICA  TURTLEDOVE, HARRY  C/P:2,P
THEY  MANNES, MARYA  C/P,P
THEY ALSO SERVE [SOCIETY OF HUMANITY/UNITY PARTY 3]  MOSCOE, MIKE  C/P,P
THEY ARE ALL AROUND US  CHAVEZ, RAY  P
THEY CAME FROM OUTER SPACE  WYNORSKI, JAMES  SFWA
THEY CAME, THEY SAW  HALLUMS, JAMES R.  SR-P

THEY DIED TWICE [DOC SAVAGE #105]/ SCREAMING MAN, THE [DOC SAVAGE 106] ROBESON, KENNETH

@P

THEY FLY AT CIRON  DELANY, SAMUEL R.  P
THEY FOUND ATLANTIS  WHEATLEY, DENNIS  SR-P,C/P,H
THEY KNEW TOO MUCH ABOUT FLYING SAUCERS  BARKER, GRAY  [C/RAN-H]
THEY NEVER COME BACK  BRETT, LEO  SR-P
THEY SHALL HAVE STARS  BLISH, JAMES  C/P:2,P
THEY THIRST  MCCAMMON, ROBERT R.  C/P,P
THEY USED DARK FORCES  WHEATLEY, DENNIS  SR-P,C/P
THEY WALKED LIKE MEN  SIMAK, CLIFFORD D.  C/P:2,P
THEY’D RATHER BE RIGHT  CLIFTON, MARK/RILEY, FRANK  H
"THEY’RE HERE..." [INVASION OF THE BODY SNATCHERS]  MCCARTHY, KEVIN/GORMAN, ED  R/REF-P
THIEF OF ALWAYS, THE  BARKER, CLIVE  H,P
THIEF OF DREAMS [@STAR REQUIEM #2]  COLE, ADRIAN  P,SFWA-P
THIEF OF HERMES, THE [XENA]  EMERSON, RU  P
THIEF OF KALIMAR, THE  DIAMOND, GRAHAM  C/P:2,P
THIEF OF KALIMAR, THE (PROOF)  DIAMOND, GRAHAM  SR-LP
THIEF OF LIVES [NOBLE DEAD SERIES 1 2]  HENDEE, BARB/HENDEE, J. C.  P
THIEF OF LLARN [LLARN 2]  FOX, GARDNER F.  C/P:2,P
THIEF OF SOULS [STAR SHARDS 2]  SHUSTERMAN, NEAL  P
THIEF OF THOTH, THE  CARTER, LIN  SR-DB0244,SR-DB0809
THIEF OF TIME [DISCWORLD 26]  PRATCHETT, TERRY  H:2,P
THIEF WITH NO SHADOW  GEE, EMILY  P
THIEF’S GAMBLE, THE [EINARINN 1]  MCKENNA, JULIET E.  C/P:2,P
THIEF’S TALE  ANDREWS, MICHAEL  LP
THIEVES’ CARNIVAL  HABER, KAREN  TD22,SFWA-TD22
THIEVES’ GUILD, THE [@DRAGONLANCE CROSSROADS]  CROOK, JEFF  P
THIEVES OF BLOOD [@EBERRON BLADE OF THE FLAME #1]  WAGGONER, TIM  P
THIEVES OF LIGHT [PHOTON]  HUDSON, MICHAEL  P
THIEVES’ WORLD [@THIEVES’ WORLD #1]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
THIEVES’ WORLD (BRITISH) [THIEVES’ WORLD 1]  ASPRIN, ROBERT LYNN  PA

THIN AIR [@STAR TREK NEW EARTH #5] RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY

C/P:2,P

THING, THE  CAMPBELL, JOHN W. JR.  SR-P
THING, THE  FOSTER, ALAN DEAN  P
THING, THE  MADISON, J. J.  P
THING FROM ANOTHER WORLD, THE  CAMPBELL, JOHN W. JR.  B
THING IN THE BROOK, THE  STORME, PETER  H
THINGS  HOWARD, IVAN  PA,C/PA
THINGS BEYOND MIDNIGHT  NOLAN, WILLIAM F.  H
THINGS HUNTING MEN [STARHUNTERS 2]  DRAKE, DAVID  PA,SFWA-P
THINGS INVISIBLE TO SEE  SCHIMEL, LAWRENCE  LPA
THINGS INVISIBLE TO SEE  WILLARD, NANCY  C/P,P
THINGS THAT ARE NOT THERE, THE  HENDERSON, C. J.  LP
THINGS TO COME (FILM SCRIPT WITH NOTES)  WELLS, H. G.  H
THINGS WITH CLAWS  BURNETT, WHIT/BURNETT, HALLIE  PA
THINK LIKE A DINOSAUR AND OTHER STORIES  KELLY, JAMES PATRICK  H
THINKING MACHINE, THE  FUTRELLE, JACQUES  LP
THINKING MACHINE AFFAIR, THE [@MAN FROM UNCLE #21]  BERNARD, JOEL  C/P,P

THINKING MACHINES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

HA

The MITSFS Titledex  September 24, 2007  Page 367
TITLE  AUTHOR  CODE
THINKING SEAT, THE  TATE, PETER  C/P,P
THINNER  BACHMAN, RICHARD  H
THINNER  KING, STEPHEN  C/P:2,P
THIRD BODY, THE  DANN, SAM  P
THIRD BOOK OF SWORDS, THE [@SWORDS #3]  SABERHAGEN, FRED  C/P,SFWA-P,P,H
THIRD DOCTOR, THE [DOCTOR WHO]  HOWE, DAVID J./WALKER, STEPHEN JAMES  R/REF-P
THIRD EAGLE, THE  MACAVOY, R. A.  C/P,P,H
THIRD EAR, THE  SIODMAK, CURT  C/P,P,H:2
THIRD ENCOUNTER [@STARSEA INVADERS 3]  STINE, G. HARRY  P
THIRD EYE, THE  COGSWELL, THEODORE R.  C/P,P
THIRD FORCE, THE  LAIDLAW, MARC  LP
THIRD FROM THE SUN  MATHESON, RICHARD  C/P:2,P
THIRD GHOST BOOK, THE  ASQUITH, CYNTHIA  PA
THIRD GRAVE, THE  CASE, DAVID  H
THIRD INDUSTRIAL REVOLUTION, THE  STINE, G. HARRY  [C/RAN-P:2]
THIRD LEVEL, THE  FINNEY, JACK  C/P,P,H
THIRD MAGIC, THE [FOREVER KING 3]  COCHRAN, MOLLY  P
THIRD MILLENNIUM, THE  STABLEFORD, BRIAN M./LANGFORD, DAVID R/REF-H
THIRD OMNI BOOK OF SCIENCE FICTION, THE  DATLOW, ELLEN  C/PA,PA
THIRD PLANET FROM ALTAIR, THE  PACKARD, EDWARD  P
THIRD WORLD WAR: AUGUST 1985, THE  HACKETT, JOHN, ET AL.  P
THIRDSPACE [@BABYLON 5]  DAVID, PETER  P
THIRST QUENCHERS, THE  RAPHAEL, RICK  H
THIRTEEN  MORGAN, RICHARD K.  H
13 ABOVE THE NIGHT  CONKLIN, GROFF  PA,C/PA
THIRTEEN BRACELETS, THE  LORY, ROBERT  C/P:2,P

13 CRIMES OF SCIENCE FICTION, THE ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

HA

THIRTEEN FRENCH SCIENCE FICTION STORIES  KNIGHT, DAMON  PA,C/PA
13 GREAT STORIES OF SCIENCE FICTION  CONKLIN, GROFF  PA,C/PA

13 HORRORS OF HALLOWEEN WAUGH, CAROL-LYNN ROSSEL/

GREENBERG, MARTIN H./ASIMOV, ISAAC

PA

THIRTEEN O’CLOCK AND OTHER ZERO HOURS  KORNBLUTH, C. M.  C/P:2,P
THIRTEEN SECONDS THAT ROCKED THE WORLD, OR THE MENTATOR  MEYER, JOHN J.  H
THIRTEEN TALES OF TERROR  LONDON, JACK  P
THIRTEEN UNCANNY TALES  GREEN, ROGER LANCELYN  HA
THIRTEENTH HOUSE, THE [TWELVE HOUSES]  SHINN, SHARON  H
13TH IMMORTAL, THE  SILVERBERG, ROBERT  DD223
13TH LEGION [WARHAMMER 40K LAST CHANCERS 1]  THORPE, GAV  P
THIRTEENTH MAJESTRAL, THE  PEIRCE, HAYFORD  C/P:2,P
THIRTEENTH ORPHAN, THE  CHAUNDLER, CHRISTINE  H
THIRTEENTH SCROLL, THE [AGHAMORE 2]  NEASON, REBECCA  P
30 DAY WONDER  WILSON, RICHARD  C/P:2,P
THIRTY-FIRST OF FEBRUARY, THE  BOND, NELSON  SR-H,H
THIRTY-FIRST OF JUNE, THE  PRIESTLEY, J. B.  H
34TH RULE, THE [STAR TREK DS9]  SHIMERMAN, ARMIN/GEORGE III, DAVID R.  C/P,P
THIRTY STRANGE STORIES  WELLS, H. G.  H:2
THIS ALIEN SHORE  FRIEDMAN, C. S.  C/P,P,H
THIS BAFFLING WORLD 1  GODWIN, JOHN  [C/RAN-P:2]
THIS BAFFLING WORLD 3  GODWIN, JOHN  [C/RAN-P]
THIS BUSINESS OF BOMFOG  DUKE, MADELAINE  C/P,P,H
THIS CROWDED EARTH  BLOCH, ROBERT  SR-DB0080
THIS DARKENING UNIVERSE [JAN DARZEK]  BIGGLE, LLOYD JR.  H,SFWA
THIS DAY ALL GODS DIE [GAP 5]  DONALDSON, STEPHEN R.  H:2,P
THIS FORTRESS WORLD  GUNN, JAMES E.  H:2,P
THIS FORTRESS WORLD (ABRIDGED)  GUNN, JAMES E.  DD223
THIS GRAY SPIRIT [STAR TREK DS9 MISSION GAMMA #2]  JARMAN, HEATHER  C/P,P
THIS HIGH MAN  LEHMAN, MILTON  [C/RAN-P]
THIS IMMORTAL  ZELAZNY, ROGER  SFWA-P:2,H:2,P
THIS IS MY FUNNIEST  RESNICK, MIKE  LPA
THIS IS OUTER SPACE  MOTZ, LLOYD  R/REF-P
THIS IS THE WAY THE WORLD BEGINS  MCINTOSH, J. T.  P
THIS IS THE WAY THE WORLD ENDS  MORROW, JAMES  H,SFWA-P
THIS IS YOUR LIFE, BHODILI [@PHOTON #4]  PETERS, DAVID  P
THIS ISLAND EARTH  JONES, RAYMOND F.  H:2
THIS OTHER EDEN  ELTON, BEN  LP
THIS PERFECT DAY  LEVIN, IRA  C/P,P
THIS PLACE HAS NO ATMOSPHERE  DANZIGER, PAULA  P
THIS PLANET FOR SALE  HAY, GEORGE  SR-P

THIS ROUGH MAGIC [SHADOW OF THE LION 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H,P

THIS ROUGH MAGIC (PROOF) [SHADOW OF THE LION 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

SR-LP

THIS SCEPTER’D ISLE [SCEPTER’D ISLE 1]  LACKEY, MERCEDES/GELLIS, ROBERTA  P

TITLE  AUTHOR  CODE
THIS SENTIENT EARTH  HOYLE, TREVOR  P
THIS SHINING SEA  PHILLIES, GEORGE  LP
THIS SHOAL OF SPACE  ARGO, JOHN  LP
THIS SIDE OF INFINITY  CARR, TERRY  PA,C/PA
THIS SIDE OF JUDGEMENT  DUNN, J. R.  H,P
THIS STAR SHALL ABIDE [NOREN 1]  ENGDAHL, SYLVIA LOUISE  H
THIS STRANGE TOMORROW  LONG, FRANK BELKNAP  C/P:2,P
THIS TIME OF DARKNESS  HOOVER, H. M.  H
THIS TOWN WILL NEVER LET US GO [@FACTION PARADOX 1]  MILES, LAWRENCE  LP
THIS WITCH  TUCKER, WILSON  H
THIS WORLD IS TABOO [MED 3]  LEINSTER, MURRAY  C/P:2,P
THOMAS AND THE WARLOCK  HUNTER, MOLLIE  LP

THOMAS BURNETT SWANN:

A BRIEF CRITICAL BIOGRAPHY AND ANNOTATED BIBLIOGRAPHY COLLINS, ROBERT A.

R/REF-P

THOMAS BURNETT SWANN BIBLIOGRAPHY, A  ROEHM, BOB  R/IND-F
THOMAS THE RHYMER  KUSHNER, ELLEN  H:2,P
THONGOR AGAINST THE GODS  CARTER, LIN  C/P,P
THONGOR AND THE DRAGON CITY  CARTER, LIN  C/P:2,P
THONGOR AND THE WIZARD OF LEMURIA  CARTER, LIN  C/P,P
THONGOR AT THE END OF TIME  CARTER, LIN  P
THONGOR FIGHTS THE PIRATES OF TARAKUS  CARTER, LIN  C/P,P
THONGOR IN THE CITY OF MAGICIANS  CARTER, LIN  P
THONGOR OF LEMURIA  CARTER, LIN  SR-P,C/P
THORA’S SACRIFICE [@PERRY RHODAN #70]  ACE  PA
THORGAL: CHILD OF THE STARS  VAN HAMME, JEAN/ROSINSKI, GRZEGORZ  CX
THORGAL: THE SORCERESS BETRAYED  VAN HAMME, JEAN/ROSINSKI, GRZEGORZ  CX
THORN [DRACULA]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
THORN AND NEEDLE  THOMPSON, PAUL B.  P
THORN OGRES OF HAGWOOD [HAGWOOD 1]  JARVIS, ROBIN  H
THORNE SMITH 3-DECKER, THE  SMITH, THORNE  H
THORNHOLD [@FORGOTTEN REALMS HARPERS]  CUNNINGHAM, ELAINE  P
THORNS  SILVERBERG, ROBERT  C/P:2,P
THORNYHOLD  STEWART, MARY  P
THOR’S HAMMER [@FUTURE AT WAR #1]  BRETNOR, REGINALD  SFWA-P,C/PA,PA
THOR’S HAMMER  WHITEFORD, WYNNE  SR-P,C/P
THOSE GENTLE VOICES  EFFINGER, GEORGE ALEC  P
THOSE IDIOTS FROM EARTH  WILSON, RICHARD  C/P,P
THOSE WHO CAN  WILSON, ROBIN SCOTT  C/PA,PA
THOSE WHO FALL FROM THE SUN  STONE, JOSEPHINE RECTOR  H
THOSE WHO FAVOR FIRE  RANDALL, MARTA  C/P,P
THOSE WHO HUNT THE NIGHT [LONDON VAMPIRES 1]  HAMBLY, BARBARA  C/P,P,H:2
THOSE WHO SURVIVE (PROOF)  BULYCHEV, KIR  H,SR-LP
THOSE WHO WALK IN DARKNESS  RIDLEY, JOHN  H
THOSE WHO WATCH  SILVERBERG, ROBERT  C/P:2,P
THOUGHTS OF GOD  KANALY, MICHAEL  H,LP
THOUGHTS THROUGH SPACE  WILKIN, SIR HUBERT/SHERMAN, HAROLD M. H
THOUSAND CITIES, THE [TIME OF TROUBLES 3]  TURTLEDOVE, HARRY  P
THOUSAND DEATHS, A  EFFINGER, GEORGE ALEC  H
THOUSAND-HEADED MAN, THE [DOC SAVAGE #2]  ROBESON, KENNETH  C/P:2,@P
THOUSAND ORCS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #1]  SALVATORE, R. A.  P
THOUSAND SHRINE WARRIOR [TOMOE GOZEN 3]  SALMONSON, JESSICA AMANDA  C/P:2,P
THOUSAND WORDS FOR STRANGER, A [TRADE PACT 1]  CZERNEDA, JULIE E.  H:2,P
THOUSAND-YEAR PLAN, THE  ASIMOV, ISAAC  C/P,DD110,P
THOUSAND YEARS A MINUTE, A  CLAUDY, CARL H.  H
1000 YEARS ON, A  MULLER, JOHN E.  SR-P
THOUSANDFOLD THOUGHT, THE [@PRINCE OF NOTHING #3]  BAKKER, R. SCOTT  H
THOUSANDSTAR [CLUSTER #4]  ANTHONY, PIERS  C/P:2,@P
THRALL AND THE DRAGON’S HEART, THE [ALFAR 3]  BOYER, ELIZABETH H.  C/P:2,P
THRALL OF HYPNO, THE [@PERRY RHODAN #20]  ACE  C/PA,PA
THRAN, THE [@MAGIC: THE GATHERING]  KING, J. ROBERT  P
THRAXAS [THRAXAS]  SCOTT, MARTIN  SR-P,C/P
THREAD THAT BINDS THE BONES, THE  HOFFMAN, NINA KIRIKI  P,SFWA-P
THREADS OF AMBITION [@BATTLETECH #44/CAPELLAN SOLUTION 1]  COLEMAN, LOREN L.  P
THREADS OF TIME  SILVERBERG, ROBERT  SR-HA,C/PA,PA
THREATS ... AND OTHER PROMISES  VINGE, VERNOR  C/P:2,SFWA-P,P
THREE [@STAR TREK TNG STARGAZER #3]  FRIEDMAN, MICHAEL JAN  P
THREE: 1971  RANDOM HOUSE  HA
3 ADVENTURE NOVELS  HAGGARD, H. RIDER  LP
THREE AGAINST THE WITCH WORLD [WITCH WORLD ESTCARP 1]  NORTON, ANDRE  C/P:2,P
THREE-BLADED DOOM  HOWARD, ROBERT E.  P
THREE BOOKS OF KNOWN SPACE [KNOWN SPACE]  NIVEN, LARRY  LP
3 BY FINNEY  FINNEY, JACK  H,LP
THREE BY HEINLEIN  HEINLEIN, ROBERT A.  H:2
THREE COMPLETE NOVELS [KELVIN HACKLEBERRY 1,2,3]  ANTHONY, PIERS/MARGROFF, ROBERT E.  H
THREE-CORNERED WAR, THE [REGIMENT 5]  DALMAS, JOHN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 369
TITLE  AUTHOR  CODE
THREE CORNERS TO NOWHERE  CAIDIN, MARTIN  C/P,P
THREE DAYS TO NEVER  POWERS, TIM  H
THREE-EYES  GORDON, STUART  C/P,P
THREE FACES OF SCIENCE FICTION  LOWNDES, ROBERT A. W.  R/REF-H
THREE FACES OF TIME, THE  LONG, FRANK BELKNAP  P
THREE FACES OF TIME  MERWIN, SAM JR.  DD121
3 FOR SPACE [SAM SPACE]  NOLAN, WILLIAM F.  LP
THREE FOR TOMORROW  DELL BOOKS  PA,C/PA
THREE FOR TOMORROW  MEREDITH PRESS  HA
THREE FROM OUT THERE  MARGULIES, LEO  PA,C/PA
THREE FROM THE LEGION  WILLIAMSON, JACK  H:2,P
THREE GO BACK  MITCHELL, J. LESLIE  GN3.15,C/P:2
THREE GOLD CROWNS [@AVENGER #14]  ROBESON, KENNETH  C/P:2,P
THREE HAINISH NOVELS [HAINISH 1,2,3]  LE GUIN, URSULA K.  H:2
THREE HANDS FOR SCORPIO  NORTON, ANDRE  H
THREE HEARTS AND THREE LIONS  ANDERSON, POUL  H:2,P
334  DISCH, THOMAS M.  C/P:2,P
THREE IMPOSTORS, THE  MACHEN, ARTHUR  C/P,P
THREE IN ONE  MARGULIES, LEO  PA,C/PA
THREE-LEGGED HOOTCH DANCER, THE [GALACTIC MIDWAY 2]  RESNICK, MIKE  C/P,P
3 MARTIAN NOVELS [BARSOOM #4,5,6]  BURROUGHS, EDGAR RICE  H:2,LP
THREE-MINUTE UNIVERSE, THE [STAR TREK]  PAUL, BARBARA  C/P:2,P
THREE NOVELS  KNIGHT, DAMON  H:2,P
THREE NOVELS  SILVERBERG, ROBERT  P
THREE NOVELS OF THE FUTURE  WELLS, H. G.  H
THREE OF SWORDS, THE [FAFHRD 1,2,3]  LEIBER, FRITZ  H
THREE PROPHETIC NOVELS  WELLS, H. G.  H:2,LP
THREE PROPHETIC SCIENCE FICTION NOVELS OF H.G. WELLS  WELLS, H. G.  H,LP
THREE QUARTERS  DE TIMMS, GRAEME  SR-P
THREE-RING PSYCHUS  SHIRLEY, JOHN  P
THREE-SEATED SPACE SHIP, THE  SLOBODKIN, LOUIS  C/P,P,H
THREE SISTERS, THE  LOCKSLEY, REBECCA  P
THREE STIGMATA OF PALMER ELDRITCH, THE  DICK, PHILIP K.  H:2,P
THREE STORIES  DOUBLEDAY BOOKS  HA
THREE SUNS OF AMARA, THE  TEMPLE, WILLIAM F.  D76380,DF129
THREE SURVIVED  SILVERBERG, ROBERT  H
3001: THE FINAL ODYSSEY [@SPACE ODYSSEY #4]  CLARKE, ARTHUR C.  H:2,P
3001: THE FINAL ODYSSEY (PROOF) [@SPACE ODYSSEY #4]  CLARKE, ARTHUR C.  SR-LP
THREE THOUSAND YEARS  MCCLARY, THOMAS CALVERT  SR-H,P,DD176

3000 YEARS OF FANTASY AND SCIENCE FICTION DE CAMP, L. SPRAGUE/

DE CAMP, CATHERINE CROOK

HA

THREE TIME-PLAYS  PRIESTLEY, J. B.  SR-P
THREE TIMES INFINITY  MARGULIES, LEO  PA,C/PA
THREE TO CONQUER  RUSSELL, ERIC FRANK  DD215,P,C/P
THREE TO DORSAI [CHILDE CYCLE]  DICKSON, GORDON R.  H:2
3 TO THE HIGHEST POWER  NOLAN, WILLIAM F.  PA,C/PA
THREE TRIPS IN TIME AND SPACE  DELL BOOKS  PA,C/PA
THREE TRIPS IN TIME AND SPACE  HAWTHORN BOOKS  HA
THREE TRUMPS SOUNDING [@TWILIGHT OF THE GODS #3]  SCHMIDT, DENNIS  C/P,SFWA-P,P

THREE WILD MEN, THE [DOC SAVAGE #121]/ FIERY MENACE, THE [DOC SAVAGE 122] ROBESON, KENNETH

@P

THREE WORLDS OF FUTURITY  SAINT CLAIR, MARGARET  DM105
THREE WORLDS TO CONQUER  ANDERSON, POUL  C/P:2,P
THRESHOLD  DOUGLASS, SARA  H
THRESHOLD  KIERNAN, CAITLIN R.  LP
THRESHOLD  LE GUIN, URSULA K.  P
THRESHOLD  MORRIS, JANET/MORRIS, CHRIS  H:2,SFWA-P:2,P
THRESHOLD  PALMER, DAVID R.  C/P:2,P
THRESHOLD  SINGER, JUDITH  P
THRESHOLD OF ETERNITY  BRUNNER, JOHN  DD335
THRESHOLD OF THE STARS  BERNA, PAUL  SR-P
THRESHOLD SHIFT  BROWN, ERIC  H
THRICE BOUND  GELLIS, ROBERTA  P
THRICE UPON A TIME  HOGAN, JAMES P.  C/P:2,P
THRICE UPON A TIME (PROOF)  HOGAN, JAMES P.  SR-LP
THROME OF THE ERRIL OF SHERILL, THE  MCKILLIP, PATRICIA A.  C/P,P
THRONE OF BHAAL [@FORGOTTEN REALMS BALDUR’S GATE II]  KARPYSHYN, DREW  P
THRONE OF BONES, THE  MCNAUGHTON, BRIAN  H
THRONE OF FOOLS [@OMARAN #2]  COLE, ADRIAN  P
THRONE OF ISIS  TARR, JUDITH  P
THRONE OF JADE [TEMERAIRE 2]  NOVIK, NAOMI  C/P,P
THRONE OF MADNESS, THE [INQUESTOR 2]  SUCHARITKUL, SOMTOW  C/P:2,P
THRONE OF SATAN [MARK HOOD 6]  DARK, JAMES  C/P,P
THRONE OF SATURN, THE  DRURY, ALLEN  H:2
THRONE OF SCONE, THE [KELTIAD 2]  KENNEALY-MORRISON, PATRICIA  C/P,P,H

TITLE  AUTHOR  CODE
THRONES OF KRONOS, THE [EXORDIUM 5]  SMITH, SHERWOOD/TROWBRIDGE, DAVE  P
THROUGH A BRAZEN MIRROR  SHERMAN, DELIA  P
THROUGH A GLASS, CLEARLY  ASIMOV, ISAAC  SR-P,C/P
THROUGH ALIEN EYES  THOMSON, AMY  C/P,LP
THROUGH CAVERNS INFINITE  WEYRICH, BECKY LEE  SR-P
THROUGH DARKEST AMERICA  BARRETT, NEAL JR.  H
THROUGH ELEGANT EYES  LAFFERTY, R. A.  SR-H
THROUGH EYES OF WONDER  BOVA, BEN  R/REF-H
THROUGH SPACE TO MARS  ROCKWOOD, ROY  H
THROUGH THE ALIMENTARY CANAL WITH GUN AND CAMERA  CHAPPELL, GEORGE  H
THROUGH THE BARRIER  TORRO, PEL  SR-P
THROUGH THE BEHRING STRAIT  VERNE, JULES  H
THROUGH THE BREACH [REACHES 2]  DRAKE, DAVID  C/P,P,H
THROUGH THE DARK CURTAIN [GUARDIANS (SAXON) 4]  SAXON, PETER  P
THROUGH THE DARKNESS [INTO THE DARKNESS 3]  TURTLEDOVE, HARRY  C/P,P,H
THROUGH THE EARTH  FEZANDIE, CLEMENT  SR-LP
THROUGH THE EYE OF A NEEDLE  CLEMENT, HAL  SR-P,C/P:2
THROUGH THE EYE OF TIME [Q 2]  HOYLE, TREVOR  C/P,P
THROUGH THE HEART  GRANT, RICHARD  C/P:2,P
THROUGH THE ICE  ANTHONY, PIERS/KORNWISE, ROBERT  C/P:2,P
THROUGH THE LOOKING GLASS  CARROLL, LEWIS  H

THROUGH THE LOOKING GLASS [@ANDROMEDA (GENE RODDENBERRY’S) #4] SHERMAN, JOSEPHA

P

THROUGH THE REALITY WARP  PFEIL, DONALD J.  C/P,P
THROUGH VIOLET EYES [VIOLETS 1]  WOODWORTH, STEPHEN  H
THROUGH WOLF’S EYES [FIREKEEPER 1]  LINDSKOLD, JANE  P
THROWBACKS, THE  SARAC, ROGER  P
THROY [CADWAL 3]  VANCE, JACK  H
THROY (VIE) [@VANCE INTEGRAL EDITION #41/CADWAL 3]  VANCE, JACK  SR-H
THUD! [DISCWORLD 33/DISCWORLD WATCHMEN 7]  PRATCHETT, TERRY  H,P
THUMBPRINTS  SARGENT, PAMELA  H
THUNDER AND ROSES [COMPLETE STORIES OF THEODORE STURGEON 4]  STURGEON, THEODORE  H
THUNDER MOUNTAIN [@FIREBRATS #3]  SIEGEL, BARBARA/SIEGEL, SCOTT  P
THUNDER OF HELL [GUARDIANS 3]  AUSTIN, RICHARD  P
THUNDER OF STARS, A  MORGAN, DAN/KIPPAX, JOHN  C/P:2,P
THUNDER OF THE CAPTAINS [BARD’S TALE 6]  LISLE, HOLLY/ALLSTON, AARON  P
THUNDER OF TIME [FOOTPRINTS OF THUNDER 2]  DAVID, JAMES F.  H
THUNDER ON NEPTUNE, A  EKLUND, GORDON  H:2
THUNDER RIFT  FARRELL, MATTHEW  P
THUNDER ROAD  CURRY, CHRIS  P
THUNDER STRIKE!  MCCOLLUM, MICHAEL  C/P:2,P
THUNDERBALL [JAMES BOND 9]  FLEMING, IAN  C/P:2,P
THUNDERBIRD FALLS [WALKER 2]  MURPHY, C. E.  H:2,LP
THUNDERBOLT MEN, THE  MACARTHUR, DAVID S.  H
THUNDERWORLD  HUGHES, ZACH  P
THURB REVOLUTION, THE [ANTHONY VILLIERS 2]  PANSHIN, ALEXEI  SR-P,C/P:2
THUVIA, MAID OF MARS [BARSOOM #4]  BURROUGHS, EDGAR RICE  H:2,@P

THUVIA, MAID OF MARS [BARSOOM 4]/ CHESSMEN OF MARS, THE [BARSOOM 5] BURROUGHS, EDGAR RICE

H:2

THX 1138  BOVA, BEN  C/P:2,P
TICKET TO NOWHERE, A  BECHER, DON  SR-P
TICKING IS IN YOUR HEAD, THE  DAVENTRY, LEONARD  C/P,P
TIDE  HUGHES, ZACH  H,P
TIDE OF VICTORY, THE [BELISARIUS 5]  FLINT, ERIC/DRAKE, DAVID  H
TIDE WENT OUT, THE  MAINE, CHARLES ERIC  C/P:2,P
TIDES  MACKAY, SCOTT  H
TIDES OF BLOOD [@DRAGONLANCE MINOTAUR WARS #2]  KNAAK, RICHARD A.  P
TIDES OF GOD, THE  REYNOLDS, TED  C/P,P
TIDES OF KREGEN, THE [DRAY PRESCOT 12]  AKERS, ALAN BURT  P
TIDES OF LIGHT [GALACTIC CENTER 4]  BENFORD, GREGORY  C/P:2,SFWA-P,P,H
TIDES OF LUST, THE  DELANY, SAMUEL R.  SR-L
TIDES OF TIME, THE  BRUNNER, JOHN  C/P,P,H
TIDES OF TREACHERY, THE [@LEGENDS OF LONE WOLF #3]  DEVER, JOE/GRANT, JOHN  P
TIES OF BLOOD AND SILVER  ROSENBERG, JOEL  C/P:2,P
TIES OF POWER [@TRADE PACT #2]  CZERNEDA, JULIE E.  C/P,P,H
TIGANA  KAY, GUY GAVRIEL  C/P,SFWA-P,P,H:3

TIGER BURNING BRIGHT BRADLEY, MARION ZIMMER/

NORTON, ANDRE/LACKEY, MERCEDES

C/P,P,H:2

TIGER BY THE TAIL [@MAN FROM PLANET X #2]  ADAMS, HUNTER  SR-P
TIGER BY THE TAIL AND OTHER SCIENCE FICTION STORIES  NOURSE, ALAN E.  H:2,P
TIGER IN THE SKY [@OUT OF TIME #2]  FINCH, SHEILA  P
TIGER IN THE STARS  HUGHES, ZACH  SR-P
TIGER RIVER  FRIEL, ARTHUR O.  C/P,P,H
TIGERS AND TRAITORS  VERNE, JULES  H,P

TITLE  AUTHOR  CODE
TIGERS OF THE SEA  HOWARD, ROBERT E.  SR-H,C/P,P
TIGRIS LEAPS, THE [@PERRY RHODAN #63]  ACE  PA,C/PA
TIJDCAPSULE, DE  BIGGLE, LLOYD JR.  FP
TIK-TOK  SLADEK, JOHN  C/P,P
TIK-TOK OF OZ [OZ #8]  BAUM, L. FRANK  C/P,@P,H
TILL WE HAVE FACES  LEWIS, C. S.  LP
TILTANGLE  MACKELWORTH, R. W.  C/P:2,P
TIMBERJAK  DEBRANDT, DON H.  P
TIME: 110100  KELLEY, LEO P.  H,SFWA
TIME AFTER TIME  ALEXANDER, KARL  P
TIME AFTER TIME  APPEL, ALLEN  H
TIME AFTER TIME  LITTLE, DENISE  PA
TIME AND AGAIN [TIME AND AGAIN 1]  FINNEY, JACK  H:2,LP
TIME AND AGAIN  SIMAK, CLIFFORD D.  C/P,SFWA,P,H
TIME AND LIGHT  BORNEFELD, WILLIAM  P
TIME AND RELATIVE [DOCTOR WHO]  NEWMAN, KIM  SR-H
TIME AND SPACE  LE PAGE, RAND  SR-P
TIME AND STARS  ANDERSON, POUL  C/P,P,H,SR-H
TIME AND THE GODS  DUNSANY, LORD  LP
TIME AND THE RANI [DOCTOR WHO]  BAKER, JANE/BAKER, PIP  P
TIME AND TIMOTHY GRENVILLE  GREENHOUGH, TERRY  SR-P
TIME AND TOMORROW  MACDONALD, JOHN D.  H:2
TIME APPOINTED, A  STATTEN, VARGO  SR-P
TIME AT THE TOP  ORMONDROYD, EDWARD  P
TIME AXIS, THE  KUTTNER, HENRY  C/P:2,P
TIME BENDER, THE [LAFAYETTE O’LEARY 1]  LAUMER, KEITH  C/P:2,P
TIME BEYOND TIME  GREEN, I. G.  P
TIME BLENDER  DORN, MICHAEL  C/P,P
TIME BOMB  TUCKER, WILSON  H:2
TIME BOMB  ZAHN, TIMOTHY  C/P:2,SFWA-P,P
TIME BRIDGE, THE  STATTEN, VARGO  SR-P,C/P
TIME CAMERA  LEE, TERRENCE  LP
TIME CAPSULE  BERMAN, MITCH  P
TIME CONNECTION, THE  MONTELEONE, THOMAS F.  C/P,P
TIME CURVE, THE  MOSKOWITZ, SAM/ELWOOD, ROGER  PA,C/PA
TIME DISSOLVER, THE  SOHL, JERRY  C/P:2,P
TIME DWELLER, THE  MOORCOCK, MICHAEL  C/P:2,P
TIME ECHO  LIONEL, ROBERT  C/P:2,P
TIME ENOUGH FOR LOVE  HEINLEIN, ROBERT A.  H:2,P
TIME EXILE  MILLERS, REINHOLD  H
TIME FAMINE  OLSEN, LANCE  LP
TIME FIGHTERS, THE [@SPACE: 1999 YEAR 2 #5]  BUTTERWORTH, MICHAEL  P
TIME-FLIGHT [DOCTOR WHO]  GRIMWADE, PETER  SR-P,C/P
TIME FOR DRAGONS, A [DRAGON WARS 1]  GENTILE, GARY  C/P:2,P
TIME FOR THE STARS  HEINLEIN, ROBERT A.  H:2,P
TIME FOR WAR, A [@TIMESHARE 3]  DANN, JOSHUA  P
TIME FOR WAR, A TIME FOR PEACE, A [STAR TREK TNG TIME TO 9]  DECANDIDO, KEITH R. A.  P
TIME FOR YESTERDAY [STAR TREK]  CRISPIN, A. C.  C/P:2,P
TIME FUTURE [COMMANDER HALLEY 1]  MCARTHUR, MAXINE  P
TIME GARDEN, THE  EAGER, EDWARD  P
TIME GATE  JAKES, JOHN  C/P,P,H
TIME GATE [@TIME GATE #1]  SILVERBERG, ROBERT/FAWCETT, BILL  PA,C/PA
TIME GLADIATOR  REYNOLDS, MACK  C/P,P
TIME HOPPERS, THE  SILVERBERG, ROBERT  H:2,P
TIME IN ADVANCE  TENN, WILLIAM  C/P:2,P
TIME IN ECLIPSE  GARNETT, DAVID S.  SR-P
TIME IN MIND  STARBUCK, KATHLYN S.  P
TIME IS A SHADOW  BEST, STAN  P
TIME IS THE SIMPLEST THING  SIMAK, CLIFFORD D.  H:2,SFWA,P
TIME-JUMP  BRUNNER, JOHN  C/P:2,P
TIME KEEPER, THE [TIME KEEPER 1]  BARTHOLOMEW, BARBARA  P
TIME LIKE AN EVER-ROLLING STREAM  MOFFETT, JUDITH  H
TIME LOTTERY [TIME LOTTERY 1]  MOSER, NANCY  LP
TIME MACHINE, THE  WELLS, H. G.  C/P:2,P
TIME MACHINE, THE/WAR OF THE WORLDS, THE  WELLS, H. G.  P
TIME MACHINE AND OTHER STORIES, THE  WELLS, H. G.  P
TIME MACHINE II  PAL, GEORGE/MORHAIM, JOE  C/P,P
TIME MACHINES  ADLER, BILL JR.  HA,LPA
TIME MACHINES, THE  ASHLEY, MIKE  R/REF-H
TIME MASTERS, THE (ORIGINAL)  TUCKER, WILSON  C/P:2,P
TIME MASTERS, THE (REVISED)  TUCKER, WILSON  H:2,P
TIME MEDDLER, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,P
TIME MERCENARIES, THE  HIGH, PHILIP E.  DH059
TIME MUST HAVE A STOP  HUXLEY, ALDOUS  H

The MITSFS Titledex  September 24, 2007  Page 372
TITLE  AUTHOR  CODE
TIME NOMADS [@DEATHLANDS]  AXLER, JAMES  P
TIME OF CHANGES, A  SILVERBERG, ROBERT  H:2,P
TIME OF DYING, A [@RAVEN #5]  KIRK, RICHARD  SR-P,C/P
TIME OF EXILE, A [DEVERRY 5/WESTLANDS 1]  KERR, KATHARINE  C/P,P,H
TIME OF GHOSTS, A [@RAVEN 2]  KIRK, RICHARD  SR-P
TIME OF JUSTICE, A [DEVERRY 7/WESTLANDS 3]  KERR, KATHARINE  P
TIME OF MADNESS, THE [@MAGELORD 2]  MARTIN, THOMAS K.  P
TIME OF OMENS, A [DEVERRY 6/WESTLANDS 2]  KERR, KATHARINE  H:2,P

TIME OF PASSAGE OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

HA

TIME OF THE ANNIHILATOR, THE  MORRESSY, JOHN  P
TIME OF THE DARK, THE [DARWATH 1]  HAMBLY, BARBARA  C/P:2,P
TIME OF THE FOURTH HORSEMAN  YARBRO, CHELSEA QUINN  C/P,P,H,SFWA
TIME OF THE FOX [TIME WARRIOR 1]  COSTELLO, MATTHEW J.  C/P,P
TIME OF THE GHOST, THE  JONES, DIANA WYNNE  C/P,P
TIME OF THE GREAT FREEZE  SILVERBERG, ROBERT  C/P:2,SFWA-P,P

TIME OF THE HAWKLORDS, THE MOORCOCK, MICHAEL/

BUTTERWORTH, MICHAEL

C/P:2,P

TIME OF THE HEDRONS, THE  ECKSTROM, JACK DENNIS  H
TIME OF THE TRANSFERENCE, THE [SPELLSINGER 6]  FOSTER, ALAN DEAN  C/P,P,H
TIME OF THE TWINS [@DRAGONLANCE LEGENDS #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
TIME OF THE VAMPIRES, THE  ELROD, P. N./GREENBERG, MARTIN H.  C/PA,PA
TIME OF WAR, A [DEVERRY]  KERR, KATHARINE  P
TIME OF YOUR LIFE [DOCTOR WHO MISSING ADVENTURES]  LYONS, STEVE  C/P:2,P
TIME ON MY HANDS  DELACORTE, PETER  H
TIME OUT  ELY, DAVID  H
TIME OUT FOR TOMORROW  WILSON, RICHARD  C/P:2,P
TIME OUT OF JOINT  DICK, PHILIP K.  C/P:2,P
TIME OUT OF MIND  BOULLE, PIERRE  C/P,P
TIME OUT OF MIND  COWPER, RICHARD  SR-P
TIME OUT OF MIND  LORD, GRAHAM  P
TIME OUT OF MIND  MAXIM, JOHN R.  C/P,SFWA-P,P,H
TIME PAST [COMMANDER HALLEY 2]  MCARTHUR, MAXINE  P
TIME PATROL, THE  ANDERSON, POUL  H:2
TIME PATROLMAN  ANDERSON, POUL  C/P:2,P
TIME PIECES  BISHOP, MICHAEL  LP
TIME PIPER  HUDDY, DELIA  P
TIME PRESSURE [MINDKILLER 2]  ROBINSON, SPIDER  C/P,SFWA-P,H,P,SFWA
TIME PROBE  CLARKE, ARTHUR C.  PA,C/PA
TIME ROGUE  KELLEY, LEO P.  P
TIME SAFARI  DRAKE, DAVID  P,SFWA-P
TIME SCOUT [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  C/P:2,P
TIME SHIFTERS, THE  MERWIN, SAM JR.  P
TIME SHIPS, THE  BAXTER, STEPHEN  C/P:2,P
TIME SLAVE  NORMAN, JOHN  C/P:2,P
TIME-SLIP  MARTIN, GRAHAM DUNSTAN  P
TIME SNAKE AND SUPERCLOWN  KING, VINCENT  SR-P
TIME SNATCHERS  HINES, LEONARD F.  LP
TIME SPIRAL [@MAGIC: THE GATHERING TIME SPIRAL #1]  MCGOUGH, SCOTT  P
TIME STATION BERLIN [TIME STATION 3]  EVANS, DAVID  P
TIME STATION LONDON [TIME STATION 1]  EVANS, DAVID  C/P,P
TIME STATION PARIS [TIME STATION 2]  EVANS, DAVID  P
TIME STORM  DICKSON, GORDON R.  SFWA-P,H:2,P
TIME STORY  GORDON, STUART  C/P:2,P
TIME STREAM, THE  TAINE, JOHN  SR-H
TIME STREAM, THE/GREATEST ADVENTURE, THE/PURPLE SAPPHIRE, THE  TAINE, JOHN  H:2,LP
TIME STREAMS [@MAGIC: THE GATHERING ARTIFACTS 3]  KING, J. ROBERT  P
TIME-SWEPT CITY, THE  MONTELEONE, THOMAS F.  P
TIME: THE SEMI-FINAL FRONTIER  FENN, LIONEL  P
TIME THIEVES  KOONTZ, DEAN R.  D00990
TIME TO BE BORN, A [@STAR TREK TNG TIME TO 1]  VORNHOLT, JOHN  P
TIME TO COME  DERLETH, AUGUST  PA,C/PA
TIME TO DIE, A [@STAR TREK TNG TIME TO 2]  VORNHOLT, JOHN  P
TIME TO HARVEST, A [@STAR TREK TNG TIME TO 4]  WARD, DAYTON/DILMORE, KEVIN  P
TIME TO HATE, A [@STAR TREK TNG TIME TO 6]  GREENBERGER, ROBERT  P
TIME TO HEAL, A [STAR TREK TNG TIME TO 8]  MACK, DAVID  P
TIME TO KILL, A [@STAR TREK TNG TIME TO 7]  MACK, DAVID  P
TIME TO LIVE  RACKHAM, JOHN  DG606
TIME TO LOVE, A [@STAR TREK TNG TIME TO 5]  GREENBERGER, ROBERT  P
TIME TO SOW, A [@STAR TREK TNG TIME TO 3]  WARD, DAYTON/DILMORE, KEVIN  P
TIME TO TELEPORT  DICKSON, GORDON R.  C/P,DD449
TIME TO TELEPORT/DELUSION WORLD  DICKSON, GORDON R.  P
TIME TO TIME [ASHTON FORD 6]  PENDLETON, DON  P
TIME TRADERS, THE [TIME TRADERS 1]  NORTON, ANDRE  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 373
TITLE  AUTHOR  CODE

TIME TRADERS (REVISED) [TIME TRADERS 1,2]/ GALACTIC DERELICT (REVISED) [TIME TRADERS 1,2] NORTON, ANDRE

P

TIME TRADERS III [TIME TRADERS 3,4]  NORTON, ANDRE/SMITH, SHERWOOD  H
TIME TRANSFER  SELLINGS, ARTHUR  SR-P
TIME TRAP [TIME TRAP 1]  DALTON, SEAN  P
TIME TRAP, THE  KELLEY, LEO P.  SR-P
TIME TRAP  LAUMER, KEITH  SR-H,C/P:2,P
TIME TRAP  PHILLIPS, ROG  SR-P
TIME TRAP, THE  STATTEN, VARGO  SR-P
TIME TRAP GAMBIT, THE [@AGENT OF TERRA #4]  MADDOCK, LARRY  C/P:2,P
TIME TRAP OF MING XIII, THE [FLASH GORDON 4]  STEFFANSON, CON  C/P,P
TIME TRAVELERS  DOZOIS, GARDNER  PA

TIME TRAVELERS, THE SILVERBERG, ROBERT/

GREENBERG, MARTIN H.

C/PA,PA

TIME TRAVELERS ACADEMY, THE  WILLIAMS, REGINALD  LP
TIME TRAVELERS STRICTLY CASH [CALLAHAN’S CROSSTIME SALOON]  ROBINSON, SPIDER  C/P:2,P
TIME TRAVELER’S WIFE, THE  NIFFENEGGER, AUDREY  H,LP
TIME TRAVELLERS, THE [DOCTOR WHO PAST DOCTORS]  GUERRIER, SIMON  P
TIME TRIP, THE  SWIGART, ROB  H
TIME TUNNEL  LEINSTER, MURRAY  C/P,P
TIME TUNNEL, THE [TIME TUNNEL 1]  LEINSTER, MURRAY  C/P:2,P
TIME TWISTER [@ALIEN TRACE #2]  MAJOR, H. M.  P
TIME TWISTER, THE  PETAJA, EMIL  C/P:2,P
TIME TWISTERS, THE  HOLLY, JOAN HUNTER  C/P,P
TIME TWISTERS  RABE, JEAN/GREENBERG, MARTIN H.  PA
TIME WANDERERS, THE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H:2,P
TIME WAR CARTER, LIN C/P,P

TIME WARS WAUGH, CHARLES G./

GREENBERG, MARTIN H.

C/PA,PA,SFWA-P

TIME WHEN, A  MCCAFFREY, ANNE  SR-H
TIME: X  TUCKER, WILSON  C/P:2,P
TIME ZERO [DOCTOR WHO EIGHTH DOCTOR]  RICHARDS, JUSTIN  P
TIMEDIVER’S DAWN [TIMEDIVERS 2]  MODESITT, L. E. JR.  C/P:2,P
TIMEFALL  KAHN, JAMES  H,P
TIMEGATES  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
TIMEGOD, THE [TIMEDIVERS 1]  MODESITT, L. E. JR.  P
TIMEGODS’ WORLD [TIMEDIVERS 1,2]  MODESITT, L. E. JR.  H
TIMEJUMPER  GREENLEAF, WILLIAM  P
TIMEKEEPER CONSPIRACY, THE [TIMEWARS 2]  HAWKE, SIMON  C/P:2,P
TIMELASH [DOCTOR WHO]  MCCOY, GLEN  SR-P,C/P
TIMELESS [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
TIMELESS LOVE  AVON BOOKS  SFWA-P
TIMELESS PASSION  O’DAY-FLANNERY, CONSTANCE  P
TIMELESS STORIES FOR TODAY AND TOMORROW  BRADBURY, RAY  SR-PA,C/PA
TIMELIKE DIPLOMACY [SINGULARITY SKY 1,2]  STROSS, CHARLES  H
TIMELIKE INFINITY [XEELEE]  BAXTER, STEPHEN  C/P,P
TIMELINE  CRICHTON, MICHAEL  H:2,P
TIMELINER  MAINE, CHARLES ERIC  H:2,P

TIMELOST GIARRATANO, JOSEPH C./

ANDREWS, ARLAN KEITH/ ANDREWS, KRIS AUSTEN

SCX

TIMEMASTER  FORWARD, ROBERT L.  C/P,P,H
TIMEPIECE  BALL, BRIAN N.  C/P:2,P
TIMEPIVOT  BALL, BRIAN N.  C/P:2,P
TIMEQUAKE  VONNEGUT, KURT JR.  LP
TIMEQUEST  NELSON, RAY FARADAY  P,SFWA-P
TIME’S CHILD  ORE, REBECCA  LP
TIME’S CHILD (PROOF)  ORE, REBECCA  SR-LP
TIME’S DARK LAUGHTER  KAHN, JAMES  C/P:2,P
TIME’S ENEMY [STAR TREK INVASION! 3/STAR TREK DS9]  GRAF, L. A.  C/P,P
TIME’S EYE (CD) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  SR-MM
TIME’S EYE (HARDCOVER AND CD) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  H
TIME’S EYE (PROOF) [TIME ODYSSEY 1]  CLARKE, ARTHUR C./BAXTER, STEPHEN  SR-LP
TIME’S FOOL  CARRINGTON, GRANT  H,SFWA
TIME’S LAST GIFT  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
TIME’S LONELY ONE [@PERRY RHODAN #42]  ACE  PA,C/PA
TIMES OF TROUBLE [@TERMINATOR 2 NEW JOHN CONNOR 3]  BLACKFORD, RUSSELL  P
TIMES WITHOUT NUMBER  BRUNNER, JOHN  DF161
TIMES WITHOUT NUMBER (REVISED)  BRUNNER, JOHN  C/P:2,P
TIMESCALE  CALDER, NIGEL  [C/RAN-H]
TIMESCAPE  BENFORD, GREGORY  C/P:2,P,H
TIMESCOOP  BRUNNER, JOHN  C/P:2,P
TIMESERVERS, THE  GRIFFIN, RUSSELL M.  P
TIMESHADOW RIDER  MAXWELL, ANN  C/P:2,SFWA-P,P

The MITSFS Titledex  September 24, 2007  Page 374
TITLE  AUTHOR  CODE
TIMESHARE [@TIMESHARE 1]  DANN, JOSHUA  P
TIMESLIDE [USS SKIPJACK 2]  LANG, SIMON  P
TIMESLIP! [TIME TUNNEL 2]  LEINSTER, MURRAY  C/P:2,P
TIMESPELL [AELWYN 1]  CHARRETTE, ROBERT N.  P
TIMESTOP!  FARMER, PHILIP JOSE  C/P,P
TIMESWEPT LOVERS  O’DAY-FLANNERY, CONSTANCE  P
TIMETIPPING  DANN, JACK  H,SFWA
TIMETRACKS  LAUMER, KEITH  C/P:2,P
TIMETRAP [STAR TREK]  DVORKIN, DAVID  C/P:2,P
TIMEVIEW [DOCTOR WHO]  BELLAMY, FRANK/BELLAMY, DAVID  A
TIMEWYRM: APOCALYPSE [DOCTOR WHO NEW ADVENTURES]  ROBINSON, NIGEL  C/P,P
TIMEWYRM: EXODUS [DOCTOR WHO NEW ADVENTURES]  DICKS, TERRANCE  C/P,P
TIMEWYRM: GENESYS [DOCTOR WHO NEW ADVENTURES]  PEEL, JOHN  C/P:2,P
TIMEWYRM: REVELATION [DOCTOR WHO NEW ADVENTURES]  CORNELL, PAUL  C/P:2,P
TIN ANGEL, THE  GOULART, RON  C/P:2,P
TIN MEN, THE  FRAYN, MICHAEL  H:2,P
TIN SOLDIER  VINGE, JOAN D.  TD23

TIN STARS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

TIN WOODMAN  BISCHOFF, DAVID F./BAILEY, DENNIS R.  C/P,SFWA,P,H
TIN WOODMAN OF OZ, THE [OZ #12]  BAUM, L. FRANK  H,@P,@C/P
TINKER [TINKER 1]  SPENCER, WEN  P
TINTAGEL  COOK, PAUL  C/P,P
TINTED VENUS: A FARCICAL ROMANCE, THE  ANSTEY, F.  H
TINY ALICE  ALBEE, EDWARD  P
TITAN  BAXTER, STEPHEN  C/P:2,P
TITAN  BOVA, BEN  H
TITAN, THE  MILLER, P. SCHUYLER  H:2
TITAN [GAEA 1]  VARLEY, JOHN  H:2,P
TITAN A.E.  PERRY, STEVE/PERRY, DAL  P
TITAN A.E.: THE JUNIOR NOVELIZATION  DALY, C. R.  LP
TITAN OF TWILIGHT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #3]  DENNING, TROY  P
TITAN’S DAUGHTER  BLISH, JAMES  C/P:2,P
TITAN’S MOON  CHARLES, NEIL  SR-P
TITANS OF CHAOS [ORPHANS OF CHAOS 3]  WRIGHT, JOHN C.  H
TITANS OF THE UNIVERSE  HARVEY, JAMES  P
TITUS ALONE [GORMENGHAST 3]  PEAKE, MERVYN  C/P,P,H
TITUS CROW VOLUME THREE [TITUS CROW 5,6]  LUMLEY, BRIAN  LP
TITUS GROAN [GORMENGHAST 1]  PEAKE, MERVYN  C/P,P,H

T’NT: TELZEY & TRIGGER [TELZEY] SCHMITZ, JAMES H./FLINT, ERIC/

GORDON, GUY

C/P,P

TO A HIGHLAND NATION  ROWLEY, CHRISTOPHER  C/P,P
TO ARKON! [@PERRY RHODAN #30]  ACE  PA,C/PA
TO BATTLE THE GODS [JALAV 5]  GREEN, SHARON  P
TO BE CONTINUED  SILVERBERG, ROBERT  H
TO BIGOTRY, NO SANCTION  DURST, MOSE  [C/RAN-H]
TO CHALLENGE CHAOS [CHALLENGE CHAOS 1]  STABLEFORD, BRIAN M.  C/P:2,P
TO CHARLES FORT, WITH LOVE  KIERNAN, CAITLIN R.  H
TO CONQUER CHAOS  BRUNNER, JOHN  C/P:2,P
TO CONQUER THE THRONE  SULLIVAN, TIM  P,SFWA-P
TO CONTROL THE STARS  HOSKINS, ROBERT  C/P:2,P
TO CRUSH THE MOON [QUEENDOM OF SOL 4]  MCCARTHY, WIL  P
TO DEMONS BOUND [SWORDS OF RAEMLLYN 1]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P,P
TO DIE IN ITALBAR  ZELAZNY, ROGER  H:2,P,SFWA
TO DREAM IN THE CITY OF SORROWS [@BABYLON 5 #9]  DRENNAN, KATHRYN M.  C/P,P

TO DREAM OF DREAMERS LOST [@WORLD OF DARKNESS GRAILS COVENANT #3] WILSON, DAVID NIALL

P

TO DREAM OF EVIL...  COMER, RALPH  P
TO ESCAPE THE STARS  HOSKINS, ROBERT  C/P:2,P
2XS [SHADOWRUN 4]  FINDLEY, NIGEL D.  C/P:2,P
TO FALL LIKE STARS  ASIRE, NANCY  P
TO FEAR THE LIGHT  BOVA, BEN/AUSTIN, A. J.  C/P,P
TO FOLLOW A STAR: NINE SCIENCE FICTION STORIES ABOUT CHRISTMAS  CARR, TERRY  HA
TO FREE ATLANTIS [FANTASTIC FOUR]  COLLINS, NANCY A.  P
TO GREEN ANGEL TOWER [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  H:2,SFWA:2
TO GREEN ANGEL TOWER (PART ONE) [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  C/P,P
TO GREEN ANGEL TOWER (PART TWO) [MEMORY,SORROW AND THORN 3]  WILLIAMS, TAD  C/P:2,P
TO HERE AND THE EASEL  STURGEON, THEODORE  C/P,P
TO HOLD INFINITY  MEANEY, JOHN  P
TO KEEP THE SHIP [GRIMES]  CHANDLER, A. BERTRAM  P
TO KILL A SHADOW [@TRAVELER #4]  DRUMM, D. B.  P
TO LIGHT A CANDLE [OBSIDIAN 2]  LACKEY, MERCEDES/MALLORY, JAMES  H
TO LIVE AGAIN  SILVERBERG, ROBERT  H:2,P
TO LIVE FOREVER  VANCE, JACK  C/P:2,P

TITLE  AUTHOR  CODE
TO LUNA WITH LOVE  BEVIS, H. U.  H
TO MAKE DEATH LOVE US  FALCONER, SOVEREIGN  H,SFWA
TO MARRY MEDUSA  STURGEON, THEODORE  H,SFWA-P,P
TO OPEN THE SKY  SILVERBERG, ROBERT  C/P:2,P
TO OUTLIVE ETERNITY  ANDERSON, POUL  LP
TO OUTRUN DOOMSDAY  BULMER, KENNETH  C/P:2,P
TO PRIME THE PUMP [GRIMES]  CHANDLER, A. BERTRAM  C/P:2,P
TO REIGN IN HELL  BRUST, STEVEN  SR-H,C/P:3,P
TO REIGN IN HELL [STAR TREK]  COX, GREG  P
TO RENEW THE AGES  COULSON, ROBERT  SR-P
TO RIDE A RATHORN [JAMETHIEL DREAM-WEAVER 4]  HODGELL, P. C.  H
TO RIDE HELL’S CHASM  WURTS, JANNY  H
TO RIDE PEGASUS [PEGASUS 1]  MCCAFFREY, ANNE  C/P:2,P
TO RULE IN AMBER [ROGER ZELAZNY’S DAWN OF AMBER 3]  BETANCOURT, JOHN GREGORY  H
TO SAIL BEYOND THE SUNSET  HEINLEIN, ROBERT A.  C/P,SFWA-P,P,H
TO SAIL THE CENTURY SEA [TIME STREAM 2]  EDMONDSON, G. C.  C/P:2,P
TO SAVE A WORLD [DARKOVER]  BRADLEY, MARION ZIMMER  P
TO SAVE THE SUN  BOVA, BEN/AUSTIN, A. J.  C/P:2,P
TO SAY NOTHING OF THE DOG [DUNWORTHY & FINCH 2]  WILLIS, CONNIE  H:2,P
TO SERVE AND SUBMIT  WRIGHT, SUSAN  P

TO SIFT THROUGH BITTER ASHES [@WORLD OF DARKNESS GRAILS COVENANT #1] WILSON, DAVID NIALL

P

TO SLEEP WITH EVIL [RAVENLOFT]  CARDARELLE, ANDRIA  P

TO SPEAK IN LIFELESS TONGUES [@WORLD OF DARKNESS GRAILS COVENANT #2] WILSON, DAVID NIALL

P

TO STAND BENEATH THE SUN  STRICKLAND, BRAD  P
TO STORM HEAVEN [STAR TREK TNG]  FRIESNER, ESTHER M.  C/P,P
TO THE DARK TOWER  LONG, LYDA BELKNAP  P
TO THE DEVIL A DAUGHTER  WHEATLEY, DENNIS  C/P,P
TO THE END OF TIME AND OTHER STORIES  WILLIAMS, ROBERT MOORE  DD427
TO THE HAUNTED MOUNTAINS  EMERSON, RU  P
TO THE HIGH REDOUBT  YARBRO, CHELSEA QUINN  C/P,P
TO THE KING A DAUGHTER [OAK,YEW,ASH,AND ROWAN 1]  NORTON, ANDRE/MILLER, SASHA  H:2
TO THE LAND OF THE ELECTRIC ANGEL  ROTSLER, WILLIAM  P
TO THE LAND OF THE LIVING [GILGAMESH]  SILVERBERG, ROBERT  C/P,P
TO THE MAGIC BORN [WARLOCK 0,1]  STASHEFF, CHRISTOPHER  H:2
TO THE RESURRECTION STATION  ARNASON, ELEANOR  P
TO THE SLAUGHTER [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
TO THE STARS [TO THE STARS 1,2,3]  HARRISON, HARRY  H:2,P
TO THE STARS  SILVERBERG, ROBERT  HA
TO THE STARS  TAKEI, GEORGE  R/REF-H,C/REF-H

TO THE SUN? (ROTH TRANSLATION)/ OFF ON A COMET! (ROTH TRANSLATION)

VERNE, JULES

LP

TO THE TOMBAUGH STATION  TUCKER, WILSON  DD479
TO THE ULTIMATE  STATTEN, VARGO  SR-P
TO THE VANISHING POINT  FOSTER, ALAN DEAN  H:2,P
TO TRADE THE STARS [@TRADE PACT #3]  CZERNEDA, JULIE E.  P
TO VENUS! TO VENUS!  GRINNELL, DAVID  D81610
TO VISIT THE QUEEN [FELINE WIZARDS 2]  DUANE, DIANE  C/P,LP
TO WALK THE NIGHT  SLOANE, WILLIAM  C/P,P,H
TO WARM THE EARTH  BELDEN, DAVID  C/P,P
TO WATERS’ END [@GAMMALAW 4]  DALEY, BRIAN  C/P,P

TO WEAVE A WEB OF MAGIC DELACROIX, CLAIRE/KURLAND, LYNN/ MCKILLIP, PATRICIA A./SHINN, SHARON

LP

TO YITH AND BEYOND  RIMEL, DUANE W.  SR-VLH
TO YOUR SCATTERED BODIES GO [RIVERWORLD 1]  FARMER, PHILIP JOSE  C/P,P,H
TOAST  STROSS, CHARLES  H
TOCHTER DES ROBOTERS, DIE [@PERRY RHODAN #6]  MAHR, KURT  FPR
TODAY AND TOMORROW AND ...  ASIMOV, ISAAC  SFWA
TODAY WE CHOOSE FACES  ZELAZNY, ROGER  C/P:2,P
TODAY WE CHOOSE FACES/BRIDGE OF ASHES  ZELAZNY, ROGER  C/P,P
TOKEN OF DRAGONSBLOOD  COLE, DEMARIS  C/P,P
TOLKIEN: A LOOK BEHIND THE LORD OF THE RINGS  CARTER, LIN  R/REF-P,C/REF-P
TOLKIEN BESTIARY, A  DAY, DAVID  R/REF-H
TOLKIEN COMPANION, THE  TYLER, J. E. A.  R/REF-H
TOLKIEN COMPASS, A  LOBDELL, JARED  R/REF-P
TOLKIEN CRITICISM: AN ANNOTATED CHECKLIST  WEST, RICHARD C.  R/IND
TOLKIEN MISCELLANY, A  TOLKIEN, J. R. R.  H
TOLKIEN QUIZ BOOK, THE  ANDREWS, BART  R/REF-P
TOLKIEN READER, THE  TOLKIEN, J. R. R.  C/P:2,P
TOLKIEN RELATION, THE  READY, WILLIAM  R/REF-H
TOLKIEN SCRAPBOOK, THE  BECKER, ALIDA  R/REF-H
TOLKIEN: THE ILLUSTRATED ENCYCLOPAEDIA  DAY, DAVID  R/REF-H
TOLKIEN’S WORLD  HELMS, RANDEL  R/REF-H

TITLE  AUTHOR  CODE
TOM CORBETT: A TRIP TO THE MOON [TOM CORBETT]  MARTIN, MARCIA  BLB
TOM O’BEDLAM  SILVERBERG, ROBERT  C/P,SFWA,P,H
TOM O’BEDLAM’S NIGHT OUT  SCHWEITZER, DARRELL  H
TOM PAINE MARU [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P:2,P
TOM SWIFT AMONG THE DIAMOND MAKERS [@TOM SWIFT #7]  APPLETON, VICTOR  H
TOM SWIFT AMONG THE FIRE FIGHTERS [@TOM SWIFT #24]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AERIAL WARSHIP [@TOM SWIFT #18]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AIR GLIDER [@TOM SWIFT #12]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AIR SCOUT [@TOM SWIFT #22]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AIRLINE EXPRESS [@TOM SWIFT #29]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS AIRSHIP [@TOM SWIFT #3]  APPLETON, VICTOR  H
TOM SWIFT AND HIS AQUATOMIC TRACKER [@TOM SWIFT JR. #23]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS ATOMIC EARTH BLASTER [@TOM SWIFT JR. #5]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS BIG DIRIGIBLE [@TOM SWIFT #33]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS BIG TUNNEL [@TOM SWIFT #19]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS CHEST OF SECRETS [@TOM SWIFT #28]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS COSMOTRON EXPRESS [@TOM SWIFT JR. #32]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS DEEP-SEA HYDRODOME [@TOM SWIFT JR. #11]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS DIVING SEACOPTER [@TOM SWIFT JR. #7]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS DYNA-4 CAPSULE [@TOM SWIFT JR. #31]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS ELECTRIC LOCOMOTIVE [@TOM SWIFT #25]  APPLETON, VICTOR  H
TOM SWIFT AND HIS ELECTRIC RIFLE [@TOM SWIFT #10]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS ELECTRIC RUNABOUT [@TOM SWIFT #5]  APPLETON, VICTOR  H
TOM SWIFT AND HIS ELECTRONIC RETROSCOPE [@TOM SWIFT JR. #14]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS FLYING BOAT [@TOM SWIFT #26]  APPLETON, VICTOR  H
TOM SWIFT AND HIS FLYING LAB [@TOM SWIFT JR. #1]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS G-FORCE INVERTER [@TOM SWIFT JR. #30]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS GIANT CANNON [@TOM SWIFT #16]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GIANT MAGNET [@TOM SWIFT #35]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GIANT ROBOT [@TOM SWIFT JR. #4]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS GIANT TELESCOPE  APPLETON, VICTOR  BLB
TOM SWIFT AND HIS GREAT OIL GUSHER [@TOM SWIFT #27]  APPLETON, VICTOR  H
TOM SWIFT AND HIS GREAT SEARCHLIGHT [@TOM SWIFT #15]  APPLETON, VICTOR  H
TOM SWIFT AND HIS HOUSE ON WHEELS [@TOM SWIFT #32]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS JETMARINE [@TOM SWIFT JR. #2]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS MAGNETIC SILENCER  APPLETON, VICTOR  BLB
TOM SWIFT AND HIS MEGASCOPE SPACE PROBER [@TOM SWIFT JR. #20]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS MOTOR-BOAT [@TOM SWIFT #2]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS MOTOR-CYCLE [@TOM SWIFT #1]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS OCEAN AIRPORT [@TOM SWIFT #37]  APPLETON, VICTOR  H
TOM SWIFT AND HIS OUTPOST IN SPACE [@TOM SWIFT JR. #6]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS PHOTO TELEPHONE [@TOM SWIFT #17]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS PLANET STONE [@TOM SWIFT #38]  APPLETON, VICTOR  H
TOM SWIFT AND HIS POLAR-RAY DYNASPHERE [@TOM SWIFT JR. #25]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS REPELATRON SKYWAY [@TOM SWIFT JR. #22]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ROCKET SHIP [@TOM SWIFT JR. #3]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS SKY RACER [@TOM SWIFT #9]  APPLETON, VICTOR  H
TOM SWIFT AND HIS SKY TRAIN [@TOM SWIFT #34]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS SONIC BOOM TRAP [@TOM SWIFT JR. #26]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS SPACE SOLARTRON [@TOM SWIFT JR. #13]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS SPECTROMARINE SELECTOR [@TOM SWIFT JR. #15]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS SUBMARINE BOAT [@TOM SWIFT #4]  APPLETON, VICTOR  H
TOM SWIFT AND HIS SUBOCEAN GEOTRON [@TOM SWIFT JR. #27]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS TALKING PICTURES [@TOM SWIFT #31]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS TELEVISION DETECTOR [@TOM SWIFT #36]  APPLETON, VICTOR  H:2
TOM SWIFT AND HIS 3-D TELEJECTOR [@TOM SWIFT JR. #24]  APPLETON, VICTOR II  H
TOM SWIFT AND HIS TRIPHIBIAN ATOMICAR [@TOM SWIFT JR. #19]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS ULTRASONIC CYCLOPLANE [@TOM SWIFT JR. #10]  APPLETON, VICTOR II  H:2
TOM SWIFT AND HIS UNDERSEA SEARCH [@TOM SWIFT #23]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WAR TANK [@TOM SWIFT #21]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WIRELESS MESSAGE [@TOM SWIFT #6]  APPLETON, VICTOR  H
TOM SWIFT AND HIS WIZARD CAMERA [@TOM SWIFT #14]  APPLETON, VICTOR  H
TOM SWIFT AND THE ASTEROID PIRATES [@TOM SWIFT JR. #21]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE CAPTIVE PLANETOID [@TOM SWIFT JR. #29]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE COSMIC ASTRONAUTS [@TOM SWIFT JR. #16]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE ELECTRONIC HYDROLUNG [@TOM SWIFT JR. #18]  APPLETON, VICTOR II  H:2
TOM SWIFT AND THE GALAXY GHOSTS [@TOM SWIFT JR. #33]  APPLETON, VICTOR II  H
TOM SWIFT AND THE MYSTERY COMET [@TOM SWIFT JR. #28]  APPLETON, VICTOR II  H
TOM SWIFT AND THE VISITOR FROM PLANET X [@TOM SWIFT JR. #17]  APPLETON, VICTOR II  H:2
TOM SWIFT CIRCLING THE GLOBE [@TOM SWIFT #30]  APPLETON, VICTOR  H
TOM SWIFT IN CAPTIVITY [@TOM SWIFT #13]  APPLETON, VICTOR  H
TOM SWIFT IN THE CAVES OF ICE [@TOM SWIFT #8]  APPLETON, VICTOR  H
TOM SWIFT IN THE CAVES OF NUCLEAR FIRE [@TOM SWIFT JR. #8]  APPLETON, VICTOR II  H:2

The MITSFS Titledex  September 24, 2007  Page 377
TITLE  AUTHOR  CODE
TOM SWIFT IN THE CITY OF GOLD [@TOM SWIFT #11]  APPLETON, VICTOR  H
TOM SWIFT IN THE LAND OF WONDERS [@TOM SWIFT #20]  APPLETON, VICTOR  H
TOM SWIFT IN THE RACE TO THE MOON [@TOM SWIFT JR. #12]  APPLETON, VICTOR II  H:2
TOM SWIFT ON THE PHANTOM SATELLITE [@TOM SWIFT JR. #9]  APPLETON, VICTOR II  H:2
TOMB, THE [ADVERSARY 2/REPAIRMAN JACK]  WILSON, F. PAUL  C/P:2,P
TOMB AND OTHER TALES, THE  LOVECRAFT, H. P.  C/P:2,P
TOMB OF HORRORS, THE [@GREYHAWK]  STROHM, KEITH FRANCIS  P
TOMB OF VALDEMAR [DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  P
TOMBS  KRAMER, EDWARD E./CROWTHER, PETER  PA
TOMBS OF ATUAN, THE [EARTHSEA 2]  LE GUIN, URSULA K.  C/P,P,H
TOMBS OF KOBOL, THE [@BATTLESTAR GALACTICA 1978 #3]  LARSON, GLEN A./THURSTON, ROBERT  P
TOMMYKNOCKERS, THE  KING, STEPHEN  H:2,P
TOMOE GOZEN [TOMOE GOZEN 1]  SALMONSON, JESSICA AMANDA  C/P:2,P
TOMORROW  ELWOOD, ROGER  SR-HA
TOMORROW  TUBB, E. C.  GRD18
TOMORROW!  WYLIE, PHILIP  C/P,P,H
TOMORROW 1  HOSKINS, ROBERT  PA,C/PA
TOMORROW AND BEYOND  SUMMERS, IAN  A
TOMORROW AND TOMORROW  COLLINS, HUNT  C/P:2,P
TOMORROW AND TOMORROW  SHEFFIELD, CHARLES  C/P,P,H
TOMORROW BITES  COX, GREG/WEISSKOPF, T. K. F.  C/PA,PA
TOMORROW CAME  COOPER, EDMUND  P
TOMORROW FILE, THE  SANDERS, LAWRENCE  C/P:2,P
TOMORROW HAPPENS  BRIN, DAVID  SR-H

TOMORROW, INC.

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

HA

TOMORROW IS TOO FAR  WHITE, JAMES  C/P:2,P
TOMORROW KNIGHT  KURLAND, MICHAEL  C/P,P
TOMORROW LIES IN AMBUSH  SHAW, BOB  C/P:2,P
TOMORROW LOG, THE  LEE, SHARON/MILLER, STEVE  LP
TOMORROW MAY BE EVEN WORSE  BRUNNER, JOHN  B
TOMORROW MIDNIGHT  BRADBURY, RAY  SR-P,C/P
TOMORROW MIGHT BE DIFFERENT  REYNOLDS, MACK  C/P:2,P
TOMORROW PEOPLE, THE  MERRIL, JUDITH  C/P:2,P
TOMORROW PLUS X  TUCKER, WILSON  C/P:2,P
TOMORROW SUCKS  COX, GREG/WEISSKOPF, T. K. F.  PA
TOMORROW TESTAMENT, THE  LONGYEAR, BARRY B.  C/P,P
TOMORROW, THE STARS  HEINLEIN, ROBERT A.  C/PA,PA
TOMORROW TIMES SEVEN  POHL, FREDERIK  C/P:3,P
TOMORROW TODAY  ZEBROWSKI, GEORGE  SR-LPA
TOMORROW WINDOWS, THE [DOCTOR WHO EIGHTH DOCTOR]  MORRIS, JONATHAN  P
TOMORROW X 4  KNIGHT, DAMON  PA,C/PA
TOMORROW’S ALTERNATIVES [FRONTIERS 1]  ELWOOD, ROGER  HA,PA
TOMORROW’S CHILDREN  ASIMOV, ISAAC  HA,PA
TOMORROW’S COMET  SOWDEN, LEWIS  H
TOMORROW’S CRIMES  WESTLAKE, DONALD E.  H
TOMORROW’S GIFT  COOPER, EDMUND  C/P:2,P
TOMORROW’S HERITAGE [@CHILDREN OF THE STARS #1]  COULSON, JUANITA  C/P,P,H
TOMORROW’S MAGIC [MAGIC’S RETURN 2]  SERVICE, PAMELA F.  H,P
TOMORROW’S SON  HOSKINS, ROBERT  H,SFWA
TOMORROW’S SPHINX  BELL, CLARE  H
TOMORROW’S WORLDS  SILVERBERG, ROBERT  HA,PA
TONGUES OF THE MOON  FARMER, PHILIP JOSE  C/P:2,P
TONGUING THE ZEITGEIST  OLSEN, LANCE  LP
TONIGHT WE STEAL THE STARS  JAKES, JOHN  D81680
TONO-BUNGAY  WELLS, H. G.  H
TONY AND THE WONDERFUL DOOR  FAST, HOWARD  H
TOO CLOSE FOR COMFORT [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P:2,P
2 L8 4 U (PROOF) [@WICKED DEAD #1]  PETRUCHA, STEFAN/PENDLETON, THOMAS  SR-LP
TOO LONG A SACRIFICE  BROXON, MILDRED DOWNEY  H:2,P
TOO MANY MAGICIANS [LORD DARCY 2]  GARRETT, RANDALL  SR-P,C/P:2
TOO TOO SOLID FLESH  O’DONOHOE, NICK  P
TOOL OF THE TRADE  HALDEMAN, JOE  SR-H,C/P:2,P
TOOLMAKER KOAN  MCLOUGHLIN, JOHN C.  C/P,SFWA,SFWA-P,P,H
TOOTH AND CLAW [STAR TREK TNG]  DURGIN, DORANNA  P
TOOTH AND CLAW [DEATH DEALER 3]  SILKE, JAMES  C/P,SFWA-P,P
TOOTH AND CLAW  WALTON, JO  C/P,P
TOOTH FAIRY, THE  JOYCE, GRAHAM  LP
TOP DOG [INGERSOL 1]  CARROLL, JERRY JAY  C/P,P
TOPPER  SMITH, THORNE  C/P,P,H
TOPPER TAKES A TRIP  SMITH, THORNE  C/P,P,H
TOR HORROR SAMPLER  TOR  SR-PA,C/PA
TOR SF SAMPLER, THE  TOR  P
TORCH, THE  BECHDOLT, JACK  SR-H,H

The MITSFS Titledex  September 24, 2007  Page 378
TITLE  AUTHOR  CODE
TORCH OF FEAR [HARRY BORG 4]  HAWKINS, WARD  P
TORCH OF HONOR, THE [TORCH OF HONOR 1]  ALLEN, ROGER MACBRIDE  C/P:2,P
TOREADOR [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  WIECK, STEWART  P
TORIAN PEARLS, THE [@BLADE #25]  LORD, JEFFREY  P
TORMENT [@PLANE SCAPE]  VALLESE, RAY/VALLESE, VALERIE  P
TORMENTS [MANSE 2]  CANTRELL, LISA W.  SFWA-P
TORNADO ALLEY  TUNING, WILLIAM  C/P,P
TORRENT OF FACES, A  BLISH, JAMES/KNIGHT, NORMAN  H:2,P
TORTURE MACHINE, THE  TABORI, PAUL  C/P,P
TORTURED PLANET, THE (PERELANDRA 3 ABRIDGED) [PERELANDRA 3]  LEWIS, C. S.  C/P:2,P
TORTURER, THE  SAXON, PETER  P
TOTAL ECLIPSE  BRUNNER, JOHN  H:2,P,SFWA
TOTAL RECALL  ANTHONY, PIERS  H
TOTAL RECALL (ABRIDGED)  ANTHONY, PIERS  C/P:2,P
TOTAL WAR, THE [@SURVIVALIST #1]  AHERN, JERRY  P
TOTAL WAR [@FLEET #5]  DRAKE, DAVID/FAWCETT, BILL  PA,SFWA-P:2
TOTEM, THE  MORRELL, DAVID  P
TOUCH, THE  KEYES, DANIEL  P
TOUCH, THE [@NECROSCOPE 14]  LUMLEY, BRIAN  H
TOUCH, THE [ADVERSARY 3]  WILSON, F. PAUL  P
TOUCH NOT THE CAT  STEWART, MARY  P
TOUCH OF DEATH, THE  CREASEY, JOHN  SR-P
TOUCH OF ETERNITY, A [@PERRY RHODAN #57]  ACE  C/PA,PA
TOUCH OF INFINITY, A  ELLISON, HARLAN  DD413
TOUCH OF INFINITY, A  FAST, HOWARD  C/P,P,H
TOUCH OF STRANGE, A  STURGEON, THEODORE  C/P,P,H

TOUCH OF YOUR SHADOW, THE WHISPER OF YOUR NAME, THE [@BABYLON 5 #5] BARRETT, NEAL JR.

P

TOUCH THE DARK [CASSANDRA PALMER]  CHANCE, KAREN  P
TOUCH THE STARS: EMERGENCE  DALMAS, JOHN/MARTIN, CARL  P
TOUCHED BY MAGIC [REANDN 1]  DURGIN, DORANNA  P
TOUCHED BY THE GODS  WATT-EVANS, LAWRENCE  H,P
TOUCHED BY VENOM [@DRAGON TEMPLE #1]  CROSS, JANINE  LP
TOUGH GUIDE TO FANTASYLAND, THE  JONES, DIANA WYNNE  R/REF-H,C/REF-P
TOUR DU MONDE EN 80 JOURS, LE/RAYON-VERT, LE  VERNE, JULES  FH
TOUR DU MONDE EN QUATRE-VINGTS JOURS, LE  VERNE, JULES  FH
TOURISM  HARRISON, M. JOHN  B
TOURMALINE, THE [ROUMANIA 2]  PARK, PAUL  H
TOURNAMENT OF THORNS, THE  SWANN, THOMAS BURNETT  C/P:2,P
TOURS OF THE BLACK CLOCK  ERICKSON, STEVE  LP
TOWARD INFINITY  KNIGHT, DAMON  HA
TOWER, THE [SPIDER WORLD 2]  WILSON, COLIN  P
TOWER AND THE HIVE, THE [ROWAN 4]  MCCAFFREY, ANNE  H,P
TOWER AT STONY WOOD, THE  MCKILLIP, PATRICIA A.  H
TOWER AT THE EDGE OF TIME  CARTER, LIN  C/P,P
TOWER OF BABEL [MAGE ASCENSION]  STEELE, JOHN H.  P
TOWER OF BEOWULF, THE  GODWIN, PARKE  P
TOWER OF DEATH, THE [CORMAC 2]  OFFUTT, ANDREW J./TAYLOR, KEITH  P
TOWER OF DESPAIR [KILLSTAR 2]  KELLY, DAVID J.  SR-P,C/P
TOWER OF DOOM [@RAVENLOFT]  ANTHONY, MARK  P
TOWER OF DREAMS  NASIR, JAMIL  P
TOWER OF FEAR, THE  COOK, GLEN  C/P,SFWA-P,P,H
TOWER OF GLASS  SILVERBERG, ROBERT  C/P,P,H
TOWER OF RAVENS, THE [@RHIANNON’S RIDE #1]  FORSYTH, KATE  P
TOWER OF SHADOWS, THE  BOWLING, DREW C.  H
TOWER OF THE ELEPHANT, THE  HOWARD, ROBERT E.  VLP
TOWER OF THE GODS [ORGANIC FUTURE]  EASTON, THOMAS A.  P
TOWER OF THE KING’S DAUGHTER [@OUTREMER #2]  BRENCHLEY, CHAZ  C/P,P
TOWER OF THE KING’S DAUGHTER (COMPLETE) [OUTREMER 1,2]  BRENCHLEY, CHAZ  P
TOWER OF THE MEDUSA  CARTER, LIN  D42900
TOWER OF ZANID, THE [KRISHNA]  DE CAMP, L. SPRAGUE  C/P:2,P
TOWER ON THE RIFT, THE [@VIEW FROM THE MIRROR #2]  IRVINE, IAN  C/P,P
TOWER TO THE SKY  JENNINGS, PHILLIP C.  C/P,SFWA-P,P
TOWERS OF DARKOVER [DARKOVER]  BRADLEY, MARION ZIMMER  C/PA,PA
TOWERS OF MELNON, THE [@BLADE #15]  LORD, JEFFREY  P
TOWERS OF REXOR, THE [@DRAGONTALES #4]  VILOTT, RHONDI  P
TOWERS OF THE SUNSET, THE [RECLUCE 2]  MODESITT, L. E. JR.  C/P:2,P
TOWERS OF TORON, THE [FALL OF THE TOWERS 2]  DELANY, SAMUEL R.  C/P,DF261,P
TOWERS OF UTOPIA, THE  REYNOLDS, MACK  C/P:2,P
TOWING JEHOVAH [CORPUS DEI 1]  MORROW, JAMES  H:2,LP
TOXICOLOGY  AYLETT, STEVE  LP
TOY SOLDIERS [DOCTOR WHO NEW ADVENTURES]  LEONARD, PAUL  C/P:2,P
TOYBOX, THE [@IMMORTAL EYES #1]  CASSADA, JACKIE  P
TOYMAKER, THE  JONES, RAYMOND F.  SR-LP

TITLE  AUTHOR  CODE
TOYMAN [DUMAREST 3]  TUBB, E. C.  D23140
TOYNBEE CONVECTOR, THE  BRADBURY, RAY  C/P,P
TRACE OF DREAMS, A  EKLUND, GORDON  C/P:2,P
TRACE OF MEMORY, A  LAUMER, KEITH  C/P:2,SFWA-P,P
TRACES  BAXTER, STEPHEN  P
TRACK OF THE WHITE WOLF [CHEYSULI 4]  ROBERSON, JENNIFER  C/P,P
TRACKER  STILLMAN, RON  P
TRACKERS [BIRTHRIGHT PROJECT 2]  MACKEL, KATHRYN  LP
TRADER [NEWFORD]  DE LINT, CHARLES  C/P:2,P
TRADER REDUX [@DEATHLANDS]  AXLER, JAMES  P
TRADER TO THE STARS  ANDERSON, POUL  P,H:2,SR-H
TRADER’S WORLD  SHEFFIELD, CHARLES  C/P:2,P
TRADING FUTURES [DOCTOR WHO EIGHTH DOCTOR]  PARKIN, LANCE  P
TRADING IN DANGER [VATTA’S WAR 1]  MOON, ELIZABETH  H
TRAGEDY DAY [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P
TRAGEDY OF THE MOON, THE  ASIMOV, ISAAC  SFWA
TRAIL BLAZERS OF THE SKIES  LANGLEY, JOHN PRENTICE  H
TRAIL OF BOHU, THE [IMARO 3]  SAUNDERS, CHARLES R.  P
TRAIL OF CTHULHU, THE  DERLETH, AUGUST  P
TRAIL OF DARKNESS  BOLESNY, DARLENE D.  P
TRAIL OF FU MANCHU, THE  ROHMER, SAX  C/P,P
TRAIL OF THE BLACK WYRM [@DRAGONLANCE TALADAS #2]  PIERSON, CHRIS  P
TRAIL OF THE SEAHAWKS  MAYHAR, ARDATH/FORTIER, RON  P
TRAILS IN DARKNESS  HOWARD, ROBERT E.  P
TRAILS OF PERIL [@WARRIOR OF VENGEANCE #2]  COE, ROSS ANTON  P
TRAITOR [@STAR WARS NEW JEDI ORDER]  STOVER, MATTHEW WOODRING  P
TRAITOR AND THE CHALICE, THE [LYREMOUTH 2]  FLETCHER, JANE  LP
TRAITOR GENERAL [@WARHAMMER 40K GAUNT’S GHOSTS]  ABNETT, DAN  P
TRAITOR TO THE BLOOD [NOBLE DEAD SERIES 1 4]  HENDEE, BARB/HENDEE, J. C.  H
TRAITOR TO THE LIVING  FARMER, PHILIP JOSE  C/P:2,SFWA-P,P
TRAITOR WINDS [STAR TREK]  GRAF, L. A.  C/P:2,P
TRAITOR’S HAND, THE [@WARHAMMER 40K CIAPHAS CAIN 3]  MITCHELL, SANDY  P
TRAITOR’S MOON [NIGHTRUNNER 3]  FLEWELLING, LYNN  C/P,P
TRAITOR’S SUN [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
TRAMONTANE [KALEVALA 4]  PETAJA, EMIL  DH036
TRAMP ROYALE  HEINLEIN, ROBERT A.  SFWA:2
TRAMPLE AN EMPIRE DOWN  REYNOLDS, MACK  C/P,P
TRAMPOLINE  LINK, KELLY  LPA
TRAN [JANISSARIES]  POURNELLE, JERRY/GREEN, ROLAND  P
TRANQUILITY ALTERNATIVE, THE  STEELE, ALLEN  P
TRANQUILITY WARS, THE  LEE, GENTRY  H,P
TRANS-MERCURIAN  LANG, KING  SR-P
TRANSATLANTIC TUNNEL, HURRAH!, A  HARRISON, HARRY  C/P:2,P
TRANSCENDENCE [DEMONWARS II 2]  SALVATORE, R. A.  H
TRANSCENDENCE [HERITAGE UNIVERSE 3]  SHEFFIELD, CHARLES  H:2,P
TRANSCENDENT [DESTINY’S CHILDREN 3]  BAXTER, STEPHEN  H
TRANSCENDENT MAN, THE  SOHL, JERRY  C/P:2,P

TRANSCENDENTAL TALES FROM ISAAC ASIMOV’S SCIENCE FICTION MAGAZINE DOZOIS, GARDNER

LPA

TRANSCENSION  BRODERICK, DAMIEN  H
TRANSFER TO YESTERDAY  HAIBLUM, ISIDORE  C/P:2,SFWA,P
TRANSFERENCE, THE  SMITH, ELLA  P
TRANSFIGURATIONS  BISHOP, MICHAEL  H:2,SFWA,P
TRANSFIGURATIONS (PROOF)  BISHOP, MICHAEL  SR-LP
TRANSFINITE MAN  KAPP, COLIN  C/P:2,P
TRANSFINITE: THE ESSENTIAL A.E. VAN VOGT  VAN VOGT, A. E.  H
TRANSFORMATION [SEYONNE 1]  BERG, CAROL  P
TRANSFORMATION: THE BREAKTHROUGH  STRIEBER, WHITLEY  [C/RAN-P]
TRANSFORMATIONS  ASHLEY, MIKE  R/REF-H
TRANSFORMATIONS  ROSELLE, DANIEL  PA
TRANSFORMATIONS (CASSETTE) [STAR TREK]  STERN, DAVE  SR-MM
TRANSFORMATIONS II  ROSELLE, DANIEL  PA,C/PA
TRANSFORMER [MORPHODITE 2]  FOSTER, M. A.  C/P:2,P
TRANSFORMING EARTH  WILLIAMSON, JACK  P
TRANSGALACTIC GUIDE TO SOLAR SYSTEM M-17, THE  ROVIN, JEFF  H,LP
TRANSGALAXIS KATALOG 1  BINGENHEIMER, HEINZ  R/IND
TRANSIT [DOCTOR WHO NEW ADVENTURES]  AARONOVITCH, BEN  C/P,P
TRANSIT  COOPER, EDMUND  C/P:2,P
TRANSIT OF EARTH  PLAYBOY BOOKS  PA,C/PA
TRANSIT TO SCORPIO [DRAY PRESCOT 1]  AKERS, ALAN BURT  C/P:2,P
TRANSITION [STARFARERS 2]  MCINTYRE, VONDA N.  C/P:2,P
TRANSITION OF H.P. LOVECRAFT, THE  LOVECRAFT, H. P.  LP
TRANSITION OF TITUS CROW, THE [TITUS CROW 1]  LUMLEY, BRIAN  C/P,P
TRANSLATION  MARLOWE, STEPHEN  P

The MITSFS Titledex  September 24, 2007  Page 380
TITLE  AUTHOR  CODE
TRANSMIGRATION  MCINTOSH, J. T.  C/P:2,P
TRANSMIGRATION OF SOULS, THE  BARTON, WILLIAM  C/P:2,P
TRANSMIGRATION OF TIMOTHY ARCHER, THE  DICK, PHILIP K.  H,SFWA,P
TRANSMISSION ERROR  KURLAND, MICHAEL  C/P,P
TRANSMUTATIONS  PANSHIN, ALEXEI  SR-LP
TRANSPLANT  JONES, MARGARET  P
TRANSPOSED HEADS, THE  MANN, THOMAS  H,P
TRANSPOSED MAN, THE  SWAIN, DWIGHT V.  DD113
TRANSVECTION MACHINE, THE  HOCH, EDWARD D.  H,SFWA,P
TRANSVERGENCE [HERITAGE UNIVERSE (REVISED) 3,4]  SHEFFIELD, CHARLES  C/P,P
TRAPPED [LEAGUE OF PEOPLES 5]  GARDNER, JAMES ALAN  H,P
TRAPPED! [@TIME POLICE #2]  NORWOOD, WARREN  P
TRAPPED IN SPACE  WILLIAMSON, JACK  C/P,LP
TRAPPED IN THE ASHES [ASHES 9]  JOHNSTONE, WILLIAM W.  P
TRAPPERS OF VENUS [@DIG ALLEN #4]  GREENE, JOSEPH  H
TRASH SEX MAGIC  STEVENSON, JENNIFER  H
TRAUMA  SHAUGHNESSY, P. J.  P
TRAVEL WITHOUT THE TARDIS [DOCTOR WHO]  AIREY, JEAN/HALDEMAN, LINDA  R/REF-P
TRAVELER, THE [FOURTH REALM 1]  TWELVE HAWKS, JOHN  LP
TRAVELER IN BLACK, THE  BRUNNER, JOHN  C/P:2,P
TRAVELER IN TIME, A  UTTLEY, ALISON  H
TRAVELERS IN TIME  STERN, PHILIP VAN DOREN  HA
TRAVELING SOUL, THE  RAE, HUGH C.  P
TRAVELING TIDE, THE  LOVE, ROSALEEN  LP
TRAVELING WITH THE DEAD [LONDON VAMPIRES 2]  HAMBLY, BARBARA  C/P,P,H
TRAVELLER IN TIME, A  UTTLEY, ALISON  P
TRAVELLERS IN MAGIC  GOLDSTEIN, LISA  H,LP
TRAVELLING CIRCUS, THE  VERNE, JULES  H
TRAVELLING TOWARDS EPSILON  JAKUBOWSKI, MAXIM  SR-PA
TRAVELS OF A REPUBLICAN RADICAL IN SEARCH OF HOT WATER  WELLS, H. G.  [C/RAN-P]
TREACHERY AND TREASON  GILMAN, LAURA ANNE/HEDDLE, JENNIFER  PA
TREACHERY IN OUTER SPACE [TOM CORBETT 6]  ROCKWELL, CAREY  H
TREACHERY OF KINGS, THE  BARRETT, NEAL JR.  P
TREACHERY ON THE DARK SIDE [MATTHEWS 6]  BURGAUER, STEVEN  LP
TREACHERY’S WAKE [@DUNGEONS & DRAGONS]  LAIN, T. H.  P
TREAD SOFTLY  LAYMON, RICHARD  SFWA-P
TREASON  CARD, ORSON SCOTT  C/P,P,H
TREASON IN TIME [@FREEDOM’S RANGERS #4]  ANDREWS, KEITH WILLIAM  P
TREASON KEEP [HYTHRUN 2]  FALLON, JENNIFER  H
TREASON OF DORTREAN, THE [@ARRANDIN #2]  HERNIMAN, MARCUS  P
TREASON OF ISENGARD [HISTORY OF MIDDLE-EARTH 7]  TOLKIEN, J. R. R.  H
TREASURE [DIRK PITT]  CUSSLER, CLIVE  P
TREASURE BOX  CARD, ORSON SCOTT  C/P,P,H
TREASURE HUNTERS, THE  WETANSON, BURT/HOOBLER, THOMAS  P
TREASURE IN THE HEART OF THE MAZE, THE [RABELAIS 2]  CARR, JAYGE  H,SFWA
TREASURE OF ATLANTIS, THE  DUNN, J. ALLEN  P
TREASURE OF GREEN KNOWE [GREEN KNOWE 2]  BOSTON, L. M.  H,LP
TREASURE OF LIGHT [GAMANT 2]  O’NEAL, KATHLEEN N.  P
TREASURE OF TAU CETI  RACKHAM, JOHN  D23775
TREASURE OF THE BLACK FALCON  BURROUGHS, JOHN COLEMAN  C/P:2,P
TREASURE OF THE STARS [@BLADE #29]  LORD, JEFFREY  P
TREASURE OF WONDERWHAT, THE [FARSTAR & SON 2]  STARR, BILL  C/P:2,P
TREASURED ONE, THE [@DREAMERS #2]  EDDINGS, DAVID/EDDINGS, LEIGH  H
TREASURES OF FANTASY  WEIS, MARGARET/HICKMAN, TRACY  C/PA,LPA
TREASURY OF FANTASY, A  WILKINS, CARY  HA,LPA
TREASURY OF GREAT GHOST STORIES, A  PECK, IRA  PA
TREASURY OF GREAT SCIENCE FICTION 1, A  BOUCHER, ANTHONY  HA
TREASURY OF GREAT SCIENCE FICTION 2, A  BOUCHER, ANTHONY  HA
TREASURY OF MODERN FANTASY, A  CARR, TERRY/GREENBERG, MARTIN H.  LPA
TREASURY OF SCIENCE FICTION, A  CONKLIN, GROFF  PA,C/PA
TREASURY OF SCIENCE FICTION CLASSICS, THE  KUEBLER, HAROLD W.  SR-HA,HA
TREATY AT DOONA [DOONA]  MCCAFFREY, ANNE/NYE, JODY LYNN  P

TREATY’S LAW [STAR TREK DAY OF HONOR 4] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

C/P,P

TREE AND LEAF  TOLKIEN, J. R. R.  R/REF-H,C/REF-H:2
TREE AND LEAF/MYTHOPOEIA  TOLKIEN, J. R. R.  SR-H:2
TREE BY TOLKIEN  WILSON, COLIN  R/REF-P
TREE LORD OF IMETEN, THE  PURDOM, TOM  DM139
TREE OF HEAVEN, THE  CHAMBERS, ROBERT W.  H
TREE OF SWORDS AND JEWELS, THE [ARAFEL 2]  CHERRYH, C. J.  C/P:3,P
TREE PEOPLE, THE  STOKES, NAOMI M.  P
TREES OF ZHARKA, THE  MACKENROTH, NANCY  P
TREK OR TREAT  FLANAGAN, TERRY/EHRHARDT, ELEANOR  [C/RAN-P]
TREK: THE NEXT GENERATION REVISED EDITION [STAR TREK TNG]  VAN HISE, JAMES  R/REF-H

The MITSFS Titledex  September 24, 2007  Page 381
TITLE  AUTHOR  CODE
TREK: THE UNAUTHORIZED A–Z [STAR TREK]  SCHUSTER, HAL/RATHBONE, WENDY  R/REF-P,C/REF-P
TREK TO KRAGGEN-COR [@SILVER CALL #1/MITHGAR 10]  MCKIERNAN, DENNIS L.  C/P,SFWA,P,H
TREK TO MADWORLD [STAR TREK]  GOLDIN, STEPHEN  C/P:2,P
TREKMASTER  JOHNSON, JAMES B.  C/P:2,P
TREKS NOT TAKEN [STAR TREK TNG]  BOYETT, STEVEN R.  LP
TREKTOONS [STAR TREK]  LISTER, MARK  SCX
TRELLISANE CONFRONTATION, THE [STAR TREK]  DVORKIN, DAVID  C/P:3,P
TREMOR IN THE BITTER EARTH, A [TIELMARAN 2]  REIMANN, KATYA  H,P
TRENCH AND BEYOND, THE  ROWE, KEN  LP
TRENCH SOLDIER, THE [@CASCA #21]  SADLER, BARRY  P
TREY OF SWORDS [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P,P,H
TRI-PLANET  KELLAR, VON  SR-P
TRIAD  FINCH, SHEILA  C/P:2,P
TRIAD [ILLUMINATION 3]  MCGARRY, TERRY  H
TRIAD  VAN VOGT, A. E.  H:2
TRIAD WORLDS, THE  BUSBY, F. M.  P
TRIAL BY CHAOS [@MECHWARRIOR DARK AGE/BATTLETECH]  YORK, J. STEVEN  P
TRIAL BY ERROR [STAR TREK DS9]  GARLAND, MARK A.  C/P,P
TRIAL BY FIRE [GUARDIANS 2]  AUSTIN, RICHARD  P
TRIAL BY FIRE [@STARGATE SG-1 #1]  BAUER, SABINE C.  P
TRIAL OF TERRA, THE  WILLIAMSON, JACK  C/P:2,P
TRIAL OF THE SEVENTH CARRIER [SEVENTH CARRIER 4]  ALBANO, PETER  P
TRIALS AND TRIBBLE-ATIONS [STAR TREK DS9]  CAREY, DIANE  P
TRIANGLE [GALACTIC EMPIRE 1,2,3]  ASIMOV, ISAAC  H
TRIANGLE [STAR TREK]  MARSHAK, SONDRA/CULBREATH, MYRNA  C/P:3,P
TRIAX  SILVERBERG, ROBERT  C/PA,PA
TRIBE, THE  WOOD, BARI  P
TRIBESMEN OF GOR [GOR 10]  NORMAN, JOHN  C/P:2,P
TRIBULATION FORCE [@LEFT BEHIND #2]  LAHAYE, TIM/JENKINS, JERRY B.  LP
TRIBULATIONS OF A CHINESE GENTLEMAN, THE  VERNE, JULES  H
TRICENTENNIAL REPORT: LETTERS FROM AMERICA, THE  ATLANTIC RICHFIELD CO.  R/REF-H
TRICK TOP HAT, THE [@SCHRODINGER’S CAT #2]  WILSON, ROBERT ANTON  C/P:2,P
TRICKSTER [SILENT EMPIRE 3]  HARPER, STEVEN  P
TRICKSTERS, THE  MAHY, MARGARET  SFWA
TRICKSTERS [DAUGHTER OF THE LIONESS 1,2]  PIERCE, TAMORA  H
TRICKSTER’S TOUCH [CONTRARYWISE 2]  GREENHALGH, ZOHRA  C/P,P
TRIGAN EMPIRE, THE  CHARTWELL BOOKS, INC.  CX

TRIGGER, THE CLARKE, ARTHUR C./

KUBE-MCDOWELL, MICHAEL P.

H

TRIGGER & FRIENDS [TELZEY]  SCHMITZ, JAMES H./FLINT, ERIC  C/P,P
TRIKON DECEPTION, THE  BOVA, BEN/POGUE, BILL  C/P,P,H

TRILL: UNJOINED [STAR TREK DS9 WORLDS 2]/ BAJOR: FRAGMENTS AND OMENS [STAR TREK DS9 WORLDS 2] MARTIN, MICHAEL A./MANGELS, ANDY/

KYM, J. NOAH

C/P,P

TRINITY  KRESS, NANCY  H
TRINITY AND OTHER STORIES  KRESS, NANCY  P
TRINITY PARADOX, THE  ANDERSON, KEVIN J./BEASON, DOUG  C/P:2,P
TRINITY VECTOR, THE  PERRY, STEVE  C/P:2,P
TRIO FOR LUTE, A [DAMIANO 1,2,3]  MACAVOY, R. A.  H:2,P
TRIP TO MARS, A  ASH, FENTON  SR-H
TRIPLANETARY [LENSMEN 1]  SMITH, E. E.  H:2,P
TRIPLE DETENTE  ANTHONY, PIERS  C/P:2,SFWA-P,P
TRIPLE MAN, THE  FANTHORPE, R. L.  SR-P
TRIPLE ZERO [@STAR WARS REPUBLIC COMMANDO 2]  TRAVISS, KAREN  P
TRIPLET  ZAHN, TIMOTHY  C/P:2,P
TRIPLICITY  DISCH, THOMAS M.  H:2
TRIPOINT [UNION/ALLIANCE]  CHERRYH, C. J.  H:2,P
TRIPS IN TIME  SILVERBERG, ROBERT  HA
TRIS’S BOOK [@CIRCLE OF MAGIC #2]  PIERCE, TAMORA  P
TRISTAN  VON STASSBURG, GOTTFRIED  P
TRITON  DELANY, SAMUEL R.  C/P:2,P
TRITONIAN RING, THE  DE CAMP, L. SPRAGUE  H:2,P
TRIUMPH  BOVA, BEN  H,P
TRIUMPH  WYLIE, PHILIP  C/P:2,P
TRIUMPH OF SOULS, A [JOURNEYS OF THE CATECHIST 3]  FOSTER, ALAN DEAN  H
TRIUMPH OF THE DARKSWORD [@DARKSWORD #3]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
TRIUMPH OF THE DRAGON [BROTHERS OF THE DRAGON 3]  BAILEY, ROBIN WAYNE  C/P,P
TRIUMPH OF TIME, THE  BLISH, JAMES  C/P:2,P
TRIUMPHANT, THE [@BOLOS #3]  WEBER, DAVID/EVANS, LINDA  P
TRIUNE MAN, THE  LUPOFF, RICHARD A.  C/P,P,H
TRIVIA MANIA: STAR TREK  EINSTEIN, XAVIER  [C/RAN-P]
TROIKA [@INDIGO #5]  COOPER, LOUISE  C/P:2,P
TROIKA INCIDENT: A PROMISE, THE  BROWN, JAMES COOKE  H
TROJAN GOLD  PETERS, ELIZABETH  SFWA-P:2
TROJAN ORBIT  REYNOLDS, MACK/ING, DEAN  C/P,P

The MITSFS Titledex  September 24, 2007  Page 382
TITLE  AUTHOR  CODE
TROLL FELL  LANGRISH, KATHERINE  H,P
TROLL FELL (PROOF)  LANGRISH, KATHERINE  SR-LP
TROLL-QUEST  ESTES, ROSE  P
TROLL-TAKEN  ESTES, ROSE  P
TROLL’S GRINDSTONE, THE [WIZARD’S WAR 1]  BOYER, ELIZABETH H.  C/P,P
TROLLSLAYER [WARHAMMER GOTREK & FELIX 1]  KING, WILLIAM  P
TRON  DALEY, BRIAN  C/P,P,H
TRON: THE STORYBOOK  WEINBERG, LAWRENCE  H
TROPHY, THE [HELMSMAN 3]  BALDWIN, BILL  C/P:2,P
TROPIC OF CREATION  KENYON, KAY  P
TROPIC OF NIGHT  GRUBER, MICHAEL  H

TROS OF SAMOTHRACE [TROS OF SAMOTHRACE 1]/ LUD OF LUNDUN [TROS OF SAMOTHRACE 1] MUNDY, TALBOT

P

TROTSKY’S RUN  HOYT, RICHARD  SFWA-P
TROUBLE AND HER FRIENDS  SCOTT, MELISSA  H,P
TROUBLE MAGNET [HUMANX FLINX]  FOSTER, ALAN DEAN  H
TROUBLE ON CLOUD CITY [@STAR WARS YOUNG JEDI KNIGHTS]  ANDERSON, KEVIN J./MOESTA, REBECCA  P
TROUBLE ON TITAN  NOURSE, ALAN E.  C/P,P,H
TROUBLE ON TRITON  DELANY, SAMUEL R.  LP
TROUBLE PLANET  STEEL, MARK  SR-P
TROUBLE TWISTERS, THE  ANDERSON, POUL  C/P:2,P
TROUBLE WITH ALIENS, THE  ANVIL, CHRISTOPHER/FLINT, ERIC  H
TROUBLE WITH JENNY’S EAR, THE  BUTTERWORTH, OLIVER  H
TROUBLE WITH LICHEN (AMERICAN)  WYNDHAM, JOHN  C/P,P,H
TROUBLE WITH LICHEN (BRITISH)  WYNDHAM, JOHN  SR-P
TROUBLE WITH MAGIC, THE [@BEWITCHING MYSTERY 1]  ALT, MADELYN  P
TROUBLE WITH MISS SWITCH, THE  WALLACE, BARBARA BROOKS  P
TROUBLE WITH TRIBBLES, THE [STAR TREK]  GERROLD, DAVID  R/REF-P,C/REF-P:2
TROUBLE WITH TYCHO, THE  SIMAK, CLIFFORD D.  DD517,C/P:2,P
TROUBLE WITH YOU EARTH PEOPLE, THE  MACLEAN, KATHERINE  LP
TROUBLED STAR  SMITH, GEORGE O.  SR-P,C/P
TROUBLED WATERS [@MEROVINGEN NIGHTS #3]  CHERRYH, C. J.  PA,C/PA
TROUBLEMAKERS  ELLISON, HARLAN  LP
TROUBLING ALONG THE BORDER, A  AAMODT, DONALD  C/P:2,P
TRUCE AT BAKURA, THE [@STAR WARS]  TYERS, KATHY  C/P,P,H
TRUCE WITH TIME, A  GODWIN, PARKE  C/P,SR-LP,H
TRUE BELIEVER, THE  HOFFER, ERIC  [C/RAN-P:2]
TRUE BLOOD  WADDELL, PATRICIA  P
TRUE GAME, THE [TRUE GAME 1,2,3]  TEPPER, SHERI S.  H:2
TRUE JAGUAR  NORWOOD, WARREN  C/P,P
TRUE KNIGHT, THE [WARHORSE OF ESDRAGON 3]  DEXTER, SUSAN  P
TRUE MEN, THE  STEELE, MARY Q.  H
TRUE NAMES  VINGE, VERNOR  BS5
TRUE NAMES ... AND OTHER DANGERS  VINGE, VERNOR  P,SFWA-P
TRUE NAMES AND THE OPENING OF THE CYBERSPACE FRONTIER  FRENKEL, JAMES  LPA
TRUE SON OF THE BEAST!  BROWN, CARTER  P
TRUE TALES OF THE UNKNOWN: THE UNINVITED  JARVIS, SHARON  PA
TRUEST POWER, THE [AGHAMORE 1]  NEASON, REBECCA  P
TRULLION: ALASTOR 2262 [ALASTOR 1]  VANCE, JACK  C/P:2,P

TRULLION: ALASTOR 2262 (VIE) [@VANCE INTEGRAL EDITION #29/ ALASTOR 1] VANCE, JACK

SR-H

TRUMPETS OF TAGAN, THE [USS SKIPJACK 1]  LANG, SIMON  P
TRUMPS OF DOOM [AMBER 6]  ZELAZNY, ROGER  C/P,H,@P,SFWA
TRUST TERRITORY  MORRIS, JANET/MORRIS, CHRIS  H:2,P
TRUTH, THE [DISCWORLD 25]  PRATCHETT, TERRY  H:2,P
TRUTH ABOUT DRAGONS, THE  ADAMS, HAZARD  H
TRUTH AND CONSEQUENCES [@CHARMED]  DOKEY, CAMERON  P

TRUTH AND SHADOWS [@MECHWARRIOR DARK AGE PROVING GROUNDS 2/BATTLETECH] DELRIO, MARTIN

P

TRUTH IS OUT THERE [X-FILES]  LOWRY, BRIAN  R/REF-P,C/REF-H
TRUTH MACHINE, THE  HALPERIN, JAMES L.  H,P
TRUTH UNTIL PARADOX  KRAUSE, STALEY/WIECK, STEWART  PA
TRUTHFUL HARP, THE  ALEXANDER, LLOYD  LP
TRUTHSAYER’S APPRENTICE, THE [LOREGIVER 1]  CHRISTIAN, DEBORAH  LP
TRUXTON KING  MCCUTCHEON, GEORGE B.  H
TRY ANOTHER WORLD  MEYER, JOHN J.  H
TRYST  THANE, ELSWYTH  P
TSADDIK OF THE SEVEN WONDERS, THE  HAIBLUM, ISIDORE  P,SFWA
TSATHOGGUA CYCLE, THE  PRICE, ROBERT M.  LPA
TSCHAI [@VANCE INTEGRAL EDITION #21]  VANCE, JACK  SR-H
TSUNAMI [@MARQ’SSAN #3]  DUCHAMP, L. TIMMEL  LP
TUBULAR ANDROID SUPERHEROES  GILDEN, MEL  C/P,P
TUCK EVERLASTING  BABBITT, NATALIE  P
TUF VOYAGING  MARTIN, GEORGE R. R.  C/P,P,H

The MITSFS Titledex  September 24, 2007  Page 383
TITLE  AUTHOR  CODE
TUG OF THE DWARF STAR  RUSK, JAMES JR.  P
TULKU  DICKINSON, PETER  P
TULPA, THE  WILLIAMSON, J. N.  P
TUMBLING AFTER  WITCOVER, PAUL  H
TUNE IN YESTERDAY  BETHANCOURT, T. ERNESTO  P
TUNED FOR MURDER [@AVENGER #9]  ROBESON, KENNETH  C/P:2,P
TUNESMITH, THE  BIGGLE, LLOYD JR.  SFWA-TD27,TD27
TUNNEL, THE  BYRNE, ROBERT  P
TUNNEL, THE  ELWOOD, ROGER  SR-HA
TUNNEL AT THE END OF THE LIGHT, THE [@TIME HUNTER 2]  PETRUCHA, STEFAN  LP
TUNNEL FROM CALAIS  RAME, DAVID  H
TUNNEL IN THE SKY  HEINLEIN, ROBERT A.  C/P:2,P
TUNNEL TERROR [DOC SAVAGE #93]  ROBESON, KENNETH  @P
TUNNEL THROUGH THE DEEPS  HARRISON, HARRY  H,P

TUNNEL THROUGH THE STARS [@STAR TREK DOMINION WAR #3/ STAR TREK TNG] VORNHOLT, JOHN

P

TUNNEL THROUGH TIME  DEL REY, LESTER  C/P,LP,H
TUNNEL THRU THE AIR, THE  GANN, W. D.  H
TUPOLEV TOO FAR, A  ALDISS, BRIAN W.  H
TURING OPTION, THE  HARRISON, HARRY/MINSKY, MARVIN  H:2,P
TURING TEST, THE [DOCTOR WHO EIGHTH DOCTOR]  LEONARD, PAUL  C/P,P
TURLOUGH AND THE EARTHLINK DILEMMA [DOCTOR WHO]  ATTWOOD, TONY  SR-P
TURN LEFT AT THURSDAY  POHL, FREDERIK  C/P:2,P
TURN OF THE CARDS [WILD CARDS 12]  MILAN, VICTOR  C/P:2,P
TURN THE OTHER CHICK  FRIESNER, ESTHER M.  HA,PA
TURNABOUT  SMITH, THORNE  C/P:2,P
TURNCOAT [NOEL KEDRAN 2]  DALTON, SEAN  P
TURNING ON  KNIGHT, DAMON  C/P:2,P
TURNING ON THE GIRLS  BERNARD, CHERYL  H
TURNING PLACE, THE  KARL, JEAN E.  H,P
TURNING POINT [SHOLAN ALLIANCE 1]  NORMAN, LISANNE  C/P,P
TURNING POINTS [THIEVES’ WORLD]  ABBEY, LYNN  HA
TURNING POINTS (PROOF) [THIEVES’ WORLD]  ABBEY, LYNN  SR-LPA
TURNING POINTS: ESSAYS ON THE ART OF SCIENCE FICTION  KNIGHT, DAMON  SFWA,R/REF-H
TURNING THE STORM [ELIANA 2]  KRITZER, NAOMI  P
TURQUOISE, THE  SETON, ANYA  P

TV: 2000 ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

PA

TWEEDLIOOP  SCHMIDT, STANLEY  H,SFWA,SFWA-P,P
12TH PLANET, THE  SITCHIN, ZECHARIA  [C/RAN-P]

TWELVE CRIMES OF CHRISTMAS, THE WAUGH, CAROL-LYNN ROSSEL/

GREENBERG, MARTIN H./ASIMOV, ISAAC

PA

TWELVE FAIR KINGDOMS [OZARK 1]  ELGIN, SUZETTE HADEN  C/P,SFWA,P

TWELVE FRIGHTS OF CHRISTMAS, THE ASIMOV, ISAAC/WAUGH, CHARLES G./

GREENBERG, MARTIN H.

C/PA,PA

12 GREAT CLASSICS OF SCIENCE FICTION  CONKLIN, GROFF  PA,C/PA
TWELVE MAIDENS, THE  FARRAR, STEWART  P
12 MONKEYS  HAND, ELIZABETH  P
12 STORIES THEY WOULDN’T LET ME DO ON TV  HITCHCOCK, ALFRED  PA
TWELVE TALES OF SUSPENSE AND THE SUPERNATURAL  GRUBB, DAVIS  P
TWELVE TIMES ZERO  BROWNE, HOWARD  GRD15
20TH CENTURY COMPUTERS AND HOW THEY WORKED [STAR TREK TNG]  FLYNN, JENNIFER  R/REF-H
TWENTIETH CENTURY DISCOVERY  ASIMOV, ISAAC  [C/RAN-P:2]
TWENTIETH SON OF ORNON, THE  SIROTA, MIKE  C/P,P
TWENTY-EIGHT SCIENCE FICTION STORIES OF H. G. WELLS  WELLS, H. G.  H:2
2018 A.D. OR THE KING KONG BLUES  LUNDWALL, SAM J.  C/P:2,P
TWENTY-FIFTH HOUR, THE  BEST, HERBERT  H
TWENTY-FIRST CENTURY SUB  HERBERT, FRANK  C/P:2,P
20 MILLION MILES TO EARTH  SLESAR, HENRY  SR-P,C/P
TWENTY-ONE BILLIONTH PARADOX  DAVENTRY, LEONARD  H,SFWA
2150 A.D.  ALEXANDER, THEA  C/P,P
TWENTY-SECOND CENTURY, THE  CHRISTOPHER, JOHN  C/P:2,P
TWENTY-SEVENTH DAY, THE  MANTLEY, JOHN  H:2,P
2076: THE AMERICAN TRICENTENNIAL  BRYANT, EDWARD  PA
20,000 LEAGUES UNDER THE SEA/AROUND THE MOON  VERNE, JULES  H
TWENTY THOUSAND LEAGUES UNDER THE SEA  VERNE, JULES  C/P,P,H,LP
TWENTY THOUSAND LEAGUES UNDER THE SEA (ANNOTATED)  VERNE, JULES  LP
TWENTY THOUSAND LEAGUES UNDER THE SEA (BONNER TRANSLATION)  VERNE, JULES  P
2020 VISION  POURNELLE, JERRY  PA,C/PA
20/20 VISION  WEST, PAMELA  P
TWENTY-TWO GOBLINS (RYDER TRANSLATION)  ANONYMOUS  H
20 YEARS OF EXCELLENT ADVENTURES (CD)  BAEN PUBLISHING ENTERPRISES  SR-MM:2
TWENTY YEARS OF THE MAGAZINE OF FANTASY AND SCIENCE FICTION  FERMAN, EDWARD L./MILLS, ROBERT P.  C/PA,PA
TWICE 22  BRADBURY, RAY  H:2

The MITSFS Titledex  September 24, 2007  Page 384
TITLE  AUTHOR  CODE
TWICE DEAD THINGS  ATTANASIO, A. A.  LP
TWICE IN TIME  WELLMAN, MANLY WADE  SR-P,C/P:2,SFWA-P
TWICE SEVEN  BOVA, BEN  P
TWICE UPON A TIME  APPEL, ALLEN  P
TWICE UPON A TIME  FONTENAY, CHARLES L.  DD266
TWICE UPON A TIME  LITTLE, DENISE  C/PA,PA
TWILIGHT [STAR TREK DS9 MISSION GAMMA #1]  GEORGE, DAVID R. III  C/P,P
TWILIGHT  STEMBER, NICHOLAS S.  LP
TWILIGHT AT THE WELL OF SOULS [WELL WORLD 5]  CHALKER, JACK L.  C/P:2,P
TWILIGHT CANDELABRA  CRADDOCK, WILLIAM J.  LP
TWILIGHT CHILDREN [@DEATHLANDS]  AXLER, JAMES  P
TWILIGHT EYES  KOONTZ, DEAN R.  P
TWILIGHT FALLING [@FORGOTTEN REALMS EREVIS CALE #1]  KEMP, PAUL S.  P
TWILIGHT JOURNEY  DAVIES, L. P.  H:2
TWILIGHT MAN, THE  MOORCOCK, MICHAEL  C/P:2,P
TWILIGHT OF BRIAREUS, THE  COWPER, RICHARD  H:2,P
TWILIGHT OF MAGIC, THE  LOFTING, HUGH  H
TWILIGHT OF REASON  BURKE, JONATHAN  SR-P
TWILIGHT OF THE BASILISKS  TRANSUE, JACOB  P
TWILIGHT OF THE CITY  PLATT, CHARLES  H,P
TWILIGHT OF THE EMPIRE [DEATHSTALKER]  GREEN, SIMON R.  C/P,P
TWILIGHT OF THE GODS [DOCTOR WHO MISSING ADVENTURES]  BULIS, CHRISTOPHER  C/P,P

TWILIGHT OF THE GODS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] CLAPHAM, MARK/MILLER, JON DE BURGH

P

TWILIGHT OF THE MIND [@ALPHA CENTAURI #3]  ELY, MICHAEL  P
TWILIGHT OF THE SERPENT  TIMLETT, PETER VALENTINE  C/P,P
TWILIGHT REALM, THE  CARPENTER, CHRISTOPHER  C/P,P
TWILIGHT RIVER, THE  EKLUND, GORDON  BS2
TWILIGHT WATCH [NIGHTWATCH 3]  LUKYANENKO, SERGEI  LP
TWILIGHT WORLD  ANDERSON, POUL  H:2,P
TWILIGHT ZONE  BADGER BOOKS  SR-PA
TWILIGHT ZONE COMPANION, THE (FIRST EDITION)  ZICREE, MARC SCOTT  [C/RAN-H]
TWILIGHT ZONE COMPANION, THE (SECOND EDITION)  ZICREE, MARC SCOTT  R/REF-H
TWILIGHT ZONE THE MOVIE  BLOCH, ROBERT  C/P:2,P

TWILIGHT ZONE: THE ORIGINAL STORIES, THE GREENBERG, MARTIN H./ MATHESON, RICHARD/ WAUGH, CHARLES G.

HA,LPA

TWILIGHT’S END [STAR TREK]  OLTION, JERRY  C/P,P
TWILIGHT’S KINGDOMS  ASIRE, NANCY  P,SFWA-P
TWIN CITIES RUN [@ENDWORLD #3]  ROBBINS, DAVID  P
TWIN DILEMMA, THE [DOCTOR WHO]  SAWARD, ERIC  SR-P,C/P
TWIN PLANETS  HIGH, PHILIP E.  C/P:2,P
TWIN WORLDS [PROFESSOR JAMESON 4]  JONES, NEIL R.  C/P:2,P
TWINKLE TWINKLE LITTLE STAR  BARZMAN, BEN  H:2
TWIST, THE  CALDER, RICHARD  P
TWIST OF FAITH [STAR TREK DS9]  POCKET BOOKS  LPA
TWIST OF FATE: THE LOCKET [@BLACKSTONE #2]  SAUL, JOHN  P
TWISTED  CONKLIN, GROFF  PA,C/PA
TWISTED CROSS, THE [WINGMAN 5]  MALONEY, MACK  P
TWISTED DRAGON [ELFWOOD]  STEIN, KEVIN  P
TWISTED IMAGES  D’AMMASSA, DON  B
TWISTED MEN, THE  VAN VOGT, A. E.  DF253
TWISTING THE ROPE  MACAVOY, R. A.  C/P:2,P
TWISTOR  CRAMER, JOHN  C/P,P,H
TWISTS IN TIME  LEINSTER, MURRAY  C/P:2,P
TWO [GERALD KNAVE]  JANIFER, LAURENCE M.  LP
TWO-BIT HEROES [IVORY 2]  EGAN, DORIS  C/P,P
TWO CAME CALLING  DORER, NANCY/DORER, FRANCES  P
TWO CROWNS FOR AMERICA  KURTZ, KATHERINE  H,P
TWO DAYS OF TERROR  SHELDON, ROY  SR-P
TWO DOCTORS, THE [DOCTOR WHO]  HOLMES, ROBERT  SR-P,C/P
TWO DOZEN DRAGON EGGS  WOLLHEIM, DONALD A.  SR-P
TWO-EDGED SWORD, A [TWO-EDGED SWORD 1]  MARTIN, THOMAS K.  P
TWO-EYES  GORDON, STUART  P
TWO FACES OF TOMORROW, THE  HOGAN, JAMES P.  C/P:2,P
TWO-FRONT WAR, THE [@STAR TREK NEW FRONTIER #3]  DAVID, PETER  C/P,P
TWO GEORGES, THE  DREYFUSS, RICHARD/TURTLEDOVE, HARRY C/P,P,H
TWO-HANDED ENGINE  KUTTNER, HENRY/MOORE, C. L.  H
TWO HAWKS FROM EARTH  FARMER, PHILIP JOSE  C/P:2,P
TWO HOURS TO DARKNESS  TREW, ANTHONY  P
TWO HUNDRED MILLION A.D.  VAN VOGT, A. E.  C/P:2,P
200 YEARS TO CHRISTMAS  MCINTOSH, J. T.  DF113
TWO IN HIDING, THE [@NIGHT-THREADS #2]  EMERSON, RU  C/P,P
TWO IN THE FIELD [SAMUEL CLEMENS FOWLER 2]  BROCK, DARRYL  LP

The MITSFS Titledex  September 24, 2007  Page 385
TITLE  AUTHOR  CODE
TWO KINGDOMS, THE [ISLANDIA]  SAXTON, MARK  H:2
TWO OF MINDS [STAR*DRIVE]  KEITH, WILLIAM H. JR.  P
TWO OF THEM, THE  RUSS, JOANNA  H,P
TWO PLANETS  LASSWITZ, KURD  C/P,P,H
TWO QUEENS OF LOCHRIN  CREIGHTON, LEE  P
TWO-SPACE WAR, THE  GROSSMAN, DAVE/FRANKOWSKI, LEO  H
TWO SWORDS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #3]  SALVATORE, R. A.  H
TWO TALES AND EIGHT TOMORROWS  HARRISON, HARRY  C/P:2,P
2061: ODYSSEY THREE [@SPACE ODYSSEY #3]  CLARKE, ARTHUR C.  H:2,P
2001: A SPACE ODYSSEY [@SPACE ODYSSEY #1]  CLARKE, ARTHUR C.  C/P,SFWA-P,H:2,SR-H,P
2001: A SPACE ODYSSEY (REVISED) [@SPACE ODYSSEY #1]  CLARKE, ARTHUR C.  P

2001: A SPACE ODYSSEY THE FINAL CHAPTERS (LP RECORD)

[SPACE ODYSSEY] CLARKE, ARTHUR C.

SR-MM

2010: ODYSSEY TWO [@SPACE ODYSSEY #2]  CLARKE, ARTHUR C.  H:2,P
2010: ODYSSEY TWO (PROOF) [@SPACE ODYSSEY #2]  CLARKE, ARTHUR C.  SR-LP
2000 YEARS ON  STATTEN, VARGO  SR-P
TWO-TIMERS, THE  SHAW, BOB  C/P:2,P
TWO TINY CLAWS  DAVIS, BRETT  P
TWO TO CONQUER [DARKOVER]  BRADLEY, MARION ZIMMER  C/P:2,P
TWO TOWERS, THE [@LORD OF THE RINGS #2]  TOLKIEN, J. R. R.  H:3,SR-H,P
TWO TOWERS, THE (ACE) [@LORD OF THE RINGS #2]  TOLKIEN, J. R. R.  SR-P

TWO VIEWS OF WONDER SCORTIA, THOMAS N./

YARBRO, CHELSEA QUINN

C/PA,PA

TWO WORLDS  LORRAINE, PAUL  SR-P,H
TYMORA’S LUCK [@FORGOTTEN REALMS LOST GODS]  NOVAK, KATE/GRUBB, JEFF  P
TYPEWRITER IN THE SKY  HUBBARD, L. RON  SR-P
TYPEWRITER IN THE SKY/FEAR  HUBBARD, L. RON  H:2
TYPHON’S CHILDREN [TYPHON 1]  ANZETTI, TONI  P
TYRANNOSAUR  DRAKE, DAVID  C/P,P
TYRANNOSAURUS REX [AGE OF DINOSAURS 1]  RIVKIN, J. F.  P,SFWA-P
TYRANNY OF THE NIGHT, THE [@INSTRUMENTALITIES OF THE NIGHT #1]  COOK, GLEN  H
TYRANOPOLIS  VAN VOGT, A. E.  SR-P
TYRANT, THE [GENERAL]  FLINT, ERIC/DRAKE, DAVID  P
TYRANT OF HADES, THE [CAGEWORLD 3]  KAPP, COLIN  C/P,P
TYRANT OF TIME  ESHBACH, LLOYD ARTHUR  SR-H
TYRANT’S TEST [@STAR WARS BLACK FLEET CRISIS 3]  KUBE-MCDOWELL, MICHAEL P.  C/P,P
TZIMISCE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  GRIFFIN, ERIC  P

The MITSFS Titledex  September 24, 2007  Page 387
TITLE  AUTHOR  CODE
UBER UNS DAS NICHTS  SCHEER, K. H.  FP
UBIK  DICK, PHILIP K.  H:3,P
U.F.O. 517  FANE, BRON  SR-P
UFO ABDUCTIONS  ROGO, D. SCOTT  [C/RAN-P]
UFO CRASH AT ROSWELL  RANDLE, KEVIN D./SCHMITT, DONALD R.  [C/RAN-P]
UFO DEFENSE [@X-COM]  DUANE, DIANE  P
UFO ENIGMA, THE  MENZEL, DONALD H./TAVES, ERNEST H.  SFWA
UFO FILES, THE  GREENBERG, MARTIN H.  PA
UFO HANDBOOK, THE  HENDRY, ALLAN  SFWA
UGGLIANS, THE  FALLAW, L. M.  H
UGLY LITTLE BOY, THE  ASIMOV, ISAAC  SFWA-TD9,TD9
UGLY LITTLE BOY, THE  ASIMOV, ISAAC/SILVERBERG, ROBERT  C/P,P,H
UGLY SWANS, THE  STRUGATSKY, ARKADY/STRUGATSKY, BORIS H:2,LP
UHURA’S SONG [STAR TREK]  KAGAN, JANET  C/P:3,SFWA-P,P
UITERSTE MOGELIJKHEDEN VAN DE NATUURWETENSCHAP, DE  PYKE, MAGNUS  FP
UITVINDINGEN VAN MORGEN  BERGIER, JACQUES/DE LATIL, PIERRE  FP
ULLER UPRISING  PIPER, H. BEAM  C/P:2,P
ULTIMA GENESI (ITALIAN) [XENOGENESIS 1]  BUTLER, OCTAVIA E.  FP
ULTIMA ONE  STOUGHTON, RICHARD P.  P
ULTIMATE, THE  FEAR, W. H.  SR-P,C/P
ULTIMATE ALIEN, THE  PREISS, BYRON  LPA
ULTIMATE BIKE PATH, THE [BICYCLING THROUGH SPACE AND TIME 2]  SIROTA, MIKE  C/P,SFWA-P,P
ULTIMATE DINOSAUR, THE  PREISS, BYRON/SILVERBERG, ROBERT  SR-VLHA
ULTIMATE DRACULA, THE  PREISS, BYRON  HA,LPA
ULTIMATE DRAGON, THE  PREISS, BYRON  LPA
ULTIMATE DRAGON, THE (PROOF)  PREISS, BYRON  SR-LPA

ULTIMATE EGOIST, THE [@COMPLETE STORIES OF THEODORE STURGEON 1] STURGEON, THEODORE

LP

ULTIMATE ENCYCLOPEDIA OF SCIENCE FICTION, THE  PRINGLE, DAVID  R/REF-H
ULTIMATE ENEMY, THE [BERSERKER 4]  SABERHAGEN, FRED  C/P:2,SFWA-P,P
ULTIMATE FOE, THE [DOCTOR WHO]  BAKER, PIP/BAKER, JANE  P
ULTIMATE FRANKENSTEIN, THE  PREISS, BYRON  HA,LPA
ULTIMATE GUIDE TO SCIENCE FICTION, THE  PRINGLE, DAVID  SFWA
ULTIMATE HALLOWEEN, THE  KAYE, MARVIN  PA
ULTIMATE HELM, THE [@SPELLJAMMER CLOAKMASTER #6]  HOWARD, RUSS T.  C/P,P
ULTIMATE HITCHHIKER’S GUIDE, THE [HITCHHIKERS 1,2,3,4,5]  ADAMS, DOUGLAS  H,LP
ULTIMATE HULK, THE [MARVEL COMICS]  LEE, STAN/DAVID, PETER  LPA
ULTIMATE INVADER AND OTHER SCIENCE FICTION, THE  WOLLHEIM, DONALD A.  SR-DD044
ULTIMATE MALLWORLD, THE [MALLWORLD]  SOMTOW, S. P.  LP
ULTIMATE PRIZE, THE [@CYBORG COMMANDO #3]  O’NEILL, PAMELA/MOHAN, KIM  P
ULTIMATE SILVER SURFER, THE [MARVEL COMICS]  LEE, STAN  LPA
ULTIMATE SOLUTION, THE  NORDEN, ERIC  P
ULTIMATE SPIDER-MAN, THE [MARVEL COMICS]  LEE, STAN  LPA
ULTIMATE SUPER-VILLAINS, THE [MARVEL COMICS]  LEE, STAN  LPA
ULTIMATE THRESHOLD, THE  GINSBURG, MIRRA  C/PA,PA
ULTIMATE TREASURE, THE [DOCTOR WHO PAST DOCTORS]  BULIS, CHRISTOPHER  C/P:2,P
ULTIMATE UNAUTHORIZED STAR TREK QUIZ BOOK, THE [STAR TREK]  BLY, ROBERT W.  [C/RAN-P]
ULTIMATE WEAPON, THE  CAMPBELL, JOHN W. JR.  C/P:2,DG585
ULTIMATE WEREWOLF, THE  PREISS, BYRON  HA,LPA
ULTIMATE WORLD  GERNSBACK, HUGO  H,SFWA,LP
ULTIMATUM  MACCLURE, VICTOR  H
ULTIMATUM  ROHMER, RICHARD  P
ULTIMATUM IN 2050 A.D.  SHARKEY, JACK  DM117
ULTIMATUM: PU 94  DAN, URI/MANN, PETER  P
ULTIMAX MAN, THE  LAUMER, KEITH  H:2,P
ULTRA SPECTRUM  STATTEN, VARGO  SR-P

ULTRAMARINES OMNIBUS, THE [@WARHAMMER 40K ULTRAMARINES 1,2,3] MCNEILL, GRAHAM

LP

UMBRELLA CONSPIRACY, THE [@RESIDENT EVIL #1]  PERRY, S. D.  P
UN LUN DUN  MIEVILLE, CHINA  H
UN-MAN AND OTHER NOVELLAS  ANDERSON, POUL  DF139
UNACCOMPANIED SONATA  CARD, ORSON SCOTT  B,H:2,P,SFWA
UNAUTHORIZED AUTOBIOGRAPHIES  RESNICK, MIKE  B
UNAUTHORIZED HISTORY OF TREK, THE  VAN HISE, JAMES  R/REF-P

UNAUTHORIZED TREKKERS’ GUIDE TO THE NEXT GENERATION AND DEEP SPACE NINE, THE VAN HISE, JAMES

R/REF-P,C/REF-P

UNBEGOTTEN, THE  CREASEY, JOHN  P
UNBEHEADED KING, THE [RELUCTANT KING 3]  DE CAMP, L. SPRAGUE  C/P,SFWA,P,H
UNBIDDEN, THE  CHETWYND-HAYES, R.  P
UNBLESSED, THE  RICHARDS, PAUL  P
UNBORN, THE  SHOBIN, DAVID  H,P
UNBROKEN WEB, THE  ADAMS, RICHARD  SR-VLH
UNCANNY, THE  KLAVAN, ANDREW  H
UNCENSORED MAN, THE  SELLINGS, ARTHUR  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 388
TITLE  AUTHOR  CODE
UNCERTAINTY PRINCIPLE, THE  BILENKIN, DMITRI  H,LP
UNCHARTED STARS [JERN MURDOCK 2]  NORTON, ANDRE  H:2,P
UNCHARTED TERRITORY  WILLIS, CONNIE  C/P:2,P
UNCLE SAM’S CHURCH  BOUTON, JOHN BELL  [C/RAN-H]
UNCLEAN [@FORGOTTEN REALMS HAUNTED LANDS 1]  BYERS, RICHARD LEE  P
UNCOLLECTED PROSE AND POETRY  LOVECRAFT, H. P.  SR-LP

UNCOLLECTED STARS ANTHONY, PIERS/MALZBERG, BARRY N./

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

PA,C/PA

UNCONFINED, THE  FANTHORPE, R. L.  SR-P
UNCONQUERABLE, THE [@BOLOS #2]  FAWCETT, BILL  PA,C/PA
UNCONQUERED COUNTRIES  RYMAN, GEOFF  H
UNCONQUERED COUNTRY, THE  RYMAN, GEOFF  P
UNCROWNED KING, THE [@SUN SWORD #2]  WEST, MICHELLE  P
UNDEAD AND UNAPPRECIATED [BETSY TAYLOR 3]  DAVIDSON, MARYJANICE  H,P
UNDEAD AND UNEASY [BETSY TAYLOR 6]  DAVIDSON, MARYJANICE  H
UNDEAD AND UNEMPLOYED [BETSY TAYLOR 2]  DAVIDSON, MARYJANICE  P
UNDEAD AND UNPOPULAR [BETSY TAYLOR 5]  DAVIDSON, MARYJANICE  H:2
UNDEAD AND UNRETURNABLE [BETSY TAYLOR 4]  DAVIDSON, MARYJANICE  H
UNDEFEATED, THE  LAUMER, KEITH  C/P:2,P
UNDER A CALCULATING STAR  MORRESSY, JOHN  C/P,SFWA,P,H
UNDER A KILLING MOON [TEX MURPHY]  CONNERS, AARON  P
UNDER AFRICAN SKIES [@FUTURE EARTHS]  RESNICK, MIKE/DOZOIS, GARDNER  C/PA,PA
UNDER ALIEN STARS  SERVICE, PAMELA F.  C/P,SFWA,P
UNDER COVER OF DARKNESS  CZERNEDA, JULIE E./PANICCIA, JANA  PA

UNDER FALLEN STARS [@FORGOTTEN REALMS THREAT FROM THE SEA #2] ODOM, MEL

P

UNDER FIRE [@K-9 CORPS]  VON GUNDEN, KENNETH  C/P,P
UNDER HEAVEN’S BRIDGE  BISHOP, MICHAEL/WATSON, IAN  C/P:2,P
UNDER OLD EARTH  SMITH, CORDWAINER  P
UNDER-PEOPLE, THE  NORMAN, ERIC  [C/RAN-P]
UNDER PLUM LAKE  DAVIDSON, LIONEL  C/P,P,H
UNDER PRESSURE  HERBERT, FRANK  C/P:2,P
UNDER SOUTH AMERICAN SKIES [@FUTURE EARTHS]  RESNICK, MIKE/DOZOIS, GARDNER  PA
UNDER THE ANDES  STOUT, REX  C/P,P,H
UNDER THE CANOPY  PAUL, BARBARA  C/P,P
UNDER THE CITY OF ANGELS  BROWN, JERRY EARL  C/P,P
UNDER THE EYE OF GOD [PHAESTOR 1]  GERROLD, DAVID  C/P:2,P
UNDER THE EYE OF NIGHT  MILLS, ROBERT E.  P
UNDER THE GREEN STAR [GREEN STAR 1]  CARTER, LIN  C/P:2,P
UNDER THE HEALING SIGN [BJ VAUGHAN 2]  O’DONOHOE, NICK  P
UNDER THE INFLUENCE  KERR, GEOFFREY  C/P:2,P
UNDER THE MOONS OF MARS  MOSKOWITZ, SAM  HA,R/REF-H
UNDER THE OCEAN TO THE SOUTH POLE  ROCKWOOD, ROY  H
UNDER THE OVERTREE  MOORE, JAMES A.  LP
UNDER THE SEA WIND  CARSON, RACHEL L.  [C/RAN-P]
UNDER THE SHADOW [@MOONRUNNER 1]  TOOMBS, JANE  C/P:2,P
UNDER THE STARS OF DRUUFON [@PERRY RHODAN #68]  ACE  C/PA,PA
UNDER THE TRIPLE SUNS  COBLENTZ, STANTON A.  SR-H,H
UNDER THE WHEEL [ALIEN STARS 3]  MITCHELL, ELIZABETH  PA
UNDER THE YOKE [DRAKA 2]  STIRLING, S. M.  C/P,SFWA-P,P
UNDER TWIN SUNS [@SPACEWAYS #8]  CLEVE, JOHN  P
UNDERGROUND MAN  DE TARDE, GABRIEL  LP
UNDERGROUND PICNIC, THE [@FLYING NUN #5]  JOHNSTON, WILLIAM  P
UNDERLAND  FARREN, MICK  H
UNDERPEOPLE, THE [INSTRUMENTALITY OF MANKIND]  SMITH, CORDWAINER  C/P,P
UNDERSEA [FINNBRANCH 2]  HAZEL, PAUL  P
UNDERSEA CITY [UNDERSEA 3]  POHL, FREDERIK/WILLIAMSON, JACK  H:2,P
UNDERSEA FLEET [UNDERSEA 2]  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
UNDERSEA QUEST [UNDERSEA 1]  POHL, FREDERIK/WILLIAMSON, JACK  C/P:2,P
UNDERSEA TRILOGY, THE [UNDERSEA 1,2,3]  POHL, FREDERIK/WILLIAMSON, JACK  SFWA-P
UNDERSTANDING MU  SANTESSON, HANS STEFAN  [C/RAN-P:2]
UNDERSTANDING TOLKIEN AND THE LORD OF THE RINGS  READY, WILLIAM  R/REF-P,C/REF-P
UNDERWATER MENACE, THE [DOCTOR WHO]  ROBINSON, NIGEL  SR-P,C/P
UNDERWORLD [@UNDERWORLD 1]  COX, GREG  P
UNDERWORLD [@RESIDENT EVIL #4]  PERRY, S. D.  P
UNDERWORLD EVOLUTION [@UNDERWORLD 2]  COX, GREG  P
UNDERWORLD OF ZELLO  DEEGAN, JON J.  SR-P
UNDESIRED PRINCESS, THE/ENCHANTED BUNNY, THE  DE CAMP, L. SPRAGUE/DRAKE, DAVID  C/P:2,SFWA-P,P
UNDISCOVERED NURSING HOME, THE [@STAR WRECK #5]  REWOLINSKI, LEAH  P
UNDYING FIRE, THE  PRATT, FLETCHER  C/P,P,H
UNDYING FIRE, THE  WELLS, H. G.  H
UNDYING LAND, THE  GILMOUR, WILLIAM  H
UNDYING MONSTER, THE  KERRUISH, JESSIE DOUGLAS  P

The MITSFS Titledex  September 24, 2007  Page 389
TITLE  AUTHOR  CODE
UNDYING WIZARD, THE  OFFUTT, ANDREW J.  C/P:2,P
UNDYING WORLD [@BLADE #8]  LORD, JEFFREY  C/P,P
UNEARTH PEOPLE, THE  NEVILLE, KRIS  C/P:2,P
UNEARTHED  MCCONNELL, ASHLEY  P
UNEARTHLY, THE  HICHENS, ROBERT  H
UNEARTHLY, THE  SINGER, KURT  [C/RAN-P]
UNEARTHLY NEIGHBORS  OLIVER, CHAD  H:2,P
UNEASY ALLIANCES [@THIEVES’ WORLD #11]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
UNENDING NIGHT, THE  SMITH, GEORGE H.  C/P:2,P
UNENDLICHE RAUM, DER  SCHEER, K. H.  FP
UNEXPECTED!, THE  CERF, BENNETT  SR-PA,C/PA
UNEXPECTED, THE  MARGULIES, LEO  C/PA,PA
UNEXPECTED APPRENTICE, AN  NYE, JODY LYNN  H
UNEXPECTED DIMENSION, THE  BUDRYS, ALGIS  C/P:2,P
UNEXPECTED DRAGON, THE [UNEXPECTED DRAGON 1,2,3]  BROWN, MARY  H
UNEXPECTED MAGIC  JONES, DIANA WYNNE  H
UNFAIR FARE AFFAIR, THE [@MAN FROM UNCLE #18]  LESLIE, PETER  C/P,P
UNFAMILIAR TERRITORY  SILVERBERG, ROBERT  C/P,SFWA,P,H
UNFINISHED TALES  TOLKIEN, J. R. R.  H:2,P
UNFORSAKEN HIERO, THE  LANIER, STERLING E.  C/P,P,H
UNFORTUNATE FURSEY, THE  WALL, MERVYN  H
UNFROZEN, THE  DREYFUSS, ERNST  P
UNHANDSOME PRINCE, THE  MOORE, JOHN  H
UNHOLY CITY, THE  FINNEY, CHARLES G.  P
UNHOLY TRINITY  RUSSELL, RAY  P
UNICORN, THE [FIVE RINGS CLAN WAR 2]  LASSIEUR, A. L.  P
UNICORN AFFAIR, THE  FRITZHAND, JAMES/GLICKSMAN, FRANK  P
UNICORN AND DRAGON [@UNICORN AND DRAGON #1]  ABBEY, LYNN  H:2,LP
UNICORN CREED, THE [ARGONIA 2]  SCARBOROUGH, ELIZABETH ANN  C/P:2,P
UNICORN DANCER, THE  SALSITZ, R. A. V.  P
UNICORN DILEMMA, THE [STRAND 2]  LEE, JOHN  P
UNICORN GAMBIT, THE [GAMING MAGI 3]  BISCHOFF, DAVID F.  C/P,P
UNICORN GIRL, THE  KURLAND, MICHAEL  C/P:2,P
UNICORN HIGHWAY  JONES, DAVID LEE  P
UNICORN MOUNTAIN  BISHOP, MICHAEL  H,P
UNICORN PEACE, THE [STRAND 4]  LEE, JOHN  P
UNICORN POINT [APPRENTICE ADEPT #6]  ANTHONY, PIERS  C/P,SFWA-P,@P,H
UNICORN QUEST, THE [STRAND 1]  LEE, JOHN  C/P,P
UNICORN SOLUTION, THE [STRAND 3]  LEE, JOHN  C/P:2,SFWA-P,P
UNICORN SONATA, THE  BEAGLE, PETER S.  H
UNICORN TRADE, THE  ANDERSON, POUL/ANDERSON, KAREN  C/P:2,P
UNICORN U. [GNOME MAN’S LAND 3]  FRIESNER, ESTHER M.  C/P,SFWA-P:2,P
UNICORN VARIATIONS  ZELAZNY, ROGER  H:2,P
UNICORN WAR, THE [STRAND 5]  LEE, JOHN  P
UNICORNS!  DANN, JACK/DOZOIS, GARDNER  C/PA,PA
UNICORNS 2  DANN, JACK/DOZOIS, GARDNER  PA
UNICORNS IN THE RAIN  COHEN, BARBARA  H
UNIDENTIFIED, THE  BARKER, GRAY  [C/RAN-P]
UNIDENTIFIED FLYING ODDBALL  CRUME, VIC  P
UNIFICATION [STAR TREK TNG]  TAYLOR, JERI  C/P:3,P
UNIFIED FIELD, THE  LEON, MARK  P
UNIFYING FORCE, THE [@STAR WARS NEW JEDI ORDER]  LUCENO, JAMES  P

UNIFYING FORCE, THE (HARDCOVER AND CD)

[@STAR WARS NEW JEDI ORDER] LUCENO, JAMES

H

UNINVITED, THE  FARRIS, JOHN  SFWA-P
UNINVITED, THE  MACARDLE, DOROTHY  H
UNINVITED, THE  MULLER, JOHN E.  SR-P
UNION CLUB MYSTERIES, THE  ASIMOV, ISAAC  C/P:2,P
UNION FIRES [@TIMERAIDER]  BARNES, JOHN  P
UNION FOREVER [@LOST REGIMENT 2]  FORSTCHEN, WILLIAM R.  C/P:2,P
UNISAVE  MADSEN, AXEL  P
UNITED PLANETS, THE  WADEY, VICTOR  H
UNITED STATES AND CHINA, THE  FAIRBANK, JOHN KING  [C/RAN-P]
UNITY [@BATTLESTAR GALACTICA 2003]  HARPER, STEVEN  LP
UNITY [@STAR TREK DS9]  PERRY, S. D.  P
UNIVERS 01  SADOUL, JACQUES  FP
UNIVERS 15  SADOUL, JACQUES  FP
UNIVERSAL BASEBALL ASSOCIATION, INC.: J. HENRY WAUGH, PROP., THE  COOVER, ROBERT  C/P,P
UNIVERSAL DATABANK, THE [DOCTOR WHO]  LOFFICIER, JEAN MARC  R/REF-P
UNIVERSAL PREY, THE  SWYCAFFER, JEFFERSON P.  P
UNIVERSAL STATION  BROWN, BETH  H
UNIVERSE, THE  ASIMOV, ISAAC  [C/RAN-P]
UNIVERSE  HEINLEIN, ROBERT A.  C/P,B,TD35,P
UNIVERSE, THE  PREISS, BYRON  SR-VLHA

TITLE  AUTHOR  CODE
UNIVERSE 1  CARR, TERRY  PA,C/PA
UNIVERSE 1 (SECOND SERIES)  SILVERBERG, ROBERT/HABER, KAREN  PA
UNIVERSE 2  CARR, TERRY  PA,C/PA
UNIVERSE 2 (SECOND SERIES)  SILVERBERG, ROBERT/HABER, KAREN  HA,LPA
UNIVERSE 3  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 3 (SECOND SERIES)  SILVERBERG, ROBERT/HABER, KAREN  C/PA,PA
UNIVERSE 4  CARR, TERRY  C/PA,PA
UNIVERSE 5  CARR, TERRY  PA,C/PA:2
UNIVERSE 6  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 7  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 8  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 9  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 10  CARR, TERRY  SFWA,PA,HA
UNIVERSE 11  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 12  CARR, TERRY  C/PA,PA,SFWA
UNIVERSE 13  CARR, TERRY  C/PA,PA,SFWA-P,SFWA
UNIVERSE 14  CARR, TERRY  C/PA,SFWA-P,PA,SFWA
UNIVERSE 15  CARR, TERRY  HA,PA,SFWA-P,SFWA
UNIVERSE 16  CARR, TERRY  HA,SFWA
UNIVERSE 17  CARR, TERRY  HA
UNIVERSE AGAINST HER, THE [TELZEY 1]  SCHMITZ, JAMES H.  C/P:2,P
UNIVERSE AND DOCTOR EINSTEIN, THE  BARNETT, LINCOLN  [C/RAN-P:2]
UNIVERSE BETWEEN, THE  NOURSE, ALAN E.  C/P:2,P
UNIVERSE DAY  O’DONNELL, K. M.  C/P:2,P
UNIVERSE MAKER, THE  VAN VOGT, A. E.  SR-DD031,P,C/P:2
UNIVERSE MAKERS, THE  WOLLHEIM, DONALD A.  R/REF-H
UNIVERSE NEXT DOOR, THE [@SCHRODINGER’S CAT 1]  WILSON, ROBERT ANTON  C/P,P
UNIVERSES OF E. E. SMITH, THE  ELLIK, RON/EVANS, BILL  R/REF-H
UNKINDNESS OF RAVENS, AN [LADY BRANWEN 1]  MEANEY, DEE MORRISON  P
UNKNOWN, THE  BENSEN, DONALD R.  PA,C/PA
UNKNOWN, THE  LEINSTER, MURRAY  B
UNKNOWN  SCHMIDT, STANLEY  PA,SFWA-P
UNKNOWN DANGER [LAND OF THE GIANTS 3]  LEINSTER, MURRAY  P
UNKNOWN DESTINY  FANE, BRON  SR-P
UNKNOWN FIVE, THE  BENSEN, DONALD R.  PA,C/PA
UNKNOWN GUEST, THE  MAETERLINCK, MAURICE  H
UNKNOWN — IS IT NEARER?, THE  DINGWALL, ERIC J./LANGDON-DAVIES, JOHN [C/RAN-P]
UNKNOWN LEWIS CARROLL, THE  CARROLL, LEWIS  LP
UNKNOWN REGIONS  HOLDSTOCK, ROBERT P.  P
UNKNOWN SECTOR: MILKY WAY [@PERRY RHODAN #45]  ACE  PA,C/PA
UNKNOWN SHORE, THE  MALCOLM, DONALD  SR-P
UNKNOWN SOLDIER, THE  REICHERT, MICKEY ZUCKER  C/P,SFWA-P,P
UNKNOWN WORLDS  SCHMIDT, STANLEY/GREENBERG, MARTIN H. HA,LPA
UNLEASH THE NIGHT [@DARK-HUNTER]  KENYON, SHERRILYN  P
UNLEASHED IN SPACE [WISHBONE 3]  STEELE, ALEXANDER  LP
UNLEASHED POWERS [@PERRY RHODAN #90]  ACE  C/PA,PA
UNLESS SHE BURN [@AREIA DARENGA #2]  MEZO, FRANCINE  P
UNLIKELY GHOSTS, THE  TURNER, JAMES  SR-PA,HA
UNLIKELY ONES, THE [UNEXPECTED DRAGON]  BROWN, MARY  H:2,SFWA-P,P
UNLIMITED DREAM COMPANY, THE  BALLARD, J. G.  C/P,P,H
UNLOCKING THE AIR AND OTHER STORIES  LE GUIN, URSULA K.  LP
UNLOVED, THE  SAUL, JOHN  P
UNMAN, THE/KOVRIGIN’S CHRONICLES  SHEFNER, VADIM  H,LP
UNMAN, THE/KOVRIGIN’S CHRONICLES (PROOF)  SHEFNER, VADIM  SR-LP
UNNATURAL DIPLOMACY [@ISAAC’S UNIVERSE #3]  GREENBERG, MARTIN H.  PA,SFWA-P
UNNATURAL HISTORY [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P,P
UNNATURAL SELECTION [BUFFY]  ODOM, MEL  P
UNNATURAL TALENT [DARK FORCES 7]  LOGAN, LES  P
UNO CONTRO HERCULUM  SOHL, JERRY  FP
UNPLEASANT PROFESSION OF JONATHAN HOAG, THE  HEINLEIN, ROBERT A.  C/P:2,P
UNPOPULAR PLANET  SMITH, EVELYN E.  C/P:2,P
UNQUENCHABLE FIRE [LIVING WORLD 1]  POLLACK, RACHEL  H,LP
UNQUIET GRAVE, THE [@DURHAM RED]  EVANS, PETER J.  P
UNQUIET SPIRIT, THE  STEEN, MARGUERITE  C/P,P
UNREAL PEOPLE, THE  SIEGEL, MARTIN  C/P,P
UNREASONING MASK, THE  FARMER, PHILIP JOSE  C/P,P,H
UNSAFE AT ANY SPEED  NADER, RALPH  [C/RAN-P]
UNSCHOOLED WIZARD, THE [SUNWOLF AND STARHAWK 1,2]  HAMBLY, BARBARA  H:2
UNSEEN ASSASSIN, THE  JANSON, HANK  SR-P
UNSEEN KING, THE  BLUE, TYSON  R/REF-P
UNSEEN TORMENT  KIMBROUGH, KATHERYN  P
UNSHAPELY THINGS [CONNOR GREY 1]  DEL FRANCO, MARK  P
UNSILENT NIGHT  LEE, TANITH  SR-H
UNSLEEP, THE  GILLON, DIANA/GILLON, MEIR  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 391
TITLE  AUTHOR  CODE
UNSLEEPING EYE, THE  COMPTON, D. G.  C/P,P
UNSPEAKABLE PEOPLE, THE  HAINING, PETER  PA
UNTELEPORTED MAN, THE  DICK, PHILIP K.  C/P:2,DG602,D15697,P

UNTHREATENED BY THE MORNING LIGHT [@AUTHOR’S CHOICE MONTHLY #2] WAGNER, KARL EDWARD

SR-H

UNTIL RELIEVED [LUCKY 13TH]  SHELLEY, RICK  P
UNTIL THE CELEBRATION [GREEN-SKY 3]  SNYDER, ZILPHA KEATLEY  P
UNTO THE BEAST  MONACO, RICHARD  P
UNTO THE LAST GENERATION  COULSON, JUANITA  SR-P
UNTO THE LAST GENERATION  EASTON, THOMAS A.  LP
UNTO ZEOR, FOREVER [SIME/GEN 2]  LICHTENBERG, JACQUELINE  C/P,P,H,SFWA
UNTOLD LEGEND OF THE BATMAN, THE  WEIN, LEN  SFWA-P
UNTOUCHABLE [WEB,THE]  BROWN, ERIC  P
UNTOUCHED BY HUMAN HANDS  SHECKLEY, ROBERT  C/P:2,P
UNTRODDEN STREETS OF TIME  FALKNER, JOHN  SR-P
UNWILLING DICTATOR, THE  VERNE, JULES  H
UNWILLING WARLORD, THE [ETHSHAR]  WATT-EVANS, LAWRENCE  P
UNWILLINGLY TO EARTH  ASHWELL, PAULINE  P
UNWISE CHILD  GARRETT, RANDALL  H:2
UNWOUND WAY, THE  ADAMS, BILL/BROOKS, CECIL  P
UP ABOVE THE WORLD  BOWLES, PAUL  P
UP IN A HEAVAL [XANTH #26]  ANTHONY, PIERS  H
UP IN THE HILLS  DUNSANY, LORD  H
UP THE LINE  SILVERBERG, ROBERT  C/P:2,P
UP THE WALL  FRIESNER, ESTHER M.  LP
UP THE WALLS OF THE WORLD  TIPTREE, JAMES JR.  C/P,H:2,P
UP THERE AND OTHER STRANGE DIRECTIONS  WOLLHEIM, DONALD A.  SR-H
UP TO THE SKY IN SHIPS  CHANDLER, A. BERTRAM  SR-DNESFA2
UP-WINGERS  ESFANDIARY, S. M.  [C/RAN-P]
UP YOUR ASTEROID!  COOPER, C. EVERETT  SR-LP
UPLAND OUTLAWS [HANDFUL OF MEN 2]  DUNCAN, DAVE  C/P,P,H
UPLIFT WAR, THE [UPLIFT 3]  BRIN, DAVID  C/P:2,P
UPLINK  FANCHER, JANE S.  C/P,P
UPON THE WINDS OF YESTERDAY  BARR, GEORGE  A
UPSIDE DOWNSIDE  GOULART, RON  P
UPTIME, DOWNTIME  PEEL, JOHN  H
URANIUM 235  MULLER, JOHN E.  SR-P

URBAN NIGHTMARES SHERMAN, JOSEPHA/

DECANDIDO, KEITH R. A.

PA

URBAN SHAMAN [WALKER 1]  MURPHY, C. E.  LP
URCHIN OF THE RIDING STARS [@MISTMANTLE #1]  MCALLISTER, M. I.  H
URSHURAK  HILDEBRANDT, BROTHERS/NICHOLS, JERRY H,SR-LP

URSULA K. LE GUIN OLANDER, JOSEPH D./

GREENBERG, MARTIN H.

R/REF-H

URSULA, THE PROUD [@PHENWICK #34]  KIMBROUGH, KATHERYN  P
URSULA’S GIFT  DISILVESTRO, ROGER L.  H
URSUN’S TEETH [@WARHAMMER VON VELTEN 2]  MCNEILL, GRAHAM  P
URSUS  DVORKIN, DAVID  H
URSUS OF ULTIMA THULE  DAVIDSON, AVRAM  C/P:2,P
URTH OF THE NEW SUN, THE [NEW SUN 5]  WOLFE, GENE  C/P,SFWA-P,P,H
USE OF WEAPONS [CULTURE]  BANKS, IAIN M.  H:2,P
USER FRIENDLY  ROBINSON, SPIDER  C/P:2,P
USHER’S PASSING  MCCAMMON, ROBERT R.  P
U.S.S. ENTERPRISE BRIDGE BLUEPRINTS  MCMASTER, MICHAEL  A
U.S.S.A.  MADSEN, DAVID  H:2
USSR: MISSILES, ROCKETS AND THE SPACE EFFORT  ARMY, DEPARTMENT OF THE  [C/RAN-H]
USUAL LUNACY, A  COMPTON, D. G.  LP
USURPER, THE [@KINGDOMS #2]  WELLS, ANGUS  C/P:2,P
USURPER’S CROWN, THE [ISAVALTA 2]  ZETTEL, SARAH  H
UTHER [CAMULOD]  WHYTE, JACK  P
UTMOST ISLAND, THE  MYERS, HENRY  H
UTOPIA [CALIBAN 3]  ALLEN, ROGER MACBRIDE  H,P
UTOPIA  FLUGGE, FRANK  FP
UTOPIA  MORE, SIR THOMAS  C/P,P
UTOPIA 3  EFFINGER, GEORGE ALEC  P
UTOPIA 14  VONNEGUT, KURT JR.  C/P:2,P
UTOPIA AFFAIR, THE [@MAN FROM UNCLE #15]  MCDANIEL, DAVID  C/P,P
UTOPIA HUNTERS [INQUESTOR 3]  SUCHARITKUL, SOMTOW  C/P:2,P
UTOPIA MINUS X  GORDON, REX  C/P:2,P

TITLE  AUTHOR  CODE
V [V]  CRISPIN, A. C.  C/P:2,P
V FOR VENDETTA  MOORE, ALAN/LLOYD, DAVID  CX
VACANT THRONE, THE [BAND OF FOUR]  GREENWOOD, ED  H,P
VACATION IN SPACE  KELLEY, LEO P.  SR-P
VACATION STORIES  RAMON Y CAJAL, SANTIAGO  H
VACCINATOR, THE  SMITH, MICHAEL MARSHALL  MB002
VACUUM DIAGRAMS  BAXTER, STEPHEN  C/P:2,LP,H
VACUUM FLOWERS  SWANWICK, MICHAEL  C/P,SFWA,P,H:2
VAGABOND OF SPACE [@PERRY RHODAN #93]  ACE  C/PA:2,PA
VAGABONDS OF GOR [GOR 24]  NORMAN, JOHN  P
VAGUES D’INVASION  RANDA, PETER  FP
VAIN COMMAND [@STARCRUISER SHENANDOAH #4]  GREEN, ROLAND J.  P
VAINQUEURS DE VEGA, LES (FRENCH) [@PERRY RHODAN #5]  SCHEER, K. H./DARLTON, CLARK  FPR
VALDAR THE OFT-BORN  GRIFFITH, GEORGE  H
VALDEMAR COMPANION, THE [VALDEMAR]  HELFERS, JOHN/LITTLE, DENISE  R/REF-H
VALDEZ HORSES, THE  HOFFMAN, LEE  P
VALE OF THE VOLE [@XANTH #10]  ANTHONY, PIERS  C/P:2,P
VALENTINA: SOUL IN SAPPHIRE  DELANEY, JOSEPH H./STIEGLER, MARC  C/P,P
VALENTINE PONTIFEX [MAJIPOOR 3]  SILVERBERG, ROBERT  H:2,P
VALENTINE’S EXILE [@VAMPIRE EARTH #5]  KNIGHT, E. E.  H
VALENTINE’S RISING [@VAMPIRE EARTH #4]  KNIGHT, E. E.  P
VALHALLA [STAR TREK DS9]  ARCHER, NATHAN  C/P,P
VALHALLA  HOLT, TOM  P
VALIANT, THE [STAR TREK TNG STARGAZER]  FRIEDMAN, MICHAEL JAN  P
VALIS  DICK, PHILIP K.  C/P:2,P
VALLEE DES MAMMOUTHS, LA  PEYRAMAURE, MICHEL  FH
VALLEE ENSEVELIE, LA  MAGOG, H. -J.  FP
VALLEY BEYOND TIME  SILVERBERG, ROBERT  C/P:2,P
VALLEY OF CREATION, THE  HAMILTON, EDMOND  C/P:2,P
VALLEY OF HORSES, THE [EARTH’S CHILDREN 2]  AUEL, JEAN M.  C/P,P,H
VALLEY OF LIGHTS  GALLAGHER, STEPHEN  SFWA-P
VALLEY OF OGRUM, THE [@ORON #4]  SMITH, DAVID C.  P
VALLEY OF PRETENDERS  CLIVE, DENNIS  B
VALLEY OF SILENCE [CIRCLE 3]  ROBERTS, NORA  H
VALLEY OF SPIDERS, THE  WELLS, H. G.  P
VALLEY OF TERROR  REY, RUSSELL  SR-P
VALLEY OF THE FLAME  KUTTNER, HENRY  C/P:2,P
VALLEY OF THE WORM  HOWARD, ROBERT E.  H
VALLEY OF THUNDER, THE [DUNGEON 3]  DE LINT, CHARLES  C/P:2,P
VALLEY SO LOW, THE  WELLMAN, MANLY WADE  SFWA
VALLEY WHERE TIME STOOD STILL, THE [MAN WHO LOVED MARS 2]  CARTER, LIN  H,P
VALNIR’S BANE [@WARHAMMER BLACK HEARTS 1]  LONG, NATHAN  P
VALOR IN THE ASHES  JOHNSTONE, WILLIAM W.  P
VALORIAN [DARK HORSE 0]  HERBERT, MARY H.  P
VALORIAN’S CHILDREN [DARK HORSE 1,2]  HERBERT, MARY H.  H
VALORIAN’S LEGACY [DARK HORSE 3,4]  HERBERT, MARY H.  H
VALOR’S CHOICE [TORIN KERR 1]  HUFF, TANYA  P
VAMPHYRI! [NECROSCOPE 2]  LUMLEY, BRIAN  SFWA-P:2
VAMPIRE [@WORLD OF DARKNESS PREDATOR & PREY #1]  BOWEN, CARL  P
VAMPIRE$  STEAKLEY, JOHN  P,LP
VAMPIRE, THE  VOLTA, ORNELLA  [C/RAN-P]
VAMPIRE AFFAIR, THE [@MAN FROM UNCLE #6]  MCDANIEL, DAVID  C/P:2,P
VAMPIRE ARMAND, THE [LESTAT 6]  RICE, ANNE  H,LP
VAMPIRE BLOOD  GRIFFITH, KATHRYN MEYER  P
VAMPIRE DETECTIVES  GREENBERG, MARTIN H.  C/PA,PA
VAMPIRE IN MOSCOW [ST. JOHN THE PURSUER]  HENRICK, RICHARD  P
VAMPIRE JOURNALS, THE  BRIERY, TRACI  P
VAMPIRE JUNCTION [TIMOTHY VALENTINE 1]  SOMTOW, S. P.  C/P:2,P
VAMPIRE LESTAT, THE [LESTAT 2]  RICE, ANNE  C/P,P,H
VAMPIRE LESTAT, THE (GRAPHIC NOVEL) [LESTAT 2]  RICE, ANNE/PEROZICH, FAYE  CX
VAMPIRE MASTER, THE  HAMILTON, EDMOND  H
VAMPIRE OF CURITIBA AND OTHER STORIES, THE  TREVISAN, DALTON  H
VAMPIRE OF THE MISTS [@RAVENLOFT]  GOLDEN, CHRISTIE  C/P:2,P
VAMPIRE PAPERS, THE  ROMKEY, MICHAEL  P
VAMPIRE PRINCESS, THE  ROMKEY, MICHAEL  P
VAMPIRE SCIENCE [DOCTOR WHO EIGHTH DOCTOR]  BLUM, JONATHAN/ORMAN, KATE  C/P:2,P
VAMPIRE SEXTETTE, THE  KAYE, MARVIN  HA,PA
VAMPIRE TAPES, THE  RANDOLPHE, ARABELLA  P
VAMPIRE TAPESTRY, THE  CHARNAS, SUZY MCKEE  C/P,SFWA-P,P,H
VAMPIRE WHO LOVED ME, THE  MEDEIROS, TERESA  H
VAMPIRE WOMEN, THE  SAMUELS, VICTOR  C/P,P
VAMPIRES  RYAN, ALAN  HA
VAMPIRE’S APPRENTICE, THE  BYERS, RICHARD LEE  P
VAMPIRE’S MOON  SAXON, PETER  P

The MITSFS Titledex  September 24, 2007  Page 394
TITLE  AUTHOR  CODE
VAMPIRES OF FINISTERE, THE [GUARDIANS (SAXON) 4]  SAXON, PETER  P
VAMPIRES OF NIGHTWORLD  BISCHOFF, DAVID F.  C/P,P
VAMPIRES OF VERMONT [@NEWS FROM THE EDGE 3]  SUMNER, MARK  P
VAMPIRES, SPIES, AND ALIEN BEINGS  AUSTIN, R. G.  P
VAMPIRES, WEREWOLVES AND OTHER MONSTERS  ELWOOD, ROGER  PA
VAMPIRESLAYER [WARHAMMER GOTREK & FELIX 6]  KING, WILLIAM  P
VAMPRIRE ODYSSEY, THE [DANIELLE WALTHERS 1]  CIENCIN, SCOTT  P

VAMPS GREENBERG, MARTIN H./ WAUGH, CHARLES G.

C/PA,PA

VAMPS AND THE CITY  SPARKS, KERRELYN  P
VAMPYR  JENNINGS, JAN  P
VAMPYR-SS [@KNIGHTS OF THE BLOOD 1]  MACMILLAN, SCOTT  C/P,P
VAN HELSING  RYAN, KEVIN  P
VAN MOLECUUL TOT MENS  SLUYSER, M.  FH
VANDALS OF THE VOID  VANCE, JACK  H:2
VANDERDEKEN’S CHILDREN [DOCTOR WHO EIGHTH DOCTOR]  BULIS, CHRISTOPHER  C/P,P
VANEGLORY [ETHICAL SOCIETY 2]  TURNER, GEORGE  P
VANG: THE BATTLEMASTER, THE [VANG]  ROWLEY, CHRISTOPHER  P
VANG: THE MILITARY FORM, THE [VANG]  ROWLEY, CHRISTOPHER  C/P,P
VANGUARD [@BATTLESTATION #2]  DRAKE, DAVID/FAWCETT, BILL  C/PA,PA,SFWA-P:2
VANGUARD FROM ALPHA  ALDISS, BRIAN W.  DD369
VANGUARD TO NEPTUNE  WALSH, J. M.  SR-P
VANGUARD TO VENUS  CASTLE, JEFFERY LLOYD  H
VANISHED [@TIME POLICE #1]  NORWOOD, WARREN  C/P,P
VANISHER, THE [DOC SAVAGE #52]  ROBESON, KENNETH  C/P:2,@P
VANISHING, THE [@LAST ON EARTH #1]  KAYE, MARILYN  P
VANISHING ACTS  DATLOW, ELLEN  HA,LPA
VANISHING ISLAND, THE  O’CONNELL, CHARLES C.  H
VANISHING POINT [DOCTOR WHO EIGHTH DOCTOR]  COLE, STEPHEN  P
VANISHING POINT  ROESSNER, MICHAELA  H:2,P
VANISHING TOWER, THE [ELRIC 4]  MOORCOCK, MICHAEL  C/P,P
VANITAS [TIMOTHY VALENTINE 3]  SOMTOW, S. P.  H
VANITY’S BROOD [@FORGOTTEN REALMS HOUSE OF SERPENTS #3]  SMEDMAN, LISA  P
VAR THE STICK [BATTLE CIRCLE #2]  ANTHONY, PIERS  C/P:2,@P
VARGR-MOON [HATHER LAMBISSON 2]  KING, BERNARD  H
VARIABLE MAN AND OTHER STORIES, THE  DICK, PHILIP K.  C/P:2,P
VARIABLE STAR  HEINLEIN, ROBERT A./ROBINSON, SPIDER  H
VARIATIONS THREE:  LEE, SHARON  SR-LP
VARJAK PAW  SAID, S. F.  LP
VARKAUS CONSPIRACY, THE  DALMAS, JOHN  C/P,SFWA-P,P
VARNEY THE VAMPYRE I [VARNEY #1]  RYMER, JAMES  LP
VARNEY THE VAMPYRE II [VARNEY #2]  RYMER, JAMES  LP
VAST  NAGATA, LINDA  C/P,P
VATHEK  BECKFORD, WILLIAM  C/P,P
VAULT, THE [LOST IN SPACE]  DEWEESE, GENE  P
VAULT OF HORROR, THE  OLECK, JACK  P
VAULT OF THE AGES  ANDERSON, POUL  H:2,SR-H,P
VAULT OF THE AGES (JAPANESE)  ANDERSON, POUL  FH
VAZKOR, SON OF VAZKOR  LEE, TANITH  C/P,P
VECTOR  SUTTON, HENRY  C/P,P
VECTOR ANALYSIS  HALDEMAN, JACK C. II  C/P,P,H
VECTOR PRIME [@STAR WARS NEW JEDI ORDER]  SALVATORE, R. A.  P
VECTOR PRIME (CD) [STAR WARS]  LUCAS BOOKS  SR-MM
VECTORS  KUBE-MCDOWELL, MICHAEL P.  P
VECTORS  SHEFFIELD, CHARLES  C/P,P

VECTORS [@STAR TREK TNG DOUBLE HELIX 2] SMITH, DEAN WESLEY/

RUSCH, KRISTINE KATHRYN

P

VEGA  HUNT, GILL  SR-P
VEGA SECTOR, THE [@PERRY RHODAN #5]  ACE  PA,C/PA
VEGAS VAMPIRE [@CHILL #4]  SHERMAN, JORY  P
VEIL  CHESBRO, GEORGE C.  C/P,P
VEIL OF A THOUSAND TEARS, THE [PEARL 2]  LUSTBADER, ERIC VAN  H
VEIL OF A THOUSAND TEARS, THE (PROOF) [PEARL 2]  LUSTBADER, ERIC VAN  H:2,SR-LP
VEIL OF SHADOW [ELVEN KINGDOMS 2]  COOKE, CATHERINE  P,SFWA-P
VEIL OF YEARS, THE [SORCERER’S TALE 2]  DOUGLAS, L. WARREN  H,P

VEILED ALLIANCES [@SEVEN SUNS] ANDERSON, KEVIN J./TERANISHI, ROBERT/

BROOME, WENDY

CX

VEILED DRAGON, THE [@FORGOTTEN REALMS HARPERS #12]  DENNING, TROY  P
VEILED KNOWLEDGE  JAMES, EDWIN  B
VEILED RAIDERS, THE [@RICK BRANT #20]  BLAINE, JOHN  H
VEILED WEB, THE  ASARO, CATHERINE  P
VEILED WORLD, THE  LORY, ROBERT  P,D31755
VEILS OF AZLAROC, THE  SABERHAGEN, FRED  C/P:2,SFWA-P,P
VELIERO DEI GHIACCI, IL  MOORCOCK, MICHAEL  FP

TITLE  AUTHOR  CODE
VELLUM [BOOK OF ALL HOURS 1]  DUNCAN, HAL  LP
VELOCITY GOSPEL, THE [ACCOMPLICE 2]  AYLETT, STEVE  LP
VELVET SHADOWS  NORTON, ANDRE  P
VELVETEEN RABBIT, THE  WILLIAMS, MARGERY  SFWA-P
VEN Y ENLOQUECE (SPANISH)  BROWN, FREDRIC  FH
VENDETTA [STAR TREK TNG]  DAVID, PETER  C/P:2,P
VENDETTA  MURDOCK, M. S.  P
VENETIAN COURT, THE  HARNESS, CHARLES L.  C/P:2,P
VENGEANCE [@STAR TREK DS9]  AB HUGH, DAFYDD  P
VENGEANCE [@ANGEL]  CIENCIN, SCOTT/JOLLEY, DAN  P
VENGEANCE FANTASTIC  LITTLE, DENISE  PA
VENGEANCE OF DRAGONS [SECRET TEXTS 2]  LISLE, HOLLY  C/P,P
VENGEANCE OF GWA  WRIGHT, S. FOWLER  H
VENGEANCE OF HERA, THE [@HERCULES]  BETANCOURT, JOHN GREGORY  P
VENGEANCE OF MASKS [CHILDERIC THE SHATTERER 1]  EDGHILL, ROSEMARY  H
VENGEANCE OF ORION [ORION 2]  BOVA, BEN  H:2,P
VENGEANCE OF SIVA  MULLER, JOHN E.  SR-P
VENGEANCE OF THE DANCING GODS [DANCING GODS 3]  CHALKER, JACK L.  C/P:2,P
VENGEANCE ON VAROS [DOCTOR WHO]  MARTIN, PHILIP  SR-P,C/P
VENGING, THE  BEAR, GREG  P
VENISS UNDERGROUND  VANDERMEER, JEFF  LP
VENOM OF ARGUS, THE [EXPENDABLES 4]  AVERY, RICHARD  C/P:2,P
VENOM-SEEKERS, THE  BERRY, BRYAN  SR-P
VENOM TREES OF SUNGA, THE  DE CAMP, L. SPRAGUE  P
VENOM’S TASTE [@FORGOTTEN REALMS HOUSE OF SERPENTS #1]  SMEDMAN, LISA  P
VENTURE INWARD  CAYCE, EDGAR  [C/RAN-P]
VENTUS  SCHROEDER, KARL  H,P
VENUS  BOVA, BEN  H
VENUS 13  HITCHCOCK, RAYMOND  SR-P
VENUS BELT, THE [NORTH AMERICAN CONFEDERACY]  SMITH, L. NEIL  C/P,P
VENUS BOY  SUTTON, LEE  H
VENUS DEVELOPMENT  BERGAMINI, DAVID  C/P:2,P
VENUS EQUILATERAL  SMITH, GEORGE O.  C/P,SR-H,P,H
VENUS FACTOR, THE  GHIDALIA, VIC/ELWOOD, ROGER  C/PA,PA
VENUS IN DANGER [@PERRY RHODAN #14]  ACE  PA,C/PA
VENUS, INC.  POHL, FREDERIK/KORNBLUTH, C. M.  H:2
VENUS OF DREAMS [VENUS OF DREAMS 1]  SARGENT, PAMELA  C/P:2,P
VENUS OF SHADOWS [VENUS OF DREAMS 2]  SARGENT, PAMELA  C/P,P,H
VENUS ON THE HALF-SHELL  FARMER, PHILIP JOSE  P
VENUS ON THE HALF-SHELL  TROUT, KILGORE  C/P:2,P
VENUS PLUS X  STURGEON, THEODORE  C/P:2,P
VENUS PRESERVED [@VENUS #4]  LEE, TANITH  LP
VENUS PROBE, THE  SAINT JOHN, DAVID  P
VENUS TRAP, THE [@PERRY RHODAN #17]  ACE  PA
VENUS VENTURE, THE  MULLER, JOHN E.  SR-P
VENUSIAN LULLABY [DOCTOR WHO MISSING ADVENTURES]  LEONARD, PAUL  C/P,P
VERBOTENE SONNENSYSTEM, DAS [@PERRY RHODAN #58]  DARLTON, CLARK  FPR
VERDANT PASSAGE, THE [@DARK SUN PRISM #1]  DENNING, TROY  C/P,P
VERDIGRIS [DOCTOR WHO PAST DOCTORS]  MAGRS, PAUL  C/P,P
VERGIL IN AVERNO  DAVIDSON, AVRAM  H,SFWA
VERMILION SANDS  BALLARD, J. G.  C/P:2,P
VERMINE DU LION, LA  CARSAC, FRANCIS  FP
VERMINTIDE [@WARHAMMER]  LEE, BRUNO  P
VERTIGO  SHAW, BOB  C/P,P
VERUCHIA [DUMAREST 8]  TUBB, E. C.  C/P:2,P
VERY BIG BANG!, A [SLUGGY FREELANCE]  ABRAMS, PETE  CX
VERY FAR AWAY FROM ANYWHERE ELSE  LE GUIN, URSULA K.  P
VERY PRIVATE LIFE, A  FRAYN, MICHAEL  C/P,P
VERY SLOW TIME MACHINE, THE  WATSON, IAN  C/P:2,P
VERY SNEAKY, ANDY CAPP  SMYTHE  [C/RAN-P]
VERY STRANGE TRIP, THE (PROOF)  WOLVERTON, DAVE  SR-LP
VESMIRNE HRY  STOPY  FH
VESTIGES OF TIME [ERIC MATHERS 3]  MEREDITH, RICHARD C.  H,SFWA,P
VETTIUS AND HIS FRIENDS  DRAKE, DAVID  C/P:2,SFWA-P,P
V.I.  DEBRANDT, DON  C/P,P
VIATOR  SHEPARD, LUCIUS  H
VICAR OF CHRIST, THE  MURPHY, WALTER F.  P
VICARION, THE  HUNTING, GARDNER  H
VICE VERSA  ANSTEY, F.  P
VICE’S PLEASURE TOY, THE  WHIP, LEATHER  WRD
VICKERS  FARREN, MICK  C/P:2,P
VICTIM PRIME  SHECKLEY, ROBERT  C/P,P
VICTORIA WINTERS  ROSS, MARILYN  P
VICTORIAN ALBUM, THE  BERCKMAN, EVELYN  H

The MITSFS Titledex  September 24, 2007  Page 396
TITLE  AUTHOR  CODE
VICTORIAN CHAISE LONGUE, THE  LASKI, MARGHANITA  C/P,P
VICTORIAN GHOST STORIES  DALBY, RICHARD  LPA

VICTORIOUS OPPOSITION, THE [CONFEDERATE STATES AMERICAN EMPIRE 3] TURTLEDOVE, HARRY

H

VICTORY [@FEY 5]  RUSCH, KRISTINE KATHRYN  P
VICTORY FOR KREGEN, A [DRAY PRESCOT 22]  AKERS, ALAN BURT  C/P,P
VICTORY ON JANUS [JANUS 2]  NORTON, ANDRE  C/P,P,H
VIDEODROME  MARTIN, JACK  P
VIDESSOS BESIEGED [VIDESSOS 5]  TURTLEDOVE, HARRY  P
VIE TRES PRIVEE, UNE  FRAYN, MICHAEL  FP
VIETNAM AND OTHER ALIEN WORLDS  HALDEMAN, JOE  SR-H
VIEW FROM A DISTANT STAR, THE  SHAPLEY, HARLOW  [C/RAN-P:2]
VIEW FROM A HEIGHT  ASIMOV, ISAAC  [C/RAN-P:2]
VIEW FROM ANOTHER SHORE  ROTTENSTEINER, FRANZ  C/PA,PA,SFWA
VIEW FROM SERENDIP, THE  CLARKE, ARTHUR C.  SFWA
VIEW FROM THE EDGE, THE  TURNER, GEORGE  SR-LPA
VIEW FROM THE STARS, THE  MILLER, WALTER M. JR.  C/P:2,P
VIEWPOINT  BOVA, BEN  R/REF-H
VIEWS FROM THE OLDEST HOUSE  GRANT, RICHARD  C/P,LP
VIGIL  MASELLO, ROBERT  P
VIGILANT [LEAGUE OF PEOPLES 2]  GARDNER, JAMES ALAN  C/P:2,P
VIGILANTE, 21ST CENTURY  WILLIAMS, ROBERT MOORE  C/P:2,P
VILE VILLAGE, THE [@LAST RANGER #7]  SARGENT, CRAIG  P
VILLAGE IN THE TREETOPS, THE  VERNE, JULES  H
VILLAGE OF STARS  STANTON, PAUL  C/P,P,H
VILLAGE OF THE DAMNED  WYNDHAM, JOHN  C/P:2,P
VILLAGE OF VAMPIRES  OTFINOSKI, STEVEN  SR-P
VILLAINS [@RIFTWORLD 2]  MCCAY, BILL  C/P,P
VILLAINS BY NECESSITY  FORWARD, EVE  H,P
VILLAINS OF VOLTURNUS [ENDLESS QUEST 8]  BLASHFIELD, JEAN  P
VILLAINS VICTORIOUS  GREENBERG, MARTIN H./HELFERS, JOHN  C/PA,PA
VILLAINY VICTORIOUS [@MISSION EARTH #9]  HUBBARD, L. RON  H
VINAS SOLAMNUS [@DRAGONLANCE LOST LEGENDS #1]  KING, J. ROBERT  P
VINCALIS THE AGITATOR  LISLE, HOLLY  C/P,LP
VINGT MILLE LIEUES SOUS LES MERS  VERNE, JULES  FP
VINTAGE ANTHOLOGY OF SCIENCE FANTASY, THE  CERF, CHRISTOPHER  PA,C/PA
VINTAGE BRADBURY, THE  BRADBURY, RAY  C/P:2,P
VINTAGE SCIENCE FICTION  HAINING, PETER  LPA
VINTAGE SEASON [VINTAGE SEASON 1]  MOORE, C. L.  SFWA-TD18,TD18
VIOLATIONS [STAR TREK VOYAGER]  WRIGHT, SUSAN  C/P:2,P
VIOLENT MAN, THE  VAN VOGT, A. E.  P
VIOLENT STARS [FLESH AND GOLD 2]  GOTLIEB, PHYLLIS  H
VIOLET APPLE, THE/WITCH, THE  LINDSAY, DAVID  LP
VIPERHAND [@FORGOTTEN REALMS MAZTICA #2]  NILES, DOUGLAS  C/P:2,P
VIPER’S KISS [@FORGOTTEN REALMS HOUSE OF SERPENTS #2]  SMEDMAN, LISA  P
VIPER’S SPAWN [@TIMECOP 1]  PARKINSON, DAN  P
VIRAVAX  RANSOM, BILL  P,SFWA
VIRGIL FINLAY  GRANT, DONALD M.  A
VIRGIN, THE [LIVING PLANET 3]  COON, SUSAN  C/P,P
VIRGIN  PATTERSON, JAMES  H,P
VIRGIN AND THE DINOSAUR, THE  GARCIA Y ROBERTSON, R.  P
VIRGIN AND THE VAMPIRE, THE  MYERS, ROBERT J.  P
VIRGIN AND THE WHEELS, THE  DE CAMP, L. SPRAGUE  C/P:2,P
VIRGIN IN FLAMES  ROHMER, SAX  H
VIRGIN PLANET  ANDERSON, POUL  SR-P,C/P:2
VIRGINS AND MARTYRS  MAGINN, SIMON  P
VIRGINTOOTH  IVANHOE, MARK  P
VIRICONIUM [VIRICONIUM]  HARRISON, M. JOHN  LP
VIRICONIUM NIGHTS [VIRICONIUM]  HARRISON, M. JOHN  P
VIRILITY FACTOR, THE  MERLE, ROBERT  H
VIRILITY GENE, THE [@HOOK #4]  ZETFORD, TULLY  C/P,P
VIRTU, THE  MONETTE, SARAH  H
VIRTUAL DEATH  AARON, SHALE  C/P:2,P
VIRTUAL DESTRUCTION [CRAIG KREIDENT 1]  ANDERSON, KEVIN J./BEASON, DOUG  C/P,P
VIRTUAL GIRL  THOMSON, AMY  C/P,P
VIRTUAL LIGHT  GIBSON, WILLIAM  H:2,P
VIRTUAL LIGHT (EXCERPT)  GIBSON, WILLIAM  SR-LP,LP
VIRTUAL MODE [MODE 1]  ANTHONY, PIERS  C/P:2,SFWA-P:2,P
VIRTUAL SOUL, A  TEIXEIRA, KEVIN  P
VIRTUAL UNREALITIES  BESTER, ALFRED  H,LP
VIRTUAL ZEN  NELSON, RAY  P
VIRTUOSITY  BISSON, TERRY  P
VIRUS  BUCHANAN, BILL  P
VIRUS  PERRY, S. D.  P

TITLE  AUTHOR  CODE
VIRUS CLANS  KANALY, MICHAEL  P
VIRUS DESTRUIDOR, O  MAINE, CHARLES ERIC  FP
VISCOUS CIRCLE [CLUSTER #5]  ANTHONY, PIERS  C/P:2,@P
VISIBLE LIGHT  CHERRYH, C. J.  C/P,P,H
VISIBLE MAN, THE  DOZOIS, GARDNER  P
VISION, THE  KOONTZ, DEAN R.  H,P
VISION OF STEPHEN, THE  BURFORD, LOLAH  C/P:2,P
VISION OF TAROT [TAROT #2]  ANTHONY, PIERS  C/P:2,@P
VISION OF THE FUTURE [STAR WARS]  ZAHN, TIMOTHY  C/P,P
VISION QUEST  SERVICE, PAMELA F.  C/P,P
VISIONARY IN RESIDENCE  STERLING, BRUCE  LP
VISIONS 2002 CALENDAR  SCIENCE FICTION BOOK CLUB  CAL
VISIONS AND VENTURERS  STURGEON, THEODORE  C/P:2,P
VISIONS FROM NOWHERE [@RETURN TO THE PLANET OF THE APES #1]  ARROW, WILLIAM  P
VISIONS IN DEATH [EVE DALLAS]  ROBB, J. D.  H
VISIONS OF FANTASY  ASIMOV, ISAAC/GREENBERG, MARTIN H.  HA
VISIONS OF FEAR  HARTWELL, DAVID G.  PA
VISIONS OF LIBERTY  TIER, MARK/GREENBERG, MARTIN H.  PA
VISIONS OF TERROR  KATZ, WILLIAM  P
VISIONS OF THE DAMNED  MARTEN, JACQUELINE  P
VISIONS OF TOMORROW  ELWOOD, ROGER  PA,C/PA
VISIONS OF TOMORROW: GREAT SCIENCE FICTION FROM THE MOVIES  EDELSON, EDWARD  R/REF-H
VISIONS OF WONDER  BOYER, ROBERT H./ZAHORSKI, KENNETH J. PA
VISIONS OF WONDER  HARTWELL, DAVID G./WOLF, MILTON T.  LPA
VISIT TO A SMALL PLANET  VIDAL, GORE  C/P:2,P
VISITING MOON, THE  HAYES, RALPH  H
VISITOR, THE  CUNNINGHAM, JERE  P
VISITOR, THE  TEPPER, SHERI S.  H
VISITOR FROM 50,000 YEARS AGO, THE (CHINESE)  PEOPLE’S LITERATURE  FP
VISITOR FROM OUTER SPACE, A  FOREIGN LANGUAGES PUB.  SR-PA
VISITORS [BUFFY]  GILMAN, LAURA ANNE/SHERMAN, JOSEPHA P
VISITORS, THE  SIMAK, CLIFFORD D.  H:2,P
VISITORS FROM OZ, THE  BAUM, L. FRANK  H
VISITORS FROM PLANET VETA, THE  GOLL, REINHOLD W.  H
VITALS  BEAR, GREG  H
VITTORIO THE VAMPIRE [NEW TALES 2]  RICE, ANNE  H:2,P
VIXEN 03  CUSSLER, CLIVE  P
VIZIER OF THE TWO-HORNED ALEXANDER, THE  STOCKTON, FRANK R.  H
VIZIER’S SECOND DAUGHTER, THE  YOUNG, ROBERT F.  P
VMR THEORY, THE  FREZZA, ROBERT  P
VOICE COMMANDS, THE  CLIVE, DENNIS  B
VOICE FOR PRINCESS, A [KEDRIGERN 1]  MORRESSY, JOHN  C/P,P
VOICE OF CEPHEUS, THE  APPLEBY, KEN  C/P:2,P
VOICE OF DASHIN, THE  GANPAT  H
VOICE OF OUR SHADOW  CARROLL, JONATHAN  P
VOICE OF THE DEMON [@ELITA #2]  JACOBY, KATE  P
VOICE OF THE DOLPHINS AND OTHER STORIES, THE  SZILARD, LEO  LP
VOICE OF THE DOLPHINS AND OTHER STORIES, THE (BRITISH)  SZILARD, LEO  SR-P
VOICE OF THE GODS [AGE OF THE FIVE 3]  CANAVAN, TRUDI  P
VOICE OF THE MOUNTAIN, THE [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  H,SFWA
VOICE OF THE PLANET  TOBIAS, MICHAEL  C/P:2,P
VOICE OF THE VISITOR  SLONAKER, LARRY  C/P:2,P
VOICE OF THE WHIRLWIND  WILLIAMS, WALTER JON  C/P:2,SFWA-P,P,H
VOICE OUT OF RAMAH, A  KILLOUGH, LEE  C/P:2,P
VOICES [WESTERN SHORE 2]  LE GUIN, URSULA K.  H
VOICES [@BABYLON 5 #1]  VORNHOLT, JOHN  C/P:2,P
VOICES FOR THE FUTURE 1  CLARESON, THOMAS D.  R/REF-H
VOICES FOR THE FUTURE 3  CLARESON, THOMAS D./WYMER, THOMAS  R/REF-P
VOICES FROM THE SKY  CLARKE, ARTHUR C.  [C/RAN-P:2]
VOICES IN THE DARK [TWILIGHT 2]  HAYNES, JAMES  P
VOICES IN THE FOG [HOLDERLY HALL 3]  CAMERON, KATE  P
VOICES OF CHAOS [STARBRIDGE 7]  CRISPIN, A. C./EMERSON, RU  P
VOICES OF DOOM  COPPER, BASIL  H
VOICES OF HEAVEN, THE  POHL, FREDERIK  H:2,P
VOICES OF HOPE [NICHOLAS SEAFORT 5]  FEINTUCH, DAVID  C/P:2,P
VOICES OF TIME, THE  BALLARD, J. G.  C/P:2,P
VOID BEYOND AND OTHER STORIES, THE  WILLIAMS, ROBERT MOORE  DD322
VOID CAPTAIN’S TALE, THE  SPINRAD, NORMAN  H:2,P
VOIDFARER [MOONWORLDS]  MCMULLEN, SEAN  H
VOIMA  BRITTAIN, C. DALE  C/P,P
VOLCANO OGRE, THE [ZARKON 3]  CARTER, LIN  C/P,SFWA,H
VOLKHAVAAR  LEE, TANITH  C/P:2,P
VOLSUNGA SAGA: THE STORY OF THE VOLSUNGS AND NIBELUNGS  ANONYMOUS  P
VOLTEFACE [INTERFACE 2]  ADLARD, MARK  P

TITLE  AUTHOR  CODE
VOLUNTEER FOR FRAGO [PERRY RHODAN 122]  BRAND, KURT  SR-LP
VON BEK (EXPANDED OMNIBUS EDITION) [VON BEK]  MOORCOCK, MICHAEL  LP
VON BEK (OMNIBUS EDITION) [VON BEK]  MOORCOCK, MICHAEL  P
VON NEUMANN’S WAR  RINGO, JOHN/TAYLOR, TRAVIS S.  H
VONNEGUT STATEMENT, THE  KLINKOWITZ, JEROME/SOMER, JOHN  R/REF-P
VOODOO CHILD  REAVES, MICHAEL  H
VOODOO MOON [@CHARMED]  STAUB, WENDY CORSI  P
VOODOO PLANET  NORTH, ANDREW  DD345
VOORLOPER  NORTON, ANDRE  H:2,P
VOR  BLISH, JAMES  C/P:2,P
VOR GAME, THE [MILES VORKOSIGAN 4]  BUJOLD, LOIS MCMASTER  C/P:2,P
VORKOSIGAN’S GAME [MILES VORKOSIGAN 3,4]  BUJOLD, LOIS MCMASTER  H
VORTEX [STEN 7]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
VORTEX  LAND, JON  P
VORTEX BLASTER, THE  SMITH, E. E.  H:2
VORTEX BLASTERS, THE  MOSKOWITZ, SAM  PA,C/PA
VOSOFF AND NIMMITZ  CASTRO, ADAM-TROY  H
VOYAGE  BAXTER, STEPHEN  C/P,LP,H
VOYAGE, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P,P,H
VOYAGE AU CENTRE DE LA TERRE  VERNE, JULES  FP
VOYAGE BEGUN, THE  BOND, NANCY  H
VOYAGE DANS LA LUNE  DE BERGERAC, CYRANO  FP
VOYAGE FROM YESTERYEAR  HOGAN, JAMES P.  C/P,P,H
VOYAGE INTO SPACE  VAN LODEN, ERLE  SR-P
VOYAGE OF ARGO, THE  RHODES, APOLLONIUS OF  P
VOYAGE OF THE "DAWN TREADER", THE [NARNIA 3]  LEWIS, C. S.  C/P:2,P
VOYAGE OF THE FOX RIDER [MITHGAR 2]  MCKIERNAN, DENNIS L.  C/P,P
VOYAGE OF THE LUNA 1, THE  CRAIGIE, DAVID  H:2
VOYAGE OF THE MOURNING DAWN [@EBERRON HEIRS OF ASH #1]  WULF, RICH  P
VOYAGE OF THE PLANETSLAYER [@CONCORDAT #2]  SWYCAFFER, JEFFERSON P.  C/P:2,P
VOYAGE OF THE SABLE KEECH, THE [POLITY SKINNER 2]  ASHER, NEAL  H
VOYAGE OF THE SHADOWMOON [MOONWORLDS 1]  MCMULLEN, SEAN  H
VOYAGE OF THE SPACE BEAGLE, THE  VAN VOGT, A. E.  C/P:2,SFWA,P
VOYAGE OF THE STAR WOLF [STAR WOLF 1]  GERROLD, DAVID  C/P,P,H
VOYAGE OF THE TRIGON  DE BOLT, ADRIANA  P
VOYAGE OF VENGEANCE [@MISSION EARTH #7]  HUBBARD, L. RON  C/P,P,H
VOYAGE TO A FORGOTTEN SUN  PFEIL, DONALD J.  C/P:2,P
VOYAGE TO ARCTURUS, A  LINDSAY, DAVID  C/P,P,H
VOYAGE TO DARI, A  WALLACE, IAN  C/P:2,P
VOYAGE TO DESTINY  COOK, LEONARD R.  H
VOYAGE TO ENEH [SEAS OF KILMOYN 1]  GREEN, ROLAND J.  H,P
VOYAGE TO FAREMIDO/CAPILLARIA  KARINTHY, FRIGYES  SR-P
VOYAGE TO THE BOTTOM OF THE SEA  JONES, RAYMOND F.  H:2
VOYAGE TO THE BOTTOM OF THE SEA  STURGEON, THEODORE  C/P:2,P
VOYAGE TO THE CITY OF THE DEAD [HUMANX]  FOSTER, ALAN DEAN  C/P:2,P
VOYAGE TO THE RED PLANET  BISSON, TERRY  C/P:2,SFWA,P,H
VOYAGER [OUTLANDER (GABALDON) 3]  GABALDON, DIANA  C/P,P,H
VOYAGER  HARPERCOLLINS  B
VOYAGER IN NIGHT  CHERRYH, C. J.  C/P,P,H
VOYAGERS [@VOYAGERS #1]  BOVA, BEN  C/P:2,SFWA-P,P,H
VOYAGERS IN TIME  SILVERBERG, ROBERT  PA,C/PA
VOYAGES  SAUER, ROB  C/PA,PA
VOYAGES BY STARLIGHT  MACLEOD, IAN R.  H
VOYAGES IN SPACE  LOCKE, GEORGE  R/IND
VOYAGES OF DOCTOR DOLITTLE, THE [DOCTOR DOLITTLE]  LOFTING, HUGH  C/P,P,H
VOYAGES OF IMAGINATION [STAR TREK]  AYERS, JEFF  R/REF-H
VULCAN! [STAR TREK]  SKY, KATHLEEN  C/P,P
VULCAN ACADEMY MURDERS, THE [STAR TREK]  LORRAH, JEAN  C/P:2,P
VULCAN’S FORGE [@STAR TREK]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
VULCAN’S GLORY [STAR TREK]  FONTANA, D. C.  C/P:2,P
VULCAN’S HAMMER  DICK, PHILIP K.  DD457
VULCAN’S HEART [STAR TREK]  SHERMAN, JOSEPHA/SHWARTZ, SUSAN  P
VULGMASTER, THE [@TALES FROM THE ONE-EYED CROW]  MCKIERNAN, DENNIS L.  C/P:2,SCX,SFWA-P
VULNERABLE  LANE, AMY  LP
VULTURES OF THE WHITE DEATH [@G-8 #5]  HOGAN, ROBERT J.  SR-P
VULTURES OF WHAPETON, THE  HOWARD, ROBERT E.  P
VURT  NOON, JEFF  LP

TITLE  AUTHOR  CODE
W3: WOMEN IN DEEP TIME  BEAR, GREG  P
WAGERED WORLD, THE  JANIFER, LAURENCE M./TREIBICH, S. J.  D81680
WAGERS OF SIN [TIME SCOUT]  ASPRIN, ROBERT LYNN/EVANS, LINDA  P
WAGES OF SIN, THE  MACLAINE, JENNA  LP
WAGES OF SIN, THE [DOCTOR WHO PAST DOCTORS]  MCINTEE, DAVID A.  C/P:2,P
WAGNER, THE WEHR-WOLF  REYNOLDS, G. W. M.  VLP
WAGON TRAIN TO THE STARS [@STAR TREK NEW EARTH #1]  CAREY, DIANE  P
WAILING ASTEROID, THE  LEINSTER, MURRAY  C/P:2,P
WAILING OCTOPUS, THE [@RICK BRANT #11]  BLAINE, JOHN  H
WAITERS ON THE DANCE [LEMMUS 1]  SAVARIN, JULIAN JAY  H
WAITING FOR GODALMING [TRILOGY THAT DARE NOT SPEAK ITS NAME 3] RANKIN, ROBERT  P
WAITING FOR THE GALACTIC BUS [BARION AND COYUL 1]  GODWIN, PARKE  H:2,P
WAITING ROOM, THE  WRIGHT, T. M.  SFWA-P
WAKE, THE [@SANDMAN #10]  GAIMAN, NEIL  CX
WAKE UP SCREAMING  WRIGHT, LEE/SHEEHAN, RICHARD G.  PA
WAKING BEAUTY  WITCOVER, PAUL  H
WAKING IN DREAMLAND  NYE, JODY LYNN  P
WAKING NIGHTMARES  CAMPBELL, RAMSEY  H
WAKING THE MOON  HAND, ELIZABETH  H,P
WALDEN TWO  SKINNER, B. F.  C/P,P,H
WALDO/MAGIC, INC.  HEINLEIN, ROBERT A.  C/P,P,H
WALDO: GENIUS IN ORBIT  HEINLEIN, ROBERT A.  P
WALK IN HELL [CONFEDERATE STATES: GREAT WAR 2]  TURTLEDOVE, HARRY  C/P,P,H
WALK IN WOLF WOOD, A  STEWART, MARY  P
WALK OUT OF THE WORLD, A  NICHOLS, RUTH  P
WALK THE MOONS ROAD  AIKIN, JIM  P
WALK THROUGH TO-MORROW  ZEIGFREID, KARL  SR-P,C/P
WALK TO THE END OF THE WORLD [HOLDFAST 1]  CHARNAS, SUZY MCKEE  C/P:2,P
WALK UP THE SKY  WILLIAMS, ROBERT MOORE  H
WALKABOUT WOMAN  ROESSNER, MICHAELA  C/P,P
WALKAWAY CLAUSE, THE  DALMAS, JOHN  C/P,SFWA-P,P
WALKER BETWEEN THE WORLDS  DESROCHERS, DIANE  C/P,P
WALKER IN SHADOWS, THE  MICHAELS, BARBARA  P
WALKER OF WORLDS [@KING’S TRAMP #1]  DE HAVEN, TOM  C/P:2,LP
WALKER-THROUGH-WALLS, THE  AYME, MARCEL  P
WALKERS  MASTERTON, GRAHAM  P
WALKERS ON THE SKY  LAKE, DAVID J.  P
WALKING IN ETERNITY [DOCTOR WHO]  EALES, JULIAN  SR-LPA
WALKING ON GLASS  BANKS, IAIN M.  P
WALKING SHADOW, THE  STABLEFORD, BRIAN M.  P
WALKING THE LABYRINTH  GOLDSTEIN, LISA  LP

WALKING TO BABYLON [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)] ORMAN, KATE

P,PA

WALL, THE [@QUANTUM LEAP]  MCCONNELL, ASHLEY  C/P,SFWA-P:2,P
WALL, THE  TUBB, E. C.  LP
WALL AROUND A STAR  POHL, FREDERIK/WILLIAMSON, JACK  C/P,P
WALL AROUND EDEN, THE  SLONCZEWSKI, JOAN  C/P,P,H
WALL AROUND THE WORLD, THE  COGSWELL, THEODORE R.  C/P:2,P
WALL AT THE EDGE OF THE WORLD, THE  AIKIN, JIM  P,SFWA-P:2
WALL OF SERPENTS [INCOMPLETE ENCHANTER]  DE CAMP, L. SPRAGUE/PRATT, FLETCHER  SR-H,P
WALL OF THE SKY, THE WALL OF THE EYE, THE  LETHEM, JONATHAN  LP
WALL OF YEARS, THE  STEPHENSON, ANDREW M.  C/P,P
WALLET OF KAI LUNG, THE [KAI LUNG]  BRAMAH, ERNEST  H
WALLS OF AIR, THE [DARWATH 2]  HAMBLY, BARBARA  C/P:2,P
WALLS OF WESTERNFORT, THE [CELAENO]  FLETCHER, JANE  LP
WALLS WITHIN WALLS  TOFTE, ARTHUR  SR-P,C/P
WALPURGIS III  RESNICK, MIKE  P
WAMPETERS, FOMA, AND GRANFALLOONS (OPINIONS)  VONNEGUT, KURT JR.  R/REF-P
WANDERER, THE  LEIBER, FRITZ  C/P:2,SFWA-P,P
WANDERER [WARRIOR 2]  MCQUINN, DONALD E.  P
WANDERER  SCHMIDT, DENNIS  C/P,P
WANDERER OF SPACE  STATTEN, VARGO  SR-P
WANDERER’S NECKLACE, THE  HAGGARD, H. RIDER  P
WANDERERS OF TIME  WYNDHAM, JOHN  SR-P
WANDERER’S RETURN, A [WINDAMEIR CIRCLE 3]  HANCOCK, NIEL  P
WANDERER’S TALE, THE [LINDORMYN 1]  BILSBOROUGH, DAVID  H
WANDERGROUND, THE  GEARHARDT, SALLY MILLER  H,P
WANDERING FIRE, THE [FIONAVAR TAPESTRY 2]  KAY, GUY GAVRIEL  H:2,SFWA,SFWA-P,P
WANDERING STARS  DANN, JACK  HA,PA,SFWA
WANDERING TELLURIAN, THE  SCHWARTZ, ALAN  DH020
WANDERING UNICORN, THE  MUJICA LAINEZ, MANUEL  H,P
WANDERING VARIABLES, THE  TRIMBLE, LOUIS  C/P:2,P
WANDERINGS OF WUNTVOR, THE [WUNTVOR 1,2,3]  GARDNER, CRAIG SHAW  H:2
WANDERLUST [@DRAGONLANCE MEETINGS 2]  KIRCHOFF, MARY/WINTER, STEVE  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 400
TITLE  AUTHOR  CODE
WANDL THE INVADER  CUMMINGS, RAY  DD497
WANDOR’S FLIGHT [WANDOR 4]  GREEN, ROLAND J.  P
WANDOR’S JOURNEY [WANDOR 2]  GREEN, ROLAND J.  P
WANDOR’S RIDE [WANDOR 1]  GREEN, ROLAND J.  P
WANDOR’S VOYAGE [WANDOR 3]  GREEN, ROLAND J.  C/P,P
WANTED: A FAIRY GODFATHER [@BARNABY #1]  JOHNSON, CROCKETT  SCX
WANTED BY THE INTERGALACTIC SECURITY BUREAU  SEIDMAN, ERIC  A
WANTED: DEAD OR ALIVE [@SPIDERMAN]  GARDNER, CRAIG SHAW  H
WANTING FACTOR, THE  DEWEESE, GENE  P
WANTING SEED, THE  BURGESS, ANTHONY  C/P:2,P
WAR, THE [@BLACKWATER #6]  MCDOWELL, MICHAEL  P
WAR [@ALIENS VS. PREDATOR 3]  PERRY, S. D.  C/P,P

WAR AGAINST THE CHTORR: INVASION, THE [WAR AGAINST THE CHTORR (ABRIDGED) 1,2] GERROLD, DAVID

H:2

WAR AGAINST THE RULL, THE  VAN VOGT, A. E.  H:2,P
WAR AMONGST THE ANGELS, THE [SECOND ETHER 3]  MOORCOCK, MICHAEL  H,LP
WAR BETWEEN THE PITIFUL TEACHERS AND THE SPLENDID KIDS, THE  KIESEL, STANLEY  C/P,P
WAR DOGS OF THE GOLDEN HORDE [@MARS ATTACKS]  MURILL, RAY W.  C/P,P
WAR DRAGONS [STAR TREK CAPTAIN’S TABLE 1]  GRAF, L. A.  C/P:2,P
WAR DRUMS [STAR TREK TNG]  VORNHOLT, JOHN  C/P:2,P
WAR FOR ETERNITY, THE  ROWLEY, CHRISTOPHER  C/P:2,P
WAR FOR THE OAKS  BULL, EMMA  C/P:2,P
WAR-GAMERS’ WORLD [MAGIRA 1]  WALKER, HUGH  P
WAR GAMES  HANSEN, KARL  C/P:2,P
WAR GAMES OF ZELOS, THE [EXPENDABLES 3]  AVERY, RICHARD  C/P:2,P
WAR GOD’S OWN, THE [HRADANI 1]  WEBER, DAVID  C/P,P,H
WAR HOUND AND THE WORLD’S PAIN, THE [VON BEK 1]  MOORCOCK, MICHAEL  C/P,P,H
WAR IN 2020, THE  PETERS, RALPH  H:2,P
WAR IN HEAVEN, THE [ETERNAL WARRIORS 1]  BEALE, THEODORE  P
WAR IN HEAVEN [@HORIZON WAR #3/MAGE: THE ASCENSION]  WEINBERG, ROBERT  P
WAR IN HEAVEN  WILLIAMS, CHARLES  H,LP
WAR IN HEAVEN [REQUIEM FOR HOMO SAPIENS 4]  ZINDELL, DAVID  P
WAR IN HELL [HEROES IN HELL]  MORRIS, JANET  C/PA,PA,SFWA-P
WAR IN OUTER SPACE, THE [@NEW TOM SWIFT #4]  APPLETON, VICTOR  H,LP
WAR IN TETHYR [FORGOTTEN REALMS NOBLES 2]  MILAN, VICTOR  P
WAR IN THE AIR, THE  WELLS, H. G.  SR-P
WAR IN THE DESERT [@ID4]  MOLSTAD, STEPHEN  C/P,P
WAR IN THE MODERN WORLD  ROPP, THEODORE  [C/RAN-P]
WAR LORDS OF SPACE  HUGHES, DENIS  SR-P
WAR MACHINE, THE [CRISIS OF EMPIRE 3]  ALLEN, ROGER MACBRIDE/DRAKE, DAVID  P,SFWA-P:2
WAR MACHINES OF KALINTH, THE [DIRSHAN 2]  LANCOUR, GENE  H:2,SFWA
WAR MINSTRELS, THE  HABER, KAREN  P
WAR OF 1938, THE  WRIGHT, S. FOWLER  SR-H
WAR OF ARGOS  LE PAGE, RAND  SR-P
WAR OF DREAMS, THE  CARTER, ANGELA  H,P
WAR OF HONOR (HARDCOVER AND CD) [HONOR HARRINGTON 10]  WEBER, DAVID  H:2
WAR OF NERVES [@ATTAR #2]  GRAHAM, ROBERT  P
WAR OF OMISSION  O’DONNELL, KEVIN JR.  C/P:2,P
WAR OF SHADOWS, A  CHALKER, JACK L.  C/P:2,P
WAR OF THE CYBERNAUTS, THE [FLASH GORDON 6]  BINGHAM, CARSON  C/P:2,P
WAR OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  PEEL, JOHN  C/P:2,P
WAR OF THE FLOWERS, THE  WILLIAMS, TAD  H
WAR OF THE GODS  ANDERSON, POUL  H:2,P
WAR OF THE GODS [@BATTLESTAR GALACTICA 1978 #7]  LARSON, GLEN A./YERMAKOV, NICHOLAS  P
WAR OF THE GURUS  RHEINGOLD, HOWARD  SR-P
WAR OF THE JEWELS, THE [HISTORY OF MIDDLE-EARTH 11]  TOLKIEN, J. R. R.  H:2
WAR OF THE LANCE, THE [@DRAGONLANCE TALES II #3]  TSR  C/PA,PA
WAR OF THE MAELSTROM [CHANGEWINDS 3]  CHALKER, JACK L.  C/P:3,SFWA-P,P

WAR OF THE PROPHETS, THE [@STAR TREK DS9 MILLENIUM #2] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

C/P,P

WAR OF THE RING, THE [HISTORY OF MIDDLE-EARTH 8]  TOLKIEN, J. R. R.  H:2
WAR OF THE TWINS [@DRAGONLANCE LEGENDS #2]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P
WAR OF THE WING-MEN  ANDERSON, POUL  C/P:2,P,DD303
WAR OF THE WIZARDS [@LONE WOLF #4]  PAGE, IAN/DEVER, JOE  P
WAR OF THE WORLDS, THE  WELLS, H. G.  C/P,SFWA-P,P,H,LP
WAR OF THE WORLDS, THE/TIME MACHINE, THE  WELLS, H. G.  C/P:2,P
WAR OF THE WORLDS: GLOBAL DISPATCHES  ANDERSON, KEVIN J.  PA
WAR OF THE WORLDS, THE (CASSETTE)  WELLS, H. G.  SR-MM

WAR OF THE WORLDS, THE TIME MACHINE, AND SELECTED SHORT STORIES, THE WELLS, H. G.

H:2

WAR OF THREE WATERS [WATERSHED 3]  NILES, DOUGLAS  H,LP
WAR OF TWO WORLDS, THE  ANDERSON, POUL  C/P,DD335
WAR ON ALEPH [@RACK #2]  JAMES, LAURENCE  P
WAR SURF  BUCKNER, M. M.  P

The MITSFS Titledex  September 24, 2007  Page 401
TITLE  AUTHOR  CODE
WAR WITH EARTH, THE  FRANKOWSKI, LEO/GROSSMAN, DAVE  P
WAR WITH THE GIZMOS  LEINSTER, MURRAY  C/P:2,P
WAR WITH THE NEWTS  CAPEK, KAREL  C/P:2,P,H

WAR WITH THE ROBOTS ASIMOV, ISAAC/WARRICK, PATRICIA S./

GREENBERG, MARTIN H.

HA

WAR WITH THE ROBOTS  HARRISON, HARRY  C/P:2,P
WAR ZONE [GUARDIANS 6]  AUSTIN, RICHARD  P
WARAUTUMN [ERON 4]  DEITZ, TOM  P
WARBLADE [WARHAMMER KONRAD 3]  FERRING, DAVID  P
WARBOTS [@WARBOTS #1]  STINE, G. HARRY  C/P,P
WARCHILD [WARCHILD 1]  BOWES, RICHARD  C/P,P
WARCHILD [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
WARCHILD [STAR TREK DS9]  FRIESNER, ESTHER M.  C/P:2,P
WARCHILD  LOWACHEE, KARIN  C/P,P
WARCHILD (PROOF) [STAR TREK DS9]  FRIESNER, ESTHER M.  SR-VLP

WARCRAFT ARCHIVE, THE [WARCRAFT] KNAAK, RICHARD A./GOLDEN, CHRISTIE/

GRUBB, JEFF/METZEN, CHRIS

H

WARDAY  STRIEBER, WHITLEY/KUNETKA, JAMES  C/P:2,P
WARDEN OF HORSES, THE [@SLOW WORLD #2]  RIPLEY, KAREN  P
WARDING OF WITCH WORLD, THE [WITCH WORLD SECRETS]  NORTON, ANDRE  H:2,P
WARDOVE, THE  SMITH, L. NEIL  C/P,P
’WARE HAWK [WITCH WORLD ESTCARP]  NORTON, ANDRE  H:2,P
WARGAMES  BISCHOFF, DAVID F.  C/P:2,P
WARHAVEN  HARVEY, M. ELAYN  H
WARHAWK [@BATTLESTAR GALACTICA 1978]  HATCH, RICHARD/GOLDEN, CHRISTOPHER  H
WARHORN  KRAMER, DANA  SFWA-P
WARHORSE  ZAHN, TIMOTHY  C/P:2,SFWA-P,P

WARHOST OF VASTMARK [WARS OF LIGHT AND SHADOW 2/ SHIPS OF MERIOR 2] WURTS, JANNY

C/P,P

WARLOCK [DOCTOR WHO NEW ADVENTURES]  CARTMEL, ANDREW  C/P:2,P
WARLOCK [@DARKWAR #2]  COOK, GLEN  C/P,P
WARLOCK  GARTON, RAY  P
WARLOCK  KOONTZ, DEAN R.  C/P,P
WARLOCK AND SON [WARLOCK 11]  STASHEFF, CHRISTOPHER  C/P,SFWA-P:2
WARLOCK AT THE WHEEL  JONES, DIANA WYNNE  H
WARLOCK ENLARGED, THE [WARLOCK 2,3,4]  STASHEFF, CHRISTOPHER  H:2
WARLOCK ENRAGED, THE [WARLOCK 4]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK HERETICAL, THE [WARLOCK 7]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK IN SPITE OF HIMSELF, THE [WARLOCK 1]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK INSANE, THE [WARLOCK 9]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK IS MISSING, THE [WARLOCK 6]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK OF RHADA, THE [RHADA 2]  GILMAN, ROBERT CHAM  C/P,P
WARLOCK OF THE WITCH WORLD [WITCH WORLD ESTCARP 2]  NORTON, ANDRE  C/P:2,P
WARLOCK ROCK, THE [WARLOCK 10]  STASHEFF, CHRISTOPHER  C/P,SFWA-P:2,P
WARLOCK UNLOCKED, THE [WARLOCK 3]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK WANDERING, THE [WARLOCK 5]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCKS AND WARRIORS  DE CAMP, L. SPRAGUE  C/PA,PA
WARLOCK’S BLADE, A  PERRY, MARK C./PERRY, MAGAHN  P,SFWA-P
WARLOCK’S COMPANION, THE [WARLOCK 8]  STASHEFF, CHRISTOPHER  C/P:2,P
WARLOCK’S GIFT  MAYHAR, ARDATH  H,SFWA
WARLOCK’S LAST RIDE, THE [WARLOCK 12]  STASHEFF, CHRISTOPHER  P
WARLOCK’S NIGHT OUT, THE [WARLOCK 5,6]  STASHEFF, CHRISTOPHER  H:2
WARLORD, THE [@WARLORD #1]  FROST, JASON  P
WARLORD!  MORRIS, JANET  C/P,P
WARLORD, THE [@CASCA #3]  SADLER, BARRY  C/P,P
WARLORD  STIRLING, S. M./DRAKE, DAVID  H
WARLORD OF ANTARES [DRAY PRESCOT 37]  AKERS, ALAN BURT  P
WARLORD OF GHANDOR  DOW DELL, DEL  P
WARLORD OF HEAVEN [@STAR REQUIEM #3]  COLE, ADRIAN  P,SFWA-P
WARLORD OF KOR  CARR, TERRY  DF177
WARLORD OF MARS, THE [BARSOOM #3]  BURROUGHS, EDGAR RICE  C/P,@P,H:2
WARLORD OF THE AIR, THE [OSWALD BASTABLE 1]  MOORCOCK, MICHAEL  C/P:2,P
WARLORD’S DOMAIN, THE [BOOK OF YEARS 4]  MORWOOD, PETER  P
WARLORDS OF GAIKON [@BLADE #18]  LORD, JEFFREY  P
WARLORDS OF NIN, THE [@DRAGON KING #2]  LAWHEAD, STEPHEN R.  C/P,P
WARLORDS OF UTOPIA [@FACTION PARADOX 3]  PARKIN, LANCE  LP
WARLORDS OF XUMA  LAKE, DAVID J.  P
WARLORD’S WORLD  ANVIL, CHRISTOPHER  P
WARM WORLDS AND OTHERWISE  TIPTREE, JAMES JR.  C/P:2,P

WARMASTERS, THE [HONOR HARRINGTON/HAMMER’S SLAMMERS/ BELISARIUS] FAWCETT, BILL

HA

WARMONGER [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  P
WARP ANGEL  HOPEN, STUART  H,P
WARP SPEED  TAYLOR, TRAVIS S.  H

The MITSFS Titledex  September 24, 2007  Page 402
TITLE  AUTHOR  CODE
WARPATH  DANIEL, TONY  C/P,P
WARPATH [@STAR TREK DS9]  MACK, DAVID  P
WARPED [STAR TREK DS9]  JETER, K. W.  H,P
WARPSWORD [@WARHAMMER MALUS DARKBLADE]  ABNETT, DAN/LEE, MIKE  P
WARREN WITCHES, THE [CHARMED]  SIMON SPOTLIGHT ENTERTAINMENT  LPA
WARRING STATES [@FACTION PARADOX 4]  HALLIDAY, MAGS L.  LP
WARRIOR, THE [HAMMER’S SLAMMERS]  DRAKE, DAVID  C/P:2,P
WARRIOR [HYTHRUN WOLFBLADE 2]  FALLON, JENNIFER  H
WARRIOR [WARRIOR 1]  MCQUINN, DONALD E.  C/P,H,LP
WARRIOR! [@THERE WILL BE WAR #5]  POURNELLE, JERRY  C/PA,PA,SFWA-P
WARRIOR, THE [@CASCA #17]  SADLER, BARRY  P
WARRIOR [@ROBOTS IN TIME 3]  WU, WILLIAM F.  P
WARRIOR AND THE WITCH, THE  MILLER, CARL  P
WARRIOR AND WITCH [MIREI 2]  BRENNAN, MARIE  P
WARRIOR BLOOD [@WARHAMMER 40K DEATHWATCH 1]  GOTO, C. S.  P
WARRIOR CHALLENGED, THE [TERRILIAN 4]  GREEN, SHARON  C/P:2,P
WARRIOR: COUPE [@BATTLETECH #43/WARRIOR 3]  STACKPOLE, MICHAEL A.  C/P:2,P
WARRIOR COVEN [@WARHAMMER 40K DEATHWATCH 2]  GOTO, C. S.  P
WARRIOR: EN GARDE [@BATTLETECH #41/WARRIOR 1]  STACKPOLE, MICHAEL A.  C/P,P
WARRIOR ENCHAINED, THE [TERRILIAN 2]  GREEN, SHARON  C/P:2,P

WARRIOR ENCHANTRESSES MASSIE-FRENCH, KATHLEEN M./ GREENBERG, MARTIN H.

C/PA,PA

WARRIOR FANTASTIC  GREENBERG, MARTIN H./HELFERS, JOHN  C/PA,PA
WARRIOR KING, THE [DAMASTES 3]  BUNCH, CHRIS  P
WARRIOR LIVES, THE [GUARDIANS OF THE FLAME 5]  ROSENBERG, JOEL  C/P,P,H
WARRIOR OF LLARN [LLARN 1]  FOX, GARDNER F.  C/P:2,P
WARRIOR OF MARS [CLAYTON DREW]  FEARN, JOHN RUSSELL  SR-P
WARRIOR OF SCORPIO [DRAY PRESCOT 3]  AKERS, ALAN BURT  C/P,P
WARRIOR OF WORLD’S END, THE [GONDWANE 1]  CARTER, LIN  C/P,P
WARRIOR PLANET  WISMER, DON  P

WARRIOR PRINCESSES SCARBOROUGH, ELIZABETH ANN/

GREENBERG, MARTIN H.

C/PA,PA

WARRIOR-PROPHET, THE [@PRINCE OF NOTHING #2]  BAKKER, R. SCOTT  H
WARRIOR REARMED, THE [TERRILIAN 3]  GREEN, SHARON  C/P:2,P
WARRIOR RETURNS, THE [@ANTEROS]  COLE, ALLAN  H,P
WARRIOR: RIPOSTE [@BATTLETECH #42/WARRIOR 2]  STACKPOLE, MICHAEL A.  C/P:2,P
WARRIOR VICTORIOUS, THE [TERRILIAN 5]  GREEN, SHARON  P
WARRIOR WHO CARRIED LIFE, THE  RYMAN, GEOFF  P
WARRIOR WITCH OF HEL [HEL 1]  DRAKE, ASA  C/P,P
WARRIOR WITHIN, THE [TERRILIAN 1]  GREEN, SHARON  C/P:2,P
WARRIOR WOMAN  BRADLEY, MARION ZIMMER  C/P:2,P
WARRIORS, THE [GUARDIANS OF THE FLAME 1,2,3]  ROSENBERG, JOEL  H:2
WARRIOR’S APPRENTICE, THE [MILES VORKOSIGAN 1]  BUJOLD, LOIS MCMASTER  C/P:2,P
WARRIOR’S APPRENTICE, THE (REVISED) [MILES VORKOSIGAN 1]  BUJOLD, LOIS MCMASTER  H
WARRIOR’S BLOOD [BUCK ROGERS (NIVEN & POURNELLE’S) 3]  MCENROE, RICHARD S.  P
WARRIOR’S BLOOD, THE [EINARINN 4]  MCKENNA, JULIET E.  P
WARRIORS FOR THE WORKING DAY [STARCRUISER SHENANDOAH 6]  GREEN, ROLAND J.  P
WARRIOR’S HEART [@DRAGONLANCE GOODLUND #1]  SULLIVAN, STEPHEN D.  LP
WARRIOR’S JOURNEY, A [@DRAGONLANCE ERGOTH #1]  THOMPSON, PAUL B./COOK, TONYA C.  C/P,P
WARRIORS OF BLOOD AND DREAM  ZELAZNY, ROGER  C/PA,PA
WARRIORS OF DAWN, THE [LER 2]  FOSTER, M. A.  C/P:2,P
WARRIORS OF DAY, THE  BLISH, JAMES  SR-P,C/P:2,GN3.16,P
WARRIORS OF ISIS [ISIS 3]  STEWART, JEAN  LP
WARRIORS OF LATAN [@BLADE #37]  LORD, JEFFREY  P
WARRIORS OF MARS [MICHAEL KANE 1]  BRADBURY, EDWARD P.  SR-P
WARRIORS OF MARS [MICHAEL KANE 1]  MOORCOCK, MICHAEL  P
WARRIORS OF SPIDER, THE [WARRIORS OF SPIDER 1]  GEAR, W. MICHAEL  C/P:2,P
WARRIORS OF TERRA, THE  FAUCETTE, JOHN M.  C/P,P
WARRIORS OF THE DEEP [DOCTOR WHO]  DICKS, TERRANCE  SR-P,C/P:2
WARRIORS OF THE RAINBOW  HARVEY, A. D.  LP
WARRIORS OF THE STORM [@RINGS OF THE MASTER #3]  CHALKER, JACK L.  C/P:2,P
WARRIORS OF THE WASTELAND [HUNTSMAN 2]  HILL, DOUGLAS  H
WARRIORS OF THLASSA MEY [THLASSA MEY 2]  MCCARTY, DENNIS  P
WARRIORS OF ULTRAMAR [WARHAMMER 40K ULTRAMARINES 2]  MCNEILL, GRAHAM  P
WARRIOR’S STAR, A  KENNETT, RICK  SR-P
WARRIOR’S TALE, THE  COLE, ALLAN/BUNCH, CHRIS  H,P
WARRIOR’S WOMAN  LINDSEY, JOHANNA  P
WARRIOR’S WORLD [BUCK ROGERS (NIVEN & POURNELLE’S) 4]  MCENROE, RICHARD S.  P
WARRIORWARDS [JIANA 2]  AB HUGH, DAFYDD  C/P:3,P
WARS OF VIS, THE [VIS 1,2]  LEE, TANITH  H
WARSLAYER, THE  EDGHILL, ROSEMARY  P
WARSPRITE  SWYCAFFER, JEFFERSON P.  P
WARSTALKER’S TRACK  DEITZ, TOM  P
WARSTRIDER [@WARSTRIDER 1]  KEITH, WILLIAM H. JR.  P,SFWA-P

The MITSFS Titledex  September 24, 2007  Page 403
TITLE  AUTHOR  CODE
WARSTRIDER: BATTLEMIND [@WARSTRIDER 6]  KEITH, WILLIAM H. JR.  P
WARSTRIDER: JACKERS [@WARSTRIDER 3]  KEITH, WILLIAM H. JR.  P
WARSTRIDER: NETLINK [@WARSTRIDER 5]  KEITH, WILLIAM H. JR.  C/P,P
WARSTRIDER: REBELLION [@WARSTRIDER 2]  KEITH, WILLIAM H. JR.  P,SFWA-P
WARSTRIDER: SYMBIONTS [@WARSTRIDER 4]  KEITH, WILLIAM H. JR.  P
WARTIDE [@TIMERAIDER]  BARNES, JOHN  P
WASHED BY A WAVE OF WIND  BELL, M. SHAYNE  LP
WASHINGTON’S DIRIGIBLE [@TIMELINE WARS #2]  BARNES, JOHN  C/P:2,P
WASP  RUSSELL, ERIC FRANK  H:2,P
WASP (UNABRIDGED)  RUSSELL, ERIC FRANK  C/P:2,P
WASTE LANDS, THE [@DARK TOWER #3]  KING, STEPHEN  C/P,SFWA,LP,H
WASTELAND OF FLINT [TIME OF THE SIXTH SUN 1]  HARLAN, THOMAS  H
WAT IS UW I.Q.  EYSENCK, H. J.  FP
WATCH, THE  DANVERS, DENNIS  H,LP
WATCH BELOW, THE  WHITE, JAMES  C/P:2,P
WATCH ON THE RHINE [POSLEEN]  RINGO, JOHN/KRATMAN, TOM  H
WATCH, THE (PROOF)  DANVERS, DENNIS  SR-LP
WATCH THE NORTHWIND RISE  GRAVES, ROBERT  C/P,P
WATCHER [@WORLD OF DARKNESS WEREWOLF]  GRANT, CHARLES L.  C/P,P
WATCHER, THE  PALMER, JANE  SR-P
WATCHER IN THE WOODS, THE  RANDALL, FLORENCE ENGEL  P
WATCHERS, THE  FOWLER, RAYMOND E.  [C/RAN-H]
WATCHERS  KOONTZ, DEAN R.  C/P,SFWA-P,P,H
WATCHERS AT THE WELL, THE [WELL WORLD 6,7,8]  CHALKER, JACK L.  H:2
WATCHER’S MASK, THE  MARKS, LAURIE J.  P
WATCHERS OF THE DARK [JAN DARZEK]  BIGGLE, LLOYD JR.  H:2
WATCHERS OUT OF TIME AND OTHERS, THE  LOVECRAFT, H. P./DERLETH, AUGUST  SR-H,H
WATCHERS: THE COMING OF THE KING  MEIKLE, WILLIAM  LP
WATCHING TREES GROW  HAMILTON, PETER F.  GZD07305
WATCHING WORLD, THE  FANTHORPE, R. L.  SR-P
WATCHMEN, THE  BOVA, BEN  SFWA-P
WATCHSTAR  SARGENT, PAMELA  H:2,P
WATCHTOWER [TORNOR 1]  LYNN, ELIZABETH A.  C/P,P,H,SFWA
WATER  MCKINLEY, ROBIN/DICKINSON, PETER  LP
WATER BABIES, THE  KINGSLEY, CHARLES  H
WATER GHOST AND OTHERS, THE  BANGS, JOHN KENDRICK  H
WATER KING’S LAUGHTER, THE [TWELVE KINGDOMS 4]  FRIESNER, ESTHER M.  C/P:2,P
WATER LOGIC [ELEMENTAL LOGIC 3]  MARKS, LAURIE J.  LP
WATER OF THE WONDROUS ISLES, THE  MORRIS, WILLIAM  C/P:2,P
WATER OF THOUGHT, THE  SABERHAGEN, FRED  DM127
WATER OF THOUGHT, THE (REVISED)  SABERHAGEN, FRED  C/P:2,SFWA-P,P

WATER SLEEPS [@BLACK COMPANY GLITTERING STONE 3/ BLACK COMPANY 8] COOK, GLEN

H

WATER SPRITES, THE [FAIRY REALM 8]  RODDA, EMILY  H
WATER WITCH  FELICE, CYNTHIA/WILLIS, CONNIE  C/P:2,P
WATER WORLD  APPEL, WILLIAM  P
WATERBORN, THE [CHILDREN OF THE CHANGELING 1]  KEYES, J. GREGORY  H,P
WATERDANCE  LOGSTON, ANNE  P
WATERDEEP [FORGOTTEN REALMS AVATAR 3]  AWLINSON, RICHARD  C/P:2,P
WATERDEEP [@FORGOTTEN REALMS AVATAR #3]  DENNING, TROY  P
WATERS OF CENTAURUS, THE  BROWN, ROSEL GEORGE  H:2,P
WATERS OF DEATH  GREENFIELD, IRVING A.  C/P:2,P
WATERSHIP DOWN [WATERSHIP DOWN 1]  ADAMS, RICHARD  C/P,P
WATERSLEEP [@DEATHLANDS]  AXLER, JAMES  P
WATERSPOUT UP  ETKIN, ANNE  B
WATERWORLD  COLLINS, MAX ALLAN  P
WATERY GRAVE [TWILIGHT #18]  TRAINOR, JOSEPH  P
WAVE, THE  MOSLEY, WALTER  H

WAVE AND THE FLAME, THE [LEAR’S DAUGHTERS 1] KELLOGG, MARJORIE BRADLEY/

ROSSOW, WILLIAM B.

P

WAVE RIDER  SCHENCK, HILBERT  P
WAVE RIDER (PROOF)  SCHENCK, HILBERT  SR-LP
WAVE WITHOUT A SHORE  CHERRYH, C. J.  C/P:2,P
WAVES  FOSTER, M. A.  C/P:2,P
WAY BACK, THE [GRIMES]  CHANDLER, A. BERTRAM  P
WAY BENEATH, THE [@KINGDOMS #3]  WELLS, ANGUS  C/P,P
WAY BETWEEN THE WORLDS, THE [@VIEW FROM THE MIRROR #4]  IRVINE, IAN  P
WAY-FARER  SCHMIDT, DENNIS  C/P:2,P
WAY HOME, A  STURGEON, THEODORE  C/P:2,P
WAY OF LIGHT, THE [MAGRAVANDIAS 3]  CONSTANTINE, STORM  H:2
WAY OF SPIDER, THE [WARRIORS OF SPIDER 2]  GEAR, W. MICHAEL  C/P:2,P
WAY OF STARS: A ROMANCE OF REINCARNATION, THE  BECK, L. ADAMS  H
WAY OF THE CLANS [@BATTLETECH #1/JADE PHOENIX 1]  THURSTON, ROBERT  C/P:2,SFWA-P:2,P
WAY OF THE DEAD [@WARHAMMER]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA

The MITSFS Titledex  September 24, 2007  Page 404
TITLE  AUTHOR  CODE
WAY OF THE GODS  KUTTNER, HENRY  B
WAY OF THE PILGRIM  DICKSON, GORDON R.  H:2,P,SFWA
WAY OF THE ROSE, THE [EVERIEN 3]  LEITH, VALERY  LP
WAY OF THE WARRIOR, THE [STAR TREK DS9]  CAREY, DIANE  P
WAY OF THE WARRIOR, THE (PART 2) [@YAMATO #2]  KATO, KEN  P
WAY OF THE WOLF [@DEATHLANDS]  AXLER, JAMES  P
WAY OF THE WOLF [@VAMPIRE EARTH #1]  KNIGHT, E. E.  P
WAY OUT  ELWOOD, ROGER  SR-HA
WAY OUT  HOWARD, IVAN  PA,C/PA
WAY OUT WORLD, THE  NEBEL, LONG JOHN  [C/RAN-P]
WAY STATION  SIMAK, CLIFFORD D.  H:2,P,SFWA

WAY THE FUTURE WAS, THE POHL, FREDERIK R/REF-P,C/REF-H:2,

C/REF-P

WAY TO DAWNWORLD, THE [FARSTAR & SON 1]  STARR, BILL  C/P:2,P
WAY TO GLORY, THE [RCN 4]  DRAKE, DAVID  H
WAY TO GROUND ZERO, THE  BARTTER, MARTHA A.  R/REF-H
WAYFARER REDEMPTION, THE [WAYFARER REDEMPTION 1]  DOUGLASS, SARA  P
WAYFINDER’S STORY [@LOST SWORDS #7]  SABERHAGEN, FRED  H:2
WAYLANDER [DRENAI WAYLANDER 1]  GEMMELL, DAVID  C/P,P
WAYS OF MAGIC, THE [@ELVEN WAYS #1]  CIENCIN, SCOTT  P
WAYS OF MAGIC, THE [MENNENKALTENEI 1,2]  TURNER, DELIA MARSHALL  H
WAYWARD KNIGHTS, THE [@DRAGONLANCE WARRIORS #7]  GREEN, ROLAND J.  P
WAYWARD MOON [@ASTROLOGER 2]  DEMARTINO, DENNY  P

WAYWARD WIZARD, THE [@DRAGONLANCE NEW ADVENTURES SUNCATCHER #1] SAMPSON, JEFF

LP

WE (GREGORY ZILBOORG TRANSLATION)  ZAMIATIN, EUGENE  C/P:2,P
WE (MIRRA GINSBURG TRANSLATION)  ZAMYATIN, YEVGENY  C/P,P,H
WE ALL DIED AT BREAKAWAY STATION  MEREDITH, RICHARD C.  C/P:2,P
WE ARE ALL LEGENDS  SCHWEITZER, DARRELL  H,LP
WE ARE NOT ALONE  SULLIVAN, WALTER  [C/RAN-P:2]
WE ARE NOT THE FIRST  TOMAS, ANDREW  [C/RAN-P]
WE CAN BUILD YOU  DICK, PHILIP K.  C/P:2,P
WE CLAIM THESE STARS!  ANDERSON, POUL  C/P:2,DD407,P
WE FEW [MARCH UPCOUNTRY 4]  WEBER, DAVID/RINGO, JOHN  H
WE FIGHT FOR THE FUTURE  MATHEWS, BASIL  [C/RAN-P]
WE HAVE ALWAYS LIVED IN THE CASTLE  JACKSON, SHIRLEY  C/P,P
WE, IN SOME STRANGE POWER’S EMPLOY, MOVE ON A RIGOROUS LINE  DELANY, SAMUEL R.  TD21,SFWA-TD21
WE OPEN ON VENUS [@STARSHIP TROUPERS #2]  STASHEFF, CHRISTOPHER  C/P:2,P
WE REACH THE MOON  WILFORD, JOHN NOBLE  [C/RAN-P]
WE, THE VENUSIANS  RACKHAM, JOHN  DM127
WE WHO ARE ABOUT TO...  RUSS, JOANNA  C/P,P
WE WHO SURVIVED THE FIFTH ICE AGE  NOEL, STERLING  C/P:2,P
WEALDWIFE’S TALE, THE  HAZEL, PAUL  P,SFWA
WEALTH OF THE VOID  STATTEN, VARGO  SR-P
WEANS, THE  NATHAN, ROBERT  H
WEAPON  MASON, ROBERT  C/P:2,P
WEAPON, THE [FREEHOLD WAR 2]  WILLIAMSON, MICHAEL Z.  H
WEAPON FROM BEYOND, THE [@STARWOLF #1]  HAMILTON, EDMOND  C/P:2,P
WEAPON MAKERS, THE  VAN VOGT, A. E.  C/P,P,H
WEAPON SHOPS OF ISHER, THE  VAN VOGT, A. E.  SR-DD053,C/P:2,P
WEAPONS OF CHOICE [AXIS OF TIME 1]  BIRMINGHAM, JOHN  LP
WEAR THE BUTCHERS’ MEDAL  BRUNNER, JOHN  P
WEASEL’S LUCK [@DRAGONLANCE HEROES 3]  WILLIAMS, MICHAEL  C/P:2,P
WEATHER WAR  LEOKUM, LEONARD/POSNICK, PAUL  P
WEATHERMAKERS, THE  BOVA, BEN  C/P:2,P
WEATHERMONGER, THE [CHANGES 3]  DICKINSON, PETER  C/P,P,H
WEAVERS OF SARAMYR [BRAIDED PATH 1]  WOODING, CHRIS  LP
WEAVERS OF WAR [WINDS OF THE FORELANDS 5]  COE, DAVID B.  H
WEAVEWORLD  BARKER, CLIVE  C/P,P,H
WEB, THE [@SURVIVALIST #5]  AHERN, JERRY  P
WEB, THE [ALIEN SPEEDWAY 3]  WYLDE, THOMAS  C/P:2,P
WEB  WYNDHAM, JOHN  SR-P
WEB 2027  MILLENNIUM  PA
WEB BETWEEN THE WORLDS, THE  SHEFFIELD, CHARLES  C/P:2,LP
WEB OF ANGELS  FORD, JOHN M.  C/P,P
WEB OF DARKNESS [WEB 2]  BRADLEY, MARION ZIMMER  C/P:2,LP
WEB OF DEFEAT [QUEST FOR THE WHITE DUCK 2]  FENN, LIONEL  C/P,P
WEB OF EVERYWHERE  BRUNNER, JOHN  C/P:2,P
WEB OF FEAR, THE [DOCTOR WHO]  DICKS, TERRANCE  P
WEB OF FUTURES  SWYCAFFER, JEFFERSON P.  C/P,P
WEB OF LIGHT [WEB 1]  BRADLEY, MARION ZIMMER  C/P:2,P
WEB OF SAND [DUMAREST 20]  TUBB, E. C.  C/P:2,P
WEB OF SPIDER, THE [WARRIORS OF SPIDER 3]  GEAR, W. MICHAEL  C/P:2,P
WEB OF THE CHOZEN, THE  CHALKER, JACK L.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 405
TITLE  AUTHOR  CODE
WEB OF THE CITY  ELLISON, HARLAN  P
WEB OF THE ROMULANS [STAR TREK]  MURDOCK, M. S.  C/P:3,P
WEB OF THE SPIDER [WAR OF THE WIZARDS 3]  OFFUTT, ANDREW J./LYON, RICHARD K.  P
WEB OF THE WITCH WORLD [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P:2,P
WEB OF WIND [SILVERGLASS 2]  RIVKIN, J. F.  P
WEB OF WIZARDRY, THE  COULSON, JUANITA  C/P:2,P
WEB SITE STORY  RANKIN, ROBERT  P
WEBMAGE  MCCULLOUGH, KELLY  P
WEBS  BAKER, SCOTT  SFWA-P
WEBSTER’S NEW WORLD THESAURUS  LAIRD, CHARLTON  DICT
WEDGE’S GAMBLE [@STAR WARS X-WING 2]  STACKPOLE, MICHAEL A.  C/P,P
WEE FREE MEN, THE [DISCWORLD 29/DISCWORLD TIFFANY ACHING 1]  PRATCHETT, TERRY  H,P
WEEPING MAY TARRY  JONES, RAYMOND F./DEL REY, LESTER  P
WEIGHER, THE  VINICOFF, ERIC/MARTIN, MARCIA  C/P,P
WEIGHTLESS IN GAZA  SHANNON, FRED  P
WEIRD GATHERING AND OTHER TALES, THE  CURRAN, RONALD  C/PA,PA
WEIRD HENRY BERG  SARGENT, SARAH  H
WEIRD HEROES 1  PREISS, BYRON  C/PA,PA
WEIRD HEROES 2  PREISS, BYRON  PA
WEIRD HEROES 6  PREISS, BYRON  PA
WEIRD HEROES 8  PREISS, BYRON  C/PA,PA
WEIRD MENACE, THE  JONES, BOB  R/IND
WEIRD MENACE #1  WEINBERG, ROBERT  SR-LPA
WEIRD OF THE WHITE WOLF, THE [ELRIC 3]  MOORCOCK, MICHAEL  C/P:2,P
WEIRD ONES, THE  BELMONT BOOKS  PA,C/PA

WEIRD ROMANCE (CD)

BRENNERT, ALAN/MENKEN, ALAN/ SPENCER, DAVID

SR-MM

WEIRD SHADOW OVER INNSMOUTH, THE  LOVECRAFT, H. P.  SR-P
WEIRD SHADOWS FROM BEYOND  CARNELL, JOHN  PA,C/PA
WEIRD SHOW  PLAYBOY BOOKS  C/PA,PA
WEIRD STORIES  RIDDELL, J. H.  H
WEIRD TALES  MARGULIES, LEO  C/PA,PA
WEIRD TALES #1  CARTER, LIN  SR-PA
WEIRD TALES 1  HAINING, PETER  SR-PA
WEIRD TALES #2  CARTER, LIN  SR-PA
WEIRD TALES #3  CARTER, LIN  SR-PA
WEIRD TALES #4  CARTER, LIN  SR-PA
WEIRD TALES FROM SHAKESPEARE  KERR, KATHARINE/GREENBERG, MARTIN H. PA,C/PA
WEIRD TALES OF TERROR AND DETECTION  HEARD, H. F.  H
WEIRD TALES: THE MAGAZINE THAT NEVER DIES  KAYE, MARVIN  HA
WEIRDSTONE OF BRISINGAMEN, THE [ALDERLY 1]  GARNER, ALAN  C/P:2,P
WEIRWOODS, THE  SWANN, THOMAS BURNETT  C/P:2,P
WEISSENBAUM’S EYE  STETTEN, GEORGE  C/P,P
WELCOME, CHAOS  WILHELM, KATE  C/P,SFWA,P,H
WELCOME HOME, JAIME [BIONIC WOMAN 1]  LOTTMAN, EILEEN  P
WELCOME TO MARS  BLISH, JAMES  H,P
WELCOME TO THE MONKEY HOUSE  VONNEGUT, KURT JR.  C/P:2,P
WELL, THE  SIROTA, MIKE  P
WELL AT THE WORLD’S END, THE  MORRIS, WILLIAM  P
WELL AT THE WORLD’S END 1, THE  MORRIS, WILLIAM  C/P:2,P
WELL AT THE WORLD’S END 2, THE  MORRIS, WILLIAM  C/P:2,P
WELL-FAVORED MAN, THE [LANDUC 1]  WILLEY, ELIZABETH  H,P
WELL-MANNERED WAR, THE [DOCTOR WHO MISSING ADVENTURES]  ROBERTS, GARETH  C/P:2,P
WELL OF DARKNESS, THE [GANDALARA 4]  GARRETT, RANDALL/HEYDRON, VICKI ANN C/P:2,P
WELL OF DARKNESS [SOVEREIGN STONE 1]  WEIS, MARGARET/HICKMAN, TRACY  H
WELL OF ETERNITY, THE [@WARCRAFT WAR OF THE ANCIENTS #1]  KNAAK, RICHARD A.  P
WELL OF LOST PLOTS, THE [THURSDAY NEXT 3]  FFORDE, JASPER  H
WELL OF SHADES, THE [BRIDEI 3]  MARILLIER, JULIET  H
WELL OF SHIUAN [MORGAINE 2]  CHERRYH, C. J.  C/P:2,P
WELL OF SOULS [@STAR TREK LOST ERA (2336 AD)]  BICK, ILSA J.  P
WELL OF STARS, THE [MARROW 2]  REED, ROBERT  H
WELL OF THE UNICORN, THE  FLETCHER, GEORGE U.  H
WELL OF THE UNICORN, THE  PRATT, FLETCHER  C/P:2,P
WELL OF THE WORLDS, THE  KUTTNER, HENRY  C/P:2,P
WELL OF THE WORLDS  PADGETT, LEWIS  C/P:2,GN3.17
WELL-TIMED ENCHANTMENT, A  VELDE, VIVIAN VANDE  H
WELL-WISHERS, THE [MAGIC OR NOT? 2]  EAGER, EDWARD  LP
WELLENLAUFER, DIE  MEYER, KAI  FH
WELLSPRING OF CHAOS [RECLUCE 12]  MODESITT, L. E. JR.  H
WELLSPRINGS OF LIFE, THE  ASIMOV, ISAAC  [C/RAN-P]
WELLSTONE, THE [QUEENDOM OF SOL 2]  MCCARTHY, WIL  P
WELSH CELTIC MYTH IN MODERN FANTASY  SULLIVAN, C. W. III  R/REF-H
WELTEN IM BRAND  BINGS, HENRY  FP
WEREBEASTS OF HEL [HEL 3]  DRAKE, ASA  P

The MITSFS Titledex  September 24, 2007  Page 406
TITLE  AUTHOR  CODE
WEREBLOOD  IVERSON, ERIC  P
WEREHUNTER  LACKEY, MERCEDES  C/P:2,P
WERENIGHT  IVERSON, ERIC  P
WERENIGHT [GERIN THE FOX 1]  TURTLEDOVE, HARRY  P,SFWA-P
WEREWOLF [@WORLD OF DARKNESS PREDATOR & PREY #3]  FLEMING, GHERBOD  P
WEREWOLF  HOUSEMAN, CLEMENCE  VLP
WEREWOLF!  PRONZINI, BILL  HA,PA
WEREWOLF AMONG US, A  KOONTZ, DEAN R.  C/P:2,P
WEREWOLF IN LEGEND, FACT, AND ART, THE  COPPER, BASIL  R/REF-H
WEREWOLF OF LONDON, THE  DREADSTONE, CARL  P
WEREWOLF OF PARIS, THE  ENDORE, GUY  SR-P,H:2
WEREWOLF OF PONKERT, THE  MUNN, H. WARNER  P
WEREWOLF PRINCIPLE, THE  SIMAK, CLIFFORD D.  H:2,P
WEREWOLF WALKS TONIGHT, THE [@SATAN SLEUTH #2]  AVALLONE, MICHAEL  P
WEREWOLF’S KISS, THE [MARLEY FAMILY WEREWOLF 1]  SCOTCH, CHERI  P
WEREWOLF’S REVENGE, THE [JAMES UNDERHILL 2]  JACCOMA, RICHARD  P
WEREWOLF’S TALE, THE [JAMES UNDERHILL 1]  JACCOMA, RICHARD  P
WEREWOLF’S TOUCH, THE [MARLEY FAMILY WEREWOLF 2]  SCOTCH, CHERI  P
WEREWOLVES OF KREGEN [DRAY PRESCOT 33]  AKERS, ALAN BURT  P
WEREWOLVES OF LONDON, THE [DAVID LYDYARD 1]  STABLEFORD, BRIAN M.  C/P,P,H
WEST OF EDEN [EDEN 1]  HARRISON, HARRY  C/P,P,H
WEST OF HONOR  POURNELLE, JERRY  P
WEST OF HONOR (REVISED)  POURNELLE, JERRY  C/P:2,P
WEST OF JANUARY  DUNCAN, DAVE  C/P:2,P
WEST OF THE SUN  PANGBORN, EDGAR  H:2,P
WESTERCON 37 PROGRAM BOOK  CROSS, DEBBIE  R/REF-P
WESTERN KING, THE [RUNE BLADE 2]  MARSTON, ANN  P
WESTERN SHORE [@ALDABRESHIN COMPASS #3]  MCKENNA, JULIET E.  P
WESTERN WIZARD, THE [RENSHAI 2]  REICHERT, MICKEY ZUCKER  C/P,P
WESTMARK [WESTMARK 1]  ALEXANDER, LLOYD  C/P,P
WESTWORLD  CRICHTON, MICHAEL  C/P:2,P
WETWARE  NOVA, CRAIG  LP
WETWARE [WARE 2]  RUCKER, RUDY  C/P:2,P
WHALE SINGER  KATZ, WELWYN WILTON  SFWA
WHALES ON STILTS [THRILLING TALES 1]  ANDERSON, M. T.  H
WHALESONG  SIEGEL, ROBERT  C/P,LP
WHAT DID I DO TOMORROW?  DAVIES, L. P.  H
WHAT DID MISS DARLINGTON SEE?  SALMONSON, JESSICA AMANDA  LP
"WHAT DO YOU CARE WHAT OTHER PEOPLE THINK?"  FEYNMAN, RICHARD P.  [C/RAN-H]
WHAT DREAMS MAY COME  MATHESON, RICHARD  P
WHAT DREAMS MAY COME [JOHN THUNSTONE]  WELLMAN, MANLY WADE  H,SFWA
WHAT ENTROPY MEANS TO ME  EFFINGER, GEORGE ALEC  C/P,SFWA,P,H
WHAT HAPPENED TO EMILY GOODE AFTER THE GREAT EXHIBITION  MOORE, RAYLYN  H,LP
WHAT HO, MAGIC!  HUFF, TANYA  LP
WHAT IF? 2  COWLEY, ROBERT  HA
WHAT IF? VOLUME 1  LUPOFF, RICHARD A.  PA
WHAT LAW THERE WAS  DEMPSEY, AL  SFWA
WHAT LAY BEYOND [STAR TREK GATEWAYS #7]  PARAMOUNT PICTURES  HA,PA
WHAT MAD UNIVERSE  BROWN, FREDRIC  C/P,P,H

WHAT MIGHT HAVE BEEN I AND II [WHAT MIGHT HAVE BEEN 1,2] BENFORD, GREGORY/

GREENBERG, MARTIN H.

HA

WHAT MIGHT HAVE BEEN III AND IV [WHAT MIGHT HAVE BEEN 3,4] BENFORD, GREGORY/

GREENBERG, MARTIN H.

HA

WHAT PRICE HONOR? [STAR TREK ENTERPRISE]  STERN, DAVE  P
WHAT PRICE VICTORY [WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
WHAT ROUGH BEAST  WATKINS, WILLIAM JOHN  C/P:2,P
WHAT SAVAGE BEAST [HULK]  DAVID, PETER  P
WHAT STRANGE STARS AND SKIES  DAVIDSON, AVRAM  C/P:2,P
WHAT YOU LEAVE BEHIND [@STAR TREK DS9]  CAREY, DIANE  P
WHAT YOU MAKE IT  SMITH, MICHAEL MARSHALL  LP
WHAT YOUR AURA TELLS ME  STANFORD, RAY  SFWA
WHATDUNITS  RESNICK, MIKE  SFWA-P
WHAT’S A GHOUL TO DO? [@GHOST HUNTER]  LAURIA, VICTORIA  P
WHAT’S BECOME OF SCREWLOOSE? AND OTHER INQUIRIES  GOULART, RON  H:2,SFWA,P
WHAT’S BIG AND YELLOW AND PUTS OUT FLAMES?  BANANA, PLASTIC  C/P,P
WHAT’S IT LIKE OUT THERE? AND OTHER STORIES  HAMILTON, EDMOND  P
WHEEL IN THE SKY, THE  BERNARD, RAFE  SR-P
WHEEL OF DREAMS  TYSON, SALINDRA  C/P,P
WHEEL OF FORTUNE  ZELAZNY, ROGER  C/PA,PA
WHEEL OF STARS  NORTON, ANDRE  C/P,SFWA-P,LP,H
WHEEL OF THE INFINITE  WELLS, MARTHA  H,P
WHEEL OF THE WINDS  ENGH, M. J.  C/P,SFWA-P,P,H
WHEEL OF TIME, THE (SYNOPSIS) [WHEEL OF TIME 1,2]  JORDAN, ROBERT  SR-L
WHEELERS  STEWART, IAN/COHEN, JAMES  H:2,P

The MITSFS Titledex  September 24, 2007  Page 407
TITLE  AUTHOR  CODE
WHEELS OF FIRE [SERRATED EDGE 2]  LACKEY, MERCEDES/SHEPHERD, MARK  C/P:2,SFWA-P,P

WHEELS OF IF, THE DE CAMP, L. SPRAGUE C/P,SFWA-TD20,H,P,

TD20

WHEELS OF PHYRE, THE [ARBITER 2]  DOUGLAS, L. WARREN  P
WHEELS WITHIN WHEELS [REPAIRMAN JACK]  WILSON, F. PAUL  C/P,SFWA,P,H
WHEELWORLD [TO THE STARS 2]  HARRISON, HARRY  C/P:2,P
WHEN CHURCHYARDS YAWN  ASQUITH, CYNTHIA  SR-PA
WHEN DARKNESS FALLS [OBSIDIAN 3]  LACKEY, MERCEDES/MALLORY, JAMES  H
WHEN DEATH BIRDS FLY [CORMAC 3]  OFFUTT, ANDREW J./TAYLOR, KEITH  C/P,P
WHEN DEMONS WALK  BRIGGS, PATRICIA  C/P,P
WHEN DRAGONS RAGE [DRAGONCROWN WAR 2]  STACKPOLE, MICHAEL A.  C/P,LP
WHEN DREAMERS CEASE TO DREAM [TIME KEEPER 3]  BARTHOLOMEW, BARBARA  P
WHEN ELEPHANTS LAST IN THE DOORYARD BLOOMED  BRADBURY, RAY  H
WHEN GOD LAUGHS  LONDON, JACK  H
WHEN GRAVITY FAILS [MARID AUDRAN 1]  EFFINGER, GEORGE ALEC  C/P,SFWA,P,H
WHEN HARLIE WAS ONE  GERROLD, DAVID  C/P,P,H
WHEN HARLIE WAS ONE, RELEASE 2.0  GERROLD, DAVID  C/P,P
WHEN HEAVEN FELL  BARTON, WILLIAM  P
WHEN HELL LAUGHS [@RED SONJA #3]  SMITH, DAVID C./TIERNEY, RICHARD L.  P
WHEN NEXT I WAKE  DORN, FRANK  P
WHEN THE BIRDS FLY SOUTH  COBLENTZ, STANTON A.  LP
WHEN THE BLACK LOTUS BLOOMS  SAUNDERS, ELIZABETH A.  SR-LPA
WHEN THE BOUGH BREAKS [SERRATED EDGE 3]  LACKEY, MERCEDES/LISLE, HOLLY  C/P:2,SFWA-P,P
WHEN THE CHANGEWINDS BLOW [CHANGEWINDS 1]  CHALKER, JACK L.  C/P:2,P
WHEN THE DEVIL DANCES [POSLEEN 3]  RINGO, JOHN  H
WHEN THE EARTH DIED  MANNHEIM, KARL  P
WHEN THE GODS ARE SILENT  LINDSKOLD, JANE  P
WHEN THE GODS CAME  ADAMS, JOHN  SR-P
WHEN THE GODS RETURNED  BEAMER, CHARLES  P
WHEN THE GREEN STAR CALLS [GREEN STAR 2]  CARTER, LIN  C/P:2,P
WHEN THE IDOLS WALKED [BRAK 4]  JAKES, JOHN  P
WHEN THE KING COMES HOME  STEVERMER, CAROLINE  P
WHEN THE KISSING HAD TO STOP  FITZGIBBON, CONSTANTINE  C/P,P,H
WHEN THE MOON DIED  SAVAGE, RICHARD  SR-P
WHEN THE MOON RAN WILD  AINSBURY, RAY  SR-P
WHEN THE MUSIC’S OVER  SHINER, LEWIS  PA
WHEN THE RED KING WOKE  KELLEAM, JOSESPH E.  H
WHEN THE SKY BURNED  BOVA, BEN  H,SFWA,P
WHEN THE SLEEPER WAKES  WELLS, H. G.  C/P:2,P
WHEN THE STAR KINGS DIE  JAKES, JOHN  C/P:2,P
WHEN THE TRIPODS CAME [TRIPODS 0]  CHRISTOPHER, JOHN  C/P,SFWA-P,P
WHEN THE WAKER SLEEPS  GOULART, RON  C/P:2,P
WHEN THE WHITES WENT  BATEMAN, ROBERT  SR-P
WHEN THE WIND BLOWS  SAUL, JOHN  H,P

WHEN THE WORLD SCREAMED AND OTHER STORIES [PROF. CHALLENGER 2] DOYLE, ARTHUR CONAN

LP

WHEN THE WORLD SHOOK  HAGGARD, H. RIDER  C/P,P
WHEN THEY COME FROM SPACE  CLIFTON, MARK  C/P,P,H
WHEN TIME STOOD STILL  ORKOW, BEN  P
WHEN TROUBLE BECKONS [MATHEW SWAIN 2]  MCQUAY, MIKE  P
WHEN TRUE NIGHT FALLS [COLDFIRE 2]  FRIEDMAN, C. S.  H:2,P
WHEN TWO WORLDS MEET  WILLIAMS, ROBERT MOORE  P
WHEN VOIHA WAKES  CHANT, JOY  C/P:2,P
WHEN WE DEAD AWAKEN  HOLT, JOHN R.  P
WHEN WE WERE GOOD  SKAL, DAVID J.  P
WHEN WE WERE REAL  BARTON, WILLIAM  P
WHEN WILL YOU RAGE? [@WORLD OF DARKNESS WEREWOLF]  WIECK, STEWART  PA
WHEN WOMEN RULE  MOSKOWITZ, SAM  HA
WHEN WORLDS COLLIDE [WHEN WORLDS COLLIDE 1]  WYLIE, PHILIP/BALMER, EDWIN  C/P,P,H:2
WHEN WORLDS RAGE [ORIANUS CREATION 2]  BURTON, D. SHANE  LP
WHENABOUTS OF BURR, THE  KURLAND, MICHAEL  C/P,P
WHERE ANGELS FEAR [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  LEVENE, REBECCA/WINSTONE, SIMON  C/P,P
WHERE ARE THE RUSSIANS?  BENTLEY, JOHN  P
WHERE DO WE GO FROM HERE?  ASIMOV, ISAAC  C/PA,PA,SFWA
WHERE DRAGONS LIE  SALSITZ, R. A. V.  P
WHERE IS EVERYBODY?  WEBB, STEPHEN  R/REF-H
WHERE IS THE BIRD OF FIRE?  SWANN, THOMAS BURNETT  C/P:2,P
WHERE LATE THE SWEET BIRDS SANG  WILHELM, KATE  SFWA-P,H:2,P,SFWA
WHERE NO MAN HAS GONE BEFORE [STAR TREK]  PEEPLES, SAMUEL A.  SR-S
WHERE NO MAN HAS TROD  DORER, NANCY/DORER, FRANCES  P
WHERE NO SUN SHINES  KELLEY, LEO P.  SR-P
WHERE SEA MEETS SKY [STAR TREK CAPTAIN’S TABLE 6]  OLTION, JERRY  C/P:2,P
WHERE SHE BRUSHED HER HAIR AND OTHER STORIES  STEELE, MAX  H
WHERE TERROR STALKED  BIRKIN, CHARLES  SR-P

The MITSFS Titledex  September 24, 2007  Page 408
TITLE  AUTHOR  CODE
WHERE THE EVIL DWELLS  SIMAK, CLIFFORD D.  H:2,P
WHERE THE NI-LACH  BENNETT, MARCIA J.  C/P:2,P
WHERE THE PAVEMENT ENDS  RUSSELL, JOHN  H
WHERE THE SHIPS DIE  DIETZ, WILLIAM C.  C/P,P
WHERE TIME WINDS BLOW  HOLDSTOCK, ROBERT P.  H:2,P
WHERE WERE YOU LAST PLUTERDAY?  VAN HERCK, PAUL  C/P:2,P
WHERE’S MY COW? [DISCWORLD]  PRATCHETT, TERRY  H
WHETTED BRONZE  NORVIL, MANNING  P
WHICH WAY TO THE FUTURE?  SCHMIDT, STANLEY  R/REF-H
WHIFF OF DEATH, A  ASIMOV, ISAAC  C/P:3,P
WHIFF OF MADNESS, A  GOULART, RON  P
WHILE THE CAT’S AWAY [@SABRINA THE TEENAGE WITCH #25]  BATRAE, MARGOT  P
WHIM OF THE DRAGON, THE [SECRET COUNTRY 3]  DEAN, PAMELA C.  C/P,P
WHIMS OF CREATION, THE  HAWKE, SIMON  C/P:2,P
WHIPPING STAR  HERBERT, FRANK  C/P,P,H
WHIRLIGIG OF TIME, THE [JAN DARZEK]  BIGGLE, LLOYD JR.  H
WHIRLPOOL OF STARS [@HOOK #1]  ZETFORD, TULLY  P
WHIRLWIND [X-FILES]  GRANT, CHARLES L.  C/P,P
WHIRLWIND ALCHEMY [CORWYN THE ALCHEMIST 2]  CLARK, DOUGLAS W.  C/P,SFWA-P,P
WHIRLWIND BENEATH THE SEA [AQUANAUTS 6]  STANTON, KEN  P

WHISKER OF HERCULES, THE [DOC SAVAGE #103]/ MAN WHO WAS SCARED, THE [DOC SAVAGE 104] ROBESON, KENNETH

@P

WHISKEY AND WATER [PROMETHEAN AGE 2]  BEAR, ELIZABETH  LP
WHISPER FROM THE STARS  SUTTON, JEFF  C/P,P
WHISPER OF BLOOD, A  DATLOW, ELLEN  PA
WHISPER OF TIME, A  DOWNING, PAULA E.  P
WHISPER OF WAVES [@FORGOTTEN REALMS WATERCOURSE #1]  ATHANS, PHILIP  P
WHISPERER AND OTHER STORIES, THE  LUMLEY, BRIAN  LP
WHISPERING BOX MYSTERY, THE [@RICK BRANT #5]  BLAINE, JOHN  H:2
WHISPERING GORILLA, THE  REED, DAVID V.  SR-P
WHISPERING NICKEL IDOLS [GARRETT 11]  COOK, GLEN  P
WHISPERING WOODS [MAGIC: THE GATHERING]  EMERY, CLAYTON  C/P:2,P
WHISPERS  KOONTZ, DEAN R.  P
WHISPERS, THE [GATES OF TIME 1]  PARKINSON, DAN  P
WHISPERS  SCHIFF, STUART DAVID  HA,PA,SFWA
WHISPERS FROM SPACE  MACVEY, JOHN W.  SFWA
WHISPERS FROM THE COTTON TREE ROOT  HOPKINSON, NALO  LP
WHISPERS FROM THE DARK SIDE OF TOMORROW  PETERS, OTHELLO  P
WHISPERS FROM THE PAST [@CHARMED]  NOONAN, ROSILAND  P
WHISPERS II  SCHIFF, STUART DAVID  HA,SFWA
WHISPERS III  SCHIFF, STUART DAVID  HA,SFWA
WHISPERS IV  SCHIFF, STUART DAVID  HA,SFWA
WHISPERS V  SCHIFF, STUART DAVID  HA,SFWA
WHISPERS VI  SCHIFF, STUART DAVID  HA,SFWA
WHISTLING WRAITH, THE [DOC SAVAGE]  ROBESON, KENNETH  C/P,P
WHITE ABACUS, THE  BRODERICK, DAMIEN  C/P,LP,H
WHITE APPLES  CARROLL, JONATHAN  LP
WHITE AS SNOW  LEE, TANITH  H:2
WHITE BULL, THE  SABERHAGEN, FRED  C/P:2,SFWA-P,P
WHITE DARKNESS [DOCTOR WHO NEW ADVENTURES]  MCINTEE, DAVID A.  C/P:2,P
WHITE DEVILS  MCAULEY, PAUL J.  P
WHITE DRAGON, THE [PERN DRAGONRIDERS 3]  MCCAFFREY, ANNE  C/P,P,H
WHITE DRAGON, THE [IN FIRE FORGED 1]  RESNICK, LAURA  H,P
WHITE EYES [DOC SAVAGE]  ROBESON, KENNETH  P
WHITE FIRE, THE [@JADE DEMONS #4]  VARDEMAN, ROBERT E.  P
WHITE FOG, THE  DENT, ROXANNE  P
WHITE GOLD WIELDER [THOMAS COVENANT 6]  DONALDSON, STEPHEN R.  C/P,P,H
WHITE GRYPHON, THE [MAGE WARS 2]  LACKEY, MERCEDES/DIXON, LARRY  C/P,P,H
WHITE GUARDIAN, THE [ETERNAL GUARDIANS 3]  CROSS, RONALD ANTHONY  H,LP
WHITE HART, THE [BOOKS OF THE ISLE 1]  SPRINGER, NANCY  H:2,P
WHITE HORSE, DARK DRAGON  FLEET, ROBERT C.  C/P,SFWA-P:2,P
WHITE HORSE WITH WINGS, A  DAVIES, ANTHEA  H
WHITE HOUSE HORRORS  GREENBERG, MARTIN H.  PA
WHITE JADE FOX, THE  NORTON, ANDRE  H:2,P
WHITE JENNA [ALTA 2]  YOLEN, JANE  C/P,P,H
WHITE LIES [ARCANE SOCIETY]  KRENTZ, JAYNE ANN  H
WHITE LIGHT  BARTON, WILLIAM/CAPOBIANCO, MICHAEL C/P,P,H
WHITE LIGHT  RUCKER, RUDY  P
WHITE LOTUS  HERSEY, JOHN  H:2
WHITE MARE, THE [DALRIADA 1]  WATSON, JULES  LP
WHITE MARE, RED STALLION  PAXSON, DIANA L.  P
WHITE MARE’S DAUGHTER [HORSE GODDESS PEOPLE 1]  TARR, JUDITH  H
WHITE MARS  ALDISS, BRIAN W./PENROSE, ROGER  H
WHITE MISTS OF POWER, THE  RUSCH, KRISTINE KATHRYN  C/P:2,P

TITLE  AUTHOR  CODE
WHITE MOON, RED DRAGON [@CHUNG KUO #6]  WINGROVE, DAVID  C/P:2,LP
WHITE MOUNTAIN, THE [@CHUNG KUO #3]  WINGROVE, DAVID  C/P,P,H
WHITE MOUNTAINS, THE [TRIPODS 1]  CHRISTOPHER, JOHN  C/P,P,H
WHITE NIGHT [DRESDEN FILES 9]  BUTCHER, JIM  H:2
WHITE NINJA  LUSTBADER, ERIC VAN  C/P,P
WHITE ORDER, THE [RECLUCE 8]  MODESITT, L. E. JR.  H:2,P
WHITE PAPERS, THE  WHITE, JAMES  SR-H
WHITE PIPES, THE  KRESS, NANCY  C/P,P,H
WHITE PLAGUE, THE  HERBERT, FRANK  C/P,P,H
WHITE PLUME MOUNTAIN [@GREYHAWK]  KIDD, PAUL  P
WHITE QUEEN [ALEUTIAN 1]  JONES, GWYNETH  H
WHITE RAVEN, THE  PAXSON, DIANA L.  P
WHITE REGIMENT, THE [REGIMENT 2]  DALMAS, JOHN  C/P:2,SFWA-P,P
WHITE ROSE, THE [BLACK COMPANY 3]  COOK, GLEN  C/P:2,SFWA-P,P
WHITE SILENCE [HIGHLANDER]  BUCHANAN, GINJER  P
WHITE SYBIL AND OTHER STORIES, THE  SMITH, CLARK ASHTON  H
WHITE TIME  LANAGAN, MARGO  H
WHITE TRIBUNAL, THE  VOLSKY, PAULA  C/P:2,LP
WHITE TYGER, THE  PARK, PAUL  H
WHITE VOYAGE, THE  CHRISTOPHER, JOHN  P
WHITE WIDOWS, THE  MERWIN, SAM JR.  P
WHITE WING  KENDALL, GORDON  C/P,SFWA-P,P
WHITE WOLF [DRENAI]  GEMMELL, DAVID  P
WHITEOUT  WALKER, SAGE  H,LP
WHITLEY STRIEBER’S ALIENS  STRIEBER, WHITLEY  PA
WHO?  BUDRYS, ALGIS  C/P:2,P
WHO CAN REPLACE A MAN?  ALDISS, BRIAN W.  C/P:2,P
WHO CENSORED ROGER RABBIT? [ROGER RABBIT 1]  WOLF, GARY K.  C/P:2,P
WHO FEARS THE DEVIL? [JOHN THE BALLADEER]  WELLMAN, MANLY WADE  SR-P,C/P
WHO GOES HERE?  SHAW, BOB  C/P,P
WHO GOES THERE?  CAMPBELL, JOHN W. JR.  SR-P,C/P,H
WHO HUNTS THE HUNTER [@SHADOWRUN 16]  SMITH, NYX  C/P,P
WHO IS JULIA?  HARRIS, BARBARA S.  P
WHO KILLED JAMES DEAN?  BEATH, WARREN NEWTON  SFWA-P
WHO KILLED KENNEDY [DOCTOR WHO]  STEVENS, JAMES/BISHOP, DAVID  C/P,P
WHO KNOCKS?  DERLETH, AUGUST  SR-HA
...WHO NEEDS ENEMIES? [HUMANX FLINX]  FOSTER, ALAN DEAN  C/P:2,P
WHO P-P-P-PLUGGED ROGER RABBIT? [ROGER RABBIT 2]  WOLF, GARY K.  H
WHO SPEAKS OF CONQUEST?  WRIGHT, LAN  DD205
WHO WAS THAT MONOLITH I SAW YOU WITH?  GOODWIN, MICHAEL  CX
WHODUNNIT [@SMALLVILLE]  SMITH, DEAN WESLEY  P
WHOLE (ALMOST) PROZINE CATALOG  SPECULATOR  SRL/IND
WHOLE MAN, THE  BRUNNER, JOHN  C/P:2,SFWA-P,P
WHOLE WIDE WORLD  MCAULEY, PAUL J.  H
WHOM THE GODS WOULD SLAY  JORGENSEN, IVAR  C/P,P
WHOPPER SPACE STORIES  CHILDREN’S PRESS  HA
WHO’S AFRAID OF BEOWULF?  HOLT, TOM  H:2,P
WHO’S WHO IN HORROR AND FANTASY FICTION  ASHLEY, MIKE  R/REF-H
WHO’S WHO IN SCIENCE FICTION  ASH, BRIAN  R/REF-H
WHO’S WHO IN SCIENCE FICTION FANDOM 1961  BROYLES, L. D.  R/REF-P
WHOSE SONG IS SUNG  SCHAEFER, FRANK  H,P
WHY CALL THEM BACK FROM HEAVEN?  SIMAK, CLIFFORD D.  C/P,P,H
WHY DO BIRDS  KNIGHT, DAMON  H
WHY DOES EVERYBODY THINK I’M NUTTY?  COLLIER, JAMES LINCOLN  H
WHY NOT YOU AND I?  WAGNER, KARL EDWARD  SFWA-P
WHY SHOULD I CUT YOUR THROAT?  VANDERMEER, JEFF  R/REF-H
WHY YOU SHOULD NEVER BEAM DOWN IN A RED SHIRT [STAR TREK]  BLY, ROBERT W.  R/REF-P,C/REF-P
WICKED  MAGUIRE, GREGORY  LP
WICKED CYBORG, THE  GOULART, RON  P
WICKED DAY, THE [ARTHURIAN 4]  STEWART, MARY  H:2,LP,P
WICKED ENCHANTMENT, THE  BENARY-ISBERT, MARGOT  P
WIDDERBURN HORROR, THE  WARNER-CROZETTI, R.  P
WIDDERSHINS [NEWFORD]  DE LINT, CHARLES  H
WIDE ANGLE LENS  FENNER, PHYLLIS R.  HA
WIDGET, THE WADGET, AND BOFF, THE  STURGEON, THEODORE  SFWA-TD9,TD9
WIDOWMAKER, THE [WIDOWMAKER 1]  RESNICK, MIKE  C/P,P
WIDOWMAKER REBORN, THE [WIDOWMAKER 2]  RESNICK, MIKE  C/P,P
WIDOWMAKER UNLEASHED, THE [WIDOWMAKER 3]  RESNICK, MIKE  P
WIDOWMAKERS, THE [WIDOWMAKER 1,2,3]  RESNICK, MIKE  H:2
WIDOW’S SON, THE [@HISTORICAL ILLUMINATUS #2]  WILSON, ROBERT ANTON  C/P,P
WIELAND: OR, THE TRANSFORMATION  BROWN, CHARLES BROCKDEN  P
WIELDING A RED SWORD [INCARNATIONS OF IMMORTALITY 4]  ANTHONY, PIERS  H:2,P,SFWA
WIELKOSC UROJONA  LEM, STANISLAW  FP
WILD, THE  STRIEBER, WHITLEY  SFWA-P

The MITSFS Titledex  September 24, 2007  Page 410
TITLE  AUTHOR  CODE
WILD, THE [REQUIEM FOR HOMO SAPIENS 3]  ZINDELL, DAVID  P
WILD ALIEN TAMER, THE [GALACTIC MIDWAY 3]  RESNICK, MIKE  C/P,P
WILD ANGEL [MAX MERRIWELL 2]  MURPHY, PAT  P
WILD ANIMALS  STRAUB, PETER  H
WILD BLOOD  LOGSTON, ANNE  C/P,P
WILD BLUE AND THE GRAY, THE  SANDERS, WILLIAM  P
WILD BOY, THE  ROCHELLE, WARREN  H
WILD BOYS, THE  BURROUGHS, WILLIAM S.  LP
WILD CARD RUN [KURTIS:P385XL47:RUTH 1]  STAMEY, SARA  P,SFWA-P
WILD CARDS [@WILD CARDS #1]  MARTIN, GEORGE R. R.  HA,PA
WILD COUNTRY [SYSTEMIC SHOCK 3]  ING, DEAN  C/P,SFWA-P,P
WILD DOG AND LONE, A [NIMNESTL 3]  CRAWFORD, DAN  P
WILD GALAXY  NOLAN, WILLIAM F.  H
WILD HUNT: CHILD OF FIRE, THE [WILD HUNT 2]  FOXE, JOCELIN  P
WILD HUNT OF THE GHOST HOUNDS, THE  LIVELY, PENELOPE  P
WILD HUNT: VENGEANCE MOON, THE [WILD HUNT 1]  FOXE, JOCELIN  P
WILD JACK  CHRISTOPHER, JOHN  H,SFWA-P,P
WILD KINGDOMS [@WARHAMMER]  EARL, ROBERT  C/P,P
WILD MACHINES, THE [@ASH #3]  GENTLE, MARY  C/P,P
WILD MAGIC [WILD MAGIC 1]  CLAYTON, JO  C/P:2,P
WILD MAGIC [FOOL’S GOLD 2]  FISHER, JUDE  H
WILD MAGIC [@IMMORTALS (PIERCE) #1]  PIERCE, TAMORA  P
WILD MAGIC [@GODWARS #3]  WELLS, ANGUS  C/P:2,P
WILD ONES, THE [GRIMES]  CHANDLER, A. BERTRAM  C/P,P
WILD REEL, THE  BRANDON, PAUL  LP
WILD ROAD, THE [WILD ROAD 1]  KING, GABRIEL  C/P,P
WILD ROBERT  JONES, DIANA WYNNE  H
WILD SEA, THE [@BARD #3]  TAYLOR, KEITH  P
WILD SEED  BUTLER, OCTAVIA E.  H:2,SFWA,P
WILD SHORE, THE [THREE CALIFORNIAS]  ROBINSON, KIM STANLEY  C/P:2,LP
WILD SWANS, THE  KERR, PEG  P
WILD TALENT  TUCKER, WILSON  C/P,P,H
WILD TALENTS  FORT, CHARLES  [C/RAN-P:2]
WILD THINGS [@PROWLERS 4]  GOLDEN, CHRISTOPHER  P

WILD THYME AND VIOLETS, OTHER UNPUBLISHED WORKS, AND ADDENDA [@VANCE INTEGRAL EDITION #44] VANCE, JACK

SR-H

WILD WILD WEST, THE  WORMSER, RICHARD  C/P,P
WILD WOOD, THE  DE LINT, CHARLES  LP
WILDCAT  THOMAS, CRAIG  P
WILDEBLOOD’S EMPIRE [DAEDALUS 3]  STABLEFORD, BRIAN M.  C/P,P
WILDER CURSE, THE [@AVENGER #23]  ROBESON, KENNETH  C/P,P
WILDERNESS  DANVERS, DENNIS  C/P,P,H
WILDERNESS OF ICE, THE  VERNE, JULES  H
WILDERNESS OF STARS, A  NOLAN, WILLIAM F.  PA,C/PA
WILDFIRE [@WILD MAGIC #2]  CLAYTON, JO  C/P:2,SFWA-P,P

WILDFIRE [@STAR TREK S.C.E. #6] DECANDIDO, KEITH R. A./MACK, DAVID/ YORK, J. STEVEN/YORK, CHRISTINA F.

P

WILDING, THE  FRIEDMAN, C. S.  H:2
WILDING, THE (PROOF)  TEM, MELANIE  SR-LP
WILDINGS, THE [DANIELLE WALTHERS 2]  CIENCIN, SCOTT  P
WILDINGS OF WESTRON, THE  LAKE, DAVID J.  C/P,P
WILDLIFE  KELLY, JAMES PATRICK  P
WILDRAITH’S LAST BATTLE  KARR, PHYLLIS ANN  C/P:2,P
WILDSIDE  GOULD, STEVEN C.  H,P
WILDSMITH  GOULART, RON  C/P:2,P
WILDWOOD  FARRIS, JOHN  SFWA-P
WILDWOOD DANCING  MARILLIER, JULIET  H
WILK ARE AMONG US, THE  HAIBLUM, ISIDORE  C/P,P,H
WILL I THINK OF YOU?  NIMOY, LEONARD  [C/RAN-P]
WILL-O-THE-WISP  SWANN, THOMAS BURNETT  C/P,P
WILL OF THE EMPRESS, THE [CIRCLE OF MAGIC]  PIERCE, TAMORA  H
WILL OF THE GODS, THE [JALAV 4]  GREEN, SHARON  P
WILL OF THE WANDERER, THE [@ROSE OF THE PROPHET #1]  WEIS, MARGARET/HICKMAN, TRACY  C/P:2,P

WILL THE LAST PERSON TO LEAVE THE PLANET PLEASE SHUT OFF THE SUN RESNICK, MIKE

H

WILLARD  GILBERT, STEPHEN  P

WILLIAM SHATNER’S PERSONAL ACCOUNT OF THE MAKING OF STAR TREK 5 [STAR TREK] SHATNER, LISABETH

R/REF-P,C/REF-H

WILLING SPIRIT, THE  ANTHONY, PIERS/TELLA, ALFRED  H,P
WILLOW  DREW, WAYLAND  C/P,P
WILLOW FILES, THE, VOL. 1 [@BUFFY]  NAVARRO, YVONNE  P
WILLOW FILES, THE, VOL. 2 [@BUFFY]  NAVARRO, YVONNE  P
WILLOW GARDEN, THE [TEN THOUSAND WILLOWS 3]  GRANT, KATHRYN  C/P,P
WIN, LOSE, DRAW [KURTIS:P385XL47:RUTH 2]  STAMEY, SARA  P

The MITSFS Titledex  September 24, 2007  Page 411
TITLE  AUTHOR  CODE
WIN, LOSE, OR DIE [JAMES BOND 22]  GARDNER, JOHN (BRITISH, 1926-)  C/P,P
WIND AFTER TIME, THE [@SHADOW WARRIOR #1]  BUNCH, CHRIS  C/P,P
WIND CHILD  MELUCH, R. M.  C/P,P
WIND CRYSTAL, THE [WESTRIA 6]  PAXSON, DIANA L.  C/P,SFWA-P,P
WIND DANCERS  MELUCH, R. M.  C/P,P
WIND FROM A BURNING WOMAN, THE  BEAR, GREG  C/P:2,P
WIND FROM A FOREIGN SKY [TIELMARAN 1]  REIMANN, KATYA  H,P
WIND FROM NOWHERE, THE  BALLARD, J. G.  C/P:2,P
WIND FROM THE ABYSS [SILISTRA 3]  MORRIS, JANET  C/P:3,P
WIND FROM THE SUN, THE  CLARKE, ARTHUR C.  C/P,H:2,P
WIND IN CAIRO, A  TARR, JUDITH  C/P,P
WIND IN THE ASHES [ASHES 6]  JOHNSTONE, WILLIAM W.  C/P,P
WIND IN THE DOOR, A [MEG MURRY 2]  L’ENGLE, MADELEINE  C/P,LP,H
WIND IN THE ROSE-BUSH, THE  FREEMAN, MARY WILKINS  P
WIND IN THE STONE  NORTON, ANDRE  C/P,P,H
WIND IN THE WILLOWS, THE  GRAHAME, KENNETH  P
WIND OF HONOR [FIVE RINGS FOUR WINDS #1]  SOESBEE, REE  P
WIND OF JUSTICE [@FIVE RINGS FOUR WINDS #3]  WULFF, RICH  P
WIND OF LIBERTY, THE  BULMER, KENNETH  SR-P
WIND OF TRUTH [@FIVE RINGS FOUR WINDS #4]  SOESBEE, REE  P
WIND OF WAR [FIVE RINGS FOUR WINDS #2]  LEBOW, JESS  P
WIND OVER STONEHENGE  DORRE, PAMELA  SR-P
WIND RIDER’S OATH (HARDCOVER AND CD) [HRADANI 3]  WEBER, DAVID  H:2
WIND RIDER’S OATH DISK, THE (CD) [HRADANI]  BAEN PUBLISHING ENTERPRISES  SR-MM
WIND SINGER, THE [WIND ON FIRE 1]  NICHOLSON, WILLIAM  H
WIND WHALES OF ISHMAEL, THE  FARMER, PHILIP JOSE  C/P:2,P
WIND WHISPERS, SHADOW SHOUTS [ALEXIA AND TIRAN 3]  GREEN, SHARON  C/P,P
WIND-WITCH, THE [WARHORSE OF ESDRAGON 2]  DEXTER, SUSAN  C/P:2,P
WINDFALL [WEATHER WARDEN 4]  CAINE, RACHEL  P
WINDHAVEN  MARTIN, GEORGE R. R./TUTTLE, LISA  H:2,P
WINDMASTER’S BANE [OTHERWORLD 1]  DEITZ, TOM  C/P:2,P

WINDOW OF OPPORTUNITY GINGRICH, NEWT/DRAKE, DAVID/

GINGRICH, MARRIANNE

SFWA-P

WINDOW ON THE FUTURE  HILL, DOUGLAS  SR-PA
WINDOWS  COMPTON, D. G.  H,SFWA,P
WINDOWS INTO TOMORROW  SILVERBERG, ROBERT  HA,PA
WINDOWS ON A LOST WORLD [STAR TREK]  MITCHELL, V. E.  C/P:2,P
WINDRIDER, THE  HOPPE, STEPHANIE T.  P
WINDS OF ALTAIR, THE  BOVA, BEN  C/P:2,P,LP,H
WINDS OF CHANGE [MAGE WINDS 2]  LACKEY, MERCEDES  C/P,P,H
WINDS OF CHANGE [@BLOOD OF TEN CHIEFS #3]  PINI, RICHARD  PA,C/PA,SFWA-P,SFWA
WINDS OF CHANGE AND OTHER STORIES, THE  ASIMOV, ISAAC  C/P,P,H:2
WINDS OF DARKOVER, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H,D89250
WINDS OF FATE [MAGE WINDS 1]  LACKEY, MERCEDES  C/P,P,H
WINDS OF FURY [MAGE WINDS 3]  LACKEY, MERCEDES  H:2,SFWA
WINDS OF GATH, THE [DUMAREST 1]  TUBB, E. C.  C/P,DH027,D89301
WINDS OF GATH, THE (WITH "FRESH GUY") [DUMAREST 1]  TUBB, E. C.  LP
WINDS OF LIMBO, THE  MOORCOCK, MICHAEL  C/P:2,P
WINDS OF THE HELIOPOLIS  FRETLAND, DON J.  SR-P
WINDS OF THE WILD SEA [@AGE OF CONAN MARAUDERS #2]  MARIOTTE, JEFF  P
WINDS OF TIME, THE  OLIVER, CHAD  H:2,P
WIND’S TWELVE QUARTERS, THE  LE GUIN, URSULA K.  H:2,P,SFWA
WINDSINGERS, THE [KI AND VANDIEN 2]  LINDHOLM, MEGAN  C/P,P
WINDSOUND  VALLEJO, DORIS  C/P,P
WINDWALKER [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 3]  CUNNINGHAM, ELAINE  P
WINE OF DREAMS, THE [WARHAMMER]  CRAIG, BRIAN  P
WINE OF THE DREAMERS  MACDONALD, JOHN D.  C/P,P,H
WINE OF VIOLENCE, THE  MORROW, JAMES  C/P,P,H
WINE, WOMEN, AND WAR [@SIX MILLION DOLLAR MAN #1]  JAHN, MICHAEL  C/P,P
WING COMMANDER [WING COMMANDER]  TELEP, PETER  P
WINGED ASSASSIN, THE [ARRIS 1]  COOKE, CATHERINE  P
WINGED ISIS [ISIS 4]  STEWART, JEAN  LP
WINGED MAGIC [DARK HORSE 4]  HERBERT, MARY H.  P
WINGED MAN, THE  VAN VOGT, A. E./HULL, E. MAYNE  C/P,P,H
WINGED PHARAOH  GRANT, JOAN  H,P
WINGMAN [WINGMAN 1]  MALONEY, MACK  P
WINGMASTER  HOUSTON, DAVID  P
WINGS IN THE NIGHT  HOWARD, ROBERT E.  LP
WINGS OF FLAME  SPRINGER, NANCY  C/P,P,H
WINGS OF MERLIN [LOST YEARS OF MERLIN 5]  BARRON, T. A.  P
WINGS OF OMEN [@THIEVES’ WORLD #6]  ASPRIN, ROBERT LYNN/ABBEY, LYNN  PA,C/PA
WINGS OF PEGASUS, THE [PEGASUS 1,2]  MCCAFFREY, ANNE  H
WINGS OF POWER  CARL, LILLIAN STEWART  C/P:2,P
WINGS OF THE EAGLE, THE  DORER, NANCY/DORER, FRANCES  P

The MITSFS Titledex  September 24, 2007  Page 412
TITLE  AUTHOR  CODE
WINGS OVER THE ROCKIES  NEWCOMB, AMBROSE  SR-H
WINGS TO THE KINGDOM [EDEN MOORE]  PRIEST, CHERIE  LP
WINNER TAKES ALL [HAWK & FISHER 2]  GREEN, SIMON R.  P
WINNER TAKES ALL [DOCTOR WHO NINTH DOCTOR]  RAYNER, JACQUELINE  H
WINNERS  ANDERSON, POUL  P
WINNIE ILLE PU  MILNE, A. A.  FH
WINNIE THE POOH  MILNE, A. A.  H
WINNING COLORS [HERIS SERRANO 3]  MOON, ELIZABETH  C/P:2,SFWA-P,P
WINNING SCHEHEREZADE  GOROG, JUDITH  SFWA
WINNING SIDE, THE [@TIME HUNTER 1]  PARKIN, LANCE  LP

WINNOWING, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #2] FLEMING, GHERBOD

P

WINTER HAWK  THOMAS, CRAIG  P
WINTER IN EDEN [EDEN 2]  HARRISON, HARRY  H:2,P
WINTER KING, THE  CARL, LILLIAN STEWART  C/P:2,P
WINTER KNIGHT [@BLACK OAK 3]  GRANT, CHARLES L.  P

WINTER MOON LACKEY, MERCEDES/LEE, TANITH/

MURPHY, C. E.

LP

WINTER OAK, THE [MAUREEN PIERCE 2]  HETLEY, JAMES A.  LP
WINTER OF FIRE  JORDAN, SHERRYL  P
WINTER OF MAGIC’S RETURN [MAGIC’S RETURN 1]  SERVICE, PAMELA F.  H,P
WINTER OF THE WORLD, THE  ANDERSON, POUL  H:2,P
WINTER PRINCE, THE  WEIN, ELIZABETH E.  P
WINTER QUEEN, THE  CARY, DEVIN  P
WINTER RECKONING  BRENNAN, NOEL-ANNE  SR-H
WINTER ROSE  MCKILLIP, PATRICIA A.  P
WINTER TIDES  BLAYLOCK, JAMES P.  P
WINTER WAKE  HAUTALA, RICK  P
WINTER WARRIORS [DRENAI]  GEMMELL, DAVID  P
WINTER WISH AND OTHER POEMS, A  LOVECRAFT, H. P.  SR-H
WINTER WORLD [WINTER WORLD 1]  MILLS, C. J.  C/P,P
WINTERHEIM [@DRAGONLANCE ICEWALL #3]  NILES, DOUGLAS  P
WINTERKING [FINNBRANCH 3]  HAZEL, PAUL  H,P
WINTERLONG [WINTERLONG 1]  HAND, ELIZABETH  C/P:2,P
WINTERMIND  KAYE, MARVIN/GODWIN, PARKE  H:2,SFWA,P
WINTER’S DAUGHTER  WHITMORE, CHARLES  P
WINTER’S HEART [WHEEL OF TIME 9]  JORDAN, ROBERT  H:2,P
WINTER’S TALE  HELPRIN, MARK  C/P,P
WINTER’S TALES  DINESEN, ISAK  H
WINTERSMITH [DISCWORLD TIFFANY ACHING 3]  PRATCHETT, TERRY  H:2
WINTERTIDE  SINCLAIR, LINNEA  P
WIPEOUT [@I SPY 4]  TIGER, JOHN  P
WIR FANDEN MENSCHEN  WORNER, HANS  FH
WIRED FOR CHAOS  RENWICK, BRETT L.  LP
WIRED HARD  TAN, CECILIA  SR-L
WISDOM OF THE FOX [GERIN THE FOX (REVISED) 1,2]  TURTLEDOVE, HARRY  C/P,P
WISDOM OF WAR, THE [@BUFFY]  GOLDEN, CHRISTOPHER  C/P,P
WISDOM TEETH [TEETH #1,2,3]  HUCKELBERY, TIMOTHY LEE  WRD
WISDOM’S DAUGHTER [AYESHA 3]  HAGGARD, H. RIDER  P
WISE WOMAN, THE  MACDONALD, GEORGE  P
WISH YOU WERE HERE  HOLT, TOM  P
WISHBRINGER  GARDNER, CRAIG SHAW  C/P:2,P
WISHING GARDEN, THE  YORKE, CHRISTY  P
WISHING SEASON  FRIESNER, ESTHER M.  C/P:2,P
WISHSONG OF SHANNARA, THE [SHANNARA 3]  BROOKS, TERRY  H:2,P
WITCH [WARRIOR 3]  MCQUINN, DONALD E.  C/P,LP
WITCH AND THE CATHEDRAL, THE [WIZARD OF YURT 4]  BRITTAIN, C. DALE  P
WITCH AND WOMBAT  CUSHMAN, CAROLYN  C/P:2,P
WITCH-BAITER, THE  BIRKIN, CHARLES  PA
WITCH BANE  NEILL, ROBERT  P
WITCH DOCTOR, THE [WIZARD IN RHYME 3]  STASHEFF, CHRISTOPHER  C/P,P,H
WITCH FAMILY, THE  ESTES, ELEANOR  LP
WITCH FANTASTIC  RESNICK, MIKE/GREENBERG, MARTIN H.  C/PA,PA
WITCH FINDER [@WARHAMMER WITCH HUNTER 1]  WERNER, C. L.  P
WITCH FIRE  BAST, ANYA  P
WIT’CH FIRE [BANNED AND THE BANISHED 1]  CLEMENS, JAMES  C/P:2,LP
WIT’CH GATE [BANNED AND THE BANISHED 4]  CLEMENS, JAMES  P
WITCH GODDESS, THE [HORSECLANS 9]  ADAMS, ROBERT  C/P:2,P
WITCH HILL [INHERITOR 2]  BRADLEY, MARION ZIMMER  C/P:2,SFWA-P,P
WITCH HOUSE  WALTON, EVANGELINE  SR-P,C/P,H
WITCH HUNTER [@WARHAMMER]  WERNER, C. L.  P
WITCH HUNTERS, THE [DOCTOR WHO PAST DOCTORS]  LYONS, STEVE  C/P:2,P
WITCH IN HEAT  WILLIAMS, J. X.  P
WITCH KILLER [WARHAMMER]  WERNER, C. L.  P

The MITSFS Titledex  September 24, 2007  Page 413
TITLE  AUTHOR  CODE
WITCH MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1,2] ARMSTRONG, KELLEY  H
WITCH OF RHOSTSHYL [SILVERGLASS 3]  RIVKIN, J. F.  C/P,P
WITCH OF THE DARK GATE  JAKES, JOHN  C/P:2,P
WITCH OF THE INDIES, THE [BLACK VULMEA]  SMITH, DAVID C.  P
WITCH POWER  FOREST, SALAMBO  SR-L
WITCH QUEEN, THE [FERN CHAPEL 3]  SIEGEL, JAN  H
WITCH QUEEN OF LOCHLANN [ARWIN 2]  SMITH, GEORGE H.  C/P,P
WITCH QUEEN OF MONGO, THE [FLASH GORDON 5]  BINGHAM, CARSON  C/P:2,P
WIT’CH STAR [BANNED AND THE BANISHED 5]  CLEMENS, JAMES  P
WIT’CH STORM [BANNED AND THE BANISHED 2]  CLEMENS, JAMES  C/P,LP
WIT’CH WAR [BANNED AND THE BANISHED 3]  CLEMENS, JAMES  H,LP
WITCH WEEK [CHRESTOMANCI 2]  JONES, DIANA WYNNE  H,LP
WITCH, WITCH, BURNING BRIGHT [SIXTH SENSE 1]  BLOCH, JOHN W.  C/P,P
WITCH-WOMAN: A TRILOGY ABOUT HER, THE  CABELL, JAMES BRANCH  H
WITCH WOOD  BUCHAN, JOHN  H,P
WITCH WORLD [WITCH WORLD ESTCARP]  NORTON, ANDRE  C/P:2,P
WITCHBLOOD  SHETTERLY, WILL  C/P,P
WITCHCRAFT  SEABROCK, WILLIAM  [C/RAN-P]
WITCHCRAFT READER, THE  HAINING, PETER  HA
WITCHDAME  SKY, KATHLEEN  C/P:2,P
WITCHERY [@GHOSTS OF ALBION]  BENSON, AMBER/GOLDEN, CHRISTOPHER  LP

WITCHES ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

PA,C/PA

WITCHES, THE  DAHL, ROALD  P

WITCHES & WIZARDS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

HA

WITCHES ABROAD [DISCWORLD 12/DISCWORLD WITCHES 3]  PRATCHETT, TERRY  C/P,SFWA-P,P
WITCHES’ BREW [LANDOVER 5]  BROOKS, TERRY  H:2,LP
WITCHES OF CHISWICK, THE [WITCHES OF CHISWICK 1]  RANKIN, ROBERT  P
WITCHES OF EASTWICK, THE  UPDIKE, JOHN  C/P,P,H
WITCHES OF EILEANAN, THE [WITCHES OF EILEANAN 1]  FORSYTH, KATE  C/P:2,P
WITCHES OF KARRES, THE [WITCHES OF KARRES 1]  SCHMITZ, JAMES H.  H:2,P
WITCHES OF KREGEN [DRAY PRESCOT 34]  AKERS, ALAN BURT  P
WITCHES OF TURNSTONE BAY, THE  BURNS, ALMA  P
WITCHES OF WENSHAR, THE [SUNWOLF AND STARHAWK 2]  HAMBLY, BARBARA  C/P:2,P
WITCHES, WRAITHS AND WARLOCKS  CURRAN, RONALD  C/PA,PA
WITCHFINDERS, THE  COMER, RALPH  SR-P
WITCHFIRES OF LETH, THE  GREENBERG, DAN  SFWA-P
WITCHING HOUR, THE  GUNN, JAMES E.  P
WITCHING HOUR, THE [MAYFAIR WITCHES 1]  RICE, ANNE  H:2,P
WITCHING OF DRACULA, THE [@DRACULA #6]  LORY, ROBERT  P
WITCHLIGHT  BRADLEY, MARION ZIMMER  H:2
WITCHLING [OTHERWORLD]  GALENORN, YASMINE  C/P,P
WITCHOPOLY [@SABRINA THE TEENAGE WITCH #22]  VORNHOLT, JOHN  P
WITCH’S BUSINESS  JONES, DIANA WYNNE  LP
WITCH’S COMPANION, THE [MAYFAIR WITCHES]  RAMSLAND, KATHERINE  R/REF-H
WITCH’S DAUGHTER, THE [YNIS AIELLE 2]  SALVATORE, R. A.  P
WITCH’S EYE [@PRIVATE SCHOOL #3]  CHARLES, STEVEN  P
WITCH’S HEAD, THE  HAGGARD, H. RIDER  H
WITCH’S SUCKLING  HALL, GIMONE  P
WITCH’S WELCOME, A [ERIC BRIGHTEYES 2]  SKALDASPILLIR, SIGFRIOUR  P
WITCHSTONE, THE  GRAHAM, VICTORIA  P
WITCHWOOD  LUKEMAN, TIM  H:2
WITCHWOOD CRADLE, THE [TWELVE KINGDOMS 3]  FRIESNER, ESTHER M.  P
WITH A FINGER IN MY I  GERROLD, DAVID  C/P:2,P
WITH A SINGLE SPELL [ETHSHAR]  WATT-EVANS, LAWRENCE  C/P:2,P
WITH A STRANGE DEVICE  RUSSELL, ERIC FRANK  P
WITH A TANGLED SKEIN [INCARNATIONS OF IMMORTALITY 3]  ANTHONY, PIERS  H:2,P,SFWA
WITH FATE CONSPIRE [@DESTINY MAKERS 1]  SHUPP, MIKE  C/P:2,P
WITH FRIENDS LIKE THESE...  FOSTER, ALAN DEAN  C/P:2,P
WITH FULL HONORS [LANNAT 1]  MCQUINN, DONALD E.  P
WITH HER BODY  GRIFFITH, NICOLA  LP

WITH MERCY TOWARD NONE [KING WITHOUT A THRONE 2/ RAGNARSON 2] COOK, GLEN

C/P:2,P

WITH RED HANDS [VIOLETS 2]  WOODWORTH, STEPHEN  H
WITH STARS IN MY EYES  WESTON, PETER  R/REF-H
WITH THE LIGHTNINGS [RCN 1]  DRAKE, DAVID  H:2,P
WITHER  PASSARELLA, J. G.  H
WITHERWING  JARRETT, DAVID  C/P:2,P
WITHOUT A TRACE  BERLITZ, CHARLES  [C/RAN-P]
WITHOUT ABSOLUTION  CASIL, AMY STERLING  H,LP
WITHOUT FEATHERS  ALLEN, WOODY  [C/RAN-P]
WITHOUT REMORSE  CLANCY, TOM  P
WITHOUT WARNING  MICHAELS, FERN  P

The MITSFS Titledex  September 24, 2007  Page 414
TITLE  AUTHOR  CODE
WITLING, THE  VINGE, VERNOR  C/P:2,P
WITNESS OF GOR [GOR 26]  NORMAN, JOHN  H
WIZ BIZ, THE [WIZ ZUMWALT 1,2]  COOK, RICK  C/P,P
WIZARD [GAEA 2]  VARLEY, JOHN  H:2,P,SFWA
WIZARD, THE [WIZARD KNIGHT 2]  WOLFE, GENE  H
WIZARD ABROAD, A [YOUNG WIZARDS 4]  DUANE, DIANE  C/P,P
WIZARD ALONE, A [YOUNG WIZARDS 6]  DUANE, DIANE  H
WIZARD AND A WARLORD, A [ROGUE WIZARD 7]  STASHEFF, CHRISTOPHER  H
WIZARD AND GLASS [@DARK TOWER #4]  KING, STEPHEN  C/P,LP,H
WIZARD AND THE FLOATING CITY, THE  ROWLEY, CHRISTOPHER  P
WIZARD AND THE WAR MACHINE, THE [WAR SURPLUS 2]  WATT-EVANS, LAWRENCE  C/P,P
WIZARD AND THE WARLORD, THE [ALFAR 4]  BOYER, ELIZABETH H.  C/P:2,P
WIZARD AT HOME, THE [@SEVEN TOWERS]  SHELLEY, RICK  P
WIZARD AT LARGE [LANDOVER 3]  BROOKS, TERRY  C/P,H:2,P
WIZARD AT MECQ, THE [@SEVEN TOWERS]  SHELLEY, RICK  P
WIZARD FANTASTIC  GREENBERG, MARTIN H.  PA
WIZARD HUNTERS, THE [FALL OF ILE-RIEN 1]  WELLS, MARTHA  H:2
WIZARD IN A FEUD, A [ROGUE WIZARD 9]  STASHEFF, CHRISTOPHER  H,P
WIZARD IN ABSENTIA, A [WARLOCK’S HEIRS 1]  STASHEFF, CHRISTOPHER  C/P:2,SFWA-P:2,P
WIZARD IN BEDLAM, A [ROGUE WIZARD 2]  STASHEFF, CHRISTOPHER  H:2,SFWA,P
WIZARD IN CHAOS, A [ROGUE WIZARD 5]  STASHEFF, CHRISTOPHER  H
WIZARD IN MIDGARD, A [ROGUE WIZARD 6]  STASHEFF, CHRISTOPHER  H
WIZARD IN MIND, A [ROGUE WIZARD 1]  STASHEFF, CHRISTOPHER  P
WIZARD IN PEACE, A [ROGUE WIZARD 4]  STASHEFF, CHRISTOPHER  P
WIZARD IN THE WAY, A [ROGUE WIZARD 8]  STASHEFF, CHRISTOPHER  H
WIZARD IN WAITING, THE [PELMEN THE POWERSHAPER 2]  HUGHES, ROBERT DON  C/P:2,P
WIZARD IN WAR, A [ROGUE WIZARD 3]  STASHEFF, CHRISTOPHER  H,P
WIZARD KING, THE [CAITHAN CRUSADE 4]  SMITH, JULIE DEAN  P
WIZARD LORD, THE [CHOSEN 1]  WATT-EVANS, LAWRENCE  H
WIZARD OF AMBERMERE, THE [AMBERMERE 3]  PIERCE, J. CALVIN  P
WIZARD OF ANHARITTE, THE  KAPP, COLIN  C/P:2,P
WIZARD OF BONES [AELWYN 3]  CHARRETTE, ROBERT N.  P
WIZARD OF CAMELOT, THE [WYRDRUNE 0]  HAWKE, SIMON  C/P:2,P
WIZARD OF EARTHSEA, A [EARTHSEA 1]  LE GUIN, URSULA K.  C/P,P,H
WIZARD OF 4TH STREET, THE [WYRDRUNE 1]  HAWKE, SIMON  C/P:2,P
WIZARD OF ISIS [ISIS 5]  STEWART, JEAN  LP

WIZARD OF KARRES, THE [WITCHES OF KARRES 2] LACKEY, MERCEDES/FLINT, ERIC/

FREER, DAVE

H

WIZARD OF LEMURIA, THE  CARTER, LIN  SR-P,C/P:2
WIZARD OF LINN, THE  VAN VOGT, A. E.  C/P:2,P
WIZARD OF LONDON, THE [ELEMENTAL MASTERS 4]  LACKEY, MERCEDES  H:2
WIZARD OF LOVECRAFT’S CAFE, THE [WYRDRUNE 7]  HAWKE, SIMON  C/P:2,P
WIZARD OF OZ, THE [OZ #1]  BAUM, L. FRANK  H,@P,@C/P:2
WIZARD OF RENTORO [@BLADE #28]  LORD, JEFFREY  P
WIZARD OF RUE MORGUE, THE [WYRDRUNE 4]  HAWKE, SIMON  C/P:2,P
WIZARD OF SANTA FE, THE [WYRDRUNE 6]  HAWKE, SIMON  C/P,P
WIZARD OF STARSHIP POSEIDON, THE  BULMER, KENNETH  DF209
WIZARD OF STORMS  VAN ARNAM, DAVE  C/P,P
WIZARD OF SUNSET STRIP, THE [WYRDRUNE 3]  HAWKE, SIMON  C/P:2,SFWA-P,P
WIZARD OF THE PIGEONS  LINDHOLM, MEGAN  C/P,P
WIZARD OF THE WINDS [TIMURA 1]  COLE, ALLAN  LP
WIZARD OF TIZARE [GUARDIANS OF THE THREE 3]  COSTELLO, MATTHEW J.  C/P:2,P
WIZARD OF VENUS, THE [VENUS #5]  BURROUGHS, EDGAR RICE  @P
WIZARD OF VENUS, THE/PIRATE BLOOD  BURROUGHS, EDGAR RICE  C/P:2
WIZARD OF WHITECHAPEL, THE [WYRDRUNE 2]  HAWKE, SIMON  C/P:2,P
WIZARD OF ZAO, THE  CARTER, LIN  C/P:2,P
WIZARD SPAWN [SWORD OF KNOWLEDGE 2]  CHERRYH, C. J./ASIRE, NANCY  C/P:2,SFWA-P:2,P
WIZARD WAR [@WIZARD WAR #1]  COOK, HUGH  C/P:2,P
WIZARD WORLD [POL DETSON 1,2]  ZELAZNY, ROGER  C/P:2,P
WIZARDBORN [RUNELORDS 3]  FARLAND, DAVID  P
WIZARDRY COMPILED, THE [WIZ ZUMWALT 2]  COOK, RICK  C/P:2,SFWA-P:2,P
WIZARDRY CONSULTED, THE [WIZ ZUMWALT 4]  COOK, RICK  C/P:2,P
WIZARDRY CURSED, THE [WIZ ZUMWALT 3]  COOK, RICK  C/P:2,SFWA-P,P
WIZARDRY QUESTED, THE [WIZ ZUMWALT 5]  COOK, RICK  P

WIZARDS ASIMOV, ISAAC/GREENBERG, MARTIN H./

WAUGH, CHARLES G.

C/PA,PA

WIZARDS  DANN, JACK/DOZOIS, GARDNER  HA
WIZARDS  WILLIS, JENNIFER SCHWAMM  LPA
WIZARDS & WARRIORS [WORLDS OF POWER 5]  MILES, ELLEN  P
WIZARDS AND WARLOCKS  GHIDALIA, VIC  C/PA,PA
WIZARDS AT WAR [YOUNG WIZARDS 8]  DUANE, DIANE  H
WIZARD’S BANE [WIZ ZUMWALT 1]  COOK, RICK  C/P:2,SFWA-P,P

WIZARD’S BETRAYAL [@DRAGONLANCE NEW ADVENTURES TRINISTYR #2] SAMPSON, JEFF LP

The MITSFS Titledex  September 24, 2007  Page 415
TITLE  AUTHOR  CODE
WIZARDS’ CONCLAVE [@DRAGONLANCE AGE OF MORTALS]  NILES, DOUGLAS  P
WIZARD’S CURSE [@DRAGONLANCE NEW ADVENTURES TRINISTYR #1]  WOODS, CHRISTINA  LP
WIZARD’S DAUGHTER, THE  MICHAELS, BARBARA  P
WIZARD’S DILEMMA, THE [YOUNG WIZARDS 5]  DUANE, DIANE  H
WIZARD’S ELEVEN [TRUE GAME 3]  TEPPER, SHERI S.  C/P,P
WIZARD’S FATE, THE [@DRAGONLANCE ERGOTH #2]  THOMPSON, PAUL B./COOK, TONYA C.  P
WIZARD’S FIRST RULE [SWORD OF TRUTH 1]  GOODKIND, TERRY  C/P:2,SFWA-P,P
WIZARD’S FUNERAL [@RED PAVILIONS #2]  HUNTER, KIM  P
WIZARD’S HEIR [FANUILH 2]  HOOD, DANIEL  P
WIZARD’S HOLIDAY [YOUNG WIZARDS 7]  DUANE, DIANE  H
WIZARD’S ISLE [@COLLECTED STORIES JACK WILLIAMSON #3]  WILLIAMSON, JACK  H
WIZARD’S MOLE  STRICKLAND, BRAD  P
WIZARDS OF ODD, THE  HAINING, PETER  C/PA,PA
WIZARDS OF SENCHURIA, THE [KEYS TO THE DIMENSIONS 4]  BULMER, KENNETH  D12140
WIZARD’S RETURN [@DRAGONLANCE NEW ADVENTURES TRINISTYR #3]  WILLIS, DAN  LP
WIZARD’S SHADOW, THE  DEXTER, SUSAN  P
WIZARD’S TREASURE, THE [DRAGON NIMBUS 4]  RADFORD, IRENE  P
WIZARD’S WAR  HILSON, PHOENIX  SR-P
WIZARD’S WARD, THE [MAURA WOODBURY 1]  HALE, DEBORAH  LP
WIZARDS’ WORLDS  NORTON, ANDRE  C/P,P,H
WIZARDWAR, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #3]  CUNNINGHAM, ELAINE  P
WIZENBEAK [WIZENBEAK 1]  GILLILAND, ALEXIS A.  C/P:2,LP
WOAD TO WUIN, THE [SIR APROPOS OF NOTHING 2]  DAVID, PETER  H,P
WOKEN FURIES [TAKESHI KOVACS 3]  MORGAN, RICHARD  H
WOLF AND IRON  DICKSON, GORDON R.  H:2,P
WOLF AND RAVEN [@SHADOWRUN 32]  STACKPOLE, MICHAEL A.  P
WOLF AND THE CROWN, THE  ATTANASIO, A. A.  C/P,LP
WOLF AND THE RAVEN, THE [WODAN’S CHILDREN 1]  PAXSON, DIANA L.  H,P
WOLF AT THE DOOR [OTHERS (WARREN) 1]  WARREN, CHRISTINE  P
WOLF BELL, THE [YNELL 2]  MURPHY, SHIRLEY ROUSSEAU  C/P,P,H
WOLF BROTHER [ANCIENT DARKNESS 1]  PAVER, MICHELLE  H
WOLF CAPTURED [FIREKEEPER 4]  LINDSKOLD, JANE  H
WOLF-DREAMS [BLOODFANG 1]  WEAVER, MICHAEL D.  P
WOLF HUNTERS [@MECHWARRIOR DARK AGE/BATTLETECH]  KILLIANY, KEVIN  P
WOLF HUNTING [FIREKEEPER 5]  LINDSKOLD, JANE  H
WOLF IN NIGHT [WOLFWALKER 7]  HARPER, TARA K.  P
WOLF IN SHADOW [@STONES OF POWER #3]  GEMMELL, DAVID  P
WOLF IN THE FOLD [HAWK & FISHER 4]  GREEN, SIMON R.  C/P:2,P
WOLF IN THE GARDEN, THE  BILL, ALFRED H.  C/P,P
WOLF JUSTICE [REANDN 2]  DURGIN, DORANNA  P
WOLF KING, THE [WOLF RICE 3]  BORCHARDT, ALICE  P
WOLF MOON  DE LINT, CHARLES  C/P:2,P
WOLF MOON  HOLT, JOHN R.  P
WOLF OF SHADOWS  STRIEBER, WHITLEY  C/P,P,H
WOLF OF WINTER, THE  VOLSKY, PAULA  C/P,LP,H
WOLF PACK [BATTLETECH 4]  CHARRETTE, ROBERT N.  C/P,P
WOLF RIDERS [WARHAMMER]  PRINGLE, DAVID  SR-LP
WOLF-SPEAKER [@IMMORTALS (PIERCE) #2]  PIERCE, TAMORA  P
WOLF STAR [@TOUR OF THE MERRIMACK #2]  MELUCH, R. M.  H
WOLF TIME  WALKER, LARS  P
WOLF WHO RULES [TINKER 2]  SPENCER, WEN  H:2
WOLF WORLDS, THE [STEN 2]  COLE, ALLAN/BUNCH, CHRIS  C/P:2,P
WOLFBANE  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
WOLFBANE (REVISED)  POHL, FREDERIK/KORNBLUTH, C. M.  P
WOLFBLADE [HYTHRUN WOLFBLADE 1]  FALLON, JENNIFER  H
WOLFBLADE [@WARHAMMER 40K SPACE WOLF 4]  KING, WILLIAM  P
WOLFEN, THE  STRIEBER, WHITLEY  P
WOLFHEAD  HARNESS, CHARLES L.  C/P:2,P
WOLFHELM [DRAGONREALM 3]  KNAAK, RICHARD A.  P
WOLFKING [WOLFKINGS]  WOOD, BRIDGET  C/P,LP
WOLFLING  DICKSON, GORDON R.  C/P,P,H
WOLFMAN, THE  DREADSTONE, CARL  P
WOLF’S BANE [WOLFWALKER 4]  HARPER, TARA K.  C/P,P
WOLF’S BLOOD [FIREKEEPER 6]  LINDSKOLD, JANE  H
WOLF’S BROTHER  LINDHOLM, MEGAN  C/P,P
WOLF’S HEAD, WOLF’S HEART [FIREKEEPER 2]  LINDSKOLD, JANE  P
WOLF’S HOUR, THE  MCCAMMON, ROBERT R.  SR-P,C/P:2
WOLFSBANE  JOHNSTONE, WILLIAM W.  P
WOLFSBANE [DOCTOR WHO PAST DOCTORS]  RAYNER, JACQUELINE  P
WOLFSKIN  MARILLIER, JULIET  H
WOLFSKLAUW  POHL, FREDERIK/KORNBLUTH, C. M.  FP

WOLFSONG [@BLOOD OF TEN CHIEFS #2] PINI, RICHARD/ASPRIN, ROBERT LYNN/

ABBEY, LYNN

SFWA-P,C/PA,LPA

WOLFWALKER [WOLFWALKER 1]  HARPER, TARA K.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 416
TITLE  AUTHOR  CODE
WOLFWINTER  SWANN, THOMAS BURNETT  C/P:2,P
WOLLHEIM’S WORLD’S BEST SF SERIES ONE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA
WOLLHEIM’S WORLD’S BEST SF SERIES FOUR  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA
WOLLHEIM’S WORLD’S BEST SF SERIES NINE  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA
WOLVERINE  CLAREMONT, CHRIS/MILLER, FRANK  CX
WOLVERINE’S DAUGHTER  DURGIN, DORANNA  P
WOLVES OF AAM, THE  CURRY, JANE LOUISE  H
WOLVES OF AUTUMN, THE  CIENCIN, SCOTT  P
WOLVES OF DARKNESS [@COLLECTED STORIES JACK WILLIAMSON #2]  WILLIAMSON, JACK  H

WOLVES OF GOD AND OTHER FEY STORIES, THE BLACKWOOD, ALGERNON/ WILSON, WILFRED

SR-H

WOLVES OF MEMORY, THE  EFFINGER, GEORGE ALEC  H:2,P
WOLVES OF THE CALLA [@DARK TOWER #5]  KING, STEPHEN  H:2
WOLVES OF THE GODS [TIMURA 2]  COLE, ALLAN  P
WOLVES ON THE BORDER [BATTLETECH 25]  CHARRETTE, ROBERT N.  P
WOMAN A DAY, A  FARMER, PHILIP JOSE  SR-P,C/P:2
WOMAN ALIVE  ERTZ, SUSAN  H
WOMAN BETWEEN THE WORLDS, THE  MACINTYRE, F. GWYNPLAINE  LP
WOMAN CLOTHED WITH THE SUN AND OTHER STORIES, THE  LUCAS, F. L.  H
WOMAN FROM ANOTHER PLANET  LONG, FRANK BELKNAP  SR-L
WOMAN OF DESTINY, A  CARD, ORSON SCOTT  P
WOMAN OF FLOWERS, THE [HEIRS TO BYZANTIUM 2]  SHWARTZ, SUSAN  C/P,P
WOMAN OF THE FUTURE, A  IRELAND, DAVID  P
WOMAN OF THE HORSECLANS, A [HORSECLANS 12]  ADAMS, ROBERT  P
WOMAN OF THE IRON PEOPLE, A [WOMAN OF THE IRON PEOPLE 1,2]  ARNASON, ELEANOR  H
WOMAN ON THE EDGE OF TIME  PIERCY, MARGE  C/P,P
WOMAN SPACE  NEW VICTORIA PUBLISHERS  SR-LPA
WOMAN WHO LOVED REINDEER, THE  PIERCE, MEREDITH ANN  C/P,SFWA-P,P
WOMAN WHO LOVED THE MOON AND OTHER STORIES, THE  LYNN, ELIZABETH A.  C/P:2,P
WOMAN WHO RIDES LIKE A MAN, THE [SONG OF THE LIONESS 3]  PIERCE, TAMORA  H,P
WOMAN WITHOUT A SHADOW  HABER, KAREN  P

WOMEN AT WAR BUJOLD, LOIS MCMASTER/ GREEN, ROLAND J.

LPA

WOMEN OF DARKNESS  PTACEK, KATHRYN  HA,PA,SFWA-P
WOMEN OF DARKNESS II  PTACEK, KATHRYN  SR-HA
WOMEN OF WONDER  SARGENT, PAMELA  C/PA:2,PA
WOMEN OF WONDER: THE CLASSIC YEARS  SARGENT, PAMELA  LPA
WOMEN OF WONDER: THE CONTEMPORARY YEARS  SARGENT, PAMELA  HA,LPA
WOMEN WHO WALK THROUGH FIRE, THE  STURGIS, SUSANNA J.  LPA
WOMEN WRITING SCIENCE FICTION AS MEN  RESNICK, MIKE  PA
WONDER EFFECT, THE  POHL, FREDERIK/KORNBLUTH, C. M.  C/P:2,P
WONDER TALES  DUNSANY, LORD  LP
WONDER WAR, THE  JANIFER, LAURENCE M.  C/P:2,P
WONDERBOOK AND TANGLEWOOD TALES  HAWTHORNE, NATHANIEL  H
WONDERFLOWER OF UTIK [@PERRY RHODAN #105]  ACE  PA
WONDERFUL ADVENTURES OF PHRA THE PHOENICIAN, THE  ARNOLD, EDWIN LESTER  H,LP
WONDERFUL FLIGHT TO THE MUSHROOM PLANET, THE  CAMERON, ELEANOR  P
WONDERFUL ICE CREAM SUIT AND OTHER PLAYS, THE  BRADBURY, RAY  C/P,P
WONDERFUL VISIT, THE  WELLS, H. G.  H
WONDERFUL WIZARD OF OZ, THE [OZ 1]  BAUM, L. FRANK  H:3,LP
WONDERFUL WORLD OF ROBERT SHECKLEY, THE  SHECKLEY, ROBERT  P
WONDERLAND [DOCTOR WHO]  CHADBOURN, MARK  SR-H
WONDERMAKERS  HOSKINS, ROBERT  C/PA,PA
WONDERMAKERS 2  HOSKINS, ROBERT  PA
WONDER’S CHILD: MY LIFE IN SCIENCE FICTION  WILLIAMSON, JACK  SFWA
WONDROUS BEGINNINGS  SILVER, STEVEN H./GREENBERG, MARTIN H. C/PA,PA
WONDROUS STRANGE  SPRIGGS, ROBIN  LP
WONG’S LOST AND FOUND EMPORIUM AND OTHER ODDITIES  WU, WILLIAM F.  SR-H
WOOD BEYOND THE WORLD, THE  MORRIS, WILLIAM  C/P,P,H
WOOD BOY, THE/BURNING MAN, THE  FEIST, RAYMOND E./WILLIAMS, TAD  CX
WOOD NYMPH AND THE CRANKY SAINT, THE [WIZARD OF YURT 2]  BRITTAIN, C. DALE  C/P,SFWA-P,P
WOOD WIFE, THE  WINDLING, TERRI  P
WOODEN HEART [@DOCTOR WHO]  DAY, MARTIN  H
WOODEN SEA, THE  CARROLL, JONATHAN  H
WOODEN SPACESHIPS, THE [TOLLER MARAQUINE 2]  SHAW, BOB  C/P:2,SFWA-P,P
WOODEN STAR, THE  TENN, WILLIAM  C/P:2,P
WOODEN SWORD, THE [WOODEN SWORD 1]  ABBEY, LYNN  SFWA-P:2
WOODS OUT BACK, THE [SPEARWIELDER 1]  SALVATORE, R. A.  C/P:2,SFWA-P,P
WOODSMAN [ORGANIC FUTURE]  EASTON, THOMAS A.  C/P:2,P
WORD-BRINGER  LLEWELLYN, EDWARD  P
WORD FOR WORLD IS FOREST, THE  LE GUIN, URSULA K.  C/P:2,P
WORD ORIGINS  HUNT, CECIL  [C/RAN-H]
WORDCHANGER  HAYNES, MARY  P
WORDS OF BLOOD [WARHAMMER 40K]  GASCOIGNE, MARC/DUNN, CHRISTIAN  PA
WORDS OF SCIENCE  ASIMOV, ISAAC  [C/RAN-P]

The MITSFS Titledex  September 24, 2007  Page 417
TITLE  AUTHOR  CODE
WORDS WITHOUT PICTURES  NILES, STEVE  LPA
WORDWRIGHT [SOULSMITH 3]  DEITZ, TOM  P,SFWA-P
WORF’S FIRST ADVENTURE [@STAR TREK TNG STARFLEET ACADEMY #1]  DAVID, PETER  H:3,LP
WORK OF JACK WILLIAMSON, THE  HAUPTMANN, RICHARD A.  R/REF-H
WORK OF THE SUN, THE [CELYDONN GREEN LION 3]  EDGERTON, TERESA  P,SFWA-P:2
WORKING FOR THE DEVIL [DANTE VALENTINE 1]  SAINTCROW, LILITH  P
WORKING OF STARS, A [MAGEWORLDS 7]  DOYLE, DEBRA/MACDONALD, JAMES D.  H,P
WORKS OF EDGAR ALLAN POE 1, THE  POE, EDGAR ALLAN  SR-H
WORKS OF EDGAR ALLAN POE 2, THE  POE, EDGAR ALLAN  SR-H
WORKS OF EDGAR ALLAN POE 3, THE  POE, EDGAR ALLAN  SR-H
WORKS OF EDGAR ALLAN POE 4, THE  POE, EDGAR ALLAN  SR-H
WORKS OF EDGAR ALLAN POE 5, THE  POE, EDGAR ALLAN  SR-H
WORKS OF H. G. WELLS  WELLS, H. G.  H
WORKS OF JULES VERNE  VERNE, JULES  H
WORKS OF M. P. SHIEL, THE  MORSE, A. REYNOLDS  R/REF-H
WORLD AFLAME  BULMER, H. K.  SR-P
WORLD AFLAME  ENGEL, LEONARD/PILLER, EMANUEL S.  SR-H
WORLD AFLAME, A [@CAP KENNEDY #13]  KERN, GREGORY  P
WORLD AND THORINN, THE  KNIGHT, DAMON  H:2,P,SFWA
WORLD ASUNDER, THE  WALLACE, IAN  C/P,P
WORLD AT BAY, THE  CAPON, PAUL  H:2
WORLD AT BAY  TUBB, E. C.  SR-P
WORLD AT THE END OF TIME, THE  POHL, FREDERIK  H:2,LP
WORLD BEFORE, THE [WESS’HAR WARS 3]  TRAVISS, KAREN  P
WORLD BELOW, THE  WRIGHT, S. FOWLER  SR-H,GN1.05
WORLD BELOW, THE (WITH "THE AMPHIBIANS")  WRIGHT, S. FOWLER  LP
WORLD BETWEEN, A  SPINRAD, NORMAN  C/P:2,P
WORLD BETWEEN AND OTHER STORIES, THE  VANCE, JACK  DM125
WORLD BEYOND THE HILL, THE  PANSHIN, ALEXEI/PANSHIN, CORY  R/REF-H
WORLD BY THE TALE, A  CAMPBELL, JOHN W. JR.  C/PA,PA
WORLD CALLED CAMELOT, A [CAMELOT 1]  LANDIS, ARTHUR H.  C/P,P
WORLD CALLED SOLITUDE, A  GOLDIN, STEPHEN  H,SFWA,P
WORLD ENDS IN HICKORY HOLLOW, THE  MAYHAR, ARDATH  H,SFWA
WORLD ENOUGH, AND TIME  KAHN, JAMES  C/P:2,P
WORLD FALL [SEVEN CIRCLES 2]  NILES, DOUGLAS  LP
WORLD FANTASY AWARDS 2, THE  SCHIFF, STUART DAVID/LEIBER, FRITZ  SFWA
WORLD GAME [DOCTOR WHO PAST DOCTORS]  DICKS, TERRANCE  P
WORLD GONE MAD, A [@PERRY RHODAN #29]  ACE  C/PA,PA
WORLD GRABBERS, THE  FAIRMAN, PAUL W.  P
WORLD I LEFT BEHIND ME, THE  WALLING, WILLIAM H.  H
WORLD IN 2030, THE  BIRKENHEAD, EARL OF  SR-H
WORLD IN AMBER, THE [WORLD IN AMBER 1]  ORR, A.  H:2,SFWA-P,P
WORLD IN DARKNESS, THE  ROBERTS, MURRAY  B
WORLD IN ECLIPSE  DEXTER, WILLIAM  C/P:2,P
WORLD IN SPACE, THE  MARSHACK, ALEXANDER  [C/RAN-P:2]
WORLD IN WINTER, THE  CHRISTOPHER, JOHN  P
WORLD INSIDE, THE  SILVERBERG, ROBERT  H:2,SFWA,P
WORLD INVISIBLE, THE  OPPENHEIM, SHULAMITH  C/P:2,P
WORLD IS ROUND, THE  ROTHMAN, TONY  C/P:2,P
WORLD JONES MADE, THE  DICK, PHILIP K.  C/P:2,P,DD150
WORLD KILLERS [@SENTINELS #4/ROBOTECH 16]  MCKINNEY, JACK  C/P,P
WORLD LOST, A  JOHNSON, JAMES B.  P
WORLD MAKERS, THE  MAXWELL, JOHN C.  H
WORLD MENDERS, THE [CULTURAL SURVEY 2]  BIGGLE, LLOYD JR.  C/P:2,SFWA,P
WORLD NAMED CLEOPATRA, A  ELWOOD, ROGER  PA,C/PA
WORLD NEXT DOOR, THE  FERGUSON, BRAD  C/P:2,SFWA-P,P
WORLD OF A THOUSAND COLORS  SILVERBERG, ROBERT  H:2,P
WORLD OF DIFFERENCE, A  CONQUEST, ROBERT  C/P:2,P
WORLD OF DIFFERENCE, A  TURTLEDOVE, HARRY  C/P:2,P
WORLD OF FANZINES: A SPECIAL FORM OF COMMUNICATION, THE  WERTHAM, FREDRIC  R/REF-H
WORLD OF FLYING SAUCERS, THE  MENZEL, DONALD H./BOYD, LYLE G.  [C/RAN-H]
WORLD OF GOL  CHARLES, NEIL  SR-P
WORLD OF HURT, A [@STARFIST 10]  SHERMAN, DAVID/CRAGG, DAN  P
WORLD OF HURT, A (PROOF) [@STARFIST 10]  SHERMAN, DAVID/CRAGG, DAN  SR-LP
WORLD OF IF  PHILLIPS, ROG  SR-P
WORLD OF MAZES [CENOTAPH ROAD 3]  VARDEMAN, ROBERT E.  C/P,P
WORLD OF NULL-A, THE  VAN VOGT, A. E.  SR-DD031,P,H,C/P:2
WORLD OF PROMISE [DUMAREST 23]  TUBB, E. C.  C/P,P
WORLD OF PTAVVS [KNOWN SPACE]  NIVEN, LARRY  C/P:2,P

WORLD OF ROBERT JORDAN’S THE WHEEL OF TIME, THE [WHEEL OF TIME] JORDAN, ROBERT/PATTERSON, TERESA

R/REF-H

WORLD OF SCIENCE FICTION 1926–1976:

THE HISTORY OF A SUBCULTURE, THE DEL REY, LESTER

R/REF-H,C/REF-H

WORLD OF STAR TREK, THE  GERROLD, DAVID  R/REF-P,C/REF-P:2

The MITSFS Titledex  September 24, 2007  Page 418
TITLE  AUTHOR  CODE
WORLD OF THE DARK CRYSTAL, THE  FROUD, BRIAN  A
WORLD OF THE MASTERMINDS  WILLIAMS, ROBERT MOORE  DD427
WORLD OF THE SLEEPER  WAYMAN, TONY RUSSELL  DH021
WORLD OF THE STARWOLVES [@STARWOLF #3]  HAMILTON, EDMOND  C/P,P
WORLD OF TIERS 1, THE [WORLD OF TIERS 1,2]  FARMER, PHILIP JOSE  H:2
WORLD OF TIERS 2, THE [WORLD OF TIERS 3,4,5]  FARMER, PHILIP JOSE  H:2
WORLD OF TOMORROW  ZEIGFREID, KARL  SR-P
WORLD OF TROUBLE, A  TOOMEY, ROBERT E. JR.  C/P:2,P
WORLD OF WONDER  PRATT, FLETCHER  HA
WORLD OUT OF MIND  MCINTOSH, J. T.  H:2,P
WORLD OUT OF TIME, A  NIVEN, LARRY  C/P,P,H
WORLD SET FREE, THE  WELLS, H. G.  H,LP
WORLD SHUFFLER, THE [LAFAYETTE O’LEARY 2]  LAUMER, KEITH  H:2,P
WORLD SOUL  EMTSEV, MIKHAIL/PARNOV, EREMEI  H:2,LP
WORLD SPIRITS [KHYREN]  BOUCHER-KAPLAN, ALINE  C/P:2,SFWA-P,P
WORLD SWAPPERS, THE  BRUNNER, JOHN  C/P:2,DD391,P
WORLD THAT COULDN’T BE, THE  GOLD, HORACE L.  PA,C/PA
WORLD THAT NEVER WAS, THE  ZEIGFREID, KARL  SR-P

WORLD THINKER AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #2] VANCE, JACK

SR-H

WORLD TREASURY OF SCIENCE FICTION, THE  HARTWELL, DAVID G.  HA:2,LPA
WORLD TREE, THE [DAUGHTERS OF BAST 3]  ISIDORE, SARAH  P
WORLD TURNED UPSIDE DOWN, THE  DRAKE, DAVID/FLINT, ERIC/BAEN, JIM  HA
WORLD UNKNOWN, A  CLAGETT, JOHN  C/P,P
WORLD WAR III  STANLEY, JOHN  P
WORLD WAR Z  BROOKS, MAX  H
WORLD WEAVER, THE  ETCHISON, CRAIG  P
WORLD WELL LOST  AIKEN, JOHN  H,SFWA
WORLD WITHIN, THE  LUKENS, ADAM  H

WORLD WITHOUT CHILDREN/EARTH QUARTER, THE (TWO DAMON KNIGHT SCIENCE FICTION NOVELS)

KNIGHT, DAMON

C/P:2,P

WORLD WITHOUT END  COCHRAN, MOLLY/MURPHY, WARREN  H,P
WORLD WITHOUT END [STAR TREK]  HALDEMAN, JOE  C/P:2,P
WORLD WITHOUT END  MACKENZIE, NIGEL  H
WORLD WITHOUT END [MOONTIDE AND MAGIC RISE 1]  RUSSELL, SEAN  C/P,P
WORLD WITHOUT MEN  MAINE, CHARLES ERIC  C/P:2,P
WORLD WITHOUT MERCY [PERRY RHODAN 131]  VOLTZ, WILLIAM  SR-LP
WORLD WITHOUT STARS  ANDERSON, POUL  C/P:2,P
WORLD WITHOUT WOMEN  KEENE, DAY/PRUYN, LEONARD G.  C/P,P
WORLD WITHOUT WOMEN, THE  MARTINI, VIRGILIO  H
WORLD WRECKERS, THE [DARKOVER]  BRADLEY, MARION ZIMMER  C/P,P,H
WORLDCON GUEST OF HONOR SPEECHES  RESNICK, MIKE/SICLARI, JOE  R/REF-H
WORLDMAKER  ELLIS, A. C.  C/P,P
WORLDMAKERS  DOZOIS, GARDNER  LPA
WORLDS [WORLDS 1]  HALDEMAN, JOE  C/P,SFWA,P,H

WORLDS APART DECARNIN, CAMILLA/GARBER, ERIC/

PALEO, LYN

LPA

WORLDS APART [WORLDS 2]  HALDEMAN, JOE  C/P,P,SR-H,H
WORLDS APART  KELLEY, LEO P.  SR-P
WORLDS APART  LOCKE, GEORGE  HA
WORLDS APART  MCINTOSH, J. T.  C/P:2,P
WORLDS AT WAR  TEMPEST BOOKS  SR-PA
WORLDS AWAY!  BEEMISH, CRAGG  SR-P
WORLDS BEGINNING  ARDREY, ROBERT  H
WORLD’S BEST SCIENCE FICTION 1965  WOLLHEIM, DONALD A./CARR, TERRY  PA,C/PA
WORLD’S BEST SCIENCE FICTION 1966  WOLLHEIM, DONALD A./CARR, TERRY  PA,C/PA
WORLD’S BEST SCIENCE FICTION 1967  WOLLHEIM, DONALD A./CARR, TERRY  PA,C/PA
WORLD’S BEST SCIENCE FICTION 1968  WOLLHEIM, DONALD A./CARR, TERRY  PA,C/PA
WORLD’S BEST SCIENCE FICTION 1969  WOLLHEIM, DONALD A./CARR, TERRY  HA,PA
WORLD’S BEST SCIENCE FICTION 1970  WOLLHEIM, DONALD A./CARR, TERRY  C/PA,PA
WORLD’S BEST SCIENCE FICTION 1971  WOLLHEIM, DONALD A./CARR, TERRY  C/PA,PA
WORLD’S BEST SCIENCE FICTION FIRST SERIES  WOLLHEIM, DONALD A./CARR, TERRY  C/PA,PA
WORLD’S BEST SF 7  WOLLHEIM, DONALD A./SAHA, ARTHUR W.  PA
WORLDS BEYOND  RIDPATH, IAN  SFWA-VLP

WORLDS BEYOND THE WORLD:

THE FANTASTIC VISION OF WILLIAM MORRIS MATHEWS, RICHARD

R/REF-P

WORLD’S DESIRE, THE  HAGGARD, H. RIDER/LANG, ANDREW  P
WORLD’S END [AGE OF MISRULE 1]  CHADBOURN, MARK  P
WORLDS’ END [@SANDMAN #8]  GAIMAN, NEIL  H,CX
WORLD’S END  SINCLAIR, UPTON  H
WORLD’S END [SNOW QUEEN 2]  VINGE, JOAN D.  SFWA-P,H:2,P,SFWA
WORLDS ENOUGH & TIME  SIMMONS, DAN  H,LP
WORLDS ENOUGH AND TIME [WORLDS 3]  HALDEMAN, JOE  C/P,P,H
WORLD’S FAIR, 1992  SILVERBERG, ROBERT  C/P,HA,P
WORLD’S FAIR GOBLIN [DOC SAVAGE #39]  ROBESON, KENNETH  C/P,@P

The MITSFS Titledex  September 24, 2007  Page 419
TITLE  AUTHOR  CODE
WORLDS FOR THE TAKING  BULMER, KENNETH  C/P:2,P
WORLD’S GREATEST ATHLETE, THE  GARDNER, GERALD/CARUSO, DEE  P

WORLDS IN COLLISION [STAR TREK] REEVES-STEVENS, JUDITH/

REEVES-STEVENS, GARFIELD

LP

WORLDS IN COLLISION  VELIKOVSKY, IMMANUEL  [C/RAN-H,C/RAN-P]
WORLDS NEAR AND FAR  CARR, TERRY  HA
WORLDS OF A. E. VAN VOGT, THE  VAN VOGT, A. E.  C/P:3,P
WORLDS OF CLIFFORD SIMAK, THE  SIMAK, CLIFFORD D.  H:2,P
WORLDS OF EXILE AND ILLUSION [HAINISH 1,2,3]  LE GUIN, URSULA K.  LP
WORLDS OF FRANK HERBERT, THE  HERBERT, FRANK  C/P:2,P
WORLDS OF FRITZ LEIBER, THE  LEIBER, FRITZ  C/P:2,P
WORLDS OF GEORGE O.  SMITH, GEORGE O.  C/P,P
WORLDS OF H. BEAM PIPER, THE  PIPER, H. BEAM  C/P,P
WORLDS OF HONOR [HONOR HARRINGTON/WORLDS OF HONOR 2]  WEBER, DAVID  C/PA,HA,PA

WORLDS OF IF POHL, FREDERIK/GREENBERG, MARTIN H./

OLANDER, JOSEPH D.

HA

WORLDS OF JACK VANCE, THE  VANCE, JACK  C/P,P
WORLDS OF MAYBE  SILVERBERG, ROBERT  PA
WORLDS OF POUL ANDERSON, THE  ANDERSON, POUL  C/P:2,P
WORLDS OF ROBERT A. HEINLEIN, THE  HEINLEIN, ROBERT A.  C/P:2,P
WORLDS OF ROBERT F. YOUNG, THE  YOUNG, ROBERT F.  SR-P,H
WORLDS OF SCIENCE FICTION, THE  HIPPLE, THEODORE W./WRIGHT, ROBERT G. LPA
WORLDS OF SCIENCE FICTION, THE  MILLS, ROBERT P.  PA,C/PA
WORLDS OF SHADOW [THREE WORLDS 1,2,3]  WATT-EVANS, LAWRENCE  H
WORLDS OF THE GOLDEN QUEEN [GOLDEN QUEEN 1,2]  FARLAND, DAVID  LP
WORLDS OF THE IMPERIUM [IMPERIUM 1]  LAUMER, KEITH  C/P:2,DF127,SFWA-P,P
WORLDS OF THE IMPERIUM (EXPANDED) [IMPERIUM 1]  LAUMER, KEITH  C/P:2,P
WORLDS OF THE WALL  MACAPP, C. C.  C/P:2,P
WORLDS OF THEODORE STURGEON, THE  STURGEON, THEODORE  C/P:2,P
WORLDS OF TOMORROW  DERLETH, AUGUST  PA,C/PA
WORLDS OF WEIRD  MARGULIES, LEO  C/PA,PA
WORLDS OF WHEN  CONKLIN, GROFF  PA,C/PA
WORLDS OF WOMEN  TAN, CECILIA  SR-L
WORLDS OF WONDER  HARRISON, HARRY  HA

WORLDS THAT WEREN’T TURTLEDOVE, HARRY/STIRLING, S. M./

GENTLE, MARY/WILLIAMS, WALTER JON

H,LP

WORLDS TO COME  KNIGHT, DAMON  C/PA,PA
WORLDS TO CONQUER  STATTEN, VARGO  SR-P
WORLDS VAST AND VARIOUS  BENFORD, GREGORY  LP
WORLDS WITHIN  PHILLIPS, ROG  SR-P
WORLDS WITHIN WORLDS  STARMONT HOUSE, INC.  SR-LPA
WORLDS WITHOUT END  SIMAK, CLIFFORD D.  C/P:2,P
WORLDS WITHOUT END [@SHADOWRUN 18/IMMORTALS 3]  SPECTOR, CAROLINE  C/P,P
WORLD’S WORST WASP JOKES, THE  PERMAN, DONALD/LEIGHT  [C/RAN-P]
WORLDSTONE  STRAUSS, VICTORIA  C/P,P
WORLDWAR: IN THE BALANCE [@WORLDWAR 1]  TURTLEDOVE, HARRY  C/P,P,H:2
WORLDWAR: STRIKING THE BALANCE [@WORLDWAR 4]  TURTLEDOVE, HARRY  C/P,P,H
WORLDWAR: TILTING THE BALANCE [@WORLDWAR 2]  TURTLEDOVE, HARRY  H:2,P
WORLDWAR: UPSETTING THE BALANCE [@WORLDWAR 3]  TURTLEDOVE, HARRY  C/P,P
WORLDWIRED [JENNY CASEY 3]  BEAR, ELIZABETH  P
WORM OUROBOROS, THE [ZIMIAMVIA]  EDDISON, E. R.  H:3,P
WORM RE-TURNS, THE  MCCONNELL, JAMES V.  [C/RAN-H]
WORMS, THE  SARRANTONIO, AL  SFWA
WORMS OF THE EARTH  HOWARD, ROBERT E.  SR-H,P
WORMWOOD (PROOF)  TAYLOR, G. P.  SR-LP
WORSE THINGS WAITING  MCNAUGHTON, BRIAN  LP
WORSE THINGS WAITING  WELLMAN, MANLY WADE  SR-H
WORSHIP OF THE GENERATIVE POWERS DURING THE MIDDLE AGES, THE  WRIGHT, THOMAS  [C/RAN-P]
WORSHIP THE NIGHT  VIGLIANTE, MARY  P
WORTHING CHRONICLE, THE [WORTHING]  CARD, ORSON SCOTT  C/P:2,P
WORTHING SAGA, THE [WORTHING]  CARD, ORSON SCOTT  C/P:2,SFWA-P,P
WOUNDED [LITTLE GODDESS 2]  LANE, AMY  LP
WOUNDED HAWK, THE [@CRUCIBLE 2]  DOUGLASS, SARA  H

WOUNDED KING, THE [@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #3] BOULLE, PHILIPPE

P

WOUNDED LAND, THE [THOMAS COVENANT 4]  DONALDSON, STEPHEN R.  C/P,P,H
WOUNDED PLANET, THE  ELWOOD, ROGER/KIDD, VIRGINIA  PA,C/PA
WOUNDED SKY, THE [STAR TREK]  DUANE, DIANE  C/P:2,P
WOUNDHEALER’S STORY [@LOST SWORDS #1]  SABERHAGEN, FRED  H:2,P
WRACK AND ROLL  DENTON, BRADLEY  C/P,P
WRAETHTU [WRAETHTHU 1,2,3]  CONSTANTINE, STORM  LP
WRAITH [@ZOE MARTINIQUE]  WELDON, PHAEDRA  LP
WRAITH BOARD [GAMING MAGI 2]  BISCHOFF, DAVID F.  C/P,P
WRAITH SQUADRON [@STAR WARS X-WING 5]  ALLSTON, AARON  P

The MITSFS Titledex  September 24, 2007  Page 420
TITLE  AUTHOR  CODE
WRAITH: THE OBLIVION (2ND EDITION) [WORLD OF DARKNESS WRAITH]  DANSKY, RICHARD E.  [C/RAN-H]
WRAITHS OF TIME  NORTON, ANDRE  H:2,P
WRAITHS OF WILL AND PLEASURE, THE [WRAETHTHU HISTORIES 1]  CONSTANTINE, STORM  H
WRAPT IN CRYSTAL  SHINN, SHARON  H,P
WRATH OF ASHAR [@KINGDOMS #1]  WELLS, ANGUS  C/P:2,P
WRATH OF FU MANCHU, THE  ROHMER, SAX  C/P,P
WRATH OF GOD  GLEASON, ROBERT  H,LP
WRATH OF KHAN, THE (PHOTOSTORY) [STAR TREK]  ANOBILE, RICHARD J.  P
WRATH OF POSEIDON, THE [@HERCULES]  BETANCOURT, JOHN GREGORY  P
WRATH OF THE PRINCES [BARD’S TALE 7]  LISLE, HOLLY/ALLSTON, AARON  P

WRATH OF THE PROPHETS [STAR TREK DS9] DAVID, PETER/FRIEDMAN, MICHAEL JAN/

GREENBERGER, ROBERT

C/P,P

WREATH OF STARS, A  SHAW, BOB  C/P,SFWA,SFWA-P,P,H
WRECK OF HEAVEN, THE [WORLD GATES 2]  LISLE, HOLLY  P
WRECK OF THE RIVER OF STARS, THE  FLYNN, MICHAEL  H

WRECKAGE OF PARADISE, THE [@WORLD OF DARKNESS DEMON THE FALLEN #3] STOLZE, GREG

P

WRECKS OF TIME, THE  MOORCOCK, MICHAEL  DH036
WRESTLER’S CRUEL STUDY, THE  DOBYNS, STEPHEN  LP
WRINKLE IN THE SKIN, A  CHRISTOPHER, JOHN  P
WRINKLE IN TIME, A [MEG MURRY 1]  L’ENGLE, MADELEINE  C/P,P,H:2
WRINKLE IN TIME, A (BRITISH) [MEG MURRY 1]  L’ENGLE, MADELEINE  SR-P
WRIT IN BLOOD [SAINT-GERMAIN 11]  YARBRO, CHELSEA QUINN  H
WRITER, THE  HUBBARD, L. RON  R/REF-H
WRITING AND SELLING SCIENCE FICTION  GRANT, CHARLES L.  R/REF-H

WRITING SCIENCE FICTION AND FANTASY DOZOIS, GARDNER/LEE, TINA/

SCHMIDT, STANLEY/ STROCK, IAN RANDALL/ WILLIAMS, SHEILA

R/REF-P

WRITING SCIENCE FICTION AND FANTASY  KILIAN, CRAWFORD  R/REF-H
WRITTEN IN VENOM  TILTON, LOIS  H,LP
WRONG END OF TIME, THE  BRUNNER, JOHN  H:2,P,SFWA
WRONG REFLECTION, THE  BRADSHAW, GILLIAN  P
WRONG SIDE OF THE MOON  ASHTON, FRANCIS/ASHTON, STEPHEN  SR-P
WT50: A TRIBUTE TO WEIRD TALES  WEINBERG, ROBERT  R/REF-H
WULFSYARN  MANN, PHILLIP  C/P,SFWA,SFWA-P,P,H
WUNDER WAR, THE [@MAN-KZIN WARS #10]  COLEBATCH, HAL  H
WYLDER’S HAND  LE FANU, J. SHERIDAN  LP
WYRD SISTERS [DISCWORLD 6/DISCWORLD WITCHES 2]  PRATCHETT, TERRY  C/P,P,H
WYRLDMAKER  BISSON, TERRY  C/P:2,P
WYRM  FABI, MARK  LP
WYRM WOLF [WORLD OF DARKNESS]  VAN BELKOM, EDO  P
WYRMHOLE  CASELBERG, JAY  P
WYRMS  CARD, ORSON SCOTT  SFWA-P,H:2,P
WYSARD  KEPHART, CAROLYN  LP
WYST: ALASTOR 1716 [ALASTOR 3]  VANCE, JACK  C/P:2,P
WYST: ALASTOR 1716 (VIE) [@VANCE INTEGRAL EDITION #31/ALASTOR 3]  VANCE, JACK  SR-H
WYVERN  ATTANASIO, A. A.  H,P
WYVERN’S SPUR, THE [FORGOTTEN REALMS FINDER’S STONE 2]  NOVAK, KATE/GRUBB, JEFF  C/P,P

The MITSFS Titledex  September 24, 2007  Page 421
TITLE  AUTHOR  CODE
X  SUDAK, EUDICE  C/P:2,P
X FACTOR, THE  NORTON, ANDRE  C/P,P,H
X FILES FIGHT THE FUTURE, THE [X-FILES]  CARTER, CHRIS/HAND, ELIZABETH  H,P
X-MACHINE, THE  MULLER, JOHN E.  SR-P

X-MEN [X-MEN] RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY

P

X-MEN THE BROOD SAGA, PART 1 [X-MEN]  CLAREMONT, CHRIS  SCX
X PRESIDENT, THE  BARUTH, PHILIP E.  LP
X PRESIDENT, THE (PROOF)  BARUTH, PHILIP E.  SR-LP

X-TREME POSSIBILITIES [X-FILES] CORNELL, PAUL/DAY, MARTIN/

TOPPING, KEITH

R/REF-P

XANADU  YOLEN, JANE  SR-HA
XANADU 2  YOLEN, JANE  HA,LPA
XANADU 3  YOLEN, JANE  HA,LPA
XANADU TALISMAN, THE [@MODESTY BLAISE]  O’DONNELL, PETER  P,SFWA-P
XANDER YEARS, THE, VOL. 1 [BUFFY]  DECANDIDO, KEITH R. A.  P
XANDER YEARS, THE, VOL. 2 [BUFFY]  MARIOTTE, JEFF  P
XANTHE AND THE ROBOTS  MACLEOD, SHEILA  SR-P
XENOCIDE [ENDER 3]  CARD, ORSON SCOTT  H:3,P
XENOCIDE (PROOF) [ENDER 3]  CARD, ORSON SCOTT  SR-LP
XENOGENESIS [XENOGENESIS 1,2,3]  BUTLER, OCTAVIA E.  H:2
XENOGENESIS  DEFORD, MIRIAM ALLEN  C/P:2,P
XENOS [WARHAMMER 40K EISENHORN 1]  ABNETT, DAN  P
XICCARPH  SMITH, CLARK ASHTON  SR-P,C/P:2
XONE OF CONTENTION [XANTH #23]  ANTHONY, PIERS  H
XORANDOR  BROOKE-ROSE, CHRISTINE  C/P,P
XT CALLED STANLEY, AN  TREBOR, ROBERT  C/P:2,P
XUAN AND THE GIRL FROM THE OTHER SIDE  BERGIN, PAUL A.  C/P:2,P

The MITSFS Titledex  September 24, 2007  Page 423
TITLE  AUTHOR  CODE
YANK AT VALHALLA, A  HAMILTON, EDMOND  D93900
YANKED! [@OUT OF TIME #1]  KRESS, NANCY  P
YARBOROUGH BRAND, THE  HOFFMAN, LEE  C/P,P
YARGO  SUSANN, JACQUELINE  H,P
YARIL’S CHILDREN  BENNETT, MARCIA J.  C/P:2,P
YARROW  DE LINT, CHARLES  C/P:2,P
YEAR 2000, THE  HARRISON, HARRY  C/PA,PA
YEAR 2018!  BLISH, JAMES  C/P,P

YEAR AFTER TOMORROW, THE DEL REY, LESTER/MATSCHAT, CECILE/

CARMER, CARL

HA

YEAR BEFORE YESTERDAY, THE  ALDISS, BRIAN W.  H,SFWA,P
YEAR OF CONSENT  CROSSEN, KENDELL FOSTER  C/P:2,P
YEAR OF INTELLIGENT TIGERS, THE [DOCTOR WHO EIGHTH DOCTOR]  ORMAN, KATE  P
YEAR OF OUR WAR, THE [FOURLANDS 1]  SWAINSTON, STEPH  H
YEAR OF THE CLOUD, THE  THOMAS, THEODORE L./WILHELM, KATE  H,P
YEAR OF THE COMET, THE  CHRISTOPHER, JOHN  LP
YEAR OF THE GRIFFIN [DERKHOLM 2]  JONES, DIANA WYNNE  P
YEAR OF THE LAST EAGLE, THE  RIENOW, LEONA TRAIN/RIENOW, ROBERT  C/P,P
YEAR OF THE LUCY, THE  MCCAFFREY, ANNE  P
YEAR OF THE QUIET SUN, THE  TUCKER, WILSON  C/P:2,SFWA-P,P
YEAR OF THE RANSOM, THE  ANDERSON, POUL  H
YEAR OF THE RATS, THE  RICCI, BARBARA GUIGNON  H
YEAR OF THE UNICORN [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
YEAR OF THE WARRIOR, THE  WALKER, LARS  P
YEAR THE CLOUD FELL, THE [FALLEN CLOUD 1]  GIAMBASTIANI, KURT R. A.  P
YEAR WHEN STARDUST FELL, THE  JONES, RAYMOND F.  H:2
YEAR ZERO  LONG, JEFF  P
YEAR’S BEST FANTASY  HARTWELL, DAVID G.  PA
YEAR’S BEST FANTASY 2  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST FANTASY 3  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
YEAR’S BEST FANTASY 4  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
YEAR’S BEST FANTASY 5  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
YEAR’S BEST FANTASY 6  HARTWELL, DAVID G./CRAMER, KATHRYN  LPA
YEAR’S BEST FANTASY 7  HARTWELL, DAVID G./CRAMER, KATHRYN  LPA
YEAR’S BEST FANTASY, FIRST ANNUAL COLLECTION, THE  DATLOW, ELLEN/WINDLING, TERRI  HA,LPA
YEAR’S BEST FANTASY, SECOND ANNUAL COLLECTION, THE  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, THIRD ANNUAL COLLECTION, THE  DATLOW, ELLEN/WINDLING, TERRI  LPA

YEAR’S BEST FANTASY AND HORROR, FOURTH ANNUAL COLLECTION, THE DATLOW, ELLEN/WINDLING, TERRI

LPA

YEAR’S BEST FANTASY AND HORROR, FIFTH ANNUAL COLLECTION, THE  DATLOW, ELLEN/WINDLING, TERRI  LPA

YEAR’S BEST FANTASY AND HORROR, SEVENTH ANNUAL COLLECTION, THE DATLOW, ELLEN/WINDLING, TERRI

LPA

YEAR’S BEST FANTASY AND HORROR, EIGHTH ANNUAL COLLECTION, THE DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, NINTH ANNUAL COLLECTION, THE  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, TENTH ANNUAL COLLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, ELEVENTH ANNUAL COLLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, TWELFTH ANNUAL COLLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, THIRTEENTH ANNUAL COLLECTION DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, FOURTEENTH ANNUAL COLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, FIFTEENTH ANNUAL COLLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA
YEAR’S BEST FANTASY AND HORROR, SIXTEENTH ANNUAL COLLECTION  DATLOW, ELLEN/WINDLING, TERRI  LPA

YEAR’S BEST FANTASY AND HORROR, SEVENTEENTH ANNUAL COLLECTION, THE DATLOW, ELLEN/LINK, KELLY/

GRANT, GAVIN J.

LPA

YEAR’S BEST FANTASY AND HORROR, EIGHTEENTH ANNUAL COLLECTION, THE DATLOW, ELLEN/LINK, KELLY/

GRANT, GAVIN J.

LPA

YEAR’S BEST FANTASY AND HORROR, NINETEENTH ANNUAL COLLECTION, THE DATLOW, ELLEN/LINK, KELLY/

GRANT, GAVIN J.

LPA

YEAR’S BEST FANTASY STORIES, THE  CARTER, LIN  PA
YEAR’S BEST FANTASY STORIES 2, THE  CARTER, LIN  C/PA,PA
YEAR’S BEST FANTASY STORIES 3, THE  CARTER, LIN  C/PA,PA
YEAR’S BEST FANTASY STORIES 4, THE  CARTER, LIN  PA
YEAR’S BEST FANTASY STORIES 5, THE  CARTER, LIN  PA
YEAR’S BEST FANTASY STORIES 6, THE  CARTER, LIN  PA
YEAR’S BEST FANTASY STORIES 7, THE  SAHA, ARTHUR W.  PA
YEAR’S BEST FANTASY STORIES 8, THE  SAHA, ARTHUR W.  C/PA,PA
YEAR’S BEST FANTASY STORIES 9, THE  SAHA, ARTHUR W.  PA
YEAR’S BEST FANTASY STORIES 10, THE  SAHA, ARTHUR W.  PA
YEAR’S BEST FANTASY STORIES 11, THE  SAHA, ARTHUR W.  PA
YEAR’S BEST FANTASY STORIES 12, THE  SAHA, ARTHUR W.  PA
YEAR’S BEST FANTASY STORIES 13, THE  SAHA, ARTHUR W.  HA,PA
YEAR’S BEST FANTASY STORIES 14, THE  SAHA, ARTHUR W.  C/PA,PA
YEAR’S BEST HORROR STORIES I, THE  DAVIS, RICHARD  C/PA,PA
YEAR’S BEST HORROR STORIES II, THE  DAVIS, RICHARD  PA
YEAR’S BEST HORROR STORIES #2, THE (BRITISH)  DAVIS, RICHARD  SR-PA
YEAR’S BEST HORROR STORIES III, THE  DAVIS, RICHARD  PA

The MITSFS Titledex  September 24, 2007  Page 424
TITLE  AUTHOR  CODE
YEAR’S BEST HORROR STORIES V, THE  PAGE, GERALD W.  PA
YEAR’S BEST HORROR STORIES VI, THE  PAGE, GERALD W.  PA
YEAR’S BEST HORROR STORIES VII, THE  PAGE, GERALD W.  PA
YEAR’S BEST HORROR STORIES VIII, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES IX, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES X, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES XI, THE  WAGNER, KARL EDWARD  C/PA,PA
YEAR’S BEST HORROR STORIES XII, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES XIII, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES XIV, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES XVI, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST HORROR STORIES XVII, THE  WAGNER, KARL EDWARD  PA
YEAR’S BEST S-F 5, THE  MERRIL, JUDITH  PA,C/PA
YEAR’S BEST S-F 6, THE  MERRIL, JUDITH  PA,C/PA
YEAR’S BEST S-F 7, THE  MERRIL, JUDITH  C/PA,PA
YEAR’S BEST S-F 8, THE  MERRIL, JUDITH  PA,C/PA
YEAR’S BEST S-F 9, THE  MERRIL, JUDITH  C/PA,PA
YEAR’S BEST S-F 10, THE  MERRIL, JUDITH  HA,PA
YEAR’S BEST S-F 11, THE  MERRIL, JUDITH  C/PA,PA
SF 12  MERRIL, JUDITH  PA,C/PA
YEAR’S BEST SCIENCE FICTION 1, THE  DOZOIS, GARDNER  LPA
YEAR’S BEST SCIENCE FICTION 2, THE  DOZOIS, GARDNER  LPA
YEAR’S BEST SCIENCE FICTION 3, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 4, THE  DOZOIS, GARDNER  LPA
YEAR’S BEST SCIENCE FICTION 5, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 6, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 7, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 8, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION #8, THE  HARRISON, HARRY/ALDISS, BRIAN W.  SR-PA
YEAR’S BEST SCIENCE FICTION 9, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 10, THE  DOZOIS, GARDNER  HA:2,LPA
YEAR’S BEST SCIENCE FICTION 11, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 12, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 13, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 14, THE  DOZOIS, GARDNER  HA:2,LPA
YEAR’S BEST SCIENCE FICTION 15, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 16, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 17, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 18, THE  DOZOIS, GARDNER  HA,LPA
YEAR’S BEST SCIENCE FICTION 19, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 20, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 21, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 22, THE  DOZOIS, GARDNER  HA
YEAR’S BEST SCIENCE FICTION 23, THE  DOZOIS, GARDNER  LPA
YEAR’S BEST SCIENCE FICTION 24, THE  DOZOIS, GARDNER  LPA

YEAR’S BEST SCIENCE FICTION AND FANTASY FOR TEENS, FIRST ANNUAL COLLECTION, THE YOLEN, JANE/HAYDEN, PATRICK NIELSEN

LPA

YEAR’S BEST SCIENCE FICTION NOVELS 1952  BLEILER, EVERETT F./DIKTY, T. E.  HA
YEAR’S BEST SCIENCE FICTION NOVELS 1954  BLEILER, EVERETT F./DIKTY, T. E.  HA
YEAR’S BEST SF (1995)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 2 (1996)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 3 (1997)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 4 (1998)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 5 (1999)  HARTWELL, DAVID G.  PA,C/PA
YEAR’S BEST SF 6 (2000)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 7 (2001)  HARTWELL, DAVID G.  C/PA,PA
YEAR’S BEST SF 8 (2002)  HARTWELL, DAVID G./CRAMER, KATHRYN  HA,PA
YEAR’S BEST SF 9 (2003)  HARTWELL, DAVID G./CRAMER, KATHRYN  HA
YEAR’S BEST SF 10 (2004)  HARTWELL, DAVID G./CRAMER, KATHRYN  HA
YEAR’S BEST SF 11 (2005)  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
YEAR’S BEST SF 12 (2006)  HARTWELL, DAVID G./CRAMER, KATHRYN  PA
YEAR’S FINEST FANTASY, THE  CARR, TERRY  PA,C/PA
YEAR’S FINEST FANTASY 2, THE  CARR, TERRY  C/PA,PA
YEARS IN THE MAKING  DE CAMP, L. SPRAGUE  H
YEARS OF LIGHT  COLOMBO, JOHN ROBERT  R/REF-P
YEARS OF RICE AND SALT, THE  ROBINSON, KIM STANLEY  H,P
YEARS OF THE CITY, THE  POHL, FREDERIK  C/P,P,H
YEARWOOD [FINNBRANCH 1]  HAZEL, PAUL  C/P:2,P
YELLOW CLAW, THE  ROHMER, SAX  C/P,P
YELLOW CLOUD, THE [DOC SAVAGE #59]  ROBESON, KENNETH  C/P:2,@P
YELLOW EYES [POSLEEN]  RINGO, JOHN/KRATMAN, TOM  H
YELLOW FOG [DON SEBASTIAN 4]  DANIELS, LES  SR-H,SFWA-P
YELLOW FRACTION, THE  GORDON, REX  C/P:2,P
YELLOW HOARD, THE [@AVENGER #2]  ROBESON, KENNETH  C/P:2,P
YELLOW PERIL, THE  SHIEL, M. P.  H

The MITSFS Titledex  September 24, 2007  Page 425
TITLE  AUTHOR  CODE
YELLOW PLANET, THE  BROWNE, GEORGE SHELDON  SR-P
YELLOW SCOURGE, THE [OPERATOR 5 #3]  STEELE, CURTIS  SR-P
YELLOW SILK, THE [@FORGOTTEN REALMS ROGUES]  BASSINGTHWAITE, DON  P
YENDI [DRAGAERA VLAD TALTOS]  BRUST, STEVEN  C/P:3,P
YESTERDAY AND TOMORROW  VERNE, JULES  H,P
YESTERDAY WE SAW MERMAIDS  FRIESNER, ESTHER M.  H
YESTERDAY’S CHILDREN [STAR WOLF 0]  GERROLD, DAVID  C/P:2,P
YESTERDAY’S CHILDREN (REVISED, PROOF) [STAR WOLF 0]  GERROLD, DAVID  SR-LP
YESTERDAY’S CHILDREN (REVISED) [STAR WOLF 0]  GERROLD, DAVID  P
YESTERDAY’S KINGS  WELLS, ANGUS  P
YESTERDAY’S MEN [ETHICAL SOCIETY 3]  TURNER, GEORGE  P
YESTERDAY’S PAWN  QUICK, W. T.  P
YESTERDAY’S SON [STAR TREK]  CRISPIN, A. C.  C/P:2,P
YESTERDAY’S TOMORROWS  POHL, FREDERIK  HA,LPA
YET MORE PENGUIN SCIENCE FICTION  ALDISS, BRIAN W.  SR-PA
YIPPY SKIPPY, THE EVIL! [SLUGGY FREELANCE 5]  ABRAMS, PETE  SCX
YLANA OF CALLISTO [JANDAR 7]  CARTER, LIN  C/P,P
YNGLING, THE [YNGLING 1]  DALMAS, JOHN  C/P:2,P
YNGLING AND THE CIRCLE OF POWER, THE [YNGLING 2]  DALMAS, JOHN  C/P:2,SFWA-P,P
YNGLING IN YAMATO, THE [YNGLING]  DALMAS, JOHN  C/P,P
YOKE, THE  MILLER, ELIZABETH  H
YOKE OF MAGIC, A [SWORDS OF RAEMLLYN 2]  VARDEMAN, ROBERT E./PROCTOR, GEO. W.  C/P,P
YOKE OF SHEN, THE [@SPACEWAYS #10]  CLEVE, JOHN  P
YON ILL WIND [XANTH #20]  ANTHONY, PIERS  H,@P,@C/P
YONDER  BEAUMONT, CHARLES  P
YONDER COMES THE OTHER END OF TIME [COYOTE JONES 5]  ELGIN, SUZETTE HADEN  C/P:2,P
YORATH THE WOLF [RULERS OF HYLOR 2]  WILDER, CHERRY  H,SFWA,P
YOR’S REVENGE [CAPTAINS OUTRAGEOUS 2]  YOUNG, ROY V.  P
YOU AND I  NIMOY, LEONARD  [C/RAN-P]

YOU BET YOUR PLANET GREENBERG, MARTIN H./ KOREN, BRITTANY A.

PA

YOU CAN BE THE STAINLESS STEEL RAT [STAINLESS STEEL RAT]  HARRISON, HARRY  P
YOU ONLY LIVE TWICE [JAMES BOND 12]  FLEMING, IAN  C/P:2,P
YOU SANE MEN  JANIFER, LAURENCE M.  C/P:2,P
YOU SHALL KNOW THEM  VERCORS  H:2,P
YOU SLAY ME [AISLING GREY 4]  MACALISTER, KATIE  C/P,P
YOU WILL NEVER BE THE SAME  SMITH, CORDWAINER  C/P:2,P
YOUNG ADULTS  PINKWATER, DANIEL MANUS  LP
YOUNG BLEYS [CHILDE CYCLE]  DICKSON, GORDON R.  H:2,P
YOUNG DEMONS  ELWOOD, ROGER/GHIDALIA, VIC  PA,C/PA
YOUNG FRANKENSTEIN  PEARLMAN, GILBERT  C/P:2,P
YOUNG LEGIONARY [LAST LEGIONARY 0]  HILL, DOUGLAS  H
YOUNG MAGICIANS, THE  CARTER, LIN  PA
YOUNG RISSA [HULZEIN DYNASTY]  BUSBY, F. M.  C/P:2,P
YOUNG UNICORNS, THE  L’ENGLE, MADELEINE  P
YOUNG VISITOR TO MARS  ELAM, RICHARD M. JR.  H
YOUNG WARRIORS, THE [@BATTLESTAR GALACTICA 1978 #4]  LARSON, GLEN A./THURSTON, ROBERT  P
YOUNGER GODS, THE [@DREAMERS #4]  EDDINGS, DAVID/EDDINGS, LEIGH  H
YOUR COMPLETE GUIDE TO FATE AND FORTUNE  BARONET  [C/RAN-P]
YOUR NEXT FIFTY YEARS  PREHODA, ROBERT W.  R/REF-H,C/REF-P
YOU’RE ALL ALONE  LEIBER, FRITZ  C/P:2,P
YOU’RE HISTORY [@SABRINA THE TEENAGE WITCH SALEM #3]  STRICKLAND, BRAD/STRICKLAND, BARBARA P
YOURS TRULY, FROM HELL  SMITH, TERRENCE LORE  P
YOUTH MONOPOLY, THE  WOBIG, ELLEN  DH048
YSABEL  KAY, GUY GAVRIEL  H
YULETIDE UNIVERSE, A (PROOF)  THOMSEN, BRIAN M.  SR-LPA
YURTH BURDEN  NORTON, ANDRE  C/P:2,P
YVGENIE  CHERRYH, C. J.  H:2,P
YVONNE, THE CONFIDENT [@PHENWICK #26]  KIMBROUGH, KATHERYN  P

The MITSFS Titledex  September 24, 2007  Page 427
TITLE  AUTHOR  CODE
Z EFFECT, THE  LAURENS, MARSHALL  C/P:2,P
Z FOR ZACHARIAH  O’BRIEN, ROBERT C.  C/P,P
Z-LENSMAN [LENSMEN 9]  KYLE, DAVID A.  C/P:2,P
Z RAY, THE  SNELL, EDMUND  H
Z-STING  WALLACE, IAN  C/P,P
ZACHERLEY’S MIDNIGHT SNACKS  ZACHERLEY  PA,C/PA
ZACHERLEY’S VULTURE STEW  ZACHERLEY  PA,C/PA

ZACHERLY FOR PRESIDENT WATSON, DICK, ELLIOTT, JOHN/

STROME, DICK

[C/RAN-H]

ZAMPER [DOCTOR WHO NEW ADVENTURES]  ROBERTS, GARETH  C/P:2,P
ZANDRA  ROTSLER, WILLIAM  H,SFWA

ZANDRU’S FORGE [@CLINGFIRE #2] BRADLEY, MARION ZIMMER/

ROSS, DEBORAH J.

H

ZANTHAR AT MOON’S MADNESS  WILLIAMS, ROBERT MOORE  C/P,P
ZANTHAR AT THE EDGE OF NEVER  WILLIAMS, ROBERT MOORE  P
ZANTHAR AT TRIP’S END  WILLIAMS, ROBERT MOORE  P
ZANTHAR OF THE MANY WORLDS  WILLIAMS, ROBERT MOORE  C/P,P
ZANTHODON [ERIC CARSTAIRS 2]  CARTER, LIN  P
ZANTI MISFITS, THE [@OUTER LIMITS]  PEEL, JOHN  LP
ZANZIBAR CAT, THE  RUSS, JOANNA  C/P,P
ZAP DAY  LINDBLAD, JOHN  P
ZAP GUN, THE  DICK, PHILIP K.  C/P:2,P
ZARAGOZ [WARHAMMER ORFEO 1]  CRAIG, BRIAN  LP
ZARDOZ  BOORMAN, JOHN/STAIR, BILL  C/P,P
ZARSTHOR’S BANE [WITCH WORLD HIGH HALLACK]  NORTON, ANDRE  C/P:2,P
ZAVANT [WARHAMMER ZAVANT KONNIGER]  RENNIE, GORDON  P
ZB4  JACKSON, ANITA  SR-P
ZEALOT [@HIGHLANDER]  LETTOW, DONNA  P
ZEITGEIST  STERLING, BRUCE  H
ZEITSTURM  CUNIS, REINMAR  FP
ZELDE M’TANA [HULZEIN DYNASTY]  BUSBY, F. M.  C/P,P
ZELEROD’S DOOM [SIME/GEN 8]  LICHTENBERG, JACQUELINE/LORRAH, JEAN C/P:2,P
ZEN GUN, THE  BAYLEY, BARRINGTON J.  P
ZENDA VENDETTA, THE [TIMEWARS 4]  HAWKE, SIMON  C/P:2,P
ZENITH ANGLE, THE  STERLING, BRUCE  H,P
ZENITH-D  LORRAINE, PAUL  SR-P
ZENTRAEDI REBELLION, THE [@ROBOTECH #19]  MCKINNEY, JACK  P
ZENYA [DUMAREST 11]  TUBB, E. C.  C/P:2,P
ZERO CITY [@DEATHLANDS]  AXLER, JAMES  P
ZERO COOL  LANGE, JOHN  C/P,P
ZERO FIELD  HUNT, GILL  SR-P
ZERO GRAVITY SWAP  PYGASTER, CAL I.  SR-L
ZERO HOUR  MILLER, BENJAMIN E.  P
ZERO HOUR [@RESIDENT EVIL #0]  PERRY, S. D.  P
ZERO HOUR  STATTEN, VARGO  SR-P
ZERO MINUS X  ZEIGFREID, KARL  SR-P
ZERO POINT [STAR*DRIVE]  BAKER, RICHARD  P
ZERO POINT  LE PAGE, RAND  SR-P
ZERO STONE, THE [JERN MURDOCK 1]  NORTON, ANDRE  H:2,P
ZERO-SUM GAMES [@PLANET BUILDERS #5]  TALLIS, ROBYN  P
ZETA MAJOR [DOCTOR WHO PAST DOCTORS]  MESSINGHAM, SIMON  C/P,P
ZEUS AND CO.  JONES, DAVID LEE  P,SFWA-P
ZHORANI: MASTER OF THE UNIVERSE  MARAS, KARL  SR-P
ZILOV BOMBS, THE  BARRON, D. G.  H
ZIMA BLUE  REYNOLDS, ALASTAIR  H
ZIMIAMVIA [ZIMIAMVIA 1,2,3]  EDDISON, E. R.  LP
ZIP-ZIP AND HIS FLYING SAUCER  SCHEALER, JOHN M.  H
ZO LEVEN WIJ MORGEN  DE LATIL, PIERRE  FP
ZOBOA  CAIDIN, MARTIN  P
ZOD WALLOP  SPENCER, WILLIAM BROWNING  P
ZODIAC  LEES, DAN  SFWA
ZODIAC  STEPHENSON, NEAL  C/P:2,LP
ZODIAC FANTASTIC  GREENBERG, MARTIN H./MORLAN, A. R.  PA
ZOLOTOV AFFAIR, THE  RIMMER, ROBERT H.  C/P,P
ZOMBIE LOVER [XANTH #22]  ANTHONY, PIERS  H
ZOMBIE SURF COMMANDOS FROM MARS! [WEIRD ZONE 1]  ABBOTT, TONY  LP
ZOMBIES OF THE GENE POOL  MCCRUMB, SHARYN  P
ZONE NULL  FRANKE, HERBERT W.  SFWA
ZONE SOLDIERS  RANSOM, DANIEL  C/P,P
ZONE YELLOW [IMPERIUM 4]  LAUMER, KEITH  C/P,P
ZOO 2000  YOLEN, JANE  HA
ZOO DES ASTORS, LE  RAYJEAN, MAX-ANDRE  FP
ZORACHUS [ZORACHUS 1]  ROGERS, MARK E.  P
ZORK CHRONICLES, THE [ZORK]  EFFINGER, GEORGE ALEC  C/P:2,P

TITLE  AUTHOR  CODE
ZOROASTER/KHALED  CRAWFORD, F. MARION  H
ZOTHIQUE [ZOTHIQUE]  SMITH, CLARK ASHTON  SR-P,C/P:2
ZOTZ!  KARIG, WALTER  H:2
ZULU HEART [LION’S BLOOD 2]  BARNES, STEVEN  H,P
ZWISCHEN SCHOPFUNG UND UNTERGANG  NOWAK, JURGEN  FP